The MITSFS Pinkdex  September 24, 2007  Page 1
AUTHOR  TITLE  CODE
"A", DOCTOR  SENSUOUS DIRTY OLD MAN, THE  [C/RAN-H,C/RAN-P]
AAMODT, DONALD  NAME TO CONJURE WITH, A  C/P:2,P
AAMODT, DONALD  TROUBLING ALONG THE BORDER, A  C/P:2,P
AARON, CHESTER  OUT OF SIGHT, OUT OF MIND  C/P,P
AARON, SHALE  VIRTUAL DEATH  C/P:2,P
AARONOVITCH, BEN  ALSO PEOPLE, THE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
AARONOVITCH, BEN  REMEMBRANCE OF THE DALEKS [DOCTOR WHO]  C/P,P
AARONOVITCH, BEN/ORMAN, KATE  SO VILE A SIN [DOCTOR WHO NEW ADVENTURES]  C/P,P
AARONOVITCH, BEN  TRANSIT [DOCTOR WHO NEW ADVENTURES]  C/P,P
AARONS, EDWARD S.  ASSIGNMENT — MOON GIRL  P
AB HUGH, DAFYDD  ARTHUR WAR LORD [ARTHUR WAR LORD 1]  P
AB HUGH, DAFYDD  BALANCE OF POWER [STAR TREK TNG]  C/P:2,P
AB HUGH, DAFYDD  CONQUERED, THE [@STAR TREK DS9 REBELS #1]  P
AB HUGH, DAFYDD  COURAGEOUS, THE [@STAR TREK DS9 REBELS #2]  C/P,P
AB HUGH, DAFYDD/LINAWEAVER, BRAD  ENDGAME [@DOOM #4]  C/P:2,P
AB HUGH, DAFYDD  FALLEN HEROES [STAR TREK DS9]  C/P:2,P
AB HUGH, DAFYDD  FALLEN HEROES (PROOF) [STAR TREK DS9]  SR-VLP
AB HUGH, DAFYDD  FAR BEYOND THE WAVE [@ARTHUR WAR LORD #2]  P
AB HUGH, DAFYDD  FINAL FURY, THE [STAR TREK INVASION! 4/STAR TREK VOYAGER]  C/P,P
AB HUGH, DAFYDD/LINAWEAVER, BRAD  HELL ON EARTH [@DOOM #2]  C/P,P
AB HUGH, DAFYDD  HEROING [JIANA 1]  C/P,P
AB HUGH, DAFYDD/LINAWEAVER, BRAD  INFERNAL SKY [@DOOM #3]  C/P:2,P
AB HUGH, DAFYDD/LINAWEAVER, BRAD  KNEE-DEEP IN THE DEAD [@DOOM #1]  C/P,P
AB HUGH, DAFYDD  LIBERATED, THE [@STAR TREK DS9 REBELS #3]  P
AB HUGH, DAFYDD  VENGEANCE [@STAR TREK DS9]  P
AB HUGH, DAFYDD  WARRIORWARDS [JIANA 2]  C/P:3,P
ABANTE, ABRAHAM  TEN SCI-FI SHORTS  H,LP
ABBEY, LYNN  BEHIND TIME [EMMA MERRIGAN 2]  C/P,P
ABBEY, LYNN  BENEATH THE WEB [WOODEN SWORD 2]  P
ABBEY, LYNN  BLACK FLAME, THE  C/P,LP,H
ABBEY, LYNN  BRAZEN GAMBIT, THE [@DARK SUN ATHAS #1]  P
ABBEY, LYNN/ASPRIN, ROBERT LYNN  CATWOMAN  P
ABBEY, LYNN  CINNABAR SHADOWS [@DARK SUN ATHAS #4]  P
ABBEY, LYNN  CONQUEST [@UNICORN AND DRAGON #2]  LP
ABBEY, LYNN  DAUGHTER OF THE BRIGHT MOON  C/P:2,LP
ABBEY, LYNN  DOWN TIME [EMMA MERRIGAN 4]  P
ABBEY, LYNN  FORGE OF VIRTUE, THE [ULTIMA]  C/P,P
ABBEY, LYNN  GUARDIANS, THE  C/P,P
ABBEY, LYNN  JERLAYNE  C/P,P
ABBEY, LYNN  NETHER SCROLL, THE [@FORGOTTEN REALMS LOST EMPIRES]  P
ABBEY, LYNN  OUT OF TIME [EMMA MERRIGAN 1]  C/P,P
ABBEY, LYNN  PLANESWALKER [MAGIC: THE GATHERING ARTIFACTS 2]  P
ABBEY, LYNN  RIFKIND’S CHALLENGE  H
ABBEY, LYNN  RISE AND FALL OF A DRAGON KING, THE [@DARK SUN ATHAS #5]  P
ABBEY, LYNN  SANCTUARY [THIEVES’ WORLD]  H
ABBEY, LYNN  SIEGE OF SHADOWS  C/P,P
ABBEY, LYNN  SIMBUL’S GIFT, THE [@FORGOTTEN REALMS NOBLES #6]  P
ABBEY, LYNN  TAKING TIME [EMMA MERRIGAN 3]  P
ABBEY, LYNN  TEMPER OF WISDOM, THE [ULTIMA]  C/P,P
ABBEY, LYNN  TURNING POINTS [THIEVES’ WORLD]  HA
ABBEY, LYNN  TURNING POINTS (PROOF) [THIEVES’ WORLD]  SR-LPA
ABBEY, LYNN  UNICORN AND DRAGON [@UNICORN AND DRAGON #1]  H:2,LP
ABBEY, LYNN  WOODEN SWORD, THE [WOODEN SWORD 1]  SFWA-P:2
ABBOTT, CARL  FRONTIERS PAST AND FUTURE  R/REF-H
ABBOTT, EDWIN A.  FLATLAND [FLATLAND]  H:2,LP
ABBOTT, TONY  ZOMBIE SURF COMMANDOS FROM MARS! [WEIRD ZONE 1]  LP
ABE, KOBO  INTER ICE AGE 4  H:2,P
ABERLY, RACHEL/ENGEL, VOLKER  MAKING OF INDEPENDENCE DAY, THE  R/REF-H
ABNETT, DAN  ARMOUR OF CONTEMPT, THE [@WARHAMMER 40K]  H
ABNETT, DAN/LEE, MIKE  BLOODSTORM [@WARHAMMER MALUS DARKBLADE 2]  P
ABNETT, DAN  BORN OF BLOOD [WARHAMMER DARKBLADE #1]  CX
ABNETT, DAN  BROTHERS OF THE SNAKE [@WARHAMMER 40K]  H
ABNETT, DAN/LEE, MIKE  DAEMON’S CURSE, THE [@WARHAMMER MALUS DARKBLADE 1]  P
ABNETT, DAN  DOUBLE EAGLE [@WARHAMMER 40K]  P
ABNETT, DAN  EISENHORN [@WARHAMMER 40K EISENHORN 1,2,3]  H,LP
ABNETT, DAN  FELL CARGO [@WARHAMMER]  P
ABNETT, DAN  FIRST AND ONLY [WARHAMMER 40K GAUNT’S GHOSTS 1]  P
ABNETT, DAN  FOUNDING, THE [@WARHAMMER 40K GAUNT’S GHOSTS]  LP
ABNETT, DAN  GHOSTMAKER [WARHAMMER 40K GAUNT’S GHOSTS 2]  P
ABNETT, DAN/VINCENT, NIK  GILEAD’S BLOOD [WARHAMMER GILEAD]  P
ABNETT, DAN  GUNS OF TANITH, THE [@WARHAMMER 40K GAUNT’S GHOSTS #5]  P
ABNETT, DAN/VINCENT, NIK/WALLIS, JAMES HAMMERS OF ULRIC [WARHAMMER]  P
ABNETT, DAN  HERETICUS [WARHAMMER 40K EISENHORN 3]  P

The MITSFS Pinkdex  September 24, 2007  Page 2
AUTHOR  TITLE  CODE
ABNETT, DAN  HIS LAST COMMAND [@WARHAMMER 40K GAUNT’S GHOSTS]  H
ABNETT, DAN  HONOUR GUARD [WARHAMMER 40K GAUNT’S GHOSTS 4]  P
ABNETT, DAN  HORUS RISING [@WARHAMMER 40K HORUS HERESY 1]  P
ABNETT, DAN  HUNT FOR DEFAY, THE [WARHAMMER INQUISITOR ASCENDANT]  CX
ABNETT, DAN  MALLEUS [WARHAMMER 40K EISENHORN 2]  P
ABNETT, DAN  NECROPOLIS [WARHAMMER 40K GAUNT’S GHOSTS 3]  P
ABNETT, DAN  RAVENOR [@WARHAMMER 40K RAVENOR 1]  P
ABNETT, DAN  RAVENOR RETURNED [@WARHAMMER 40K RAVENOR 2]  H
ABNETT, DAN  RAVENOR ROGUE [@WARHAMMER 40K RAVENOR 3]  H
ABNETT, DAN/LEE, MIKE  REAPER OF SOULS [@WARHAMMER MALUS DARKBLADE]  P
ABNETT, DAN  RIDERS OF THE DEAD [@WARHAMMER]  H
ABNETT, DAN  SABBAT MARTYR [@WARHAMMER 40K GAUNT’S GHOSTS]  P
ABNETT, DAN  STRAIGHT SILVER [@WARHAMMER 40K GAUNT’S GHOSTS]  P
ABNETT, DAN  TRAITOR GENERAL [@WARHAMMER 40K GAUNT’S GHOSTS]  P
ABNETT, DAN/LEE, MIKE  WARPSWORD [@WARHAMMER MALUS DARKBLADE]  P
ABNETT, DAN  XENOS [WARHAMMER 40K EISENHORN 1]  P
ABRAHAM, DANIEL  SHADOW IN SUMMER, A [LONG PRICE 1]  H
ABRAMOV, A./ABRAMOV, S.  HORSEMEN FROM NOWHERE  H
ABRAMS, PETE  VERY BIG BANG!, A [SLUGGY FREELANCE]  CX
ABRAMS, PETE  YIPPY SKIPPY, THE EVIL! [SLUGGY FREELANCE 5]  SCX
ABRAMS, R. VAUGHN  PARA  H
ACE  ACTION: DIVISION 3 [@PERRY RHODAN #94]  C/PA,PA
ACE  AGAIN: ATLAN! [@PERRY RHODAN #46]  PA,C/PA
ACE  AMBASSADORS FROM AURIGEL, THE [@PERRY RHODAN #64]  PA,C/PA
ACE  ATLAN IN DANGER [@PERRY RHODAN #82]  C/PA,PA
ACE  ATOM HELL OF GRAUTIER, THE [@PERRY RHODAN #71]  PA
ACE  ATTACK FROM THE UNSEEN [@PERRY RHODAN #50]  PA,C/PA
ACE  BEWARE THE MICROBOTS [@PERRY RHODAN #35]  PA,C/PA
ACE  BLAZING SUN [@PERRY RHODAN #86]  C/PA:2,PA
ACE  BLOCKADE: LEPSO [@PERRY RHODAN #101]  PA
ACE  BLUE DWARFS, THE [@PERRY RHODAN #54]  PA,C/PA
ACE  BLUE SYSTEM, THE [@PERRY RHODAN #99]  C/PA,PA
ACE  BONDS OF ETERNITY, THE [@PERRY RHODAN #69]  PA
ACE  CALLER FROM ETERNITY [@PERRY RHODAN #106]  PA
ACE  CAVES OF THE DRUUFS [@PERRY RHODAN #72]  C/PA,PA
ACE  CHALLENGE OF THE UNKNOWN [@PERRY RHODAN #32]  C/PA,PA
ACE  CHECKMATE: UNIVERSE [@PERRY RHODAN #74]  C/PA,PA
ACE  COLUMBUS AFFAIR, THE [@PERRY RHODAN #80]  PA
ACE  CONFLICT CENTER: NAATOR [@PERRY RHODAN #77]  C/PA,PA
ACE  COSMIC DECOY, THE [@PERRY RHODAN #21]  PA,C/PA
ACE  COSMIC TRAITOR [@PERRY RHODAN #26]  PA,C/PA
ACE  CRIMSON UNIVERSE [@PERRY RHODAN #67]  PA
ACE  CRYSTAL PRINCE, THE [@ATLAN #4]/WAR OF THE GHOSTS [ATLAN 5]  PA
ACE  DEAD LIVE, THE [@PERRY RHODAN #48]  PA,C/PA
ACE  DEATH WAITS IN SEMISPACE [@PERRY RHODAN #61]  C/PA,PA
ACE  DESERT OF DEATH’S DOMAIN [@PERRY RHODAN #100]  C/PA,PA
ACE  DIMENSION SEARCH [@PERRY RHODAN #60]  PA
ACE  DUEL UNDER THE DOUBLE SUN [@PERRY RHODAN #108]  PA
ACE  EARTH DIES, THE [@PERRY RHODAN #41]  PA,C/PA
ACE  EMPEROR AND THE MONSTER, THE [@PERRY RHODAN #107]  PA
ACE  ENEMY IN THE DARK [@PERRY RHODAN #85]  PA
ACE  ENTERPRISE STARDUST [@PERRY RHODAN #1]  PA,C/PA
ACE  EPIDEMIC CENTER: ARALON [@PERRY RHODAN #37]  C/PA,PA
ACE  ERNST ELLERT RETURNS! [@PERRY RHODAN #83]  PA
ACE  ESCAPE TO VENUS [@PERRY RHODAN #15]  PA,C/PA
ACE  FALSE FRONT [@PERRY RHODAN #103]  PA
ACE  FLEET OF THE SPRINGERS, THE [@PERRY RHODAN #22]  PA,C/PA

ACE FLIGHT FROM TARKIHL [@ATLAN #2]/

MENACE OF ATOMIGEDDON (SPECIAL) [PERRY RHODAN]

PA

ACE  FORTRESS ATLANTIS [@PERRY RHODAN #52]  C/PA:2,PA
ACE  FORTRESS OF THE SIX MOONS [@PERRY RHODAN #7]  PA,C/PA
ACE  FRIEND TO MANKIND [@PERRY RHODAN #91]  PA
ACE  GALACTIC ALARM [@PERRY RHODAN #3]  PA,C/PA
ACE  GALACTIC RIDDLE, THE [@PERRY RHODAN #8]  PA,C/PA
ACE  GHOSTS OF GOL, THE [@PERRY RHODAN #10]  PA,C/PA
ACE  GIANT’S PARTNER, THE [@PERRY RHODAN #33]  PA,C/PA
ACE  GUARDIANS, THE [@PERRY RHODAN #58]  C/PA:2,PA

ACE HERITAGE OF THE LIZARD PEOPLE [@PERRY RHODAN #113]/

DEATH’S DEMAND [PERRY RHODAN 114]

C/PA,PA

ACE  HORN: GREEN [@PERRY RHODAN #96]  C/PA,PA
ACE  HORROR, THE [@PERRY RHODAN #66]  C/PA,PA
ACE  IDOL FROM PASSA, THE [@PERRY RHODAN #98]  PA
ACE  IMMORTAL UNKNOWN, THE [@PERRY RHODAN #13]  PA,C/PA

The MITSFS Pinkdex  September 24, 2007  Page 3
AUTHOR  TITLE  CODE
ACE  IN THE CENTER OF THE GALAXY (NOVEL) [PERRY RHODAN]  PA
ACE  INFINITY FLIGHT [@PERRY RHODAN #24]  PA,C/PA
ACE  INTERLUDE ON SILIKO 5 [@PERRY RHODAN #59]  C/PA,PA
ACE  INVASION FROM SPACE [@PERRY RHODAN #4]  PA,C/PA
ACE  LAST DAYS OF ATLANTIS, THE [@PERRY RHODAN #62]  PA
ACE  LIFE HUNT [@PERRY RHODAN #43]  PA,C/PA
ACE  MAN AND MONSTER [@PERRY RHODAN #36]  PA,C/PA
ACE  MAN WITH TWO FACES, THE [@PERRY RHODAN #104]  C/PA,PA
ACE  MENACE OF THE MUTANT MASTER [@PERRY RHODAN #18]  C/PA,PA
ACE  MICRO-TECHS, THE [@PERRY RHODAN #55]  C/PA,PA
ACE  MUTANTS VS. MUTANTS [@PERRY RHODAN #19]  PA
ACE  MYSTERY OF THE ANTI, THE [@PERRY RHODAN #88]  C/PA,PA

ACE PALE COUNTRY PURSUIT [@ATLAN #3]/

ROBOT THREAT: NEW YORK (SPECIAL) [PERRY RHODAN]

PA

ACE  PERIL ON ICE PLANET [@PERRY RHODAN #23]  PA,C/PA
ACE  PHANTOM FLEET [@PERRY RHODAN #97]  PA
ACE  PLAGUE OF OBLIVION, THE [@PERRY RHODAN #28]  PA,C/PA
ACE  PLANET OF THE DYING SUN, THE [@PERRY RHODAN #11]  PA,C/PA
ACE  PLANET OF THE GODS [@PERRY RHODAN #27]  PA,C/PA
ACE  PLANET TOPIDE, PLEASE REPLY! [@PERRY RHODAN #75]  C/PA,PA
ACE  PLASMA MONSTER, THE [@PERRY RHODAN #95]  C/PA,PA
ACE  POWER KEY [@PERRY RHODAN #78]  PA
ACE  POWER’S PRICE [@PERRY RHODAN #89]  C/PA,PA
ACE  PRISONER OF TIME [@PERRY RHODAN #56]  C/PA:2,PA
ACE  PROJECT: EARTHSAVE [@PERRY RHODAN #38]  PA,C/PA
ACE  PSEUDO ONE, THE [@PERRY RHODAN #44]  PA,C/PA
ACE  PUCKY’S GREATEST HOUR [@PERRY RHODAN #81]  PA
ACE  QUEST THROUGH SPACE AND TIME [@PERRY RHODAN #9]  PA,C/PA
ACE  RADIANT DOME, THE [@PERRY RHODAN #2]  PA,C/PA
ACE  REALM OF THE TRI-PLANETS [@PERRY RHODAN #31]  PA,C/PA
ACE  REBELS OF TUGLAN, THE [@PERRY RHODAN #12]  PA,C/PA
ACE  RECRUITS FOR ARKON [@PERRY RHODAN #76]  C/PA,PA
ACE  RED EYE OF BETELGEUSE [@PERRY RHODAN #40]  PA,C/PA
ACE  RENEGADES OF THE FUTURE [@PERRY RHODAN #65]  C/PA,PA
ACE  RETURN FROM THE VOID [@PERRY RHODAN #51]  PA,C/PA

ACE SABOTEURS IN A-1 [@PERRY RHODAN #115]/ PSYCHO DUEL, THE [PERRY RHODAN 116]

C/PA,PA

ACE SAVIOR OF THE EMPIRE [@PERRY RHODAN #117]/ SHADOWS ATTACK, THE [PERRY RHODAN 118]

PA

ACE  SECRET BARRIER X [@PERRY RHODAN #16]  C/PA,PA
ACE  SECRET MISSION: MOLUK [@PERRY RHODAN #84]  PA
ACE  SECRET OF THE TIME VAULT, THE [@PERRY RHODAN #6]  PA,C/PA

ACE SEEDS OF RUIN [@PERRY RHODAN #111]/

PLANET MECHANICA [PERRY RHODAN 112]

C/PA,PA

ACE  SHADOW OF THE MUTANT MASTER [@PERRY RHODAN #47]  PA,C/PA
ACE  SILENCE OF GOM, THE [@PERRY RHODAN #39]  PA,C/PA
ACE  SLEEPERS, THE [@PERRY RHODAN #79]  PA
ACE  SNOWMAN IN FLAMES [@PERRY RHODAN #25]  PA,C/PA
ACE  SOLAR ASSASSINS [@PERRY RHODAN #49]  PA,C/PA
ACE  SOS: SPACESHIP TITAN! [@PERRY RHODAN #34]  PA,C/PA
ACE  SPACESHIP OF ANCESTORS [@PERRY RHODAN #73]  PA,C/PA

ACE SPIDER DESERT [@ATLAN #1]/

WASP MEN ATTACK, THE (SPECIAL) [PERRY RHODAN]

PA

ACE  SPOOR OF THE ANTIS [@PERRY RHODAN #102]  C/PA,PA
ACE  SPYBOT! [@PERRY RHODAN #53]  PA,C/PA
ACE  STARLESS REALM, THE [@PERRY RHODAN #87]  PA

ACE STOLEN SPACEFLEET, THE [@PERRY RHODAN #109]/

SGT. ROBOT [PERRY RHODAN 110]

PA

ACE  TARGET STAR, THE [@PERRY RHODAN #92]  C/PA:2,PA
ACE  THORA’S SACRIFICE [@PERRY RHODAN #70]  PA
ACE  THRALL OF HYPNO, THE [@PERRY RHODAN #20]  C/PA,PA
ACE  TIGRIS LEAPS, THE [@PERRY RHODAN #63]  PA,C/PA
ACE  TIME’S LONELY ONE [@PERRY RHODAN #42]  PA,C/PA
ACE  TO ARKON! [@PERRY RHODAN #30]  PA,C/PA
ACE  TOUCH OF ETERNITY, A [@PERRY RHODAN #57]  C/PA,PA
ACE  UNDER THE STARS OF DRUUFON [@PERRY RHODAN #68]  C/PA,PA
ACE  UNKNOWN SECTOR: MILKY WAY [@PERRY RHODAN #45]  PA,C/PA
ACE  UNLEASHED POWERS [@PERRY RHODAN #90]  C/PA,PA
ACE  VAGABOND OF SPACE [@PERRY RHODAN #93]  C/PA:2,PA
ACE  VEGA SECTOR, THE [@PERRY RHODAN #5]  PA,C/PA
ACE  VENUS IN DANGER [@PERRY RHODAN #14]  PA,C/PA
ACE  VENUS TRAP, THE [@PERRY RHODAN #17]  PA
ACE  WONDERFLOWER OF UTIK [@PERRY RHODAN #105]  PA

The MITSFS Pinkdex  September 24, 2007  Page 4
AUTHOR  TITLE  CODE
ACE  WORLD GONE MAD, A [@PERRY RHODAN #29]  C/PA,PA
ACHILLI, JUSTIN  GIOVANNI [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
ACKER, KATHY  EMPIRE OF THE SENSELESS  H,LP
ACKERMAN, FORREST J.  ACKERMANTHOLOGY!  LPA
ACKERMAN, FORREST J.  BEST SCIENCE FICTION FOR 1973  PA
ACKERMAN, FORREST J.  GOSH! WOW! (SENSE OF WONDER)  C/PA,PA
ACKERMAN, FORREST J.  SCIENCE-FICTION CLASSICS  LPA
ACKROYD, PETER  FIRST LIGHT  LP
ACKROYD, PETER  PLATO PAPERS, THE  LP
ACRES, MARK  DARK DIVIDE [RUNESWORD 5]  C/P,P
ACRES, MARK  DRAGON WAR [LAND BETWEEN THE RIVERS 2]  P
ACRES, MARK  DRAGONSPAWN [LAND BETWEEN THE RIVERS 1]  P
ACRES, MARK  INDOMITABLE THUNDER [@WORLD OF ADEN THUNDERSCAPE]  P
ACRES, MARK  LORD OF LANCES [JANISSARIES]  P
ADAMS, BILL/BROOKS, CECIL  END OF FAME, THE  P
ADAMS, BILL/BROOKS, CECIL  UNWOUND WAY, THE  P
ADAMS, DOUGLAS  ANHALTER DURCH DIE GALAXIS, PER  FP
ADAMS, DOUGLAS  DIRK GENTLY’S HOLISTIC DETECTIVE AGENCY [DIRK GENTLY 1]  H:2,P
ADAMS, DOUGLAS  HITCHHIKERS’ GUIDE TO THE GALAXY, THE [HITCHHIKERS 1]  C/P,H:2,P

ADAMS, DOUGLAS HITCHHIKERS’ GUIDE TO THE GALAXY, THE (5/6 CASSETTES)

[HITCHHIKERS 1]

SR-MM

ADAMS, DOUGLAS HITCHHIKERS’ GUIDE TO THE GALAXY, THE (6 CASSETTES)

[HITCHHIKERS 1]

SR-MM:2

ADAMS, DOUGLAS HITCHHIKERS’ GUIDE TO THE GALAXY, THE (WITH CASABLANCA)

[HITCHHIKERS]

SR-MM

ADAMS, DOUGLAS  ILLUSTRATED HITCHHIKER’S GUIDE TO THE GALAXY, THE [HITCHHIKERS] A
ADAMS, DOUGLAS  LIFE, THE UNIVERSE AND EVERYTHING [HITCHHIKERS 3]  C/P,P,H:2
ADAMS, DOUGLAS  LONG DARK TEA-TIME OF THE SOUL, THE [DIRK GENTLY 2]  C/P,P,H
ADAMS, DOUGLAS/LLOYD, JOHN  MEANING OF LIFF, THE  [C/RAN-P]
ADAMS, DOUGLAS  MOSTLY HARMLESS [HITCHHIKERS 5]  H:2
ADAMS, DOUGLAS  ORIGINAL HITCHHIKER RADIO SCRIPTS, THE [HITCHHIKERS]  H,LP
ADAMS, DOUGLAS  RESTAURANT AT THE END OF THE UNIVERSE, THE [HITCHHIKERS 2]  C/P,P,H

ADAMS, DOUGLAS RESTAURANT AT THE END OF THE UNIVERSE, THE (1 CASSETTE)

[HITCHHIKERS 2]

SR-MM

ADAMS, DOUGLAS  SALMON OF DOUBT, THE  H
ADAMS, DOUGLAS  SHADA [DOCTOR WHO]  SR-P
ADAMS, DOUGLAS  SO LONG, AND THANKS FOR ALL THE FISH [HITCHHIKERS 4]  C/P,P,H
ADAMS, DOUGLAS  ULTIMATE HITCHHIKER’S GUIDE, THE [HITCHHIKERS 1,2,3,4,5]  H,LP
ADAMS, EUSTACE L.  FLYING WINDMILL, THE  H
ADAMS, HAZARD  TRUTH ABOUT DRAGONS, THE  H
ADAMS, HUNTER  SHE-BEAST, THE [@MAN FROM PLANET X #1]  SR-P
ADAMS, HUNTER  TIGER BY THE TAIL [@MAN FROM PLANET X #2]  SR-P
ADAMS, JOHN  WHEN THE GODS CAME  SR-P

ADAMS, KATHLEEN/ ATCHINSON, FRANCES ELIZABETH BOOK OF ENCHANTMENT, A

H

ADAMS, RICHARD  GIRL IN A SWING, THE  C/P,P
ADAMS, RICHARD  MAIA  C/P,P,H
ADAMS, RICHARD  PLAGUE DOGS, THE  H,P
ADAMS, RICHARD  SHARDIK  H:2,P
ADAMS, RICHARD  TALES FROM WATERSHIP DOWN [WATERSHIP DOWN 2]  C/P,P
ADAMS, RICHARD  UNBROKEN WEB, THE  SR-VLH
ADAMS, RICHARD  WATERSHIP DOWN [WATERSHIP DOWN 1]  C/P,P
ADAMS, ROBERT  ALTERNATIVES  SFWA-P

ADAMS, ROBERT/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

BARBARIANS

PA

ADAMS, ROBERT  BILI THE AXE [HORSECLANS 10]  C/P:2,P
ADAMS, ROBERT  CASTAWAYS IN TIME [@CASTAWAYS IN TIME #1]  C/P,LP
ADAMS, ROBERT  CAT OF SILVERY HUE, A [HORSECLANS 4]  C/P:2,P
ADAMS, ROBERT  CHAMPION OF THE LAST BATTLE [HORSECLANS 11]  C/P:2,P
ADAMS, ROBERT  CLAN OF THE CATS, THE [HORSECLANS 18]  P
ADAMS, ROBERT  COMING OF THE HORSECLANS, THE [HORSECLANS 1]  C/P:2,P
ADAMS, ROBERT  DEATH OF A LEGEND, THE [HORSECLANS 8]  C/P:2,P
ADAMS, ROBERT  HORSECLANS ODYSSEY [HORSECLANS 7]  C/P:2,P
ADAMS, ROBERT  HORSES OF THE NORTH [HORSECLANS 13]  P
ADAMS, ROBERT  MADMAN’S ARMY [HORSECLANS 17]  P
ADAMS, ROBERT  MAN CALLED MILO MORAI, A [HORSECLANS 14]  C/P,P
ADAMS, ROBERT  MEMORIES OF MILO MORAI, THE [HORSECLANS 15]  C/P,P
ADAMS, ROBERT  MONSTERS AND MAGICIANS [STAIRWAY TO FOREVER 2]  C/P:2,P
ADAMS, ROBERT  OF BEGINNINGS AND ENDINGS [@CASTAWAYS IN TIME #6]  P
ADAMS, ROBERT  OF CHIEFS AND CHAMPIONS [@CASTAWAYS IN TIME #4]  P
ADAMS, ROBERT  OF MYTHS AND MONSTERS [@CASTAWAYS IN TIME #5]  P
ADAMS, ROBERT  OF QUESTS AND KINGS [@CASTAWAYS IN TIME #3]  C/P:2,P
ADAMS, ROBERT  PATRIMONY, THE [HORSECLANS 6]  C/P:2,P
ADAMS, ROBERT  PHANTOM REGIMENTS  C/PA,PA,SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 5
AUTHOR  TITLE  CODE
ADAMS, ROBERT  REVENGE OF THE HORSECLANS [HORSECLANS 3]  C/P:2,P

ADAMS, ROBERT/GREENBERG, MARTIN H./ ADAMS, PAMELA CRIPPEN ROBERT ADAMS’ BOOK OF ALTERNATE WORLDS

C/PA,PA

ADAMS, ROBERT/GREENBERG, MARTIN H./ ADAMS, PAMELA CRIPPEN ROBERT ADAMS’ BOOK OF SOLDIERS

C/PA,PA

ADAMS, ROBERT  SAVAGE MOUNTAINS, THE [HORSECLANS 5]  C/P:2,P
ADAMS, ROBERT  SEVEN MAGICAL JEWELS OF IRELAND, THE [@CASTAWAYS IN TIME #2]  C/P,P
ADAMS, ROBERT  STAIRWAY TO FOREVER [STAIRWAY TO FOREVER 1]  C/P,SFWA-P,P
ADAMS, ROBERT  SWORDS OF THE HORSECLANS [HORSECLANS 2]  C/P:2,P
ADAMS, ROBERT  TALES OF THE HORSECLANS [HORSECLANS 1,2,3]  H
ADAMS, ROBERT  WITCH GODDESS, THE [HORSECLANS 9]  C/P:2,P
ADAMS, ROBERT  WOMAN OF THE HORSECLANS, A [HORSECLANS 12]  P
ADAMS, TERRY A.  MASTER OF CHAOS, THE [SENTIENCE 2]  C/P:2,P
ADAMS, TERRY A.  SENTIENCE [SENTIENCE 1]  C/P:2,P
ADDAMS, CHARLES  MY CROWD  CX
ADDISCOMBE, JOHN  IRON ARMY, THE  B
ADDISON, JOSEPH  TESSERACT  C/P:2,P
ADKINS, PATRICK H.  LORD OF THE CROOKED PATHS [TITANS 1]  P
ADKINS, PATRICK H.  MASTER OF THE FEARFUL DEPTHS [TITANS 2]  P
ADKINS, PATRICK H.  SONS OF THE TITANS [TITANS 3]  P
ADLARD, MARK  INTERFACE [INTERFACE 1]  C/P,P
ADLARD, MARK  MULTIFACE [INTERFACE 3]  P
ADLARD, MARK  VOLTEFACE [INTERFACE 2]  P
ADLER, ALLEN  MACH 1  H:2
ADLER, ALLEN  TERROR ON PLANET IONUS  C/P:2,P
ADLER, BILL  LETTERS TO THE AIR FORCE ON UFO’S  [C/RAN-P:2]
ADLER, BILL JR.  TIME MACHINES  HA,LPA
ADLER, C. S.  EDDIE’S BLUE-WINGER DRAGON  P
ADLER, C. S.  MAGICAL COACH, THE  P
ADLER, IRVING  SUN AND ITS FAMILY, THE  [C/RAN-P]
ADLER, PAUL  SAUCER HILL  P
ADRIAN, JACK  PILGRIMAGE TO HELL [@DEATHLANDS]  C/P,P
ADVENT BOOKS  SCIENCE FICTION NOVEL, THE  R/REF-H
AESOP  AESOP’S FABLES  H
AGEL, JEROME  MAKING OF KUBRICK’S 2001, THE  R/REF-P,C/REF-P:2
AGINS, PHYLLIS CAROL  SUISAN  P,SFWA-P
AGUIRRE, FORREST/VANDERMEER, JEFF  LEVIATHAN 3  LPA
AHERN, JERRY  ARSENAL, THE [@SURVIVALIST #16]  P
AHERN, JERRY  AWAKENING, THE [@SURVIVALIST #10]  P
AHERN, JERRY  DOOMSAYER, THE [@SURVIVALIST #4]  P
AHERN, JERRY  EARTH FIRE [@SURVIVALIST #9]  P
AHERN, JERRY  END IS COMING, THE [@SURVIVALIST #8]  P
AHERN, JERRY/AHERN, SHARON  FREEMAN, THE  P
AHERN, JERRY/AHERN, SHARON  GOLDEN SHIELD OF IBF, THE  P
AHERN, JERRY  MID-WAKE [@SURVIVALIST]  P
AHERN, JERRY  NIGHTMARE BEGINS, THE [@SURVIVALIST #2]  P
AHERN, JERRY  ORDEAL, THE [@SURVIVALIST #17]  P
AHERN, JERRY  OVERLORD [@SURVIVALIST #15]  P
AHERN, JERRY  PROPHET, THE [@SURVIVALIST #7]  P
AHERN, JERRY  PURSUIT [@SURVIVALIST #13]  P
AHERN, JERRY  QUEST, THE [@SURVIVALIST #3]  P
AHERN, JERRY  REBELLION, THE [@SURVIVALIST #12]  P
AHERN, JERRY  REPRISAL, THE [@SURVIVALIST #11]  P
AHERN, JERRY  SAVAGE HORDE, THE [@SURVIVALIST #6]  P
AHERN, JERRY  TAKERS, THE  C/P,P
AHERN, JERRY  TERROR, THE [@SURVIVALIST #14]  P
AHERN, JERRY  TOTAL WAR, THE [@SURVIVALIST #1]  P
AHERN, JERRY  WEB, THE [@SURVIVALIST #5]  P
AHMED, ROLLO  COMPLETE BOOK OF WITCHCRAFT, THE  [C/RAN-P]
AICKMAN, ROBERT  COLD HAND IN MINE  H,P
AICKMAN, ROBERT  FONTANA BOOK OF GREAT GHOST STORIES, THE  PA
AICKMAN, ROBERT  3RD FONTANA BOOK OF GREAT GHOST STORIES, THE  SR-PA,C/PA
AICKMAN, ROBERT  4TH FONTANA BOOK OF GREAT GHOST STORIES, THE  PA
AICKMAN, ROBERT  8TH FONTANA BOOK OF GREAT GHOST STORIES, THE  SR-PA
AICKMAN, ROBERT  MODEL, THE  H
AICKMAN, ROBERT  PAINTED DEVILS  H,SFWA
AIKEN, JOAN  COCKATRICE BOYS, THE  H,LP
AIKEN, JOAN  CUCKOO TREE, THE  P
AIKEN, JOAN  DIDO AND PA  H
AIKEN, JOAN  GREEN FLASH, THE  H,P
AIKEN, JOAN  NECKLACE OF RAINDROPS, A  P
AIKEN, JOHN  WORLD WELL LOST  H,SFWA
AIKIN, JIM  WALK THE MOONS ROAD  P
AIKIN, JIM  WALL AT THE EDGE OF THE WORLD, THE  P,SFWA-P:2

The MITSFS Pinkdex  September 24, 2007  Page 6
AUTHOR  TITLE  CODE
AINSBURY, RAY  WHEN THE MOON RAN WILD  SR-P
AIREY, JEAN  DOCTOR AND THE ENTERPRISE, THE [DOCTOR WHO/STAR TREK]  SR-LP
AIREY, JEAN/HALDEMAN, LINDA  TRAVEL WITHOUT THE TARDIS [DOCTOR WHO]  R/REF-P
AKERS, ALAN BURT  ALLIES OF ANTARES [DRAY PRESCOT 26]  P
AKERS, ALAN BURT  ARENA OF ANTARES [DRAY PRESCOT 7]  P
AKERS, ALAN BURT  ARMADA OF ANTARES [DRAY PRESCOT 11]  C/P,P
AKERS, ALAN BURT  AVENGER OF ANTARES [DRAY PRESCOT 10]  P
AKERS, ALAN BURT  BEASTS OF ANTARES [DRAY PRESCOT 23]  P
AKERS, ALAN BURT  BLADESMAN OF ANTARES [DRAY PRESCOT 9]  P
AKERS, ALAN BURT  CAPTIVE SCORPIO [DRAY PRESCOT 17]  P
AKERS, ALAN BURT  DELIA OF VALLIA [DRAY PRESCOT 28]  P
AKERS, ALAN BURT  FIRES OF SCORPIO [DRAY PRESCOT 29]  P
AKERS, ALAN BURT  FLIERS OF ANTARES [DRAY PRESCOT 8]  P
AKERS, ALAN BURT  FORTUNE FOR KREGEN, A [DRAY PRESCOT 21]  P
AKERS, ALAN BURT  GOLDEN SCORPIO [DRAY PRESCOT 18]  P
AKERS, ALAN BURT  KROZAIR OF KREGEN [DRAY PRESCOT 14]  P
AKERS, ALAN BURT  LEGIONS OF ANTARES [DRAY PRESCOT 25]  P
AKERS, ALAN BURT  LIFE FOR KREGEN, A [DRAY PRESCOT 19]  P
AKERS, ALAN BURT  MANHOUNDS OF ANTARES [DRAY PRESCOT 6]  C/P,P
AKERS, ALAN BURT  MASKS OF SCORPIO [DRAY PRESCOT 31]  C/P,P
AKERS, ALAN BURT  MAZES OF SCORPIO [DRAY PRESCOT 27]  P
AKERS, ALAN BURT  OMENS OF KREGEN [DRAY PRESCOT 36]  P
AKERS, ALAN BURT  PRINCE OF SCORPIO [DRAY PRESCOT 5]  C/P,P
AKERS, ALAN BURT  REBEL OF ANTARES [DRAY PRESCOT 24]  P
AKERS, ALAN BURT  RENEGADE OF KREGEN [DRAY PRESCOT 13]  P
AKERS, ALAN BURT  SAVAGE SCORPIO [DRAY PRESCOT 16]  C/P:2,P
AKERS, ALAN BURT  SECRET SCORPIO [DRAY PRESCOT 15]  P
AKERS, ALAN BURT  SEG THE BOWMAN [DRAY PRESCOT 32]  C/P,P
AKERS, ALAN BURT  STORM OVER VALLIA [DRAY PRESCOT 35]  P
AKERS, ALAN BURT  SUNS OF SCORPIO, THE [DRAY PRESCOT 2]  C/P,P
AKERS, ALAN BURT  SWORD FOR KREGEN, A [DRAY PRESCOT 20]  P
AKERS, ALAN BURT  SWORDSHIPS OF SCORPIO [DRAY PRESCOT 4]  C/P,P
AKERS, ALAN BURT  TALONS OF SCORPIO [DRAY PRESCOT 30]  P
AKERS, ALAN BURT  TIDES OF KREGEN, THE [DRAY PRESCOT 12]  P
AKERS, ALAN BURT  TRANSIT TO SCORPIO [DRAY PRESCOT 1]  C/P:2,P
AKERS, ALAN BURT  VICTORY FOR KREGEN, A [DRAY PRESCOT 22]  C/P,P
AKERS, ALAN BURT  WARLORD OF ANTARES [DRAY PRESCOT 37]  P
AKERS, ALAN BURT  WARRIOR OF SCORPIO [DRAY PRESCOT 3]  C/P,P
AKERS, ALAN BURT  WEREWOLVES OF KREGEN [DRAY PRESCOT 33]  P
AKERS, ALAN BURT  WITCHES OF KREGEN [DRAY PRESCOT 34]  P
ALAMA, PAULINE J.  EYE OF NIGHT, THE  C/P,P
ALANIZ, ALEX  BEYOND FUTURE SHOCK  LP
ALBAN, ANTONY  CATHARSIS CENTRAL  C/P,P
ALBAN, ANTONY  DAY OF THE SHIELD, THE  C/P:2,P
ALBANO, PETER  TRIAL OF THE SEVENTH CARRIER [SEVENTH CARRIER 4]  P
ALBEE, EDWARD  TINY ALICE  P
ALDANI, LINO  FANTASCIENZA, LA  R/REF-P
ALDISS, BRIAN W.  ALL ABOUT VENUS  C/PA,PA
ALDISS, BRIAN W.  BAREFOOT IN THE HEAD  C/P:2,P

ALDISS, BRIAN W.

BILLION YEAR SPREE SFWA,R/REF-H, C/REF-P:2

ALDISS, BRIAN W.  BOOK OF BRIAN ALDISS, THE  C/P:2,P
ALDISS, BRIAN W.  BOW DOWN TO NUL  C/P:2,DD443,P
ALDISS, BRIAN W.  CANOPY OF TIME, THE  C/P,P
ALDISS, BRIAN W.  COMMON CLAY  H:2
ALDISS, BRIAN W.  CRYPTOZOIC!  H:2,P
ALDISS, BRIAN W.  CULTURAL BREAKS  H
ALDISS, BRIAN W.  DARK LIGHT-YEARS, THE  C/P:2,P
ALDISS, BRIAN W./HARRISON, HARRY  DECADE: THE 1940’S  HA
ALDISS, BRIAN W./HARRISON, HARRY  DECADE: THE 1950’S  HA
ALDISS, BRIAN W.  DETACHED RETINA, THE  R/REF-P
ALDISS, BRIAN W.  DRACULA UNBOUND  C/P,P
ALDISS, BRIAN W.  EARTHWORKS  H:2,P
ALDISS, BRIAN W.  EIGHTY MINUTE HOUR, THE  C/P,SFWA,P,H
ALDISS, BRIAN W.  ENEMIES OF THE SYSTEM  P,SFWA
ALDISS, BRIAN W.  EQUATOR AND SEGREGATION  P
ALDISS, BRIAN W.  EVIL EARTHS  C/PA,PA
ALDISS, BRIAN W.  FORGOTTEN LIFE  P
ALDISS, BRIAN W.  FRANKENSTEIN UNBOUND  C/P,SFWA,P,H
ALDISS, BRIAN W.  GALACTIC EMPIRES 1  C/PA,PA
ALDISS, BRIAN W.  GALACTIC EMPIRES 2  PA,C/PA
ALDISS, BRIAN W.  GALAXIES LIKE GRAINS OF SAND  C/P:2,P
ALDISS, BRIAN W.  GREYBEARD  C/P:2,P
ALDISS, BRIAN W.  HELLICONIA SPRING [HELLICONIA 1]  H:2,P,SFWA:2

AUTHOR  TITLE  CODE
ALDISS, BRIAN W.  HELLICONIA SUMMER [HELLICONIA 2]  H:2,P,SFWA
ALDISS, BRIAN W.  HELLICONIA WINTER [HELLICONIA 3]  C/P,SFWA-P,H:2,P,SFWA
ALDISS, BRIAN W./HARRISON, HARRY  HELL’S CARTOGRAPHERS  SFWA,R/REF-H
ALDISS, BRIAN W.  HOTHOUSE  C/P:2,P
ALDISS, BRIAN W.  ISLAND CALLED MOREAU, AN  C/P,P,H
ALDISS, BRIAN W.  LAST ORDERS  C/P:2,P
ALDISS, BRIAN W.  LONG AFTERNOON OF EARTH, THE  C/P:2,P
ALDISS, BRIAN W.  MALACIA TAPESTRY, THE  C/P,SFWA,P,H
ALDISS, BRIAN W.  MALE RESPONSE, THE  SR-P
ALDISS, BRIAN W.  MAN IN HIS TIME  C/P,P,H
ALDISS, BRIAN W.  MOMENT OF ECLIPSE  H,SFWA
ALDISS, BRIAN W.  NEANDERTHAL PLANET  H:2,P
ALDISS, BRIAN W./HARRISON, HARRY  NEBULA AWARD STORIES 2  HA,PA
ALDISS, BRIAN W.  NEW ARRIVALS, OLD ENCOUNTERS  H,SFWA,P
ALDISS, BRIAN W.  NO TIME LIKE TOMORROW  C/P:2,P
ALDISS, BRIAN W.  NON-STOP  H:2,P
ALDISS, BRIAN W.  PENGUIN SCIENCE FICTION  SR-PA
ALDISS, BRIAN W.  PENGUIN SCIENCE FICTION OMNIBUS, THE  C/PA,PA
ALDISS, BRIAN W.  PERILOUS PLANETS  C/PA,PA
ALDISS, BRIAN W.  PRIMAL URGE, THE  C/P:2,P
ALDISS, BRIAN W.  REPORT ON PROBABILITY A  C/P,P
ALDISS, BRIAN W.  ROMANCE OF THE EQUATOR, A  H:2
ALDISS, BRIAN W.  SALIVA TREE, THE  SR-P,TD3,SFWA-TD3
ALDISS, BRIAN W.  SEASONS IN FLIGHT  H
ALDISS, BRIAN W.  SEX AND THE BLACK MACHINE  B
ALDISS, BRIAN W.  SHAPE OF FURTHER THINGS, THE  SFWA,R/REF-H
ALDISS, BRIAN W.  SOMEWHERE EAST OF LIFE  H
ALDISS, BRIAN W.  SPACE ODYSSEYS  C/PA,PA,SFWA
ALDISS, BRIAN W.  SPACE OPERA  C/PA,PA
ALDISS, BRIAN W.  SPACE, TIME AND NATHANIEL  P
ALDISS, BRIAN W.  STARSHIP  C/P:2,P
ALDISS, BRIAN W.  STARSWARM  C/P:2,P
ALDISS, BRIAN W.  SUPER-STATE  H
ALDISS, BRIAN W.  SUPERTOYS LAST ALL SUMMER LONG  LP
ALDISS, BRIAN W.  TUPOLEV TOO FAR, A  H
ALDISS, BRIAN W.  VANGUARD FROM ALPHA  DD369
ALDISS, BRIAN W./PENROSE, ROGER  WHITE MARS  H
ALDISS, BRIAN W.  WHO CAN REPLACE A MAN?  C/P:2,P
ALDISS, BRIAN W.  YEAR BEFORE YESTERDAY, THE  H,SFWA,P
ALDISS, BRIAN W.  YET MORE PENGUIN SCIENCE FICTION  SR-PA
ALDRED, SOPHIE/TUCKER, MIKE  ACE! [DOCTOR WHO]  R/REF-H
ALDRIDGE, RAY  EMPEROR OF EVERYTHING, THE [@EMANCIPATOR #2]  C/P,P
ALDRIDGE, RAY  ORPHEUS MACHINE, THE [@EMANCIPATOR #3]  C/P,P
ALDRIDGE, RAY  PHARAOH CONTRACT, THE [@EMANCIPATOR #1]  C/P:2,P
ALDRIN, BUZZ/BARNES, JOHN  ENCOUNTER WITH TIBER  P
ALDRIN, BUZZ/BARNES, JOHN  RETURN, THE  H:2
ALEXANDER, ALMA  CHANGER OF DAYS [ANGHARA 2]  P
ALEXANDER, ALMA  HIDDEN QUEEN, THE [ANGHARA 1]  P
ALEXANDER, BRANDON  KISS OF DARKNESS [@CHARMED]  P
ALEXANDER, C. M.  ALIEN ATLAS  P
ALEXANDER, C. M.  BIODROIDS 2300  P
ALEXANDER, DAVID  DARK MESSIAH [@PHOENIX #1]  P
ALEXANDER, DAVID  METALSTORM [@PHOENIX #4]  P
ALEXANDER, DAVID  NOMAD  P
ALEXANDER, DAVID  STAR TREK CREATOR  SFWA
ALEXANDER, DAVID M.  FANE  C/P,P
ALEXANDER, JAN  DARKWATER  P
ALEXANDER, KARL  TIME AFTER TIME  P
ALEXANDER, KATHERINE  DREAM OF THINGS THAT WERE, A  P
ALEXANDER, LLOYD  ARKADIANS, THE (PROOF)  SR-LP
ALEXANDER, LLOYD  BEGGAR QUEEN, THE [WESTMARK 3]  P
ALEXANDER, LLOYD  BLACK CAULDRON, THE [PRYDAIN 2]  C/P:2,LP,H
ALEXANDER, LLOYD  BOOK OF THREE, THE [PRYDAIN 1]  C/P:2,LP,H
ALEXANDER, LLOYD  CASTLE OF LLYR, THE [PRYDAIN 3]  C/P:3,LP
ALEXANDER, LLOYD  COLL AND HIS WHITE PIG  LP
ALEXANDER, LLOYD  FIRST TWO LIVES OF LUK AS-KASHA, THE  P
ALEXANDER, LLOYD  FOUNDLING, THE [PRYDAIN 6]  H,P
ALEXANDER, LLOYD  HIGH KING, THE [PRYDAIN 5]  C/P:2,LP
ALEXANDER, LLOYD  ILLYRIAN ADVENTURE, THE  P
ALEXANDER, LLOYD  KESTREL, THE [WESTMARK 2]  P
ALEXANDER, LLOYD  PRYDAIN CHRONICLES, THE [PRYDAIN 1,2,3,4,5,6]  H
ALEXANDER, LLOYD  TARAN WANDERER [PRYDAIN 4]  C/P:3,LP
ALEXANDER, LLOYD  TRUTHFUL HARP, THE  LP
ALEXANDER, LLOYD  WESTMARK [WESTMARK 1]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 8
AUTHOR  TITLE  CODE
ALEXANDER, MARC  MIST LIZARD, THE  P
ALEXANDER, PAUL  RED DWARF LOG NO. 1996 [RED DWARF]  [C/RAN-H]
ALEXANDER, THEA  2150 A.D.  C/P,P
ALGOL PRESS  EXPLORING CORDWAINER SMITH  R/REF-P
ALICKE, WALTER  MOON IN SCIENCE & FICTION, THE  R/REF-H
ALIGHIERI, DANTE  PURGATORIO, THE (CIARDI TRANSLATION)  P
ALKON, PAUL K.  SCIENCE FICTION BEFORE 1900  R/REF-H
ALLAN, MARGARET  SISTER OF THE SKY  P
ALLAWAY, HOWARD  SPACE SHUTTLE AT WORK, THE  [C/RAN-H]
ALLEN, CARLTON C.  SPACE STATION EIGHT  H
ALLEN, DICK  SCIENCE FICTION: THE FUTURE  LP
ALLEN, HENRY WILSON  GENESIS FIVE  C/P:2,P
ALLEN, L. DAVID  BALLANTINE TEACHERS’ GUIDE TO SCIENCE FICTION, THE  R/REF-P,C/REF-P
ALLEN, L. DAVID  SCIENCE FICTION: AN INTRODUCTION  R/REF-H,C/REF-H
ALLEN, L. DAVID  SCIENCE FICTION READER’S GUIDE  R/REF-P,C/REF-P:2
ALLEN, ROGER MACBRIDE  ALLIES AND ALIENS  P
ALLEN, ROGER MACBRIDE  AMBUSH AT CORELLIA [STAR WARS CORELLIAN 1]  C/P:2,P
ALLEN, ROGER MACBRIDE  ASSAULT AT SELONIA [STAR WARS CORELLIAN 2]  C/P:2,P
ALLEN, ROGER MACBRIDE  CALIBAN [CALIBAN 1]  H,SFWA:2,P
ALLEN, ROGER MACBRIDE  CAUSE OF DEATH, THE [BUREAU OF SPECIAL INVESTIGATIONS]  P
ALLEN, ROGER MACBRIDE  DEPTHS OF TIME, THE [SOLACE 1]  H,LP
ALLEN, ROGER MACBRIDE  FARSIDE CANNON  C/P:2,SFWA-P,P
ALLEN, ROGER MACBRIDE  GAME OF WORLDS, THE [@OUT OF TIME #3]  P
ALLEN, ROGER MACBRIDE  INFERNO [CALIBAN 2]  P
ALLEN, ROGER MACBRIDE  MODULAR MAN, THE  C/P,P,H
ALLEN, ROGER MACBRIDE  OCEAN OF YEARS, THE [SOLACE 2]  P
ALLEN, ROGER MACBRIDE  ORPHAN OF CREATION  P,SFWA-P
ALLEN, ROGER MACBRIDE  RING OF CHARON, THE [HUNTED EARTH 1]  C/P,SFWA-P,P,H
ALLEN, ROGER MACBRIDE  ROGUE POWERS [TORCH OF HONOR 2]  C/P,P
ALLEN, ROGER MACBRIDE  SHATTERED SPHERE, THE [HUNTED EARTH 2]  H,P
ALLEN, ROGER MACBRIDE  SHORES OF TOMORROW, THE [SOLACE 3]  P
ALLEN, ROGER MACBRIDE  SHOWDOWN AT CENTERPOINT [STAR WARS CORELLIAN 3]  C/P:2,P
ALLEN, ROGER MACBRIDE/KOTANI, ERIC  SUPERNOVA  C/P,P
ALLEN, ROGER MACBRIDE  TORCH OF HONOR, THE [TORCH OF HONOR 1]  C/P:2,P
ALLEN, ROGER MACBRIDE  UTOPIA [CALIBAN 3]  H,P
ALLEN, ROGER MACBRIDE/DRAKE, DAVID  WAR MACHINE, THE [CRISIS OF EMPIRE 3]  P,SFWA-P:2
ALLEN, VERGIE  JOURNEYS  P
ALLEN, WOODY  WITHOUT FEATHERS  [C/RAN-P]
ALLINGHAM, CEDRIC  FLYING SAUCER FROM MARS  [C/RAN-H]
ALLISON, CLYDE  FOR YOUR SIGHS ONLY  SR-L
ALLISON, CLYDE  NAUTIPUSS  SR-L

ALLISON, LEONARD/JENKIN, LEONARD/ PERRAULT, ROBERT SURVIVAL PRINTOUT

C/PA,PA

ALLSTON, AARON  BETRAYAL [STAR WARS LEGACY OF THE FORCE 1]  H
ALLSTON, AARON  DOC SIDHE [DOC SIDHE 1]  P
ALLSTON, AARON  DOUBLE JEOPARDY [CAR WARRIORS 2]  P
ALLSTON, AARON  ENEMY LINES [@STAR WARS NEW JEDI ORDER ENEMY LINES 1,2]  H
ALLSTON, AARON  EXILE [@STAR WARS LEGACY OF THE FORCE 4]  H,P
ALLSTON, AARON  GALATEA IN 2-D  P,SFWA-P
ALLSTON, AARON  IRON FIST [@STAR WARS X-WING 6]  C/P,P
ALLSTON, AARON  REBEL DREAM [@STAR WARS NEW JEDI ORDER ENEMY LINES 1]  P
ALLSTON, AARON  REBEL STAND [@STAR WARS NEW JEDI ORDER ENEMY LINES 2]  P
ALLSTON, AARON  SIDHE DEVIL [DOC SIDHE 2]  P
ALLSTON, AARON  SOLO COMMAND [@STAR WARS X-WING 7]  P
ALLSTON, AARON  STARFIGHTERS OF ADUMAR [@STAR WARS X-WING 9]  C/P,P
ALLSTON, AARON  TERMINATOR DREAMS [TERMINATOR 3]  P
ALLSTON, AARON  WRAITH SQUADRON [@STAR WARS X-WING 5]  P
ALLUM, TOM  BOY BEYOND THE MOON  H
ALPER, GERALD A.  MY NAME IS VLADIMIR SLOIFOISKI  P
ALPERS, FUCHS/HAHN/JESCHKE  LEXICON DER SCIENCE FICTION LITERATUR  R/REF-P
ALT, MADELYN  CHARMED DEATH, A [@BEWITCHING MYSTERY 2]  P
ALT, MADELYN  TROUBLE WITH MAGIC, THE [@BEWITCHING MYSTERY 1]  P
ALTEN, STEVE  DOMAIN  H
ALTEN, STEVE  GOLIATH  H
ALTER, ROBERT EDMOND  PATH TO SAVAGERY  C/P,P
ALTMAN, STEVEN-ELLIOT  DEPRIVERS  LP
ALTON, ANDREA I.  DEMON OF UNDOING  P

ALTOV, GENRIKH/ ZHURAVLYOVA, VALENTINA BALLAD OF THE STARS

H

AMADO, JORGE  DONA FLOR AND HER TWO HUSBANDS  P
AMBROSE, DAVID  MAN WHO TURNED INTO HIMSELF, THE  H
AMBUEHL, JAMES  HARDBOILED CTHULHU  LPA
AMES, CLYDE  GORGONZOLA, WON’T YOU PLEASE COME HOME?  P
AMES, MILDRED  ANNA TO THE INFINITE POWER  P

AUTHOR  TITLE  CODE
AMIES, CHRIS  DEAD GROUND  LP
AMIS, KINGSLEY  ALTERATION, THE  H,P
AMIS, KINGSLEY  COLONEL SUN [JAMES BOND 14]  P
AMIS, KINGSLEY  GREEN MAN, THE  H,P
AMIS, KINGSLEY  NEW MAPS OF HELL  R/REF-P,C/REF-H:2
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM  C/PA,PA
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM 2  C/PA,PA
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM 3  PA,C/PA
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM 4  C/PA,PA
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM 5  C/PA,PA
AMIS, KINGSLEY/CONQUEST, ROBERT  SPECTRUM (SELECTIONS)  C/PA,PA
AMOSOFF, N.  NOTES FROM THE FUTURE  H
AMPER, DRAX  FAR BEYOND THE BLUE  SR-P
AMTHOR, TERRY K.  SPY IN ISENGARD, A  P
ANCHORS, WILLIAM E. JR./MITCHELL, FLINT LOST IN SPACE SCRAPBOOK VOL. TWO, THE [LOST IN SPACE]  VLP
ANDERS, LOU  FAST FORWARD 1  LPA
ANDERS, LOU  FUTURESHOCKS  LPA
ANDERS, LOU  LIVE WITHOUT A NET  LPA
ANDERS, LOU  OUTSIDE THE BOX  LPA
ANDERSEN, HANS CHRISTIAN  ANDERSEN’S FAIRY TALES  H,P
ANDERSON, BARTH  PATRON SAINT OF PLAGUES, THE  LP
ANDERSON, CHESTER  BUTTERFLY KID, THE  C/P:2,P
ANDERSON, CHESTER/KURLAND, MICHAEL  TEN YEARS TO DOOMSDAY  C/P:2,P
ANDERSON, COLIN  MAGELLAN  H,P
ANDERSON, DENNIS LEE  ARTHUR, KING  C/P,P
ANDERSON, DOUGLAS A.  TALES BEFORE TOLKIEN  HA
ANDERSON, JACK  MILLENIUM  LP
ANDERSON, JACK  MILLENNIUM  P
ANDERSON, JANI  BRINGING DOWN THE MOON  SR-LP
ANDERSON, KEVIN J./MOESTA, REBECCA  AKIMA’S STORY [@TITAN A.E.]  P
ANDERSON, KEVIN J.  ANTIBODIES [X-FILES]  H
ANDERSON, KEVIN J./BEASON, DOUG  ASSEMBLERS OF INFINITY  C/P,P,H
ANDERSON, KEVIN J.  BLINDFOLD  P
ANDERSON, KEVIN J./MOESTA, REBECCA  CALE’S STORY [@TITAN A.E.]  P
ANDERSON, KEVIN J.  CAPTAIN NEMO  H,P
ANDERSON, KEVIN J.  CHAMPIONS OF THE FORCE [@STAR WARS JEDI ACADEMY 3]  C/P:2,P
ANDERSON, KEVIN J.  CLIMBING OLYMPUS  P
ANDERSON, KEVIN J./MOESTA, REBECCA  CRISIS AT CRYSTAL REEF [@STAR WARS YOUNG JEDI KNIGHTS]  P
ANDERSON, KEVIN J.  DARK APPRENTICE [@STAR WARS JEDI ACADEMY 2]  C/P:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  DARKEST KNIGHT [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J.  DARKSABER [@STAR WARS]  H:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  DELUSIONS OF GRANDEUR [@STAR WARS YOUNG JEDI KNIGHTS]  P
ANDERSON, KEVIN J./MOESTA, REBECCA  DIVERSITY ALLIANCE [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J./MOESTA, REBECCA  EMPEROR’S PLAGUE, THE [@STAR WARS YOUNG JEDI KNIGHTS]  P
ANDERSON, KEVIN J./BEASON, DOUG  FALLOUT [CRAIG KREIDENT 2]  P
ANDERSON, KEVIN J.  FANTASTIC VOYAGE MICROCOSM  P
ANDERSON, KEVIN J.  FOREST OF STARS, A [@SEVEN SUNS #2]  H,P
ANDERSON, KEVIN J.  FOREST OF STARS, A (PROOF) [@SEVEN SUNS #2]  SR-LP
ANDERSON, KEVIN J.  GAMEARTH [GAMEARTH 1]  P
ANDERSON, KEVIN J.  GAMEPLAY [GAMEARTH 2]  P
ANDERSON, KEVIN J.  GROUND ZERO [X-FILES]  P
ANDERSON, KEVIN J./MOESTA, REBECCA  HEIRS OF THE FORCE [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J.  HIDDEN EMPIRE [SEVEN SUNS 1]  C/P,P
ANDERSON, KEVIN J.  HOPSCOTCH  H
ANDERSON, KEVIN J.  HORIZON STORMS [@SEVEN SUNS #3]  H
ANDERSON, KEVIN J./BEASON, DOUG  IGNITION  H,P
ANDERSON, KEVIN J./BEASON, DOUG  ILL WIND  H:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  JABBA’S PALACE POP-UP BOOK [STAR WARS]  SR-MM
ANDERSON, KEVIN J./MOESTA, REBECCA  JEDI BOUNTY [@STAR WARS YOUNG JEDI KNIGHTS]  P
ANDERSON, KEVIN J.  JEDI SEARCH [@STAR WARS JEDI ACADEMY 1]  C/P:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  JEDI UNDER SIEGE [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J./BEASON, DOUG  LETHAL EXPOSURE [CRAIG KREIDENT 3]  C/P,P
ANDERSON, KEVIN J./BEASON, DOUG  LIFELINE  C/P:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  LIGHTSABERS [@STAR WARS YOUNG JEDI KNIGHTS]  C/P:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  LOST ONES, THE [@STAR WARS YOUNG JEDI KNIGHTS]  C/P:2,P
ANDERSON, KEVIN J.  RESURRECTION, INC.  C/P,P
ANDERSON, KEVIN J./MOESTA, REBECCA  RETURN TO ORD MANTELL [@STAR WARS YOUNG JEDI KNIGHTS]  P
ANDERSON, KEVIN J.  RUINS [X-FILES]  H,P
ANDERSON, KEVIN J.  SCATTERED SUNS [SEVEN SUNS 4]  H
ANDERSON, KEVIN J./MOESTA, REBECCA  SHADOW ACADEMY [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J./MOESTA, REBECCA  SHARDS OF ALDERAAN [@STAR WARS YOUNG JEDI KNIGHTS]  C/P,P
ANDERSON, KEVIN J.  TALES FROM JABBA’S PALACE [STAR WARS]  PA,C/PA
ANDERSON, KEVIN J.  TALES FROM THE MOS EISLEY CANTINA [STAR WARS]  C/PA,PA
ANDERSON, KEVIN J.  TALES OF THE BOUNTY HUNTERS [STAR WARS]  PA

The MITSFS Pinkdex  September 24, 2007  Page 10
AUTHOR  TITLE  CODE
ANDERSON, KEVIN J./BEASON, DOUG  TRINITY PARADOX, THE  C/P:2,P
ANDERSON, KEVIN J./MOESTA, REBECCA  TROUBLE ON CLOUD CITY [@STAR WARS YOUNG JEDI KNIGHTS]  P

ANDERSON, KEVIN J./TERANISHI, ROBERT/ BROOME, WENDY VEILED ALLIANCES [@SEVEN SUNS]

CX

ANDERSON, KEVIN J./BEASON, DOUG  VIRTUAL DESTRUCTION [CRAIG KREIDENT 1]  C/P,P
ANDERSON, KEVIN J.  WAR OF THE WORLDS: GLOBAL DISPATCHES  PA
ANDERSON, M. T.  WHALES ON STILTS [THRILLING TALES 1]  H
ANDERSON, MARGARET J.  IN THE CIRCLE OF TIME  P
ANDERSON, MARGARET J.  IN THE KEEP OF TIME  P
ANDERSON, MARGARET J.  MISTS OF TIME, THE  H
ANDERSON, POUL  AFTER DOOMSDAY  C/P:2,P

ANDERSON, POUL AGENT OF THE TERRAN EMPIRE (ORIGINAL) (BRITISH)

[DOMINIC FLANDRY]

SR-P

ANDERSON, POUL  AGENT OF THE TERRAN EMPIRE (REVISED) [DOMINIC FLANDRY]  C/P:2,P
ANDERSON, POUL  ALIGHT IN THE VOID  C/P,P
ANDERSON, POUL  ALL ONE UNIVERSE  LP
ANDERSON, POUL  ANNALS OF THE TIME PATROL  H:2
ANDERSON, POUL  ARMIES OF ELFLAND, THE  C/P:2,P
ANDERSON, POUL  AVATAR, THE  C/P,P,H
ANDERSON, POUL  AVONTUUR IN HET VERLEDEN  FP
ANDERSON, POUL  BEST OF POUL ANDERSON, THE  C/P,P
ANDERSON, POUL  BEYOND THE BEYOND  SR-H,H:2,P
ANDERSON, POUL  BOAT OF A MILLION YEARS, THE  H:2,SFWA-P,P
ANDERSON, POUL  BOOK OF POUL ANDERSON, THE  P
ANDERSON, POUL  BRAIN WAVE  C/P:2,P
ANDERSON, POUL  BROKEN SWORD, THE  C/P:2,SFWA-P,P
ANDERSON, POUL  BYWORLDER, THE  C/P:2,SFWA-P,P
ANDERSON, POUL  CIRCUS OF HELLS, A [DOMINIC FLANDRY]  C/P:2,P
ANDERSON, POUL  COLD VICTORY  C/P,P
ANDERSON, POUL  CONAN THE REBEL [BANTAM CONAN 6]  C/P,P
ANDERSON, POUL  CONFLICT  C/P:2,P
ANDERSON, POUL  CORRIDORS OF TIME, THE  C/P,H:2,P
ANDERSON, POUL/ANDERSON, KAREN  DAHUT [@KING OF YS 3]  C/P,SFWA-P,P
ANDERSON, POUL  DANCER FROM ATLANTIS, THE  H:2,P
ANDERSON, POUL  DARK BETWEEN THE STARS, THE  H,P
ANDERSON, POUL  DAY OF THEIR RETURN, THE  H:2,P

ANDERSON, POUL/ BROXON, MILDRED DOWNEY DEMON OF SCATTERY, THE

H:2,LP

ANDERSON, POUL  DEVIL’S GAME, THE  C/P:2,P
ANDERSON, POUL  DIALOGUE WITH DARKNESS  C/P:2,P
ANDERSON, POUL/ANDERSON, KAREN  DOG AND THE WOLF, THE [@KING OF YS 4]  P,SFWA-P
ANDERSON, POUL  EARTH BOOK OF STORMGATE, THE  H:2,P
ANDERSON, POUL  EARTHMAN, GO HOME!  DD479,P
ANDERSON, POUL/DICKSON, GORDON R.  EARTHMAN’S BURDEN  C/P:2,P
ANDERSON, POUL  ENEMY STARS, THE  SR-H,H:2,P
ANDERSON, POUL  ENSIGN FLANDRY [DOMINIC FLANDRY]  C/P:2,P
ANDERSON, POUL  EXPLORATIONS  C/P:2,P
ANDERSON, POUL  FANTASY  C/P:2,P
ANDERSON, POUL  FIRE TIME  H:2,P,SFWA:2
ANDERSON, POUL  FLANDRY [DOMINIC FLANDRY]  P,SFWA-P
ANDERSON, POUL  FLANDRY OF TERRA [DOMINIC FLANDRY]  C/P,P,H
ANDERSON, POUL  FLEET OF STARS, THE [HARVEST OF STARS 4]  C/P:2,P
ANDERSON, POUL  FOR LOVE AND GLORY  H:2
ANDERSON, POUL/ANDERSON, KAREN  GALLICENAE [@KING OF YS 2]  C/P,SFWA-P,P
ANDERSON, POUL  GAME OF EMPIRE, THE [DOMINIC FLANDRY]  C/P,P,H
ANDERSON, POUL  GENESIS  C/P,P
ANDERSON, POUL  GEVANGENEN VAN ATLANTIS, DE  FP
ANDERSON, POUL  GODS LAUGHED, THE  C/P:2,P
ANDERSON, POUL  GOING FOR INFINITY  H
ANDERSON, POUL  GOLDEN HORN, THE [@LAST VIKING #1]  P
ANDERSON, POUL  GOLDEN SLAVE, THE  C/P,P
ANDERSON, POUL  GUARDIANS OF TIME  C/P:2,P
ANDERSON, POUL  HARVEST OF STARS [HARVEST OF STARS 2]  H:2,P
ANDERSON, POUL  HARVEST THE FIRE [HARVEST OF STARS 1]  H:2,P
ANDERSON, POUL  HIGH CRUSADE, THE  H:2,SFWA-P,P
ANDERSON, POUL/DICKSON, GORDON R.  HOKA! [HOKA 1]  C/P,LP,H
ANDERSON, POUL/DICKSON, GORDON R.  HOKA! HOKA! HOKA! [HOKA 2]  P
ANDERSON, POUL/DICKSON, GORDON R.  HOKAS POKAS! [HOKA 3]  P
ANDERSON, POUL  HOMEBREW  SR-H
ANDERSON, POUL  HOMEWARD AND BEYOND  C/P:2,SFWA,P
ANDERSON, POUL  HORN OF TIME, THE  C/P:2,P
ANDERSON, POUL  HROLF KRAKI’S SAGA  C/P:2,SFWA-P,P
ANDERSON, POUL  IMPERIAL STARS, THE [DOMINIC FLANDRY]  H
ANDERSON, POUL  INCONSTANT STAR [MAN-KZIN WARS]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 11
AUTHOR  TITLE  CODE
ANDERSON, POUL/ANDERSON, KAREN  KING OF YS, THE, VOLUME I [KING OF YS 1,2]  H
ANDERSON, POUL/ANDERSON, KAREN  KING OF YS, THE, VOLUME II [KING OF YS 3,4]  H
ANDERSON, POUL  KINSHIP WITH THE STARS  C/P:2,P
ANDERSON, POUL  KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY]  H:2,P
ANDERSON, POUL  LET THE SPACEMEN BEWARE!  DF209
ANDERSON, POUL  LONG NIGHT, THE  C/P,P
ANDERSON, POUL  LONG WAY HOME, THE  C/P:2,P
ANDERSON, POUL  LONGEST VOYAGE, THE  SFWA-TD30,TD30
ANDERSON, POUL  MAKESHIFT ROCKET, THE  DF139
ANDERSON, POUL  MAN WHO COUNTS, THE  C/P:2,P
ANDERSON, POUL  MAURAI AND KITH  C/P:2,P
ANDERSON, POUL  MAYDAY ORBIT  P,DF104
ANDERSON, POUL  MERMAN’S CHILDREN, THE  C/P,P,H,SFWA
ANDERSON, POUL  MIDSUMMER TEMPEST, A  C/P,SFWA-P,H,P,SFWA
ANDERSON, POUL  MIRKHEIM  H:2,P
ANDERSON, POUL  MOTHER OF KINGS  H
ANDERSON, POUL  NEBULA AWARD STORIES 4  C/PA,PA
ANDERSON, POUL  NEW AMERICA [RUSTUM]  C/P:2,P
ANDERSON, POUL  NIGHT FACE, THE  C/P:2,P
ANDERSON, POUL/ANDERSON, KAREN  NIGHT FANTASTIC, THE  PA
ANDERSON, POUL  NO TRUCE WITH KINGS  TD5
ANDERSON, POUL  NO WORLD OF THEIR OWN  P,DD110
ANDERSON, POUL  OPERATION CHAOS [OPERATION CHAOS 1]  C/P:2,SFWA,P
ANDERSON, POUL  OPERATION LUNA [OPERATION CHAOS 2]  H,P
ANDERSON, POUL  ORBIT UNLIMITED [RUSTUM]  C/P:2,P
ANDERSON, POUL  ORION SHALL RISE  H:2,LP
ANDERSON, POUL  PAST TIMES  C/P:2,SFWA-P,P
ANDERSON, POUL  PEOPLE OF THE WIND, THE  C/P:2,SFWA-P,P
ANDERSON, POUL  PEOPLE OF THE WIND, THE/DAY OF THEIR RETURN, THE  P
ANDERSON, POUL  PEREGRINE, THE  C/P:2,P
ANDERSON, POUL  PLANET OF NO RETURN  DD199
ANDERSON, POUL  PSYCHOTECHNIC LEAGUE, THE  C/P:2,P
ANDERSON, POUL  QUEEN OF AIR AND DARKNESS AND OTHER STORIES, THE  C/P:2,P
ANDERSON, POUL  QUESTION AND ANSWER  C/P:2,P
ANDERSON, POUL  REBEL WORLDS, THE [DOMINIC FLANDRY]  C/P:2,P

ANDERSON, POUL REBEL WORLDS, THE [DOMINIC FLANDRY]/

KNIGHT OF GHOSTS AND SHADOWS, A [DOMINIC FLANDRY]

P

ANDERSON, POUL  ROAD OF THE SEA HORSE, THE [@LAST VIKING #2]  C/P:2,P
ANDERSON, POUL  ROGUE SWORD  P
ANDERSON, POUL/ANDERSON, KAREN  ROMA MATER [@KING OF YS 1]  C/P:2,P
ANDERSON, POUL  SATAN’S WORLD  SR-H,H:2,P
ANDERSON, POUL  SATURN GAME, THE  H,TD14,SFWA-TD14
ANDERSON, POUL  SEVEN CONQUESTS  C/P,SR-H,H,P
ANDERSON, POUL  SHIELD  C/P:2,P
ANDERSON, POUL  SHIELD OF TIME, THE  H:2,P
ANDERSON, POUL  SIGN OF THE RAVEN, THE [@LAST VIKING #3]  C/P,P
ANDERSON, POUL  SNOWS OF GANYMEDE, THE  DD303
ANDERSON, POUL/DICKSON, GORDON R.  SOUND AND THE FURRY, THE [HOKA 2,3]  H
ANDERSON, POUL  SPACE FOLK  C/P:2,SFWA-P,P
ANDERSON, POUL  STAR FOX, THE  SR-H,H:2,P
ANDERSON, POUL/DICKSON, GORDON R.  STAR PRINCE CHARLIE  C/P:2,P
ANDERSON, POUL  STAR WAYS  C/P:2,P,DD255
ANDERSON, POUL  STARFARERS  H:2
ANDERSON, POUL  STARS ARE ALSO FIRE, THE [HARVEST OF STARS 3]  H,P
ANDERSON, POUL  STARSHIP  C/P:2,P
ANDERSON, POUL  STONE IN HEAVEN, A [DOMINIC FLANDRY]  H:2,P
ANDERSON, POUL  STRANGERS FROM EARTH  SR-P,C/P:2
ANDERSON, POUL  TALES OF THE FLYING MOUNTAINS  C/P,P,H
ANDERSON, POUL  TAU ZERO  H:2,P
ANDERSON, POUL  THERE WILL BE TIME  H:2,P
ANDERSON, POUL  THERMONUCLEAR WARFARE  [C/RAN-P]
ANDERSON, POUL  THREE HEARTS AND THREE LIONS  H:2,P
ANDERSON, POUL  THREE WORLDS TO CONQUER  C/P:2,P
ANDERSON, POUL  TIME AND STARS  C/P,P,H,SR-H
ANDERSON, POUL  TIME PATROL, THE  H:2
ANDERSON, POUL  TIME PATROLMAN  C/P:2,P
ANDERSON, POUL  TO OUTLIVE ETERNITY  LP
ANDERSON, POUL  TRADER TO THE STARS  P,H:2,SR-H
ANDERSON, POUL  TROUBLE TWISTERS, THE  C/P:2,P
ANDERSON, POUL  TWILIGHT WORLD  H:2,P
ANDERSON, POUL  UN-MAN AND OTHER NOVELLAS  DF139
ANDERSON, POUL/ANDERSON, KAREN  UNICORN TRADE, THE  C/P:2,P
ANDERSON, POUL  VAULT OF THE AGES  H:2,SR-H,P
ANDERSON, POUL  VAULT OF THE AGES (JAPANESE)  FH

AUTHOR  TITLE  CODE
ANDERSON, POUL  VIRGIN PLANET  SR-P,C/P:2
ANDERSON, POUL  WAR OF THE GODS  H:2,P
ANDERSON, POUL  WAR OF THE WING-MEN  C/P:2,P,DD303
ANDERSON, POUL  WAR OF TWO WORLDS, THE  C/P,DD335
ANDERSON, POUL  WE CLAIM THESE STARS!  C/P:2,DD407,P
ANDERSON, POUL  WINNERS  P
ANDERSON, POUL  WINTER OF THE WORLD, THE  H:2,P
ANDERSON, POUL  WORLD WITHOUT STARS  C/P:2,P
ANDERSON, POUL  WORLDS OF POUL ANDERSON, THE  C/P:2,P
ANDERSON, POUL  YEAR OF THE RANSOM, THE  H
ANDERSON, WILLIAM C.  ADAM M-1  H
ANDERSON, WILLIAM C.  PENELOPE  C/P,P
ANDERSSON, C. DEAN  I AM FRANKENSTEIN  P
ANDERSSON, C. DEAN  RAW PAIN MAX  P
ANDERSSON, LEIF  INDEX TILL...HAPNA!  R/IND
ANDREEV, K.  AL’MANAKH NAUCHNOI FANTASTIKI 1964, VYPUSK 2  FH
ANDREWS, ALLEN  PIG PLANTAGENET, THE  C/P,P
ANDREWS, AUSTEN  MACHINATIONS [@ULTIMA TECHNOCRAT WAR #1]  P
ANDREWS, AUSTEN  MAELSTROM [@ULTIMA TECHNOCRAT WAR #3]  P
ANDREWS, AUSTEN  MASQUERADE [@ULTIMA TECHNOCRAT WAR #2]  P

ANDREWS, BART FROM THE BLOB TO STAR WARS:

THE SCIENCE FICTION MOVIE QUIZ BOOK

R/REF-P,C/REF-P

ANDREWS, BART/DUNNING, BRAD  STAR TREK QUIZ BOOK  [C/RAN-P:2]
ANDREWS, BART  TOLKIEN QUIZ BOOK, THE  R/REF-P
ANDREWS, CHRIS  MACHINATIONS  LPA
ANDREWS, DONNA  ACCESS DENIED [TURING HOPPER 3]  P
ANDREWS, DONNA  CLICK HERE FOR MURDER [TURING HOPPER 2]  P
ANDREWS, DONNA  DELETE ALL SUSPECTS [TURING HOPPER 4]  H
ANDREWS, ILONA  MAGIC BITES  P
ANDREWS, J. S.  BELL OF NENDRUM, THE  P
ANDREWS, KEITH WILLIAM  FREEDOM’S RANGERS [@FREEDOM’S RANGERS 1]  C/P,P
ANDREWS, KEITH WILLIAM  RAIDERS OF THE REVOLUTION [@FREEDOM’S RANGERS 2]  P
ANDREWS, KEITH WILLIAM  SEARCH AND DESTROY [@FREEDOM’S RANGERS #3]  P
ANDREWS, KEITH WILLIAM  SINK THE ARMADA! [@FREEDOM’S RANGERS #5]  P
ANDREWS, KEITH WILLIAM  SNOW KILL [@FREEDOM’S RANGERS #6]  P
ANDREWS, KEITH WILLIAM  TREASON IN TIME [@FREEDOM’S RANGERS #4]  P
ANDREWS, MICHAEL  THIEF’S TALE  LP
ANDREZEL, PIERRE  ANGELIC AVENGERS, THE  P
ANGHELIDES, PETER/COLE, STEPHEN  ANCESTOR CELL, THE [DOCTOR WHO EIGHTH DOCTOR]  P
ANGHELIDES, PETER  FRONTIER WORLDS [DOCTOR WHO EIGHTH DOCTOR]  P
ANGHELIDES, PETER  KURSAAL [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
ANIOLOWSKI, SCOTT DAVID  MADE IN GOATSWOOD  LP
ANIOLOWSKI, SCOTT DAVID  RETURN TO LOVECRAFT COUNTRY  LPA
ANOBILE, RICHARD J.  OUTLAND: THE MOVIE NOVEL  VLP
ANOBILE, RICHARD J.  WRATH OF KHAN, THE (PHOTOSTORY) [STAR TREK]  P
ANONYMOUS  ADULT VERSION OF THE ESCAPADES OF CLEOPATRA, THE  SR-L
ANONYMOUS  CITADELS ON EARTH [@FLASH GORDON #6]  C/P,P
ANONYMOUS  CITADELS UNDER ATTACK [@FLASH GORDON #5]  P
ANONYMOUS  CRISIS ON CITADEL II [@FLASH GORDON #3]  P
ANONYMOUS  FORCES FROM THE FEDERATION [@FLASH GORDON #4]  P
ANONYMOUS  INDEX TO FANTASY AND SCIENCE FICTION IN MUNSEY PUBLICATIONS  R/IND
ANONYMOUS  JAUFRY THE KNIGHT AND THE FAIR BRUNISSENDE  LP
ANONYMOUS  MABINOGION, THE  P
ANONYMOUS  MAGIA NEGRA: AMULETOS Y TALISMANES, LA  FP
ANONYMOUS  MAHABHARATA  H
ANONYMOUS  MASSACRE IN THE 22ND CENTURY [@FLASH GORDON #1]  C/P:2,P
ANONYMOUS  NO-FRILLS BOOK: SCIENCE FICTION  C/P,P
ANONYMOUS  SONG OF THE VOLSUNGS AND THE NIBELUNGS  P
ANONYMOUS  STAR TREK (VIEWMASTER STEREO PICTURES) [STAR TREK]  SR-MM
ANONYMOUS  STRANGE DEATH OF ADOLF HITLER, THE  H
ANONYMOUS  TALES FROM THE ARABIAN NIGHTS (BURTON TRANSLATION)  HA
ANONYMOUS  TWENTY-TWO GOBLINS (RYDER TRANSLATION)  H
ANONYMOUS  VOLSUNGA SAGA: THE STORY OF THE VOLSUNGS AND NIBELUNGS  P
ANSARA, MICHAEL  FRONTIER  H
ANSCOMBE, RODERICK  SECRET LIFE OF LASZLO, COUNT DRACULA, THE  H
ANSON, JAY  AMITYVILLE HORROR, THE  [C/RAN-P:2]
ANSTEY, F.  BRASS BOTTLE, THE  H
ANSTEY, F.  TINTED VENUS: A FARCICAL ROMANCE, THE  H
ANSTEY, F.  VICE VERSA  P
ANTCZAK, STEPHEN L.  DAYDREAMS UNDERTAKEN  LP
ANTHONY, MARK  BEYOND THE PALE [LAST RUNE 1]  P
ANTHONY, MARK  BLOOD OF MYSTERY [LAST RUNE 4]  P
ANTHONY, MARK  CRYPT OF THE SHADOWKING [@FORGOTTEN REALMS HARPERS #6]  P
ANTHONY, MARK  CURSE OF THE SHADOWMAGE [@FORGOTTEN REALMS HARPERS #11]  P
ANTHONY, MARK  DARK REMAINS, THE [LAST RUNE 3]  P

The MITSFS Pinkdex  September 24, 2007  Page 13
AUTHOR  TITLE  CODE
ANTHONY, MARK  ESCAPE FROM UNDERMOUNTAIN [@FORGOTTEN REALMS NOBLES #3]  P
ANTHONY, MARK  FIRST STONE, THE [LAST RUNE 6]  P
ANTHONY, MARK  GATES OF WINTER, THE [LAST RUNE 5]  P
ANTHONY, MARK  GATES OF WINTER, THE (PROOF) [LAST RUNE 5]  SR-LP
ANTHONY, MARK  KEEP OF FIRE, THE [LAST RUNE 2]  P
ANTHONY, MARK/PORATH, ELLEN  KINDRED SPIRITS [DRAGONLANCE MEETINGS 1]  C/P:2,P
ANTHONY, MARK  TOWER OF DOOM [@RAVENLOFT]  P
ANTHONY, PATRICIA  BROTHER TERMITE  C/P,P,H
ANTHONY, PATRICIA  COLD ALLIES  C/P,P,H
ANTHONY, PATRICIA  CONSCIENCE OF THE BEAGLE  C/P,SR-LP
ANTHONY, PATRICIA  CRADLE OF SPLENDOR  P
ANTHONY, PATRICIA  EATING MEMORIES  C/P,LP
ANTHONY, PATRICIA  GOD’S FIRES  C/P:2,P
ANTHONY, PATRICIA  HAPPY POLICEMAN  C/P,P
ANTHONY, PIERS  ALIEN PLOT  P
ANTHONY, PIERS  AND ETERNITY [INCARNATIONS OF IMMORTALITY 7]  C/P:2,P,H
ANTHONY, PIERS  ANTHONOLOGY  C/P,SFWA-P,P,H
ANTHONY, PIERS  BALOOK  P
ANTHONY, PIERS  BATTLE CIRCLE [BATTLE CIRCLE 1,2,3]  C/P:2,P
ANTHONY, PIERS  BEARING AN HOURGLASS [INCARNATIONS OF IMMORTALITY 2]  C/P,P,H
ANTHONY, PIERS  BEING A GREEN MOTHER [INCARNATIONS OF IMMORTALITY 5]  C/P,P,H

ANTHONY, PIERS BIO OF AN OGRE R/REF-P,SFWA,R/REF-H, C/REF-P:2,SFWA-P

ANTHONY, PIERS  BLUE ADEPT [APPRENTICE ADEPT #2]  C/P,@P,H
ANTHONY, PIERS  BUT WHAT OF EARTH?  C/P:2,SFWA-P,P
ANTHONY, PIERS/COULSON, ROBERT  BUT WHAT OF EARTH?  SR-P
ANTHONY, PIERS  CASTLE ROOGNA [@XANTH #3]  C/P:2,P
ANTHONY, PIERS  CASTLE ROOGNA (PROOF) [XANTH 3]  SR-LP
ANTHONY, PIERS/FARMER, PHILIP JOSE  CATERPILLAR’S QUESTION, THE  H,SFWA:2,P
ANTHONY, PIERS  CENTAUR AISLE [@XANTH #4]  C/P,P,H
ANTHONY, PIERS  CHAINING THE LADY [CLUSTER #2]  C/P:2,@P
ANTHONY, PIERS  CHAOS MODE [MODE 3]  C/P,P,H
ANTHONY, PIERS/MARGROFF, ROBERT E.  CHIMAERA’S COPPER [KELVIN HACKLEBERRY 3]  C/P,P,H
ANTHONY, PIERS  CHTHON [CHTHON #1]  C/P:2,@P
ANTHONY, PIERS  CLUSTER [CLUSTER #1]  C/P:2,@P
ANTHONY, PIERS  COLOR OF HER PANTIES, THE [@XANTH #15]  C/P,SFWA-P,P,H
ANTHONY, PIERS  CREWEL LYE [@XANTH #8]  C/P:2,P
ANTHONY, PIERS  CURRANT EVENTS [@XANTH #28]  H
ANTHONY, PIERS  DASTARD, THE [XANTH #24]  C/P,P,H
ANTHONY, PIERS/FUENTES, ROBERTO  DEAD MORN  C/P:2,SFWA-P,P
ANTHONY, PIERS  DEMONS DON’T DREAM [@XANTH #16]  H:2,P
ANTHONY, PIERS  DOOON MODE [MODE 4]  H
ANTHONY, PIERS  DOUBLE EXPOSURE [APPRENTICE ADEPT 1,2,3]  H:2
ANTHONY, PIERS  DRAGON ON A PEDESTAL [@XANTH #7]  C/P,P,H
ANTHONY, PIERS/MARGROFF, ROBERT E.  DRAGON’S GOLD [KELVIN HACKLEBERRY 1]  C/P:2,SFWA-P,P
ANTHONY, PIERS/BRADY, JULIE  DREAM A LITTLE DREAM  P
ANTHONY, PIERS/MARGROFF, ROBERT E.  E.S.P. WORM, THE  C/P:4,SFWA-P,P
ANTHONY, PIERS  EXECUTIVE [@BIO OF A SPACE TYRANT #4]  C/P:2,P
ANTHONY, PIERS  FAITH OF TAROT [TAROT #3]  C/P:2,@P
ANTHONY, PIERS  FAUN & GAMES [XANTH #21]  H,@P,@C/P
ANTHONY, PIERS  FIREFLY  H,P
ANTHONY, PIERS  FOR LOVE OF EVIL [INCARNATIONS OF IMMORTALITY 6]  C/P:2,P
ANTHONY, PIERS  FRACTAL MODE [MODE 2]  H:2,SFWA-P:2,P
ANTHONY, PIERS  GEIS OF THE GARGOYLE [@XANTH #18]  C/P,P
ANTHONY, PIERS  GHOST  C/P,SFWA-P,P,H
ANTHONY, PIERS  GOD OF TAROT [TAROT #1]  C/P:2,@P
ANTHONY, PIERS  GOLEM IN THE GEARS [@XANTH #9]  C/P,P,H
ANTHONY, PIERS/LEMING, RON  GUTBUCKET QUEST, THE  H
ANTHONY, PIERS  HARD SELL  C/P:2,P
ANTHONY, PIERS  HARPY THYME [@XANTH #17]  C/P,P,H
ANTHONY, PIERS  HASAN  C/P:2,SFWA-P,P
ANTHONY, PIERS  HEAVEN CENT [@XANTH #11]  H:2,P
ANTHONY, PIERS  HOPE OF EARTH [GEODYSSEY 3]  H
ANTHONY, PIERS  HOW PRECIOUS WAS THAT WHILE  R/REF-H,C/REF-H
ANTHONY, PIERS/LACKEY, MERCEDES  IF I PAY THEE NOT IN GOLD  H,SFWA,P
ANTHONY, PIERS  IRON MAIDEN, THE [BIO OF A SPACE TYRANT 6]  LP
ANTHONY, PIERS  ISLE OF VIEW [@XANTH #13]  H:2,SFWA,P
ANTHONY, PIERS  ISLE OF WOMAN [GEODYSSEY 1]  H:2,P
ANTHONY, PIERS  JUXTAPOSITION [APPRENTICE ADEPT #3]  C/P,@P,H
ANTHONY, PIERS  KEY TO CHROMA [CHROMAGIC 2]  H
ANTHONY, PIERS  KEY TO DESTINY [CHROMAGIC 3]  H
ANTHONY, PIERS  KEY TO HAVOC [CHROMAGIC 1]  LP
ANTHONY, PIERS  KEY TO LIBERTY [CHROMAGIC 4]  H
ANTHONY, PIERS  KILLOBYTE  C/P,SFWA-P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 14
AUTHOR  TITLE  CODE
ANTHONY, PIERS  KIRLIAN QUEST [CLUSTER #3]  C/P:2,@P
ANTHONY, PIERS  MACROSCOPE  C/P:2,P
ANTHONY, PIERS  MAGIC OF XANTH, THE [XANTH 1,2,3]  H
ANTHONY, PIERS  MAN FROM MUNDANIA [@XANTH #12]  C/P:2,P
ANTHONY, PIERS  MERCENARY [@BIO OF A SPACE TYRANT #2]  C/P:2,P
ANTHONY, PIERS  MERCYCLE  C/P:2,SFWA-P:2,P
ANTHONY, PIERS/FUENTES, ROBERTO  MISTRESS OF DEATH  P
ANTHONY, PIERS  MUSE OF ART [GEODYSSEY 4]  H
ANTHONY, PIERS  MUTE  C/P:2,P
ANTHONY, PIERS  NIGHT MARE [@XANTH #6]  C/P,P,H
ANTHONY, PIERS  OGRE, OGRE [@XANTH #5]  C/P:2,P
ANTHONY, PIERS  OMNIVORE [OF MAN AND MANTA 1]  C/P,P,H
ANTHONY, PIERS  ON A PALE HORSE [INCARNATIONS OF IMMORTALITY 1]  H:2,P
ANTHONY, PIERS/MARGROFF, ROBERT E.  ORC’S OPAL [KELVIN HACKLEBERRY 4]  C/P:2,P
ANTHONY, PIERS  ORN [OF MAN AND MANTA 2]  C/P,P,H
ANTHONY, PIERS  OUT OF PHAZE [APPRENTICE ADEPT #4]  C/P:2,@P
ANTHONY, PIERS  OX [OF MAN AND MANTA 3]  C/P,P,H
ANTHONY, PIERS  PET PEEVE [@XANTH #29]  P
ANTHONY, PIERS  PHAZE DOUBT [APPRENTICE ADEPT #7]  C/P,@P,SFWA-P:2,H
ANTHONY, PIERS  PHTHOR [CHTHON #2]  C/P:2,@P
ANTHONY, PIERS/NYE, JODY LYNN  PIERS ANTHONY’S VISUAL GUIDE TO XANTH [XANTH]  R/REF-P,C/REF-H:2
ANTHONY, PIERS  POLITICIAN [@BIO OF A SPACE TYRANT #3]  C/P:2,P
ANTHONY, PIERS  PORNUCOPIA  SR-L
ANTHONY, PIERS/HALL, FRANCES  PRETENDER  C/P:2,SFWA-P,LP
ANTHONY, PIERS  PROSTHO PLUS  C/P:2,SFWA-P,P

ANTHONY, PIERS/RICHEY, JAMES/ RIGGS, ALAN QUEST FOR THE FALLEN STAR

H

ANTHONY, PIERS  QUESTION QUEST [@XANTH #14]  C/P:2,P
ANTHONY, PIERS  RACE AGAINST TIME  C/P:2,SFWA-P,P
ANTHONY, PIERS  REALTY CHECK  LP
ANTHONY, PIERS  REFUGEE [@BIO OF A SPACE TYRANT #1]  C/P:2,P
ANTHONY, PIERS/MARGROFF, ROBERT E.  RING, THE  C/P:2,SFWA-P,P
ANTHONY, PIERS  RINGS OF ICE  C/P:2,P
ANTHONY, PIERS  ROBOT ADEPT [APPRENTICE ADEPT #5]  C/P:2,@P,SFWA-P
ANTHONY, PIERS  ROC AND A HARD PLACE [@XANTH #19]  C/P:2,P
ANTHONY, PIERS/TAEUSCH, JO ANNE  SECRET OF SPRING, THE  H:2,P
ANTHONY, PIERS/MARGROFF, ROBERT E.  SERPENT’S SILVER [KELVIN HACKLEBERRY 2]  H:2,P
ANTHONY, PIERS  SHADE OF THE TREE  C/P,SFWA-P,P,H
ANTHONY, PIERS  SHAME OF MAN [GEODESSEY 2]  H,P
ANTHONY, PIERS  SOS THE ROPE [BATTLE CIRCLE 1]  C/P,@P,H
ANTHONY, PIERS  SOURCE OF MAGIC, THE [@XANTH #2]  C/P:2,P
ANTHONY, PIERS  SPELL FOR CHAMELEON, A [@XANTH #1]  C/P:2,P
ANTHONY, PIERS  SPELL FOR CHAMELEON, A (ABRIDGED; CASSETTE)  SR-MM
ANTHONY, PIERS/PICKOVER, CLIFFORD A.  SPIDER LEGS  H,P
ANTHONY, PIERS  SPLIT INFINITY [APPRENTICE ADEPT #1]  C/P:2,@P
ANTHONY, PIERS  STATESMAN [@BIO OF A SPACE TYRANT #5]  C/P:2,P
ANTHONY, PIERS  STEPPE  H:2,P
ANTHONY, PIERS  STORK NAKED [@XANTH #30]  H
ANTHONY, PIERS  SWELL FOOP [XANTH #25]  H,P
ANTHONY, PIERS  TAROT [TAROT 1,2,3]  LP
ANTHONY, PIERS  TATHAM MOUND  H,SFWA-P,P
ANTHONY, PIERS  THOUSANDSTAR [CLUSTER #4]  C/P:2,@P
ANTHONY, PIERS/MARGROFF, ROBERT E.  THREE COMPLETE NOVELS [KELVIN HACKLEBERRY 1,2,3]  H
ANTHONY, PIERS/KORNWISE, ROBERT  THROUGH THE ICE  C/P:2,P
ANTHONY, PIERS  TOTAL RECALL  H
ANTHONY, PIERS  TOTAL RECALL (ABRIDGED)  C/P:2,P
ANTHONY, PIERS  TRIPLE DETENTE  C/P:2,SFWA-P,P

ANTHONY, PIERS/MALZBERG, BARRY N./ GREENBERG, MARTIN H./ WAUGH, CHARLES G.

UNCOLLECTED STARS

PA,C/PA

ANTHONY, PIERS  UNICORN POINT [APPRENTICE ADEPT #6]  C/P,SFWA-P,@P,H
ANTHONY, PIERS  UP IN A HEAVAL [XANTH #26]  H
ANTHONY, PIERS  VALE OF THE VOLE [@XANTH #10]  C/P:2,P
ANTHONY, PIERS  VAR THE STICK [BATTLE CIRCLE #2]  C/P:2,@P
ANTHONY, PIERS  VIRTUAL MODE [MODE 1]  C/P:2,SFWA-P:2,P
ANTHONY, PIERS  VISCOUS CIRCLE [CLUSTER #5]  C/P:2,@P
ANTHONY, PIERS  VISION OF TAROT [TAROT #2]  C/P:2,@P
ANTHONY, PIERS  WIELDING A RED SWORD [INCARNATIONS OF IMMORTALITY 4]  H:2,P,SFWA
ANTHONY, PIERS/TELLA, ALFRED  WILLING SPIRIT, THE  H,P
ANTHONY, PIERS  WITH A TANGLED SKEIN [INCARNATIONS OF IMMORTALITY 3]  H:2,P,SFWA
ANTHONY, PIERS  XONE OF CONTENTION [XANTH #23]  H
ANTHONY, PIERS  YON ILL WIND [XANTH #20]  H,@P,@C/P
ANTHONY, PIERS  ZOMBIE LOVER [XANTH #22]  H
ANTIEAU, KIM  COYOTE COWGIRL  H

AUTHOR  TITLE  CODE
ANTIEAU, KIM  GAIA WEBSTERS, THE  LP
ANTIEAU, KIM  JIGSAW WOMAN, THE  LP
ANTILL, KEITH  MOON IN THE GROUND  H
ANTON, UWE  BRENNENDEN SCHIFFE, DIE [@PERRY RHODAN ANDROMEDA 1]  FPR
ANTWORTH, TERRANCE  DEATH FROM THE SEA  LP
ANVIL, CHRISTOPHER  DAY THE MACHINES STOPPED, THE  P
ANVIL, CHRISTOPHER  INTERSTELLAR PATROL [INTERSTELLAR PATROL]  P

ANVIL, CHRISTOPHER INTERSTELLAR PATROL II: THE FEDERATION OF HUMANITY

[INTERSTELLAR PATROL]

H

ANVIL, CHRISTOPHER  PANDORA’S LEGIONS  P
ANVIL, CHRISTOPHER  PANDORA’S PLANET  C/P,SFWA,P,H
ANVIL, CHRISTOPHER  STEEL, THE MIST, AND THE BLAZING SUN, THE  P
ANVIL, CHRISTOPHER  STRANGERS IN PARADISE  C/P:2,P
ANVIL, CHRISTOPHER/FLINT, ERIC  TROUBLE WITH ALIENS, THE  H
ANVIL, CHRISTOPHER  WARLORD’S WORLD  P
ANZETTI, TONI  RIDERS OF LEVIATHAN [TYPHON 2]  P
ANZETTI, TONI  TYPHON’S CHILDREN [TYPHON 1]  P
APPEL, ALLEN  TIME AFTER TIME  H
APPEL, ALLEN  TWICE UPON A TIME  P
APPEL, BENJAMIN  DEATH MASTER, THE  P
APPEL, BENJAMIN  DEVIL AND W. KASPAR, THE  C/P,P
APPEL, BENJAMIN  FANTASTIC MIRROR, THE  R/REF-H
APPEL, BENJAMIN  FUNHOUSE, THE  C/P,P
APPEL, WILLIAM  WATER WORLD  P
APPLE, A. E.  MR. CHANG’S CRIME RAY  H
APPLEBY, KEN  VOICE OF CEPHEUS, THE  C/P:2,P
APPLEGATE, K. A.  ENTER THE ENCHANTED [@EVERWORLD #3]  P
APPLETON, VICTOR  ALIEN PROBE, THE [@NEW TOM SWIFT #3]  H,LP
APPLETON, VICTOR  AQUATECH WARRIORS [@TOM SWIFT (ARCHWAY) #6]  P
APPLETON, VICTOR  ARK TWO [@NEW TOM SWIFT #7]  H,LP
APPLETON, VICTOR  ASTRAL FORTRESS, THE [@NEW TOM SWIFT #5]  C/P,LP
APPLETON, VICTOR  BLACK DRAGON, THE [@TOM SWIFT (ARCHWAY) #1]  P
APPLETON, VICTOR  CITY IN THE STARS, THE [@NEW TOM SWIFT #1]  H,LP
APPLETON, VICTOR  CRATER OF MYSTERY [@NEW TOM SWIFT #8]  LP
APPLETON, VICTOR  CYBORG KICKBOXER [@TOM SWIFT (ARCHWAY) #3]  P
APPLETON, VICTOR  DEATH QUAKE [@TOM SWIFT (ARCHWAY) #12]  P
APPLETON, VICTOR  DNA DISASTER, THE [@TOM SWIFT (ARCHWAY) #4]  P
APPLETON, VICTOR  FIRE BIKER [@TOM SWIFT (ARCHWAY) #9]  P
APPLETON, VICTOR  GATEWAY TO DOOM [@NEW TOM SWIFT #9]  LP
APPLETON, VICTOR  INVISIBLE FORCE, THE [@NEW TOM SWIFT #10]  LP
APPLETON, VICTOR  MICROBOTS, THE [@TOM SWIFT (ARCHWAY) #8]  P
APPLETON, VICTOR  MIND GAMES [@TOM SWIFT (ARCHWAY) #10]  P
APPLETON, VICTOR  MONSTER MACHINE [@TOM SWIFT (ARCHWAY) #5]  P
APPLETON, VICTOR  MOONSTALKER [@TOM SWIFT (ARCHWAY) #7]  P
APPLETON, VICTOR  MUTANT BEACH [@TOM SWIFT (ARCHWAY) #11]  P
APPLETON, VICTOR  NEGATIVE ZONE, THE [@TOM SWIFT (ARCHWAY) #2]  P
APPLETON, VICTOR  PLANET OF NIGHTMARES [@NEW TOM SWIFT #11]  LP
APPLETON, VICTOR  RESCUE MISSION, THE [@NEW TOM SWIFT #6]  C/P,LP
APPLETON, VICTOR  TERROR ON THE MOONS OF JUPITER [@NEW TOM SWIFT #2]  LP
APPLETON, VICTOR  TOM SWIFT AMONG THE DIAMOND MAKERS [@TOM SWIFT #7]  H
APPLETON, VICTOR  TOM SWIFT AMONG THE FIRE FIGHTERS [@TOM SWIFT #24]  H
APPLETON, VICTOR  TOM SWIFT AND HIS AERIAL WARSHIP [@TOM SWIFT #18]  H
APPLETON, VICTOR  TOM SWIFT AND HIS AIR GLIDER [@TOM SWIFT #12]  H
APPLETON, VICTOR  TOM SWIFT AND HIS AIR SCOUT [@TOM SWIFT #22]  H
APPLETON, VICTOR  TOM SWIFT AND HIS AIRLINE EXPRESS [@TOM SWIFT #29]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS AIRSHIP [@TOM SWIFT #3]  H
APPLETON, VICTOR  TOM SWIFT AND HIS BIG DIRIGIBLE [@TOM SWIFT #33]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS BIG TUNNEL [@TOM SWIFT #19]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS CHEST OF SECRETS [@TOM SWIFT #28]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS ELECTRIC LOCOMOTIVE [@TOM SWIFT #25]  H
APPLETON, VICTOR  TOM SWIFT AND HIS ELECTRIC RIFLE [@TOM SWIFT #10]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS ELECTRIC RUNABOUT [@TOM SWIFT #5]  H
APPLETON, VICTOR  TOM SWIFT AND HIS FLYING BOAT [@TOM SWIFT #26]  H
APPLETON, VICTOR  TOM SWIFT AND HIS GIANT CANNON [@TOM SWIFT #16]  H
APPLETON, VICTOR  TOM SWIFT AND HIS GIANT MAGNET [@TOM SWIFT #35]  H
APPLETON, VICTOR  TOM SWIFT AND HIS GIANT TELESCOPE  BLB
APPLETON, VICTOR  TOM SWIFT AND HIS GREAT OIL GUSHER [@TOM SWIFT #27]  H
APPLETON, VICTOR  TOM SWIFT AND HIS GREAT SEARCHLIGHT [@TOM SWIFT #15]  H
APPLETON, VICTOR  TOM SWIFT AND HIS HOUSE ON WHEELS [@TOM SWIFT #32]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS MAGNETIC SILENCER  BLB
APPLETON, VICTOR  TOM SWIFT AND HIS MOTOR-BOAT [@TOM SWIFT #2]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS MOTOR-CYCLE [@TOM SWIFT #1]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS OCEAN AIRPORT [@TOM SWIFT #37]  H
APPLETON, VICTOR  TOM SWIFT AND HIS PHOTO TELEPHONE [@TOM SWIFT #17]  H:2

AUTHOR  TITLE  CODE
APPLETON, VICTOR  TOM SWIFT AND HIS PLANET STONE [@TOM SWIFT #38]  H
APPLETON, VICTOR  TOM SWIFT AND HIS SKY RACER [@TOM SWIFT #9]  H
APPLETON, VICTOR  TOM SWIFT AND HIS SKY TRAIN [@TOM SWIFT #34]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS SUBMARINE BOAT [@TOM SWIFT #4]  H
APPLETON, VICTOR  TOM SWIFT AND HIS TALKING PICTURES [@TOM SWIFT #31]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS TELEVISION DETECTOR [@TOM SWIFT #36]  H:2
APPLETON, VICTOR  TOM SWIFT AND HIS UNDERSEA SEARCH [@TOM SWIFT #23]  H
APPLETON, VICTOR  TOM SWIFT AND HIS WAR TANK [@TOM SWIFT #21]  H
APPLETON, VICTOR  TOM SWIFT AND HIS WIRELESS MESSAGE [@TOM SWIFT #6]  H
APPLETON, VICTOR  TOM SWIFT AND HIS WIZARD CAMERA [@TOM SWIFT #14]  H
APPLETON, VICTOR  TOM SWIFT CIRCLING THE GLOBE [@TOM SWIFT #30]  H
APPLETON, VICTOR  TOM SWIFT IN CAPTIVITY [@TOM SWIFT #13]  H
APPLETON, VICTOR  TOM SWIFT IN THE CAVES OF ICE [@TOM SWIFT #8]  H
APPLETON, VICTOR  TOM SWIFT IN THE CITY OF GOLD [@TOM SWIFT #11]  H
APPLETON, VICTOR  TOM SWIFT IN THE LAND OF WONDERS [@TOM SWIFT #20]  H
APPLETON, VICTOR  WAR IN OUTER SPACE, THE [@NEW TOM SWIFT #4]  H,LP
APPLETON, VICTOR II  TOM SWIFT AND HIS AQUATOMIC TRACKER [@TOM SWIFT JR. #23]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS ATOMIC EARTH BLASTER [@TOM SWIFT JR. #5]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS COSMOTRON EXPRESS [@TOM SWIFT JR. #32]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS DEEP-SEA HYDRODOME [@TOM SWIFT JR. #11]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS DIVING SEACOPTER [@TOM SWIFT JR. #7]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS DYNA-4 CAPSULE [@TOM SWIFT JR. #31]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS ELECTRONIC RETROSCOPE [@TOM SWIFT JR. #14]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS FLYING LAB [@TOM SWIFT JR. #1]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS G-FORCE INVERTER [@TOM SWIFT JR. #30]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS GIANT ROBOT [@TOM SWIFT JR. #4]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS JETMARINE [@TOM SWIFT JR. #2]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS MEGASCOPE SPACE PROBER [@TOM SWIFT JR. #20]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS OUTPOST IN SPACE [@TOM SWIFT JR. #6]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS POLAR-RAY DYNASPHERE [@TOM SWIFT JR. #25]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS REPELATRON SKYWAY [@TOM SWIFT JR. #22]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS ROCKET SHIP [@TOM SWIFT JR. #3]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS SONIC BOOM TRAP [@TOM SWIFT JR. #26]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS SPACE SOLARTRON [@TOM SWIFT JR. #13]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS SPECTROMARINE SELECTOR [@TOM SWIFT JR. #15]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS SUBOCEAN GEOTRON [@TOM SWIFT JR. #27]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS 3-D TELEJECTOR [@TOM SWIFT JR. #24]  H
APPLETON, VICTOR II  TOM SWIFT AND HIS TRIPHIBIAN ATOMICAR [@TOM SWIFT JR. #19]  H:2
APPLETON, VICTOR II  TOM SWIFT AND HIS ULTRASONIC CYCLOPLANE [@TOM SWIFT JR. #10]  H:2
APPLETON, VICTOR II  TOM SWIFT AND THE ASTEROID PIRATES [@TOM SWIFT JR. #21]  H:2
APPLETON, VICTOR II  TOM SWIFT AND THE CAPTIVE PLANETOID [@TOM SWIFT JR. #29]  H:2
APPLETON, VICTOR II  TOM SWIFT AND THE COSMIC ASTRONAUTS [@TOM SWIFT JR. #16]  H:2
APPLETON, VICTOR II  TOM SWIFT AND THE ELECTRONIC HYDROLUNG [@TOM SWIFT JR. #18]  H:2
APPLETON, VICTOR II  TOM SWIFT AND THE GALAXY GHOSTS [@TOM SWIFT JR. #33]  H
APPLETON, VICTOR II  TOM SWIFT AND THE MYSTERY COMET [@TOM SWIFT JR. #28]  H
APPLETON, VICTOR II  TOM SWIFT AND THE VISITOR FROM PLANET X [@TOM SWIFT JR. #17]  H:2
APPLETON, VICTOR II  TOM SWIFT IN THE CAVES OF NUCLEAR FIRE [@TOM SWIFT JR. #8]  H:2
APPLETON, VICTOR II  TOM SWIFT IN THE RACE TO THE MOON [@TOM SWIFT JR. #12]  H:2
APPLETON, VICTOR II  TOM SWIFT ON THE PHANTOM SATELLITE [@TOM SWIFT JR. #9]  H:2
ARATYR, RISA  HUNTER OF THE LIGHT  P
ARCH, E. L.  BRIDGE TO YESTERDAY  H
ARCH, E. L.  MAN WITH THREE EYES, THE  H
ARCHER, GEOFFREY  SKYDANCER  P,SFWA-P
ARCHER, NATHAN  COLD WAR [@PREDATOR 2]  C/P,P
ARCHER, NATHAN  CONCRETE JUNGLE [@PREDATOR 1]  P
ARCHER, NATHAN  MARTIAN DEATHTRAP [@MARS ATTACKS]  C/P,P,H
ARCHER, NATHAN  RAGNAROK [STAR TREK VOYAGER]  P
ARCHER, NATHAN  VALHALLA [STAR TREK DS9]  C/P,P
ARCHER, PETER  GREAT WHITE WYRM, THE [@DRAGONLANCE CHAMPIONS]  P
ARCHER, PETER  RATH AND STORM [@MAGIC: THE GATHERING]  PA
ARDEN, TOM  EMPRESS OF THE ENDLESS DREAM [@OROKON #5]  P
ARDEN, TOM  HARLEQUIN’S DANCE [@OROKON #1]  P
ARDEN, TOM  KING AND QUEEN OF SWORDS, THE [@OROKON #2]  P
ARDEN, TOM  NIGHTDREAMERS [DOCTOR WHO]  SR-H
ARDEN, TOM  SHADOW BLACK  LP
ARDEN, TOM  SISTERHOOD OF THE BLUE STORM [@OROKON #4]  H,P
ARDEN, TOM  SULTAN OF THE MOON AND STARS [@OROKON #3]  P
ARDIES, TOM  THEIR MAN IN THE WHITE HOUSE  SFWA
ARDREY, ROBERT  WORLDS BEGINNING  H
ARETON, EMIL COHEN  NEW ATLANTEANS, THE  LP
ARGO, JOHN  LANTERN ROAD  LP
ARGO, JOHN  PIONEERS  LP
ARGO, JOHN  THIS SHOAL OF SPACE  LP

The MITSFS Pinkdex  September 24, 2007  Page 17
AUTHOR  TITLE  CODE
ARIOSTO, LUDOVICO  RING OF ANGELICA, THE [@ORLANDO FURIOSO 1]  C/P,P
ARISIA 1998 CONVENTION COMMITTEE  ARISIA ’98 (PROGRAM BOOK)  VLPA
ARISIA 2000 CONVENTION COMMITTEE  ARISIA ’00 (PROGRAM BOOK)  VLPA
ARISS, BRUCE  FULL CIRCLE  H
ARLEN, MICHAEL  MAN’S MORTALITY  H
ARMSTRONG, JENNIFER/BUTCHER, NANCY  KILN, THE [FIRE-US]  H
ARMSTRONG, KELLEY  BITTEN [WOMEN OF THE OTHERWORLD 1]  P
ARMSTRONG, KELLEY  BROKEN [WOMEN OF THE OTHERWORLD]  P

ARMSTRONG, KELLEY DIME STORE MAGIC

[WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1]

C/P:2,P

ARMSTRONG, KELLEY  HAUNTED [WOMEN OF THE OTHERWORLD]  P

ARMSTRONG, KELLEY INDUSTRIAL MAGIC

[WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 2]

C/P,P

ARMSTRONG, KELLEY  NO HUMANS INVOLVED [WOMEN OF THE OTHERWORLD]  H
ARMSTRONG, KELLEY  STOLEN [WOMEN OF THE OTHERWORLD 2]  P
ARMSTRONG, KELLEY  WITCH MAGIC [WOMEN OF THE OTHERWORLD PAIGE WINTERBOURNE 1,2] H
ARMSTRONG, MICHAEL  AFTER THE ZAP  C/P,P
ARMSTRONG, MICHAEL  AGVIQ  C/P,P
ARMSTRONG, MICHAEL  HIDDEN WAR, THE  P
ARMSTRONG, SARAH  BLOOD RED ROSES [TWILIGHT 8]  P
ARMY, DEPARTMENT OF THE  USSR: MISSILES, ROCKETS AND THE SPACE EFFORT  [C/RAN-H]
ARNASON, ELEANOR  CHANGING WOMEN [@WOMAN OF THE IRON PEOPLE #2]  P,SFWA-P
ARNASON, ELEANOR  DAUGHTER OF THE BEAR KING  P
ARNASON, ELEANOR  IN THE LIGHT OF SIGMA DRACONIS [@WOMAN OF THE IRON PEOPLE #1]  P
ARNASON, ELEANOR  ORDINARY PEOPLE  LP
ARNASON, ELEANOR  RING OF SWORDS  H
ARNASON, ELEANOR  SWORD SMITH, THE  P
ARNASON, ELEANOR  TO THE RESURRECTION STATION  P
ARNASON, ELEANOR  WOMAN OF THE IRON PEOPLE, A [WOMAN OF THE IRON PEOPLE 1,2]  H
ARNESON, D. J./TALLARICO, TONY  BLACK STAR CHRONICLES  P
ARNOLD, ALAN  ONCE UPON A GALAXY  R/REF-P,C/REF-P
ARNOLD, EDWIN LESTER  GULLIVER OF MARS  C/P:2,P
ARNOLD, EDWIN LESTER  WONDERFUL ADVENTURES OF PHRA THE PHOENICIAN, THE  H,LP
ARONICA, LOU/MCCARTHY, SHAWNA  FULL SPECTRUM  PA,C/PA

ARONICA, LOU/MCCARTHY, SHAWNA/ STOUT, AMY/LOBRUTTO, PAT FULL SPECTRUM 2

C/PA,PA

ARONICA, LOU/STOUT, AMY/ MITCHELL, BETSY FULL SPECTRUM 3

C/PA,PA,SR-LPA

ARONICA, LOU/STOUT, AMY/ MITCHELL, BETSY FULL SPECTRUM 4

C/PA,LPA

ARONICA, LOU/STOUT, AMY/ MITCHELL, BETSY FULL SPECTRUM 4 (PROOF)

SR-LPA

ARROW, WILLIAM ESCAPE FROM TERROR LAGOON

[@RETURN TO THE PLANET OF THE APES #2]

P

ARROW, WILLIAM  MAN, THE HUNTED ANIMAL [@RETURN TO THE PLANET OF THE APES #3]  P
ARROW, WILLIAM  VISIONS FROM NOWHERE [@RETURN TO THE PLANET OF THE APES #1]  P
ARROWSMITH, NANCY/MOORE, GEORGE  FIELD GUIDE TO THE LITTLE PEOPLE, A  R/REF-P
ARTHUR, KERI  DANGEROUS GAMES [RILEY JENSON 4]  P
ARTHUR, KERI  FULL MOON RISING [RILEY JENSON 1]  P
ARTHUR, KERI  KISSING SIN [RILEY JENSON 2]  P
ARTHUR, KERI  TEMPTING EVIL [RILEY JENSON 3]  P
ARTHUR, ROBERT  MONSTER MIX  PA
ASARO, CATHERINE  ALPHA [CHARON 2]  H
ASARO, CATHERINE  ASCENDANT SUN [SKOLIAN EMPIRE 5]  H,P
ASARO, CATHERINE  CATCH THE LIGHTNING [SKOLIAN EMPIRE 2]  C/P:2,P
ASARO, CATHERINE  CHARMED SPHERE, THE  LP
ASARO, CATHERINE  DAWN STAR, THE [MEL DAWNFIELD 2]  LP
ASARO, CATHERINE  FINAL KEY, THE [SKOLIAN TRIAD 2]  H
ASARO, CATHERINE  FIRE OPAL, THE [LOST CONTINENT 2]  LP
ASARO, CATHERINE  IRRESISTIBLE FORCES  LPA
ASARO, CATHERINE  LAST HAWK, THE [SKOLIAN EMPIRE 3]  C/P,P,H
ASARO, CATHERINE  MISTED CLIFFS, THE [LOST CONTINENT 1]  LP
ASARO, CATHERINE  MOON’S SHADOW, THE [SKOLIAN EMPIRE 8]  H
ASARO, CATHERINE  PHOENIX CODE, THE  P
ASARO, CATHERINE  PRIMARY INVERSION [SKOLIAN EMPIRE 1]  C/P,P
ASARO, CATHERINE  QUANTUM ROSE, THE [SKOLIAN EMPIRE 6]  H,P
ASARO, CATHERINE  RADIANT SEAS, THE [SKOLIAN EMPIRE 4]  C/P,P,H
ASARO, CATHERINE  SCHISM [@TRIAD #1]  H
ASARO, CATHERINE  SKYFALL [SKOLIAN EMPIRE 9]  P
ASARO, CATHERINE  SPHERICAL HARMONIC [SKOLIAN EMPIRE 7]  H
ASARO, CATHERINE  SUNRISE ALLEY [CHARON 1]  H
ASARO, CATHERINE  VEILED WEB, THE  P
ASBURY, HERBERT  DEVIL OF PEI-LING, THE  H
ASH, BRIAN  FACES OF THE FUTURE  R/REF-H

AUTHOR  TITLE  CODE
ASH, BRIAN  WHO’S WHO IN SCIENCE FICTION  R/REF-H
ASH, CONSTANCE  HORSEGIRL, THE [HORSEGIRL 1]  C/P,P
ASH, CONSTANCE  NOT OF WOMAN BORN  PA
ASH, CONSTANCE  STALKING HORSE, THE [HORSEGIRL 2]  C/P,P
ASH, CONSTANCE  STALLION QUEEN, THE  C/P,SFWA-P,P
ASH, FENTON  TRIP TO MARS, A  SR-H
ASH, SARAH  CHILDREN OF THE SERPENT GATE [TEARS OF ARTAMON 3]  H
ASH, SARAH  LORD OF SNOW AND SHADOWS [TEARS OF ARTAMON 1]  C/P,P
ASH, SARAH  LORD OF SNOW AND SHADOWS (PROOF) [TEARS OF ARTAMON 1]  SR-LP
ASH, SARAH  PRISONER OF THE IRON TOWER [TEARS OF ARTAMON 2]  H
ASHBY, RICHARD  ACT OF GOD  SR-P,C/P
ASHE, GEOFFREY  DISCOVERY OF KING ARTHUR, THE  SFWA
ASHER, MARTY  SHELTER  H,SFWA,LP
ASHER, NEAL  AFRICA ZERO  LP
ASHER, NEAL  BRASS MAN [POLITY GRIDLINKED 3]  P
ASHER, NEAL  COWL  H
ASHER, NEAL  ENGINEER, THE  LP
ASHER, NEAL  ENGINEER RECONDITIONED, THE  LP
ASHER, NEAL  GRIDLINKED [POLITY GRIDLINKED 1]  C/P,P,H
ASHER, NEAL  LINE OF POLITY, THE [POLITY GRIDLINKED 2]  C/P,P
ASHER, NEAL  PARASITE, THE  LP
ASHER, NEAL  POLITY AGENT [POLITY GRIDLINKED 4]  H
ASHER, NEAL  PRADOR MOON [POLITY]  LP
ASHER, NEAL  RUNCIBLE TALES [POLITY]  SR-LP
ASHER, NEAL  SKINNER, THE [POLITY SKINNER 1]  H:2,LP
ASHER, NEAL  VOYAGE OF THE SABLE KEECH, THE [POLITY SKINNER 2]  H
ASHERMAN, ALLAN  MAKING OF STAR TREK II: THE WRATH OF KHAN, THE [STAR TREK]  R/REF-P
ASHERMAN, ALLAN  STAR TREK INTERVIEW BOOK, THE  R/REF-P
ASHERMAN, ALLEN  STAR TREK COMPENDIUM, THE (REVISED 1986) [STAR TREK]  R/REF-H
ASHERMAN, ALLEN  STAR TREK COMPENDIUM, THE (REVISED 1993) [STAR TREK]  R/REF-H
ASHLEY, JENNIFER  CALLING, THE [@IMMORTALS]  P
ASHLEY, MIKE  ALGERNON BLACKWOOD  R/IND
ASHLEY, MIKE  ALMAS DE METAL  FP
ASHLEY, MIKE  CAMELOT CHRONICLES, THE  LPA
ASHLEY, MIKE  COMPLETE INDEX TO ASTOUNDING/ANALOG, THE  R/IND
ASHLEY, MIKE  GATEWAYS TO FOREVER  R/REF-H
ASHLEY, MIKE/LOWNDES, ROBERT A. W.  GERNSBACK DAYS, THE  R/REF-H
ASHLEY, MIKE  ILLUSTRATED BOOK OF SCIENCE FICTION LISTS, THE  R/REF-H
ASHLEY, MIKE  MAMMOTH BOOK OF ARTHURIAN LEGENDS, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF AWESOME COMIC FANTASY, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF COMIC FANTASY, THE  HA,LPA
ASHLEY, MIKE  MAMMOTH BOOK OF COMIC FANTASY II, THE  HA
ASHLEY, MIKE  MAMMOTH BOOK OF EXTREME SCIENCE FICTION, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF FANTASY, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF NEW COMIC FANTASY, THE  HA
ASHLEY, MIKE/BROWN, ERIC  MAMMOTH BOOK OF NEW JULES VERNE ADVENTURES, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF SCIENCE FICTION, THE  LPA
ASHLEY, MIKE  MAMMOTH BOOK OF SORCERERS’ TALES, THE  LPA
ASHLEY, MIKE  MERLIN CHRONICLES, THE  PA
ASHLEY, MIKE  PENDRAGON CHRONICLES, THE  HA
ASHLEY, MIKE  RANDOM HOUSE BOOK OF FANTASY STORIES, THE  LPA
ASHLEY, MIKE  RANDOM HOUSE BOOK OF SCIENCE FICTION STORIES, THE  LPA
ASHLEY, MIKE  SOULS IN METAL  HA,PA
ASHLEY, MIKE  TIME MACHINES, THE  R/REF-H
ASHLEY, MIKE  TRANSFORMATIONS  R/REF-H
ASHLEY, MIKE  WHO’S WHO IN HORROR AND FANTASY FICTION  R/REF-H
ASHLEY, STEVEN  LOVE OUT OF TIME  P
ASHMAN, HOWARD  MANDRAKE THE MAGICIAN  C/P,P
ASHTON, FRANCIS  ALAS, THAT GREAT CITY  H
ASHTON, FRANCIS  BREAKING OF THE SEALS, THE  H,LP
ASHTON, FRANCIS/ASHTON, STEPHEN  WRONG SIDE OF THE MOON  SR-P
ASHTON, MARVIN  PEOPLE OF ASA  SR-P
ASHWELL, PAULINE  PROJECT FARCRY  LP
ASHWELL, PAULINE  PROJECT FARCRY (PROOF)  SR-LP
ASHWELL, PAULINE  UNWILLINGLY TO EARTH  P
ASIMOV, ISAAC  ADDING A DIMENSION  [C/RAN-P]
ASIMOV, ISAAC  ALTERNATE ASIMOVS, THE  H,P
ASIMOV, ISAAC  ASIMOV CHRONICLES 1, THE  C/P,P
ASIMOV, ISAAC  ASIMOV CHRONICLES 2, THE  P
ASIMOV, ISAAC  ASIMOV CHRONICLES 3, THE  P
ASIMOV, ISAAC  ASIMOV CHRONICLES 4, THE  C/P:2,P
ASIMOV, ISAAC  ASIMOV CHRONICLES 5, THE  C/P,P
ASIMOV, ISAAC  ASIMOV ON NUMBERS  [C/RAN-H]
ASIMOV, ISAAC  ASIMOV SELECCION 1  FP

The MITSFS Pinkdex  September 24, 2007  Page 19
AUTHOR  TITLE  CODE
ASIMOV, ISAAC  ASIMOV SELECCION 3  FP
ASIMOV, ISAAC  ASIMOV’S CHRONOLOGY OF SCIENCE AND DISCOVERY  R/REF-H
ASIMOV, ISAAC  ASIMOV’S MYSTERIES  H:2,P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

ATLANTIS

PA,C/PA

ASIMOV, ISAAC  AZAZEL  C/P,P,H:2
ASIMOV, ISAAC  BANQUETS OF THE BLACK WIDOWERS  C/P:2,P
ASIMOV, ISAAC  BEFORE THE GOLDEN AGE  SFWA,HA
ASIMOV, ISAAC  BEFORE THE GOLDEN AGE 1  PA,C/PA
ASIMOV, ISAAC  BEFORE THE GOLDEN AGE 2  PA,C/PA
ASIMOV, ISAAC  BEFORE THE GOLDEN AGE 3  PA,C/PA
ASIMOV, ISAAC  BEGINNING AND THE END, THE  SFWA,R/REF-P
ASIMOV, ISAAC  BEST MYSTERIES OF ISAAC ASIMOV, THE  C/P:2,SFWA,P
ASIMOV, ISAAC  BEST OF ISAAC ASIMOV, THE  C/P,P,H
ASIMOV, ISAAC  BEST SCIENCE FICTION OF ISAAC ASIMOV, THE  SFWA
ASIMOV, ISAAC  BICENTENNIAL MAN AND OTHER STORIES, THE  H:2,P,SFWA
ASIMOV, ISAAC  BLOTE ZON, DE  FP
ASIMOV, ISAAC  BUY JUPITER AND OTHER STORIES  H:2,P,SFWA
ASIMOV, ISAAC  CAILLOUX DANS LE CIEL  FP
ASIMOV, ISAAC  CASEBOOK OF THE BLACK WIDOWERS  C/P:2,P
ASIMOV, ISAAC  CAVES OF STEEL, THE [LIJE BALEY 1]  C/P:2,P,H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

COMETS

PA

ASIMOV, ISAAC  COMPLETE ADVENTURES OF LUCKY STARR, THE [LUCKY STARR 1,2,3,4,5,6] H
ASIMOV, ISAAC  COMPLETE ROBOT, THE  H,SFWA
ASIMOV, ISAAC  COMPLETE STORIES, THE, VOLUME 1  LP
ASIMOV, ISAAC  COMPLETE STORIES, THE, VOLUME 2  H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

COSMIC KNIGHTS

C/PA,PA

ASIMOV, ISAAC  COUNTING THE EONS  SFWA
ASIMOV, ISAAC  CURRENTS OF SPACE, THE [GALACTIC EMPIRE 2]  C/P:2,P,H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

CURSES

PA

ASIMOV, ISAAC  DAVID STARR, SPACE RANGER [LUCKY STARR 1]  C/P:2,P

ASIMOV, ISAAC DAVID STARR, SPACE RANGER [LUCKY STARR 1,2]/

LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 1,2]

P

ASIMOV, ISAAC  DEATH DEALERS, THE  P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

DEVILS

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

DRAGON TALES

PA

ASIMOV, ISAAC  EARLY ASIMOV, THE  H:2,SFWA
ASIMOV, ISAAC  EARLY ASIMOV 1, THE  C/P:2,P
ASIMOV, ISAAC  EARLY ASIMOV 2, THE  C/P:2,P
ASIMOV, ISAAC  EARTH IS ROOM ENOUGH  C/P,P,H
ASIMOV, ISAAC  EARTH: OUR CROWDED SPACESHIP  [C/RAN-P]
ASIMOV, ISAAC  EDGE OF TOMORROW, THE  H:2,SFWA-P,P
ASIMOV, ISAAC  EIGHT STORIES FROM THE REST OF THE ROBOTS  C/P:3,P
ASIMOV, ISAAC/GREENBERG, MARTIN H.  ELECTION DAY 2084  HA
ASIMOV, ISAAC  END OF ETERNITY, THE  C/P:2,P
ASIMOV, ISAAC  ES, ROBOTS (LATVIAN)  FH
ASIMOV, ISAAC  FACT AND FANCY  R/REF-H,C/REF-P:2
ASIMOV, ISAAC  FACT AND FANCY (ARABIC)  FP

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

FAERIES

PA

ASIMOV, ISAAC  FANTASTIC VOYAGE  H:3,P
ASIMOV, ISAAC  FANTASTIC VOYAGE II: DESTINATION BRAIN  C/P,P,H,SFWA
ASIMOV, ISAAC  FAR AS HUMAN EYE COULD SEE  SFWA
ASIMOV, ISAAC  FAR ENDS OF TIME AND EARTH, THE  H
ASIMOV, ISAAC/CONKLIN, GROFF  FIFTY SHORT SCIENCE FICTION TALES  PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

FLYING SAUCERS

C/PA,PA

ASIMOV, ISAAC  FONDAZIONE E TERRA (ITALIAN) [FOUNDATION 5]  FH
ASIMOV, ISAAC  FORWARD THE FOUNDATION [FOUNDATION]  C/P,P,H:2
ASIMOV, ISAAC  FOUNDATION [FOUNDATION 1]  C/P:2,P
ASIMOV, ISAAC  FOUNDATION (REVISED) [FOUNDATION 1]  P
ASIMOV, ISAAC  FOUNDATION AND EARTH [FOUNDATION 5]  H:3,SFWA,P
ASIMOV, ISAAC  FOUNDATION AND EMPIRE [FOUNDATION 2]  C/P:2,P,SR-H,H
ASIMOV, ISAAC  FOUNDATION TRILOGY, THE [FOUNDATION 1,2,3]  H:2,LP
ASIMOV, ISAAC  FOUNDATION’S EDGE [FOUNDATION 4]  C/P,P,H:2,SFWA
ASIMOV, ISAAC  FROM EARTH TO HEAVEN  [C/RAN-P:2]
ASIMOV, ISAAC  FUNDACION E IMPERIO  FP

ASIMOV, ISAAC/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

FUTURE I, THE C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 20
AUTHOR  TITLE  CODE

ASIMOV, ISAAC/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

FUTURE IN QUESTION, THE

PA

ASIMOV, ISAAC  GELIEBTER ROBOTER  FP

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

GHOSTS

C/PA,PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

GIANTS

C/PA,PA

ASIMOV, ISAAC  GODS THEMSELVES, THE  C/P:2,H:2,P,SFWA
ASIMOV, ISAAC  GOLD: THE FINAL SCIENCE FICTION COLLECTION  C/P,P,H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

GREAT SCIENCE FICTION STORIES BY THE WORLD’S GREATEST SCIENTISTS

C/PA,PA

ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 1, THE (1939)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 2, THE (1940)  PA,C/PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 3, THE (1941)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 4, THE (1942)  PA,C/PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 5, THE (1943)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 6, THE (1944)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 7, THE (1945)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 8, THE (1946)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 9, THE (1947)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 10, THE (1948)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 11, THE (1949)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 12, THE (1950)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 13, THE (1951)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 14, THE (1952)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 15, THE (1953)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 16, THE (1954)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 17, THE (1955)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 18, THE (1956)  C/PA,PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 19, THE (1957)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 20, THE (1958)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 21, THE (1959)  PA
ASIMOV, ISAAC/GREENBERG, MARTIN H.  GREAT SF STORIES 23, THE (1961)  PA

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

GREAT TALES OF CLASSIC SCIENCE FICTION

HA

ASIMOV, ISAAC  HAVE YOU SEEN THESE?  SR-H
ASIMOV, ISAAC  HEAVENLY HOST, THE  C/P,P
ASIMOV, ISAAC  HOMBRE DEL BICENTENARIO, EL  FP
ASIMOV, ISAAC  HUGO WINNERS 1, THE  PA,C/PA
ASIMOV, ISAAC  HUGO WINNERS 1 AND 2, THE  HA
ASIMOV, ISAAC  HUGO WINNERS 2, THE  HA,SFWA
ASIMOV, ISAAC  HUGO WINNERS 3, THE  SFWA,HA
ASIMOV, ISAAC  HUGO WINNERS 3 BOOK 1, THE  C/PA,PA
ASIMOV, ISAAC  HUGO WINNERS 3 BOOK 2, THE  C/PA,PA
ASIMOV, ISAAC  HUGO WINNERS 4, THE  HA,SFWA
ASIMOV, ISAAC  HUGO WINNERS 5, THE  HA
ASIMOV, ISAAC  HUMAN BODY, THE  [C/RAN-P]
ASIMOV, ISAAC  I. ASIMOV  R/REF-H
ASIMOV, ISAAC  I, ROBOT  C/P,SR-P,P,H:3
ASIMOV, ISAAC  IK, ROBOT  FP
ASIMOV, ISAAC  IN JOY STILL FELT (AUTOBIOGRAPHY 1954–1978)  R/REF-H,C/REF-H
ASIMOV, ISAAC  IN MEMORY YET GREEN (AUTOBIOGRAPHY 1920–1954)  R/REF-H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

INTERGALACTIC EMPIRES

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

INVASIONS

PA

ASIMOV, ISAAC  IS ANYONE THERE?  [C/RAN-P:2]

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

ISAAC ASIMOV PRESENTS:

THE BEST SCIENCE FICTION OF THE 19TH CENTURY

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H.  ISAAC ASIMOV PRESENTS THE GOLDEN YEARS OF SCIENCE FICTION  HA:3
ASIMOV, ISAAC  ISAAC ASIMOV’S SCIENCE FICTION SERIE 7 (NORWEGIAN)  FP
ASIMOV, ISAAC  KONETS VECHNOSTI  FH

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

LAST MAN ON EARTH, THE

C/PA,PA

ASIMOV, ISAAC  LECHEROUS LIMERICKS  [C/RAN-P]
ASIMOV, ISAAC  LEFT HAND OF THE ELECTRON, THE  [C/RAN-P]
ASIMOV, ISAAC  LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 4]  C/P:2,P
ASIMOV, ISAAC  LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5]  C/P:2,P

ASIMOV, ISAAC LUCKY STARR AND THE MOONS OF JUPITER [LUCKY STARR 5,6]/ LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 5,6]

P

ASIMOV, ISAAC  LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3]  C/P,P,H

ASIMOV, ISAAC LUCKY STARR AND THE OCEANS OF VENUS [LUCKY STARR 3,4]/

LUCKY STARR AND THE BIG SUN OF MERCURY [LUCKY STARR 3,4]

P

ASIMOV, ISAAC  LUCKY STARR AND THE PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 21
AUTHOR  TITLE  CODE
ASIMOV, ISAAC  LUCKY STARR AND THE RINGS OF SATURN [LUCKY STARR 6]  C/P:2,P
ASIMOV, ISAAC  MAGIC: THE FINAL FANTASY COLLECTION  P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

MAGICAL WISHES

C/PA,PA

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

MAMMOTH BOOK OF GOLDEN AGE SCIENCE FICTION, THE

LPA

ASIMOV, ISAAC  MAN WHO UPSET THE UNIVERSE, THE  C/P:2,P
ASIMOV, ISAAC  MARTIAN WAY, THE  SR-H
ASIMOV, ISAAC  MARTIAN WAY AND OTHER STORIES, THE  H:2,P
ASIMOV, ISAAC  MARTIAN WAY, THE (AUTOGRAPHED)  SR-P
ASIMOV, ISAAC  MEASURE OF THE UNIVERSE, THE  R/REF-H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

MICROCOSMIC TALES

HA

ASIMOV, ISAAC  MODERNE NATUURWETENSCHAPPEN 1, DE  FP
ASIMOV, ISAAC  MODERNE NATUURWETENSCHAPPEN 2, DE  FP

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

MONSTERS

C/PA,PA

ASIMOV, ISAAC  MORE STORIES FROM THE HUGO WINNERS II  PA,C/PA
ASIMOV, ISAAC  MORE TALES OF THE BLACK WIDOWERS  C/P:2,P
ASIMOV, ISAAC  MURDER AT THE ABA  H:2,P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

MUTANTS

PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

MYTHICAL BEASTIES

C/PA,PA

ASIMOV, ISAAC  NAKED SUN, THE [LIJE BALEY 2]  C/P:3,P,H
ASIMOV, ISAAC  NEBULA AWARD STORIES 8  C/PA,PA
ASIMOV, ISAAC  NEMESIS  C/P,P,H
ASIMOV, ISAAC  NEW HUGO WINNERS, THE  HA,PA
ASIMOV, ISAAC  NEW HUGO WINNERS 2, THE  HA,PA,SFWA-P
ASIMOV, ISAAC/SILVERBERG, ROBERT  NIGHTFALL  C/P:2,P,H
ASIMOV, ISAAC  NIGHTFALL AND OTHER STORIES  H:2,P
ASIMOV, ISAAC  NINE TOMORROWS  H:2,P
ASIMOV, ISAAC  OF TIME AND SPACE AND OTHER THINGS  [C/RAN-P:2]

ASIMOV, ISAAC/CARR, TERRY/ GREENBERG, MARTIN H.

100 GREAT FANTASY SHORT SHORT STORIES

PA,SFWA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

100 GREAT SCIENCE FICTION SHORT SHORT STORIES

PA,SFWA

ASIMOV, ISAAC  OPUS 100  [C/RAN-H,C/RAN-P:2]
ASIMOV, ISAAC  OPUS 200  [C/RAN-P:2]
ASIMOV, ISAAC  OPUS 300  [C/RAN-H]
ASIMOV, ISAAC  OPUS 300 (PROOF)  SR-LP
ASIMOV, ISAAC  OUT OF THE EVERYWHERE  R/REF-H
ASIMOV, ISAAC  PEBBLE IN THE SKY [GALACTIC EMPIRE 3]  C/P,P,H,GN3.14
ASIMOV, ISAAC  PIRATES OF THE ASTEROIDS [LUCKY STARR 2]  C/P,P
ASIMOV, ISAAC  PLANET THAT WASN’T, THE  SFWA
ASIMOV, ISAAC/SILVERBERG, ROBERT  POSITRONIC MAN, THE  C/P,P,H:2
ASIMOV, ISAAC  PRELUDE TO FOUNDATION [FOUNDATION 0]  C/P,P,H
ASIMOV, ISAAC  QUICK AND EASY MATH  R/REF-H
ASIMOV, ISAAC  REBELLIOUS STARS, THE  DD084
ASIMOV, ISAAC  REST OF THE ROBOTS, THE  H:2
ASIMOV, ISAAC  ROBOT DREAMS  C/P,SFWA-P:2,P
ASIMOV, ISAAC  ROBOT NOVELS, THE  H:3
ASIMOV, ISAAC  ROBOT VISIONS  C/P,SFWA-P:2,P,H
ASIMOV, ISAAC  ROBOT VISIONS (PROOF)  SR-LP
ASIMOV, ISAAC  ROBOTS AND EMPIRE  C/P:2,P,H,SFWA
ASIMOV, ISAAC  ROBOTS, LES  FP
ASIMOV, ISAAC  ROBOTS OF DAWN, THE [LIJE BALEY 3]  C/P:2,P,H

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SCIENCE FICTION

HA

ASIMOV, ISAAC  SCIENCE FICTION BY ASIMOV  SR-P
ASIMOV, ISAAC  SCIENCE FICTION MASTERPIECES  HA

ASIMOV, ISAAC/MARTIN, GEORGE R. R./ GREENBERG, MARTIN H.

SCIENCE FICTION WEIGHT LOSS BOOK, THE

HA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SCIENCE FICTIONAL OLYMPICS, THE

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SCIENCE FICTIONAL SOLAR SYSTEM, THE

PA,SFWA

ASIMOV, ISAAC  SCIENCE, NUMBERS, AND I  [C/RAN-P:2]
ASIMOV, ISAAC  SCIENCE PAST — SCIENCE FUTURE  SFWA
ASIMOV, ISAAC  SECOND FOUNDATION [FOUNDATION 3]  C/P:2,P,H:3,SR-P,SR-H
ASIMOV, ISAAC  SECONDE FONDATION [FOUNDATION 3]  FP

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

7 CARDINAL VIRTUES OF SCIENCE FICTION, THE

C/PA,PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

7 DEADLY SINS AND CARDINAL VIRTUES OF SCIENCE FICTION, THE HA

The MITSFS Pinkdex  September 24, 2007  Page 22
AUTHOR  TITLE  CODE

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

7 DEADLY SINS OF SCIENCE FICTION, THE

HA,PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SHERLOCK HOLMES THROUGH TIME AND SPACE

LPA

ASIMOV, ISAAC  SOLAR SYSTEM AND BACK, THE  [C/RAN-P]
ASIMOV, ISAAC  SPACE DICTIONARY  [C/RAN-P]

ASIMOV, ISAAC/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

SPACE MAIL

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SPACE MAIL II

C/PA,PA

ASIMOV, ISAAC  SPACE RANGER [LUCKY STARR 1]  P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SPACE SHUTTLES

C/PA,PA

ASIMOV, ISAAC/LAURANCE, ALICE  SPECULATIONS  SR-HA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SPELLS

C/PA,PA

ASIMOV, ISAAC  STARS IN THEIR COURSES, THE  [C/RAN-P]
ASIMOV, ISAAC  STARS, LIKE DUST, THE [GALACTIC EMPIRE 1]  C/P:2,P,H

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

STARSHIPS

C/PA,PA

ASIMOV, ISAAC  STORIES FROM THE HUGO WINNERS II  C/PA,PA
ASIMOV, ISAAC  SUN SHINES BRIGHT, THE  [C/RAN-P]
ASIMOV, ISAAC  SUPER HUGOS, THE  PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

SUPERMEN

PA,C/PA

ASIMOV, ISAAC  TALES OF THE BLACK WIDOWERS  C/P:2,P
ASIMOV, ISAAC  TAUSENDJAHRES PLAN, DER  FP

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

THINKING MACHINES

HA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

13 CRIMES OF SCIENCE FICTION, THE

HA

ASIMOV, ISAAC  THOUSAND-YEAR PLAN, THE  C/P,DD110,P
ASIMOV, ISAAC  THROUGH A GLASS, CLEARLY  SR-P,C/P

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

TIN STARS

C/PA,PA

ASIMOV, ISAAC  TODAY AND TOMORROW AND ...  SFWA
ASIMOV, ISAAC  TOMORROW’S CHILDREN  HA,PA
ASIMOV, ISAAC  TRAGEDY OF THE MOON, THE  SFWA
ASIMOV, ISAAC  TRIANGLE [GALACTIC EMPIRE 1,2,3]  H

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

TV: 2000

PA

ASIMOV, ISAAC/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

TWELVE FRIGHTS OF CHRISTMAS, THE

C/PA,PA

ASIMOV, ISAAC  TWENTIETH CENTURY DISCOVERY  [C/RAN-P:2]
ASIMOV, ISAAC  UGLY LITTLE BOY, THE  SFWA-TD9,TD9
ASIMOV, ISAAC/SILVERBERG, ROBERT  UGLY LITTLE BOY, THE  C/P,P,H
ASIMOV, ISAAC  UNION CLUB MYSTERIES, THE  C/P:2,P
ASIMOV, ISAAC  UNIVERSE, THE  [C/RAN-P]
ASIMOV, ISAAC  VIEW FROM A HEIGHT  [C/RAN-P:2]
ASIMOV, ISAAC/GREENBERG, MARTIN H.  VISIONS OF FANTASY  HA

ASIMOV, ISAAC/WARRICK, PATRICIA S./ GREENBERG, MARTIN H.

WAR WITH THE ROBOTS

HA

ASIMOV, ISAAC  WELLSPRINGS OF LIFE, THE  [C/RAN-P]
ASIMOV, ISAAC  WHERE DO WE GO FROM HERE?  C/PA,PA,SFWA
ASIMOV, ISAAC  WHIFF OF DEATH, A  C/P:3,P
ASIMOV, ISAAC  WINDS OF CHANGE AND OTHER STORIES, THE  C/P,P,H:2

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

WITCHES

PA,C/PA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

WITCHES & WIZARDS

HA

ASIMOV, ISAAC/GREENBERG, MARTIN H./ WAUGH, CHARLES G.

WIZARDS

C/PA,PA

ASIMOV, ISAAC  WORDS OF SCIENCE  [C/RAN-P]
ASIMOV, JANET  MIND TRANSFER  C/P,SFWA-P,P,H
ASIMOV, JANET  MURDER AT THE GALACTIC WRITERS’ SOCIETY  C/P,P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY AND THE COURT JESTER [NORBY]  P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY AND THE OLDEST DRAGON [NORBY]  SFWA-P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY AND YOBO’S GREAT ADVENTURE [NORBY 4]  C/P,SFWA-P:2,P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY CHRONICLES, THE [NORBY 1]  C/P:2,P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY DOWN TO EARTH [NORBY 5]  P,SFWA-P:2
ASIMOV, JANET/ASIMOV, ISAAC  NORBY: ROBOT FOR HIRE [NORBY 2]  C/P:3,P
ASIMOV, JANET/ASIMOV, ISAAC  NORBY THROUGH TIME AND SPACE [NORBY]  C/P:2,P
ASIMOV, JANET  PSHRINKS ANONYMOUS  C/P,P
ASIRE, NANCY  TEARS OF TIME  P,SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 23
AUTHOR  TITLE  CODE
ASIRE, NANCY  TO FALL LIKE STARS  P
ASIRE, NANCY  TWILIGHT’S KINGDOMS  P,SFWA-P
ASKEGREN, PIERCE  AFTERIMAGE [@BUFFY]  P

ASKEGREN, PIERCE AVENGERS AND THE THUNDERBOLTS, THE

[AVENGERS (MARVEL COMICS CHARACTERS)]

C/P,P

ASKEGREN, PIERCE  EXIT STRATEGY [INCONSTANT MOON 3]  P
ASKEGREN, PIERCE  FALL GIRL [INCONSTANT MOON 2]  C/P,P
ASKEGREN, PIERCE  GATEWAY TO THE STARS [@TRAVELLER 1]  P
ASKEGREN, PIERCE  HUMAN RESOURCE [INCONSTANT MOON 1]  P
ASKEGREN, PIERCE/FINGEROTH, DANNY  SABOTAGE [@DOOM’S DAY #2]  C/P,P
ASNIN, SCOTT  COLD WIND FROM ORION, A  C/P,P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  AFTERMATH [@THIEVES’ WORLD #10]  PA,C/PA
ASPRIN, ROBERT LYNN  ANOTHER FINE MYTH... [MYTH ADVENTURES 1]  C/P:2,LP
ASPRIN, ROBERT LYNN/ABBEY, LYNN  BLOOD TIES [@THIEVES’ WORLD #9]  PA,C/PA
ASPRIN, ROBERT LYNN  BUG WARS, THE  H,SFWA-P:2,P
ASPRIN, ROBERT LYNN/NYE, JODY LYNN  CLASS DIS-MYTHED [MYTH ADVENTURES]  H
ASPRIN, ROBERT LYNN  COLD CASH WAR, THE [COLD CASH WAR 1]  C/P,SFWA-P:2,P,H
ASPRIN, ROBERT LYNN/FAWCETT, BILL  COLD CASH WARRIOR [COLD CASH WAR 2]  P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  CROSS-CURRENTS  HA
ASPRIN, ROBERT LYNN/ABBEY, LYNN  DEAD OF WINTER, THE [@THIEVES’ WORLD #7]  PA,C/PA

ASPRIN, ROBERT LYNN/ FRIESNER, ESTHER M.

E.GODZ

H:2

ASPRIN, ROBERT LYNN/ABBEY, LYNN  FACE OF CHAOS, THE [@THIEVES’ WORLD #5]  PA,C/PA
ASPRIN, ROBERT LYNN/EVANS, LINDA  FOR KING & COUNTRY [TIME SCOUT]  H,P
ASPRIN, ROBERT LYNN  HIT OR MYTH [MYTH ADVENTURES 4]  C/P:2,LP
ASPRIN, ROBERT LYNN/EVANS, LINDA  HOUSE THAT JACK BUILT, THE [TIME SCOUT]  P
ASPRIN, ROBERT LYNN/NYE, JODY LYNN  LICENSE INVOKED  C/P,P
ASPRIN, ROBERT LYNN  LITTLE MYTH MARKER [MYTH ADVENTURES 6]  C/P,SFWA-P,P,H
ASPRIN, ROBERT LYNN/TAKEI, GEORGE  MIRROR FRIEND, MIRROR FOE  C/P:2,P
ASPRIN, ROBERT LYNN  MYTH ADVENTURES [MYTH ADVENTURES 1,2,3,4]  H:2
ASPRIN, ROBERT LYNN  MYTH ALLIANCES (OMNIBUS EDITION) [MYTH ADVENTURES 5,6,7]  H:2
ASPRIN, ROBERT LYNN/NYE, JODY LYNN  MYTH ALLIANCES (SINGLE NOVEL) [MYTH ADVENTURES]  H,P
ASPRIN, ROBERT LYNN  MYTH CONCEPTIONS [MYTH ADVENTURES 2]  C/P,P,H
ASPRIN, ROBERT LYNN  MYTH DIRECTIONS [MYTH ADVENTURES 3]  C/P,LP,H

ASPRIN, ROBERT LYNN MYTH DIRECTIONS [MYTH ADVENTURES 3,4]/ HIT OR MYTH [MYTH ADVENTURES 3,4]

P

ASPRIN, ROBERT LYNN/NYE, JODY LYNN  MYTH-GOTTEN GAINS [MYTH ADVENTURES]  LP
ASPRIN, ROBERT LYNN  M.Y.T.H. INC. IN ACTION [MYTH INC 3]  C/P:2,SFWA-P:2,P,H
ASPRIN, ROBERT LYNN  M.Y.T.H. INC. LINK [MYTH ADVENTURES 7/MYTH INC 1]  C/P,SFWA-P,H,LP
ASPRIN, ROBERT LYNN  MYTH-ING PERSONS [MYTH ADVENTURES 5]  C/P,P,H
ASPRIN, ROBERT LYNN  MYTH-ION IMPROBABLE [MYTH ADVENTURES]  C/P,P
ASPRIN, ROBERT LYNN  MYTH-NOMERS AND IM-PERVECTIONS [MYTH INC 2]  C/P:2,SFWA-P,P,H
ASPRIN, ROBERT LYNN/NYE, JODY LYNN  MYTH-TAKEN IDENTITY [MYTH ADVENTURES]  LP
ASPRIN, ROBERT LYNN/NYE, JODY LYNN  MYTH-TOLD TALES [MYTH ADVENTURES]  P
ASPRIN, ROBERT LYNN/HECK, PETER J.  NO PHULE LIKE AN OLD PHULE [PHULE 5]  P
ASPRIN, ROBERT LYNN/HECK, PETER J.  PHULE AND HIS MONEY, A [PHULE 3]  C/P,P
ASPRIN, ROBERT LYNN/HECK, PETER J.  PHULE ME TWICE [PHULE 4]  C/P:2,P
ASPRIN, ROBERT LYNN  PHULE’S COMPANY [PHULE 1]  C/P:2,SFWA-P:2,P
ASPRIN, ROBERT LYNN/HECK, PETER J.  PHULE’S ERRAND [PHULE 6]  C/P,P
ASPRIN, ROBERT LYNN  PHULE’S PARADISE [PHULE 2]  C/P:2,SFWA-P:2,P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  PRICE OF VICTORY, THE [THIEVES’ WORLD 10,11,12]  H:2
ASPRIN, ROBERT LYNN/DEL CARLO, ERIC  RESURRECTION [@WARTORN 1]  P
ASPRIN, ROBERT LYNN/EVANS, LINDA  RIPPING TIME [TIME SCOUT]  P
ASPRIN, ROBERT LYNN  SANCTUARY  HA
ASPRIN, ROBERT LYNN/ABBEY, LYNN  SHADOWS OF SANCTUARY [@THIEVES’ WORLD #3]  PA,C/PA
ASPRIN, ROBERT LYNN  SHADOWS OF SANCTUARY (BRITISH) [THIEVES’ WORLD 3]  PA
ASPRIN, ROBERT LYNN/ABBEY, LYNN  SHATTERED SPHERE, THE  HA
ASPRIN, ROBERT LYNN  SOMETHING M.Y.T.H. INC. [MYTH ADVENTURES]  C/P,P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  SOUL OF THE CITY [@THIEVES’ WORLD #8]  PA,C/PA
ASPRIN, ROBERT LYNN/ABBEY, LYNN  STEALER’S SKY [@THIEVES’ WORLD #12]  PA,C/PA,SFWA-P
ASPRIN, ROBERT LYNN  STORM SEASON [@THIEVES’ WORLD #4]  PA,C/PA
ASPRIN, ROBERT LYNN  SWEET MYTH-TERY OF LIFE [MYTH ADVENTURES]  C/P:2,P,H
ASPRIN, ROBERT LYNN/ABBEY, LYNN  TALES FROM THE VULGAR UNICORN [@THIEVES’ WORLD #2]  PA,C/PA
ASPRIN, ROBERT LYNN  TAMBU  H:2,LP
ASPRIN, ROBERT LYNN/ABBEY, LYNN  THIEVES’ WORLD [@THIEVES’ WORLD #1]  PA,C/PA
ASPRIN, ROBERT LYNN  THIEVES’ WORLD (BRITISH) [THIEVES’ WORLD 1]  PA
ASPRIN, ROBERT LYNN/EVANS, LINDA  TIME SCOUT [TIME SCOUT]  C/P:2,P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  UNEASY ALLIANCES [@THIEVES’ WORLD #11]  PA,C/PA
ASPRIN, ROBERT LYNN/EVANS, LINDA  WAGERS OF SIN [TIME SCOUT]  P
ASPRIN, ROBERT LYNN/ABBEY, LYNN  WINGS OF OMEN [@THIEVES’ WORLD #6]  PA,C/PA
ASQUITH, CYNTHIA  GHOST BOOK, THE  PA
ASQUITH, CYNTHIA  SECOND GHOST BOOK, THE  SR-PA
ASQUITH, CYNTHIA  THIRD GHOST BOOK, THE  PA

AUTHOR  TITLE  CODE
ASQUITH, CYNTHIA  WHEN CHURCHYARDS YAWN  SR-PA
ATACK, CHRIS  PROJECT MALDON  P
ATHANS, PHILIP  ANNIHILATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #5]  P
ATHANS, PHILIP  BALDUR’S GATE [@FORGOTTEN REALMS BALDUR’S GATE I]  P

ATHANS, PHILIP HALLS OF STORMWEATHER, THE

[@FORGOTTEN REALMS SEMBIA GATEWAY TO THE REALMS #1]

PA

ATHANS, PHILIP  REALMS OF MYSTERY [@FORGOTTEN REALMS]  PA
ATHANS, PHILIP  REALMS OF THE DEEP [@FORGOTTEN REALMS THREAT FROM THE SEA]  PA

ATHANS, PHILIP REALMS OF THE DRAGONS

[@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS]

PA

ATHANS, PHILIP REALMS OF THE DRAGONS II

[@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS]

PA

ATHANS, PHILIP  REALMS OF THE ELVES [@FORGOTTEN REALMS LAST MYTHAL]  PA
ATHANS, PHILIP  SCREAM OF STONE [@FORGOTTEN REALMS WATERCOURSE #3]  P
ATHANS, PHILIP  SHADOWS OF AMN [@FORGOTTEN REALMS BALDUR’S GATE II]  P
ATHANS, PHILIP  WHISPER OF WAVES [@FORGOTTEN REALMS WATERCOURSE #1]  P
ATHELING, WILLIAM JR.  ISSUE AT HAND, THE  R/REF-H,C/REF-H
ATHELING, WILLIAM JR.  MORE ISSUES AT HAND  R/REF-H
ATHERTON, GERTRUDE  BLACK OXEN  SR-H
ATKEY, BERTRAM  ESCAPES OF MR. HONEY, THE  H
ATLANTIC RICHFIELD CO.  TRICENTENNIAL REPORT: LETTERS FROM AMERICA, THE  R/REF-H
ATTANASIO, A. A.  ARC OF THE DREAM [RADIX 3]  C/P,P
ATTANASIO, A. A.  BEASTMARKS  H
ATTANASIO, A. A.  DRAGON AND THE UNICORN, THE  C/P:2,LP
ATTANASIO, A. A.  EAGLE AND THE SWORD, THE  LP
ATTANASIO, A. A.  HELLBOUND [@CROW]  H,LP
ATTANASIO, A. A.  HUNTING THE GHOST DANCER  H,P
ATTANASIO, A. A.  IN OTHER WORLDS [RADIX 2]  C/P,P
ATTANASIO, A. A.  LAST LEGENDS OF EARTH, THE [RADIX 4]  C/P:2,SFWA,LP
ATTANASIO, A. A.  MOON’S WIFE, THE  H
ATTANASIO, A. A.  RADIX [RADIX 1]  C/P:2,P
ATTANASIO, A. A.  SERPENT AND THE GRAIL, THE  P
ATTANASIO, A. A.  SOLIS  H,P
ATTANASIO, A. A.  TWICE DEAD THINGS  LP
ATTANASIO, A. A.  WOLF AND THE CROWN, THE  C/P,LP
ATTANASIO, A. A.  WYVERN  H,P
ATTIAS, DIANA/SMITH, LINDSAY  EMPIRE STRIKES BACK NOTEBOOK, THE  R/REF-H
ATTWOOD, TONY  TERRY NATION’S BLAKE’S 7 THE PROGRAMME GUIDE [BLAKE’S 7]  R/REF-P
ATTWOOD, TONY  TURLOUGH AND THE EARTHLINK DILEMMA [DOCTOR WHO]  SR-P
ATWOOD, MARGARET  CAT’S EYE  LP
ATWOOD, MARGARET  HANDMAID’S TALE, THE  C/P,P
ATWOOD, MARGARET  ORYX AND CRAKE  H
AUBIN, ETIENNE  DRACULA AND THE VIRGINS OF THE UNDEAD  SR-P
AUDEN, RENEE  HIGH THRUST  SR-L
AUDLEY, ANSELM  HERESY [AQUASILVA 1]  H
AUDLEY, ANSELM  INQUISITION [@AQUASILVA #2]  P
AUEL, JEAN M.  CLAN OF THE CAVE BEAR, THE [EARTH’S CHILDREN 1]  C/P:2,P
AUEL, JEAN M.  MAMMOTH HUNTERS, THE [EARTH’S CHILDREN 3]  C/P,P,H
AUEL, JEAN M.  PLAINS OF PASSAGE, THE [EARTH’S CHILDREN 4]  P
AUEL, JEAN M.  VALLEY OF HORSES, THE [EARTH’S CHILDREN 2]  C/P,P,H
AUGUSTUS, ALBERT JR.  SLAVES OF LOMOORO, THE  SR-P
AUSTIN, ALICIA  ALICIA AUSTIN’S AGE OF DREAMS  A
AUSTIN, J. D.  BOBBY’S GIRL  P
AUSTIN, J. D.  SECOND CONTACT  P
AUSTIN, R. G.  VAMPIRES, SPIES, AND ALIEN BEINGS  P
AUSTIN, RICHARD  BRUTE FORCE [GUARDIANS 7]  P
AUSTIN, RICHARD  DESOLATION ROAD [GUARDIANS 8]  P
AUSTIN, RICHARD  GUARDIANS, THE [GUARDIANS 1]  P
AUSTIN, RICHARD  NIGHT OF THE PHOENIX [GUARDIANS 4]  P
AUSTIN, RICHARD  PLAGUE YEARS [GUARDIANS 12]  P
AUSTIN, RICHARD  THUNDER OF HELL [GUARDIANS 3]  P
AUSTIN, RICHARD  TRIAL BY FIRE [GUARDIANS 2]  P
AUSTIN, RICHARD  WAR ZONE [GUARDIANS 6]  P
AVALLONE, MICHAEL  BENEATH THE PLANET OF THE APES  C/P:2,P
AVALLONE, MICHAEL  BIRDS-OF-A-FEATHER AFFAIR, THE [@GIRL FROM UNCLE #1]  P
AVALLONE, MICHAEL  BLAZING AFFAIR, THE [@GIRL FROM UNCLE #2]  P
AVALLONE, MICHAEL  LIVING BOMB, THE  P
AVALLONE, MICHAEL  MAN FROM U.N.C.L.E., THE [@MAN FROM UNCLE #1]  C/P:2,P
AVALLONE, MICHAEL  TALES OF THE FRIGHTENED  C/P,P
AVALLONE, MICHAEL  WEREWOLF WALKS TONIGHT, THE [@SATAN SLEUTH #2]  P
AVERY, FIONA  CROWN ROSE, THE  H
AVERY, RICHARD  DEATHWORMS OF KRATOS, THE [EXPENDABLES 1]  C/P:2,P
AVERY, RICHARD  RINGS OF TANTALUS, THE [EXPENDABLES 2]  C/P:2,P
AVERY, RICHARD  VENOM OF ARGUS, THE [EXPENDABLES 4]  C/P:2,P

AUTHOR  TITLE  CODE
AVERY, RICHARD  WAR GAMES OF ZELOS, THE [EXPENDABLES 3]  C/P:2,P
AVON  AVON GHOST READER  SR-PA
AVON  GIRL WITH THE HUNGRY EYES, THE  SR-PA
AVON BOOKS  TIMELESS LOVE  SFWA-P
AVONOVA BOOKS  FUTURE QUARTET  HA,PA
AWLINSON, RICHARD  SHADOWDALE [FORGOTTEN REALMS AVATAR 1]  C/P:2,P
AWLINSON, RICHARD  TANTRAS [FORGOTTEN REALMS AVATAR 2]  C/P:2,P
AWLINSON, RICHARD  WATERDEEP [FORGOTTEN REALMS AVATAR 3]  C/P:2,P
AXLER, JAMES  AURORA QUEST [@EARTHBLOOD]  P
AXLER, JAMES  BITTER FRUIT [@DEATHLANDS]  P
AXLER, JAMES  BLACK HARVEST [@DEATHLANDS]  P
AXLER, JAMES  COLD ASYLUM [@DEATHLANDS]  P
AXLER, JAMES  CRATER LAKE [@DEATHLANDS]  P
AXLER, JAMES  CROSSWAYS [@DEATHLANDS]  P
AXLER, JAMES  CRUCIBLE OF TIME [@DEATHLANDS]  C/P,P
AXLER, JAMES  DARK EMBLEM [@DEATHLANDS]  P
AXLER, JAMES  DECTRA CHAIN [@DEATHLANDS]  P
AXLER, JAMES  DEEP TREK [@EARTHBLOOD]  P
AXLER, JAMES  DEMONS OF EDEN [@DEATHLANDS]  P
AXLER, JAMES  EARTHBLOOD [@EARTHBLOOD]  P
AXLER, JAMES  ECLIPSE AT NOON [@DEATHLANDS]  P
AXLER, JAMES  EMERALD FIRE [@DEATHLANDS]  P
AXLER, JAMES  ENCOUNTER [@DEATHLANDS]  P
AXLER, JAMES  EQUINOX ZERO [@OUTLANDERS]  P
AXLER, JAMES  EVIL ABYSS [@OUTLANDERS]  P
AXLER, JAMES  FREEDOM LOST [@DEATHLANDS]  P
AXLER, JAMES  GROUND ZERO [@DEATHLANDS]  P
AXLER, JAMES  HELL RISING [@OUTLANDERS]  P
AXLER, JAMES  HOMEWARD BOUND [@DEATHLANDS]  P
AXLER, JAMES  ICE AND FIRE [@DEATHLANDS]  P
AXLER, JAMES  ICEBLOOD [@OUTLANDERS]  P
AXLER, JAMES  KEEPERS OF THE SUN [@DEATHLANDS]  P
AXLER, JAMES  LATITUDE ZERO [@DEATHLANDS]  P
AXLER, JAMES  MARS ARENA, THE [@DEATHLANDS]  C/P,P
AXLER, JAMES  MOON FATE [@DEATHLANDS]  P
AXLER, JAMES  NEUTRON SOLSTICE [@DEATHLANDS]  P
AXLER, JAMES  NIGHTMARE PASSAGE [@DEATHLANDS]  P
AXLER, JAMES  NORTHSTAR RISING [DEATHLANDS]  P
AXLER, JAMES  PARALLAX RED [@OUTLANDERS]  P
AXLER, JAMES  PONY SOLDIERS [@DEATHLANDS]  P
AXLER, JAMES  RAT KING [@DEATHLANDS]  P
AXLER, JAMES  RED HOLOCAUST [@DEATHLANDS]  P
AXLER, JAMES  ROAD WARS [@DEATHLANDS]  P
AXLER, JAMES  SEA OF PLAGUE [@OUTLANDERS HEART OF THE WORLD 2]  P
AXLER, JAMES  SHADOW WORLD [@DEATHLANDS]  P
AXLER, JAMES  SHADOWFALL [@DEATHLANDS]  P
AXLER, JAMES  SHAKING EARTH [@DEATHLANDS]  P
AXLER, JAMES  SKYDARK SPAWN [@DEATHLANDS]  P
AXLER, JAMES  STONEFACE [@DEATHLANDS]  P
AXLER, JAMES  SUN LORD [@OUTLANDERS]  P
AXLER, JAMES  TALON AND FANG [OUTLANDERS HEART OF THE WORLD 1]  P
AXLER, JAMES  TIME NOMADS [@DEATHLANDS]  P
AXLER, JAMES  TRADER REDUX [@DEATHLANDS]  P
AXLER, JAMES  TWILIGHT CHILDREN [@DEATHLANDS]  P
AXLER, JAMES  WATERSLEEP [@DEATHLANDS]  P
AXLER, JAMES  WAY OF THE WOLF [@DEATHLANDS]  P
AXLER, JAMES  ZERO CITY [@DEATHLANDS]  P
AXTON, DAVID  PRISON OF ICE  P
AYCLIFFE, JONATHAN  LOST, THE  H,P
AYCOCK, DALE  STAR DRIFTER  P
AYCOCK, DALE  STARSPINNER  P
AYERS, JEFF  VOYAGES OF IMAGINATION [STAR TREK]  R/REF-H
AYLESWORTH, JOHN  FEE, FEI, FO, FUM  C/P:2,P
AYLETT, STEVE  ATOM  H
AYLETT, STEVE  INFLATABLE VOLUNTEER, THE  LP
AYLETT, STEVE  TOXICOLOGY  LP
AYLETT, STEVE  VELOCITY GOSPEL, THE [ACCOMPLICE 2]  LP
AYME, MARCEL  ACROSS PARIS AND OTHER STORIES  H
AYME, MARCEL  WALKER-THROUGH-WALLS, THE  P
AYRTON, MICHAEL  MAZE MAKER, THE  P

AUTHOR  TITLE  CODE
BABBITT, NATALIE  TUCK EVERLASTING  P
BACH, RICHARD  JONATHAN LIVINGSTON SEAGULL  P
BACH, RICHARD  ONE  H
BACHMAN, RICHARD  REGULATORS, THE  H
BACHMAN, RICHARD  RUNNING MAN, THE  SR-L
BACHMAN, RICHARD  THINNER  H
BACON, JONATHAN/TROYANOVICH, STEVE  OMNIUMGATHUM  SR-VLPA
BACON-SMITH, CAMILLE  DAEMONS, INC. [KEVIN BRADLEY AND EVAN DAVIS 1,2]  H
BACON-SMITH, CAMILLE  EYE OF THE DAEMON [KEVIN BRADLEY AND EVAN DAVIS 1]  C/P,P
BACON-SMITH, CAMILLE  EYES OF THE EMPRESS [KEVIN BRADLEY AND EVAN DAVIS 2]  P
BACON-SMITH, CAMILLE  FACE OF TIME, THE  P
BADGER BOOKS  MENACE FROM MERCURY  SR-PA
BADGER BOOKS  SPACE-BORNE  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #36  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #41  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #43  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #47  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #49  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #51  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #53  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #55  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #57  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #59  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #61  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #63  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #65  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #67  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #69  SR-PA,C/PA
BADGER BOOKS  SUPERNATURAL STORIES #71  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #73  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #75  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #77  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #79  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #81  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #83  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #85  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #87  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #89  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #91  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #93  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #95  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #97  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #99  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #103  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #105  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #107  SR-PA
BADGER BOOKS  SUPERNATURAL STORIES #109  SR-PA
BADGER BOOKS  TWILIGHT ZONE  SR-PA
BAEN BOOKS  FOREVER AFTER  C/PA:2,PA
BAEN BOOKS  GODS OF WAR, THE  PA,SFWA-P
BAEN, JAMES  BEST FROM GALAXY III, THE  C/PA:2,PA
BAEN, JAMES  BEST FROM GALAXY IV, THE  PA,C/PA
BAEN, JAMES  BEST FROM IF III, THE  C/PA,PA
BAEN, JAMES  BEST OF DESTINIES 1980, THE  PA
BAEN, JAMES  DESTINIES 1 #1  PA,C/PA
BAEN, JAMES  DESTINIES 1 #2  C/PA,PA
BAEN, JAMES  DESTINIES 1 #3  C/PA,PA
BAEN, JAMES  DESTINIES 1 #4  C/PA,PA
BAEN, JAMES  DESTINIES 1 #5  PA,C/PA
BAEN, JAMES  DESTINIES 2 #1  PA
BAEN, JAMES  DESTINIES 2 #2  PA
BAEN, JAMES  DESTINIES 2 #3  C/PA,PA
BAEN, JAMES  DESTINIES 2 #4  PA
BAEN, JAMES  DESTINIES 3 #1  C/PA,PA
BAEN, JAMES  DESTINIES 3 #2  PA,C/PA
BAEN, JAMES  GALAXY: THE BEST OF MY YEARS  C/PA,PA
BAEN, JAMES  NEW DESTINIES 1  PA,C/PA
BAEN, JAMES  NEW DESTINIES 2  PA,C/PA
BAEN, JAMES  NEW DESTINIES 3  C/PA,PA,SFWA-P
BAEN, JAMES  NEW DESTINIES 4  PA,C/PA,SFWA-P
BAEN, JAMES  NEW DESTINIES 6  PA,C/PA,SFWA-P
BAEN, JAMES  NEW DESTINIES 7  PA,C/PA,SFWA-P
BAEN, JAMES  NEW DESTINIES 8  C/PA,PA,SFWA-P:2
BAEN, JAMES  NEW DESTINIES 9  C/PA,PA
BAEN PUBLISHING ENTERPRISES  AT ALL COSTS DISK, THE (CD) [HONOR HARRINGTON]  SR-MM
BAEN PUBLISHING ENTERPRISES  DRAGONS DISK, THE (CD)  SR-MM
BAEN PUBLISHING ENTERPRISES  HONORVERSE DISK, THE (CD) [HONOR HARRINGTON]  SR-MM
BAEN PUBLISHING ENTERPRISES  POSLEEN DISK, THE (CD) [POSLEEN]  SR-MM

AUTHOR  TITLE  CODE
BAEN PUBLISHING ENTERPRISES  20 YEARS OF EXCELLENT ADVENTURES (CD)  SR-MM:2
BAEN PUBLISHING ENTERPRISES  WIND RIDER’S OATH DISK, THE (CD) [HRADANI]  SR-MM
BAHNSON, AGNEW H. JR.  STARS ARE TOO HIGH, THE  C/P:2,P
BAHRENBURG, BRUCE  CREATION OF DINO DE LAURENTIIS’ KING KONG, THE  R/REF-P,C/REF-P
BAILEY, DALE  FALLEN, THE  H
BAILEY, DALE  RESURRECTION MAN’S LEGACY, THE  H
BAILEY, DAVID  SHORT TRIPS: SEVEN DEADLY SINS [DOCTOR WHO]  HA
BAILEY, GERALD EARL  SWORD OF POYANA [THORGRIM 2]  P
BAILEY, GERALD EARL  SWORD OF THE NURLINGAS [THORGRIM 1]  P
BAILEY, HILARY  NEW WORLDS 9  SR-PA
BAILEY, HILARY  NEW WORLDS 10  SR-PA
BAILEY, J. O.  PILGRIMS THROUGH SPACE AND TIME  R/REF-H,C/REF-H
BAILEY, LYNNE  SPLENDID YOU  P
BAILEY, ROBIN WAYNE  ARCHITECTS OF DREAMS  HA,LPA
BAILEY, ROBIN WAYNE  BLOODSONGS [FROST 3]  C/P:2,SFWA-P,P
BAILEY, ROBIN WAYNE  BROTHERS OF THE DRAGON [BROTHERS OF THE DRAGON 1]  C/P:2,P
BAILEY, ROBIN WAYNE  DRAGONKIN  H
BAILEY, ROBIN WAYNE  ENCHANTER  C/P:2,P
BAILEY, ROBIN WAYNE  FLAMES OF THE DRAGON [BROTHERS OF THE DRAGON 2]  C/P:2,SFWA-P,P
BAILEY, ROBIN WAYNE  FROST [FROST 1]  C/P:2,SFWA-P,P
BAILEY, ROBIN WAYNE  LAKE OF FIRE, THE [DUNGEON 4]  C/P:2,SFWA-P,P
BAILEY, ROBIN WAYNE  LOST CITY OF ZORK, THE [ZORK]  C/P:2,P
BAILEY, ROBIN WAYNE  NIGHT’S ANGEL  LP
BAILEY, ROBIN WAYNE  NIGHTWATCH [GREYHAWK]  C/P,P
BAILEY, ROBIN WAYNE  SHADOWDANCE  H,P
BAILEY, ROBIN WAYNE  SKULL GATE [FROST 2]  C/P,SFWA-P,P
BAILEY, ROBIN WAYNE  SWORDS AGAINST THE SHADOWLAND [FAFHRD]  H,LP
BAILEY, ROBIN WAYNE  TRIUMPH OF THE DRAGON [BROTHERS OF THE DRAGON 3]  C/P,P
BAIN, F. W.  DRAUGHT OF THE BLUE, A  H
BAIN, F. W.  LIVERY OF EVE, THE  H
BAIRD, ALISON  ARCHONS OF THE STARS, THE [@DRAGON THRONE #3]  LP
BAIRD, ALISON  EMPIRE OF THE STARS, THE [@DRAGON THRONE #2]  LP
BAIRD, ALISON  STONE OF STARS, THE [@DRAGON THRONE #1]  LP
BAIRD, WILHELMINA  CHAOS COME AGAIN  P
BAIRD, WILHELMINA  CLIPJOINT [CRASHCOURSE 2]  P
BAIRD, WILHELMINA  CRASHCOURSE [CRASHCOURSE 1]  C/P:2,SFWA-P,P
BAIRD, WILHELMINA  PSYKOSIS  P
BAKER, DENYS VAL  FACE IN THE MIRROR, THE  H
BAKER, ERIC T.  CHECKMATE  C/P,P
BAKER, FRED  PTOLIA [PTOLIA 1]  P
BAKER, JANE/BAKER, PIP  TIME AND THE RANI [DOCTOR WHO]  P
BAKER, KAGE  ANVIL OF THE WORLD, THE  H
BAKER, KAGE  BLACK PROJECTS, WHITE KNIGHTS [COMPANY]  H
BAKER, KAGE  CHILDREN OF THE COMPANY, THE [COMPANY 6]  H:2
BAKER, KAGE  DARK MONDAYS  H
BAKER, KAGE  GODS AND PAWNS [COMPANY 8]  H
BAKER, KAGE  GRAVEYARD GAME, THE [COMPANY 4]  LP
BAKER, KAGE  IN BAD COMPANY [COMPANY 3,4]  H
BAKER, KAGE  IN THE GARDEN OF IDEN [COMPANY 1]  C/P:2,P
BAKER, KAGE  LIFE OF THE WORLD TO COME, THE [COMPANY 5]  H
BAKER, KAGE  MACHINE’S CHILD, THE [COMPANY 7]  H
BAKER, KAGE  MENDOZA IN HOLLYWOOD [COMPANY 3]  P
BAKER, KAGE  ON COMPANY TIME [COMPANY 1,2]  H
BAKER, KAGE  RUDE MECHANICALS [COMPANY]  SR-H
BAKER, KAGE  SKY COYOTE [COMPANY 2]  P
BAKER, KAGE  SONS OF HEAVEN, THE [COMPANY 9]  H
BAKER, KEITH  CITY OF TOWERS, THE [@EBERRON DREAMING DARK #1]  P
BAKER, KEITH  GATES OF NIGHT, THE [@EBERRON DREAMING DARK #3]  P
BAKER, KEITH  SHATTERED LAND, THE [@EBERRON DREAMING DARK #2]  P
BAKER, LINDA P.  IRDA, THE [DRAGONLANCE LOST HISTORIES 2]  C/P,P
BAKER, PIP/BAKER, JANE  MARK OF THE RANI, THE [DOCTOR WHO]  P
BAKER, PIP/BAKER, JANE  RACE AGAINST TIME [DOCTOR WHO]  [C/RAN-P]
BAKER, PIP/BAKER, JANE  TERROR OF THE VERVOIDS [DOCTOR WHO]  SR-P,C/P
BAKER, PIP/BAKER, JANE  ULTIMATE FOE, THE [DOCTOR WHO]  P
BAKER, RICHARD  CITY OF RAVENS, THE [@FORGOTTEN REALMS CITIES]  P
BAKER, RICHARD  CONDEMNATION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #3]  H,P
BAKER, RICHARD  FARTHEST REACH [@FORGOTTEN REALMS LAST MYTHAL #2]  P
BAKER, RICHARD  FINAL GATE [@FORGOTTEN REALMS LAST MYTHAL #3]  P
BAKER, RICHARD  FORSAKEN HOUSE [@FORGOTTEN REALMS LAST MYTHAL #1]  P
BAKER, RICHARD  SHADOW STONE [@FORGOTTEN REALMS]  P
BAKER, RICHARD  ZERO POINT [STAR*DRIVE]  P
BAKER, ROBERT A.  STRESS ANALYSIS OF A STRAPLESS EVENING GOWN, A  [C/RAN-H]

BAKER, RUSSELL OUR NEXT PRESIDENT:

THE INCREDIBLE STORY OF WHAT HAPPENED IN THE 1968 ELECTIONS P

AUTHOR  TITLE  CODE
BAKER, SCOTT  ANCESTRAL HUNGERS  LP
BAKER, SCOTT  DHAMPIRE  C/P:2,P
BAKER, SCOTT  DRINK THE FIRE FROM THE FLAMES [ASHLU]  P,SFWA-P
BAKER, SCOTT  FIREDANCE [ASHLU]  C/P,P
BAKER, SCOTT  NIGHT CHILD  H,SFWA,P
BAKER, SCOTT  SYMBIOTE’S CROWN  P
BAKER, SCOTT  WEBS  SFWA-P
BAKER, SHARON  BURNING TEARS OF SASSURUM [NAPHAR 3]  P
BAKER, SHARON  JOURNEY TO MEMBLIAR [NAPHAR 2]  C/P,P
BAKER, SHARON  QUARRELING, THEY MET THE DRAGON [NAPHAR 1]  P
BAKER, WILL  SHADOW HUNTER  P
BAKIS, KIRSTEN  LIVES OF THE MONSTER DOGS  H
BAKKER, R. SCOTT  DARKNESS THAT COMES BEFORE, THE [@PRINCE OF NOTHING #1]  H
BAKKER, R. SCOTT  THOUSANDFOLD THOUGHT, THE [@PRINCE OF NOTHING #3]  H
BAKKER, R. SCOTT  WARRIOR-PROPHET, THE [@PRINCE OF NOTHING #2]  H
BAKKER, ROBERT T.  RAPTOR RED  C/P,P
BAKST, HAROLD  STRANGE VOYAGE OF KIAN THE MARINER, THE  P
BALCHIN, NIGEL  KINGS OF INFINITE SPACE  SR-P,H
BALDWIN, BILL  CANBY’S LEGION  C/P:2,P
BALDWIN, BILL  DEFENDERS, THE [HELMSMAN 5]  C/P:2,P
BALDWIN, BILL  DEFIANCE, THE [HELMSMAN 7]  P
BALDWIN, BILL  GALACTIC CONVOY [HELMSMAN 2]  C/P:2,P
BALDWIN, BILL  HELMSMAN, THE [HELMSMAN 1]  C/P:2,P
BALDWIN, BILL  MERCENARIES, THE [HELMSMAN 4]  C/P:2,P
BALDWIN, BILL  SIEGE, THE [HELMSMAN 6]  C/P,P
BALDWIN, BILL  TROPHY, THE [HELMSMAN 3]  C/P:2,P

BALDWIN, LIZZ REALMS OF SHADOW

[@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS]

PA

BALFOUR, BRUCE  DIGITAL DEAD, THE  P
BALFOUR, BRUCE  FORGE OF MARS, THE  P
BALFOUR, BRUCE  PROMETHEUS ROAD  P
BALFOUR, BRUCE  STAR CRUSADER  C/P,P
BALIZET, CAROL  SEVEN LAST YEARS, THE  P
BALL, BRIAN N.  PLANET PROBABILITY  C/P:2,P
BALL, BRIAN N.  PROBABILITY MAN, THE  C/P:2,P
BALL, BRIAN N.  REGIMENTS OF NIGHT, THE  C/P:2,P
BALL, BRIAN N.  SINGULARITY STATION  C/P,P
BALL, BRIAN N.  SPACE GUARDIANS, THE [SPACE: 1999 3]  C/P:2,P
BALL, BRIAN N.  SUNDOG  C/P:2,P
BALL, BRIAN N.  SURVIVAL [@SPACE: 1999]  P
BALL, BRIAN N.  TALES OF SCIENCE FICTION  SR-PA,C/PA
BALL, BRIAN N.  TIMEPIECE  C/P:2,P
BALL, BRIAN N.  TIMEPIVOT  C/P:2,P
BALL, MARGARET  CHANGEWEAVER  C/P:2,SFWA-P,P
BALL, MARGARET  DISAPPEARING ACT  H
BALL, MARGARET  FLAMEWEAVER  C/P,SFWA-P,P
BALL, MARGARET  LOST IN TRANSLATION  C/P:2,P
BALL, MARGARET  MATHEMAGICS [CHICKS IN CHAINMAIL 4]  C/P,P
BALL, MARGARET  NO EARTHLY SUNNE  C/P,P
BALL, MARGARET  SHADOW GATE, THE  C/P,P
BALLANTINE BOOKS  DOUBLE PHOENIX  PA,C/PA
BALLANTINE BOOKS  SOMETIME, NEVER  SR-PA,C/PA
BALLANTINE BOOKS  STAR WARS BLUEPRINTS [STAR WARS]  A
BALLANTINE BOOKS  STAR WARS TRILOGY, THE [STAR WARS]  LP
BALLANTINE BOOKS  TALES OF THE INCREDIBLE  SCX
BALLANTYNE, TONY  CAPACITY [RECURSION 2]  P
BALLANTYNE, TONY  DIVERGENCE [RECURSION 3]  P
BALLANTYNE, TONY  RECUSRION [RECURSION 1]  P
BALLARD, J. G.  BEST SCIENCE FICTION OF J. G. BALLARD  P
BALLARD, J. G.  BEST SHORT STORIES OF J. G. BALLARD, THE  H,P
BALLARD, J. G.  BILLENIUM  C/P:2,P
BALLARD, J. G.  BURNING WORLD, THE  C/P:2,P
BALLARD, J. G.  CATACLISMO SOLAR  FP
BALLARD, J. G.  CHRONOPOLIS AND OTHER STORIES  H:2
BALLARD, J. G.  CHRONOPOLIS: THE SCIENCE FICTION OF J. G. BALLARD  C/P:2,P
BALLARD, J. G.  CONCRETE ISLAND  P
BALLARD, J. G.  CRASH  P
BALLARD, J. G.  CRYSTAL WORLD, THE  H:2,P
BALLARD, J. G.  DAY OF CREATION, THE  H,P
BALLARD, J. G.  DISASTER AREA, THE  SR-P
BALLARD, J. G.  DROUGHT, THE  C/P,P
BALLARD, J. G.  DROWNED WORLD, THE  C/P:2,VLP
BALLARD, J. G.  DROWNED WORLD, THE/WIND FROM NOWHERE, THE  H:2
BALLARD, J. G.  FOUR-DIMENSIONAL NIGHTMARE, THE  SR-P
BALLARD, J. G.  HELLO AMERICA  P

AUTHOR  TITLE  CODE
BALLARD, J. G.  HIGH-RISE  H,P
BALLARD, J. G.  IMPOSSIBLE MAN, THE  C/P:2,P
BALLARD, J. G.  LOVE AND NAPALM: EXPORT U.S.A.  H
BALLARD, J. G.  LOW-FLYING AIRCRAFT  SR-LP
BALLARD, J. G.  MEMORIES OF THE SPACE AGE  H
BALLARD, J. G.  MUNDO DE CRISTAL, O  FP
BALLARD, J. G.  PASSPORT TO ETERNITY  C/P:2,P
BALLARD, J. G.  SUPER-CANNES  H
BALLARD, J. G.  TERMINAL BEACH  C/P,P
BALLARD, J. G.  TERMINAL BEACH, THE  SR-P
BALLARD, J. G.  UNLIMITED DREAM COMPANY, THE  C/P,P,H
BALLARD, J. G.  VERMILION SANDS  C/P:2,P
BALLARD, J. G.  VOICES OF TIME, THE  C/P:2,P
BALLARD, J. G.  WIND FROM NOWHERE, THE  C/P:2,P
BALLINGER, W. A.  DRUMS OF THE DARK GODS  P
BALLOTTI, MATTHEW  BLACK BLADE, THE  LP
BALLOU, ARTHUR W.  MAROONED IN ORBIT  H
BALMER, EDWIN/WYLIE, PHILIP  AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  H
BAMBER, GEORGE  SEA IS BOILING HOT, THE  C/P:2,P
BANANA COMPANY  BANANA 9000 OPERATING SYSTEM (MACINTOSH)  SR-MM
BANANA, PLASTIC  WHAT’S BIG AND YELLOW AND PUTS OUT FLAMES?  C/P,P
BANCHI, CHARLOTTE A  OUR PLACE IN TIME  LP
BANERJEE, ANJALI  SILVER SPELL, THE [@KNIGHTS OF THE SILVER DRAGON #8]  LP
BANGS, JOHN KENDRICK  HOUSE-BOAT ON THE STYX, A  H
BANGS, JOHN KENDRICK  MR. MUNCHAUSEN  SR-H
BANGS, JOHN KENDRICK  WATER GHOST AND OTHERS, THE  H
BANISTER, MANLY  CONQUEST OF EARTH  C/P:2,P
BANKER, ASHOK K.  PRINCE OF AYODHYA [@RAMAYANA #1]  P
BANKER, ASHOK K.  PRINCE OF AYODHYA (PROOF) [RAMAYANA 1]  SR-LP
BANKS, DAVID  DOCTOR WHO CYBERMEN [DOCTOR WHO]  R/REF-H
BANKS, DAVID  ICEBERG [DOCTOR WHO NEW ADVENTURES]  C/P,P
BANKS, IAIN M.  AGAINST A DARK BACKGROUND  C/P:2,P
BANKS, IAIN M.  ALGEBRAIST, THE  H
BANKS, IAIN M.  BRIDGE, THE  C/P,P,H
BANKS, IAIN M.  COMPLICITY  P
BANKS, IAIN M.  CONSIDER PHLEBAS [CULTURE]  H,P
BANKS, IAIN M.  EXCESSION [CULTURE]  C/P,P,H
BANKS, IAIN M.  FEERSUM ENDJINN  C/P,LP,H
BANKS, IAIN M.  INVERSIONS  H:2
BANKS, IAIN M.  LOOK TO WINDWARD [CULTURE]  H
BANKS, IAIN M.  PLAYER OF GAMES, THE [CULTURE]  C/P:3,P,H
BANKS, IAIN M.  SONG OF STONE, A  P
BANKS, IAIN M.  STATE OF THE ART, THE [CULTURE]  H,P
BANKS, IAIN M.  USE OF WEAPONS [CULTURE]  H:2,P
BANKS, IAIN M.  WALKING ON GLASS  P
BANKS, L. A.  HUNTED, THE [VAMPIRE HUNTRESS 3]  LP
BANKS, L. A.  MINION [VAMPIRE HUNTRESS 1]  LP
BANKS, L. A.  MINION (SPECIAL HUNTRESS EDITION) [VAMPIRE HUNTRESS]  P
BANKS, LYNNE REID  INDIAN IN THE CUPBOARD, THE [INDIAN IN THE CUPBOARD 1]  C/P:2,P
BANKS, MICHAEL/LAMBE, DEAN R.  ODYSSEUS SOLUTION, THE  C/P,P
BANKS, RAYMOND E.  MOONRAPERS, THE  SR-L
BANNISTER, PAUL  STRANGE HAPPENINGS  [C/RAN-H]
BANTAM SPECTRA  SPECIAL EDITIONS SAMPLER 1992  B
BANTAM SPECTRA  SPECTRA SAMPLER 1991  B
BANTAM SPECTRA  SPECTRA TO THE STARS!  B
BANTAM SPECTRA  SPECTRA TO THE STARS: SPECTRA SAMPLER 1993  B
BARBET, PIERRE  BAPHOMET’S METEOR  C/P,P
BARBET, PIERRE  CAVERNICOLES DE WOLF, LES  FP
BARBET, PIERRE  COMPANHEIROS DO INFINITO, OS  FP
BARBET, PIERRE  COSMIC CRUSADERS  C/P:2,P
BARBET, PIERRE  EMPEROR OF ERIDANUS, THE  P
BARBET, PIERRE  ENCHANTED PLANET, THE  P
BARBET, PIERRE  ETOILE DU NEANT, L  FP
BARBET, PIERRE  GAMES PSYBORGS PLAY  C/P,P
BARBET, PIERRE  JOAN-OF-ARC REPLAY, THE  C/P:2,P
BARBET, PIERRE  LIMIERS DE L’INFINI, LES  FP
BARBET, PIERRE  NAPOLEONS OF ERIDANUS, THE  C/P,P
BARBET, PIERRE  SECRET DES QUASARS, LE  FP
BARBOUR, DAVID/RALEIGH, RICHARD  SHADOWS BEND  LP
BARBREE, JAY  PILOT ERROR [@SIX MILLION DOLLAR MAN #4]  P
BARCLAY, JAMES  CHRONICLES OF THE RAVEN 1 [CHRONICLES OF THE RAVEN 1,2]  H
BARCLAY, JAMES  DAWNTHIEF [CHRONICLES OF THE RAVEN 1]  C/P:2,P
BARCLAY, JAMES  DEMONSTORM [LEGENDS OF THE RAVEN 3]  P
BARCLAY, JAMES  ELFSORROW [LEGENDS OF THE RAVEN 1]  LP
BARCLAY, JAMES  NIGHTCHILD [CHRONICLES OF THE RAVEN 3]  LP

The MITSFS Pinkdex  September 24, 2007  Page 31
AUTHOR  TITLE  CODE
BARCLAY, JAMES  NOONSHADE [CHRONICLES OF THE RAVEN 2]  LP
BARCLAY, JAMES  SHADOWHEART [LEGENDS OF THE RAVEN 2]  LP
BARING-GOULD, SABINE  BOOK OF WEREWOLVES, THE  R/REF-H,C/REF-H
BARING-GOULD, WILLIAM S.  ANNOTATED SHERLOCK HOLMES 1, THE  SR-VLH
BARING-GOULD, WILLIAM S.  ANNOTATED SHERLOCK HOLMES 2, THE  SR-VLH
BARJAVEL, RENE  ASHES, ASHES  H,P
BARJAVEL, RENE  FUTURE TIMES THREE  C/P:2,P
BARJAVEL, RENE  ICE PEOPLE, THE  H:2,P
BARJAVEL, RENE  IMMORTALS, THE  C/P,P
BARKER, ALBERT  APOLLO LEGACY, THE  P
BARKER, CLIVE  CABAL  H
BARKER, CLIVE  CLIVE BARKER’S BOOKS OF BLOOD 1  P
BARKER, CLIVE  CLIVE BARKER’S BOOKS OF BLOOD 2  P
BARKER, CLIVE  CLIVE BARKER’S BOOKS OF BLOOD 3  P
BARKER, CLIVE  DAMNATION GAME, THE  H:2,P
BARKER, CLIVE  DAYS OF MAGIC, NIGHTS OF WAR [ABARAT 2]  H
BARKER, CLIVE  ESSENTIAL CLIVE BARKER, THE  H
BARKER, CLIVE  EVERVILLE [ART 2]  H,LP
BARKER, CLIVE  FORMS OF HEAVEN  H
BARKER, CLIVE  GALILEE  H
BARKER, CLIVE  GREAT AND SECRET SHOW, THE [ART 1]  H,P
BARKER, CLIVE  IMAJICA  H:2,LP
BARKER, CLIVE  IN THE FLESH  H:2,P
BARKER, CLIVE  INCARNATIONS  H,LP
BARKER, CLIVE  INHUMAN CONDITION, THE  H:2,P
BARKER, CLIVE  SACRAMENT  H
BARKER, CLIVE  THIEF OF ALWAYS, THE  H,P
BARKER, CLIVE  WEAVEWORLD  C/P,P,H
BARKER, GRAY  THEY KNEW TOO MUCH ABOUT FLYING SAUCERS  [C/RAN-H]
BARKER, GRAY  UNIDENTIFIED, THE  [C/RAN-P]
BARKER, M. A. R.  FLAMESONG [TEKUMEL 2]  C/P,P
BARKER, M. A. R.  MAN OF GOLD, THE [TEKUMEL 1]  C/P,P
BARKER, THOMAS W.  FIVE FOR INFINITY  P
BARLOUGH, JEFFREY E.  DARK SLEEPER [WESTERN LIGHTS 1]  LP
BARLOUGH, JEFFREY E.  HOUSE IN THE HIGH WOOD, THE [WESTERN LIGHTS 2]  LP
BARLOUGH, JEFFREY E.  STRANGE CARGO [WESTERN LIGHTS 3]  LP
BARLOW, JAMES  ONE HALF OF THE WORLD  H

BARLOWE, WAYNE DOUGLAS/ SUMMERS, IAN/MEACHAM, BETH BARLOWE’S GUIDE TO EXTRATERRESTRIALS

R/REF-H

BARLOWE, WAYNE DOUGLAS  EXPEDITION  VLP
BARNES, ARTHUR K.  INTERPLANETARY HUNTER  H:2,P
BARNES, JOHN  APOSTROPHES & APOCALYPSES  H,LP
BARNES, JOHN  ARMIES OF MEMORY, THE [THOUSAND CULTURES 4]  H
BARNES, JOHN  CAESAR’S BICYCLE [@TIMELINE WARS #3]  C/P,P
BARNES, JOHN  CANDLE  H,P
BARNES, JOHN  DUKE OF URANIUM, THE [JAK JINNAKA 1]  C/P:2,P
BARNES, JOHN  EARTH MADE OF GLASS [THOUSAND CULTURES 2]  C/P,P,H
BARNES, JOHN  FINITY  C/P,P,H
BARNES, JOHN  IN THE HALL OF THE MARTIAN KING [JAK JINNAKA 3]  C/P,P
BARNES, JOHN  KALEIDOSCOPE CENTURY  H,P
BARNES, JOHN  MAN WHO PULLED DOWN THE SKY, THE  C/P:2,P
BARNES, JOHN  MERCHANTS OF SOULS, THE [THOUSAND CULTURES 3]  H,P
BARNES, JOHN  MILLION OPEN DOORS, A [THOUSAND CULTURES 1]  C/P,P,H
BARNES, JOHN  MOTHER OF STORMS  C/P,P
BARNES, JOHN  ONE FOR THE MORNING GLORY  C/P,P,H
BARNES, JOHN  ORBITAL RESONANCE  C/P,SFWA,P,H
BARNES, JOHN  PATTON’S SPACESHIP [@TIMELINE WARS #1]  C/P:2,P
BARNES, JOHN  PRINCESS OF THE AERIE, A [JAK JINNAKA 2]  P
BARNES, JOHN  SIN OF ORIGIN  P
BARNES, JOHN  SKY SO BIG AND BLACK, THE  H:2
BARNES, JOHN  UNION FIRES [@TIMERAIDER]  P
BARNES, JOHN  WARTIDE [@TIMERAIDER]  P
BARNES, JOHN  WASHINGTON’S DIRIGIBLE [@TIMELINE WARS #2]  C/P:2,P
BARNES, STEVEN  BLOOD BROTHERS  P
BARNES, STEVEN  CESTUS DECEPTION, THE [@STAR WARS CLONE WARS]  P
BARNES, STEVEN  CHARISMA  H,P
BARNES, STEVEN  FAR BEYOND THE STARS [@STAR TREK DS9]  P
BARNES, STEVEN  FIREDANCE [AUBRY KNIGHT]  H
BARNES, STEVEN  GORGON CHILD [AUBRY KNIGHT 2]  C/P:2,SFWA-P,P
BARNES, STEVEN  IRON SHADOWS  H
BARNES, STEVEN  KUNDALINI EQUATION, THE  C/P,SFWA-P,P
BARNES, STEVEN  LION’S BLOOD [LION’S BLOOD 1]  C/P,P
BARNES, STEVEN  STREETLETHAL [AUBRY KNIGHT 1]  C/P:2,SFWA-P,P
BARNES, STEVEN  ZULU HEART [LION’S BLOOD 2]  H,P

The MITSFS Pinkdex  September 24, 2007  Page 32
AUTHOR  TITLE  CODE
BARNES-SVARNEY, PATRICIA  LOYALTIES [@STAR TREK TNG STARFLEET ACADEMY #10]  LP
BARNES-SVARNEY, PATRICIA  QUARANTINE [@STAR TREK VOYAGER STARFLEET ACADEMY #3]  LP
BARNES-SVARNEY, PATRICIA  TEACHER’S PET [@SABRINA THE TEENAGE WITCH SALEM #2]  P
BARNETT, LINCOLN  UNIVERSE AND DOCTOR EINSTEIN, THE  [C/RAN-P:2]
BARNITZ, CHARLES  DEEPEST SEA, THE  P
BARNWELL, WILLIAM  BLESSING PAPERS, THE  P
BARNWELL, WILLIAM  IMRAM  P
BARNWELL, WILLIAM  SIGMA CURVE, THE  P
BARON, MIKE  DEMONS [@WITCHBLADE]  P
BARON, NICK  GLORY’S END [TIME TOURS 2]  P
BARONET  YOUR COMPLETE GUIDE TO FATE AND FORTUNE  [C/RAN-P]
BARR, DENSIL NEVE  MAN WITH ONLY ONE HEAD, THE  SR-P
BARR, DONALD  PLANET IN ARMS, A  C/P:2,P
BARR, DONALD  SPACE RELATIONS  C/P,P,H
BARR, GEORGE  UPON THE WINDS OF YESTERDAY  A
BARR, MARLENE S.  FUTURE FEMALES: A CRITICAL ANTHOLOGY  R/REF-P
BARR, MIKE W.  GEMINI [STAR TREK]  P
BARR, TYRONE C.  LAST FOURTEEN, THE  SR-P,C/P
BARR, TYRONE C.  SPLIT WORLDS  SR-P
BARRETT, MICHAEL DENNIS  ASYLUM AND CIRCUS  P
BARRETT, NEAL JR.  ALDAIR, ACROSS THE MISTY SEA [ALDAIR 3]  P
BARRETT, NEAL JR.  ALDAIR IN ALBION [ALDAIR 1]  P
BARRETT, NEAL JR.  ALDAIR, MASTER OF SHIPS [ALDAIR 2]  C/P,P
BARRETT, NEAL JR.  ALDAIR: THE LEGION OF BEASTS [ALDAIR 4]  P
BARRETT, NEAL JR.  C’ERA UNA VOLTA L’AMERICA  FP
BARRETT, NEAL JR.  DAWN’S UNCERTAIN LIGHT  P
BARRETT, NEAL JR.  DUNGEONS & DRAGONS THE MOVIE [@DUNGEONS & DRAGONS]  P
BARRETT, NEAL JR.  GATES OF TIME, THE  D27400
BARRETT, NEAL JR.  HEREAFTER GANG, THE  H
BARRETT, NEAL JR.  HIGHWOOD  D33710
BARRETT, NEAL JR.  JUDGE DREDD [JUDGE DREDD]  P
BARRETT, NEAL JR.  KARMA CORPS, THE  C/P:2,P
BARRETT, NEAL JR.  KELWIN  C/P,P
BARRETT, NEAL JR.  LEAVES OF TIME, THE  C/P:2,P
BARRETT, NEAL JR.  PERPETUITY BLUES AND OTHER STORIES  H
BARRETT, NEAL JR.  PROPHECY MACHINE, THE  C/P,P
BARRETT, NEAL JR.  STRESS PATTERN  C/P,P
BARRETT, NEAL JR.  THROUGH DARKEST AMERICA  H

BARRETT, NEAL JR.

TOUCH OF YOUR SHADOW, THE WHISPER OF YOUR NAME, THE [@BABYLON 5 #5]

P

BARRETT, NEAL JR.  TREACHERY OF KINGS, THE  P
BARRETT, WILLIAM E.  FOOLS OF TIME, THE  C/P,P,H
BARRIE, JAMES M.  PETER PAN  P
BARRIE, MONICA  QUEEN OF KNIGHTS  P
BARRINGER, LESLIE  GERFALCON [NEUSTRIAN CYCLE 1]  LP
BARRINGER, LESLIE  JORIS OF THE ROCK [NEUSTRIAN CYCLE 2]  LP
BARRINGER, LESLIE  SHY LEOPARDESS [NEUSTRIAN CYCLE 3]  LP
BARRON, D. G.  ZILOV BOMBS, THE  H
BARRON, LAIRD  IMAGO SEQUENCE, THE  H
BARRON, NEIL  ANATOMY OF WONDER: SCIENCE FICTION  R/REF-P,C/REF-H
BARRON, T. A.  ANCIENT ONE, THE  P
BARRON, T. A.  CHILD OF THE DARK PROPHECY [@GREAT TREE OF AVALON 1]  P
BARRON, T. A.  ETERNAL FLAME [@GREAT TREE OF AVALON #3]  H
BARRON, T. A.  HEARTLIGHT  P
BARRON, T. A.  LOST YEARS OF MERLIN, THE [LOST YEARS OF MERLIN 1]  P
BARRON, T. A.  MERLIN EFFECT, THE  C/P,P
BARRON, T. A.  MIRROR OF MERLIN, THE [LOST YEARS OF MERLIN 4]  P
BARRON, T. A.  SEVEN SONGS OF MERLIN, THE [LOST YEARS OF MERLIN 2]  P
BARRON, T. A.  SHADOWS ON THE STARS [GREAT TREE OF AVALON 2]  P
BARRON, T. A.  WINGS OF MERLIN [LOST YEARS OF MERLIN 5]  P
BARRY, DAVE/PEARSON, RIDLEY  PETER AND THE SHADOW THIEVES [PETER 2]  H
BARRY, DAVE/PEARSON, RIDLEY  PETER AND THE STARCATCHERS [PETER 1]  H
BARRY, MAX  JENNIFER GOVERNMENT  H:2
BART BOOKS  BART SCIENCE FICTION TRIPLET #1  PA
BARTH, JOHN  CHIMERA  C/P,P
BARTH, JOHN  GILES GOAT-BOY  P
BARTHELME, DONALD  KING, THE  H
BARTHOLOMEW, BARBARA  CHILD OF TOMORROW [TIME KEEPER 2]  P
BARTHOLOMEW, BARBARA  FLIGHT INTO THE UNKNOWN  P
BARTHOLOMEW, BARBARA  TIME KEEPER, THE [TIME KEEPER 1]  P
BARTHOLOMEW, BARBARA  WHEN DREAMERS CEASE TO DREAM [TIME KEEPER 3]  P
BARTLETT, R. E.  PERSONIFIED PROJECT, THE  LP
BARTON, ERLE  PLANET SEEKERS, THE  SR-P
BARTON, JAMES  AFTERMATH [WASTEWORLD 1]  P

AUTHOR  TITLE  CODE
BARTON, JAMES  RESURRECTION [WASTEWORLD 2]  P
BARTON, WILLIAM  ACTS OF CONSCIENCE  P
BARTON, WILLIAM/CAPOBIANCO, MICHAEL ALPHA CENTAURI  C/P,LP,H
BARTON, WILLIAM DARK SKY LEGION C/P:3,SFWA-P,P
BARTON, WILLIAM/CAPOBIANCO, MICHAEL FELLOW TRAVELER  C/P:2,P
BARTON, WILLIAM HUNTING ON KUNDERER D48245
BARTON, WILLIAM/CAPOBIANCO, MICHAEL IRIS  C/P,SFWA-P,P,H
BARTON, WILLIAM  PLAGUE OF ALL COWARDS, A  C/P:2,P
BARTON, WILLIAM  TRANSMIGRATION OF SOULS, THE  C/P:2,P
BARTON, WILLIAM  WHEN HEAVEN FELL  P
BARTON, WILLIAM  WHEN WE WERE REAL  P
BARTON, WILLIAM/CAPOBIANCO, MICHAEL WHITE LIGHT  C/P,P,H
BARTTER, MARTHA A.  WAY TO GROUND ZERO, THE  R/REF-H
BARUTH, PHILIP E.  X PRESIDENT, THE  LP
BARUTH, PHILIP E.  X PRESIDENT, THE (PROOF)  SR-LP
BARZMAN, BEN  ECHO X  C/P:2,P
BARZMAN, BEN  TWINKLE TWINKLE LITTLE STAR  H:2
BASILE, GLORIA VITANZA  EYE OF THE EAGLE [GLOBAL 2000 1]  P
BASILE, GLORIA VITANZA  JACKAL HELIX, THE [GLOBAL 2000 2]  P
BASILE, GLORIA VITANZA  STING OF THE SCORPION, THE [GLOBAL 2000 3]  P
BASS, MILTON R.  FORCE RED  H,P
BASS, T. J.  GODWHALE, THE  C/P:2,P
BASS, T. J.  HALF PAST HUMAN  C/P:2,P
BASS, THOMAS  EUDAEMONIC PIE, THE  [C/RAN-H]
BASSETT, JAMES C.  LIVING REAL  C/P,P
BASSINGTHWAITE, DON  BINDING STONE, THE [@EBERRON DRAGON BELOW #1]  P
BASSINGTHWAITE, DON  BREATHE DEEPLY  C/P,P
BASSINGTHWAITE, DON  GRIEVING TREE, THE [@EBERRON DRAGON BELOW #2]  P
BASSINGTHWAITE, DON  IF WHISPERS CALL [DARK MATTER 2]  P
BASSINGTHWAITE, DON  KILLING SONG, THE [@EBERRON DRAGON BELOW #3]  P
BASSINGTHWAITE, DON/GROSS, DAVE  MISTRESS OF THE NIGHT [@FORGOTTEN REALMS PRIESTS]  P
BASSINGTHWAITE, DON  POMEGRANATES FULL AND FINE [@WORLD OF DARKNESS]  P
BASSINGTHWAITE, DON  SUCH PAIN [@WORLD OF DARKNESS MAGE 1]  C/P,P
BASSINGTHWAITE, DON  YELLOW SILK, THE [@FORGOTTEN REALMS ROGUES]  P
BAST, ANYA  WITCH FIRE  P
BASTRAW, MICHAEL  FIFTY EXTREMELY SHORT FICTION STORIES  SR-LPA
BATCHELOR, JOHN CALVIN  BIRTH OF THE PEOPLE’S REPUBLIC OF ANTARCTICA, THE  H,SFWA,LP
BATCHELOR, JOHN CALVIN  FURTHER ADVENTURES OF HALLEY’S COMET, THE  P
BATEMAN, ROBERT  WHEN THE WHITES WENT  SR-P

BATES, ANDREW HERALDS OF THE STORM

[@WORLD OF DARKNESS YEAR OF THE SCARAB #1]

P

BATES, ANDREW  LAND OF THE DEAD [@WORLD OF DARKNESS YEAR OF THE SCARAB #3]  P

BATES, ANDREW LAY DOWN WITH LIONS

[@WORLD OF DARKNESS YEAR OF THE SCARAB #2]

P

BATO, JOSEPH  SORCERER, THE  H
BATRAE, MARGOT  WHILE THE CAT’S AWAY [@SABRINA THE TEENAGE WITCH #25]  P
BAUDINO, GAEL  BRANCH AND CROWN [@WATER! #3]  C/P,P
BAUDINO, GAEL  DOVE LOOKED IN, THE [@WATER! #2]  P
BAUDINO, GAEL  DRAGON DEATH [DRAGONSWORD 3]  C/P,SFWA-P,P
BAUDINO, GAEL  DRAGONSWORD [DRAGONSWORD 1]  C/P:2,P
BAUDINO, GAEL  DUEL OF DRAGONS [DRAGONSWORD 2]  C/P,SFWA-P:2,P
BAUDINO, GAEL  GOSSAMER AXE  C/P:2,P
BAUDINO, GAEL  MAZE OF MOONLIGHT [STRANDS OF STARLIGHT 2]  C/P,P
BAUDINO, GAEL  O GREENEST BRANCH! [@WATER! #1]  P
BAUDINO, GAEL  SHROUD OF SHADOW [STRANDS OF STARLIGHT 3]  C/P:2,P
BAUDINO, GAEL  SPIRES OF SPIRIT [STRANDS OF STARLIGHT]  P
BAUDINO, GAEL  STRANDS OF STARLIGHT [STRANDS OF STARLIGHT 1]  C/P:2,P
BAUDINO, GAEL  STRANDS OF SUNLIGHT [STRANDS OF STARLIGHT 4]  C/P,P
BAUER, SABINE C.  SURVIVAL OF THE FITTEST [@STARGATE SG-1 #7]  P
BAUER, SABINE C.  TRIAL BY FIRE [@STARGATE SG-1 #1]  P
BAUER, STEVEN  SATYRDAY  H,P
BAUM, L. FRANK  ANNOTATED WIZARD OF OZ, THE  SR-H,H
BAUM, L. FRANK  DOROTHY AND THE WIZARD IN OZ [OZ #4]  H,@P,@C/P
BAUM, L. FRANK  EMERALD CITY OF OZ, THE [OZ #6]  H:2,@P,@C/P:2
BAUM, L. FRANK  GLINDA OF OZ [OZ #14]  C/P,@P,H
BAUM, L. FRANK  JOHN DOUGH AND THE CHERUB  LP
BAUM, L. FRANK  LAND OF OZ, THE [OZ #2]  H,@P,@C/P:2
BAUM, L. FRANK  LIFE AND ADVENTURES OF SANTA CLAUS, THE  C/P,P,H
BAUM, L. FRANK  LOST PRINCESS OF OZ, THE [OZ #11]  H,@P
BAUM, L. FRANK  MAGIC OF OZ, THE [OZ #13]  H,@P
BAUM, L. FRANK  MAGICAL MONARCH OF MO, THE  H,LP
BAUM, L. FRANK  MARVELOUS LAND OF OZ, THE [OZ 2]  H:2,LP
BAUM, L. FRANK  MASTER KEY: AN ELECTRICAL FAIRY TALE, THE  H,LP
BAUM, L. FRANK  OZMA OF OZ [OZ #3]  H:2,@P,@C/P

The MITSFS Pinkdex  September 24, 2007  Page 34
AUTHOR  TITLE  CODE
BAUM, L. FRANK  PATCHWORK GIRL OF OZ, THE [OZ #7]  H,@P
BAUM, L. FRANK  QUEEN ZIXI OF IX; OR, THE STORY OF THE MAGIC CLOAK  H,LP
BAUM, L. FRANK  RINKITINK IN OZ [OZ #10]  C/P,@P,H
BAUM, L. FRANK  ROAD TO OZ, THE [OZ #5]  C/P,@P,@C/P,H
BAUM, L. FRANK  SCARECROW OF OZ, THE [OZ #9]  H:2,@P
BAUM, L. FRANK  SKY ISLAND  H
BAUM, L. FRANK  TIK-TOK OF OZ [OZ #8]  C/P,@P,H
BAUM, L. FRANK  TIN WOODMAN OF OZ, THE [OZ #12]  H,@P,@C/P
BAUM, L. FRANK  VISITORS FROM OZ, THE  H
BAUM, L. FRANK  WIZARD OF OZ, THE [OZ #1]  H,@P,@C/P:2
BAUM, L. FRANK  WONDERFUL WIZARD OF OZ, THE [OZ 1]  H:3,LP
BAUM, THOMAS  IT LOOKS ALIVE TO ME!  LP
BAUMBACH, LAURA  DETAILS OF THE HUNT  LP
BAXENDALE, TREVOR  COLDHEART [DOCTOR WHO EIGHTH DOCTOR]  P
BAXENDALE, TREVOR  DEADSTONE MEMORIAL, THE [@DOCTOR WHO EIGHTH DOCTOR]  P
BAXENDALE, TREVOR  EATER OF WASPS [DOCTOR WHO EIGHTH DOCTOR]  P
BAXENDALE, TREVOR  FEAR OF THE DARK [DOCTOR WHO PAST DOCTORS]  P
BAXENDALE, TREVOR  JANUS CONJUNCTION, THE [DOCTOR WHO EIGHTH DOCTOR]  P
BAXTER, JOHN  HERMES FALL, THE  C/P,P
BAXTER, JOHN  OFF-WORLDERS, THE  DG588
BAXTER, JOHN  PACIFIC BOOK OF SCIENCE FICTION, THE  HA
BAXTER, JOHN  SCIENCE FICTION IN THE CINEMA  [C/RAN-P]
BAXTER, JOHN  SECOND PACIFIC BOOK OF SCIENCE FICTION, THE  HA
BAXTER, LORNA  EGGCHILD, THE  P
BAXTER, STEPHEN  ANTI-ICE  C/P,P
BAXTER, STEPHEN  COALESCENT [DESTINY’S CHILDREN 1]  H,P
BAXTER, STEPHEN  DEEP FUTURE  [C/RAN-P:2]
BAXTER, STEPHEN  EMPEROR [@TIME’S TAPESTRY #1]  H
BAXTER, STEPHEN  EVOLUTION  H:2
BAXTER, STEPHEN  EXULTANT [DESTINY’S CHILDREN 2]  H
BAXTER, STEPHEN  FLUX [XEELEE]  C/P,P,H
BAXTER, STEPHEN  GULLIVERZONE [WEB,THE]  P
BAXTER, STEPHEN  ICEBONES [MAMMOTHS]  H,LP
BAXTER, STEPHEN  LONGTUSK [MAMMOTHS 2]  C/P,P
BAXTER, STEPHEN  MANIFOLD: ORIGIN [REID MALENFANT 3]  C/P,P,H
BAXTER, STEPHEN  MANIFOLD: SPACE [REID MALENFANT 2]  C/P,P,H:2
BAXTER, STEPHEN  MANIFOLD: TIME [REID MALENFANT 1]  C/P,P,H
BAXTER, STEPHEN  MOONSEED  H,P
BAXTER, STEPHEN  PHASE SPACE [REID MALENFANT]  H
BAXTER, STEPHEN  RAFT [XEELEE]  C/P:2,SFWA-P,P
BAXTER, STEPHEN  REALITY DUST  GZD07306
BAXTER, STEPHEN  RIDING THE ROCK [XEELEE]  SR-LP
BAXTER, STEPHEN  RING [XEELEE]  C/P:2,P
BAXTER, STEPHEN  SILVERHAIR [MAMMOTHS 1]  C/P,P
BAXTER, STEPHEN  TIME SHIPS, THE  C/P:2,P
BAXTER, STEPHEN  TIMELIKE INFINITY [XEELEE]  C/P,P
BAXTER, STEPHEN  TITAN  C/P:2,P
BAXTER, STEPHEN  TRACES  P
BAXTER, STEPHEN  TRANSCENDENT [DESTINY’S CHILDREN 3]  H
BAXTER, STEPHEN  VACUUM DIAGRAMS  C/P:2,LP,H
BAXTER, STEPHEN  VOYAGE  C/P,LP,H
BAY, AUSTIN  PRISM  P
BAYLEY, BARRINGTON J.  ANNIHILATION FACTOR  D33710
BAYLEY, BARRINGTON J.  COLLISION COURSE  C/P:2,P
BAYLEY, BARRINGTON J.  COLLISION WITH CHRONOS  P
BAYLEY, BARRINGTON J.  EMPIRE OF TWO WORLDS  P
BAYLEY, BARRINGTON J.  EYE OF TERROR [WARHAMMER 40K]  P
BAYLEY, BARRINGTON J.  FALL OF CHRONOPOLIS, THE  C/P:2,P
BAYLEY, BARRINGTON J.  FOREST OF PELDAN, THE  P
BAYLEY, BARRINGTON J.  GARMENTS OF CAEAN, THE (U.S. EDITION)  C/P,SFWA,P,H
BAYLEY, BARRINGTON J.  GRAND WHEEL, THE  P
BAYLEY, BARRINGTON J.  PILLARS OF ETERNITY, THE  P
BAYLEY, BARRINGTON J.  SOUL OF THE ROBOT  C/P,SFWA,P,H
BAYLEY, BARRINGTON J.  STAR VIRUS, THE  D78400
BAYLEY, BARRINGTON J.  STAR WINDS  C/P,P
BAYLEY, BARRINGTON J.  ZEN GUN, THE  P
BAYLUS, ROBERT F.  PEOPLE EXCHANGE, THE  SR-P,SFWA-P
BBC BOOKS  SHOOTING SCRIPTS, THE [DOCTOR WHO TENTH DOCTOR]  R/REF-H
BBC TV  DOCTOR WHO ANNUAL 1985 [DOCTOR WHO]  HA
BBC TV  DOCTOR WHO ANNUAL 1986 [DOCTOR WHO]  HA

BBC TV DOCTOR WHO ANNUAL (YEAR UNKNOWN, COPYRIGHT 1981)

[DOCTOR WHO]

HA

BBC TV  DOCTOR WHO: JOURNEY THROUGH TIME [DOCTOR WHO]  H
BEAGLE, PETER S.  FANTASY WORLDS OF PETER S. BEAGLE, THE  H:2,LP

The MITSFS Pinkdex  September 24, 2007  Page 35
AUTHOR  TITLE  CODE
BEAGLE, PETER S.  FINE AND PRIVATE PLACE, A  C/P:2,P
BEAGLE, PETER S.  FOLK OF THE AIR, THE  C/P:2,SFWA,P
BEAGLE, PETER S.  INNKEEPER’S SONG, THE  H,LP
BEAGLE, PETER S.  LAST UNICORN, THE  C/P:2,P
BEAGLE, PETER S.  LINE BETWEEN, THE  LP
BEAGLE, PETER S./BERLINER, JANET  PETER S. BEAGLE’S IMMORTAL UNICORN  HA
BEAGLE, PETER S./BERLINER, JANET  PETER S. BEAGLE’S IMMORTAL UNICORN, VOLUME 1  PA
BEAGLE, PETER S./BERLINER, JANET  PETER S. BEAGLE’S IMMORTAL UNICORN, VOLUME 2  PA

BEAGLE, PETER S.

RHINOCEROS WHO QUOTED NIETZCHE AND OTHER ODD ACQUAINTANCES, THE

LP

BEAGLE, PETER S.  TAMSIN  H:2
BEAGLE, PETER S.  UNICORN SONATA, THE  H
BEALE, THEODORE  WAR IN HEAVEN, THE [ETERNAL WARRIORS 1]  P
BEALS, CARLETON  DAWN OVER THE AMAZON  H
BEAMER, CHARLES  WHEN THE GODS RETURNED  P
BEAR, DAVID  KEEPING TIME  H,P
BEAR, ELIZABETH  BLOOD AND IRON [PROMETHEAN AGE 1]  LP
BEAR, ELIZABETH  CARNIVAL  P
BEAR, ELIZABETH  CHAINS THAT YOU REFUSE, THE  LP
BEAR, ELIZABETH  HAMMERED [JENNY CASEY 1]  C/P,P
BEAR, ELIZABETH  NEW AMSTERDAM  H
BEAR, ELIZABETH  SCARDOWN [JENNY CASEY 2]  P
BEAR, ELIZABETH  WHISKEY AND WATER [PROMETHEAN AGE 2]  LP
BEAR, ELIZABETH  WORLDWIRED [JENNY CASEY 3]  P
BEAR, GREG  ANVIL OF STARS [FORGE OF GOD 2]  H:2,P
BEAR, GREG  BEAR’S FANTASIES  SR-H
BEAR, GREG  BEYOND HEAVEN’S RIVER  P,SFWA-P
BEAR, GREG  BLOOD MUSIC  H:2,P,SFWA
BEAR, GREG  COLLECTED STORIES OF GREG BEAR, THE  H
BEAR, GREG  CORONA [STAR TREK]  C/P:3,P
BEAR, GREG  DARWIN’S CHILDREN [SHEVA 2]  H
BEAR, GREG  DARWIN’S RADIO [SHEVA 1]  H:2,P
BEAR, GREG  DINOSAUR SUMMER  P
BEAR, GREG  EARLY HARVEST  SR-H
BEAR, GREG  EON [EON 1]  C/P,SFWA-P:2,H:2,P
BEAR, GREG  ETERNITY [EON 2]  SFWA-P,H:2,P
BEAR, GREG  FORGE OF GOD, THE [FORGE OF GOD 1]  H:2,SFWA-P,P
BEAR, GREG  FOUNDATION AND CHAOS [SECOND FOUNDATION 2]  C/P,P,H
BEAR, GREG  HARDFOUGHT  TD2,SFWA-TD2
BEAR, GREG  HEADS [QUEEN OF ANGELS 2]  C/P,P,H
BEAR, GREG  HEGIRA  C/P,SFWA-P,P
BEAR, GREG  INFINITY CONCERTO, THE [MICHAEL PERRIN 1]  C/P:2,P
BEAR, GREG  LEGACY [EON 0]  H:2,P
BEAR, GREG  MOVING MARS [QUEEN OF ANGELS 3]  H:2,P
BEAR, GREG  PSYCHLONE  C/P:2,SFWA-P,P
BEAR, GREG  QUEEN OF ANGELS [QUEEN OF ANGELS 1]  C/P:2,P
BEAR, GREG  ROGUE PLANET [@STAR WARS]  P
BEAR, GREG  SERPENT MAGE, THE [MICHAEL PERRIN 2]  C/P:2,P
BEAR, GREG  SISTERS  B
BEAR, GREG  SLANT [QUEEN OF ANGELS 4]  C/P:2,P
BEAR, GREG  SONGS OF EARTH & POWER [MICHAEL PERRIN (REVISED) 1,2]  P
BEAR, GREG  STRENGTH OF STONES  C/P,P
BEAR, GREG  TANGENTS  H:2,SFWA,P
BEAR, GREG  VENGING, THE  P
BEAR, GREG  VITALS  H
BEAR, GREG  W3: WOMEN IN DEEP TIME  P
BEAR, GREG  WIND FROM A BURNING WOMAN, THE  C/P:2,P
BEATH, WARREN NEWTON  WHO KILLED JAMES DEAN?  SFWA-P
BEATTY, JEROME JR.  MARIA LOONEY AND THE COSMIC CIRCUS  LP
BEATTY, JEROME JR.  MARIA LOONEY AND THE REMARKABLE ROBOT  LP
BEATTY, JEROME JR.  MARIA LOONEY ON THE RED PLANET  LP
BEATTY, JEROME JR.  MATTHEW LOONEY AND THE SPACE PIRATES  C/P,LP
BEATTY, JEROME JR.  MATTHEW LOONEY IN THE OUTBACK  LP
BEATTY, JEROME JR.  MATTHEW LOONEY’S INVASION OF THE EARTH  C/P,LP
BEATTY, JEROME JR.  MATTHEW LOONEY’S VOYAGE TO THE EARTH  H,LP
BEAUMONT, CHARLES  BEST OF BEAUMONT  P
BEAUMONT, CHARLES  EDGE, THE  SR-P
BEAUMONT, CHARLES  HUNGER AND OTHER STORIES, THE  C/P,P
BEAUMONT, CHARLES  MAGIC MAN, THE  C/P:2,P
BEAUMONT, CHARLES  NIGHT RIDE AND OTHER JOURNEYS  P
BEAUMONT, CHARLES  SHADOW PLAY  SR-P
BEAUMONT, CHARLES  YONDER  P
BEAUSEIGNEUR, JAMES  ACTS OF GOD [@CHRIST CLONE #3]  H
BEAUSEIGNEUR, JAMES  BIRTH OF AN AGE [@CHRIST CLONE #2]  P

The MITSFS Pinkdex  September 24, 2007  Page 36
AUTHOR  TITLE  CODE
BEAUSEIGNEUR, JAMES  BIRTH OF AN AGE (PROOF) [CHRIST CLONE 2]  SR-LP
BEAUSEIGNEUR, JAMES  IN HIS IMAGE [@CHRIST CLONE #1]  P
BEBRIS, CARRIE A.  POOL OF RADIANCE [@FORGOTTEN REALMS RUINS OF MYTH DRANNOR]  P
BECHDOLT, JACK  TORCH, THE  SR-H,H
BECHER, DON  TICKET TO NOWHERE, A  SR-P
BECK, CALVIN  FRANKENSTEIN READER, THE  SR-PA,C/PA
BECK, L. ADAMS  DREAMS AND DELIGHTS  H
BECK, L. ADAMS  HOUSE OF FULFILMENT: THE ROMANCE OF A SOUL, THE  H
BECK, L. ADAMS  WAY OF STARS: A ROMANCE OF REINCARNATION, THE  H
BECK, ROBERT E.  LITERATURE OF THE SUPERNATURAL  LP
BECKER, ALIDA  TOLKIEN SCRAPBOOK, THE  R/REF-H
BECKER, HYAM YONA  TEMPLE OF HASHEM, THE  C/P,P
BECKFORD, WILLIAM  HISTORY OF THE CALIPH WATHEK, THE  LP
BECKFORD, WILLIAM  VATHEK  C/P,P
BECKNER, MORTON  MONEY PLAYS  H
BEDARD, MICHAEL  DARKER MAGIC, A  P
BEDARD, MICHAEL  REDWORK  SFWA-P
BEDDOR, FRANK  LOOKING GLASS WARS, THE  H
BEDFORD-JONES, H./ROBERTSON, W. C.  TEMPLE OF THE TEN, THE  SR-H
BEDFORD, K. A.  HYDROGEN STEEL  LP
BEDWELL-GRIME, STEPHANIE  GUARDIAN ANGEL  LP
BEECHING, JACK  DAKOTA PROJECT, THE  P
BEEDING, FRANCIS  HIDDEN KINGDOM, THE  H
BEEDING, FRANCIS  SPELLBOUND  H
BEEMISH, CRAGG  WORLDS AWAY!  SR-P
BEESE, P. J./HAMILTON, TODD CAMERON  GUARDSMAN, THE  C/P,P
BEHN, HARRY  FARAWAY LURS, THE  P
BEHR, IRA STEVEN  FERENGI RULES OF ACQUISITION, THE [STAR TREK DS9]  R/REF-P,C/REF-P
BELDEN, DAVID  CHILDREN OF ARABLE  C/P,P
BELDEN, DAVID  TO WARM THE EARTH  C/P,P
BELDEN, WILANNE SCHNEIDER  MIND-CALL  H
BELDEN, WILANNE, SCHNEIDER  MIND-HOLD  H:2
BELIAEV, ALEXANDER  AMPHIBIAN, THE  H
BELIAEV, ALEXANDER  PROFESSOR DOWELL’S HEAD  H:2,LP

BELL, ANDREA L./ MOLINA-GAVILAN, YOLANDA COSMOS LATINOS

LPA

BELL, CLARE  CLAN GROUND [RATHA 2]  H
BELL, CLARE  JAGUAR PRINCESS, THE  P
BELL, CLARE  PEOPLE OF THE SKY [RATHA 5]  C/P,SFWA-P,P,H
BELL, CLARE  RATHA’S CREATURE [RATHA 1]  C/P,SFWA,P
BELL, CLARE  TOMORROW’S SPHINX  H
BELL, DOUGLAS  MOJO AND THE PICKLE JAR  C/P,P
BELL, GORDON BENNETT  GOLDEN TROUBADOUR, THE  H
BELL, GORDON BENNETT  GOLDEN TROUBADOUR, THE (PROOF)  SR-LP
BELL, HILARI  FALL OF A KINGDOM [@FARSALA #1]  P
BELL, HILARI  NAVOHAR  P
BELL, HILARI  RISE OF A HERO [@FARSALA #2]  H
BELL, JADRIEN  A.D. 999  P
BELL, M. SHAYNE  INUIT  P
BELL, M. SHAYNE  NICOJI  C/P,SFWA-P,P
BELL, M. SHAYNE  WASHED BY A WAVE OF WIND  LP
BELL, NEAL  GONE TO BE SNAKES NOW  C/P,P
BELL, ROBERT  IN REALMS UNKNOWN  H
BELL, THORNTON  CHAOS  SR-P
BELL, THORNTON  SPACE TRAP  SR-P,H
BELL, WILLIAM DIXON  MOON COLONY, THE  H
BELL, WILLIAM DIXON  SACRED SCIMITER, THE  H
BELLAIRS, JOHN  CURSE OF THE BLUE FIGURINE, THE  H,P
BELLAIRS, JOHN  FACE IN THE FROST, THE  C/P,SFWA-P,P,H
BELLAIRS, JOHN  FIGURE IN THE SHADOWS, THE  LP
BELLAIRS, JOHN  HOUSE WITH A CLOCK IN ITS WALLS, THE  LP
BELLAIRS, JOHN  LAMP FROM THE WARLOCK’S TOMB, THE  P
BELLAIRS, JOHN  LETTER, THE WITCH, AND THE RING, THE  H,LP
BELLAIRS, JOHN  MUMMY, THE WILL, AND THE CRYPT, THE  LP
BELLAMY, EDWARD  EQUALITY  SR-H
BELLAMY, EDWARD  LOOKING BACKWARD 2000–1887  H,SR-H,SR-L,P
BELLAMY, FRANCIS RUFUS  ATTA  DD079,P
BELLAMY, FRANK/BELLAMY, DAVID  TIMEVIEW [DOCTOR WHO]  A
BELLEMARE, PIERRE  LES DOSSIERS D’INTERPOL  FP
BELMONT BOOKS  MASTERS OF SCIENCE FICTION  PA,C/PA
BELMONT BOOKS  NOW AND BEYOND  PA,C/PA
BELMONT BOOKS  WEIRD ONES, THE  PA,C/PA
BEMMANN, HANS  BROKEN GODDESS, THE  P,SFWA,SFWA-P
BEMMANN, HANS  STONE AND THE FLUTE, THE  H,LP
BENARY-ISBERT, MARGOT  WICKED ENCHANTMENT, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 37
AUTHOR  TITLE  CODE
BENCHLEY, PETER  BEAST  P
BENCHLEY, PETER  JAWS  P
BENEFIELD, BARRY  EDDIE AND THE ARCHANGEL MIKE  H
BENFORD, GREGORY  ACROSS THE SEA OF SUNS (ORIGINAL) [GALACTIC CENTER 2]  H:2

BENFORD, GREGORY ACROSS THE SEA OF SUNS (REVISED: EXTRA CHAPTER AT END)

[GALACTIC CENTER 2]

C/P,P

BENFORD, GREGORY  AGAINST INFINITY  C/P:2,P

BENFORD, GREGORY/ GREENBERG, MARTIN H.

ALTERNATE AMERICAS [@WHAT MIGHT HAVE BEEN #4]

C/PA,PA

BENFORD, GREGORY/ GREENBERG, MARTIN H.

ALTERNATE EMPIRES [@WHAT MIGHT HAVE BEEN #1]

C/PA,PA,SFWA-P

BENFORD, GREGORY/ GREENBERG, MARTIN H.

ALTERNATE HEROES [@WHAT MIGHT HAVE BEEN #2]

C/PA,PA

BENFORD, GREGORY  ARTIFACT  C/P:2,SFWA-P:2,P,H
BENFORD, GREGORY  BEYOND INFINITY [FALL OF NIGHT]  H
BENFORD, GREGORY  COSM  C/P:2,P
BENFORD, GREGORY  DEEPER THAN THE DARKNESS  P
BENFORD, GREGORY  EATER  H,P
BENFORD, GREGORY  FAR FUTURES  HA
BENFORD, GREGORY/EKLUND, GORDON  FIND THE CHANGELING  C/P,P
BENFORD, GREGORY  FOUNDATION’S FEAR [SECOND FOUNDATION 1]  H:2,P
BENFORD, GREGORY  FURIOUS GULF [GALACTIC CENTER 5]  C/P,P,H
BENFORD, GREGORY  GREAT SKY RIVER [GALACTIC CENTER 3]  C/P,SFWA,P,H
BENFORD, GREGORY/BRIN, DAVID  HEART OF THE COMET  C/P,SFWA,P,H

BENFORD, GREGORY/ GREENBERG, MARTIN H.

HITLER VICTORIOUS

HA,PA,SFWA-P

BENFORD, GREGORY/CARTER, PAUL A.  ICEBORN  TD14,SFWA-TD14
BENFORD, GREGORY/EKLUND, GORDON  IF THE STARS ARE GODS  C/P:2,P,H
BENFORD, GREGORY  IN ALIEN FLESH  C/P,SFWA-P,P,H
BENFORD, GREGORY  IN THE OCEAN OF NIGHT [GALACTIC CENTER 1]  C/P,P,H
BENFORD, GREGORY  JUPITER PROJECT  C/P:2,P
BENFORD, GREGORY  MARTIAN RACE, THE  H
BENFORD, GREGORY  MATTER’S END  C/P:2,P
BENFORD, GREGORY  MICROCOSMS  PA
BENFORD, GREGORY  NEBULA AWARDS SHOWCASE 2000  LPA
BENFORD, GREGORY  SAILING BRIGHT ETERNITY [GALACTIC CENTER 6]  C/P:2,P
BENFORD, GREGORY/ROTSLER, WILLIAM  SHIVA DESCENDING  C/P:2,P
BENFORD, GREGORY  STARS IN SHROUD, THE  H:3,P
BENFORD, GREGORY  SUNBORN, THE  H
BENFORD, GREGORY  TIDES OF LIGHT [GALACTIC CENTER 4]  C/P:2,SFWA-P,P,H
BENFORD, GREGORY  TIMESCAPE  C/P:2,P,H

BENFORD, GREGORY/ GREENBERG, MARTIN H.

WHAT MIGHT HAVE BEEN I AND II [WHAT MIGHT HAVE BEEN 1,2]

HA

BENFORD, GREGORY/ GREENBERG, MARTIN H.

WHAT MIGHT HAVE BEEN III AND IV [WHAT MIGHT HAVE BEEN 3,4]

HA

BENFORD, GREGORY  WORLDS VAST AND VARIOUS  LP
BENJAMIN, CURT  GATES OF HEAVEN, THE [SEVEN BROTHERS 3]  H
BENJAMIN, CURT  PRINCE OF DREAMS, THE [SEVEN BROTHERS 2]  P
BENJAMIN, CURT  PRINCE OF SHADOW, THE [SEVEN BROTHERS 1]  C/P,P
BENNET, ROBERT AMES  BOWL OF BAAL, THE  H
BENNETT, ANNA ELIZABETH  LITTLE WITCH  C/P,P
BENNETT, CHRISTOPHER L.  BURIED AGE, THE [@STAR TREK TNG]  P
BENNETT, CHRISTOPHER L.  EX MACHINA [@STAR TREK]  P
BENNETT, CHRISTOPHER L.  ORION’S HOUNDS [STAR TREK TITAN]  P
BENNETT, GARY L.  STAR SAILORS, THE  H:2
BENNETT, GEORGE  GREAT TALES OF ACTION AND ADVENTURE  C/PA,PA
BENNETT, MARCIA J.  BEYOND THE DRAAK’S TEETH  C/P:2,P
BENNETT, MARCIA J.  SEEKING THE DREAM BROTHER  C/P:2,P
BENNETT, MARCIA J.  SHADOW SINGER  C/P:2,P
BENNETT, MARCIA J.  WHERE THE NI-LACH  C/P:2,P
BENNETT, MARCIA J.  YARIL’S CHILDREN  C/P:2,P
BENNETT, MIKE  CORDWAINER SMITH CHECKLIST, A  R/IND
BENNETT, NIGEL/ELROD, P. N.  HIS FATHER’S SON  P
BENNETT, NIGEL/ELROD, P. N.  KEEPER OF THE KING  C/P,P
BENNI, STEFANO  TERRA!  H,LP
BENOIT, HENDRA  HENDRA’S BOOK [@PSI PATROL 2]  P
BENOIT, PIERRE  ATLANTIDA  P
BENSEN, DONALD R.  AND HAVING WRIT...  C/P,P,H
BENSEN, DONALD R.  UNKNOWN, THE  PA,C/PA
BENSEN, DONALD R.  UNKNOWN FIVE, THE  PA,C/PA
BENSON, AMBER/GOLDEN, CHRISTOPHER  ACCURSED [@GHOSTS OF ALBION]  LP
BENSON, AMBER/GOLDEN, CHRISTOPHER  INITIATION [@GHOSTS OF ALBION]  SR-H
BENSON, AMBER/GOLDEN, CHRISTOPHER  WITCHERY [@GHOSTS OF ALBION]  LP
BENSON, ANN  PLAGUE TALES, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 38
AUTHOR  TITLE  CODE
BENSON, E. F.  DAVID BLAIZE AND THE BLUE DOOR  H
BENSON, ROBERT HUGH  NECROMANCERS, THE  LP
BENT, JORJ  SOMETIME AFTER THE EQUINOX  P
BENTLEY, JOHN  WHERE ARE THE RUSSIANS?  P
BENTLY, NORMAN S.  ARMADA OF THE AIR  H
BERBERICK, NANCY VARIAN  CHILD OF ELVISH, A [JEWELS OF ELVISH 2]  C/P:2,P
BERBERICK, NANCY VARIAN  DALAMAR THE DARK [@DRAGONLANCE CLASSICS]  P
BERBERICK, NANCY VARIAN  INHERITANCE, THE [@DRAGONLANCE CLASSICS]  P
BERBERICK, NANCY VARIAN  JEWELS OF ELVISH, THE [JEWELS OF ELVISH 1]  P
BERBERICK, NANCY VARIAN  LIONESS, THE [@DRAGONLANCE AGE OF MORTALS]  P
BERBERICK, NANCY VARIAN  PANTHER’S HOARD, THE  P,SFWA-P:2
BERBERICK, NANCY VARIAN  PRISONER OF HAVEN [@DRAGONLANCE AGE OF MORTALS]  P
BERBERICK, NANCY VARIAN  SHADOW OF THE SEVENTH MOON  C/P,SFWA-P:2,P
BERBERICK, NANCY VARIAN  STORMBLADE [DRAGONLANCE HEROES 2]  C/P:2,P
BERCKMAN, EVELYN  VICTORIAN ALBUM, THE  H
BERESFORD, J. D.  HAMPDENSHIRE WONDER, THE  H
BERESFORD, LESLIE  FLYING FISH, THE  B
BERG, CAROL  DAUGHTER OF ANCIENTS [BRIDGE OF D’ARNATH 4]  P
BERG, CAROL  FLESH AND SPIRIT  LP
BERG, CAROL  GUARDIANS OF THE KEEP [BRIDGE OF D’ARNATH 2]  P
BERG, CAROL  RESTORATION [SEYONNE 3]  P
BERG, CAROL  REVELATION [SEYONNE 2]  P
BERG, CAROL  SON OF AVONAR [BRIDGE OF D’ARNATH 1]  P
BERG, CAROL  SONG OF THE BEAST  P
BERG, CAROL  SOUL WEAVER, THE [BRIDGE OF D’ARNATH 3]  P
BERG, CAROL  TRANSFORMATION [SEYONNE 1]  P
BERGAMINI, DAVID  VENUS DEVELOPMENT  C/P:2,P
BERGAUST, ERIK/BELLER, WILLIAM  SATELLITE!  [C/RAN-H,C/RAN-P:2]
BERGER, ERIC  BEST SHORT STORIES  C/PA,PA
BERGER, HAROLD L.  SCIENCE FICTION AND THE NEW DARK AGE  R/REF-P
BERGER, THOMAS  ARTHUR REX  P
BERGER, THOMAS  CHANGING THE PAST  H
BERGER, THOMAS  REGIMENT OF WOMEN  P
BERGIER, JACQUES/DE LATIL, PIERRE  UITVINDINGEN VAN MORGEN  FP
BERGIN, PAUL A.  XUAN AND THE GIRL FROM THE OTHER SIDE  C/P:2,P
BERGLUND, E. P.  SHARDS OF DARKNESS  LP
BERGLUND, EDWARD P.  DISCIPLES OF CTHULHU, THE  C/PA,PA
BERGSTROM, ELAINE  BARONESS OF BLOOD [RAVENLOFT]  P
BERGSTROM, ELAINE  BLOOD RITES  P
BERGSTROM, ELAINE  BLOOD TO BLOOD  P
BERGSTROM, ELAINE  DAUGHTER OF THE NIGHT  P
BERGSTROM, ELAINE  DOOR THROUGH WASHINGTON SQUARE, THE  P
BERGSTROM, ELAINE  SHATTERED GLASS  C/P,P
BERGSTROM, ELAINE  TAPESTRY OF DARK SOULS [@RAVENLOFT]  P
BERK, HOWARD  SUN GROWS COLD, THE  H,P
BERKE, HELEN  CAPTAIN MIDNIGHT AND SHEIK JOMAK KHAN  BLB
BERKLEY BOOKS  BEYOND  PA,C/PA
BERLINER, JANET/GUTHRIDGE, GEORGE  CHILD OF THE JOURNEY [MADAGASCAR MANIFESTO 2]  P
BERLINER, JANET/GUTHRIDGE, GEORGE  CHILD OF THE LIGHT [MADAGASCAR MANIFESTO 1]  P
BERLINER, JANET/GUTHRIDGE, GEORGE  CHILDREN OF THE DUSK [MADAGASCAR MANIFESTO 3]  P
BERLINER, JANET/GUTHRIDGE, GEORGE  MADAGASCAR MANIFESTO, THE [MADAGASCAR MANIFESTO 1,2,3]  LP
BERLITZ, CHARLES  WITHOUT A TRACE  [C/RAN-P]
BERLYN, MICHAEL  CRYSTAL PHOENIX  P
BERLYN, MICHAEL  ETERNAL ENEMY, THE  H,SFWA
BERLYN, MICHAEL  INTEGRATED MAN, THE  C/P,P
BERMAN, JUDITH  BEAR DAUGHTER  LP
BERMAN, MITCH  TIME CAPSULE  P
BERNA, PAUL  THRESHOLD OF THE STARS  SR-P
BERNARD, CHERYL  TURNING ON THE GIRLS  H
BERNARD, CHRISTINE  FONTANA BOOK OF GREAT HORROR STORIES, THE  SR-PA
BERNARD, CHRISTINE  SECOND FONTANA BOOK OF GREAT HORROR STORIES, THE  C/PA,PA
BERNARD, CHRISTINE  THIRD FONTANA BOOK OF GREAT HORROR STORIES, THE  PA
BERNARD, CHRISTINE  FOURTH FONTANA BOOK OF GREAT HORROR STORIES, THE  PA
BERNARD, D. V.  GOD IN THE IMAGE OF WOMAN  LP
BERNARD, JOEL  THINKING MACHINE AFFAIR, THE [@MAN FROM UNCLE #21]  C/P,P
BERNARD, RAFE  ARMY OF THE UNDEAD [INVADERS 3]  P
BERNARD, RAFE  WHEEL IN THE SKY, THE  SR-P
BERNARD, RAYMOND  HOLLOW EARTH, THE  [C/RAN-P:2]
BERNE, ERIC  GAMES PEOPLE PLAY  [C/RAN-H]

BERNHARD, HUBERT J./ BENNETT, DOROTHY A./RICE, HUGH S.

NEW HANDBOOK OF THE HEAVENS

[C/RAN-P:2]

BERNING, FRANK W.  KAMPF IM UTOPIA  FP
BERNING, FRANK W.  ...NORDPOL RUFT ATLANTIS...  FP
BERNING, FRANK W.  SYSTEM 2000 — 3. PLANET  FP

The MITSFS Pinkdex  September 24, 2007  Page 39
AUTHOR  TITLE  CODE
BERNSTEIN, MOREY  SEARCH FOR BRIDEY MURPHY, THE  [C/RAN-P]
BERRY, ADRIAN  GIANT LEAP, THE  R/REF-H
BERRY, ADRIAN  LABYRINTH OF LIES  H
BERRY, ADRIAN  NEXT TEN THOUSAND YEARS, THE  P
BERRY, BRYAN  AND THE STARS REMAIN  SR-P
BERRY, BRYAN  BORN IN CAPTIVITY  H
BERRY, BRYAN  DREAD VISITOR  SR-P
BERRY, BRYAN  FROM WHAT FAR STAR?  H
BERRY, BRYAN  RETURN TO EARTH  SR-P
BERRY, BRYAN  VENOM-SEEKERS, THE  SR-P
BERRY, JAMES R.  GALACTIC INVADERS  SR-P
BERRY, JAMES R.  QUAS STARBRITE  C/P,P
BERRY, STEPHEN AMES  AI WAR, THE [BIOFAB 3]  SFWA-P
BERRY, STEPHEN AMES  BATTLE FOR TERRA TWO, THE [BIOFAB 2]  C/P,P
BERRY, STEPHEN AMES  BIOFAB WAR, THE [BIOFAB 1]  P
BERRY, STEPHEN AMES  FINAL ASSAULT [BIOFAB 4]  P,SFWA-P
BERTASO, JOSE  ANTOLOGIA DE FICCAO CIENTIFICA 1  FP
BERTASO, JOSE  ANTOLOGIA DE FICCAO CIENTIFICA 2  FP
BERTASO, JOSE  ANTOLOGIA DE FICCAO CIENTIFICA 3  FP
BERTIN, JACK  BROOD OF HELIOS  H
BERTIN, JACK  INTERPLANETARY ADVENTURERS, THE  H
BERTIN, JACK  PYRAMIDS FROM SPACE, THE  C/P,P
BERTIN, JOANNE  DRAGON AND PHOENIX [DRAGONLORD LINDEN 2]  H:2
BERTIN, JOANNE  DRAGON AND PHOENIX (PROOF) [DRAGONLORD LINDEN 2]  SR-LP
BERTIN, JOANNE  LAST DRAGONLORD, THE [DRAGONLORD LINDEN 1]  H,P
BERTONI, ALFIO/MISSIAJA, GIANLUIGI  CATALOGO GENERALE DELLA FANTASCIENZA  R/IND
BES SHAHAR, ELUKI  ARCHANGEL BLUES [@HELLFLOWER #3]  P
BES SHAHAR, ELUKI  BUTTERFLY AND HELLFLOWER [BUTTERFLY ST. CYR 1,2,3]  H
BES SHAHAR, ELUKI  DARKTRADERS [HELLFLOWER 2]  C/P:2,P
BES SHAHAR, ELUKI  HELLFLOWER [HELLFLOWER 1]  C/P:2,SFWA-P:2,P
BESAW, VICTOR  ALIEN, THE  P
BESAW, VICTOR  SWORD OF SHANDAR, THE  C/P,P
BESHER, ALEXANDER  CHI [VIRTUAL REALITY 3]  H
BESHER, ALEXANDER  MIR [VIRTUAL REALITY 2]  H
BESHER, ALEXANDER  RIM [VIRTUAL REALITY 1]  P
BESSON, LUC  ARTHUR AND THE FORBIDDEN CITY  LP
BEST, HERBERT  TWENTY-FIFTH HOUR, THE  H
BEST, STAN  TIME IS A SHADOW  P
BESTER, ALFRED  COMPUTER CONNECTION, THE  C/P,P,H
BESTER, ALFRED  DARK SIDE OF THE EARTH, THE  C/P:2,P
BESTER, ALFRED  DECEIVERS, THE  H:2,SFWA-P,LP
BESTER, ALFRED  DEMOLISHED MAN, THE  H:2,LP
BESTER, ALFRED  EXTRO  P
BESTER, ALFRED  GOLEM 100  C/P,P,H
BESTER, ALFRED  HOMEM DEMOLIDO, O  FH
BESTER, ALFRED/ZELAZNY, ROGER  PSYCHOSHOP  H:2,LP
BESTER, ALFRED  REDEMOLISHED  P
BESTER, ALFRED  STAR LIGHT, STAR BRIGHT  H
BESTER, ALFRED  STARBURST  C/P:2,P
BESTER, ALFRED  STARLIGHT  H:3,P
BESTER, ALFRED  STARS MY DESTINATION, THE  C/P,H:2,P,SFWA
BESTER, ALFRED  VIRTUAL UNREALITIES  H,LP
BETANCOURT, JOHN GREGORY  BLIND ARCHER, THE  C/P,P
BETANCOURT, JOHN GREGORY  CHAOS AND AMBER [ROGER ZELAZNY’S DAWN OF AMBER 2]  H
BETANCOURT, JOHN GREGORY  DAWN OF AMBER, THE [ROGER ZELAZNY’S DAWN OF AMBER 1]  H:2
BETANCOURT, JOHN GREGORY  DRAGON SORCERER, THE  P
BETANCOURT, JOHN GREGORY  DRAGONS OF KOMAKO, THE [DR. BONES]  P
BETANCOURT, JOHN GREGORY  GATES OF HADES, THE  P
BETANCOURT, JOHN GREGORY  HAG’S CONTRACT [BIRTHRIGHT 4]  P
BETANCOURT, JOHN GREGORY  HEART OF THE WARRIOR, THE [STAR TREK DS9]  C/P:2,P
BETANCOURT, JOHN GREGORY  INCIDENT AT ARBUK [STAR TREK VOYAGER]  P
BETANCOURT, JOHN GREGORY  INFECTION [@STAR TREK TNG DOUBLE HELIX 1]  P
BETANCOURT, JOHN GREGORY  JOHNNY ZED  C/P,P

BETANCOURT, JOHN GREGORY/ BUSHYAGER, LINDA E.

PACIFICA

LP

BETANCOURT, JOHN GREGORY  PLAYING IN WONDERLAND  LP
BETANCOURT, JOHN GREGORY  REMEMORY  C/P,P
BETANCOURT, JOHN GREGORY  ROGUE PIRATE  P
BETANCOURT, JOHN GREGORY  SHADOWS OF AMBER [@ROGER ZELAZNY’S DAWN OF AMBER 4]  H
BETANCOURT, JOHN GREGORY  SWASHBUCKLING EDITOR STORIES  LP
BETANCOURT, JOHN GREGORY  TO RULE IN AMBER [ROGER ZELAZNY’S DAWN OF AMBER 3]  H
BETANCOURT, JOHN GREGORY  VENGEANCE OF HERA, THE [@HERCULES]  P
BETANCOURT, JOHN GREGORY  WRATH OF POSEIDON, THE [@HERCULES]  P
BETHANCOURT, T. ERNESTO  DOG DAYS OF ARTHUR CANE, THE  P

AUTHOR  TITLE  CODE
BETHANCOURT, T. ERNESTO  MORTAL INSTRUMENTS, THE  P
BETHANCOURT, T. ERNESTO  TUNE IN YESTERDAY  P
BETHKE, BRUCE  HEADCRASH  H,P
BETHKE, BRUCE  MAVERICK [ROBOTS AND ALIENS 5]  P,SFWA-P:2
BETHKE, BRUCE/DAY, VOX  REBEL MOON  C/P,P
BEVIS, H. U.  ALIEN ABDUCTORS, THE  H
BEVIS, H. U.  TO LUNA WITH LOVE  H
BEYER, KIRSTEN  FUSION [STAR TREK VOYAGER STRING THEORY 2]  P
BEYNON, JOHN  STOWAWAY TO MARS  SR-P,C/P:2
BHAKTIVEDANTA BOOK TRUST  COMING BACK: THE SCIENCE OF REINCARNATION  [C/RAN-P:2]
BHATT, JACK  ODYSSEY OF THE DAMNED  P
BICK, ILSA J.  BLOOD AVATAR [@MECHWARRIOR DARK AGE/BATTLETECH]  P
BICK, ILSA J.  DAUGHTER OF THE DRAGON [@MECHWARRIOR DARK AGE/BATTLETECH]  P
BICK, ILSA J.  DRAGON RISING [@MECHWARRIOR DARK AGE/BATTLETECH]  P
BICK, ILSA J.  WELL OF SOULS [@STAR TREK LOST ERA (2336 AD)]  P
BICKERTON, DEREK  KING OF THE SEA  P
BICKHAM, JACK M.  ARIEL  P,SFWA-P
BICKHAM, JACK M.  DAY SEVEN  P
BIDMEAD, CHRISTOPHER H.  CASTROVALVA [DOCTOR WHO]  SR-P,C/P
BIDMEAD, CHRISTOPHER H.  FRONTIOS [DOCTOR WHO]  SR-P,P
BIDMEAD, CHRISTOPHER H.  LOGOPOLIS [DOCTOR WHO]  SR-P,C/P:2
BIEMILLER, CARL L.  FOLLOW THE WHALES  SFWA
BIEMILLER, CARL L.  HYDRONAUTS, THE  H
BIERCE, AMBROSE  COMPLETE SHORT STORIES OF AMBROSE BIERCE, VOL. 1, THE  P
BIERCE, AMBROSE  COMPLETE SHORT STORIES OF AMBROSE BIERCE, VOL. 2, THE  P
BIERCE, AMBROSE  DEVIL’S DICTIONARY, THE  H:2,LP
BIG FINISH  DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME TWO [DOCTOR WHO]  HA
BIG FINISH  DOCTOR WHO: THE AUDIO SCRIPTS, VOLUME THREE [DOCTOR WHO]  HA
BIGGLE, LLOYD JR.  ALL THE COLORS OF DARKNESS [JAN DARZEK]  C/P,P,H:2
BIGGLE, LLOYD JR.  ANGRY ESPERS, THE  DD485
BIGGLE, LLOYD JR.  FURY OUT OF TIME, THE  C/P:2,P
BIGGLE, LLOYD JR.  GALAXY OF STRANGERS, A  H,SFWA
BIGGLE, LLOYD JR.  LIGHT THAT NEVER WAS, THE  C/P:2,SFWA,P,H
BIGGLE, LLOYD JR.  METALLIC MUSE, THE  H:2,SFWA,P
BIGGLE, LLOYD JR.  MONUMENT  H:2,P,SFWA
BIGGLE, LLOYD JR.  NEBULA AWARD STORIES 7  C/PA,PA
BIGGLE, LLOYD JR.  RULE OF THE DOOR AND OTHER FANCIFUL REGULATIONS, THE  C/P,P,H
BIGGLE, LLOYD JR.  SILENCE IS DEADLY [JAN DARZEK]  H:2,SFWA
BIGGLE, LLOYD JR.  SILENT SKY, THE  P
BIGGLE, LLOYD JR.  STILL, SMALL VOICE OF TRUMPETS, THE [CULTURAL SURVEY 1]  C/P,P,H:2
BIGGLE, LLOYD JR.  THIS DARKENING UNIVERSE [JAN DARZEK]  H,SFWA
BIGGLE, LLOYD JR.  TIJDCAPSULE, DE  FP
BIGGLE, LLOYD JR.  TUNESMITH, THE  SFWA-TD27,TD27
BIGGLE, LLOYD JR.  WATCHERS OF THE DARK [JAN DARZEK]  H:2
BIGGLE, LLOYD JR.  WHIRLIGIG OF TIME, THE [JAN DARZEK]  H
BIGGLE, LLOYD JR.  WORLD MENDERS, THE [CULTURAL SURVEY 2]  C/P:2,SFWA,P
BIGHAM, MADGE A.  MERRY ANIMAL TALES  H
BILENKIN, DMITRI  UNCERTAINTY PRINCIPLE, THE  H,LP
BILL, ALFRED H.  WOLF IN THE GARDEN, THE  C/P,P
BILLIAS, STEPHEN  AMERICAN BOOK OF THE DEAD, THE  P
BILLIAS, STEPHEN  DERYNI CHALLENGE [DERYNI]  P,SFWA-P
BILLIAS, STEPHEN  HOLO MEN [CYBERPUNK 2.0.2.0.]  P
BILLIAS, STEPHEN  HORRIBLE HUMES [RUNESWORD 4]  C/P,P
BILLIAS, STEPHEN  QUEST FOR THE 36, THE  C/P,P
BILLIAS, STEPHEN  RAVENGERS, THE [CYBERPUNK 2.0.2.0.]  P
BILLOT, MARCEL/WELLS, F. V.  PERFUMERY TECHNOLOGY  [C/RAN-H]
BILLS, RANDALL N.  HERETIC’S FAITH [@MECHWARRIOR DARK AGE/BATTLETECH]  P
BILLS, RANDALL N.  HUNTERS OF THE DEEP [@MECHWARRIOR DARK AGE/BATTLETECH]  P
BILLS, RANDALL N.  IMMINENT CRISIS [@MECHWARRIOR]  P
BILLS, RANDALL N.  PATH OF GLORY [@BATTLETECH 49]  P
BILSBOROUGH, DAVID  WANDERER’S TALE, THE [LINDORMYN 1]  H
BILYEU, RICHARD  TANELORN ARCHIVES, THE  R/IND:2
BINDER, EANDO  ADAM LINK — ROBOT  C/P:2,P
BINDER, EANDO  ADAO E EVA ROBOTS  FP
BINDER, EANDO  ANTON YORK, IMMORTAL  C/P:2,P
BINDER, EANDO  ANTON YORK, INMORTAL (SPANISH)  FP
BINDER, EANDO  DOUBLE MAN, THE  C/P:2,P
BINDER, EANDO  ENSLAVED BRAINS  H
BINDER, EANDO  FIVE STEPS TO TOMORROW  C/P,P
BINDER, EANDO  GET OFF MY WORLD  C/P,P
BINDER, EANDO  IMPOSSIBLE WORLD, THE  C/P:2,P
BINDER, EANDO  LORDS OF CREATION  C/P:2,P
BINDER, EANDO  MENACE OF THE SAUCERS  C/P,P
BINDER, EANDO  NIGHT OF THE SAUCERS  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 41
AUTHOR  TITLE  CODE
BINDER, EANDO  PUZZLE OF THE SPACE PYRAMIDS  P
BINDER, EANDO  SECRET OF THE RED SPOT  P

BINDER, OTTO AVENGERS BATTLE THE EARTH-WRECKER, THE [AVENGERS (MARVEL COMICS CHARACTERS)]

C/P,P

BINDER, OTTO  HOSPITAL HORROR, THE  P
BINDER, OTTO  PLANETS  [C/RAN-H]
BINGENHEIMER, HEINZ  TRANSGALAXIS KATALOG 1  R/IND
BINGHAM, CARSON  WAR OF THE CYBERNAUTS, THE [FLASH GORDON 6]  C/P:2,P
BINGHAM, CARSON  WITCH QUEEN OF MONGO, THE [FLASH GORDON 5]  C/P:2,P
BINGS, HENRY  WELTEN IM BRAND  FP
BINNS, JOHN  SHORT TRIPS: A UNIVERSE OF TERRORS [DOCTOR WHO]  HA
BINNS, JOHN  SHORT TRIPS: LIFE SCIENCES [DOCTOR WHO]  HA
BINNS, JOHN  SHORT TRIPS: STEEL SKIES [DOCTOR WHO]  HA
BINNS, JOHN  SHORT TRIPS: 2040 [DOCTOR WHO]  HA
BIRKENHEAD, EARL OF  WORLD IN 2030, THE  SR-H
BIRKIN, CHARLES  MY NAME IS DEATH  P
BIRKIN, CHARLES  SPAWN OF SATAN  P
BIRKIN, CHARLES  WHERE TERROR STALKED  SR-P
BIRKIN, CHARLES  WITCH-BAITER, THE  PA
BIRMINGHAM, JOHN  DESIGNATED TARGETS [AXIS OF TIME 2]  LP
BIRMINGHAM, JOHN  FINAL IMPACT [AXIS OF TIME 3]  LP
BIRMINGHAM, JOHN  WEAPONS OF CHOICE [AXIS OF TIME 1]  LP
BIRTOLO, DYLAN  SHADOW CHASER, THE  LP
BISCHOFF, DAVID F.  ALIEN ISLAND [@SPACE PRECINCT #3]  C/P,P
BISCHOFF, DAVID F.  ANCIENT, THE [@SEAQUEST DSV 3]  C/P,P
BISCHOFF, DAVID F.  BLOB, THE  P
BISCHOFF, DAVID F.  CRUNCH BUNCH, THE  P

BISCHOFF, DAVID F./ MONTELEONE, THOMAS F.

DAY OF THE DRAGONSTAR [DRAGONSTAR 1]

C/P:2,P

BISCHOFF, DAVID F.  DEITY-FATHER, THE [@SPACE PRECINCT #1]  P
BISCHOFF, DAVID F.  DEMON WING [@SPACE PRECINCT #2]  P
BISCHOFF, DAVID F.  DESTINY DICE, THE [GAMING MAGI 1]  C/P:2,P

BISCHOFF, DAVID F./ MONTELEONE, THOMAS F.

DRAGONSTAR DESTINY [DRAGONSTAR 3]

C/P,P

BISCHOFF, DAVID F.  GALACTIC WARRIORS [STARHOUNDS 2]  C/P:2,P
BISCHOFF, DAVID F.  GENOCIDE [@ALIENS 4]  P
BISCHOFF, DAVID F.  GREMLINS 2 [GREMLINS]  P
BISCHOFF, DAVID F.  GROUNDED [STAR TREK TNG]  C/P:2,P
BISCHOFF, DAVID F.  HACKERS  P
BISCHOFF, DAVID F.  H.P. LOVECRAFT INSTITUTE, THE  H
BISCHOFF, DAVID F.  HUNTER’S PLANET [@ALIENS VS. PREDATOR 2]  C/P,P
BISCHOFF, DAVID F.  INFINITE BATTLE, THE [STARHOUNDS 1]  C/P:2,P
BISCHOFF, DAVID F.  JACK LONDON, STAR WARRIOR  LP
BISCHOFF, DAVID F.  J.R.R. TOLKIEN UNIVERSITY  LP
BISCHOFF, DAVID F.  MACROCOSMIC CONFLICT, THE [STARHOUNDS 3]  C/P,P
BISCHOFF, DAVID F.  MANDALA  C/P:2,P

BISCHOFF, DAVID F./ MONTELEONE, THOMAS F.

NIGHT OF THE DRAGONSTAR [DRAGONSTAR 2]

C/P:2,P

BISCHOFF, DAVID F.  NIGHT OF THE LIVING SHARK [MELVINGE 1]  P
BISCHOFF, DAVID F.  NIGHTWORLD  C/P:2,P

BISCHOFF, DAVID F./BROWN, RICH/ RICHARDSON, LINDA PERSONAL DEMON, A

C/P:2,P

BISCHOFF, DAVID F.  PHILIP K. DICK HIGH  LP
BISCHOFF, DAVID F.  SEARCH FOR DINOSAURS [@TIME MACHINE #2]  P

BISCHOFF, DAVID F./ LAMPTON, CHRISTOPHER SEEKER, THE

SR-P,C/P

BISCHOFF, DAVID F.  SHIP OF GHOSTS [@FARSCAPE]  C/P,P
BISCHOFF, DAVID F.  STAR FALL  C/P:2,P
BISCHOFF, DAVID F.  STAR SPRING  C/P:2,P
BISCHOFF, DAVID F./BAILEY, DENNIS R.  TIN WOODMAN  C/P,SFWA,P,H
BISCHOFF, DAVID F.  UNICORN GAMBIT, THE [GAMING MAGI 3]  C/P,P
BISCHOFF, DAVID F.  VAMPIRES OF NIGHTWORLD  C/P,P
BISCHOFF, DAVID F.  WARGAMES  C/P:2,P
BISCHOFF, DAVID F.  WRAITH BOARD [GAMING MAGI 2]  C/P,P
BISHOP, ANNE  BELLADONNA [EPHEMERA]  H
BISHOP, ANNE  BLACK JEWELS TRILOGY, THE [BLACK JEWELS 1,2,3]  LP
BISHOP, ANNE  DAUGHTER OF THE BLOOD [BLACK JEWELS 1]  C/P,P
BISHOP, ANNE  DREAMS MADE FLESH  LP
BISHOP, ANNE  HEIR TO THE SHADOWS [BLACK JEWELS 2]  C/P,P
BISHOP, ANNE  HOUSE OF GAIAN [TIR ALAINN 3]  P
BISHOP, ANNE  INVISIBLE RING, THE [BLACK JEWELS]  C/P,P
BISHOP, ANNE  PILLARS OF THE WORLD, THE [TIR ALAINN 1]  C/P,P
BISHOP, ANNE  QUEEN OF THE DARKNESS [BLACK JEWELS 3]  C/P,P
BISHOP, ANNE  SEBASTIAN [EPHEMERA]  H

AUTHOR  TITLE  CODE
BISHOP, ANNE  SHADOWS AND LIGHT [TIR ALAINN 2]  C/P,P
BISHOP, DAVID  AMORALITY TALE [DOCTOR WHO PAST DOCTORS]  P
BISHOP, DAVID  BAD MOON RISING [@JUDGE DREDD]  P
BISHOP, DAVID  DOMINO EFFECT, THE [DOCTOR WHO EIGHTH DOCTOR]  P
BISHOP, DAVID  EMPIRE OF DEATH [DOCTOR WHO PAST DOCTORS]  P
BISHOP, DAVID  KINGDOM OF THE BLIND [@JUDGE DREDD]  P
BISHOP, DAVID  MURDER IN MARIENBURG, A [@WARHAMMER]  P
BISHOP, DAVID  SAVAGE AMUSEMENT, THE [JUDGE DREDD]  P
BISHOP, DAVID  STRANGELOVE GAMBIT, THE [@NIKOLAI DANTE]  P
BISHOP, GEORGE  SHUTTLE PEOPLE, THE  P
BISHOP, GERALD  NEW SCIENCE FICTION BOOKS PUBLISHED IN GREAT BRITAIN 1968, 1969  R/IND-F
BISHOP, K. J.  ETCHED CITY, THE  LP
BISHOP, MICHAEL  ANCIENT OF DAYS  H,P
BISHOP, MICHAEL  AND STRANGE AT ECBATAN THE TREES  H:2,SFWA,SFWA-TD16
BISHOP, MICHAEL  AT THE CITY LIMITS OF FATE  SR-LP
BISHOP, MICHAEL  BENEATH THE SHATTERED MOONS  P
BISHOP, MICHAEL  BLOODED ON ARACHNE  C/P,P
BISHOP, MICHAEL  BRITTLE INNINGS  H,P
BISHOP, MICHAEL  CATACOMB YEARS  H,SFWA,P
BISHOP, MICHAEL/WATSON, IAN  CHANGES  PA
BISHOP, MICHAEL  CLOSE ENCOUNTERS WITH THE DEITY  LP
BISHOP, MICHAEL  COUNT GEIGER’S BLUES  H,LP
BISHOP, MICHAEL  EYES OF FIRE  C/P,P
BISHOP, MICHAEL  EYES OF FIRE (PROOF)  SR-LP
BISHOP, MICHAEL  FUNERAL FOR THE EYES OF FIRE, A  C/P:2,P
BISHOP, MICHAEL  LITTLE KNOWLEDGE, A  H,P
BISHOP, MICHAEL  NEBULA AWARDS 23  HA
BISHOP, MICHAEL  NEBULA AWARDS 24  SFWA,LPA
BISHOP, MICHAEL  NEBULA AWARDS 25  LPA
BISHOP, MICHAEL  NO ENEMY BUT TIME  SFWA-P,H:2,P,SFWA
BISHOP, MICHAEL  PHILIP K. DICK IS DEAD, ALAS  LP
BISHOP, MICHAEL  SECRET ASCENSION, THE  C/P,SFWA-P,P,H
BISHOP, MICHAEL  STOLEN FACES  P
BISHOP, MICHAEL  TIME PIECES  LP
BISHOP, MICHAEL  TRANSFIGURATIONS  H:2,SFWA,P
BISHOP, MICHAEL  TRANSFIGURATIONS (PROOF)  SR-LP
BISHOP, MICHAEL/WATSON, IAN  UNDER HEAVEN’S BRIDGE  C/P:2,P
BISHOP, MICHAEL  UNICORN MOUNTAIN  H,P
BISHOP, TOBY  AIRS BENEATH THE MOON  P
BISSON, TERRY  BEARS DISCOVER FIRE  H
BISSON, TERRY  DEAR ABBEY  SR-LP
BISSON, TERRY  FIFTH ELEMENT, THE  P
BISSON, TERRY  FIRE ON THE MOUNTAIN  H,P
BISSON, TERRY  GALAXY QUEST  P
BISSON, TERRY  GREETINGS  H
BISSON, TERRY  IN THE UPPER ROOM AND OTHER LIKELY STORIES  H
BISSON, TERRY  JOHNNY MNEMONIC  P
BISSON, TERRY  NUMBERS DON’T LIE  LP
BISSON, TERRY  PICKUP ARTIST, THE  H
BISSON, TERRY  PIRATES OF THE UNIVERSE  H,LP
BISSON, TERRY  TALKING MAN  H,SFWA,P
BISSON, TERRY  VIRTUOSITY  P
BISSON, TERRY  VOYAGE TO THE RED PLANET  C/P:2,SFWA,P,H
BISSON, TERRY  WYRLDMAKER  C/P:2,P
BIXBY, E. REW  FIRE THRONE MOUNTAIN  P
BIXBY, JEROME  CALL FOR AN EXORCIST  SR-L
BIXBY, JEROME  MIRROR, MIRROR [STAR TREK]  SR-S
BIXBY, JEROME  SPACE BY THE TALE  C/P:2,P
BJORKLUND, CHARLES R.  STAR WARS INTERGALACTIC PASSPORT, THE  R/REF-P
BLACK, CAMPBELL  ASTERISK DESTINY  P
BLACK, CAMPBELL  BRAINFIRE  P
BLACK, CAMPBELL  RAIDERS OF THE LOST ARK  C/P,P
BLACK, IAN STUART  DOCTOR WHO: THE WAR MACHINES [DOCTOR WHO]  P
BLACK, IAN STUART  MACRA TERROR, THE [DOCTOR WHO]  SR-P
BLACK, IAN STUART  SAVAGES, THE [DOCTOR WHO]  SR-P
BLACK, SHEILA  STARGATE (ADAPTED FOR YOUNGER READERS) [STARGATE]  LP
BLACKBURN, JOHN  BURY HIM DARKLY  H
BLACKBURN, JOHN  CHILDREN OF THE NIGHT  P
BLACKBURN, JOHN  FOR FEAR OF LITTLE MEN  SR-P
BLACKBURN, JOHN  RELUCTANT SPY, THE  P
BLACKBURN, JOHN  SCENT OF NEW-MOWN HAY, A  H:2
BLACKFORD, RUSSELL  DARK FUTURES [@TERMINATOR 2 NEW JOHN CONNOR 1]  P
BLACKFORD, RUSSELL  EVIL HOUR, AN [@TERMINATOR 2 NEW JOHN CONNOR #2]  P
BLACKFORD, RUSSELL  TEMPTING OF THE WITCH KING, THE  SR-P

The MITSFS Pinkdex  September 24, 2007  Page 43
AUTHOR  TITLE  CODE
BLACKFORD, RUSSELL  TIMES OF TROUBLE [@TERMINATOR 2 NEW JOHN CONNOR 3]  P
BLACKHALL, RAY  STEALTH PLANET  H,LP
BLACKHALL, RAY  STEALTH PLANET (PROOF)  H,SR-LP
BLACKSTON, DANIEL E.  LORDS OF SWORDS  LPA
BLACKSTON, DANIEL E.  SAGES & SWORDS  LPA
BLACKWOOD, ALGERNON  ANCIENT SORCERIES  VLP
BLACKWOOD, ALGERNON  BEST GHOST STORIES OF ALGERNON BLACKWOOD  LP
BLACKWOOD, ALGERNON  BEST SUPERNATURAL TALES OF ALGERNON BLACKWOOD, THE  H
BLACKWOOD, ALGERNON  DOLL AND ONE OTHER, THE  H
BLACKWOOD, ALGERNON  DUDLEY AND GILDEROY  C/P,P,H
BLACKWOOD, ALGERNON  EPISODES BEFORE THIRTY  SR-H
BLACKWOOD, ALGERNON  EXTRA DAY, THE  SR-H
BLACKWOOD, ALGERNON  HUMAN CHORD, THE  SR-H
BLACKWOOD, ALGERNON  PRISONER IN FAIRYLAND, A  SR-H
BLACKWOOD, ALGERNON  TALES OF THE MYSTERIOUS AND MACABRE  H
BLACKWOOD, ALGERNON  TALES OF THE UNCANNY AND SUPERNATURAL  H

BLACKWOOD, ALGERNON/ WILSON, WILFRED WOLVES OF GOD AND OTHER FEY STORIES, THE

SR-H

BLACKWOOD, GARY L.  BEYOND THE DOOR  SFWA
BLAINE, JOHN  BLUE GHOST MYSTERY, THE [@RICK BRANT #15]  H
BLAINE, JOHN  CAVES OF FEAR, THE [@RICK BRANT #8]  H
BLAINE, JOHN  DANGER BELOW! [@RICK BRANT #23]  H
BLAINE, JOHN  DEADLY DUTCHMAN, THE [@RICK BRANT #22]  H
BLAINE, JOHN  EGYPTIAN CAT MYSTERY, THE [@RICK BRANT #16]  H:2
BLAINE, JOHN  ELECTRONIC MIND READER, THE [@RICK BRANT #12]  H:2
BLAINE, JOHN  FLAMING MOUNTAIN, THE [@RICK BRANT #17]  H
BLAINE, JOHN  FLYING STINGAREE, THE [@RICK BRANT #18]  H
BLAINE, JOHN  GOLDEN SKULL, THE [@RICK BRANT #10]  H:2
BLAINE, JOHN  LOST CITY, THE [@RICK BRANT #2]  H:2
BLAINE, JOHN  100 FATHOMS UNDER [@RICK BRANT #4]  H
BLAINE, JOHN  PHANTOM SHARK, THE [@RICK BRANT #6]  H:2
BLAINE, JOHN  PIRATES OF SHAN, THE [@RICK BRANT #14]  H
BLAINE, JOHN  ROCKET JUMPER [@RICK BRANT #21]  H
BLAINE, JOHN  ROCKET’S SHADOW, THE [@RICK BRANT #1]  H:2
BLAINE, JOHN  RUBY RAY MYSTERY, THE [@RICK BRANT #19]  H
BLAINE, JOHN  SCARLET LAKE MYSTERY, THE [@RICK BRANT #13]  H
BLAINE, JOHN  SEA GOLD [@RICK BRANT #3]  H
BLAINE, JOHN  SMUGGLER’S REEF [@RICK BRANT #7]  H
BLAINE, JOHN  STAIRWAY TO DANGER [@RICK BRANT #9]  H:2
BLAINE, JOHN  VEILED RAIDERS, THE [@RICK BRANT #20]  H
BLAINE, JOHN  WAILING OCTOPUS, THE [@RICK BRANT #11]  H
BLAINE, JOHN  WHISPERING BOX MYSTERY, THE [@RICK BRANT #5]  H:2
BLAIR, ADRIAN  COSMIC CONQUEST  SR-P
BLAIR, ANNETTE  SCOT, THE WITCH AND THE WARDROBE, THE  P
BLAIR, CYNTHIA  CURSE, THE [@DARK MOON #1]  P
BLAIR, JOHN  BRIGHT ANGEL  P
BLAIR, JOHN  LANDSCAPE OF DARKNESS, A  C/P,P
BLAIR, KARIN  MEANING IN STAR TREK  R/REF-P,C/REF-P:2
BLAIR, WILLIAM  FIRE! SURVIVAL AND PREVENTION  [C/RAN-P]
BLAKE, KATHERINE  INTERIOR LIFE, THE  C/P:2,P
BLAKE, STERLING  CHILLER  H
BLAKE, SUSAN  HAUNTED DOLLHOUSE, THE [TWILIGHT #22]  P
BLAKENEY, JAY D.  CHILDREN OF ANTHI, THE [ANTHI]  P
BLAKENEY, JAY D.  GODA WAR, THE  C/P,P
BLAKENEY, JAY D.  REQUIEM FOR ANTHI [ANTHI]  P
BLANK, CLAIR  ADVENTURE GIRLS IN THE AIR, THE  H
BLANKENSHIP, WILLIAM D.  HELIX FILE, THE  H
BLANPIED, PAMELA WHARTON  DRAGONS  C/P,P
BLASHFIELD, JEAN  VILLAINS OF VOLTURNUS [ENDLESS QUEST 8]  P
BLATTY, WILLIAM PETER  EXORCIST, THE [EXORCIST 1]  C/P:2,P
BLATTY, WILLIAM PETER  LEGION [EXORCIST 2]  H:2
BLAYLOCK, JAMES P.  ALL THE BELLS ON EARTH  P
BLAYLOCK, JAMES P.  DIGGING LEVIATHAN, THE  P
BLAYLOCK, JAMES P.  DISAPPEARING DWARF, THE  C/P,P
BLAYLOCK, JAMES P.  ELFIN SHIP, THE  C/P:2,P
BLAYLOCK, JAMES P.  HOMUNCULUS  P
BLAYLOCK, JAMES P.  LAND OF DREAMS  C/P,SFWA,P,H
BLAYLOCK, JAMES P.  LAST COIN, THE  C/P,SFWA-P,P,H
BLAYLOCK, JAMES P.  LORD KELVIN’S MACHINE  P,SFWA-P:2
BLAYLOCK, JAMES P.  NIGHT RELICS  SFWA
BLAYLOCK, JAMES P.  PAPER DRAGONS  B
BLAYLOCK, JAMES P.  PAPER GRAIL, THE  C/P:2,SFWA,P,H
BLAYLOCK, JAMES P.  STONE GIANT, THE  SFWA-P
BLAYLOCK, JAMES P.  WINTER TIDES  P

The MITSFS Pinkdex  September 24, 2007  Page 44
AUTHOR  TITLE  CODE
BLEILER, EVERETT F./DIKTY, T. E.  BEST SCIENCE FICTION STORIES, 1949, THE  HA
BLEILER, EVERETT F./DIKTY, T. E.  BEST SCIENCE FICTION STORIES, 1951, THE  HA
BLEILER, EVERETT F./DIKTY, T. E.  BEST SCIENCE FICTION STORIES, 1952, THE  HA
BLEILER, EVERETT F.  CHECKLIST OF FANTASTIC LITERATURE, THE  SRL/IND

BLEILER, EVERETT F.

CHECKLIST OF SCIENCE-FICTION AND SUPERNATURAL FICTION 1800–1948, THE

R/IND

BLEILER, EVERETT F.  EIGHT DIME NOVELS  VLPA
BLEILER, EVERETT F.  FIVE VICTORIAN GHOST NOVELS  LPA
BLEILER, EVERETT F./DIKTY, T. E.  FRONTIERS IN SPACE  PA,C/PA
BLEILER, EVERETT F.  GUIDE TO SUPERNATURAL FICTION, THE  R/IND
BLEILER, EVERETT F./DIKTY, T. E.  IMAGINATION UNLIMITED  C/PA,PA
BLEILER, EVERETT F.  SCIENCE-FICTION THE EARLY YEARS  R/IND
BLEILER, EVERETT F.  SCIENCE-FICTION THE GERNSBACK YEARS  R/REF-H
BLEILER, EVERETT F./DIKTY, T. E.  YEAR’S BEST SCIENCE FICTION NOVELS 1952  HA
BLEILER, EVERETT F./DIKTY, T. E.  YEAR’S BEST SCIENCE FICTION NOVELS 1954  HA
BLISH, JAMES  AND ALL THE STARS A STAGE  C/P:2,P,H
BLISH, JAMES  ANYWHEN  H:2,P
BLISH, JAMES  BEST OF JAMES BLISH, THE  C/P,P
BLISH, JAMES  BLACK EASTER  C/P,P,H
BLISH, JAMES  CASE OF CONSCIENCE, A  H:2,P
BLISH, JAMES  CITIES IN FLIGHT [CITIES IN FLIGHT 1,2,3,4]  C/P,H:2,P
BLISH, JAMES  CITIES IN FLIGHT, VOL. 1 [CITIES IN FLIGHT 1,2]  C/P:2,SFWA-P,P
BLISH, JAMES  CITIES IN FLIGHT, VOL. 2 [CITIES IN FLIGHT 3,4]  P
BLISH, JAMES  COMMON TIME  B
BLISH, JAMES  DAY AFTER JUDGMENT, THE  C/P,SFWA,SR-P,H
BLISH, JAMES  DEVIL’S DAY, THE  C/P:2,SFWA-P,P
BLISH, JAMES  DOCTOR MIRABILIS  SR-H,P
BLISH, JAMES/LOWNDES, ROBERT A. W.  DUPLICATED MAN, THE  C/P:2,P
BLISH, JAMES  EARTHMAN, COME HOME  H:2,P
BLISH, JAMES  ESPER  C/P:2,P
BLISH, JAMES  FALLEN STAR  SR-P,C/P:2,H
BLISH, JAMES  FROZEN YEAR, THE  C/P,P,H
BLISH, JAMES  GALACTIC CLUSTER  C/P:3,P
BLISH, JAMES  JACK OF EAGLES  C/P,GN3.19,P,H
BLISH, JAMES  LIFE FOR THE STARS, A  C/P:2,P
BLISH, JAMES  MIDSUMMER CENTURY  H:2,P,SFWA
BLISH, JAMES  MISSION TO THE HEART STARS  C/P:2,P
BLISH, JAMES  NEBULA AWARD STORIES 5  C/PA,PA
BLISH, JAMES  NEW DREAMS THIS MORNING  PA,C/PA
BLISH, JAMES  NIGHT SHAPES, THE  C/P:2,P
BLISH, JAMES  PAIR FROM SPACE: GIANTS IN THE EARTH, A  SR-DB0612
BLISH, JAMES  QUINCUNX OF TIME, THE  C/P:2,P
BLISH, JAMES  SEEDLING STARS, THE  C/P:2,P,H
BLISH, JAMES  SO CLOSE TO HOME  C/P:2,P
BLISH, JAMES  SPOCK MUST DIE! [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR DWELLERS, THE  C/P:2,P
BLISH, JAMES  STAR TREK [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 2 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 3 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 4 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 5 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 6 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 7 [STAR TREK]  C/P:3,P
BLISH, JAMES  STAR TREK 8 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 9 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 10 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK 11 [STAR TREK]  C/P,P,H
BLISH, JAMES/LAWRENCE, J. A.  STAR TREK 12 [STAR TREK]  C/P:2,P
BLISH, JAMES  STAR TREK READER, THE [STAR TREK]  H:2
BLISH, JAMES  STAR TREK READER II, THE [STAR TREK]  H
BLISH, JAMES  STAR TREK READER III, THE [STAR TREK]  H
BLISH, JAMES  STAR TREK READER IV, THE [STAR TREK]  H
BLISH, JAMES/LAWRENCE, J. A.  STAR TREK: THE CLASSIC EPISODES 1 [STAR TREK]  P
BLISH, JAMES  STAR TREK: THE CLASSIC EPISODES 2 [STAR TREK]  P
BLISH, JAMES/LAWRENCE, J. A.  STAR TREK: THE CLASSIC EPISODES 3 [STAR TREK]  P
BLISH, JAMES  TALE THAT WAGS THE GOD, THE  R/REF-H
BLISH, JAMES  TESTAMENT OF ANDROS, THE  P
BLISH, JAMES  THEY SHALL HAVE STARS  C/P:2,P
BLISH, JAMES  TITAN’S DAUGHTER  C/P:2,P
BLISH, JAMES/KNIGHT, NORMAN  TORRENT OF FACES, A  H:2,P
BLISH, JAMES  TRIUMPH OF TIME, THE  C/P:2,P
BLISH, JAMES  VOR  C/P:2,P
BLISH, JAMES  WARRIORS OF DAY, THE  SR-P,C/P:2,GN3.16,P
BLISH, JAMES  WELCOME TO MARS  H,P

AUTHOR  TITLE  CODE
BLISH, JAMES  YEAR 2018!  C/P,P
BLISHEN, EDWARD  SCIENCE FICTION STORIES  LPA
BLOCH, ARTHUR  MURPHY’S LAW  [C/RAN-P]
BLOCH, JOHN W.  WITCH, WITCH, BURNING BRIGHT [SIXTH SENSE 1]  C/P,P
BLOCH, ROBERT  AMERICAN GOTHIC  SFWA-P
BLOCH, ROBERT  ATOMS AND EVIL  C/P:2,P
BLOCH, ROBERT  BEST OF ROBERT BLOCH, THE  C/P:2,P
BLOCH, ROBERT  BITTER ENDS  PA
BLOCH, ROBERT  BOGEY MEN  C/P:2,P
BLOCH, ROBERT  COLD CHILLS  C/P,SFWA,P,H
BLOCH, ROBERT  DRAGONS AND NIGHTMARES  C/P:2,P
BLOCH, ROBERT  EIGHTH STAGE OF FANDOM, THE  R/REF-P
BLOCH, ROBERT  FEAR AND TREMBLING  SFWA-P
BLOCH, ROBERT  FEAR PLANET, THE  SR-H
BLOCH, ROBERT  FEAR TODAY — GONE TOMORROW  P
BLOCH, ROBERT  FINAL RECKONINGS  LP
BLOCH, ROBERT  FIREBUG  SFWA-P
BLOCH, ROBERT  FLOWERS FROM THE MOON AND OTHER LUNACIES  H
BLOCH, ROBERT  HORROR-7  P
BLOCH, ROBERT  IT’S ALL IN YOUR MIND  SR-P
BLOCH, ROBERT/NORTON, ANDRE  JEKYLL LEGACY, THE  P
BLOCH, ROBERT  KIDNAPPER, THE  P,SFWA-P
BLOCH, ROBERT  LADIES’ DAY  SR-DB0080
BLOCH, ROBERT  LORI  SFWA-P
BLOCH, ROBERT  LOST IN TIME AND SPACE WITH LEFTY FEEP  H:2,LP
BLOCH, ROBERT  MIDNIGHT PLEASURES  H,SFWA
BLOCH, ROBERT  MORE NIGHTMARES  P
BLOCH, ROBERT  MYSTERIES OF THE WORM  C/P,P
BLOCH, ROBERT  NIGHT OF THE RIPPER, THE  H,SFWA
BLOCH, ROBERT  NIGHT-WORLD  P,SFWA-P
BLOCH, ROBERT  NIGHTMARES  P
BLOCH, ROBERT  ONCE AROUND THE BLOCH  R/REF-H
BLOCH, ROBERT  OPENER OF THE WAY, THE  SR-H,H
BLOCH, ROBERT  OUT OF MY HEAD  SR-H
BLOCH, ROBERT  PLEASANT DREAMS  P
BLOCH, ROBERT  PSYCHO [PSYCHO 1]  SFWA-P
BLOCH, ROBERT  PSYCHO HOUSE [PSYCHO 3]  SFWA-P
BLOCH, ROBERT  PSYCHO II [PSYCHO 2]  SFWA-P
BLOCH, ROBERT  PSYCHOPATHS  SFWA
BLOCH, ROBERT  SCARF, THE  SR-P
BLOCH, ROBERT  SNEAK PREVIEW  C/P:2,P
BLOCH, ROBERT  STRANGE EONS  P
BLOCH, ROBERT  SUCH STUFF AS SCREAMS ARE MADE OF  C/P,P
BLOCH, ROBERT  TALES IN A JUGULAR VEIN  P
BLOCH, ROBERT  THIS CROWDED EARTH  SR-DB0080
BLOCH, ROBERT  TWILIGHT ZONE THE MOVIE  C/P:2,P
BLOCK, FRANCESCA LIA  ECSTASIA  SFWA
BLOCK, FRANCESCA LIA  PRIMAVERA  P
BLOCK, THOMAS H.  MAYDAY  P
BLOCK, THOMAS H.  ORBIT  H,P
BLOODSTONE, JOHN  GODMAN!  SR-P
BLOOM, BRITTON  MATRIX CUBED [BUCK ROGERS INNER PLANETS 3]  C/P,P
BLOOM, HAROLD P.  FLIGHT TO LUCIFER, THE  H,LP
BLUE, TYSON  UNSEEN KING, THE  R/REF-P
BLUM, JONATHAN/ORMAN, KATE  FALLEN GODS [DOCTOR WHO]  SR-H
BLUM, JONATHAN/BOOTH, RUPERT  PRISONER’S DILEMMA, THE [PRISONER]  LP
BLUM, JONATHAN/ORMAN, KATE  SEEING I [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
BLUM, JONATHAN/ORMAN, KATE  UNNATURAL HISTORY [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
BLUM, JONATHAN/ORMAN, KATE  VAMPIRE SCIENCE [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
BLUM, RALPH  SIMULTANEOUS MAN, THE  C/P:2,P
BLUM, ROBERT S.  GIRL FROM THE EMERALINE ISLAND, THE  C/P:3,P
BLUMGARTEN, JAMES  ASTRONAUT, THE  C/P,P
BLUMLEIN, MICHAEL  MOVEMENT OF MOUNTAINS, THE  H,P
BLUMLEIN, MICHAEL, M. D.  HEALER, THE  H
BLUMRICH, JOSEF  SPACESHIPS OF EZEKIEL, THE  [C/RAN-P:2]
BLY, ROBERT W.  ULTIMATE UNAUTHORIZED STAR TREK QUIZ BOOK, THE [STAR TREK]  [C/RAN-P]
BLY, ROBERT W.  WHY YOU SHOULD NEVER BEAM DOWN IN A RED SHIRT [STAR TREK]  R/REF-P,C/REF-P
BLYTHE, DANIEL  DIMENSION RIDERS, THE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
BLYTHE, DANIEL  INFINITE REQUIEM [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
BOARDMAN, TOM  ABC OF SCIENCE FICTION, AN  PA,C/PA
BOARDMAN, TOM  CONNOISSEUR’S SCIENCE FICTION  SR-PA,C/PA
BOARDMAN, TOM JR.  SCIENCE FICTION STORIES  HA
BODE, N. E.  SLIPPERY MAP, THE (PROOF)  SR-LP
BODE, VAUGHN  BODE’S CARTOON CONCERT  SCX
BODELSEN, ANDERS  FREEZING DOWN  H:2,P

The MITSFS Pinkdex  September 24, 2007  Page 46
AUTHOR  TITLE  CODE
BOGDANOV, ALEXANDER  RED STAR: THE FIRST SOVIET UTOPIA  LP
BOGEN, K. B.  GO QUEST, YOUNG MAN  P
BOGGS, TIMOTHY  EYE OF THE RAM, THE [@HERCULES THE LEGENDARY JOURNEYS]  C/P,P
BOGNER, NORMAN  SNOWMAN  P
BOHNHOFF, MAYA KAATHRYN  CRYSTAL ROSE, THE [MERI 3]  P
BOHNHOFF, MAYA KAATHRYN  MERI, THE [MERI 1]  C/P,SFWA-P,P
BOHNHOFF, MAYA KAATHRYN  SPIRIT GATE, THE  P
BOHNHOFF, MAYA KAATHRYN  TAMINY [MERI 2]  P,SFWA-P
BOILEAU, PIERRE/NARCEJAC, THOMAS  CHOICE CUTS  P
BOK, HANNES  BEYOND THE GOLDEN STAIR  P
BOK, HANNES  SORCERER’S SHIP, THE  C/P:2,P
BOK, HANNES  SPINNER OF SILVER AND THISTLE  SR-H
BOLAND, JOHN  NO REFUGE  SR-P
BOLESNY, DARLENE D.  TRAIL OF DARKNESS  P
BOLIN, M. C.  ARMAGEDDON  P
BOLME, EDWARD  ALABASTER STAFF, THE [@FORGOTTEN REALMS ROGUES]  P
BOLME, EDWARD  BOUND BY IRON [@EBERRON INQUISITIVES]  P
BOLME, EDWARD  ORB OF XORIAT, THE [@EBERRON WAR-TORN #2]  P
BOLME, EDWARD  STEEL THRONE, THE [FIVE RINGS FOUR WINDS 0]  P
BOLTON, JOHANNA M.  ALIEN WITHIN, THE  P
BOLTON, JOHANNA M.  MISSION: TORI  P
BONANNO, MARGARET WANDER  BURNING DREAMS [STAR TREK]  P
BONANNO, MARGARET WANDER  CATALYST OF SORROWS [@STAR TREK LOST ERA (2360 AD)]  P
BONANNO, MARGARET WANDER  DWELLERS IN THE CRUCIBLE [STAR TREK]  C/P:3,P
BONANNO, MARGARET WANDER  OTHERS, THE [OTHERS (BONANNO) 1]  H,SFWA
BONANNO, MARGARET WANDER  OTHERWHERE [OTHERS (BONANNO) 2]  H
BONANNO, MARGARET WANDER  OTHERWISE [OTHERS (BONANNO) 3]  H
BONANNO, MARGARET WANDER  PRETERNATURAL [PRETERNATURAL 1]  C/P,P
BONANNO, MARGARET WANDER  PRETERNATURAL 3 [PRETERNATURAL 3]  H:2
BONANNO, MARGARET WANDER  PRETERNATURAL TOO: GYRE [PRETERNATURAL 2]  H,LP
BONANNO, MARGARET WANDER  PROBE [STAR TREK]  C/P,P,H
BONANNO, MARGARET WANDER  STRANGERS FROM THE SKY [STAR TREK]  C/P:2,P
BONANSINGA, JAY  OBLIVION  SR-H
BOND, EVELYN  DARK SONATA [IRA VEDDER 4]  P
BOND, EVELYN  GIRL FROM NOWHERE, THE [IRA VEDDER 2]  P
BOND, J. HARVEY  OTHER WORLD, THE (BRITISH)  SR-P,C/P
BOND, JONATHAN E./KOKE, JAK  TERMINUS EXPERIMENT, THE [@SHADOWRUN 34]  P
BOND, NANCY  VOYAGE BEGUN, THE  H
BOND, NELSON  EXILES OF TIME  C/P:2,P
BOND, NELSON  MONSTER, THE  B
BOND, NELSON  NIGHTMARES AND DAYDREAMS  SR-H
BOND, NELSON  NO TIME LIKE THE FUTURE  C/P:2,P
BOND, NELSON  REMARKABLE EXPLOITS OF LANCELOT BIGGS: SPACEMAN, THE  H,LP
BOND, NELSON  THIRTY-FIRST OF FEBRUARY, THE  SR-H,H
BOND, SIMON  BIZARRE SIGHTS & ODD VISIONS  [C/RAN-P]
BONE, J. F.  CONFEDERATION MATADOR  LP
BONE, J. F./MYERS, ROY  GIFT OF THE MANTI  SR-P
BONE, J. F.  LANI PEOPLE, THE  C/P,P
BONE, J. F.  LEGACY  SR-P
BONE, J. F.  MEDDLERS, THE  SR-P
BONFIGLIOLI, KYRIL  SF REPRISE 4  PA
BONFIGLIOLI, KYRIL  SF REPRISE 6  PA
BONHAM, FRANK  FOREVER FORMULA, THE  C/P,P,H
BONHAM, FRANK  MISSING PERSONS LEAGUE, THE  P
BONNER, RICHARD  BOY INVENTORS’ FLYING SHIP  H
BONSELLS, WALDEMAR  MAYA THE BEE  H
BONTA, VANNA  FLIGHT  LP
BONTLY, THOMAS  CELESTIAL CHESS  P
BOOK-OF-THE-MONTH CLUB  RADICAL UTOPIAS  HA
BOORMAN, JOHN/STAIR, BILL  ZARDOZ  C/P,P
BOOTHBY, GUY  DR. NIKOLA RETURNS  LP
BOOTHBY, GUY  FAREWELL, NIKOLA  H
BORCHARDT, ALICE  DRAGON QUEEN, THE [TALES OF GUINEVERE 1]  H,P
BORCHARDT, ALICE  NIGHT OF THE WOLF [WOLF RICE 1]  P
BORCHARDT, ALICE  RAVEN WARRIOR, THE [TALES OF GUINEVERE 2]  H
BORCHARDT, ALICE  SILVER WOLF, THE [WOLF RICE 2]  C/P:2,P
BORCHARDT, ALICE  WOLF KING, THE [WOLF RICE 3]  P
BORGES, JORGE LUIS  ALEPH AND OTHER STORIES, THE (1933–1969)  H,P
BORGES, JORGE LUIS  BOOK OF FANTASY, THE  LPA

BORGES, JORGE LUIS/ GUERRERO, MARGARITA BOOK OF IMAGINARY BEINGS, THE

C/P,LP

BORGES, JORGE LUIS  BOOK OF SAND, THE  LP
BORGES, JORGE LUIS  LABYRINTHS  LP
BORGES, JORGE LUIS  LIBRO DE ARENA, EL  FP

The MITSFS Pinkdex  September 24, 2007  Page 47
AUTHOR  TITLE  CODE
BORISOFF, NORMAN  BEWITCHED AND BEWILDERED  P
BORN, FRANZ  JULES VERNE: THE MAN WHO INVENTED THE FUTURE  R/REF-P
BORNEFELD, WILLIAM  TIME AND LIGHT  P
BORODIN, GEORGE  SPURIOUS SUN  H
BORRELLO, ALFRED  H. G. WELLS: AUTHOR IN AGONY  R/REF-H
BORSCH, FRANK  ARK OF THE STARS [@PERRY RHODAN LEMURA #1]  P
BOSSE, MALCOLM  MISTER TOUCH  H
BOSTON, JOHN  NAKED CAME THE SASQUATCH  P
BOSTON, L. M.  CHILDREN OF GREEN KNOWE, THE [GREEN KNOWE 1]  H,LP
BOSTON, L. M.  ENEMY AT GREEN KNOWE, AN [GREEN KNOWE 5]  H,LP
BOSTON, L. M.  RIVER AT GREEN KNOWE, THE [GREEN KNOWE 3]  H,LP
BOSTON, L. M.  STRANGER AT GREEN KNOWE, A [GREEN KNOWE 4]  H,LP
BOSTON, L. M.  TREASURE OF GREEN KNOWE [GREEN KNOWE 2]  H,LP
BOTT, CLAIRE  CLOCKWORK WOMAN, THE [@TIME HUNTER 3]  LP
BOUCHER, ANTHONY/MCCOMAS, J. FRANCIS BEST FROM FANTASY AND SCIENCE FICTION 3, THE  C/PA,PA
BOUCHER, ANTHONY  BEST FROM FANTASY AND SCIENCE FICTION 4, THE  HA,PA
BOUCHER, ANTHONY  BEST FROM FANTASY AND SCIENCE FICTION 5, THE  HA,PA
BOUCHER, ANTHONY  BEST FROM FANTASY AND SCIENCE FICTION 6, THE  C/PA,PA
BOUCHER, ANTHONY  BEST FROM FANTASY AND SCIENCE FICTION 7, THE  C/PA,PA
BOUCHER, ANTHONY  BEST FROM FANTASY AND SCIENCE FICTION 8, THE  C/PA,PA
BOUCHER, ANTHONY  CASE OF THE SOLID KEY, THE  P
BOUCHER, ANTHONY  COMPLEAT BOUCHER, THE  H
BOUCHER, ANTHONY  COMPLEAT WEREWOLF, THE  H:2,P
BOUCHER, ANTHONY  FAR AND AWAY  C/P:2,P
BOUCHER, ANTHONY  NINE TIMES NINE  P
BOUCHER, ANTHONY  ROCKET TO THE MORGUE  SR-P,C/P:2
BOUCHER, ANTHONY  TREASURY OF GREAT SCIENCE FICTION 1, A  HA
BOUCHER, ANTHONY  TREASURY OF GREAT SCIENCE FICTION 2, A  HA
BOUCHER, CHRIS  CORPSE MARKER [DOCTOR WHO PAST DOCTORS]  P
BOUCHER, CHRIS  LAST MAN RUNNING [DOCTOR WHO PAST DOCTORS]  C/P:2,P
BOUCHER, CHRIS  MATCH OF THE DAY [@DOCTOR WHO PAST DOCTORS]  P
BOUCHER, CHRIS  PSI-ENCE FICTION [DOCTOR WHO]  P
BOUCHER-KAPLAN, ALINE  WORLD SPIRITS [KHYREN]  C/P:2,SFWA-P,P
BOULLE, PHILIPPE  HAUNTING THE DEAD [@WORLD OF DARKNESS ORPHEUS]  PA

BOULLE, PHILIPPE MADNESS OF PRIESTS, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #2]

P

BOULLE, PHILIPPE/GRIFFIN, ERIC RED TALONS [WORLD OF DARKNESS WEREWOLF 3]/ FIANNA [WORLD OF DARKNESS WEREWOLF 3]

PA

BOULLE, PHILIPPE WOUNDED KING, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE VICTORIAN AGE #3]

P

BOULLE, PIERRE  DESPERATE GAMES  H
BOULLE, PIERRE  GARDEN ON THE MOON  C/P,P,H
BOULLE, PIERRE  PLANET OF THE APES  C/P,P,H
BOULLE, PIERRE  TIME OUT OF MIND  C/P,P
BOULT, S. KYE  SOLO KILL  C/P,P
BOUNDS, SYDNEY J.  BEST OF SYDNEY J. BOUNDS, VOLUME 1, THE  LP
BOUNDS, SYDNEY J.  BEST OF SYDNEY J. BOUNDS, VOLUME 2, THE  LP
BOUNDS, SYDNEY J.  MOON RAIDERS, THE  SR-P
BOUNDS, SYDNEY J.  ROBOT BRAINS, THE  C/P,P
BOUTON, JOHN BELL  UNCLE SAM’S CHURCH  [C/RAN-H]
BOVA, BEN  ALIEN WITHIN, THE [@VOYAGERS #2]  C/P,SFWA-P,P,H:2
BOVA, BEN  ANALOG 9  HA
BOVA, BEN  ANALOG ANNUAL  PA,C/PA
BOVA, BEN  ANALOG YEARBOOK  HA,LPA
BOVA, BEN  AS ON A DARKLING PLAIN  C/P:2,SFWA-P,H,P,SFWA
BOVA, BEN  ASTEROID WARS, THE [ASTEROID WARS 1,2,3]  H
BOVA, BEN  ASTRAL MIRROR, THE  C/P:2,SFWA-P,P
BOVA, BEN  BATTLE STATION  C/P,P
BOVA, BEN  BEST OF ANALOG, THE  C/PA,LPA
BOVA, BEN/MYRUS, DON  BEST OF OMNI SCIENCE FICTION, THE  VLPA
BOVA, BEN/MYRUS, DON  BEST OF OMNI SCIENCE FICTION NO. 4, THE  VLPA
BOVA, BEN  BROTHERS  H,P
BOVA, BEN  CHALLENGES  H,P
BOVA, BEN  CITY OF DARKNESS  P
BOVA, BEN  COLONY  C/P:2,SFWA-P,P
BOVA, BEN  CYBERBOOKS  C/P,SFWA-P,P,H
BOVA, BEN  DEATH DREAM  C/P,P,H
BOVA, BEN  DUELING MACHINE, THE  SR-H,C/P:2,P
BOVA, BEN  EMPIRE BUILDERS [PRIVATEERS 2]  P
BOVA, BEN  ESCAPE!  SR-H,P
BOVA, BEN  ESCAPE PLUS  C/P,P
BOVA, BEN  EXILED FROM EARTH  H:2,LP
BOVA, BEN  EXILES TRILOGY, THE  C/P,P
BOVA, BEN  FAINT ECHOES, DISTANT STARS  R/REF-H

The MITSFS Pinkdex  September 24, 2007  Page 48
AUTHOR  TITLE  CODE
BOVA, BEN  FLIGHT OF EXILES  H
BOVA, BEN  FORWARD IN TIME  C/P,P,H,SFWA
BOVA, BEN  FUTURE CRIME  C/P:2,P
BOVA, BEN/DICKSON, GORDON R.  GREMLINS GO HOME  C/P:2,P
BOVA, BEN  JUPITER  H,P
BOVA, BEN  KINSMAN  H:2,P
BOVA, BEN  KINSMAN SAGA, THE  C/P,SFWA-P,P,H
BOVA, BEN  MANY WORLDS OF SCIENCE FICTION, THE  SR-HA,HA
BOVA, BEN  MARS [JAMIE WATERMAN 1]  C/P:2,P,H
BOVA, BEN  MAXWELL’S DEMONS  C/P,P,H
BOVA, BEN  MERCURY  H
BOVA, BEN  MILLENNIUM  H:2,SFWA,P
BOVA, BEN  MOONRISE [MOONBASE 1]  C/P,P
BOVA, BEN  MOONWAR [MOONBASE 2]  C/P,P
BOVA, BEN  MULTIPLE MAN, THE  C/P,SFWA-P,P,H
BOVA, BEN  NOTES TO A SCIENCE FICTION WRITER  R/REF-H
BOVA, BEN  ORION [ORION 1]  C/P,SFWA-P:2,P,H
BOVA, BEN  ORION AMONG THE STARS [ORION 4]  C/P,SFWA,P,H
BOVA, BEN  ORION AND THE CONQUEROR [ORION 3]  C/P,P
BOVA, BEN  ORION IN THE DYING TIME [ORION 2]  C/P:2,P,H
BOVA, BEN  OUT OF THE SUN  SR-H,C/P,P
BOVA, BEN  PEACEKEEPERS  C/P:2,SFWA-P,P
BOVA, BEN  POWERSAT  H
BOVA, BEN  PRECIPICE, THE [ASTEROID WARS 1]  H,P
BOVA, BEN  PRIVATEERS [PRIVATEERS 1]  C/P:2,SFWA-P,P,H
BOVA, BEN  PROMETHEANS  C/P:2,P
BOVA, BEN  RETURN TO MARS [JAMIE WATERMAN 2]  C/P,P,H
BOVA, BEN  ROCK RATS, THE [ASTEROID WARS 2]  P
BOVA, BEN  SAM GUNN FOREVER [SAM GUNN]  P
BOVA, BEN  SAM GUNN OMNIBUS, THE [SAM GUNN]  H
BOVA, BEN  SAM GUNN, UNLIMITED [SAM GUNN]  P
BOVA, BEN  SATURN  H
BOVA, BEN  SCIENCE FICTION HALL OF FAME 2A, THE  SFWA,PA,HA
BOVA, BEN  SCIENCE FICTION HALL OF FAME 2B, THE  C/PA,PA,SFWA
BOVA, BEN  SILENT WAR, THE [ASTEROID WARS 3]  P
BOVA, BEN/LEWIS, ANTHONY R.  SPACE TRAVEL  R/REF-H
BOVA, BEN  STAR BROTHERS [@VOYAGERS #3]  C/P:2,SFWA-P,P,H
BOVA, BEN  STAR PEACE: ASSURED SURVIVAL  SFWA
BOVA, BEN  STAR WATCHMAN  C/P:2,P
BOVA, BEN  STARCROSSED, THE  C/P:2,SFWA-P,P
BOVA, BEN  TALES OF THE GRAND TOUR  H
BOVA, BEN  TEST OF FIRE  C/P,P
BOVA, BEN  THROUGH EYES OF WONDER  R/REF-H
BOVA, BEN  THX 1138  C/P:2,P
BOVA, BEN  TITAN  H
BOVA, BEN/AUSTIN, A. J.  TO FEAR THE LIGHT  C/P,P
BOVA, BEN/AUSTIN, A. J.  TO SAVE THE SUN  C/P:2,P
BOVA, BEN/POGUE, BILL  TRIKON DECEPTION, THE  C/P,P,H
BOVA, BEN  TRIUMPH  H,P
BOVA, BEN  TWICE SEVEN  P
BOVA, BEN  VENGEANCE OF ORION [ORION 2]  H:2,P
BOVA, BEN  VENUS  H
BOVA, BEN  VIEWPOINT  R/REF-H
BOVA, BEN  VOYAGERS [@VOYAGERS #1]  C/P:2,SFWA-P,P,H
BOVA, BEN  WATCHMEN, THE  SFWA-P
BOVA, BEN  WEATHERMAKERS, THE  C/P:2,P
BOVA, BEN  WHEN THE SKY BURNED  H,SFWA,P
BOVA, BEN  WINDS OF ALTAIR, THE  C/P:2,P,LP,H
BOWEN, CARL  DAY DARK AS NIGHT, A [@EXALTED #1]  P
BOWEN, CARL  MAGE [@WORLD OF DARKNESS PREDATOR & PREY #5]  P
BOWEN, CARL  PILLAR OF THE SUN [@EXALTED #4]  P

BOWEN, CARL/FLEMING, GHERBOD SILENT STRIDERS [WORLD OF DARKNESS WEREOLF 2]/ BLACK FURIES [WORLD OF DARKNESS WEREOLF 2]

PA

BOWEN, CARL  VAMPIRE [@WORLD OF DARKNESS PREDATOR & PREY #1]  P
BOWEN, CHARLES  HUMANOIDS, THE  [C/RAN-P]
BOWEN, GARY  FLOATING WORLDS  B
BOWEN, GARY  GREEN ECHO  B
BOWEN, GARY  QUEER DESTINIES  SR-L
BOWEN, JOHN  AFTER THE RAIN  C/P,P,H
BOWEN, MARJORIE  KECKSIES AND OTHER TWILIGHT TALES  H
BOWER, B. M.  ADAM CHASERS, THE  H
BOWES, RICHARD  FERAL CELL  P
BOWES, RICHARD  FROM THE FILES OF THE TIME RANGERS  H
BOWES, RICHARD  GOBLIN MARKET [WARCHILD 2]  P

AUTHOR  TITLE  CODE
BOWES, RICHARD  MINIONS OF THE MOON  H
BOWES, RICHARD  WARCHILD [WARCHILD 1]  C/P,P
BOWKER, RICHARD  DOVER BEACH  P
BOWKER, RICHARD  FORBIDDEN SANCTUARY  C/P:2,P
BOWKER, RICHARD  MARLBOROUGH STREET  C/P:2,SFWA,P
BOWKER, RICHARD  REPLICA  C/P,P
BOWKETT, STEPHEN  DREAMCASTLE [WEB,THE]  P
BOWLES, PAUL  UP ABOVE THE WORLD  P
BOWLING, DREW C.  TOWER OF SHADOWS, THE  H
BOXLEITNER, BRUCE  FRONTIER EARTH [@FRONTIER EARTH 1]  H
BOXLEITNER, BRUCE  SEARCHER [@FRONTIER EARTH 2]  P
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1977  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1978  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1979  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1980  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1981  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1982  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1983  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SCIENCE FICTION MAGAZINES 1984  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SEMI-PROFESSIONAL FANTASY MAGAZINES 1982  R/IND
BOYAJIAN, JERRY/JOHNSON, KENNETH R.  INDEX TO THE SEMI-PROFESSIONAL FANTASY MAGAZINES 1983  R/IND
BOYCE, CHRIS  CATCHWORLD  H,SFWA,P
BOYD, DONNA  ALCHEMIST, THE  LP
BOYD, JOHN  ANDROMEDA GUN  C/P,P,H
BOYD, JOHN  BARNARD’S PLANET  C/P,P,H
BOYD, JOHN  DOOMSDAY GENE, THE  H
BOYD, JOHN  GIRL WITH THE JADE GREEN EYES, THE  SR-P
BOYD, JOHN  GORGON FESTIVAL, THE  C/P,P,H
BOYD, JOHN  I.Q. MERCHANT, THE  H
BOYD, JOHN  LAST STARSHIP FROM EARTH, THE  H:2,P
BOYD, JOHN  ORGAN BANK FARM, THE  C/P,P,H
BOYD, JOHN  POLLINATORS OF EDEN, THE  H:2,P
BOYD, JOHN  RAKEHELLS OF HEAVEN, THE  C/P,P,H
BOYD, JOHN  SEX AND THE HIGH COMMAND  C/P:2,P,H
BOYD, MALCOLM  FANTASY WORLDS OF PETER STONE AND OTHER FABLES, THE  P
BOYER, ELIZABETH H.  BLACK LYNX, THE [SKYLA 2]  P
BOYER, ELIZABETH H.  CLAN OF THE WARLORD, THE [SKYLA 1]  P
BOYER, ELIZABETH H.  CURSE OF SLAGFID, THE [WIZARD’S WAR 2]  P
BOYER, ELIZABETH H.  DRAGON’S CARBUNCLE, THE [WIZARD’S WAR 3]  C/P,P
BOYER, ELIZABETH H.  ELVES AND THE OTTERSKIN, THE [ALFAR 2]  C/P:2,P
BOYER, ELIZABETH H.  KEEPER OF CATS [SKYLA 3]  C/P,P
BOYER, ELIZABETH H.  LORD OF CHAOS, THE [WIZARD’S WAR 4]  C/P:2,P
BOYER, ELIZABETH H.  SWORD AND THE SATCHEL, THE [ALFAR 1]  C/P:2,P
BOYER, ELIZABETH H.  THRALL AND THE DRAGON’S HEART, THE [ALFAR 3]  C/P:2,P
BOYER, ELIZABETH H.  TROLL’S GRINDSTONE, THE [WIZARD’S WAR 1]  C/P,P
BOYER, ELIZABETH H.  WIZARD AND THE WARLORD, THE [ALFAR 4]  C/P:2,P
BOYER, RICHARD L.  GIANT RAT OF SUMATRA, THE  P
BOYER, ROBERT H./ZAHORSKI, KENNETH J.  DARK IMAGININGS  HA,LPA
BOYER, ROBERT H./ZAHORSKI, KENNETH J.  FANTASISTS ON FANTASY  R/REF-P
BOYER, ROBERT H./ZAHORSKI, KENNETH J.  FANTASTIC IMAGINATION, THE  C/PA,PA
BOYER, ROBERT H./ZAHORSKI, KENNETH J.  PHOENIX TREE, THE  C/PA,PA
BOYER, ROBERT H./ZAHORSKI, KENNETH J.  VISIONS OF WONDER  PA
BOYETT, STEVEN R.  ARCHITECT OF SLEEP, THE  P
BOYETT, STEVEN R.  ARIEL  C/P:2,P
BOYETT, STEVEN R.  TREKS NOT TAKEN [STAR TREK TNG]  LP
BOYLAN, GRACE DUFFIE  KISS OF GLORY, THE  H
BOYLL, RANDALL  DARKMAN  P
BOYLL, RANDALL  DEMON KNIGHT  P
BOYLL, RANDALL  HANGMAN, THE [@DARKMAN #1]  P
BOYLL, RANDALL  IN THE FACE OF DEATH [@DARKMAN #4]  P
BOYLL, RANDALL  PRICE OF FEAR, THE [@DARKMAN #2]  P
BOYNTON, SANDRA  CHOCOLATE, THE CONSUMING PASSION  [C/RAN-H]
BOYS’ LIFE EDITORS  BOYS’ LIFE BOOK OF OUTER SPACE STORIES, THE  HA
BRACK, VEKTIS  CASTAWAY FROM SPACE  SR-P
BRACK, VEKTIS  ODYSSEY IN SPACE  SR-P
BRACKETT, LEIGH  ALPHA CENTAURI — OR DIE!  C/P:2,DF187
BRACKETT, LEIGH  ARK OF MARS, THE  B
BRACKETT, LEIGH  BEST OF LEIGH BRACKETT, THE  H:2,P
BRACKETT, LEIGH  BEST OF PLANET STORIES #1, THE  PA,C/PA
BRACKETT, LEIGH  BIG JUMP, THE  C/P:2,DD103,P
BRACKETT, LEIGH  BOOK OF SKAITH, THE [SKAITH 1,2,3]  H:2
BRACKETT, LEIGH  COMING OF THE TERRANS, THE  C/P:2,P
BRACKETT, LEIGH/KASDAN, LAWRENCE  EMPIRE STRIKES BACK THE SCREENPLAY, THE [STAR WARS MOVIE 5]  SR-S
BRACKETT, LEIGH  ERIC JOHN STARK: OUTLAW OF MARS  P

The MITSFS Pinkdex  September 24, 2007  Page 50
AUTHOR  TITLE  CODE
BRACKETT, LEIGH  GALACTIC BREED, THE  DD099
BRACKETT, LEIGH  GINGER STAR, THE [SKAITH 1]  C/P:2,P
BRACKETT, LEIGH  HALFLING AND OTHER STORIES, THE  C/P:2,P
BRACKETT, LEIGH  HOUNDS OF SKAITH, THE [SKAITH 2]  C/P:2,P
BRACKETT, LEIGH  JEWEL OF BAS, THE  TD22,SFWA-TD22
BRACKETT, LEIGH  LONG TOMORROW, THE  H:2,P
BRACKETT, LEIGH  MARTIAN QUEST  H
BRACKETT, LEIGH  NEMESIS FROM TERRA, THE  DF123,P,SFWA-TD8,TD8
BRACKETT, LEIGH  PEOPLE OF THE TALISMAN  DM101
BRACKETT, LEIGH  REAVERS OF SKAITH, THE [SKAITH 3]  C/P:2,P
BRACKETT, LEIGH  SECRET OF SINHARAT, THE  DM101
BRACKETT, LEIGH  STARMEN, THE  H
BRACKETT, LEIGH  STARMEN OF LLYRDIS, THE  C/P:2,P
BRACKETT, LEIGH  SWORD OF RHIANNON, THE  C/P:2,SR-DD036,P
BRADBURY, EDWARD P.  BARBARIANS OF MARS [MICHAEL KANE 3]  SR-P,C/P:2
BRADBURY, EDWARD P.  BLADES OF MARS [MICHAEL KANE 2]  SR-P,C/P:2
BRADBURY, EDWARD P.  WARRIORS OF MARS [MICHAEL KANE 1]  SR-P
BRADBURY, RAY  ABISMO DE CHICAGO, O  FP
BRADBURY, RAY  BRADBURY SPEAKS  R/REF-H
BRADBURY, RAY  BRADBURY STORIES  LP
BRADBURY, RAY  CAT’S PAJAMAS, THE  H
BRADBURY, RAY  CIRCUS OF DOCTOR LAO, THE  C/PA,PA
BRADBURY, RAY  COMPLETE POEMS OF RAY BRADBURY, THE  [C/RAN-P]
BRADBURY, RAY  CRONICI MARTIENE  FP
BRADBURY, RAY  DANDELION WINE  C/P,P,H
BRADBURY, RAY  DEATH IS A LONELY BUSINESS  P
BRADBURY, RAY  DRIVING BLIND  P
BRADBURY, RAY  FAHRENHEIT 451  SR-P,H:2,P
BRADBURY, RAY  FROM THE DUST RETURNED  H:2
BRADBURY, RAY  GOLDEN APPLES OF THE SUN, THE  C/P:2,P,H
BRADBURY, RAY  GREEN SHADOWS, WHITE WHALE  P
BRADBURY, RAY  HALLOWEEN TREE, THE  C/P,P,H
BRADBURY, RAY  I SING THE BODY ELECTRIC!  C/P:2,P,H
BRADBURY, RAY  ILLUSTRATED MAN, THE  C/P,H,SR-P
BRADBURY, RAY  LONG AFTER MIDNIGHT  C/P,P,H
BRADBURY, RAY  MACHINERIES OF JOY, THE  C/P:3,P
BRADBURY, RAY  MAQUINAS DA ALEGRIA, AS  FP
BRADBURY, RAY  MARTIAN CHRONICLES, THE  H:2,SFWA,SR-P
BRADBURY, RAY  MEDICINE FOR MELANCHOLY, A  C/P,P
BRADBURY, RAY  OCTOBER COUNTRY, THE  C/P:2,P
BRADBURY, RAY  ONE MORE FOR THE ROAD  H
BRADBURY, RAY  PILLAR OF FIRE AND OTHER PLAYS  P
BRADBURY, RAY  QUICKER THAN THE EYE  P
BRADBURY, RAY  R IS FOR ROCKET  C/P:2,P
BRADBURY, RAY  S IS FOR SPACE  C/P,P,H
BRADBURY, RAY  SILVER LOCUSTS, THE  SR-P
BRADBURY, RAY  SOMETHING WICKED THIS WAY COMES  C/P,P,H
BRADBURY, RAY  STORIES OF RAY BRADBURY, THE  H
BRADBURY, RAY  SWITCH ON THE NIGHT  H
BRADBURY, RAY  TIMELESS STORIES FOR TODAY AND TOMORROW  SR-PA,C/PA
BRADBURY, RAY  TOMORROW MIDNIGHT  SR-P,C/P
BRADBURY, RAY  TOYNBEE CONVECTOR, THE  C/P,P
BRADBURY, RAY  TWICE 22  H:2
BRADBURY, RAY  VINTAGE BRADBURY, THE  C/P:2,P
BRADBURY, RAY  WHEN ELEPHANTS LAST IN THE DOORYARD BLOOMED  H
BRADBURY, RAY  WONDERFUL ICE CREAM SUIT AND OTHER PLAYS, THE  C/P,P
BRADDOCK, HANOVI  ASHES OF THE SUN [MAGIC: THE GATHERING]  C/P,P
BRADFORD, MATTHEW C.  INVASION FROM SPACE  SR-P

BRADLEY, MARION ZIMMER/ ROSS, DEBORAH J.

ALTON GIFT, THE [DARKOVER]

H

BRADLEY, MARION ZIMMER  BEST OF MARION ZIMMER BRADLEY, THE  C/P,P
BRADLEY, MARION ZIMMER  BEST OF MARION ZIMMER BRADLEY’S FANTASY MAGAZINE, THE  PA

BRADLEY, MARION ZIMMER/ WATERS, ELISABETH BEST OF MARION ZIMMER BRADLEY’S FANTASY MAGAZINE, THE, VOL. II

C/PA,PA

BRADLEY, MARION ZIMMER/MAY, JULIAN/ NORTON, ANDRE BLACK TRILLIUM [TRILLIUM 1]

C/P,P,H

BRADLEY, MARION ZIMMER  BLOODY SUN, THE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  BLOODY SUN, THE (EXPANDED) [DARKOVER]  C/P,P,H
BRADLEY, MARION ZIMMER  BRASS DRAGON, THE  C/P:2,P,D37250
BRADLEY, MARION ZIMMER  CASTLE TERROR  C/P,P
BRADLEY, MARION ZIMMER  CATCH TRAP, THE  P
BRADLEY, MARION ZIMMER  CHILDREN OF HASTUR [DARKOVER]  H:2
BRADLEY, MARION ZIMMER  CITY OF SORCERY [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  COLORS OF SPACE, THE  C/P:2,P
BRADLEY, MARION ZIMMER  DARK INTRUDER AND OTHER STORIES, THE  DF273

The MITSFS Pinkdex  September 24, 2007  Page 51
AUTHOR  TITLE  CODE
BRADLEY, MARION ZIMMER  DARK SATANIC  C/P:2,SFWA-P,P
BRADLEY, MARION ZIMMER  DARKOVER LANDFALL [DARKOVER]  H:2,P
BRADLEY, MARION ZIMMER  DOMAINS OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  DOOR THROUGH SPACE, THE  C/P:2,DF117,P
BRADLEY, MARION ZIMMER  DRUMS OF DARKNESS  P
BRADLEY, MARION ZIMMER  ENDLESS UNIVERSE  C/P:2,P
BRADLEY, MARION ZIMMER  ENDLESS VOYAGE  C/P:2,P
BRADLEY, MARION ZIMMER  EXILE’S SONG [DARKOVER]  H:2,P
BRADLEY, MARION ZIMMER  FALCONS OF NARABEDLA  P,DF273
BRADLEY, MARION ZIMMER  FALL OF ATLANTIS, THE [WEB 1,2]  C/P:2,P

BRADLEY, MARION ZIMMER/ ROSS, DEBORAH J.

FALL OF NESKAYA, THE [@CLINGFIRE #1/DARKOVER]

H,P

BRADLEY, MARION ZIMMER  FIREBRAND, THE  H:2,P

BRADLEY, MARION ZIMMER/ ROSS, DEBORAH J.

FLAME IN HALI, A [@CLINGFIRE #3]

H

BRADLEY, MARION ZIMMER  FORBIDDEN TOWER, THE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  FOREST HOUSE, THE [MISTS OF AVALON 2]  H,LP
BRADLEY, MARION ZIMMER  FOUR MOONS OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  FREE AMAZONS OF DARKOVER [DARKOVER]  PA,C/PA
BRADLEY, MARION ZIMMER  GHOSTLIGHT  H:2,LP
BRADLEY, MARION ZIMMER/LISLE, HOLLY  GLENRAVEN [GLENRAVEN 1]  C/P,P
BRADLEY, MARION ZIMMER  GRATITUDE OF KINGS, THE  H:2
BRADLEY, MARION ZIMMER  GRAVELIGHT  H:2
BRADLEY, MARION ZIMMER  GREYHAVEN  PA,C/PA
BRADLEY, MARION ZIMMER  HAWKMISTRESS! [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  HEARTLIGHT  LP
BRADLEY, MARION ZIMMER  HEIRS OF HAMMERFELL, THE [DARKOVER]  H:2,P
BRADLEY, MARION ZIMMER  HERITAGE OF HASTUR, THE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  HOUSE BETWEEN THE WORLDS, THE  H:2,P,SFWA
BRADLEY, MARION ZIMMER  HUNTERS OF THE RED MOON  C/P:2,P
BRADLEY, MARION ZIMMER/LISLE, HOLLY  IN THE RIFT [GLENRAVEN 2]  H:2,P
BRADLEY, MARION ZIMMER  INHERITOR, THE [INHERITOR 1]  C/P:2,LP
BRADLEY, MARION ZIMMER  JEWEL OF ARWEN, THE  B
BRADLEY, MARION ZIMMER  KEEPER’S PRICE, THE [DARKOVER]  PA,C/PA
BRADLEY, MARION ZIMMER  LADY OF AVALON  H
BRADLEY, MARION ZIMMER  LADY OF THE TRILLIUM [TRILLIUM 4]  C/P,P,H
BRADLEY, MARION ZIMMER  LERONI OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  LYTHANDE [THIEVES’ WORLD]  C/P:2,P
BRADLEY, MARION ZIMMER  MARION ZIMMER BRADLEY’S DARKOVER [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  MISTS OF AVALON, THE [MISTS OF AVALON 1]  H:2,LP
BRADLEY, MARION ZIMMER  NIGHT’S DAUGHTER  C/P,P
BRADLEY, MARION ZIMMER  OATH OF THE RENUNCIATES [DARKOVER]  H
BRADLEY, MARION ZIMMER  OTHER SIDE OF THE MIRROR, THE [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  PARTING OF ARWEN, THE  B
BRADLEY, MARION ZIMMER  PLANET SAVERS, THE [DARKOVER]  DF153

BRADLEY, MARION ZIMMER PLANET SAVERS, THE [DARKOVER]/

SWORD OF ALDONES, THE [DARKOVER]

C/P,P

BRADLEY, MARION ZIMMER  PLANET SAVERS, THE (INCLUDES THE WATERFALL) [DARKOVER]  C/P,P

BRADLEY, MARION ZIMMER/ PAXSON, DIANA L.

PRIESTESS OF AVALON [MISTS OF AVALON 3]

H

BRADLEY, MARION ZIMMER  RED SUN OF DARKOVER [DARKOVER]  C/PA,PA

BRADLEY, MARION ZIMMER/ LACKEY, MERCEDES REDISCOVERY [DARKOVER]

H:2,P

BRADLEY, MARION ZIMMER  RENUNCIATES OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  RUINS OF ISIS, THE  C/P,P,H
BRADLEY, MARION ZIMMER  SAGA OF THE RENUNCIATES, THE [DARKOVER]  P
BRADLEY, MARION ZIMMER  SEVEN FROM THE STARS  C/P,DF127,P
BRADLEY, MARION ZIMMER  SHADOW MATRIX, THE [DARKOVER]  H:2,P
BRADLEY, MARION ZIMMER  SHARRA’S EXILE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  SHATTERED CHAIN, THE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  SNOWS OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  SOUVENIR OF MONIQUE  C/P,P
BRADLEY, MARION ZIMMER  SPELL SWORD, THE [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  SPELLS OF WONDER  C/PA,PA
BRADLEY, MARION ZIMMER  STAR OF DANGER [DARKOVER]  C/P,P,H
BRADLEY, MARION ZIMMER  STORMQUEEN! [DARKOVER]  C/P,P,H
BRADLEY, MARION ZIMMER  SURVEY SHIP  C/P:2,LP

BRADLEY, MARION ZIMMER/ ZIMMER, PAUL EDWIN SURVIVORS, THE

C/P:2,P

BRADLEY, MARION ZIMMER  SWORD AND SORCERESS  PA,C/PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS II  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS III  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS IV  PA,C/PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS V  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 52
AUTHOR  TITLE  CODE
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS VI  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS VII  PA,C/PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS VIII  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS IX  C/PA,PA,SFWA-P
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS X  PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XI  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XII  PA,C/PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XIII  PA,C/PA

BRADLEY, MARION ZIMMER/ HOLMAN, RACHEL E.

SWORD AND SORCERESS XIV

PA,C/PA

BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XV  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XVI  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XVII  C/PA,PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XVIII  PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XIX  PA
BRADLEY, MARION ZIMMER  SWORD AND SORCERESS XX  PA
BRADLEY, MARION ZIMMER  SWORD OF ALDONES, THE [DARKOVER]  C/P,DF153,P
BRADLEY, MARION ZIMMER  SWORD OF CHAOS [DARKOVER]  PA,C/PA
BRADLEY, MARION ZIMMER  THENDARA HOUSE [DARKOVER]  C/P:2,P

BRADLEY, MARION ZIMMER/ NORTON, ANDRE/LACKEY, MERCEDES TIGER BURNING BRIGHT

C/P,P,H:2

BRADLEY, MARION ZIMMER  TO SAVE A WORLD [DARKOVER]  P
BRADLEY, MARION ZIMMER  TOWERS OF DARKOVER [DARKOVER]  C/PA,PA
BRADLEY, MARION ZIMMER  TRAITOR’S SUN [DARKOVER]  C/P,P,H
BRADLEY, MARION ZIMMER  TWO TO CONQUER [DARKOVER]  C/P:2,P
BRADLEY, MARION ZIMMER  WARRIOR WOMAN  C/P:2,P
BRADLEY, MARION ZIMMER  WEB OF DARKNESS [WEB 2]  C/P:2,LP
BRADLEY, MARION ZIMMER  WEB OF LIGHT [WEB 1]  C/P:2,P
BRADLEY, MARION ZIMMER  WINDS OF DARKOVER, THE [DARKOVER]  C/P,P,H,D89250
BRADLEY, MARION ZIMMER  WITCH HILL [INHERITOR 2]  C/P:2,SFWA-P,P
BRADLEY, MARION ZIMMER  WITCHLIGHT  H:2
BRADLEY, MARION ZIMMER  WORLD WRECKERS, THE [DARKOVER]  C/P,P,H

BRADLEY, MARION ZIMMER/ ROSS, DEBORAH J.

ZANDRU’S FORGE [@CLINGFIRE #2]

H

BRADLEY, REBECCA  LADY IN GIL [LADY IN GIL 1]  C/P,P
BRADLEY, REBECCA  LADY PAIN [LADY IN GIL 3]  P
BRADLEY, REBECCA  SCION’S LADY [LADY IN GIL 2]  P
BRADLEY, WILL  ARK LIBERTY  C/P,SFWA-P,P
BRADSHAW, GILLIAN  BEYOND THE NORTH WIND  H
BRADSHAW, GILLIAN  HAWK OF MAY [GWALCHMAI 1]  H:2,P
BRADSHAW, GILLIAN  HAWK OF MAY (PROOF) [GWALCHMAI 1]  SR-LP
BRADSHAW, GILLIAN  HORSES OF HEAVEN  C/P,P,H
BRADSHAW, GILLIAN  IN WINTER’S SHADOW  H:2,P
BRADSHAW, GILLIAN  KINGDOM OF SUMMER [GWALCHMAI 2]  H:2,P
BRADSHAW, GILLIAN  WRONG REFLECTION, THE  P
BRAGA, BRANDON  BIRTHRIGHT [STAR TREK TNG]  SR-S
BRAKE, COLIN  COLONY OF LIES, THE [DOCTOR WHO PAST DOCTORS]  P
BRAKE, COLIN  ESCAPE VELOCITY [DOCTOR WHO EIGHTH DOCTOR]  P
BRAKE, COLIN  PRICE OF PARADISE, THE [DOCTOR WHO TENTH DOCTOR]  H
BRAMAH, ERNEST  KAI LUNG UNROLLS HIS MAT [KAI LUNG]  C/P,P
BRAMAH, ERNEST  KAI LUNG’S GOLDEN HOURS [KAI LUNG]  H:2,P
BRAMAH, ERNEST  WALLET OF KAI LUNG, THE [KAI LUNG]  H
BRAMSTRUP, M. A.  COURIER  SR-P
BRAMSTRUP, M. A.  STAR DESTROYER  SR-P
BRAND, KURT  AGENTS OF DESTRUCTION [PERRY RHODAN 134]  SR-LP
BRAND, KURT  BLITZKRIEG GALACTICA [PERRY RHODAN 129]  SR-LP
BRAND, KURT  PERIL UNLIMITED [PERRY RHODAN 130]  SR-LP
BRAND, KURT  SCHATZKAMMER DER STERNE [@PERRY RHODAN #3]  FPR
BRAND, KURT  SINISTER POWER, THE [PERRY RHODAN 124]  SR-LP
BRAND, KURT  VOLUNTEER FOR FRAGO [PERRY RHODAN 122]  SR-LP
BRAND, LARRY  BIRTHPYRE  P
BRAND, MAX  DAN BARRY’S DAUGHTER  P
BRAND, MAX  GARDEN OF EDEN, THE  P
BRAND, MAX  NIGHT HORSEMAN, THE  P
BRANDAO, IGNACIO DELOYOLA  AND STILL THE EARTH  P
BRANDEL, MARC  MAN WHO LIKED WOMEN, THE  P
BRANDNER, GARY  AARDVARK AFFAIR, THE [@BIG BRAIN #1]  C/P:2,P
BRANDNER, GARY  BEELZEBUB BUSINESS, THE [@BIG BRAIN #2]  P
BRANDNER, GARY  ENERGY ZERO [@BIG BRAIN #3]  P
BRANDNER, GARY  HELLBORN  P
BRANDNER, GARY  HOWLING, THE [@HOWLING #1]  P
BRANDNER, GARY  HOWLING II, THE [@HOWLING #2]  P
BRANDNER, GARY  HOWLING III, THE [@HOWLING #3]  P
BRANDON, PAUL  SWIM THE MOON  LP

The MITSFS Pinkdex  September 24, 2007  Page 53
AUTHOR  TITLE  CODE
BRANDON, PAUL  WILD REEL, THE  LP
BRANNEY, SEAN/LEMAN, ANDREW  SHOGGOTH ON THE ROOF, A: THE DOCUMENTARY (DVD)  SR-MM
BRANNEY, SEAN/LEMAN, ANDREW  SHOGGOTH ON THE ROOF, A: THE MUSICAL  SR-VLP
BRAUN, MATTHEW  SECOND COMING OF LUCAS BROKAW, THE  P
BRAUNBECK, GARY A.  IN HOLLOW HOUSES [DARK MATTER 1]  P
BRAUTIGAN, RICHARD  HAWKLINE MONSTER, THE  P
BRAVER, GARY  GRAY MATTER  H
BRAY, PATRICIA  DEVLIN’S HONOR [SWORD OF CHANGE 2]  C/P,P
BRAY, PATRICIA  DEVLIN’S HONOR (PROOF) [SWORD OF CHANGE 2]  SR-P
BRAY, PATRICIA  DEVLIN’S JUSTICE [SWORD OF CHANGE 3]  C/P,P
BRAY, PATRICIA  DEVLIN’S LUCK [SWORD OF CHANGE 1]  C/P,P
BRAY, PATRICIA  FIRST BETRAYAL, THE  P
BREARLEY, JOHN  NIGHT HAWK, THE  B
BREBNER, PERCY JAMES  IVORY DISC, THE  H
BREDE, ARNOLD  SISTER EARTH  SR-P
BREDENBERG, JEFF  DREAM COMPASS, THE  C/P,P
BREDENBERG, JEFF  DREAM VESSEL, THE  P,SFWA-P
BREDENBERG, JEFF  MAN IN THE MOON MUST DIE, THE  C/P,SFWA-P,P
BREDESON, LENORE/NELSON, DICK  SPACE JOKES, COSMIC CARTOONS, AND MARTIAN LAUGHS  [C/RAN-P]
BREEN, WALTER  DARKOVER CONCORDANCE, THE  R/REF-H
BREGGIN, PETER  AFTER THE GOOD WAR  P
BREIG, JOSEPH A.  DEVIL YOU SAY, THE  H
BRENCHLEY, CHAZ  DARK WAY TO GLORY, A [@OUTREMER #3]  P
BRENCHLEY, CHAZ  DEVIL IN THE DUST, THE [@OUTREMER #1]  P
BRENCHLEY, CHAZ  END OF ALL ROADS, THE [@OUTREMER #6]  P
BRENCHLEY, CHAZ  FEAST OF THE KING’S SHADOW [@OUTREMER #4]  P
BRENCHLEY, CHAZ  FEAST OF THE KING’S SHADOW (COMPLETE) [OUTREMER 3,4]  P
BRENCHLEY, CHAZ  HAND OF THE KING’S EVIL [@OUTREMER #5]  P
BRENCHLEY, CHAZ  KEYS TO D’ESPERANCE, THE  B
BRENCHLEY, CHAZ  TOWER OF THE KING’S DAUGHTER [@OUTREMER #2]  C/P,P
BRENCHLEY, CHAZ  TOWER OF THE KING’S DAUGHTER (COMPLETE) [OUTREMER 1,2]  P
BRENNAN, C. M.  GENESIS WEB, THE [BUCK ROGERS INVADERS OF CHARON 1]  P
BRENNAN, CAITLIN  MOUNTAIN’S CALL, THE [AURELIAN EMPIRE 1]  LP
BRENNAN, CAITLIN  SHATTERED DANCE [AURELIAN EMPIRE 3]  LP
BRENNAN, CAITLIN  SONG OF UNMAKING [AURELIAN EMPIRE 2]  LP
BRENNAN, FREDERICK H.  ONE OF OUR H BOMBS IS MISSING  P
BRENNAN, HERBIE  FAERIE WARS  LP
BRENNAN, J. H.  SHIVA ACCUSED  SFWA

BRENNAN, JOSEPH PAYNE/ GRANT, DONALD M.

ACT OF PROVIDENCE

SR-H

BRENNAN, JOSEPH PAYNE  BORDERS JUST BEYOND, THE  SR-H
BRENNAN, JOSEPH PAYNE  CASEBOOK OF LUCIUS LEFFING, THE  SR-H
BRENNAN, JOSEPH PAYNE  NINE HORRORS AND A DREAM  P
BRENNAN, JOSEPH PAYNE  SHAPES OF MIDNIGHT, THE  P
BRENNAN, JOSEPH PAYNE  STORIES OF DARKNESS AND DREAD  SR-H
BRENNAN, MARIE  DOPPELGANGER [MIREI 1]  P
BRENNAN, MARIE  WARRIOR AND WITCH [MIREI 2]  P
BRENNAN, NOEL-ANNE  BLOOD OF THE LAND, THE [RILSIN 2]  P
BRENNAN, NOEL-ANNE  DAUGHTER OF THE DESERT  P
BRENNAN, NOEL-ANNE  SWORD OF THE LAND, THE [RILSIN 1]  P
BRENNAN, NOEL-ANNE  WINTER RECKONING  SR-H
BRENNER, MAYER ALAN  CATASTROPHE’S SPELL [DANCE OF GODS 1]  C/P,P
BRENNER, MAYER ALAN  SPELL OF APOCALYPSE [DANCE OF GODS 4]  P
BRENNER, MAYER ALAN  SPELL OF FATE [DANCE OF GODS 3]  P
BRENNER, MAYER ALAN  SPELL OF INTRIGUE [DANCE OF GODS 2]  P
BRENNERT, ALAN  HER PILGRIM SOUL  H,SFWA
BRENNERT, ALAN  KINDRED SPIRITS  P
BRENNERT, ALAN  MA QUI AND OTHER PHANTOMS  SR-LP

BRENNERT, ALAN/MENKEN, ALAN/ SPENCER, DAVID WEIRD ROMANCE (CD)

SR-MM

BRENT, JEREMY  PLASTIC MAN, THE  P
BRENT, MADELEINE  MERLIN’S KEEP  P
BRETNOR, REGINALD  CRAFT OF SCIENCE FICTION, THE  SFWA
BRETNOR, REGINALD  GILPIN’S SPACE  P
BRETNOR, REGINALD  MODERN SCIENCE FICTION: ITS MEANING AND ITS FUTURE  R/REF-H
BRETNOR, REGINALD  ORION’S SWORD [@FUTURE AT WAR #3]  PA,C/PA,SFWA-P
BRETNOR, REGINALD  SCHIMMELHORN FILE: MEMOIRS OF A DIRTY OLD GENIUS, THE  C/P:2,P
BRETNOR, REGINALD  SCHIMMELHORN’S GOLD  C/P:2,SFWA-P,P
BRETNOR, REGINALD  SCIENCE FICTION, TODAY AND TOMORROW  SFWA,R/REF-P,C/REF-H
BRETNOR, REGINALD  SPEAR OF MARS, THE [@FUTURE AT WAR #2]  PA,C/PA,SFWA-P
BRETNOR, REGINALD  THOR’S HAMMER [@FUTURE AT WAR #1]  SFWA-P,C/PA,PA
BRETON, THIERRY/BENEICH, DENIS  SOFTWAR  H,SR-LP
BRETT, LEO  ALIEN ONES, THE  SR-P,C/P
BRETT, LEO  BLACK INFINITY  SR-P

The MITSFS Pinkdex  September 24, 2007  Page 54
AUTHOR  TITLE  CODE
BRETT, LEO  EXIT HUMANITY  SR-P
BRETT, LEO  FACE IN THE NIGHT  SR-P
BRETT, LEO  FACELESS PLANET  SR-P
BRETT, LEO  FORBIDDEN, THE  SR-P
BRETT, LEO  FROM REALMS BEYOND  SR-P
BRETT, LEO  IMMORTALS, THE  SR-P
BRETT, LEO  MIND FORCE  H
BRETT, LEO  NIGHTMARE  SR-P
BRETT, LEO  POWER SPHERE  SR-P
BRETT, LEO  THEY NEVER COME BACK  SR-P
BREWER, GENE  K-PAX [K-PAX 1]  P
BREWER, GENE  ON A BEAM OF LIGHT [K-PAX 2]  P
BREWER-GIORGIO, GAIL  IS ELVIS ALIVE?  [C/RAN-P]
BREWSTER, DENNIS  I DREAM OF JEANNIE  LP
BRIARTON, GRENDEL  COLLECTED FEGHOOT, THE  LP
BRIARTON, GRENDEL  COMPLEAT FEGHOOT, THE  SR-P,C/P:2

BRIDGES, BILL/ACHILLI, JUSTIN BONE GNAWERS [WORLD OF DARKNESS WEREOLF 4]/ STARGAZERS [WORLD OF DARKNESS WEREOLF 4]

PA

BRIDGES, LAURIE  ASHTON HORROR, THE [@DARK FORCES 12]  P
BRIDGES, LAURIE/ALEXANDER, PAUL  DEVIL WIND [@DARK FORCES 4]  P
BRIDGES, LAURIE/ALEXANDER, PAUL  MAGIC SHOW [@DARK FORCES 2]  P
BRIDGES, WILLIAM  SILVER CROWN, THE [WORLD OF DARKNESS RAGE]  C/P,P
BRIERY, TRACI  VAMPIRE JOURNALS, THE  P
BRIGGS, NICHOLAS/GUERRIER, SIMON  SHORT TRIPS: DALEK EMPIRE [DOCTOR WHO]  HA
BRIGGS, PATRICIA  BLOOD BOUND [MERCY THOMPSON 2]  C/P,P
BRIGGS, PATRICIA  DRAGON BLOOD [WARD OF HUROG 2]  C/P,P
BRIGGS, PATRICIA  DRAGON BONES [WARD OF HUROG 1]  C/P,P
BRIGGS, PATRICIA  HOB’S BARGAIN, THE  P
BRIGGS, PATRICIA  MASQUES  C/P,P
BRIGGS, PATRICIA  MOON CALLED [MERCY THOMPSON 1]  P
BRIGGS, PATRICIA  RAVEN’S SHADOW [SERAPH 1]  P
BRIGGS, PATRICIA  RAVEN’S STRIKE [SERAPH 2]  P
BRIGGS, PATRICIA  STEAL THE DRAGON  P
BRIGGS, PATRICIA  WHEN DEMONS WALK  C/P,P
BRIMMICOMBE-WOOD, LEE  ALIENS COLONIAL MARINES TECHNICAL MANUAL [@ALIENS]  VLP
BRIN, DAVID  BRIGHTNESS REEF [UPLIFT 4]  H:2,P
BRIN, DAVID  EARTH  H:2,P
BRIN, DAVID  EARTHCLAN [UPLIFT 2,3]  H:2
BRIN, DAVID  FOUNDATION’S TRIUMPH [SECOND FOUNDATION 3]  C/P,P,H
BRIN, DAVID  GLORY SEASON  C/P:2,P,H
BRIN, DAVID  GLORY SEASON (EXCERPT)  SR-LP
BRIN, DAVID  HEAVEN’S REACH [UPLIFT 6]  H:2,P
BRIN, DAVID  INFINITY’S SHORE [UPLIFT 5]  H:2,P
BRIN, DAVID  KILN PEOPLE  H:2,P
BRIN, DAVID  OTHERNESS  C/P:2,P
BRIN, DAVID  POSTMAN, THE  C/P:2,SFWA,P
BRIN, DAVID  PRACTICE EFFECT, THE  C/P,SR-P,H
BRIN, DAVID  RIVER OF TIME, THE  C/P:2,P,H
BRIN, DAVID  STARTIDE RISING [UPLIFT 2]  C/P:2,P
BRIN, DAVID  STARTIDE RISING (REVISED) [UPLIFT 2]  C/P,P,H
BRIN, DAVID  SUNDIVER [UPLIFT 1]  C/P:2,P
BRIN, DAVID  TOMORROW HAPPENS  SR-H
BRIN, DAVID  UPLIFT WAR, THE [UPLIFT 3]  C/P:2,P
BRINDEL, JUNE RACHUY  PHAEDRA  H
BRINEY, ROBERT/WOOD, EDWARD  SF BIBLIOGRAPHIES  R/IND
BRINK, CAROL RYRIE  ANDY BUCKRAM’S TIN MEN  P
BRINKLEY, WILLIAM  LAST SHIP, THE  H
BRINLEY, B. R., CAPT.  ROCKET MANUAL FOR AMATEURS  [C/RAN-P]
BRINLEY, BERTRAND R.  MAD SCIENTISTS’ CLUB, THE  C/P:2,P
BRINLEY, BERTRAND R.  NEW ADVENTURES OF THE MAD SCIENTISTS’ CLUB, THE  P
BRINTON, HENRY  PURPLE-6  H,P
BRISCO, PAT  OTHER PEOPLE, THE  P
BRITAIN, DAN  CIVIL WAR II  P
BRITAIN, DAN  GODMAKERS, THE  C/P:2,P
BRITAIN, KRISTEN  FIRST RIDER’S CALL [GREEN RIDER 2]  H,P
BRITAIN, KRISTEN  GREEN RIDER [GREEN RIDER 1]  H:2,P
BRITE, POPPY Z.  LAZARUS HEART, THE [@CROW]  LP
BRITE, POPPY Z.  LOST SOULS  C/P,P
BRITE, POPPY Z.  LOVE IN VEIN  C/PA,LPA
BRITE, POPPY Z.  LOVE IN VEIN II  LPA
BRITTAIN, C. DALE  BAD SPELL IN YURT, A [WIZARD OF YURT 1]  P,SFWA-P
BRITTAIN, C. DALE/BOUCHARD, ROBERT A.  COUNT SCAR  C/P,P
BRITTAIN, C. DALE  DAUGHTER OF MAGIC [WIZARD OF YURT 5]  P
BRITTAIN, C. DALE  IS THIS APOCALYPSE NECESSARY?  LP

The MITSFS Pinkdex  September 24, 2007  Page 55
AUTHOR  TITLE  CODE
BRITTAIN, C. DALE  MAGE QUEST [WIZARD OF YURT 3]  C/P,SFWA-P,P
BRITTAIN, C. DALE  VOIMA  C/P,P
BRITTAIN, C. DALE  WITCH AND THE CATHEDRAL, THE [WIZARD OF YURT 4]  P
BRITTAIN, C. DALE  WOOD NYMPH AND THE CRANKY SAINT, THE [WIZARD OF YURT 2]  C/P,SFWA-P,P
BRITTON, DAVID/BUTTERWORTH, MICHAEL  SAVOY BOOK, THE  SR-PA
BRIZZI, MARY T.  PHILIP JOSE FARMER  R/REF-P
BRIZZOLARA, JOHN  EMPIRE’S HORIZON  P
BROCK, BETTY  NO FLYING IN THE HOUSE  P
BROCK, DARRYL  IF I NEVER GET BACK [SAMUEL CLEMENS FOWLER 1]  H
BROCK, DARRYL  TWO IN THE FIELD [SAMUEL CLEMENS FOWLER 2]  LP
BRODERICK, DAMIEN  BLACK GRAIL, THE  C/P:2,P
BRODERICK, DAMIEN  DREAMING DRAGONS, THE  C/P,P
BRODERICK, DAMIEN  GODPLAYERS [PLAYERS IN THE CONTEST OF WORLDS 1]  LP
BRODERICK, DAMIEN  JUDAS MANDALA, THE  C/P:2,P
BRODERICK, DAMIEN  K-MACHINES [PLAYERS IN THE CONTEST OF WORLDS 2]  LP
BRODERICK, DAMIEN  SORCERER’S WORLD  C/P:2,P
BRODERICK, DAMIEN  STRIPED HOLES  P
BRODERICK, DAMIEN  TRANSCENSION  H
BRODERICK, DAMIEN  WHITE ABACUS, THE  C/P,LP,H

BRODEUR, GREG/ CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ GALANTER, DAVE/JOLLEY, DAN/ ROSENBERG, AARON SOME ASSEMBLY REQUIRED [@STAR TREK S.C.E. #3]

P

BRODIE-INNES, J. W.  DEVIL’S MISTRESS, THE  P

BRODKIN, SYLVIA Z./ PEARSON, ELIZABETH J.

SCIENCE FICTION

HA

BROECKER, RANDY  BROECKER SAMPLER, THE  B
BROLL, W. W.  PHANTOME UNTER UNS  FP
BROM  PLUCKER, THE  H
BRONDFIELD, JEROME  NIGHT IN FUNLAND  PA
BROOKE, GUN  PROTECTOR OF THE REALM [@SUPREME CONSTELLATIONS #1]  LP
BROOKE, GUN  REBEL’S QUEST [@SUPREME CONSTELLATIONS #2]  LP
BROOKE, KEITH  EXPATRIA  LP
BROOKE, KEITH  GENETOPIA  H
BROOKE-ROSE, CHRISTINE  XORANDOR  C/P,P
BROOKES, OWEN  INHERITANCE  P
BROOKS, MAX  WORLD WAR Z  H
BROOKS, NED/MARTIN, DON  HANNES BOK ILLUSTRATION INDEX  R/IND,R/IND-F
BROOKS, NED  INTERIM: HANNES BOK ILLUSTRATION INDEX (SUMMER 1967)  SRL/IND
BROOKS, NED/BEAHM, GEORGE W.  INTERIM: VAUGHN F. BODE ILLUSTRATION INDEX (SPRING 1971)  R/IND-F
BROOKS, TERRY  ANGEL FIRE EAST [VOID & WORD 3]  H
BROOKS, TERRY  ANTRAX [VOYAGE OF THE JERLE SHANNARA 2/SHANNARA JERLE]  H,P
BROOKS, TERRY  ARMAGEDDON’S CHILDREN  H
BROOKS, TERRY  BLACK UNICORN, THE [@LANDOVER 2]  C/P,P,H:2
BROOKS, TERRY  DRUID OF SHANNARA, THE [SHANNARA HERITAGE 2]  H:2,P
BROOKS, TERRY  ELF QUEEN OF SHANNARA, THE [SHANNARA HERITAGE 3]  H:2,P
BROOKS, TERRY  ELFSTONES OF SHANNARA, THE [SHANNARA 2]  C/P,LP,H
BROOKS, TERRY  FIRST KING OF SHANNARA [SHANNARA 0]  H:2,P
BROOKS, TERRY  HOOK  C/P,P,H
BROOKS, TERRY  ILSE WITCH [@VOYAGE OF THE JERLE SHANNARA #1/SHANNARA JERLE]  H
BROOKS, TERRY  JARKA RUUS [HIGH DRUID OF SHANNARA 1]  H
BROOKS, TERRY  JARKA RUUS (PROOF) [HIGH DRUID OF SHANNARA 1]  SR-LP
BROOKS, TERRY  KNIGHT OF THE WORD, A [VOID & WORD 2]  H,P
BROOKS, TERRY  MAGIC KINGDOM FOR SALE — SOLD! [LANDOVER 1]  C/P,H:2,P
BROOKS, TERRY  MORGAWR [@VOYAGE OF THE JERLE SHANNARA 3/SHANNARA JERLE]  H
BROOKS, TERRY  PHANTOM MENACE, THE [STAR WARS MOVIE 1]  H:2,P
BROOKS, TERRY  RUNNING WITH THE DEMON [VOID & WORD 1]  H,P
BROOKS, TERRY  SCIONS OF SHANNARA, THE [SHANNARA HERITAGE 1]  H:2,P
BROOKS, TERRY  SOMETIMES THE MAGIC WORKS  R/REF-H
BROOKS, TERRY  STRAKEN [HIGH DRUID OF SHANNARA 3]  H
BROOKS, TERRY  SWORD OF SHANNARA, THE [SHANNARA 1]  C/P,LP,H,SFWA:2
BROOKS, TERRY  TALISMANS OF SHANNARA, THE [SHANNARA HERITAGE 4]  H:2
BROOKS, TERRY  TANEQUIL [HIGH DRUID OF SHANNARA 2]  H
BROOKS, TERRY  TANGLE BOX, THE [LANDOVER 4]  C/P,P,H
BROOKS, TERRY  WISHSONG OF SHANNARA, THE [SHANNARA 3]  H:2,P
BROOKS, TERRY  WITCHES’ BREW [LANDOVER 5]  H:2,LP
BROOKS, TERRY  WIZARD AT LARGE [LANDOVER 3]  C/P,H:2,P
BROOKS, WALTER  ORIGINAL MR. ED, THE  P
BROOKS, WALTER R.  FREDDY AND THE FLYING SAUCER PLANS  H
BROSNAN, JOHN  FALL OF THE SKY LORDS, THE [SKY LORDS 3]  P
BROSNAN, JOHN  FUTURE TENSE: THE CINEMA OF SCIENCE FICTION  R/REF-H
BROSNAN, JOHN  OPOPONAX INVASION, THE  LP
BROSNAN, JOHN  SKY LORDS, THE [SKY LORDS 1]  SFWA
BROTHERS, ROBERTSKI  MCATRIX DERIDED, THE  H

The MITSFS Pinkdex  September 24, 2007  Page 56
AUTHOR  TITLE  CODE
BROTHERTON, MIKE  STAR DRAGON  H
BROWER, BROCK  LATE GREAT CREATURE, THE  P
BROWER, KENNETH  STARSHIP AND THE CANOE, THE  [C/RAN-P]
BROWN, BETH  UNIVERSAL STATION  H
BROWN, CARTER  GIRL WHO WAS POSSESSED, THE  P
BROWN, CARTER  TRUE SON OF THE BEAST!  P
BROWN, CHARLES BROCKDEN  WIELAND: OR, THE TRANSFORMATION  P
BROWN, CHARLES N./STRAHAN, JONATHAN LOCUS AWARDS, THE  LPA

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1984

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1986

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1987

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1988

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1989

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1990

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, AND HORROR: 1991

R/IND

BROWN, CHARLES N./ CONTENTO, WILLIAM G.

SCIENCE FICTION IN PRINT: 1985

R/IND

BROWN, DALE  DAY OF THE CHEETAH [MAC MCLANAHAN 2]  H
BROWN, DALE  FLIGHT OF THE OLD DOG [MAC MCLANAHAN 1]  P
BROWN, DALE  SILVER TOWER  H,P
BROWN, DAN  DA VINCI CODE, THE  H
BROWN, DENNIS PHILLIP  FOXGLOVE HOLLOW  P
BROWN, ERIC  DEEP FUTURE  H
BROWN, ERIC  HELIX  P
BROWN, ERIC  PENUMBRA  P
BROWN, ERIC  THRESHOLD SHIFT  H
BROWN, ERIC  UNTOUCHABLE [WEB,THE]  P
BROWN, FREDRIC  AND THE GODS LAUGHED  H
BROWN, FREDRIC  ANGELS AND SPACESHIPS  H
BROWN, FREDRIC  BEST OF FREDRIC BROWN, THE  H:2,P
BROWN, FREDRIC  DAYMARES  C/P:2,P
BROWN, FREDRIC  FROM THESE ASHES  H
BROWN, FREDRIC  HIS NAME WAS DEATH  P
BROWN, FREDRIC  HONEYMOON IN HELL  C/P:2,P
BROWN, FREDRIC  LIGHTS IN THE SKY ARE STARS, THE  C/P:2,P,H
BROWN, FREDRIC  LUNA DE MIEL EL INFIERNO (SPANISH)  FH
BROWN, FREDRIC  MARTIANS AND MADNESS  H:2
BROWN, FREDRIC  MARTIANS, GO HOME  C/P,P,H
BROWN, FREDRIC  MIND THING, THE  C/P:2,P
BROWN, FREDRIC  MURDERERS, THE  P
BROWN, FREDRIC  NIGHTMARES AND GEEZENSTACKS  P
BROWN, FREDRIC  ONE FOR THE ROAD  P
BROWN, FREDRIC  PARADOX LOST  H:2,SFWA,P
BROWN, FREDRIC  ROGUE IN SPACE  C/P:2,P
BROWN, FREDRIC/REYNOLDS, MACK  SCIENCE-FICTION CARNIVAL  PA,C/PA
BROWN, FREDRIC  SPACE ON MY HANDS  SR-P,C/P:2
BROWN, FREDRIC  STAR SHINE  C/P:2,P
BROWN, FREDRIC  VEN Y ENLOQUECE (SPANISH)  FH
BROWN, FREDRIC  WHAT MAD UNIVERSE  C/P,P,H
BROWN, GEORGE SHELDON  DESTINATION MARS  SR-P
BROWN, GEORGE SHELDON  PLANETOID PERIL, THE  SR-P
BROWN, HIMAN  STRANGE TALES FROM CBS RADIO MYSTERY THEATER  PA
BROWN, JAMES COOKE  TROIKA INCIDENT: A PROMISE, THE  H
BROWN, JERRY EARL  DARKHOLD  C/P,P
BROWN, JERRY EARL  EARTHFALL  C/P,P
BROWN, JERRY EARL  UNDER THE CITY OF ANGELS  C/P,P
BROWN, MARY  DRAGONNE’S EG [UNEXPECTED DRAGON 3]  P
BROWN, MARY  MASTER OF MANY TREASURES [UNEXPECTED DRAGON 2]  P
BROWN, MARY  PIGS DON’T FLY [UNEXPECTED DRAGON 1]  C/P:2,P
BROWN, MARY  STRANGE DELIVERANCE  P
BROWN, MARY  UNEXPECTED DRAGON, THE [UNEXPECTED DRAGON 1,2,3]  H
BROWN, MARY  UNLIKELY ONES, THE [UNEXPECTED DRAGON]  H:2,SFWA-P,P
BROWN, MOLLY  BAD TIMING AND OTHER STORIES  LP
BROWN, ROSEL GEORGE  GALACTIC SIBYL SUE BLUE  C/P:2,P
BROWN, ROSEL GEORGE  HANDFUL OF TIME, A  C/P:2,P
BROWN, ROSEL GEORGE  SIBYL SUE BLUE  H:2
BROWN, ROSEL GEORGE  WATERS OF CENTAURUS, THE  H:2,P
BROWN, SIMON  DAUGHTER OF INDEPENDENCE [KYDAN 3]  P

The MITSFS Pinkdex  September 24, 2007  Page 57
AUTHOR  TITLE  CODE
BROWN, SIMON  EMPIRE’S DAUGHTER [KYDAN 1]  C/P,P
BROWN, SIMON  FIRE AND SWORD [@KEYS OF POWER #2]  P
BROWN, SIMON  INHERITANCE [@KEYS OF POWER #1]  P
BROWN, SIMON  RIVAL’S SON [KYDAN 2]  P
BROWN, SIMON  SOVEREIGN [@KEYS OF POWER #3]  P
BROWN, SLATER  SPACEWARD BOUND  H
BROWN, STAN  CRAB, THE [FIVE RINGS CLAN WAR 5]  P

BROWN, STAN DRAGON DAY [@DRAGONLANCE NEW ADVENTURES #6/

DRAGONLANCE DRAGON QUARTET 2]

LP

BROWN, TIMOTHY  DARK KNIGHT OF KARAMEIKOS [MYSTARA]  P
BROWN, WENZELL  POSSESS AND CONQUER  P
BROWNE, C. J.  ESCAPE CARTHUS  P
BROWNE, GEORGE SHELDON  YELLOW PLANET, THE  SR-P
BROWNE, GERALD A.  HAZARD  P
BROWNE, GERALD A.  STONE 588  P
BROWNE, HOWARD  CARBON-COPY KILLER  GRD15
BROWNE, HOWARD  TWELVE TIMES ZERO  GRD15
BROWNE, ROBERT  NEW ATOMS’ BOMBSHELL, THE  C/P,P
BROXON, MILDRED DOWNEY  TOO LONG A SACRIFICE  H:2,P
BROYLES, L. D.  WHO’S WHO IN SCIENCE FICTION FANDOM 1961  R/REF-P
BRUCHAC, JOSEPH  BEARWALKER (PROOF)  SR-LP
BRUCKNER, KARL  DEUX ROBOTS, LES  FH
BRUGALETTE, PHILLIP  NINE GATES, THE  P
BRUMMELS, J. V.  DEUS EX MACHINA  C/P:2,P
BRUNN, ROBERT  INITIATION, THE [TWILIGHT #3]  P
BRUNNER, JOHN  AGE OF MIRACLES  C/P:2,P
BRUNNER, JOHN  ALTAR ON ASCONEL, THE  DM123
BRUNNER, JOHN  ASTRONAUTS MUST NOT LAND, THE  DF227
BRUNNER, JOHN  ATLANTIC ABOMINATION, THE  C/P:2,DD465,P
BRUNNER, JOHN  AVENGERS OF CARRIG, THE  C/P:2,P
BRUNNER, JOHN  BEDLAM PLANET  C/P:2,P
BRUNNER, JOHN  BEST OF JOHN BRUNNER, THE  P
BRUNNER, JOHN  BLACK IS THE COLOR  C/P:2,P
BRUNNER, JOHN  BOOK OF JOHN BRUNNER, THE  C/P,P
BRUNNER, JOHN  BORN UNDER MARS  C/P:2,P
BRUNNER, JOHN  BOTSCHAFT AUS DEM ALL  FP
BRUNNER, JOHN  CASTAWAYS’ WORLD  DF242
BRUNNER, JOHN  CATCH A FALLING STAR  C/P:2,P
BRUNNER, JOHN  CHILDREN OF THE THUNDER  C/P:2,P
BRUNNER, JOHN  COMPLEAT TRAVELLER IN BLACK, THE  C/P,LP
BRUNNER, JOHN  CRUCIBLE OF TIME, THE  C/P:2,P,H
BRUNNER, JOHN  DAY OF THE STAR CITIES, THE  C/P:2,P
BRUNNER, JOHN  DEVIL’S WORK, THE  H
BRUNNER, JOHN  DOADORES DE MUNDOS  FP
BRUNNER, JOHN  DOUBLE, DOUBLE  C/P:2,P
BRUNNER, JOHN  DRAMATURGES OF YAN, THE  C/P:2,P
BRUNNER, JOHN  DREAMING EARTH, THE  C/P:3,P
BRUNNER, JOHN  ECHO IN THE SKULL  DD385
BRUNNER, JOHN  ENDLESS SHADOW  DF299
BRUNNER, JOHN  ENIGMA FROM TANTALUS  DM115
BRUNNER, JOHN  ENTRY TO ELSEWHEN  C/P:2,P
BRUNNER, JOHN  EVIL THAT MEN DO, THE  SR-DB1010
BRUNNER, JOHN  FATHER OF LIES  SR-DB0081
BRUNNER, JOHN  FROM THIS DAY FORWARD  C/P:2,SFWA,P,H:2
BRUNNER, JOHN  GAUDY SHADOWS, THE  SR-P,C/P:2
BRUNNER, JOHN  GIVE WARNING TO THE WORLD  C/P,P
BRUNNER, JOHN  100TH MILLENNIUM, THE  DD362
BRUNNER, JOHN  INFINITIVE OF GO, THE  C/P:2,P
BRUNNER, JOHN  INTERSTELLAR EMPIRE  C/P:2,P
BRUNNER, JOHN  INTO THE SLAVE NEBULA  C/P:2,P
BRUNNER, JOHN  JAGGED ORBIT, THE  C/P:2,P,H
BRUNNER, JOHN  LISTEN! THE STARS!  DF215
BRUNNER, JOHN  LONG RESULT, THE  C/P:2,P
BRUNNER, JOHN  MANSHAPE  C/P:2,P
BRUNNER, JOHN  MAZE OF STARS, A  C/P:2,P,H
BRUNNER, JOHN  MEETING AT INFINITY  C/P:2,DD507,P
BRUNNER, JOHN  MORE THINGS IN HEAVEN  C/P:2,P
BRUNNER, JOHN  MUDDLE EARTH  C/P:2,P
BRUNNER, JOHN  NEW SETTLEMENT OF OLD SCORES, A  VLP
BRUNNER, JOHN  NO FUTURE IN IT  C/P,SR-P,H
BRUNNER, JOHN  NOW THEN!  C/P:2,P
BRUNNER, JOHN  OUT OF MY MIND  C/P:2,P
BRUNNER, JOHN  PLANET OF YOUR OWN, A  DG592
BRUNNER, JOHN  PLAYERS AT THE GAME OF PEOPLE  H:2,P
BRUNNER, JOHN  POLYMATH  C/P:2,P

AUTHOR  TITLE  CODE
BRUNNER, JOHN  PRODUCTIONS OF TIME, THE  C/P:2,P
BRUNNER, JOHN  QUICKSAND  H:2,P
BRUNNER, JOHN  REPAIRMEN OF CYCLOPS, THE  C/P:2,DM115
BRUNNER, JOHN  RITES OF OHE, THE  DF242
BRUNNER, JOHN  SANCTUARY IN THE SKY  DD471
BRUNNER, JOHN  SECRET AGENT OF TERRA  DF133
BRUNNER, JOHN  SHEEP LOOK UP, THE  H:2,P
BRUNNER, JOHN  SHIFT KEY, THE  P
BRUNNER, JOHN  SHOCKWAVE RIDER, THE  H:2,P
BRUNNER, JOHN  SKYNAPPERS, THE  DD457
BRUNNER, JOHN  SLAVERS OF SPACE  DD421
BRUNNER, JOHN  SOCIETA DEL TEMPO, LA  FP
BRUNNER, JOHN  SPACE-TIME JUGGLER, THE  DF227
BRUNNER, JOHN  SQUARES OF THE CITY, THE  C/P:2,SFWA-P,P
BRUNNER, JOHN  STAND ON ZANZIBAR  H:2,P
BRUNNER, JOHN  STARDROPPERS, THE  C/P:2,P
BRUNNER, JOHN  STONE THAT NEVER CAME DOWN, THE  C/P,SFWA,P,H:2
BRUNNER, JOHN  SUPER BARBARIANS, THE  C/P:2,P
BRUNNER, JOHN  THRESHOLD OF ETERNITY  DD335
BRUNNER, JOHN  TIDES OF TIME, THE  C/P,P,H
BRUNNER, JOHN  TIME-JUMP  C/P:2,P
BRUNNER, JOHN  TIMES WITHOUT NUMBER  DF161
BRUNNER, JOHN  TIMES WITHOUT NUMBER (REVISED)  C/P:2,P
BRUNNER, JOHN  TIMESCOOP  C/P:2,P
BRUNNER, JOHN  TO CONQUER CHAOS  C/P:2,P
BRUNNER, JOHN  TOMORROW MAY BE EVEN WORSE  B
BRUNNER, JOHN  TOTAL ECLIPSE  H:2,P,SFWA
BRUNNER, JOHN  TRAVELER IN BLACK, THE  C/P:2,P
BRUNNER, JOHN  WEAR THE BUTCHERS’ MEDAL  P
BRUNNER, JOHN  WEB OF EVERYWHERE  C/P:2,P
BRUNNER, JOHN  WHOLE MAN, THE  C/P:2,SFWA-P,P
BRUNNER, JOHN  WORLD SWAPPERS, THE  C/P:2,DD391,P
BRUNNER, JOHN  WRONG END OF TIME, THE  H:2,P,SFWA
BRUSH, KAREN A.  DEMON PIG  C/P,P
BRUSH, KAREN A.  PIG, THE PRINCE, AND THE UNICORN, THE  P
BRUSS, B. R.  CREATURE EPARSE, LA  FP
BRUSS, B. R.  ESPACE NOIR, L  FP
BRUSS, B. R.  ETRANGE PLANETE ORGA, L  FP
BRUSS, B. R.  MYSTERE DES SUPS, LE  FP
BRUSSOLO, SERGE  SEMINATORI DI ABISSI, I (ITALIAN)  FP
BRUST, STEVEN  AGYAR  H,P
BRUST, STEVEN  ATHYRA [DRAGAERA VLAD TALTOS]  C/P:3,SFWA-P:2,P
BRUST, STEVEN  BOOK OF ATHYRA, THE [DRAGAERA VLAD TALTOS]  LP
BRUST, STEVEN  BOOK OF JHEREG, THE [DRAGAERA VLAD TALTOS]  H
BRUST, STEVEN  BOOK OF TALTOS, THE [DRAGAERA VLAD TALTOS]  H:2
BRUST, STEVEN  BROKEDOWN PALACE [DRAGAERA FENARIO]  C/P:3,P
BRUST, STEVEN  COWBOY FENG’S SPACE BAR AND GRILLE  C/P:2,P
BRUST, STEVEN  DRAGON [DRAGAERA VLAD TALTOS]  C/P:2,P,H
BRUST, STEVEN  DZUR [DRAGAERA VLAD TALTOS]  H
BRUST, STEVEN  FIVE HUNDRED YEARS AFTER [DRAGAERA KHAAVREN 2]  C/P,P,H
BRUST, STEVEN/BULL, EMMA  FREEDOM & NECESSITY  C/P:2,P
BRUST, STEVEN/LINDHOLM, MEGAN  GYPSY  H:2,P
BRUST, STEVEN  ISSOLA [DRAGAERA VLAD TALTOS]  C/P,P,H
BRUST, STEVEN  JHEREG [DRAGAERA VLAD TALTOS]  C/P:2,P
BRUST, STEVEN  LORD OF CASTLE BLACK, THE [DRAGAERA ADRILANKHA 2]  H
BRUST, STEVEN  ORCA [DRAGAERA VLAD TALTOS]  C/P:3,P
BRUST, STEVEN  PATHS OF THE DEAD, THE [DRAGAERA ADRILANKHA 1]  H:2,P
BRUST, STEVEN  PHOENIX [DRAGAERA VLAD TALTOS]  C/P:2,SFWA-P:2,P
BRUST, STEVEN  PHOENIX GUARDS, THE [DRAGAERA KHAAVREN 1]  C/P,P,H
BRUST, STEVEN  SETHRA LAVODE [DRAGAERA ADRILANKHA 3]  H
BRUST, STEVEN  SUN, THE MOON, AND THE STARS, THE  H:2,SFWA,P
BRUST, STEVEN  TALTOS [DRAGAERA VLAD TALTOS]  C/P:3,P
BRUST, STEVEN  TECKLA [DRAGAERA VLAD TALTOS]  C/P:3,P
BRUST, STEVEN  TO REIGN IN HELL  SR-H,C/P:3,P
BRUST, STEVEN  YENDI [DRAGAERA VLAD TALTOS]  C/P:3,P
BRYAN, JESSICA  ACROSS A WINE-DARK SEA  P
BRYAN, JESSICA  BENEATH A SAPPHIRE SEA  P
BRYAN, KATHLEEN  SERPENT AND THE ROSE, THE [WAR OF THE ROSE 1]  LP
BRYANT, DOROTHY  KIN OF ATA ARE WAITING FOR YOU, THE  SFWA
BRYANT, EDWARD  AMONG THE DEAD  H,SFWA,P
BRYANT, EDWARD  CINNABAR  C/P,P,H
BRYANT, EDWARD  PARTICLE THEORY  C/P,P
BRYANT, EDWARD/ELLISON, HARLAN  PHOENIX WITHOUT ASHES  C/P:2,P
BRYANT, EDWARD  2076: THE AMERICAN TRICENTENNIAL  PA

AUTHOR  TITLE  CODE
BRYANT, PETER  RED ALERT  C/P:2,P
BRYCE, ROBERT  COLCHICINE FACTOR, THE  P
BSFS  BALTICON 33 PROGRAM BOOK: DAVID WEBER GOH  HA,VLPA
BUCHAN, JOHN  BLANKET OF THE DARK, THE  H
BUCHAN, JOHN  DANCING FLOOR, THE  H
BUCHAN, JOHN  WITCH WOOD  H,P
BUCHANAN, BILL  VIRUS  P
BUCHANAN, GINJER  WHITE SILENCE [HIGHLANDER]  P
BUCHEIM, ROBERT  NEW SPACE HANDBOOK  [C/RAN-P]
BUCHER-JONES, SIMON  DEATH OF ART, THE [DOCTOR WHO NEW ADVENTURES]  C/P,P
BUCHER-JONES, SIMON  GHOST DEVICES [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
BUCHER-JONES, SIMON/HALE, KELLY  GRIMM REALITY [DOCTOR WHO EIGHTH DOCTOR]  P
BUCHER-JONES, SIMON/CLAPHAM, MARK  TAKING OF PLANET 5, THE [DOCTOR WHO EIGHTH DOCTOR]  P
BUCHWALD, ART  ESTABLISHMENT IS ALIVE & WELL IN WASHINGTON, THE  [C/RAN-P]
BUCHWALD, ART  "I AM NOT A CROOK"  [C/RAN-P]
BUCK, CHARLES H.  MASTER CURE, THE  P
BUCKELL, TOBIAS S.  CRYSTAL RAIN  H
BUCKELL, TOBIAS S.  RAGAMUFFIN  H
BUCKLAND, RAYMOND  COMMITTEE, THE  P
BUCKLEY, CHRISTOPHER  LITTLE GREEN MEN  H
BUCKNER, M. M.  HYPERTHOUGHT  P
BUCKNER, M. M.  NEUROLINK  P
BUCKNER, M. M.  WAR SURF  P
BUCKNER, ROBERT  STARFIRE  C/P:2,P
BUD, BUSTER  ECHEC AUX FUSEES  FP
BUDDHA, THE  TEACHINGS OF THE COMPASSIONATE BUDDHA, THE  PA
BUDRYS, ALGIS  AMSIRS AND THE IRON THORN, THE  C/P:2,P
BUDRYS, ALGIS  BLOOD AND BURNING  C/P,P
BUDRYS, ALGIS  BUDRYS’ INFERNO  C/P:2,P
BUDRYS, ALGIS  ENTERTAINMENT  LP
BUDRYS, ALGIS  FALLING TORCH, THE  C/P,P
BUDRYS, ALGIS  FALLING TORCH (COMPLETE)  C/P,P
BUDRYS, ALGIS  FALSE NIGHT  SR-P,C/P
BUDRYS, ALGIS  HARD LANDING  P
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE  C/PA,PA
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 2  C/PA,PA
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 3  C/PA,PA
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 4  C/PA,PA
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 5  PA,SFWA-P
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 8  C/PA,PA
BUDRYS, ALGIS  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 18  PA
BUDRYS, ALGIS  MAN OF EARTH  C/P:2,P
BUDRYS, ALGIS  MICHAELMAS  H:2,P
BUDRYS, ALGIS  ROGUE MOON  C/P,SR-P,H
BUDRYS, ALGIS  SOME WILL NOT DIE  C/P,P
BUDRYS, ALGIS  SOME WILL NOT DIE (REVISED)  LP
BUDRYS, ALGIS  UNEXPECTED DIMENSION, THE  C/P:2,P
BUDRYS, ALGIS  WHO?  C/P:2,P
BUDZ, MARK  CLADE  C/P,P
BUDZ, MARK  CLADE (PROOF)  H,SR-LP
BUDZ, MARK  CRACHE  P
BUDZ, MARK  CRACHE (PROOF)  SR-P
BUDZ, MARK  IDOLIN  P
BUETTNER, ROBERT  ORPHANAGE [JASON WANDER 1]  C/P,P
BUETTNER, ROBERT  ORPHAN’S DESTINY [JASON WANDER 2]  P
BUJOLD, LOIS MCMASTER  BARRAYAR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 2]  C/P:2,SFWA-P,P,H
BUJOLD, LOIS MCMASTER  BEGUILEMENT [@SHARING KNIFE 1]  H
BUJOLD, LOIS MCMASTER  BORDERS OF INFINITY [MILES VORKOSIGAN 3]  C/P:2,SFWA-P,P
BUJOLD, LOIS MCMASTER  BROTHERS IN ARMS [MILES VORKOSIGAN 2]  C/P:2,P
BUJOLD, LOIS MCMASTER  CETAGANDA [MILES VORKOSIGAN 6]  C/P,P,H
BUJOLD, LOIS MCMASTER  CIVIL CAMPAIGN, A [MILES VORKOSIGAN 9]  H,P
BUJOLD, LOIS MCMASTER  CORDELIA’S HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1,2]  LP
BUJOLD, LOIS MCMASTER  CURSE OF CHALION, THE [CHALION 1]  H:2
BUJOLD, LOIS MCMASTER  DIPLOMATIC IMMUNITY [MILES VORKOSIGAN 10]  H,P
BUJOLD, LOIS MCMASTER  DREAMWEAVER’S DILEMMA  H
BUJOLD, LOIS MCMASTER  ETHAN OF ATHOS  C/P,P,H
BUJOLD, LOIS MCMASTER  FALLING FREE  C/P:2,SFWA-P,P
BUJOLD, LOIS MCMASTER  HALLOWED HUNT, THE [CHALION 3]  H:2
BUJOLD, LOIS MCMASTER  KOMARR [MILES VORKOSIGAN 8]  H:2,P
BUJOLD, LOIS MCMASTER  LEGACY [@SHARING KNIFE 2]  H
BUJOLD, LOIS MCMASTER  MEMORY [MILES VORKOSIGAN 7]  H:2,P
BUJOLD, LOIS MCMASTER  MIRROR DANCE [MILES VORKOSIGAN 5]  H,SFWA,P
BUJOLD, LOIS MCMASTER  PALADIN OF SOULS [CHALION 2]  H
BUJOLD, LOIS MCMASTER  SHARDS OF HONOR [MILES VORKOSIGAN 0/CORDELIA NAISMITH 1]  C/P:2,P,H

The MITSFS Pinkdex  September 24, 2007  Page 60
AUTHOR  TITLE  CODE
BUJOLD, LOIS MCMASTER  SPIRIT RING, THE  H:2,P
BUJOLD, LOIS MCMASTER  VOR GAME, THE [MILES VORKOSIGAN 4]  C/P:2,P
BUJOLD, LOIS MCMASTER  VORKOSIGAN’S GAME [MILES VORKOSIGAN 3,4]  H
BUJOLD, LOIS MCMASTER  WARRIOR’S APPRENTICE, THE [MILES VORKOSIGAN 1]  C/P:2,P
BUJOLD, LOIS MCMASTER  WARRIOR’S APPRENTICE, THE (REVISED) [MILES VORKOSIGAN 1]  H

BUJOLD, LOIS MCMASTER/ GREEN, ROLAND J.

WOMEN AT WAR

LPA

BULFINCH, THOMAS  AGE OF CHIVALRY, THE/LEGENDS OF CHARLEMAGNE, THE  [C/RAN-P]
BULFINCH, THOMAS  BULFINCH’S MYTHOLOGY: THE AGE OF FABLE  R/REF-H
BULGAKOV, MIKHAIL  HEART OF A DOG, THE  LP
BULGAKOV, MIKHAIL  MASTER AND MARGARITA, THE  P
BULIS, CHRISTOPHER  CITY AT WORLD’S END [DOCTOR WHO PAST DOCTORS]  P
BULIS, CHRISTOPHER  DEVICE OF DEATH, A [DOCTOR WHO MISSING ADVENTURES]  C/P,P
BULIS, CHRISTOPHER  EYE OF THE GIANT, THE [DOCTOR WHO MISSING ADVENTURES]  C/P,P
BULIS, CHRISTOPHER  IMPERIAL MOON [DOCTOR WHO PAST DOCTORS]  C/P,P
BULIS, CHRISTOPHER  PALACE OF THE RED SUN [DOCTOR WHO PAST DOCTORS]  P
BULIS, CHRISTOPHER  SHADOWMIND [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
BULIS, CHRISTOPHER  SORCERER’S APPRENTICE, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
BULIS, CHRISTOPHER  STATE OF CHANGE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
BULIS, CHRISTOPHER  TEMPEST [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
BULIS, CHRISTOPHER  TWILIGHT OF THE GODS [DOCTOR WHO MISSING ADVENTURES]  C/P,P
BULIS, CHRISTOPHER  ULTIMATE TREASURE, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
BULIS, CHRISTOPHER  VANDERDEKEN’S CHILDREN [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
BULL, EMMA  BONE DANCE  H:2,SFWA-P:2,P
BULL, EMMA/SHETTERLY, WILL  DOUBLE FEATURE  SR-H,H
BULL, EMMA  FALCON  C/P:2,SFWA-P,P
BULL, EMMA  FINDER [BORDERLANDS 4]  C/P:2,P
BULL, EMMA  WAR FOR THE OAKS  C/P:2,P
BULL, R. C.  PERTURBED SPIRITS  PA
BULLETT, GERALD  MR. GODLY BESIDE HIMSELF  H
BULMER, H. K.  CHALLENGE  SR-P
BULMER, H. K.  EMPIRE OF CHAOS  SR-P
BULMER, H. K.  GALACTIC INTRIGUE  SR-P
BULMER, H. K.  STARS ARE OURS, THE  H
BULMER, H. K.  WORLD AFLAME  SR-P
BULMER, KENNETH  BEHOLD THE STARS  DM131
BULMER, KENNETH  BEYOND THE SILVER SKY  DD507
BULMER, KENNETH  BLAZON  P
BULMER, KENNETH  CHANGELING WORLDS, THE  DD369
BULMER, KENNETH  CHARIOTS OF RA, THE [KEYS TO THE DIMENSIONS 7]  D10293
BULMER, KENNETH  CITY UNDER THE SEA  DD255
BULMER, KENNETH  CYCLE OF NEMESIS  C/P:2,P
BULMER, KENNETH  DEMONS, THE  SR-P
BULMER, KENNETH  DEMONS’ WORLD  DF289
BULMER, KENNETH  DIAMOND CONTESSA, THE [KEYS TO THE DIMENSIONS 8]  P
BULMER, KENNETH  DOOMSDAY MEN, THE  H:2,P
BULMER, KENNETH  EARTH GODS ARE COMING, THE  DD453
BULMER, KENNETH  ELECTRIC SWORD-SWALLOWERS, THE  D05595
BULMER, KENNETH  HUNTERS OF JUNDAGAI, THE [KEYS TO THE DIMENSIONS 6]  D68310
BULMER, KENNETH  INSANE CITY, THE  C/P:2,P
BULMER, KENNETH  KANDAR  C/P,P
BULMER, KENNETH  KEY TO IRUNIUM, THE [KEYS TO THE DIMENSIONS 2]  DH020
BULMER, KENNETH  KEY TO VENUDINE, THE [KEYS TO THE DIMENSIONS 3]  DH065
BULMER, KENNETH  LAND BEYOND THE MAP [KEYS TO THE DIMENSIONS 1]  DM111
BULMER, KENNETH  MILLION YEAR HUNT, THE  DF285
BULMER, KENNETH  NEW WRITINGS IN SF 22  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 23  SR-PA,C/PA
BULMER, KENNETH  NEW WRITINGS IN SF 24  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 25  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 26  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 27  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 28  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 29  SR-PA
BULMER, KENNETH  NEW WRITINGS IN SF 30  SR-PA
BULMER, KENNETH  NO MAN’S WORLD  DF104
BULMER, KENNETH  ON THE SYMB-SOCKET CIRCUIT  C/P:2,P
BULMER, KENNETH  ROLLER COASTER WORLD  P
BULMER, KENNETH  SECRET OF ZI, THE  DD331
BULMER, KENNETH  SHIPS OF DUROSTORUM, THE [KEYS TO THE DIMENSIONS 5]  D76096
BULMER, KENNETH  STAR VENTURERS, THE  D22600
BULMER, KENNETH  SWORDS OF THE BARBARIANS  C/P,P
BULMER, KENNETH  TO OUTRUN DOOMSDAY  C/P:2,P
BULMER, KENNETH  WIND OF LIBERTY, THE  SR-P
BULMER, KENNETH  WIZARD OF STARSHIP POSEIDON, THE  DF209

AUTHOR  TITLE  CODE
BULMER, KENNETH  WIZARDS OF SENCHURIA, THE [KEYS TO THE DIMENSIONS 4]  D12140
BULMER, KENNETH  WORLDS FOR THE TAKING  C/P:2,P
BULYCHEV, KIR  HALF A LIFE  H,LP
BULYCHEV, KIR  THOSE WHO SURVIVE (PROOF)  H,SR-LP
BUNCH, CHRIS  CORSAIR  P
BUNCH, CHRIS  DARKNESS OF GOD, THE [@SHADOW WARRIOR #3]  P
BUNCH, CHRIS  DEMON KING, THE [DAMASTES 2]  LP
BUNCH, CHRIS  DOG FROM HELL, THE [@STAR RISK,LTD. 4]  P
BUNCH, CHRIS  DOUBLECROSS PROGRAM, THE [STAR RISK,LTD. 3]  P
BUNCH, CHRIS  EMPIRE STONE, THE  P
BUNCH, CHRIS  FIREMASK [LAST LEGION 2]  C/P,P
BUNCH, CHRIS  HOMEFALL [LAST LEGION 4]  C/P:2,P
BUNCH, CHRIS  HUNT THE HEAVENS [@SHADOW WARRIOR #2]  P
BUNCH, CHRIS  KNIGHTHOOD OF THE DRAGON [@DRAGONMASTER #2]  LP
BUNCH, CHRIS  LAST BATTLE, THE [@DRAGONMASTER #3]  LP
BUNCH, CHRIS  LAST LEGION, THE [LAST LEGION 1]  C/P,P
BUNCH, CHRIS  SCOUNDREL WORLDS, THE [STAR RISK,LTD. 2]  P
BUNCH, CHRIS  SEER KING, THE [DAMASTES 1]  C/P,P
BUNCH, CHRIS  STAR RISK, LTD. [STAR RISK,LTD. 1]  P
BUNCH, CHRIS  STORM FORCE [LAST LEGION 3]  C/P,P
BUNCH, CHRIS  STORM OF WINGS [@DRAGONMASTER #1]  LP
BUNCH, CHRIS  WARRIOR KING, THE [DAMASTES 3]  P
BUNCH, CHRIS  WIND AFTER TIME, THE [@SHADOW WARRIOR #1]  C/P,P
BUNCH, DAVID R.  MODERAN  C/P:2,P
BURCH, RALPH  DUPLICATE LOVERS  SR-L
BURDEKIN, KATHARINE  BURNING RING, THE  H
BURDEKIN, KATHERINE  END OF THIS DAY’S BUSINESS, THE  LP
BURDICK, EUGENE/WHEELER, HARVEY  FAIL-SAFE  C/P:2,P
BURFORD, LOLAH  VISION OF STEPHEN, THE  C/P:2,P
BURGAUER, STEVEN  BRAZEN RULE, THE [MATTHEWS 1]  H:2,LP
BURGAUER, STEVEN  FORNAX [MATTHEWS 3]  H,LP
BURGAUER, STEVEN  IN THE SHADOW OF OMEN [MATTHEWS 4]  LP
BURGAUER, STEVEN  LAST AMERICAN, THE [MATTHEWS 2]  H,LP
BURGAUER, STEVEN  TREACHERY ON THE DARK SIDE [MATTHEWS 6]  LP
BURGER, DIONYS  SPHERELAND  H,LP
BURGER, JOANNE  SCIENCE FICTION BOOKS PUBLISHED IN 1967  R/IND
BURGER, JOANNE  SF PUBLISHED IN 1968  R/IND
BURGER, JOANNE  SF PUBLISHED IN 1969  R/IND:2
BURGER, JOANNE  SF PUBLISHED IN 1970  R/IND
BURGER, JOANNE  SF PUBLISHED IN 1971  R/IND
BURGER, JOANNE  SF PUBLISHED IN 1972  R/IND
BURGER, JOANNE  SF PUBLISHED IN 1973  R/IND
BURGESS, ANTHONY  CLOCKWORK ORANGE, A  C/P:2,LP
BURGESS, ANTHONY  DEVIL’S MODE, THE  P
BURGESS, ANTHONY  END OF THE WORLD NEWS, THE  H
BURGESS, ANTHONY  HONEY FOR THE BEARS  C/P,P
BURGESS, ANTHONY  1985  H
BURGESS, ANTHONY  NOTHING LIKE THE SUN  C/P,P
BURGESS, ANTHONY  WANTING SEED, THE  C/P:2,P
BURGESS, MELVIN  BLOODTIDE  H,P
BURGO, JOSEPH  LIGHTS OF BARBRIN, THE  P
BURKA, LAUREN P.  MATE  SR-L
BURKE, JOHN  MOON ZERO TWO  C/P,P
BURKE, JOHN  TALES OF UNEASE  PA
BURKE, JONATHAN  ECHOING WORLDS, THE  SR-P
BURKE, JONATHAN  HOTEL COSMOS  SR-P
BURKE, JONATHAN  REVOLT OF THE HUMANS  SR-P
BURKE, JONATHAN  TWILIGHT OF REASON  SR-P
BURKE, RUSTY  SHORT BIOGRAPHY OF ROBERT E. HOWARD, A  R/REF-H
BURKE, THOMAS  MORE LIMEHOUSE NIGHTS  H
BURKETT, WILLIAM R. JR.  BLOOD LINES  P
BURKETT, WILLIAM R. JR.  BLOODSPORT  P
BURKETT, WILLIAM R. JR.  SLEEPING PLANET  C/P,P,H
BURKHOLZ, HERBERT  SENSITIVES, THE  P
BURKS, ARTHUR J.  BLACK MEDICINE  SR-H
BURKS, ARTHUR J.  LOOK BEHIND YOU!  SR-H
BURLAND, HARRIS  DACOBRA, OR THE WHITE PRIESTS OF AHRIMAN  H
BURNES, CAROLINE  BEWITCHING FAMILIAR  P
BURNETT, WHIT/BURNETT, HALLIE  19 TALES OF TERROR  PA
BURNETT, WHIT/BURNETT, HALLIE  THINGS WITH CLAWS  PA
BURNFORD, SHEILA  MR. NOAH AND THE SECOND FLOOD  P
BURNS, ALMA  WITCHES OF TURNSTONE BAY, THE  P
BURNS, LAURA J./METZ, MELINDA  APOCALYPSE MEMORIES [@BUFFY]  C/P,P
BURNS, LAURA J.  INHERIT THE WITCH [@CHARMED]  P

The MITSFS Pinkdex  September 24, 2007  Page 62
AUTHOR  TITLE  CODE
BURNS, LAURA J.  SWEET TALKIN’ DEMON [@CHARMED]  P
BURNS, STEPHEN L.  CALL FROM A DISTANT SHORE  P
BURNS, STEPHEN L.  FLESH AND SILVER  C/P:2,P
BURRISS, ELI EDWARD  TABOO, MAGIC, SPIRITS  [C/RAN-H]
BURROUGHS, EDGAR RICE  AT THE EARTH’S CORE [PELLUCIDAR #1]  H:2,@P

BURROUGHS, EDGAR RICE AT THE EARTH’S CORE [PELLUCIDAR 1,2,3]/

PELLUCIDAR [PELLUCIDAR 1,2,3]/

TANAR OF PELLUCIDAR [PELLUCIDAR 1,2,3]

LP

BURROUGHS, EDGAR RICE  BACK TO THE STONE AGE [PELLUCIDAR #5]  C/P,@P,H
BURROUGHS, EDGAR RICE  BANDIT OF HELL’S BEND, THE  P
BURROUGHS, EDGAR RICE  BEASTS OF TARZAN, THE [TARZAN #3]  C/P,@P,H
BURROUGHS, EDGAR RICE  BEYOND THE FARTHEST STAR  C/P:2,P
BURROUGHS, EDGAR RICE  CARSON OF VENUS [VENUS #3]  C/P:2,@P
BURROUGHS, EDGAR RICE  CAVE GIRL, THE  H:2,SR-P,P
BURROUGHS, EDGAR RICE  CHESSMEN OF MARS, THE [BARSOOM #5]  C/P:2,@P
BURROUGHS, EDGAR RICE  DEPUTY SHERIFF OF COMMANCHE COUNTY, THE  P
BURROUGHS, EDGAR RICE  ESCAPE ON VENUS [VENUS #4]  C/P:2,@P
BURROUGHS, EDGAR RICE  ETERNAL SAVAGE, THE  C/P:2,P
BURROUGHS, EDGAR RICE  FIGHTING MAN OF MARS, A [BARSOOM #7]  C/P,@P,H
BURROUGHS, EDGAR RICE  GIRL FROM FARRIS’S, THE  SR-VLP
BURROUGHS, EDGAR RICE  GODS OF MARS, THE [BARSOOM #2]  C/P,@P,H
BURROUGHS, EDGAR RICE  GODS OF MARS, THE [BARSOOM 2]/WARLORD OF MARS, THE [BARSOOM 3] H:2
BURROUGHS, EDGAR RICE  I AM A BARBARIAN  C/P,P,H
BURROUGHS, EDGAR RICE  JOHN CARTER OF MARS [BARSOOM #11]  C/P:2,@P
BURROUGHS, EDGAR RICE  JOHN CARTER OF MARS (ORIGINAL) [BARSOOM 11]  SR-L
BURROUGHS, EDGAR RICE  JUNGLE TALES OF TARZAN [TARZAN #6]  H:2,@P
BURROUGHS, EDGAR RICE  LAD AND THE LION, THE  C/P:2,P
BURROUGHS, EDGAR RICE  LAND OF HIDDEN MEN, THE  C/P:2,P
BURROUGHS, EDGAR RICE  LAND OF TERROR [PELLUCIDAR #6]  C/P:2,@P
BURROUGHS, EDGAR RICE  LAND THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 1]  C/P:2,P
BURROUGHS, EDGAR RICE  LAND THAT TIME FORGOT, THE/MOON MAID, THE  LP

BURROUGHS, EDGAR RICE LAND THAT TIME FORGOT, THE (OMNIBUS EDITION)

[LAND THAT TIME FORGOT 1,2,3]

H:2

BURROUGHS, EDGAR RICE  LLANA OF GATHOL [BARSOOM #10]  C/P:2,@P
BURROUGHS, EDGAR RICE  LLANA OF GATHOL [BARSOOM 10]/JOHN CARTER OF MARS [BARSOOM 11] H:2
BURROUGHS, EDGAR RICE  LOST CONTINENT, THE  C/P:2,P
BURROUGHS, EDGAR RICE  LOST ON VENUS [VENUS #2]  C/P:2,@P
BURROUGHS, EDGAR RICE  MAD KING, THE  C/P,P
BURROUGHS, EDGAR RICE  MAN-EATER, THE  SR-LP
BURROUGHS, EDGAR RICE  MASTERMIND OF MARS, THE [BARSOOM #6]  C/P,@P,H

BURROUGHS, EDGAR RICE MASTERMIND OF MARS, THE [BARSOOM 6]/ FIGHTING MAN OF MARS, A [BARSOOM 7]

H:2

BURROUGHS, EDGAR RICE  MONSTER MEN, THE  C/P:2,P
BURROUGHS, EDGAR RICE  MOON MAID, THE  C/P:2,P
BURROUGHS, EDGAR RICE  MOON MEN, THE  C/P:2,P
BURROUGHS, EDGAR RICE  MUCKER, THE  H
BURROUGHS, EDGAR RICE  OAKDALE AFFAIR, THE  C/P:2,P
BURROUGHS, EDGAR RICE  OUT OF TIME’S ABYSS [LAND THAT TIME FORGOT 3]  C/P:2,P
BURROUGHS, EDGAR RICE  OUTLAW OF TORN, THE  C/P,P,H
BURROUGHS, EDGAR RICE  PELLUCIDAR [PELLUCIDAR #2]  C/P,@P,H
BURROUGHS, EDGAR RICE  PEOPLE THAT TIME FORGOT, THE [LAND THAT TIME FORGOT 2]  C/P:2,P
BURROUGHS, EDGAR RICE  PIRATES OF VENUS [VENUS #1]  C/P:2,@P
BURROUGHS, EDGAR RICE  PRINCESS OF MARS, A [BARSOOM #1]  H:3,@P,SR-H

BURROUGHS, EDGAR RICE PRINCESS OF MARS, A [BARSOOM 1]/

FIGHTING MAN OF MARS, A [BARSOOM 7]

LP

BURROUGHS, EDGAR RICE  RETURN OF TARZAN, THE [TARZAN #2]  H:2,@P
BURROUGHS, EDGAR RICE  RIDER, THE  C/P:2,P
BURROUGHS, EDGAR RICE  SAVAGE PELLUCIDAR [PELLUCIDAR #7]  C/P:2,@P
BURROUGHS, EDGAR RICE  SCIENCE FICTION CLASSICS [BARSOOM 4,5,6/PELLUCIDAR 2,3]  H
BURROUGHS, EDGAR RICE  SON OF TARZAN, THE [TARZAN #4]  C/P,@P,H
BURROUGHS, EDGAR RICE  SWORDS OF MARS [BARSOOM #8]  C/P:2,@P
BURROUGHS, EDGAR RICE  SWORDS OF MARS [BARSOOM 8]/SYNTHETIC MEN OF MARS [BARSOOM 9] H:2
BURROUGHS, EDGAR RICE  SYNTHETIC MEN OF MARS [BARSOOM #9]  C/P:2,@P
BURROUGHS, EDGAR RICE  TANAR OF PELLUCIDAR [PELLUCIDAR #3]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE ANT-MEN [TARZAN #10]  H:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE CASTAWAYS [TARZAN #24]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE CITY OF GOLD [TARZAN #16]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE FORBIDDEN CITY [TARZAN #20]  C/P,@P,H
BURROUGHS, EDGAR RICE  TARZAN AND "THE FOREIGN LEGION" [TARZAN #22]  C/P,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE GOLDEN LION [TARZAN #9]  H:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE JEWELS OF OPAR [TARZAN #5]  H:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE LEOPARD MEN [TARZAN #18]  C/P,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE LION MAN [TARZAN #17]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN AND THE LOST EMPIRE [TARZAN #12]  C/P,@P,H

The MITSFS Pinkdex  September 24, 2007  Page 63
AUTHOR  TITLE  CODE
BURROUGHS, EDGAR RICE  TARZAN AND THE MADMAN [TARZAN #23]  C/P,@P
BURROUGHS, EDGAR RICE  TARZAN AT THE EARTH’S CORE [TARZAN #13/PELLUCIDAR 4]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN, LORD OF THE JUNGLE [TARZAN #11]  C/P,@P,H:2
BURROUGHS, EDGAR RICE  TARZAN OF THE APES [TARZAN #1]  H:2,@P
BURROUGHS, EDGAR RICE  TARZAN THE INVINCIBLE [TARZAN #14]  C/P:2,@P
BURROUGHS, EDGAR RICE  TARZAN THE MAGNIFICENT [TARZAN #21]  C/P,@P
BURROUGHS, EDGAR RICE  TARZAN THE TERRIBLE [TARZAN #8]  C/P,@P,H
BURROUGHS, EDGAR RICE  TARZAN THE UNTAMED [TARZAN #7]  H:2,@P
BURROUGHS, EDGAR RICE  TARZAN TRIUMPHANT [TARZAN #15]  C/P,@P,H
BURROUGHS, EDGAR RICE  TARZAN TWINS, THE [TARZAN]  H
BURROUGHS, EDGAR RICE  TARZAN’S QUEST [TARZAN #19]  @P
BURROUGHS, EDGAR RICE  3 MARTIAN NOVELS [BARSOOM #4,5,6]  H:2,LP
BURROUGHS, EDGAR RICE  THUVIA, MAID OF MARS [BARSOOM #4]  H:2,@P

BURROUGHS, EDGAR RICE THUVIA, MAID OF MARS [BARSOOM 4]/ CHESSMEN OF MARS, THE [BARSOOM 5]

H:2

BURROUGHS, EDGAR RICE  WARLORD OF MARS, THE [BARSOOM #3]  C/P,@P,H:2
BURROUGHS, EDGAR RICE  WIZARD OF VENUS, THE [VENUS #5]  @P
BURROUGHS, EDGAR RICE  WIZARD OF VENUS, THE/PIRATE BLOOD  C/P:2
BURROUGHS, JOHN COLEMAN  TREASURE OF THE BLACK FALCON  C/P:2,P
BURROUGHS, WILLIAM S.  CITIES OF THE RED NIGHT  LP
BURROUGHS, WILLIAM S.  EXTERMINATOR!  LP
BURROUGHS, WILLIAM S.  NOVA EXPRESS  H
BURROUGHS, WILLIAM S.  WILD BOYS, THE  LP
BURTIS, THOMSON  HAUNTED AIRWAYS  H
BURTON, D. SHANE  BLACK STORM RISING [ORIANUS CREATION 1]  LP
BURTON, D. SHANE  WHEN WORLDS RAGE [ORIANUS CREATION 2]  LP
BURTON, JOSH MICHAEL  JOURNEY INTO THE GATEWAY OF DIMENSIONS [TAVAN 1]  LP
BURTON, LEVAR  AFTERMATH  C/P,P
BURY, STEPHEN  INTERFACE  C/P,LP
BUSBY, F. M.  ALIEN DEBT, THE [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  ALL THESE EARTHS  C/P:2,P
BUSBY, F. M.  ARROW FROM EARTH  P
BUSBY, F. M.  BREEDS OF MAN, THE  C/P:2,P
BUSBY, F. M.  CAGE A MAN [DEMU 1]  H:2,P
BUSBY, F. M.  DEMU TRILOGY, THE [DEMU 1,2,3]  C/P:2,P
BUSBY, F. M.  GETTING HOME  C/P,P
BUSBY, F. M.  ISLANDS OF TOMORROW  P
BUSBY, F. M.  LONG VIEW, THE  C/P,P,H
BUSBY, F. M.  PROUD ENEMY, THE [DEMU 2]  C/P:2,P
BUSBY, F. M.  REBEL DYNASTY 2, THE [HULZEIN DYNASTY]  P
BUSBY, F. M.  REBEL’S QUEST [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  REBELS’ SEED [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  RISSA AND TREGARE [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  RISSA KERGUELEN [HULZEIN DYNASTY]  C/P,P,H
BUSBY, F. M.  SINGULARITY PROJECT, THE  C/P,P
BUSBY, F. M.  SLOW FREIGHT  C/P:2,P
BUSBY, F. M.  STAR REBEL [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  TRIAD WORLDS, THE  P
BUSBY, F. M.  YOUNG RISSA [HULZEIN DYNASTY]  C/P:2,P
BUSBY, F. M.  ZELDE M’TANA [HULZEIN DYNASTY]  C/P,P
BUSH, ANNE KELLEHER  CHILDREN OF ENCHANTMENT [POWER AND THE PATTERN 2]  P
BUSH, ANNE KELLEHER  DAUGHTER OF PROPHECY [POWER AND THE PATTERN 1]  C/P,P
BUSH, ANNE KELLEHER  KNIGHT, THE HARP AND THE MAIDEN, THE  C/P,P
BUSH, ANNE KELLEHER  MISBEGOTTEN KING, THE  C/P,P
BUSHYAGER, LINDA E.  MASTER OF HAWKS  C/P,P
BUSHYAGER, LINDA E.  SPELLSTONE OF SHALTUS, THE  C/P:2,P
BUTCHER, JIM  ACADEM’S FURY [@CODEX ALERA #2]  H
BUTCHER, JIM  BLOOD RITES [@DRESDEN FILES #6]  C/P,P
BUTCHER, JIM  CURSOR’S FURY [@CODEX ALERA #3]  H:2
BUTCHER, JIM  DEAD BEAT [DRESDEN FILES 7]  H
BUTCHER, JIM  DEATH MASKS [@DRESDEN FILES #5]  C/P,P
BUTCHER, JIM  FOOL MOON [@DRESDEN FILES #2]  C/P:2,P
BUTCHER, JIM  FURIES OF CALDERON [CODEX ALERA 1]  H
BUTCHER, JIM  GRAVE PERIL [@DRESDEN FILES #3]  C/P,P
BUTCHER, JIM  PROVEN GUILTY [DRESDEN FILES 8]  H
BUTCHER, JIM  STORM FRONT [@DRESDEN FILES #1]  C/P,P
BUTCHER, JIM  SUMMER KNIGHT [@DRESDEN FILES #4]  C/P,P
BUTCHER, JIM  WHITE NIGHT [DRESDEN FILES 9]  H:2
BUTLER, IVAN  HORROR IN THE CINEMA  R/REF-P
BUTLER, JACK  NIGHTSHADE  SFWA
BUTLER, OCTAVIA E.  ADULTHOOD RITES [XENOGENESIS 2]  H:2,P
BUTLER, OCTAVIA E.  BLOOD CHILD  H
BUTLER, OCTAVIA E.  CLAY’S ARK  C/P,SFWA,P,H
BUTLER, OCTAVIA E.  DAWN [XENOGENESIS 1]  H:2,P

The MITSFS Pinkdex  September 24, 2007  Page 64
AUTHOR  TITLE  CODE
BUTLER, OCTAVIA E.  FLEDGLING  H
BUTLER, OCTAVIA E.  IMAGO [XENOGENESIS 3]  C/P,P,H
BUTLER, OCTAVIA E.  KINDRED  C/P,P,H
BUTLER, OCTAVIA E.  MIND OF MY MIND  C/P,SFWA,P,H
BUTLER, OCTAVIA E.  PARABLE OF THE SOWER  C/P,P,H
BUTLER, OCTAVIA E.  PARABLE OF THE TALENTS  C/P,LP,H
BUTLER, OCTAVIA E.  PATTERNMASTER  C/P,SFWA,P,H
BUTLER, OCTAVIA E.  SURVIVOR  C/P,SFWA,P,H
BUTLER, OCTAVIA E.  ULTIMA GENESI (ITALIAN) [XENOGENESIS 1]  FP
BUTLER, OCTAVIA E.  WILD SEED  H:2,SFWA,P
BUTLER, OCTAVIA E.  XENOGENESIS [XENOGENESIS 1,2,3]  H:2
BUTLER, SAMUEL  EREWHON  P
BUTTERWORTH, MICHAEL  EDGE OF THE INFINITE, THE [@SPACE: 1999 YEAR 2 #6]  P
BUTTERWORTH, MICHAEL  MIND-BREAKS OF SPACE [@SPACE: 1999 YEAR 2 #2]  C/P,P
BUTTERWORTH, MICHAEL  PLANETS OF PERIL [@SPACE: 1999 YEAR 2 #1]  C/P,P
BUTTERWORTH, MICHAEL  PSYCHOMORPH, THE [@SPACE: 1999 YEAR 2 #4]  C/P,P
BUTTERWORTH, MICHAEL  QUEENS OF DELIRIA  SR-P
BUTTERWORTH, MICHAEL  SPACE: 1999 YEAR TWO [SPACE: 1999 YEAR 2]  SR-VLH
BUTTERWORTH, MICHAEL  SPACE-JACKERS, THE [@SPACE: 1999 YEAR 2 #3]  P
BUTTERWORTH, MICHAEL  TIME FIGHTERS, THE [@SPACE: 1999 YEAR 2 #5]  P
BUTTERWORTH, OLIVER  ENORMOUS EGG, THE  P
BUTTERWORTH, OLIVER  TROUBLE WITH JENNY’S EAR, THE  H
BYATT, A. S.  LITTLE BLACK BOOK OF STORIES  LP
BYERS, EDWARD A.  BABYLON GATE, THE  C/P,P
BYERS, EDWARD A.  LONG FORGETTING, THE  C/P:2,P
BYERS, RICHARD LEE  BLACK BOUQUET, THE [@FORGOTTEN REALMS ROGUES]  P
BYERS, RICHARD LEE  CARAVAN OF SHADOWS [WORLD OF DARKNESS WRAITH]  P

BYERS, RICHARD LEE/PETRUCHA, STEFAN CHILDREN OF GAIA [WORLD OF DARKNESS WEREOLF 5]/

UKTENA [WORLD OF DARKNESS WEREOLF 5]

PA

BYERS, RICHARD LEE  DISSOLUTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #1]  P
BYERS, RICHARD LEE  ENEMY WITHIN, THE [@WARHAMMER]  P
BYERS, RICHARD LEE  FORBIDDEN [SWORD & SORCERY DEAD GOD 3]  P
BYERS, RICHARD LEE  FORSAKEN [SWORD & SORCERY DEAD GOD 1]  P
BYERS, RICHARD LEE  FORSWORN [SWORD & SORCERY DEAD GOD 2]  P

BYERS, RICHARD LEE/EMERY, CLAYTON/ GREENWOOD, ED/GROSS, DAVE/ KEMP, PAUL S./SMEDMAN, LISA/ WHITNEY-ROBINSON, VORONICA HALLS OF STORMWEATHER, THE [@FORGOTTEN REALMS SEMBIA]

P

BYERS, RICHARD LEE  NETHERWORLD [WORLD OF DARKNESS VAMPIRE]  P
BYERS, RICHARD LEE  ON A DARKLING PLAIN  P
BYERS, RICHARD LEE  QUEEN OF THE DEPTHS [@FORGOTTEN REALMS PRIESTS]  P
BYERS, RICHARD LEE  RAGE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #1]  P
BYERS, RICHARD LEE  RITE, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #2]  P
BYERS, RICHARD LEE  RUIN, THE [@FORGOTTEN REALMS YEAR OF ROGUE DRAGONS #3]  P
BYERS, RICHARD LEE  SHATTERED MASK, THE [@FORGOTTEN REALMS SEMBIA]  P
BYERS, RICHARD LEE  UNCLEAN [@FORGOTTEN REALMS HAUNTED LANDS 1]  P
BYERS, RICHARD LEE  VAMPIRE’S APPRENTICE, THE  P
BYFIELD, BARBARA NINDE  BOOK OF WEIRD, THE  R/REF-H
BYLINSKY, GENE  LIFE IN DARWIN’S UNIVERSE  SFWA
BYRNE, ROBERT  TUNNEL, THE  P
BYRNE, STUART J.  ALPHA TRAP, THE  P

AUTHOR  TITLE  CODE
CABELL, JAMES BRANCH  CHIVALRY  H
CABELL, JAMES BRANCH  CREAM OF THE JEST, THE  P
CABELL, JAMES BRANCH  DOMNEI  C/P:2,P
CABELL, JAMES BRANCH  FIGURES OF EARTH  C/P,P,H
CABELL, JAMES BRANCH  HIGH PLACE, THE  C/P:2,P
CABELL, JAMES BRANCH  JURGEN  H:2,SR-P
CABELL, JAMES BRANCH  LETTERS OF JAMES BRANCH CABELL, THE  R/REF-H
CABELL, JAMES BRANCH  LINE OF LOVE, THE  LP
CABELL, JAMES BRANCH  SILVER STALLION, THE  C/P,P,H
CABELL, JAMES BRANCH  SOMETHING ABOUT EVE  C/P,P
CABELL, JAMES BRANCH  WITCH-WOMAN: A TRILOGY ABOUT HER, THE  H
CACEK, P. D.  CANYONS  H
CADIGAN, PAT  DERVISH IS DIGITAL [KONSTANTIN 2]  H
CADIGAN, PAT  DIRTY WORK  H
CADIGAN, PAT  FOOLS  C/P,P
CADIGAN, PAT  MINDPLAYERS  C/P:2,P
CADIGAN, PAT  MINDPLAYERS (PROOF)  SR-LP
CADIGAN, PAT  PATTERNS  LP
CADIGAN, PAT  PROMISED LAND [LOST IN SPACE]  P
CADIGAN, PAT  SYNNERS  C/P:2,P
CADIGAN, PAT  TEA FROM AN EMPTY CUP [KONSTANTIN 1]  H,P
CADNUM, MICHAEL  JUDAS GLASS, THE  H
CADNUM, MICHAEL  SAINT PETER’S WOLF  H
CADY, JACK  NIGHT WE BURIED ROAD DOG, THE  H
CADY, JACK  OFF SEASON, THE  H
CAHILL, ROBERT ELLIS  NEW ENGLAND’S MARVELOUS MONSTERS  [C/RAN-H]
CAHILL, ROBERT ELLIS  NEW ENGLAND’S THINGS THAT GO BUMP IN THE NIGHT  [C/RAN-H]
CAHILL, ROBERT ELLIS  NEW ENGLAND’S WITCHES AND WIZARDS  [C/RAN-H]
CAIDIN, MARTIN  AQUARIUS MISSION  C/P,P
CAIDIN, MARTIN  BEAMRIDERS!  C/P:2,SFWA-P,P
CAIDIN, MARTIN  CYBORG [@CYBORG #1]  C/P:2,P
CAIDIN, MARTIN  CYBORG IV [SIX MILLION DOLLAR MAN 6]  H,@P
CAIDIN, MARTIN  DARK MESSIAH [MESSIAH STONE 2]  C/P:2,P
CAIDIN, MARTIN  ENCOUNTER THREE  C/P,P
CAIDIN, MARTIN  EXIT EARTH  C/P:2,P
CAIDIN, MARTIN  FINAL COUNTDOWN, THE  P
CAIDIN, MARTIN  FOUR CAME BACK  C/P:2,SFWA-P,P,H
CAIDIN, MARTIN  GOD MACHINE, THE  C/P:2,SFWA-P,P
CAIDIN, MARTIN  HIGH CRYSTAL [SIX MILLION DOLLAR MAN 3]  H:2,@P
CAIDIN, MARTIN  INDIANA JONES AND THE SKY PIRATES [INDIANA JONES]  C/P,P
CAIDIN, MARTIN  INDIANA JONES AND THE WHITE WITCH [INDIANA JONES]  P
CAIDIN, MARTIN  KILLER STATION  P
CAIDIN, MARTIN  LAST FATHOM, THE  P
CAIDIN, MARTIN  LIFE IN THE FUTURE, A [BUCK ROGERS]  H
CAIDIN, MARTIN  MANFAC  C/P,P
CAIDIN, MARTIN  MAROONED (ORIGINAL)  SR-P
CAIDIN, MARTIN  MAROONED (REVISED)  C/P,P
CAIDIN, MARTIN  MENDELOV CONSPIRACY, THE  P
CAIDIN, MARTIN  MESSIAH STONE, THE [MESSIAH STONE 1]  C/P,P
CAIDIN, MARTIN  NO MAN’S WORLD  H
CAIDIN, MARTIN  OPERATION NUKE [@CYBORG #2]  P
CAIDIN, MARTIN  PERDIDO NO ESPACO 1  FP
CAIDIN, MARTIN  PERDIDO NO ESPACO 2  FP
CAIDIN, MARTIN  PRISON SHIP  C/P:2,SFWA-P,P
CAIDIN, MARTIN  STAR BRIGHT  C/P:2,SFWA-P:2,P
CAIDIN, MARTIN  THREE CORNERS TO NOWHERE  C/P,P
CAIDIN, MARTIN  ZOBOA  P
CAIN, ROBERT  CAPO’S REVENGE [@CYBERNARC #4]  P
CAINE, GEOFFREY  CURSE OF THE VAMPIRE  P
CAINE, RACHEL  CHILL FACTOR [WEATHER WARDEN 3]  C/P,P
CAINE, RACHEL  DEAD GIRLS’ DANCE, THE [@MORGANVILLE VAMPIRES 2]  P
CAINE, RACHEL  FIRESTORM [WEATHER WARDEN 5]  P
CAINE, RACHEL  GLASS HOUSES [@MORGANVILLE VAMPIRES 1]  P
CAINE, RACHEL  HEAT STROKE [WEATHER WARDEN 2]  P
CAINE, RACHEL  ILL WIND [WEATHER WARDEN 1]  P
CAINE, RACHEL  WINDFALL [WEATHER WARDEN 4]  P
CAIRD, JANET  LOCH, THE  P
CALDECOTT, MOYRA  SHADOW ON THE STONES [SACRED STONES 3]  P
CALDECOTT, MOYRA  TALL STONES, THE [SACRED STONES 1]  P
CALDECOTT, MOYRA  TEMPLE OF THE SUN, THE [SACRED STONES 2]  H,P
CALDER, NIGEL  SPACESHIPS OF THE MIND  R/REF-H
CALDER, NIGEL  TIMESCALE  [C/RAN-H]
CALDER, RICHARD  CYTHERA  H
CALDER, RICHARD  DEAD GIRLS [DEAD 1]  P

The MITSFS Pinkdex  September 24, 2007  Page 66
AUTHOR  TITLE  CODE
CALDER, RICHARD  DEAD GIRLS DEAD BOYS DEAD THINGS [DEAD 1,2,3]  LP
CALDER, RICHARD  IMPAKTO  P
CALDER, RICHARD  TWIST, THE  P
CALDWELL, STEVEN  ALIENS IN SPACE  A
CALDWELL, STEVEN  STAR QUEST  A
CALDWELL, TAYLOR  DIALOGUES WITH THE DEVIL  C/P,P
CALDWELL, TAYLOR/STEARN, JESS  ROMANCE OF ATLANTIS, THE  P
CALENDAR  CHRONICLES OF NARNIA CALENDAR, THE (1982)  CAL
CALENDAR  DRAGONLANCE AND OTHER WORLDS (1996)  CAL
CALENDAR  EXPLORING THE UNIVERSE (1987)  CAL
CALENDAR  FRANK FRAZETTA CALENDAR, THE (1977)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1973)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1975)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1976)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1979)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1980)  CAL
CALENDAR  J. R. R. TOLKIEN CALENDAR, THE (1981)  CAL
CALENDAR  J.R.R. TOLKIEN CALENDAR 1998  CAL
CALENDAR  MICHAEL WHELAN’S YEAR OF WONDER (1989)  CAL
CALENDAR  MISS PIGGY COVER GIRL FANTASY CALENDAR (1981)  CAL
CALENDAR  STAR TREK (STARDATE) CALENDAR (1976)  CAL
CALENDAR  STAR TREK CALENDAR 1991  CAL
CALENDAR  STAR TREK DEEP SPACE NINE CALENDAR (1994) [STAR TREK DS9]  CAL
CALENDAR  STAR WARS CALENDAR (1978)  CAL
CALENDAR  STAR WARS POSTER ART CALENDAR (1980)  CAL
CALENDAR  STAR WARS RETURN OF THE JEDI (1984)  CAL
CALENDAR  STAR WARS THE EMPIRE STRIKES BACK CALENDAR (1981)  CAL

CALENDAR TERRY PRATCHETT 2000 DISCWORLD COLLECTOR’S EDITION CALENDAR,

THE

CAL

CALIF, R. C.  RUST  P
CALL, DEBORAH  ART OF STAR WARS: EPISODE V, THE [STAR WARS]  A
CALLAHAN, JAY  FOOTPRINTS OF THE DEAD [TWILIGHT 14]  P
CALLAHAN, JAY  NIGHT OF THE WOLF  P
CALLAHAN, WILLIAM  MACHINE THAT THOUGHT, THE  B
CALLANDER, DON  AEROMANCER [MANCER 4]  P
CALLANDER, DON  AQUAMANCER [MANCER 2]  P,SFWA-P
CALLANDER, DON  DRAGON COMPANION [DRAGON COMPANION 1]  C/P:2,P
CALLANDER, DON  DRAGON RESCUE [DRAGON COMPANION 2]  C/P,P
CALLANDER, DON  DRAGON TEMPEST  P
CALLANDER, DON  GEOMANCER [MANCER 3]  C/P,SFWA-P:2,P
CALLANDER, DON  MARBLEHEART  P
CALLANDER, DON  PYROMANCER [MANCER 1]  P
CALLENBACH, ERNEST  ECOTOPIA  C/P:2,LP
CALLENBACH, ERNEST  ECOTOPIA EMERGING  P
CALLIN, GRANT  LION ON THARTHEE, A  C/P:2,SFWA-P,P
CALLIN, GRANT  SATURNALIA  C/P:2,P
CALVINO, ITALO  CASTLE OF CROSSED DESTINIES, THE  LP
CALVINO, ITALO  COSMICOMICS  C/P:2,P
CALVINO, ITALO  T ZERO  C/P,P,H
CALVO, FRANCESCO  THERION CHRONICLES, THE  LP
CAMERON, BERL  BLACK INFINITY  SR-P
CAMERON, BERL  DESTINATION ALPHA  SR-P
CAMERON, BERL  LOST AEONS  SR-P
CAMERON, BERL  PHOTOMESIS  SR-P
CAMERON, BERL  SPHERO NOVA  SR-P
CAMERON, ELEANOR  BEYOND SILENCE  P
CAMERON, ELEANOR  MR. BASS’S PLANETOID  H
CAMERON, ELEANOR  MYSTERY FOR MR. BASS, A  H
CAMERON, ELEANOR  SPELL IS CAST, A  P
CAMERON, ELEANOR  STOWAWAY TO THE MUSHROOM PLANET  P
CAMERON, ELEANOR  WONDERFUL FLIGHT TO THE MUSHROOM PLANET, THE  P
CAMERON, IAN  ISLAND AT THE TOP OF THE WORLD  P
CAMERON, IAN  MOUNTAINS AT THE BOTTOM OF THE WORLD, THE  P
CAMERON, J. D.  BLOOD TIDE [@OMEGA SUB #4]  P
CAMERON, J. D.  CITY OF FEAR [@OMEGA SUB #3]  P
CAMERON, J. D.  COMMAND DECISION [@OMEGA SUB #2]  P
CAMERON, J. D.  DEATH DIVE [@OMEGA SUB #5]  P
CAMERON, J. D.  OMEGA SUB [@OMEGA SUB #1]  P
CAMERON, J. D.  RAVEN RISING [@OMEGA SUB #6]  P
CAMERON, KATE  AWAKENING DREAM, THE [WHISPERING HILLS 5]  P
CAMERON, KATE  CURSE OF WHISPERING HILLS, THE [WHISPERING HILLS 2]  P
CAMERON, KATE  ECHOES OF EVIL [WHISPERING HILLS 6]  P
CAMERON, KATE  LEGEND OF HOLDERLY HALL, THE [HOLDERLY HALL 1]  P
CAMERON, KATE  MUSIC FROM THE PAST [HOLDERLY HALL 5]  P

The MITSFS Pinkdex  September 24, 2007  Page 67
AUTHOR  TITLE  CODE
CAMERON, KATE  PORTRAITS OF THE PAST [HOLDERLY HALL 6]  P
CAMERON, KATE  SHADOWS OF THE PAST [HOLDERLY HALL 2]  P
CAMERON, KATE  SHADOWS ON THE MOON [WHISPERING HILLS 3]  P
CAMERON, KATE  VOICES IN THE FOG [HOLDERLY HALL 3]  P
CAMERON, LOU  CYBERNIA  C/P:2,P
CAMERON, SKYLA DAWN  RIVER  LP
CAMPBELL, ALAN  SCAR NIGHT [DEEPGATE 1]  H
CAMPBELL, BARBARA  HEARTWOOD [TRICKSTER’S GAME 1]  P
CAMPBELL, BILL  SUNSHINE PATRIOTS  LP

CAMPBELL, ELIZABETH MONTGOMERY/ SOLOMON, DAVID SEARCH FOR MORAG, THE

SFWA

CAMPBELL, H. J.  ANOTHER SPACE — ANOTHER TIME  SR-P
CAMPBELL, H. J.  BEYOND THE VISIBLE  SR-P,H
CAMPBELL, H. J.  BRAIN ULTIMATE  SR-P
CAMPBELL, H. J.  DIFERENTE ESPACIO, DIFERENTE TIEMPO  FP
CAMPBELL, H. J.  ONCE UPON A SPACE  SR-P
CAMPBELL, H. J.  PLANETA ROJO, EL  FP
CAMPBELL, JACK  DAUNTLESS [@LOST FLEET 1]  P
CAMPBELL, JACK  FEARLESS [@LOST FLEET 2]  P
CAMPBELL, JEFFREY  HOMING, THE  H

CAMPBELL, JIM/SPURRIER, SIMON/ RENNIE, GORDON/WALKER, KEV HERETIC SAINT [@WARHAMMER 40K DAEMONIFUGE]

SCX

CAMPBELL, JOHN W. JR.  AARN MUNRO IL GIOVIANO  FH
CAMPBELL, JOHN W. JR.  ANALOG 1  PA,HA
CAMPBELL, JOHN W. JR.  ANALOG 2  C/PA,PA,HA
CAMPBELL, JOHN W. JR.  ANALOG 3  HA
CAMPBELL, JOHN W. JR.  ANALOG 4  HA
CAMPBELL, JOHN W. JR.  ANALOG 5  HA
CAMPBELL, JOHN W. JR.  ANALOG 6  C/PA,PA
CAMPBELL, JOHN W. JR.  ANALOG 7  HA,PA
CAMPBELL, JOHN W. JR.  ANALOG 8  HA,SFWA
CAMPBELL, JOHN W. JR.  ANTHOLOGY: THREE NOVELS  SFWA
CAMPBELL, JOHN W. JR.  ASTOUNDING SCIENCE FICTION ANTHOLOGY, THE  PA,HA
CAMPBELL, JOHN W. JR.  ASTOUNDING TALES OF SPACE AND TIME  PA,C/PA
CAMPBELL, JOHN W. JR.  ATOMIC STORY, THE  R/REF-H
CAMPBELL, JOHN W. JR.  BEST OF JOHN W. CAMPBELL, THE  C/P:2,P,H
CAMPBELL, JOHN W. JR.  BLACK STAR PASSES, THE  SR-H,C/P:2,P
CAMPBELL, JOHN W. JR.  CLOAK OF AESIR  C/P:2,P
CAMPBELL, JOHN W. JR.  COLLECTED EDITORIALS FROM ANALOG  R/REF-H
CAMPBELL, JOHN W. JR.  COUNTERCOMMANDMENT  C/PA,PA
CAMPBELL, JOHN W. JR.  DEAD KNOWLEDGE  B
CAMPBELL, JOHN W. JR.  ELIMINATION  B
CAMPBELL, JOHN W. JR.  FROM UNKNOWN WORLDS  HA
CAMPBELL, JOHN W. JR.  INCREDIBLE PLANET, THE  SR-H
CAMPBELL, JOHN W. JR.  INVADERS FROM THE INFINITE  H:2,P
CAMPBELL, JOHN W. JR.  ISLANDS OF SPACE  SR-H,C/P:2,P
CAMPBELL, JOHN W. JR.  JOHN W. CAMPBELL LETTERS 1, THE  R/REF-P
CAMPBELL, JOHN W. JR.  MIGHTIEST MACHINE, THE  SR-H,C/P,P,H
CAMPBELL, JOHN W. JR.  MOON IS HELL!, THE (INCLUDES THE ELDER GODS)  SR-H,C/P,SR-LP,H
CAMPBELL, JOHN W. JR.  NEW DAWN, A  H
CAMPBELL, JOHN W. JR.  PERMANENT IMPLOSION, THE  C/PA,PA
CAMPBELL, JOHN W. JR.  PLANETEERS, THE  DG585
CAMPBELL, JOHN W. JR.  PROLOGUE TO ANALOG  HA
CAMPBELL, JOHN W. JR.  SPACE BEYOND, THE  C/P:2,P
CAMPBELL, JOHN W. JR.  THING, THE  SR-P
CAMPBELL, JOHN W. JR.  THING FROM ANOTHER WORLD, THE  B
CAMPBELL, JOHN W. JR.  ULTIMATE WEAPON, THE  C/P:2,DG585
CAMPBELL, JOHN W. JR.  WHO GOES THERE?  SR-P,C/P,H
CAMPBELL, JOHN W. JR.  WORLD BY THE TALE, A  C/PA,PA
CAMPBELL, JOSEPH  ARABIAN NIGHTS  PA
CAMPBELL, RAMSEY  ANCIENT IMAGES  SFWA-P
CAMPBELL, RAMSEY  COLD PRINT  SFWA-P
CAMPBELL, RAMSEY  DARK COMPANIONS  H,P
CAMPBELL, RAMSEY  DEMONS BY DAYLIGHT  SR-H,P
CAMPBELL, RAMSEY  DOLL WHO ATE HIS MOTHER, THE  H,P
CAMPBELL, RAMSEY  FACE THAT MUST DIE, THE  C/P,SR-P,H
CAMPBELL, RAMSEY  FINE FRIGHTS  SFWA-P
CAMPBELL, RAMSEY  HEIGHT OF THE SCREAM, THE  SR-H
CAMPBELL, RAMSEY  HUNGRY MOON, THE  SFWA-P
CAMPBELL, RAMSEY  LONG LOST, THE  P
CAMPBELL, RAMSEY  NAMELESS, THE  H,P
CAMPBELL, RAMSEY  NEEDING GHOSTS  H,LP
CAMPBELL, RAMSEY  NEW TALES OF THE CTHULHU MYTHOS  SR-HA,HA
CAMPBELL, RAMSEY  NEW TERRORS 1  SR-PA
CAMPBELL, RAMSEY  NEW TERRORS 2  SR-PA

The MITSFS Pinkdex  September 24, 2007  Page 68
AUTHOR  TITLE  CODE
CAMPBELL, RAMSEY  OBSESSION  P
CAMPBELL, RAMSEY  PARASITE, THE  H,SFWA-P,P
CAMPBELL, RAMSEY  SCARED STIFF  LP
CAMPBELL, RAMSEY  SCARED STIFF (PROOF)  SR-LP:2
CAMPBELL, RAMSEY  SUPERHORROR  SR-HA
CAMPBELL, RAMSEY  WAKING NIGHTMARES  H
CAMPBELL, TYREE  GEMSTONES  B
CANAVAN, TRUDI  HIGH LORD, THE [BLACK MAGICIAN 3]  C/P,P
CANAVAN, TRUDI  LAST OF THE WILDS [AGE OF THE FIVE 2]  P
CANAVAN, TRUDI  MAGICIANS’ GUILD, THE [BLACK MAGICIAN 1]  C/P,P
CANAVAN, TRUDI  NOVICE, THE [BLACK MAGICIAN 2]  C/P,P
CANAVAN, TRUDI  PRIESTESS OF THE WHITE [AGE OF THE FIVE 1]  P
CANAVAN, TRUDI  VOICE OF THE GODS [AGE OF THE FIVE 3]  P
CANNING, JOHN  50 GREAT GHOST STORIES  C/PA
CANNING, VICTOR  CRIMSON CHALICE, THE  P
CANNON, P. H.  SCREAM FOR JEEVES  H,SR-LP
CANTER, MARK  EMBER FROM THE SUN  H
CANTOR, HAL  GHOSTS AND THINGS  PA
CANTOR, JAY  KRAZY KAT  H:2
CANTRELL, LISA W.  MANSE, THE [MANSE 1]  SFWA-P
CANTRELL, LISA W.  RIDGE, THE  SFWA-P
CANTRELL, LISA W.  TORMENTS [MANSE 2]  SFWA-P
CANTRIL, HADLEY  INVASION FROM MARS, THE  R/REF-P
CAPEK, KAREL  ABSOLUTE AT LARGE, THE  LP
CAPEK, KAREL  KRAKATIT  H
CAPEK, KAREL  R. U. R.  P
CAPEK, KAREL  WAR WITH THE NEWTS  C/P:2,P,H
CAPOBIANCO, MICHAEL  BURSTER  P
CAPON, PAUL  DOWN TO EARTH  H
CAPON, PAUL  INTO THE TENTH MILLENNIUM  SR-P
CAPON, PAUL  LOST: A MOON  H
CAPON, PAUL  WORLD AT BAY, THE  H:2
CAPPELLI, OSCAR  RHAPSODY FOR A UNICORN  LP
CAPPS, CARROLL M.  SECRET OF THE SUNLESS WORLD  C/P,P
CARAKER, MARY  FACES OF CETI, THE  H
CARAKER, MARY  SEVEN WORLDS  C/P:2,P
CARAKER, MARY  SNOWS OF JASPRE, THE  H,P
CARBERRY, C. A.  CELESTIAL PURSUITS IN THE HUB OF THE UNIVERSE  LP
CARD, ORSON SCOTT  ABYSS, THE  C/P:2,P
CARD, ORSON SCOTT  ALVIN JOURNEYMAN [ALVIN MAKER 4]  C/P:2,P
CARD, ORSON SCOTT  ALVIN WANDERING [ALVIN MAKER 4,5]  H
CARD, ORSON SCOTT/FERRELL, KEITH  BLACK MIST  PA
CARD, ORSON SCOTT  CALL OF EARTH, THE [HOMECOMING 2]  C/P,P,H
CARD, ORSON SCOTT  CAPITOL  C/P:2,LP
CARD, ORSON SCOTT  CHANGED MAN, THE  C/P:2,P
CARD, ORSON SCOTT  CHILDREN OF THE MIND [ENDER 4]  H:3,P
CARD, ORSON SCOTT  CRUEL MIRACLES  P
CARD, ORSON SCOTT  CRYSTAL CITY, THE [ALVIN MAKER 6]  H
CARD, ORSON SCOTT  DRAGONS OF DARKNESS  C/PA,LPA
CARD, ORSON SCOTT  DRAGONS OF LIGHT  C/PA,PA,SR-LPA
CARD, ORSON SCOTT  EARTHBORN [HOMECOMING 5]  C/P,P,H:2
CARD, ORSON SCOTT  EARTHFALL [HOMECOMING 4]  H:2,P
CARD, ORSON SCOTT/OLEXA, KEITH  EMPIRE OF DREAMS AND MIRACLES  HA,LPA
CARD, ORSON SCOTT  ENCHANTMENT  H
CARD, ORSON SCOTT  ENDER’S GAME (ORIGINAL) [ENDER 1]  P
CARD, ORSON SCOTT  ENDER’S GAME (REVISED 1991 EDITION) [ENDER 1]  C/P:3,SFWA-P,P
CARD, ORSON SCOTT  ENDER’S SHADOW [ENDER 1.1]  H:2,P

CARD, ORSON SCOTT ENDER’S WAR (CONTAINS ORIGINAL NOVEL VERSION OF ENDER’S GAME)

[ENDER 1,2]

H:2

CARD, ORSON SCOTT  EYE FOR EYE  SFWA-TD27,TD27
CARD, ORSON SCOTT  FIRST MEETINGS IN THE ENDERVERSE [ENDER]  H
CARD, ORSON SCOTT  FLUX  P
CARD, ORSON SCOTT  FOLK OF THE FRINGE, THE  C/P:2,P,H
CARD, ORSON SCOTT  FUTURE ON FIRE  C/PA,PA,SFWA-P
CARD, ORSON SCOTT  FUTURE ON ICE  LPA
CARD, ORSON SCOTT  HART’S HOPE  C/P:2,P
CARD, ORSON SCOTT  HATRACK RIVER [ALVIN MAKER 1,2,3]  H
CARD, ORSON SCOTT  HEARTFIRE [ALVIN MAKER 5]  H,P

CARD, ORSON SCOTT/OLEXA, KEITH/ O’TOOLE, CHRISTIAN HITTING THE SKIDS IN PIXELTOWN

LPA

CARD, ORSON SCOTT  HOMEBODY  P
CARD, ORSON SCOTT  HOMECOMING: EARTH [HOMECOMING 4,5]  H:2
CARD, ORSON SCOTT  HOMECOMING: HARMONY [HOMECOMING 1,2,3]  H:2
CARD, ORSON SCOTT  HOT SLEEP: THE WORTHING CHRONICLE [WORTHING]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 69
AUTHOR  TITLE  CODE
CARD, ORSON SCOTT  HOW TO WRITE SCIENCE FICTION AND FANTASY  R/REF-H
CARD, ORSON SCOTT  LOST BOYS  H:2,P
CARD, ORSON SCOTT/KIDD, KATHRYN H.  LOVELOCK [MAYFLOWER 1]  H,P
CARD, ORSON SCOTT  MAGIC STREET  H
CARD, ORSON SCOTT  MAPS IN A MIRROR  H:2,SFWA
CARD, ORSON SCOTT  MASTERPIECES  HA
CARD, ORSON SCOTT  MEMORY OF EARTH, THE [HOMECOMING 1]  C/P,P,H
CARD, ORSON SCOTT  MONKEY SONATAS  P
CARD, ORSON SCOTT  PASTWATCH  C/P,P,H
CARD, ORSON SCOTT  PLANET CALLED TREASON, A  H:2,P
CARD, ORSON SCOTT  PRENTICE ALVIN [ALVIN MAKER 3]  C/P,SFWA-P,P,H
CARD, ORSON SCOTT  RED PROPHET [ALVIN MAKER 2]  H:2,SFWA-P,P
CARD, ORSON SCOTT  SAINTS  SFWA-P
CARD, ORSON SCOTT  SARAH [WOMEN OF GENESIS 1]  H
CARD, ORSON SCOTT  SEVENTH SON [ALVIN MAKER 1]  SFWA-P,H:2,P
CARD, ORSON SCOTT  SHADOW OF THE GIANT [ENDER 1.4]  H:2,P
CARD, ORSON SCOTT  SHADOW OF THE HEGEMON [ENDER 1.2]  C/P,P,H
CARD, ORSON SCOTT  SHADOW PUPPETS [ENDER 1.3]  H:2,P
CARD, ORSON SCOTT  SHIPS OF EARTH, THE [HOMECOMING 3]  H:2,P
CARD, ORSON SCOTT  SONGMASTER  C/P,SFWA,P,H
CARD, ORSON SCOTT  SPEAKER FOR THE DEAD [ENDER 2]  C/P:2,SFWA,SFWA-P,P
CARD, ORSON SCOTT  TREASON  C/P,P,H
CARD, ORSON SCOTT  TREASURE BOX  C/P,P,H
CARD, ORSON SCOTT  UNACCOMPANIED SONATA  B,H:2,P,SFWA
CARD, ORSON SCOTT  WOMAN OF DESTINY, A  P
CARD, ORSON SCOTT  WORTHING CHRONICLE, THE [WORTHING]  C/P:2,P
CARD, ORSON SCOTT  WORTHING SAGA, THE [WORTHING]  C/P:2,SFWA-P,P
CARD, ORSON SCOTT  WYRMS  SFWA-P,H:2,P
CARD, ORSON SCOTT  XENOCIDE [ENDER 3]  H:3,P
CARD, ORSON SCOTT  XENOCIDE (PROOF) [ENDER 3]  SR-LP
CARDARELLE, ANDRIA  TO SLEEP WITH EVIL [RAVENLOFT]  P
CARELMAN, JACQUES  CATALOG OF FANTASTIC THINGS  [C/RAN-H]
CAREY, DIANE  ANCIENT BLOOD [STAR TREK DAY OF HONOR 1/STAR TREK TNG]  C/P:2,P
CAREY, DIANE  BATTLESTATIONS! [STAR TREK]  C/P:2,P
CAREY, DIANE  BEST DESTINY [STAR TREK]  C/P,P,H
CAREY, DIANE  BROKEN BOW [@STAR TREK ENTERPRISE]  P
CAREY, DIANE  CADET KIRK [@STAR TREK STARFLEET ACADEMY #3]  LP
CAREY, DIANE  CALL TO ARMS... [@STAR TREK DOMINION WAR #2/STAR TREK DS9]  C/P,P
CAREY, DIANE  CHAINMAIL [STAR TREK GATEWAYS #2/STAR TREK CHALLENGER]  P
CAREY, DIANE  CHALLENGER [@STAR TREK NEW EARTH #6]  C/P,P
CAREY, DIANE  DESCENT [STAR TREK TNG]  C/P,P
CAREY, DIANE  DREADNOUGHT! [STAR TREK]  C/P:3,P
CAREY, DIANE  ENDGAME [STAR TREK VOYAGER]  LP
CAREY, DIANE  EQUINOX [STAR TREK VOYAGER]  P
CAREY, DIANE  FINAL FRONTIER [STAR TREK]  C/P:2,P
CAREY, DIANE  FIRE SHIP [STAR TREK CAPTAIN’S TABLE 4/STAR TREK VOYAGER]  C/P,P
CAREY, DIANE/KIRKLAND, DR. JAMES I.  FIRST FRONTIER [STAR TREK]  C/P:2,P
CAREY, DIANE  FIRST STRIKE [STAR TREK INVASION! 1]  C/P,P
CAREY, DIANE  FLASHBACK [STAR TREK VOYAGER]  C/P:2,P
CAREY, DIANE  GHOST SHIP [STAR TREK TNG]  C/P:2,P
CAREY, DIANE  GREAT STARSHIP RACE, THE [STAR TREK]  C/P:3,P
CAREY, DIANE  RED SECTOR [@STAR TREK TNG DOUBLE HELIX 3]  P

CAREY, DIANE ...SACRIFICE OF ANGELS [@STAR TREK DOMINION WAR #4/

STAR TREK DS9]

P

CAREY, DIANE  SEARCH, THE [STAR TREK DS9]  C/P,P
CAREY, DIANE  SHIP OF THE LINE [STAR TREK TNG]  P
CAREY, DIANE  STARFLEET ACADEMY [STAR TREK]  C/P,P
CAREY, DIANE  STATION RAGE [STAR TREK DS9]  C/P,P
CAREY, DIANE  TRIALS AND TRIBBLE-ATIONS [STAR TREK DS9]  P
CAREY, DIANE  WAGON TRAIN TO THE STARS [@STAR TREK NEW EARTH #1]  P
CAREY, DIANE  WAY OF THE WARRIOR, THE [STAR TREK DS9]  P
CAREY, DIANE  WHAT YOU LEAVE BEHIND [@STAR TREK DS9]  P
CAREY, JACQUELINE  BANEWREAKER [@SUNDERING #1]  H
CAREY, JACQUELINE  GODSLAYER [@SUNDERING #2]  H
CAREY, JACQUELINE  KUSHIEL’S AVATAR [KUSHIEL 3]  H,P
CAREY, JACQUELINE  KUSHIEL’S CHOSEN [KUSHIEL 2]  H,P
CAREY, JACQUELINE  KUSHIEL’S DART [KUSHIEL 1]  H,P
CAREY, JACQUELINE  KUSHIEL’S JUSTICE [KUSHIEL 5]  H
CAREY, JACQUELINE  KUSHIEL’S SCION [KUSHIEL 4]  H
CAREY, PETER  ILLYWHACKER  H
CARKEET, DAVID  I BEEN THERE BEFORE  H
CARL, LILLIAN STEWART  SABAZEL  P
CARL, LILLIAN STEWART  SHADOW DANCERS  C/P,SFWA-P,P
CARL, LILLIAN STEWART  WINGS OF POWER  C/P:2,P

AUTHOR  TITLE  CODE
CARL, LILLIAN STEWART  WINTER KING, THE  C/P:2,P
CARLILE, CLANCY  SPORE 7  P
CARLIN, MIKE  BEGINNINGS [STAR TREK TNG]  CX
CARLISLE, ANNE  LIQUID SKY  LP
CARLOCK, MICHAELA  PLANET DREAMS  LP
CARLSEN, CHRIS  BULL CHIEF, THE [@BERSERKER SWIFTAXE 2]  P
CARLSEN, CHRIS  HORNED WARRIOR, THE [@BERSERKER SWIFTAXE 3]  P
CARLSEN, CHRIS  SHADOW OF THE WOLF [@BERSERKER SWIFTAXE 1]  P
CARLSON, DALE  HUMAN APES, THE  H
CARLSON, DALE  PLANT PEOPLE, THE  P
CARLSON, DALE/CARLSON, DANNY  SHINING POOL, THE  H
CARLSON, WILLIAM K.  ELYSIUM  H,SFWA
CARLSON, WILLIAM K.  SUNRISE WEST  H
CARLTON, CLARK  FACE/OFF  P
CARLYON, RICHARD  DARK LORD OF PENGERSICK, THE  P
CARMER, CARL  HURRICANE’S CHILDREN, THE  H
CARMODY, ISOBELLE  ASHLING [OBERNEWTYN 3]  H
CARMODY, ISOBELLE  FARSEEKERS, THE [OBERNEWTYN 2]  H,P
CARMODY, ISOBELLE  OBERNEWTYN [OBERNEWTYN 1]  P
CARMODY, ISOBELLE  OBERNEWTYN & THE FARSEEKERS [OBERNEWTYN 1,2]  H
CARNELL, JOHN  LAMBDA I AND OTHER STORIES  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 1  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 2  SR-PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 3  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 4  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 5  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 6  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 7  SR-HA,PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 8  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 9  PA,C/PA
CARNELL, JOHN  NEW WRITINGS IN SF 10  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 12  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 13  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 14  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 15  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 16  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 17  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 18  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 19  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 20  SR-PA
CARNELL, JOHN  NEW WRITINGS IN SF 21  SR-PA
CARNELL, JOHN  ROADKILL OF MIDDLE EARTH, THE  LP
CARNELL, JOHN  WEIRD SHADOWS FROM BEYOND  PA,C/PA
CARO, DENNIS R.  DEVINE WAR  H,SFWA
CARO, DENNIS R.  MAN IN THE DARKSUIT, THE  C/P,P
CARPENTER, CHRISTOPHER  TWILIGHT REALM, THE  C/P,P
CARPENTER, ELMER J.  MOONSPIN  P
CARPENTER, HUMPHREY  INKLINGS, THE  R/REF-H
CARPENTER, HUMPHREY  LETTERS OF J.R.R. TOLKIEN, THE  R/REF-H,C/REF-H
CARPENTER, LEONARD  CONAN OF THE RED BROTHERHOOD [CONAN]  P
CARPENTER, LEONARD  CONAN, SCOURGE OF THE BLOODY COAST [CONAN]  P
CARPENTER, LEONARD  CONAN THE GLADIATOR [CONAN]  P
CARPENTER, LEONARD  CONAN THE GREAT [CONAN]  P,SFWA-P
CARPENTER, LEONARD  CONAN THE HERO [CONAN]  C/P,P
CARPENTER, LEONARD  CONAN THE OUTCAST [CONAN]  P
CARPENTER, LEONARD  CONAN THE RAIDER [CONAN]  P,SFWA,SFWA-P,LP
CARPENTER, LEONARD  CONAN THE RENEGADE [CONAN]  P,SFWA-P
CARPENTER, LEONARD  CONAN THE SAVAGE [CONAN]  P
CARPENTER, LEONARD  CONAN THE WARLORD [CONAN]  P,SFWA-P
CARPENTER, LEONARD  LORD OF THE BLACK RIVER [CONAN]  P
CARPOZI, GEORGE JR.  SUNSTRIKE  C/P,P
CARR, CHARLES  COLONISTS OF SPACE  SR-P
CARR, CHARLES  SALAMANDER WAR  SR-P
CARR, JAYGE  LEVIATHAN’S DEEP  H:2,SFWA,P
CARR, JAYGE  NAVIGATOR’S SINDROME [RABELAIS 1]  H,SFWA
CARR, JAYGE  RABELAISIAN REPRISE [RABELAIS 3]  H,SFWA
CARR, JAYGE  TREASURE IN THE HEART OF THE MAZE, THE [RABELAIS 2]  H,SFWA
CARR, JOHN DICKSON  BURNING COURT, THE  P
CARR, JOHN DICKSON  DEVIL IN VELVET, THE  P
CARR, JOHN DICKSON  FIRE, BURN!  C/P,P
CARR, JOHN F.  OPHIDIAN CONSPIRACY, THE  P
CARR, JOHN F.  PAIN GAIN  P
CARR, NICK  AMERICA’S SECRET SERVICE ACE  R/REF-P
CARR, ROBERT SPENCER  BEYOND INFINITY  C/P:2,P

AUTHOR  TITLE  CODE
CARR, ROBERT SPENCER  INVADERS, THE  B
CARR, TERRY  BEST FROM UNIVERSE, THE  SFWA
CARR, TERRY  BEST SCIENCE FICTION NOVELLAS OF THE YEAR #1, THE  C/PA,PA
CARR, TERRY  BEST SCIENCE FICTION NOVELLAS OF THE YEAR #2, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #2, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #3, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #4, THE  C/PA:2,PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #5, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #6, THE  PA,C/PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #7, THE  PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #8, THE  HA,PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #9, THE  PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #10, THE  C/PA,PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #11, THE  C/PA,PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #12, THE  C/PA,PA
CARR, TERRY  BEST SCIENCE FICTION OF THE YEAR #13, THE  C/PA,PA
CARR, TERRY  BETWEEN TWO WORLDS  SR-DNESFA1
CARR, TERRY  BEYOND REALITY  HA
CARR, TERRY  CIRQUE  C/P,SFWA,P,H
CARR, TERRY  CLASSIC SCIENCE FICTION: THE FIRST GOLDEN AGE  SFWA
CARR, TERRY  CREATURES FROM BEYOND  HA
CARR, TERRY  DREAM’S EDGE  HA,LPA
CARR, TERRY  EXALTATION OF STARS, AN  PA,HA
CARR, TERRY  FANTASY ANNUAL III  PA
CARR, TERRY  FANTASY ANNUAL IV  HA,PA
CARR, TERRY  FANTASY ANNUAL V  PA
CARR, TERRY  FELLOWSHIP OF THE STARS  SR-HA,SFWA,HA
CARR, TERRY  IDES OF TOMORROW, THE  SR-HA,HA
CARR, TERRY  INFINITE ARENA, THE  HA
CARR, TERRY  INTO THE UNKNOWN  HA
CARR, TERRY  LIGHT AT THE END OF THE UNIVERSE, THE  C/P,P
CARR, TERRY  NEW WORLDS OF FANTASY  C/PA,PA
CARR, TERRY  NEW WORLDS OF FANTASY #2  PA,C/PA
CARR, TERRY  NEW WORLDS OF FANTASY #3  PA
CARR, TERRY  ON OUR WAY TO THE FUTURE  PA,C/PA
CARR, TERRY  OTHERS, THE  PA,C/PA
CARR, TERRY  PLANETS OF WONDER  HA
CARR, TERRY  SCIENCE FICTION FOR PEOPLE WHO HATE SCIENCE FICTION  HA,PA
CARR, TERRY  SCIENCE FICTION HALL OF FAME 4, THE  HA,PA
CARR, TERRY  TERRY CARR’S BEST SCIENCE FICTION OF THE YEAR (#14)  C/PA,PA,SFWA-P
CARR, TERRY  TERRY CARR’S BEST SCIENCE FICTION OF THE YEAR #15  PA,SFWA-P
CARR, TERRY  TERRY CARR’S BEST SCIENCE FICTION AND FANTASY OF THE YEAR #16  HA
CARR, TERRY  THIS SIDE OF INFINITY  PA,C/PA
CARR, TERRY  TO FOLLOW A STAR: NINE SCIENCE FICTION STORIES ABOUT CHRISTMAS HA
CARR, TERRY/GREENBERG, MARTIN H.  TREASURY OF MODERN FANTASY, A  LPA
CARR, TERRY  UNIVERSE 1  PA,C/PA
CARR, TERRY  UNIVERSE 2  PA,C/PA
CARR, TERRY  UNIVERSE 3  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 4  C/PA,PA
CARR, TERRY  UNIVERSE 5  PA,C/PA:2
CARR, TERRY  UNIVERSE 6  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 7  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 8  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 9  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 10  SFWA,PA,HA
CARR, TERRY  UNIVERSE 11  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 12  C/PA,PA,SFWA
CARR, TERRY  UNIVERSE 13  C/PA,PA,SFWA-P,SFWA
CARR, TERRY  UNIVERSE 14  C/PA,SFWA-P,PA,SFWA
CARR, TERRY  UNIVERSE 15  HA,PA,SFWA-P,SFWA
CARR, TERRY  UNIVERSE 16  HA,SFWA
CARR, TERRY  UNIVERSE 17  HA
CARR, TERRY  WARLORD OF KOR  DF177
CARR, TERRY  WORLDS NEAR AND FAR  HA
CARR, TERRY  YEAR’S FINEST FANTASY, THE  PA,C/PA
CARR, TERRY  YEAR’S FINEST FANTASY 2, THE  C/PA,PA
CARRADINE, F. J.  RUNES & RIDDLES [EMPIRE OF ELWOLF]  LP
CARRAU, BOB  MONSTERS AND ALIENS FROM GEORGE LUCAS  H
CARRELL, CHRISTOPHER  BEYOND THIS HORIZON  VLPA
CARRERE, EMMANUEL  I AM ALIVE AND YOU ARE DEAD  R/REF-H
CARRIGAN, RICHARD/CARRIGAN, NANCY  SIREN STARS, THE  C/P,P
CARRINGTON, GRANT  TIME’S FOOL  H,SFWA
CARROLL, GLADYS HASTY  MAN ON THE MOUNTAIN  P
CARROLL, JERRY JAY  DOG EAT DOG [INGERSOL 2]  H,LP

The MITSFS Pinkdex  September 24, 2007  Page 72
AUTHOR  TITLE  CODE
CARROLL, JERRY JAY  INHUMAN BEINGS  C/P,LP
CARROLL, JERRY JAY  TOP DOG [INGERSOL 1]  C/P,P
CARROLL, JONATHAN  BLACK COCKTAIL  SFWA,LP
CARROLL, JONATHAN  BONES OF THE MOON  H:2,P
CARROLL, JONATHAN  CHILD ACROSS THE SKY, A  P
CARROLL, JONATHAN  FROM THE TEETH OF ANGELS  P
CARROLL, JONATHAN  KISSING THE BEEHIVE  LP
CARROLL, JONATHAN  LAND OF LAUGHS, THE  C/P,P
CARROLL, JONATHAN  MARRIAGE OF STICKS, THE  LP
CARROLL, JONATHAN  OUTSIDE THE DOG MUSEUM  C/P,P
CARROLL, JONATHAN  SLEEPING IN FLAME  P,LP
CARROLL, JONATHAN  VOICE OF OUR SHADOW  P
CARROLL, JONATHAN  WHITE APPLES  LP
CARROLL, JONATHAN  WOODEN SEA, THE  H

CARROLL, LEWIS/LE GALLIENNE, EVA/ FRIEBUS, FLORIDA ALICE IN WONDERLAND (STAGE ADAPTATION)

SR-P

CARROLL, LEWIS  ANNOTATED ALICE, THE  SR-VLH,H
CARROLL, LEWIS  AVENTURES D’ALICE AU PAYS DES MERVEILLES  FP
CARROLL, LEWIS  THROUGH THE LOOKING GLASS  H
CARROLL, LEWIS  UNKNOWN LEWIS CARROLL, THE  LP
CARSAC, FRANCIS  VERMINE DU LION, LA  FP
CARSON, JOHN F.  BOYS WHO VANISHED, THE  H
CARSON, RACHEL L.  UNDER THE SEA WIND  [C/RAN-P]
CARSON, ROBIN  PAWN OF TIME  H
CARTER, ANGELA  FIREWORKS  C/P,LP
CARTER, ANGELA  HEROES AND VILLAINS  C/P,P
CARTER, ANGELA  INFERNAL DESIRE MACHINES OF DOCTOR HOFFMAN, THE  LP
CARTER, ANGELA  MAGIC TOYSHOP, THE  P
CARTER, ANGELA  NIGHTS AT THE CIRCUS  H
CARTER, ANGELA  PASSION OF NEW EVE, THE  H,LP
CARTER, ANGELA  WAR OF DREAMS, THE  H,P
CARTER, CARMEN  CHILDREN OF HAMLIN, THE [STAR TREK TNG]  C/P:3,P
CARTER, CARMEN  DEVIL’S HEART, THE [STAR TREK TNG]  C/P,P,H

CARTER, CARMEN/ DAVID, PETER/FRIEDMAN, MICHAEL JAN/ GREENBERGER, ROBERT DOOMSDAY WORLD [STAR TREK TNG]

C/P:2,P

CARTER, CARMEN  DREAMS OF THE RAVEN [STAR TREK]  C/P:2,P
CARTER, CHRIS/HAND, ELIZABETH  X FILES FIGHT THE FUTURE, THE [X-FILES]  H,P
CARTER, DEE  BLUE CORDON  SR-P
CARTER, DEE  CHLOROPLASM  SR-P
CARTER, DEE  PURPLE ISLANDS  SR-P
CARTER, LIN  AS THE GREEN STAR RISES [GREEN STAR 4]  C/P,P
CARTER, LIN  BARBARIAN OF WORLD’S END, THE [GONDWANE 4]  P
CARTER, LIN  BEYOND THE GATES OF DREAM  C/P:2,P
CARTER, LIN  BLACK LEGION OF CALLISTO [JANDAR 2]  C/P:2,P
CARTER, LIN  BLACK STAR, THE  C/P:2,P
CARTER, LIN  BY THE LIGHT OF THE GREEN STAR [GREEN STAR 3]  C/P:2,P
CARTER, LIN  CALLIPYGIA [TERRA MAGICA 4]  P
CARTER, LIN  CITY OUTSIDE THE WORLD, THE [MAN WHO LOVED MARS 3]  C/P,P
CARTER, LIN  DARYA OF THE BRONZE AGE [ERIC CARSTAIRS 4]  P
CARTER, LIN  DOWN TO A SUNLESS SEA [MAN WHO LOVED MARS 4]  C/P,P
CARTER, LIN  DRAGONROUGE [TERRA MAGICA 2]  C/P,P
CARTER, LIN  DRAGONS, ELVES, AND HEROES  PA
CARTER, LIN  EARTH-SHAKER, THE [ZARKON 4]  H,SFWA
CARTER, LIN  ENCHANTRESS OF WORLD’S END, THE [GONDWANE 2]  C/P,P
CARTER, LIN  ERIC OF ZANTHODON [ERIC CARSTAIRS 5]  P
CARTER, LIN  FLAME OF IRIDAR, THE  SR-DB0759
CARTER, LIN  FLASHING SWORDS! #1  PA,C/PA
CARTER, LIN  FLASHING SWORDS! #2  C/PA,PA
CARTER, LIN  FLASHING SWORDS! #3: WARRIORS AND WIZARDS  PA,C/PA
CARTER, LIN  FLASHING SWORDS! #4: BARBARIANS AND BLACK MAGICIANS  PA,C/PA
CARTER, LIN  FLASHING SWORDS! #5: DEMONS AND DAGGERS  PA,HA
CARTER, LIN  FOUND WANTING  C/P:2,P
CARTER, LIN  GIANT OF WORLD’S END [GONDWANE]  P
CARTER, LIN  GOLDEN CITIES, FAR  SR-PA,C/PA
CARTER, LIN  HORROR WEARS BLUE  SFWA
CARTER, LIN  HUROK OF THE STONE AGE [ERIC CARSTAIRS 3]  P
CARTER, LIN  IMAGINARY WORLDS  R/REF-P,C/REF-P
CARTER, LIN  IMMORTAL OF WORLD’S END, THE [GONDWANE 3]  C/P,P
CARTER, LIN  IN THE GREEN STAR’S GLOW [GREEN STAR 5]  P
CARTER, LIN  INVISIBLE DEATH [ZARKON 2]  H:2,SFWA,P
CARTER, LIN  JANDAR OF CALLISTO [JANDAR 1]  C/P:2,P
CARTER, LIN  JOURNEY TO THE UNDERGROUND WORLD [ERIC CARSTAIRS 1]  P
CARTER, LIN  KELLORY THE WARLOCK  SFWA

AUTHOR  TITLE  CODE
CARTER, LIN  KESRICK [TERRA MAGICA 1]  C/P:2,P
CARTER, LIN  KINGDOMS OF SORCERY  HA,SFWA
CARTER, LIN  LANKAR OF CALLISTO [JANDAR 6]  C/P,P
CARTER, LIN  LOST WORLD OF TIME  P
CARTER, LIN  LOST WORLDS  P
CARTER, LIN  LOVECRAFT: A LOOK BEHIND THE CTHULHU MYTHOS  R/REF-P,C/REF-P
CARTER, LIN  MAD EMPRESS OF CALLISTO [JANDAR 4]  C/P,P
CARTER, LIN  MAGIC OF ATLANTIS, THE  C/PA,PA
CARTER, LIN  MAN WHO LOVED MARS, THE [MAN WHO LOVED MARS 1]  C/P:2,P
CARTER, LIN  MAN WITHOUT A PLANET, THE [GREAT IMPERIUM]  DG606
CARTER, LIN  MANDRICARDO [TERRA MAGICA 3]  P
CARTER, LIN  MIND WIZARDS OF CALLISTO [JANDAR 5]  C/P,P
CARTER, LIN  NEMESIS OF EVIL, THE [ZARKON 1]  P,SFWA
CARTER, LIN  NEW WORLDS FOR OLD  SR-PA
CARTER, LIN  OUTWORLDER [GREAT IMPERIUM]  C/P:2,P
CARTER, LIN  PIRATE OF WORLD’S END, THE [GONDWANE 5]  P
CARTER, LIN  PURLOINED PLANET, THE  SR-DB1010
CARTER, LIN  QUEST OF KADJI, THE  C/P:2,P
CARTER, LIN  REALMS OF WIZARDRY  HA,SFWA
CARTER, LIN  RENEGADE OF CALLISTO [JANDAR 8]  P
CARTER, LIN  SKY PIRATES OF CALLISTO [JANDAR 3]  C/P,P
CARTER, LIN  SPAWN OF CTHULHU, THE  PA,C/PA
CARTER, LIN  STAR MAGICIANS, THE  DG588
CARTER, LIN  STAR ROGUE [GREAT IMPERIUM]  C/P:2,P
CARTER, LIN  TARA OF THE TWILIGHT  SR-P
CARTER, LIN  THIEF OF THOTH, THE  SR-DB0244,SR-DB0809
CARTER, LIN  THONGOR AGAINST THE GODS  C/P,P
CARTER, LIN  THONGOR AND THE DRAGON CITY  C/P:2,P
CARTER, LIN  THONGOR AND THE WIZARD OF LEMURIA  C/P,P
CARTER, LIN  THONGOR AT THE END OF TIME  P
CARTER, LIN  THONGOR FIGHTS THE PIRATES OF TARAKUS  C/P,P
CARTER, LIN  THONGOR IN THE CITY OF MAGICIANS  P
CARTER, LIN  THONGOR OF LEMURIA  SR-P,C/P
CARTER, LIN  TIME WAR  C/P,P
CARTER, LIN  TOLKIEN: A LOOK BEHIND THE LORD OF THE RINGS  R/REF-P,C/REF-P
CARTER, LIN  TOWER AT THE EDGE OF TIME  C/P,P
CARTER, LIN  TOWER OF THE MEDUSA  D42900
CARTER, LIN  UNDER THE GREEN STAR [GREEN STAR 1]  C/P:2,P
CARTER, LIN  VALLEY WHERE TIME STOOD STILL, THE [MAN WHO LOVED MARS 2]  H,P
CARTER, LIN  VOLCANO OGRE, THE [ZARKON 3]  C/P,SFWA,H
CARTER, LIN  WARRIOR OF WORLD’S END, THE [GONDWANE 1]  C/P,P
CARTER, LIN  WEIRD TALES #1  SR-PA
CARTER, LIN  WEIRD TALES #2  SR-PA
CARTER, LIN  WEIRD TALES #3  SR-PA
CARTER, LIN  WEIRD TALES #4  SR-PA
CARTER, LIN  WHEN THE GREEN STAR CALLS [GREEN STAR 2]  C/P:2,P
CARTER, LIN  WIZARD OF LEMURIA, THE  SR-P,C/P:2
CARTER, LIN  WIZARD OF ZAO, THE  C/P:2,P
CARTER, LIN  YEAR’S BEST FANTASY STORIES, THE  PA
CARTER, LIN  YEAR’S BEST FANTASY STORIES 2, THE  C/PA,PA
CARTER, LIN  YEAR’S BEST FANTASY STORIES 3, THE  C/PA,PA
CARTER, LIN  YEAR’S BEST FANTASY STORIES 4, THE  PA
CARTER, LIN  YEAR’S BEST FANTASY STORIES 5, THE  PA
CARTER, LIN  YEAR’S BEST FANTASY STORIES 6, THE  PA
CARTER, LIN  YLANA OF CALLISTO [JANDAR 7]  C/P,P
CARTER, LIN  YOUNG MAGICIANS, THE  PA
CARTER, LIN  ZANTHODON [ERIC CARSTAIRS 2]  P
CARTER, M. L.  CURSE OF THE UNDEAD, THE  PA
CARTER, M. L.  DEMON LOVERS AND STRANGE SEDUCTIONS  PA,C/PA
CARTER, PAUL A.  CREATION OF TOMORROW, THE  R/REF-H
CARTER, RAPHAEL  FORTUNATE FALL, THE  LP
CARTIER, EDD  KNOWN AND THE UNKNOWN, THE  A
CARTLAND, BARBARA  LOST LOVE  P
CARTMEL, ANDREW  ATOM BOMB BLUES [DOCTOR WHO PAST DOCTORS]  P
CARTMEL, ANDREW  CAT’S CRADLE: WARHEAD [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CARTMEL, ANDREW  FOREIGN DEVILS [DOCTOR WHO]  SR-H,H
CARTMEL, ANDREW  WARCHILD [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CARTMEL, ANDREW  WARLOCK [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CARTMILL, CLEVE  SPACE SCAVENGERS, THE  P
CARVER, JEFFREY A.  CLYPSIS [@ALIEN SPEEDWAY 1]  C/P:2,P
CARVER, JEFFREY A.  DOWN THE STREAM OF STARS [CHANGELING STAR 2]  C/P:2,P
CARVER, JEFFREY A.  DRAGON RIGGER [STAR DRAGONS 2]  C/P,P,H
CARVER, JEFFREY A.  DRAGONS IN THE STARS [STAR DRAGONS 1]  C/P,P
CARVER, JEFFREY A.  ETERNITY’S END  H:2

The MITSFS Pinkdex  September 24, 2007  Page 74
AUTHOR  TITLE  CODE
CARVER, JEFFREY A.  FROM A CHANGELING STAR [CHANGELING STAR 1]  H:2,P
CARVER, JEFFREY A.  INFINITE SEA, THE [CHAOS CHRONICLES 3]  H,P
CARVER, JEFFREY A.  INFINITY LINK, THE  C/P,SFWA-P:2,P,H:2
CARVER, JEFFREY A.  NEPTUNE CROSSING [CHAOS CHRONICLES 1]  C/P:2,P,H
CARVER, JEFFREY A.  PANGLOR  P
CARVER, JEFFREY A.  RAPTURE EFFECT, THE  H:2,SFWA-P,P
CARVER, JEFFREY A.  SEAS OF ERNATHE  SR-P
CARVER, JEFFREY A.  STAR RIGGER’S WAY  C/P,SR-P,H:2
CARVER, JEFFREY A.  STRANGE ATTRACTORS [CHAOS CHRONICLES 2]  C/P,P,H
CARWYN, GILES/FAHNESTOCK, TODD  HEIR OF AUTUMN [OHNDARIEN 1]  H
CARWYN, GILES/FAHNESTOCK, TODD  MISTRESS OF WINTER [OHNDARIEN 2]  H
CARY, DEVIN  WINTER QUEEN, THE  P
CARY, JOYCE  AFRICAN WITCH, THE  H
CASE, DAVID  CELL, THE  H
CASE, DAVID  THIRD GRAVE, THE  H
CASE, GEORGE  SILENCE DESCENDS  LP

CASE, RICHARD D./GESS, NICHOLAS M./ ROGERSON, EDWARD G./ HIRSCHLER, PHILIP/ SALTONSTALL, WILLIAM L. JR.

DAIMON (LATIN)

FH

CASELBERG, JAY  WYRMHOLE  P
CASEWIT, CURTIS W.  PEACEMAKERS, THE  C/P,P
CASH, STEVE  MEQ, THE  LP
CASIL, AMY STERLING  WITHOUT ABSOLUTION  H,LP
CASPIAN, JONATHA ARIADNE  NIGHTMARE DREAM, THE [@TORG POSSIBILITY WARS #3]  C/P,P
CASSADA, JACKIE  SHADOWS ON THE HILL [@IMMORTAL EYES #2]  P
CASSADA, JACKIE  TOYBOX, THE [@IMMORTAL EYES #1]  P
CASSEDY, SYLVIA  BEHIND THE ATTIC WALL  LP
CASSUTT, MICHAEL  DRAGON SEASON  C/P,P
CASSUTT, MICHAEL  STAR COUNTRY, THE  H,SFWA
CAST, P. C.  ELPHAME’S CHOICE  LP
CASTANEDA, CARLOS  JOURNEY TO IXTLAN  [C/RAN-P]
CASTANEDA, CARLOS  SEPARATE REALITY, A  [C/RAN-P]
CASTANEDA, CARLOS  TALES OF POWER  [C/RAN-P]
CASTANEDA, CARLOS  TEACHINGS OF DON JUAN: A YAQUI WAY OF KNOWLEDGE, THE  [C/RAN-P:2]
CASTEVANO, ROMAN  RITES OF THE DEMON [@DEATHWALKER #1]  P
CASTLE, DAMON  STARBRIGHT  P
CASTLE, JEFFERY LLOYD  SATELLITE E ONE  H:2,P
CASTLE, JEFFERY LLOYD  VANGUARD TO VENUS  H
CASTRO, ADAM-TROY  ALIEN DARKNESS, AN  LP
CASTRO, ADAM-TROY  DESPERATE, DECAYING DARKNESS, A  LP
CASTRO, ADAM-TROY  GATHERING OF THE SINISTER SIX, THE [SPIDER-MAN]  P
CASTRO, ADAM-TROY  VOSOFF AND NIMMITZ  H
CAT, IVAN  BURNING HEART OF NIGHT, THE  P
CAT, IVAN  EYES OF LIGHT AND DARKNESS, THE  P
CATALOGUE, THE  INDEX OF POSSIBILITIES, AN  SFWA-VLP
CATTON, JOHN PAUL  KITSUNE [@TIME HUNTER 4]  LP
CAUDETT, FRANK  A LA LUNA EN UN SEGUNDO [ENVIADO SECRETO 81]  FP
CAVE, HUGH B.  DISCIPLES OF DREAD  P
CAVE, HUGH B.  EVIL, THE  P
CAVE, HUGH B.  LEGION OF THE DEAD  C/P,P
CAVE, HUGH B.  MURGUNSTRUMM  SR-H
CAVE, HUGH B.  NEBULON HORROR, THE  P
CAVE, PETER  LAST ECHO, THE [INVASION: EARTH]  LP
CAVELOS, JEANNE  CASTING SHADOWS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #1]  C/P,P
CAVELOS, JEANNE  INVOKING DARKNESS [@BABYLON 5 PASSING OF THE TECHNO-MAGES #3]  C/P,P
CAVELOS, JEANNE  MANY FACES OF VAN HELSING, THE  LPA
CAVELOS, JEANNE  SHADOW WITHIN, THE [@BABYLON 5 #7]  C/P,P
CAVELOS, JEANNE  SUMMONING LIGHT [@BABYLON 5 PASSING OF THE TECHNO-MAGES #2]  C/P,P
CAWTHORN, JAMES/MOORCOCK, MICHAEL  FANTASY THE BEST 100 BOOKS  R/REF-H
CAYCE, EDGAR  EDGAR CAYCE ON ATLANTIS  [C/RAN-P]
CAYCE, EDGAR  VENTURE INWARD  [C/RAN-P]
CECILIONE, MICHAEL  DOMINATION  P
CENTAUR PRESS  SWORDSMEN AND SUPERMEN  PA
CENTAURI COMPANY  CENTAURI EXPRESS 1 #3 (CASSETTE)  SR-MM
CERASINI, MARC  AVP: ALIEN VS. PREDATOR [ALIENS/PREDATOR]  P
CERASINI, MARC  GODZILLA AT WORLD’S END  P
CERF, BENNETT  FAMOUS GHOST STORIES  PA
CERF, BENNETT  UNEXPECTED!, THE  SR-PA,C/PA
CERF, CHRISTOPHER  VINTAGE ANTHOLOGY OF SCIENCE FANTASY, THE  PA,C/PA
CHABON, MICHAEL  MCSWEENEY’S ENCHANTED CHAMBER OF ATONISHING STORIES  LPA
CHABON, MICHAEL  MCSWEENEY’S MAMMOTH TREASURY OF THRILLING TALES  LPA
CHADBOURN, MARK  ALWAYS FOREVER [AGE OF MISRULE 3]  LP
CHADBOURN, MARK  DARKEST HOUR [AGE OF MISRULE 2]  P

The MITSFS Pinkdex  September 24, 2007  Page 75
AUTHOR  TITLE  CODE
CHADBOURN, MARK  WONDERLAND [DOCTOR WHO]  SR-H
CHADBOURN, MARK  WORLD’S END [AGE OF MISRULE 1]  P
CHAFE, PAUL C.  DEATH OF THE PHOENIX [MISSION CRITICAL]  P
CHAFE, PAUL C.  DESTINY’S FORGE [@MAN-KZIN WARS]  H
CHALKER, JACK L.  AND THE DEVIL WILL DRAG YOU UNDER  C/P:2,P
CHALKER, JACK L.  AND THE DEVIL WILL DRAG YOU UNDER (PROOF)  SR-LP
CHALKER, JACK L.  BALSHAZZAR’S SERPENT [THREE KINGS 1]  H
CHALKER, JACK L.  BIRTH OF FLUX AND ANCHOR, THE [@SOUL RIDER #4]  C/P:2,SFWA-P,P
CHALKER, JACK L.  CERBERUS: A WOLF IN THE FOLD [@FOUR LORDS OF THE DIAMOND #2]  C/P:2,P
CHALKER, JACK L.  CHANGEWINDS, THE [CHANGEWINDS 1,2,3]  LP
CHALKER, JACK L.  CHARON: A DRAGON AT THE GATE [@FOUR LORDS OF THE DIAMOND #3]  C/P:2,P
CHALKER, JACK L.  CHILDREN OF FLUX AND ANCHOR [@SOUL RIDER #5]  C/P:2,SFWA-P,P
CHALKER, JACK L.  CYBERNETIC WALRUS, THE [@WONDERLAND GAMBIT #1]  C/P,SFWA-P,LP,H:2
CHALKER, JACK L.  DANCE BAND ON THE TITANIC  C/P:2,P
CHALKER, JACK L.  DANCERS IN THE AFTERGLOW  C/P:2,P
CHALKER, JACK L.  DANZATORI DEL CREPUSCOLO, I (ITALIAN)  FP

CHALKER, JACK L.

DEMONS AT RAINBOW BRIDGE, THE [QUINTARA 1] C/P,SFWA-P:2,H:2,P,

SFWA

CHALKER, JACK L.  DEMONS OF THE DANCING GODS [DANCING GODS 2]  C/P:2,P
CHALKER, JACK L.  DOWNTIMING THE NIGHT SIDE  C/P:2,SFWA-P,P

CHALKER, JACK L.

ECHOES OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 1]

C/P,LP,H

CHALKER, JACK L.  EMPIRES OF FLUX AND ANCHOR [@SOUL RIDER #2]  C/P:2,P
CHALKER, JACK L.  EXILES AT THE WELL OF SOULS [WELL WORLD 2]  C/P:2,P

CHALKER, JACK L.

FOUR LORDS OF THE DIAMOND, THE [FOUR LORDS OF THE DIAMOND 1,2,3,4]

H:2

CHALKER, JACK L.  FUNF ZAUBERSTEIN ZU BINDEN FUNF VER-SCHIEDNE WELTEN  FP
CHALKER, JACK L.  GHOST OF THE WELL OF SOULS [WELL WORLD]  P
CHALKER, JACK L.  GODS OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 3] H:2,LP
CHALKER, JACK L.  HORRORS OF THE DANCING GODS [DANCING GODS 5]  C/P:2,P
CHALKER, JACK L.  HOT-WIRED DODO, THE [@WONDERLAND GAMBIT #3]  H:2,P
CHALKER, JACK L.  HOTEL ANDROMEDA  C/PA,PA,SFWA-P
CHALKER, JACK L.  IDENTITY MATRIX, THE  C/P:2,P
CHALKER, JACK L.  INFORMAL BIOGRAPHY OF SCROOGE MCDUCK, AN  R/REF-P
CHALKER, JACK L.  JUNGLE OF STARS, A  C/P:2,P
CHALKER, JACK L.  KASPAR’S BOX [THREE KINGS 3]  P
CHALKER, JACK L.  KASPAR’S BOX (PROOF) [THREE KINGS 3]  SR-LP
CHALKER, JACK L.  LABYRINTH OF DREAMS, THE [@G.O.D. INC. #1]  C/P:2,SFWA-P,P
CHALKER, JACK L.  LILITH: A SNAKE IN THE GRASS [@FOUR LORDS OF THE DIAMOND #1]  C/P:2,P
CHALKER, JACK L.  LORDS OF THE MIDDLE DARK [@RINGS OF THE MASTER #1]  C/P:2,P
CHALKER, JACK L.  MARCH HARE NETWORK, THE [@WONDERLAND GAMBIT #2]  H:2,P
CHALKER, JACK L.  MASKS OF THE MARTYRS [@RINGS OF THE MASTER #4]  C/P:2,P
CHALKER, JACK L.  MASTERS OF FLUX AND ANCHOR [@SOUL RIDER #3]  C/P:2,SFWA-P,P
CHALKER, JACK L.  MAZE IN THE MIRROR, THE [@G.O.D. INC. #3]  C/P:2,SFWA-P,P
CHALKER, JACK L.  MEDUSA: A TIGER BY THE TAIL [@FOUR LORDS OF THE DIAMOND #4]  C/P:2,P
CHALKER, JACK L.  MELCHIOR’S FIRE [THREE KINGS 2]  H
CHALKER, JACK L.  MESSIAH CHOICE, THE  H:2,SFWA-P,P
CHALKER, JACK L.  MIDNIGHT AT THE WELL OF SOULS [WELL WORLD 1]  C/P:3,P
CHALKER, JACK L.  MOREAU FACTOR, THE  P
CHALKER, JACK L.  NINETY TRILLION FAUSTS, THE [QUINTARA 3]  H:2,SFWA-P,P
CHALKER, JACK L.  PIRATES OF THE THUNDER [@RINGS OF THE MASTER #2]  C/P:2,P
CHALKER, JACK L.  PRIAM’S LENS  P
CHALKER, JACK L.  QUEST FOR THE WELL OF SOULS [WELL WORLD 3]  C/P:3,P

CHALKER, JACK L./RESNICK, MIKE/ EFFINGER, GEORGE ALEC RED TAPE WAR, THE

C/P,P,H,SFWA

CHALKER, JACK L.  RETURN OF NATHAN BRAZIL, THE [WELL WORLD 4]  C/P:2,P
CHALKER, JACK L.  RIDERS OF THE WINDS [CHANGEWINDS 2]  C/P:2,SFWA-P,P
CHALKER, JACK L.  RIVER OF DANCING GODS, THE [DANCING GODS 1]  C/P:2,P
CHALKER, JACK L.  RUN TO CHAOS KEEP, THE [QUINTARA 2]  SFWA-P:2,H:2,P,SFWA:2
CHALKER, JACK L.  SEA IS FULL OF STARS, THE [WELL WORLD 6]  C/P,P
CHALKER, JACK L.  SHADOW DANCERS, THE [@G.O.D. INC. #2]  C/P:2,SFWA-P,P

CHALKER, JACK L.

SHADOW OF THE WELL OF SOULS [WELL WORLD WATCHERS AT THE WELL 2]

H:2,LP

CHALKER, JACK L.  SONGS OF THE DANCING GODS [DANCING GODS 4]  C/P:2,P
CHALKER, JACK L.  SPIRITS OF FLUX AND ANCHOR [@SOUL RIDER #1]  C/P:2,P
CHALKER, JACK L.  TWILIGHT AT THE WELL OF SOULS [WELL WORLD 5]  C/P:2,P
CHALKER, JACK L.  VENGEANCE OF THE DANCING GODS [DANCING GODS 3]  C/P:2,P
CHALKER, JACK L.  WAR OF SHADOWS, A  C/P:2,P
CHALKER, JACK L.  WAR OF THE MAELSTROM [CHANGEWINDS 3]  C/P:3,SFWA-P,P
CHALKER, JACK L.  WARRIORS OF THE STORM [@RINGS OF THE MASTER #3]  C/P:2,P
CHALKER, JACK L.  WATCHERS AT THE WELL, THE [WELL WORLD 6,7,8]  H:2
CHALKER, JACK L.  WEB OF THE CHOZEN, THE  C/P:2,P
CHALKER, JACK L.  WHEN THE CHANGEWINDS BLOW [CHANGEWINDS 1]  C/P:2,P
CHAMBERLIN, ANN  MERLIN OF ST. GILLES’ WELL, THE [JOAN OF ARC 1]  LP

The MITSFS Pinkdex  September 24, 2007  Page 76
AUTHOR  TITLE  CODE
CHAMBERLIN, ANN  MERLIN OF THE OAK WOOD, THE  P
CHAMBERS, ANDY  SURVIVAL INSTINCT [@NECROMUNDA]  P
CHAMBERS, HOWARD V.  FACTS ON THE FLYING SAUCER CONTROVERSY, THE  [C/RAN-H]
CHAMBERS, JAMES  ISAAC ASIMOV’S HISTORY OF I-BOTICS  HA
CHAMBERS, JANE  BURNING  P
CHAMBERS, ROBERT W.  GREEN MOUSE, THE  H
CHAMBERS, ROBERT W.  IN SEARCH OF THE UNKNOWN  LP
CHAMBERS, ROBERT W.  KING IN YELLOW, THE  SR-H,C/P:3,P
CHAMBERS, ROBERT W.  MAKER OF MOONS, THE  H
CHAMBERS, ROBERT W.  SLAYER OF SOULS, THE  H
CHAMBERS, ROBERT W.  TALKERS, THE  H
CHAMBERS, ROBERT W.  TREE OF HEAVEN, THE  H
CHAMBERS, STEPHEN  HOPE’S END  C/P,P
CHAMBERS, WHITMAN  INVASION!  H
CHAMPETIER, JOEL  DRAGON’S EYE, THE  LP
CHANCE, KAREN  CLAIMED BY SHADOW [CASSANDRA PALMER]  P
CHANCE, KAREN  TOUCH THE DARK [CASSANDRA PALMER]  P
CHANDLER, A. BERTRAM  ALTERNATE MARTIANS, THE  DM129
CHANDLER, A. BERTRAM  ALTERNATE ORBITS [GRIMES]  D13783
CHANDLER, A. BERTRAM  ANARCH LORDS, THE [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  BEYOND THE GALACTIC RIM [GRIMES]  DF237
CHANDLER, A. BERTRAM  BIG BLACK MARK, THE [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  BRING BACK YESTERDAY [GRIMES]  DD517
CHANDLER, A. BERTRAM  BROKEN CYCLE, THE [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  CATCH THE STAR WINDS  C/P:2,P
CHANDLER, A. BERTRAM  COILS OF TIME, THE  DM107,P
CHANDLER, A. BERTRAM  COMMODORE AT SEA, THE [GRIMES]/SPARTAN PLANET [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  CONTRABAND FROM OTHERSPACE [GRIMES]  DG609
CHANDLER, A. BERTRAM  DARK DIMENSIONS, THE [GRIMES]  D13783
CHANDLER, A. BERTRAM  DARK DIMENSIONS, THE [GRIMES]/RIM GODS, THE [GRIMES]  P
CHANDLER, A. BERTRAM  EMPRESS OF OUTER SPACE  DM129
CHANDLER, A. BERTRAM  FAR TRAVELER, THE [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  FRONTIER OF THE DARK  C/P,P
CHANDLER, A. BERTRAM  GATEWAY TO NEVER, THE [GRIMES]  D37062
CHANDLER, A. BERTRAM  GLORY PLANET  H:2
CHANDLER, A. BERTRAM  HAMELIN PLAGUE, THE  C/P:2,P
CHANDLER, A. BERTRAM  HARD WAY UP, THE [GRIMES]  D31755
CHANDLER, A. BERTRAM  INHERITORS, THE [GRIMES]  D37062
CHANDLER, A. BERTRAM  INHERITORS, THE [GRIMES]/GATEWAY TO NEVER [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  INTO THE ALTERNATE UNIVERSE [GRIMES]  DM107

CHANDLER, A. BERTRAM INTO THE ALTERNATE UNIVERSE [GRIMES]/ CONTRABAND FROM OTHERSPACE [GRIMES]

C/P,P

CHANDLER, A. BERTRAM  JOHN GRIMES: LIEUTENANT OF THE SURVEY SERVICE [GRIMES]  H
CHANDLER, A. BERTRAM  JOHN GRIMES: RESERVE COMMODORE [GRIMES]  H:2
CHANDLER, A. BERTRAM  JOHN GRIMES: RIM COMMANDER [GRIMES]  H
CHANDLER, A. BERTRAM  JOHN GRIMES: RIM RUNNER [GRIMES]  H
CHANDLER, A. BERTRAM  JOHN GRIMES: SURVEY CAPTAIN [GRIMES]  H
CHANDLER, A. BERTRAM  JOHN GRIMES: TRAMP CAPTAIN [GRIMES]  H
CHANDLER, A. BERTRAM  KELLY COUNTRY  C/P,P
CHANDLER, A. BERTRAM  LAST AMAZON, THE [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  MATILDA’S STEPCHILDREN [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  NEBULA ALERT  DG632
CHANDLER, A. BERTRAM  RENDEZVOUS ON A LOST WORLD  DF117
CHANDLER, A. BERTRAM  RIM GODS, THE [GRIMES]  D72400
CHANDLER, A. BERTRAM  RIM OF SPACE, THE [GRIMES]  C/P,DF133,P,H
CHANDLER, A. BERTRAM  RIM OF SPACE, THE [GRIMES]/SHIP FROM OUTSIDE, THE [GRIMES]  C/P,P
CHANDLER, A. BERTRAM  ROAD TO THE RIM, THE [GRIMES]  DH029
CHANDLER, A. BERTRAM  ROAD TO THE RIM, THE [GRIMES]/HARD WAY UP, THE [GRIMES]  P
CHANDLER, A. BERTRAM  SEA BEASTS, THE  P
CHANDLER, A. BERTRAM  SHIP FROM OUTSIDE, THE [GRIMES]  DF237
CHANDLER, A. BERTRAM  SPACE MERCENARIES  DM133
CHANDLER, A. BERTRAM  SPARTAN PLANET [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  STAR COURIER [GRIMES]  P
CHANDLER, A. BERTRAM  STAR LOOT [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  TO KEEP THE SHIP [GRIMES]  P
CHANDLER, A. BERTRAM  TO PRIME THE PUMP [GRIMES]  C/P:2,P
CHANDLER, A. BERTRAM  UP TO THE SKY IN SHIPS  SR-DNESFA2
CHANDLER, A. BERTRAM  WAY BACK, THE [GRIMES]  P
CHANDLER, A. BERTRAM  WILD ONES, THE [GRIMES]  C/P,P
CHANDLER, EVAN  DYING LIGHT  P
CHANT, JOY  GREY MANE OF MORNING, THE  H:2,LP,P
CHANT, JOY  HIGH KINGS, THE  P
CHANT, JOY  RED MOON AND BLACK MOUNTAIN  H:2,P
CHANT, JOY  WHEN VOIHA WAKES  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 77
AUTHOR  TITLE  CODE
CHAPMAN, D. D./TARZAN, DELORIS LEHMAN RED TIDE  P
CHAPMAN, GORDON  REX COLE JR. AND THE GRINNING GHOST  H
CHAPMAN, VERA  BLAEDUD THE BIRDMAN  C/P:2,P
CHAPMAN, VERA  GREEN KNIGHT, THE [THREE DAMOSELS 1]  C/P,P
CHAPMAN, VERA  KING ARTHUR’S DAUGHTER [THREE DAMOSELS 3]  C/P:2,P
CHAPMAN, VERA  KING’S DAMOSEL, THE [THREE DAMOSELS 2]  C/P,P
CHAPPELL, GEORGE  THROUGH THE ALIMENTARY CANAL WITH GUN AND CAMERA  H
CHARBONNEAU, LOUIS  BARRIER WORLD  C/P:2,P
CHARBONNEAU, LOUIS  CORPUS EARTHLING  SR-P,C/P
CHARBONNEAU, LOUIS  DOWN TO EARTH  C/P:2,P
CHARBONNEAU, LOUIS  EMBRYO  P
CHARBONNEAU, LOUIS  NO PLACE ON EARTH  H:2,P
CHARBONNEAU, LOUIS  PSYCHEDELIC-40  C/P:2,P
CHARBONNEAU, LOUIS  SENSITIVES, THE  P
CHARBONNEAU, LOUIS  SENTINEL STARS, THE  C/P:2,P
CHARKIN, PAUL  LIGHT OF MARS  SR-P
CHARKIN, PAUL  OTHER SIDE OF NIGHT, THE  SR-P
CHARLES, NEIL  BEYOND ZOASTER  SR-P
CHARLES, NEIL  LAND OF ESA, THE  SR-P
CHARLES, NEIL  PARA-ROBOT  SR-P
CHARLES, NEIL  PLANET THA  SR-P
CHARLES, NEIL  PRE-GARGANTUA  SR-P
CHARLES, NEIL  TITAN’S MOON  SR-P
CHARLES, NEIL  WORLD OF GOL  SR-P
CHARLES, ROBERT  COMET, THE  P,SFWA-P
CHARLES, STEVEN  ACADEMY OF TERROR [@PRIVATE SCHOOL #2]  P
CHARLES, STEVEN  ENEMY WITHIN, THE [@PRIVATE SCHOOL #5]  P
CHARLES, STEVEN  LAST ALIEN, THE [@PRIVATE SCHOOL #6]  P
CHARLES, STEVEN  NIGHTMARE SESSION [@PRIVATE SCHOOL #1]  P
CHARLES, STEVEN  SKELETON KEY [@PRIVATE SCHOOL #4]  P
CHARLES, STEVEN  WITCH’S EYE [@PRIVATE SCHOOL #3]  P
CHARNAS, SUZY MCKEE  BRONZE KING, THE [VALENTINE MARSH 1]  H,SFWA
CHARNAS, SUZY MCKEE  CONQUEROR’S CHILD, THE [HOLDFAST 4]  H,LP
CHARNAS, SUZY MCKEE  DOROTHEA DREAMS  P,SFWA,SFWA-P
CHARNAS, SUZY MCKEE  FURIES, THE [HOLDFAST 3]  C/P,P,H
CHARNAS, SUZY MCKEE  GOLDEN THREAD, THE [VALENTINE MARSH 3]  LP
CHARNAS, SUZY MCKEE  KINGDOM OF KEVIN MALONE, THE  H
CHARNAS, SUZY MCKEE  MOTHERLINES [HOLDFAST 2]  H,P
CHARNAS, SUZY MCKEE  SILVER GLOVE, THE [VALENTINE MARSH 2]  H
CHARNAS, SUZY MCKEE  SLAVE AND THE FREE, THE [HOLDFAST 1,2]  LP
CHARNAS, SUZY MCKEE  VAMPIRE TAPESTRY, THE  C/P,SFWA-P,P,H
CHARNAS, SUZY MCKEE  WALK TO THE END OF THE WORLD [HOLDFAST 1]  C/P:2,P

CHARRETTE, ROBERT N.

CHOOSE YOUR ENEMIES CAREFULLY [SHADOWRUN 2/ SECRETS OF POWER 2]

C/P:2,P

CHARRETTE, ROBERT N.  EYE OF THE SERPENT [AELWYN 2]  C/P,P
CHARRETTE, ROBERT N.  FIND YOUR OWN TRUTH [SHADOWRUN 3/SECRETS OF POWER 3]  C/P,P
CHARRETTE, ROBERT N.  HEIR TO THE DRAGON [@BATTLETECH 28]  P
CHARRETTE, ROBERT N.  INITIATION TO WAR [@MECHWARRIOR/BATTLETECH]  P
CHARRETTE, ROBERT N.  JUST COMPENSATION [SHADOWRUN 19]  C/P:2,P
CHARRETTE, ROBERT N.  KING BENEATH THE MOUNTAIN, A [AWAKENED KING 2]  P
CHARRETTE, ROBERT N.  KNIGHT AMONG KNAVES, A [AWAKENED KING 3]  P
CHARRETTE, ROBERT N.  NEVER DEAL WITH A DRAGON [SHADOWRUN 1/SECRETS OF POWER 1]  C/P:2,P
CHARRETTE, ROBERT N.  NEVER TRUST AN ELF [SHADOWRUN 6]  P
CHARRETTE, ROBERT N.  PRINCE AMONG MEN, A  P
CHARRETTE, ROBERT N.  TIMESPELL [AELWYN 1]  P
CHARRETTE, ROBERT N.  WIZARD OF BONES [AELWYN 3]  P
CHARRETTE, ROBERT N.  WOLF PACK [BATTLETECH 4]  C/P,P
CHARRETTE, ROBERT N.  WOLVES ON THE BORDER [BATTLETECH 25]  P
CHARTERIS, LESLIE  SAINT AND THE LAST HERO, THE  SR-P
CHARTERIS, LESLIE  FANTASTIC SAINT, THE  SFWA
CHARTERIS, LESLIE  SAINT ERRANT  P
CHARTERIS, LESLIE  SAINT STEPS IN, THE  H
CHARTWELL BOOKS, INC.  TRIGAN EMPIRE, THE  CX
CHASE, CAROL  HAWK’S FLIGHT  C/P,P
CHASE, ROBERT R.  CRUCIBLE [FOX 2]  P
CHASE, ROBERT R.  GAME OF FOX AND LION, THE [FOX 1]  P
CHASE, ROBERT R.  SHAPERS  P
CHAUNDLER, CHRISTINE  THIRTEENTH ORPHAN, THE  H
CHAUVIN, CY  MULTITUDE OF VISIONS, A  R/REF-P
CHAVEZ, RAY  THEY ARE ALL AROUND US  P
CHAYEFSKY, PADDY  ALTERED STATES  C/P,P,H
CHEETHAM, ANTHONY  SCIENCE AGAINST MAN  PA,C/PA
CHEETHAM, ERIKA  PROPHECIES OF NOSTRADAMUS, THE  [C/RAN-P]
CHEPAITIS, B. A.  FEAR OF GOD, THE [FEAR 2]  C/P,P

AUTHOR  TITLE  CODE
CHEPAITIS, B. A.  FEAR PRINCIPLE, THE [FEAR 1]  P
CHEPAITIS, B. A.  LEARNING FEAR [FEAR 3]  P
CHERNENKO, DAN  BASTARD KING, THE [@SCEPTER OF MERCY #1]  LP
CHERNENKO, DAN  CHERNAGOR PIRATES, THE [@SCEPTER OF MERCY #2]  LP
CHERNENKO, DAN  SCEPTER’S RETURN, THE [@SCEPTER OF MERCY #3]  LP
CHERRYH, C. J.  ANGEL WITH THE SWORD  H:2,P
CHERRYH, C. J.  ARAFEL’S SAGA [ARAFEL 1,2]  H:2
CHERRYH, C. J.  AT THE EDGE OF SPACE [HANAN REBELLION 1,2]  P
CHERRYH, C. J.  BOOK OF MORGAINE, THE [MORGAINE 1,2,3]  H:2
CHERRYH, C. J.  BROTHERS OF EARTH [HANAN REBELLION 1]  H:2,P
CHERRYH, C. J.  CHANUR’S HOMECOMING [CHANUR 4]  H:2,P
CHERRYH, C. J.  CHANUR’S LEGACY [CHANUR 5]  C/P,SFWA,P,H:2
CHERRYH, C. J.  CHANUR’S VENTURE [CHANUR 2]  H:2,P
CHERRYH, C. J.  CHERNEVOG  H:2,P
CHERRYH, C. J.  CLOUD’S RIDER  C/P,P,H
CHERRYH, C. J.  COLLECTED SHORT FICTION OF C.J. CHERRYH, THE  H
CHERRYH, C. J.  CUCKOO’S EGG  H:2,P
CHERRYH, C. J.  CYTEEN [UNION/ALLIANCE CYTEEN 1,2,3]  H
CHERRYH, C. J.  CYTEEN: THE BETRAYAL [UNION/ALLIANCE CYTEEN 1]  C/P:2,P
CHERRYH, C. J.  CYTEEN: THE REBIRTH [UNION/ALLIANCE CYTEEN 2]  C/P:2,P
CHERRYH, C. J.  CYTEEN: THE VINDICATION [UNION/ALLIANCE CYTEEN 3]  C/P:2,P
CHERRYH, C. J.  DEFENDER [ATEVI 5]  H:2,P
CHERRYH, C. J.  DELIVERER [ATEVI 9]  H
CHERRYH, C. J.  DESTROYER [ATEVI 7]  H:2
CHERRYH, C. J./FISH, LESLIE  DIRGE FOR SABIS, A [SWORD OF KNOWLEDGE #1]  C/P:2,SFWA-P,P
CHERRYH, C. J.  DIVINE RIGHT [@MEROVINGEN NIGHTS #5]  PA
CHERRYH, C. J.  DOWNBELOW STATION [UNION/ALLIANCE]  H:2,P
CHERRYH, C. J.  DREAMING TREE, THE [ARAFEL (REVISED) 1,2]  P
CHERRYH, C. J.  DREAMSTONE, THE [ARAFEL 1]  C/P:2,P
CHERRYH, C. J.  ENDGAME [@MEROVINGEN NIGHTS #7]  C/PA,PA
CHERRYH, C. J.  EXILE’S GATE [MORGAINE 4]  H:2,P
CHERRYH, C. J.  EXPLORER [ATEVI 6]  H:2
CHERRYH, C. J.  FADED SUN, THE [FADED SUN 1,2,3]  C/P:2,P
CHERRYH, C. J.  FADED SUN: KESRITH, THE [FADED SUN 1]  H:2,P
CHERRYH, C. J.  FADED SUN: KUTATH, THE [FADED SUN 3]  H:2,P
CHERRYH, C. J.  FADED SUN: SHON’JIR, THE [FADED SUN 2]  H:2,P
CHERRYH, C. J.  FAERY IN SHADOW  C/P,P,H
CHERRYH, C. J.  FESTIVAL MOON [@MEROVINGEN NIGHTS #1]  C/PA,PA
CHERRYH, C. J.  FEVER SEASON [@MEROVINGEN NIGHTS #2]  C/PA,PA
CHERRYH, C. J.  FINITY’S END [UNION/ALLIANCE]  H:2,P
CHERRYH, C. J.  FIRES OF AZEROTH [MORGAINE 3]  C/P:2,P
CHERRYH, C. J.  FLOOD TIDE [@MEROVINGEN NIGHTS #6]  C/PA,PA
CHERRYH, C. J.  FOREIGNER [ATEVI 1]  H:2,P
CHERRYH, C. J.  FORGE OF HEAVEN [GENE WARS 2]  H:2
CHERRYH, C. J.  FORTRESS IN THE EYE OF TIME [TRISTEN 1]  H:2,P
CHERRYH, C. J.  FORTRESS IN THE EYE OF TIME (PROOF) [TRISTEN 1]  SR-LP
CHERRYH, C. J.  FORTRESS OF DRAGONS [TRISTEN 4]  C/P,P,H
CHERRYH, C. J.  FORTRESS OF EAGLES [TRISTEN 2]  H:2,P
CHERRYH, C. J.  FORTRESS OF ICE [TRISTEN 5]  H
CHERRYH, C. J.  FORTRESS OF OWLS [TRISTEN 3]  C/P,P,H
CHERRYH, C. J.  FORTY THOUSAND IN GEHENNA [UNION/ALLIANCE]  H:2,P
CHERRYH, C. J.  GATE OF IVREL [MORGAINE 1]  C/P:2,P
CHERRYH, C. J./MORRIS, JANET  GATES OF HELL, THE [HEROES IN HELL]  C/P,P,H
CHERRYH, C. J.  GLASS AND AMBER  SR-H
CHERRYH, C. J.  GOBLIN MIRROR, THE  H:2
CHERRYH, C. J.  HAMMERFALL [GENE WARS 1]  H:2,P
CHERRYH, C. J.  HEAVY TIME  C/P,P,H:2
CHERRYH, C. J.  HELLBURNER  H:2,P
CHERRYH, C. J.  HESTIA  C/P:2,P
CHERRYH, C. J.  HUNTER OF WORLDS [HANAN REBELLION 2]  C/P,P,H
CHERRYH, C. J.  INHERITOR [ATEVI 3]  H:2,P
CHERRYH, C. J.  INVADER [ATEVI 2]  H:2
CHERRYH, C. J.  KIF STRIKE BACK, THE [CHANUR 3]  C/P,P,H
CHERRYH, C. J./MORRIS, JANET  KINGS IN HELL [HEROES IN HELL]  C/P,P
CHERRYH, C. J.  LEGIONS OF HELL [HEROES IN HELL]  C/P,P
CHERRYH, C. J.  LOIS & CLARK [LOIS & CLARK]  H
CHERRYH, C. J.  MERCHANTER’S LUCK [UNION/ALLIANCE]  C/P,P,H
CHERRYH, C. J.  PALADIN, THE  H:2,SFWA-P,P
CHERRYH, C. J.  PORT ETERNITY  H:2,P
CHERRYH, C. J.  PRECURSOR [ATEVI 4]  H:2,P
CHERRYH, C. J.  PRETENDER [ATEVI 8]  H
CHERRYH, C. J.  PRIDE OF CHANUR, THE [CHANUR 1]  C/P,P,H
CHERRYH, C. J.  PUERTA DE IVREL, LA [MORGAINE 1]  FP

The MITSFS Pinkdex  September 24, 2007  Page 79
AUTHOR  TITLE  CODE
CHERRYH, C. J./LACKEY, MERCEDES  REAP THE WHIRLWIND [SWORD OF KNOWLEDGE 3]  C/P:2,SFWA-P:2,P
CHERRYH, C. J.  RIDER AT THE GATE  H:2,P
CHERRYH, C. J.  RIMRUNNERS [UNION/ALLIANCE]  C/P,P,H
CHERRYH, C. J.  RUSALKA  C/P,P,H
CHERRYH, C. J.  SERPENT’S REACH  C/P,P,H
CHERRYH, C. J.  SMUGGLER’S GOLD [@MEROVINGEN NIGHTS #4]  PA,C/PA
CHERRYH, C. J.  SUNFALL  C/P,P,H

CHERRYH, C. J./LACKEY, MERCEDES/ ASIRE, NANCY/FISH, LESLIE SWORD OF KNOWLEDGE, THE [SWORD OF KNOWLEDGE 1,2,3]

LP

CHERRYH, C. J.  TREE OF SWORDS AND JEWELS, THE [ARAFEL 2]  C/P:3,P
CHERRYH, C. J.  TRIPOINT [UNION/ALLIANCE]  H:2,P
CHERRYH, C. J.  TROUBLED WATERS [@MEROVINGEN NIGHTS #3]  PA,C/PA
CHERRYH, C. J.  VISIBLE LIGHT  C/P,P,H
CHERRYH, C. J.  VOYAGER IN NIGHT  C/P,P,H
CHERRYH, C. J.  WAVE WITHOUT A SHORE  C/P:2,P
CHERRYH, C. J.  WELL OF SHIUAN [MORGAINE 2]  C/P:2,P
CHERRYH, C. J./ASIRE, NANCY  WIZARD SPAWN [SWORD OF KNOWLEDGE 2]  C/P:2,SFWA-P:2,P
CHERRYH, C. J.  YVGENIE  H:2,P
CHESBRO, GEORGE C.  AFFAIR OF SORCERERS, AN [MONGO FREDERICKSON 3]  H:2,P
CHESBRO, GEORGE C.  BEASTS OF VALHALLA, THE [MONGO FREDERICKSON 4]  H,P
CHESBRO, GEORGE C.  CITY OF WHISPERING STONE [MONGO FREDERICKSON 2]  C/P,P,H
CHESBRO, GEORGE C.  SHADOW OF A BROKEN MAN [MONGO FREDERICKSON 1]  H,P
CHESBRO, GEORGE C.  VEIL  C/P,P
CHESTER, DEBORAH  CHALICE, THE [@SWORD,RING AND CHALICE 3]  P
CHESTER, DEBORAH  CRIMSON CLAW, THE [@ALIEN CHRONICLES #2]  P
CHESTER, DEBORAH  CRYSTAL EYE, THE [@ALIEN CHRONICLES #3]  P
CHESTER, DEBORAH  GOLDEN ONE, THE [@ALIEN CHRONICLES #1]  C/P,P
CHESTER, DEBORAH  KING BETRAYED, THE [SWORD,RING AND CHALICE PHERESA 3]  P
CHESTER, DEBORAH  KING IMPERILLED, THE [SWORD,RING AND CHALICE PHERESA 4]  P
CHESTER, DEBORAH  QUEEN’S GAMBIT, THE [SWORD,RING AND CHALICE PHERESA 1]  C/P,P
CHESTER, DEBORAH  QUEEN’S KNIGHT, THE [SWORD,RING AND CHALICE PHERESA 2]  P
CHESTER, DEBORAH  REALM OF LIGHT [REIGN OF SHADOWS 3]  P
CHESTER, DEBORAH  REIGN OF SHADOWS [REIGN OF SHADOWS 1]  C/P,P
CHESTER, DEBORAH  RING, THE [@SWORD,RING AND CHALICE 2]  P
CHESTER, DEBORAH  SHADOW WAR [REIGN OF SHADOWS 2]  P
CHESTER, DEBORAH  SWORD, THE [@SWORD,RING AND CHALICE 1]  P
CHESTER, MICHAEL  MYSTERY OF THE LOST MOON, THE  H
CHESTER, WILLIAM L.  HAWK OF THE WILDERNESS [KIOGA 1]  P
CHESTER, WILLIAM L.  KIOGA OF THE UNKNOWN LAND [KIOGA 4]  C/P,P
CHESTER, WILLIAM L.  KIOGA OF THE WILDERNESS [KIOGA 2]  P
CHESTER, WILLIAM L.  ONE AGAINST A WILDERNESS [KIOGA 3]  P
CHESTERTON, G. K.  MAN WHO WAS THURSDAY, THE  C/P,P
CHETWIN, GRACE  ATHELING, THE  H,P
CHETWIN, GRACE  OUT OF THE DARK WORLD  P
CHETWIN, GRACE  RIDDLE AND THE RUNE, THE [TALES OF GOM]  P
CHETWIN, GRACE  STARSTONE, THE [TALES OF GOM]  P
CHETWYND-HAYES, R.  AND LOVE SURVIVED  P
CHETWYND-HAYES, R.  GRANGE, THE  P,SFWA-P
CHETWYND-HAYES, R.  OTHER SIDE, THE  SFWA-P
CHETWYND-HAYES, R.  UNBIDDEN, THE  P
CHEUNG, H. K.  LIFE TRANSFORM MIRROR (CHINESE)  FP
CHEUNG, H. K.  SHOTLAND PUZZLE, THE (CHINESE)  FP
CHI CHING CHOU  FLYING TOWARDS PLUTO (CHINESE)  FP
CHIANG, TED  STORIES OF YOUR LIFE AND OTHERS  H
CHIBA, MILAN  NOONBLAZE  P
CHILD, RICHARD WASHBURN  BLUE WALL, THE  H
CHILDREN’S PRESS  WHOPPER SPACE STORIES  HA

CHILDRESS, DAVID HATCHER/ SHAVER, RICHARD LOST CONTINENTS & THE HOLLOW EARTH

[C/RAN-H]

CHILSON, ROBERT  ANTES DEL FIN  FP
CHILSON, ROBERT  AS THE CURTAIN FALLS  C/P:2,P
CHILSON, ROBERT  BLACK AS BLOOD  P
CHILSON, ROBERT  MEN LIKE RATS  P
CHILSON, ROBERT  REFUGE [ROBOT CITY 5]  C/P:2,SFWA-P,P
CHILSON, ROBERT  ROUNDED WITH SLEEP  P
CHILSON, ROBERT  SHORES OF KANSAS, THE  C/P,P
CHILSON, ROBERT  STAR-CROWNED KINGS, THE  C/P:2,P
CHILTON, CHARLES  JOURNEY INTO SPACE  SR-P,C/P
CHILTON, CHARLES  RED PLANET, THE  SR-P
CHILTON, I. M.  NIGHTMARE  P
CHILTON, IRMA  TAKE AWAY THE FLOWERS  H
CHIPPINDALE, PETER  LAPTOP OF THE GODS  LP
CHIZMAR, RICHARD T.  BEST OF CEMETARY DANCE, THE, VOLUME ONE  LPA
CHIZMAR, RICHARD T.  BEST OF CEMETARY DANCE, THE, VOLUME TWO  LPA

The MITSFS Pinkdex  September 24, 2007  Page 80
AUTHOR  TITLE  CODE
CHIZMAR, RICHARD T.  EARTH STRIKES BACK, THE  PA
CHIZMAR, RICHARD T./MORRISH, ROBERT  OCTOBER DREAMS  LPA
CHOLFIN, BRYAN  BEST OF CRANK!, THE  HA,LPA
CHOPRA, DEPAK  RETURN OF MERLIN, THE  H
CHRISMAN, ARTHUR BOWIE  SHEN OF THE SEA: CHINESE STORIES FOR CHILDREN  H
CHRISTENSEN, DON R.  SPACE GHOST/SORCERESS OF CYBA-3, THE  BLB
CHRISTENSEN, MARK  ALOHA  H
CHRISTIAN, CATHERINE  PENDRAGON, THE  P
CHRISTIAN, DEBORAH  KAR KALIM  C/P,P
CHRISTIAN, DEBORAH  MAINLINE  P
CHRISTIAN, DEBORAH  TRUTHSAYER’S APPRENTICE, THE [LOREGIVER 1]  LP
CHRISTIE, AGATHA  BIG FOUR, THE  P
CHRISTOPHER, JOHN  BEYOND THE BURNING LANDS [LUKE 2]  H,P
CHRISTOPHER, JOHN  CAVES OF NIGHT, THE  P
CHRISTOPHER, JOHN  CITY OF GOLD AND LEAD, THE [TRIPODS 2]  C/P:2,P
CHRISTOPHER, JOHN  GUARDIANS, THE  C/P,P,H
CHRISTOPHER, JOHN  LITTLE PEOPLE, THE  H:2,P
CHRISTOPHER, JOHN  LONG WINTER, THE  C/P:2,P
CHRISTOPHER, JOHN  LOTUS CAVES, THE  P,SFWA-P
CHRISTOPHER, JOHN  NO BLADE OF GRASS  H:2,LP
CHRISTOPHER, JOHN  PENDULUM  H:2,P
CHRISTOPHER, JOHN  PLANET IN PERIL  C/P:2,P
CHRISTOPHER, JOHN  POOL OF FIRE, THE [TRIPODS 3]  C/P:2,P
CHRISTOPHER, JOHN  POSSESSORS, THE  H:2,P
CHRISTOPHER, JOHN  PRINCE IN WAITING, THE [LUKE 1]  H:2,P
CHRISTOPHER, JOHN  RAGGED EDGE, THE  H:2,P
CHRISTOPHER, JOHN  SWEENEY’S ISLAND  P
CHRISTOPHER, JOHN  SWORD OF THE SPIRITS, THE [LUKE 3]  C/P,P,H
CHRISTOPHER, JOHN  TWENTY-SECOND CENTURY, THE  C/P:2,P
CHRISTOPHER, JOHN  WHEN THE TRIPODS CAME [TRIPODS 0]  C/P,SFWA-P,P
CHRISTOPHER, JOHN  WHITE MOUNTAINS, THE [TRIPODS 1]  C/P,P,H
CHRISTOPHER, JOHN  WHITE VOYAGE, THE  P
CHRISTOPHER, JOHN  WILD JACK  H,SFWA-P,P
CHRISTOPHER, JOHN  WORLD IN WINTER, THE  P
CHRISTOPHER, JOHN  WRINKLE IN THE SKIN, A  P
CHRISTOPHER, JOHN  YEAR OF THE COMET, THE  LP
CHRISTOPHER, MILBOURNE  HOUDINI: THE UNTOLD STORY  [C/RAN-P]
CHRONISTER, ALAN B.  CRY WOLF  P
CHUNK, GEORGE  FOUNDATION AND EARTH PAPER BLOCK  SFWA-VLP
CHUPP, SAM  SINS OF THE FATHERS [WORLD OF DARKNESS WRAITH]  P
CHURCH, RALPH  MORK AND MINDY  P
CHURCHWARD, JAMES  CHILDREN OF MU, THE  [C/RAN-P:2]
CHURCHWARD, JAMES  COSMIC FORCES OF MU, THE  [C/RAN-P]
CHURCHWARD, JAMES  LOST CONTINENT OF MU, THE  [C/RAN-P]
CHURCHWARD, JAMES  SACRED SYMBOLS OF MU, THE  [C/RAN-P:2]
CHURCHWARD, JAMES  SECOND BOOK OF THE COSMIC FORCES OF MU, THE  [C/RAN-P]
CIENCIN, SCOTT  ANCIENT GAMES [@ELVEN WAYS #2]  P

CIENCIN, SCOTT/DECANDIDO, KEITH R. A./ DILMORE, KEVIN/JARMAN, HEATHER/ WARD, DAYTON BREAKDOWNS [@STAR TREK S.C.E. #7]

P

CIENCIN, SCOTT  DINOVERSE  H
CIENCIN, SCOTT  HIGH SPIRTS [CHARMED]  P
CIENCIN, SCOTT  LIGHT OF THE WORLD [@CHARMED]  P
CIENCIN, SCOTT  LOTUS AND THE ROSE, THE  P
CIENCIN, SCOTT  LUCK BE A LADY [@CHARMED]  P
CIENCIN, SCOTT/CIENCIN, DENISE  MORTAL FEAR [@BUFFY]  LP
CIENCIN, SCOTT/CIENCIN, DENISE  NEMESIS [@ANGEL]  P
CIENCIN, SCOTT  NIGHT OF GLORY [@ELVEN WAYS #3]  P
CIENCIN, SCOTT  NIGHT PARADE, THE [FORGOTTEN REALMS HARPERS #4]  C/P,P
CIENCIN, SCOTT  PARLIAMENT OF BLOOD  P
CIENCIN, SCOTT  SHADOWDALE [@FORGOTTEN REALMS AVATAR #1]  P
CIENCIN, SCOTT  SWEET SIXTEEN [@BUFFY]  P
CIENCIN, SCOTT  TANTRAS [@FORGOTTEN REALMS AVATAR #2]  P
CIENCIN, SCOTT  VAMPRIRE ODYSSEY, THE [DANIELLE WALTHERS 1]  P
CIENCIN, SCOTT/JOLLEY, DAN  VENGEANCE [@ANGEL]  P
CIENCIN, SCOTT  WAYS OF MAGIC, THE [@ELVEN WAYS #1]  P
CIENCIN, SCOTT  WILDINGS, THE [DANIELLE WALTHERS 2]  P
CIENCIN, SCOTT  WOLVES OF AUTUMN, THE  P
CLADINGBOEL, NEIL C.  TALE SPIN  SR-P
CLAGETT, JOHN  ORANGE R, THE  C/P,P
CLAGETT, JOHN  WORLD UNKNOWN, A  C/P,P
CLANCY, TOM  CARDINAL OF THE KREMLIN, THE [JACK RYAN]  P
CLANCY, TOM  HUNT FOR RED OCTOBER, THE  C/P:2,P
CLANCY, TOM  RED STORM RISING  C/P:2,P

AUTHOR  TITLE  CODE
CLANCY, TOM  WITHOUT REMORSE  P
CLAPHAM, MARK  HOPE [DOCTOR WHO EIGHTH DOCTOR]  P

CLAPHAM, MARK/MILLER, JON DE BURGH TWILIGHT OF THE GODS

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

P

CLARE, JOHN  PASSIONATE INVADERS, THE  H
CLAREMONT, CHRIS  DEBT OF HONOR [STAR TREK]  CX
CLAREMONT, CHRIS  FIRSTFLIGHT [NICOLE SHEA 1]  C/P:2,SFWA-P,P
CLAREMONT, CHRIS  GROUNDED! [NICOLE SHEA 2]  C/P:2,SFWA-P:2,P
CLAREMONT, CHRIS  HIGH FRONTIER  H
CLAREMONT, CHRIS/LUCAS, GEORGE  SHADOW DAWN [SHADOW WAR 2]  H:2
CLAREMONT, CHRIS/LUCAS, GEORGE  SHADOW STAR [SHADOW WAR 3]  P
CLAREMONT, CHRIS  SUNDOWNER [NICOLE SHEA 3]  P
CLAREMONT, CHRIS/MILLER, FRANK  WOLVERINE  CX
CLAREMONT, CHRIS  X-MEN THE BROOD SAGA, PART 1 [X-MEN]  SCX
CLARESON, THOMAS D.  MANY FUTURES, MANY WORLDS  R/REF-H
CLARESON, THOMAS D.  MUNDOS CREADOS I  FP
CLARESON, THOMAS D.  MUNDOS CREADOS II  FP
CLARESON, THOMAS D.  SCIENCE FICTION CRITICISM: AN ANNOTATED CHECKLIST  R/IND
CLARESON, THOMAS D.  SF: THE OTHER SIDE OF REALISM  R/REF-H,C/REF-H
CLARESON, THOMAS D.  SPECTRUM OF WORLDS, A  HA
CLARESON, THOMAS D.  VOICES FOR THE FUTURE 1  R/REF-H
CLARESON, THOMAS D./WYMER, THOMAS  VOICES FOR THE FUTURE 3  R/REF-P
CLARK, CURT  ANARCHAOS  C/P:2,P
CLARK, DOUGLAS W.  ALCHEMY UNLIMITED [CORWYN THE ALCHEMIST 1]  C/P,P
CLARK, DOUGLAS W.  REHEARSAL FOR A RENAISSANCE  C/P,P
CLARK, DOUGLAS W.  SAVING SOLACE [@DRAGONLANCE CHAMPIONS]  P
CLARK, DOUGLAS W.  WHIRLWIND ALCHEMY [CORWYN THE ALCHEMIST 2]  C/P,SFWA-P,P
CLARK, IAN  PRODIGY  P
CLARK, JAN  EARTH HERALD  P
CLARK, LYDIA BENSON  DEMON CAT  P
CLARK, LYDIA BENSON  SEANCE FOR SUSAN  P
CLARK, MARY  STOP THAT WITCH!  P
CLARK, RONALD  BOMB THAT FAILED, THE  H
CLARK, SIMON  DALEK FACTOR, THE [DOCTOR WHO]  SR-H
CLARK, SIMON  NIGHT OF THE TRIFFIDS, THE [TRIFFIDS 2]  P
CLARK, VIRGINIA ROYCE  GOING TOWARD JASON  LP
CLARK, WILLIAM J.  AUTHOR INDEX TO THE DOC SAVAGE MAGAZINE, AN  R/IND
CLARKE, A. V./BULMER, H. K.  CYBERNETIC CONTROLLER  SR-P
CLARKE, A. V./BULMER, H. K.  SPACE TREASON  SR-P
CLARKE, ARTHUR C.  AGAINST THE FALL OF NIGHT  C/P:2,P
CLARKE, ARTHUR C.  ANDERE SEITE DES HIMMELS, DIE  FP
CLARKE, ARTHUR C.  ARTHUR C. CLARKE’S JULY 20, 2019  H:2
CLARKE, ARTHUR C.  ASTOUNDING DAYS  R/REF-P,C/REF-H
CLARKE, ARTHUR C.  BEST OF ARTHUR C. CLARKE: 1937–1955, THE  SR-P,C/P
CLARKE, ARTHUR C./BENFORD, GREGORY  BEYOND THE FALL OF NIGHT  H:2,P
CLARKE, ARTHUR C.  CHALLENGE OF THE SPACESHIP, THE  [C/RAN-P:2]
CLARKE, ARTHUR C.  CHILDHOOD’S END  H:3,SR-H,P
CLARKE, ARTHUR C.  CITY AND THE STARS, THE  H:2,P
CLARKE, ARTHUR C.  COAST OF CORAL, THE  [C/RAN-P]
CLARKE, ARTHUR C.  COLLECTED STORIES OF ARTHUR C. CLARKE  H
CLARKE, ARTHUR C.  COMING OF THE SPACE AGE, THE  [C/RAN-H]
CLARKE, ARTHUR C./LEE, GENTRY  CRADLE  C/P,P,H
CLARKE, ARTHUR C.  DEEP RANGE, THE  C/P,SR-H,P,H
CLARKE, ARTHUR C.  DOLFIJNENEILAND  FP
CLARKE, ARTHUR C.  DOLPHIN ISLAND  C/P:2,P
CLARKE, ARTHUR C.  EARTHLIGHT  C/P,P,H
CLARKE, ARTHUR C.  EXPEDITION TO EARTH  C/P,P,H
CLARKE, ARTHUR C.  EXPLORATION OF SPACE, THE  [C/RAN-H,C/RAN-P:2]
CLARKE, ARTHUR C.  FALL OF MOONDUST, A  C/P,P,H
CLARKE, ARTHUR C.  FOUNTAINS OF PARADISE, THE  H:2,P
CLARKE, ARTHUR C.  FROM THE OCEAN, FROM THE STARS  SR-H,H
CLARKE, ARTHUR C./LEE, GENTRY  GARDEN OF RAMA, THE [RAMA 3]  H:2,P
CLARKE, ARTHUR C.  GHOST FROM THE GRAND BANKS, THE  C/P,P,H
CLARKE, ARTHUR C.  GLIDE PATH  SR-H,C/P:2,P
CLARKE, ARTHUR C.  GOING INTO SPACE  [C/RAN-H]
CLARKE, ARTHUR C.  GREETINGS, CARBON-BASED BIPEDS!  R/REF-H
CLARKE, ARTHUR C.  HAMMER OF GOD, THE  C/P,P,H
CLARKE, ARTHUR C.  ILE DES DAUPHINS, L’  FH
CLARKE, ARTHUR C.  IMPERIAL EARTH  C/P,P,H,SR-H
CLARKE, ARTHUR C.  ISLANDS IN THE SKY  C/P,P,H
CLARKE, ARTHUR C.  ISOLA NEL CIELO, L  FP
CLARKE, ARTHUR C./BAXTER, STEPHEN  LIGHT OF OTHER DAYS, THE  C/P,P,H
CLARKE, ARTHUR C.  LION OF COMARRE, THE/AGAINST THE FALL OF NIGHT  H:2,P
CLARKE, ARTHUR C.  LOST WORLDS OF 2001, THE  C/P:2,P
CLARKE, ARTHUR C.  MAKING OF A MOON, THE  R/REF-H

The MITSFS Pinkdex  September 24, 2007  Page 82
AUTHOR  TITLE  CODE
CLARKE, ARTHUR C.  MASTER OF SPACE  C/P,P
CLARKE, ARTHUR C.  MEETING WITH MEDUSA, A  TD1
CLARKE, ARTHUR C.  MORE THAN ONE UNIVERSE  C/P,P
CLARKE, ARTHUR C.  NINE BILLION NAMES OF GOD, THE  C/P:2,P
CLARKE, ARTHUR C.  1984: SPRING — A CHOICE OF FUTURES  [C/RAN-P]
CLARKE, ARTHUR C./HYAMS, PETER  ODYSSEY FILE, THE  R/REF-P
CLARKE, ARTHUR C.  OTHER SIDE OF THE SKY, THE  C/P,P,H
CLARKE, ARTHUR C.  PRELUDE TO MARS  H:2,SR-H
CLARKE, ARTHUR C.  PRELUDE TO SPACE  C/P:2,GN1.03,SR-P,P
CLARKE, ARTHUR C.  PROFILES OF THE FUTURE  [C/RAN-P:2]
CLARKE, ARTHUR C.  PROJECT SOLAR SAIL  C/PA,PA
CLARKE, ARTHUR C.  PROMISE OF SPACE, THE  [C/RAN-P]
CLARKE, ARTHUR C./LEE, GENTRY  RAMA II [RAMA 2]  C/P,P,H
CLARKE, ARTHUR C./LEE, GENTRY  RAMA REVEALED [RAMA 4]  H:2,P
CLARKE, ARTHUR C.  REACH FOR TOMORROW  SR-H,C/P:2,P
CLARKE, ARTHUR C.  RENDEZVOUS WITH RAMA [RAMA 1]  H:2,P
CLARKE, ARTHUR C.  REPORT ON PLANET THREE AND OTHER SPECULATIONS  [C/RAN-H,C/RAN-P:2]
CLARKE, ARTHUR C./MCQUAY, MIKE  RICHTER 10  C/P:2,P
CLARKE, ARTHUR C.  S. O. S. LUA  FP
CLARKE, ARTHUR C.  SANDS OF MARS, THE  H:2,SR-H,P
CLARKE, ARTHUR C./PROCTOR, GEORGE W.  SCIENCE FICTION HALL OF FAME 3, THE  HA
CLARKE, ARTHUR C.  SENTINEL, THE  H:2,P
CLARKE, ARTHUR C.  SONGS OF DISTANT EARTH, THE  C/P,P,H,SFWA
CLARKE, ARTHUR C.  SPACE DREAMERS, THE  P
CLARKE, ARTHUR C./BAXTER, STEPHEN  SUNSTORM [TIME ODYSSEY 2]  H
CLARKE, ARTHUR C.  TALES FROM PLANET EARTH  H:2,LP,P
CLARKE, ARTHUR C.  TALES FROM THE "WHITE HART"  H:2,P
CLARKE, ARTHUR C.  TALES OF TEN WORLDS  C/P:2,P
CLARKE, ARTHUR C.  TASTEN IN DE TOEKOMST  FP
CLARKE, ARTHUR C.  3001: THE FINAL ODYSSEY [@SPACE ODYSSEY #4]  H:2,P
CLARKE, ARTHUR C.  3001: THE FINAL ODYSSEY (PROOF) [@SPACE ODYSSEY #4]  SR-LP
CLARKE, ARTHUR C.  TIME PROBE  PA,C/PA
CLARKE, ARTHUR C./BAXTER, STEPHEN  TIME’S EYE (CD) [TIME ODYSSEY 1]  SR-MM
CLARKE, ARTHUR C./BAXTER, STEPHEN  TIME’S EYE (HARDCOVER AND CD) [TIME ODYSSEY 1]  H
CLARKE, ARTHUR C./BAXTER, STEPHEN  TIME’S EYE (PROOF) [TIME ODYSSEY 1]  SR-LP

CLARKE, ARTHUR C./ KUBE-MCDOWELL, MICHAEL P.

TRIGGER, THE

H

CLARKE, ARTHUR C.  2061: ODYSSEY THREE [@SPACE ODYSSEY #3]  H:2,P
CLARKE, ARTHUR C.  2001: A SPACE ODYSSEY [@SPACE ODYSSEY #1]  C/P,SFWA-P,H:2,SR-H,P
CLARKE, ARTHUR C.  2001: A SPACE ODYSSEY (REVISED) [@SPACE ODYSSEY #1]  P

CLARKE, ARTHUR C.

2001: A SPACE ODYSSEY THE FINAL CHAPTERS (LP RECORD) [SPACE ODYSSEY]

SR-MM

CLARKE, ARTHUR C.  2010: ODYSSEY TWO [@SPACE ODYSSEY #2]  H:2,P
CLARKE, ARTHUR C.  2010: ODYSSEY TWO (PROOF) [@SPACE ODYSSEY #2]  SR-LP
CLARKE, ARTHUR C.  VIEW FROM SERENDIP, THE  SFWA
CLARKE, ARTHUR C.  VOICES FROM THE SKY  [C/RAN-P:2]
CLARKE, ARTHUR C.  WIND FROM THE SUN, THE  C/P,H:2,P
CLARKE, JOAN  HAPPY PLANET, THE  H
CLARKE, LINDSAY  CHYMICAL WEDDING, THE  P
CLARKE, ROBERT  LESS THAN HUMAN  C/P:2,P
CLARKE, SUSANNA  JONATHAN STRANGE & MR. NORRELL  H
CLARO, JOE  SPACECAMP  P
CLASON, CLYDE B.  ARK OF VENUS  H
CLAUDY, CARL H.  BLUE GROTTO TERROR, THE  H:2
CLAUDY, CARL H.  LAND OF NO SHADOW, THE  H:2
CLAUDY, CARL H.  MYSTERY MEN OF MARS, THE  H:2
CLAUDY, CARL H.  THOUSAND YEARS A MINUTE, A  H
CLAYTON, JO  BAIT OF DREAMS, A  C/P:2,P
CLAYTON, JO  BLUE MAGIC [SOUL DRINKER 2]  C/P:2,P
CLAYTON, JO  BURNING GROUND, THE [SHADOWSONG 2]  C/P:2,P
CLAYTON, JO  CHANGER’S MOON [DUEL OF SORCERY 3]  C/P:2,P
CLAYTON, JO  CRYSTAL HEAT [@SHADOWSONG #3]  C/P,P
CLAYTON, JO  DANCE DOWN THE STARS [@DANCER #3]  C/P:2,P
CLAYTON, JO  DANCER’S RISE [@DANCER #1]  C/P:2,P
CLAYTON, JO  DIADEM FROM THE STARS [DIADEM 1]  C/P:3,P
CLAYTON, JO  DRINKER OF SOULS [SOUL DRINKER 1]  C/P:2,P
CLAYTON, JO  DRUM CALLS [DRUMS OF CHAOS 2]  P
CLAYTON, JO/MURPHY, KEVIN ANDREW  DRUM INTO SILENCE [DRUMS OF CHAOS 3]  H,P
CLAYTON, JO  DRUM WARNING [DRUMS OF CHAOS 1]  C/P,P
CLAYTON, JO  FIRE IN THE SKY [@SHADOWSONG #1]  C/P:2,P
CLAYTON, JO  GATHERING OF STONES, A [SOUL DRINKER 3]  C/P:2,P
CLAYTON, JO  GHOSTHUNT [DIADEM 7]  C/P:2,P
CLAYTON, JO  IRSUD [DIADEM 3]  C/P:2,P
CLAYTON, JO  LAMARCHOS [DIADEM 2]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 83
AUTHOR  TITLE  CODE
CLAYTON, JO  MAEVE [DIADEM 4]  C/P:2,P
CLAYTON, JO  MAGIC WARS, THE [@WILD MAGIC #3]  C/P,P
CLAYTON, JO  MOONGATHER [DUEL OF SORCERY 1]  C/P:2,P
CLAYTON, JO  MOONSCATTER [DUEL OF SORCERY 2]  C/P:2,P
CLAYTON, JO  NOWHERE HUNT, THE [DIADEM 6]  C/P:2,P
CLAYTON, JO  QUESTER’S ENDGAME [DIADEM 9]  C/P:2,P
CLAYTON, JO  SERPENT WALTZ [@DANCER #2]  C/P:2,P
CLAYTON, JO  SHADOW OF THE WARMASTER  C/P:2,P
CLAYTON, JO  SHADOWKILL [SHADITH’S QUEST 3/DIADEM 12]  C/P,P
CLAYTON, JO  SHADOWPLAY [SHADITH’S QUEST 1/DIADEM 10]  C/P:2,P
CLAYTON, JO  SHADOWSPEER [SHADITH’S QUEST 2/DIADEM 11]  C/P:2,P
CLAYTON, JO  SKEEN’S LEAP [SKEEN 1]  C/P:2,P
CLAYTON, JO  SKEEN’S RETURN [SKEEN 2]  C/P:2,P
CLAYTON, JO  SKEEN’S SEARCH [SKEEN 3]  C/P:2,P
CLAYTON, JO  SNARES OF IBEX, THE [DIADEM 8]  C/P:2,P
CLAYTON, JO  SOUL DRINKER, THE [SOUL DRINKER 1,2,3]  H:2
CLAYTON, JO  STAR HUNTERS [DIADEM 5]  C/P:2,P
CLAYTON, JO  WILD MAGIC [WILD MAGIC 1]  C/P:2,P
CLAYTON, JO  WILDFIRE [@WILD MAGIC #2]  C/P:2,SFWA-P,P
CLEE, MONA  BRANCH POINT  C/P:2,P
CLEE, MONA  OVERSHOOT  P
CLEGG, BARBARA  ENLIGHTENMENT [DOCTOR WHO]  SR-P,C/P:2
CLEGG, DOUGLAS  LADY OF THE SERPENTS, THE [@VAMPYRICON 2]  H
CLEGG, DOUGLAS  MORDRED, BASTARD SON  H

CLEM, RALPH/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

CITY: 2000 A.D.: URBAN LIFE THROUGH SCIENCE FICTION, THE

C/PA,PA

CLEMENS, JAMES  HINTERLAND [GODSLAYER 2]  H
CLEMENS, JAMES  SHADOWFALL [GODSLAYER 1]  H:2
CLEMENS, JAMES  WIT’CH FIRE [BANNED AND THE BANISHED 1]  C/P:2,LP
CLEMENS, JAMES  WIT’CH GATE [BANNED AND THE BANISHED 4]  P
CLEMENS, JAMES  WIT’CH STAR [BANNED AND THE BANISHED 5]  P
CLEMENS, JAMES  WIT’CH STORM [BANNED AND THE BANISHED 2]  C/P,LP
CLEMENS, JAMES  WIT’CH WAR [BANNED AND THE BANISHED 3]  H,LP
CLEMENS, RODGERS  PRESENCE, THE  P
CLEMENT, AERON  COLD MOONS, THE  LP
CLEMENT, HAL  BEST OF HAL CLEMENT, THE  C/P:2,P
CLEMENT, HAL  BOTSCHAFTER VON DEN STERNEN  FP
CLEMENT, HAL  CLOSE TO CRITICAL  C/P:2,P
CLEMENT, HAL  CYCLE OF FIRE  C/P:2,P
CLEMENT, HAL  ESSENTIAL HAL CLEMENT 1, THE: TRIO FOR SLIDE RULE & TYPEWRITER  H
CLEMENT, HAL  ESSENTIAL HAL CLEMENT 2, THE: MUSIC OF MANY SPHERES  H

CLEMENT, HAL ESSENTIAL HAL CLEMENT 3, THE:

VARIATIONS ON A THEME BY SIR ISAAC NEWTON

H

CLEMENT, HAL  FOSSIL [ISAAC’S UNIVERSE]  C/P:2,P
CLEMENT, HAL  FROM OUTER SPACE  C/P:2,P
CLEMENT, HAL  HALF LIFE  H:2,P
CLEMENT, HAL  HEAVY PLANET  LP
CLEMENT, HAL  ICEWORLD  SR-H,C/P:2,P
CLEMENT, HAL  INTUIT  H
CLEMENT, HAL  MICROBE DETECTIVE, LE  FP
CLEMENT, HAL  MISSION OF GRAVITY  C/P,P,H
CLEMENT, HAL  NATIVES OF SPACE  C/P:2,P
CLEMENT, HAL  NEEDLE  C/P,P,H
CLEMENT, HAL  NITROGEN FIX, THE  C/P,P,H
CLEMENT, HAL  NOISE  H
CLEMENT, HAL  OCEAN ON TOP  C/P:2,P
CLEMENT, HAL  PERSECUCION COSMICA  FP
CLEMENT, HAL  SMALL CHANGES  H
CLEMENT, HAL  SPACE LASH  C/P:2,P
CLEMENT, HAL  STAR LIGHT  C/P:2,P
CLEMENT, HAL  STILL RIVER  H:2,P
CLEMENT, HAL  SYMBIOSE  FH
CLEMENT, HAL  THROUGH THE EYE OF A NEEDLE  SR-P,C/P:2
CLEMENT, HENRY  SHE WAITS  P
CLEVE, JOHN  ASSIGNMENT: HELLHOLE [@SPACEWAYS #14]  P
CLEVE, JOHN  CARNADYNE HORDE, THE [@SPACEWAYS #17]  P
CLEVE, JOHN  CORUNDUM’S WOMAN [@SPACEWAYS #2]  P
CLEVE, JOHN  ESCAPE FROM MACHO [@SPACEWAYS #3]  C/P:2,P
CLEVE, JOHN  ICEWORLD CONNECTION, THE [@SPACEWAYS #11]  P
CLEVE, JOHN  IN QUEST OF QALARA [@SPACEWAYS #9]  P
CLEVE, JOHN  JONUTA RISING! [@SPACEWAYS #13]  P
CLEVE, JOHN  KING OF THE SLAVERS [@SPACEWAYS #19]  C/P,P
CLEVE, JOHN  MANHUNTRESS, THE [@SPACEWAYS #7]  P
CLEVE, JOHN  MANLIB  SR-L

The MITSFS Pinkdex  September 24, 2007  Page 84
AUTHOR  TITLE  CODE
CLEVE, JOHN  MASTER OF MISFIT [@SPACEWAYS #5]  P
CLEVE, JOHN  OF ALIEN BONDAGE [@SPACEWAYS #1]  P
CLEVE, JOHN  PLANET MURDERER, THE [@SPACEWAYS #16]  P
CLEVE, JOHN  PURRFECT PLUNDER [@SPACEWAYS #6]  P
CLEVE, JOHN  RACE ACROSS THE STARS [@SPACEWAYS #18]  P
CLEVE, JOHN  SATANA ENSLAVED [@SPACEWAYS #4]  P
CLEVE, JOHN  SEXORCIST, THE  SR-L
CLEVE, JOHN  STAR SLAVER [@SPACEWAYS #12]  P
CLEVE, JOHN  STARSHIP SAPPHIRE [@SPACEWAYS #15]  P
CLEVE, JOHN  UNDER TWIN SUNS [@SPACEWAYS #8]  P
CLEVE, JOHN  YOKE OF SHEN, THE [@SPACEWAYS #10]  P
CLIFFORD, SARAH  ADAM AND HIS WOMEN  P
CLIFTON, MARK  EIGHT KEYS TO EDEN  H:2,P
CLIFTON, MARK/RILEY, FRANK  FOREVER MACHINE, THE  SR-P,C/P:2
CLIFTON, MARK  SCIENCE FICTION OF MARK CLIFTON, THE  H
CLIFTON, MARK/RILEY, FRANK  THEY’D RATHER BE RIGHT  H
CLIFTON, MARK  WHEN THEY COME FROM SPACE  C/P,P,H
CLINE, LEONARD  DARK CHAMBER, THE  P
CLINE, LEONARD  GOD HEAD  H
CLINE, LEONARD  LISTEN, MOON!  H
CLINE, LINDA  MIRACLE SEASON, THE  P
CLINGERMAN, MILDRED  CUPFUL OF SPACE, A  C/P:2,P
CLINTON, JEFF  KANE’S ODYSSEY  SR-P
CLIVE, DENNIS  VALLEY OF PRETENDERS  B
CLIVE, DENNIS  VOICE COMMANDS, THE  B
CLOUGH, BRENDA W.  CRYSTAL CROWN, THE [AVERIDAN 1]  C/P:2,P
CLOUGH, BRENDA W.  DOORS OF DEATH AND LIFE [PEARL OF IMMORTALITY 2]  H
CLOUGH, BRENDA W.  DRAGON OF MISHBIL, THE [AVERIDAN 4]  C/P:2,P
CLOUGH, BRENDA W.  HOW LIKE A GOD [PEARL OF IMMORTALITY 1]  H,P
CLOUGH, BRENDA W.  NAME OF THE SUN, THE [AVERIDAN 3]  C/P:2,P
CLOUGH, BRENDA W.  REALM BENEATH, THE [AVERIDAN 2]  C/P:2,P
CLOW, MARTHA DEMEY  STARBREED  C/P:2,P
CLOWES, CAROLYN  PANDORA PRINCIPLE, THE [STAR TREK]  C/P:3,P
CLUTE, JOHN  APPLESEED  LP
CLUTE, JOHN/GRANT, JOHN  ENCYCLOPEDIA OF FANTASY, THE  R/REF-H
CLUTE, JOHN/NICHOLLS, PETER  ENCYCLOPEDIA OF SCIENCE FICTION, THE  R/REF-H

CLUTE, JOHN/GREENLAND, COLIN/ PRINGLE, DAVID INTERZONE: THE FIRST ANTHOLOGY [INTERZONE 1]

HA

CLUTE, JOHN/PRINGLE, DAVID/ OUNSLEY, SIMON INTERZONE: THE SECOND ANTHOLOGY [INTERZONE 2]

PA

CLUTE, JOHN  SCIENCE FICTION: THE ILLUSTRATED ENCYCLOPEDIA  R/REF-H
CLUTE, JOHN/DORSEY, CANDAS JANE  TESSERACTS 8  LPA

CLYMER, THEODORE/RUTH, LEO/ EVANECHKO, PETER/HIGGS, JULIA SPECULATIONS

LP

COATES, JOHN  HERE TODAY  H
COATSWORTH, ELIZABETH  CAT WHO WENT TO HEAVEN, THE  P
COATSWORTH, ELIZABETH  ENCHANTED, THE  P
COBLENTZ, STANTON A.  AFTER 12,000 YEARS  SR-H,SR-LP
COBLENTZ, STANTON A.  ANIMAL PEOPLE, THE  C/P:2,P
COBLENTZ, STANTON A.  CRIMSON CAPSULE, THE  H:2
COBLENTZ, STANTON A.  DAY THE WORLD STOPPED, THE  H:2
COBLENTZ, STANTON A.  HIDDEN WORLD  C/P,P
COBLENTZ, STANTON A.  INTO PLUTONIAN DEPTHS  SR-P
COBLENTZ, STANTON A.  ISLAND PEOPLE, THE  P
COBLENTZ, STANTON A.  LORD OF TRANERICA  LP
COBLENTZ, STANTON A.  MOON PEOPLE, THE  C/P,P
COBLENTZ, STANTON A.  NEXT DOOR TO THE SUN  LP
COBLENTZ, STANTON A.  PLANET OF YOUTH, THE  H
COBLENTZ, STANTON A.  SUNKEN WORLD, THE  SR-LP
COBLENTZ, STANTON A.  UNDER THE TRIPLE SUNS  SR-H,H
COBLENTZ, STANTON A.  WHEN THE BIRDS FLY SOUTH  LP
COBLEY, MICHAEL  SHADOWKINGS [@SHADOWKINGS #1]  LP
COCHRAN, LIN NIELSEN/HANCOCK, SCOTT  SOUTHERN FRIED SCI-FI AND JAMBALAYA GENRES  SR-L,LPA
COCHRAN, MOLLY/MURPHY, WARREN  BROKEN SWORD, THE [FOREVER KING 2]  C/P,P,H
COCHRAN, MOLLY/MURPHY, WARREN  FOREVER KING, THE [FOREVER KING 1]  H:2,P
COCHRAN, MOLLY  THIRD MAGIC, THE [FOREVER KING 3]  P
COCHRAN, MOLLY/MURPHY, WARREN  WORLD WITHOUT END  H,P
COCHRANE, WILLIAM E.  CLASS SIX CLIMB  P
COCKCROFT, T. G. L.  INDEX TO FICTION IN RADIO NEWS AND OTHER MAGAZINES  R/IND-F
COCKCROFT, T. G. L.  INDEX TO THE WEIRD FICTION MAGAZINES — AUTHOR  SRL/IND
COCKCROFT, T. G. L.  INDEX TO THE WEIRD FICTION MAGAZINES — TITLE  SRL/IND
COCKRELL, MARIAN  SHADOW CASTLE  P

COCKRUM, KURT/LEVACK, DANIEL J. H./ UNDERWOOD, TIM FANTASMS 2.1:

A BIBLIOGRAPHY OF THE PUBLISHED FICTION OF JACK VANCE R/IND

The MITSFS Pinkdex  September 24, 2007  Page 85
AUTHOR  TITLE  CODE
COE, DAVID B.  BONDS OF VENGEANCE [WINDS OF THE FORELANDS 3]  H
COE, DAVID B.  CHILDREN OF AMARID [LON TOBYN 1]  C/P:2,P
COE, DAVID B.  EAGLE-SAGE [LON TOBYN 3]  H,P
COE, DAVID B.  OUTLANDERS, THE [LON TOBYN 2]  C/P:2,P
COE, DAVID B.  RULES OF ASCENSION [WINDS OF THE FORELANDS 1]  H
COE, DAVID B.  SEEDS OF BETRAYAL [WINDS OF THE FORELANDS 2]  P
COE, DAVID B.  SHAPERS OF DARKNESS [WINDS OF THE FORELANDS 4]  H
COE, DAVID B.  WEAVERS OF WAR [WINDS OF THE FORELANDS 5]  H
COE, ROSS ANTON  SORCERER’S BLOOD [@WARRIOR OF VENGEANCE #1]  P
COE, ROSS ANTON  TRAILS OF PERIL [@WARRIOR OF VENGEANCE #2]  P
COFFEY, FRANK  MODERN MASTERS OF HORROR  HA
COFFEY, FRANK  SHAMAN, THE  H,P
COFFMAN, VIRGINIA  CHALET DIABOLIQUE [LUCIFER COVE 5]  P
COFFMAN, VIRGINIA  DEVIL’S MISTRESS, THE [LUCIFER COVE 1]  P
COFFMAN, VIRGINIA  DEVIL’S VIRGIN, THE [LUCIFER COVE 3]  P
COFFMAN, VIRGINIA  FROM SATAN, WITH LOVE [LUCIFER COVE 6]  P
COFFMAN, VIRGINIA  MASQUE OF SATAN [LUCIFER COVE 4]  P
COFFMAN, VIRGINIA  PRIESTESS OF THE DAMNED [LUCIFER COVE 2]  P
COGGINS, JACK/PRATT, FLETCHER  BY SPACE SHIP TO THE MOON  [C/RAN-H]
COGGINS, JACK/PRATT, FLETCHER  ROCKETS, JETS, GUIDED MISSILES, AND SPACE SHIPS  [C/RAN-H]
COGSWELL, THEODORE R.  PROCEEDINGS OF THE INSTITUTE FOR TWENTY-FIRST CENTURY STUDIES  R/REF-H

COGSWELL, THEODORE R./ SPANO, CHARLES A.

SPOCK, MESSIAH! [STAR TREK]

C/P:3,P

COGSWELL, THEODORE R.  THIRD EYE, THE  C/P,P
COGSWELL, THEODORE R.  WALL AROUND THE WORLD, THE  C/P:2,P
COHEN, BARBARA  UNICORNS IN THE RAIN  H
COHEN, BARNEY  BLOOD ON THE MOON  P
COHEN, BARNEY/BAEN, JAMES  TAKING OF SATCON STATION, THE  P
COHEN, DANIEL  MONSTERS OF STAR TREK, THE  R/REF-P
COHEN, DANIEL  SUPERMONSTERS  R/REF-P
COHEN, GARY G.  CIVILIZATION’S LAST HURRAH  P
COHEN, JAMES/RUBIE, PETER  MINDBENDER  C/P,P
COHEN, JOEL H.  SIX MILLION DOLLAR MAN AND THE BIONIC WOMAN, THE  [C/RAN-P]
COHEN, MICHAEL  JOHN BIRCH COLORING BOOK, THE  CX
COLANDER, VALERIE N.  NEENA GATHERING  C/P:2,P
COLE, ADRIAN  BANE OF NIGHTMARES [@DREAM LORDS #3]  P
COLE, ADRIAN  BLOOD RED ANGEL  P
COLE, ADRIAN  CRIMSON TALISMAN, THE [@EBERRON WAR-TORN #1]  P
COLE, ADRIAN  GODS IN ANGER, THE [@OMARAN #4]  P
COLE, ADRIAN  KING OF LIGHT AND SHADOWS, THE [@OMARAN #3]  P
COLE, ADRIAN  LABYRINTH OF WORLDS [@STAR REQUIEM #4]  C/P,SFWA-P,P
COLE, ADRIAN  LORD OF NIGHTMARES [@DREAM LORDS #2]  P
COLE, ADRIAN  MOTHER OF STORMS [@STAR REQUIEM #1]  P,SFWA-P
COLE, ADRIAN  PLACE AMONG THE FALLEN, A [@OMARAN #1]  H,SFWA,P
COLE, ADRIAN  PLAGUE OF NIGHTMARES, A [@DREAM LORDS #1]  C/P:2,P
COLE, ADRIAN  THIEF OF DREAMS [@STAR REQUIEM #2]  P,SFWA-P
COLE, ADRIAN  THRONE OF FOOLS [@OMARAN #2]  P
COLE, ADRIAN  WARLORD OF HEAVEN [@STAR REQUIEM #3]  P,SFWA-P
COLE, ALLAN/BUNCH, CHRIS  COURT OF A THOUSAND SUNS, THE [STEN 3]  C/P:2,P
COLE, ALLAN/BUNCH, CHRIS  EMPIRE’S END [STEN 8]  C/P:2,P
COLE, ALLAN/BUNCH, CHRIS  FAR KINGDOMS, THE  C/P,P
COLE, ALLAN/BUNCH, CHRIS  FLEET OF THE DAMNED [STEN 4]  C/P:2,P
COLE, ALLAN  GODS AWAKEN, THE [TIMURA 3]  P
COLE, ALLAN/BUNCH, CHRIS  KINGDOMS OF THE NIGHT [ANTEROS]  C/P,P
COLE, ALLAN/BUNCH, CHRIS  RETURN OF THE EMPEROR, THE [STEN 6]  C/P:2,P
COLE, ALLAN/BUNCH, CHRIS  REVENGE OF THE DAMNED [STEN 5]  C/P:2,P
COLE, ALLAN/BUNCH, CHRIS  STEN [STEN 1]  C/P:2,P
COLE, ALLAN/BUNCH, CHRIS  VORTEX [STEN 7]  C/P:2,P
COLE, ALLAN  WARRIOR RETURNS, THE [@ANTEROS]  H,P
COLE, ALLAN/BUNCH, CHRIS  WARRIOR’S TALE, THE  H,P
COLE, ALLAN  WIZARD OF THE WINDS [TIMURA 1]  LP
COLE, ALLAN/BUNCH, CHRIS  WOLF WORLDS, THE [STEN 2]  C/P:2,P
COLE, ALLAN  WOLVES OF THE GODS [TIMURA 2]  P
COLE, BURT  FUNCO FILE, THE  H:2,P
COLE, BURT  QUICK, THE  H
COLE, DEMARIS  TOKEN OF DRAGONSBLOOD  C/P,P
COLE, E. B.  PHILOSOPHICAL CORPS, THE  H:2
COLE, KRESLEY  HUNGRE LIKE NO OTHER, A [IMMORTALS AFTER DARK 1]  H
COLE, KRESLEY  NO REST FOR THE WICKED [IMMORTALS AFTER DARK 2]  H
COLE, STEPHEN  ART OF DESTRUCTION, THE [@DOCTOR WHO]  H
COLE, STEPHEN  FEAST OF THE DROWNED, THE [DOCTOR WHO TENTH DOCTOR]  H
COLE, STEPHEN  MONSTERS INSIDE, THE [DOCTOR WHO NINTH DOCTOR]  H
COLE, STEPHEN  MORE SHORT TRIPS [DOCTOR WHO]  PA

The MITSFS Pinkdex  September 24, 2007  Page 86
AUTHOR  TITLE  CODE

COLE, STEPHEN PROFESOR BERNICE SUMMERFIELD AND THE GODS OF THE

UNDERWORLD [BERNICE SUMMERFIELD 2]

C/P,P

COLE, STEPHEN  SHORT TRIPS [DOCTOR WHO]  C/PA,PA
COLE, STEPHEN/RAYNER, JACQUELINE  SHORT TRIPS AND SIDE STEPS [DOCTOR WHO]  PA
COLE, STEPHEN  TEN LITTLE ALIENS [DOCTOR WHO PAST DOCTORS]  P
COLE, STEPHEN  TIMELESS [DOCTOR WHO EIGHTH DOCTOR]  P
COLE, STEPHEN  TO THE SLAUGHTER [DOCTOR WHO EIGHTH DOCTOR]  P
COLE, STEPHEN  VANISHING POINT [DOCTOR WHO EIGHTH DOCTOR]  P
COLE, W. R.  CHECKLIST OF SCIENCE FICTION ANTHOLOGIES, A  SRL/IND
COLEBATCH, HAL  WUNDER WAR, THE [@MAN-KZIN WARS #10]  H
COLEMAN, JAMES A.  MODERNE THEORIEEN OVER HET HEELAL  FP
COLEMAN, JAMES A.  RELATIVITY FOR THE LAYMAN  [C/RAN-P]
COLEMAN, JAMES NELSON  NULL-FREQUENCY IMPULSER, THE  C/P:2,P
COLEMAN, JAMES NELSON  SEEKER FROM THE STARS  C/P:2,P
COLEMAN, LOREN L.  BINDING FORCE [@BATTLETECH 32]  P
COLEMAN, LOREN L.  BLOOD OF THE ISLE [@MECHWARRIOR DARK AGE/BATTLETECH]  P
COLEMAN, LOREN L.  BLOOD OF WOLVES [@AGE OF CONAN KERN #1]  P
COLEMAN, LOREN L.  BLOODLINES [@MAGIC: THE GATHERING ARTIFACTS 4]  C/P,P

COLEMAN, LOREN L.

BY TEMPTATIONS AND BY WAR [@MECHWARRIOR DARK AGE/ BATTLETECH]

LP

COLEMAN, LOREN L.

BY TEMPTATIONS AND BY WAR (PROOF) [@MECHWARRIOR DARK AGE/ BATTLETECH]

SR-LP

COLEMAN, LOREN L.  CALL TO ARMS, A [MECHWARRIOR DARK AGE/BATTLETECH]  P
COLEMAN, LOREN L.  CIMMERIAN RAGE [@AGE OF CONAN KERN #2]  P
COLEMAN, LOREN L.  DOUBLE-BLIND [@BATTLETECH 31]  P
COLEMAN, LOREN L.  ENDGAME [@BATTLETECH]  P
COLEMAN, LOREN L.  FLASHPOINT [@BATTLETECH 30]  P
COLEMAN, LOREN L.  FORTRESS REPUBLIC [@MECHWARRIOR DARK AGE/BATTLETECH]  P
COLEMAN, LOREN L.  ILLUSIONS OF VICTORY [@BATTLETECH 47]  P
COLEMAN, LOREN L.  KILLING FIELDS, THE [@BATTLETECH #45/CAPELLAN SOLUTION 2]  P
COLEMAN, LOREN L.  PATRIOTS AND TYRANTS [@BATTLETECH 52]  P
COLEMAN, LOREN L.  SONGS OF VICTORY [@AGE OF CONAN KERN #3]  P
COLEMAN, LOREN L.  STORMS OF FATE [@BATTLETECH]  C/P,P
COLEMAN, LOREN L.  SWORD OF SEDITION [@MECHWARRIOR DARK AGE/BATTLETECH]  P
COLEMAN, LOREN L.  THREADS OF AMBITION [@BATTLETECH #44/CAPELLAN SOLUTION 1]  P
COLEMAN, WIM/PERRIN, PAT  JAMAIS VU PAPERS, THE (PROOF)  SFWA
COLERIDGE, JOHN  MARTIAN MARTYRS  B
COLERIDGE, JOHN  NEW LIFE, THE  B
COLFER, EOIN  ARTEMIS FOWL [ARTEMIS FOWL 1]  C/P,LP
COLFER, EOIN  ARTEMIS FOWL: THE ARCTIC INCIDENT [ARTEMIS FOWL 2]  LP
COLFER, EOIN  ARTEMIS FOWL: THE ETERNITY CODE [ARTEMIS FOWL 3]  H
COLFER, EOIN  ARTEMIS FOWL: THE LOST COLONY [ARTEMIS FOWL 5]  H
COLFER, EOIN  ARTEMIS FOWL: THE OPAL DECEPTION [ARTEMIS FOWL 4]  H
COLLIER BOOKS  MORE SOVIET SCIENCE FICTION  PA,C/PA
COLLIER BOOKS  NEW SOVIET SCIENCE FICTION  HA,LPA
COLLIER BOOKS  SOVIET SCIENCE FICTION  PA,C/PA
COLLIER, JAMES LINCOLN  WHY DOES EVERYBODY THINK I’M NUTTY?  H
COLLIER, JOHN  FANCIES AND GOODNIGHTS  C/P,SR-P,LP,H
COLLIER, JOHN  JOHN COLLIER READER, THE  H:2
COLLIER, MICHAEL  LONGEST DAY [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
COLLIER, MICHAEL  TAINT, THE [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
COLLIGNON, JEFF  HER MONSTER  P
COLLINS, BARNABAS/COLLINS, QUENTIN  DARK SHADOWS BOOK OF VAMPIRES AND WEREWOLVES, THE  PA
COLLINS, CHARLES M.  FRIGHT  PA
COLLINS, ERROL  MARINERS OF SPACE  H
COLLINS, HELEN  MUTAGENESIS  C/P:2,P
COLLINS, HUNT  TOMORROW AND TOMORROW  C/P:2,P
COLLINS, LARRY/LAPIERRE, DOMINIQUE  FIFTH HORSEMAN, THE  C/P,P
COLLINS, MABEL  IDYLL OF THE WHITE LOTUS, THE  P
COLLINS, MAX ALLAN  AFTER THE DARK [@DARK ANGEL]  P
COLLINS, MAX ALLAN  BEFORE THE DAWN [@DARK ANGEL]  C/P,P
COLLINS, MAX ALLAN  SKIN GAME [@DARK ANGEL]  C/P,P
COLLINS, MAX ALLAN  WATERWORLD  P
COLLINS, MICHAEL  LUKAN WAR  C/P,P
COLLINS, MICHAEL  PLANETS OF DEATH, THE  C/P:2,P
COLLINS, MICHAEL (ASTRONAUT)  CARRYING THE FIRE  R/REF-P

COLLINS, MIKE/ EDGINTON, IAN/GREENBARGER, ROBERT/ HAUMAN, GLENN/MARIOTTE, JEFF NO SURRENDER [@STAR TREK S.C.E. #4]

P

COLLINS, NANCY A.  DARKEST HEART [SONJA BLUE]  H
COLLINS, NANCY A.  DEAD MAN’S HAND  LP
COLLINS, NANCY A.  DOZEN BLACK ROSES, A [SONJA BLUE]  LP
COLLINS, NANCY A.  GAHAN WILSON’S THE ULTIMATE HAUNTED HOUSE  LPA
COLLINS, NANCY A.  IN THE BLOOD  P

AUTHOR  TITLE  CODE
COLLINS, NANCY A.  MIDNIGHT BLUE [SONJA BLUE 1,2,3]  H
COLLINS, NANCY A.  SONJA BLUE COLLECTION, THE [SONJA BLUE 1,2,3]  LP
COLLINS, NANCY A.  SUNGLASSES AFTER DARK [SONJA BLUE 1]  C/P,P
COLLINS, NANCY A.  TO FREE ATLANTIS [FANTASTIC FOUR]  P
COLLINS, PAUL  ALIEN WORLDS  SR-HA
COLLINS, PAUL  DISTANT WORLDS  SR-HA
COLLINS, PAUL  EARTHBORN, THE [EARTHBORN WARS 1]  H
COLLINS, PAUL  ENVISAGED WORLDS  SR-HA
COLLINS, PAUL  FRONTIER WORLDS  SR-PA
COLLINS, PAUL  OTHER WORLDS  SR-HA

COLLINS, ROBERT A.

THOMAS BURNETT SWANN:

A BRIEF CRITICAL BIOGRAPHY AND ANNOTATED BIBLIOGRAPHY

R/REF-P

COLLINS, WILKIE  TALES OF TERROR AND THE SUPERNATURAL  LP
COLLODI, CARLO  PINOCCHIO  H
COLOMB, P.  GREAT WAR OF 189–, THE  SR-H
COLOMBO, JOHN ROBERT  YEARS OF LIGHT  R/REF-P
COLVIN, JAMES  DEEP FIX, THE  SR-P
COMBES, SHARON  CALY  P
COMBES, SHARON  CHERRON  P
COMER, RALPH  TO DREAM OF EVIL...  P
COMER, RALPH  WITCHFINDERS, THE  SR-P
COMFORT, ALEX  PROCESS OF AGEING, THE  [C/RAN-P]
COMFORT, ALEX  TETRARCH  H
COMFORT, ALEX  TETRARCH (PROOF)  SR-LP

COMMITTEE ON AERONAUTICAL AND SPACE SCIENCES SOLAR POWER FROM SATELLITES

[C/RAN-H]

COMPTON, D. G.  ASCENDANCIES  H,SFWA,P
COMPTON, D. G.  CHRONOCULES  C/P:2,P
COMPTON, D. G.  CONTINUOUS KATHERINE MORTENHOE, THE  P
COMPTON, D. G.  ELECTRIC CROCODILE, THE  P
COMPTON, D. G.  FAREWELL, EARTH’S BLISS  C/P:2,P
COMPTON, D. G.  MISSIONARIES, THE  C/P:2,P
COMPTON, D. G.  QUALITY OF MERCY, THE  P
COMPTON, D. G.  SILENT MULTITUDE, THE  C/P:2,P
COMPTON, D. G.  STEEL CROCODILE, THE  C/P:2,P
COMPTON, D. G.  SYNTHAJOY  C/P:2,P
COMPTON, D. G.  UNSLEEPING EYE, THE  C/P,P
COMPTON, D. G.  USUAL LUNACY, A  LP
COMPTON, D. G.  WINDOWS  H,SFWA,P
COMSTOCK, JARROD  LOVE MACHINE, THE [@THESE LAWLESS WORLDS #1]  P
COMSTOCK, JARROD  SCALES OF JUSTICE [@THESE LAWLESS WORLDS #2]  P
CONAWAY, J. C.  MAGICIAN’S SLEEVE, THE  P
CONDE, NICHOLAS  RELIGION, THE  P
CONDON, EDWARD U.  SCIENTIFIC STUDY OF UNIDENTIFIED FLYING OBJECTS  [C/RAN-P]
CONDON, RICHARD  MANCHURIAN CANDIDATE, THE  P
CONDON, RICHARD  OLDEST CONFESSION, THE  P
CONE, MOLLY  MISHMASH AND THE ROBOT  P
CONEY, MICHAEL G.  BRONTOMEK!  SR-P,C/P
CONEY, MICHAEL G.  CAT KARINA  C/P:2,P
CONEY, MICHAEL G.  CELESTIAL STEAM LOCOMOTIVE, THE [SONG OF EARTH 1]  H:2,SFWA
CONEY, MICHAEL G.  CHARISMA  SR-P,C/P
CONEY, MICHAEL G.  FANG THE GNOME  H
CONEY, MICHAEL G.  FRIENDS COME IN BOXES  C/P,P,H
CONEY, MICHAEL G.  GIRL WITH A SYMPHONY IN HER FINGERS, THE  H
CONEY, MICHAEL G.  GODS OF THE GREATAWAY [SONG OF EARTH 2]  H:2
CONEY, MICHAEL G.  GODS OF THE GREATAWAY (PROOF) [SONG OF EARTH 2]  SR-LP
CONEY, MICHAEL G.  HELLO SUMMER, GOODBYE  H
CONEY, MICHAEL G.  HERO OF DOWNWAYS, THE  C/P:2,P
CONEY, MICHAEL G.  JAWS THAT BITE, THE CLAWS THAT CATCH, THE  P
CONEY, MICHAEL G.  KING OF THE SCEPTER’D ISLE  H:2,P
CONEY, MICHAEL G.  MIRROR IMAGE  C/P,P
CONEY, MICHAEL G.  MONITOR FOUND IN ORBIT  P
CONEY, MICHAEL G.  NEPTUNE’S CAULDRON  P
CONEY, MICHAEL G.  RAX  P
CONEY, MICHAEL G.  SYZYGY  C/P:2,P
CONGDON, DON/CONGDON, MICHAEL  ALONE BY NIGHT  C/PA,PA
CONGDON, DON  STORIES FOR THE DEAD OF NIGHT  PA
CONGDON, DON  TALES OF LOVE AND HORROR  PA
CONGDON, S. PERRY II  DRAMA READER, THE  LP
CONKLIN, GROFF  ANOTHER PART OF THE GALAXY  PA,C/PA
CONKLIN, GROFF  BEST OF SCIENCE FICTION, THE (1946 EDITION; COMPLETE)  HA
CONKLIN, GROFF  BIG BOOK OF SCIENCE FICTION, THE  PA,C/PA
CONKLIN, GROFF  BRRRR!  PA
CONKLIN, GROFF  CLASSIC BOOK OF SCIENCE FICTION, THE  HA
CONKLIN, GROFF  CROSSROADS IN TIME  C/PA,PA

AUTHOR  TITLE  CODE
CONKLIN, GROFF  DIMENSION 4  PA,C/PA
CONKLIN, GROFF  ELSEWHERE AND ELSEWHEN  PA,C/PA
CONKLIN, GROFF  FIVE-ODD  PA,C/PA
CONKLIN, GROFF  5 UNEARTHLY VISIONS  PA,C/PA
CONKLIN, GROFF  4 FOR THE FUTURE  PA,C/PA
CONKLIN, GROFF  GIANTS UNLEASHED  PA,C/PA
CONKLIN, GROFF  GOLDEN AGE OF SCIENCE FICTION, THE  HA
CONKLIN, GROFF  GRAVEYARD READER, THE  PA,C/PA
CONKLIN, GROFF/FABRICANT, NOAH D.  GREAT SCIENCE FICTION ABOUT DOCTORS  PA,C/PA
CONKLIN, GROFF  GREAT SCIENCE FICTION BY SCIENTISTS  PA,C/PA
CONKLIN, GROFF  GREAT STORIES OF SPACE TRAVEL  PA,C/PA
CONKLIN, GROFF  IN THE GRIP OF TERROR  PA
CONKLIN, GROFF  INVADERS OF EARTH  HA
CONKLIN, GROFF  INVADERS OF EARTH (15 STORIES)  PA,C/PA
CONKLIN, GROFF  INVADERS OF EARTH (17 STORIES)  PA,C/PA
CONKLIN, GROFF  MINDS UNLEASHED  C/PA
CONKLIN, GROFF  OMNIBUS OF SCIENCE FICTION  HA
CONKLIN, GROFF  OPERATION FUTURE  PA,C/PA
CONKLIN, GROFF  POSSIBLE WORLDS OF SCIENCE FICTION  PA,C/PA
CONKLIN, GROFF  SCIENCE FICTION ADVENTURES IN DIMENSION  PA,C/PA
CONKLIN, GROFF  SCIENCE FICTION ADVENTURES IN MUTATION  C/PA,PA
CONKLIN, GROFF  SCIENCE FICTION GALAXY, THE  SR-PA
CONKLIN, GROFF  SCIENCE FICTION ODDITIES  PA,C/PA
CONKLIN, GROFF  SCIENCE FICTION OMNIBUS  SR-PA,C/PA
CONKLIN, GROFF  SCIENCE FICTION TERROR TALES  C/PA,PA
CONKLIN, GROFF  SCIENCE FICTION THINKING MACHINES  PA,C/PA
CONKLIN, GROFF  SEVEN COME INFINITY  PA,C/PA
CONKLIN, GROFF  SEVEN TRIPS THROUGH TIME AND SPACE  PA,C/PA
CONKLIN, GROFF  17 X INFINITY  PA,C/PA
CONKLIN, GROFF  6 GREAT SHORT NOVELS OF SCIENCE FICTION  PA,C/PA
CONKLIN, GROFF  SIX GREAT SHORT SCIENCE FICTION NOVELS  PA,C/PA
CONKLIN, GROFF  SUPERNATURAL READER, THE  HA,PA
CONKLIN, GROFF  13 ABOVE THE NIGHT  PA,C/PA
CONKLIN, GROFF  13 GREAT STORIES OF SCIENCE FICTION  PA,C/PA
CONKLIN, GROFF  TREASURY OF SCIENCE FICTION, A  PA,C/PA
CONKLIN, GROFF  12 GREAT CLASSICS OF SCIENCE FICTION  PA,C/PA
CONKLIN, GROFF  TWISTED  PA,C/PA
CONKLIN, GROFF  WORLDS OF WHEN  PA,C/PA
CONLY, JANE LESLIE  R-T, MARGARET, AND THE RATS OF NIMH [NIMH 3]  H
CONNELL, ADAM  COUNTERFEIT KINGS  LP
CONNER, MICHAEL  ARCHANGEL  P
CONNER, MIGUEL  QUEEN OF DARKNESS, THE  P
CONNER, MIKE  EYE OF THE SUN  P
CONNER, MIKE  GROUPMIND  P
CONNER, MIKE  HOUDINI DIRECTIVE, THE  P
CONNER, MIKE  I AM NOT THE OTHER HOUDINI  H,SFWA,P
CONNERS, AARON  PANDORA DIRECTIVE, THE [@TEX MURPHY]  P
CONNERS, AARON  UNDER A KILLING MOON [TEX MURPHY]  P
CONNOLLY, FLYNN  RISING OF THE MOON, THE  P
CONNOLLY, JOHN  BOOK OF LOST THINGS, THE  H
CONNORS, WILLIAM W./BEBRIS, CARRIE A.  SHADOWBORN [@RAVENLOFT]  P
CONQUEST, ROBERT  WORLD OF DIFFERENCE, A  C/P:2,P
CONRAD, ALLAN  GLAHMIAN SHOCK  P
CONRAD, EARL  DA VINCI MACHINE, THE  C/P,P,H
CONRAD, JOSEPH  TALES OF HEARSAY  H
CONROY, RICK  MISSION FROM MARS  SR-P
CONROY, ROBERT  1901  P
CONSOLMAGNO, GUY  BROTHER ASTRONOMER  R/REF-H
CONSTANTINE, LARRY  INFINITE LOOP  HA
CONSTANTINE, STORM  BEWITCHMENTS OF LOVE AND HATE, THE [WRAETHTHU 2]  P
CONSTANTINE, STORM  BURYING THE SHADOW  LP
CONSTANTINE, STORM  CROWN OF SILENCE, THE [MAGRAVANDIAS 2]  H:2
CONSTANTINE, STORM  ENCHANTMENTS OF FLESH AND SPIRIT, THE [WRAETHTHU 1]  SFWA-P
CONSTANTINE, STORM  FULFILMENTS OF FATE AND DESIRE, THE [WRAETHTHU 3]  P,SFWA-P
CONSTANTINE, STORM  HERMETECH  C/P,P
CONSTANTINE, STORM  SEA DRAGON HEIR [MAGRAVANDIAS 1]  H:2
CONSTANTINE, STORM  WAY OF LIGHT, THE [MAGRAVANDIAS 3]  H:2
CONSTANTINE, STORM  WRAETHTU [WRAETHTHU 1,2,3]  LP
CONSTANTINE, STORM  WRAITHS OF WILL AND PLEASURE, THE [WRAETHTHU HISTORIES 1]  H
CONTENTO, WILLIAM G.  INDEX TO SCIENCE FICTION ANTHOLOGIES AND COLLECTIONS  R/IND
CONTENTO, WILLIAM G.  INDEX TO SCIENCE FICTION ANTHOLOGIES AND COLLECTIONS 1977–1983  R/IND
CONVERTITO, BILL  ROMBELLA SHUTTLE, THE  C/P,P
CONVISIER, JOSH  ECHELON  LP
CONWAY, D. J.  DREAM WARRIOR, THE  LP
CONWAY, GERARD F.  MIDNIGHT DANCERS, THE  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 89
AUTHOR  TITLE  CODE
CONWAY, GERARD F.  MINDSHIP  C/P:2,P
CONWAY, J. C.  QUARREL WITH THE MOON  SFWA-P
CONWAY, NORMAN  OMEGA OPERATION, THE [HUNTER 1]  P
COOK, DAVID  BEYOND THE MOONS [@SPELLJAMMER CLOAKMASTER #1]  C/P:2,P
COOK, DAVID  HORSELORDS [FORGOTTEN REALMS EMPIRES 1]  C/P:2,P
COOK, DAVID  KING PINCH [@FORGOTTEN REALMS NOBLES #1]  P
COOK, DAVID  SOLDIERS OF ICE [@FORGOTTEN REALMS HARPERS #7]  P
COOK, DAWN  DECOY PRINCESS, THE [DECOY PRINCESS 1]  C/P,P
COOK, DAWN  FIRST TRUTH [FIRST TRUTH 1]  P
COOK, DAWN  FORGOTTEN TRUTH [FIRST TRUTH 3]  P
COOK, DAWN  HIDDEN TRUTH [FIRST TRUTH 2]  P
COOK, DAWN  LOST TRUTH [FIRST TRUTH 4]  P
COOK, DAWN  PRINCESS AT SEA [DECOY PRINCESS 2]  C/P,P
COOK, GLEN  ALL DARKNESS MET [DREAD EMPIRE 3/RAGNARSON 5]  C/P:2,P
COOK, GLEN  ANGRY LEAD SKIES [GARRETT 10]  C/P,P
COOK, GLEN  ANNALS OF THE BLACK COMPANY [BLACK COMPANY 1,2,3]  H:2
COOK, GLEN  BITTER GOLD HEARTS [GARRETT 2]  C/P:2,P
COOK, GLEN  BLACK COMPANY, THE [BLACK COMPANY 1]  C/P:2,SFWA-P,P

COOK, GLEN BLACK COMPANY GOES SOUTH, THE [BLACK COMPANY 4,5/

BLACK COMPANY SOUTH 1,2]

H

COOK, GLEN BLACK COMPANY, THE: GLITTERING STONE I [BLACK COMPANY GLITTERING STONE 1,2]

H

COOK, GLEN BLACK COMPANY, THE: GLITTERING STONE II [BLACK COMPANY GLITTERING STONE 3,4]

H

COOK, GLEN BLEAK SEASONS [@BLACK COMPANY GLITTERING STONE 1/

BLACK COMPANY 6]

H,P

COOK, GLEN  CEREMONY [@DARKWAR #3]  P
COOK, GLEN  COLD COPPER TEARS [GARRETT 3]  C/P:2,P
COOK, GLEN  DEADLY QUICKSILVER LIES [GARRETT 7]  P,SFWA-P
COOK, GLEN  DOOMSTALKER [@DARKWAR #1]  P
COOK, GLEN  DRAGON NEVER SLEEPS, THE  C/P:2,P
COOK, GLEN  DREAD BRASS SHADOWS [GARRETT 5]  C/P,P
COOK, GLEN  DREAMS OF STEEL [BLACK COMPANY 5/BLACK COMPANY SOUTH 2]  C/P:2,SFWA-P,P
COOK, GLEN  FADED STEEL HEAT [GARRETT 9]  C/P,P
COOK, GLEN  FIRE IN HIS HANDS, THE [KING WITHOUT A THRONE 1/RAGNARSON 1]  C/P,P
COOK, GLEN  GARRETT FILES, THE [GARRETT 1,2,3]  H
COOK, GLEN  GARRETT INVESTIGATES [GARRETT 7,8,9]  H
COOK, GLEN  GARRETT ON THE CASE [GARRETT 10,11]  H
COOK, GLEN  GARRETT, P.I. [GARRETT 4,5,6]  H
COOK, GLEN  HEIRS OF BABYLON, THE  C/P:2,P
COOK, GLEN  ILL FATE MARSHALLING, AN [DREAD EMPIRE 4/RAGNARSON 6]  C/P,SFWA-P,P

COOK, GLEN LORD OF THE SILENT KINGDOM

[@INSTRUMENTALITIES OF THE NIGHT #2]

H

COOK, GLEN  MATTER OF TIME, A  C/P,P
COOK, GLEN  OCTOBER’S BABY [DREAD EMPIRE 2/RAGNARSON 4]  C/P:2,P
COOK, GLEN  OLD TIN SORROWS [GARRETT 4]  C/P,P
COOK, GLEN  PASSAGE AT ARMS [STARFISHERS]  H,P
COOK, GLEN  PETTY PEWTER GODS [GARRETT 8]  C/P:2,P
COOK, GLEN  REAP THE EAST WIND [DREAD EMPIRE 5/RAGNARSON 7]  C/P:2,SFWA-P,P
COOK, GLEN  RED IRON NIGHTS [GARRETT 6]  P,SFWA-P:2
COOK, GLEN  SHADOW GAMES [BLACK COMPANY 4/BLACK COMPANY SOUTH 1]  C/P,SFWA-P,P
COOK, GLEN  SHADOW OF ALL NIGHT FALLING, A [DREAD EMPIRE 1/RAGNARSON 3]  C/P:2,P
COOK, GLEN  SHADOWLINE [STARFISHERS 1]  P
COOK, GLEN  SHADOWS LINGER [BLACK COMPANY 2]  SR-P,C/P:2

COOK, GLEN SHE IS THE DARKNESS [@BLACK COMPANY GLITTERING STONE 2/

BLACK COMPANY 7]

H,P

COOK, GLEN  SILVER SPIKE, THE [BLACK COMPANY]  C/P:2,SFWA-P,P

COOK, GLEN SOLDIERS LIVE [@BLACK COMPANY GLITTERING STONE 4/

BLACK COMPANY 9]

H:2

COOK, GLEN  STARFISHERS [STARFISHERS 2]  C/P,P
COOK, GLEN  STARS’ END [STARFISHERS 3]  C/P,P
COOK, GLEN  SUNG IN BLOOD  SR-H,H
COOK, GLEN  SWEET SILVER BLUES [GARRETT 1]  C/P:2,P
COOK, GLEN  SWORDBEARER, THE  C/P:2,SFWA-P,P
COOK, GLEN  TOWER OF FEAR, THE  C/P,SFWA-P,P,H
COOK, GLEN  TYRANNY OF THE NIGHT, THE [@INSTRUMENTALITIES OF THE NIGHT #1]  H
COOK, GLEN  WARLOCK [@DARKWAR #2]  C/P,P

COOK, GLEN WATER SLEEPS [@BLACK COMPANY GLITTERING STONE 3/

BLACK COMPANY 8]

H

COOK, GLEN  WHISPERING NICKEL IDOLS [GARRETT 11]  P
COOK, GLEN  WHITE ROSE, THE [BLACK COMPANY 3]  C/P:2,SFWA-P,P

COOK, GLEN WITH MERCY TOWARD NONE [KING WITHOUT A THRONE 2/

RAGNARSON 2] C/P:2,P

AUTHOR  TITLE  CODE
COOK, HUGH  HERO’S RETURN, THE [@WIZARD WAR #3]  C/P:2,P
COOK, HUGH  LORDS OF THE SWORD [@WIZARD WAR]  P,SFWA-P:2
COOK, HUGH  ORACLE, THE [@WIZARD WAR #4]  P
COOK, HUGH  QUESTING HERO, THE [@WIZARD WAR #2]  C/P:2,P
COOK, HUGH  WIZARD WAR [@WIZARD WAR #1]  C/P:2,P
COOK, LEONARD R.  VOYAGE TO DESTINY  H
COOK, MONTE  GLASS PRISON, THE [@FORGOTTEN REALMS]  P
COOK, MONTE  OF AGED ANGELS [DARK MATTER 4]  P
COOK, PAUL  ALEJANDRA VARIATIONS, THE  P
COOK, PAUL  DUENDE MEADOW  P
COOK, PAUL  ENGINES OF DAWN, THE  C/P,P
COOK, PAUL  FORTRESS ON THE SUN  C/P,P
COOK, PAUL  HALO  P
COOK, PAUL  ON THE RIM OF THE MANDALA  P
COOK, PAUL  TINTAGEL  C/P,P
COOK, RICK  LIMBO SYSTEM  C/P:2,SFWA-P,P
COOK, RICK  MALL PURCHASE NIGHT  C/P:2,SFWA-P,P
COOK, RICK  WIZ BIZ, THE [WIZ ZUMWALT 1,2]  C/P,P
COOK, RICK  WIZARDRY COMPILED, THE [WIZ ZUMWALT 2]  C/P:2,SFWA-P:2,P
COOK, RICK  WIZARDRY CONSULTED, THE [WIZ ZUMWALT 4]  C/P:2,P
COOK, RICK  WIZARDRY CURSED, THE [WIZ ZUMWALT 3]  C/P:2,SFWA-P,P
COOK, RICK  WIZARDRY QUESTED, THE [WIZ ZUMWALT 5]  P
COOK, RICK  WIZARD’S BANE [WIZ ZUMWALT 1]  C/P:2,SFWA-P,P
COOK, ROBIN  BRAIN  H,P
COOK, ROBIN  CHROMOSOME 6  P
COOK, ROBIN  COMA  C/P:2,P
COOK, ROBIN  FEVER  C/P,P
COOK, ROBIN  GODPLAYER  P
COOK, ROBIN  INVASION  P
COOK, ROBIN  MINDBEND  P
COOK, ROBIN  MORTAL FEAR  C/P,P
COOK, ROBIN  MUTATION  H,P
COOK, ROBIN  OUTBREAK  H,P
COOK, SUE WEINLEIN  CHILDREN OF THE RUNE [MONTE COOK’S ARCANA EVOLVED]  C/PA,PA
COOK, SUE WEINLEIN  DRAGON’S RETURN, THE [MONTE COOK’S ARCANA EVOLVED]  PA
COOK, WALLACE WILLIAM  ADRIFT IN THE UNKNOWN  SR-P
COOK, WILLIAM W.  ROUND TRIP TO THE YEAR 2000, A  LP
COOKE, CATHERINE  CRIMSON GODDESS, THE [ARRIS 3]  C/P:2,P
COOKE, CATHERINE  HIDDEN TEMPLE, THE [ELVEN KINGDOMS 3]  C/P,SFWA-P,P
COOKE, CATHERINE  MASK OF THE WIZARD [ELVEN KINGDOMS 1]  C/P,SFWA-P,P
COOKE, CATHERINE  REALM OF THE GODS [ARRIS 2]  P
COOKE, CATHERINE  VEIL OF SHADOW [ELVEN KINGDOMS 2]  P,SFWA-P
COOKE, CATHERINE  WINGED ASSASSIN, THE [ARRIS 1]  P
COOKE, JOHN PAYTON  LAKE, THE  P
COOL, TOM  INFECTRESS  P
COOL, TOM  SECRET REALMS  H,P
COOMBS, CHARLES  MYSTERY OF SATELLITE 7  C/P,P
COON, GENE L.  DEVIL IN THE DARK, THE [STAR TREK]  SR-S
COON, GENE L.  ERRAND OF MERCY [STAR TREK]  SR-S
COON, GENE L./RODDENBERRY, GENE  SPACE SEED [STAR TREK]  SR-S
COON, HORACE  43,000 YEARS LATER  C/P:2,P
COON, SUSAN  CASSILEE [LIVING PLANET 2]  C/P,P
COON, SUSAN  CHIY-UNE [LIVING PLANET 4]  P
COON, SUSAN  RAHNE [LIVING PLANET 1]  P
COON, SUSAN  VIRGIN, THE [LIVING PLANET 3]  C/P,P
COONEY, CAROLINE B.  OUT OF TIME  P
COONTS, STEPHEN  SAUCER  P
COONTZ, OTTO  ISLE OF THE SHAPESHIFTERS  P
COOPER, BRENDA  SILVER SHIP AND THE SEA, THE  H
COOPER, C. EVERETT  UP YOUR ASTEROID!  SR-LP
COOPER, EDMUND  ALL FOOLS’ DAY  P
COOPER, EDMUND  CLOUD WALKER, THE  C/P,P
COOPER, EDMUND  DEADLY IMAGE  C/P:2,P
COOPER, EDMUND  FAR SUNSET, A  C/P:2,P
COOPER, EDMUND  FIVE TO TWELVE  H:2,P
COOPER, EDMUND  GENDER GENOCIDE  P
COOPER, EDMUND  KRONK  P
COOPER, EDMUND  LAST CONTINENT, THE  C/P,P
COOPER, EDMUND  NEWS FROM ELSEWHERE  C/P:2,P
COOPER, EDMUND  OVERMAN CULTURE, THE  H:2,P
COOPER, EDMUND  PRISONER OF FIRE  SR-P
COOPER, EDMUND  SEA HORSE IN THE SKY  H:2,P
COOPER, EDMUND  SEED OF LIGHT  C/P:3,P
COOPER, EDMUND  SLAVES OF HEAVEN, THE  H:2,SFWA,P

AUTHOR  TITLE  CODE
COOPER, EDMUND  TENTH PLANET, THE  C/P,P,H
COOPER, EDMUND  TOMORROW CAME  P
COOPER, EDMUND  TOMORROW’S GIFT  C/P:2,P
COOPER, EDMUND  TRANSIT  C/P:2,P
COOPER, HENRY S. F. JR.  SEARCH FOR LIFE ON MARS, THE  [C/RAN-H]
COOPER, HUGHES  SEXMAX  P
COOPER, JEFFREY  NIGHTMARES ON ELM STREET, THE [ELM STREET 1,2,3]  P
COOPER, LOUISE  AISLING [INDIGO 8]  C/P:2,P
COOPER, LOUISE  AVATAR [@INDIGO #6]  C/P:2,P
COOPER, LOUISE  AVENGER, THE [CHAOS GATE 3]  C/P:2,P
COOPER, LOUISE  BOOK OF PARADOX, THE  C/P,P
COOPER, LOUISE  CROWN OF HORN  P
COOPER, LOUISE  DECEIVER, THE [CHAOS GATE 1]  C/P,P
COOPER, LOUISE  HEART OF GLASS [DARK ENCHANTMENT]  P
COOPER, LOUISE  IN MEMORY OF SARAH BAILEY  SR-P
COOPER, LOUISE  INFANTA [@INDIGO #3]  C/P,SFWA-P,P
COOPER, LOUISE  INFERNO [@INDIGO #2]  C/P:2,SFWA-P,P
COOPER, LOUISE  INITIATE, THE [TIME MASTER 1]  C/P:2,SFWA-P,P
COOPER, LOUISE  MASTER, THE [TIME MASTER 3]  C/P:2,SFWA-P,P
COOPER, LOUISE  MIRAGE  C/P:2,SFWA-P,P
COOPER, LOUISE  NEMESIS [@INDIGO #1]  C/P,SFWA-P,P
COOPER, LOUISE  NOCTURNE [@INDIGO #4]  C/P,SFWA-P,P
COOPER, LOUISE  OUTCAST, THE [TIME MASTER 2]  C/P:2,SFWA-P,P
COOPER, LOUISE  PRETENDER, THE [CHAOS GATE 2]  P
COOPER, LOUISE  REVENANT [@INDIGO #7]  C/P,P
COOPER, LOUISE  RIP TIDE [DOCTOR WHO]  SR-H
COOPER, LOUISE  SACRAMENT OF NIGHT  H
COOPER, LOUISE  SLEEP OF STONE, THE  C/P,SFWA,P
COOPER, LOUISE  STAR ASCENDANT [TIME MASTER 0]  C/P,P,H
COOPER, LOUISE  TROIKA [@INDIGO #5]  C/P:2,P
COOPER, MORTON  MUNSTERS, THE  P
COOPER, PARLEY J.  FEMINISTS, THE  P
COOPER, PARLEY J.  MY LADY EVIL  P
COOPER, PARLEY J.  REVEREND MAMA  P
COOPER, SONNI  BLACK FIRE [STAR TREK]  C/P:3,P
COOPER, SUSAN  BOGGART, THE  SFWA
COOPER, SUSAN  DARK IS RISING, THE [DARK IS RISING 2]  C/P,LP,H:2
COOPER, SUSAN  GREENWITCH [DARK IS RISING 3]  H:3,LP
COOPER, SUSAN  GREY KING, THE [DARK IS RISING 4]  C/P,LP,H:2
COOPER, SUSAN  MANDRAKE  SR-P,C/P
COOPER, SUSAN  MATTHEW’S DRAGON  SFWA
COOPER, SUSAN  OVER SEA, UNDER STONE [DARK IS RISING 1]  C/P,SR-P,H:2
COOPER, SUSAN  SEAWARD  C/P,SFWA,P
COOPER, SUSAN  SILVER ON THE TREE [DARK IS RISING 5]  C/P,P,H:2
COOVER, ROBERT  UNIVERSAL BASEBALL ASSOCIATION, INC.: J. HENRY WAUGH, PROP., THE  C/P,P
COPPEL, ALFRED  BURNING MOUNTAIN, THE  H
COPPEL, ALFRED  DARK DECEMBER  C/P:2,P
COPPEL, ALFRED  GLORY [GOLDENWING 1]  H
COPPEL, ALFRED  GLORY’S PEOPLE [GOLDENWING 3]  H,P
COPPEL, ALFRED  GLORY’S WAR [GOLDENWING 2]  P
COPPER, BASIL  AND AFTERWARD, THE DARK  H
COPPER, BASIL  GREAT WHITE SPACE, THE  C/P,P
COPPER, BASIL  HERE BE DAEMONS  H
COPPER, BASIL  NECROPOLIS  H
COPPER, BASIL  PHANTOM AND THE SCORPIA MENACE, THE [PHANTOM 3]  P
COPPER, BASIL  VOICES OF DOOM  H
COPPER, BASIL  WEREWOLF IN LEGEND, FACT, AND ART, THE  R/REF-H
COPPERFIELD, DAVID  DAVID COPPERFIELD’S BEYOND IMAGINATION  PA
COPPERFIELD, DAVID/BERLINER, JANET  DAVID COPPERFIELD’S TALES OF THE IMPOSSIBLE  HA,PA
CORBEN, RICHARD  NEVERWHERE  CX
CORBETT, SCOTT  HAIRY HORROR TRICK, THE  P
CORBY, ADAM  DIVINE QUEEN, THE [DOOM-QUEST OF ARA-KARN 2]  P
CORBY, ADAM  FORMER KING, THE [DOOM-QUEST OF ARA-KARN 1]  C/P,P
CORDELL, ALEXANDER  IF YOU BELIEVE THE SOLDIERS  H,SFWA
CORDELL, BRUCE R.  DARKVISION [@FORGOTTEN REALMS WIZARDS]  P
CORDELL, BRUCE R.  LADY OF POISON [@FORGOTTEN REALMS]  P
CORELLI, MARIE  MIGHTY ATOM, THE  P
CORELLI, MARIE  ROMANCE OF TWO WORLDS, A  H
CORELLI, MARIE  SORROWS OF SATAN, THE  H
COREY, PAUL  PLANET OF THE BLIND, THE  C/P,P
CORLETT, WILLIAM  STEPS UP THE CHIMNEY, THE [MAGICIAN’S HOUSE 1]  P
CORLEY, DONALD  HAUNTED JESTER, THE  H
CORLEY, EDWIN  JESUS FACTOR, THE  P
CORLEY, EDWIN  SARGASSO  C/P,P

AUTHOR  TITLE  CODE
CORLEY, EDWIN  SIEGE  P
CORMAN, AVERY  OH, GOD!  C/P,P
CORMIER, ROBERT  FADE  H,P
CORNELL, PAUL  BRITISH SUMMERTIME  LP
CORNELL, PAUL  GOTH OPERA [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
CORNELL, PAUL  HAPPY ENDINGS [DOCTOR WHO NEW ADVENTURES]  P
CORNELL, PAUL  HUMAN NATURE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CORNELL, PAUL  LICENCE DENIED [DOCTOR WHO]  R/REF-P
CORNELL, PAUL  LIFE DURING WARTIME [BERNICE SUMMERFIELD]  HA
CORNELL, PAUL  LIFE OF SURPRISES, A [BERNICE SUMMERFIELD]  HA
CORNELL, PAUL  LOVE AND WAR [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CORNELL, PAUL  NO FUTURE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
CORNELL, PAUL  OH NO IT ISN’T! [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P

CORNELL, PAUL PROFESOR BERNICE SUMMERFIELD AND THE DEAD MEN DIARIES

[BERNICE SUMMERFIELD]

PA

CORNELL, PAUL  SCREAM OF THE SHALKA [@DOCTOR WHO NINTH (ANIMATED) DOCTOR]  P
CORNELL, PAUL  SHADOWS OF AVALON, THE [DOCTOR WHO]  P
CORNELL, PAUL  SHORT TRIPS: A CHRISTMAS TREASURY [DOCTOR WHO]  HA
CORNELL, PAUL  SOMETHING MORE  LP
CORNELL, PAUL  TIMEWYRM: REVELATION [DOCTOR WHO NEW ADVENTURES]  C/P:2,P

CORNELL, PAUL/DAY, MARTIN/ TOPPING, KEITH X-TREME POSSIBILITIES [X-FILES]

R/REF-P

CORNISH, EDWARD  1999: THE WORLD OF TOMORROW  [C/RAN-H]
CORNMARKET REPRINTS  BATTLE OF DORKING CONTROVERSY, THE  HA
CORNWELL, SUE/KOTT, MIKE  OFFICIAL PRICE GUIDE: STAR TREK AND STAR WARS COLLECTIBLES, THE R/REF-P
CORRAN, MARY  DARKFELL  P
CORREY, LEE  ABODE OF LIFE, THE [STAR TREK]  C/P:3,P
CORREY, LEE  CONTRABAND ROCKET  DD146
CORREY, LEE  MANNA  C/P:2,P
CORREY, LEE  MATTER OF METALAW, A  C/P,P
CORREY, LEE  SHUTTLE DOWN  C/P:2,P
CORREY, LEE  SPACE DOCTOR  C/P:2,P
CORREY, LEE  STAR DRIVER  C/P:2,P
CORWIN  FELINE FETISHES  SR-L
CORWIN, NORMAN  DOG IN THE SKY  H
CORY, DAVID  ICEBERG EXPRESS, THE  H
CORY, HOWARD L.  MIND MONSTERS, THE  DG602
CORY, HOWARD L.  SWORD OF LANKOR, THE  C/P:2,P
COSTELLO, MATTHEW J.  FATE’S TRICK  P,SFWA-P
COSTELLO, MATTHEW J.  FIRE BELOW [@SEAQUEST DSV 2]  P
COSTELLO, MATTHEW J.  HOUR OF THE SCORPION [@TIME WARRIOR #2]  P
COSTELLO, MATTHEW J.  REVOLT ON MAJIPOOR  P,SFWA-P
COSTELLO, MATTHEW J.  TIME OF THE FOX [TIME WARRIOR 1]  C/P,P
COSTELLO, MATTHEW J.  WIZARD OF TIZARE [GUARDIANS OF THE THREE 3]  C/P:2,P
COSTELLO, PETER  JULES VERNE: INVENTOR OF SCIENCE FICTION  SFWA
COSTIKYAN, GREG  ANOTHER DAY, ANOTHER DUNGEON [CUPS AND SORCERY 1]  P,SFWA-P
COSTIKYAN, GREG  BY THE SWORD [MAGIC OF THE PLAINS #1]  C/P,@H:2,P
COSTIKYAN, GREG  FIRST CONTRACT  H
COSTIKYAN, GREG  ONE QUEST, HOLD THE DRAGONS [CUPS AND SORCERY 2]  P
COTTINGTON, LADY  LADY COTTINGTON’S PRESSED FAIRY CALENDAR 1997  CAL
COTTON, DONALD  GUNFIGHTERS, THE [DOCTOR WHO]  SR-P
COTTON, DONALD  MYTH MAKERS, THE [DOCTOR WHO]  C/P,P
COULSON, JUANITA  CRISIS ON CHEIRON  DH027
COULSON, JUANITA  DARK PRIESTESS  P
COULSON, JUANITA  DEATH GOD’S CITADEL, THE  C/P:2,P
COULSON, JUANITA  LEGACY OF EARTH [@CHILDREN OF THE STARS #3]  C/P,P
COULSON, JUANITA  OUTWARD BOUND [@CHILDREN OF THE STARS #2]  C/P,P,H
COULSON, JUANITA  PAST OF FOREVER, THE [@CHILDREN OF THE STARS #4]  C/P,P
COULSON, JUANITA  SINGING STONES, THE  DH077
COULSON, JUANITA  SPACE TRAP  SR-P
COULSON, JUANITA  STAR SISTER  C/P,P
COULSON, JUANITA  TOMORROW’S HERITAGE [@CHILDREN OF THE STARS #1]  C/P,P,H
COULSON, JUANITA  UNTO THE LAST GENERATION  SR-P
COULSON, JUANITA  WEB OF WIZARDRY, THE  C/P:2,P
COULSON, ROBERT/DE WEESE, GENE  GATES OF THE UNIVERSE  SR-P,C/P:2
COULSON, ROBERT  TO RENEW THE AGES  SR-P
COUNSELMAN, MARY ELIZABETH  HALF IN SHADOW  SR-H
COUNTER, BEN  BLEEDING CHALICE, THE [@WARHAMMER 40K SOUL DRINKERS 2]  P
COUNTER, BEN  CHAPTER WAR [@WARHAMMER 40K]  P
COUNTER, BEN  CRIMSON TEARS [@WARHAMMER 40K SOUL DRINKERS 3]  P
COUNTER, BEN  DAEMON WORLD [@WARHAMMER 40K]  P
COUNTER, BEN  DARK ADEPTUS [@WARHAMMER 40K GREY KNIGHTS 2]  P
COUNTER, BEN  GALAXY IN FLAMES [@WARHAMMER 40K HORUS HERESY 3]  P
COUNTER, BEN  GREY KNIGHTS [@WARHAMMER 40K GREY KNIGHTS 1]  P

AUTHOR  TITLE  CODE
COUNTER, BEN  SOUL DRINKER [@WARHAMMER 40K SOUL DRINKERS 1]  P
COUPER, STEPHEN  DYING OF PARADISE  SR-P
COUPPEY, MADELEINE  RUMOUR IN THE FOREST, THE  H
COURT, NATHAN  MATHEMATICS IN FUN AND IN EARNEST  [C/RAN-P]
COURTNEY, VINCENT  HARVEST OF BLOOD  P
COVER, ARTHUR BYRON  AUTUMN ANGELS  C/P,P
COVER, ARTHUR BYRON  BORN IN FIRE [RISING STARS 1]  P
COVER, ARTHUR BYRON  EAST WIND COMING, AN  C/P,P
COVER, ARTHUR BYRON  FLASH GORDON  P
COVER, ARTHUR BYRON  NIGHT OF THE LIVING RERUN [BUFFY]  P
COVER, ARTHUR BYRON  PLANETFALL [PLANETFALL 1]  C/P:2,P
COVER, ARTHUR BYRON  PLATYPUS OF DOOM AND OTHER NIHILISTS, THE  P
COVER, ARTHUR BYRON  PRODIGY [ROBOT CITY 4]  C/P:2,SFWA-P,P
COVER, ARTHUR BYRON  RED STAR, THE  LP
COVER, ARTHUR BYRON  SOUND OF WINTER, THE  P
COVER, ARTHUR BYRON  STATIONFALL [PLANETFALL 2]  C/P,P
COVILLE, BRUCE  ALIENS ATE MY HOMEWORK [ROD ALLBRIGHT 1]  P
COVILLE, BRUCE  BRUCE COVILLE’S BOOK OF ALIENS  LPA
COVILLE, BRUCE  BRUCE COVILLE’S BOOK OF MAGIC  LPA
COVILLE, BRUCE  DARK ABYSS, THE [DUNGEON 2]  C/P:2,P
COVILLE, BRUCE  DINOSAUR THAT FOLLOWED ME HOME, THE [CAMP HAUNTED HILL 3]  LP
COVILLE, BRUCE  DRAGONSLAYERS, THE  LP
COVILLE, BRUCE  EYES OF THE TAROT [DARK FORCES 9]  P
COVILLE, BRUCE  I LEFT MY SNEAKERS IN DIMENSION X [ROD ALLBRIGHT 2]  LP
COVILLE, BRUCE  INTO THE LAND OF THE UNICORNS  LP
COVILLE, BRUCE  JENNIFER MURDLEY’S TOAD [MAGIC SHOP 1]  LP
COVILLE, BRUCE  JEREMY THATCHER, DRAGON HATCHER [MAGIC SHOP 2]  LP
COVILLE, BRUCE  MONSTER’S RING, THE [MAGIC SHOP 3]  C/P,LP
COVILLE, BRUCE  MY TEACHER FLUNKED THE PLANET [MY TEACHERS 4]  LP
COVILLE, BRUCE  MY TEACHER FRIED MY BRAINS [MY TEACHERS 2]  P
COVILLE, BRUCE  MY TEACHER GLOWS IN THE DARK [MY TEACHERS 3]  P
COVILLE, BRUCE  MY TEACHER IS AN ALIEN [MY TEACHERS 1]  P
COVILLE, BRUCE  SPACE STATION ICE-3  C/P,P
COWAN, DALE  DEADLY SLEEP [TWILIGHT 1]  P
COWE-SPIGAI, HERETH/NEIGHLY, PATRICK  LIFEDEATH [DOCTOR WHO]  SR-PA
COWLEY, ROBERT  WHAT IF? 2  HA
COWLEY, STEWART/HERRIDGE, CHARLES  GREAT SPACE BATTLES  A
COWLEY, STEWART  SPACECRAFT 2000 TO 2100 AD  A
COWLEY, STEWART  SPACEWRECK  A
COWPER, RICHARD  BREAKTHROUGH  C/P:2,P
COWPER, RICHARD  CLONE  H:2,P,SFWA
COWPER, RICHARD  DREAM OF KINSHIP, A [BIRD OF KINSHIP 2]  C/P,P
COWPER, RICHARD  KULDESAK  SR-P,H,SFWA
COWPER, RICHARD  OUT THERE WHERE THE BIG SHIPS GO  C/P,P
COWPER, RICHARD  PHOENIX  H:2,P
COWPER, RICHARD  PROFUNDIS  P
COWPER, RICHARD  ROAD TO CORLAY, THE [BIRD OF KINSHIP 1]  C/P,P,H
COWPER, RICHARD  TAPESTRY OF TIME, A [BIRD OF KINSHIP 3]  P
COWPER, RICHARD  TIME OUT OF MIND  SR-P
COWPER, RICHARD  TWILIGHT OF BRIAREUS, THE  H:2,P
COX, DONALD/STOIKO, MICHAEL  SPACEPOWER  [C/RAN-H]
COX, ERLE  OUT OF THE SILENCE  H
COX, F. BRET/DUNCAN, ANDY  CROSSROADS  HA
COX, GREG  ASSIGNMENT: ETERNITY [STAR TREK]  C/P,P
COX, GREG  BATTLE ON! [XENA]  R/REF-P
COX, GREG  BLACK SHORE, THE [STAR TREK VOYAGER]  C/P:2,P
COX, GREG  BLOOD ENEMY [@UNDERWORLD 0]  P
COX, GREG  DAREDEVIL  H
COX, GREG/BETANCOURT, JOHN GREGORY  DEVIL IN THE SKY [STAR TREK DS9]  C/P:2,P
COX, GREG  EUGENICS WARS, VOLUME 1 [STAR TREK]  P
COX, GREG  EUGENICS WARS, VOLUME 2 [STAR TREK]  H,P
COX, GREG  IRON MAN: THE ARMOR TRAP  C/P,P
COX, GREG  LOST AND FOUND [@GAMMA QUEST #1]  P
COX, GREG  Q-SPACE [STAR TREK TNG]  C/P,P
COX, GREG  Q-STRIKE [@STAR TREK TNG]  C/P,P
COX, GREG  Q-ZONE [STAR TREK TNG]  C/P:2,P
COX, GREG  TO REIGN IN HELL [STAR TREK]  P
COX, GREG/WEISSKOPF, T. K. F.  TOMORROW BITES  C/PA,PA
COX, GREG/WEISSKOPF, T. K. F.  TOMORROW SUCKS  PA
COX, GREG  UNDERWORLD [@UNDERWORLD 1]  P
COX, GREG  UNDERWORLD EVOLUTION [@UNDERWORLD 2]  P
COX, IRVING JR.  OF SUCH AS THESE  B
COX, JOAN  MINDSONG  C/P:2,P
COX, JOAN  STAR WEB  C/P,P

AUTHOR  TITLE  CODE
COYNE, JAMES P.  STRIKE EAGLES  LP
COYNE, JOHN  FURY  SFWA
COYNE, JOHN  HOBGOBLIN  H,P
COYNE, JOHN  PIERCING, THE  P
COYNE, JOHN  SEARING, THE  H
COYNE, JOHN  SHROUD, THE  P
CRADDOCK, WILLIAM J.  TWILIGHT CANDELABRA  LP
CRADOCK, PHYLLIS  GATEWAY TO REMEMBRANCE  H
CRAIG, BRIAN  PAWNS OF CHAOS [WARHAMMER 40K]  P
CRAIG, BRIAN  PLAGUE DAEMON [WARHAMMER ORFEO 2]  LP
CRAIG, BRIAN  STORM WARRIORS [WARHAMMER ORFEO 3]  H,P
CRAIG, BRIAN  WINE OF DREAMS, THE [WARHAMMER]  P
CRAIG, BRIAN  ZARAGOZ [WARHAMMER ORFEO 1]  LP
CRAIG, WILLIAM  TASHKENT CRISIS, THE  P
CRAIGIE, DAVID  DARK ATLANTIS  H
CRAIGIE, DAVID  VOYAGE OF THE LUNA 1, THE  H:2
CRAMER, JOHN  EINSTEIN’S BRIDGE  C/P,LP
CRAMER, JOHN  TWISTOR  C/P,P,H
CRAMER, KATHRYN/PAUTZ, PETER D.  ARCHITECTURE OF FEAR, THE  SFWA
CRAMER, KATHRYN/HARTWELL, DAVID G.  CHRISTMAS GHOSTS  SFWA
CRANDALL, MELISSA  EARTH 2 [EARTH 2]  C/P,P
CRANDALL, MELISSA  SEARCH AND RESCUE [QUANTUM LEAP]  C/P,P
CRANDALL, MELISSA  SHELL GAME [STAR TREK]  C/P:2,P
CRANE, NATHALIA  SUNKEN GARDEN, THE  H
CRANE, ROBERT  HERO’S WALK  C/P:2,P
CRAVENS, GWYNETH/MARR, JOHN S.  BLACK DEATH, THE  C/P,P
CRAWFORD, BETTY ANNE  BUSHIDO INCIDENT, THE  C/P:2,P
CRAWFORD, DAN  ROUSE A SLEEPING CAT [NIMNESTL 1]  P
CRAWFORD, DAN  SURE DEATH OF A MOUSE, THE [NIMNESTL 2]  C/P,P
CRAWFORD, DAN  WILD DOG AND LONE, A [NIMNESTL 3]  P
CRAWFORD, F. MARION  KHALED  C/P:2,P
CRAWFORD, F. MARION  ZOROASTER/KHALED  H
CRAWFORD PUBLICATIONS  GARDEN OF FEAR, THE  SR-PA
CRAWLEY, RAYBURN  CHATTERING GODS  H
CRAWSHAY-WILLIAMS, ELIOT  MAN WHO MET HIMSELF AND OTHER STORIES, THE  H
CRAY, ROBERTA  SWORD AND THE LION, THE  P,SFWA-P:2
CREAN, MICHELLE SHIREY  DANCER OF THE SIXTH  P
CREASEY, JOHN  CHILDREN OF HATE, THE  H
CREASEY, JOHN  DEPTHS, THE  P
CREASEY, JOHN  FAMINE, THE  H
CREASEY, JOHN  FLOOD, THE  P
CREASEY, JOHN  INFERNO, THE  P
CREASEY, JOHN  INSULATORS, THE  P
CREASEY, JOHN  PLAGUE OF SILENCE, THE  P
CREASEY, JOHN  PROPHET OF FIRE, THE  H
CREASEY, JOHN  SLEEP, THE  P
CREASEY, JOHN  SMOG, THE  P
CREASEY, JOHN  TERROR, THE  C/P,P
CREASEY, JOHN  TOUCH OF DEATH, THE  SR-P
CREASEY, JOHN  UNBEGOTTEN, THE  P
CREIGHTON, LEE  TWO QUEENS OF LOCHRIN  P
CRENSHAW, NADINE  BALOR OF THE EVIL EYE  P
CRICHTON, MICHAEL  AIRFRAME  H:2
CRICHTON, MICHAEL  ANDROMEDA STRAIN, THE  C/P,P,H
CRICHTON, MICHAEL  CONGO  C/P,P,H
CRICHTON, MICHAEL  DISCLOSURE  H:2
CRICHTON, MICHAEL  EATERS OF THE DEAD  C/P:2,P
CRICHTON, MICHAEL  JURASSIC PARK [JURASSIC PARK 1]  H:2,P
CRICHTON, MICHAEL  LOST WORLD, THE [JURASSIC PARK 2]  H:2,P
CRICHTON, MICHAEL  PREY  H:2
CRICHTON, MICHAEL  SPHERE  H:2,P
CRICHTON, MICHAEL  TERMINAL MAN, THE  C/P,P,H
CRICHTON, MICHAEL  TIMELINE  H:2,P
CRICHTON, MICHAEL  WESTWORLD  C/P:2,P
CRILLEY, PAUL  NIGHT OF THE LONG SHADOWS [@EBERRON INQUISITIVES]  P
CRISPIN, A. C./O’MALLEY, KATHLEEN  ALIEN RESURRECTION [ALIENS]  P
CRISPIN, A. C./KING, T. JACKSON  ANCESTOR’S WORLD [STARBRIDGE 6]  P
CRISPIN, A. C./MARSHALL, DEBORAH A.  DEATH TIDE [V]  C/P,P
CRISPIN, A. C.  EYES OF THE BEHOLDERS, THE [STAR TREK TNG]  C/P:2,P
CRISPIN, A. C.  HUTT GAMBIT, THE [@STAR WARS HAN SOLO 2]  P
CRISPIN, A. C.  PARADISE SNARE, THE [@STAR WARS HAN SOLO 1]  C/P,P
CRISPIN, A. C.  REBEL DAWN [@STAR WARS HAN SOLO 3]  P
CRISPIN, A. C.  SAREK [STAR TREK]  H:2,P
CRISPIN, A. C./MARSHALL, DEBORAH A.  SERPENT’S GIFT [@STARBRIDGE #4]  C/P,SFWA-P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 95
AUTHOR  TITLE  CODE
CRISPIN, A. C./ELLIOTT, JANNEAN  SHADOW WORLD [@STARBRIDGE #3]  C/P:2,SFWA-P:2,P
CRISPIN, A. C./O’MALLEY, KATHLEEN  SILENT DANCES [@STARBRIDGE #2]  C/P:2,SFWA-P:2,P
CRISPIN, A. C./O’MALLEY, KATHLEEN  SILENT SONGS [@STARBRIDGE #5]  P,SFWA-P
CRISPIN, A. C.  STARBRIDGE [@STARBRIDGE #1]  C/P:2,SFWA-P:2,P
CRISPIN, A. C.  STORMS OF DESTINY [EXILES OF BOQ’URAIN 1]  P
CRISPIN, A. C.  TIME FOR YESTERDAY [STAR TREK]  C/P:2,P
CRISPIN, A. C.  V [V]  C/P:2,P
CRISPIN, A. C./EMERSON, RU  VOICES OF CHAOS [STARBRIDGE 7]  P
CRISPIN, A. C.  YESTERDAY’S SON [STAR TREK]  C/P:2,P
CRISPIN, EDMUND  BEST SF THREE  SR-LPA
CRISTABEL  CRUACHAN AND THE KILLANE, THE  P
CRISTABEL  GOLDEN OLIVE, THE  P
CRISTABEL  MANALACOR OF VELTAKIN  P
CRISTABEL  MORTAL IMMORTALS, THE  H,SFWA
CROCCO, KYLE  HEROES INC.  C/P,P
CROCCO, KYLE  HEROES WANTED  C/P,P
CROMPTON, ANNE ELIOT  MERLIN’S HARP  P
CROMPTON, ANNE ELIOT  PERCIVAL’S ANGEL  P
CRONIN, THOMAS W.  AS IT IS ON MARS (REVISED SECOND EDITION) [AS IT IS ON MARS 1]  H
CRONIN, THOMAS W.  GIVE US THIS MARS [AS IT IS ON MARS 2]  H
CRONIN, THOMAS W.  GLORY BE TO MARS [AS IT IS ON MARS 3]  H

CROOK, JEFF/GRUBBER, JOHN/ HERBERT, MARY H./ BERBERICK, NANCY VARIAN BERTREM’S GUIDE TO THE WAR OF SOULS, VOL. ONE

[@DRAGONLANCE WAR OF SOULS]

P

CROOK, JEFF  CONUNDRUM [@DRAGONLANCE AGE OF MORTALS]  P
CROOK, JEFF  DARK THANE [@DRAGONLANCE AGE OF MORTALS]  P
CROOK, JEFF  THIEVES’ GUILD, THE [@DRAGONLANCE CROSSROADS]  P
CROSS, DEBBIE  WESTERCON 37 PROGRAM BOOK  R/REF-P
CROSS, JANINE  FORGED BY FIRE [@DRAGON TEMPLE #3]  LP
CROSS, JANINE  SHADOWED BY WINGS [@DRAGON TEMPLE #2]  LP
CROSS, JANINE  TOUCHED BY VENOM [@DRAGON TEMPLE #1]  LP
CROSS, JOHN KEIR  ANGRY PLANET, THE  H:2
CROSS, JOHN KEIR  OTHER PASSENGER, THE  C/P,P
CROSS, RONALD ANTHONY  FOURTH GUARDIAN, THE [@ETERNAL GUARDIANS #1]  LP
CROSS, RONALD ANTHONY  LOST GUARDIAN, THE [@ETERNAL GUARDIANS #2]  H,P
CROSS, RONALD ANTHONY  PRISONERS OF PARADISE  H
CROSS, RONALD ANTHONY  WHITE GUARDIAN, THE [ETERNAL GUARDIANS 3]  H,LP
CROSS, TOM PEETE/SLOVER, CLARK HARRIS ANCIENT IRISH TALES  HA
CROSSEN, KENDELL FOSTER  ADVENTURES IN TOMORROW  C/PA,PA
CROSSEN, KENDELL FOSTER  FUTURE TENSE  SR-HA
CROSSEN, KENDELL FOSTER  YEAR OF CONSENT  C/P:2,P
CROWCROFT, PETER  MONSTER  P
CROWDER, HERBERT  AMBUSH AT OSIRAK  P
CROWLEY, ALEISTER  MOONCHILD  C/P,P
CROWLEY, JOHN  AEGYPT [AEGYPT 1]  H,LP
CROWLEY, JOHN  BEASTS  H:2,SFWA,P
CROWLEY, JOHN  DAEMONMANIA  H
CROWLEY, JOHN  DEEP, THE  C/P,SFWA,P,H
CROWLEY, JOHN  ENDLESS THINGS [AEGYPT]  H
CROWLEY, JOHN  ENGINE SUMMER  C/P,SFWA,P,H
CROWLEY, JOHN  GREAT WORK OF TIME  C/P,P
CROWLEY, JOHN  LITTLE, BIG  C/P,LP,H,SFWA
CROWLEY, JOHN  LOVE AND SLEEP [AEGYPT 2]  LP
CROWLEY, JOHN  NOVELTIES & SOUVENIRS  LP
CROWLEY, JOHN  NOVELTY  SR-LP
CROWTHER, PETER  BLUE MOTEL  PA
CROWTHER, PETER  CITIES  HA
CROWTHER, PETER  CONSTELLATIONS  PA
CROWTHER, PETER/KRAMER, EDWARD E.  DANTE’S DISCIPLES  PA
CROWTHER, PETER  DESTINATION UNKNOWN  LPA
CROWTHER, PETER  FORBIDDEN PLANETS  PA
CROWTHER, PETER  FOURSIGHT  HA
CROWTHER, PETER  FUTURES  PA
CROWTHER, PETER  HEAVEN SENT  PA
CROWTHER, PETER  INFINITIES  HA
CROWTHER, PETER  MARS PROBES  PA
CROWTHER, PETER  MOON SHOTS  C/PA,PA
CROWTHER, PETER  TALES IN SPACE  HA,LPA
CROZIER, OUIDA  SHADOWS AFTER DARK  LP
CRUMB, R.  R. CRUMB’S CARLOAD O’ COMICS  CX
CRUME, VIC  SHAGGY D.A., THE  P
CRUME, VIC  UNIDENTIFIED FLYING ODDBALL  P
CULLEN, SEAMUS  ASTRA AND FLONDRIX  C/P,LP
CUMMINGS, M. A.  EXILE AND OTHER TALES OF FANTASY  P

AUTHOR  TITLE  CODE
CUMMINGS, RAY  BEYOND THE STARS  C/P:2,P
CUMMINGS, RAY  BEYOND THE VANISHING POINT  DD331
CUMMINGS, RAY  BRAND NEW WORLD, A  C/P:2,P
CUMMINGS, RAY  BRIGANDS OF THE MOON  C/P,P,H
CUMMINGS, RAY  EXILE OF TIME, THE  C/P:2,P
CUMMINGS, RAY  EXPLORERS INTO INFINITY  H
CUMMINGS, RAY  GIRL IN THE GOLDEN ATOM, THE  H,LP
CUMMINGS, RAY  INSECT INVASION, THE  H
CUMMINGS, RAY  MAN WHO MASTERED TIME, THE  H,DD173
CUMMINGS, RAY  PRINCESS OF THE ATOM, THE  C/P,P
CUMMINGS, RAY  SHADOW GIRL, THE  C/P:2,P
CUMMINGS, RAY  TAMA OF THE LIGHT COUNTRY  C/P:2,P
CUMMINGS, RAY  TAMA, PRINCESS OF MERCURY  C/P,P
CUMMINGS, RAY  TARRANO THE CONQUEROR  H
CUMMINGS, RAY  WANDL THE INVADER  DD497
CUNIS, REINMAR  ZEITSTURM  FP
CUNNINGHAM, ELAINE  BEST OF THE REALMS BOOK III, THE [@FORGOTTEN REALMS]  P
CUNNINGHAM, ELAINE  DARK JOURNEY [@STAR WARS NEW JEDI ORDER]  P

CUNNINGHAM, ELAINE DAUGHTER OF THE DROW

[@FORGOTTEN REALMS STARLIGHT AND SHADOWS 1]

P

CUNNINGHAM, ELAINE  DREAM SPHERES [@FORGOTTEN REALMS]  P
CUNNINGHAM, ELAINE  ELFSHADOW [@FORGOTTEN REALMS HARPERS #2]  C/P:2,P
CUNNINGHAM, ELAINE  ELFSONG [@FORGOTTEN REALMS HARPERS #8]  C/P,P
CUNNINGHAM, ELAINE  EVERMEET: ISLAND OF ELVES [FORGOTTEN REALMS]  P
CUNNINGHAM, ELAINE  FLOODGATE, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #2]  P
CUNNINGHAM, ELAINE  MAGEHOUND, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #1]  P
CUNNINGHAM, ELAINE  RADIANT DRAGON, THE [SPELLJAMMER CLOAKMASTER 4]  C/P,P
CUNNINGHAM, ELAINE  SHADOWS IN THE DARKNESS [CHANGELING DETECTIVE 1]  H
CUNNINGHAM, ELAINE  SHADOWS IN THE STARLIGHT [CHANGELING DETECTIVE 2]  H
CUNNINGHAM, ELAINE  SILVER SHADOWS [@FORGOTTEN REALMS HARPERS #13]  P
CUNNINGHAM, ELAINE  TANGLED WEBS [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 2]  P
CUNNINGHAM, ELAINE  THORNHOLD [@FORGOTTEN REALMS HARPERS]  P
CUNNINGHAM, ELAINE  WINDWALKER [@FORGOTTEN REALMS STARLIGHT AND SHADOWS 3]  P
CUNNINGHAM, ELAINE  WIZARDWAR, THE [@FORGOTTEN REALMS COUNSELORS & KINGS #3]  P
CUNNINGHAM, JERE  VISITOR, THE  P
CUNNINGHAM, MICHAEL  SPECIMEN DAYS  LP
CURRAN, RONALD  WEIRD GATHERING AND OTHER TALES, THE  C/PA,PA
CURRAN, RONALD  WITCHES, WRAITHS AND WARLOCKS  C/PA,PA
CURRAN, TIM  HIVE  LP
CURRY, CHRIS  THUNDER ROAD  P
CURRY, JANE LOUISE  WOLVES OF AAM, THE  H
CURTIS, GENE  SEVENTH MOUNTAIN, THE [MAGI 1]  LP
CURTIS, MYRON  GOSPEL PROBE, THE  LP
CURTIS, RICHARD  FOOL FOR AN AGENT  R/REF-H
CURTIS, RICHARD  FUTURE TENSE  PA,C/PA
CUSHMAN, CAROLYN  WITCH AND WOMBAT  C/P:2,P
CUSICK, RICHIE TANKERSLEY  BUFFY THE VAMPIRE SLAYER [BUFFY]  P
CUSICK, RICHIE TANKERSLEY  EVIL ON THE BAYOU [TWILIGHT #21]  P
CUSICK, RICHIE TANKERSLEY  HARVEST, THE [BUFFY]  P
CUSSLER, CLIVE  CYCLOPS  P
CUSSLER, CLIVE  NIGHT PROBE!  P
CUSSLER, CLIVE  RAISE THE TITANIC!  C/P,P
CUSSLER, CLIVE  TREASURE [DIRK PITT]  P
CUSSLER, CLIVE  VIXEN 03  P
CUSTER, FRED  MORTE DELLA TERRA  FP
CUTTER, LEAH R.  CAVES OF BUDA, THE  P
CUTTER, LEAH R.  JAGUAR AND THE WOLF, THE  P
CUTTER, LEAH R.  PAPER MAGE  C/P,P
CYMRI, CHRYS  DRAGON REFORGED [GONARD 2]  P
CYMRI, CHRYS  DRAGONS CAN ONLY RUST [GONARD 1]  P
CZERNEDA, JULIE E.  BEHOLDER’S EYE [WEB SHIFTERS 1]  C/P,P
CZERNEDA, JULIE E.  CHANGING VISION [@WEB SHIFTERS #2]  C/P:2,P
CZERNEDA, JULIE E.  EXPLORER  LPA
CZERNEDA, JULIE E.  HIDDEN IN SIGHT [WEB SHIFTERS #3]  P
CZERNEDA, JULIE E.  IN THE COMPANY OF OTHERS  H,P
CZERNEDA, JULIE E.  MIGRATION [@SPECIES IMPERATIVE #2]  H
CZERNEDA, JULIE E.  ORBITER  LPA
CZERNEDA, JULIE E.  REGENERATION [@SPECIES IMPERATIVE #3]  H
CZERNEDA, JULIE E./SZPINDEL, ISAAC  REVISIONS  C/PA:2,PA
CZERNEDA, JULIE E.  SPACE INC.  C/PA,PA
CZERNEDA, JULIE E.  STARDUST  LPA
CZERNEDA, JULIE E.  SURVIVAL [@SPECIES IMPERATIVE #1]  H
CZERNEDA, JULIE E.  THOUSAND WORDS FOR STRANGER, A [TRADE PACT 1]  H:2,P
CZERNEDA, JULIE E.  TIES OF POWER [@TRADE PACT #2]  C/P,P,H

AUTHOR  TITLE  CODE
CZERNEDA, JULIE E.  TO TRADE THE STARS [@TRADE PACT #3]  P
CZERNEDA, JULIE E./PANICCIA, JANA  UNDER COVER OF DARKNESS  PA

The MITSFS Pinkdex  September 24, 2007  Page 99
AUTHOR  TITLE  CODE
DA CRUZ, DANIEL  AYES OF TEXAS, THE [TEXAS 1]  C/P,P
DA CRUZ, DANIEL  F-CUBED  C/P,P
DA CRUZ, DANIEL  GROTTO OF THE FORMIGANS, THE  C/P:2,P
DA CRUZ, DANIEL  MIXED DOUBLES  C/P,P
DA CRUZ, DANIEL  TEXAS ON THE ROCKS [TEXAS 2]  C/P,P
DA CRUZ, DANIEL  TEXAS TRIUMPHANT [TEXAS 3]  C/P,P
DA VINCI, LEONARDO  DELUGE, THE  SR-P
DAGMAR, PETER  SANDS OF TIME  H
DAHL, ROALD  JAMES AND THE GIANT PEACH  P
DAHL, ROALD  KISS KISS  SR-P
DAHL, ROALD  MATILDA  H
DAHL, ROALD  SOME TIME NEVER  H
DAHL, ROALD  SWITCH BITCH  P
DAHL, ROALD  WITCHES, THE  P
DAIN, ALEX  BANE OF KANTHOS, THE  D42800
DALBY, RICHARD  MAMMOTH BOOK OF GHOST STORIES, THE  LPA
DALBY, RICHARD  TALES OF WITCHCRAFT  HA
DALBY, RICHARD  VICTORIAN GHOST STORIES  LPA
DALE, FLOYD D.  HUNTER’S FIRE, A  P
D’ALEMBERT, ARTHUR  SONG OF THE SWAN, THE  LP
D’ALEMBERT, ARTHUR  SONG OF THE SWAN, PART 2  LP
DALEY, BRIAN  BROKEN COUNTRY, THE [@GAMMALAW 3]  C/P,P
DALEY, BRIAN  DOOMFARERS OF CORAMONDE, THE [CORAMONDE 1]  C/P:2,P

DALEY, BRIAN EMPIRE STRIKES BACK, THE:

NATIONAL PUBLIC RADIO DRAMATIZATION, THE [STAR WARS]

LP

DALEY, BRIAN  FALL OF THE WHITE SHIP AVATAR [HOBART FLOYT]  C/P:2,P
DALEY, BRIAN  HAN SOLO ADVENTURES, THE [STAR WARS HAN SOLO]  C/P,P
DALEY, BRIAN  HAN SOLO AND THE LOST LEGACY [STAR WARS HAN SOLO]  C/P,P,H
DALEY, BRIAN  HAN SOLO AT STARS’ END [STAR WARS HAN SOLO]  C/P,P,H
DALEY, BRIAN  HAN SOLO’S REVENGE [STAR WARS HAN SOLO]  C/P,P,H
DALEY, BRIAN  JINX ON A TERRAN INHERITANCE [HOBART FLOYT 2]  C/P:2,P
DALEY, BRIAN  REQUIEM FOR A RULER OF WORLDS [HOBART FLOYT 1]  C/P:2,P
DALEY, BRIAN  SCREAMING ACROSS THE SKY, A [@GAMMALAW 2]  C/P,P
DALEY, BRIAN  SMOKE ON THE WATER [@GAMMALAW 1]  C/P,P

DALEY, BRIAN STAR WARS: NATIONAL PUBLIC RADIO DRAMATIZATION, THE

[STAR WARS]

LP

DALEY, BRIAN  STARFOLLOWERS OF CORAMONDE, THE [CORAMONDE 2]  C/P:2,P
DALEY, BRIAN  TAPESTRY OF MAGICS, A  C/P,P
DALEY, BRIAN  TO WATERS’ END [@GAMMALAW 4]  C/P,P
DALEY, BRIAN  TRON  C/P,P,H
DALKEY, KARA  ASCENSION [@WATER #1]  LP
DALKEY, KARA  BHAGAVATI [BLOOD OF THE GODDESS 3]  H
DALKEY, KARA  BIJAPUR [BLOOD OF THE GODDESS 2]  P
DALKEY, KARA  CRYSTAL SAGE  C/P,P
DALKEY, KARA  CURSE OF SAGAMORE, THE [SAGAMORE 1]  P
DALKEY, KARA  EURYALE  P
DALKEY, KARA  GENPEI  H
DALKEY, KARA  GOA [BLOOD OF THE GODDESS 1]  P
DALKEY, KARA  NIGHTINGALE, THE  H,SFWA,P
DALKEY, KARA  REUNION [@WATER #2]  LP
DALKEY, KARA  STEEL ROSE  C/P,P
DALKEY, KARA  SWORD OF SAGAMORE, THE [SAGAMORE 2]  C/P,P
DALLAIRE, NATALIE/COLE, STEPHEN  PARALLEL 59 [DOCTOR WHO EIGHTH DOCTOR]  P
DALMAS, JOHN  BAVARIAN GATE, THE [CURTIS MACURDY 2]  P
DALMAS, JOHN  FANGLITH  C/P,P
DALMAS, JOHN  GENERAL’S PRESIDENT, THE  C/P,SFWA-P,P
DALMAS, JOHN  HELVERTI INVASION, THE [LIZARD WAR 2]  P
DALMAS, JOHN  HOMECOMING  P
DALMAS, JOHN  KALIF’S WAR, THE [REGIMENT 3]  C/P,SFWA-P,P
DALMAS, JOHN  LANTERN OF GOD, THE  C/P,SFWA-P,P
DALMAS, JOHN  LION OF FARSIDE, THE [CURTIS MACURDY 1]  C/P,P
DALMAS, JOHN  LION RETURNS, THE [CURTIS MACURDY 3]  C/P,P
DALMAS, JOHN  LIZARD WAR, THE [LIZARD WAR 1]  C/P:2,P
DALMAS, JOHN  ORC WARS, THE [@YNGLING #1,2]  C/P:2,SFWA-P,P
DALMAS, JOHN/MARTIN, ROD  PLAYMASTERS, THE  C/P,P
DALMAS, JOHN  PUPPET MASTER, THE  P
DALMAS, JOHN  REALITY MATRIX, THE  C/P:2,P
DALMAS, JOHN  REGIMENT, THE [REGIMENT 1]  C/P:2,P
DALMAS, JOHN  REGIMENT’S WAR, THE [REGIMENT 4]  C/P,SFWA-P,P
DALMAS, JOHN  RETURN TO FANGLITH  C/P:2,P
DALMAS, JOHN  SCROLL OF MAN, THE  P,SFWA-P
DALMAS, JOHN  SOLDIERS  C/P,P
DALMAS, JOHN  THREE-CORNERED WAR, THE [REGIMENT 5]  C/P,P
DALMAS, JOHN/MARTIN, CARL  TOUCH THE STARS: EMERGENCE  P

AUTHOR  TITLE  CODE
DALMAS, JOHN  VARKAUS CONSPIRACY, THE  C/P,SFWA-P,P
DALMAS, JOHN  WALKAWAY CLAUSE, THE  C/P,SFWA-P,P
DALMAS, JOHN  WHITE REGIMENT, THE [REGIMENT 2]  C/P:2,SFWA-P,P
DALMAS, JOHN  YNGLING, THE [YNGLING 1]  C/P:2,P
DALMAS, JOHN  YNGLING AND THE CIRCLE OF POWER, THE [YNGLING 2]  C/P:2,SFWA-P,P
DALMAS, JOHN  YNGLING IN YAMATO, THE [YNGLING]  C/P,P
DALOS, GYORGY  1985  H:2,LP
DALTON, CLAIRE  SECOND LIFE OF CECILY PRIDE  P
DALTON, JOHN J.  CATTLE MUTILATORS, THE  P
DALTON, SEAN  BEYOND THE VOID [@STARHAWKS #3]  C/P,P
DALTON, SEAN  CODE NAME PEREGRINE [@STARHAWKS #2]  C/P,P
DALTON, SEAN  DESTINATION: MUTINY [@STARHAWKS #5]  C/P:2,P
DALTON, SEAN  PIECES OF EIGHT [TIME TRAP 3]  P
DALTON, SEAN  PUZZLE [EARTH 2]  C/P,P
DALTON, SEAN  RESTORATION [NOEL KEDRAN 1]  P
DALTON, SEAN  ROSTMA LURE, THE [@STARHAWKS #4]  C/P,P
DALTON, SEAN  SALUKAN GAMBIT [@STARHAWKS #6]  P
DALTON, SEAN  SHOWDOWN [TIME TRAP 2]  P
DALTON, SEAN  SPACE HAWKS [@STARHAWKS #1]  C/P:2,P
DALTON, SEAN  TERMINATION [NOEL KEDRAN 3]  C/P,P
DALTON, SEAN  TIME TRAP [TIME TRAP 1]  P
DALTON, SEAN  TURNCOAT [NOEL KEDRAN 2]  P
DALY, C. R.  TITAN A.E.: THE JUNIOR NOVELIZATION  LP
D’AMELIO, DAN  SILVABAMBA  SR-P
D’AMMASSA, DON  TWISTED IMAGES  B
DAN, URI/MANN, PETER  ULTIMATUM: PU 94  P
DANBY, MARY  FIFTH FONTANA BOOK OF GREAT HORROR STORIES, THE  C/PA,PA
DANE, CHRISTOPHER  GALACTIC ARENA, THE  P
DANE, CHRISTOPHER  RIDERS OF THE DRAGON  P
DANE, CLEMENCE  ARROGANT HISTORY OF WHITE BEN, THE  H
DANGERFIELD, PAUL  ISLAND OF THE VOODOO DOLLS  SR-L
DANIEL, COLIN  DEMON TREE [TWILIGHT 9]  P
DANIEL, JERRY C.  SPACE MACHINE, THE  H
DANIEL, TONY  EARTHLING  H,LP
DANIEL, TONY  METAPLANETARY [METAPLANETARY 1]  H
DANIEL, TONY  SUPERLUMINAL [METAPLANETARY 2]  H
DANIEL, TONY  WARPATH  C/P,P
DANIELEWSKI, MARK Z.  HOUSE OF LEAVES  LP
DANIELL, TINA  COMPANIONS, THE [@DRAGONLANCE MEETINGS #6]  C/P,P
DANIELL, TINA  DARK HEART [@DRAGONLANCE MEETINGS #3]  C/P:2,P
DANIELL, TINA  MAQUESTA KAR-THON [@DRAGONLANCE WARRIORS #2]  P
DANIELS, CASEY  DON OF THE DEAD [PEPPER MARTIN]  P
DANIELS, DOROTHY  GHOST SONG  P
DANIELS, JONATHAN  CLASH OF ANGELS  H
DANIELS, LES  BLACK CASTLE, THE [DON SEBASTIAN 1]  H,P
DANIELS, LES  CITIZEN VAMPIRE [DON SEBASTIAN 3]  P
DANIELS, LES  NO BLOOD SPILLED [DON SEBASTIAN 5]  P,SFWA-P
DANIELS, LES  SILVER SKULL, THE [DON SEBASTIAN 2]  P
DANIELS, LES  YELLOW FOG [DON SEBASTIAN 4]  SR-H,SFWA-P
DANIELS, MAX  OFFWORLD  P
DANIELS, MAX  SPACE GUARDIAN, THE  P
DANIELS, NORMAN  MAGNETIC MAN, THE [AVENGERS #8]  C/P:2,P
DANIELS, NORMAN  MOON EXPRESS [AVENGERS #9]  C/P,P
DANN, JACK/DOZOIS, GARDNER  A.I.S  PA
DANN, JACK/DOZOIS, GARDNER  ALIENS AMONG US  C/PA:2,PA
DANN, JACK/DOZOIS, GARDNER  ANGELS!  PA
DANN, JACK/DOZOIS, GARDNER  ARMAGEDDONS  PA
DANN, JACK/DOZOIS, GARDNER  BESTIARY!  C/PA,PA
DANN, JACK/DOZOIS, GARDNER  BEYOND FLESH  PA
DANN, JACK/DOZOIS, GARDNER  BEYOND SINGULARITY  PA
DANN, JACK/DOZOIS, GARDNER  CLONES  PA
DANN, JACK/DOZOIS, GARDNER  DEMONS!  PA
DANN, JACK/DOZOIS, GARDNER  DINOSAURS!  C/PA,PA
DANN, JACK/DOZOIS, GARDNER  DINOSAURS II  PA
DANN, JACK/DOZOIS, GARDNER  DOGTALES!  PA
DANN, JACK/DOZOIS, GARDNER  DRAGONS!  PA
DANN, JACK/WEBB, JANEEN  DREAMING DOWN UNDER  HA,LPA
DANN, JACK/DOZOIS, GARDNER  ESCAPE FROM EARTH  HA
DANN, JACK/ZEBROWSKI, GEORGE  FASTER THAN LIGHT  C/PA,PA,SFWA
DANN, JACK  FICTION FACTORY, THE  H
DANN, JACK/DOZOIS, GARDNER  FUTURE CRIMES  PA
DANN, JACK/DOZOIS, GARDNER  FUTURE POWER  SR-HA,SFWA,HA
DANN, JACK/DOZOIS, GARDNER  FUTURE SPORTS  PA
DANN, JACK/DOZOIS, GARDNER  FUTURE WAR  PA

AUTHOR  TITLE  CODE
DANN, JACK/DOZOIS, GARDNER  FUTURES PAST  PA
DANN, JACK/DOZOIS, GARDNER  GENOMETRY  PA
DANN, JACK/DOZOIS, GARDNER  HACKERS  PA
DANN, JACK/DOZOIS, GARDNER  HORSES!  PA
DANN, JACK  IMMORTAL  HA,SFWA
DANN, JACK/DOZOIS, GARDNER  IMMORTALS  PA
DANN, JACK/DOZOIS, GARDNER  INVADERS!  PA
DANN, JACK  JUBILEE  H
DANN, JACK  JUNCTION  P
DANN, JACK/DOZOIS, GARDNER  LITTLE PEOPLE!  PA
DANN, JACK/DOZOIS, GARDNER  MAGICATS!  C/PA,PA
DANN, JACK/DOZOIS, GARDNER  MAGICATS 2  C/PA,PA
DANN, JACK  MAN WHO MELTED, THE  H,P
DANN, JACK/DOZOIS, GARDNER  MERMAIDS!  C/PA,PA
DANN, JACK  MORE WANDERING STARS  SFWA
DANN, JACK/DOZOIS, GARDNER  NANOTECH  C/PA,PA
DANN, JACK  NEBULA AWARDS 32  LPA
DANN, JACK  NEBULA AWARDS SHOWCASE 2005, THE  LPA
DANN, JACK/DOZOIS, GARDNER  SEASERPENTS!  PA
DANN, JACK/DOZOIS, GARDNER  SORCERERS!  C/PA:2,PA
DANN, JACK/DOZOIS, GARDNER  SPACE SOLDIERS  C/PA,PA
DANN, JACK  STARHIKER  H:2,SFWA
DANN, JACK/DOZOIS, GARDNER  TIMEGATES  C/PA,PA
DANN, JACK  TIMETIPPING  H,SFWA
DANN, JACK/DOZOIS, GARDNER  UNICORNS!  C/PA,PA
DANN, JACK/DOZOIS, GARDNER  UNICORNS 2  PA
DANN, JACK  WANDERING STARS  HA,PA,SFWA
DANN, JACK/DOZOIS, GARDNER  WIZARDS  HA
DANN, JEANNE VAN BUREN/DANN, JACK  IN THE FIELD OF FIRE  PA,SFWA-P,SFWA
DANN, JOSHUA  SECOND TIME AROUND [@TIMESHARE 2]  P
DANN, JOSHUA  TIME FOR WAR, A [@TIMESHARE 3]  P
DANN, JOSHUA  TIMESHARE [@TIMESHARE 1]  P
DANN, SAM  THIRD BODY, THE  P
DANSKY, RICHARD E.  BELOVED OF THE DEAD [@EXALTED SECOND AGE #2]  P
DANSKY, RICHARD E.  CHILDREN OF THE DRAGON [@EXALTED SECOND AGE #3]  P
DANSKY, RICHARD E.  CHOSEN THE SUN [@EXALTED SECOND AGE #1]  P
DANSKY, RICHARD E.  WRAITH: THE OBLIVION (2ND EDITION) [WORLD OF DARKNESS WRAITH]  [C/RAN-H]
DANVERS, DENNIS  CIRCUIT OF HEAVEN [CIRCUIT OF HEAVEN 1]  H,P
DANVERS, DENNIS  END OF DAYS [CIRCUIT OF HEAVEN 2]  C/P,P
DANVERS, DENNIS  FOURTH WORLD, THE  H
DANVERS, DENNIS  WATCH, THE  H,LP
DANVERS, DENNIS  WATCH, THE (PROOF)  SR-LP
DANVERS, DENNIS  WILDERNESS  C/P,P,H
DANZIGER, PAULA  THIS PLACE HAS NO ATMOSPHERE  P
D’ARC, BIANCA  ICE DRAGON, THE [@DRAGON KNIGHTS #3]  LP
DARK HARVEST  NIGHT VISIONS 4  SR-HA
DARK HARVEST  NIGHT VISIONS 6  SR-HA
DARK HARVEST  NIGHT VISIONS 8  SR-HA
DARK HARVEST  NIGHT VISIONS 9  SR-HA
DARK, JAMES  INVISIBLES, THE [MARK HOOD 9]  P
DARK, JAMES  OPERATION OCTOPUS [MARK HOOD 8]  P
DARK, JAMES  SWORD OF GENGHIS KHAN, THE [MARK HOOD 7]  P
DARK, JAMES  THRONE OF SATAN [MARK HOOD 6]  C/P,P
DARLTON, CLARK  ATOM FIRE ON MECHANICA [PERRY RHODAN 121]  SR-LP
DARLTON, CLARK  DEADMEN SHOULDN’T DIE [PERRY RHODAN 132]  SR-LP
DARLTON, CLARK  GUCKY UND DIE MORDWESPEN [@PERRY RHODAN #28]  FPR
DARLTON, CLARK  MANNER FUR LACERTAE [@PERRY RHODAN #64]  FPR
DARLTON, CLARK  SENTINELS OF SOLITUDE [PERRY RHODAN 127]  SR-LP
DARLTON, CLARK  SIE KAMEN VON KENDARKOS [@PERRY RHODAN #142]  FPR
DARLTON, CLARK  VERBOTENE SONNENSYSTEM, DAS [@PERRY RHODAN #58]  FPR
DARNAY, ARSEN  HOSTAGE FOR HINTERLAND, A  C/P,P
DARNAY, ARSEN  KARMA  H,P
DARNAY, ARSEN  KARMA AFFAIR, THE  P
DARNAY, ARSEN  PURGATORY ZONE, THE  C/P,P
DARNAY, ARSEN  SIEGE OF FALTARA, THE  P
DARNAY, ARSEN  SPLENDID FREEDOM, THE  C/P,P
DARNTON, JOHN  EXPERIMENT, THE  H
DARNTON, JOHN  NEANDERTHAL  C/P,P,H
DART-THORNTON, CECILIA  BATTLE OF EVERNIGHT, THE [BITTERBYNDE 3]  C/P,P
DART-THORNTON, CECILIA  BATTLE OF EVERNIGHT, THE (PROOF) [BITTERBYNDE 3]  SR-LP
DART-THORNTON, CECILIA  ILL-MADE MUTE, THE [BITTERBYNDE 1]  C/P:2,P
DART-THORNTON, CECILIA  IRON TREE, THE [CROWTHISTLE 1]  H,P
DART-THORNTON, CECILIA  LADY OF THE SORROWS, THE [BITTERBYNDE 2]  C/P,P
DARVILL-EVANS, PETER  ASYLUM [DOCTOR WHO PAST DOCTORS]  P
DARVILL-EVANS, PETER  DECEIT [DOCTOR WHO NEW ADVENTURES]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 102
AUTHOR  TITLE  CODE
DARVILL-EVANS, PETER  INDEPENDENCE DAY [DOCTOR WHO PAST DOCTORS]  C/P,P
DATLOW, ELLEN  ALIEN SEX  SR-HA,C/PA,PA
DATLOW, ELLEN/WINDLING, TERRI  BLACK HEART, IVORY BONES [COMPLEMENTARY FAIRY COLORS 6]  LPA
DATLOW, ELLEN/WINDLING, TERRI  BLACK SWAN, WHITE RAVEN [COMPLEMENTARY FAIRY COLORS 4]  HA
DATLOW, ELLEN/WINDLING, TERRI  BLACK THORN, WHITE ROSE [COMPLEMENTARY FAIRY COLORS 2]  C/PA,PA
DATLOW, ELLEN  BLOOD IS NOT ENOUGH  C/PA,PA
DATLOW, ELLEN  DARK, THE  LPA
DATLOW, ELLEN/WINDLING, TERRI  FAERY REEL, THE  HA
DATLOW, ELLEN  FIRST OMNI BOOK OF SCIENCE FICTION, THE  C/PA,PA
DATLOW, ELLEN  FOURTH OMNI BOOK OF SCIENCE FICTION, THE  PA
DATLOW, ELLEN  LITTLE DEATHS  PA
DATLOW, ELLEN  OFF LIMITS  C/PA,PA
DATLOW, ELLEN  OMNI BEST SCIENCE FICTION TWO  LPA
DATLOW, ELLEN  OMNI VISIONS TWO  LPA
DATLOW, ELLEN/WINDLING, TERRI  RUBY SLIPPERS, GOLDEN TEARS [COMPLEMENTARY FAIRY COLORS 3]  HA,PA
DATLOW, ELLEN/WINDLING, TERRI  SALON FANTASTIQUE  LPA
DATLOW, ELLEN  SECOND OMNI BOOK OF SCIENCE FICTION, THE  PA
DATLOW, ELLEN/WINDLING, TERRI  SILVER BIRCH, BLOOD MOON [COMPLEMENTARY FAIRY COLORS 5]  HA,LPA
DATLOW, ELLEN/WINDLING, TERRI  SIRENS  PA
DATLOW, ELLEN/WINDLING, TERRI  SNOW WHITE, BLOOD RED [COMPLEMENTARY FAIRY COLORS 1]  SR-HA,PA,C/PA,SFWA
DATLOW, ELLEN/WINDLING, TERRI  SWAN SISTER  LPA
DATLOW, ELLEN  THIRD OMNI BOOK OF SCIENCE FICTION, THE  C/PA,PA
DATLOW, ELLEN  VANISHING ACTS  HA,LPA
DATLOW, ELLEN  WHISPER OF BLOOD, A  PA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY, FIRST ANNUAL COLLECTION, THE  HA,LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY, SECOND ANNUAL COLLECTION, THE  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, THIRD ANNUAL COLLECTION, THE  LPA

DATLOW, ELLEN/WINDLING, TERRI YEAR’S BEST FANTASY AND HORROR, FOURTH ANNUAL COLLECTION,

THE

LPA

DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, FIFTH ANNUAL COLLECTION, THE  LPA

DATLOW, ELLEN/WINDLING, TERRI YEAR’S BEST FANTASY AND HORROR, SEVENTH ANNUAL COLLECTION,

THE

LPA

DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, EIGHTH ANNUAL COLLECTION, THE LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, NINTH ANNUAL COLLECTION, THE  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, TENTH ANNUAL COLLECTION  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, ELEVENTH ANNUAL COLLECTION  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, TWELFTH ANNUAL COLLECTION  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, THIRTEENTH ANNUAL COLLECTION LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, FOURTEENTH ANNUAL COLECTION LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, FIFTEENTH ANNUAL COLLECTION  LPA
DATLOW, ELLEN/WINDLING, TERRI  YEAR’S BEST FANTASY AND HORROR, SIXTEENTH ANNUAL COLLECTION  LPA

DATLOW, ELLEN/LINK, KELLY/ GRANT, GAVIN J.

YEAR’S BEST FANTASY AND HORROR, SEVENTEENTH ANNUAL COLLECTION, THE

LPA

DATLOW, ELLEN/LINK, KELLY/ GRANT, GAVIN J.

YEAR’S BEST FANTASY AND HORROR, EIGHTEENTH ANNUAL COLLECTION, THE

LPA

DATLOW, ELLEN/LINK, KELLY/ GRANT, GAVIN J.

YEAR’S BEST FANTASY AND HORROR, NINETEENTH ANNUAL COLLECTION, THE

LPA

DAVENPORT, BASIL  DEALS WITH THE DEVIL  PA,C/PA
DAVENPORT, BASIL  INVISIBLE MEN  PA,C/PA
DAVENPORT, BASIL  TALES TO BE TOLD IN THE DARK  PA
DAVENTRY, LEONARD  MAN OF DOUBLE DEED, A  C/P,P,H
DAVENTRY, LEONARD  REFLECTIONS IN A MIRAGE  C/P,P
DAVENTRY, LEONARD  REFLECTIONS IN A MIRAGE/TICKING IS IN YOUR HEAD, THE  H
DAVENTRY, LEONARD  TICKING IS IN YOUR HEAD, THE  C/P,P
DAVENTRY, LEONARD  TWENTY-ONE BILLIONTH PARADOX  H,SFWA
DAVID, JAMES F.  FOOTPRINTS OF THUNDER [FOOTPRINTS OF THUNDER 1]  P
DAVID, JAMES F.  SHIP OF THE DAMNED  P
DAVID, JAMES F.  THUNDER OF TIME [FOOTPRINTS OF THUNDER 2]  H
DAVID, JOHN  BECOMING DEATH [@ENEMIES OF SOCIETY #1]  LP
DAVID, JOHN  DESTROYING WORLDS [@ENEMIES OF SOCIETY #2]  LP
DAVID, JOHN  PASSION FOR DEAD LEAVES [ENEMIES OF SOCIETY 3]  LP
DAVID, JOHN  SEND IN THE SWORDS [ENEMIES OF SOCIETY 4]  LP
DAVID, PETER  AFTER THE FALL [STAR TREK NEW FRONTIER 15]  H
DAVID, PETER  ARMIES OF LIGHT AND DARK [@BABYLON 5 LEGIONS OF FIRE #2]  C/P,P
DAVID, PETER  BATMAN FOREVER  P
DAVID, PETER  BODY AND SOUL [ALIEN NATION 3]  C/P:2,P
DAVID, PETER  CAPTAIN’S DAUGHTER, THE [STAR TREK]  C/P:2,P
DAVID, PETER  CHAOS KID, THE [@PSI-MAN #4]  C/P,P
DAVID, PETER  COLD WARS [STAR TREK GATEWAYS #6/STAR TREK NEW FRONTIER]  P
DAVID, PETER  DARK ALLIES [@STAR TREK NEW FRONTIER #8]  P
DAVID, PETER  DEATHSCAPE [@PSI-MAN #2]  P

DAVID, PETER/FRIEDMAN, MICHAEL JAN/ GREENBERGER, ROBERT DISINHERITED, THE [STAR TREK]

C/P:2,P

DAVID, PETER  DOUBLE OR NOTHING [@STAR TREK TNG DOUBLE HELIX 5]  P
DAVID, PETER  END GAME [@STAR TREK NEW FRONTIER #4]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 103
AUTHOR  TITLE  CODE
DAVID, PETER  EXCALIBUR: RENAISSANCE [@STAR TREK NEW FRONTIER #10]  P
DAVID, PETER  EXCALIBUR: REQUIEM [@STAR TREK NEW FRONTIER #9]  P
DAVID, PETER  EXCALIBUR: RESTORATION [@STAR TREK NEW FRONTIER #11]  P
DAVID, PETER  FIRE ON HIGH [@STAR TREK NEW FRONTIER #6]  P
DAVID, PETER  GODS ABOVE [@STAR TREK NEW FRONTIER 13]  P
DAVID, PETER  HAVEN [@PSI-MAN #6]  P
DAVID, PETER  HOUSE OF CARDS [@STAR TREK NEW FRONTIER #1]  C/P,P
DAVID, PETER  HOWLING MAD  C/P,P
DAVID, PETER  HULK [HULK]  C/P,P
DAVID, PETER  IMZADI [STAR TREK TNG]  H:2,P
DAVID, PETER  IMZADI FOREVER [STAR TREK TNG]  LP
DAVID, PETER  IN THE BEGINNING [@BABYLON 5]  C/P,P
DAVID, PETER  INTO THE VOID [@STAR TREK NEW FRONTIER #2]  C/P,P
DAVID, PETER  KNIGHT LIFE [ARTHUR PENN 1]  P,SFWA-P
DAVID, PETER  LINE OF FIRE [@STAR TREK TNG STARFLEET ACADEMY #2]  H,LP

DAVID, PETER LONG NIGHT OF CENTAURI PRIME, THE [@BABYLON 5 LEGIONS OF FIRE #1]

P

DAVID, PETER  MAIN STREET D.O.A. [@PSI-MAN #3]  P
DAVID, PETER  MARTYR [@STAR TREK NEW FRONTIER #5]  P
DAVID, PETER  MIND FORCE WARRIOR [@PSI-MAN #1]  P
DAVID, PETER  MISSING IN ACTION [STAR TREK NEW FRONTIER 16]  H
DAVID, PETER  NO LIMITS [@STAR TREK NEW FRONTIER]  LPA

DAVID, PETER ONCE BURNED [STAR TREK CAPTAIN’S TABLE 5/

STAR TREK NEW FRONTIER]

C/P,P

DAVID, PETER  ONE KNIGHT ONLY [ARTHUR PENN 2]  P
DAVID, PETER  OUT OF THE DARKNESS [@BABYLON 5 LEGIONS OF FIRE #3]  C/P,P
DAVID, PETER  Q-IN-LAW [STAR TREK TNG]  C/P:2,P
DAVID, PETER  Q-SQUARED [STAR TREK TNG]  H:2,P
DAVID, PETER  QUIET PLACE, THE [@STAR TREK NEW FRONTIER #7]  P
DAVID, PETER  RETURN OF SWAMP THING, THE  P
DAVID, PETER  RIFT, THE [STAR TREK]  C/P:3,P
DAVID, PETER  ROCK AND A HARD PLACE, A [STAR TREK TNG]  C/P:2,P
DAVID, PETER  ROCKETEER, THE  P
DAVID, PETER  SAGITTARIUS IS BLEEDING [@BATTLESTAR GALACTICA 2003]  LP
DAVID, PETER  SIEGE, THE [STAR TREK DS9]  C/P:2,P
DAVID, PETER  SIR APROPOS OF NOTHING [SIR APROPOS OF NOTHING 1]  C/P,P
DAVID, PETER  SPIDER-MAN [SPIDER-MAN]  P
DAVID, PETER  SPIDER-MAN 2  H
DAVID, PETER  SPIKE VS. DRACULA [BUFFY SPIKE]  CX
DAVID, PETER  STALKER [@PSI-MAN #5]  P
DAVID, PETER  STONE AND ANVIL [@STAR TREK NEW FRONTIER 14]  P
DAVID, PETER  STRIKE ZONE [STAR TREK TNG]  C/P:3,P
DAVID, PETER  SURVIVAL [@STAR TREK TNG STARFLEET ACADEMY #3]  H,LP
DAVID, PETER  THIRDSPACE [@BABYLON 5]  P
DAVID, PETER  TWO-FRONT WAR, THE [@STAR TREK NEW FRONTIER #3]  C/P,P
DAVID, PETER  VENDETTA [STAR TREK TNG]  C/P:2,P
DAVID, PETER  WHAT SAVAGE BEAST [HULK]  P
DAVID, PETER  WOAD TO WUIN, THE [SIR APROPOS OF NOTHING 2]  H,P
DAVID, PETER  WORF’S FIRST ADVENTURE [@STAR TREK TNG STARFLEET ACADEMY #1]  H:3,LP

DAVID, PETER/FRIEDMAN, MICHAEL JAN/ GREENBERGER, ROBERT WRATH OF THE PROPHETS [STAR TREK DS9]

C/P,P

DAVIDS, PAUL/DAVIDS, HOLLACE  STAR WARS BOOK ONE [STAR WARS]  H
DAVIDSON, AVRAM  ADVENTURES IN UNHISTORY  H
DAVIDSON, AVRAM  AVRAM DAVIDSON TREASURY, THE  H
DAVIDSON, AVRAM  BEST FROM FANTASY AND SCIENCE FICTION 12, THE  HA,PA
DAVIDSON, AVRAM  BEST FROM FANTASY AND SCIENCE FICTION 13, THE  C/PA,PA
DAVIDSON, AVRAM  BEST FROM FANTASY AND SCIENCE FICTION 14, THE  C/PA,PA
DAVIDSON, AVRAM  BEST OF AVRAM DAVIDSON, THE  H,SFWA
DAVIDSON, AVRAM/DAVIS, GRANIA  BOSS IN THE WALL, A TREATISE ON THE HOUSE DEVIL, THE  SR-LP
DAVIDSON, AVRAM  CLASH OF STAR-KINGS  DG576
DAVIDSON, AVRAM  COLLECTED FANTASIES  P
DAVIDSON, AVRAM  ENEMY OF MY ENEMY, THE  C/P:2,P
DAVIDSON, AVRAM  ENQUIRIES OF DOCTOR ESZTERHAZY, THE  P
DAVIDSON, AVRAM  ISLAND UNDER THE EARTH, THE  C/P:2,P
DAVIDSON, AVRAM  KAR-CHEE REIGN, THE  DG574
DAVIDSON, AVRAM  KAR-CHEE REIGN, THE/ROGUE DRAGON  C/P,P
DAVIDSON, AVRAM  MAGIC FOR SALE  C/PA,PA
DAVIDSON, AVRAM/DAVIS, GRANIA  MARCO POLO AND THE SLEEPING BEAUTY  P,SFWA-P
DAVIDSON, AVRAM  MASTERS OF THE MAZE  C/P:2,P
DAVIDSON, AVRAM  MUTINY IN SPACE  C/P:2,P
DAVIDSON, AVRAM  OR ALL THE SEAS WITH OYSTERS  C/P:2,P
DAVIDSON, AVRAM  OTHER NINETEENTH CENTURY, THE  LP
DAVIDSON, AVRAM  PEREGRINE: PRIMUS  H:2,SFWA,P
DAVIDSON, AVRAM  PEREGRINE: SECUNDUS  P

AUTHOR  TITLE  CODE
DAVIDSON, AVRAM  PHOENIX AND THE MIRROR, THE  C/P,P,H
DAVIDSON, AVRAM  REDWARD EDWARD PAPERS, THE  H,SFWA
DAVIDSON, AVRAM  ROGUE DRAGON  C/P:2,P
DAVIDSON, AVRAM  RORK!  C/P:2,P
DAVIDSON, AVRAM  STRANGE SEAS AND SHORES  H:2,P
DAVIDSON, AVRAM  URSUS OF ULTIMA THULE  C/P:2,P
DAVIDSON, AVRAM  VERGIL IN AVERNO  H,SFWA
DAVIDSON, AVRAM  WHAT STRANGE STARS AND SKIES  C/P:2,P
DAVIDSON, FRANK P./COX, JOHN STUART  MACRO  [C/RAN-H]
DAVIDSON, LIONEL  UNDER PLUM LAKE  C/P,P,H
DAVIDSON, MARYJANICE  DEAD AND LOVING IT  LP
DAVIDSON, MARYJANICE  DERIK’S BANE [WYNDHAM WEREWOLF]  P
DAVIDSON, MARYJANICE  HELLO, GORGEOUS  LP
DAVIDSON, MARYJANICE  SLEEPING WITH THE FISHES  P
DAVIDSON, MARYJANICE  UNDEAD AND UNAPPRECIATED [BETSY TAYLOR 3]  H,P
DAVIDSON, MARYJANICE  UNDEAD AND UNEASY [BETSY TAYLOR 6]  H
DAVIDSON, MARYJANICE  UNDEAD AND UNEMPLOYED [BETSY TAYLOR 2]  P
DAVIDSON, MARYJANICE  UNDEAD AND UNPOPULAR [BETSY TAYLOR 5]  H:2
DAVIDSON, MARYJANICE  UNDEAD AND UNRETURNABLE [BETSY TAYLOR 4]  H
DAVIDSON, MICHAEL  DAUGHTER OF IS  C/P:2,P
DAVIDSON, MICHAEL  KARMA MACHINE, THE  C/P,P
DAVIE, GEORGE E.  MICKEY MOUSE ADVENTURE IN OUTER SPACE  BLB
DAVIES, ANTHEA  WHITE HORSE WITH WINGS, A  H
DAVIES, FREDRIC  CROSS OF GOLD AFFAIR, THE [MAN FROM UNCLE 14]  C/P:2,P
DAVIES, L. P.  ALIEN, THE  H,SFWA
DAVIES, L. P.  ARTIFICIAL MAN, THE  H:2,P
DAVIES, L. P.  DIMENSION A  C/P:2,P
DAVIES, L. P.  GENESIS TWO  C/P,P,H
DAVIES, L. P.  GRAVE MATTER, A  H
DAVIES, L. P.  PAPER DOLLS, THE  H,P
DAVIES, L. P.  PSYCHOGEIST  H:2,P
DAVIES, L. P.  TWILIGHT JOURNEY  H:2
DAVIES, L. P.  WHAT DID I DO TOMORROW?  H
DAVIES, PETE  DOLLARVILLE  H
DAVIES, RUSSELL T.  DAMAGED GOODS [DOCTOR WHO NEW ADVENTURES]  C/P,P
DAVIES, VALENTINE  MIRACLE ON 34TH STREET  P
DAVIS, BART  BLIND PROPHET  H
DAVIS, BRETT  BONE WARS  C/P,P
DAVIS, BRETT  FAERY CONVENTION, THE  P
DAVIS, BRETT  HAIR OF THE DOG  C/P,P
DAVIS, BRETT  TWO TINY CLAWS  P
DAVIS, CAROL/REESE, ESTHER D.  MIRROR’S EDGE [@QUANTUM LEAP]  P
DAVIS, CAROL  OBSESSIONS [@QUANTUM LEAP]  P
DAVIS, DON/DAVIS, JAY  SINS OF THE FLESH  SFWA-P
DAVIS, ELIZABETH  MY SOUL TO KEEP  P
DAVIS, GERRY/BINGEMAN, ALISON  CELESTIAL TOYMAKER, THE [DOCTOR WHO]  SR-P
DAVIS, GERRY  DOCTOR WHO AND THE CYBERMEN [DOCTOR WHO]  SR-P,C/P:2
DAVIS, GERRY  DOCTOR WHO AND THE TENTH PLANET [DOCTOR WHO]  SR-P,C/P:2
DAVIS, GERRY  DOCTOR WHO AND THE TOMB OF THE CYBERMEN [DOCTOR WHO]  SR-P,C/P:2
DAVIS, GRAEME  BLOOD AND HONOR [@EBERRON WAR-TORN #4]  P
DAVIS, GRANIA  MOONBIRD  H
DAVIS, GRANIA  RAINBOW ANNALS, THE  P
DAVIS, JAMES P  BLOODWALK [FORGOTTEN REALMS WIZARDS 5]  P
DAVIS, MAGGIE  FORBIDDEN OBJECTS  P
DAVIS, MARGARET  MINDLIGHT  C/P,P
DAVIS, MARGARET  MINDS APART  P
DAVIS PUBLICATIONS  ASTRONAUTS AND ANDROIDS  C/PA,PA
DAVIS PUBLICATIONS  ISAAC ASIMOV’S TOMORROW’S VOICES  HA,SFWA
DAVIS PUBLICATIONS  SIX DECADES  SR-LPA
DAVIS, RICHARD  YEAR’S BEST HORROR STORIES I, THE  C/PA,PA
DAVIS, RICHARD  YEAR’S BEST HORROR STORIES II, THE  PA
DAVIS, RICHARD  YEAR’S BEST HORROR STORIES #2, THE (BRITISH)  SR-PA
DAVIS, RICHARD  YEAR’S BEST HORROR STORIES III, THE  PA
DAVIS, TOM/AYKROYD, DAN  CONEHEADS: THE LIFE AND TIMES OF BELDAR CONEHEAD  LP
DAVISON, PETER  PETER DAVISON’S BOOK OF ALIEN PLANETS  SR-PA
DAWSON, ROXANN/GRAHAM, DANIEL  ENTERING TENEBREA [@TENEBREA #1]  P
DAWSON, ROXANN/GRAHAM, DANIEL  TENEBREA RISING [@TENEBREA #3]  P
DAWSON, ROXANN/GRAHAM, DANIEL  TENEBREA’S HOPE [@TENEBREA #2]  P
DAY, ALYSSA  ATLANTIS RISING  P
DAY, BRADFORD M.  COMPLETE CHECKLIST OF SCIENCE-FICTION MAGAZINES, THE  R/IND:2
DAY, BRADFORD M.  EDGAR RICE BURROUGHS: A BIBLIOGRAPHY  R/IND
DAY, CHET  HACKER, THE  SFWA-P
DAY, DAVID  HOBBIT COMPANION, THE [LORD OF THE RINGS]  R/REF-H
DAY, DAVID  TOLKIEN BESTIARY, A  R/REF-H

The MITSFS Pinkdex  September 24, 2007  Page 105
AUTHOR  TITLE  CODE
DAY, DAVID  TOLKIEN: THE ILLUSTRATED ENCYCLOPAEDIA  R/REF-H
DAY, DONALD B.  INDEX TO THE SCIENCE FICTION MAGAZINES, 1926–1950  R/IND
DAY, LIONEL  BURIED WORLD, THE  B

DAY, MARTIN/BEECH, LEN ANOTHER GIRL, ANOTHER PLANET

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

C/P,P

DAY, MARTIN  BUNKER SOLDIERS [DOCTOR WHO PAST DOCTORS]  P
DAY, MARTIN  MENAGERIE, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
DAY, MARTIN  SLEEP OF REASON, THE [@DOCTOR WHO EIGHTH DOCTOR]  P
DAY, MARTIN  WOODEN HEART [@DOCTOR WHO]  H
DAYTON, ARWEN ELYS  RESURRECTION  P
DAYTON, GAIL  BARBED ROSE, THE [ONE ROSE 2]  LP
DAYTON, GAIL  COMPASS ROSE, THE [ONE ROSE 1]  LP
DAZAI, OSAMU  BLUE BAMBOO  P
DE ANGELES, LY  QUICKENING, THE  LP
DE ANGELES, LY  SHINING ISLE  LP
DE BERGERAC, CYRANO  VOYAGE DANS LA LUNE  FP
DE BIBIENA, JEAN-GALLI  AMOROUS PHILANDRE  SR-P
DE BIE, ERIK SCOTT  DEPTHS OF MADNESS [@FORGOTTEN REALMS DUNGEONS]  P
DE BOLT, ADRIANA  ALIEN WITHIN, THE  P
DE BOLT, ADRIANA  VOYAGE OF THE TRIGON  P
DE CAMP, L. SPRAGUE  ARROWS OF HERCULES, THE  P
DE CAMP, L. SPRAGUE  BEST OF L. SPRAGUE DE CAMP, THE  H:2,P
DE CAMP, L. SPRAGUE  BLADE OF CONAN, THE [CONAN]  R/REF-P,C/REF-P
DE CAMP, L. SPRAGUE  BLOND BARBARIANS AND NOBLE SAVAGES  R/REF-P

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK BONES OF ZORA, THE [KRISHNA]

H,P

DE CAMP, L. SPRAGUE  BRONZE GOD OF RHODES, THE  C/P,P
DE CAMP, L. SPRAGUE/PRATT, FLETCHER  CARNELIAN CUBE, THE  C/P:2,P
DE CAMP, L. SPRAGUE/PRATT, FLETCHER  CASTELLO D’ACCIAIO, IL  FP
DE CAMP, L. SPRAGUE/PRATT, FLETCHER  CASTLE OF IRON, THE  C/P,P,H
DE CAMP, L. SPRAGUE  CLOCKS OF IRAZ, THE [RELUCTANT KING 2]  C/P:2,P

DE CAMP, L. SPRAGUE/PRATT, FLETCHER COMPLEAT ENCHANTER: THE MAGICAL MISADVENTURES OF HAROLD

SHEA, THE [INCOMPLETE ENCHANTER]

C/P,P,H

DE CAMP, L. SPRAGUE/PRATT, FLETCHER  COMPLETE COMPLEAT ENCHANTER, THE [INCOMPLETE ENCHANTER]  C/P:2,SFWA-P,P
DE CAMP, L. SPRAGUE  CONAN AND THE SPIDER GOD [BANTAM CONAN 5]  C/P,P
DE CAMP, L. SPRAGUE/CARTER, LIN  CONAN OF AQUILONIA [CONAN 11]  C/P:2,P
DE CAMP, L. SPRAGUE/CARTER, LIN  CONAN OF THE ISLES [CONAN 12]  C/P:2,P
DE CAMP, L. SPRAGUE  CONAN READER, THE [CONAN]  R/REF-H
DE CAMP, L. SPRAGUE/CARTER, LIN  CONAN THE BARBARIAN [CONAN]  C/P:2,P
DE CAMP, L. SPRAGUE/CARTER, LIN  CONAN THE BUCCANEER [CONAN 6]  C/P:2,P
DE CAMP, L. SPRAGUE/CARTER, LIN  CONAN THE LIBERATOR [BANTAM CONAN 2]  C/P:2,P

DE CAMP, L. SPRAGUE/CARTER, LIN/ NYBERG, BJORN CONAN THE SWORDSMAN [BANTAM CONAN 1]

C/P,P

DE CAMP, L. SPRAGUE  CONTINENT MAKERS AND OTHER TALES OF THE VIAGENS, THE  H:2,P
DE CAMP, L. SPRAGUE  COSMIC MANHUNT  SR-DD061

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK/ GRIFFIN, JANE WHITTINGTON DARK VALLEY DESTINY: THE LIFE OF ROBERT E. HOWARD

R/REF-H

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK DAY OF THE DINOSAUR, THE

[C/RAN-P]

DE CAMP, L. SPRAGUE  DIVIDE AND RULE  SR-H,C/P:2,P
DE CAMP, L. SPRAGUE  DRAGON OF THE ISHTAR GATE, THE  C/P:2,P
DE CAMP, L. SPRAGUE  ELEPHANT FOR ARISTOTLE, AN  P

DE CAMP, L. SPRAGUE/PRATT, FLETCHER ENCHANTER COMPLEATED, THE (ORIGINALLY WALL OF SERPENTS)

[INCOMPLETE ENCHANTER]

SR-P

DE CAMP, L. SPRAGUE/ STASHEFF, CHRISTOPHER ENCHANTER REBORN, THE [INCOMPLETE ENCHANTER]

PA,C/PA,SFWA-P

DE CAMP, L. SPRAGUE  ENGINES  [C/RAN-P]

DE CAMP, L. SPRAGUE/ STASHEFF, CHRISTOPHER EXOTIC ENCHANTER, THE [INCOMPLETE ENCHANTER]

PA

DE CAMP, L. SPRAGUE  FALLIBLE FIEND, THE  C/P:2,P
DE CAMP, L. SPRAGUE  FANTASTIC SWORDSMEN, THE  C/PA,PA
DE CAMP, L. SPRAGUE/MILLER, P. SCHUYLER GENUS HOMO  SR-H,C/P:2,P
DE CAMP, L. SPRAGUE  GLORY THAT WAS, THE  C/P,P,H
DE CAMP, L. SPRAGUE  GOBLIN TOWER, THE [RELUCTANT KING 1]  C/P:2,P
DE CAMP, L. SPRAGUE  GOLDEN WIND, THE  H,P
DE CAMP, L. SPRAGUE  GREAT FETISH, THE  C/P,SFWA,P,H
DE CAMP, L. SPRAGUE  GUN FOR DINOSAUR, A  H:2,P
DE CAMP, L. SPRAGUE  HAND OF ZEI, THE [KRISHNA]  C/P,P,H,DF249
DE CAMP, L. SPRAGUE  HONORABLE BARBARIAN, THE [RELUCTANT KING]  H:2,P
DE CAMP, L. SPRAGUE  HOSTAGE OF ZIR, THE [KRISHNA]  H:2,P
DE CAMP, L. SPRAGUE  HOSTAGE OF ZIR, THE (PROOF) [KRISHNA]  SR-LP
DE CAMP, L. SPRAGUE/PRATT, FLETCHER  INCOMPLETE ENCHANTER, THE [INCOMPLETE ENCHANTER]  SR-H,C/P:2,P

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK INCORPORATED KNIGHT, THE H,SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 106
AUTHOR  TITLE  CODE
DE CAMP, L. SPRAGUE  LEST DARKNESS FALL  GN4.24,P,H:2
DE CAMP, L. SPRAGUE/DRAKE, DAVID  LEST DARKNESS FALL/TO BRING THE LIGHT  P
DE CAMP, L. SPRAGUE  LITERARY SWORDSMEN AND SORCERERS  R/REF-H
DE CAMP, L. SPRAGUE  LOST CONTINENTS  [C/RAN-P]
DE CAMP, L. SPRAGUE  LOVECRAFT: A BIOGRAPHY  R/REF-H
DE CAMP, L. SPRAGUE  MACCHINA DEL TEMPO, LA  FP

DE CAMP, L. SPRAGUE MISCAST BARBARIAN: A BIOGRAPHY OF ROBERT E. HOWARD, THE

(1906–1936)

R/REF-H

DE CAMP, L. SPRAGUE  PHANTOMS AND FANCIES  SR-H

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK PIXILATED PEERESS, THE

C/P,P,H

DE CAMP, L. SPRAGUE  PRISONER OF ZHAMANAK, THE [KRISHNA]  H,P
DE CAMP, L. SPRAGUE  PURPLE PTERODACTYLS, THE  SR-H,C/P:2,P
DE CAMP, L. SPRAGUE  QUEEN OF ZAMBA, THE [KRISHNA]  C/P:2,P
DE CAMP, L. SPRAGUE  RELUCTANT KING, THE [RELUCTANT KING 1,2,3]  H:2
DE CAMP, L. SPRAGUE  RELUCTANT SHAMAN AND OTHER FANTASTIC TALES, THE  C/P:2,P
DE CAMP, L. SPRAGUE  RIVERS OF TIME  P
DE CAMP, L. SPRAGUE  ROGUE QUEEN  H:2,P

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK SCIENCE FICTION HANDBOOK (REVISED)

R/REF-H,C/REF-H

DE CAMP, L. SPRAGUE  SCRIBBLINGS  SR-H,SR-L
DE CAMP, L. SPRAGUE  SEARCH FOR ZEI, THE [KRISHNA]  DF249,H
DE CAMP, L. SPRAGUE  SIR HAROLD AND THE GNOME KING [INCOMPLETE ENCHANTER]  LP
DE CAMP, L. SPRAGUE  SPELL OF SEVEN, THE  PA,C/PA

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK STONES OF NOMURU, THE

C/P:2,SFWA-P,P

DE CAMP, L. SPRAGUE  SWORDS AND SORCERY  PA,C/PA

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK SWORDS OF ZINJABAN, THE [KRISHNA]

P

DE CAMP, L. SPRAGUE/PRATT, FLETCHER  TALES FROM GAVAGAN’S BAR (EXPANDED)  C/P,SR-H,P,H

DE CAMP, L. SPRAGUE/ DE CAMP, CATHERINE CROOK 3000 YEARS OF FANTASY AND SCIENCE FICTION

HA

DE CAMP, L. SPRAGUE  TOWER OF ZANID, THE [KRISHNA]  C/P:2,P
DE CAMP, L. SPRAGUE  TRITONIAN RING, THE  H:2,P
DE CAMP, L. SPRAGUE  UNBEHEADED KING, THE [RELUCTANT KING 3]  C/P,SFWA,P,H
DE CAMP, L. SPRAGUE/DRAKE, DAVID  UNDESIRED PRINCESS, THE/ENCHANTED BUNNY, THE  C/P:2,SFWA-P,P
DE CAMP, L. SPRAGUE  VENOM TREES OF SUNGA, THE  P
DE CAMP, L. SPRAGUE  VIRGIN AND THE WHEELS, THE  C/P:2,P
DE CAMP, L. SPRAGUE/PRATT, FLETCHER  WALL OF SERPENTS [INCOMPLETE ENCHANTER]  SR-H,P
DE CAMP, L. SPRAGUE  WARLOCKS AND WARRIORS  C/PA,PA

DE CAMP, L. SPRAGUE WHEELS OF IF, THE C/P,SFWA-TD20,H,P,

TD20

DE CAMP, L. SPRAGUE  YEARS IN THE MAKING  H
DE CHANCIE, JOHN  BRIDE OF THE CASTLE [CASTLE 8]  C/P,P
DE CHANCIE, JOHN  CASTLE DREAMS [CASTLE 6]  C/P:2,P
DE CHANCIE, JOHN/GREENBERG, MARTIN H. CASTLE FANTASTIC  PA
DE CHANCIE, JOHN  CASTLE FOR RENT [CASTLE 2]  C/P:2,SFWA-P,P
DE CHANCIE, JOHN  CASTLE KIDNAPPED [CASTLE 3]  C/P:2,SFWA-P,P
DE CHANCIE, JOHN  CASTLE MURDERS [CASTLE 5]  C/P:2,P
DE CHANCIE, JOHN  CASTLE PERILOUS [CASTLE 1]  C/P:2,P
DE CHANCIE, JOHN  CASTLE SPELLBOUND [CASTLE 7]  C/P,P
DE CHANCIE, JOHN  CASTLE WAR! [CASTLE 4]  C/P:2,SFWA-P,P
DE CHANCIE, JOHN/BISCHOFF, DAVID F.  DR. DIMENSION [DOCTOR DIMENSION]  C/P:2,P
DE CHANCIE, JOHN/BISCHOFF, DAVID F.  DR. DIMENSION: MASTERS OF SPACETIME [DOCTOR DIMENSION]  P
DE CHANCIE, JOHN  FROM PRUSSIA WITH LOVE [CASTLE FALKENSTEIN]  P
DE CHANCIE, JOHN  INNERVERSE  P
DE CHANCIE, JOHN  KRUTON INTERFACE, THE [U.S.S. REPULSE 1]  C/P,P
DE CHANCIE, JOHN  LIVING WITH ALIENS  C/P,P
DE CHANCIE, JOHN  MAGICNET  C/P,P,H
DE CHANCIE, JOHN  MASTERMINDS OF FALKENSTEIN [@CASTLE FALKENSTEIN]  P
DE CHANCIE, JOHN  PARADOX ALLEY [SKYWAY 3]  C/P,P
DE CHANCIE, JOHN  RED LIMIT FREEWAY [SKYWAY 2]  C/P:3,P
DE CHANCIE, JOHN  STARRIGGER [SKYWAY 1]  C/P:2,P
DE CHANCIE, JOHN  TALONS [@WITCHBLADE]  P
DE CHANCIE, JOHN/MILLER, JOHN J.  TALONS [@WITCHBLADE]/TERRIBLE BEAUTY, A [@WITCHBLADE]  P
DE CLES, JON  PARTICOLORED UNICORN, THE  P
DE CONEAU, ALEXANDER  MONK’S MAGIC  H
DE FELITTA, FRANK  AUDREY ROSE  H,P
DE FELITTA, FRANK  ENTITY, THE  P
DE FELITTA, FRANK  GOLGOTHA FALLS  H
DE GOURMONT, REMY  MR. ANTIPHILOS, SATYR  H
DE HAAN, TOM  MIRROR FOR PRINCES, A  H
DE HAVEN, TOM  END-OF-EVERYTHING MAN, THE [@KING’S TRAMP #2]  C/P,P,H
DE HAVEN, TOM  FREAKS’ AMOUR  LP

The MITSFS Pinkdex  September 24, 2007  Page 107
AUTHOR  TITLE  CODE
DE HAVEN, TOM  JOE GOSH  H
DE HAVEN, TOM  LAST HUMAN, THE [@KING’S TRAMP #3]  C/P,LP
DE HAVEN, TOM  LAST HUMAN, THE (PROOF) [KING’S TRAMP 3]  SR-LP
DE HAVEN, TOM  WALKER OF WORLDS [@KING’S TRAMP #1]  C/P:2,LP
DE KRUIF, PAUL  MICROBE HUNTERS  [C/RAN-P]
DE LA HIRE, JEAN  MYSTERE JAUNE, LE  FP
DE LA MARE, WALTER J.  BEST STORIES OF WALTER DE LA MARE  SR-H
DE LA MARE, WALTER J.  HENRY BROCKEN  SR-H
DE LA MARE, WALTER J.  RETURN, THE  LP
DE LA REE, GERRY  BOK  A
DE LA REE, GERRY  BOOK OF VIRGIL FINLAY, THE  A
DE LANCIE, JOHN/DAVID, PETER  I, Q [STAR TREK TNG]  P
DE LANCIE, JOHN/COOL, TOM  SOLDIER OF LIGHT  H,P
DE LARRABEITI, MICHAEL  ACROSS THE DARK METROPOLIS [BORRIBLES 3]  P
DE LARRABEITI, MICHAEL  BORRIBLES, THE [BORRIBLES 1]  C/P,P
DE LARRABEITI, MICHAEL  BORRIBLES GO FOR BROKE, THE [BORRIBLES 2]  P
DE LATIL, PIERRE  ZO LEVEN WIJ MORGEN  FP
DE LILLO, DON  RATNER’S STAR  H,P
DE LINT, CHARLES  ANGEL OF DARKNESS  H,LP
DE LINT, CHARLES  BLUE GIRL, THE  H
DE LINT, CHARLES  DREAMS UNDERFOOT  C/P,P,H
DE LINT, CHARLES  DRINK DOWN THE MOON [JACKY ROWAN 2]  C/P:2,SFWA-P:2,P
DE LINT, CHARLES  FAIR IN EMAIN MACHA, THE  SFWA-TD19,TD19
DE LINT, CHARLES  FORESTS OF THE HEART [NEWFORD]  H
DE LINT, CHARLES  GREENMANTLE  C/P:2,P
DE LINT, CHARLES  HARP OF THE GREY ROSE, THE  C/P:2,LP
DE LINT, CHARLES  HIDDEN CITY, THE [DUNGEON 5]  C/P:2,P
DE LINT, CHARLES  INTO THE GREEN  C/P,P,H
DE LINT, CHARLES  IVORY AND THE HORN, THE [NEWFORD]  C/P,P,H
DE LINT, CHARLES  JACK OF KINROWAN [JACKY ROWAN 1,2]  P
DE LINT, CHARLES  JACK, THE GIANT-KILLER [JACKY ROWAN 1]  C/P,SFWA-P,P,H
DE LINT, CHARLES  JACK THE GIANT KILLER (PROOF) [JACKY ROWAN 1]  SR-LP
DE LINT, CHARLES  LITTLE COUNTRY, THE  H:2,LP
DE LINT, CHARLES  MEMORY AND DREAM [NEWFORD]  H,P
DE LINT, CHARLES  MOONHEART [TAMSON HOUSE 1]  C/P:2,P
DE LINT, CHARLES  MOONLIGHT AND VINES [NEWFORD]  C/P,P,H
DE LINT, CHARLES  MULENGRO  C/P,P
DE LINT, CHARLES  ONION GIRL, THE [NEWFORD]  H,LP
DE LINT, CHARLES  RIDDLE OF THE WREN, THE  C/P,P
DE LINT, CHARLES  SOMEPLACE TO BE FLYING [NEWFORD]  C/P:2,P
DE LINT, CHARLES  SPIRITS IN THE WIRES  H
DE LINT, CHARLES  SPIRITWALK [TAMSON HOUSE 2]  C/P,P,H
DE LINT, CHARLES  SVAHA  C/P:2,SFWA-P,P
DE LINT, CHARLES  TAPPING THE DREAM TREE [NEWFORD]  H
DE LINT, CHARLES  TRADER [NEWFORD]  C/P:2,P
DE LINT, CHARLES  VALLEY OF THUNDER, THE [DUNGEON 3]  C/P:2,P
DE LINT, CHARLES  WIDDERSHINS [NEWFORD]  H
DE LINT, CHARLES  WILD WOOD, THE  LP
DE LINT, CHARLES  WOLF MOON  C/P:2,P
DE LINT, CHARLES  YARROW  C/P:2,P
DE MARINIS, RICK  CINDER  P
DE MARINO, LAWRENCE  ODYSSEY PROJECT, THE  P
DE MAUPASSANT, GUY  SELECTED TALES OF GUY DE MAUPASSANT  SR-VLH
DE MILLE, JAMES  STRANGE MANUSCRIPT FOUND IN A COPPER CYLINDER, A  H
DE PIERRES, MARIANNE  CODE NOIR  P
DE PIERRES, MARIANNE  CRASH DELUXE [PARRISH PLESSIS 3]  P
DE PIERRES, MARIANNE  NYLON ANGEL [PARRISH PLESSIS 1]  P
DE ROSA, PETER  BEST OF ALL POSSIBLE WORLDS, THE  SR-P
DE TARDE, GABRIEL  UNDERGROUND MAN  LP
DE TIMMS, GRAEME  THREE QUARTERS  SR-P
DE VET, CHARLES V./MACLEAN, KATHERINE COSMIC CHECKMATE  DF149
DE VET, CHARLES V./MACLEAN, KATHERINE SECOND GAME  C/P,P
DE VET, CHARLES V. SPECIAL FEATURE C/P,P

DE VORE, HOWARD HISTORY OF THE HUGO, NEBULA,

AND INTERNATIONAL FANTASY AWARDS, A

SRL/IND

DE VRIES, TJOMME  MENS IN HET UITDIJENDE HEELAL 1, DE  FP
DE VRIES, TJOMME  MENS IN HET UITDIJENDE HEELAL 2, DE  FP
DEAKINS, JOHN  BARROW  P
DEAN, PAMELA  DUBIOUS HILLS, THE  H,P
DEAN, PAMELA  JUNIPER, GENTIAN, AMD ROSEMARY  LP
DEAN, PAMELA C.  HIDDEN LAND, THE [SECRET COUNTRY 2]  C/P,P
DEAN, PAMELA C.  SECRET COUNTRY, THE [SECRET COUNTRY 1]  P
DEAN, PAMELA C.  TAM LIN  H,P
DEAN, PAMELA C.  WHIM OF THE DRAGON, THE [SECRET COUNTRY 3]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 108
AUTHOR  TITLE  CODE
DEANE, HAMILTON/BALDERSTON, JOHN L.  DRACULA: THE VAMPIRE PLAY  H
DEAREN, PATRICK  STARFLIGHT TO FAROUL  P
DEBIE, ERIK SCOTT  GHOSTWALKER [@FORGOTTEN REALMS FIGHTERS]  P
DEBRANDT, DON  SHAKEDOWN [@ANGEL]  P
DEBRANDT, DON  V.I.  C/P,P
DEBRANDT, DON H.  QUICKSILVER SCREEN, THE  P
DEBRANDT, DON H.  STEELDRIVER  P
DEBRANDT, DON H.  TIMBERJAK  P
DECANDIDO, KEITH R. A.  ART OF THE IMPOSSIBLE, THE [@STAR TREK LOST ERA (2328-46 AD)]  P
DECANDIDO, KEITH R. A.  ARTICLES OF THE FEDERATION [STAR TREK]  P
DECANDIDO, KEITH R. A.  BLACKOUT [@BUFFY]  P
DECANDIDO, KEITH R. A.  BRAVE AND THE BOLD, BOOK ONE [@STAR TREK]  P
DECANDIDO, KEITH R. A.  BRAVE AND THE BOLD, BOOK TWO [@STAR TREK]  P
DECANDIDO, KEITH R. A.  CYCLE OF HATRED [WORLD OF WARCRAFT]  P
DECANDIDO, KEITH R. A.  DEATHLESS, THE [@BUFFY]  P

DECANDIDO, KEITH R. A.

DEMONS OF AIR AND DARKNESS [STAR TREK GATEWAYS #4/ STAR TREK DS9]

P

DECANDIDO, KEITH R. A.  DESTRUCTION OF ILLUSIONS [@ANDROMEDA (GENE RODDENBERRY’S) #1]  P
DECANDIDO, KEITH R. A.  DIPLOMATIC IMPLAUSIBILITY [STAR TREK TNG]  P
DECANDIDO, KEITH R. A.  DRAGON PRECINCT  P
DECANDIDO, KEITH R. A.  ENEMY TERRITORY [@STAR TREK I.K.S. GORKON 3]  P

DECANDIDO, KEITH R. A./ GEORGE, DAVID R. III FERENGINAR: SATISFACTION IS NOT GUARANTEED

[STAR TREK DS9 WORLDS 3]/

DOMINION, THE: OLYMPUS DESCENDING [STAR TREK DS9 WORLDS 3]

C/P,P

DECANDIDO, KEITH R. A.  GOOD DAY TO DIE, A [@STAR TREK I.K.S. GORKON 1]  P

DECANDIDO, KEITH R. A./DILMORE, KEVIN/ GOLDEN, CHRISTIE/SMITH, DEAN WESLEY/ WARD, DAYTON HAVE TECH, WILL TRAVEL [@STAR TREK S.C.E. #1]

P

DECANDIDO, KEITH R. A.  HONOR BOUND [@STAR TREK I.K.S. GORKON 2]  P
DECANDIDO, KEITH R. A.  HOUSE OF CARDS [@FARSCAPE]  C/P,P
DECANDIDO, KEITH R. A.  IMAGININGS  LPA

DECANDIDO, KEITH R. A./DILMORE, KEVIN/ MACK, DAVID/WARD, DAYTON MIRACLE WORKERS [@STAR TREK S.C.E. #2]

P

DECANDIDO, KEITH R. A.  NOVA [@STARCRAFT GHOST]  P
DECANDIDO, KEITH R. A.  SERENITY  P
DECANDIDO, KEITH R. A.  TALES FROM THE CAPTAIN’S TABLE [STAR TREK]  LPA
DECANDIDO, KEITH R. A.  TALES OF THE DOMINION WAR [@STAR TREK]  LPA
DECANDIDO, KEITH R. A.  TIME FOR WAR, A TIME FOR PEACE, A [STAR TREK TNG TIME TO 9]  P

DECANDIDO, KEITH R. A./MACK, DAVID/ YORK, J. STEVEN/YORK, CHRISTINA F.

WILDFIRE [@STAR TREK S.C.E. #6]

P

DECANDIDO, KEITH R. A.  XANDER YEARS, THE, VOL. 1 [BUFFY]  P
DECARLO, ELISA  DEVIL YOU SAY, THE  C/P,P
DECARLO, ELISA  STRONG SPIRITS  P

DECARNIN, CAMILLA/GARBER, ERIC/ PALEO, LYN WORLDS APART

LPA

DEDMAN, STEPHEN  ART OF ARROW CUTTING, THE [MAGE AND TAKUMO 1]  LP
DEDMAN, STEPHEN  FISTFUL OF DATA, A [@SHADOWRUN]  P
DEDMAN, STEPHEN  FOREIGN BODIES  H
DEDMAN, STEPHEN  SHADOWS BITE [MAGE AND TAKUMO 2]  H,LP
DEE, ROGER  EARTH GONE MAD, AN  DD084
DEE, RON  BLOOD  P
DEE, RON  BLOOD LUST  P
DEEGAN, JON J.  AMATEURS IN ALCHEMY  SR-P
DEEGAN, JON J.  ANTRO, THE LIFE-GIVER  SR-P,H
DEEGAN, JON J.  EXILES IN TIME  SR-P
DEEGAN, JON J.  GREAT ONES, THE  H
DEEGAN, JON J.  UNDERWORLD OF ZELLO  SR-P
DEFILIPPIS  FOREKNOWLEDGE [QUANTUM LEAP]  P
DEFONTENAY, C. I.  STAR (PSI CASSIOPEIA)  P
DEFORD, MIRIAM ALLEN  ELSEWHERE, ELSEWHEN, ELSEHOW  H
DEFORD, MIRIAM ALLEN  SPACE, TIME, AND CRIME  PA,C/PA
DEFORD, MIRIAM ALLEN  XENOGENESIS  C/P:2,P
DEGRAEFF, ALLEN  HUMAN AND OTHER BEINGS  PA
DEIGHTON, LEN  BILLION DOLLAR BRAIN, THE  P
DEIGHTON, LEN  SS-GB  H,SFWA,P
DEITZ, TOM  ABOVE THE LOWER SKY  C/P,P,H
DEITZ, TOM  BLOODWINTER [ERON 1]  LP
DEITZ, TOM  DARKTHUNDER’S WAY [OTHERWORLD 3]  C/P:2,P
DEITZ, TOM  DEMONS IN THE GREEN, THE  P
DEITZ, TOM  DREAMBUILDER [SOULSMITH 2]  P,SFWA-P
DEITZ, TOM  DREAMSEEKER’S ROAD [OTHERWORLD 6]  H,P
DEITZ, TOM  FIRESHAPER’S DOOM [OTHERWORLD 2]  C/P:2,P
DEITZ, TOM  GHOSTCOUNTRY’S WRATH  P
DEITZ, TOM  GRYPHON KING, THE [OTHERWORLD]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 109
AUTHOR  TITLE  CODE
DEITZ, TOM  LANDSLAYER’S LAW [OTHERWORLD]  P
DEITZ, TOM  SOULSMITH [SOULSMITH 1]  P
DEITZ, TOM  SPRINGWAR [ERON 2]  LP
DEITZ, TOM  STONESKIN’S REVENGE [OTHERWORLD 5]  C/P,P
DEITZ, TOM  SUMMERBLOOD [ERON 3]  P
DEITZ, TOM  SUNSHAKER’S WAR [OTHERWORLD 4]  C/P:2,P
DEITZ, TOM  WARAUTUMN [ERON 4]  P
DEITZ, TOM  WARSTALKER’S TRACK  P
DEITZ, TOM  WINDMASTER’S BANE [OTHERWORLD 1]  C/P:2,P
DEITZ, TOM  WORDWRIGHT [SOULSMITH 3]  P,SFWA-P
DEL FRANCO, MARK  UNSHAPELY THINGS [CONNOR GREY 1]  P
DEL MARTIA, ASTRON  DAWN OF DARKNESS  SR-P
DEL MARTIA, ASTRON  INTERSTELLAR ESPIONAGE  SR-P
DEL MARTIA, ASTRON  ONE AGAINST TIME  C/P:2,P
DEL MARTIA, ASTRON  SPACE PIRATES  SR-P
DEL REY, JUDY-LYNN  STELLAR #1  PA,C/PA
DEL REY, JUDY-LYNN  STELLAR #2  PA,C/PA
DEL REY, JUDY-LYNN  STELLAR #3  C/PA,PA
DEL REY, JUDY-LYNN  STELLAR #4  C/PA,PA
DEL REY, JUDY-LYNN  STELLAR #5  PA,C/PA,SFWA-P
DEL REY, JUDY-LYNN  STELLAR #6  PA
DEL REY, JUDY-LYNN  STELLAR #7  PA,C/PA
DEL REY, JUDY-LYNN  STELLAR SHORT NOVELS  PA,C/PA
DEL REY, LESTER  ALARM IN DE ATOOMCENTRALE  FP
DEL REY, LESTER  "...AND SOME WERE HUMAN"  C/P:2,P
DEL REY, LESTER  ATTACK FROM ATLANTIS  H:2,P
DEL REY, LESTER  BADGE OF INFAMY  D76960
DEL REY, LESTER  BEST OF LESTER DEL REY, THE  C/P,P,H
DEL REY, LESTER  BEST SCIENCE FICTION STORIES OF THE YEAR 1  HA,PA
DEL REY, LESTER  BEST SCIENCE FICTION STORIES OF THE YEAR 2  HA,PA
DEL REY, LESTER  BEST SCIENCE FICTION STORIES OF THE YEAR 3  HA,PA
DEL REY, LESTER  BEST SCIENCE FICTION STORIES OF THE YEAR 4  C/PA,PA
DEL REY, LESTER  BEST SCIENCE FICTION STORIES OF THE YEAR 5  C/PA,PA
DEL REY, LESTER  CAVE OF SPEARS, THE  H:2
DEL REY, LESTER  DAY OF THE GIANTS  C/P,P,H
DEL REY, LESTER  DAY THE SUN STOOD STILL, THE  PA,C/PA
DEL REY, LESTER  EARLY DEL REY  H:2,SFWA
DEL REY, LESTER  EARLY DEL REY 1, THE  C/P,P
DEL REY, LESTER  EARLY DEL REY 2, THE  C/P:2,P
DEL REY, LESTER  ELEVENTH COMMANDMENT, THE  C/P:2,P
DEL REY, LESTER  FANTASTIC SCIENCE FICTION ART 1926–1954  A
DEL REY, LESTER  GODS AND GOLEMS  C/P:2,P
DEL REY, LESTER  MAN WITHOUT A PLANET, THE  P
DEL REY, LESTER  MAROONED ON MARS  H:2,P
DEL REY, LESTER  MOON OF MUTINY  H:2,P
DEL REY, LESTER  MORTALS AND MONSTERS  C/P:2,P
DEL REY, LESTER  MYSTERIOUS PLANET, THE  C/P,P
DEL REY, LESTER  NERVES  C/P:2,P
DEL REY, LESTER/KESSLER, RISA  ONCE UPON A TIME  SR-HA,HA
DEL REY, LESTER  OUTPOST OF JUPITER  C/P,P,H
DEL REY, LESTER/VAN LHIN, ERIK  POLICE YOUR PLANET  C/P:2,P
DEL REY, LESTER  PRISONERS OF SPACE  H
DEL REY, LESTER  PSTALEMATE  C/P:2,P,H
DEL REY, LESTER  ROBOTS AND CHANGELINGS  C/P:2,P
DEL REY, LESTER  ROCKET JOCKEY  C/P:2,P
DEL REY, LESTER  RUNAWAY ROBOT, THE  LP
DEL REY, LESTER  SCHEME OF THINGS, THE  P
DEL REY, LESTER  SIEGE PERILOUS  C/P:2,P
DEL REY, LESTER  SKY IS FALLING, THE  D76960
DEL REY, LESTER  SKY IS FALLING, THE/BADGE OF INFAMY  SR-P,C/P:2
DEL REY, LESTER  SPACE FLIGHT  [C/RAN-H]
DEL REY, LESTER  STEP TO THE STARS  H:2,P
DEL REY, LESTER  TUNNEL THROUGH TIME  C/P,LP,H

DEL REY, LESTER WORLD OF SCIENCE FICTION 1926–1976: THE HISTORY OF A SUBCULTURE, THE

R/REF-H,C/REF-H

DEL REY, LESTER/MATSCHAT, CECILE/ CARMER, CARL YEAR AFTER TOMORROW, THE

HA

DELACORTE, PETER  TIME ON MY HANDS  H
DELACORTE PRESS  PATH INTO THE UNKNOWN: THE BEST OF SOVIET SCIENCE FICTION  HA

DELACROIX, CLAIRE/KURLAND, LYNN/ MCKILLIP, PATRICIA A./SHINN, SHARON TO WEAVE A WEB OF MAGIC

LP

DELANEY, JOSEPH H.  IN THE FACE OF MY ENEMY  C/P:2,P
DELANEY, JOSEPH H.  LORDS TEMPORAL  C/P,P
DELANEY, JOSEPH H./STIEGLER, MARC  VALENTINA: SOUL IN SAPPHIRE  C/P,P
DELANY, SAMUEL R.  ATLANTIS: THREE TALES  H

AUTHOR  TITLE  CODE
DELANY, SAMUEL R.  BABEL-17  C/P:2,P
DELANY, SAMUEL R.  BABEL-17 (ITALIAN)  FP
DELANY, SAMUEL R.  BALLAD OF BETA-2, THE  DM121,C/P:2,P
DELANY, SAMUEL R.  BALLAD OF BETA-2, THE/EMPIRE STAR  C/P:2,P
DELANY, SAMUEL R.  BALLADE DE BETA 2, LA  FP
DELANY, SAMUEL R.  BRIDGE OF LOST DESIRE, THE  C/P,P,H
DELANY, SAMUEL R.  CAPTIVES OF THE FLAME  DF199
DELANY, SAMUEL R.  CITY OF A THOUSAND SUNS [FALL OF THE TOWERS 3]  C/P:2,P
DELANY, SAMUEL R.  DHALGREN  C/P,P,H
DELANY, SAMUEL R.  DISTANT STARS  H,LP
DELANY, SAMUEL R.  DRIFTGLASS  H:2,P
DELANY, SAMUEL R.  EINSTEIN INTERSECTION, THE  C/P,P,H
DELANY, SAMUEL R.  EMPIRE STAR  DM139
DELANY, SAMUEL R.  FALL OF THE TOWERS, THE [FALL OF THE TOWERS 1,2,3]  C/P:2,P
DELANY, SAMUEL R.  FLIGHT FROM NEVERYON  P
DELANY, SAMUEL R.  HEAVENLY BREAKFAST: AN ESSAY ON THE WINTER OF LOVE  R/REF-P
DELANY, SAMUEL R.  JEWEL-HINGED JAW, THE  R/REF-P,C/REF-H:2
DELANY, SAMUEL R.  JEWELS OF APTOR, THE  DF173
DELANY, SAMUEL R.  JEWELS OF APTOR, THE (RESTORED)  C/P:2,P
DELANY, SAMUEL R.  MOTION OF LIGHT IN WATER, THE  R/REF-H
DELANY, SAMUEL R.  NEBULA WINNERS 13  C/PA,PA,SFWA
DELANY, SAMUEL R.  NEVERYONA  H:2,LP
DELANY, SAMUEL R.  NOVA  H:2,P
DELANY, SAMUEL R.  OUT OF THE DEAD CITY [FALL OF THE TOWERS 1]  P
DELANY, SAMUEL R./HACKER, MARILYN  QUARK 1  PA,C/PA
DELANY, SAMUEL R./HACKER, MARILYN  QUARK 2  C/PA,PA
DELANY, SAMUEL R./HACKER, MARILYN  QUARK 3  C/PA,PA
DELANY, SAMUEL R./HACKER, MARILYN  QUARK 4  PA,C/PA
DELANY, SAMUEL R.  SILENT INTERVIEWS  R/REF-H
DELANY, SAMUEL R.  STAR PIT, THE  SFWA-TD4,TD4
DELANY, SAMUEL R.  STARS IN MY POCKET LIKE GRAINS OF SAND  C/P,P,H
DELANY, SAMUEL R.  TALES OF NEVERYON  C/P:2,P
DELANY, SAMUEL R.  THEY FLY AT CIRON  P
DELANY, SAMUEL R.  TIDES OF LUST, THE  SR-L
DELANY, SAMUEL R.  TOWERS OF TORON, THE [FALL OF THE TOWERS 2]  C/P,DF261,P
DELANY, SAMUEL R.  TRITON  C/P:2,P
DELANY, SAMUEL R.  TROUBLE ON TRITON  LP
DELANY, SAMUEL R.  WE, IN SOME STRANGE POWER’S EMPLOY, MOVE ON A RIGOROUS LINE  TD21,SFWA-TD21
DELL BOOKS  PATH INTO THE UNKNOWN: THE BEST OF SOVIET SCIENCE FICTION  PA,C/PA
DELL BOOKS  THREE FOR TOMORROW  PA,C/PA
DELL BOOKS  THREE TRIPS IN TIME AND SPACE  PA,C/PA
DELL, JEFFREY  NEWS FOR HEAVEN  SR-P
DELRIO, MARTIN  MIDNIGHT JUSTICE [SPIDER-MAN]  C/P,P
DELRIO, MARTIN  MORTAL KOMBAT [MORTAL KOMBAT 1]  P

DELRIO, MARTIN SERVICE FOR THE DEAD

[@MECHWARRIOR DARK AGE PROVING GROUNDS 3/BATTLETECH]

P

DELRIO, MARTIN SILENCE IN THE HEAVENS, A

[@MECHWARRIOR DARK AGE PROVING GROUNDS 1/BATTLETECH]

P

DELRIO, MARTIN TRUTH AND SHADOWS

[@MECHWARRIOR DARK AGE PROVING GROUNDS 2/BATTLETECH]

P

DEMARTINO, DENNY  HEART OF STONE [@ASTROLOGER 1]  P
DEMARTINO, DENNY  WAYWARD MOON [@ASTROLOGER 2]  P
DEMPSEY, AL  MISS FINNEY KILLS NOW AND THEN  SFWA-P
DEMPSEY, AL  WHAT LAW THERE WAS  SFWA
DENAERDE, STEFAN  OPERATION SURVIVAL EARTH  P
DENGELEGI, PAUL  CASCA THE DEFIANT [CASCA]  P
DENGELEGI, PAUL  CASCA THE LIBERATOR [CASCA]  P
DENGROVE, RICHARD A.  BIBLIOGRAPH OF A. E. VAN VOGT  SRL/IND
DENHAM, ALICE  GHOST AND MRS. MUIR, THE  P
DENIRO, ALAN  SKINNY DIPPING IN THE LAKE OF THE DEAD  LP
DENNING, TROY  AMBER ENCHANTRESS, THE [@DARK SUN PRISM #3]  C/P,P
DENNING, TROY  BEYOND THE HIGH ROAD [@FORGOTTEN REALMS CORMYR 2]  P
DENNING, TROY  CERULEAN STORM, THE [@DARK SUN PRISM #5]  C/P,P
DENNING, TROY  CRIMSON LEGION, THE [@DARK SUN PRISM #2]  C/P,P

DENNING, TROY CRUCIBLE: THE TRIAL OF CYRIC THE MAD [@FORGOTTEN REALMS AVATAR 5]

P

DENNING, TROY  DARK NEST [@STAR WARS DARK NEST 1,2,3]  H
DENNING, TROY  DRAGONWALL [FORGOTTEN REALMS EMPIRES 2]  C/P:2,P
DENNING, TROY  FACES OF DECEPTION [@FORGOTTEN REALMS LOST EMPIRES]  P
DENNING, TROY  GIANT AMONG US, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #2]  P
DENNING, TROY  OATH OF STONEKEEP, THE  P
DENNING, TROY  OBSIDIAN ORACLE, THE [@DARK SUN PRISM #4]  P
DENNING, TROY  OGRE’S PACT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #1]  P
DENNING, TROY  PARCHED SEA, THE [@FORGOTTEN REALMS HARPERS #1]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 111
AUTHOR  TITLE  CODE
DENNING, TROY  SIEGE, THE [@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #2]  P

DENNING, TROY SORCERER, THE

[@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #3]

P

DENNING, TROY  STAR BY STAR [@STAR WARS NEW JEDI ORDER]  P

DENNING, TROY SUMMONING, THE

[@FORGOTTEN REALMS RETURN OF THE ARCHWIZARDS #1]

P

DENNING, TROY  TATOOINE GHOST [@STAR WARS]  P
DENNING, TROY  TEMPEST [@STAR WARS LEGACY OF THE FORCE 3]  P
DENNING, TROY  TITAN OF TWILIGHT, THE [@FORGOTTEN REALMS TWILIGHT GIANTS #3]  P
DENNING, TROY  VEILED DRAGON, THE [@FORGOTTEN REALMS HARPERS #12]  P
DENNING, TROY  VERDANT PASSAGE, THE [@DARK SUN PRISM #1]  C/P,P
DENNING, TROY  WATERDEEP [@FORGOTTEN REALMS AVATAR #3]  P
DENNIS, CAROL L.  DRAGON’S KNIGHT  P
DENNIS, CAROL L.  DRAGON’S PAWN  C/P,P
DENNIS, CAROL L.  DRAGON’S QUEEN  P
DENT, LESTER  DOC SAVAGE FILES, THE [DOC SAVAGE]  R/REF-P
DENT, ROXANNE  WHITE FOG, THE  P
DENTON, BRADLEY  BUDDY HOLLY IS ALIVE AND WELL ON GANYMEDE  C/P,P
DENTON, BRADLEY  CALVIN COOLIDGE HOME FOR DEAD COMEDIANS, THE  LP
DENTON, BRADLEY  CONFLAGRATION ARTIST, A  LP
DENTON, BRADLEY  LUNATICS  LP
DENTON, BRADLEY  WRACK AND ROLL  C/P,P
DERESKE, JO  LONE SENTINEL, THE  SFWA
DERLETH, AUGUST  BEACHHEADS IN SPACE  C/PA,PA
DERLETH, AUGUST  BEYOND TIME AND SPACE  PA,C/PA
DERLETH, AUGUST  COUNTRY MATTERS  LP
DERLETH, AUGUST  CTHULHU MYTHOS, THE  H
DERLETH, AUGUST  DARK MIND, DARK HEART  SR-PA,C/PA
DERLETH, AUGUST  DARK OF THE MOON  HA
DERLETH, AUGUST  DARK THINGS  SR-HA
DERLETH, AUGUST  HARRIGAN’S FILE  SR-H
DERLETH, AUGUST  MASK OF CTHULHU, THE  C/P:2,P
DERLETH, AUGUST  MR. FAIRLIE’S FINAL JOURNEY  SR-H
DERLETH, AUGUST  NIGHT’S YAWNING PEAL  PA
DERLETH, AUGUST  NOT LONG FOR THIS WORLD  H,P
DERLETH, AUGUST  OTHER SIDE OF THE MOON, THE  PA,C/PA
DERLETH, AUGUST  OUTER REACHES, THE  HA,PA
DERLETH, AUGUST  OVER THE EDGE  SR-PA
DERLETH, AUGUST  PORTALS OF TOMORROW  HA
DERLETH, AUGUST  RETURN TO SAC PRAIRIE  LP
DERLETH, AUGUST  SLEEP NO MORE  HA
DERLETH, AUGUST  SOMEONE IN THE DARK  C/P,P
DERLETH, AUGUST  SOMETHING NEAR  SR-H
DERLETH, AUGUST  STRANGE PORTS OF CALL  C/PA,PA
DERLETH, AUGUST  TALES OF THE CTHULHU MYTHOS 1  PA,C/PA
DERLETH, AUGUST  TALES OF THE CTHULHU MYTHOS 2  PA,C/PA
DERLETH, AUGUST  TIME TO COME  PA,C/PA
DERLETH, AUGUST  TRAIL OF CTHULHU, THE  P
DERLETH, AUGUST  WHO KNOCKS?  SR-HA
DERLETH, AUGUST  WORLDS OF TOMORROW  PA,C/PA
DERMOTT, VERN  PLANET FINDERS  C/P,P,H
DESJARDIN, MARIE  FOR THE TIME BEING  LP
DESLEAL, ALVARO MENEN  ILUSTRE FAMILIA ANDROIDE, LA  FP
DESMOND, HUGH  TERRIBLE AWAKENING, THE  H
DESMOND, WILLIAM H.  SCIENCE-FICTION MAGAZINE CHECKLIST 1961–1972, THE  R/IND:3
DESROCHERS, DIANE  WALKER BETWEEN THE WORLDS  C/P,P
DETWILLER, DENNIS  DELTA GREEN: DENIED TO THE ENEMY [DELTA GREEN]  LP
DETZER, DIANE  PLANET OF FEAR  H
DEUTCHMAN, DEBORAH  SIGNALS  H
DEUTSCHMAN, DEBORAH  SIGNALS  P
DEVENPORT, EMILY  EGGHEADS  C/P,P
DEVENPORT, EMILY  GODHEADS  C/P,P
DEVENPORT, EMILY  KRONOS CONDITION, THE  P
DEVENPORT, EMILY  LARISSA  C/P,P
DEVENPORT, EMILY  SCORPIANNE  C/P,P
DEVENPORT, EMILY  SHADE  P
DEVER, JOE  CAPTIVES OF RAAG, THE [LONE WOLF 14]  P
DEVER, JOE  CASTLE DEATH [LONE WOLF 7]  P
DEVER, JOE  CAULDRON OF FEAR, THE [LONE WOLF 9]  P
DEVER, JOE/CHALK, GARY  CAVERNS OF KALTE, THE [LONE WOLF 3]  C/P,P
DEVER, JOE  CHASM OF DOOM, THE [LONE WOLF 4]  P
DEVER, JOE/GRANT, JOHN  DARK DOOR OPENS, THE [@LEGENDS OF LONE WOLF #2]  P
DEVER, JOE  DARKE CRUSADE, THE [LONE WOLF 15]  P
DEVER, JOE  DAWN OF THE DRAGONS [LONE WOLF]  P

The MITSFS Pinkdex  September 24, 2007  Page 112
AUTHOR  TITLE  CODE
DEVER, JOE  DEATHLORD OF IXIA, THE [LONE WOLF 17]  P
DEVER, JOE  DUNGEONS OF TORGAR, THE [LONE WOLF 10]  P
DEVER, JOE/GRANT, JOHN  ECLIPSE OF THE KAI [@LEGENDS OF LONE WOLF #1]  P
DEVER, JOE  FIRE ON THE WATER [LONE WOLF 2]  P
DEVER, JOE/CHALK, GARY  FLIGHT FROM THE DARK [LONE WOLF 1]  P
DEVER, JOE  JUNGLE OF HORRORS, THE [LONE WOLF 8]  P
DEVER, JOE  KINGDOMS OF TERROR, THE [LONE WOLF 6]  P
DEVER, JOE  LEGACY OF VASHNA, THE [LONE WOLF 16]  P
DEVER, JOE  MASTERS OF DARKNESS, THE [LONE WOLF 12]  C/P,P
DEVER, JOE  PLAGUE LORDS OF RUEL, THE [LONE WOLF 13]  P
DEVER, JOE  PRISONERS OF TIME, THE [LONE WOLF 11]  P
DEVER, JOE  SHADOW ON THE SAND [LONE WOLF 5]  P
DEVER, JOE/GRANT, JOHN  TIDES OF TREACHERY, THE [@LEGENDS OF LONE WOLF #3]  P
DEVEREAUX, ROBERT  SANTA STEPS OUT  P
DEVIN, FLANNA  ALIEN ENCOUNTER  P

DEVLIN, DEAN/EMMERICH, ROLAND/ MOLSTAD, STEPHEN INDEPENDENCE DAY [@ID4]

P

DEVLIN, DEAN/EMMERICH, ROLAND  STARGATE [STARGATE]  P
DEVORE, HOWARD  HUGO, NEBULA AND WORLD FANTASY AWARDS, THE  R/REF-P
DEVOS, ELISABETH  SERAPHIM RISING, THE  P
DEWDNEY, A. K.  PLANIVERSE, THE  LP
DEWEESE, GENE  CHAIN OF ATTACK [STAR TREK]  C/P:3,P
DEWEESE, GENE/COULSON, ROBERT  CHARLES FORT NEVER MENTIONED WOMBATS  H,SFWA
DEWEESE, GENE  ENGINES OF DESTINY [@STAR TREK]  P
DEWEESE, GENE  FINAL NEXUS, THE [STAR TREK]  C/P:2,P
DEWEESE, GENE  INTO THE NEBULA [STAR TREK TNG]  C/P:2,P
DEWEESE, GENE  JEREMY CASE  SR-P,C/P
DEWEESE, GENE  KING OF THE DEAD [@RAVENLOFT]  P
DEWEESE, GENE  LORD OF THE NECROPOLIS [RAVENLOFT]  P
DEWEESE, GENE/COULSON, ROBERT  NOW YOU SEE IT/HIM/THEM...  H,SFWA
DEWEESE, GENE  PEACEKEEPERS, THE [STAR TREK TNG]  C/P:3,P
DEWEESE, GENE  RENEGADE [STAR TREK]  C/P:2,P
DEWEESE, GENE  VAULT, THE [LOST IN SPACE]  P
DEWEESE, GENE  WANTING FACTOR, THE  P
DEWEESE, JEAN  DOLL WITH OPAL EYES, THE  P
DEWOLF, PARKER  LEFT HAND OF DEATH, THE [@EBERRON LANTERNLIGHT #1]  P
DEXTER, JOHN  SINNERS OF HWANG, THE  SR-L
DEXTER, SUSAN  MOUNTAINS OF CHANNADRAN, THE [WINTER KING’S WAR 3]  C/P,P
DEXTER, SUSAN  PRINCE OF ILL LUCK, THE [WARHORSE OF ESDRAGON 1]  P
DEXTER, SUSAN  RING OF ALLAIRE, THE [WINTER KING’S WAR 1]  C/P:2,P
DEXTER, SUSAN  SWORD OF CALANDRA, THE [WINTER KING’S WAR 2]  C/P,P
DEXTER, SUSAN  TRUE KNIGHT, THE [WARHORSE OF ESDRAGON 3]  P
DEXTER, SUSAN  WIND-WITCH, THE [WARHORSE OF ESDRAGON 2]  C/P:2,P
DEXTER, SUSAN  WIZARD’S SHADOW, THE  P
DEXTER, WILLIAM  CHILDREN OF THE VOID  C/P:2,P
DEXTER, WILLIAM  WORLD IN ECLIPSE  C/P:2,P
DEYO, DAVID D. JR.  ALL THE DEVILS ARE HERE  SR-LPA
DI FILIPPO, PAUL  BABYLON SISTERS  LP
DI FILIPPO, PAUL  EMPEROR OF GONDWANALAND, THE  LP
DI FILIPPO, PAUL  FRACTAL PAISLEYS  H
DI FILIPPO, PAUL  LOST PAGES  LP
DI FILIPPO, PAUL  MOUTHFUL OF TONGUES, A  H
DI FILIPPO, PAUL  NEUTRINO DRAG  H,LP
DI FILIPPO, PAUL  RIBOFUNK  P
DI FILIPPO, PAUL  SHUTEYE FOR THE TIMEBROKER  LP
DI FILIPPO, PAUL  STEAMPUNK TRILOGY, THE  H
DI FILIPPO, PAUL  STRANGE TRADES  H
DIAMOND, GRAHAM  BEASTS OF HADES, THE [EMPIRE PRINCESS 4]  P
DIAMOND, GRAHAM  DUNGEONS OF KUBA [EMPIRE PRINCESS 2]  P
DIAMOND, GRAHAM  FALCON OF EDEN, THE [EMPIRE PRINCESS 3]  P
DIAMOND, GRAHAM  HAVEN, THE [EMPIRE PRINCESS 0]  P
DIAMOND, GRAHAM  LADY OF THE HAVEN [EMPIRE PRINCESS 1]  P
DIAMOND, GRAHAM  SAMARKAND  P
DIAMOND, GRAHAM  SAMARKAND DAWN  P
DIAMOND, GRAHAM  THIEF OF KALIMAR, THE  C/P:2,P
DIAMOND, GRAHAM  THIEF OF KALIMAR, THE (PROOF)  SR-LP
DIAZ-PLAJA, FERNANDO  DESFILE DE LA VICTORIA, EL  FP
DIBBLE, BIRNEY  PAN  P
DIBDIN, MICHAEL  LAST SHERLOCK HOLMES STORY, THE  P
DIBELL, ANSEN  CIRCLE, CRESCENT, STAR [KANTMORIE 2]  C/P:2,P
DIBELL, ANSEN  PURSUIT OF THE SCREAMER [KANTMORIE 1]  C/P:2,P
DIBELL, ANSEN  SUMMERFAIR [KANTMORIE 3]  P
DICHIARA, ROBERT  DEAL WITH THE DEVIL  P
DICK, PHILIP K.  BEST OF PHILIP K. DICK, THE  P

AUTHOR  TITLE  CODE
DICK, PHILIP K.  BLADE RUNNER  C/P:2,P
DICK, PHILIP K.  BOOK OF PHILIP K. DICK, THE  C/P,P
DICK, PHILIP K.  CLANS OF THE ALPHANE MOON  C/P:2,P
DICK, PHILIP K.  COLLECTED STORIES OF PHILIP K. DICK VOLUME 1, THE  LP
DICK, PHILIP K.  COLLECTED STORIES OF PHILIP K. DICK VOLUME 2, THE  LP
DICK, PHILIP K.  COLLECTED STORIES OF PHILIP K. DICK VOLUME 3, THE  LP
DICK, PHILIP K.  COLLECTED STORIES OF PHILIP K. DICK VOLUME 5, THE  LP
DICK, PHILIP K.  CONFESSIONS OF A CRAP ARTIST  P
DICK, PHILIP K.  COSMIC PUPPETS, THE  DD249
DICK, PHILIP K.  COUNTER-CLOCK WORLD  C/P:2,P
DICK, PHILIP K.  CRACK IN SPACE, THE  C/P:2,P
DICK, PHILIP K./ZELAZNY, ROGER  DEUS IRAE  C/P,SFWA,P,H
DICK, PHILIP K.  DIVINE INVASION, THE  C/P,P,H
DICK, PHILIP K.  DO ANDROIDS DREAM OF ELECTRIC SHEEP?  C/P:3,P
DICK, PHILIP K.  DR. BLOODMONEY, OR HOW WE GOT ALONG AFTER THE BOMB  C/P:2,P
DICK, PHILIP K.  DR. FUTURITY  DD421,D15697
DICK, PHILIP K.  EYE IN THE SKY  C/P:2,P
DICK, PHILIP K.  FLOW MY TEARS, THE POLICEMAN SAID  C/P,P,H,SFWA:2
DICK, PHILIP K.  GALACTIC POT-HEALER  H:2,P
DICK, PHILIP K.  GAME-PLAYERS OF TITAN, THE  C/P,P,H
DICK, PHILIP K./NELSON, RAY FARADAY  GANYMEDE TAKEOVER, THE  C/P:2,P
DICK, PHILIP K.  GOLDEN MAN, THE  C/P:2,P,H
DICK, PHILIP K.  I HOPE I SHALL ARRIVE SOON  H,SFWA,P
DICK, PHILIP K.  LIES, INC.  LP
DICK, PHILIP K.  MAN IN THE HIGH CASTLE, THE  C/P,P,H:2
DICK, PHILIP K.  MAN WHO JAPED, THE  DD193
DICK, PHILIP K.  MARTIAN TIME-SLIP  C/P:2,P
DICK, PHILIP K.  MARY AND THE GIANT  H:2
DICK, PHILIP K.  MAZE OF DEATH, A  C/P,P,H
DICK, PHILIP K.  MINORITY REPORT, THE [COLLECTED STORIES OF PHILIP K. DICK 4]  LP
DICK, PHILIP K.  MR. LARS, SOGNATORE D’ARMI  FP
DICK, PHILIP K.  NOW WAIT FOR LAST YEAR  H:2,P
DICK, PHILIP K.  OUR FRIENDS FROM FROLIX 8  H:2,P
DICK, PHILIP K.  PAYCHECK  LP
DICK, PHILIP K.  PENULTIMATE TRUTH, THE  C/P:2,P
DICK, PHILIP K.  PRESERVING MACHINE, THE  C/P,P,H
DICK, PHILIP K.  RADIO FREE ALBEMUTH  H:2,P
DICK, PHILIP K.  SCANNER DARKLY, A  C/P,H:3,LP,SFWA
DICK, PHILIP K.  SECOND VARIETY  B
DICK, PHILIP K.  SIMULACRA, THE  C/P:2,P
DICK, PHILIP K.  SOLAR LOTTERY  C/P,DD103,SFWA-P,P
DICK, PHILIP K.  THREE STIGMATA OF PALMER ELDRITCH, THE  H:2,P
DICK, PHILIP K.  TIME OUT OF JOINT  C/P:2,P
DICK, PHILIP K.  TRANSMIGRATION OF TIMOTHY ARCHER, THE  H,SFWA,P
DICK, PHILIP K.  UBIK  H:3,P
DICK, PHILIP K.  UNTELEPORTED MAN, THE  C/P:2,DG602,D15697,P
DICK, PHILIP K.  VALIS  C/P:2,P
DICK, PHILIP K.  VARIABLE MAN AND OTHER STORIES, THE  C/P:2,P
DICK, PHILIP K.  VULCAN’S HAMMER  DD457
DICK, PHILIP K.  WE CAN BUILD YOU  C/P:2,P
DICK, PHILIP K.  WORLD JONES MADE, THE  C/P:2,P,DD150
DICK, PHILIP K.  ZAP GUN, THE  C/P:2,P
DICK, R. A.  GHOST AND MRS. MUIR, THE  H
DICKENS, MONICA  MESSENGER, THE  P
DICKINSON, CHARLES  SHORTCUT IN TIME, A  LP
DICKINSON, PETER  BLUE HAWK, THE  C/P:2,P
DICKINSON, PETER  BOX OF NOTHING, A  H,P
DICKINSON, PETER  CHANGES, THE [CHANGES 1,2,3]  H,LP
DICKINSON, PETER  DEVIL’S CHILDREN, THE [CHANGES 1]  H
DICKINSON, PETER  GLASS-SIDED ANTS’ NEST, THE  C/P,P
DICKINSON, PETER  GREEN GENE, THE  C/P,P,H
DICKINSON, PETER  HEARTSEASE [CHANGES 2]  H:2
DICKINSON, PETER  POISON ORACLE, THE  SFWA
DICKINSON, PETER  TULKU  P
DICKINSON, PETER  WEATHERMONGER, THE [CHANGES 3]  C/P,P,H
DICKINSON, WILLIAM GROFT  BORROBIL  P
DICKS, TERRANCE  ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  H

DICKS, TERRANCE ADVENTURES OF K9 AND OTHER MECHANICAL CREATURES

[DOCTOR WHO]

P

DICKS, TERRANCE  AMBASSADORS OF DEATH, THE [DOCTOR WHO]  SR-P
DICKS, TERRANCE  ARC OF INFINITY [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  BLOOD HARVEST [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
DICKS, TERRANCE  CATASTROPHEA [DOCTOR WHO PAST DOCTORS]  C/P,P
DICKS, TERRANCE  CAVES OF ANDROZANI, THE [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE/LETTS, BARRY  DEADLY REUNION [DOCTOR WHO PAST DOCTORS]  P

AUTHOR  TITLE  CODE
DICKS, TERRANCE  DEATH TO THE DALEKS [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND AN UNEARTHLY CHILD [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE ABOMINABLE SNOWMEN [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE ANDROID INVASION [DOCTOR WHO]  C/P,P
DICKS, TERRANCE  DOCTOR WHO AND THE ANDROIDS OF TARA [DOCTOR WHO]  SR-P,P
DICKS, TERRANCE  DOCTOR WHO AND THE ARMAGEDDON FACTOR [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE AUTON INVASION [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE BRAIN OF MORBIUS [DOCTOR WHO]  SR-P,H,P
DICKS, TERRANCE  DOCTOR WHO AND THE CARNIVAL OF MONSTERS [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE CLAWS OF AXOS [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE DALEK INVASION OF EARTH [DOCTOR WHO]  SR-P,C/P:3
DICKS, TERRANCE  DOCTOR WHO AND THE DAY OF THE DALEKS [DOCTOR WHO]  C/P:2,P
DICKS, TERRANCE  DOCTOR WHO AND THE DEADLY ASSASSIN [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE DESTINY OF THE DALEKS [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE FACE OF EVIL [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE GENESIS OF THE DALEKS [DOCTOR WHO]  C/P:2,P
DICKS, TERRANCE  DOCTOR WHO AND THE GIANT ROBOT [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE HAND OF FEAR [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE HORNS OF NIMON [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE HORROR OF FANG ROCK [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE IMAGE OF THE FENDAHL [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE INVASION OF TIME [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE INVISIBLE ENEMY [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE KEEPER OF TRAKEN [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE LOCH NESS MONSTER [DOCTOR WHO]  C/P,P
DICKS, TERRANCE  DOCTOR WHO AND THE MONSTER OF PELADON [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE MUTANTS [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE NIGHTMARE OF EDEN [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE PLANET OF EVIL [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE PLANET OF THE DALEKS [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE PLANET OF THE SPIDERS [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE POWER OF KROLL [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE PYRAMIDS OF MARS [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE REVENGE OF THE CYBERMEN [DOCTOR WHO]  C/P:2,P
DICKS, TERRANCE  DOCTOR WHO AND THE ROBOTS OF DEATH [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE STATE OF DECAY [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE STONES OF BLOOD [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE SUNMAKERS [DOCTOR WHO]  SR-P
DICKS, TERRANCE  DOCTOR WHO AND THE TALONS OF WENG-CHIANG [DOCTOR WHO]  C/P:2,P
DICKS, TERRANCE  DOCTOR WHO AND THE TERROR OF THE AUTONS [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE THREE DOCTORS [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  DOCTOR WHO AND THE TIME MONSTER, THE [DOCTOR WHO]  P
DICKS, TERRANCE  DOCTOR WHO AND THE TIME WARRIOR [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE UNDERWORLD [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO AND THE WEB OF FEAR [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  DOCTOR WHO: THE WHEEL IN SPACE [DOCTOR WHO]  P
DICKS, TERRANCE  EIGHT DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
DICKS, TERRANCE  ENDGAME [DOCTOR WHO EIGHTH DOCTOR]  P
DICKS, TERRANCE  FACELESS ONES, THE [@DOCTOR WHO]  P
DICKS, TERRANCE  FIVE DOCTORS, THE [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  FOUR TO DOOMSDAY [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  FURTHER ADVENTURES OF DOCTOR WHO, THE [DOCTOR WHO]  H
DICKS, TERRANCE  INFERNO [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  KINDA [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  KROTONS, THE [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE/HULKE, MALCOLM  MAKING OF DOCTOR WHO, THE  R/REF-P,C/REF-P:2
DICKS, TERRANCE  MEAN STREETS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  P
DICKS, TERRANCE  MEGLOS [DOCTOR WHO]  SR-P
DICKS, TERRANCE  MIND OF EVIL, THE [DOCTOR WHO]  SR-P,P
DICKS, TERRANCE  MYSTERIOUS PLANET, THE [DOCTOR WHO]  P
DICKS, TERRANCE  PLAYERS [DOCTOR WHO PAST DOCTORS]  C/P,P
DICKS, TERRANCE  SEEDS OF DEATH, THE [DOCTOR WHO]  SR-P
DICKS, TERRANCE  SHAKEDOWN [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
DICKS, TERRANCE  SMUGGLERS, THE [DOCTOR WHO]  SR-P,C/P
DICKS, TERRANCE  SNAKEDANCE [DOCTOR WHO]  SR-P
DICKS, TERRANCE  SPACE PIRATES, THE [DOCTOR WHO]  P
DICKS, TERRANCE  TIMEWYRM: EXODUS [DOCTOR WHO NEW ADVENTURES]  C/P,P
DICKS, TERRANCE  WARMONGER [DOCTOR WHO PAST DOCTORS]  P
DICKS, TERRANCE  WARRIORS OF THE DEEP [DOCTOR WHO]  SR-P,C/P:2
DICKS, TERRANCE  WEB OF FEAR, THE [DOCTOR WHO]  P
DICKS, TERRANCE  WORLD GAME [DOCTOR WHO PAST DOCTORS]  P
DICKSON, GORDON R.  ALIEN ART  H:2,P
DICKSON, GORDON R.  ALIEN ART/ARCTURUS LANDING  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 115
AUTHOR  TITLE  CODE
DICKSON, GORDON R.  ALIEN FROM ARCTURUS  DD139
DICKSON, GORDON R.  ALIEN WAY, THE  C/P:2,SFWA-TD31,P
DICKSON, GORDON R.  ANCIENT, MY ENEMY  H:2,P,SFWA
DICKSON, GORDON R./WIXON, DAVID W.  ANTAGONIST [CHILDE CYCLE]  H
DICKSON, GORDON R.  ARCTURUS LANDING  C/P:2,SFWA-P,P
DICKSON, GORDON R.  BEGINNINGS  C/P:2,SFWA-P,P
DICKSON, GORDON R.  BEYOND THE DAR AL-HARB  C/P:2,SFWA-P,P
DICKSON, GORDON R.  BLOOD AND WAR [@HARRIERS #2]  C/PA,PA
DICKSON, GORDON R.  BOOK OF GORDON DICKSON, THE  C/P:2,P
DICKSON, GORDON R.  CHANTRY GUILD, THE [CHILDE CYCLE]  C/P,SFWA-P,P,H
DICKSON, GORDON R.  COMBAT SF  C/PA,PA,SFWA
DICKSON, GORDON R.  DELUSION WORLD  DF119
DICKSON, GORDON R.  DICKSON!  SR-H
DICKSON, GORDON R.  DORSAI! [CHILDE CYCLE]  C/P:2,P
DICKSON, GORDON R.  DORSAI COMPANION, THE [CHILDE CYCLE]  H,LP

DICKSON, GORDON R./DENNING, TROY/ GLABERSON, CORY DORSAI’S COMMAND [@COMBAT COMMAND CHILDE CYCLE]

P

DICKSON, GORDON R.  DRAGON AND THE DJINN, THE [DRAGON KNIGHT]  C/P,P
DICKSON, GORDON R.  DRAGON AND THE FAIR MAID OF KENT, THE [DRAGON KNIGHT]  H
DICKSON, GORDON R.  DRAGON AND THE GEORGE, THE [DRAGON KNIGHT 1]  C/P,P,H:2
DICKSON, GORDON R.  DRAGON AND THE GNARLY KING, THE [DRAGON KNIGHT]  C/P,P,H
DICKSON, GORDON R.  DRAGON AT WAR, THE [DRAGON KNIGHT 4]  H,SFWA,SFWA-P:2,P
DICKSON, GORDON R.  DRAGON IN LYONESSE, THE [DRAGON KNIGHT]  H
DICKSON, GORDON R.  DRAGON KNIGHT, THE [DRAGON KNIGHT 2]  H:2,SFWA,P
DICKSON, GORDON R.  DRAGON ON THE BORDER, THE [DRAGON KNIGHT 3]  H:2,SFWA-P:2,P
DICKSON, GORDON R.  DRAGON, THE EARL, AND THE TROLL, THE [DRAGON KNIGHT 5]  H,P
DICKSON, GORDON R.  EARTH LORDS, THE  C/P:2,SFWA-P,P
DICKSON, GORDON R.  ENDS  C/P:2,SFWA-P,P
DICKSON, GORDON R.  FAR CALL, THE  H:2,SFWA-P,P
DICKSON, GORDON R.  FINAL ENCYCLOPEDIA, THE [CHILDE CYCLE]  C/P,P,H
DICKSON, GORDON R.  FOREVER MAN, THE  C/P,SFWA-P,P,H
DICKSON, GORDON R.  FORWARD!  C/P:2,P
DICKSON, GORDON R.  GENETIC GENERAL, THE [CHILDE CYCLE]  C/P,DD449,P
DICKSON, GORDON R.  GORDON R. DICKSON’S SF BEST  P
DICKSON, GORDON R./BOVA, BEN  GREMLINS GO HOME  H
DICKSON, GORDON R.  GUIDED TOUR  P,SFWA-P
DICKSON, GORDON R.  HOME FROM THE SHORE  C/P:2,LP
DICKSON, GORDON R./BOVA, BEN  HOUR OF THE GREMLINS  P
DICKSON, GORDON R.  HOUR OF THE HORDE  C/P,P,H
DICKSON, GORDON R.  HUMAN EDGE, THE  P
DICKSON, GORDON R.  IN IRON YEARS  C/P,P,H,SFWA
DICKSON, GORDON R.  IN THE BONE  P,SFWA-P
DICKSON, GORDON R.  INVADERS!  C/P,P
DICKSON, GORDON R./GREEN, ROLAND  JAMIE THE RED  C/P:2,P
DICKSON, GORDON R.  LAST DREAM, THE  C/P,P
DICKSON, GORDON R.  LAST MASTER, THE  C/P:2,P
DICKSON, GORDON R./HARRISON, HARRY  LIFESHIP, THE  C/P:2
DICKSON, GORDON R.  LOST DORSAI [CHILDE CYCLE]  C/P,LP,H
DICKSON, GORDON R.  LOVE NOT HUMAN  C/P:2,SFWA-P,P
DICKSON, GORDON R.  L’UOMO ETERNO (ITALIAN)  FP
DICKSON, GORDON R.  MAGNIFICENT WILF, THE  H,P
DICKSON, GORDON R.  MAN FROM EARTH, THE  C/P:2,P
DICKSON, GORDON R.  MAN THE WORLDS REJECTED, THE  C/P:2,SFWA-P,P
DICKSON, GORDON R.  MANKIND ON THE RUN  DD164
DICKSON, GORDON R.  MASTERS OF EVERON  C/P,H,LP
DICKSON, GORDON R.  MINDSPAN  C/P:2,P
DICKSON, GORDON R.  MISSION TO UNIVERSE  C/P:2,SFWA-P,P
DICKSON, GORDON R.  MISSION TO UNIVERSE (REVISED)  C/P:2,P
DICKSON, GORDON R.  MUTANTS  C/P:2,P
DICKSON, GORDON R.  NAKED TO THE STARS  C/P:2,SFWA-TD31,P
DICKSON, GORDON R.  NEBULA WINNERS 12  C/PA,PA,SFWA
DICKSON, GORDON R.  NECROMANCER [CHILDE CYCLE]  C/P,P,H
DICKSON, GORDON R./BOVA, BEN  NO MAS DUENDES  FP
DICKSON, GORDON R.  NO ROOM FOR MAN [CHILDE CYCLE]  C/P:2,P
DICKSON, GORDON R.  NONE BUT MAN  C/P,P,H
DICKSON, GORDON R.  NONE BUT MAN (WITH ADDITIONAL STORY)  P
DICKSON, GORDON R.  OF WAR AND HONOR [@HARRIERS #1]  PA,C/PA,SFWA-P
DICKSON, GORDON R.  ON THE RUN  C/P:2,SFWA-P,P
DICKSON, GORDON R.  OTHER [CHILDE CYCLE]  C/P,P,H
DICKSON, GORDON R.  OUTPOSTER, THE  C/P,SFWA-P:2,P,H
DICKSON, GORDON R.  PRITCHER MASS, THE  C/P,SFWA-P,P,H,SFWA
DICKSON, GORDON R.  PRO  C/P:2,P
DICKSON, GORDON R.  R-MASTER, THE  H:2,P
DICKSON, GORDON R.  SECRET UNDER THE SEA  LP

AUTHOR  TITLE  CODE
DICKSON, GORDON R.  SECRETS OF THE DEEP  P
DICKSON, GORDON R.  SLEEPWALKER’S WORLD  C/P,P,H
DICKSON, GORDON R.  SOLDIER, ASK NOT [CHILDE CYCLE]  C/P:2,P
DICKSON, GORDON R.  SPACE DOGFIGHTS  C/PA,PA
DICKSON, GORDON R.  SPACE SWIMMERS, THE  C/P:2,P
DICKSON, GORDON R.  SPACE WINNERS  C/P,SFWA-P,P,H
DICKSON, GORDON R.  SPACEPAW  C/P,SFWA-P,P,H
DICKSON, GORDON R.  SPACIAL DELIVERY  DF119,C/P,SFWA-P
DICKSON, GORDON R.  SPIRIT OF DORSAI, THE [CHILDE CYCLE]  C/P,LP,H
DICKSON, GORDON R.  STAR ROAD, THE  H:2,SFWA,P
DICKSON, GORDON R.  STEEL BROTHER  C/P:2,SFWA-P,P
DICKSON, GORDON R.  STRANGER, THE  C/P,SFWA-P,P
DICKSON, GORDON R.  SURVIVAL!  C/P:2,P
DICKSON, GORDON R.  TACTICS OF MISTAKE [CHILDE CYCLE]  C/P,P,H,SFWA
DICKSON, GORDON R.  THREE TO DORSAI [CHILDE CYCLE]  H:2
DICKSON, GORDON R.  TIME STORM  SFWA-P,H:2,P
DICKSON, GORDON R.  TIME TO TELEPORT  C/P,DD449
DICKSON, GORDON R.  TIME TO TELEPORT/DELUSION WORLD  P
DICKSON, GORDON R.  WAY OF THE PILGRIM  H:2,P,SFWA
DICKSON, GORDON R.  WOLF AND IRON  H:2,P
DICKSON, GORDON R.  WOLFLING  C/P,P,H
DICKSON, GORDON R.  YOUNG BLEYS [CHILDE CYCLE]  H:2,P
DIETRICH, WILLIAM  GETTING BACK  H
DIETZ, WILLIAM C.  ALIEN BOUNTY [SAM MCCADE 3]  C/P,P
DIETZ, WILLIAM C.  BODYGUARD  C/P:2,P
DIETZ, WILLIAM C.  BY BLOOD ALONE [LEGION OF THE DAMNED 3]  C/P,P
DIETZ, WILLIAM C.  BY FORCE OF ARMS [LEGION OF THE DAMNED 4]  P
DIETZ, WILLIAM C./DRAKE, DAVID  CLUSTER COMMAND [CRISIS OF EMPIRE 2]  P,SFWA-P
DIETZ, WILLIAM C.  DEATHDAY [DEATHDAY 1]  C/P,P
DIETZ, WILLIAM C.  DRIFTER [DRIFTER 1]  C/P:2,SFWA-P:2,P
DIETZ, WILLIAM C.  DRIFTER’S RUN [DRIFTER 2]  C/P,P
DIETZ, WILLIAM C.  DRIFTER’S WAR [DRIFTER 3]  C/P:2,SFWA-P,P
DIETZ, WILLIAM C.  EARTHRISE [DEATHDAY 2]  P
DIETZ, WILLIAM C.  FINAL BATTLE, THE [LEGION OF THE DAMNED 2]  C/P,P
DIETZ, WILLIAM C.  FOR MORE THAN GLORY [LEGION OF THE DAMNED 5]  H,P
DIETZ, WILLIAM C.  FOR THOSE WHO FELL [LEGION OF THE DAMNED 6]  H
DIETZ, WILLIAM C.  FREEHOLD  P
DIETZ, WILLIAM C.  GALACTIC BOUNTY [SAM MCCADE 1]  P
DIETZ, WILLIAM C.  HALO: THE FLOOD [@HALO]  P
DIETZ, WILLIAM C.  IMPERIAL BOUNTY [SAM MCCADE 2]  P
DIETZ, WILLIAM C.  LEGION OF THE DAMNED [LEGION OF THE DAMNED 1]  C/P:2,SFWA-P,P
DIETZ, WILLIAM C.  LOGOS RUN [JAK REBO 2]  H
DIETZ, WILLIAM C.  MARS PRIME  P
DIETZ, WILLIAM C.  MATRIX MAN  C/P:2,P
DIETZ, WILLIAM C.  MCCADE ON THE RUN [SAM MCCADE 3,4]  LP
DIETZ, WILLIAM C.  MCCADE’S BOUNTY [SAM MCCADE 4]  P
DIETZ, WILLIAM C.  PRISON PLANET  C/P,P
DIETZ, WILLIAM C.  REBEL AGENT [@STAR WARS DARK FORCES 2]  H
DIETZ, WILLIAM C.  RUNNER [JAK REBO 1]  P
DIETZ, WILLIAM C.  SOLDIER FOR THE EMPIRE [@STAR WARS DARK FORCES 1]  H
DIETZ, WILLIAM C.  STEELHEART  P
DIETZ, WILLIAM C.  WHERE THE SHIPS DIE  C/P,P
D’IGNAZIO, FRED  ESCAPE FROM ROBOTROPOLIS [ROBOT ODYSSEY 1]  H:2
DIKTY, T. E.  BEST SCIENCE FICTION STORIES AND NOVELS 1956, THE  HA
DIKTY, T. E.  BEST SCIENCE FICTION STORIES AND NOVELS, NINTH SERIES, THE  HA
DIKTY, T. E.  5 TALES FROM TOMORROW  PA,C/PA
DIKTY, T. E.  6 FROM WORLDS BEYOND  PA,C/PA
DILLARD, J. M.  BLOODTHIRST [STAR TREK]  C/P:2,P
DILLARD, J. M.  DEMONS [STAR TREK]  C/P:3,P
DILLARD, J. M.  EMISSARY [STAR TREK DS9]  C/P:2,P
DILLARD, J. M.  EMISSARY (2 CASSETTES) [STAR TREK DS9]  SR-MM
DILLARD, J. M.  ENTERPRISE: THE EXPANSE [STAR TREK ENTERPRISE]  LP
DILLARD, J. M.  LOST YEARS, THE [STAR TREK]  C/P,P,H:2
DILLARD, J. M.  MINDSHADOW [STAR TREK]  C/P:2,P
DILLARD, J. M./O’MALLEY, KATHLEEN  POSSESSION [STAR TREK TNG]  C/P:2,P
DILLARD, J. M.  RECOVERY [STAR TREK]  C/P,P
DILLARD, J. M.  RESURRECTION, THE [WAR OF THE WORLDS 1]  C/P,P
DILLARD, J. M.  STAR TREK FEDERATION PASSPORT [STAR TREK]  B
DILLARD, J. M.  STAR TREK: FIRST CONTACT [STAR TREK TNG]  H:2
DILLARD, J. M.  STAR TREK GENERATIONS [STAR TREK]  H
DILLARD, J. M.  STAR TREK INSURRECTION [STAR TREK TNG]  H
DILLARD, J. M.  STAR TREK: NEMESIS [STAR TREK TNG]  H
DILLARD, J. M.  STAR TREK V: THE FINAL FRONTIER [STAR TREK]  C/P:2,P
DILLARD, J. M.  STAR TREK VI: THE UNDISCOVERED COUNTRY [STAR TREK]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 117
AUTHOR  TITLE  CODE
DILLARD, J. M.  SURAK’S SOUL [STAR TREK ENTERPRISE]  P

DILLE, ROBERT C.

COLLECTED WORKS OF BUCK ROGERS IN THE 25TH CENTURY, THE [BUCK ROGERS]

CX

DILLON, PETER C.  ACE ALPHA-NUMERIC ANNOTATED SCIENCE FICTION CHECKLIST, AN  R/IND
DIMARCO, JENNIFER  ESCAPE TO THE WIND  LP
DINESEN, ISAK  SEVEN GOTHIC TALES  P
DINESEN, ISAK  WINTER’S TALES  H
DINGWALL, ERIC J./LANGDON-DAVIES, JOHN UNKNOWN — IS IT NEARER?, THE  [C/RAN-P]
DIONE, R. L.  GOD DRIVES A FLYING SAUCER  [C/RAN-P]
DIPPER, ALAN  GOLDEN VIRGIN, THE  P
DISCH, THOMAS M.  BAD MOON RISING  HA,SFWA
DISCH, THOMAS M.  BRAVE LITTLE TOASTER GOES TO MARS, THE  H
DISCH, THOMAS M.  BUSINESSMAN, THE  H,SFWA
DISCH, THOMAS M.  CAMP CONCENTRATION  C/P,P,H
DISCH, THOMAS M.  CAMPO ARCHIMEDE  FP
DISCH, THOMAS M.  ECHO ROUND HIS BONES  C/P:2,P
DISCH, THOMAS M.  FUN WITH YOUR NEW HEAD  H:2,SFWA,P
DISCH, THOMAS M.  FUNDAMENTAL DISCH  C/P:2,P
DISCH, THOMAS M.  GENOCIDES, THE  C/P:2,P
DISCH, THOMAS M.  GETTING INTO DEATH  H
DISCH, THOMAS M.  GETTING INTO DEATH AND OTHER STORIES  C/P,P
DISCH, THOMAS M.  MAN WHO HAD NO IDEA, THE  C/P:2,P
DISCH, THOMAS M.  MANKIND UNDER THE LEASH  DG597
DISCH, THOMAS M.  M.D., THE  C/P,P
DISCH, THOMAS M./NAYLOR, CHARLES  NEW CONSTELLATIONS  HA,SFWA
DISCH, THOMAS M.  NEW IMPROVED SUN, THE  SFWA
DISCH, THOMAS M.  ON WINGS OF SONG  C/P,P,H
DISCH, THOMAS M.  ONE HUNDRED AND TWO H-BOMBS  C/P:2,P
DISCH, THOMAS M.  PRISONER, THE [PRISONER]  C/P,P
DISCH, THOMAS M.  RUINS OF EARTH, THE  PA,C/PA
DISCH, THOMAS M./NAYLOR, CHARLES  STRANGENESS  C/PA,PA
DISCH, THOMAS M.  334  C/P:2,P
DISCH, THOMAS M.  TRIPLICITY  H:2
DISILVESTRO, ROGER L.  LIVING WITH THE REPTILES  H
DISILVESTRO, ROGER L.  URSULA’S GIFT  H
DITCHBURN, ELIZABETH/GROVE, HEATHER  RISEN, THE [WORLD OF DARKNESS WRAITH]  SR-VLH
DIVONO, SHARMAN  BLOOD MOON  P
DIXON, DOUGAL  AFTER MAN  A
DIXON, DOUGAL  NEW DINOSAURS, THE  A
DIXON, JAMES  IT LIVES AGAIN!  P
DIXON, RICHARD  DESTINATION: AMALTHEIA  SR-PA,C/PA
DIXON, ROGER  NOAH II  C/P,P,H
DIXON, THOMAS  FALL OF A NATION, THE  H
DOBYNS, STEPHEN  WRESTLER’S CRUEL STUDY, THE  LP
DOCKERY, KEVIN/NILES, DOUGLAS  TASK FORCE MARS [@STARSTRIKE]  P
DOCTOROW, CORY  DOWN AND OUT IN THE MAGIC KINGDOM  H
DOCTOROW, CORY  EASTERN STANDARD TRIBE  H
DOCTOROW, CORY  OVERCLOCKED  LP
DOCTOROW, CORY  PLACE SO FOREIGN, A  H,LP
DOCTOROW, CORY  SOMEONE COMES TO TOWN, SOMEONE LEAVES TOWN  H
DODDERIDGE, ESME  NEW GULLIVER, THE  LP
DODGE, MICHAEL J.  STAR TREK: VOYAGE TO ADVENTURE  P
DOGBOLT, BARNABY  EVE’S SECOND APPLE  H
DOHERTY, ROBERT  AREA 51 [AREA 51 1]  P
DOHERTY, ROBERT  AREA 51: EXCALIBUR [AREA 51 6]  P
DOHERTY, ROBERT  AREA 51: LEGEND [AREA 51 9]  C/P,P
DOHERTY, ROBERT  AREA 51: NOSFERATU [@AREA 51 8]  P
DOHERTY, ROBERT  AREA 51: NOSFERATU (PROOF) [AREA 51 8]  SR-P
DOHERTY, ROBERT  AREA 51: THE GRAIL [AREA 51 5]  P
DOHERTY, ROBERT  AREA 51: THE MISSION [AREA 51 3]  P
DOHERTY, ROBERT  AREA 51: THE REPLY [AREA 51 2]  P
DOHERTY, ROBERT  AREA 51: THE SPHINX [AREA 51 4]  P
DOHERTY, ROBERT  AREA 51: THE TRUTH [@AREA 51 7]  P
DOHERTY, ROBERT  PROJECT AURA [PSYCHIC WARRIOR]  P
DOHERTY, ROBERT  ROCK, THE  P
DOKEY, CAMERON  HAUNTED BY DESIRE [@CHARMED]  P
DOKEY, CAMERON  HERE BE MONSTERS [BUFFY]  P
DOKEY, CAMERON  PICTURE PERFECT [@CHARMED]  P
DOKEY, CAMERON  SUMMONED, THE [@ANGEL]  P
DOKEY, CAMERON  TRUTH AND CONSEQUENCES [@CHARMED]  P
DOLAN, BILL  AFRIKORPS [@AFRIKORPS #1]  P
DOLAN, BILL  IRON HORSE [@AFRIKORPS #2]  P
DOLAN, BILL  LION MOUNTAIN [@AFRIKORPS #5]  P
DOLAN, BILL  SEA STALLION [@AFRIKORPS #4]  P

The MITSFS Pinkdex  September 24, 2007  Page 118
AUTHOR  TITLE  CODE
DOLAN, MIKE  SANTANA MORNING  SR-P,C/P
DOLBIER, MAURICE  MAGIC SHOP, THE  H
DOLINSKY, MIKE  MIND ONE  P
DOLLEY, CHRIS  RESONANCE  H
DOMATILLA, JOHN  LAST CRIME, THE  H,SFWA
DOMINIK, HANS  LAND AUS FEUER UND WASSER  FP
DONALDSON, STEPHEN R.  CHAOS AND ORDER [GAP 4]  H:2,P
DONALDSON, STEPHEN R.  DARK AND HUNGRY GOD ARISES, A [GAP 3]  H:2,@P
DONALDSON, STEPHEN R.  DAUGHTER OF REGALS AND OTHER TALES  C/P,P,H
DONALDSON, STEPHEN R.  FORBIDDEN KNOWLEDGE [GAP 2]  H:2,P
DONALDSON, STEPHEN R.  GILDEN-FIRE  H:2,SR-H
DONALDSON, STEPHEN R.  ILLEARTH WAR, THE [THOMAS COVENANT 2]  C/P,H,LP
DONALDSON, STEPHEN R.  LORD FOUL’S BANE [THOMAS COVENANT 1]  C/P,H,LP
DONALDSON, STEPHEN R.  MAN RIDES THROUGH, A [MORDANT’S NEED 2]  H:2,P
DONALDSON, STEPHEN R.  MIRROR OF HER DREAMS, THE [MORDANT’S NEED 1]  H:2,P,SFWA
DONALDSON, STEPHEN R.  ONE TREE, THE [THOMAS COVENANT 5]  C/P,P,H
DONALDSON, STEPHEN R.  POWER THAT PRESERVES, THE [THOMAS COVENANT 3]  C/P,LP,H:2
DONALDSON, STEPHEN R.  REAL STORY, THE [GAP 1]  H:2,P
DONALDSON, STEPHEN R.  REAVE THE JUST AND OTHER TALES  H
DONALDSON, STEPHEN R.  RUNES OF THE EARTH, THE [LAST THOMAS COVENANT 1]  H
DONALDSON, STEPHEN R.  STRANGE DREAMS  HA,LPA
DONALDSON, STEPHEN R.  THIS DAY ALL GODS DIE [GAP 5]  H:2,P
DONALDSON, STEPHEN R.  WHITE GOLD WIELDER [THOMAS COVENANT 6]  C/P,P,H
DONALDSON, STEPHEN R.  WOUNDED LAND, THE [THOMAS COVENANT 4]  C/P,P,H
DONEGAN, GREG  ASSAULT ON ATLANTIS [ATLANTIS 5]  P
DONEGAN, GREG  ATLANTIS [ATLANTIS 1]  C/P,P
DONEGAN, GREG  ATLANTIS: DEVIL’S SEA [ATLANTIS 3]  P
DONEGAN, GREG  BATTLE FOR ATLANTIS [ATLANTIS 6]  P
DONIS, MILES  FALL OF NEW YORK, THE  P
DONNELLY, MARCOS  PROPHETS FOR THE END OF TIME  C/P,P
DONOHUE, KEITH  STOLEN CHILD, THE  LP
DONOHUE, TREVOR  SAVAGE TOMORROW  SR-P,C/P
DONOVAN, DALE/JOHNS, LINDA  SIGN OF THE SHAPESHIFTER [@KNIGHTS OF THE SILVER DRAGON #3]  LP
DOOHAN, JAMES/STIRLING, S. M.  INDEPENDENT COMMAND, THE [FLIGHT ENGINEER 3]  H,P
DOOHAN, JAMES/STIRLING, S. M.  PRIVATEER, THE [FLIGHT ENGINEER 2]  C/P,P
DOOHAN, JAMES/STIRLING, S. M.  RISING, THE [FLIGHT ENGINEER 1]  P
DORER, NANCY/DORER, FRANCES  BY DAYBREAK THE EAGLE  P
DORER, NANCY/DORER, FRANCES  RETURN OF THE EAGLE  P
DORER, NANCY/DORER, FRANCES  TWO CAME CALLING  P
DORER, NANCY/DORER, FRANCES  WHERE NO MAN HAS TROD  P
DORER, NANCY/DORER, FRANCES  WINGS OF THE EAGLE, THE  P
DORMAN, SONYA  PLANET PATROL  H
DORN, FRANK  APPOINTMENT WITH YESTERDAY  P
DORN, FRANK  SUNWATCH  P
DORN, FRANK  WHEN NEXT I WAKE  P
DORN, MICHAEL  TIME BLENDER  C/P,P
DORRE, PAMELA  WIND OVER STONEHENGE  SR-P
DORRINGTON, ALBERT  RADIUM TERRORS, THE  SR-H
DORSEY, CANDAS JANE  BLACK WINE  H,LP
DORSEY, CANDAS JANE  MACHINE SEX  LP
DORSEY, CANDAS JANE  PARADIGM OF EARTH, A  LP
DORSEY, CANDAS JANE/TRUSCOTT, GERRY  TESSERACTS 3  C/PA,PA
DOUBLEDAY BOOKS  BEST-IN-BOOKS (CONTAINS "DANDELION WINE" BY RAY BRADBURY)  HA
DOUBLEDAY BOOKS  BEST-IN-BOOKS (CONTAINS "THE FAR TRAVELLER" BY MANNING COLES)  HA
DOUBLEDAY BOOKS  FIVE FATES  HA
DOUBLEDAY BOOKS  KILLERBOWL T-SHIRT  SFWA-VLP
DOUBLEDAY BOOKS  NEW VISIONS: A COLLECTION OF MODERN SCIENCE FICTION ART, THE  SFWA
DOUBLEDAY BOOKS  PRELUDE TO FOUNDATION PAPER BLOCK  SFWA-VLP
DOUBLEDAY BOOKS  SCIENCE FICTION CALENDAR 1980  SFWA-VLP
DOUBLEDAY BOOKS  THREE STORIES  HA
DOUGLAS, CAROLE NELSON  COUNTERPROBE [PROBE 2]  H:2,SFWA-P,P
DOUGLAS, CAROLE NELSON  CUP OF CLAY [TALISWOMAN 1]  H:2,P
DOUGLAS, CAROLE NELSON  EXILES OF THE RYNTH  C/P:2,P
DOUGLAS, CAROLE NELSON  HEIR OF RENGARTH [SWORD AND CIRCLET 2]  H:2,SFWA-P,P
DOUGLAS, CAROLE NELSON  KEEPERS OF EDANVANT [SWORD AND CIRCLET 1]  H:2,SFWA-P,P

DOUGLAS, CAROLE NELSON PROBE [PROBE 1] C/P,SFWA-P:2,H,LP,

SFWA

DOUGLAS, CAROLE NELSON  SEED UPON THE WIND [TALISWOMAN 2]  H:2
DOUGLAS, CAROLE NELSON  SEVEN OF SWORDS [SWORD AND CIRCLET 3]  C/P,SFWA-P,P,H
DOUGLAS, CAROLE NELSON  SIX OF SWORDS  C/P:2,P
DOUGLAS, CHARLOTTE/KEARNEY, SUSAN  BATTLE OF BETAZED, THE [STAR TREK TNG]  C/P,P
DOUGLAS, IAN  BATTLESPACE [@LEGACY #2]  P
DOUGLAS, IAN  EUROPA STRIKE [@HERITAGE #3]  C/P,P
DOUGLAS, IAN  LUNA MARINE [@HERITAGE #2]  C/P,P
DOUGLAS, IAN  SEMPER MARS [@HERITAGE #1]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 119
AUTHOR  TITLE  CODE
DOUGLAS, IAN  STAR CORPS [@LEGACY #1]  P
DOUGLAS, IAN  STAR MARINES [@LEGACY #3]  P
DOUGLAS, L. WARREN  BRIGHT ISLANDS IN A DARK SEA  P
DOUGLAS, L. WARREN  CANNON’S ORB  P
DOUGLAS, L. WARREN  GLAICE [ARBITER 1]  P
DOUGLAS, L. WARREN  ISLE BEYOND TIME, THE  P
DOUGLAS, L. WARREN  PLAGUE OF CHANGE, A  P
DOUGLAS, L. WARREN  SACRED POOL, THE [SORCERER’S TALE 1]  H
DOUGLAS, L. WARREN  SIMPLY HUMAN  P
DOUGLAS, L. WARREN  STEPWATER [ARBITER]  P
DOUGLAS, L. WARREN  VEIL OF YEARS, THE [SORCERER’S TALE 2]  H,P
DOUGLAS, L. WARREN  WHEELS OF PHYRE, THE [ARBITER 2]  P
DOUGLAS, NORMAN  IN THE BEGINNING  H
DOUGLASS, SARA  BEYOND THE HANGING WALL  H
DOUGLASS, SARA  CRUSADER [WAYFARER REDEMPTION 6]  H
DOUGLASS, SARA  DARKWITCH RISING [TROY GAME 3]  H
DOUGLASS, SARA  DRUID’S SWORD [TROY GAME 4]  H
DOUGLASS, SARA  ENCHANTER [WAYFARER REDEMPTION 2]  H
DOUGLASS, SARA  GODS’ CONCUBINE [@TROY GAME #2]  P
DOUGLASS, SARA  HADES’ DAUGHTER [TROY GAME #1]  H:2,@P
DOUGLASS, SARA  NAMELESS DAY, THE [CRUCIBLE 1]  P
DOUGLASS, SARA  PILGRIM [WAYFARER REDEMPTION 5]  H
DOUGLASS, SARA  SERPENT BRIDE, THE [DARKGLASS MOUNTAIN 1]  H
DOUGLASS, SARA  SERPENT BRIDE, THE (PROOF) [DARKGLASS MOUNTAIN 1]  SR-LP
DOUGLASS, SARA  SINNER [WAYFARER REDEMPTION 4]  P
DOUGLASS, SARA  STARMAN [WAYFARER REDEMPTION 3]  H:2
DOUGLASS, SARA  THRESHOLD  H
DOUGLASS, SARA  WAYFARER REDEMPTION, THE [WAYFARER REDEMPTION 1]  P
DOUGLASS, SARA  WOUNDED HAWK, THE [@CRUCIBLE 2]  H
DOW DELL, DEL  SPEARMEN OF ARN  P
DOW DELL, DEL  WARLORD OF GHANDOR  P
DOWD, TOM  BURNING BRIGHT [@SHADOWRUN 15]  P
DOWD, TOM  NIGHT’S PAWN [@SHADOWRUN 10]  C/P,P
DOWLING, TERRY  RYNOSSEROS  H:2
DOWNER, ANN  BOOKS OF THE KEEPERS, THE [SPELLKEY 3]  H
DOWNER, ANN  SPELLKEY TRILOGY, THE [SPELLKEY (REVISED) 1,2,3]  P
DOWNING, PAULA E.  FALLWAY  P
DOWNING, PAULA E.  FLARE STAR  P
DOWNING, PAULA E.  RINN’S STAR  C/P,P
DOWNING, PAULA E.  WHISPER OF TIME, A  P
DOXEY, WILLIAM S.  ESPIONAGE  P

DOYLE, ADRIAN CONAN/ CARR, JOHN DICKSON EXPLOITS OF SHERLOCK HOLMES, THE

P

DOYLE, ARTHUR CONAN  ADVENTURES OF SHERLOCK HOLMES, THE  C/P,P
DOYLE, ARTHUR CONAN  CHIEN DES BASKERVILLE, LE  FP
DOYLE, ARTHUR CONAN  GREAT KEINPLATZ EXPERIMENT, THE  SR-H
DOYLE, ARTHUR CONAN  HOUND OF THE BASKERVILLES, THE  P
DOYLE, ARTHUR CONAN  LOST WORLD, THE  C/P:2,P

DOYLE, ARTHUR CONAN LOST WORLD, THE [PROF. CHALLENGER 1]/ POISON BELT, THE [PROF. CHALLENGER 1]

SR-H

DOYLE, ARTHUR CONAN  MARACOT DEEP, THE  SR-H,C/P,P
DOYLE, ARTHUR CONAN  POISON BELT, THE  C/P:2,P
DOYLE, ARTHUR CONAN  RETURN OF SHERLOCK HOLMES, THE  H
DOYLE, ARTHUR CONAN  STUDY IN SCARLET, A/SIGN OF FOUR, THE  P

DOYLE, ARTHUR CONAN WHEN THE WORLD SCREAMED AND OTHER STORIES

[PROF. CHALLENGER 2]

LP

DOYLE, DEBRA/MACDONALD, JAMES D.  BY HONOR BETRAY’D [MAGEWORLDS 3]  P
DOYLE, DEBRA/MACDONALD, JAMES D.  GATHERING FLAME, THE [MAGEWORLDS 4]  P
DOYLE, DEBRA/MACDONALD, JAMES D.  JUDGMENT NIGHT  P
DOYLE, DEBRA/MACDONALD, JAMES D.  LAND OF MIST AND SNOW  P
DOYLE, DEBRA/MACDONALD, JAMES D.  LONG HUNT, THE [MAGEWORLDS 5]  P
DOYLE, DEBRA/MACDONALD, JAMES D.  NIGHT OF THE LIVING RAT [MELVINGE 2]  C/P,P
DOYLE, DEBRA/MACDONALD, JAMES D.  PRICE OF THE STARS, THE [MAGEWORLDS 1]  C/P:2,P
DOYLE, DEBRA/MACDONALD, JAMES D.  REQUIEM FOR BOONE [@EARTH FINAL CONFLICT]  C/P,LP
DOYLE, DEBRA/MACDONALD, JAMES D.  STARPILOT’S GRAVE [MAGEWORLDS 2]  C/P,P
DOYLE, DEBRA/MACDONALD, JAMES D.  STARS ASUNDER, THE [MAGEWORLDS 6]  H:2
DOYLE, DEBRA/MACDONALD, JAMES D.  WORKING OF STARS, A [MAGEWORLDS 7]  H,P
DOZOIS, GARDNER/DANN, JACK  ALIENS!  C/PA,PA
DOZOIS, GARDNER  ANOTHER WORLD  HA
DOZOIS, GARDNER  BEST OF THE BEST, THE  LPA
DOZOIS, GARDNER  BEST OF THE BEST 2, THE  LPA
DOZOIS, GARDNER  BEST SCIENCE FICTION STORIES OF THE YEAR 6  C/PA,PA
DOZOIS, GARDNER  BEST SCIENCE FICTION STORIES OF THE YEAR 7  HA,PA
DOZOIS, GARDNER  BEST SCIENCE FICTION STORIES OF THE YEAR 8  HA,PA

The MITSFS Pinkdex  September 24, 2007  Page 120
AUTHOR  TITLE  CODE
DOZOIS, GARDNER  BEST SCIENCE FICTION STORIES OF THE YEAR 9  HA
DOZOIS, GARDNER  DAY IN THE LIFE, A  HA,PA
DOZOIS, GARDNER  EXPLORERS  LPA
DOZOIS, GARDNER  FURTHEST HORIZON, THE  LPA
DOZOIS, GARDNER  GALILEO’S CHILDREN  HA
DOZOIS, GARDNER  GEODESIC DREAMS  P
DOZOIS, GARDNER  GOOD NEW STUFF, THE  LPA
DOZOIS, GARDNER  GOOD OLD STUFF, THE  LPA
DOZOIS, GARDNER  ISAAC ASIMOV’S ALIENS  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S CAMELOT  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S CHRISTMAS  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S CYBERDREAMS  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S DETECTIVES  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S EARTH  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S FATHER’S DAY  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S GHOSTS  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S HALLOWEEN  PA
DOZOIS, GARDNER  ISAAC ASIMOV’S MARS  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S MOONS  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S MOTHER’S DAY  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S ROBOTS  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S SKIN DEEP  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S SOLAR SYSTEM  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S UTOPIAS  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S VALENTINES  C/PA,PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S VAMPIRES  PA
DOZOIS, GARDNER/WILLIAMS, SHEILA  ISAAC ASIMOV’S WEREWOLVES  PA
DOZOIS, GARDNER  KILLING ME SOFTLY  PA,LPA
DOZOIS, GARDNER  MODERN CLASSIC SHORT NOVELS OF SCIENCE FICTION  HA
DOZOIS, GARDNER  MODERN CLASSICS OF FANTASY  LPA
DOZOIS, GARDNER  MODERN CLASSICS OF SCIENCE FICTION  HA
DOZOIS, GARDNER  MORNING CHILD AND OTHER STORIES  LP
DOZOIS, GARDNER  NEBULA AWARDS SHOWCASE 2006  LPA
DOZOIS, GARDNER/STRAHAN, JONATHAN  NEW SPACE OPERA, THE  SR-LPA

DOZOIS, GARDNER/ EFFINGER, GEORGE ALEC NIGHTMARE BLUE

C/P,P

DOZOIS, GARDNER  ONE MILLION A.D.  HA
DOZOIS, GARDNER/CASPER, SUSAN  RIPPER!  SFWA-P
DOZOIS, GARDNER/SCHMIDT, STANLEY  ROADS NOT TAKEN  PA
DOZOIS, GARDNER  STRANGE DAYS  H
DOZOIS, GARDNER  STRANGERS  C/P,P,H
DOZOIS, GARDNER  SUPERMEN  LPA
DOZOIS, GARDNER  TIME TRAVELERS  PA

DOZOIS, GARDNER TRANSCENDENTAL TALES FROM ISAAC ASIMOV’S SCIENCE FICTION

MAGAZINE

LPA

DOZOIS, GARDNER  VISIBLE MAN, THE  P
DOZOIS, GARDNER  WORLDMAKERS  LPA

DOZOIS, GARDNER/LEE, TINA/ SCHMIDT, STANLEY/ STROCK, IAN RANDALL/ WILLIAMS, SHEILA WRITING SCIENCE FICTION AND FANTASY

R/REF-P

DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 1, THE  LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 2, THE  LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 3, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 4, THE  LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 5, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 6, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 7, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 8, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 9, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 10, THE  HA:2,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 11, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 12, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 13, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 14, THE  HA:2,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 15, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 16, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 17, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 18, THE  HA,LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 19, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 20, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 21, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 22, THE  HA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 23, THE  LPA
DOZOIS, GARDNER  YEAR’S BEST SCIENCE FICTION 24, THE  LPA

The MITSFS Pinkdex  September 24, 2007  Page 121
AUTHOR  TITLE  CODE
DRACHMAN, THEODORE S.  CRY PLAGUE!  SR-DD013
DRAGO, HENRY  OUT OF THE SILENT NORTH  H
DRAGO, TY  PHOBOS  H
DRAKE, ASA  DEATH RIDERS OF HEL [HEL 2]  P
DRAKE, ASA  WARRIOR WITCH OF HEL [HEL 1]  C/P,P
DRAKE, ASA  WEREBEASTS OF HEL [HEL 3]  P
DRAKE, DAVID  ALL THE WAY TO THE GALLOWS  C/P,P
DRAKE, DAVID/MORRIS, JANET  ARC RIDERS [ARC RIDERS]  C/P:2,P
DRAKE, DAVID/MOSIMAN, BILLIE SUE  ARMAGEDDON  C/PA,PA
DRAKE, DAVID  AT ANY PRICE [HAMMER’S SLAMMERS]  C/P:2,P
DRAKE, DAVID  BALEFIRES  H
DRAKE, DAVID/FAWCETT, BILL  BATTLESTATION [@BATTLESTATION #1]  C/PA,PA,SFWA-P:2
DRAKE, DAVID  BIRDS OF PREY  C/P,SFWA-P,P,H
DRAKE, DAVID  BLUEBLOODS [STARHUNTERS 3]  C/PA,PA,SFWA-P
DRAKE, DAVID/FAWCETT, BILL  BREAKTHROUGH [@FLEET #3]  C/PA,PA,SFWA-P
DRAKE, DAVID  BRIDGEHEAD  C/P:2,SFWA-P,P
DRAKE, DAVID  BUTCHER’S BILL, THE [HAMMER’S SLAMMERS]  C/P:2,P
DRAKE, DAVID  CAUGHT IN THE CROSSFIRE [HAMMER’S SLAMMERS]  C/P,P

DRAKE, DAVID COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 1

[HAMMER’S SLAMMERS]

H

DRAKE, DAVID COMPLETE HAMMER’S SLAMMERS, THE, VOLUME 2

[HAMMER’S SLAMMERS]

H

DRAKE, DAVID/FAWCETT, BILL  COUNTERATTACK [@FLEET #2]  PA,C/PA,SFWA-P
DRAKE, DAVID  COUNTING THE COST [HAMMER’S SLAMMERS]  P,SFWA-P
DRAKE, DAVID/FAWCETT, BILL  CRISIS [@FLEET #6]  C/PA,PA,SFWA-P:2
DRAKE, DAVID  CROSS THE STARS  C/P:2,P
DRAKE, DAVID  DAGGER [THIEVES’ WORLD]  C/P,SFWA-P,P,H
DRAKE, DAVID  DOGS OF WAR  PA
DRAKE, DAVID  DRAGON LORD, THE  C/P,SFWA-P,H,P,SFWA
DRAKE, DAVID  DRAGON LORD, THE (PROOF)  SR-LP
DRAKE, DAVID  ETERNAL CITY, THE  P,SFWA-P:2
DRAKE, DAVID/MORRIS, JANET  EXPLORERS IN HELL [HEROES IN HELL]  P,SFWA-P
DRAKE, DAVID  FAR SIDE OF THE STARS, THE [RCN 3]  H:2
DRAKE, DAVID  FAR SIDE OF THE STARS, THE (PROOF) [RCN 3]  SR-LP
DRAKE, DAVID  FIRESHIPS [REACHES 3]  C/P,P,H
DRAKE, DAVID/FAWCETT, BILL  FLEET, THE [@FLEET #1]  C/PA,PA,SFWA-P
DRAKE, DAVID  FOREIGN LEGIONS [RANKS OF BRONZE]  HA,PA
DRAKE, DAVID  FORLORN HOPE, THE  C/P:2,SFWA-P:2,P
DRAKE, DAVID  FORTRESS  C/P:2,SFWA,SFWA-P,P,H
DRAKE, DAVID  FORTRESS OF GLASS, THE [CROWN OF THE ISLES 1]  H
DRAKE, DAVID/MORRIS, JANET  FOURTH ROME, THE [ARC RIDERS]  C/P,P
DRAKE, DAVID  FROM THE HEART OF DARKNESS  P
DRAKE, DAVID  GODDESS OF THE ICE REALM [LORD OF THE ISLES 5]  H,P
DRAKE, DAVID  GRIMMER THAN HELL  H,P
DRAKE, DAVID  HAMMER’S SLAMMERS [HAMMER’S SLAMMERS]  C/P:2,P
DRAKE, DAVID  HAMMER’S SLAMMERS (INCLUDES THE TANK LORDS)  C/P:2,P
DRAKE, DAVID/MIESEL, SANDRA  HEADS TO THE STORM  PA,SFWA-P:2
DRAKE, DAVID/THOMAS, THOMAS T.  HONORABLE DEFENSE, AN [CRISIS OF EMPIRE 1]  C/P:2,SFWA-P,P
DRAKE, DAVID/KJELGAARD, JIM  HUNTER RETURNS, THE  C/P:2,SFWA-P,P
DRAKE, DAVID  IGNITING THE REACHES [REACHES 1]  C/P,P,H
DRAKE, DAVID  JUNGLE, THE  H:2,P
DRAKE, DAVID/WAGNER, KARL EDWARD  KILLER  C/P:2,SFWA-P,P
DRAKE, DAVID  LACEY AND HIS FRIENDS  C/P:2,P
DRAKE, DAVID  LT. LEARY, COMMANDING [RCN 2]  C/P,P,H
DRAKE, DAVID  LORD OF THE ISLES [LORD OF THE ISLES 1]  C/P:2,P,H
DRAKE, DAVID  MASTER OF THE CAULDRON [LORD OF THE ISLES 6]  P
DRAKE, DAVID  MEN HUNTING THINGS [STARHUNTERS 1]  SFWA-P,C/PA,PA
DRAKE, DAVID  MILITARY DIMENSION, THE  C/P:2,P
DRAKE, DAVID  MILITARY DIMENSION: MARK II, THE  P
DRAKE, DAVID  MISTRESS OF THE CATACOMBS [LORD OF THE ISLES 4]  C/P:2,P,H

DRAKE, DAVID/FLINT, ERIC/SPOOR, RYK E./ KUTTNER, HENRY MOUNTAIN MAGIC

C/PA,PA

DRAKE, DAVID  NORTHWORLD [NORTHWORLD 1]  C/P:2,SFWA-P:2,P
DRAKE, DAVID  NORTHWORLD JUSTICE [NORTHWORLD 3]  C/P:2,SFWA-P:2,P
DRAKE, DAVID  NORTHWORLD TRILOGY [NORTHWORLD 1,2,3]  C/P,P
DRAKE, DAVID  NORTHWORLD VENGEANCE [NORTHWORLD 2]  C/P,SFWA-P:2,P
DRAKE, DAVID  OLD NATHAN  C/P:2,SFWA-P,P
DRAKE, DAVID  OTHER TIMES THAN PEACE  H
DRAKE, DAVID  PATRIOTS  C/P:2,P
DRAKE, DAVID  PAYING THE PIPER [HAMMER’S SLAMMERS]  P
DRAKE, DAVID  QUEEN OF DEMONS [LORD OF THE ISLES 2]  C/P,P,H:2
DRAKE, DAVID  RANKS OF BRONZE [RANKS OF BRONZE]  C/P:2,P
DRAKE, DAVID  REDLINERS  C/P:2,P
DRAKE, DAVID  ROLLING HOT [HAMMER’S SLAMMERS]  C/P:2,SFWA-P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 122
AUTHOR  TITLE  CODE
DRAKE, DAVID  SEA HAG, THE [CRYSTAL WALLS 1]  C/P,SFWA-P,P
DRAKE, DAVID/MIESEL, SANDRA  SEPARATE STAR, A  C/PA,PA,SFWA-P
DRAKE, DAVID  SERVANT OF THE DRAGON [LORD OF THE ISLES 3]  C/P:3,P,H
DRAKE, DAVID  SHARP END, THE [HAMMER’S SLAMMERS]  C/P,SFWA,P,H
DRAKE, DAVID  SKYRIPPER  C/P,P
DRAKE, DAVID  SOME GOLDEN HARBOR [RCN 5]  H

DRAKE, DAVID/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SPACE DREADNOUGHTS

PA,C/PA

DRAKE, DAVID/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SPACE GLADIATORS

C/PA,PA

DRAKE, DAVID/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SPACE INFANTRY

PA,C/PA

DRAKE, DAVID  SQUARE DEAL, THE [@CAR WARRIORS #1]  LP
DRAKE, DAVID  STARLINER  C/P,P
DRAKE, DAVID  SURFACE ACTION  C/P:2,SFWA-P:2,P
DRAKE, DAVID/FAWCETT, BILL  SWORN ALLIES [@FLEET #4]  PA,SFWA-P:2
DRAKE, DAVID  TANK LORDS, THE [HAMMER’S SLAMMERS]  C/P,P
DRAKE, DAVID  THINGS HUNTING MEN [STARHUNTERS 2]  PA,SFWA-P
DRAKE, DAVID  THROUGH THE BREACH [REACHES 2]  C/P,P,H
DRAKE, DAVID  TIME SAFARI  P,SFWA-P
DRAKE, DAVID/FAWCETT, BILL  TOTAL WAR [@FLEET #5]  PA,SFWA-P:2
DRAKE, DAVID  TYRANNOSAUR  C/P,P
DRAKE, DAVID/FAWCETT, BILL  VANGUARD [@BATTLESTATION #2]  C/PA,PA,SFWA-P:2
DRAKE, DAVID  VETTIUS AND HIS FRIENDS  C/P:2,SFWA-P,P
DRAKE, DAVID  VOYAGE, THE [HAMMER’S SLAMMERS]  C/P,P,H
DRAKE, DAVID  WARRIOR, THE [HAMMER’S SLAMMERS]  C/P:2,P
DRAKE, DAVID  WAY TO GLORY, THE [RCN 4]  H
DRAKE, DAVID  WITH THE LIGHTNINGS [RCN 1]  H:2,P
DRAKE, DAVID/FLINT, ERIC/BAEN, JIM  WORLD TURNED UPSIDE DOWN, THE  HA
DRAKE, EMILY  CURSE OF ARKADY, THE [@MAGICKERS #2]  P
DRAKE, EMILY  DRAGON GUARD, THE [@MAGICKERS #3]  P
DRAKE, EMILY  MAGICKERS, THE [MAGICKERS 1]  P
DRAKE, H. L.  NULL-A WORLDS OF A. E. VAN VOGT, THE  R/REF-P
DRAULANS, DIRK  RED QUEEN, THE  H
DREADSTONE, CARL  BRIDE OF FRANKENSTEIN, THE  P
DREADSTONE, CARL  CREATURE FROM THE BLACK LAGOON  P
DREADSTONE, CARL  DRACULA’S DAUGHTER  P
DREADSTONE, CARL  MUMMY, THE  P
DREADSTONE, CARL  WEREWOLF OF LONDON, THE  P
DREADSTONE, CARL  WOLFMAN, THE  P
DREIFUSS, JEROME  FURLOUGH FROM HEAVEN  H
DREIFUSS, KURT  OTHER SIDE OF THE UNIVERSE, THE  H
DRENNAN, KATHRYN M.  TO DREAM IN THE CITY OF SORROWS [@BABYLON 5 #9]  C/P,P
DREW, WAYLAND  DRAGONSLAYER  H:2,P
DREW, WAYLAND  ERTHRING CYCLE, THE [ERTHRING 1,2,3]  H:2
DREW, WAYLAND  GAIAN EXPEDIENT [ERTHRING 2]  C/P:2,P
DREW, WAYLAND  MASTER OF NORRIYA, THE [ERTHRING 3]  P
DREW, WAYLAND  MEMOIRS OF ALCHERINGIA, THE [ERTHRING 1]  C/P:2,P
DREW, WAYLAND  WILLOW  C/P,P
DREW, WEYLAND  *BATTERIES NOT INCLUDED  P
DREYFUSS, ERNST  UNFROZEN, THE  P
DREYFUSS, RICHARD/TURTLEDOVE, HARRY TWO GEORGES, THE  C/P,P,H
DRIZARI, NELO  ALBANIAN-ENGLISH AND ENGLISH-ALBANIAN DICTIONARY  DICT
DRUMM, D. B.  BORDER WAR [@TRAVELER #6]  P
DRUMM, D. B.  CHILDREN’S CRUSADE, THE [@TRAVELER #11]  P
DRUMM, D. B.  FIRST, YOU FIGHT [@TRAVELER #1]  P
DRUMM, D. B.  GHOST DANCERS [@TRAVELER #13]  P
DRUMM, D. B.  HELL ON EARTH [@TRAVELER #10]  P
DRUMM, D. B.  KINGDOM COME [@TRAVELER #2]  C/P,P
DRUMM, D. B.  PREY, THE [@TRAVELER #12]  P
DRUMM, D. B.  ROAD GHOST, THE [@TRAVELER #7]  P
DRUMM, D. B.  ROAD WAR [@TRAVELER #5]  P
DRUMM, D. B.  STALKERS, THE [@TRAVELER #3]  P
DRUMM, D. B.  STALKING TIME, THE [@TRAVELER #9]  P
DRUMM, D. B.  TERMINAL ROAD [@TRAVELER #8]  P
DRUMM, D. B.  TO KILL A SHADOW [@TRAVELER #4]  P
DRURY, ALLEN  THRONE OF SATURN, THE  H:2
DU BRUL, JACK  DEEP FIRE RISING (PROOF)  SR-LP
DU MAURIER, DAPHNE  DAPHNE DU MAURIER’S CLASSICS OF THE MACABRE  SFWA
DU MAURIER, GEORGE  MARTIAN, THE  H
DUANE, DIANE  BOOK OF NIGHT WITH MOON, THE [FELINE WIZARDS 1]  C/P,LP
DUANE, DIANE  DARK MIRROR [STAR TREK TNG]  H:2,P
DUANE, DIANE  DEEP WIZARDRY  H
DUANE, DIANE  DOCTOR’S ORDERS [STAR TREK]  C/P:3,P

AUTHOR  TITLE  CODE
DUANE, DIANE  DOOR INTO FIRE, THE [TALE OF THE FIVE 1]  P
DUANE, DIANE  DOOR INTO FIRE, THE (REVISED) [TALE OF THE FIVE 1]  C/P,SFWA-P,LP
DUANE, DIANE  DOOR INTO SHADOW, THE [TALE OF THE FIVE 2]  H:2,SFWA-P,LP
DUANE, DIANE  DOOR INTO SUNSET, THE [TALE OF THE FIVE 3]  SR-P,H
DUANE, DIANE  EMPTY CHAIR, THE [STAR TREK RIHANNSU 5]  P
DUANE, DIANE/MORWOOD, PETER  HIGH MOON [@SPACE COPS 3]  C/P:2,P
DUANE, DIANE  HIGH WIZARDRY [YOUNG WIZARDS 3]  H:2,P
DUANE, DIANE  HONOR BLADE [STAR TREK RIHANNSU 4]  C/P,P
DUANE, DIANE  INTELLIVORE [STAR TREK TNG]  C/P:2,P
DUANE, DIANE/MORWOOD, PETER  KILL STATION [@SPACE COPS 1]  C/P:2,P
DUANE, DIANE/MORWOOD, PETER  MINDBLAST [@SPACE COPS 2]  C/P,P
DUANE, DIANE  MY ENEMY, MY ALLY [STAR TREK RIHANNSU 1]  C/P:3,P
DUANE, DIANE  NIGHTFALL AT ALGEMRON [STAR*DRIVE HARBINGER 3]  P
DUANE, DIANE  OCTOPUS AGENDA, THE [@SPIDERMAN]  H
DUANE, DIANE  RIHANNSU [STAR TREK RIHANNSU 1,2,3,4]  H
DUANE, DIANE/MORWOOD, PETER  ROMULAN WAY, THE [STAR TREK RIHANNSU 2]  C/P:2,P
DUANE, DIANE/MORWOOD, PETER  SEAQUEST DSV: THE NOVEL [@SEAQUEST DSV 1]  P
DUANE, DIANE  SO YOU WANT TO BE A WIZARD [YOUNG WIZARDS 1]  C/P,P,H
DUANE, DIANE  SPIDER-MAN: THE VENOM FACTOR [SPIDER-MAN]  P
DUANE, DIANE  SPOCK’S WORLD [STAR TREK]  H:2,P
DUANE, DIANE  STARRISE AT CORRIVALE [STAR*DRIVE HARBINGER 1]  P
DUANE, DIANE  STEALING THE ELF-KING’S ROSES  C/P,P
DUANE, DIANE  STORM AT ELDALA [STAR*DRIVE HARBINGER 2]  P
DUANE, DIANE  SUPPORT YOUR LOCAL WIZARD [YOUNG WIZARDS 1,2,3]  H
DUANE, DIANE  SWORDHUNT [STAR TREK RIHANNSU 3]  C/P:2,P
DUANE, DIANE  TO VISIT THE QUEEN [FELINE WIZARDS 2]  C/P,LP
DUANE, DIANE  UFO DEFENSE [@X-COM]  P
DUANE, DIANE  WIZARD ABROAD, A [YOUNG WIZARDS 4]  C/P,P
DUANE, DIANE  WIZARD ALONE, A [YOUNG WIZARDS 6]  H
DUANE, DIANE  WIZARDS AT WAR [YOUNG WIZARDS 8]  H
DUANE, DIANE  WIZARD’S DILEMMA, THE [YOUNG WIZARDS 5]  H
DUANE, DIANE  WIZARD’S HOLIDAY [YOUNG WIZARDS 7]  H
DUANE, DIANE  WOUNDED SKY, THE [STAR TREK]  C/P:2,P
DUBOIS, BRENDAN  RESURRECTION DAY  H,P
DUBOIS, MARGUERITE-MARIE, ET AL.  LAROUSSE’S FRENCH-ENGLISH ENGLISH-FRENCH DICTIONARY  DICT
DUBOIS, THEODORA  SARAH HALL’S SEA GOD  P
DUBOIS, THEODORA  SOLUTION T-25  SR-P,C/P:2
DUBOWSKI, CATHY EAST  DOG’S LIFE, A [@SABRINA THE TEENAGE WITCH #9]  P
DUBOWSKI, CATHY EAST  PSYCHIC KITTY [@SABRINA THE TEENAGE WITCH SALEM #6]  P
DUBOWSKI, CATHY EAST  SANTA’S LITTLE HELPER [@SABRINA THE TEENAGE WITCH #5]  P
DUBOWSKI, MARK  CAT TV [@SABRINA THE TEENAGE WITCH SALEM #1]  P
DUBOWSKI, MARK  KING OF CATS, THE [@SABRINA THE TEENAGE WITCH SALEM #4]  P
DUBREUIL, LINDA  EVIL, EVIL  P
DUBREUIL, LINDA  NIGHTMARE BABY  P
DUCHAMP, L. TIMMEL  ALANYA TO ALANYA [@MARQ’SSAN #1]  LP
DUCHAMP, L. TIMMEL  LOVE’S BODY, DANCING IN TIME  LP
DUCHAMP, L. TIMMEL  RED ROSE RAGES (BLEEDING), THE  LP
DUCHAMP, L. TIMMEL  RENEGADE [@MARQ’SSAN #2]  LP
DUCHAMP, L. TIMMEL  TALKING BACK  LPA
DUCHAMP, L. TIMMEL  TSUNAMI [@MARQ’SSAN #3]  LP
DUDLEY, ROY C.  GALACTIC GAMBIT  H
DUDLEY, TERRENCE  BLACK ORCHID [DOCTOR WHO]  SR-P,C/P
DUDLEY, TERRENCE  K-9 AND COMPANY [DOCTOR WHO]  SR-P,C/P
DUDLEY, TERRENCE  KING’S DEMONS, THE [DOCTOR WHO]  SR-P
DUE, TANANARIVE  BETWEEN, THE  LP
DUE, TANANARIVE  MY SOUL TO KEEP  LP
DUIGON, LEE  PRECOG  P
DUKE, CYNTHIA J.  ELEMENT 115  P
DUKE, MADELAINE  CLARET, SANDWICHES AND SIN  P
DUKE, MADELAINE  THIS BUSINESS OF BOMFOG  C/P,P,H
DUMAS, ALEXANDRE  HORROR AT FONTENAY  SR-P
DUNCAN, ANDY  BELUTHAHATCHIE AND OTHER STORIES  H
DUNCAN, BRUCE  MIRROR IMAGE  SR-DB0081
DUNCAN, DAVE  ALCHEMIST’S APPRENTICE, THE  LP
DUNCAN, DAVE  CHILDREN OF CHAOS [CHILDREN OF CHAOS 1]  H
DUNCAN, DAVE  COMING OF WISDOM, THE [SEVENTH SWORD 2]  C/P:2,P
DUNCAN, DAVE  CROOKED HOUSE, THE  C/P,P
DUNCAN, DAVE  CURSED, THE  P
DUNCAN, DAVE  CUTTING EDGE, THE [HANDFUL OF MEN 1]  C/P:2,P
DUNCAN, DAVE  DESTINY OF THE SWORD, THE [SEVENTH SWORD 3]  C/P:2,P
DUNCAN, DAVE  EMPEROR AND CLOWN [MAN OF HIS WORD 4]  C/P,P
DUNCAN, DAVE  FAERY LANDS FORLORN [MAN OF HIS WORD 2]  C/P:2,P
DUNCAN, DAVE  FUTURE INDEFINITE [@GREAT GAME #3]  P
DUNCAN, DAVE  GILDED CHAIN, THE [KING’S BLADES]  H,P

AUTHOR  TITLE  CODE
DUNCAN, DAVE  HERO!  C/P:2,P
DUNCAN, DAVE  HUNTERS’ HAUNT, THE [OMAR 2]  C/P:2,P
DUNCAN, DAVE  IMPOSSIBLE ODDS [KING’S BLADES]  H,P
DUNCAN, DAVE  JAGUAR KNIGHTS, THE [KING’S BLADES]  H
DUNCAN, DAVE  LIVING GOD, THE [HANDFUL OF MEN 4]  C/P,P,H
DUNCAN, DAVE  LORD OF THE FIRE LANDS [KING’S BLADES]  P
DUNCAN, DAVE  MAGIC CASEMENT [MAN OF HIS WORD 1]  C/P,P
DUNCAN, DAVE  MOTHER OF LIES [CHILDREN OF CHAOS 2]  H
DUNCAN, DAVE  PARAGON LOST [@KING’S BLADES]  P
DUNCAN, DAVE  PAST IMPERATIVE [@GREAT GAME #1]  H,P
DUNCAN, DAVE  PERILOUS SEAS [MAN OF HIS WORD 3]  C/P,P
DUNCAN, DAVE  PRESENT TENSE [@GREAT GAME #2]  H
DUNCAN, DAVE  REAVER ROAD, THE  H,P
DUNCAN, DAVE  RELUCTANT SWORDSMAN, THE [SEVENTH SWORD 1]  C/P:2,P
DUNCAN, DAVE  ROSE-RED CITY, A  P
DUNCAN, DAVE  SHADOW  P
DUNCAN, DAVE  SIR STALWART [KING’S DAGGERS 1]  P
DUNCAN, DAVE  SKY OF SWORDS [KING’S BLADES]  H
DUNCAN, DAVE  STRICKEN FIELD, THE [HANDFUL OF MEN 3]  C/P,P,H
DUNCAN, DAVE  STRINGS  C/P,P
DUNCAN, DAVE  UPLAND OUTLAWS [HANDFUL OF MEN 2]  C/P,P,H
DUNCAN, DAVE  WEST OF JANUARY  C/P:2,P
DUNCAN, DAVID  BEYOND EDEN  C/P,P
DUNCAN, DAVID  DARK DOMINION  C/P:2,P
DUNCAN, DAVID  OCCAM’S RAZOR  C/P,P,H
DUNCAN, HAL  INK [BOOK OF ALL HOURS 2]  LP
DUNCAN, HAL  VELLUM [BOOK OF ALL HOURS 1]  LP
DUNCAN, LOIS  GIFT OF MAGIC, A  C/P:2,P
DUNCAN, LOIS  LOCKED IN TIME  P
DUNLOP, EILEEN  ROBINSHEUGH  P
DUNMORE, SPENCER  SOUND OF WINGS, THE  C/P,SFWA-P,P
DUNN, J. ALLEN  TREASURE OF ATLANTIS, THE  P
DUNN, J. R.  DAYS OF CAIN  C/P,P
DUNN, J. R.  FULL TIDE OF NIGHT  H,P
DUNN, J. R.  THIS SIDE OF JUDGEMENT  H,P
DUNN, JACK/HALDEMAN, JACK C. II  ECHOES OF THUNDER  TD32
DUNN, PHILIP  BLACK MOON, THE [CABAL 2]  P
DUNN, PHILIP  CABAL, THE [CABAL 1]  P
DUNN, PHILIP  EVANGELIST, THE [CABAL 3]  P
DUNN, SAUL  COMING OF STEELEYE, THE [STEELEYE 1]  SR-P
DUNN, SAUL  STEELEYE: WATERSPACE [STEELEYE]  SR-P
DUNNE, J. W.  EXPERIMENT WITH TIME, AN  [C/RAN-P]
DUNNING, LAWRENCE  NEUTRON TWO IS CRITICAL  P
DUNSANY, LORD  AT THE EDGE OF THE WORLD  SR-P
DUNSANY, LORD  BEYOND THE FIELDS WE KNOW  SR-P
DUNSANY, LORD  BOOK OF WONDER, THE  SR-L,LP
DUNSANY, LORD  CHARWOMAN’S SHADOW, THE  C/P,P,H
DUNSANY, LORD  COLLECTED JORKENS, THE, VOLUME TWO [JORKENS]  H
DUNSANY, LORD  COLLECTED JORKENS, THE, VOLUME THREE [JORKENS]  H
DUNSANY, LORD  COMPLETE PEGANA, THE  LP
DUNSANY, LORD  DON RODRIGUEZ: CHRONICLES OF SHADOW VALLEY  P
DUNSANY, LORD  DREAMER’S TALES AND OTHER STORIES, A  SR-P,H
DUNSANY, LORD  CURSE OF THE WISE WOMAN, THE  SR-P
DUNSANY, LORD  FIFTY-ONE TALES  LP
DUNSANY, LORD  FOOD OF DEATH: FIFTY-ONE TALES, THE  LP
DUNSANY, LORD  FOURTH BOOK OF JORKENS, THE [JORKENS]  H
DUNSANY, LORD  GHOSTS OF THE HEAVISIDE LAYER AND OTHER FANTASMS, THE  H
DUNSANY, LORD  GODS, MEN AND GHOSTS  H,LP
DUNSANY, LORD  GODS OF PEGANA, THE  LP
DUNSANY, LORD  KING OF ELFLAND’S DAUGHTER, THE  C/P:2,P
DUNSANY, LORD  PLAYS OF GODS AND MEN  H
DUNSANY, LORD  TALES OF THREE HEMISPHERES  SR-H
DUNSANY, LORD  TIME AND THE GODS  LP
DUNSANY, LORD  UP IN THE HILLS  H
DUNSANY, LORD  WONDER TALES  LP
DUNTEMANN, JEFF  CUNNING BLOOD, THE  H
DUPREY, RICHARD/O’LEARY, BRIAN  SPACESHIP TITANIC  H
DURBIN, FREDERIC C.  DRAGONFLY  H
DURGIN, DORANNA  BARRENLANDS [DUN LADY’S JESS 4]  P
DURGIN, DORANNA  CHANGESPELL [DUN LADY’S JESS 2]  P
DURGIN, DORANNA  CHANGESPELL LEGACY [DUN LADY’S JESS 3]  P
DURGIN, DORANNA  DUN LADY’S JESS [DUN LADY’S JESS 1]  P
DURGIN, DORANNA  FEARLESS [@ANGEL]  P
DURGIN, DORANNA  FERAL DARKNESS, A  P

The MITSFS Pinkdex  September 24, 2007  Page 125
AUTHOR  TITLE  CODE
DURGIN, DORANNA  HERITAGE [EARTH FINAL CONFLICT]  H
DURGIN, DORANNA  IMPRESSIONS [@ANGEL]  P
DURGIN, DORANNA  SEER’S BLOOD  P
DURGIN, DORANNA  TOOTH AND CLAW [STAR TREK TNG]  P
DURGIN, DORANNA  TOUCHED BY MAGIC [REANDN 1]  P
DURGIN, DORANNA  WOLF JUSTICE [REANDN 2]  P
DURGIN, DORANNA  WOLVERINE’S DAUGHTER  P
DURHAM, DAVID ANTHONY  ACACIA [WAR WITH THE MEIN 1]  H
DURHAM, DAVID ANTHONY  ACACIA (PROOF) [WAR WITH THE MEIN 1]  SR-LP
DURHAM, VICTOR  SUBMARINE BOYS’ TRIAL TRIP  H
DURRELL, LAWRENCE  NUNQUAM  C/P,P
DURST, MOSE  TO BIGOTRY, NO SANCTION  [C/RAN-H]
DURWOOD, THOMAS  CLOSE ENCOUNTERS OF THE THIRD KIND: A DOCUMENT OF THE FILM  R/REF-H,C/REF-H
DUTOURD, JEAN  DOG’S HEAD, A  SR-P
DVORKIN, DAVID  BUDSPY  H,SFWA
DVORKIN, DAVID/DVORKIN, DANIEL  CAPTAIN’S HONOR, THE [STAR TREK TNG]  C/P:2,P
DVORKIN, DAVID  CENTRAL HEAT  C/P:2,P
DVORKIN, DAVID  CHILDREN OF SHINY MOUNTAIN, THE  P
DVORKIN, DAVID  GREEN GOD, THE  P
DVORKIN, DAVID  PIT PLANET  H
DVORKIN, DAVID  SEEKERS, THE  H:2
DVORKIN, DAVID  TIMETRAP [STAR TREK]  C/P:2,P
DVORKIN, DAVID  TRELLISANE CONFRONTATION, THE [STAR TREK]  C/P:3,P
DVORKIN, DAVID  URSUS  H
DWIGHT, JEFFRY  AGE OF WONDERS, THE  LPA
DWIGHT, JEFFRY  BETWEEN THE DARKNESS AND THE FIRE  LPA
DYE, CHARLES  PRISONER IN THE SKULL  H,P
DYMOND, ANDREW  DARK SIDE OF THE SUN [@FARSCAPE]  C/P,P

DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

100 ASTOUNDING LITTLE ALIEN STORIES

HA

DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

100 CREEPY LITTLE CREATURE STORIES

HA

DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

100 GHASTLY LITTLE GHOST STORIES

HA

DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

100 WICKED LITTLE WITCH STORIES

HA

DZIEMIANOWICZ, STEFAN R./ WEINBERG, ROBERT/ GREENBERG, MARTIN H.

100 WILD LITTLE WEIRD TALES HA,LPA

AUTHOR  TITLE  CODE
EAGER, EDWARD  HALF MAGIC  H,LP
EAGER, EDWARD  KNIGHT’S CASTLE  LP
EAGER, EDWARD  MAGIC OR NOT? [MAGIC OR NOT? 1]  H,LP
EAGER, EDWARD  TIME GARDEN, THE  P
EAGER, EDWARD  WELL-WISHERS, THE [MAGIC OR NOT? 2]  LP
EALES, JULIAN  WALKING IN ETERNITY [DOCTOR WHO]  SR-LPA
EARL, ROBERT  BURNING SHORE, THE [@WARHAMMER]  P
EARL, ROBERT  CORRUPTED, THE [@WARHAMMER]  P
EARL, ROBERT  SAVAGE CITY [@WARHAMMER]  P
EARL, ROBERT  WILD KINGDOMS [@WARHAMMER]  C/P,P
EARLE, RICHARD  FOREVER EMBER  SR-L
EARLL, TONY  MU REVEALED  [C/RAN-P]
EARLS, BILL  GLADIATOR, THE  P
EARNSHAW, BRIAN  DRAGONFALL 5 AND THE HIJACKERS  SR-P
EARNSHAW, BRIAN  DRAGONFALL 5 AND THE ROYAL BEAST  SR-P
EARNSHAW, BRIAN  DRAGONFALL 5 AND THE SPACE COWBOYS  SR-P
EARNSHAW, BRIAN  PIANETA NELL’OCCHIO DEL TEMPO, IL  FP
EARTHLING PUBLICATIONS  EARTHLING SAMPLER, THE  B
EASTERMAN, DANIEL  NIGHT OF THE APOCALYPSE  H
EASTON, EDWARD  MISCAST GENTLEMAN, THE  SR-P
EASTON, M. COLEMAN  FISHERMAN’S CURSE, THE  C/P,P
EASTON, M. COLEMAN  ISKIIR  C/P,P
EASTON, M. COLEMAN  MASTERS OF GLASS  C/P,P
EASTON, M. COLEMAN  SPIRITS OF CAVERN AND HEARTH  H
EASTON, M. COLEMAN  SWIMMERS BENEATH THE BRIGHT  C/P,P
EASTON, THOMAS A.  ALIEN RESONANCE  H,LP
EASTON, THOMAS A.  BIGFOOT STALKS THE COAST OF MAINE  LP
EASTON, THOMAS A.  ELECTRIC GENE MACHINE, THE [ORGANIC FUTURE]  LP
EASTON, THOMAS A.  GREENHOUSE [ORGANIC FUTURE]  C/P:2,P
EASTON, THOMAS A.  SEEDS OF DESTINY [ORGANIC FUTURE]  P
EASTON, THOMAS A.  SPARROWHAWK [ORGANIC FUTURE]  C/P,P
EASTON, THOMAS A.  TOWER OF THE GODS [ORGANIC FUTURE]  P
EASTON, THOMAS A.  UNTO THE LAST GENERATION  LP
EASTON, THOMAS A.  WOODSMAN [ORGANIC FUTURE]  C/P:2,P
EBBS, PAUL  BOOK OF THE STILL, THE [DOCTOR WHO EIGHTH DOCTOR]  P
EBLY, P.  CIUDAD QUE NO EXISTIA, LA  FH
EBON, MARTIN  BEYOND SPACE AND TIME: AN ESP CASEBOOK  [C/RAN-H]
EBON, MARTIN  CLONING OF MAN, THE  [C/RAN-P]
EBON, MARTIN  TEST YOUR ESP  [C/RAN-P]
ECCARIUS, J. G.  LAST DAYS OF CHRIST THE VAMPIRE  P
ECCARIUS, J. G.  RESURRECTION 2027  P
ECKERT, ALLAN W.  HAB THEORY, THE  C/P,P,H
ECKERT, WIN SCOTT  MYTHS FOR THE MODERN AGE  LPA
ECKLAR, JULIA  DIVINE INTERVENTION (CD)  SR-MM
ECKLAR, JULIA  KOBAYASHI MARU, THE [STAR TREK]  C/P:2,P
ECKLAR, JULIA  REGENESIS  C/P:2,P
ECKSTROM, JACK DENNIS  TIME OF THE HEDRONS, THE  H
ECO, UMBERTO  FOUCAULT’S PENDULUM  C/P,P,H
ECO, UMBERTO  ISLAND OF THE DAY BEFORE, THE  H,LP
ECO, UMBERTO  NAME OF THE ROSE, THE  H,P
EDDINGS, DAVID/EDDINGS, LEIGH  BELGARATH THE SORCERER [BELGARIAD/MALLOREON]  H:2,P
EDDINGS, DAVID  BELGARIAD 1, THE [BELGARIAD 1,2,3]  H:2
EDDINGS, DAVID  BELGARIAD 2, THE [BELGARIAD 4,5]  H:2
EDDINGS, DAVID  CASTLE OF WIZARDRY [BELGARIAD 4]  C/P:3,P
EDDINGS, DAVID/EDDINGS, LEIGH  CRYSTAL GORGE [@DREAMERS #3]  H
EDDINGS, DAVID  DEMON LORD OF KARANDA [MALLOREON 3]  H:2,P
EDDINGS, DAVID  DIAMOND THRONE, THE [ELENIUM 1]  H:2,P
EDDINGS, DAVID  DOMES OF FIRE [TAMULI 1]  H:2,LP
EDDINGS, DAVID/EDDINGS, LEIGH  ELDER GODS, THE [@DREAMERS #1]  H:2
EDDINGS, DAVID/EDDINGS, LEIGH  ELDER GODS, THE (PROOF) [DREAMERS 1]  SR-LP
EDDINGS, DAVID  ENCHANTERS’ END GAME [BELGARIAD 5]  C/P:2,P
EDDINGS, DAVID  GUARDIANS OF THE WEST [MALLOREON 1]  C/P,H,P,SFWA
EDDINGS, DAVID  HIDDEN CITY, THE [TAMULI 3]  H:2
EDDINGS, DAVID  KING OF THE MURGOS [MALLOREON 2]  H:2,P
EDDINGS, DAVID  LOSERS, THE  P
EDDINGS, DAVID  LOSERS, THE & HIGH HUNT  H
EDDINGS, DAVID  MAGICIAN’S GAMBIT [BELGARIAD 3]  C/P:2,P
EDDINGS, DAVID  PAWN OF PROPHECY [BELGARIAD 1]  C/P:2,P
EDDINGS, DAVID/EDDINGS, LEIGH  POLGARA THE SORCERESS [BELGARIAD/MALLOREON]  H:3
EDDINGS, DAVID  QUEEN OF SORCERY [BELGARIAD 2]  C/P:2,P
EDDINGS, DAVID/EDDINGS, LEIGH  REDEMPTION OF ALTHALUS, THE  H:2,P
EDDINGS, DAVID/EDDINGS, LEIGH  REGINA’S SONG  H
EDDINGS, DAVID/EDDINGS, LEIGH  RIVAN CODEX, THE [BELGARIAD/MALLOREON]  P
EDDINGS, DAVID  RUBY KNIGHT, THE [ELENIUM 2]  H:2,P

AUTHOR  TITLE  CODE
EDDINGS, DAVID  SAPPHIRE ROSE, THE [ELENIUM 3]  C/P,P,H
EDDINGS, DAVID  SEERESS OF KELL [MALLOREON 5]  C/P,H:2,P
EDDINGS, DAVID  SHINING ONES, THE [TAMULI 2]  H:2,P
EDDINGS, DAVID  SORCERESS OF DARSHIVA [MALLOREON 4]  H:2,P
EDDINGS, DAVID/EDDINGS, LEIGH  TREASURED ONE, THE [@DREAMERS #2]  H
EDDINGS, DAVID/EDDINGS, LEIGH  YOUNGER GODS, THE [@DREAMERS #4]  H
EDDISON, E. R.  FISH DINNER IN MEMISON, A [ZIMIAMVIA 2]  SR-H,C/P:2,P
EDDISON, E. R.  MEZENTIAN GATE, THE [ZIMIAMVIA 3]  C/P:2,P
EDDISON, E. R.  MISTRESS OF MISTRESSES [ZIMIAMVIA 1]  SR-H,C/P:2,P
EDDISON, E. R.  WORM OUROBOROS, THE [ZIMIAMVIA]  H:3,P
EDDISON, E. R.  ZIMIAMVIA [ZIMIAMVIA 1,2,3]  LP
EDDY, C. M. JR.  EXIT INTO ETERNITY  SR-LP
EDDY, MURIEL E./EDDY, C. M. JR.  GENTLEMAN FROM ANGELL STREET, THE  R/REF-P
EDELMAN, DAVID LOUIS  INFOQUAKE [JUMP 225 1]  LP
EDELSON, EDWARD  GREAT SCIENCE FICTION FROM THE MOVIES  R/REF-P
EDELSON, EDWARD  VISIONS OF TOMORROW: GREAT SCIENCE FICTION FROM THE MOVIES  R/REF-H
EDGERTON, TERESA  CASTLE OF THE SILVER WHEEL, THE [CELYDONN GWENLLIANT 1]  SFWA-P:2
EDGERTON, TERESA  CHILD OF SATURN [CELYDONN GREEN LION 1]  C/P:2,SFWA-P,P
EDGERTON, TERESA  GNOME’S ENGINE, THE [GOBLIN MOON 2]  C/P:2,SFWA-P:2,P
EDGERTON, TERESA  GOBLIN MOON [GOBLIN MOON 1]  C/P:2,SFWA-P,P
EDGERTON, TERESA  GRAIL AND THE RING, THE [CELYDONN GWENLLIANT 2]  C/P,SFWA-P,P
EDGERTON, TERESA  MOON AND THE THORN, THE [CELYDONN GWENLLIANT 3]  C/P,P
EDGERTON, TERESA  MOON IN HIDING, THE [CELYDONN GREEN LION 2]  C/P:2,SFWA-P,P
EDGERTON, TERESA  QUEEN’S NECKLACE, THE  LP
EDGERTON, TERESA  WORK OF THE SUN, THE [CELYDONN GREEN LION 3]  P,SFWA-P:2
EDGHILL, ROSEMARY  CLOAK OF NIGHT AND DAGGERS, THE [TWELVE TREASURES 3]  P
EDGHILL, ROSEMARY  CUP OF MORNING SHADOWS, THE [TWELVE TREASURES 2]  C/P:2,P
EDGHILL, ROSEMARY  MURDER BY MAGIC: TWENTY TALES OF CRIME AND THE SUPERNATURAL LPA
EDGHILL, ROSEMARY  SWORD OF MAIDEN’S TEARS, THE [TWELVE TREASURES 1]  C/P,P
EDGHILL, ROSEMARY  VENGEANCE OF MASKS [CHILDERIC THE SHATTERER 1]  H
EDGHILL, ROSEMARY  WARSLAYER, THE  P
EDGLEY, LESLIE  JUDAS GOAT, THE  SR-DD013
EDMONDSON, G. C.  ALUMINUM MAN, THE  C/P:2,P
EDMONDSON, G. C./KOTLAN, C. M.  BLACK MAGICIAN, THE  P
EDMONDSON, G. C.  BLUE FACE  C/P:2,P
EDMONDSON, G. C.  CHAPAYECA  H,SFWA
EDMONDSON, G. C./KOTLAN, C. M.  CUNNINGHAM EQUATIONS, THE  C/P:2,P
EDMONDSON, G. C.  MAN WHO CORRUPTED EARTH, THE  C/P,P
EDMONDSON, G. C./KOTLAN, C. M.  MAXIMUM EFFORT  P
EDMONDSON, G. C.  SHIP THAT SAILED THE TIME STREAM, THE [TIME STREAM 1]  DM109
EDMONDSON, G. C.  SHIP THAT SAILED THE TIME STREAM, THE (REVISED) [TIME STREAM 1]  C/P:3,P
EDMONDSON, G. C.  STRANGER THAN YOU THINK  DM109
EDMONDSON, G. C./KOTLAN, C. M.  TAKEOVER, THE  P
EDMONDSON, G. C.  T.H.E.M.  H
EDMONDSON, G. C.  TO SAIL THE CENTURY SEA [TIME STREAM 2]  C/P:2,P
EDSON, J. T.  BUNDUKI  C/P,P
EDSON, J. T.  FEARLESS MASTER OF THE JUNGLE  SR-P
EDWARDS, CLAUDIA J.  BRIGHT AND SHINING TIGER  P
EDWARDS, CLAUDIA J.  ELDRIE THE HEALER [@BASTARD PRINCESS #1]  P
EDWARDS, CLAUDIA J.  HORSEWOMAN IN GODSLAND, A  C/P,P
EDWARDS, CLAUDIA J.  TAMING THE FOREST KING  C/P,P
EDWARDS, DAVID  NEXT STOP — MARS!  H
EDWARDS, FRANK  FLYING SAUCERS — HERE AND NOW!  [C/RAN-H]
EDWARDS, FRANK  FLYING SAUCERS — SERIOUS BUSINESS  [C/RAN-H,C/RAN-P]
EDWARDS, FRANK  STRANGE WORLD  [C/RAN-P:2]
EDWARDS, FRANK  STRANGER THAN SCIENCE  [C/RAN-P]
EDWARDS, FRANK  STRANGEST OF ALL  [C/RAN-P:2]
EDWARDS, GAVIN  ’SCUSE ME WHILE I KISS THIS GUY  [C/RAN-P]
EDWARDS, GRAHAM  DRAGONCHARM [FORTUNE 1]  C/P,P
EDWARDS, GRAHAM  DRAGONSTORM [FORTUNE 2]  P
EDWARDS, GRAHAM  STONE & SEA [STONE 2]  P
EDWARDS, GRAHAM  STONE & SKY [STONE 1]  P
EDWARDS, GRAHAM  STONE & SUN [STONE 3]  P
EDWARDS, NORMAN  INVASION FROM 2500  C/P,P
EDWARDS, PAUL  DEADLY CYBORGS, THE [JOHN EAGLE EXPEDITER]  P
EDWARDS, PETER  TERMINUS  H
EFFINGER, GEORGE ALEC  BIRD OF TIME, THE  H:2
EFFINGER, GEORGE ALEC  BUDAYEEN NIGHTS [MARID AUDRAN]  H
EFFINGER, GEORGE ALEC  DEATH IN FLORENCE  H,SFWA
EFFINGER, GEORGE ALEC  DIRTY TRICKS  H,SFWA
EFFINGER, GEORGE ALEC  ESCAPE TO TOMORROW [@PLANET OF THE APES #2]  C/P,P
EFFINGER, GEORGE ALEC  EXILE KISS, THE [MARID AUDRAN 3]  C/P,LP,H
EFFINGER, GEORGE ALEC  FIRE IN THE SUN, A [MARID AUDRAN 2]  C/P,P,H
EFFINGER, GEORGE ALEC  HEROICS  H,SFWA

AUTHOR  TITLE  CODE
EFFINGER, GEORGE ALEC  IDLE PLEASURES  P
EFFINGER, GEORGE ALEC  IRRATIONAL NUMBERS  H,SFWA
EFFINGER, GEORGE ALEC  JOURNEY INTO TERROR [@PLANET OF THE APES #3]  C/P,P
EFFINGER, GEORGE ALEC  LIVE! FROM PLANET EARTH  H
EFFINGER, GEORGE ALEC  MAN THE FUGITIVE [@PLANET OF THE APES #1]  C/P:2,P
EFFINGER, GEORGE ALEC  MAUREEN BIRNBAUM, BARBARIAN SWORDSPERSON  H:2
EFFINGER, GEORGE ALEC  MIXED FEELINGS  H:2
EFFINGER, GEORGE ALEC  NICK OF TIME, THE  H,SFWA
EFFINGER, GEORGE ALEC  OLD FUNNY STUFF, THE [@AUTHOR’S CHOICE MONTHLY #1]  SR-H
EFFINGER, GEORGE ALEC  RELATIVES  H:2,SFWA,P
EFFINGER, GEORGE ALEC  SCHRODINGER’S KITTEN  B
EFFINGER, GEORGE ALEC  THOSE GENTLE VOICES  P
EFFINGER, GEORGE ALEC  THOUSAND DEATHS, A  H
EFFINGER, GEORGE ALEC  UTOPIA 3  P
EFFINGER, GEORGE ALEC  WHAT ENTROPY MEANS TO ME  C/P,SFWA,P,H
EFFINGER, GEORGE ALEC  WHEN GRAVITY FAILS [MARID AUDRAN 1]  C/P,SFWA,P,H
EFFINGER, GEORGE ALEC  WOLVES OF MEMORY, THE  H:2,P
EFFINGER, GEORGE ALEC  ZORK CHRONICLES, THE [ZORK]  C/P:2,P
EGAN, DORIS  COMPLETE IVORY, THE [IVORY 1,2,3]  P
EGAN, DORIS  GATE OF IVORY, THE [IVORY 1]  C/P:2,P
EGAN, DORIS  GUILT-EDGED IVORY [IVORY 3]  C/P:2,P
EGAN, DORIS  TWO-BIT HEROES [IVORY 2]  C/P,P
EGAN, GREG  AXIOMATIC  H
EGAN, GREG  DIASPORA  C/P,P,H
EGAN, GREG  DISTRESS  C/P,P
EGAN, GREG  LUMINOUS  H,P
EGAN, GREG  PERMUTATION CITY  C/P:2,P
EGAN, GREG  QUARANTINE  C/P,P,H
EGAN, GREG  SCHILD’S LADDER  H:2
EGAN, GREG  TERANESIA  C/P,P,H
EGAN, KEVIN  PERSEUS BREED, THE  C/P,P
EGLETON, CLIVE  JUDAS MANDATE, THE  P
EGLETON, CLIVE  LAST POST FOR A PARTISAN  P
EGLETON, CLIVE  PIECE OF RESISTANCE, A  P
EHRENREICH, BARBARA  KIPPER’S GAME  H
EHRENSVARD, GOSTA  MAN ON ANOTHER WORLD  [C/RAN-H]
EHRLICH, MAX  BIG EYE, THE  C/P,SR-P,H:2
EHRLICH, MAX  EDICT, THE  H:2,P
EHRLICH, MAX  REINCARNATION IN VENICE  P
EHRLICH, MAX  REINCARNATION OF PETER PROUD, THE  H
EHRLICH, PAUL R.  END OF AFFLUENCE, THE  [C/RAN-P]
EICHNER, HENRY M.  ATLANTEAN CHRONICLES  R/REF-H
EILERS, ROBERT  HERMES STONE, THE  P
EINSTEIN, CHARLES  DAY NEW YORK WENT DRY, THE  C/P:2,P
EINSTEIN, XAVIER  TRIVIA MANIA: STAR TREK  [C/RAN-P]
EISELEY, LOREN  IMMENSE JOURNEY, THE  [C/RAN-P:2]
EISELEY, LOREN  INVISIBLE PYRAMID, THE  [C/RAN-H]
EISENBERG, LARRY  BEST LAID SCHEMES, THE  H,P
EISENSTEIN, PHYLLIS  BORN TO EXILE  C/P,P,H
EISENSTEIN, PHYLLIS  CRYSTAL PALACE, THE [CRAY ORMORU 2]  C/P:2,P
EISENSTEIN, PHYLLIS  IN THE HANDS OF GLORY  C/P,P
EISENSTEIN, PHYLLIS  IN THE RED LORD’S REACH  C/P,SFWA-P,P
EISENSTEIN, PHYLLIS  SHADOW OF EARTH  C/P:2,P
EISENSTEIN, PHYLLIS  SORCERER’S SON [CRAY ORMORU 1]  C/P:2,P
EISENSTEIN, PHYLLIS  SORCERER’S SON (PROOF) [CRAY ORMORU 1]  SR-LP
EISLER, STEVEN  SPACE WARS: WORLDS AND WEAPONS  A
EISNER, SIMON  NAKED STORM, THE  P
EKLUND, GORDON  ALL TIMES POSSIBLE  C/P,P
EKLUND, GORDON  BEYOND THE RESURRECTION  H,SFWA
EKLUND, GORDON  BLACK KNIGHT OF THE IRON SPHERE [@LORD TEDRIC #3]  H,LP,P
EKLUND, GORDON  DANCE OF THE APOCALYPSE  SR-P
EKLUND, GORDON  DEVIL WORLD [STAR TREK]  C/P:2,P
EKLUND, GORDON  ECLIPSE OF DAWN, THE  C/P,P
EKLUND, GORDON  FALLING TOWARD FOREVER  SR-P,C/P
EKLUND, GORDON  GARDEN OF WINTER, THE  P
EKLUND, GORDON  GRAYSPACE BEAST, THE  H:2,SFWA,P
EKLUND, GORDON/ANDERSON, POUL  INHERITORS OF EARTH  C/P,P,H
EKLUND, GORDON  LORD TEDRIC [LORD TEDRIC #1]  H,@P,@C/P:2
EKLUND, GORDON  SERVING IN TIME  SR-P,C/P
EKLUND, GORDON  SPACE PIRATES [@LORD TEDRIC #2]  P
EKLUND, GORDON  STARLESS WORLD, THE [STAR TREK]  C/P:2,LP
EKLUND, GORDON  THUNDER ON NEPTUNE, A  H:2
EKLUND, GORDON  TRACE OF DREAMS, A  C/P:2,P
EKLUND, GORDON  TWILIGHT RIVER, THE  BS2

AUTHOR  TITLE  CODE
ELAM, RICHARD M. JR.  SCIENCE FICTION STORIES  C/P:2,LP
ELAM, RICHARD M. JR.  SUPER SCIENCE STORIES  LP
ELAM, RICHARD M. JR.  YOUNG VISITOR TO MARS  H
ELDER, JOSEPH  EROS IN ORBIT  HA,PA,SFWA
ELDER, JOSEPH  FARTHEST REACHES, THE  C/PA,PA
ELDER, MICHAEL  ALIEN EARTH, THE  C/P,P
ELDER, MICHAEL  FLIGHT TO TERROR  C/P,P
ELDER, MICHAEL  NOWHERE ON EARTH  C/P,P
ELDER, MICHAEL  PARADISE IS NOT ENOUGH  C/P,P
ELFLANDSSON, GALAD  BLACK WOLF, THE  SR-H
ELGIN, SUZETTE HADEN  AND THEN THERE’LL BE FIREWORKS [OZARK 3]  C/P,SFWA,P,H
ELGIN, SUZETTE HADEN  AT THE SEVENTH LEVEL [COYOTE JONES 3]  C/P,P
ELGIN, SUZETTE HADEN  COMMUNIPATH WORLDS [COYOTE JONES 1,2,3]  C/P:2,P
ELGIN, SUZETTE HADEN  COMMUNIPATHS, THE [COYOTE JONES 1]  D11560
ELGIN, SUZETTE HADEN  EARTHSONG [@NATIVE TONGUE #3]  C/P:2,P
ELGIN, SUZETTE HADEN  GRAND JUBILEE, THE [OZARK 2]  C/P,SFWA,P,H
ELGIN, SUZETTE HADEN  FURTHEST [COYOTE JONES 2]  C/P:2,P
ELGIN, SUZETTE HADEN  JUDAS ROSE, THE [NATIVE TONGUE 2]  C/P:2,P
ELGIN, SUZETTE HADEN  NATIVE TONGUE [NATIVE TONGUE 1]  C/P:2,P
ELGIN, SUZETTE HADEN  OZARK TRILOGY, THE [OZARK 1,2,3]  H:2
ELGIN, SUZETTE HADEN  STAR-ANCHORED, STAR-ANGERED [COYOTE JONES 4]  H:2,SFWA,P
ELGIN, SUZETTE HADEN  TWELVE FAIR KINGDOMS [OZARK 1]  C/P,SFWA,P
ELGIN, SUZETTE HADEN  YONDER COMES THE OTHER END OF TIME [COYOTE JONES 5]  C/P:2,P
ELIAS, ALBERT J.  BOWMAN TEST, THE  C/P,P
ELIOT, E. C.  KEMLO AND THE CRAZY PLANET  H
ELIOT, E. C.  KEMLO AND THE STAR MEN  SR-P
ELIOT, E. C.  KEMLO AND THE ZONES OF SILENCE  H
ELKIN, STANLEY  LIVING END, THE  P
ELLIK, RON/EVANS, BILL  UNIVERSES OF E. E. SMITH, THE  R/REF-H
ELLIOT, GREG  DEMON DOPPELGANGERS [@CHARMED]  P
ELLIOT, JEFFREY M.  KINDRED SPIRITS  LP
ELLIOT, JEFFREY M.  SCIENCE FICTION VOICES #4  R/REF-P
ELLIOT, LEE  BIO-MUTON  SR-P
ELLIOT, LEE  OVERLORD NEW YORK  SR-P
ELLIOTT, BRUCE  ASYLUM EARTH  C/P:2,P
ELLIOTT, BRUCE  RIVET IN GRANDFATHER’S NECK, THE  C/P:2,P
ELLIOTT, ELTON  NANODREAMS  PA,SFWA-P
ELLIOTT, H. CHANDLER  REPRIEVE FROM PARADISE  H
ELLIOTT, JOHN  DRAGON FEAST  P
ELLIOTT, KATE  BURNING STONE, THE [CROWN OF STARS 3]  C/P,P,H
ELLIOTT, KATE  CHILD OF FLAME [CROWN OF STARS 4]  H:2
ELLIOTT, KATE  CROWN OF STARS [CROWN OF STARS 7]  H
ELLIOTT, KATE  EARTHLY CROWN, AN [@SWORD OF HEAVEN #1]  C/P:2,P
ELLIOTT, KATE  GATHERING STORM, THE [CROWN OF STARS 5]  H
ELLIOTT, KATE  HIS CONQUERING SWORD [@SWORD OF HEAVEN #2]  C/P:2,SFWA-P,P
ELLIOTT, KATE  IN THE RUINS [CROWN OF STARS 6]  H
ELLIOTT, KATE  JARAN [JARAN 1]  C/P:2,SFWA-P,P
ELLIOTT, KATE  KING’S DRAGON [CROWN OF STARS 1]  C/P,P,H
ELLIOTT, KATE  LAW OF BECOMING, THE [@JARAN #4]  C/P,P
ELLIOTT, KATE  PRINCE OF DOGS [CROWN OF STARS 2]  C/P,P,H
ELLIOTT, KATE  SPIRIT GATE [CROSSROADS 1]  H
ELLIOTT, RICHARD  BURNT LANDS, THE  P
ELLIOTT, RICHARD  SWORD OF ALLAH, THE  P
ELLIS, A. C.  WORLDMAKER  C/P,P
ELLIS, ALBERT C./SLATEN, JEFF  DEATH JAG  P
ELLIS, NOVALYNE PRICE  ONE WHO WALKED ALONE: ROBERT E. HOWARD  R/REF-H
ELLIS, PETER BERRESFORD  H. RIDER HAGGARD: A VOICE FROM THE INFINITE  R/REF-H
ELLIS, PETER BERRESFORD  LAST ADVENTURER: THE LIFE OF TALBOT MUNDY, THE  R/REF-H
ELLIS, WALTER  BEWARE TOMORROW  P
ELLISON, HARLAN  AGAIN, DANGEROUS VISIONS  HA,SFWA
ELLISON, HARLAN  AGAIN, DANGEROUS VISIONS 2  PA,C/PA
ELLISON, HARLAN  ALONE AGAINST TOMORROW  C/P,P,H:2
ELLISON, HARLAN  ANGRY CANDY  H,LP
ELLISON, HARLAN  APPROACHING OBLIVION  H:2,SFWA,P
ELLISON, HARLAN  BEAST THAT SHOUTED LOVE AT THE HEART OF THE WORLD, THE  H:2,P
ELLISON, HARLAN  CITY ON THE EDGE OF FOREVER, THE [STAR TREK]  SR-S
ELLISON, HARLAN  DANGEROUS VISIONS  HA,PA
ELLISON, HARLAN  DANGEROUS VISIONS #1  PA
ELLISON, HARLAN  DANGEROUS VISIONS #2  PA
ELLISON, HARLAN  DANGEROUS VISIONS #3  PA
ELLISON, HARLAN  DEADLY STREETS, THE  C/P,P
ELLISON, HARLAN  DEATHBIRD STORIES  SR-H,C/P,SFWA-P,P
ELLISON, HARLAN  DEATHBIRD STORIES (CORRECTED)  H:2
ELLISON, HARLAN  DOOMSMAN  SR-DB0779,SR-DB0244

AUTHOR  TITLE  CODE
ELLISON, HARLAN  EARTHMAN, GO HOME!  C/P,P
ELLISON, HARLAN  EDGE IN MY VOICE, AN  R/REF-H,C/REF-H:2
ELLISON, HARLAN  ELLISON WONDERLAND  C/P,P,H
ELLISON, HARLAN  ESSENTIAL ELLISON, THE (REVISED, EXPANDED)  LP
ELLISON, HARLAN  FROM THE LAND OF FEAR  C/P:2,P
ELLISON, HARLAN  GENTLEMAN JUNKIE  P
ELLISON, HARLAN  GLASS TEAT, THE  [C/RAN-P:2]
ELLISON, HARLAN  I HAVE NO MOUTH AND I MUST SCREAM  C/P:2,P
ELLISON, HARLAN/ASIMOV, ISAAC  I, ROBOT: THE ILLUSTRATED SCREENPLAY  LP
ELLISON, HARLAN  ILLUSTRATED HARLAN ELLISON, THE  CX
ELLISON, HARLAN  LOVE AIN’T NOTHING BUT SEX MISSPELLED  C/P:2,P
ELLISON, HARLAN  MAN WITH NINE LIVES, THE  DD413
ELLISON, HARLAN  MEDEA: HARLAN’S WORLD  HA,LPA
ELLISON, HARLAN  MEMOS FROM PURGATORY  [C/RAN-P:2]
ELLISON, HARLAN  NO DOORS, NO WINDOWS  P
ELLISON, HARLAN  OTHER GLASS TEAT, THE  [C/RAN-P:2]
ELLISON, HARLAN  OVER THE EDGE  C/P,P
ELLISON, HARLAN  PAINGOD  C/P:2,P
ELLISON, HARLAN  PARTNERS IN WONDER  C/P:2,P
ELLISON, HARLAN  RUN FOR THE STARS  TD32
ELLISON, HARLAN  SHATTERDAY  C/P,P,H
ELLISON, HARLAN  SLEEPLESS NIGHTS IN THE PROCRUSTEAN BED  R/REF-P
ELLISON, HARLAN  SLIPPAGE  LP
ELLISON, HARLAN  SPIDER KISS  C/P:2,P
ELLISON, HARLAN  STALKING THE NIGHTMARE  H,P
ELLISON, HARLAN  STRANGE WINE  C/P,SFWA,P,H
ELLISON, HARLAN  TOUCH OF INFINITY, A  DD413
ELLISON, HARLAN  TROUBLEMAKERS  LP
ELLISON, HARLAN  WEB OF THE CITY  P
ELMORE, LARRY/ELMORE, ROBERT  RUNES OF AUTUMN  P
ELPHINSTONE, MARGARET  INCOMER, THE  LP
ELRICK, GEORGE S.  MAJOR MATT MASON: MOON MISSION  BLB
ELRICK, GEORGE S.  SCIENCE FICTION HANDBOOK FOR READERS AND WRITERS  R/REF-H
ELROD, P. N.  ADVENTURES OF MYHR, THE  P
ELROD, P. N.  ART IN THE BLOOD [@VAMPIRE FILES #4]  C/P,P
ELROD, P. N.  BLOOD ON THE WATER [@VAMPIRE FILES #6]  C/P:2,P
ELROD, P. N.  BLOODCIRCLE [@VAMPIRE FILES #3]  C/P,P
ELROD, P. N.  BLOODLIST [@VAMPIRE FILES #1]  C/P:2,P
ELROD, P. N.  CHILL IN THE BLOOD, A [@VAMPIRE FILES 7]  H
ELROD, P. N.  DANCE OF DEATH [JONATHAN BARRETT 4]  C/P,P
ELROD, P. N.  DARK SLEEP, THE [VAMPIRE FILES 8]  H:2
ELROD, P. N.  DEATH AND THE MAIDEN [JONATHAN BARRETT 2]  C/P:2,P
ELROD, P. N.  DEATH MASQUE [JONATHAN BARRETT 3]  C/P,P
ELROD, P. N.  DRACULA IN LONDON  LPA
ELROD, P. N.  FIRE IN THE BLOOD [@VAMPIRE FILES #5]  C/P:2,P
ELROD, P. N.  I, STRAHD [@RAVENLOFT]  C/P:2,P
ELROD, P. N.  LADY CRYMSYN [@VAMPIRE FILES 9]  H
ELROD, P. N.  LIFEBLOOD [@VAMPIRE FILES #2]  C/P:2,P
ELROD, P. N.  MY BIG FAT SUPERNATURAL WEDDING  LPA
ELROD, P. N.  QUINCEY MORRIS, VAMPIRE  C/P:2,P
ELROD, P. N.  RED DEATH [JONATHAN BARRETT 1]  P,SFWA-P:2
ELROD, P. N.  SONG IN THE DARK [@VAMPIRE FILES 11]  H
ELROD, P. N./GREENBERG, MARTIN H.  TIME OF THE VAMPIRES, THE  C/PA,PA
ELTON, BEN  THIS OTHER EDEN  LP
ELWOOD, ROGER  ADRIFT IN SPACE  SR-HA
ELWOOD, ROGER/MOSKOWITZ, SAM  ALIEN EARTH AND OTHER STORIES  C/PA,PA
ELWOOD, ROGER  ALIEN WORLDS  PA,C/PA
ELWOOD, ROGER  AND WALK NOW GENTLY THROUGH THE FIRE  SR-HA,HA
ELWOOD, ROGER/GHIDALIA, VIC  ANDROIDS, TIME MACHINES AND BLUE GIRAFFES  HA
ELWOOD, ROGER  BERSERKERS, THE  SR-HA
ELWOOD, ROGER  CHILDREN OF INFINITY  HA,SFWA

ELWOOD, ROGER CHRONICLES OF A COMER AND OTHER RELIGIOUS SCIENCE FICTION

STORIES

LPA

ELWOOD, ROGER  CONTINUUM 1  C/PA,PA,SFWA
ELWOOD, ROGER  CONTINUUM 2  HA,PA,SFWA
ELWOOD, ROGER  CONTINUUM 3  C/PA,PA,SFWA
ELWOOD, ROGER  CONTINUUM 4  C/PA,PA
ELWOOD, ROGER  CRISIS  SR-HA
ELWOOD, ROGER  DEMON KIND  PA
ELWOOD, ROGER  DYSTOPIAN VISIONS  SR-HA
ELWOOD, ROGER/SILVERBERG, ROBERT  EPOCH  C/PA,PA
ELWOOD, ROGER  FAR SIDE OF TIME, THE  SR-HA,HA
ELWOOD, ROGER  FIFTY-METER MONSTERS AND OTHER HORRORS, THE  SR-PA
ELWOOD, ROGER  FLAME TREE PLANET  SR-PA
ELWOOD, ROGER  FUTURE CITY  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 132
AUTHOR  TITLE  CODE
ELWOOD, ROGER  FUTURE CORRUPTION  PA
ELWOOD, ROGER  FUTURE QUEST  PA,C/PA
ELWOOD, ROGER  FUTURELOVE  SR-HA,HA
ELWOOD, ROGER  GIFTS OF ASTI AND OTHER STORIES OF SCIENCE FICTION, THE  HA
ELWOOD, ROGER  GRADUATED ROBOT, THE  SR-HA
ELWOOD, ROGER/GHIDALIA, VIC  HORROR HUNTERS  C/PA,PA
ELWOOD, ROGER  HORROR TALES  SR-VLHA
ELWOOD, ROGER  INVASION OF THE ROBOTS  PA,C/PA
ELWOOD, ROGER  JOURNEY TO ANOTHER STAR  SR-HA
ELWOOD, ROGER  KILLER PLANTS, THE  SR-HA
ELWOOD, ROGER  LEARNING MAZE, THE  HA
ELWOOD, ROGER/GHIDALIA, VIC  LITTLE MONSTERS, THE  C/PA,PA
ELWOOD, ROGER  LONG NIGHT OF WAITING  SR-HA
ELWOOD, ROGER  MIND ANGEL, THE  SR-HA
ELWOOD, ROGER  MISSING WORLD, THE  SR-HA
ELWOOD, ROGER  MONSTER TALES  SR-VLHA
ELWOOD, ROGER/GHIDALIA, VIC  MORE LITTLE MONSTERS  PA
ELWOOD, ROGER  MORE SCIENCE FICTION TALES  HA,SR-VLHA
ELWOOD, ROGER  NEW MIND, THE [FRONTIERS 2]  C/PA,PA
ELWOOD, ROGER  NIGHT OF THE SPHINX  SR-HA
ELWOOD, ROGER  OMEGA  C/PA,PA,SFWA
ELWOOD, ROGER  OTHER SIDE OF TOMORROW, THE  C/PA,PA
ELWOOD, ROGER  SALVATION BEHIND BARS  [C/RAN-P]
ELWOOD, ROGER/KIDD, VIRGINIA  SAVING WORLDS  HA
ELWOOD, ROGER  SCIENCE FICTION TALES  SR-VLHA
ELWOOD, ROGER  SHOWCASE  SR-HA
ELWOOD, ROGER  SIX SCIENCE FICTION PLAYS  C/PA,PA
ELWOOD, ROGER  STRANGE GODS  PA
ELWOOD, ROGER/MOSKOWITZ, SAM  STRANGE SIGNPOSTS  HA
ELWOOD, ROGER  SURVIVAL FROM INFINITY  SR-HA
ELWOOD, ROGER  TEN TOMORROWS  C/PA,PA
ELWOOD, ROGER  TOMORROW  SR-HA
ELWOOD, ROGER  TOMORROW’S ALTERNATIVES [FRONTIERS 1]  HA,PA
ELWOOD, ROGER  TUNNEL, THE  SR-HA
ELWOOD, ROGER  VAMPIRES, WEREWOLVES AND OTHER MONSTERS  PA
ELWOOD, ROGER  VISIONS OF TOMORROW  PA,C/PA
ELWOOD, ROGER  WAY OUT  SR-HA
ELWOOD, ROGER  WORLD NAMED CLEOPATRA, A  PA,C/PA
ELWOOD, ROGER/KIDD, VIRGINIA  WOUNDED PLANET, THE  PA,C/PA
ELWOOD, ROGER/GHIDALIA, VIC  YOUNG DEMONS  PA,C/PA
ELY, DAVID  SECONDS  P
ELY, DAVID  TIME OUT  H
ELY, MICHAEL  CENTAURI DAWN [@ALPHA CENTAURI #1]  P
ELY, MICHAEL  DRAGON SUN [@ALPHA CENTAURI #2]  P
ELY, MICHAEL  TWILIGHT OF THE MIND [@ALPHA CENTAURI #3]  P
ELZE, WINIFRED  CHANGELING GARDEN, THE  H
EMERSON, CAROLINE D.  MAGIC TUNNEL, THE  C/P:2,LP
EMERSON, JANE  CITY OF DIAMOND  P
EMERSON, RU  AGAINST THE GIANTS [@GREYHAWK]  P
EMERSON, RU  ART OF THE SWORD, THE [@NIGHT-THREADS 5]  C/P,P
EMERSON, RU  CALLING OF THE THREE, THE [NIGHT-THREADS 1]  SFWA-P:2
EMERSON, RU  CRAFT OF LIGHT, THE [@NIGHT-THREADS 4]  P
EMERSON, RU  EMPTY THRONE, THE [@XENA]  P
EMERSON, RU  GO QUEST, YOUNG MAN [@XENA]  P
EMERSON, RU  HUNTRESS AND THE SPHINX, THE [@XENA]  P
EMERSON, RU  IN THE CAVES OF EXILE [NEDAO 2]  P,SFWA-P
EMERSON, RU  KEEP ON THE BORDERLANDS [@GREYHAWK]  P
EMERSON, RU  MASQUES [@BEAUTY AND THE BEAST]  P
EMERSON, RU  ON THE SEAS OF DESTINY  P
EMERSON, RU  ONE LAND, ONE DUKE [@NIGHT-THREADS #3]  C/P,P
EMERSON, RU  PRINCESS OF FLAMES, THE  P
EMERSON, RU  SCIENCE OF POWER, THE [@NIGHT-THREADS 6]  C/P,P
EMERSON, RU  SPELL BOUND  P,SFWA-P:2
EMERSON, RU  THIEF OF HERMES, THE [XENA]  P
EMERSON, RU  TO THE HAUNTED MOUNTAINS  P
EMERSON, RU  TWO IN HIDING, THE [@NIGHT-THREADS #2]  C/P,P
EMERY, CLAYTON  CARD MASTER [FANTASY ADVENTURES]  P
EMERY, CLAYTON  DANGEROUS GAMES [@FORGOTTEN REALMS ARCANE AGE NETHERIL #2]  P
EMERY, CLAYTON  FINAL SACRIFICE [MAGIC: THE GATHERING]  C/P:2,P
EMERY, CLAYTON/WAJENBERG, EARL  4-D FUNHOUSE, THE  P
EMERY, CLAYTON  HAZEZON [@MAGIC: THE GATHERING LEGENDS 3]  P
EMERY, CLAYTON  JEDIT [@MAGIC: THE GATHERING LEGENDS 2]  P
EMERY, CLAYTON  JOHAN [@MAGIC: THE GATHERING LEGENDS 1]  P

EMERY, CLAYTON MORTAL CONSEQUENCES

[@FORGOTTEN REALMS ARCANE AGE NETHERIL #3] P

AUTHOR  TITLE  CODE
EMERY, CLAYTON  OUTCASTS [RUNESWORD 1]  C/P,P
EMERY, CLAYTON  SHATTERED CHAINS [MAGIC: THE GATHERING]  C/P:2,P
EMERY, CLAYTON  STAR OF CURSRAH [FORGOTTEN REALMS LOST EMPIRES]  P
EMERY, CLAYTON  SWORD PLAY [@FORGOTTEN REALMS ARCANE AGE NETHERIL #1]  P
EMERY, CLAYTON  TALES OF ROBIN HOOD  P,SFWA-P
EMERY, CLAYTON  WHISPERING WOODS [MAGIC: THE GATHERING]  C/P:2,P
EMMERSON, STEVE  CASUALTIES OF WAR [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
EMMERSON, STEVE  DARK PROGENY [DOCTOR WHO EIGHTH DOCTOR]  P
EMMS, WILLIAM  GALAXY FOUR [DOCTOR WHO]  SR-P,C/P
EMMS, WILLIAM  MISSION TO VENUS [DOCTOR WHO]  [C/RAN-P]
EMSHWILLER, CAROL  CARMEN DOG  LP
EMSHWILLER, CAROL  I LIVE WITH YOU  LP
EMSHWILLER, CAROL  MOUNT, THE  LP
EMSHWILLER, CAROL  REPORT TO THE MEN’S CLUB  LP
EMSHWILLER, CAROL  SECRET CITY, THE  LP
EMSHWILLER, PETER R.  HOST, THE  C/P,P
EMSHWILLER, PETER R.  SHORT BLADE  P
EMTSEV, MIKHAIL/PARNOV, EREMEI  WORLD SOUL  H:2,LP
ENDE, MICHAEL  MOMO  SFWA,LP
ENDE, MICHAEL  NEVERENDING STORY, THE  H:2,LP
ENDORE, GUY  METHINKS THE LADY...  H
ENDORE, GUY  WEREWOLF OF PARIS, THE  SR-P,H:2
ENEY, DICK  FANCYCLOPEDIA II  R/REF-H
ENEY, DICK  PROCEEDINGS; DISCON, THE  R/REF-P
ENGDAHL, SYLVIA LOUISE  ANYWHERE, ANYWHEN  SR-HA
ENGDAHL, SYLVIA LOUISE  BEYOND THE TOMORROW MOUNTAINS [NOREN 2]  H
ENGDAHL, SYLVIA LOUISE  DOORS OF THE UNIVERSE, THE [NOREN 3]  H
ENGDAHL, SYLVIA LOUISE  ENCHANTRESS FROM THE STARS  LP
ENGDAHL, SYLVIA LOUISE  FAR SIDE OF EVIL, THE  H:2,LP
ENGDAHL, SYLVIA LOUISE  JOURNEY BETWEEN WORLDS  H
ENGDAHL, SYLVIA LOUISE  THIS STAR SHALL ABIDE [NOREN 1]  H
ENGEL, JOEL  GENE RODDENBERRY: THE MYTH AND THE MAN BEHIND STAR TREK  R/REF-H
ENGEL, JOEL  ROD SERLING (BIOGRAPHY)  R/REF-H
ENGEL, LEONARD/PILLER, EMANUEL S.  WORLD AFLAME  SR-H
ENGH, M. J.  ARSLAN  C/P,SFWA,SFWA-P,P
ENGH, M. J.  RAINBOW MAN  LP
ENGH, M. J.  WHEEL OF THE WINDS  C/P,SFWA-P,P,H
ENGLAND, GEORGE ALLAN  AIR TRUST, THE  LP
ENGLAND, GEORGE ALLAN  DARKNESS AND DAWN  H,LP
ENGLAND, GEORGE ALLAN  GOLDEN BLIGHT, THE  H
ENGLAND, TERRY  REWIND  C/P,P
ENGLE, ELOISE  COUNTDOWN FOR CINDY  C/P:2,P
ENGLEBARDT, STANLEY L.  ELEKTRONICA  FP
ENGLEHART, STEPHEN  POINT MAN, THE  C/P,P
ENGLING, RICHARD  BODY MORTGAGE  P
ENGSTROM, ELIZABETH  BLACK AMBROSIA  SFWA-P
ENKE, BRIAN  SHADOWS OF MEDUSA  H,LP
ENSTROM, ROBERT  BETA COLONY  H,SFWA
ENSTROM, ROBERT  ENCOUNTER PROGRAM  H,SFWA
ENTWISTLE, BASIL/ROOTS, JOHN MCCOOK  MORAL RE-ARMAMENT — WHAT IS IT?  [C/RAN-P]
EOS BOOKS  EOS READER 2004-2005, THE  H:2,SR-LP
ERDMAN, PAUL E.  CRASH OF ’79, THE  P
ERENBERGER, TIMOTHY  ABACAR THE WIZAR [ABACAR 1]  LP
ERENBERGER, TIMOTHY  ABACAR THE WIZARD AND THE DARK LORD OF DRAGONS [ABACAR 2]  LP
ERICKSON, LYNN  PARADOX  P
ERICKSON, PAUL  ARK, THE [DOCTOR WHO]  SR-P,C/P
ERICKSON, STEVE  DAYS BETWEEN STATIONS  LP
ERICKSON, STEVE  TOURS OF THE BLACK CLOCK  LP
ERICSON, ERIC  SORCERER, THE  H
ERIKSON, STEVEN  BLOOD FOLLOWS [MALAZAN]  SR-LP
ERIKSON, STEVEN  BONEHUNTERS, THE [MALAZAN 6]  LP
ERIKSON, STEVEN  DEADHOUSE GATES [MALAZAN 2]  C/P,LP
ERIKSON, STEVEN  DEVIL DELIVERED, THE  SR-H
ERIKSON, STEVEN  GARDENS OF THE MOON [MALAZAN 1]  C/P,P,H:2
ERIKSON, STEVEN  HEALTHY DEAD, THE  SR-H
ERIKSON, STEVEN  HOUSE OF CHAINS [MALAZAN 4]  H
ERIKSON, STEVEN  MEMORIES OF ICE [MALAZAN 3]  H,LP
ERIKSON, STEVEN  MIDNIGHT TIDES [MALAZAN 5]  H
ERNSTING, WALTER  DAY THE GODS DIED, THE  P
ERSKINE, JIM/MORAN, GEORGE  FOLD A BANANA  [C/RAN-H]
ERSKINE, JOHN  ADAM AND EVE  H
ERTZ, SUSAN  WOMAN ALIVE  H
ESCHBACH, ANDREAS  CARPET MAKERS, THE  H
ESCHBACH, ANDREAS  JESUS VIDEO, DAS  FP

AUTHOR  TITLE  CODE
ESCHBACH, ANDREAS  SOLARSTATION & KELWITTS STERN  FP
ESENWEIN, V. BERG  ADVENTURES TO COME  SR-HA
ESFANDIARY, S. M.  OPTIMISM ONE  [C/RAN-P]
ESFANDIARY, S. M.  TELESPHERES  [C/RAN-P]
ESFANDIARY, S. M.  UP-WINGERS  [C/RAN-P]
ESHBACH, LLOYD ARTHUR  ARMLET OF THE GODS, THE [LAND BEYOND THE GATE 2]  P
ESHBACH, LLOYD ARTHUR  LAND BEYOND THE GATE, THE [LAND BEYOND THE GATE 1]  C/P,P
ESHBACH, LLOYD ARTHUR  OF WORLDS BEYOND: THE SCIENCE OF SCIENCE FICTION WRITING  R/REF-H
ESHBACH, LLOYD ARTHUR  OVER MY SHOULDER  R/REF-H
ESHBACH, LLOYD ARTHUR  SCROLL OF LUCIFER, THE [LAND BEYOND THE GATE 4]  C/P:2,P
ESHBACH, LLOYD ARTHUR  SORCERESS OF SCATH, THE [LAND BEYOND THE GATE 3]  P
ESHBACH, LLOYD ARTHUR  TYRANT OF TIME  SR-H
ESKRIDGE, KELLEY  DANGEROUS SPACE  LP
ESKRIDGE, KELLEY  SOLITAIRE  H
ESQUIVEL, LAURA  LAW OF LOVE, THE  H
ESRAC, WILLIAM  DANCE TO THE SUN  P
ESTES, ELEANOR  WITCH FAMILY, THE  LP
ESTES, ROSE  BLOOD OF THE TIGER [LOST LANDS 1]  P
ESTES, ROSE  BROTHER TO THE LION [LOST LANDS 2]  P
ESTES, ROSE  CHILDREN OF THE DRAGON  P
ESTES, ROSE  CIRCUS OF FEAR  P
ESTES, ROSE  DEMON HAND, THE [@GREYHAWK #5]  C/P:2,P
ESTES, ROSE  DUNGEON OF DREAD  [C/RAN-P:2]
ESTES, ROSE  ELFWOOD  P,SFWA-P
ESTES, ROSE  EYES HAVE IT, THE [GREYHAWK]  P
ESTES, ROSE  HUNTER, THE [HUNTER 1]  C/P,P
ESTES, ROSE  HUNTER ON ARENA, THE [HUNTER 2]  P
ESTES, ROSE  HUNTER VICTORIOUS, THE [HUNTER 3]  P
ESTES, ROSE  MASTER WOLF [@GREYHAWK #3]  C/P,P
ESTES, ROSE  MOUNTAIN OF MIRRORS  [C/RAN-P:2]
ESTES, ROSE  MOUNTAINS AND MADNESS [@IRON DRAGONS]  P,SFWA-P
ESTES, ROSE  NAME OF THE GAME, THE [@GREYHAWK #6]  C/P:2,P
ESTES, ROSE  PILLARS OF PENTEGARN  [C/RAN-P]
ESTES, ROSE  PRICE OF POWER [@GREYHAWK #4]  C/P:2,P
ESTES, ROSE  RETURN TO BROOKMERE  [C/RAN-P]
ESTES, ROSE  REVENGE OF THE RAINBOW DRAGONS  [C/RAN-P:2]
ESTES, ROSE  REVOLT OF THE DWARVES  [C/RAN-P]
ESTES, ROSE/WHAM, TOM  SKRYLING’S BLADE [RUNESWORD 2]  C/P,SFWA-P:2,P
ESTES, ROSE  SPIRIT OF THE HAWK [LOST LANDS 3]  P
ESTES, ROSE/WHAM, TOM  STONE OF TIME, THE [RUNESWORD 6]  P
ESTES, ROSE  TROLL-QUEST  P
ESTES, ROSE  TROLL-TAKEN  P
ESTEY, DALE  LOST TALE, A  P
ESTLEMAN, LOREN  SHERLOCK HOLMES VS. DRACULA  P
ETCHEMENDY, NANCY  STRANGER FROM THE STARS  LP
ETCHISON, CRAIG  WORLD WEAVER, THE  P
ETCHISON, DENNIS  CUTTING EDGE  SFWA
ETCHISON, DENNIS  DARK COUNTRY, THE  SR-H,H
ETCHISON, DENNIS  DARKSIDE  C/P:2,P
ETCHISON, DENNIS  FOG, THE  P
ETCHISON, DENNIS  MASTERS OF DARKNESS  SFWA-P
ETCHISON, DENNIS  MASTERS OF DARKNESS 2  SFWA-P
ETHERIDGE, RUTLEDGE  AGENT OF CHAOS [FLEETS OF THE GREAT BATTLE 2]  P
ETHERIDGE, RUTLEDGE  AGENT OF DESTRUCTION [FLEETS OF THE GREAT BATTLE 1]  P
ETHERIDGE, RUTLEDGE  BROTHER JOHN  P
ETHERIDGE, RUTLEDGE  FIRST DUELIST, THE  SFWA-P
ETHERIDGE, RUTLEDGE  LEGEND OF THE DUELIST  P,SFWA-P:2
ETKIN, ANNE  WATERSPOUT UP  B
EULO, KEN  HOUSE OF CAINE, THE  SFWA-P
EVANS, ARTHUR B.  JULES VERNE REDISCOVERED  R/REF-H
EVANS, CHRISTOPHER  CAPELLA’S GOLDEN EYES  P
EVANS, CHRISTOPHER  MIND AT BAY  SR-PA
EVANS, CHRISTOPHER  MIND IN CHAINS  SR-PA
EVANS, D. R.  PALINDOR [THREE LANDS 1]  LP
EVANS, D. R.  PHENDRIC [THREE LANDS 3]  LP
EVANS, D. R.  SHADOW [THREE LANDS 2]  LP
EVANS, DAVID  TIME STATION BERLIN [TIME STATION 3]  P
EVANS, DAVID  TIME STATION LONDON [TIME STATION 1]  C/P,P
EVANS, DAVID  TIME STATION PARIS [TIME STATION 2]  P
EVANS, E. EVERETT  ALIEN MINDS  SR-H
EVANS, E. EVERETT  FOOD FOR DEMONS  SR-LP
EVANS, E. EVERETT  MAN OF MANY MINDS  C/P:2,P
EVANS, HILARY/EVANS, DIK  BEYOND THE GASLIGHT: SCIENCE IN POPULAR FICTION 1895–1905  HA
EVANS, I. O.  SCIENCE FICTION THROUGH THE AGES 2  SR-PA

AUTHOR  TITLE  CODE
EVANS, IAN  STAR MAIDENS  SR-P,C/P
EVANS, JOHN  HALO IN BRASS  P
EVANS, LINDA  FAR EDGE OF DARKNESS  C/P,P
EVANS, LINDA  SLEIPNIR  P,SFWA-P
EVANS, PETER J.  UNQUIET GRAVE, THE [@DURHAM RED]  P
EVANS, ROBLEY  J. R. R. TOLKIEN  R/REF-P,C/REF-P
EWERS, H. G.  DIE MACHT DER SCHATTEN [@PERRY RHODAN #133]  FPR
EWERS, H. G.  DIE TODESMASCHINE [@PERRY RHODAN #119]  FPR
EWERS, H. G.  DIE WELT DER GLUCKSELIGEN [@PERRY RHODAN #65]  FPR
EWERS, H. G.  DIE ZEITSPRINGER [@PERRY RHODAN #7]  FPR
EWERS, H. G.  PHANTOM-STATION [@PERRY RHODAN #16]  FPR
EYRE, KATHERINE WIGMORE  LUTE AND THE GLOVE, THE  P
EYSENCK, H. J.  WAT IS UW I.Q.  FP

AUTHOR  TITLE  CODE
FABER, ARTHUR  OUTERSPACE SEX ORGY  SR-L
FABI, MARK  WYRM  LP
FADIMAN, CLIFTON  FANTASIA MATHEMATICA  H,LP
FADIMAN, CLIFTON  MATHEMATICAL MAGPIE, THE  [C/RAN-H]
FAIN, JEFF  BURNING, THE  P
FAIR, CHARLES  FROM THE JAWS OF VICTORY  [C/RAN-H]
FAIRBAIRNS, ZOE  BENEFITS  P
FAIRBANK, JOHN KING  UNITED STATES AND CHINA, THE  [C/RAN-P]
FAIRMAN, PAUL W.  BRIDGET LOVES BERNIE  P
FAIRMAN, PAUL W.  CITY UNDER THE SEA  P
FAIRMAN, PAUL W.  DIABOLIST, THE  P
FAIRMAN, PAUL W.  DOOMSDAY EXHIBIT, THE  C/P,P
FAIRMAN, PAUL W.  FRANKENSTEIN WHEEL, THE  P
FAIRMAN, PAUL W.  GIRL WITH SOMETHING EXTRA, THE  P
FAIRMAN, PAUL W.  I, THE MACHINE  C/P:2,P
FAIRMAN, PAUL W.  LOVE AMERICAN STYLE  P
FAIRMAN, PAUL W.  LOVE AMERICAN STYLE #2  P
FAIRMAN, PAUL W.  NEVER TRUST A MARTIAN  B
FAIRMAN, PAUL W.  REST IN AGONY  P
FAIRMAN, PAUL W.  SEARCH FOR A DEAD NYMPHO  P
FAIRMAN, PAUL W.  THAT GIRL  P
FAIRMAN, PAUL W.  WORLD GRABBERS, THE  P
FALCONER, SOVEREIGN  TO MAKE DEATH LOVE US  H,SFWA
FALK, LEE  KILLER’S TOWN [PHANTOM 9]  P
FALKNER, JOHN  OVERLORDS OF ANDROMEDA  SR-P
FALKNER, JOHN  UNTRODDEN STREETS OF TIME  SR-P
FALLAW, L. M.  UGGLIANS, THE  H
FALLETTA, NICHOLAS  PARADOXICON, THE  SFWA
FALLON, JENNIFER  EYE OF THE LABYRINTH [SECOND SONS 2]  P
FALLON, JENNIFER  HARSHINI [HYTHRUN 3]  H
FALLON, JENNIFER  LION OF SENET, THE [SECOND SONS 1]  P
FALLON, JENNIFER  LORD OF THE SHADOWS [SECOND SONS 3]  P
FALLON, JENNIFER  MEDALON [HYTHRUN 1]  H
FALLON, JENNIFER  TREASON KEEP [HYTHRUN 2]  H
FALLON, JENNIFER  WARRIOR [HYTHRUN WOLFBLADE 2]  H
FALLON, JENNIFER  WOLFBLADE [HYTHRUN WOLFBLADE 1]  H
FANCHER, JANE S.  GROUNDTIES  P
FANCHER, JANE S.  HARMONIES OF THE NET [GROUNDTIES 3]  C/P,P
FANCHER, JANE S.  RING OF DESTINY [DANCE OF THE RINGS 3]  P
FANCHER, JANE S.  RING OF INTRIGUE [DANCE OF THE RINGS 2]  C/P,P
FANCHER, JANE S.  RING OF LIGHTNING [DANCE OF THE RINGS 1]  P
FANCHER, JANE S.  UPLINK  C/P,P
FANE, BRON  BLUE JUGGERNAUT  H
FANE, BRON  INTRUDERS, THE  SR-P
FANE, BRON  LAST MAN ON EARTH  SR-P
FANE, BRON  MACABRE ONES, THE  SR-P
FANE, BRON  NEMESIS  SR-P
FANE, BRON  RODENT MUTATION  SR-P
FANE, BRON  SOFTLY BY MOONLIGHT  SR-P
FANE, BRON  SOMEWHERE OUT THERE  SR-P
FANE, BRON  SUSPENSION  SR-P
FANE, BRON  U.F.O. 517  SR-P
FANE, BRON  UNKNOWN DESTINY  SR-P
FANTASY FILMS  LORD OF THE RINGS, THE (MOVIE PROMOTION)  A
FANTHORPE, R. L.  ALIEN FROM THE STARS  H
FANTHORPE, R. L.  ASTEROID MAN  H
FANTHORPE, R. L.  GOLDEN CHALICE, THE  SR-P
FANTHORPE, R. L.  HAND OF DOOM  H
FANTHORPE, R. L.  HYPERSPACE  SR-P
FANTHORPE, R. L.  NEGATIVE MINUS  SR-P
FANTHORPE, R. L.  NEURON WORLD  SR-P
FANTHORPE, R. L.  OUT OF THE DARKNESS  SR-P
FANTHORPE, R. L.  SPACE FURY  SR-P,C/P:2
FANTHORPE, R. L.  TRIPLE MAN, THE  SR-P
FANTHORPE, R. L.  UNCONFINED, THE  SR-P
FANTHORPE, R. L.  WATCHING WORLD, THE  SR-P
FARADAY, ROBERT  ANYTIME RINGS, THE  LP
FARADAY, ROBERT  SAMAX, THE GLADIATOR  LP
FARCA, MARIE C.  COMPLEX MAN  H,SFWA
FARCA, MARIE C.  EARTH  H:2,SFWA
FARJEON, ELEANOR  GLASS SLIPPER, THE  P
FARJEON, J. JEFFERSON  DEATH OF A WORLD  H
FARLAND, DAVID  BROTHERHOOD OF THE WOLF [RUNELORDS 2]  H:2,P
FARLAND, DAVID  LAIR OF BONES, THE [RUNELORDS 4]  H

The MITSFS Pinkdex  September 24, 2007  Page 138
AUTHOR  TITLE  CODE
FARLAND, DAVID  RUNELORDS, THE [RUNELORDS 1]  C/P:2,P
FARLAND, DAVID  SONS OF THE OAK [RUNELORDS 5]  H
FARLAND, DAVID  WIZARDBORN [RUNELORDS 3]  P
FARLAND, DAVID  WORLDS OF THE GOLDEN QUEEN [GOLDEN QUEEN 1,2]  LP
FARLEY, RALPH M.  EARTH MAN ON VENUS, AN  SR-P
FARLEY, RALPH M.  HIDDEN UNIVERSE, THE  H
FARLEY, RALPH M.  OMNIBUS OF TIME, THE  H
FARLEY, RALPH M.  RADIO BEASTS, THE  C/P:2,P
FARLEY, RALPH M.  RADIO MAN, THE  SR-LP
FARLEY, RALPH M.  RADIO PLANET, THE  C/P:2,P
FARLEY, WALTER  ISLAND STALLION RACES, THE  H
FARMER, NANCY  HOUSE OF THE SCORPION, THE  H
FARMER, PHILIP JOSE  ADVENTURE OF THE PEERLESS PEER, THE  SR-H,C/P,P
FARMER, PHILIP JOSE  ALLEY GOD, THE  C/P:2,P
FARMER, PHILIP JOSE  BARNSTORMER IN OZ, A  LP
FARMER, PHILIP JOSE  BEHIND THE WALLS OF TERRA [WORLD OF TIERS 4]  C/P:2,P
FARMER, PHILIP JOSE  BLOWN [HERALD CHILDE 2]  SR-L
FARMER, PHILIP JOSE  BOOK OF PHILIP JOSE FARMER, THE  C/P,P
FARMER, PHILIP JOSE  CACHE, THE  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  CACHE FROM OUTER SPACE  DF165
FARMER, PHILIP JOSE  CELESTIAL BLUEPRINT AND OTHER STORIES, THE  DF165
FARMER, PHILIP JOSE  CLASSIC PHILIP JOSE FARMER 1952–1964, THE  H
FARMER, PHILIP JOSE  CLASSIC PHILIP JOSE FARMER 1964–1973, THE  H
FARMER, PHILIP JOSE  DARE  SR-P,C/P:2
FARMER, PHILIP JOSE  DARK DESIGN, THE [RIVERWORLD 3]  C/P,P,H
FARMER, PHILIP JOSE  DARK DESIGN, THE (PROOF) [RIVERWORLD 3]  SR-LP
FARMER, PHILIP JOSE  DARK HEART OF TIME, THE [@TARZAN]  P
FARMER, PHILIP JOSE  DARK IS THE SUN  H:2,P
FARMER, PHILIP JOSE  DAY OF THE TIMESTOP, THE  P
FARMER, PHILIP JOSE  DAYWORLD [DAYWORLD 1]  C/P,P,H
FARMER, PHILIP JOSE  DAYWORLD BREAKUP [DAYWORLD 3]  H:2,SFWA-P,P
FARMER, PHILIP JOSE  DAYWORLD REBEL [DAYWORLD 2]  C/P,SFWA-P,P,H

FARMER, PHILIP JOSE DOC SAVAGE: HIS APOCALYPTIC LIFE SFWA,R/REF-P,

C/REF-H:3,C/REF-P

FARMER, PHILIP JOSE  DOWN IN THE BLACK GANG  H:2,P
FARMER, PHILIP JOSE  ESCAPE FROM LOKI [DOC SAVAGE]  P
FARMER, PHILIP JOSE  FABULOUS RIVERBOAT, THE [RIVERWORLD 2]  C/P,P,H
FARMER, PHILIP JOSE  FATHER TO THE STARS  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  FEAST UNKNOWN, A  C/P,SR-L,P
FARMER, PHILIP JOSE  FLESH  C/P,SR-P,H
FARMER, PHILIP JOSE  FLESH/LORD TYGER  P
FARMER, PHILIP JOSE  FLIGHT TO OPAR  C/P,P
FARMER, PHILIP JOSE  GATE OF TIME, THE  C/P:2,P
FARMER, PHILIP JOSE  GATES OF CREATION, THE [WORLD OF TIERS 2]  C/P:2,P
FARMER, PHILIP JOSE  GODS OF RIVERWORLD [RIVERWORLD 5]  H:2,P
FARMER, PHILIP JOSE  GREATHEART SILVER  C/P:2,P
FARMER, PHILIP JOSE  GREEN ODYSSEY, THE  C/P,P,H
FARMER, PHILIP JOSE  HADON OF ANCIENT OPAR  C/P:2,P
FARMER, PHILIP JOSE  IMAGE OF THE BEAST, THE [HERALD CHILDE 1]  C/P,SR-L
FARMER, PHILIP JOSE  IMAGE OF THE BEAST [HERALD CHILDE 1,2]/BLOWN [HERALD CHILDE 1,2] C/P:2,P
FARMER, PHILIP JOSE  INSIDE OUTSIDE  H:2,P
FARMER, PHILIP JOSE  JESUS ON MARS  C/P:2,P
FARMER, PHILIP JOSE  LAVALITE WORLD, THE [WORLD OF TIERS 5]  C/P:2,P
FARMER, PHILIP JOSE  LORD OF THE TREES  D51375,P
FARMER, PHILIP JOSE  LORD OF THE TREES/MAD GOBLIN, THE  C/P,P
FARMER, PHILIP JOSE  LORD TYGER  C/P:2,P
FARMER, PHILIP JOSE  LOVERS, THE  C/P,P,H
FARMER, PHILIP JOSE  MAD GOBLIN, THE  C/P,D51375
FARMER, PHILIP JOSE  MAGIC LABYRINTH, THE [RIVERWORLD 4]  C/P,P,H,SFWA
FARMER, PHILIP JOSE  MAKER OF UNIVERSES, THE [WORLD OF TIERS 1]  C/P:2,P
FARMER, PHILIP JOSE  MNOGOYARUSNEE MIR  FH
FARMER, PHILIP JOSE  MORE THAN FIRE [WORLD OF TIERS 6]  H
FARMER, PHILIP JOSE  MOTHER WAS A LOVELY BEAST  C/PA,PA
FARMER, PHILIP JOSE  NIGHT OF LIGHT  C/P:2,P
FARMER, PHILIP JOSE  NOTHING BURNS IN HELL  P
FARMER, PHILIP JOSE  OTHER LOG OF PHILEAS FOGG, THE  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  PEARLS FROM PEORIA  H
FARMER, PHILIP JOSE  PRIVATE COSMOS, A [WORLD OF TIERS 3]  C/P:2,P
FARMER, PHILIP JOSE  PURPLE BOOK, THE  C/P:2,P
FARMER, PHILIP JOSE  QUEST TO RIVERWORLD [RIVERWORLD]  PA
FARMER, PHILIP JOSE  RED ORC’S RAGE  H,P
FARMER, PHILIP JOSE  RIDERS OF THE PURPLE WAGE  C/P,P
FARMER, PHILIP JOSE  RIVER OF ETERNITY  H
FARMER, PHILIP JOSE  RIVERWORLD AND OTHER STORIES [RIVERWORLD]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 139
AUTHOR  TITLE  CODE
FARMER, PHILIP JOSE  STATIONS OF THE NIGHTMARE  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  STONE GOD AWAKENS, THE  C/P:2,P
FARMER, PHILIP JOSE  STRANGE RELATIONS  C/P,P,H
FARMER, PHILIP JOSE  TALES OF RIVERWORLD [RIVERWORLD]  C/PA,PA
FARMER, PHILIP JOSE  TARZAN ALIVE  C/P,SFWA,P,H
FARMER, PHILIP JOSE  TIME’S LAST GIFT  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  TIMESTOP!  C/P,P
FARMER, PHILIP JOSE  TO YOUR SCATTERED BODIES GO [RIVERWORLD 1]  C/P,P,H
FARMER, PHILIP JOSE  TONGUES OF THE MOON  C/P:2,P
FARMER, PHILIP JOSE  TRAITOR TO THE LIVING  C/P:2,SFWA-P,P
FARMER, PHILIP JOSE  TWO HAWKS FROM EARTH  C/P:2,P
FARMER, PHILIP JOSE  UNREASONING MASK, THE  C/P,P,H
FARMER, PHILIP JOSE  VENUS ON THE HALF-SHELL  P
FARMER, PHILIP JOSE  WIND WHALES OF ISHMAEL, THE  C/P:2,P
FARMER, PHILIP JOSE  WOMAN A DAY, A  SR-P,C/P:2
FARMER, PHILIP JOSE  WORLD OF TIERS 1, THE [WORLD OF TIERS 1,2]  H:2
FARMER, PHILIP JOSE  WORLD OF TIERS 2, THE [WORLD OF TIERS 3,4,5]  H:2
FARRAND, PHIL  NITPICKER’S GUIDE FOR CLASSIC TREKKERS, THE [STAR TREK]  R/REF-H

FARRAND, PHIL NITPICKER’S GUIDE FOR NEXT GENERATION TREKKERS, THE

[STAR TREK TNG]

R/REF-H

FARRAR, STEWART  OMEGA  H
FARRAR, STEWART  SWORD OF ORLEY, THE  H
FARRAR, STEWART  TWELVE MAIDENS, THE  P
FARRELL, JACKSON T.  FLIGHT OF THE ENDEAVOR, THE  P
FARRELL, MATTHEW  THUNDER RIFT  P
FARRELL, S. L.  HEIR OF STONE [@CLOUDMAGES #3]  H
FARRELL, S. L.  HOLDER OF LIGHTNING [@CLOUDMAGES #1]  H
FARRELL, S. L.  HOLDER OF LIGHTNINGS (PROOF) [@CLOUDMAGES #1]  SR-LP
FARRELL, S. L.  MAGE OF CLOUDS [@CLOUDMAGES #2]  P
FARREN, MICK  ARMAGEDDON CRAZY, THE  C/P,P
FARREN, MICK  BACK FROM HELL [CAR WARRIORS 3]  P
FARREN, MICK  CITIZEN PHAID [PHAID 2]  C/P:2,P
FARREN, MICK  DARKLOST [VICTOR RENQUIST 1]  H
FARREN, MICK  FEELIES, THE  P
FARREN, MICK  KINDLING  H
FARREN, MICK  LAST STAND OF THE DNA COWBOYS, THE  C/P,P
FARREN, MICK  LONG ORBIT, THE  P
FARREN, MICK  MARS — THE RED PLANET  P
FARREN, MICK  MORE THAN MORTAL [VICTOR RENQUIST 2]  H
FARREN, MICK  NECROM  C/P,P
FARREN, MICK  PHAID THE GAMBLER [PHAID 1]  C/P:2,P
FARREN, MICK  PROTECTORATE  P
FARREN, MICK  TEXTS OF FESTIVAL, THE  C/P,P
FARREN, MICK  THEIR MASTER’S WAR  P
FARREN, MICK  UNDERLAND  H
FARREN, MICK  VICKERS  C/P:2,P
FARRER, MATTHEW  CROSSFIRE [@WARHAMMER 40K SHIRA CALPURNIA 1]  P
FARRER, MATTHEW  JUNKTION [@NECROMUNDA]  P
FARRER, MATTHEW  LEGACY [@WARHAMMER 40K SHIRA CALPURNIA 2]  P
FARRERE, CLAUDE  HOUSE OF THE SECRET, THE  SR-H
FARRINGTON, IAN  SHORT TRIPS: A DAY IN THE LIFE [DOCTOR WHO]  HA
FARRINGTON, IAN  SHORT TRIPS: MONTSERS [DOCTOR WHO]  HA
FARRINGTON, IAN  SHORT TRIPS: PAST TENSE [DOCTOR WHO]  HA
FARRINGTON, IAN  SHORT TRIPS: THE CENTENARIAN [DOCTOR WHO]  HA
FARRIS, JOHN  AXMAN COMETH, THE  SFWA-P
FARRIS, JOHN  CATACOMBS  SFWA-P
FARRIS, JOHN  FIENDS  SFWA-P
FARRIS, JOHN  FURY, THE  C/P:2,P
FARRIS, JOHN  SCARE TACTICS  SFWA-P
FARRIS, JOHN  SHARP PRACTICE  SFWA-P
FARRIS, JOHN  UNINVITED, THE  SFWA-P
FARRIS, JOHN  WILDWOOD  SFWA-P
FARSHTEY, GREG  HELL’S FEAST [BLOODSHADOWS]  LPA
FARSHTEY, GREG  RIVER OF GOD, THE [SHATTERZONE 3]  P
FARSHTEY, GREG  SHATTERED AND OTHER STORIES [SHATTERZONE]  LPA
FAST, HOWARD  AU SEUIL DU FUTUR  FP
FAST, HOWARD  EDGE OF TOMORROW, THE  C/P:2,P
FAST, HOWARD  GENERAL ZAPPED AN ANGEL, THE  C/P,P,H
FAST, HOWARD  TONY AND THE WONDERFUL DOOR  H
FAST, HOWARD  TOUCH OF INFINITY, A  C/P,P,H
FAST, JONATHAN  MORTAL GODS  P
FAST, JONATHAN  SECRETS OF SYNCHRONICITY, THE  C/P:2,P
FAST, JULIUS  LEAGUE OF GREY-EYED WOMEN, THE  H
FAST, JULIUS  OUT OF THIS WORLD  SR-PA,C/PA

The MITSFS Pinkdex  September 24, 2007  Page 140
AUTHOR  TITLE  CODE
FATE, EDITORS OF  BEYOND THE STRANGE  [C/RAN-P]
FAUCETTE, JOHN M.  AGE OF RUIN, THE  DH103
FAUCETTE, JOHN M.  CROWN OF INFINITY  DH051
FAUCETTE, JOHN M.  SIEGE OF EARTH  C/P:2,P
FAUCETTE, JOHN M.  WARRIORS OF TERRA, THE  C/P,P
FAURE, RAOUL C.  CAVE AND THE ROCK, THE  H
FAUST, JOE CLIFFORD  BODDEKKER’S DEMONS [PEMBROKE HALL 2]  P
FAUST, JOE CLIFFORD  COMPANY MAN, THE  P
FAUST, JOE CLIFFORD  DEATH OF HONOR, A  C/P,P,H
FAUST, JOE CLIFFORD  DESPERATE MEASURES [ANGEL’S LUCK 1]  C/P,P,H
FAUST, JOE CLIFFORD  ESSENCE OF EVIL, THE [ANGEL’S LUCK 3]  H,P
FAUST, JOE CLIFFORD  FERMAN’S DEVILS [PEMBROKE HALL 1]  P
FAUST, JOE CLIFFORD  HANDLING IT [PEMBROKE HALL 1,2]  H
FAUST, JOE CLIFFORD  PRECIOUS CARGO [ANGEL’S LUCK 2]  H,P
FAUST, MINISTER  COYOTE KINGS OF THE SPACE-AGE BACHELOR PAD, THE  LP
FAWCETT, BILL  CATS IN SPACE AND OTHER PLACES  PA,C/PA,SFWA-P
FAWCETT, BILL  COLD STEEL [@BOLOS #6]  PA
FAWCETT, BILL  DZURLORD  SFWA-P
FAWCETT, BILL  FAR STARS WAR, THE [@WAR YEARS #1]  PA
FAWCETT, BILL  HONOR OF THE REGIMENT [@BOLOS #1]  SFWA-P,C/PA,PA
FAWCETT, BILL  JUPITER WAR, THE [@WAR YEARS #3]  PA
FAWCETT, BILL  LAST STAND [@BOLOS #4]  C/PA,PA
FAWCETT, BILL/RANDALL, NEIL  LORD OF CRAGSCLAW [GUARDIANS OF THE THREE 1]  C/P:2,P
FAWCETT, BILL  OLD GUARD [@BOLOS #5]  PA
FAWCETT, BILL  SIEGE OF ARISTA, THE [@WAR YEARS #2]  C/PA,PA
FAWCETT, BILL  UNCONQUERABLE, THE [@BOLOS #2]  PA,C/PA

FAWCETT, BILL WARMASTERS, THE [HONOR HARRINGTON/HAMMER’S SLAMMERS/

BELISARIUS]

HA

FAWKES, RICHARD  FACE OF THE ENEMY  C/P,P
FAWKES, RICHARD  NATURE OF THE BEAST  P
FAX PUBLISHERS  FAMOUS FANTASTIC CLASSICS #1  SR-LP
FAX PUBLISHERS  FAMOUS FANTASTIC CLASSICS #2  SR-LP
FAYLE, HELEN/COX, SUE  NINE LIVES  SR-LPA
FAYLE, HELEN/EALES, JULIAN  PERFECT TIMING 2 [DOCTOR WHO]  SR-LPA
FBI  PHILIP K. DICK  R/REF-H
FEAR, W. H.  LUNAR FLIGHT  SR-P
FEAR, W. H.  OPERATION SATELLITE  SR-P
FEAR, W. H.  ULTIMATE, THE  SR-P,C/P
FEARN, JOHN RUSSELL  EMPEROR OF MARS [CLAYTON DREW]  SR-P
FEARN, JOHN RUSSELL  GODDESS OF MARS [CLAYTON DREW]  SR-P
FEARN, JOHN RUSSELL  LINERS OF TIME  H:2
FEARN, JOHN RUSSELL  MANTON’S WORLD  LP
FEARN, JOHN RUSSELL  OPERATION VENUS  SR-P
FEARN, JOHN RUSSELL  RED MEN OF MARS [CLAYTON DREW]  SR-P
FEARN, JOHN RUSSELL  SLAVES OF IJAX  SR-P
FEARN, JOHN RUSSELL  WARRIOR OF MARS [CLAYTON DREW]  SR-P
FEARON-PITMAN  GALAXY 5 (TEACHER’S GUIDE)  R/REF-P
FEARON-PITMAN  SPACE POLICE (TEACHER’S GUIDE)  R/REF-P
FEARON-PITMAN  SPECTER BOOKS (TEACHER’S GUIDE)  R/REF-P
FEDAHK, FATIMA  SEXORCIST, THE  SR-L
FEDERBUSH, ARNOLD  ICE!  P
FEDERBUSH, ARNOLD  MAN WHO LIVED IN INNER SPACE, THE  C/P,P
FEEHAN, CHRISTINE  CONSPIRACY GAME  P
FEEHAN, CHRISTINE  DARK CELEBRATION [CARPATHIAN 17]  H
FEEHAN, CHRISTINE  DARK DEMON [CARPATHIAN 16]  P
FEEHAN, CHRISTINE  DARK DESIRE [CARPATHIAN 2]  P
FEEHAN, CHRISTINE  DARK DESTINY [CARPATHIAN 13]  P
FEEHAN, CHRISTINE  DARK MAGIC [CARPATHIAN 4]  P
FEEHAN, CHRISTINE  DARK SECRET [CARPATHIAN 15]  H
FEEHAN, CHRISTINE  DARK SYMPHONY [CARPATHIAN 10]  P

FEEHAN, CHRISTINE/SHAYNE, MAGGIE/ HOLLY, EMMA/KNIGHT, ANGELA HOT BLOODED

P

FEEHAN, CHRISTINE  MIND GAME [GHOSTWALKERS 2]  P
FEEHAN, CHRISTINE  SHADOW GAME [GHOSTWALKERS 1]  H
FEELEY, GREGORY  OXYGEN BARONS, THE  C/P,P
FEIFFER, JULES  BARREL OF LAUGHS, A VALE OF TEARS, A  LP
FEIFFER, JULES  PASSIONELLA AND OTHER STORIES  SCX
FEINTUCH, DAVID  CHALLENGER’S HOPE [NICHOLAS SEAFORT 2]  C/P:2,P
FEINTUCH, DAVID  CHILDREN OF HOPE [NICHOLAS SEAFORT 7]  H
FEINTUCH, DAVID  FISHERMAN’S HOPE [NICHOLAS SEAFORT 4]  C/P:2,P
FEINTUCH, DAVID  KING, THE [RODRIGO OF CALEDON 2]  H
FEINTUCH, DAVID  MIDSHIPMAN’S HOPE [NICHOLAS SEAFORT 1]  C/P:2,P
FEINTUCH, DAVID  PATRIARCH’S HOPE [NICHOLAS SEAFORT 6]  C/P,P,H
FEINTUCH, DAVID  PRISONER’S HOPE [NICHOLAS SEAFORT 3]  C/P:2,P

AUTHOR  TITLE  CODE
FEINTUCH, DAVID  SEAFORT’S HOPE [NICHOLAS SEAFORT 1,2]  H
FEINTUCH, DAVID  STILL, THE [RODRIGO OF CALEDON 1]  C/P:2,LP
FEINTUCH, DAVID  VOICES OF HOPE [NICHOLAS SEAFORT 5]  C/P:2,P
FEIST, RAYMOND E.  DARKNESS AT SETHANON, A [RIFTWAR 4]  C/P,P,H,SFWA
FEIST, RAYMOND E./WURTS, JANNY  DAUGHTER OF THE EMPIRE [MARA 1]  C/P,P,H
FEIST, RAYMOND E.  EXILE’S RETURN [CONCLAVE OF SHADOWS 3]  H
FEIST, RAYMOND E.  FAERIE TALE  C/P,P,H
FEIST, RAYMOND E.  FAERIE TALE (BRITISH)  SR-H
FEIST, RAYMOND E.  FLIGHT OF THE NIGHTHAWKS [DARKWAR 1]  H

FEIST, RAYMOND E./ FORSTCHEN, WILLIAM R.

HONORED ENEMY [@LEGENDS OF THE RIFTWAR #1/RIFTWAR]

H

FEIST, RAYMOND E.  INTO A DARK REALM [DARKWAR 2]  H
FEIST, RAYMOND E.  KING OF FOXES [CONCLAVE OF SHADOWS 2]  H
FEIST, RAYMOND E.  KING OF FOXES (PROOF) [CONCLAVE OF SHADOWS 2]  SR-LP
FEIST, RAYMOND E.  KING’S BUCCANEER, THE [RIFTWAR]  H:2,P
FEIST, RAYMOND E.  KRONDOR: TEAR OF THE GODS [RIFTWAR LEGACY 3]  P
FEIST, RAYMOND E.  KRONDOR: THE ASSASSINS [RIFTWAR LEGACY 2]  P
FEIST, RAYMOND E.  KRONDOR: THE BETRAYAL [RIFTWAR LEGACY 1]  H:2,P
FEIST, RAYMOND E.  MAGICIAN [RIFTWAR 1,2]  H,SFWA:2
FEIST, RAYMOND E.  MAGICIAN: APPRENTICE [RIFTWAR 1]  C/P:3,P
FEIST, RAYMOND E.  MAGICIAN: MASTER [RIFTWAR 2]  C/P:2,P
FEIST, RAYMOND E./WURTS, JANNY  MISTRESS OF THE EMPIRE [MARA 3]  H:2,P
FEIST, RAYMOND E.  PRINCE OF THE BLOOD [RIFTWAR]  C/P,P,H,SFWA
FEIST, RAYMOND E.  RAGE OF A DEMON KING [SERPENTWAR 3]  C/P,P,H
FEIST, RAYMOND E.  RISE OF A MERCHANT PRINCE [SERPENTWAR 2]  H:2,P
FEIST, RAYMOND E.  RISE OF A MERCHANT PRINCE (PROOF) [SERPENTWAR 2]  SR-LP
FEIST, RAYMOND E./WURTS, JANNY  SERVANT OF THE EMPIRE [MARA 2]  H:2,P
FEIST, RAYMOND E.  SHADOW OF A DARK QUEEN [SERPENTWAR 1]  C/P,P,H
FEIST, RAYMOND E.  SHARDS OF A BROKEN CROWN [SERPENTWAR 4]  H:2,P
FEIST, RAYMOND E.  SILVERTHORN [RIFTWAR 3]  C/P,P,H,SFWA
FEIST, RAYMOND E.  TALON OF THE SILVER HAWK [CONCLAVE OF SHADOWS 1]  P
FEIST, RAYMOND E./WILLIAMS, TAD  WOOD BOY, THE/BURNING MAN, THE  CX
FELDMANN, SUSAN  STORYTELLING STONE, THE  PA
FELICE, CYNTHIA  DOUBLE NOCTURNE  H:2,P
FELICE, CYNTHIA  DOWNTIME  H:2
FELICE, CYNTHIA  ECLIPSES  C/P:2,P
FELICE, CYNTHIA  GODSFIRE  C/P:2,P
FELICE, CYNTHIA  ICEMAN  C/P:2,P
FELICE, CYNTHIA  KHAN’S PERSUASION, THE  C/P:2,P
FELICE, CYNTHIA  SUNBOUND, THE  C/P,P
FELICE, CYNTHIA/WILLIS, CONNIE  WATER WITCH  C/P:2,P
FELLOWS, OSCAR L.  OPERATION DAMOCLES  P
FENN, GEORGE MANVILLE  SEVEN FROZEN SAILORS  H
FENN, LIONEL  AGNES DAY [QUEST FOR THE WHITE DUCK 3]  C/P,SFWA-P,P
FENN, LIONEL  BLOOD RIVER DOWN [QUEST FOR THE WHITE DUCK 1]  P,SFWA-P
FENN, LIONEL  BY THE TIME I GET TO NASHVILLE [DIEGO 2]  P
FENN, LIONEL  KENT MONTANA AND THE ONCE AND FUTURE THING [KENT MONTANA 3] C/P,P

FENN, LIONEL KENT MONTANA AND THE REALLY UGLY THINGS FROM MARS

[KENT MONTANA 1]

C/P:2,SFWA-P:2,P

FENN, LIONEL KENT MONTANA AND THE REASONABLY INVISIBLE MAN

[KENT MONTANA 2]

C/P,P

FENN, LIONEL  MARK OF THE MODERATELY VICIOUS VAMPIRE, THE [KENT MONTANA 4]  P
FENN, LIONEL  ONCE UPON A TIME IN THE EAST [DIEGO 1]  P
FENN, LIONEL  SEVEN SPEARS OF THE W’DCH’CK, THE  C/P:2,SFWA-P,P
FENN, LIONEL  668: THE NEIGHBOR OF THE BEAST [KENT MONTANA 5]  C/P,P
FENN, LIONEL  TIME: THE SEMI-FINAL FRONTIER  P
FENN, LIONEL  WEB OF DEFEAT [QUEST FOR THE WHITE DUCK 2]  C/P,P
FENNER, PHYLLIS R.  WIDE ANGLE LENS  HA
FENTON, EDWARD  NINE QUESTIONS, THE  H
FERGUSON, BRAD  CRISIS ON CENTAURUS [STAR TREK]  C/P:2,P
FERGUSON, BRAD  FLAG FULL OF STARS, A [STAR TREK]  C/P:3,P
FERGUSON, BRAD  LAST STAND, THE [STAR TREK TNG]  C/P:3,P
FERGUSON, BRAD  WORLD NEXT DOOR, THE  C/P:2,SFWA-P,P
FERGUSON, NEIL  DOUBLE HELIX FALL  LP
FERGUSSON, BRUCE  MACE OF SOULS, THE [SIX KINGDOMS]  C/P,P,H
FERGUSSON, BRUCE  SHADOW OF HIS WINGS, THE [SIX KINGDOMS]  H,SFWA,P
FERMAN, EDWARD L./MALZBERG, BARRY N. ARENA: SPORTS SF  HA,SFWA
FERMAN, EDWARD L./VAN GELDER, GORDON BEST FROM FANTASY & SCIENCE FICTION, THE  HA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 15, THE  C/PA,PA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 16, THE  C/PA,PA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 17, THE  C/PA,PA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 18, THE  HA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 19, THE  C/PA:2,PA,SFWA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 20, THE  C/PA,PA,SFWA

AUTHOR  TITLE  CODE
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 22, THE  C/PA,PA,SFWA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 23, THE  PA,SFWA
FERMAN, EDWARD L.  BEST FROM FANTASY AND SCIENCE FICTION 24, THE  HA,PA,SFWA

FERMAN, EDWARD L.

BEST FROM FANTASY AND SCIENCE FICTION:

A SPECIAL 25TH ANNIVERSARY ANTHOLOGY, THE

HA,PA,SFWA

FERMAN, EDWARD L./JORDAN, ANNE BEST HORROR STORIES FORM THE MAGAZINE OF FANTASY AND SCIENCE

FICTION 1, THE

PA

FERMAN, EDWARD L./MALZBERG, BARRY N. FINAL STAGE  SR-PA,C/PA
FERMAN, EDWARD L./MALZBERG, BARRY N. GRAVEN IMAGES  SR-HA,HA

FERMAN, EDWARD L.

MAGAZINE OF FANTASY AND SCIENCE FICTION:

A 30-YEAR RETROSPECTIVE, THE

HA,SFWA

FERMAN, EDWARD L.  OI, ROBOT  [C/RAN-P]
FERMAN, EDWARD L./MILLS, ROBERT P.  TWENTY YEARS OF THE MAGAZINE OF FANTASY AND SCIENCE FICTION  C/PA,PA
FERMAN, JOSEPH W.  NO LIMITS  PA,C/PA
FERN, YVONNE  GENE RODDENBERRY: THE LAST CONVERSATION  R/REF-P
FERRIGNO, ROBERT  PRAYERS FOR THE ASSASSIN  H
FERRING, DAVID  KONRAD [WARHAMMER KONRAD 1]  C/P,LP
FERRING, DAVID  KONRAD SAGA, THE [@WARHAMMER KONRAD 1,2,3]  H
FERRING, DAVID  SHADOWBREED [WARHAMMER KONRAD 2]  LP
FERRING, DAVID  WARBLADE [WARHAMMER KONRAD 3]  P
FESSIER, MICHAEL  FULLY DRESSED AND IN HIS RIGHT MIND  SR-P,C/P
FEYNMAN, RICHARD P.  SURELY YOU’RE JOKING, MR. FEYNMAN  [C/RAN-H]
FEYNMAN, RICHARD P.  "WHAT DO YOU CARE WHAT OTHER PEOPLE THINK?"  [C/RAN-H]
FEZANDIE, CLEMENT  THROUGH THE EARTH  SR-LP
FFORDE, JASPER  EYRE AFFAIR, THE [THURSDAY NEXT 1]  LP
FFORDE, JASPER  LOST IN A GOOD BOOK [THURSDAY NEXT 2]  LP
FFORDE, JASPER  SOMETHING ROTTEN [THURSDAY NEXT 4]  H
FFORDE, JASPER  WELL OF LOST PLOTS, THE [THURSDAY NEXT 3]  H
FICHMAN, FREDERICK  SETI  C/P:2,P
FIEDLER, LESLIE A.  IN DREAMS AWAKE  PA
FIEDLER, LESLIE A.  OLAF STAPLEDON, A MAN DIVIDED  R/REF-H
FILBRUN, J. S.  GEMINI RISING  P
FINCH, SHEILA  GARDEN OF THE SHAPED, THE [SHAPER EXILE 1]  C/P:2,P
FINCH, SHEILA  INFINITY’S WEB  C/P,P
FINCH, SHEILA  SHAPER’S LEGACY [SHAPER EXILE 2]  C/P:2,P
FINCH, SHEILA  SHAPING THE DAWN [SHAPER EXILE 3]  C/P:2,SFWA-P,P
FINCH, SHEILA  TIGER IN THE SKY [@OUT OF TIME #2]  P
FINCH, SHEILA  TRIAD  C/P:2,P
FINDER, JAN HOWARD  ALIEN ENCOUNTERS  SR-HA
FINDLEY, NIGEL D.  BROKEN SPHERE, THE [@SPELLJAMMER CLOAKMASTER #5]  P
FINDLEY, NIGEL D.  HOUSE OF THE SUN [@SHADOWRUN 17]  P
FINDLEY, NIGEL D.  INTO THE VOID [@SPELLJAMMER CLOAKMASTER #2]  C/P,P
FINDLEY, NIGEL D.  LONE WOLF [@SHADOWRUN 12]  P
FINDLEY, NIGEL D.  LOST KAER [@EARTHDAWN]  P
FINDLEY, NIGEL D.  NO LIMITS [@VIRTUAL WORLD]  P
FINDLEY, NIGEL D.  SHADOWPLAY [@SHADOWRUN 9]  C/P,P
FINDLEY, NIGEL D.  2XS [SHADOWRUN 4]  C/P:2,P
FINDLEY, TIMOTHY  NOT WANTED ON THE VOYAGE  H
FINE, STEPHEN  MOLLY DEAR  H
FINGEROTH, DANNY/FEIN ERIC  RAMPAGE [@DOOM’S DAY 1]  P
FINLAY, CHARLES COLEMAN  PRODIGAL TROLL, THE  LP
FINLAY, ROGER T.  ADVENTURES ON STRANGE ISLANDS  H
FINN, RALPH L.  CAPTIVE ON THE FLYING SAUCERS  SR-P
FINN, RALPH L.  FREAKS AGAINST SUPERMEN  SR-P
FINNEY, CHARLES G.  CIRCUS OF DR. LAO, THE  C/P:2,SR-LP,LP
FINNEY, CHARLES G.  GHOSTS OF MANACLE, THE  C/P:2,P
FINNEY, CHARLES G.  UNHOLY CITY, THE  P
FINNEY, JACK  ABOUT TIME  LP
FINNEY, JACK  BODY SNATCHERS, THE  C/P:2,P
FINNEY, JACK  FROM TIME TO TIME [TIME AND AGAIN 2]  H,LP
FINNEY, JACK  I LOVE GALESBURG IN THE SPRINGTIME  SR-P
FINNEY, JACK  INVASION OF THE BODY SNATCHERS  C/P:2,LP
FINNEY, JACK  MARION’S WALL  C/P,P
FINNEY, JACK  THIRD LEVEL, THE  C/P,P,H
FINNEY, JACK  3 BY FINNEY  H,LP
FINNEY, JACK  TIME AND AGAIN [TIME AND AGAIN 1]  H:2,LP
FINTUSHEL, ELIOT  BREAKFAST WITH THE ONES YOU LOVE  LP
FIRESIGN THEATRE  FIRESIGN THEATRE’S BIG BOOK OF PLAYS, THE  VLP
FIRESIGN THEATRE  FIRESIGN THEATRE’S BIG MYSTERY JOKE BOOK, THE  VLP
FISCH, SHOLLY  GHOSTBUSTERS: THE RETURN [GHOSTBUSTERS]  P
FISCHER, LEONARD  LET OUT THE BEAST  SR-P
FISCHER, MARJORIE/HUMPHRIES, ROLFE  PAUSE TO WONDER  HA
FISH, ROBERT L.  MEMOIRS OF SCHLOCK HOMES, THE [BAGEL STREET DOZEN 2]  H
FISHER, CATHERINE  LAMMAS FIELD, THE  H,LP

The MITSFS Pinkdex  September 24, 2007  Page 143
AUTHOR  TITLE  CODE
FISHER, CATHERINE  ORACLE BETRAYED, THE [ORACLE 1]  P
FISHER, DAVID  DOCTOR WHO AND THE CREATURE FROM THE PIT [DOCTOR WHO]  SR-P,C/P
FISHER, DAVID  DOCTOR WHO AND THE LEISURE HIVE [DOCTOR WHO]  SR-P
FISHER, JAMES P.  GREAT BRAIN ROBBERY, THE  C/P:2,P
FISHER, JUDE  ROSE OF THE WORLD, THE [FOOL’S GOLD 3]  H
FISHER, JUDE  SORCERY RISING [FOOL’S GOLD 1]  C/P,P
FISHER, JUDE  WILD MAGIC [FOOL’S GOLD 2]  H
FISHER, LEONARD EVERETT  NOONAN  H
FISHER, LOU  BLUE ICE PILOT, THE  C/P,P
FISHER, LOU  SUNSTOP 8  P
FISHER, PAUL R.  ASH STAFF, THE [ASH STAFF 1]  C/P,P
FISHER, PAUL R.  HAWKS OF FELLHEATH, THE [ASH STAFF 2]  C/P,P
FISHER, PAUL R.  MONT CANT GOLD  C/P,P
FISHER, PAUL R.  PRINCESS AND THE THORN, THE [ASH STAFF 3]  P
FISHER, R. L.  PRINCE OF WHALES, THE  P,SFWA-P
FISHER, VARDIS  GOLDEN ROOMS, THE  H
FISK, NICHOLAS  RAG, A BONE, AND A HANK OF HAIR, A  H
FISK, NICHOLAS  SPACE HOSTAGES  SR-P
FITCH, ED  CASTLE OF DECEPTION  LP
FITCH, MARINA  BORDER, THE  P
FITCH, MARINA  SEVENTH HEART, THE  P
FITZGERALD, GREGORY/DILLON, JOHN  LATE GREAT FUTURE, THE  PA
FITZGERALD, GREGORY  NEUTRON STARS  PA
FITZGIBBON, CONSTANTINE  WHEN THE KISSING HAD TO STOP  C/P,P,H
FLACKES, B.  DUEL IN NIGHTMARE WORLDS  SR-P
FLAMMARION, CAMILLE  OMEGA  SR-H
FLANAGAN, GRAEME  ROBERT BLOCH: A BIO-BIBLIOGRAPHY  R/IND
FLANAGAN, TERRY/EHRHARDT, ELEANOR  TREK OR TREAT  [C/RAN-P]
FLEET, ROBERT C.  LAST MOUNTAIN  P,SFWA-P:2
FLEET, ROBERT C.  WHITE HORSE, DARK DRAGON  C/P,SFWA-P:2,P
FLEISCHER, LEONORE  HEAVEN CAN WAIT  C/P,P
FLEISCHMAN, THEO  DOUBLE EXPOSURE  H,P
FLEISHER, MICHAEL L.  GREAT SUPERMAN BOOK, THE  R/REF-H
FLEMING, GHERBOD  ASSAMITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P

FLEMING, GHERBOD DARK PROPHECY

[@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #3]

P

FLEMING, GHERBOD DEVIL’S ADVOCATE, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #1]

P

FLEMING, GHERBOD  EXECUTIONER [@WORLD OF DARKNESS PREDATOR & PREY #6]  P
FLEMING, GHERBOD  GANGREL [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
FLEMING, GHERBOD  JUDGE [@WORLD OF DARKNESS PREDATOR & PREY #2]  P
FLEMING, GHERBOD  JURY [@WORLD OF DARKNESS PREDATOR & PREY #4]  P
FLEMING, GHERBOD  NOSFERATU [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
FLEMING, GHERBOD  WEREWOLF [@WORLD OF DARKNESS PREDATOR & PREY #3]  P

FLEMING, GHERBOD WINNOWING, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE BLOOD CURSE #2]

P

FLEMING, IAN  CASINO ROYALE [JAMES BOND 1]  C/P,P
FLEMING, IAN  DIAMONDS ARE FOREVER [JAMES BOND 4]  P
FLEMING, IAN  DOCTOR NO [JAMES BOND 6]  C/P:2,P
FLEMING, IAN  FOR YOUR EYES ONLY [JAMES BOND 8]  C/P:2,P
FLEMING, IAN  FROM RUSSIA, WITH LOVE [JAMES BOND 5]  C/P,P
FLEMING, IAN  GOLDFINGER [JAMES BOND 7]  C/P,P
FLEMING, IAN  LIVE AND LET DIE [JAMES BOND 2]  C/P,P
FLEMING, IAN  MAN WITH THE GOLDEN GUN, THE [JAMES BOND 13]  C/P,P
FLEMING, IAN  MOONRAKER [JAMES BOND 3]  C/P:2,P
FLEMING, IAN  ON HER MAJESTY’S SECRET SERVICE [JAMES BOND 11]  C/P,P
FLEMING, IAN  SPY WHO LOVED ME, THE [JAMES BOND 10]  C/P:2,P
FLEMING, IAN  THUNDERBALL [JAMES BOND 9]  C/P:2,P
FLEMING, IAN  YOU ONLY LIVE TWICE [JAMES BOND 12]  C/P:2,P
FLEMING, JOAN  CHILL AND THE KILL, THE  P
FLEMING, PETER  FLYING VISIT, THE  H
FLES, BARTHOLD  SATURDAY EVENING POST FANTASY STORIES, THE  SR-PA
FLETCHER, GEORGE U.  WELL OF THE UNICORN, THE  H
FLETCHER, JANE  EMPRESS AND THE ACOLYTE, THE [LYREMOUTH 3]  LP
FLETCHER, JANE  EXILE AND THE SORCERER, THE [LYREMOUTH 1]  LP
FLETCHER, JANE  TEMPLE AT LANDFALL, THE [CELAENO]  LP
FLETCHER, JANE  TRAITOR AND THE CHALICE, THE [LYREMOUTH 2]  LP
FLETCHER, JANE  WALLS OF WESTERNFORT, THE [CELAENO]  LP
FLETCHER, SUSAN  DRAGON’S MILK  LP
FLEWELLING, LYNN  BONE DOLL’S TWIN, THE [TOBIN OF SKALA 1]  P
FLEWELLING, LYNN  HIDDEN WARRIOR [TOBIN OF SKALA 2]  C/P,P
FLEWELLING, LYNN  LUCK IN THE SHADOWS [NIGHTRUNNER 1]  C/P,P
FLEWELLING, LYNN  ORACLE’S QUEEN, THE [TOBIN OF SKALA 3]  C/P,P
FLEWELLING, LYNN  STALKING DARKNESS [NIGHTRUNNER 2]  C/P,P

AUTHOR  TITLE  CODE
FLEWELLING, LYNN  TRAITOR’S MOON [NIGHTRUNNER 3]  C/P,P
FLICKER, TED  GOOD AMERICAN, THE  LP
FLIEGER, VERLYN  SPLINTERED LIGHT  R/REF-P
FLINDERS, KARL  LOVE MACHINERY, THE  SR-L
FLINN, DENNY MARTIN  FEARFUL SUMMONS, THE [STAR TREK]  C/P:2,P
FLINT, ERIC  BEST OF JIM BAEN’S UNIVERSE, THE (2006)  HA
FLINT, ERIC/SPOOR, RYK E.  BOUNDARY  H
FLINT, ERIC/WENTWORTH, K. D.  COURSE OF EMPIRE, THE  H
FLINT, ERIC/WENTWORTH, K. D.  COURSE OF EMPIRE, THE (PROOF)  SR-LP
FLINT, ERIC/DRAKE, DAVID  DANCE OF TIME, THE [BELISARIUS 6]  H
FLINT, ERIC/DRAKE, DAVID  DESTINY’S SHIELD [BELISARIUS 3]  C/P:2,P
FLINT, ERIC/DRAKE, DAVID  FORTUNE’S STROKE [BELISARIUS 4]  H:2
FLINT, ERIC/ROACH, RICHARD  FORWARD THE MAGE [JOE’S WORLD 2]  C/P,P,H
FLINT, ERIC  GRANTVILLE GAZETTE, VOL. 1 [SIXTEEN THIRTY-TWO]  PA
FLINT, ERIC  GRANTVILLE GAZETTE, VOL. 2 [SIXTEEN THIRTY-TWO]  HA
FLINT, ERIC  GRANTVILLE GAZETTE, VOL. 3 [SIXTEEN THIRTY-TWO]  HA
FLINT, ERIC/DRAKE, DAVID  IN THE HEART OF DARKNESS [BELISARIUS 2]  C/P,P
FLINT, ERIC  MOTHER OF DEMONS  C/P:2,P
FLINT, ERIC/DRAKE, DAVID  OBLIQUE APPROACH, AN [BELISARIUS 1]  C/P,P
FLINT, ERIC  PHILOSOPHICAL STRANGLER, THE [JOE’S WORLD 1]  H,P
FLINT, ERIC/FREER, DAVE  RATS, THE BATS & THE UGLY, THE [RATS & BATS 2]  H
FLINT, ERIC  RING OF FIRE  PA
FLINT, ERIC  RING OF FIRE (PROOF) [SIXTEEN THIRTY-TWO]  SR-LPA
FLINT, ERIC  RIVERS OF WAR, THE [RIVERS OF WAR 1]  H
FLINT, ERIC/DENNIS, ANDREW  1635: THE CANNON LAW [SIXTEEN THIRTY-TWO]  H
FLINT, ERIC/DENNIS, ANDREW  1634: THE GALILEO AFFAIR [SIXTEEN THIRTY-TWO 3]  H:2
FLINT, ERIC  1632 [SIXTEEN THIRTY-TWO 1]  H
FLINT, ERIC/DEMARCE, VIRGINIA  1634: THE RAM REBELLION [SIXTEEN THIRTY-TWO]  H
FLINT, ERIC/DRAKE, DAVID  TIDE OF VICTORY, THE [BELISARIUS 5]  H
FLINT, ERIC/DRAKE, DAVID  TYRANT, THE [GENERAL]  P
FLINT, HOMER EON  DEVOLUTIONIST, THE/EMANCIPATRIX, THE  C/P:2,P
FLINT, HOMER EON  LORD OF DEATH, THE/QUEEN OF LIFE, THE  C/P,P
FLINT, KENNETH C.  CHALLENGE OF THE CLANS [FINN MACCUMHAL 1]  C/P,P
FLINT, KENNETH C.  CHAMPIONS OF THE SIDHE [SIDHE 2]  C/P:2,P
FLINT, KENNETH C.  CROMM  H,P
FLINT, KENNETH C.  DARK DRUID, THE [FINN MACCUMHAL 3]  P
FLINT, KENNETH C.  DARKENING FLOOD, THE  P
FLINT, KENNETH C.  ISLE OF DESTINY  C/P,P
FLINT, KENNETH C.  LEGENDS REBORN  C/P,P
FLINT, KENNETH C.  MASTER OF THE SIDHE [SIDHE 3]  C/P:2,P
FLINT, KENNETH C.  OTHERWORLD  C/P,P
FLINT, KENNETH C.  RIDERS OF THE SIDHE, THE [SIDHE 1]  C/P:2,P
FLINT, KENNETH C.  STORM UPON ULSTER, A  C/P:2,P
FLORESCU, RADU  IN SEARCH OF FRANKENSTEIN  R/REF-P
FLUGGE, FRANK  UTOPIA  FP
FLYNN, CASEY  ENCHANTED ISLES, THE [@GODS OF IRELAND #2]  P
FLYNN, CASEY  MOST ANCIENT SONG [GODS OF IRELAND 1]  C/P,P
FLYNN, JENNIFER  20TH CENTURY COMPUTERS AND HOW THEY WORKED [STAR TREK TNG]  R/REF-H
FLYNN, JOHN L.  FALL OF THE REPUBLIC (STORY TREATMENT) [STAR WARS]  SR-S
FLYNN, MICHAEL  EIFELHEIM  H
FLYNN, MICHAEL  FALLING STARS [STARS 4]  H
FLYNN, MICHAEL  FIRESTAR [STARS 1]  H:2,P
FLYNN, MICHAEL  FOREST OF TIME AND OTHER STORIES, THE  LP
FLYNN, MICHAEL  IN THE COUNTRY OF THE BLIND  C/P:2,SFWA-P,P
FLYNN, MICHAEL  IN THE COUNTRY OF THE BLIND (REVISED)  H,P
FLYNN, MICHAEL  LODESTAR [STARS 3]  H,P
FLYNN, MICHAEL  NANOTECH CHRONICLES, THE  C/P,SFWA-P,P
FLYNN, MICHAEL  ROGUE STAR [STARS 2]  P
FLYNN, MICHAEL  WRECK OF THE RIVER OF STARS, THE  H
FLYNN, TOM  GALACTIC RAPTURE  LP
FODOR, NANDOR  BETWEEN TWO WORLDS  [C/RAN-P]
FODOR, NANDOR  HAUNTED MIND, THE  [C/RAN-P]
FODOR, R. V./TAYLOR, G. J.  IMPACT!  P
FOGLIO, PHIL/FOGLIO, KAJA  AGATHA HETERODYNE AND THE AIRSHIP CITY [@GIRL GENIUS #2]  CX
FOGLIO, PHIL/FOGLIO, KAJA  AGATHA HETERODYNE AND THE BEETLEBURG CLANK [@GIRL GENIUS #1]  CX
FOGLIO, PHIL/FOGLIO, KAJA  AGATHA HETERODYNE AND THE CIRCUS OF DREAMS [@GIRL GENIUS #4]  CX

FOGLIO, PHIL/FOGLIO, KAJA AGATHA HETERODYNE AND THE CLOCKWORK PRINCESS

[@GIRL GENIUS #5]

CX

FOGLIO, PHIL/FOGLIO, KAJA  AGATHA HETERODYNE AND THE MONSTER ENGINE [@GIRL GENIUS #3]  CX
FOLEY, JOHN  MAN IN THE MOON  SR-P
FONSTAD, KAREN WYNN  ATLAS OF MIDDLE-EARTH, THE [LORD OF THE RINGS]  R/REF-H
FONSTAD, KAREN WYNN  ATLAS OF PERN, THE [PERN]  R/REF-H,C/REF-H:3
FONSTAD, KAREN WYNN  FORGOTTEN REALMS ATLAS, THE [FORGOTTEN REALMS]  A
FONTANA, D. C.  QUESTOR TAPES, THE  C/P:2,P

AUTHOR  TITLE  CODE
FONTANA, D. C.  VULCAN’S GLORY [STAR TREK]  C/P:2,P
FONTENAY, CHARLES L.  DAY THE OCEANS OVERFLOWED, THE  C/P,P
FONTENAY, CHARLES L.  REBELS OF THE RED PLANET  P,DF113
FONTENAY, CHARLES L.  TWICE UPON A TIME  DD266
FOOTE, KEITH D.  INTERSTELLAR INTRIGUE (PROOF) [@OUTREACH INVESTIGATIONS #1]  SR-P
FORBECK, MATT  BLOOD BOWL [@BLOOD BOWL 1]  P
FORBECK, MATT  DEAD BALL [@BLOOD BOWL 2]  P
FORBECK, MATT  DEATH MATCH [@BLOOD BOWL 3]  P
FORBECK, MATT  DRAGONS REVEALED [KNIGHTS OF THE DRAGONS REVELATIONS 2]  LP
FORBECK, MATT  MARKED FOR DEATH [@EBERRON LOST MARK #1]  P

FORBECK, MATT PROPHECY OF THE DRAGONS

[KNIGHTS OF THE DRAGONS REVELATIONS 1]

LP

FORBECK, MATT  QUEEN OF DEATH, THE [@EBERRON LOST MARK #3]  P
FORBECK, MATT  ROAD TO DEATH, THE [@EBERRON LOST MARK #2]  P

FORBECK, MATT SECRET OF THE SPIRITKEEPER, THE

[@KNIGHTS OF THE SILVER DRAGON #1]

LP

FORBES, CAROLINE  NEEDLE ON FULL, THE  LP
FORBES, DAVID  AMBER WIZARD, THE [OSSERIAN 1]  P
FORBES, EDITH  EXIT TO REALITY  H
FORD, GARRET  SCIENCE AND SORCERY  PA
FORD, JEFFREY  BEYOND, THE [WELL-BUILT CITY 3]  H
FORD, JEFFREY  EMPIRE OF ICE CREAM, THE  H
FORD, JEFFREY  FANTASY WRITER’S ASSISTANT, THE  H
FORD, JEFFREY  MEMORANDA [WELL-BUILT CITY 2]  P
FORD, JEFFREY  PHYSIOGNOMY, THE [WELL-BUILT CITY 1]  C/P:2,LP
FORD, JOHN M.  CASTING FORTUNE [LIAVEK 6]  H,SFWA-P
FORD, JOHN M.  DRAGON WAITING, THE  C/P:2,P,H
FORD, JOHN M.  FINAL REFLECTION, THE [STAR TREK]  C/P:3,P
FORD, JOHN M.  FROM THE END OF THE TWENTIETH CENTURY  SR-H,H
FORD, JOHN M.  FUGUE STATE  SFWA-TD25,TD25
FORD, JOHN M.  GROWING UP WEIGHTLESS  C/P:2,LP
FORD, JOHN M.  HOW MUCH FOR JUST THE PLANET? [STAR TREK]  C/P,P,H
FORD, JOHN M.  LAST HOT TIME, THE  H
FORD, JOHN M.  PRINCES OF THE AIR, THE  C/P:2,SFWA-P,P
FORD, JOHN M.  SCHOLARS OF NIGHT, THE  P,SFWA-P
FORD, JOHN M.  WEB OF ANGELS  C/P,P
FORD, PATRICK K.  MABINOGI, THE  LP
FORD, RICHARD  CHILDREN OF ASHGAROTH, THE [FARADAWN 3]  P
FORD, RICHARD  QUEST FOR THE FARADAWN [FARADAWN 1]  H,P
FOREIGN LANGUAGES PUB.  VISITOR FROM OUTER SPACE, A  SR-PA
FOREST, JEAN-CLAUDE  BARBARELLA  CX
FOREST, SALAMBO  WITCH POWER  SR-L
FORMAN, JAMES D.  DOOMSDAY PLUS TWELVE  H
FORREST, ELIZABETH  DARK TIDE  P
FORREST, ELIZABETH  PHOENIX FIRE  P,SFWA-P
FORREST, KATHERINE V.  DAUGHTERS OF A CORAL DAWN [MATERNAS 1]  LP
FORREST, KATHERINE V.  DAUGHTERS OF AN AMBER MOON [MATERNAS 2]  LP
FORREST, KATHERINE V.  DAUGHTERS OF AN EMERALD DUSK [MATERNAS 3]  LP
FORRESTER, JOHN  BESTIARY MOUNTAIN [BESTIARY 1]  H,P
FORRESTER, JOHN  SECRET OF THE ROUND BEAST, THE [BESTIARY 2]  P
FORSTCHEN, WILLIAM R.  ACTION STATIONS [@WING COMMANDER 6]  C/P,P
FORSTCHEN, WILLIAM R.  ALEXANDRIAN RING, THE [GAMESTER WARS 1]  C/P:2,P
FORSTCHEN, WILLIAM R.  ARENA [MAGIC: THE GATHERING]  C/P:2,P
FORSTCHEN, WILLIAM R.  ARTICLE 23 [JUSTIN BELL 2]  C/P:2,P
FORSTCHEN, WILLIAM R.  ASSASSIN GAMBIT, THE [GAMESTER WARS 2]  C/P:2,P
FORSTCHEN, WILLIAM R.  BAND OF BROTHERS, A [LOST REGIMENT #7]  P
FORSTCHEN, WILLIAM R.  BATTLE HYMN [@LOST REGIMENT 5]  P
FORSTCHEN, WILLIAM R./DEMAREST, JAKI  CATSEYE [SHATTERED LIGHT 3]  P
FORSTCHEN, WILLIAM R./MORRISON, GREG CRYSTAL SORCERORS, THE [CRYSTAL WARRIORS 2]  P
FORSTCHEN, WILLIAM R./MORRISON, GREG CRYSTAL WARRIORS, THE [CRYSTAL WARRIORS 1]  C/P:2,P
FORSTCHEN, WILLIAM R.  DARKNESS UPON THE ICE, A [ICE 3]  C/P,P
FORSTCHEN, WILLIAM R.  DOWN TO THE SEA [@LOST REGIMENT 9]  C/P,P
FORSTCHEN, WILLIAM R./KEITH, ANDREW  FALSE COLORS [@WING COMMANDER 7]  C/P,P
FORSTCHEN, WILLIAM R.  FATEFUL LIGHTNING [@LOST REGIMENT 4]  P
FORSTCHEN, WILLIAM R.  FLAME UPON THE ICE, THE [ICE 2]  C/P:2,P
FORSTCHEN, WILLIAM R.  FLEET ACTION [@WING COMMANDER 3]  P,SFWA-P
FORSTCHEN, WILLIAM R.  FORGOTTEN WAR, THE [@STAR TREK TNG]  C/P,P
FORSTCHEN, WILLIAM R./SEGRIFF, LARRY  FOUR MAGICS, THE  P
FORSTCHEN, WILLIAM R./KEITH, ANDREW  HEART OF THE TIGER [@WING COMMANDER 4]  P
FORSTCHEN, WILLIAM R.  ICE PROPHET [ICE 1]  C/P:2,P
FORSTCHEN, WILLIAM R.  INTO THE SEA OF STARS  C/P,P
FORSTCHEN, WILLIAM R.  MEN OF WAR [@LOST REGIMENT #8]  C/P,P
FORSTCHEN, WILLIAM R.  NAPOLEON WAGER, THE [GAMESTER WARS 3]  P
FORSTCHEN, WILLIAM R.  NEVER SOUND RETREAT [@LOST REGIMENT #6]  P

The MITSFS Pinkdex  September 24, 2007  Page 146
AUTHOR  TITLE  CODE
FORSTCHEN, WILLIAM R./OHLANDER, BEN  PRICE OF FREEDOM, THE [@WING COMMANDER 5]  P
FORSTCHEN, WILLIAM R.  PROMETHEUS [JUSTIN BELL 3]  C/P,P
FORSTCHEN, WILLIAM R.  RALLY CRY [@LOST REGIMENT 1]  H,P
FORSTCHEN, WILLIAM R.  STAR VOYAGER ACADEMY [JUSTIN BELL 1]  C/P,P
FORSTCHEN, WILLIAM R.  TERRIBLE SWIFT SWORD [@LOST REGIMENT 3]  C/P,P
FORSTCHEN, WILLIAM R.  UNION FOREVER [@LOST REGIMENT 2]  C/P:2,P
FORSTER, E. M.  CELESTIAL OMNIBUS AND OTHER STORIES, THE  C/P,P
FORSTER, E. M.  COLLECTED SHORT STORIES  SR-P
FORSYTH, KATE  CURSED TOWERS, THE [WITCHES OF EILEANAN 3]  C/P:2,P
FORSYTH, KATE  FATHOMLESS CAVES, THE [WITCHES OF EILEANAN 6]  C/P,P
FORSYTH, KATE  FORBIDDEN LAND, THE [WITCHES OF EILEANAN 4]  C/P,P
FORSYTH, KATE  HEART OF STARS, THE [@RHIANNON’S RIDE #3]  P
FORSYTH, KATE  POOL OF TWO MOONS, THE [WITCHES OF EILEANAN 2]  C/P:2,P
FORSYTH, KATE  SHINING CITY, THE [@RHIANNON’S RIDE #2]  P
FORSYTH, KATE  SKULL OF THE WORLD, THE [WITCHES OF EILEANAN 5]  C/P,P
FORSYTH, KATE  TOWER OF RAVENS, THE [@RHIANNON’S RIDE #1]  P
FORSYTH, KATE  WITCHES OF EILEANAN, THE [WITCHES OF EILEANAN 1]  C/P:2,P
FORT, CHARLES  BOOK OF THE DAMNED, THE  [C/RAN-P:2]
FORT, CHARLES  LO!  [C/RAN-H,C/RAN-P:2]
FORT, CHARLES  WILD TALENTS  [C/RAN-P:2]
FORTIER, RON  TERMINATOR: THE BURNING EARTH [TERMINATOR]  CX
FORTUNE, DION  SECRETS OF DOCTOR TAVERNER, THE  H
FORTUNE, JULIE  SACRIFICE MOON [@STARGATE SG-1 #2]  P
FORWARD, EVE  ANIMIST  H
FORWARD, EVE  VILLAINS BY NECESSITY  H,P
FORWARD, ROBERT L.  CAMELOT 30K  C/P:2,P
FORWARD, ROBERT L.  DRAGON’S EGG  C/P:2,SFWA,P,H
FORWARD, ROBERT L.  DRAGON’S EGG (PROOF)  SR-LP
FORWARD, ROBERT L.  FLIGHT OF THE DRAGONFLY, THE  C/P:2,SFWA,LP
FORWARD, ROBERT L.  INDISTINGUISHABLE FROM MAGIC  C/P:2,P

FORWARD, ROBERT L./ FORWARD, MARGARET DODSON MAROONED ON EDEN [ROCHEWORLD 3]

C/P,SFWA-P,P

FORWARD, ROBERT L.  MARTIAN RAINBOW  C/P,P,H

FORWARD, ROBERT L./ FORWARD, MARTHA DODSON OCEAN UNDER THE ICE [ROCHEWORLD 4]

C/P,P

FORWARD, ROBERT L./ FULLER, JULIE FORWARD RESCUED FROM PARADISE [ROCHEWORLD 5]

C/P,P

FORWARD, ROBERT L./ FULLER, JULIE FORWARD RETURN TO ROCHEWORLD [ROCHEWORLD 2]

C/P:2,SFWA-P,P

FORWARD, ROBERT L.  ROCHEWORLD [ROCHEWORLD 1]  C/P:2,SFWA-P,P
FORWARD, ROBERT L.  SATURN RUKH  C/P,P,H
FORWARD, ROBERT L.  STARQUAKE  C/P,P,H,SFWA
FORWARD, ROBERT L.  TIMEMASTER  C/P,P,H
FORWARD, SIMON A.  DRIFT [DOCTOR WHO PAST DOCTORS]  P
FORWARD, SIMON A.  EMOTIONAL CHEMISTRY [DOCTOR WHO EIGHTH DOCTOR]  P
FORWARD, SIMON A.  SHELL SHOCK [DOCTOR WHO]  SR-H
FOSS, CHRIS  SCIENCE FICTION ART  A
FOSTER, ALAN DEAN  ALIEN [ALIENS]  H:2,P
FOSTER, ALAN DEAN  ALIEN 3 [ALIENS]  C/P,P
FOSTER, ALAN DEAN  ALIEN NATION [ALIEN NATION]  C/P:2,P
FOSTER, ALAN DEAN  ALIENS [ALIENS]  P
FOSTER, ALAN DEAN  APPROACHING STORM, THE [@STAR WARS]  H,P
FOSTER, ALAN DEAN  BETCHA CAN’T READ JUST ONE  C/PA,PA
FOSTER, ALAN DEAN  BLACK HOLE, THE  C/P,P,H
FOSTER, ALAN DEAN  BLOODHYPE [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  CACHALOT [HUMANX]  H:2,P
FOSTER, ALAN DEAN  CALL TO ARMS, A [@DAMNED #1]  C/P,P,H
FOSTER, ALAN DEAN  CANDLE OF DISTANT EARTH, THE  H
FOSTER, ALAN DEAN  CARNIVORES OF LIGHT AND DARKNESS [JOURNEYS OF THE CATECHIST 1] H
FOSTER, ALAN DEAN  CAT-A-LYST  C/P,SFWA-P,P,H
FOSTER, ALAN DEAN  CHORUS SKATING [SPELLSINGER 8]  C/P,P
FOSTER, ALAN DEAN  CHRONICLES OF RIDDICK, THE  P
FOSTER, ALAN DEAN  CLASH OF THE TITANS  C/P,P
FOSTER, ALAN DEAN  CODGERSPACE  C/P,SFWA-P:2,P,H,SFWA
FOSTER, ALAN DEAN  CYBER WAY  C/P:2,SFWA-P:2,P
FOSTER, ALAN DEAN  DARK STAR  C/P:2,P
FOSTER, ALAN DEAN  DAY OF THE DISSONANCE, THE [SPELLSINGER 3]  C/P:2,P
FOSTER, ALAN DEAN  DELUGE DRIVERS, THE [HUMANX ICERIGGER 3]  C/P:2,P
FOSTER, ALAN DEAN/RUSSELL, ERIC FRANK DESIGN FOR GREAT-DAY  C/P:2,P
FOSTER, ALAN DEAN  DIG, THE  C/P,P,H
FOSTER, ALAN DEAN  DINOTOPIA LOST [DINOTOPIA]  H
FOSTER, ALAN DEAN  DIRGE [HUMANX FOUNDING COMMONWEALTH 2]  H
FOSTER, ALAN DEAN  DIUTURNITY’S DAWN [HUMANX FOUNDING COMMONWEALTH 3]  H
FOSTER, ALAN DEAN  DROWNING WORLD [HUMANX]  H

The MITSFS Pinkdex  September 24, 2007  Page 147
AUTHOR  TITLE  CODE
FOSTER, ALAN DEAN  END OF THE MATTER, THE [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  FALSE MIRROR, THE [@DAMNED #2]  C/P,P,H
FOSTER, ALAN DEAN  FLINX IN FLUX [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  FLINX’S FOLLY [HUMANX]  P
FOSTER, ALAN DEAN  FLINX’S FOLLY (PROOF) [HUMANX FLINX]  SR-LP
FOSTER, ALAN DEAN  FOR LOVE OF MOTHER-NOT [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  GLORY LANE  C/P:2,SFWA-P,P
FOSTER, ALAN DEAN  GREENTHIEVES  H:2,P
FOSTER, ALAN DEAN  HAND OF DINOTOPIA, THE [DINOTOPIA]  H
FOSTER, ALAN DEAN  HORROR ON THE BEACH, THE  SR-P
FOSTER, ALAN DEAN  HOUR OF THE GATE, THE [SPELLSINGER 2]  C/P:2,P
FOSTER, ALAN DEAN  HOWLING STONES, THE [HUMANX]  H:2,P
FOSTER, ALAN DEAN  I INSIDE, THE  C/P:2,P
FOSTER, ALAN DEAN  ICERIGGER [HUMANX ICERIGGER 1]  C/P,P,H
FOSTER, ALAN DEAN  IMPOSSIBLE PLACES  P
FOSTER, ALAN DEAN  INTERLOPERS  C/P,P
FOSTER, ALAN DEAN  INTO THE OUT OF  H:2
FOSTER, ALAN DEAN  INTO THE THINKING KINGDOMS [JOURNEYS OF THE CATECHIST 2]  P
FOSTER, ALAN DEAN  JED THE DEAD  C/P,P
FOSTER, ALAN DEAN  KINGDOMS OF LIGHT  H
FOSTER, ALAN DEAN  KRULL  C/P,P
FOSTER, ALAN DEAN  LAST STARFIGHTER, THE  C/P:2,P
FOSTER, ALAN DEAN  LIFE FORM  C/P:2,P
FOSTER, ALAN DEAN  LIGHT-YEARS BENEATH MY FEET, THE  H
FOSTER, ALAN DEAN  LOST AND FOUND [TAKEN 1]  H
FOSTER, ALAN DEAN  LUANA  P
FOSTER, ALAN DEAN  MAD AMOS  C/P,P
FOSTER, ALAN DEAN  MAN WHO USED THE UNIVERSE, THE  H:2,P
FOSTER, ALAN DEAN  MAORI  C/P,SFWA-P,P
FOSTER, ALAN DEAN  METROGNOME, THE  C/P:2,P
FOSTER, ALAN DEAN  MID-FLINX [HUMANX FLINX]  C/P,P,H
FOSTER, ALAN DEAN  MIDWORLD [HUMANX]  H:2,P
FOSTER, ALAN DEAN  MISSION TO MOULOKIN [HUMANX ICERIGGER 2]  H:2,P
FOSTER, ALAN DEAN  MISSION TO MOULOKIN (PROOF) [HUMANX ICERIGGER 2]  SR-LP
FOSTER, ALAN DEAN  MOCKING PROGRAM, THE  C/P,P,H
FOSTER, ALAN DEAN  MOMENT OF THE MAGICIAN, THE [SPELLSINGER 4]  C/P,P,H
FOSTER, ALAN DEAN  MONTEZUMA STRIP  C/P,P
FOSTER, ALAN DEAN  MOULOKIN MISSION, DIE [HUMANX ICERIGGER 2]  FP
FOSTER, ALAN DEAN  NOR CRYSTAL TEARS [HUMANX]  C/P:2,P
FOSTER, ALAN DEAN  ORPHAN STAR [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  OUTLAND  C/P:2,P
FOSTER, ALAN DEAN  PARALLELITIES  C/P,P,H
FOSTER, ALAN DEAN  PATHS OF THE PERAMBULATOR, THE [SPELLSINGER 5]  C/P,P,H
FOSTER, ALAN DEAN  PHYLOGENESIS [HUMANX FOUNDING COMMONWEALTH 1]  H
FOSTER, ALAN DEAN  QUOZL  C/P:2,SFWA-P,P
FOSTER, ALAN DEAN  REUNION [HUMANX FLINX]  H
FOSTER, ALAN DEAN  RUNNING FROM THE DEITY [HUMANX FLINX]  H
FOSTER, ALAN DEAN  SAGRAMANDA  H
FOSTER, ALAN DEAN  SEASON OF THE SPELLSONG [SPELLSINGER 1,2,3]  H:2
FOSTER, ALAN DEAN  SENTENCED TO PRISM [HUMANX]  C/P:2,P
FOSTER, ALAN DEAN  SHADOWKEEP  C/P:2,P
FOSTER, ALAN DEAN  SLIDING SCALES [HUMANX FLINX]  P
FOSTER, ALAN DEAN  SLIPT  C/P:2,P

FOSTER, ALAN DEAN/ GREENBERG, MARTIN H.

SMART DRAGONS, FOOLISH ELVES

C/PA,PA,SFWA-P

FOSTER, ALAN DEAN  SON OF SPELLSINGER [SPELLSINGER 7]  C/P:2,SFWA-P,P
FOSTER, ALAN DEAN  SPELLSINGER [SPELLSINGER 1]  C/P:2,P
FOSTER, ALAN DEAN  SPELLSINGER’S SCHERZO [SPELLSINGER 4,5,6]  H
FOSTER, ALAN DEAN  SPLINTER OF THE MIND’S EYE [STAR WARS]  H:2,P
FOSTER, ALAN DEAN  SPOILS OF WAR, THE [@DAMNED #3]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 1 [STAR TREK]  C/P:3,P
FOSTER, ALAN DEAN  STAR TREK LOG ONE LOG TWO LOG THREE [STAR TREK]  P
FOSTER, ALAN DEAN  STAR TREK LOG 2 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 3 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 4 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG FOUR LOG FIVE LOG SIX [STAR TREK]  P
FOSTER, ALAN DEAN  STAR TREK LOG 5 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 6 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 7 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG SEVEN LOG EIGHT LOG NINE [STAR TREK]  P
FOSTER, ALAN DEAN  STAR TREK LOG 8 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 9 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STAR TREK LOG 10 [STAR TREK]  C/P:2,P
FOSTER, ALAN DEAN  STARMAN  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 148
AUTHOR  TITLE  CODE
FOSTER, ALAN DEAN  TAR-AIYM KRANG, THE [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  THING, THE  P
FOSTER, ALAN DEAN  TIME OF THE TRANSFERENCE, THE [SPELLSINGER 6]  C/P,P,H
FOSTER, ALAN DEAN  TO THE VANISHING POINT  H:2,P
FOSTER, ALAN DEAN  TRIUMPH OF SOULS, A [JOURNEYS OF THE CATECHIST 3]  H
FOSTER, ALAN DEAN  TROUBLE MAGNET [HUMANX FLINX]  H
FOSTER, ALAN DEAN  VOYAGE TO THE CITY OF THE DEAD [HUMANX]  C/P:2,P
FOSTER, ALAN DEAN  ...WHO NEEDS ENEMIES? [HUMANX FLINX]  C/P:2,P
FOSTER, ALAN DEAN  WITH FRIENDS LIKE THESE...  C/P:2,P
FOSTER, GEORGE C.  CHANGE, THE  SR-P
FOSTER, M. A.  DAY OF THE KLESH, THE [LER 3]  C/P:2,P
FOSTER, M. A.  GAMEPLAYERS OF ZAN, THE [LER 1]  C/P:2,P
FOSTER, M. A.  MORPHODITE, THE [MORPHODITE 1]  C/P:2,P
FOSTER, M. A.  OWL TIME  C/P:2,P
FOSTER, M. A.  PRESERVER [MORPHODITE 3]  C/P:2,P
FOSTER, M. A.  TRANSFORMER [MORPHODITE 2]  C/P:2,P
FOSTER, M. A.  WARRIORS OF DAWN, THE [LER 2]  C/P:2,P
FOSTER, M. A.  WAVES  C/P:2,P
FOSTER, RABBI SOLOMON  SPIRITUAL TRAILS TO HAPPINESS  [C/RAN-H]
FOSTER, RICHARD  REST MUST DIE, THE  P

FOSTER, ROBERT COMPLETE GUIDE TO MIDDLE-EARTH:

FROM THE HOBBIT TO THE SILMARILLION, THE

R/REF-P,C/REF-P

FOSTER, ROBERT  GUIDE TO MIDDLE-EARTH, A  R/REF-H,C/REF-P
FOWLER, CHRISTOPHER  RED BRIDE  SFWA
FOWLER, CHRISTOPHER  ROOFWORLD  P
FOWLER, KAREN JOY  ARTIFICIAL THINGS  C/P:2,P

FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

JAMES TIPTREE AWARD ANTHOLOGY 1, THE

LPA

FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

JAMES TIPTREE AWARD ANTHOLOGY 2, THE

LPA

FOWLER, KAREN JOY/MURPHY, PAT/ NOTKIN, DEBBIE/SMITH, JEFFREY D.

JAMES TIPTREE AWARD ANTHOLOGY 3, THE

LPA

FOWLER, RAYMOND E.  WATCHERS, THE  [C/RAN-H]
FOWLES, JOHN  MAGUS, THE  P

FOX, ANDREW BRIDE OF THE FAT WHITE VAMPIRE

[FAT WHITE VAMPIRE (JULES DUCHON) 2]

LP

FOX, ANDREW  FAT WHITE VAMPIRE BLUES [FAT WHITE VAMPIRE (JULES DUCHON) 1]  LP
FOX, GARDNER F.  ARSENAL OF MIRACLES, THE  P,DF299
FOX, GARDNER F.  CONEHEAD  C/P:2,P
FOX, GARDNER F.  ESCAPE ACROSS THE COSMOS  C/P:2,P
FOX, GARDNER F.  FIVE WEEKS IN A BALLOON  P
FOX, GARDNER F.  HUNTER OUT OF TIME, THE  C/P:2,P
FOX, GARDNER F.  KOTHAR AND THE CONJURER’S CURSE [KOTHAR 4]  C/P,P
FOX, GARDNER F.  KOTHAR AND THE DEMON QUEEN [KOTHAR 3]  P
FOX, GARDNER F.  KOTHAR AND THE WIZARD SLAYER [KOTHAR 5]  C/P,P
FOX, GARDNER F.  KOTHAR — BARBARIAN SWORDSMAN [KOTHAR 1]  P
FOX, GARDNER F.  KOTHAR OF THE MAGIC SWORD [KOTHAR 2]  P
FOX, GARDNER F.  KYRIK AND THE LOST QUEEN [KYRIK 4]  P
FOX, GARDNER F.  KYRIK AND THE WIZARD’S SWORD [KYRIK 3]  P
FOX, GARDNER F.  KYRIK FIGHTS THE DEMON WORLD [KYRIK 2]  P
FOX, GARDNER F.  KYRIK: WARLOCK WARRIOR [KYRIK 1]  P
FOX, GARDNER F.  THIEF OF LLARN [LLARN 2]  C/P:2,P
FOX, GARDNER F.  WARRIOR OF LLARN [LLARN 1]  C/P:2,P
FOX, H. MUNRO  PERSONALITY OF ANIMALS, THE  [C/RAN-P]
FOX, W. RANDOLPH  AFTER THE APOCALYPSE  P
FOXE, JOCELIN  WILD HUNT: CHILD OF FIRE, THE [WILD HUNT 2]  P
FOXE, JOCELIN  WILD HUNT: VENGEANCE MOON, THE [WILD HUNT 1]  P
FOXXE, ELLEN  SEASON OF SHADOWS [@SUMMERLANDS #1]  P
FOXXE, ELLEN  SEASON OF STORMS [@SUMMERLANDS #2]  P
FOY, GEORGE  CONTRABAND  H,P
FOY, GEORGE  LAST HARBOR, THE  P
FOY, GEORGE  MEMORY OF FIRE, THE  P
FOY, GEORGE  SHIFT, THE  P
FRAKES, JONATHAN/SMITH, DEAN WESLEY  ABDUCTORS: CONSPIRACY, THE  H,P
FRAKES, RANDALL/WISHER, BILL  TERMINATOR, THE [TERMINATOR]  P
FRAKES, RANDALL  TERMINATOR 2 [TERMINATOR]  C/P,P
FRANCE, ANATOLE  PENGUIN ISLAND  H
FRANCE, ANATOLE  REVOLT OF THE ANGELS, THE  H
FRANCIS, DIANA PHARAOH  PATH OF BLOOD [REISIL 3]  C/P,P
FRANCIS, DIANA PHARAOH  PATH OF FATE [REISIL 1]  C/P,P
FRANCIS, DIANA PHARAOH  PATH OF HONOR [REISIL 2]  C/P,P
FRANCIS, DOROTHY BRENNER  BLINK OF THE MIND [TWILIGHT #5]  P
FRANCIS, H. G.  GROSSEN VON AGOPP, DIE [@PERRY RHODAN #138]  FPR
FRANCOIS, YVES REGIS  CTZ PARADIGM, THE  H,SFWA

The MITSFS Pinkdex  September 24, 2007  Page 149
AUTHOR  TITLE  CODE
FRANE, JEFF  FRITZ LEIBER  R/REF-P
FRANK, ALAN  GALACTIC ALIENS  A
FRANK, ALAN  SCI-FI NOW  R/REF-H
FRANK, JACQUELYN  GIDEON [@NIGHTWALKERS]  P
FRANK, JACQUELYN  JACOB [@NIGHTWALKERS]  P
FRANK, PAT  ALAS, BABYLON  C/P:2,P,H
FRANK, PAT  FORBIDDEN AREA  H:2,P
FRANK, PAT  MR. ADAM  H:2,P
FRANK, SUZANNE  SUNRISE OF THE MEDITERRANEAN  H
FRANKE, HERBERT W.  GEDANKENNETZ, DAS  FH
FRANKE, HERBERT W.  GRUNE KOMET, DER  FH
FRANKE, HERBERT W.  MIND NET, THE  P
FRANKE, HERBERT W.  ORCHID CAGE, THE  C/P,P
FRANKE, HERBERT W.  ORCHIDEENKAFIG, DER  FH
FRANKE, HERBERT W.  SCIENCE FICTION STORY-READER 4 (GERMAN)  FP
FRANKE, HERBERT W.  ZONE NULL  SFWA
FRANKLIN, CHERYL J.  FIRE CROSSING [NETWORK-CONSORTIUM 2]  C/P:2,P
FRANKLIN, CHERYL J.  FIRE GET [TAORMIN 1]  C/P:2,P
FRANKLIN, CHERYL J.  FIRE LORD [TAORMIN 2]  C/P:2,P
FRANKLIN, CHERYL J.  GHOST SHADOW [NETWORK-CONSORTIUM 4]  P
FRANKLIN, CHERYL J.  INQUISITOR, THE [NETWORK-CONSORTIUM 3]  C/P:2,P
FRANKLIN, CHERYL J.  LIGHT IN EXILE, THE [NETWORK-CONSORTIUM 1]  C/P,P
FRANKLIN, CHERYL J.  SABLE, SHADOW, AND ICE  C/P,P
FRANKLIN, H. BRUCE  COUNTDOWN TO MIDNIGHT  PA
FRANKLIN, H. BRUCE  FUTURE PERFECT  LPA
FRANKLIN, H. BRUCE  ROBERT A. HEINLEIN: AMERICA AS SCIENCE FICTION  R/REF-H
FRANKLIN, JAY  RAT RACE, THE  C/P,GN2.10,SR-LP
FRANKLIN, KAMERON M.  MAIDEN OF PAIN [@FORGOTTEN REALMS PRIESTS]  P
FRANKLIN, MAX  DARK, THE  P
FRANKLIN, MAX  STARSKY AND HUTCH #6: THE PSYCHIC  P
FRANKOS, STEVEN  BEYOND LICH GATE [ELIXIR OF LIFE 1]  P
FRANKOS, STEVEN  CATHEDRAL OF THORNS [ELIXIR OF LIFE 2]  P
FRANKOS, STEVEN  DARKLIGHT GRIMOIRE, THE [WHEEL 3]  P
FRANKOS, STEVEN  HEART OF SPARRILL, THE [WHEEL 2]  P
FRANKOS, STEVEN  JEWEL OF EQUILIBRANT, THE [WHEEL 1]  P
FRANKOS, STEVEN  LEGEND REBORN, A [ELIXIR OF LIFE 3]  P
FRANKOWSKI, LEO  BOY AND HIS TANK, A  H,LP
FRANKOWSKI, LEO  CONRAD’S QUEST FOR RUBBER [CONRAD STARGARD 6]  C/P,P
FRANKOWSKI, LEO  CONRAD’S TIME MACHINE [CONRAD STARGARD 0]  H
FRANKOWSKI, LEO  COPERNICK’S REBELLION  P
FRANKOWSKI, LEO  CROSS-TIME ENGINEER, THE [CONRAD STARGARD 1]  C/P:2,P
FRANKOWSKI, LEO  FATA MORGANA, THE  C/P,LP
FRANKOWSKI, LEO  FLYING WARLORD, THE [CONRAD STARGARD 4]  C/P:2,P
FRANKOWSKI, LEO  HIGH-TECH KNIGHT, THE [CONRAD STARGARD 2]  C/P:2,P
FRANKOWSKI, LEO  LORD CONRAD’S LADY [CONRAD STARGARD 5]  C/P:2,P
FRANKOWSKI, LEO  RADIANT WARRIOR, THE [CONRAD STARGARD 3]  C/P:2,P
FRANKOWSKI, LEO/GROSSMAN, DAVE  WAR WITH EARTH, THE  P

FRANSON, DONALD/DEVORE, HOWARD HISTORY OF THE HUGO, NEBULA,

AND INTERNATIONAL FANTASY AWARDS, A (1978)

R/REF-P

FRANSON, ROBERT WILFRED  SHADOW OF THE SHIP, THE  C/P:2,P
FRASER, GEORGE MACDONALD  FLASHMAN [FLASHMAN]  P
FRASER, GEORGE MACDONALD  FLASHMAN AT THE CHARGE [FLASHMAN]  P
FRAYN, MICHAEL  TIN MEN, THE  H:2,P
FRAYN, MICHAEL  VERY PRIVATE LIFE, A  C/P,P
FRAYN, MICHAEL  VIE TRES PRIVEE, UNE  FP
FRAZEE, STEVE  SKY BLOCK, THE  C/P,P
FRAZETTA, FRANK  FANTASTIC ART OF FRANK FRAZETTA, THE  A
FREAS, FRANK KELLY  FRANK KELLY FREAS: THE ART OF SCIENCE FICTION  A
FREDD, STEPHEN  SERUM 2000  P
FREDDI, CHRIS  ELDER, THE  H
FREE, COLIN  SOFT KILL, THE  C/P,P
FREEDMAN, NANCY  JOSHUA SON OF NONE  C/P,P,H
FREEMAN, BARBARA E.  BROOM-ADELAIDE  H
FREEMAN, LORNA  COVENANTS [BORDERLANDS (IVERSTERRE)]  P
FREEMAN, LORNA  KING’S OWN, THE [@BORDERLANDS (IVERSTERRE)]  P
FREEMAN, MARY WILKINS  WIND IN THE ROSE-BUSH, THE  P
FREEMANTLE, BRIAN  BETRAYALS  SFWA-P
FREEMANTLE, BRIAN  DEAKEN’S WAR  SFWA-P
FREER, DAVE  FORLORN, THE  C/P,P
FREER, DAVE  MANKIND WITCH, A [SHADOW OF THE LION 3]  H
FREER, DAVE/FLINT, ERIC  PYRAMID SCHEME  H
FREER, DAVE/FLINT, ERIC  RATS, BATS, & VATS [RATS & BATS 1]  H
FREIREICH, VALERIE J.  BEACON, THE  C/P,P
FREIREICH, VALERIE J.  BECOMING HUMAN [HARMONY OF WORLDS]  P

The MITSFS Pinkdex  September 24, 2007  Page 150
AUTHOR  TITLE  CODE
FREIREICH, VALERIE J.  IMPOSTOR  C/P,P
FREIREICH, VALERIE J.  TESTAMENT [HARMONY OF WORLDS]  C/P,P
FREMION, YVES  REVES DE SABLE, CHATEAUX DE SANG  FP
FRENCH, JOSEPH L.  TALES OF TERROR  HA
FRENCH, ROBERT  NOW, THE GODS  P
FRENCH, ROBERT  SAND OF SATURN  P
FRENCH, ROBERT  THEM  P
FRENKEL, JAMES  BANGS AND WHIMPERS  HA
FRENKEL, JAMES  TRUE NAMES AND THE OPENING OF THE CYBERSPACE FRONTIER  LPA
FRETLAND, DON J.  PERSIMMION SEQUENCE, THE  SR-P,C/P
FRETLAND, DON J.  WINDS OF THE HELIOPOLIS  SR-P
FREWER, GLYN  ADVENTURE IN FORGOTTEN VALLEY  H
FREWIN, ANTHONY  ONE HUNDRED YEARS OF SCIENCE FICTION ILLUSTRATION  A
FREZZA, ROBERT  CAIN’S LAND [VERESHCHAGIN 3]  P
FREZZA, ROBERT  FIRE IN A FARAWAY PLACE [VERESHCHAGIN 2]  P
FREZZA, ROBERT  MCLENDON’S SYNDROME  C/P,P,H
FREZZA, ROBERT  SMALL COLONIAL WAR, A [VERESHCHAGIN 1]  C/P:2,P
FREZZA, ROBERT  VMR THEORY, THE  P
FRIEDBERG, GERTRUDE  REVOLVING BOY, THE  C/P,P,H
FRIEDELL, EGON  RETURN OF THE TIME MACHINE, THE  C/P:2,P
FRIEDMAN, C. S.  BLACK SUN RISING [COLDFIRE 1]  H:2,P
FRIEDMAN, C. S.  CROWN OF SHADOWS [COLDFIRE 3]  H,P
FRIEDMAN, C. S.  FEAST OF SOULS [@MAGISTER #1]  H
FRIEDMAN, C. S.  IN CONQUEST BORN  C/P:2,P
FRIEDMAN, C. S.  IN CONQUEST BORN (REVISED)  P
FRIEDMAN, C. S.  MADNESS SEASON, THE  C/P:2,P
FRIEDMAN, C. S.  THIS ALIEN SHORE  C/P,P,H
FRIEDMAN, C. S.  WHEN TRUE NIGHT FALLS [COLDFIRE 2]  H:2,P
FRIEDMAN, C. S.  WILDING, THE  H:2
FRIEDMAN, FAVIUS  GREAT HORROR MOVIES  R/REF-P,C/REF-P
FRIEDMAN, MICHAEL JAN  ALL GOOD THINGS [STAR TREK TNG]  H:2,P
FRIEDMAN, MICHAEL JAN/DELANCIE, JOHN BEST OF STAR TREK, THE NEXT GENERATION, THE  CX
FRIEDMAN, MICHAEL JAN  CALL TO DARKNESS, A [STAR TREK TNG]  C/P:2,P
FRIEDMAN, MICHAEL JAN  CONSTITUTION [@STAR TREK MY BROTHER’S KEEPER #2]  P
FRIEDMAN, MICHAEL JAN  CROSSOVER [STAR TREK TNG]  P
FRIEDMAN, MICHAEL JAN  DEATH IN WINTER [STAR TREK TNG]  H
FRIEDMAN, MICHAEL JAN  DOUBLE, DOUBLE [STAR TREK]  C/P:2,P
FRIEDMAN, MICHAEL JAN  DUJONIAN’S HOARD [STAR TREK CAPTAIN’S TABLE 2/STAR TREK TNG]  C/P,P
FRIEDMAN, MICHAEL JAN  ENIGMA [@STAR TREK TNG STARGAZER]  P
FRIEDMAN, MICHAEL JAN  ENTERPRISE [@STAR TREK MY BROTHER’S KEEPER #3]  P
FRIEDMAN, MICHAEL JAN  FACES OF FIRE [STAR TREK]  C/P:2,P

FRIEDMAN, MICHAEL JAN/ GOLDEN, CHRISTIE FIRST VIRTUE, THE [@STAR TREK TNG DOUBLE HELIX 6]

P

FRIEDMAN, MICHAEL JAN  FORTRESS AND THE FIRE, THE  C/P:2,P
FRIEDMAN, MICHAEL JAN  FORTUNE’S LIGHT [STAR TREK TNG]  C/P:2,P
FRIEDMAN, MICHAEL JAN  GAUNTLET [@STAR TREK TNG STARGAZER #1]  P
FRIEDMAN, MICHAEL JAN  GLOVE OF MAIDEN’S HAIR, THE  P
FRIEDMAN, MICHAEL JAN  HAMMER AND THE HORN, THE  C/P,P
FRIEDMAN, MICHAEL JAN  HER KLINGON SOUL [STAR TREK DAY OF HONOR 3/STAR TREK VOYAGER] C/P:2,P
FRIEDMAN, MICHAEL JAN  KAHLESS [STAR TREK TNG]  P
FRIEDMAN, MICHAEL JAN  LEGACY [STAR TREK]  C/P:3,P
FRIEDMAN, MICHAEL JAN  MAKER [@STAR TREK TNG STARGAZER]  P

FRIEDMAN, MICHAEL JAN MYSTERY OF THE MISSING CREW

[@STAR TREK TNG STARFLEET ACADEMY #6]

H,LP

FRIEDMAN, MICHAEL JAN  OBLIVION [@STAR TREK TNG STARGAZER]  P
FRIEDMAN, MICHAEL JAN  PLANET X [@STAR TREK TNG/X-MEN]  C/P,P
FRIEDMAN, MICHAEL JAN  PROGENITOR [@STAR TREK TNG STARGAZER #2]  P
FRIEDMAN, MICHAEL JAN  RELICS [STAR TREK TNG]  C/P,P
FRIEDMAN, MICHAEL JAN  REPUBLIC [@STAR TREK MY BROTHER’S KEEPER #1]  P
FRIEDMAN, MICHAEL JAN/RYAN, KEVIN  REQUIEM [STAR TREK TNG]  C/P:3,P
FRIEDMAN, MICHAEL JAN  REUNION [STAR TREK TNG]  C/P:2,P
FRIEDMAN, MICHAEL JAN  SARATOGA [STAR TREK DS9]  P
FRIEDMAN, MICHAEL JAN  SEEKERS AND THE SWORD, THE  C/P:2,P
FRIEDMAN, MICHAEL JAN  SHADOWS ON THE SUN [STAR TREK]  C/P:2,P

FRIEDMAN, MICHAEL JAN/KRAUSE, PETER/ MARCO, PABLO STAR LOST, THE [STAR TREK,THE NEXT GENERATION]

CX

FRIEDMAN, MICHAEL JAN  STAR TREK FEDERATION TRAVEL GUIDE [STAR TREK]  B
FRIEDMAN, MICHAEL JAN  STARFLEET YEAR ONE [STAR TREK]  P
FRIEDMAN, MICHAEL JAN  THREE [@STAR TREK TNG STARGAZER #3]  P
FRIEDMAN, MICHAEL JAN  VALIANT, THE [STAR TREK TNG STARGAZER]  P
FRIEL, ARTHUR O.  PATHLESS TRAIL, THE  C/P,P
FRIEL, ARTHUR O.  TIGER RIVER  C/P,P,H
FRIEND, BEVERLY  SCIENCE FICTION: THE CLASSROOM IN ORBIT  R/REF-P
FRIEND, ED  GREEN HORNET IN THE INFERNAL LIGHT, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 151
AUTHOR  TITLE  CODE
FRIEND, OSCAR J.  KID FROM MARS, THE  H
FRIEND, OSCAR J.  STAR MEN, THE  H

FRIESNER, ESTHER M./ GREENBERG, MARTIN H.

ALIEN PREGNANT BY ELVIS

C/PA,PA

FRIESNER, ESTHER M.  CHICK IS IN THE MAIL, THE [CHICKS IN CHAINMAIL 5]  PA,C/PA
FRIESNER, ESTHER M.  CHICKS IN CHAINMAIL [CHICKS IN CHAINMAIL 1]  PA
FRIESNER, ESTHER M.  CHICKS ’N CHAINED MALES [CHICKS IN CHAINMAIL 3]  HA,PA
FRIESNER, ESTHER M.  CHILD OF THE EAGLE  C/P,P
FRIESNER, ESTHER M.  DEMON BLUES [HERE BE DEMONS 2]  P
FRIESNER, ESTHER M.  DID YOU SAY CHICKS?! [CHICKS IN CHAINMAIL 2]  C/PA,PA
FRIESNER, ESTHER M.  DRUID’S BLOOD  P
FRIESNER, ESTHER M.  ELF DEFENCE  C/P:2,P
FRIESNER, ESTHER M.  GNOME MAN’S LAND [GNOME MAN’S LAND 1]  C/P:2,SFWA-P:2,P
FRIESNER, ESTHER M.  HARLOT’S RUSE  P
FRIESNER, ESTHER M.  HARPY HIGH [GNOME MAN’S LAND 2]  C/P,P
FRIESNER, ESTHER M.  HERE BE DEMONS [HERE BE DEMONS 1]  C/P,P
FRIESNER, ESTHER M.  HOORAY FOR HELLYWOOD [HERE BE DEMONS 3]  P,SFWA-P:2
FRIESNER, ESTHER M.  MAJYK BY ACCIDENT [KENDAR GANGLE 1]  C/P,SFWA-P,P
FRIESNER, ESTHER M.  MAJYK BY DESIGN [KENDAR GANGLE]  C/P,P
FRIESNER, ESTHER M.  MAJYK BY HOOK OR CROOK [KENDAR GANGLE 2]  P,SFWA-P
FRIESNER, ESTHER M.  MEN IN BLACK II [MEN IN BLACK]  P
FRIESNER, ESTHER M.  MUSTAPHA AND HIS WISE DOG [TWELVE KINGDOMS 1]  C/P:2,P
FRIESNER, ESTHER M.  NEW YORK BY KNIGHT  C/P,P
FRIESNER, ESTHER M.  PSALMS OF HEROD, THE [BECCA 1]  P
FRIESNER, ESTHER M.  SHERWOOD GAME, THE  P
FRIESNER, ESTHER M.  SILVER MOUNTAIN, THE  C/P,P
FRIESNER, ESTHER M.  SPELLS OF MORTAL WEAVING [TWELVE KINGDOMS 2]  C/P,P
FRIESNER, ESTHER M.  SPHYNXES WILD  P
FRIESNER, ESTHER M.  SWORD OF MARY, THE [BECCA 2]  P
FRIESNER, ESTHER M.  TEMPING FATE  H
FRIESNER, ESTHER M.  TO STORM HEAVEN [STAR TREK TNG]  C/P,P
FRIESNER, ESTHER M.  TURN THE OTHER CHICK  HA,PA
FRIESNER, ESTHER M.  UNICORN U. [GNOME MAN’S LAND 3]  C/P,SFWA-P:2,P
FRIESNER, ESTHER M.  UP THE WALL  LP
FRIESNER, ESTHER M.  WARCHILD [STAR TREK DS9]  C/P:2,P
FRIESNER, ESTHER M.  WARCHILD (PROOF) [STAR TREK DS9]  SR-VLP
FRIESNER, ESTHER M.  WATER KING’S LAUGHTER, THE [TWELVE KINGDOMS 4]  C/P:2,P
FRIESNER, ESTHER M.  WISHING SEASON  C/P:2,P
FRIESNER, ESTHER M.  WITCHWOOD CRADLE, THE [TWELVE KINGDOMS 3]  P
FRIESNER, ESTHER M.  YESTERDAY WE SAW MERMAIDS  H
FRITCH, CHARLES E.  HORSES’ ASTEROID  P
FRITZHAND, JAMES/GLICKSMAN, FRANK  UNICORN AFFAIR, THE  P
FROST, GREGORY  ATTACK OF THE JAZZ GIANTS  H
FROST, GREGORY  CRIMSON SPEAR  LP
FROST, GREGORY  FITCHER’S BRIDES  LP
FROST, GREGORY  LYREC  C/P:2,P
FROST, GREGORY  PURE COLD LIGHT, THE  P,SFWA-P
FROST, GREGORY  REMSCELA  C/P,P
FROST, GREGORY  TAIN  C/P,P
FROST, JASON  BADLAND [@WARLORD #3]  P
FROST, JASON  CUTTHROAT, THE [@WARLORD #2]  P
FROST, JASON  KILLER’S KEEP [@WARLORD #6]  P
FROST, JASON  PRISONLAND [@WARLORD #4]  P
FROST, JASON  TERMINAL ISLAND [@WARLORD #5]  P
FROST, JASON  WARLORD, THE [@WARLORD #1]  P
FROST, MARK  LIST OF 7, THE  H
FROUD, BRIAN/LEE, ALAN  FAERIES (TWENTY-FIFTH ANNIVERSARY EDITION)  A
FROUD, BRIAN  WORLD OF THE DARK CRYSTAL, THE  A
FRY, STEPHEN  MAKING HISTORY  H,LP
FULLER, JOHN G.  INCIDENT AT EXETER  [C/RAN-P]
FULLER, JOHN G.  INTERRUPTED JOURNEY, THE  [C/RAN-H,C/RAN-P]
FULLER, JOHN G.  PACK, THE  P
FULLER, KIMBERLY  HOME  H,P
FULLER, ROGER  SON OF FLUBBER  P
FULLILOVE, ERIC JAMES  CIRCLE OF ONE [JENNY SIXA 1]  C/P,P
FULLILOVE, ERIC JAMES  STRANGER, THE [JENNY SIXA 2]  P
FUNKE, CORNELIA  DRAGON RIDER  H
FUNNELL, AUGUSTINE  BRANDYJACK  SR-P
FUNNELL, AUGUSTINE  REBELS OF MERKA  SR-P
FUREY, MAGGIE  AURIAN [AURIAN 1]  C/P,P
FUREY, MAGGIE  DHIAMMARA [AURIAN 4]  P
FUREY, MAGGIE  ECHO OF ETERNITY [SHADOWLEAGUE 3]  C/P,P
FUREY, MAGGIE  HARP OF WINDS [AURIAN 2]  C/P:2,P
FUREY, MAGGIE  HEART OF MYRIAL, THE [SHADOWLEAGUE 1]  P

AUTHOR  TITLE  CODE
FUREY, MAGGIE  SORCERESS [WEB,THE]  P
FUREY, MAGGIE  SPIRIT OF THE STONE, THE [SHADOWLEAGUE 2]  P
FUREY, MAGGIE  SWORD OF FLAME [AURIAN 3]  C/P:2,P
FURMAN, A. L.  GHOST STORIES  PA
FURMAN, A. L.  HAUNTED STORIES  PA
FURMAN, A. L.  MORE GHOST STORIES  PA
FURMAN, A. L.  OUTER SPACE STORIES  C/PA,PA
FURMAN, A. L.  SPACE ADVENTURES  PA
FURMAN, A. L.  TEEN-AGE GHOST STORIES  HA
FURMAN, A. L.  TEEN-AGE OUTER SPACE STORIES  HA
FURNAS, C. C./MCCARTHY, JOE  ENGINEER, THE  [C/RAN-H]

FURTH, ROBIN STEPHEN KING’S THE DARK TOWER A CONCORDANCE VOLUME I

[DARK TOWER]

R/REF-P

FUTRELLE, JACQUES  BEST "THINKING MACHINE" DETECTIVE STORIES  LP
FUTRELLE, JACQUES  THINKING MACHINE, THE  LP
FUTURE, STEVE  DOOMED NATION OF THE SKIES  SR-P
FUTURE, STEVE  SLAVE TRADERS OF THE SKY  SR-P
FYFE, H. B.  D-99  C/P:2,P
FYFE, H. B.  MOONWALK  B
FYSH  PLANET WAR  SR-P

The MITSFS Pinkdex  September 24, 2007  Page 153
AUTHOR  TITLE  CODE
GABALDON, DIANA  BREATH OF SNOW AND ASHES, A [OUTLANDER (GABALDON) 6]  H
GABALDON, DIANA  DRAGONFLY IN AMBER [OUTLANDER (GABALDON) 2]  C/P,P
GABALDON, DIANA  OUTLANDER [OUTLANDER (GABALDON) 1]  H:2,P
GABALDON, DIANA  VOYAGER [OUTLANDER (GABALDON) 3]  C/P,P,H
GADALLAH, LESLIE  CAT’S GAMBIT [ORIANI 2]  C/P,P
GADALLAH, LESLIE  CAT’S PAWN [ORIANI 1]  P
GADALLAH, LESLIE  LOREMASTERS, THE  C/P:2,P
GADDIS, VINCENT H.  MYSTERIOUS FIRES AND LIGHTS  [C/RAN-P]
GAGE, WILSON  MISS OSBORNE-THE-MOP  P
GAIL, O. W./PETRI, W.  RUIMTEVAART  FP
GAIMAN, NEIL  ABSOLUTE SANDMAN, VOL. 1 [SANDMAN]  CX
GAIMAN, NEIL  ADVENTURES IN THE DREAM TRADE  SR-H
GAIMAN, NEIL  AMERICAN GODS  C/P:2,P
GAIMAN, NEIL  ANANSI BOYS, THE  H:2
GAIMAN, NEIL  ANGELS & VISITATIONS  SR-H,H:2
GAIMAN, NEIL/MCKEAN, DAVE  BLACK ORCHID  CX
GAIMAN, NEIL  BOOKS OF MAGIC  CX
GAIMAN, NEIL  BRIEF LIVES [@SANDMAN #7]  H:2,CX
GAIMAN, NEIL  CORALINE  H:2
GAIMAN, NEIL  DEATH: THE HIGH COST OF LIVING  CX
GAIMAN, NEIL  DEATH: THE TIME OF YOUR LIFE  CX
GAIMAN, NEIL  DOLL’S HOUSE, THE [@SANDMAN #2]  H:2,CX
GAIMAN, NEIL  DON’T PANIC [HITCHHIKERS]  R/REF-P
GAIMAN, NEIL  DREAM COUNTRY [@SANDMAN #3]  H,CX
GAIMAN, NEIL  DREAM HUNTERS, THE [@SANDMAN]  CX
GAIMAN, NEIL  FABLES & REFLECTIONS [@SANDMAN #6]  H,CX
GAIMAN, NEIL  FRAGILE THINGS  H
GAIMAN, NEIL  GAME OF YOU, A [@SANDMAN #5]  H,CX
GAIMAN, NEIL/PRATCHETT, TERRY  GOOD OMENS  H:2,SFWA,LP
GAIMAN, NEIL  KINDLY ONES, THE [@SANDMAN #9]  H,CX
GAIMAN, NEIL  M IS FOR MAGIC  H
GAIMAN, NEIL  M IS FOR MAGIC (PROOF)  SR-LP
GAIMAN, NEIL/WAGNER, MATT  NEIL GAIMAN’S MIDNIGHT DAYS  H,CX
GAIMAN, NEIL  NEVERWHERE  C/P:2,P
GAIMAN, NEIL  PRELUDES & NOCTURNES [@SANDMAN #1]  H,CX
GAIMAN, NEIL/KRAMER, ED  SANDMAN BOOK OF DREAMS, THE [SANDMAN]  LPA
GAIMAN, NEIL  SEASON OF MISTS [@SANDMAN #4]  H,CX
GAIMAN, NEIL  SMOKE AND MIRRORS  H
GAIMAN, NEIL  STARDUST  C/P,P,H
GAIMAN, NEIL/STEWART, ALEX  TEMPS [@TEMPS 1]  SR-PA
GAIMAN, NEIL  WAKE, THE [@SANDMAN #10]  CX
GAIMAN, NEIL  WORLDS’ END [@SANDMAN #8]  H,CX
GALANTER, DAVE/BRODEUR, GREG  BATTLE LINES [@STAR TREK VOYAGER]  C/P:2,P
GALANTER, DAVE/BRODEUR, GREG  DEAD ZONE [@STAR TREK TNG MAXIMUM WARP #1]  P
GALANTER, DAVE/BRODEUR, GREG  FOREIGN FOES [STAR TREK TNG]  C/P:2,P
GALANTER, DAVE/BRODEUR, GREG  FOREVER DARK [@STAR TREK TNG MAXIMUM WARP #2]  P
GALASSIA  AMORE A QUATTRO DIMENSIONI  FP
GALASSIA  DESTINAZIONE UOMO  FP
GALASSIA  SEDICI MAPPE DEL NOSTRO FUTURO  FP
GALAXY  BEST FROM GALAXY 1, THE  C/PA,PA
GALAXY  BEST FROM GALAXY 2, THE  C/PA,PA
GALENORN, YASMINE  CHANGELING [OTHERWORLD]  P
GALENORN, YASMINE  WITCHLING [OTHERWORLD]  C/P,P
GALIN, MITCHELL/ALLEN, TOM  TALES FROM THE DARKSIDE 1  PA
GALLAGHER, DIANA G.  ALIEN DARK, THE  C/P,P
GALLAGHER, DIANA G.  BAD BARGAIN [@BUFFY]  P
GALLAGHER, DIANA G.  BEWARE WHAT YOU WISH [@CHARMED]  P
GALLAGHER, DIANA G.  BRIDAL BEDLAM [@SABRINA THE TEENAGE WITCH #23]  P
GALLAGHER, DIANA G.  DARK VENGEANCE [@CHARMED]  P
GALLAGHER, DIANA G.  DOG DAY AFTERNOON [@SABRINA THE TEENAGE WITCH SALEM #5]  P
GALLAGHER, DIANA G.  DOOMSDAY DECK [BUFFY]  P
GALLAGHER, DIANA G.  HALLOWEEN HAVOC [@SABRINA THE TEENAGE WITCH #4]  P
GALLAGHER, DIANA G.  HURRICANE HEX [@CHARMED]  P
GALLAGHER, DIANA G.  LOTSA LUCK [@SABRINA THE TEENAGE WITCH #10]  P
GALLAGHER, DIANA G.  MIST AND STONE [@CHARMED]  P
GALLAGHER, DIANA G.  MYSTIC KNOLL [@CHARMED]  P

GALLAGHER, DIANA G.

NOW YOU SEE HER, NOW YOU DON’T [@SABRINA THE TEENAGE WITCH #16]

P

GALLAGHER, DIANA G.  OBSIDIAN FATE [@BUFFY]  C/P,P
GALLAGHER, DIANA G.  PRIME EVIL [@BUFFY]  C/P,P
GALLAGHER, DIANA G.  SHADOWS [@SMALLVILLE]  P
GALLAGHER, DIANA G.  SHADOWS (PROOF) [@SMALLVILLE]  SR-P,C/P
GALLAGHER, DIANA G.  SHAMROCK SHENANIGANS [@SABRINA THE TEENAGE WITCH #19]  P
GALLAGHER, DIANA G.  SHOWDOWN AT THE MALL [@SABRINA THE TEENAGE WITCH #2]  P

AUTHOR  TITLE  CODE
GALLAGHER, DIANA G.  SPARK AND BURN [@BUFFY]  P
GALLAGHER, DIANA G.  SPIRIT OF THE WOLF [@CHARMED]  P
GALLAGHER, FRED  MEGATOKYO 2  SCX
GALLAGHER, STEPHEN  VALLEY OF LIGHTS  SFWA-P
GALLERY BOOKS  SPACE ODYSSEY  HA
GALLUN, RAYMOND Z.  BEST OF RAYMOND Z. GALLUN, THE  C/P,P
GALLUN, RAYMOND Z.  BIOBLAST  C/P,P
GALLUN, RAYMOND Z.  DOUBLE IDENTITY  B
GALLUN, RAYMOND Z.  EDEN CYCLE, THE  C/P,P
GALLUN, RAYMOND Z.  GUTHRIE METHOD, THE  B
GALLUN, RAYMOND Z.  PEOPLE MINUS X  DD291
GALLUN, RAYMOND Z.  PLANET STRAPPERS, THE  C/P,P
GALLUN, RAYMOND Z.  SKYCLIMBER  P
GALOUYE, DANIEL F.  DARK UNIVERSE  C/P:2,P
GALOUYE, DANIEL F.  INFINITE MAN, THE  C/P:2,P
GALOUYE, DANIEL F.  LORDS OF THE PSYCHON  C/P:2,P
GALOUYE, DANIEL F.  SCOURGE OF SCREAMERS, A  C/P:2,P
GALOUYE, DANIEL F.  SIMULACRON-3  C/P:2,P
GAMOW, GEORGE  BIOGRAPHY OF THE EARTH  [C/RAN-P]
GAMOW, GEORGE  BIRTH AND DEATH OF THE SUN, THE  [C/RAN-P]
GAMOW, GEORGE  CREATION OF THE UNIVERSE, THE  [C/RAN-P]
GAMOW, GEORGE  MR. TOMPKINS IN PAPERBACK  LP
GAMOW, GEORGE  ONE, TWO, THREE...INFINITY  [C/RAN-P:2]
GANN, ERNEST K.  BRAIN 2000  H
GANN, W. D.  TUNNEL THRU THE AIR, THE  H
GANNETT, LEWIS  LIVING ONE, THE  SFWA
GANNETT, LEWIS  MAGAZINE BEACH  H,LP
GANPAT  HARILEK  H
GANPAT  MIRROR OF DREAMS  H
GANPAT  VOICE OF DASHIN, THE  H
GANTZ, KENNETH F.  NOT IN SOLITUDE  H:2,P
GARBER, ERIC/GOMEZ, JEWELLE  SWORDS OF THE RAINBOW  LPA
GARCIA, ERIC  ANONYMOUS REX  LP
GARCIA, ROBERT T.  CHILLED TO THE BONE  SR-LPA
GARCIA Y ROBERTSON, R.  ATLANTIS FOUND  P
GARCIA Y ROBERTSON, R.  SPIRAL DANCE, THE  H,P
GARCIA Y ROBERTSON, R.  VIRGIN AND THE DINOSAUR, THE  P
GARDEN, NANCY  FOURS CROSSING  P
GARDINE, MICHAEL  LAMIA  P
GARDNER, CRAIG SHAW  BACK TO THE FUTURE PART 2 [BACK TO THE FUTURE 2]  P
GARDNER, CRAIG SHAW  BACK TO THE FUTURE PART 3 [BACK TO THE FUTURE 3]  P
GARDNER, CRAIG SHAW  BAD DAY FOR ALI BABA, A [OTHER SINBAD 2]  C/P,SFWA-P:2,P
GARDNER, CRAIG SHAW  BATMAN [BATMAN 1]  C/P:2,P
GARDNER, CRAIG SHAW  BATMAN MURDERS, THE [BATMAN 2]  P
GARDNER, CRAIG SHAW  BATMAN RETURNS  C/P,P
GARDNER, CRAIG SHAW  BRIDE OF THE SLIME MONSTER [CINEVERSE 2]  C/P:2,SFWA-P:2,P
GARDNER, CRAIG SHAW  CINEVERSE CYCLE, THE [CINEVERSE 1,2,3]  H
GARDNER, CRAIG SHAW  CYLONS’ SECRET, THE [@BATTLESTAR GALACTICA 2003]  LP
GARDNER, CRAIG SHAW  CYLON’S SECRET, THE (PROOF) [@BATTLESTAR GALACTICA 2003]  SR-LP
GARDNER, CRAIG SHAW  DARK MIRROR [@ANGEL]  C/P,P
GARDNER, CRAIG SHAW  DIFFICULTY WITH DWARVES, A [WUNTVOR 1]  C/P:2,P
GARDNER, CRAIG SHAW  DISAGREEMENT WITH DEATH, A [WUNTVOR 3]  C/P:2,SFWA-P,P
GARDNER, CRAIG SHAW  DRAGON BURNING [DRAGON CIRCLE 3]  H,@P
GARDNER, CRAIG SHAW  DRAGON SLEEPING [DRAGON CIRCLE 1]  H:2,SFWA,P
GARDNER, CRAIG SHAW  DRAGON WAKING [DRAGON CIRCLE 2]  C/P,P,H
GARDNER, CRAIG SHAW  EXCESS OF ENCHANTMENTS, AN [WUNTVOR 2]  C/P:2,P
GARDNER, CRAIG SHAW  EXPLOITS OF EBENEZUM, THE [EBENEZUM 1,2,3]  H
GARDNER, CRAIG SHAW  LAST ARABIAN NIGHT, THE [OTHER SINBAD 3]  C/P,SFWA-P,P
GARDNER, CRAIG SHAW  LOST BOYS, THE  P
GARDNER, CRAIG SHAW  MALADY OF MAGICKS, A [EBENEZUM 1]  C/P:2,P
GARDNER, CRAIG SHAW  MULTITUDE OF MONSTERS, A [EBENEZUM 2]  C/P:2,P
GARDNER, CRAIG SHAW  NIGHT IN THE NETHERHELLS, A [EBENEZUM 3]  C/P:2,SFWA-P,P
GARDNER, CRAIG SHAW  OTHER SINBAD, THE [OTHER SINBAD 1]  C/P,SFWA-P:2,P
GARDNER, CRAIG SHAW  RETURN TO CHAOS [@BUFFY]  C/P:2,P
GARDNER, CRAIG SHAW  REVENGE OF THE FLUFFY BUNNIES [CINEVERSE 3]  C/P,SFWA-P:2,P
GARDNER, CRAIG SHAW  SLAVES OF THE VOLCANO GOD [CINEVERSE 1]  C/P:2,SFWA-P,P
GARDNER, CRAIG SHAW  WANDERINGS OF WUNTVOR, THE [WUNTVOR 1,2,3]  H:2
GARDNER, CRAIG SHAW  WANTED: DEAD OR ALIVE [@SPIDERMAN]  H
GARDNER, CRAIG SHAW  WISHBRINGER  C/P:2,P
GARDNER, ERLE STANLEY  HUMAN ZERO, THE  H
GARDNER, GERALD/CARUSO, DEE  WORLD’S GREATEST ATHLETE, THE  P
GARDNER, JAMES ALAN  ASCENDING [LEAGUE OF PEOPLES 4]  H,P
GARDNER, JAMES ALAN  COMMITMENT HOUR  H,P
GARDNER, JAMES ALAN  EXPENDABLE [LEAGUE OF PEOPLES 1]  H:3,P

AUTHOR  TITLE  CODE
GARDNER, JAMES ALAN  GRAVITY WELLS  LP
GARDNER, JAMES ALAN  HUNTED [LEAGUE OF PEOPLES 3]  H:2,P
GARDNER, JAMES ALAN  MAN OF BRONZE, THE [@LARA CROFT TOMB RAIDER]  P
GARDNER, JAMES ALAN  RADIANT [LEAGUE OF PEOPLES 6]  H
GARDNER, JAMES ALAN  TRAPPED [LEAGUE OF PEOPLES 5]  H,P
GARDNER, JAMES ALAN  VIGILANT [LEAGUE OF PEOPLES 2]  C/P:2,P
GARDNER, JOHN (AMERICAN, 1933-1982)  FREDDY’S BOOK  P
GARDNER, JOHN (AMERICAN, 1933-1982)  GRENDEL  C/P:2,P
GARDNER, JOHN (BRITISH, 1926-)  BROKEN CLAW [JAMES BOND 23]  P
GARDNER, JOHN (BRITISH, 1926-)  FOR SPECIAL SERVICES [JAMES BOND 16]  C/P,P
GARDNER, JOHN (BRITISH, 1926-)  GOLGOTHA  P
GARDNER, JOHN (BRITISH, 1926-)  ICEBREAKER [JAMES BOND 17]  C/P,P
GARDNER, JOHN (BRITISH, 1926-)  LICENSE RENEWED [JAMES BOND 15]  P
GARDNER, JOHN (BRITISH, 1926-)  NO DEALS, MR. BOND [JAMES BOND 20]  C/P,P
GARDNER, JOHN (BRITISH, 1926-)  NOBODY LIVES FOREVER [JAMES BOND 19]  C/P,P
GARDNER, JOHN (BRITISH, 1926-)  ROLE OF HONOR [JAMES BOND 18]  C/P,P
GARDNER, JOHN (BRITISH, 1926-)  SCORPIUS [JAMES BOND 21]  P
GARDNER, JOHN (BRITISH, 1926-)  WIN, LOSE, OR DIE [JAMES BOND 22]  C/P,P
GARDNER, JOHN W.  ELEKTRICITEIT ZONDER DYNAMO  FP
GARDNER, MARSHALL B.  JOURNEY TO THE EARTH’S INTERIOR, A  H
GARDNER, MATT  CURSE OF QUINTANA ROO, THE  P
GARDNER, MAURICE B.  BANTAN — GOD-LIKE ISLANDER  H
GARDNER, MAURICE B.  BANTAN VALIANT  H
GARDNER, RICHARD  MANDRILL  P
GARFINKLE, RICHARD  ALL OF AN INSTANT  H,LP
GARFINKLE, RICHARD  CELESTIAL MATTERS  H,LP
GARFORTH, JOHN  FLOATING GAME, THE [@AVENGERS 1]  C/P:2,P
GARFORTH, JOHN  LAUGH WAS ON LAZARUS, THE [@AVENGERS 2]  P
GARFORTH, JOHN  PASSING OF GLORIA MUNDAY, THE [@AVENGERS 3]  C/P,P
GARIS, HOWARD R.  ROCKET RIDERS OVER THE DESERT  H
GARLAND, MARK A./MCGRAW, CHARLES G. DEMON BLADE [DEMON BLADE 1]  C/P,P
GARLAND, MARK A./MCCRAW, CHARLES G. DORELLA  C/P:2,SFWA-P,P
GARLAND, MARK A./MCGRAW, CHARLES G. FROST [DEMON BLADE 2]  P
GARLAND, MARK A./MCGRAW, CHARLES G. GHOST OF A CHANCE [STAR TREK VOYAGER]  C/P:2,P
GARLAND, MARK A.  SWORD OF THE PROPHETS  P
GARLAND, MARK A.  TRIAL BY ERROR [STAR TREK DS9]  C/P,P
GARN, JAKE/COHEN, STEPHEN PAUL  NIGHT LAUNCH  P
GARNER, ALAN  ELIDOR  C/P:2,P
GARNER, ALAN  LAD OF THE GAD, THE  H
GARNER, ALAN  MOON OF GOMRATH, THE [ALDERLY 2]  C/P:2,P
GARNER, ALAN  OWL SERVICE, THE  C/P,P,H
GARNER, ALAN  RED SHIFT  H:2,P
GARNER, ALAN  WEIRDSTONE OF BRISINGAMEN, THE [ALDERLY 1]  C/P:2,P
GARNER, ROLF  IMMORTALS, THE  SR-P
GARNER, ROLF  INDESTRUCTIBLE, THE  SR-P
GARNER, ROLF  RESURGENT DUST  SR-P
GARNETT, BILL  DOWN BOUND TRAIN  H,SFWA,P
GARNETT, DAV  MIRROR IN THE SKY  P
GARNETT, DAV  STARSEEKERS, THE  C/P:2,P
GARNETT, DAVID  BIKINI PLANET  P
GARNETT, DAVID  LADY INTO FOX  H
GARNETT, DAVID  NEW WORLDS 1  HA,SR-LPA
GARNETT, DAVID S.  TIME IN ECLIPSE  SR-P
GARRETT, EILEEN J.  ADVENTURES IN THE SUPERNORMAL  [C/RAN-P:2]
GARRETT, RANDALL  ANYTHING YOU CAN DO...  P
GARRETT, RANDALL  BEST OF RANDALL GARRETT, THE  C/P:2,P
GARRETT, RANDALL/JANIFER, LAURENCE M. BRAIN TWISTER  P
GARRETT, RANDALL/HEYDRON, VICKI ANN  BRONZE OF EDDARTA, THE [GANDALARA 3]  C/P,P
GARRETT, RANDALL  EARTH INVADER  C/P:2,P
GARRETT, RANDALL  GALLERY OF THE SAINTS, A  [C/RAN-P]
GARRETT, RANDALL/HEYDRON, VICKI ANN  GANDALARA CYCLE I, THE [GANDALARA 1,2,3]  C/P:2,P
GARRETT, RANDALL/HEYDRON, VICKI ANN  GANDALARA CYCLE II, THE [GANDALARA 4,5,6]  C/P:2,P
GARRETT, RANDALL/HEYDRON, VICKI ANN  GLASS OF DYSKORNIS, THE [GANDALARA 2]  C/P:2,P
GARRETT, RANDALL  LORD DARCY [LORD DARCY 1,2,3]  H
GARRETT, RANDALL  LORD DARCY (BAEN) [LORD DARCY 1,2,3,4]  LP
GARRETT, RANDALL  LORD DARCY INVESTIGATES [LORD DARCY 3]  C/P:2,P
GARRETT, RANDALL  MURDER AND MAGIC [LORD DARCY 1]  C/P:2,P
GARRETT, RANDALL/HARRIS, LARRY M.  PAGAN PASSIONS  SR-P
GARRETT, RANDALL/HEYDRON, VICKI ANN  RETURN TO EDDARTA [GANDALARA 6]  C/P,P
GARRETT, RANDALL/HEYDRON, VICKI ANN  RIVER WALL, THE [GANDALARA 7]  C/P:2,P
GARRETT, RANDALL/HEYDRON, VICKI ANN  SEARCH FOR KA, THE [GANDALARA 5]  C/P,P
GARRETT, RANDALL  STARSHIP DEATH  C/P,P
GARRETT, RANDALL/HEYDRON, VICKI ANN  STEEL OF RAITHSKAR, THE [GANDALARA 1]  C/P:2,P
GARRETT, RANDALL  TAKEOFF!  LP

The MITSFS Pinkdex  September 24, 2007  Page 156
AUTHOR  TITLE  CODE
GARRETT, RANDALL  TAKEOFF TOO!  LP
GARRETT, RANDALL  TOO MANY MAGICIANS [LORD DARCY 2]  SR-P,C/P:2
GARRETT, RANDALL  UNWISE CHILD  H:2
GARRETT, RANDALL/HEYDRON, VICKI ANN  WELL OF DARKNESS, THE [GANDALARA 4]  C/P:2,P
GARRETT, SUSAN M.  INTIMATIONS OF MORTALITY [FOREVER KNIGHT]  P
GARRISON, PETER  CHANGELING WAR, THE [@CHANGELING #1]  C/P,LP
GARRISON, PETER  MAGIC DEAD, THE [@CHANGELING #3]  P
GARRISON, PETER  SORCERER’S GUN, THE [@CHANGELING #2]  C/P,LP
GARSON, PAUL  GREAT QUILL, THE  H:2,SFWA
GARTH, G. G.  DRIVEN TO KILL!  P
GARTH, WILL  DR. CYCLOPS  C/P,P
GARTON, RAY  ALL THAT GLITTERS [@SABRINA THE TEENAGE WITCH #12]  P
GARTON, RAY  DARK CHANNEL  H
GARTON, RAY  INVADERS FROM MARS  P
GARTON, RAY  LIVE GIRLS  SR-L
GARTON, RAY  RESURRECTING RAVANA [BUFFY]  C/P,P
GARTON, RAY  SEDUCTIONS  SR-L
GARTON, RAY  WARLOCK  P
GARVIN, RICHARD M./ADDEO, EDMOND G.  FORTEC CONSPIRACY, THE  C/P,P
GARVIN, RICHARD M./ADDEO, EDMOND G.  TALBOTT AGREEMENT, THE  H,P
GARY, ROMAIN  GASP, THE  C/P,P
GASCOIGNE, MARC/DUNN, CHRISTIAN  BRINGERS OF DEATH [@WARHAMMER 40K]  PA

GASCOIGNE, MARC/JONES, ANDY/ DUNN, CHRISTIAN CALL OF CHAOS, THE [WARHAMMER]

CX

GASCOIGNE, MARC/DUNN, CHRISTIAN  COLD HAND OF BETRAYAL, THE [@WARHAMMER]  P
GASCOIGNE, MARC/DUNN, CHRISTIAN  CRUCIBLE OF WAR [@WARHAMMER 40K]  PA
GASCOIGNE, MARC/JONES, ANDY  DARK IMPERIUM [WARHAMMER 40K]  PA
GASCOIGNE, MARC/JONES, ANDY  INTO THE MAELSTROM [WARHAMMER 40K]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  INVASION! [@WARHAMMER]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  LET THE GALAXY BURN [@WARHAMMER 40K]  LPA
GASCOIGNE, MARC/DUNN, CHRISTIAN  LORDS OF VALOUR [WARHAMMER]  PA
GASCOIGNE, MARC/JONES, ANDY  REALM OF CHAOS [WARHAMMER]  PA
GASCOIGNE, MARC/JONES, ANDY  STATUS: DEADZONE [WARHAMMER 40K]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  SWORDS OF THE EMPIRE [@WARHAMMER]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  TALES FROM THE DARK MILLENNIUM [@WARHAMMER 40K]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  TALES FROM THE TEN-TAILED CAT [WARHAMMER]  CX
GASCOIGNE, MARC/DUNN, CHRISTIAN  TALES OF THE OLD WORLD [@WARHAMMER]  LPA
GASCOIGNE, MARC/DUNN, CHRISTIAN  WAY OF THE DEAD [@WARHAMMER]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  WHAT PRICE VICTORY [WARHAMMER 40K]  PA
GASCOIGNE, MARC/DUNN, CHRISTIAN  WORDS OF BLOOD [WARHAMMER 40K]  PA
GASKELL, JANE  ATLAN [ATLAN]  C/P,P,H
GASKELL, JANE  CITY, THE [ATLAN]  C/P,P,H
GASKELL, JANE  DRAGON, THE [ATLAN]  C/P:2,P
GASKELL, JANE  KING’S DAUGHTER [ATLAN]  C/P,P
GASKELL, JANE  SERPENT, THE [ATLAN]  C/P:2,P
GASKELL, JANE  SOME SUMMER LANDS [ATLAN]  C/P,P,H
GASKELL, JANE  STRANGE EVIL  P
GASKELL, JANE  SWEET, SWEET SUMMER, A  H
GASPERINI, JIM  SAIL WITH PIRATES [@TIME MACHINE #4]  P
GASPERINI, JIM  SECRET OF THE KNIGHTS [@TIME MACHINE #1]  C/P,P
GAT, DIMITRI V.  SHEPHERD IS MY LORD, THE  H,SFWA
GATISS, MARK  LAST OF THE GADERENE [DOCTOR WHO PAST DOCTORS]  C/P,P
GATISS, MARK  NIGHTSHADE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
GATISS, MARK  ROUNDHEADS, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
GATISS, MARK  ST ANTHONY’S FIRE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
GATLAND, KENNETH W.  MANNED SPACECRAFT  [C/RAN-H]
GATLAND, KENNETH W.  PROJECT SATELLITE  [C/RAN-H]
GAUGER, RICK  CHARON’S ARK  C/P,P
GAWRON, JEAN MARK  ALGORITHM  P
GAWRON, JEAN MARK  APOLOGY FOR RAIN, AN  H,SFWA
GAWRON, JEAN MARK  DREAM OF GLASS  H:2
GAYLE, HENRY K.  SPAWN OF THE VORTEX  H
GEAR, W. MICHAEL  ARTIFACT, THE  C/P,P
GEAR, W. MICHAEL  COUNTERMEASURES [@FORBIDDEN BORDERS #3]  C/P:2,P

GEAR, W. MICHAEL/ GEAR, KATHLEEN O’NEAL PEOPLE OF THE EARTH [PEOPLE OF THE WOLF 1]

C/P,P

GEAR, W. MICHAEL/ GEAR, KATHLEEN O’NEAL PEOPLE OF THE FIRE [PEOPLE OF THE WOLF 2]

SFWA-P

GEAR, W. MICHAEL/ GEAR, KATHLEEN O’NEAL PEOPLE OF THE SEA

H

GEAR, W. MICHAEL  RELIC OF EMPIRE [@FORBIDDEN BORDERS #2]  C/P:2,SFWA-P,P
GEAR, W. MICHAEL  REQUIEM FOR THE CONQUEROR [@FORBIDDEN BORDERS #1]  C/P:2,SFWA-P:2,P
GEAR, W. MICHAEL  STARSTRIKE  C/P:2,P
GEAR, W. MICHAEL  WARRIORS OF SPIDER, THE [WARRIORS OF SPIDER 1]  C/P:2,P

AUTHOR  TITLE  CODE
GEAR, W. MICHAEL  WAY OF SPIDER, THE [WARRIORS OF SPIDER 2]  C/P:2,P
GEAR, W. MICHAEL  WEB OF SPIDER, THE [WARRIORS OF SPIDER 3]  C/P:2,P
GEARHARDT, SALLY MILLER  WANDERGROUND, THE  H,P
GEARY, PATRICIA  LIVING IN ETHER  P
GEARY, PATRICIA  STRANGE TOYS  C/P,P
GEDGE, PAULINE  STARGATE  C/P,P,H
GEE, EMILY  THIEF WITH NO SHADOW  P
GEE, MAGGIE  ICE PEOPLE, THE  LP
GEE, MAURICE  PRIESTS OF FERRIS, THE  SR-P
GELB, JEFF/GARRETT, MICHAEL  STRANGER BY NIGHT [HOT BLOOD]  PA
GELLIS, ROBERTA  BULL GOD  P
GELLIS, ROBERTA  MORTAL BANE, A  P
GELLIS, ROBERTA  THRICE BOUND  P
GEMMELL, DAVID  AGAINST THE HORDE [DRENAI 1 (ORIGINAL)]  P
GEMMELL, DAVID  BLOODSTONE [STONES OF POWER 5]  C/P,P
GEMMELL, DAVID  DARK MOON  P
GEMMELL, DAVID  DARK PRINCE [MACEDON 2]  LP
GEMMELL, DAVID  ECHOES OF THE GREAT SONG  P
GEMMELL, DAVID  FIRST CHRONICLES OF DRUSS THE LEGEND, THE [DRENAI]  C/P,P
GEMMELL, DAVID  GHOST KING [STONES OF POWER 1]  C/P,P
GEMMELL, DAVID  HAWK ETERNAL, THE [HAWK QUEEN 2]  P
GEMMELL, DAVID  HERO IN THE SHADOWS [DRENAI]  P
GEMMELL, DAVID  IN THE REALM OF THE WOLF [DRENAI WAYLANDER 2]  P
GEMMELL, DAVID  IRONHAND’S DAUGHTER [HAWK QUEEN]  P
GEMMELL, DAVID  JERUSALEM MAN, THE [STONES OF POWER 3]  P,SFWA-P
GEMMELL, DAVID  KING BEYOND THE GATE, THE [DRENAI 2]  P
GEMMELL, DAVID  KNIGHTS OF DARK RENOWN  P
GEMMELL, DAVID  LAST GUARDIAN, THE [@STONES OF POWER #4]  P
GEMMELL, DAVID  LAST SWORD OF POWER [@STONES OF POWER #2]  P
GEMMELL, DAVID  LEGEND [DRENAI 1 (REVISED)]  C/P:2,P
GEMMELL, DAVID  LEGEND OF DEATHWALKER  P
GEMMELL, DAVID  LION OF MACEDON [MACEDON 1]  H,LP
GEMMELL, DAVID  LORD OF THE SILVER BOW [@TROY 1]  LP
GEMMELL, DAVID  MIDNIGHT FALCON [RIGANTE 2]  P
GEMMELL, DAVID  MORNINGSTAR  P
GEMMELL, DAVID  QUEST FOR LOST HEROES [DRENAI 3]  P
GEMMELL, DAVID  RAVENHEART [RIGANTE 3]  H,P
GEMMELL, DAVID  SHIELD OF THUNDER [@TROY 2]  H
GEMMELL, DAVID  STORMRIDER [RIGANTE 4]  P
GEMMELL, DAVID  SWORD IN THE STORM [RIGANTE 1]  P
GEMMELL, DAVID  SWORDS OF NIGHT AND DAY, THE [SKILGANNON THE DAMNED]  H
GEMMELL, DAVID  TALES OF THE RIGANTE [RIGANTE 1,2]  H
GEMMELL, DAVID  WAYLANDER [DRENAI WAYLANDER 1]  C/P,P
GEMMELL, DAVID  WHITE WOLF [DRENAI]  P
GEMMELL, DAVID  WINTER WARRIORS [DRENAI]  P
GEMMELL, DAVID  WOLF IN SHADOW [@STONES OF POWER #3]  P
GEMS, JONATHAN  MARS ATTACKS! [@MARS ATTACKS]  P
GENOA, CHRIS  FOOP!  LP
GENTILE, GARY  DRAGONS PAST [DRAGON WARS 2]  C/P,P
GENTILE, GARY  NO FUTURE FOR DRAGONS [DRAGON WARS 3]  P
GENTILE, GARY  TIME FOR DRAGONS, A [DRAGON WARS 1]  C/P:2,P
GENTLE, MARY  ANCIENT LIGHT [ORTHE 2]  C/P,P,H
GENTLE, MARY  ARCHITECTURE OF DESIRE, THE  P,SFWA
GENTLE, MARY  CARTHAGE ASCENDANT [@ASH #2]  C/P,P
GENTLE, MARY  GOLDEN WITCHBREED [ORTHE 1]  C/P,P,H
GENTLE, MARY  GRUNTS!  C/P:2,P
GENTLE, MARY  HAWK IN SILVER, A  C/P,P,H
GENTLE, MARY  LION’S EYE, THE [@ILARIO 1]  LP
GENTLE, MARY  LOST BURGUNDY [@ASH #4]  C/P,P
GENTLE, MARY  RATS AND GARGOYLES  C/P,SFWA,P
GENTLE, MARY  SCHOLARS AND SOLDIERS  P
GENTLE, MARY  SECRET HISTORY, A [@ASH #1]  C/P:2,P
GENTLE, MARY  WILD MACHINES, THE [@ASH #3]  C/P,P
GENTRY, CURT  LAST DAYS OF THE LATE, GREAT STATE OF CALIFORNIA, THE  C/P,P
GEORGE, DAVID R. III  FIRE AND THE ROSE, THE [STAR TREK CRUCIBLE]  P
GEORGE, DAVID R. III  PROVENANCE OF SHADOWS [STAR TREK CRUCIBLE]  P
GEORGE, DAVID R. III  SERPENTS AMONG THE RUINS [@STAR TREK LOST ERA (2311 AD)]  P
GEORGE, DAVID R. III  STAR TO EVERY WANDERING, THE [STAR TREK CRUCIBLE]  P
GEORGE, DAVID R. III  TWILIGHT [STAR TREK DS9 MISSION GAMMA #1]  C/P,P
GEORGE, JEAN CRAIGHEAD  HOLD ZERO!  H
GEORGE, PETER  COMMANDER-1  C/P:2,P

GEORGE, PETER DR. STRANGELOVE OR:

HOW I LEARNED TO STOP WORRYING AND LOVE THE BOMB

C/P,P

GEORGE, SARA  FATAL SHADOWS  P

The MITSFS Pinkdex  September 24, 2007  Page 158
AUTHOR  TITLE  CODE
GERANI, GARY/SCHULMAN, PAUL H.  FANTASTIC TELEVISION  R/REF-H
GERBER, MICHAEL  BARRY TROTTER AND THE SHAMELESS PARODY  H
GERNSBACK, HUGO  FORECAST 1955  B
GERNSBACK, HUGO  FORECAST 1956  B
GERNSBACK, HUGO  FORECAST 1957  B
GERNSBACK, HUGO  FORECAST 1958  B
GERNSBACK, HUGO  FORECAST 1959  B
GERNSBACK, HUGO  FORECAST 1960  B
GERNSBACK, HUGO  FORECAST 1961  B
GERNSBACK, HUGO  FORECAST 1963  B
GERNSBACK, HUGO  FORECAST 1964  B
GERNSBACK, HUGO  RALPH 124C41+  C/P:2,P,H
GERNSBACK, HUGO  ULTIMATE WORLD  H,SFWA,LP
GERRAND, ROB  BEST AUSTRALIAN SCIENCE FICTION WRITING, THE  LPA
GERROLD, DAVID  ALTERNATE GERROLDS  LP
GERROLD, DAVID  ALTERNITIES  C/PA,PA
GERROLD, DAVID  ASCENTS OF WONDER  C/PA,PA
GERROLD, DAVID  BATTLE FOR THE PLANET OF THE APES  C/P:2,P
GERROLD, DAVID  BLOOD AND FIRE  LP
GERROLD, DAVID  BOUNCING OFF THE MOON [DINGILLIAD 2]  H
GERROLD, DAVID  CHESS WITH A DRAGON  C/P,P,H
GERROLD, DAVID  CHILD OF EARTH [SEA OF GRASS 1]  LP
GERROLD, DAVID  COVENANT OF JUSTICE, A [PHAESTOR 2]  C/P,P
GERROLD, DAVID  DAY FOR DAMNATION, A [WAR AGAINST THE CHTORR 2]  C/P:2,P

GERROLD, DAVID DAY FOR DAMNATION, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 2]

C/P,LP

GERROLD, DAVID  DEATHBEAST  C/P:2,P
GERROLD, DAVID  ENCOUNTER AT FARPOINT [STAR TREK TNG]  C/P:2,P
GERROLD, DAVID/NIVEN, LARRY  FLYING SORCERERS, THE  C/P:2,P
GERROLD, DAVID  GALACTIC WHIRLPOOL, THE [STAR TREK]  C/P:3,P
GERROLD, DAVID  GENERATION  PA,C/PA
GERROLD, DAVID  JUMPING OFF THE PLANET [DINGILLIAD 1]  H,P
GERROLD, DAVID  LEAPING TO THE STARS [DINGILLIAD 3]  H
GERROLD, DAVID  MAN WHO FOLDED HIMSELF, THE  H:2,P,SFWA
GERROLD, DAVID  MAN WHO FOLDED HIMSELF, THE (REVISED)  LP
GERROLD, DAVID  MARTIAN CHILD, THE  LP
GERROLD, DAVID  MATTER FOR MEN, A [WAR AGAINST THE CHTORR 1]  C/P,P

GERROLD, DAVID MATTER FOR MEN, A (ABRIDGED)

[WAR AGAINST THE CHTORR (ABRIDGED) 1]

C/P,LP

GERROLD, DAVID  MIDDLE OF NOWHERE, THE [STAR WOLF 2]  C/P,P
GERROLD, DAVID  MOONSTAR ODYSSEY  C/P:2,P
GERROLD, DAVID  PROTOSTARS  PA,C/PA
GERROLD, DAVID  RAGE FOR REVENGE, A [WAR AGAINST THE CHTORR 3]  C/P:2,P
GERROLD, DAVID  SCIENCE FICTION EMPHASIS #1  PA,C/PA
GERROLD, DAVID  SEASON FOR SLAUGHTER, A [WAR AGAINST THE CHTORR 4]  C/P,P
GERROLD, DAVID  SPACE SKIMMER  C/P:2,P
GERROLD, DAVID  STARHUNT [STAR WOLF 0]  P
GERROLD, DAVID  TALES OF THE STAR WOLF [STAR WOLF 1,2,3]  H
GERROLD, DAVID  TROUBLE WITH TRIBBLES, THE [STAR TREK]  R/REF-P,C/REF-P:2
GERROLD, DAVID  UNDER THE EYE OF GOD [PHAESTOR 1]  C/P:2,P
GERROLD, DAVID  VOYAGE OF THE STAR WOLF [STAR WOLF 1]  C/P,P,H

GERROLD, DAVID WAR AGAINST THE CHTORR: INVASION, THE

[WAR AGAINST THE CHTORR (ABRIDGED) 1,2]

H:2

GERROLD, DAVID  WHEN HARLIE WAS ONE  C/P,P,H
GERROLD, DAVID  WHEN HARLIE WAS ONE, RELEASE 2.0  C/P,P
GERROLD, DAVID  WITH A FINGER IN MY I  C/P:2,P
GERROLD, DAVID  WORLD OF STAR TREK, THE  R/REF-P,C/REF-P:2
GERROLD, DAVID  YESTERDAY’S CHILDREN [STAR WOLF 0]  C/P:2,P
GERROLD, DAVID  YESTERDAY’S CHILDREN (REVISED, PROOF) [STAR WOLF 0]  SR-LP
GERROLD, DAVID  YESTERDAY’S CHILDREN (REVISED) [STAR WOLF 0]  P
GESTON, MARK S.  DAY STAR, THE  C/P:2,P
GESTON, MARK S.  LORDS OF THE STARSHIP  C/P:2,P
GESTON, MARK S.  MIRROR TO THE SKY  H:2,SFWA-P
GESTON, MARK S.  OUT OF THE MOUTH OF THE DRAGON  C/P:2,P
GESTON, MARK S.  SIEGE OF WONDER, THE  H,SFWA,P
GHIDALIA, VIC/ELWOOD, ROGER  BEWARE MORE BEASTS  PA
GHIDALIA, VIC/ELWOOD, ROGER  BEWARE THE BEASTS  PA
GHIDALIA, VIC  DEVIL’S GENERATION, THE  C/PA,PA
GHIDALIA, VIC  EIGHT STRANGE TALES  PA
GHIDALIA, VIC  FEAST OF FEAR  PA
GHIDALIA, VIC  GOOSEFLESH!  PA
GHIDALIA, VIC  MUMMY WALKS AMONG US, THE  PA
GHIDALIA, VIC  NIGHTMARE GARDEN  PA
GHIDALIA, VIC  ODDBALLS, THE  PA

The MITSFS Pinkdex  September 24, 2007  Page 159
AUTHOR  TITLE  CODE
GHIDALIA, VIC  SATAN’S PETS  PA
GHIDALIA, VIC/ELWOOD, ROGER  VENUS FACTOR, THE  C/PA,PA
GHIDALIA, VIC  WIZARDS AND WARLOCKS  C/PA,PA
GHOLSTON, HOMER N.  KOIEC COROLLARY, THE  P
GIAMBASTIANI, KURT R. A.  FROM THE HEART OF THE STORM [FALLEN CLOUD 4]  P
GIAMBASTIANI, KURT R. A.  SHADOW OF THE STORM [FALLEN CLOUD 3]  P
GIAMBASTIANI, KURT R. A.  SPIRIT OF THUNDER, THE [FALLEN CLOUD 2]  H,P
GIAMBASTIANI, KURT R. A.  YEAR THE CLOUD FELL, THE [FALLEN CLOUD 1]  P

GIARRATANO, JOSEPH C./ ANDREWS, ARLAN KEITH/ ANDREWS, KRIS AUSTEN TIMELOST

SCX

GIBBONS, FLOYD  RED NAPOLEON, THE  H,P
GIBBONS, GAVIN  ON BOARD THE FLYING SAUCERS  [C/RAN-P]
GIBSON, EDWARD  IN THE WRONG HANDS  P
GIBSON, EDWARD  REACH  H,P
GIBSON, WALTER B.  COIN OF EL DIABLO AFFAIR, THE [MAN FROM UNCLE]  VLP
GIBSON, WALTER B.  RETURN OF THE SHADOW  P
GIBSON, WALTER B.  ROD SERLING’S THE TWILIGHT ZONE  H,P
GIBSON, WALTER B.  ROD SERLING’S TWILIGHT ZONE REVISITED  H,P
GIBSON, WILLIAM  ALL TOMORROW’S PARTIES  H,LP
GIBSON, WILLIAM  BURNING CHROME  C/P,SFWA-P,P,H,SFWA
GIBSON, WILLIAM  COUNT ZERO [NEUROMANCER 2]  SFWA-P,H:2,P,SFWA
GIBSON, WILLIAM/STERLING, BRUCE  DIFFERENCE ENGINE, THE  H:2,P
GIBSON, WILLIAM  IDORU  H:2,P
GIBSON, WILLIAM  MONA LISA OVERDRIVE [NEUROMANCER 3]  C/P,SFWA-P,H:2,P
GIBSON, WILLIAM  NEUROMANCER [NEUROMANCER 1]  H:2,P
GIBSON, WILLIAM  PATTERN RECOGNITION  C/P,LP,H
GIBSON, WILLIAM  VIRTUAL LIGHT  H:2,P
GIBSON, WILLIAM  VIRTUAL LIGHT (EXCERPT)  SR-LP,LP
GIER, SCOTT G.  GENELLAN: FIRST VICTORY [@GENELLAN #3]  P
GIER, SCOTT G.  GENELLAN: IN THE SHADOW OF THE MOON [@GENELLAN #2]  P
GIER, SCOTT G.  GENELLAN: PLANETFALL [@GENELLAN #1]  P
GIFFORD, DENIS  PICTORIAL HISTORY OF HORROR MOVIES, A  R/REF-H
GIFFORD, DENIS  SCIENCE FICTION FILM  R/REF-P
GILBERT, C. M.  OZINE CONQUEST, THE  P
GILBERT, STEPHEN  LANDSLIDE, THE  H
GILBERT, STEPHEN  WILLARD  P
GILDEN, MEL  BOOGEYMEN [STAR TREK TNG]  C/P:2,P
GILDEN, MEL  HAWAIIAN UFO ALIENS  C/P,P
GILDEN, MEL  PUMPKINS OF TIME, THE  P
GILDEN, MEL  RETURN OF CAPTAIN CONQUER, THE  SFWA
GILDEN, MEL  STARSHIP TRAP, THE [STAR TREK]  C/P:2,P
GILDEN, MEL  SURFING SAMURAI ROBOTS  C/P:2,SFWA-P:2,P
GILDEN, MEL  TUBULAR ANDROID SUPERHEROES  C/P,P
GILES, RAYMOND  NIGHT OF THE VAMPIRE  P
GILFORD, C. B.  LIQUID MAN, THE  P
GILHOOLEY, LORD  MAUGIS, YE SORCERER  SR-H
GILLER, MARC D.  HAMMERJACK  LP
GILLER, MARK D.  PRODIGAL  LP
GILLESPIE, BRUCE  PHILIP K. DICK: ELECTRIC SHEPHERD  R/REF-P
GILLESPIE, MARC D.  HAMMERJACK  LP
GILLIAM, RICHARD/GREENBERG, MARTIN H. BOOK OF KINGS, THE  C/PA,PA

GILLIAM, RICHARD/GREENBERG, MARTIN H./ KRAMER, EDWARD E.

CONFEDERACY OF THE DEAD

PA

GILLIAM, RICHARD/GREENBERG, MARTIN H./ KRAMER, EDWARD E.

EXCALIBUR

HA,LPA

GILLIAM, RICHARD/GREENBERG, MARTIN H./ KRAMER, EDWARD E.

GRAILS: QUESTS OF THE DAWN

HA,LPA,SFWA

GILLIAM, RICHARD/GREENBERG, MARTIN H./ KRAMER, EDWARD E.

GRAILS: VISITATIONS OF THE NIGHT

LPA

GILLIGAN, ELIZABETH  MAGIC’S SILKEN SNARE [SILKEN MAGIC 1]  P
GILLIGAN, ELIZABETH  SILKEN SHROUD, THE [SILKEN MAGIC 2]  P
GILLILAND, ALEXIS A.  END OF THE EMPIRE, THE  C/P,P
GILLILAND, ALEXIS A.  LONG SHOT FOR ROSINANTE [ROSINANTE 2]  C/P:2,P
GILLILAND, ALEXIS A.  LORD OF THE TROLL-BATS  P
GILLILAND, ALEXIS A.  PIRATES OF ROSINANTE, THE [ROSINANTE 3]  C/P:2,P
GILLILAND, ALEXIS A.  REVOLUTION FROM ROSINANTE, THE [ROSINANTE 1]  P
GILLILAND, ALEXIS A.  SHADOW SHAIA, THE [WIZENBEAK 2]  P
GILLILAND, ALEXIS A.  WIZENBEAK [WIZENBEAK 1]  C/P:2,LP
GILLING, TOM  SOOTERKIN, THE  LP
GILLMORE, INEZ HAYNES  ANGEL ISLAND  LP
GILLON, DIANA/GILLON, MEIR  UNSLEEP, THE  C/P:2,P
GILLULY, SHEILA  BOY FROM THE BURREN, THE [PAINTER 1]  P
GILLULY, SHEILA  GREENBRIAR QUEEN  P

The MITSFS Pinkdex  September 24, 2007  Page 160
AUTHOR  TITLE  CODE
GILLULY, SHEILA  RITNYM’S DAUGHTER  C/P:2,P
GILMA, CAROLYN IVES  HALFWAY HUMAN  P
GILMAN, CAROLYN IVES  CANDLE IN A BOTTLE  LP
GILMAN, CHARLOTTE P.  HERLAND  H
GILMAN, DOROTHY  CLAIRVOYANT COUNTESS, THE  P
GILMAN, DOROTHY  MAZE IN THE HEART OF THE CASTLE, THE  P
GILMAN, GREER ILENE  MOONWISE  C/P,P
GILMAN, LAURA ANNE  BRING IT ON [RETRIEVERS]  LP
GILMAN, LAURA ANNE  BURNING BRIDGES [RETRIEVERS]  LP
GILMAN, LAURA ANNE  CURSE THE DARK [RETRIEVERS]  H,LP
GILMAN, LAURA ANNE/SHERMAN, JOSEPHA DEEP WATER [BUFFY]  P

GILMAN, LAURA ANNE/ DECANDIDO, KEITH R. A.

OTHERWERE

PA

GILMAN, LAURA ANNE  STAYING DEAD [RETRIEVERS]  H,LP
GILMAN, LAURA ANNE/HEDDLE, JENNIFER  TREACHERY AND TREASON  PA
GILMAN, LAURA ANNE/SHERMAN, JOSEPHA VISITORS [BUFFY]  P
GILMAN, ROBERT CHAM  NAVIGATOR OF RHADA, THE [RHADA 3]  C/P,P,H
GILMAN, ROBERT CHAM  REBEL OF RHADA, THE [RHADA 1]  H:2,P
GILMAN, ROBERT CHAM  STARKAHN OF RHADA, THE [RHADA 4]  C/P,P,H
GILMAN, ROBERT CHAM  WARLOCK OF RHADA, THE [RHADA 2]  C/P,P
GILMAR, XANDRA  ERROR DE LA MONTANA, EL  FP
GILMORE, ANTHONY  SPACE HAWK  H
GILMOUR, H. B.  EYES OF LAURA MARS  P
GILMOUR, WILLIAM  UNDYING LAND, THE  H
GILSON, CHARLES  CAT AND THE CURATE  H
GINGRICH, NEWT/FORSTCHEN, WILLIAM R.  GETTYSBURG [GETTYSBURG 1]  H
GINGRICH, NEWT/FORSTCHEN, WILLIAM R.  GRANT COMES EAST [GETTYSBURG 2]  H
GINGRICH, NEWT/FORSTCHEN, WILLIAM R.  NEVER CALL RETREAT [GETTYSBURG 3]  H
GINGRICH, NEWT/FORSTCHEN, WILLIAM R.  1945  H,SFWA,P

GINGRICH, NEWT/DRAKE, DAVID/ GINGRICH, MARRIANNE WINDOW OF OPPORTUNITY

SFWA-P

GINSBURG, MIRRA  AIR OF MARS, THE  HA
GINSBURG, MIRRA  LAST DOOR TO AIYA  HA
GINSBURG, MIRRA  ULTIMATE THRESHOLD, THE  C/PA,PA
GIPE, GEORGE  BACK TO THE FUTURE [BACK TO THE FUTURE 1]  C/P,P
GIPE, GEORGE  EXPLORERS  P
GIPE, GEORGE  GREMLINS [GREMLINS]  C/P,P
GIRARDEAU, EMILE  AVONTUUR DER NATUURWETENSCHAP, HET  FP
GLADNEY, HEATHER  BLOODSTORM [NAGA TEOT 2]  C/P:2,P
GLADNEY, HEATHER  TEOT’S WAR [NAGA TEOT 1]  P,SFWA-P
GLASBY, JOHN  PROJECT JOVE  D68310
GLASS, ISABEL  DAUGHTER OF EXILE  P
GLASS, ISABEL  DIVIDED CROWN, THE  H
GLASS, JAMES C.  EMPRESS OF LIGHT [SHANJI INTERSTELLAR EMPIRE 2]  C/P,P
GLASS, JAMES C.  SHANJI [SHANJI INTERSTELLAR EMPIRE 1]  P
GLEASON, ROBERT  WRATH OF GOD  H,LP

GLENN, JEROME CLAYTON/ ROBINSON, GEORGE S.

SPACE TREK

[C/RAN-P]

GLICK, RUTH  INVASION OF THE BLUE LIGHTS  P
GLOAG, JULIAN  OUR MOTHER’S HOUSE  P
GLOSS, MOLLY  DAZZLE OF DAY, THE  LP
GLOSS, MOLLY  OUTSIDE THE GATES  H
GLUT, DONALD F.  BUGGED!  P
GLUT, DONALD F.  EMPIRE STRIKES BACK, THE [STAR WARS MOVIE 5]  H:2,P
GLUT, DONALD F.  SPAWN  SR-P
GLYNN, A. A.  PLAN FOR CONQUEST  SR-P,H
GNAWING IDEAS  MYTHIC PERSPECTIVES FICTION 1998  LPA
GOBLE, LOU  KALEVIDE, THE  C/P:2,P
GOBLE, NEIL  ASIMOV ANALYZED  R/REF-H,C/REF-H
GOD  BHAGAVAD-GITA  PA
GOD  GREAT NEWS: THE NEW TESTAMENT, THE (NEW INTERNATIONAL)  PA
GOD  HOLY BIBLE, THE (GIDEON)  HA
GOD  NEW TESTAMENT, THE (KING JAMES)  PA
GOD  NEW TESTAMENT, THE (REVISED STANDARD)  HA
GODDARD, KEN  OUTER PERIMETER  P
GODEY, JOHN  MAN IN QUESTION, THE (COVER FROM LAUMER’S NIGHT OF DELUSIONS)  SR-L
GODFREY, MARTYN  LAST WAR, THE  P
GODWIN, JOHN  THIS BAFFLING WORLD 1  [C/RAN-P:2]
GODWIN, JOHN  THIS BAFFLING WORLD 3  [C/RAN-P]
GODWIN, PARKE  BELOVED EXILE  H,P
GODWIN, PARKE  FIRE WHEN IT COMES, THE  H,SFWA
GODWIN, PARKE  FIRELORD  C/P,P,H
GODWIN, PARKE  INVITATION TO CAMELOT  C/PA,PA
GODWIN, PARKE  LAST RAINBOW, THE  C/P:2,LP

The MITSFS Pinkdex  September 24, 2007  Page 161
AUTHOR  TITLE  CODE
GODWIN, PARKE  LIMBO SEARCH  C/P,P
GODWIN, PARKE  ROBIN AND THE KING [SHERWOOD 2]  C/P,P
GODWIN, PARKE  SHERWOOD [SHERWOOD 1]  P,SFWA-P
GODWIN, PARKE  SNAKE OIL WARS, THE [BARION AND COYUL 2]  H
GODWIN, PARKE  TOWER OF BEOWULF, THE  P
GODWIN, PARKE  TRUCE WITH TIME, A  C/P,SR-LP,H
GODWIN, PARKE  WAITING FOR THE GALACTIC BUS [BARION AND COYUL 1]  H:2,P
GODWIN, TOM  BEYOND ANOTHER SUN  P
GODWIN, TOM  COLD EQUATIONS, THE  P
GODWIN, TOM  SPACE BARBARIANS, THE  C/P:2,P
GODWIN, TOM  SPACE PRISON  C/P,P
GODWIN, TOM  SURVIVORS, THE  H
GOFF, GEORGENA  BLACK DOG, THE  P
GOLD, BOB  ANAL ASTRONAUT  SR-L
GOLD, HORACE L.  BODYGUARD  PA,C/PA
GOLD, HORACE L.  BODYGUARD AND FOUR OTHER SHORT NOVELS FROM GALAXY  HA
GOLD, HORACE L./KREPPS, ROBERT W.  ENORMOUS ROOM, THE  B
GOLD, HORACE L.  5 GALAXY SHORT NOVELS  C/PA,PA
GOLD, HORACE L.  SECOND GALAXY READER OF SCIENCE FICTION, THE  HA
GOLD, HORACE L.  THIRD GALAXY READER, THE  C/PA,PA
GOLD, HORACE L.  FOURTH GALAXY READER, THE  C/PA,PA
GOLD, HORACE L.  FIFTH GALAXY READER, THE  C/PA,PA
GOLD, HORACE L.  SIXTH GALAXY READER, THE  HA
GOLD, HORACE L.  MIND PARTNER  PA,C/PA
GOLD, HORACE L.  WORLD THAT COULDN’T BE, THE  PA,C/PA

GOLD KEY COMICS STAR TREK: THE KEY COLLECTION VOLUME 3

[STAR TREK KEY COLLECTION 3]

CX

GOLDBERG, GERRY/STOROSCHUK, STEPHEN/ CORBETT, FRED NIGHTTOUCH: JOURNEYING INTO THE REALMS OF NIGHTMARE

VLPA

GOLDBERG, MARSHALL/KAY, KENNETH  DISPOSABLE PEOPLE  P
GOLDBERG, MARSHALL  NERVE  P
GOLDEN BOOKS  STAR WARS THE GREATEST BATTLES (3-D)  A
GOLDEN, CHRISTIE  CLOAK AND DAGGER [@STAR TREK VOYAGER DARK MATTERS #1]  P
GOLDEN, CHRISTIE  DANCE OF THE DEAD [@RAVENLOFT]  P
GOLDEN, CHRISTIE  ENEMY OF MY ENEMY [@STAR TREK VOYAGER SPIRIT WALK 2]  P
GOLDEN, CHRISTIE  ENEMY WITHIN, THE [@RAVENLOFT]  P
GOLDEN, CHRISTIE  FARTHER SHORE, THE [@STAR TREK VOYAGER HOMECOMING 2]  P
GOLDEN, CHRISTIE  FIRSTBORN [@STARCRAFT DARK TEMPLAR #1]  P
GOLDEN, CHRISTIE  GHOST DANCE [@STAR TREK VOYAGER DARK MATTERS #2]  P
GOLDEN, CHRISTIE  HOMECOMING [@STAR TREK VOYAGER HOMECOMING 1]  P
GOLDEN, CHRISTIE  IN STONE’S GRASP [KEVLA OF ARUKAN 2]  LP
GOLDEN, CHRISTIE  INSTRUMENT OF FATE  P
GOLDEN, CHRISTIE  INVASION AMERICA [INVASION AMERICA 1]  P
GOLDEN, CHRISTIE  KING’S MAN & THIEF  P
GOLDEN, CHRISTIE  LAST ROUNDUP, THE [@STAR TREK]  P
GOLDEN, CHRISTIE  LORD OF THE CLANS [@WARCRAFT #2]  P
GOLDEN, CHRISTIE  MAROONED [STAR TREK VOYAGER]  C/P,P
GOLDEN, CHRISTIE  MURDERED SUN, THE [STAR TREK VOYAGER]  P
GOLDEN, CHRISTIE  NO MAN’S LAND [STAR TREK GATEWAYS #5/STAR TREK VOYAGER]  P
GOLDEN, CHRISTIE  OLD WOUNDS [@STAR TREK VOYAGER SPIRIT WALK 1]  P
GOLDEN, CHRISTIE  ON FIRE’S WINGS [KEVLA OF ARUKAN 1]  LP
GOLDEN, CHRISTIE  ON THE RUN [INVASION AMERICA 2]  P
GOLDEN, CHRISTIE  RISE OF THE HORDE [WORLD OF WARCRAFT]  P
GOLDEN, CHRISTIE  SEVEN OF NINE [STAR TREK VOYAGER]  C/P:2,P
GOLDEN, CHRISTIE  SHADOW OF HEAVEN [@STAR TREK VOYAGER DARK MATTERS #3]  P
GOLDEN, CHRISTIE  VAMPIRE OF THE MISTS [@RAVENLOFT]  C/P:2,P
GOLDEN, CHRISTOPHER  ANGEL SOULS AND DEVIL HEARTS [SHADOWS 2]  C/P,P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  BLOODED [BUFFY]  C/P,P
GOLDEN, CHRISTOPHER  BOYS ARE BACK IN TOWN, THE  LP
GOLDEN, CHRISTOPHER  BOYS ARE BACK IN TOWN, THE (PROOF)  SR-LP
GOLDEN, CHRISTOPHER/HOLDER, NANCY  CHILD OF THE HUNT [@BUFFY]  C/P,P
GOLDEN, CHRISTOPHER  CODENAME WOLVERINE [X-MEN]  H
GOLDEN, CHRISTOPHER  DARK TIMES [@BUFFY LOST SLAYER #2]  C/P
GOLDEN, CHRISTOPHER  GATHERING DARK, THE [SHADOW SAGA]  C/P,P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  GHOST ROADS [@BUFFY GATEKEEPER 2]  C/P:2,P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  HALLOWEEN RAIN [BUFFY]  P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  IMMORTAL [BUFFY]  C/P,P
GOLDEN, CHRISTOPHER  KING OF THE DEAD [@BUFFY LOST SLAYER #3]  C/P:2
GOLDEN, CHRISTOPHER  LOST SLAYER, THE [@BUFFY LOST SLAYER 1,2,3,4]  P

GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

MONSTER ISLAND [@BUFFY/ANGEL]

P

GOLDEN, CHRISTOPHER  MYTH HUNTERS, THE [VEIL 1]  LP

GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

NIMBLE MAN, THE [@MENAGERIE] P

AUTHOR  TITLE  CODE
GOLDEN, CHRISTOPHER  OF MASQUES AND MARTYRS [SHADOWS 3]  P
GOLDEN, CHRISTOPHER  OF SAINTS AND SHADOWS [SHADOWS 1]  C/P,P
GOLDEN, CHRISTOPHER  ORIGINAL SINS [@BUFFY LOST SLAYER #4]  C/P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  OUT OF THE MADHOUSE [@BUFFY GATEKEEPER 1]  C/P,P
GOLDEN, CHRISTOPHER  OZ: INTO THE WILD [@BUFFY]  P
GOLDEN, CHRISTOPHER  PREDATOR AND PREY [@PROWLERS 3]  P
GOLDEN, CHRISTOPHER  PREDATOR’S SMILE [@DAREDEVIL]  P
GOLDEN, CHRISTOPHER  PRETTY MAIDS ALL IN A ROW [BUFFY SPIKE & DRU]  H,P
GOLDEN, CHRISTOPHER  PROPHECIES [@BUFFY LOST SLAYER #1]  C/P
GOLDEN, CHRISTOPHER  SINS OF THE FATHER [BUFFY]  C/P,P
GOLDEN, CHRISTOPHER/HOLDER, NANCY  SONS OF ENTROPY [@BUFFY GATEKEEPER 3]  P
GOLDEN, CHRISTOPHER  STRANGEWOOD  LP

GOLDEN, CHRISTOPHER/ SNIEGOSKI, THOMAS E.

TEARS OF THE FURIES [@MENAGERIE]

P

GOLDEN, CHRISTOPHER  WILD THINGS [@PROWLERS 4]  P
GOLDEN, CHRISTOPHER  WISDOM OF WAR, THE [@BUFFY]  C/P,P
GOLDIN, STEPHEN  ALIEN CONDITION, THE  C/PA,PA
GOLDIN, STEPHEN  AND NOT MAKE DREAMS YOUR MASTER  C/P:2,P
GOLDIN, STEPHEN  ASSAULT ON THE GODS  P,SFWA
GOLDIN, STEPHEN  CARAVAN  SR-P
GOLDIN, STEPHEN  CRYSTALS OF AIR AND WATER [PARSINA 3]  P
GOLDIN, STEPHEN  ETERNITY BRIGADE, THE  C/P,P
GOLDIN, STEPHEN  ETERNITY BRIGADE, THE (PROOF)  SR-P
GOLDIN, STEPHEN  FINISH LINE  SR-P
GOLDIN, STEPHEN  HERDS  SR-P
GOLDIN, STEPHEN/MASON, MARY  JADE DARCY AND THE AFFAIR OF HONOR [@JADE DARCY #1]  C/P,P
GOLDIN, STEPHEN/MASON, MARY  JADE DARCY AND THE ZEN PIRATES [@JADE DARCY #2]  C/P:2,P
GOLDIN, STEPHEN  MINDFLIGHT  C/P:2,P
GOLDIN, STEPHEN  SCAVENGER HUNT  SR-P
GOLDIN, STEPHEN  SHRINE OF THE DESERT MAGE [PARSINA 1]  P
GOLDIN, STEPHEN  STORYTELLER AND THE JAHN, THE [PARSINA 2]  P
GOLDIN, STEPHEN  TREK TO MADWORLD [STAR TREK]  C/P:2,P
GOLDIN, STEPHEN  WORLD CALLED SOLITUDE, A  H,SFWA,P
GOLDING, MORTON J.  NIGHT MARE  C/P:2,P
GOLDING, WILLIAM  BRASS BUTTERFLY, THE  SR-P
GOLDING, WILLIAM  DARKNESS VISIBLE  P
GOLDING, WILLIAM  INHERITORS, THE  P
GOLDING, WILLIAM  LORD OF THE FLIES  C/P:2,P
GOLDMAN, WILLIAM  CONTROL  P
GOLDMAN, WILLIAM  PRINCESS BRIDE, THE  SR-P,C/P:2
GOLDMAN, WILLIAM  SILENT GONDOLIERS, THE  C/P,P
GOLDS, ARLENE  FROM DREAM TO DREAM  H
GOLDS, CASSANDRA  MICHAEL AND THE SECRET WAR  SFWA
GOLDSBOROUGH, F.  CRIMSON SPELL, THE [@CHARMED]  P
GOLDSMITH, DAVID H.  KURT VONNEGUT: FANTASIST OF FIRE AND ICE  R/REF-P
GOLDSMITH, HOWARD  INVASION: 2200 A.D.  H
GOLDSTEIN, LISA  ALCHEMIST’S DOOR, THE  H
GOLDSTEIN, LISA  DARK CITIES UNDERGROUND  H,LP
GOLDSTEIN, LISA  DREAM YEARS, THE  H:2,P
GOLDSTEIN, LISA  MASK FOR THE GENERAL, A  H,P
GOLDSTEIN, LISA  RED MAGICIAN, THE  C/P:2,SFWA,SFWA-P,LP
GOLDSTEIN, LISA  STRANGE DEVICES OF THE SUN AND MOON  H
GOLDSTEIN, LISA  SUMMER KING, WINTER FOOL  P
GOLDSTEIN, LISA  TRAVELLERS IN MAGIC  H,LP
GOLDSTEIN, LISA  WALKING THE LABYRINTH  LP
GOLDSTEIN, STAN/GOLDSTEIN, FRED  STAR TREK SPACEFLIGHT CHRONOLOGY [STAR TREK]  VLP
GOLDSTEIN, WILLIAM  DR. PHIBES  C/P,P
GOLDSTEIN, WILLIAM  DR. PHIBES RISES AGAIN!  P
GOLDSTON, ROBERT  LAST OF LAZARUS, THE  P
GOLDSTONE, LAWRENCE  OFF-LINE  H
GOLL, REINHOLD W.  VISITORS FROM PLANET VETA, THE  H
GOMERY, PERCY  END OF THE CIRCLE  H

GONICK, LARRY CARTOON HISTORY OF THE UNIVERSE 1: THE EVOLUTION OF EVERYTHING, THE

SR-VLP

GONICK, LARRY  CARTOON HISTORY OF THE UNIVERSE 2: STICKS AND STONES, THE  SR-VLP
GONICK, LARRY  CARTOON HISTORY OF THE UNIVERSE 3: RIVER REALMS, THE  SR-VLP

GONICK, LARRY CARTOON HISTORY OF THE UNIVERSE 4: PART OF THE OLD TESTAMENT, THE

SR-VLP

GONICK, LARRY  CARTOON HISTORY OF THE UNIVERSE 5: BRAINS AND BRONZE, THE  SR-VLP

GONICK, LARRY CARTOON HISTORY OF THE UNIVERSE 6: WHO ARE THESE ATHENIANS?, THE

SR-VLP

GONZALEZ, MIKE  COLLINS SPANISH-ENGLISH ENGLISH-SPANISH DICTIONARY  DICT
GOOD, I. J.  SCIENTIST SPECULATES, THE  [C/RAN-H]
GOODFELLOW, CODY  RADIANT DAWN [RADIANT DAWN 1]  LP
GOODFELLOW, CODY  RAVENOUS DUSK [RADIANT DAWN 2]  LP

AUTHOR  TITLE  CODE
GOODKIND, TERRY  BLOOD OF THE FOLD [SWORD OF TRUTH 3]  C/P:2,P,H
GOODKIND, TERRY  CHAINFIRE [SWORD OF TRUTH 9]  H
GOODKIND, TERRY  DEBT OF BONES [SWORD OF TRUTH 10]  H
GOODKIND, TERRY  FAITH OF THE FALLEN [SWORD OF TRUTH 6]  C/P,P,H
GOODKIND, TERRY  NAKED EMPIRE [SWORD OF TRUTH 8]  H
GOODKIND, TERRY  PHANTOM [SWORD OF TRUTH 11]  H
GOODKIND, TERRY  PILLARS OF CREATION, THE [SWORD OF TRUTH 7]  H:2
GOODKIND, TERRY  SOUL OF THE FIRE [SWORD OF TRUTH 5]  H:2,P
GOODKIND, TERRY  STONE OF TEARS [SWORD OF TRUTH 2]  C/P:2,P
GOODKIND, TERRY  TEMPLE OF THE WINDS [SWORD OF TRUTH 4]  H:2,P
GOODKIND, TERRY  WIZARD’S FIRST RULE [SWORD OF TRUTH 1]  C/P:2,SFWA-P,P
GOODMAN, LAURIE  SPELL OF DECEIT, A  P
GOODSTONE, TONY  PULPS, THE  R/REF-H
GOODWIN, HAROLD LELAND  SCIENCE BOOK OF SPACE TRAVEL, THE  [C/RAN-P:2]
GOODWIN, MICHAEL  WHO WAS THAT MONOLITH I SAW YOU WITH?  CX
GOONAN, KATHLEEN ANN  BONES OF TIME, THE  C/P,P
GOONAN, KATHLEEN ANN  CRESCENT CITY RHAPSODY [NANOTECH]  H,P
GOONAN, KATHLEEN ANN  IN WAR TIMES  H
GOONAN, KATHLEEN ANN  LIGHT MUSIC [NANOTECH]  H
GOONAN, KATHLEEN ANN  MISSISSIPPI BLUES [NANOTECH VERITY 2]  LP
GOONAN, KATHLEEN ANN  QUEEN CITY JAZZ [NANOTECH VERITY 1]  C/P,P
GORDEN, GREG  PROPHECY [@EARTHDAWN]  P
GORDON, DONALD  FLIGHT OF THE BAT  H
GORDON, DONALD  LEAP IN THE DARK  P
GORDON, DONALD  STAR-RAKER  H
GORDON, FRITZ  FLIGHT OF THE BAMBOO SAUCER, THE  C/P,P
GORDON, JOAN  JOE HALDEMAN  R/REF-P
GORDON, REX  FIRST ON MARS  C/P:2,P
GORDON, REX  FIRST THROUGH TIME  C/P:2,P
GORDON, REX  FIRST TO THE STARS  C/P:2,P
GORDON, REX  UTOPIA MINUS X  C/P:2,P
GORDON, REX  YELLOW FRACTION, THE  C/P:2,P
GORDON, STUART  FIRE IN THE ABYSS  C/P,P
GORDON, STUART  HIDDEN WORLD, THE [WATCHERS 2]  P
GORDON, STUART  ONE-EYE  P
GORDON, STUART  SMILE ON THE VOID  C/P,SFWA,P,H
GORDON, STUART  THREE-EYES  C/P,P
GORDON, STUART  TIME STORY  C/P:2,P
GORDON, STUART  TWO-EYES  P
GOREY, EDWARD  GASHLYCRUMB TIMES 2000 CALENDAR, THE  CAL
GORMAN, ED/GREENBERG, MARTIN H.  PREDATORS  SFWA
GORMAN, ED/GREENBERG, MARTIN H.  STALKERS  PA
GORODISCHER, ANGELICA  KALPA IMPERIAL  LP
GOROG, JUDITH  WINNING SCHEHEREZADE  SFWA
GORTON, RON  SOUL, THE  C/P,P
GOSCINNY/UDERZO  ASTERIX THE LEGIONARY  [C/RAN-H]
GOSS, THEODORA  IN THE FOREST OF FORGETTING  H
GOTLIEB, PHYLLIS  EMPEROR, SWORDS, PENTACLES  C/P:2,P
GOTLIEB, PHYLLIS  FLESH AND GOLD [FLESH AND GOLD 1]  H,LP
GOTLIEB, PHYLLIS  HEART OF RED IRON [BARRAZAN V 2]  H
GOTLIEB, PHYLLIS  JUDGMENT OF DRAGONS, A  C/P:3,P
GOTLIEB, PHYLLIS  KINGDOM OF THE CATS, THE  C/P:2,P
GOTLIEB, PHYLLIS  MINDWORLDS [FLESH AND GOLD 3]  H
GOTLIEB, PHYLLIS  O MASTER CALIBAN! [BARRAZAN V 1]  C/P,SFWA,P,H
GOTLIEB, PHYLLIS  SON OF THE MORNING AND OTHER STORIES  P
GOTLIEB, PHYLLIS  SUNBURST  C/P:2,P
GOTLIEB, PHYLLIS/BARBOUR, DOUGLAS  TESSERACTS 2  LPA
GOTLIEB, PHYLLIS  VIOLENT STARS [FLESH AND GOLD 2]  H
GOTO, C. S.  ASCENSION [@WARHAMMER 40K DAWN OF WAR 2]  P
GOTO, C. S.  ELDAR PROPHECY [@WARHAMMER 40K]  P
GOTO, C. S.  SALVATION [@NECROMUNDA]  P
GOTO, C. S.  TEMPEST [@WARHAMMER 40K DAWN OF WAR 3]  P
GOTO, C. S.  WARRIOR BLOOD [@WARHAMMER 40K DEATHWATCH 1]  P
GOTO, C. S.  WARRIOR COVEN [@WARHAMMER 40K DEATHWATCH 2]  P
GOTSCHALK, FELIX C.  GROWING UP IN TIER 3000  C/P,P
GOTTESMAN, RONALD/GEDULD, HARRY  GIRL IN THE HAIRY PAW, THE  R/REF-H
GOTTLIEB, HINKO  KEY TO THE GREAT GATE, THE  H
GOTTLIEB, SHERRY  LOVE BITE  P
GOUDGE, ELIZABETH  LITTLE WHITE HORSE, THE  P
GOULART, RON  AFTER THINGS FELL APART  C/P:3,P
GOULART, RON  BIG BANG  P
GOULART, RON  BLOOD WEDDING [@VAMPIRELLA #4]  P
GOULART, RON  BLOODSTALK [@VAMPIRELLA #1]  C/P,P
GOULART, RON  BRAINZ, INC.  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 164
AUTHOR  TITLE  CODE
GOULART, RON  BRINKMAN  H,SFWA
GOULART, RON  BROKE DOWN ENGINE  C/P,P,H
GOULART, RON  CALLING DR. PATCHWORK  P
GOULART, RON  CAPRICORN ONE  C/P,P
GOULART, RON  CHALLENGERS OF THE UNKNOWN  P
GOULART, RON  CHAMELEON CORPS AND OTHER SHAPE CHANGERS, THE  H:2,P
GOULART, RON  CHEAP THRILLS  R/REF-H
GOULART, RON  CLOCKWORK’S PIRATES  D11182
GOULART, RON  COWBOY HEAVEN  H:2,SFWA
GOULART, RON  CRACKPOT  H:2,SFWA
GOULART, RON  CURSE OF THE OBELISK, THE  C/P:2,P
GOULART, RON  CYBORG KING, THE [STAR HAWKS 2]  P
GOULART, RON  DAREDEVILS, LTD. [EXCHAMELEON 1]  C/P:2,P
GOULART, RON  DEADWALK [@VAMPIRELLA #3]  P
GOULART, RON  DEATH CELL  C/P:2,P
GOULART, RON  DEATHGAME [@VAMPIRELLA #5]  P
GOULART, RON  DR. SCOFFLAW  BS3
GOULART, RON  EMPEROR OF THE LAST DAYS, THE  P
GOULART, RON  EMPIRE 99 [STAR HAWKS 1]  C/P,P
GOULART, RON  ENORMOUS HOURGLASS, THE  P
GOULART, RON  EVERYBODY COMES TO COSMO’S [EXCHAMELEON 3]  C/P,P
GOULART, RON  EYE OF THE VULTURE [GYPSY 2/WEIRD HEROES 7]  P
GOULART, RON  FIRE-EATER, THE  C/P:2,P
GOULART, RON  FLUX  C/P:2,P
GOULART, RON  GADGET MAN  C/P,SFWA,P,H
GOULART, RON  GALAXY JANE  C/P:2,P
GOULART, RON  GHOST BREAKER  D11182
GOULART, RON  HAIL HIBBLER  C/P:2,P
GOULART, RON  HAWKSHAW  H,SFWA,P
GOULART, RON  HELLHOUND PROJECT, THE  H,SFWA
GOULART, RON  HELLO, LEMURIA HELLO  P
GOULART, RON  HELLQUAD  C/P:2,P
GOULART, RON  NEMO  C/P,P
GOULART, RON  NUTZENBOLTS AND MORE TROUBLES WITH MACHINES  H:2,SFWA
GOULART, RON  ODD JOB #101  H,SFWA
GOULART, RON  ON ALIEN WINGS [@VAMPIRELLA #2]  C/P,P
GOULART, RON  PANCHRONICON PLOT, THE  C/P,P
GOULART, RON  PLUNDER  C/P:2,P
GOULART, RON  PRISONER OF BLACKWOOD CASTLE, THE  C/P:2,P
GOULART, RON  QUEST OF THE GYPSY [GYPSY 1]  C/P,P
GOULART, RON  ROBOT IN THE CLOSET, THE  P
GOULART, RON  SAME LIE TWICE, THE  P
GOULART, RON  SHAGGY PLANET  C/P:2,P
GOULART, RON  SKYROCKET STEELE  P

GOULART, RON SKYROCKET STEELE CONQUERS THE UNIVERSE

[@AUTHOR’S CHOICE MONTHLY #11]

SR-H

GOULART, RON  SPACEHAWK, INC.  C/P:2,P
GOULART, RON  STARPIRATE’S BRAIN [EXCHAMELEON 2]  C/P:2,P
GOULART, RON  SUICIDE, INC.  C/P:2,P
GOULART, RON  SWORD SWALLOWER, THE  C/P,P,H
GOULART, RON  TALENT FOR THE INVISIBLE, A  C/P:2,P
GOULART, RON  TIN ANGEL, THE  C/P:2,P
GOULART, RON  UPSIDE DOWNSIDE  P
GOULART, RON  WHAT’S BECOME OF SCREWLOOSE? AND OTHER INQUIRIES  H:2,SFWA,P
GOULART, RON  WHEN THE WAKER SLEEPS  C/P:2,P
GOULART, RON  WHIFF OF MADNESS, A  P
GOULART, RON  WICKED CYBORG, THE  P
GOULART, RON  WILDSMITH  C/P:2,P
GOULD, CHESTER  DICK TRACY MEETS THE NIGHT CRAWLER  H
GOULD, STEVEN C.  BLIND WAVES  C/P,P,H
GOULD, STEVEN C./MIXON, LAURA J.  GREENWAR  P
GOULD, STEVEN C.  HELM  H,P
GOULD, STEVEN C.  JUMPER [JUMPER 1]  H:2,P
GOULD, STEVEN C.  REFLEX [JUMPER 2]  H
GOULD, STEVEN C.  WILDSIDE  H,P
GOWER, DANIEL H.  ORPHEUS PROCESS, THE  P
GRAAT, HEINRICH  REVENGE OF INCREASE SEWALL, THE  P
GRABINSKI, STEFAN  DARK DOMAIN, THE  LP
GRAF, L. A.  ARMAGEDDON SKY [STAR TREK DAY OF HONOR 2/STAR TREK DS9]  C/P:2,P
GRAF, L. A.  CARETAKER [STAR TREK VOYAGER]  C/P,P
GRAF, L. A.  DEATH COUNT [STAR TREK]  C/P:2,P
GRAF, L. A.  EXTREME PREJUDICE [@ALIEN NATION #7]  C/P,P
GRAF, L. A.  FIRESTORM [STAR TREK]  C/P:2,P
GRAF, L. A.  FUTURE IMPERFECT [@STAR TREK JANUS GATE #2]  P

AUTHOR  TITLE  CODE
GRAF, L. A.  ICE TRAP [STAR TREK]  C/P:2,P
GRAF, L. A.  JANUS GATE, THE [STAR TREK JANUS GATE 1,2,3]  H
GRAF, L. A.  PAST PROLOGUE [@STAR TREK JANUS GATE #3]  P
GRAF, L. A.  PRESENT TENSE [@STAR TREK JANUS GATE #1]  P
GRAF, L. A.  ROUGH TRAILS [@STAR TREK NEW EARTH #3]  C/P,P
GRAF, L. A.  TIME’S ENEMY [STAR TREK INVASION! 3/STAR TREK DS9]  C/P,P
GRAF, L. A.  TRAITOR WINDS [STAR TREK]  C/P:2,P
GRAF, L. A.  WAR DRAGONS [STAR TREK CAPTAIN’S TABLE 1]  C/P:2,P
GRAHAM, DANIEL, JR.  GATEKEEPERS, THE  H,P
GRAHAM, DAVID  DOWN TO A SUNLESS SEA  H:2,P
GRAHAM, DENISE B.  CURSE OF THE LOST GROVE [@KNIGHTS OF THE SILVER DRAGON #10]  LP
GRAHAM, DENISE B.  EYE OF FORTUNE [@KNIGHTS OF THE SILVER DRAGON #4]  LP
GRAHAM, IAN  MONUMENT  C/P,P,H
GRAHAM, IRVING  HOW TO SELL THROUGH MAIL ORDER  [C/RAN-H]
GRAHAM, KENNON  SURPRISE GUESTS, THE [LAND OF THE LOST]  H
GRAHAM, L. B.  BEYOND THE SUMMERLAND [BINDING OF THE BLADE 1]  LP
GRAHAM, L. B.  BRINGER OF STORMS [BINDING OF THE BLADE 2]  LP
GRAHAM, L. B.  FATHER OF DRAGONS [@BINDING OF THE BLADE #4]  LP
GRAHAM, L. B.  SHADOW IN THE DEEP [@BINDING OF THE BLADE #3]  LP
GRAHAM, MITCHELL  ANCIENT LEGACY, THE [MATHEW LEWIN 3]  P
GRAHAM, MITCHELL  EMERALD CAVERN, THE [MATHEW LEWIN 2]  P
GRAHAM, MITCHELL  FIFTH RING, THE [MATHEW LEWIN 1]  P
GRAHAM, ROBERT  ATTAR’S REVENGE [@ATTAR #1]  C/P,P
GRAHAM, ROBERT  WAR OF NERVES [@ATTAR #2]  P
GRAHAM, VICTORIA  WITCHSTONE, THE  P
GRAHAME, KENNETH  DREAM DAYS  LP
GRAHAME, KENNETH  GOLDEN AGE, THE  LP
GRAHAME, KENNETH  WIND IN THE WILLOWS, THE  P
GRANT, ALAN  CURSE [@SMALLVILLE]  P
GRANT, ALAN  DRAGON [@SMALLVILLE]  P
GRANT, CHARLES L.  AFTER MIDNIGHT  PA
GRANT, CHARLES L.  ASCENSION  P
GRANT, CHARLES L.  BEST OF SHADOWS, THE  HA
GRANT, CHARLES L.  BLOODWIND, THE  C/P,SFWA-P,P
GRANT, CHARLES L.  CURSE, THE  P
GRANT, CHARLES L.  DIALING THE WIND  SFWA-P
GRANT, CHARLES L.  DODD, MEAD GALLERY OF HORROR, THE  HA
GRANT, CHARLES L.  DOOM CITY [GREYSTONE BAY 2]  PA,SFWA-P
GRANT, CHARLES L.  FEARS  C/PA,PA
GRANT, CHARLES L.  FINAL SHADOWS  HA
GRANT, CHARLES L.  FOR FEAR OF THE NIGHT  SFWA-P
GRANT, CHARLES L.  GALLERY OF HORROR  PA
GRANT, CHARLES L.  GLOW OF CANDLES AND OTHER STORIES, A  P
GRANT, CHARLES L.  GOBLINS [X-FILES]  C/P:2,P
GRANT, CHARLES L.  GOBLINS [X-FILES]/WHIRLWIND [X-FILES]  H,LP
GRANT, CHARLES L.  GRAVE, THE  P,SFWA-P
GRANT, CHARLES L.  GREYSTONE BAY [GREYSTONE BAY 1]  SFWA-P
GRANT, CHARLES L.  HORRORS  PA
GRANT, CHARLES L.  HOUR OF THE OXRUN DEAD, THE  H,SFWA-P,P
GRANT, CHARLES L.  HUNTING GROUND [@BLACK OAK 4]  P
GRANT, CHARLES L.  IN A DARK DREAM  P,SFWA-P
GRANT, CHARLES L.  IN THE MOOD [MILLENNIUM QUARTET 2]  H
GRANT, CHARLES L.  LAST CALL OF MOURNING, THE  H,SFWA-P
GRANT, CHARLES L.  LEGION  P
GRANT, CHARLES L.  LONG NIGHT OF THE GRAVE, THE  H
GRANT, CHARLES L.  MIDNIGHT  PA
GRANT, CHARLES L.  NIGHT SONGS  C/P,P
GRANT, CHARLES L.  NIGHT VISIONS 2  SR-HA
GRANT, CHARLES L.  NIGHTMARE SEASONS  P,SFWA
GRANT, CHARLES L.  NIGHTMARES  PA
GRANT, CHARLES L.  ORCHARD, THE  SFWA-P
GRANT, CHARLES L.  PET, THE  P,SFWA-P
GRANT, CHARLES L.  QUIET NIGHT OF FEAR, A  C/P,P
GRANT, CHARLES L.  RAVENS OF THE MOON, THE  H,SFWA
GRANT, CHARLES L.  RIDERS IN THE SKY  H
GRANT, CHARLES L.  SEAHARP HOTEL, THE [GREYSTONE BAY 3]  SFWA-P
GRANT, CHARLES L.  SHADOW OF ALPHA, THE  C/P:2,P
GRANT, CHARLES L.  SHADOWS  HA,PA,SFWA
GRANT, CHARLES L.  SHADOWS 2  HA,PA
GRANT, CHARLES L.  SHADOWS 3  C/PA,PA,SFWA
GRANT, CHARLES L.  SHADOWS 4  HA,PA,SFWA
GRANT, CHARLES L.  SHADOWS 5  HA,PA,SFWA
GRANT, CHARLES L.  SHADOWS 6  SFWA
GRANT, CHARLES L.  SHADOWS 7  HA,PA,SFWA

AUTHOR  TITLE  CODE
GRANT, CHARLES L.  SHADOWS 8  HA,PA,SFWA
GRANT, CHARLES L.  SHADOWS 9  HA,SFWA
GRANT, CHARLES L.  SHADOWS 10  HA,SFWA
GRANT, CHARLES L.  SOFT WHISPER OF THE DEAD, THE  SR-H
GRANT, CHARLES L.  SOUND OF MIDNIGHT, THE  H,SFWA-P
GRANT, CHARLES L.  STUNTS  SFWA
GRANT, CHARLES L.  SYMPHONY [MILLENNIUM QUARTET 1]  P
GRANT, CHARLES L.  TALES FROM THE NIGHTSIDE  SR-H,SFWA-P
GRANT, CHARLES L.  WATCHER [@WORLD OF DARKNESS WEREWOLF]  C/P,P
GRANT, CHARLES L.  WHIRLWIND [X-FILES]  C/P,P
GRANT, CHARLES L.  WINTER KNIGHT [@BLACK OAK 3]  P
GRANT, CHARLES L.  WRITING AND SELLING SCIENCE FICTION  R/REF-H
GRANT, CLAY  DEMON SAMURAI, THE  P
GRANT, DONALD M.  TALBOT MUNDY: MESSENGER OF DESTINY  R/REF-H
GRANT, DONALD M.  VIRGIL FINLAY  A
GRANT, JOAN  EYES OF HORNS  P
GRANT, JOAN  RETURN TO ELYSIUM  P
GRANT, JOAN  SCARLET FEATHER  P
GRANT, JOAN  SO MOSES WAS BORN  P
GRANT, JOAN  WINGED PHARAOH  H,P
GRANT, JOHN/WILSON, COLIN  QINMEARTHA AND THE GIRL CHILD LOCHI/TOMB OF THE OLD ONES  LP
GRANT, KATHRYN  BLACK JADE ROAD, THE [TEN THOUSAND WILLOWS 2]  C/P,P
GRANT, KATHRYN  PHOENIX BELLS, THE [TEN THOUSAND WILLOWS 1]  P,SFWA-P
GRANT, KATHRYN  WILLOW GARDEN, THE [TEN THOUSAND WILLOWS 3]  C/P,P
GRANT, MARK  CHRISTMAS SLAUGHTER  P
GRANT, MARK  HOLOCAUST HORROR [@MUTANTS AMOK #4]  C/P,P
GRANT, MARK  MUTANT HELL [@MUTANTS AMOK #2]  C/P,P
GRANT, MARK  MUTANTS AMOK [@MUTANTS AMOK #1]  C/P,P
GRANT, MARK  REBEL ATTACK [@MUTANTS AMOK #3]  P
GRANT, MAXWELL  DEATH TOWER, THE  P
GRANT, MAXWELL  GANGDOM’S DOOM  P
GRANT, MAXWELL  GHOST MAKERS, THE  P
GRANT, MAXWELL  GREEN EYES  P
GRANT, MAXWELL  HIDDEN DEATH  P
GRANT, MAXWELL  LIVING SHADOW, THE  C/P,P
GRANT, MAXWELL  SHADOW: DESTINATION MOON, THE  P
GRANT, MAXWELL  SHADOW — GO MAD!  P
GRANT, MAXWELL  SHADOW’S REVENGE, THE  P
GRANT, MAXWELL  SILENT DEATH, THE  P
GRANT, RICHARD  RUMORS OF SPRING  C/P,SFWA,P,H
GRANT, RICHARD  SARABAND OF LOST TIME  C/P,P
GRANT, RICHARD  TEX AND MOLLY IN THE AFTERLIFE  P
GRANT, RICHARD  THROUGH THE HEART  C/P:2,P
GRANT, RICHARD  VIEWS FROM THE OLDEST HOUSE  C/P,LP
GRANT, ROB  COLONY  P
GRANT, ROB  INCOMPETENCE  H
GRANT, STEVEN/DUFFY, JO  PUNISHER, THE  CX
GRASS, GUNTER  FLOUNDER, THE  P
GRAVEL, GEARY  ALCHEMISTS, THE [AUTUMN WORLD MOSAIC 1]  C/P:2,P
GRAVEL, GEARY  DREAMWRIGHT [@MIGHT AND MAGIC #1]  P
GRAVEL, GEARY  DUAL TO THE DEATH [@BATMAN ANIMATED SERIES 2]  P
GRAVEL, GEARY  KEY FOR THE NONESUCH, A [FADING WORLDS 1]  C/P:2,P
GRAVEL, GEARY  MASK OF THE PHANTASM [@BATMAN]  P
GRAVEL, GEARY  PATHFINDERS, THE [AUTUMN WORLD MOSAIC 2]  P
GRAVEL, GEARY  RETURN OF THE BREAKNECK BOYS [FADING WORLDS 2]  P
GRAVEL, GEARY  SHADOWS OF THE PAST [@BATMAN ANIMATED SERIES 1]  P
GRAVEL, GEARY  SHADOWSMITH, THE [@MIGHT AND MAGIC #2]  C/P,P
GRAVERSEN, PAT  INVISIBLE FIRE  P
GRAVES, ROBERT  SEVEN DAYS IN NEW CRETE  C/P,P
GRAVES, ROBERT  WATCH THE NORTHWIND RISE  C/P,P
GRAVES, STEPHAN  SWORDFISH ISLAND COVENANT, THE  LP
GRAY, ALASDAIR  HISTORY MAKER, A  LP
GRAY, ALASDAIR  LANARK  H,LP
GRAY, ALASDAIR  TEN TALES TALL AND TRUE  H
GRAY, CURME  MURDER IN MILLENNIUM VI  H:2
GRAY, JULIA  CRYSTAL DESERT, THE [@GUARDIAN #3]  P
GRAY, JULIA  DARK MOON, THE [@GUARDIAN #1]  P
GRAY, JULIA  FIRE MUSIC [TIGUAFAYA 2]  C/P,P
GRAY, JULIA  ICE MAGE [TIGUAFAYA 1]  P
GRAY, JULIA  JASPER FOREST, THE [@GUARDIAN #2]  P
GRAY, LINDA CROCKETT  SCRYER  P,SFWA-P
GRAY, LINDA CROCKETT  TANGERINE  SFWA-P
GRAY, NICHOLAS STUART  GRIMBOLD’S OTHER WORLD  P
GRAY, NICHOLAS STUART  SEVENTH SWAN, THE  P

AUTHOR  TITLE  CODE
GRAY, ROD  LAID IN THE FUTURE  SR-L
GRAYSON, DEVIN  CITY [@SMALLVILLE]  P
GRAZIER, JAMES  RUNTS OF 61 CYGNI C  C/P,P
GRAZIUNAS, DAINA/STARLIN, JIM  PREDATORS  P
GREELEY, ANDREW M.  ANGEL FIRE  SFWA-P
GREELEY, ANDREW M.  EMERALD MAGIC  PA
GREELEY, ANDREW M.  FINAL PLANET, THE  H,SFWA-P,P
GREELEY, ANDREW M.  GOD GAME  H:2,SFWA,SFWA-P,P
GREELEY, ANDREW M.  MAGIC CUP, THE  C/P:2,P
GREELEY, ANDREW M./CASSUTT, MICHAEL  SACRED VISIONS  HA
GREEN, CELIA  OUT OF THE BODY EXPERIENCES  [C/RAN-P]
GREEN, EDWIN  AIR MONSTER  H
GREEN, I. G.  TIME BEYOND TIME  P
GREEN, JEN/LEFANU, SARAH  DESPATCHES FROM THE FRONTIERS OF THE FEMALE MIND  SR-PA
GREEN, JONATHAN  CONQUEST OF ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 2]  P
GREEN, JONATHAN  CRUSADE FOR ARMAGEDDON [@WARHAMMER 40K BLACK TEMPLARS 1]  P
GREEN, JONATHAN  DEAD AND THE DAMNED, THE [WARHAMMER]  P
GREEN, JONATHAN  IRON HANDS [@WARHAMMER 40K]  P
GREEN, JONATHAN  MAGESTORM [@WARHAMMER]  P
GREEN, JONATHAN  NECROMANCER [@WARHAMMER]  P
GREEN, JOSEPH  CONSCIENCE INTERPLANETARY  C/P,SFWA,P,H
GREEN, JOSEPH  GOLD THE MAN  P
GREEN, JOSEPH  HORDE, THE  SR-P
GREEN, JOSEPH  LOAFERS OF REFUGE, THE  C/P:2,P
GREEN, JOSEPH  MIND BEHIND THE EYE, THE  C/P:2,P
GREEN, JOSEPH  STAR PROBE  P
GREEN, MARTIN  EARTH AGAIN REDEEMED, THE  H
GREEN, ROGER LANCELYN  THIRTEEN UNCANNY TALES  HA
GREEN, ROLAND J./MURRAY, FRIEDA A.  BOOK OF KANTELA, THE [THRONE OF SHERRAN 1]  C/P,SFWA-P,LP
GREEN, ROLAND J.  CONAN AND THE DEATH LORD OF THANZA [CONAN]  P
GREEN, ROLAND J.  CONAN AND THE GODS OF THE MOUNTAIN [CONAN]  P
GREEN, ROLAND J.  CONAN AND THE MISTS OF DOOM [@CONAN]  P,SFWA-P
GREEN, ROLAND J.  CONAN AT THE DEMON’S GATE [CONAN]  LP
GREEN, ROLAND J.  CONAN THE GUARDIAN [CONAN]  P,SFWA-P
GREEN, ROLAND J.  CONAN THE RELENTLESS  P
GREEN, ROLAND J.  CONAN THE VALIANT [CONAN]  P,SFWA-P:2
GREEN, ROLAND J.  DIVISION OF THE SPOILS [STARCRUISER SHENANDOAH 2]  P
GREEN, ROLAND J./CARR, JOHN F.  GREAT KINGS’ WAR [PARATIME POLICE]  C/P:2,P
GREEN, ROLAND J.  KNIGHTS OF THE CROWN [@DRAGONLANCE WARRIORS #1]  P
GREEN, ROLAND J.  KNIGHTS OF THE ROSE [@DRAGONLANCE WARRIORS #5]  P
GREEN, ROLAND J.  KNIGHTS OF THE SWORD [@DRAGONLANCE WARRIORS #3]  P
GREEN, ROLAND J.  MOUNTAIN WALKS, THE [PEACE COMPANY]  P
GREEN, ROLAND J.  ON THE VERGE [STAR*DRIVE 1]  P
GREEN, ROLAND J.  PAINFUL FIELD, THE [STARCRUISER SHENANDOAH 5]  P
GREEN, ROLAND J.  PEACE COMPANY [PEACE COMPANY]  P
GREEN, ROLAND J.  SQUADRON ALERT [STARCRUISER SHENANDOAH 1]  C/P:2,P
GREEN, ROLAND J.  SUM OF THINGS, THE [@STARCRUISER SHENANDOAH #3]  P
GREEN, ROLAND J.  TALE OF THE COMET  P
GREEN, ROLAND J.  THESE GREEN FOREIGN HILLS [PEACE COMPANY]  C/P,P
GREEN, ROLAND J.  VAIN COMMAND [@STARCRUISER SHENANDOAH #4]  P
GREEN, ROLAND J.  VOYAGE TO ENEH [SEAS OF KILMOYN 1]  H,P
GREEN, ROLAND J.  WANDOR’S FLIGHT [WANDOR 4]  P
GREEN, ROLAND J.  WANDOR’S JOURNEY [WANDOR 2]  P
GREEN, ROLAND J.  WANDOR’S RIDE [WANDOR 1]  P
GREEN, ROLAND J.  WANDOR’S VOYAGE [WANDOR 3]  C/P,P
GREEN, ROLAND J.  WARRIORS FOR THE WORKING DAY [STARCRUISER SHENANDOAH 6]  P
GREEN, ROLAND J.  WAYWARD KNIGHTS, THE [@DRAGONLANCE WARRIORS #7]  P
GREEN, SCOTT E.  CONTEMPORARY SCIENCE FICTION, FANTASY, AND HORROR POETRY  R/REF-H
GREEN, SCOTT E.  ISAAC ASIMOV  R/REF-H
GREEN, SHARON  BETRAYALS [@BLENDING #4]  C/P,P
GREEN, SHARON  CHALLENGES [@BLENDING #3]  C/P,P
GREEN, SHARON  CHOSEN OF MIDA [JALAV 3]  C/P,P
GREEN, SHARON  COMPETITIONS [@BLENDING #2]  C/P,P
GREEN, SHARON  CONVERGENCE [@BLENDING #1]  C/P,P
GREEN, SHARON  CRYSTALS OF MIDA, THE [JALAV 1]  C/P,P
GREEN, SHARON  DARK MIRROR, DARK DREAMS [ALEXIA AND TIRAN 2/HIDDEN REALMS 2] P
GREEN, SHARON  DAWN SONG  C/P,P
GREEN, SHARON  DECEPTIONS [@BLENDING ENTHRONED #2]  P
GREEN, SHARON  DESTINY [@BLENDING ENTHRONED #3]  C/P,P
GREEN, SHARON  FAR SIDE OF FOREVER, THE [@FAR SIDE OF FOREVER #1]  C/P,P
GREEN, SHARON  GAME’S END [HIDDEN REALMS 3]  P
GREEN, SHARON  GATEWAY TO XANADU [DIANA SANTEE 2]  P
GREEN, SHARON  HELLHOUND MAGIC [@FAR SIDE OF FOREVER #2]  P
GREEN, SHARON  HIDDEN REALMS, THE [HIDDEN REALMS 1]  SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 168
AUTHOR  TITLE  CODE
GREEN, SHARON  INTRIGUES [@BLENDING ENTHRONED #1]  C/P,P
GREEN, SHARON  LADY BLADE, LORD FIGHTER  C/P,P
GREEN, SHARON  MIND GUEST [DIANA SANTEE 1]  P
GREEN, SHARON  MISTS OF THE AGES  P
GREEN, SHARON  OATH TO MIDA, AN [JALAV 2]  C/P,P
GREEN, SHARON  PROPHECY [@BLENDING #5]  C/P,P
GREEN, SHARON  REBEL PRINCE, THE  C/P,P
GREEN, SHARON  SILVER PRINCESS, GOLDEN KNIGHT [ALEXIA AND TIRAN 1]  P,SFWA-P
GREEN, SHARON  TO BATTLE THE GODS [JALAV 5]  P
GREEN, SHARON  WARRIOR CHALLENGED, THE [TERRILIAN 4]  C/P:2,P
GREEN, SHARON  WARRIOR ENCHAINED, THE [TERRILIAN 2]  C/P:2,P
GREEN, SHARON  WARRIOR REARMED, THE [TERRILIAN 3]  C/P:2,P
GREEN, SHARON  WARRIOR VICTORIOUS, THE [TERRILIAN 5]  P
GREEN, SHARON  WARRIOR WITHIN, THE [TERRILIAN 1]  C/P:2,P
GREEN, SHARON  WILL OF THE GODS, THE [JALAV 4]  P
GREEN, SHARON  WIND WHISPERS, SHADOW SHOUTS [ALEXIA AND TIRAN 3]  C/P,P
GREEN, SIMON R.  AGENTS OF LIGHT AND DARKNESS [NIGHTSIDE 2]  P
GREEN, SIMON R.  BEYOND THE BLUE MOON  C/P,P
GREEN, SIMON R.  BLOOD AND HONOR  C/P,P
GREEN, SIMON R.  BLUE MOON RISING  C/P:2,SFWA-P:2,P
GREEN, SIMON R.  BONES OF HAVEN, THE [HAWK & FISHER 6]  P
GREEN, SIMON R.  DEATHSTALKER [DEATHSTALKER 1]  C/P:2,P
GREEN, SIMON R.  DEATHSTALKER CODA [DEATHSTALKER]  H
GREEN, SIMON R.  DEATHSTALKER DESTINY [DEATHSTALKER 5]  C/P,P
GREEN, SIMON R.  DEATHSTALKER HONOR [DEATHSTALKER 4]  P
GREEN, SIMON R.  DEATHSTALKER LEGACY [DEATHSTALKER]  H
GREEN, SIMON R.  DEATHSTALKER REBELLION [DEATHSTALKER 2]  C/P,P
GREEN, SIMON R.  DEATHSTALKER RETURN [DEATHSTALKER]  P
GREEN, SIMON R.  DEATHSTALKER WAR [DEATHSTALKER 3]  C/P,P
GREEN, SIMON R.  DOWN AMONG THE DEAD MEN  C/P,P
GREEN, SIMON R.  DRINKING MIDNIGHT WINE  LP
GREEN, SIMON R.  EVERYBODY COMES TO THE NIGHTSIDE [NIGHTSIDE 1,2,3]  H
GREEN, SIMON R.  GHOSTWORLD  P
GREEN, SIMON R.  GOD KILLER, THE [HAWK & FISHER 3]  C/P:2,P
GREEN, SIMON R.  GUARD AGAINST DISHONOR [HAWK & FISHER 5]  C/P:2,P
GREEN, SIMON R.  HAWK & FISHER [HAWK & FISHER 1]  P
GREEN, SIMON R.  HELL TO PAY [NIGHTSIDE 7]  P
GREEN, SIMON R.  HELLWORLD  P
GREEN, SIMON R.  HEX AND THE CITY [NIGHTSIDE 4]  P
GREEN, SIMON R.  MAN WITH THE GOLDEN TORC, THE  H
GREEN, SIMON R.  MISTWORLD  C/P:2,P
GREEN, SIMON R.  NIGHTINGALE’S LAMENT [NIGHTSIDE 3]  P
GREEN, SIMON R.  PATHS NOT TAKEN [NIGHTSIDE 5]  P
GREEN, SIMON R.  SHADOWS FALL  C/P,P
GREEN, SIMON R.  SHARPER THAN A SERPENT’S TOOTH [NIGHTSIDE 6]  P
GREEN, SIMON R.  SOMETHING FROM THE NIGHTSIDE [NIGHTSIDE 1]  C/P,P
GREEN, SIMON R.  SWORDS OF HAVEN [@HAWK & FISHER 1,2,3]  P
GREEN, SIMON R.  TWILIGHT OF THE EMPIRE [DEATHSTALKER]  C/P,P
GREEN, SIMON R.  WINNER TAKES ALL [HAWK & FISHER 2]  P
GREEN, SIMON R.  WOLF IN THE FOLD [HAWK & FISHER 4]  C/P:2,P
GREEN, TERENCE M.  BLUE LIMBO  H,P
GREENBERG, DAN  WITCHFIRES OF LETH, THE  SFWA-P
GREENBERG, MARTIN  COMING ATTRACTIONS  HA
GREENBERG, MARTIN  FIVE SCIENCE FICTION NOVELS  HA
GREENBERG, MARTIN  MEN AGAINST THE STARS  PA,C/PA
GREENBERG, MARTIN  ROBOT AND THE MAN, THE  HA
GREENBERG, MARTIN H.  AFTER THE KING  C/PA,PA
GREENBERG, MARTIN H.  ALL HELL BREAKING LOOSE  PA
GREENBERG, MARTIN H.  AMAZING STORIES VISIONS OF OTHER WORLDS  LPA

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

ANIMAL BRIGADE 3000

PA

GREENBERG, MARTIN H./DAVIS, RUSSELL  APPRENTICE FANTASTIC  C/PA,PA

GREENBERG, MARTIN H./ POTTER, ALEXANDER ASSASSIN FANTASTIC

PA

GREENBERG, MARTIN H./SEGRIFF, LARRY  BATTLE MAGIC  C/PA,PA
GREENBERG, MARTIN H./HELFERS, JOHN  BLACK CATS AND BROKEN MIRRORS  C/PA,PA
GREENBERG, MARTIN H./GORMAN, ED  CAT CRIMES 2  PA
GREENBERG, MARTIN H.  CELEBRITY VAMPIRES  PA
GREENBERG, MARTIN H./HUGHES, KERRIE  CHILDREN OF MAGIC  PA
GREENBERG, MARTIN H.  CHRISTMAS ON GANYMEDE  C/PA,PA
GREENBERG, MARTIN H.  CIVIL WAR FANTASTIC  PA

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

COMMANDO BRIGADE 3000

PA

GREENBERG, MARTIN H./URBAN, SCOTT H.  CONSPIRACY FILES, THE  PA

The MITSFS Pinkdex  September 24, 2007  Page 169
AUTHOR  TITLE  CODE

GREENBERG, MARTIN H./GORMAN, ED/ MUNSTER, BILL DEAN KOONTZ COMPANION, THE

SFWA

GREENBERG, MARTIN H.  DIPLOMACY GUILD, THE [@ISAAC’S UNIVERSE #1]  C/PA,PA
GREENBERG, MARTIN H.  DRACULA: PRINCE OF DARKNESS  PA

GREENBERG, MARTIN H./ GREENBERG, ROSALIND M.

DRAGON FANTASTIC

C/PA,PA

GREENBERG, MARTIN H.  ELF FANTASTIC  PA
GREENBERG, MARTIN H.  ELF MAGIC  C/PA,PA
GREENBERG, MARTIN H./DAVIS, RUSSELL  FAERIE TALES  PA
GREENBERG, MARTIN H.  FANTASTIC ADVENTURES OF ROBIN HOOD, THE  C/PA,PA
GREENBERG, MARTIN H.  FANTASTIC LIVES  R/REF-H

GREENBERG, MARTIN H./ KOREN, BRITTANY A.

FANTASY GONE WRONG

PA

GREENBERG, MARTIN H./SEGRIFF, LARRY  FAR FRONTIERS  C/PA,PA
GREENBERG, MARTIN H.  FOUNDATION’S FRIENDS [FOUNDATION]  C/PA,SFWA-P,LPA
GREENBERG, MARTIN H.  FRANKENSTEIN: THE MONSTER WAKES  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF BATMAN, THE [DC COMICS]  C/PA,PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF BATMAN 2, THE [DC COMICS]  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF BATMAN 3, THE [DC COMICS]  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF SUPERMAN, THE [DC COMICS]  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF THE JOKER, THE [DC COMICS]  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF WONDER WOMAN, THE [DC COMICS]  PA
GREENBERG, MARTIN H.  FURTHER ADVENTURES OF XENA WARRIOR PRINCESS [XENA]  PA
GREENBERG, MARTIN H./HELFERS, JOHN  FUTURE CRIMES  C/PA,PA
GREENBERG, MARTIN H./SEGRIFF, LARRY  FUTURE NET  PA
GREENBERG, MARTIN H./SEGRIFF, LARRY  FUTURE WARS  PA
GREENBERG, MARTIN H.  GATEWAYS  PA
GREENBERG, MARTIN H./TIER, MARK  GIVE ME LIBERTY  PA
GREENBERG, MARTIN H./SEGRIFF, LARRY  GUARDSMEN OF TOMORROW  C/PA,PA
GREENBERG, MARTIN H./DAVIS, RUSSELL  HAUNTED HOLIDAYS  PA

GREENBERG, MARTIN H./ GREENBERG, ROSALIND M.

HORSE FANTASTIC

PA,C/PA

GREENBERG, MARTIN H./DAVIS, RUSSELL  IF I WERE AN EVIL OVERLORD  PA
GREENBERG, MARTIN H./HELFERS, JOHN  IN THE SHADOW OF EVIL  PA
GREENBERG, MARTIN H./HELFERS, JOHN  KNIGHT FANTASTIC  PA
GREENBERG, MARTIN H./HELFERS, JOHN  LITTLE RED RIDING HOOD IN THE BIG BAD CITY  PA
GREENBERG, MARTIN H.  LORD OF THE FANTASTIC  HA,LPA
GREENBERG, MARTIN H./PACK, JANET  MAGIC TAILS  PA
GREENBERG, MARTIN H./HUGHES, KERRIE  MAIDEN, MATRON, CRONE  C/PA,PA
GREENBERG, MARTIN H.  MERLIN  PA
GREENBERG, MARTIN H./DAVIS, RUSSELL  MILLENNIUM 3001  PA

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

MONSTER BRIGADE 3000

PA

GREENBERG, MARTIN H./HELFERS, JOHN  MUTANT FILES, THE  PA
GREENBERG, MARTIN H.  MY FAVORITE FANTASY STORY  PA
GREENBERG, MARTIN H.  MY FAVORITE SCIENCE FICTION STORY  PA
GREENBERG, MARTIN H.  NEBULA AWARD-WINNING NOVELLAS  LPA
GREENBERG, MARTIN H.  NEW HUGO WINNERS 4, THE  C/PA,PA

GREENBERG, MARTIN H./ ARTHURS, BRUCE D.

OLYMPUS

PA

GREENBERG, MARTIN H./SEGRIFF, LARRY  PAST IMPERFECT  PA

GREENBERG, MARTIN H./ GREENBERG, ROSALIND M.

PHANTOMS

PA

GREENBERG, MARTIN H./ KOREN, BRITTANY A.

PHAROAH FANTASTIC

PA

GREENBERG, MARTIN H.  PHASES IN CHAOS [@ISAAC’S UNIVERSE #2]  PA

GREENBERG, MARTIN H./ WARRICK, PATRICIA S.

POLITICAL SCIENCE FICTION

LPA

GREENBERG, MARTIN H./ OLANDER, JOSEPH D./ WARRICK, PATRICIA S.

RUN TO STARLIGHT

HA

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

SCIENCE FICTION OF THE 50’S

HA,LPA

GREENBERG, MARTIN H./SEGRIFF, LARRY  SILICON DREAMS  PA

GREENBERG, MARTIN H./ KOREN, BRITTANY A.

SINGLE WHITE VAMPIRE SEEKS SAME

PA

GREENBERG, MARTIN H./ POTTER, ALEXANDER SIRIUS THE DOG STAR

PA

GREENBERG, MARTIN H./HELFERS, JOHN  SLIPSTREAMS  PA
GREENBERG, MARTIN H./HELFERS, JOHN  SPACE STATIONS  PA
GREENBERG, MARTIN H./SEGRIFF, LARRY  SPELL FANTASTIC  C/PA,PA
GREENBERG, MARTIN H./HELFERS, JOHN  STAR COLONIES  C/PA,PA
GREENBERG, MARTIN H.  STARFALL [STAR*DRIVE]  PA

GREENBERG, MARTIN H./ SCHIMEL, LAWRENCE TAROT FANTASTIC C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 170
AUTHOR  TITLE  CODE
GREENBERG, MARTIN H.  TASTE FOR BLOOD, A  LPA

GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

TOMORROW, INC.

HA

GREENBERG, MARTIN H./ MATHESON, RICHARD/ WAUGH, CHARLES G.

TWILIGHT ZONE: THE ORIGINAL STORIES, THE

HA,LPA

GREENBERG, MARTIN H.  UFO FILES, THE  PA
GREENBERG, MARTIN H.  UNNATURAL DIPLOMACY [@ISAAC’S UNIVERSE #3]  PA,SFWA-P
GREENBERG, MARTIN H.  VAMPIRE DETECTIVES  C/PA,PA

GREENBERG, MARTIN H./ WAUGH, CHARLES G.

VAMPS

C/PA,PA

GREENBERG, MARTIN H./HELFERS, JOHN  VILLAINS VICTORIOUS  C/PA,PA
GREENBERG, MARTIN H./HELFERS, JOHN  WARRIOR FANTASTIC  C/PA,PA
GREENBERG, MARTIN H.  WHITE HOUSE HORRORS  PA
GREENBERG, MARTIN H.  WIZARD FANTASTIC  PA

GREENBERG, MARTIN H./ KOREN, BRITTANY A.

YOU BET YOUR PLANET

PA

GREENBERG, MARTIN H./MORLAN, A. R.  ZODIAC FANTASTIC  PA

GREENBERG, ROSALIND M./ GREENBERG, MARTIN H.

CHRISTMAS BESTIARY

PA,C/PA

GREENBERGER, ROBERT  DOORS INTO CHAOS [STAR TREK GATEWAYS #3/STAR TREK TNG]  P
GREENBERGER, ROBERT  ROMULAN STRATAGEM, THE [STAR TREK TNG]  C/P:2,P
GREENBERGER, ROBERT  TIME TO HATE, A [@STAR TREK TNG TIME TO 6]  P
GREENBERGER, ROBERT  TIME TO LOVE, A [@STAR TREK TNG TIME TO 5]  P
GREENE, JOSEPH  CAPTIVES IN SPACE [@DIG ALLEN #2]  P,@H
GREENE, JOSEPH  FORGOTTEN STAR, THE [DIG ALLEN #1]  P,@H
GREENE, JOSEPH  JOURNEY TO JUPITER [@DIG ALLEN #3]  H
GREENE, JOSEPH  ROBOTS OF SATURN [@DIG ALLEN #5]  H
GREENE, JOSEPH  TRAPPERS OF VENUS [@DIG ALLEN #4]  H
GREENE, KIRBY  BROTHERHOOD OF THE STARS  P
GREENER, LESLIE  MOON AHEAD  H
GREENER, LESLIE/HUTCHINSON, JOHN  MOON AHEAD (CONDENSED)  SR-P
GREENFIELD, IRVING A.  ANCIENT OF DAYS, THE  P
GREENFIELD, IRVING A.  FACE OF HIM, THE  P
GREENFIELD, IRVING A.  GODS’ TEMPTRESS, THE  C/P,P
GREENFIELD, IRVING A.  JULIUS CAESAR IS ALIVE AND WELL  P
GREENFIELD, IRVING A.  OTHERS, THE  C/P:2,P
GREENFIELD, IRVING A.  STAR TRIAL  P
GREENFIELD, IRVING A.  STARS WILL JUDGE, THE  P
GREENFIELD, IRVING A.  WATERS OF DEATH  C/P:2,P
GREENHALGH, ZOHRA  CONTRARYWISE [CONTRARYWISE 1]  P
GREENHALGH, ZOHRA  TRICKSTER’S TOUCH [CONTRARYWISE 2]  C/P,P
GREENHOUGH, TERRY  TIME AND TIMOTHY GRENVILLE  SR-P
GREENLAND, COLIN  HARM’S WAY  C/P:2,SFWA-P,P
GREENLAND, COLIN  MOTHER OF PLENTY [PLENTY 3]  P
GREENLAND, COLIN  SEASONS OF PLENTY [PLENTY 2]  C/P,P
GREENLAND, COLIN  TAKE BACK PLENTY [PLENTY 1]  C/P:2,P
GREENLEAF, WILLIAM  CLARION  C/P,P
GREENLEAF, WILLIAM  PANDORA STONE, THE  P
GREENLEAF, WILLIAM  STARJACKED!  P
GREENLEAF, WILLIAM  TARTARUS INCIDENT, THE  P
GREENLEAF, WILLIAM  TIMEJUMPER  P
GREENO, GAYLE  EXILES’ RETURN [@GHATTI’S TALE #3]  C/P,P
GREENO, GAYLE  FARTHEST SEEKING, THE [@GHATTENS’ GAMBIT #2]  C/P,P
GREENO, GAYLE  FINDER’S SEEKERS [@GHATTI’S TALE #1]  C/P,P
GREENO, GAYLE  MIND SNARE  C/P,P
GREENO, GAYLE  MIND-SPEAKERS’ CALL [@GHATTI’S TALE #2]  C/P,SFWA-P,P
GREENO, GAYLE  SUNDERLIES SEEKING [@GHATTENS’ GAMBIT #1]  P
GREENWOOD, ED  ALL SHADOWS FLED [@FORGOTTEN REALMS SHADOW OF AVATAR #3]  P
GREENWOOD, ED  BEST OF THE REALMS, BOOK II, THE [@FORGOTTEN REALMS]  P
GREENWOOD, ED/CUNNINGHAM, ELAINE  CITY OF SPLENDORS, THE [@FORGOTTEN REALMS CITIES]  P
GREENWOOD, ED  CLOAK OF SHADOWS [@FORGOTTEN REALMS SHADOW OF AVATAR #2]  P
GREENWOOD, ED/GRUBB, JEFF  CORMYR [@FORGOTTEN REALMS CORMYR #1]  P
GREENWOOD, ED  CROWN OF FIRE [@FORGOTTEN REALMS HARPERS #9/SHANDRIL 2]  C/P:2,P
GREENWOOD, ED/DENNING, TROY  DEATH OF THE DRAGON [@FORGOTTEN REALMS CORMYR #3]  P
GREENWOOD, ED  DRAGON’S ASCENSION, A [BAND OF FOUR]  C/P,P
GREENWOOD, ED  DRAGON’S DOOM [BAND OF FOUR]  H
GREENWOOD, ED  ELMINSTER IN HELL [@FORGOTTEN REALMS ELMINSTER]  P
GREENWOOD, ED  ELMINSTER IN MYTH DRANNOR [@FORGOTTEN REALMS ELMINSTER]  H

GREENWOOD, ED ELMINSTER: THE MAKING OF A MAGE [@FORGOTTEN REALMS ELMINSTER]

C/P,P,H

GREENWOOD, ED  ELMINSTER’S DAUGHTER [@FORGOTTEN REALMS ELMINSTER]  H
GREENWOOD, ED  HAND OF FIRE [@FORGOTTEN REALMS SHANDRIL #3]  P
GREENWOOD, ED  KINGLESS LAND, THE [BAND OF FOUR]  P

AUTHOR  TITLE  CODE
GREENWOOD, ED  SHADOWS OF DOOM [@FORGOTTEN REALMS SHADOW OF AVATAR #1]  P
GREENWOOD, ED  SILVERFALL [@FORGOTTEN REALMS]  P
GREENWOOD, ED  SPELLFIRE [FORGOTTEN REALMS SHANDRIL 1]  C/P:2,P
GREENWOOD, ED  STORMLIGHT [@FORGOTTEN REALMS HARPERS #14]  P

GREENWOOD, ED SWORDS OF EVENINGSTAR

[@FORGOTTEN REALMS KNIGHTS OF MYTH DRANNOR #1]

H

GREENWOOD, ED  TEMPTATION OF ELMINSTER, THE [@FORGOTTEN REALMS]  P
GREENWOOD, ED  VACANT THRONE, THE [BAND OF FOUR]  H,P
GREER, GERY/RUDDICK, BOB  MAX AND ME AND THE TIME MACHINE  P
GREG, PERCY  ACROSS THE ZODIAC  LP
GREGORIAN, JOYCE BALLOU  BROKEN CITADEL, THE [BROKEN CITADEL 1]  C/P,P
GREGORIAN, JOYCE BALLOU  CASTLEDOWN [BROKEN CITADEL 2]  P
GREGORIAN, JOYCE BALLOU  GREAT WHEEL, THE [BROKEN CITADEL 3]  SFWA-P
GREGORY, GUY  HEROES OF ZARA KEEP  P
GREGORY, JOHN  LEGACY OF THE STARS  P
GREGORY, STEPHEN  CORMORANT, THE  P
GRENDON, STEPHEN  MR. GEORGE AND OTHER ODD PERSONS  SR-H
GRESSMAN, THOMAS S.  DAGGER POINT [@BATTLETECH]  P
GRESSMAN, THOMAS S.  DYING TIME, THE [@MECHWARRIOR]  P
GRESSMAN, THOMAS S.  HUNTERS, THE [@BATTLETECH #35/TWILIGHT OF THE CLANS 3]  C/P,P
GRESSMAN, THOMAS S.  OPERATION SIERRA-75 [@VOR THE MAELSTROM]  P
GRESSMAN, THOMAS S.  SHADOWS OF WAR [@BATTLETECH #38/TWILIGHT OF THE CLANS 6]  P
GRESSMAN, THOMAS S.  SWORD AND FIRE [@BATTLETECH #37/TWILIGHT OF THE CLANS 5]  C/P,P
GREY, CHARLES  EXTRA MAN, THE  SR-P
GRIBBIN, JOHN/CHOWN, MARCUS  DOUBLE PLANET  C/P,P
GRIBBIN, JOHN  FATHER TO THE MAN  P,SFWA-P
GRIBBIN, JOHN  INNERVISIONS  P
GRIBBIN, JOHN/CHOWN, MARCUS  REUNION  H
GRIDBAN, VOLSTED  DEBRACY’S DRUG  SR-P
GRIDBAN, VOLSTED  DYNO-DEPRESSANT, THE  SR-P
GRIDBAN, VOLSTED  FROZEN LIMIT  SR-P
GRIDBAN, VOLSTED  LONELY ASTRONOMER, THE  SR-P
GRIDBAN, VOLSTED  MAGNETIC BRAIN  SR-P
GRIDBAN, VOLSTED  MOONS FOR SALE  SR-P
GRIDBAN, VOLSTED  PLANETOID DISPOSALS LTD.  SR-P
GRIDBAN, VOLSTED  PURPLE WIZARD, THE  SR-P
GRIFFEN, ELIZABETH L.  SHAGGY DOG, THE  P
GRIFFIN, ERIC  TZIMISCE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
GRIFFIN, P. M.  CALL TO ARMS [STAR COMMANDOS 9]  C/P,P
GRIFFIN, P. M.  COLONY IN PERIL [STAR COMMANDOS 2]  P,SFWA-P
GRIFFIN, P. M.  DEATH PLANET [STAR COMMANDOS 4]  P
GRIFFIN, P. M.  FIRE PLANET [STAR COMMANDOS 7]  P
GRIFFIN, P. M.  JUNGLE ASSAULT [STAR COMMANDOS 8]  P
GRIFFIN, P. M.  MIND SLAVER [STAR COMMANDOS 5]  P
GRIFFIN, P. M.  MISSION UNDERGROUND [STAR COMMANDOS 3]  C/P:2,P
GRIFFIN, P. M.  RETURN TO WAR [STAR COMMANDOS 6]  C/P,P
GRIFFIN, P. M.  STAR COMMANDOS [STAR COMMANDOS 1]  C/P:2,P
GRIFFIN, PENI R.  DIG IN TIME, A  SFWA
GRIFFIN, PENI R.  SWITCHING WELL  SFWA
GRIFFIN PUBLISHING  MACHINE-GOD LAUGHS, THE  HA
GRIFFIN, RUSSELL M.  BLIND MEN AND THE ELEPHANT, THE  P,SFWA-P
GRIFFIN, RUSSELL M.  CENTURY’S END  P
GRIFFIN, RUSSELL M.  MAKESHIFT GOD, THE  C/P:2,P
GRIFFIN, RUSSELL M.  TIMESERVERS, THE  P
GRIFFITH, GEORGE  ANGEL OF THE REVOLUTION, THE  LP
GRIFFITH, GEORGE  OLGA ROMANOFF  LP
GRIFFITH, GEORGE  RAID OF "LE VENGEUR", THE  SR-LP
GRIFFITH, GEORGE  STORIES OF OTHER WORLDS & HONEYMOON IN SPACE, A  LP
GRIFFITH, GEORGE  VALDAR THE OFT-BORN  H
GRIFFITH, KATHRYN MEYER  VAMPIRE BLOOD  P
GRIFFITH, NICOLA  AMMONITE  C/P:2,P
GRIFFITH, NICOLA/PAGEL, STEPHEN  BENDING THE LANDSCAPE: FANTASY  LPA
GRIFFITH, NICOLA/PAGEL, STEPHEN  BENDING THE LANDSCAPE: SCIENCE FICTION  HA,LPA
GRIFFITH, NICOLA  SLOW RIVER  H,LP
GRIFFITH, NICOLA  WITH HER BODY  LP
GRIFFITH, WILLIAM JR.  FANTASIES TWO  H
GRIFFITHS, ALAN  STRANGE NEWS FROM HEAVEN  H
GRIMES, LEE  DINOSAUR NEXUS  P
GRIMES, LEE  RETRO LIVES  P,SFWA-P
GRIMSLEY, JIM  LAST GREEN TREE, THE  H
GRIMSON, TODD  BRAND NEW CHERRY FLAVOR  H,LP
GRIMSON, TODD  STAINLESS  C/P,SFWA,LP
GRIMWADE, PETER  MAWDRYN UNDEAD [DOCTOR WHO]  SR-P,C/P:2
GRIMWADE, PETER  PLANET OF FIRE [DOCTOR WHO]  SR-P,C/P

The MITSFS Pinkdex  September 24, 2007  Page 172
AUTHOR  TITLE  CODE
GRIMWADE, PETER  TIME-FLIGHT [DOCTOR WHO]  SR-P,C/P
GRIMWOOD, JON COURTENAY  EFFENDI [EL ISKANDRYIA 2]  LP
GRIMWOOD, JON COURTENAY  FELAHEEN [EL ISKANDRYIA 3]  H
GRIMWOOD, JON COURTENAY  LUCIFER’S DRAGON  LP
GRIMWOOD, JON COURTENAY  PASHAZADE [EL ISKANDRYIA 1]  H
GRIMWOOD, JON COURTENAY  REDROBE  H
GRIMWOOD, JON COURTENAY  REMIX  LP
GRIMWOOD, KEN  REPLAY  C/P,SFWA,P,H
GRINNELL, DAVID  ACROSS TIME  DD286,P,C/P:2
GRINNELL, DAVID/CARTER, LIN  DESTINATION: SATURN  DH085
GRINNELL, DAVID  DESTINY’S ORBIT  H,DF161
GRINNELL, DAVID  EDGE OF TIME  C/P:2,DD362,P
GRINNELL, DAVID  MARTIAN MISSILE, THE  P,DD465
GRINNELL, DAVID  TO VENUS! TO VENUS!  D81610
GRODY, LEATHEL  FAITH OF THE UNFORGOTTEN [@FOUNDATIONS OF HOPE #1]  H
GROELL, ANNE LESLEY  ANVIL OF THE SUN [@CLOAK AND DAGGER #1]  P
GROELL, ANNE LESLEY  BRIDGE OF VALOR [@CLOAK AND DAGGER #2]  P
GROELL, ANNE LESLEY  CAULDRON OF INIQUITY [@CLOAK AND DAGGER #3]  P
GROSS, DAVE  BLACK WOLF [@FORGOTTEN REALMS SEMBIA]  P
GROSS, DAVE  LORD OF STORMWEATHER [@FORGOTTEN REALMS SEMBIA]  P

GROSS, DAVE OPPORTUNITY FOR PROFIT, AN

[@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #5]

P

GROSS, RICHARD  MANN AUS DEM ANDEREN JAHRTAUSEND, DER  FH
GROSSBACH, ROBERT  DEVIL AND MAX DEVLIN, THE  P
GROSSMAN, DAVE/FRANKOWSKI, LEO  TWO-SPACE WAR, THE  H
GROTEBOER, JEFFREY  EDGE OF INFINITY  P
GROTTA-KURSKA, DANIEL  J. R. R. TOLKIEN: ARCHITECT OF MIDDLE EARTH  R/REF-P,C/REF-P
GROVES, J. W.  SHELLBREAK  C/P:2,P
GRUBB, DAVIS  ANCIENT LIGHTS  LP
GRUBB, DAVIS  TWELVE TALES OF SUSPENSE AND THE SUPERNATURAL  P
GRUBB, JEFF  BROTHERS’ WAR, THE [@MAGIC: THE GATHERING ARTIFACTS 1]  P
GRUBB, JEFF  ETERNAL ICE, THE [@MAGIC: THE GATHERING ICE AGE 2]  P
GRUBB, JEFF  GATHERING DARK, THE [@MAGIC: THE GATHERING ICE AGE 1]  P
GRUBB, JEFF  LAST GUARDIAN, THE [@WARCRAFT #3]  P
GRUBB, JEFF  LIBERTY’S CRUSADE [@STARCRAFT #1]  P
GRUBB, JEFF  LORD TOEDE [DRAGONLANCE VILLAINS 5]  P
GRUBB, JEFF  SHATTERED ALLIANCE, THE [@MAGIC: THE GATHERING ICE AGE 3]  P
GRUBER, MICHAEL  TROPIC OF NIGHT  H
GRUEN, VON  MORTALS OF RENI, THE  SR-P,H
GRUNDY, STEPHAN  GILGAMESH  H
GRUNDY, STEPHAN  RHINEGOLD  H,P
GRUSSKIN, EDWARD  INVINCIBLE DOC SAVAGE, THE  B
GRYPHON, TALIA  KEY TO CONFLICT  P
GUERRIER, SIMON  LIFE WORTH LIVING, A [BERNICE SUMMERFIELD]  HA
GUERRIER, SIMON  TIME TRAVELLERS, THE [DOCTOR WHO PAST DOCTORS]  P
GUIEU, JIMMY  PLANETE EN PERIL  FP
GUIN, WYMAN  LIVING WAY OUT  C/P:2,P
GUIN, WYMAN  STANDING JOY, THE  C/P,P
GUNN, EILEEN  STABLE STRATEGIES AND OTHERS  LP
GUNN, JAMES E.  ALTERNATE WORLDS  R/REF-H,C/REF-H:3
GUNN, JAMES E.  BREAKING POINT  C/P,SFWA,P,H
GUNN, JAMES E.  BURNING, THE  C/P:2,P
GUNN, JAMES E.  CRISIS!  C/P,SFWA-P,P
GUNN, JAMES E.  DREAMERS, THE  H,SFWA
GUNN, JAMES E.  END OF THE DREAMS, THE  H:2
GUNN, JAMES E.  FUTURE IMPERFECT  C/P:2,P
GUNN, JAMES E.  FUTURO AL ROGO (ITALIAN)  FP
GUNN, JAMES E.  GIFT FROM THE STARS  LP
GUNN, JAMES E.  IMMORTAL, THE  P
GUNN, JAMES E.  IMMORTALS, THE  C/P:2,P
GUNN, JAMES E.  JOY MACHINE, THE [STAR TREK]  C/P,P
GUNN, JAMES E.  JOY MAKERS, THE  C/P:2,P
GUNN, JAMES E.  KAMPUS  P
GUNN, JAMES E.  LISTENERS, THE  H:2,P,SFWA
GUNN, JAMES E.  MAGICIANS, THE  C/P,SFWA,P,H
GUNN, JAMES E.  MIND MASTER, THE  P
GUNN, JAMES E.  NEBULA AWARD STORIES 10  C/PA,PA
GUNN, JAMES E.  NEW ENCYCLOPEDIA OF SCIENCE FICTION, THE  R/REF-H
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #1: FROM GILGAMESH TO WELLS, THE  C/PA,PA
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #2: FROM WELLS TO HEINLEIN, THE  C/PA,PA
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #3: FROM HEINLEIN TO HERE, THE  C/PA,PA
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #4: FROM HERE TO FOREVER, THE  C/PA,PA
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #5: THE BRITISH WAY, THE  LPA
GUNN, JAMES E.  ROAD TO SCIENCE FICTION #6: AROUND THE WORLD, THE  LPA

AUTHOR  TITLE  CODE
GUNN, JAMES E.  SOME DREAMS ARE NIGHTMARES  H,SFWA
GUNN, JAMES E.  STATION IN SPACE  C/P:2,P
GUNN, JAMES E.  TEMPO DI STREGHE  FP
GUNN, JAMES E.  THIS FORTRESS WORLD  H:2,P
GUNN, JAMES E.  THIS FORTRESS WORLD (ABRIDGED)  DD223
GUNN, JAMES E.  WITCHING HOUR, THE  P
GUNNARSSON, THORARINN  BATTLE OF THE RING [STARWOLVES 2]  P,SFWA-P
GUNNARSSON, THORARINN  DRAGONKING OF MYSTARA [@MYSTARA DRAGONLORD 2]  P
GUNNARSSON, THORARINN  DRAGONLORD OF MYSTARA [@MYSTARA DRAGONLORD 1]  P
GUNNARSSON, THORARINN  DRAGONMAGE OF MYSTARA [@MYSTARA DRAGONLORD 3]  P
GUNNARSSON, THORARINN  DRAGON’S DOMAIN  P
GUNNARSSON, THORARINN  DRAGONS ON THE TOWN [CALIFORNIA DRAGONS 3]  P
GUNNARSSON, THORARINN  HUMAN, BEWARE! [CALIFORNIA DRAGONS 2]  P
GUNNARSSON, THORARINN  MAKE WAY FOR DRAGONS! [CALIFORNIA DRAGONS 1]  C/P,P
GUNNARSSON, THORARINN  REVENGE OF THE VALKYRIE [SONG OF THE DWARVES 2]  C/P,P
GUNNARSSON, THORARINN  SONG OF THE DWARVES [SONG OF THE DWARVES 1]  P
GUNNARSSON, THORARINN  STARWOLVES, THE [STARWOLVES 1]  C/P,P
GUNNARSSON, THORARINN  STARWOLVES DREADNOUGHT [STARWOLVES 4]  P
GUNNARSSON, THORARINN  TACTICAL ERROR [STARWOLVES 3]  P
GUNTHER, MAX  DOOM WIND (PROOF)  SR-LP
GUON, ELLEN  BEDLAM BOYZ [ERIC BANYON]  C/P:2,SFWA-P,P
GUPTILL, PAUL/DRUMM, CHRIS  MANY WORLDS OF LARRY NIVEN, THE  R/REF-P
GURNEY, DAVID  EVIL UNDER THE WATER, THE  P
GURNEY, DAVID  ’F’ CERTIFICATE, THE  H
GURNEY, JAMES  DINOTOPIA CALENDAR, THE (1994)  CAL
GUTTENBERG, ELYSE  DAUGHTER OF THE SHAMAN  P
GUTTENBERG, ELYSE  SUMMER LIGHT  P
GUTTENBERG, ELYSE  SUNDER, ECLIPSE & SEED  C/P,P
GUTTERIDGE, LINDSAY  COLD WAR IN A COUNTRY GARDEN [MATTHEW DILKE 1]  C/P:2,P
GUTTERIDGE, LINDSAY  KILLER PINE [MATTHEW DILKE 2]  H,P
GYGAX, GARY  ANUBIS MURDERS, THE [@DANGEROUS JOURNEYS 1]  C/P,SFWA-P,P
GYGAX, GARY  ARTIFACT OF EVIL [@GREYHAWK #2]  C/P:2,P
GYGAX, GARY  CITY OF HAWKS [GORD THE ROGUE]  C/P:2,SFWA-P,P
GYGAX, GARY  COME ENDLESS DARKNESS [GORD THE ROGUE]  C/P:2,P
GYGAX, GARY  DANCE OF DEMONS [GORD THE ROGUE]  C/P,P
GYGAX, GARY  DEATH IN DELHI [@DANGEROUS JOURNEYS 3]  P
GYGAX, GARY  NIGHT ARRANT [GORD THE ROGUE]  C/P:2,P
GYGAX, GARY  SAGA OF OLD CITY [@GREYHAWK #1]  C/P:2,P
GYGAX, GARY  SAMARKAND SOLUTION, THE [@DANGEROUS JOURNEYS 2]  C/P,P
GYGAX, GARY  SEA OF DEATH [GORD THE ROGUE]  C/P:2,P

AUTHOR  TITLE  CODE
HABER, KAREN  BLESS THE BEASTS [STAR TREK VOYAGER]  P
HABER, KAREN  CROSSING INFINITY  H
HABER, KAREN/STRAHAN, JONATHAN  FANTASY THE BEST OF 2004  PA
HABER, KAREN  MUTANT LEGACY [MUTANT SEASON 4]  P
HABER, KAREN  MUTANT PRIME, THE [MUTANT SEASON 2]  C/P:2,LP
HABER, KAREN/SILVERBERG, ROBERT  MUTANT SEASON, THE [MUTANT SEASON 1]  C/P,LP,P
HABER, KAREN  MUTANT STAR [MUTANT SEASON 3]  P
HABER, KAREN/STRAHAN, JONATHAN  SCIENCE FICTION THE BEST OF 2003  PA
HABER, KAREN/STRAHAN, JONATHAN  SCIENCE FICTION THE BEST OF 2004  PA
HABER, KAREN  SISTER BLOOD  P
HABER, KAREN  THIEVES’ CARNIVAL  TD22,SFWA-TD22
HABER, KAREN  WAR MINSTRELS, THE  P
HABER, KAREN  WOMAN WITHOUT A SHADOW  P
HABER, MELISSA GLENN  BEYOND THE DRAGON PORTAL  H
HACKETT, JOHN, ET AL.  THIRD WORLD WAR: AUGUST 1985, THE  P
HADDIX, MARGARET PETERSON  RUNNING OUT OF TIME  LP
HADER, BERTA/HADER, ELMER  SKYROCKET, THE  H
HADFIELD, JOHN  CHAMBER OF HORRORS, A  HA
HADINGHAM, EVAN  CIRCLES AND STANDING STONES  SFWA
HADLEY, ARTHUR T.  JOY WAGON, THE  C/P,P
HADLEY, FRANKLIN  PLANET BIG ZERO  C/P:2,P
HAGBERG, DAVID  LAST COME THE CHILDREN  SFWA-P
HAGBERG, DAVID  TERMINATOR 3: RISE OF THE MACHINES [TERMINATOR]  P
HAGE, RICHARD  ANCIENT DESIGNS  LP
HAGGARD, H. RIDER  ALLAN’S WIFE [ALLAN QUATERMAIN]  LP
HAGGARD, H. RIDER  AYESHA: THE RETURN OF SHE [AYESHA 2]  C/P,P
HAGGARD, H. RIDER  CLEOPATRA  C/P,P
HAGGARD, H. RIDER  ERIC BRIGHTEYES [ERIC BRIGHTEYES 1]  C/P,LP
HAGGARD, H. RIDER  HEART OF THE WORLD  LP
HAGGARD, H. RIDER  KING SOLOMON’S MINES [ALLAN QUATERMAIN]  C/P,P,H
HAGGARD, H. RIDER  MARION ISLE  H
HAGGARD, H. RIDER  MONTEZUMA’S DAUGHTER  H
HAGGARD, H. RIDER  MORNING STAR  C/P,P
HAGGARD, H. RIDER  NADA THE LILY  LP
HAGGARD, H. RIDER  PEOPLE OF THE MIST, THE  C/P:2,P
HAGGARD, H. RIDER  QUEEN OF THE DAWN  H
HAGGARD, H. RIDER  QUEEN SHEBA’S RING  SR-L
HAGGARD, H. RIDER  RED EVE  H
HAGGARD, H. RIDER  SHE [AYESHA 1]  H:2,P
HAGGARD, H. RIDER  SHE AND ALLAN [AYESHA 4/ALLAN QUATERMAIN]  LP
HAGGARD, H. RIDER  SPIRIT OF BAMBATSE, THE  LP
HAGGARD, H. RIDER  TALE OF THREE LIONS, A  H
HAGGARD, H. RIDER  3 ADVENTURE NOVELS  LP
HAGGARD, H. RIDER  WANDERER’S NECKLACE, THE  P
HAGGARD, H. RIDER  WHEN THE WORLD SHOOK  C/P,P
HAGGARD, H. RIDER  WISDOM’S DAUGHTER [AYESHA 3]  P
HAGGARD, H. RIDER  WITCH’S HEAD, THE  H
HAGGARD, H. RIDER/LANG, ANDREW  WORLD’S DESIRE, THE  P
HAGGARD, WILLIAM  SLOW BURNER  H
HAGGARD, WILLIAM  TELEMANN TOUCH, THE  H
HAGGERTY, JAMES J. JR.  SPACECRAFT  [C/RAN-P:2]
HAHN, STEVE  MINDWIPE!  SR-P
HAIBLUM, ISIDORE  CRYSTAL WORLD  P
HAIBLUM, ISIDORE  HAND OF GANZ, THE  P
HAIBLUM, ISIDORE  IDENTITY PLUNDERERS, THE  P
HAIBLUM, ISIDORE  INTERWORLD  C/P,P
HAIBLUM, ISIDORE  MUTANTS ARE COMING, THE  C/P,SFWA,P
HAIBLUM, ISIDORE  NIGHTMARE EXPRESS  C/P,P
HAIBLUM, ISIDORE  OUT OF SYNC  P
HAIBLUM, ISIDORE  OUTERWORLD  BS3
HAIBLUM, ISIDORE  RETURN, THE  P
HAIBLUM, ISIDORE  SPECTERWORLD  C/P,P
HAIBLUM, ISIDORE  TRANSFER TO YESTERDAY  C/P:2,SFWA,P
HAIBLUM, ISIDORE  TSADDIK OF THE SEVEN WONDERS, THE  P,SFWA
HAIBLUM, ISIDORE  WILK ARE AMONG US, THE  C/P,P,H
HAILEY, ARTHUR  IN HIGH PLACES  H
HAINING, PETER  ANCIENT MYSTERIES READER, THE  SFWA
HAINING, PETER  BEYOND THE CURTAIN OF DARK  PA
HAINING, PETER  DEAD OF NIGHT  HA
HAINING, PETER  DEADLY NIGHTSHADE  HA
HAINING, PETER  DOCTOR WHO: A CELEBRATION  R/REF-H,C/REF-H:2
HAINING, PETER  DOCTOR WHO: THE KEY TO TIME [DOCTOR WHO]  R/REF-H
HAINING, PETER  EVIL PEOPLE, THE  PA
HAINING, PETER  FANTASTIC PULPS, THE  LP
HAINING, PETER  FLYING SORCERERS, THE  PA

The MITSFS Pinkdex  September 24, 2007  Page 176
AUTHOR  TITLE  CODE
HAINING, PETER  FRANKENSTEIN OMNIBUS, THE  HA
HAINING, PETER  FREAK SHOW, THE  SR-PA
HAINING, PETER  FUTURE MAKERS, THE  HA,PA
HAINING, PETER  GHOULS, THE  HA,PA
HAINING, PETER  JULES VERNE COMPANION, THE  R/REF-H
HAINING, PETER  KNIGHTS OF MADNESS  PA
HAINING, PETER  LUCIFER SOCIETY, THE  PA
HAINING, PETER  MAMMOTH BOOK OF 20TH CENTURY GHOST STORIES, THE  LPA
HAINING, PETER  MIDNIGHT PEOPLE, THE  PA
HAINING, PETER  NECROMANCERS, THE  HA
HAINING, PETER  NIGHTMARE READER, THE  HA,SFWA
HAINING, PETER  SATANISTS, THE  PA
HAINING, PETER  TERROR!  A
HAINING, PETER  UNSPEAKABLE PEOPLE, THE  PA
HAINING, PETER  VINTAGE SCIENCE FICTION  LPA
HAINING, PETER  WEIRD TALES 1  SR-PA
HAINING, PETER  WITCHCRAFT READER, THE  HA
HAINING, PETER  WIZARDS OF ODD, THE  C/PA,PA
HAIRSTON, ANDREA  MINDSCAPE  LP
HALACY, D. S.  RETURN FROM LUNA  H
HALDEMAN, JACK C. II  FALL OF WINTER, THE  C/P,P
HALDEMAN, JACK C. II/DANN, JACK  HIGH STEEL  H,P
HALDEMAN, JACK C. II  PERRY’S PLANET [STAR TREK]  C/P,P
HALDEMAN, JACK C. II  VECTOR ANALYSIS  C/P,P,H
HALDEMAN, JOE  ALL MY SINS REMEMBERED  H:2,P

HALDEMAN, JOE/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

BODY ARMOR: 2000

PA,C/PA

HALDEMAN, JOE  BUYING TIME  H:2,P
HALDEMAN, JOE  CAMOUFLAGE  H
HALDEMAN, JOE  COMING, THE  C/P,P,H
HALDEMAN, JOE  DEALING IN FUTURES  H:2,P
HALDEMAN, JOE  FOREVER FREE  H,P
HALDEMAN, JOE  FOREVER PEACE  C/P,P,H
HALDEMAN, JOE  FOREVER WAR, THE  C/P,P,H
HALDEMAN, JOE/GREENBERG, MARTIN H.  FUTURE WEAPONS OF WAR  HA
HALDEMAN, JOE  GUARDIAN  H
HALDEMAN, JOE  HEMINGWAY HOAX, THE  H:2,P
HALDEMAN, JOE  INFINITE DREAMS  H:2,P
HALDEMAN, JOE  MINDBRIDGE  H:2,P
HALDEMAN, JOE  NEBULA AWARD STORIES 17  C/PA,PA
HALDEMAN, JOE  1968  P
HALDEMAN, JOE  NONE SO BLIND  H
HALDEMAN, JOE  OLD TWENTIETH  H
HALDEMAN, JOE  PLANET OF JUDGMENT [STAR TREK]  C/P:2,P
HALDEMAN, JOE  SEPARATE WAR AND OTHER STORIES, A  H

HALDEMAN, JOE/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SPACEFIGHTERS

PA,C/PA

HALDEMAN, JOE  STUDY WAR NO MORE  C/PA,PA

HALDEMAN, JOE/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SUPERTANKS

PA,C/PA,SFWA-P

HALDEMAN, JOE/HALDEMAN, JACK C. II  THERE IS NO DARKNESS  C/P:2,P
HALDEMAN, JOE  TOOL OF THE TRADE  SR-H,C/P:2,P
HALDEMAN, JOE  VIETNAM AND OTHER ALIEN WORLDS  SR-H
HALDEMAN, JOE  WORLD WITHOUT END [STAR TREK]  C/P:2,P
HALDEMAN, JOE  WORLDS [WORLDS 1]  C/P,SFWA,P,H
HALDEMAN, JOE  WORLDS APART [WORLDS 2]  C/P,P,SR-H,H
HALDEMAN, JOE  WORLDS ENOUGH AND TIME [WORLDS 3]  C/P,P,H
HALDEMAN, LINDA  ESBAE: A WINTER’S TALE  P
HALDEMAN, LINDA  LASTBORN OF ELVINWOOD, THE  P
HALE, DEBORAH  DESTINED QUEEN, THE [MAURA WOODBURY 2]  LP
HALE, DEBORAH  WIZARD’S WARD, THE [MAURA WOODBURY 1]  LP
HALE, F. J.  IN THE SEA NYMPH’S LAIR [SPELL WARS 2]  P
HALE, F. J.  OGRE CASTLE [SPELL WARS 1]  P
HALE, KELLY  ERASING SHERLOCK [FACTION PARADOX 5]  LP
HALE, MICHAEL  FOLD IN THE TENT OF THE SKY, A  H
HALES, E. E. Y.  CHARIOT OF FIRE  P
HALIFAX, LORD  LORD HALIFAX’S GHOST BOOK  [C/RAN-H,C/RAN-P]
HALL, ADAM  QUEEN IN DANGER  P
HALL, ADAM  SIBLING, THE  P
HALL, ANGUS  DEVILDAY  P
HALL, ANGUS  SCARS OF DRACULA, THE  P
HALL, AUSTIN/FLINT, HOMER EON  BLIND SPOT, THE  SR-H,C/P:2,P
HALL, AUSTIN  PEOPLE OF THE COMET  SR-H,H
HALL, AUSTIN  SPOT OF LIFE, THE  C/P,P
HALL, GIMONE  WITCH’S SUCKLING  P

AUTHOR  TITLE  CODE
HALL, HAL W.  SCIENCE FICTION BOOK REVIEW INDEX 1971  R/IND
HALL, HAL W.  SCIENCE FICTION BOOK REVIEW INDEX 1972 VOLUME 3  R/IND
HALL, HAL W.  SCIENCE FICTION BOOK REVIEW INDEX 1980 VOLUME 11  R/IND
HALL, JOHN RYDER  FUTUREWORLD  P
HALL, JOHN RYDER  SINBAD AND THE EYE OF THE TIGER  P
HALL, MARK EDWARD  LOST VILLAGE, THE  H
HALL, S. C.  AMULET, THE  SR-L
HALL, SANDI  GODMOTHERS, THE  LP
HALL, STEPHANIE  QUEEN OF COINS  P
HALLAHAN, WILLIAM H.  KEEPER OF THE CHILDREN  P
HALLAHAN, WILLIAM H.  SEARCH FOR JOSEPH TULLY, THE  P
HALLIDAY, MAGS L.  HISTORY 101 [DOCTOR WHO EIGHTH DOCTOR]  P
HALLIDAY, MAGS L.  WARRING STATES [@FACTION PARADOX 4]  LP
HALLMARK  JUST FOR YOU: THE WORLD’S FIRST SCIENCE FICTION BIRTHDAY CARD  WRD
HALLOWELL, MADELYNE  ANCESTRAL BLOOD  LP
HALLUMS, JAMES R.  THEY CAME, THEY SAW  SR-P
HALPERIN, JAMES L.  FIRST IMMORTAL, THE  H,P
HALPERIN, JAMES L.  TRUTH MACHINE, THE  H,P
HALPERN, JAY  JADE UNICORN, THE  P
HAMBLY, BARBARA  ARMIES OF DAYLIGHT, THE [DARWATH 3]  C/P:2,P
HAMBLY, BARBARA  BEAUTY AND THE BEAST [BEAUTY AND THE BEAST]  C/P,P
HAMBLY, BARBARA  BRIDE OF THE RAT GOD  C/P:2,P
HAMBLY, BARBARA  CHILDREN OF THE JEDI [STAR WARS]  H:2,P
HAMBLY, BARBARA  CIRCLE OF THE MOON [SISTERS OF THE RAVEN 2]  LP
HAMBLY, BARBARA  CROSSROAD [STAR TREK]  C/P:2,P
HAMBLY, BARBARA  DARK HAND OF MAGIC, THE [SUNWOLF AND STARHAWK 3]  C/P,P,H
HAMBLY, BARBARA  DARKMAGE [WINDROSE 1,2]  H:2
HAMBLY, BARBARA  DOG WIZARD [WINDROSE 3]  C/P,P,H
HAMBLY, BARBARA  DRAGONSBANE [WINTERLANDS 1]  C/P,P,H
HAMBLY, BARBARA  DRAGONSHADOW [WINTERLANDS 2]  H
HAMBLY, BARBARA  DRAGONSTAR [WINTERLANDS 4]  H
HAMBLY, BARBARA  GHOST-WALKER [STAR TREK]  C/P:3,P
HAMBLY, BARBARA  ICEFALCON’S QUEST [DARWATH 5]  C/P:2,P,H
HAMBLY, BARBARA  ISHMAEL [STAR TREK]  C/P:3,P
HAMBLY, BARBARA  KNIGHT OF THE DEMON QUEEN [WINTERLANDS 3]  P
HAMBLY, BARBARA  LADIES OF MANDRIGYN, THE [SUNWOLF AND STARHAWK 1]  C/P:3,P
HAMBLY, BARBARA  MAGICIANS OF NIGHT, THE [SUN-CROSS 2]  C/P:2,P
HAMBLY, BARBARA  MOTHER OF WINTER [DARWATH 4]  H:2,P
HAMBLY, BARBARA  PLANET OF TWILIGHT [@STAR WARS]  P
HAMBLY, BARBARA  RAINBOW ABYSS, THE [SUN-CROSS 1]  C/P:2,P
HAMBLY, BARBARA  SILENT TOWER, THE [WINDROSE 1]  C/P:2,P
HAMBLY, BARBARA  SILICON MAGE, THE [WINDROSE 2]  C/P:2,P
HAMBLY, BARBARA/GREENBERG, MARTIN H. SISTERS OF THE NIGHT  C/PA,LPA
HAMBLY, BARBARA  SISTERS OF THE RAVEN [SISTERS OF THE RAVEN 1]  H:2,LP
HAMBLY, BARBARA  SONG OF ORPHEUS [BEAUTY AND THE BEAST]  C/P:3,P
HAMBLY, BARBARA  STRANGER AT THE WEDDING  C/P:2,P
HAMBLY, BARBARA  SUN-CROSS [SUN-CROSS 1,2]  H:2
HAMBLY, BARBARA  THOSE WHO HUNT THE NIGHT [LONDON VAMPIRES 1]  C/P,P,H:2
HAMBLY, BARBARA  TIME OF THE DARK, THE [DARWATH 1]  C/P:2,P
HAMBLY, BARBARA  TRAVELING WITH THE DEAD [LONDON VAMPIRES 2]  C/P,P,H
HAMBLY, BARBARA  UNSCHOOLED WIZARD, THE [SUNWOLF AND STARHAWK 1,2]  H:2
HAMBLY, BARBARA  WALLS OF AIR, THE [DARWATH 2]  C/P:2,P
HAMBLY, BARBARA  WITCHES OF WENSHAR, THE [SUNWOLF AND STARHAWK 2]  C/P:2,P
HAMILTON, ALEX  SPLINTERS  SR-PA
HAMILTON, EDMOND  BATTLE FOR THE STARS  P,H:2,SFWA-TD8,TD8
HAMILTON, EDMOND  BEST OF EDMOND HAMILTON, THE  C/P,P,H:2
HAMILTON, EDMOND  BEYOND THE MOON  C/P,P
HAMILTON, EDMOND  CALLING CAPTAIN FUTURE  C/P,P
HAMILTON, EDMOND  CAPTAIN FUTURE AND THE SPACE EMPEROR  P
HAMILTON, EDMOND  CAPTAIN FUTURE’S CHALLENGE  C/P,P
HAMILTON, EDMOND  CITY AT WORLD’S END  SR-P,C/P:2,GN3.18,P
HAMILTON, EDMOND  CLOSED WORLDS, THE [@STARWOLF #2]  C/P,P
HAMILTON, EDMOND  COMET KINGS, THE  P
HAMILTON, EDMOND  CRASHING SUNS  C/P:2,P
HAMILTON, EDMOND  DOOMSTAR  C/P:2,P
HAMILTON, EDMOND  FUGITIVE OF THE STARS  DM111
HAMILTON, EDMOND  GALAXY MISSION  C/P,P
HAMILTON, EDMOND  HAUNTED STARS, THE  H:2,P
HAMILTON, EDMOND  INVASIONE DELLA GALASSIA, L  FP
HAMILTON, EDMOND  LUTA INTERGALACTICA  FP
HAMILTON, EDMOND  MAGICIAN OF MARS, THE  C/P:2,P
HAMILTON, EDMOND  OUTLAW WORLD  C/P:2,P
HAMILTON, EDMOND  OUTLAWS OF THE MOON  C/P:2,P
HAMILTON, EDMOND  OUTSIDE THE UNIVERSE  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 178
AUTHOR  TITLE  CODE
HAMILTON, EDMOND  PLANETS IN PERIL  C/P,P
HAMILTON, EDMOND  QUEST BEYOND THE STARS  P
HAMILTON, EDMOND  RETURN TO THE STARS  C/P:2,P
HAMILTON, EDMOND/BRACKETT, LEIGH  STAR AND THE STAR-KINGS  H
HAMILTON, EDMOND  STAR KINGS, THE  C/P,P,H
HAMILTON, EDMOND  STAR OF LIFE, THE  H:2,P
HAMILTON, EDMOND  STARWOLF [STARWOLF 1,2,3]  C/P:2,P
HAMILTON, EDMOND  SUN SMASHER, THE  DD351,P
HAMILTON, EDMOND  VALLEY OF CREATION, THE  C/P:2,P
HAMILTON, EDMOND  VAMPIRE MASTER, THE  H
HAMILTON, EDMOND  WEAPON FROM BEYOND, THE [@STARWOLF #1]  C/P:2,P
HAMILTON, EDMOND  WHAT’S IT LIKE OUT THERE? AND OTHER STORIES  P
HAMILTON, EDMOND  WORLD OF THE STARWOLVES [@STARWOLF #3]  C/P,P
HAMILTON, EDMOND  YANK AT VALHALLA, A  D93900
HAMILTON, JESSICA  ELIZABETH  P

HAMILTON, LAURELL K./ HARRIS, CHARLAINE/ DAVIDSON, MARYJANICE/ KNIGHT, ANGELA/TAYLOR, VICKIE BITE

P

HAMILTON, LAURELL K.  BLACK MOON INN [ANITA BLAKE 7,8]  H
HAMILTON, LAURELL K.  BLOODY BONES [ANITA BLAKE 5]  C/P:2,P
HAMILTON, LAURELL K.  BLUE MOON [ANITA BLAKE 8]  C/P:2,P
HAMILTON, LAURELL K.  BURNT OFFERINGS [ANITA BLAKE 7]  C/P:2,P
HAMILTON, LAURELL K.  CARESS OF TWILIGHT, A [MEREDITH GENTRY 2]  H,P
HAMILTON, LAURELL K.  CERULEAN SINS [ANITA BLAKE 11]  C/P,H
HAMILTON, LAURELL K.  CIRCUS OF THE DAMNED [ANITA BLAKE 3]  C/P:2,P
HAMILTON, LAURELL K.  CLUB VAMPYRE [ANITA BLAKE 1,2,3]  H

HAMILTON, LAURELL K./ DAVIDSON, MARYJANICE/WILKS, EILEEN/ YORK, REBECCA CRAVINGS

H,P

HAMILTON, LAURELL K.  DANSE MACABRE [ANITA BLAKE 14]  H:2
HAMILTON, LAURELL K.  DEATH OF A DARKLORD [@RAVENLOFT]  C/P,LP
HAMILTON, LAURELL K.  GUILTY PLEASURES [ANITA BLAKE 1]  C/P:2,P
HAMILTON, LAURELL K.  HARLEQUIN, THE [ANITA BLAKE 15]  H
HAMILTON, LAURELL K.  INCUBUS DREAMS [ANITA BLAKE 12]  H
HAMILTON, LAURELL K.  KILLING DANCE, THE [ANITA BLAKE 6]  C/P:2,P
HAMILTON, LAURELL K.  KISS OF SHADOWS, A [MEREDITH GENTRY 1]  H
HAMILTON, LAURELL K.  LAUGHING CORPSE, THE [ANITA BLAKE 2]  C/P:2,P
HAMILTON, LAURELL K.  LUNATIC CAFE, THE [ANITA BLAKE 4]  C/P:2,P
HAMILTON, LAURELL K.  MACAH [ANITA BLAKE 13]  P
HAMILTON, LAURELL K.  MIDNIGHT CAFE, THE [ANITA BLAKE 4,5,6]  H
HAMILTON, LAURELL K.  MISTRAL’S KISS [MEREDITH GENTRY 5]  H
HAMILTON, LAURELL K.  NARCISSUS IN CHAINS [ANITA BLAKE 10]  C/P,P,H
HAMILTON, LAURELL K.  NIGHTSEER  C/P:2,P
HAMILTON, LAURELL K.  NIGHTSHADE [STAR TREK TNG]  C/P:2,P
HAMILTON, LAURELL K.  OBSIDIAN BUTTERFLY [ANITA BLAKE 9]  C/P,P,H
HAMILTON, LAURELL K.  SEDUCED BY MOONLIGHT [MEREDITH GENTRY 3]  H
HAMILTON, LAURELL K.  STRANGE CANDY  H
HAMILTON, LAURELL K.  STROKE OF MIDNIGHT, A [MEREDITH GENTRY 4]  H
HAMILTON, PETER F.  CONFEDERATION HANDBOOK, THE [NIGHT’S DAWN]  P
HAMILTON, PETER F.  FALLEN DRAGON  H
HAMILTON, PETER F.  JUDAS UNCHAINED [COMMONWEALTH 2]  H
HAMILTON, PETER F.  LIGHTSTORM [WEB,THE]  P
HAMILTON, PETER F.  MINDSTAR RISING [GREG MANDEL 1]  C/P:2,P
HAMILTON, PETER F.  MISSPENT YOUTH  H
HAMILTON, PETER F.  NAKED GOD, THE [NIGHT’S DAWN 3]  H
HAMILTON, PETER F.  NANO FLOWER, THE [GREG MANDEL 3]  H
HAMILTON, PETER F.  NEUTRONIUM ALCHEMIST [NIGHT’S DAWN 2]  H
HAMILTON, PETER F.  NEUTRONIUM ALCHEMIST 1: CONSOLIDATION [NIGHT’S DAWN 2]  C/P,P
HAMILTON, PETER F.  NEUTRONIUM ALCHEMIST 2: CONFLICT [NIGHT’S DAWN 2]  P
HAMILTON, PETER F.  PANDORA’S STAR [COMMONWEALTH 1]  H
HAMILTON, PETER F.  QUANTUM MURDER, A [GREG MANDEL 2]  C/P,P,H
HAMILTON, PETER F.  REALITY DYSFUNCTION [NIGHT’S DAWN 1]  H
HAMILTON, PETER F.  REALITY DYSFUNCTION 1: EMERGENCE [NIGHT’S DAWN 1]  C/P:2,P
HAMILTON, PETER F.  REALITY DYSFUNCTION 2: EXPANSION [NIGHT’S DAWN 1]  C/P,P
HAMILTON, PETER F.  SECOND CHANCE AT EDEN, A  P
HAMILTON, PETER F.  WATCHING TREES GROW  GZD07305
HAMILTON, VIRGINIA  DUSTLAND [JUSTICE 2]  C/P,P
HAMILTON, VIRGINIA  GATHERING, THE [JUSTICE 3]  C/P,P
HAMILTON, VIRGINIA  JUSTICE AND HER BROTHERS [JUSTICE 1]  C/P,P
HAMMELL, IAN  BURNING GODDESS, THE [@SHADOW WORLD 1]  P
HAMMELL, IAN  CITY OF ASSASSINS [@SHADOW WORLD 3]  P
HAMMELL, IAN  CLOCK STRIKES SWORD [@SHADOW WORLD 2]  P
HAMMELL, IAN  STORMRIDERS [@SHADOW WORLD 4]  P

The MITSFS Pinkdex  September 24, 2007  Page 179
AUTHOR  TITLE  CODE
HAMMETT, DASHIELL  CREEPS BY NIGHT  C/PA,PA
HAMMOND AND COMPANY  OUR WORLD IN SPACE ATLAS  [C/RAN-H]
HAMMOND, WARREN  KOP [JUNO 1]  H
HAMPTON, KIRK  MOONHARE, THE  LP
HANCOCK, KAREN  LIGHT OF EIDON [GUARDIAN-KING 1]  LP
HANCOCK, KAREN  RETURN OF THE GUARDIAN-KING [GUARDIAN-KING 4]  LP
HANCOCK, KAREN  SHADOW OVER KIRIATH [GUARDIAN-KING 3]  LP
HANCOCK, KAREN  SHADOW WITHIN, THE [GUARDIAN-KING 2]  LP
HANCOCK, NIEL  ACROSS THE FAR MOUNTAIN [@WILDERNESS OF FOUR #1]  C/P,P
HANCOCK, NIEL  BRIDGE OF DAWN, THE [WINDAMEIR CIRCLE 4]  P
HANCOCK, NIEL  CALIX STAY [@CIRCLE OF LIGHT #3]  C/P,P
HANCOCK, NIEL  DRAGON WINTER  C/P:2,P
HANCOCK, NIEL  FARAGON FAIRINGAY [@CIRCLE OF LIGHT #2]  C/P:2,P
HANCOCK, NIEL  FIRES OF WINDAMEIR, THE [WINDAMEIR CIRCLE 1]  P
HANCOCK, NIEL  GREYFAX GRIMWALD [@CIRCLE OF LIGHT #1]  C/P:2,P
HANCOCK, NIEL  ON THE BOUNDARIES OF DARKNESS [@WILDERNESS OF FOUR #3]  P
HANCOCK, NIEL  PLAINS OF THE SEA, THE [@WILDERNESS OF FOUR #2]  C/P,P
HANCOCK, NIEL  ROAD TO THE MIDDLE ISLANDS, THE [@WILDERNESS OF FOUR #4]  P
HANCOCK, NIEL  SEA OF SILENCE, THE [WINDAMEIR CIRCLE 2]  P
HANCOCK, NIEL  SQUARING THE CIRCLE [@CIRCLE OF LIGHT #4]  C/P:2,P
HANCOCK, NIEL  WANDERER’S RETURN, A [WINDAMEIR CIRCLE 3]  P
HAND, ELIZABETH  AESTIVAL TIDE [WINTERLONG 2]  C/P,P
HAND, ELIZABETH  BLACK LIGHT  H
HAND, ELIZABETH  FRENCHMAN, THE [@MILLENIUM]  C/P,P
HAND, ELIZABETH  GLIMMERING  H,P
HAND, ELIZABETH  ICARUS DESCENDING  C/P,P
HAND, ELIZABETH  ICARUS DESCENDING (PROOF)  SR-LP
HAND, ELIZABETH  SAFFRON AND BRIMSTONE  LP
HAND, ELIZABETH  12 MONKEYS  P
HAND, ELIZABETH  WAKING THE MOON  H,P
HAND, ELIZABETH  WINTERLONG [WINTERLONG 1]  C/P:2,P
HANDLEY, MAX  MEANWHILE  LP
HANLEY, VICTORIA  SEER AND THE SWORD, THE  P

HANLON, MICHAEL SCIENCE OF THE HITCHHIKER’S GUIDE TO THE GALAXY, THE

[HITCHHIKERS]

R/REF-H

HANRATTY, PETER  BOOK OF MORDRED, THE  P
HANRATTY, PETER  LAST KNIGHT OF ALBION, THE  LP
HANSEN, KARL  DREAM GAMES  C/P,P
HANSEN, KARL  WAR GAMES  C/P:2,P
HARBINSON, W. A.  DREAM MAKER  SFWA
HARBINSON, W. A.  GENESIS  C/P:2,P
HARBINSON, W. A.  REVELATION  P
HARBOTTLE, PHILIP  FANTASY ADVENTURES #1  LPA
HARBOTTLE, PHILIP  FANTASY ADVENTURES #2  LPA
HARBOTTLE, PHILIP  FANTASY ADVENTURES 3  LPA
HARBOTTLE, PHILIP/WALLACE, SEAN  FANTASY ANNUAL 4  LPA
HARBOTTLE, PHILIP/WALLACE, SEAN  FANTASY ANNUAL 5  LPA
HARD, ANNETTE  STRANGER AT THE CROSSROADS  P
HARD, T. W.  SUM VII  H
HARD, T. W.  SUM VII (PROOF)  SR-LP
HARDIE, RAYMOND  ABYSSOS  SFWA-P

HARDING, LEE ALTERED I:

URSULA K. LE GUIN’S SCIENCE FICTION WRITING WORKSHOP, THE

H,SR-LP

HARDING, LEE  FALLEN SPACEMAN, THE  H,LP
HARDING, LEE  FUTURE SANCTUARY  SR-P
HARDING, LEE  MISPLACED PERSONS  P
HARDING, LEE  ROOMS OF PARADISE  SR-HA,HA
HARDING, RICHARD  BAY CITY BURNOUT [@OUTRIDER #4]  P
HARDING, RICHARD  BLOOD HIGHWAY [@OUTRIDER #3]  P
HARDING, RICHARD  FIRE AND ICE [@OUTRIDER #2]  P
HARDING, RICHARD  OUTRIDER, THE [@OUTRIDER #1]  P
HARDINGE, FRANCES  FLY BY NIGHT  H
HARDWICK, MICHAEL/HARDWICK, MOLLIE  PRIVATE LIFE OF SHERLOCK HOLMES, THE  P
HARDY, JASON M.  DROPS OF CORRUPTION [SHADOWRUN]  P
HARDY, JASON M./BILLS, RANDALL N.  PRINCIPLES OF DESOLATION [@MECHWARRIOR DARK AGE/BATTLETECH]  P
HARDY, JASON M.  SCORPION JAR, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  P
HARDY, LYNDON  MASTER OF THE FIVE MAGICS  C/P:2,P
HARDY, LYNDON  RIDDLE OF THE SEVEN REALMS  C/P:2,P
HARDY, LYNDON  SECRET OF THE SIXTH MAGIC  C/P:2,P
HARDY, NAOMI  MUCH OF MAGIC AND MIRACLES  P
HARDY, PHIL  ENCYCLOPEDIA OF SCIENCE FICTION MOVIES, THE  R/REF-H
HARKON, FRANZ  SPAWN OF SPACE  SR-P
HARLAN, THOMAS  DARK LORD, THE [OATH OF EMPIRE 4]  H
HARLAN, THOMAS  GATE OF FIRE, THE [OATH OF EMPIRE 2]  H,P

AUTHOR  TITLE  CODE
HARLAN, THOMAS  HOUSE OF REEDS [TIME OF THE SIXTH SUN 2]  H
HARLAN, THOMAS  SHADOW OF ARARAT, THE [OATH OF EMPIRE 1]  C/P,P
HARLAN, THOMAS  SHADOW OF ARARAT, THE (PROOF) [OATH OF EMPIRE 1]  SR-LP
HARLAN, THOMAS  STORM OF HEAVEN, THE [OATH OF EMPIRE 3]  H
HARLAN, THOMAS  WASTELAND OF FLINT [TIME OF THE SIXTH SUN 1]  H
HARLEQUIN BOOKS S. A.  NOCTURNE (NOVEL EXCERPTS)  PA
HARLING, ROBERT  ENORMOUS SHADOW, THE  H
HARMAN, ANDREW  FAHRENHEIT 666  P
HARMAN, ANDREW  FROGS OF WAR, THE  P
HARMAN, ANDREW  IT CAME FROM ON HIGH  P
HARMAN, ANDREW  SCRYING GAME, THE  P
HARMAN, ANDREW  SORCERER’S APPENDIX, THE  P
HARMON, DAVID P.  DEADLY YEARS, THE (FOTONOVEL) [STAR TREK]  [C/RAN-P]
HARMON, DAVID P./COON, GENE L.  PIECE OF THE ACTION, A [STAR TREK]  SR-S
HARNESS, CHARLES L.  CATALYST, THE  C/P,P
HARNESS, CHARLES L.  CATALYST, THE (PROOF)  SR-LP
HARNESS, CHARLES L.  CYBELE, WITH BLUEBONNETS  H
HARNESS, CHARLES L.  FIREBIRD  H,P
HARNESS, CHARLES L.  FLIGHT INTO YESTERDAY  H
HARNESS, CHARLES L.  KRONO  C/P,P
HARNESS, CHARLES L.  LUNAR JUSTICE  P
HARNESS, CHARLES L.  LURID DREAMS  C/P,P
HARNESS, CHARLES L.  ORNAMENT TO HIS PROFESSION, AN  H
HARNESS, CHARLES L.  PARADOX MEN, THE  DD118
HARNESS, CHARLES L.  PARADOX MEN, THE (REVISED)  H
HARNESS, CHARLES L.  REDWORLD  C/P,P
HARNESS, CHARLES L.  RING OF RITORNEL, THE  C/P:2,P
HARNESS, CHARLES L.  RINGS  H
HARNESS, CHARLES L.  ROSE, THE  C/P,P
HARNESS, CHARLES L.  VENETIAN COURT, THE  C/P:2,P
HARNESS, CHARLES L.  WOLFHEAD  C/P:2,P
HARPER, GEORGE W.  GYPSY EARTH  H,SFWA
HARPER, HARRY  DAWN OF THE SPACE AGE  [C/RAN-H]
HARPER, RORY  PETROGYPSIES  P,SFWA-P
HARPER, STEVEN  DEAD MAN ON THE MOON [LCSI]  LP
HARPER, STEVEN  DREAMER [SILENT EMPIRE 1]  P
HARPER, STEVEN  NIGHTMARE [SILENT EMPIRE 2]  P
HARPER, STEVEN  OFFSPRING [SILENT EMPIRE 4]  P
HARPER, STEVEN  TRICKSTER [SILENT EMPIRE 3]  P
HARPER, STEVEN  UNITY [@BATTLESTAR GALACTICA 2003]  LP
HARPER, TARA K.  CAT SCRATCH FEVER  C/P:2,P
HARPER, TARA K.  CATARACT  C/P,P
HARPER, TARA K.  GRAYHEART [WOLFWALKER 6]  C/P:2,P
HARPER, TARA K.  LIGHTWING  C/P:2,P
HARPER, TARA K.  SHADOW LEADER [WOLFWALKER 2]  C/P,P
HARPER, TARA K.  SILVER MOONS, BLACK STEEL [WOLFWALKER 5]  P
HARPER, TARA K.  STORM RUNNER [WOLFWALKER 3]  C/P,P
HARPER, TARA K.  WOLF IN NIGHT [WOLFWALKER 7]  P
HARPER, TARA K.  WOLF’S BANE [WOLFWALKER 4]  C/P,P
HARPER, TARA K.  WOLFWALKER [WOLFWALKER 1]  C/P:2,P
HARPERCOLLINS  VOYAGER  B
HARPERPRISM  REALMS OF TOLKIEN [LORD OF THE RINGS]  A
HARPMAN, JACQUELINE  I WHO HAVE NEVER KNOWN MEN  P
HARPUR, PATRICK  SERPENT’S CIRCLE, THE  H
HARRINGTON, ALAN  PARADISE 1  H
HARRINGTON, ALAN  REVELATIONS OF DOCTOR MODESTO, THE  P
HARRIS, ANNE  ACCIDENTAL CREATURES  LP
HARRIS, ANNE  INVENTING MEMORY  LP
HARRIS, BARBARA S.  WHO IS JULIA?  P
HARRIS, CHARLAINE  ALL TOGETHER DEAD [SOUTHERN VAMPIRES 7]  H:2
HARRIS, CHARLAINE  CLUB DEAD [SOUTHERN VAMPIRES 3]  P
HARRIS, CHARLAINE  DEAD AS A DOORNAIL [SOUTHERN VAMPIRES 5]  H
HARRIS, CHARLAINE  DEAD BY DAY [SOUTHERN VAMPIRES 4,5]  H
HARRIS, CHARLAINE  DEAD IN DIXIE [SOUTHERN VAMPIRES 1,2,3]  H
HARRIS, CHARLAINE  DEAD TO THE WORLD [SOUTHERN VAMPIRES 4]  H
HARRIS, CHARLAINE  DEAD UNTIL DARK [SOUTHERN VAMPIRES 1]  P
HARRIS, CHARLAINE  DEFINITELY DEAD [SOUTHERN VAMPIRES 6]  H:2
HARRIS, CHARLAINE  GRAVE SIGHT  H
HARRIS, CHARLAINE  GRAVE SURPRISE  H:2
HARRIS, CHARLAINE  LIVING DEAD IN DALLAS [SOUTHERN VAMPIRES 2]  C/P:2,P
HARRIS, CLARE WINGER  AWAY FROM THE HERE AND NOW  H
HARRIS, CLARE WINGER/BREUER, MILES J.  DEAD WORLD, THE  B
HARRIS, DEBORAH TURNER  BURNING STONE, THE [MAGES OF GARILLON 1]  C/P,SFWA,SFWA-P,P
HARRIS, DEBORAH TURNER  CALEDON OF THE MISTS [CALEDON 1]  C/P,SFWA-P,P

The MITSFS Pinkdex  September 24, 2007  Page 181
AUTHOR  TITLE  CODE
HARRIS, DEBORAH TURNER  CITY OF EXILE, THE  C/P,P
HARRIS, DEBORAH TURNER  GAUNTLET OF MALICE, THE [MAGES OF GARILLON 2]  C/P,SFWA,SFWA-P,LP
HARRIS, DEBORAH TURNER  QUEEN OF ASHES, THE [CALEDON 2]  C/P,P
HARRIS, DEBORAH TURNER  SPIRAL OF FIRE [MAGES OF GARILLON 3]  C/P,P
HARRIS, GERALDINE  CHILDREN OF THE WIND, THE [SEVEN CITADELS 2]  P
HARRIS, GERALDINE  DEAD KINGDOM, THE [SEVEN CITADELS 3]  P
HARRIS, GERALDINE  PRINCE OF THE GODBORN [SEVEN CITADELS 1]  P
HARRIS, GERALDINE  SEVENTH GATE, THE [SEVEN CITADELS 4]  C/P,P
HARRIS, JOHN BEYNON  SECRET PEOPLE, THE  C/P:2,P
HARRIS, MARILYN  CONJURERS, THE  C/P,P
HARRIS, MARILYN  NIGHT GAMES  SFWA-P

HARRIS, MARK DOCTOR WHO TECHNICAL MANUAL, THE

(OFFICIAL ANNIVERSARY VOLUME) [DOCTOR WHO]

R/REF-H,C/REF-H

HARRIS, MARK JONATHAN  SOLAY  SFWA
HARRIS, RAYMOND  BROKEN WORLDS, THE  P
HARRIS, RAYMOND  SCHIZOGENIC MAN, THE  P
HARRIS, RAYMOND  SHADOWS OF THE WHITE SUN  H:2
HARRIS, ROBERT  FATHERLAND  H:2,P
HARRIS, ROSEMARY  SEA MAGIC  H
HARRIS, THOMAS  RED DRAGON  [C/RAN-P]
HARRISON, BARBARA  COLD NIGHT’S DEATH, A  P
HARRISON, EMMA  GARDEN OF EVIL [@CHARMED]  P
HARRISON, EMMA  PHOEBE WHO? [@CHARMED]  P
HARRISON, EMMA  QUEEN’S CURSE, THE [@CHARMED]  P
HARRISON, EMMA  SOMETHING WICCAN THIS WAY COMES [@CHARMED]  P
HARRISON, EMMA  TALE OF TWO PIPERS, A [@CHARMED]  P

HARRISON, HARRY ADVENTURES OF THE STAINLESS STEEL RAT, THE

[STAINLESS STEEL RAT 1,2,3]

C/P,P,H

HARRISON, HARRY/GORDON, THEODORE J.  AHEAD OF TIME  SFWA
HARRISON, HARRY/ALDISS, BRIAN W.  ASTOUNDING-ANALOG READER 1, THE  HA,SFWA
HARRISON, HARRY/ALDISS, BRIAN W.  ASTOUNDING-ANALOG READER 2, THE  HA,SFWA
HARRISON, HARRY  ASTOUNDING — JOHN W. CAMPBELL MEMORIAL ANTHOLOGY  C/PA,PA,SFWA
HARRISON, HARRY  BACKDROP OF STARS  PA
HARRISON, HARRY  BEST OF HARRY HARRISON, THE  C/P,P
HARRISON, HARRY  BEST SF: 1967  PA,C/PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1968  C/PA,PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1969  C/PA,PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1970  HA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1971  PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1972  C/PA,PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1973  C/PA,PA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1974  HA
HARRISON, HARRY/ALDISS, BRIAN W.  BEST SF: 1975  HA
HARRISON, HARRY  BILL, THE GALACTIC HERO [BILL 0]  C/P,P,H
HARRISON, HARRY/SHECKLEY, ROBERT  BILL, THE GALACTIC HERO ON THE PLANET OF BOTTLED BRAINS [BILL 2] C/P:2,P

HARRISON, HARRY/BISCHOFF, DAVID F.

BILL, THE GALACTIC HERO ON THE PLANET OF TASTELESS PLEASURE [BILL 3]

C/P:2,P

HARRISON, HARRY/BISCHOFF, DAVID F.

BILL, THE GALACTIC HERO ON THE PLANET OF TEN THOUSAND BARS [BILL 5]

C/P,P

HARRISON, HARRY/HALDEMAN, JACK C. II BILL, THE GALACTIC HERO ON THE PLANET OF ZOMBIE VAMPIRES

[BILL 4]

C/P:2,P

HARRISON, HARRY/HARRIS, DAVID  BILL, THE GALACTIC HERO: THE FINAL INCOHERENT ADVENTURE [BILL 6] C/P,SFWA-P,P
HARRISON, HARRY  BILL THE GALACTIC HERO: THE PLANET OF THE ROBOT SLAVES [BILL 1]  C/P,P
HARRISON, HARRY  CALIFORNIA ICEBERG, THE  H
HARRISON, HARRY  CAPTIVE UNIVERSE  C/P,P,H
HARRISON, HARRY  DALETH EFFECT, THE  C/P,P,H
HARRISON, HARRY  DEATHWORLD  C/P:2,P
HARRISON, HARRY  DEATHWORLD 2  C/P,P
HARRISON, HARRY  DEATHWORLD 3  C/P:2,P
HARRISON, HARRY  DEATHWORLD TRILOGY, THE  C/P,P,H
HARRISON, HARRY  50 IN 50  H
HARRISON, HARRY  GALACTIC DREAMS  H
HARRISON, HARRY  GALACTIC DREAMS (PROOF)  SR-LP
HARRISON, HARRY  HAMMER AND THE CROSS, THE [HAMMER AND THE CROSS 1]  C/P:2,P
HARRISON, HARRY  HOMEWORLD [TO THE STARS 1]  C/P:2,P
HARRISON, HARRY  IN OUR HANDS, THE STARS  H
HARRISON, HARRY  INVASION: EARTH  C/P,LP
HARRISON, HARRY  JUPITER LEGACY, THE  C/P:2,P
HARRISON, HARRY  JUPITER PLAGUE, THE  P,SFWA-P
HARRISON, HARRY/HOLM, JOHN  KING AND EMPEROR [HAMMER AND THE CROSS 3]  C/P,P,H
HARRISON, HARRY/DICKSON, GORDON R.  LIFESHIP, THE  P,SFWA
HARRISON, HARRY  LIGHT FANTASTIC, THE  SFWA,LPA
HARRISON, HARRY  MAKE ROOM! MAKE ROOM!  C/P,P,H
HARRISON, HARRY  MAN FROM P.I.G., THE  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 182
AUTHOR  TITLE  CODE
HARRISON, HARRY  MECHANISMO  A
HARRISON, HARRY  MEN FROM P.I.G. AND R.O.B.O.T., THE  H:2
HARRISON, HARRY  NOVA 1  C/PA,PA
HARRISON, HARRY  NOVA 2  HA,PA
HARRISON, HARRY  NOVA 3  HA,SFWA
HARRISON, HARRY  NOVA 4  C/PA,PA,SFWA
HARRISON, HARRY/HOLM, JOHN  ONE KING’S WAY [HAMMER AND THE CROSS 2]  C/P,P
HARRISON, HARRY  ONE STEP FROM EARTH  H:2,SFWA-P,P
HARRISON, HARRY  OUTDATED MAN, THE  C/PA,PA
HARRISON, HARRY  PLAGUE FROM SPACE  C/P,P,H
HARRISON, HARRY  PLANET OF NO RETURN  C/P,LP,H
HARRISON, HARRY  PLANET OF THE DAMNED  C/P:2,SFWA-P,P
HARRISON, HARRY  PRIME NUMBER  C/P:2,P
HARRISON, HARRY  QE2 IS MISSING, THE  P
HARRISON, HARRY  QUEEN VICTORIA’S REVENGE  SFWA-P
HARRISON, HARRY  REBEL IN TIME, A  C/P:2,P
HARRISON, HARRY  RETURN TO EDEN [EDEN 3]  H:2,P
HARRISON, HARRY/PUGNER, CAROL  SCIENCE FICTION READER, A  LPA
HARRISON, HARRY  SF: AUTHORS’ CHOICE  PA,C/PA
HARRISON, HARRY  SF: AUTHORS’ CHOICE 2  PA,C/PA
HARRISON, HARRY  SF: AUTHORS’ CHOICE 3  C/PA,PA
HARRISON, HARRY  SF: AUTHORS’ CHOICE 4  C/PA,PA
HARRISON, HARRY  SKYFALL  C/P,SFWA-P,P,H
HARRISON, HARRY/EDWARDS, MALCOLM  SPACECRAFT IN FACT AND FICTION  A
HARRISON, HARRY  SPACESHIP MEDIC  SR-P
HARRISON, HARRY  STAINLESS STEEL RAT, THE [STAINLESS STEEL RAT 1]  C/P,P
HARRISON, HARRY  STAINLESS STEEL RAT FOR PRESIDENT, THE [STAINLESS STEEL RAT 5]  C/P,P,H
HARRISON, HARRY  STAINLESS STEEL RAT GETS DRAFTED, THE [STAINLESS STEEL RAT 6]  H:2,P
HARRISON, HARRY  STAINLESS STEEL RAT GOES TO HELL, THE [STAINLESS STEEL RAT 8]  C/P,P,H
HARRISON, HARRY  STAINLESS STEEL RAT IS BORN, A [STAINLESS STEEL RAT 0]  H:2,P
HARRISON, HARRY  STAINLESS STEEL RAT JOINS THE CIRCUS, THE [STAINLESS STEEL RAT 9]  H

HARRISON, HARRY STAINLESS STEEL RAT JOINS THE CIRCUS, THE (PROOF)

[STAINLESS STEEL RAT 9]

SR-LP

HARRISON, HARRY  STAINLESS STEEL RAT SAVES THE WORLD, THE [STAINLESS STEEL RAT 2]  C/P,P,H
HARRISON, HARRY  STAINLESS STEEL RAT SINGS THE BLUES, THE [STAINLESS STEEL RAT 7]  H:2,P
HARRISON, HARRY  STAINLESS STEEL RAT WANTS YOU!, THE [STAINLESS STEEL RAT 4]  C/P,P,H
HARRISON, HARRY  STAINLESS STEEL RAT’S REVENGE, THE [STAINLESS STEEL RAT 3]  C/P,P,H
HARRISON, HARRY  STAINLESS STEEL VISIONS [STAINLESS STEEL RAT]  C/P,P,H
HARRISON, HARRY  STAR SMASHERS OF THE GALAXY RANGERS  C/P,P,H
HARRISON, HARRY  STARS & STRIPES FOREVER [STARS & STRIPES 1]  C/P,P,H
HARRISON, HARRY  STARS & STRIPES IN PERIL [STARS & STRIPES 2]  H,P
HARRISON, HARRY  STARS & STRIPES TRIUMPHANT [STARS & STRIPES 3]  P
HARRISON, HARRY  STARWORLD [TO THE STARS 3]  C/P:2,P
HARRISON, HARRY/STOVER, LEON E.  STONEHENGE  C/P,SFWA,P
HARRISON, HARRY  TECHNICOLOR TIME MACHINE, THE  C/P:2,SFWA-P,P
HARRISON, HARRY/MCALLISTER, BRUCE  THERE WON’T BE WAR  C/PA,PA
HARRISON, HARRY  TO THE STARS [TO THE STARS 1,2,3]  H:2,P
HARRISON, HARRY  TRANSATLANTIC TUNNEL, HURRAH!, A  C/P:2,P
HARRISON, HARRY  TUNNEL THROUGH THE DEEPS  H,P
HARRISON, HARRY/MINSKY, MARVIN  TURING OPTION, THE  H:2,P
HARRISON, HARRY  TWO TALES AND EIGHT TOMORROWS  C/P:2,P
HARRISON, HARRY  WAR WITH THE ROBOTS  C/P:2,P
HARRISON, HARRY  WEST OF EDEN [EDEN 1]  C/P,P,H
HARRISON, HARRY  WHEELWORLD [TO THE STARS 2]  C/P:2,P
HARRISON, HARRY  WINTER IN EDEN [EDEN 2]  H:2,P
HARRISON, HARRY  WORLDS OF WONDER  HA
HARRISON, HARRY  YEAR 2000, THE  C/PA,PA
HARRISON, HARRY/ALDISS, BRIAN W.  YEAR’S BEST SCIENCE FICTION #8, THE  SR-PA
HARRISON, HARRY  YOU CAN BE THE STAINLESS STEEL RAT [STAINLESS STEEL RAT]  P

HARRISON, KIM/SANDS, LYNSAY/ ARMSTRONG, KELLEY/HANDELAND, LORI DATES FROM HELL

P

HARRISON, KIM  DEAD WITCH WALKING [RACHEL MORGAN 1]  P
HARRISON, KIM  EVERY WHICH WAY BUT DEAD [RACHEL MORGAN 3]  C/P,P
HARRISON, KIM  FISTFUL OF CHARMS, A [RACHEL MORGAN 4]  C/P,P
HARRISON, KIM  FOR A FEW DEMONS MORE [RACHEL MORGAN 5]  H:2
HARRISON, KIM  GOOD, THE BAD, AND THE UNDEAD, THE [RACHEL MORGAN 2]  P
HARRISON, M. JOHN  CENTAURI DEVICE, THE  H:2,SFWA:2,SR-P
HARRISON, M. JOHN  COMMITTED MEN, THE  H,SFWA
HARRISON, M. JOHN  FLOATING GODS, THE [VIRICONIUM]  P
HARRISON, M. JOHN  LIGHT [LIGHT 1]  H
HARRISON, M. JOHN  MACHINE IN SHAFT TEN, THE  SR-P
HARRISON, M. JOHN  NOVA SWING [LIGHT 2]  H
HARRISON, M. JOHN  PASTEL CITY, THE [VIRICONIUM 1]  H:2,SFWA,P
HARRISON, M. JOHN  STORM OF WINGS, A [VIRICONIUM 2]  C/P,SFWA,P

AUTHOR  TITLE  CODE
HARRISON, M. JOHN  STORM OF WINGS, A (BRITISH) [VIRICONIUM 2]  SR-P
HARRISON, M. JOHN  TOURISM  B
HARRISON, M. JOHN  VIRICONIUM [VIRICONIUM]  LP
HARRISON, M. JOHN  VIRICONIUM NIGHTS [VIRICONIUM]  P
HARRISON, METTE IVIE  PRINCESS AND THE HOUND, THE  H
HARRISON, MICHAEL  HIGHER THINGS  H
HARRISON, PAYNE  STORMING INTREPID  C/P:2,P
HARRISON, SUE  MOTHER EARTH FATHER SKY  H
HARRISON, WILLIAM  ROLLERBALL  C/P,P
HARRY, ERIC L.  SOCIETY OF THE MIND  P
HART, JOHNNY  B.C. GREAT ZOT, I’M BEAUTIFUL  [C/RAN-P]
HART, STEVE  GOLDEN VOYAGE OF SINBAD, THE  P
HARTENFELS, JEROME  LAZARUS  P
HARTLEY, JESS  IN NORTHERN TWILIGHT [@EXALTED #3]  P
HARTLEY, L. P.  FACIAL JUSTICE  H
HARTMAN, KEITH  GUMSHOE, THE WITCH & THE VIRTUAL CORPSE, THE  H,LP
HARTMAN, WILLIAM K.  MARS UNDERGROUND  P
HARTRIDGE, JON  BINARY DIVINE  H:2,P
HARTRIDGE, JON  EARTHJACKET  H
HARTWELL, DAVID G.  AGE OF WONDERS  R/REF-H,C/REF-P
HARTWELL, DAVID G./CRAMER, KATHRYN  ASCENT OF WONDER, THE  HA
HARTWELL, DAVID G./BRODERICK, DAMIEN CENTAURUS  HA
HARTWELL, DAVID G.  CHRISTMAS FOREVER  HA
HARTWELL, DAVID G.  CHRISTMAS MAGIC  PA
HARTWELL, DAVID G.  CHRISTMAS STARS  C/PA,LPA
HARTWELL, DAVID G.  COLOR OF EVIL, THE [DARK DESCENT 1]  PA
HARTWELL, DAVID G.  DARK DESCENT, THE  LPA
HARTWELL, DAVID G.  FABULOUS FORMLESS DARKNESS, A  PA
HARTWELL, DAVID G./CRAMER, KATHRYN  HARD SF RENAISSANCE, THE  HA
HARTWELL, DAVID G.  MASTERPIECES OF FANTASY AND ENCHANTMENT  HA
HARTWELL, DAVID G.  MASTERPIECES OF FANTASY AND WONDER  HA
HARTWELL, DAVID G.  MEDUSA IN THE SHIELD, THE [DARK DESCENT 2]  PA
HARTWELL, DAVID G./GRANT, GLENN  NORTHERN STARS  HA
HARTWELL, DAVID G./GRANT, GLENN  NORTHERN SUNS  LPA
HARTWELL, DAVID G.  SCIENCE FICTION CENTURY, THE  HA
HARTWELL, DAVID G./CRAMER, KATHRYN  SPACE OPERA RENAISSANCE, THE  HA
HARTWELL, DAVID G./CRAMER, KATHRYN  SPIRITS OF CHRISTMAS  PA
HARTWELL, DAVID G.  VISIONS OF FEAR  PA
HARTWELL, DAVID G./WOLF, MILTON T.  VISIONS OF WONDER  LPA
HARTWELL, DAVID G.  WORLD TREASURY OF SCIENCE FICTION, THE  HA:2,LPA
HARTWELL, DAVID G.  YEAR’S BEST FANTASY  PA
HARTWELL, DAVID G.  YEAR’S BEST FANTASY 2  C/PA,PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST FANTASY 3  PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST FANTASY 4  PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST FANTASY 5  PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST FANTASY 6  LPA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST FANTASY 7  LPA
HARTWELL, DAVID G.  YEAR’S BEST SF (1995)  C/PA,PA
HARTWELL, DAVID G.  YEAR’S BEST SF 2 (1996)  C/PA,PA
HARTWELL, DAVID G.  YEAR’S BEST SF 3 (1997)  C/PA,PA
HARTWELL, DAVID G.  YEAR’S BEST SF 4 (1998)  C/PA,PA
HARTWELL, DAVID G.  YEAR’S BEST SF 5 (1999)  PA,C/PA
HARTWELL, DAVID G.  YEAR’S BEST SF 6 (2000)  C/PA,PA
HARTWELL, DAVID G.  YEAR’S BEST SF 7 (2001)  C/PA,PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST SF 8 (2002)  HA,PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST SF 9 (2003)  HA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST SF 10 (2004)  HA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST SF 11 (2005)  PA
HARTWELL, DAVID G./CRAMER, KATHRYN  YEAR’S BEST SF 12 (2006)  PA
HARVARD LAMPOON  BORED OF THE RINGS  C/P:3,SFWA-P,P
HARVEY, A. D.  WARRIORS OF THE RAINBOW  LP
HARVEY, FRANK  AIR FORCE!  C/P,P
HARVEY, JAMES  TITANS OF THE UNIVERSE  P
HARVEY, M. ELAYN  WARHAVEN  H
HARVEY, WILLIAM FRYER  BEAST WITH FIVE FINGERS, THE  H
HASSE, HENRY L.  STARS WILL WAIT, THE  H
HASSLER, DONALD M.  HAL CLEMENT  R/REF-P
HASSLER, KENNETH W.  DESTINATION: TERRA  H
HASSLER, KENNETH W.  DREAM SQUAD, THE  H
HASSLER, KENNETH W.  INTERGALAC AGENT  H
HASSLER, KENNETH W.  MESSAGE FROM EARTH, A  H
HASTINGS, MILO  CITY OF ENDLESS NIGHT  H,LP
HATCH, ERIC  HORSE IN THE GRAY FLANNEL SUIT, THE  P
HATCH, GEORGE  GUIGNOIR AND OTHER FURIES  LPA

The MITSFS Pinkdex  September 24, 2007  Page 184
AUTHOR  TITLE  CODE
HATCH, GEORGE  NOCTULPA NO. 4  LPA
HATCH, GEORGE  SINISTRE  LPA
HATCH, GEORGE  SOULS IN PAWN  LPA
HATCH, GERALD  DAY THE EARTH FROZE, THE  C/P:2,P
HATCH, RICHARD/GOLDEN, CHRISTOPHER  ARMAGEDDON [@BATTLESTAR GALACTICA 1978]  P
HATCH, RICHARD/LINAWEAVER, BRAD  DESTINY [@BATTLESTAR GALACTICA 1978]  P
HATCH, RICHARD/LINAWEAVER, BRAD  PARADIS [@BATTLESTAR GALACTICA 1978]  P
HATCH, RICHARD/RODGERS, ALAN  REBELLION [@BATTLESTAR GALACTICA 1978]  H
HATCH, RICHARD/LINAWEAVER, BRAD  REDEMPTION [@BATTLESTAR GALACTICA 1978]  H
HATCH, RICHARD/GOLDEN, CHRISTOPHER  WARHAWK [@BATTLESTAR GALACTICA 1978]  H
HATHAWAY-NAYNE, ANNE  THESE OUR REVELS [@FOREVER KNIGHT]  P
HAUF, MICHELE  GOSSAMYR  LP
HAUF, MICHELE  RHIANA  LP
HAUF, MICHELE  SERAPHIM  LP

HAUPT, HANNA BEATE/HESTON, LILLA/ LITTELL, JOY/SOLOTAROFF, SARAH FLIGHTS OF FANTASY

LP

HAUPTMANN, GERHART  ISLAND OF THE GREAT MOTHER, THE  H
HAUPTMANN, RICHARD A.  WORK OF JACK WILLIAMSON, THE  R/REF-H
HAUSMAN, GERALD  STARGAZER (LARGE PAPERBACK AND CASSETTE) [STARGAZER 1]  SR-MM
HAUTALA, RICK  BEYOND THE SHROUD [WORLD OF DARKNESS WRAITH]  H
HAUTALA, RICK  HIDDEN SAINT, THE [@POLTERGEIST THE LEGACY]  P
HAUTALA, RICK  WINTER WAKE  P
HAVERSINE, THEDA  ESCAPE FROM FLATLAND [FLATLAND]  SR-P
HAWDON, ROBIN  RUSTLE IN THE GRASS, A  P,SFWA,SFWA-P
HAWKE, SIMON  AMBIVALENT MAGICIAN, THE  P
HAWKE, SIMON  ARGONAUT AFFAIR, THE [TIMEWARS 7]  P
HAWKE, SIMON  BATMAN: TO STALK A SPECTER  P
HAWKE, SIMON  BLAZE OF GLORY [STAR TREK TNG]  C/P:2,P
HAWKE, SIMON  BROKEN BLADE, THE [@DARK SUN ATHAS #3]  P
HAWKE, SIMON  CLEOPATRA CRISIS, THE [TIMEWARS 11]  P
HAWKE, SIMON  DRACULA CAPER, THE [TIMEWARS 8]  C/P,P
HAWKE, SIMON  HELLFIRE REBELLION, THE [TIMEWARS 10]  C/P,P
HAWKE, SIMON  INADEQUATE ADEPT, THE [MARVIN BREWSTER 2]  C/P:2,P
HAWKE, SIMON  IRON THRONE, THE [BIRTHRIGHT 1]  P
HAWKE, SIMON  IVANHOE GAMBIT, THE [TIMEWARS 1]  C/P:2,P
HAWKE, SIMON  JASON LIVES  P
HAWKE, SIMON  KHYBER CONNECTION, THE [TIMEWARS 6]  P
HAWKE, SIMON  LAST WIZARD, THE [WYRDRUNE 8]  C/P,P
HAWKE, SIMON  LILLIPUT LEGION, THE [TIMEWARS 9]  C/P,P
HAWKE, SIMON  NAUTILUS SANCTION, THE [TIMEWARS 5]  C/P:2,P
HAWKE, SIMON  NINE LIVES OF CATSEYE GOMEZ, THE [WYRDRUNE]  C/P,P
HAWKE, SIMON  NOMAD, THE [@DARK SUN TRIBE OF ONE #3]  P
HAWKE, SIMON  OUTCAST, THE [@DARK SUN TRIBE OF ONE #1]  C/P,P
HAWKE, SIMON  PATRIAN TRANSGRESSION, THE [STAR TREK]  C/P:2,P
HAWKE, SIMON  PIMPERNEL PLOT, THE [TIMEWARS 3]  C/P:2,P
HAWKE, SIMON  PREDATOR 2 [PREDATOR]  P
HAWKE, SIMON  PSYCHODROME [PSYCHODROME 1]  C/P,SFWA-P,P
HAWKE, SIMON  RELUCTANT SORCERER, THE [MARVIN BREWSTER 1]  C/P,P
HAWKE, SIMON  ROMULAN PRIZE, THE [STAR TREK TNG]  C/P:2,P
HAWKE, SIMON  SAMURAI WIZARD, THE [WYRDRUNE 5]  C/P,P
HAWKE, SIMON  SEEKER, THE [@DARK SUN TRIBE OF ONE #2]  P
HAWKE, SIMON  SHAPECHANGER SCENARIO, THE [PSYCHODROME 2]  C/P,SFWA-P,P
HAWKE, SIMON  SIX-GUN SOLUTION, THE [TIMEWARS 12]  C/P,P
HAWKE, SIMON  TIMEKEEPER CONSPIRACY, THE [TIMEWARS 2]  C/P:2,P
HAWKE, SIMON  WHIMS OF CREATION, THE  C/P:2,P
HAWKE, SIMON  WIZARD OF CAMELOT, THE [WYRDRUNE 0]  C/P:2,P
HAWKE, SIMON  WIZARD OF 4TH STREET, THE [WYRDRUNE 1]  C/P:2,P
HAWKE, SIMON  WIZARD OF LOVECRAFT’S CAFE, THE [WYRDRUNE 7]  C/P:2,P
HAWKE, SIMON  WIZARD OF RUE MORGUE, THE [WYRDRUNE 4]  C/P:2,P
HAWKE, SIMON  WIZARD OF SANTA FE, THE [WYRDRUNE 6]  C/P,P
HAWKE, SIMON  WIZARD OF SUNSET STRIP, THE [WYRDRUNE 3]  C/P:2,SFWA-P,P
HAWKE, SIMON  WIZARD OF WHITECHAPEL, THE [WYRDRUNE 2]  C/P:2,P
HAWKE, SIMON  ZENDA VENDETTA, THE [TIMEWARS 4]  C/P:2,P
HAWKINS, JANE E.  QUANTUM GATE  C/P,P
HAWKINS, JIM  LIVING ONE, THE  C/P,P
HAWKINS, WARD  BLAZE OF WRATH [HARRY BORG 3]  C/P,P
HAWKINS, WARD  RED FLAME BURNING [HARRY BORG 1]  C/P,P
HAWKINS, WARD  SWORD OF FIRE [HARRY BORG 2]  P
HAWKINS, WARD  TORCH OF FEAR [HARRY BORG 4]  P
HAWKS, CHESTER  PYTHON MEN OF LOST CITY  SR-LP

HAWORTH PRESS SPECIAL COLLECTIONS SCIENCE/FICTION COLLECTIONS:

FANTASY, SUPERNATURAL AND WEIRD TALES

R/REF-P

HAWTHORN BOOKS  FOUR FUTURES  HA
HAWTHORN BOOKS  NO MIND OF MAN  HA

The MITSFS Pinkdex  September 24, 2007  Page 185
AUTHOR  TITLE  CODE
HAWTHORN BOOKS  THREE TRIPS IN TIME AND SPACE  HA
HAWTHORNE, JULIAN  GOLDEN FLEECE, THE  H
HAWTHORNE, JULIAN  HAUNTS AND ENIGMAS  HA
HAWTHORNE, NATHANIEL  HOUSE OF THE SEVEN GABLES, THE  C/P,P
HAWTHORNE, NATHANIEL  TANGLEWOOD TALES  P
HAWTHORNE, NATHANIEL  WONDERBOOK AND TANGLEWOOD TALES  H
HAY, GEORGE  FLIGHT OF THE HESPER  SR-P
HAY, GEORGE  PULSAR 1  SR-PA
HAY, GEORGE  STOPWATCH  SR-PA
HAY, GEORGE  THIS PLANET FOR SALE  SR-P
HAY, JACOB/KESHISHIAN, JOHN M.  AUTOPSY FOR A COSMONAUT  C/P,P
HAYDON, ELIZABETH  DESTINY [SYMPHONY OF AGES 3]  H,P
HAYDON, ELIZABETH  ELEGY FOR A LOST STAR [SYMPHONY OF AGES 5]  H,P
HAYDON, ELIZABETH  FLOATING ISLAND  H
HAYDON, ELIZABETH  PROPHECY [SYMPHONY OF AGES 2]  H,P
HAYDON, ELIZABETH  REQUIEM FOR THE SUN [SYMPHONY OF AGES 4]  H
HAYDON, ELIZABETH  REQUIEM FOR THE SUN (PROOF) [SYMPHONY OF AGES 4]  SR-LP
HAYDON, ELIZABETH  RHAPSODY [SYMPHONY OF AGES 1]  H,P
HAYES, CHARLES DOUGLAS  PORTALS IN A NORTHERN SKY  H
HAYES, RALPH  VISITING MOON, THE  H
HAYES, WILLIAM D.  PROJECT: GENIUS  LP
HAYLES, BRIAN  DOCTOR WHO AND THE CURSE OF PELADON [DOCTOR WHO]  SR-P,C/P:2
HAYLES, BRIAN  DOCTOR WHO AND THE ICE WARRIORS [DOCTOR WHO]  SR-P,C/P
HAYNES, BETSY  POWER, THE [TWILIGHT 6]  P
HAYNES, JAMES  VOICES IN THE DARK [TWILIGHT 2]  P
HAYNES, MARY  WORDCHANGER  P
HAYNES, SIMON  HAL SPACEJOCK [@HAL SPACEJOCK 1]  P
HAYNES, SIMON  HAL SPACEJOCK - SECOND COURSE [@HAL SPACEJOCK 2]  P
HAZEL, PAUL  UNDERSEA [FINNBRANCH 2]  P
HAZEL, PAUL  WEALDWIFE’S TALE, THE  P,SFWA
HAZEL, PAUL  WINTERKING [FINNBRANCH 3]  H,P
HAZEL, PAUL  YEARWOOD [FINNBRANCH 1]  C/P:2,P
HEALD, DENISE LOPES  MISTWALKER  P
HEALD, TIM  MAKING OF SPACE: 1999, THE  R/REF-P,C/REF-P
HEALY, RAYMOND J./MCCOMAS, J. FRANCIS ADVENTURES IN TIME AND SPACE  HA,SR-HA
HEALY, RAYMOND J./MCCOMAS, J. FRANCIS ADVENTURES IN TIME AND SPACE (SELECTIONS)  PA,C/PA
HEALY, RAYMOND J./MCCOMAS, J. FRANCIS MORE ADVENTURES IN TIME AND SPACE  PA,C/PA
HEALY, RAYMOND J.  NEW TALES OF SPACE AND TIME  C/PA,SR-PA
HEALY, RAYMOND J.  9 TALES OF SPACE AND TIME  SR-HA
HEARD, GERALD  IS ANOTHER WORLD WATCHING?  [C/RAN-P]
HEARD, H. F.  DOPPELGANGERS  C/P:2,P
HEARD, H. F.  WEIRD TALES OF TERROR AND DETECTION  H
HEARN, LAFCADIO  IN GHOSTLY JAPAN  P
HEARN, LIAN  ACROSS THE NIGHTINGALE FLOOR [@OTORI #1]  H:2
HEARN, LIAN  ACROSS THE NIGHTINGALE FLOOR (PROOF) [@OTORI #1]  SR-LP
HEARN, LIAN  BRILLIANCE OF THE MOON [@OTORI #3]  H
HEARN, LIAN  GRASS FOR HIS PILLOW [@OTORI #2]  H:2
HEARN, LIAN  HARSH CRY OF THE HERON, THE [@OTORI #4]  H
HEATH, PETER  ASSASSINS FROM TOMORROW [MIND BROTHERS 2]  C/P:2,P
HEATH, PETER  MEN WHO DIE TWICE [MIND BROTHERS 3]  P
HEATH, PETER  MIND BROTHERS, THE [MIND BROTHERS 1]  C/P:2,P
HEERMANN, TRAVIS  IVORY STAR, THE  P
HEINLEIN, ROBERT A.  ASSIGNMENT IN ETERNITY  SR-H,C/P:2,P
HEINLEIN, ROBERT A.  ASSIGNMENT IN ETERNITY (ABRIDGED)  SR-P
HEINLEIN, ROBERT A.  BETWEEN PLANETS  C/P,P,H,SR-LP
HEINLEIN, ROBERT A.  BEYOND THIS HORIZON  C/P,P,H
HEINLEIN, ROBERT A.  CAT WHO WALKS THROUGH WALLS, THE  C/P,P,H
HEINLEIN, ROBERT A.  CITIZEN OF THE GALAXY  C/P:2,P
HEINLEIN, ROBERT A.  DAY AFTER TOMORROW, THE  C/P:2,P
HEINLEIN, ROBERT A.  DOOR INTO SUMMER, THE  C/P,P,H
HEINLEIN, ROBERT A.  DOUBLE STAR  C/P:2,P
HEINLEIN, ROBERT A.  DOUBLE STAR (CHINESE)  FP
HEINLEIN, ROBERT A.  ESCALA NO TEMPO  FP
HEINLEIN, ROBERT A.  EXPANDED UNIVERSE  C/P,H,LP
HEINLEIN, ROBERT A.  FANTASIES OF ROBERT A. HEINLEIN, THE  LP
HEINLEIN, ROBERT A.  FARMER IN THE SKY  C/P,P,H
HEINLEIN, ROBERT A.  FARNHAM’S FREEHOLD  H:2,SFWA-P,P
HEINLEIN, ROBERT A.  FOR US, THE LIVING  H
HEINLEIN, ROBERT A.  FRIDAY  C/P,H,LP
HEINLEIN, ROBERT A.  GLORY ROAD  C/P,P,H,SFWA
HEINLEIN, ROBERT A.  GREEN HILLS OF EARTH, THE  C/P,P,H
HEINLEIN, ROBERT A.  GREEN HILLS OF EARTH, THE (DIMENSION X RADIO SHOW) (LP RECORD)  SR-MM

HEINLEIN, ROBERT A.

GRUMBLES FROM THE GRAVE R/REF-H,C/REF-H:3,

C/REF-P

AUTHOR  TITLE  CODE
HEINLEIN, ROBERT A.  HAVE SPACE SUIT — WILL TRAVEL  C/P,P,H:2
HEINLEIN, ROBERT A.  HEINLEIN TRIO, A  H:2
HEINLEIN, ROBERT A.  I WILL FEAR NO EVIL  H:2,P
HEINLEIN, ROBERT A.  JOB: A COMEDY OF JUSTICE  C/P,P,H
HEINLEIN, ROBERT A.  MAN WHO SOLD THE MOON, THE  C/P:2,B,P
HEINLEIN, ROBERT A.  MENACE FROM EARTH, THE  C/P:2,P
HEINLEIN, ROBERT A.  METHUSELAH’S CHILDREN  C/P:2,P,H
HEINLEIN, ROBERT A.  MOON IS A HARSH MISTRESS, THE  C/P:2,P
HEINLEIN, ROBERT A.  NUMBER OF THE BEAST, THE  C/P,H,LP
HEINLEIN, ROBERT A.  OFF THE MAIN SEQUENCE  H
HEINLEIN, ROBERT A.  ORPHANS OF THE SKY  H:2,P
HEINLEIN, ROBERT A.  PAST THROUGH TOMORROW, THE  C/P,P,H
HEINLEIN, ROBERT A.  PODKAYNE OF MARS  C/P,P,H
HEINLEIN, ROBERT A.  PODKAYNE OF MARS (WITH BOTH ENDINGS)  H,P
HEINLEIN, ROBERT A.  PUPPET MASTERS, THE  C/P:3,P
HEINLEIN, ROBERT A.  PUPPET MASTERS, THE (UNABRIDGED)  P
HEINLEIN, ROBERT A.  RED PLANET  C/P:2,SR-LP,P
HEINLEIN, ROBERT A.  RED PLANET (UNABRIDGED)  C/P,P
HEINLEIN, ROBERT A.  REQUIEM  C/P,P,H
HEINLEIN, ROBERT A.  REVOLT IN 2100  C/P,P,H
HEINLEIN, ROBERT A.  REVOLT IN 2100/METHUSELAH’S CHILDREN  P
HEINLEIN, ROBERT A.  REVOLTA NA LUA (VOLUME 1 OF 2)  FP
HEINLEIN, ROBERT A.  ROCKET SHIP GALILEO  C/P,P,H
HEINLEIN, ROBERT A.  ROLLING STONES, THE  C/P:2,P
HEINLEIN, ROBERT A.  6 X H  C/P:2,P
HEINLEIN, ROBERT A.  SIXTH COLUMN  C/P:2,SFWA-P,P
HEINLEIN, ROBERT A.  SPACE CADET  C/P,P,H
HEINLEIN, ROBERT A.  STAR BEAST, THE  C/P:2,P
HEINLEIN, ROBERT A.  STARMAN JONES  C/P:2,P
HEINLEIN, ROBERT A.  STARSHIP TROOPERS  C/P,P,H
HEINLEIN, ROBERT A.  STRANGER IN A STRANGE LAND  C/P,P
HEINLEIN, ROBERT A.  STRANGER IN A STRANGE LAND (UNCUT)  H:2,SFWA:2,LP
HEINLEIN, ROBERT A.  THREE BY HEINLEIN  H:2
HEINLEIN, ROBERT A.  TIME ENOUGH FOR LOVE  H:2,P
HEINLEIN, ROBERT A.  TIME FOR THE STARS  H:2,P
HEINLEIN, ROBERT A.  TO SAIL BEYOND THE SUNSET  C/P,SFWA-P,P,H
HEINLEIN, ROBERT A.  TOMORROW, THE STARS  C/PA,PA
HEINLEIN, ROBERT A.  TRAMP ROYALE  SFWA:2
HEINLEIN, ROBERT A.  TUNNEL IN THE SKY  C/P:2,P
HEINLEIN, ROBERT A.  UNIVERSE  C/P,B,TD35,P
HEINLEIN, ROBERT A.  UNPLEASANT PROFESSION OF JONATHAN HOAG, THE  C/P:2,P
HEINLEIN, ROBERT A./ROBINSON, SPIDER  VARIABLE STAR  H
HEINLEIN, ROBERT A.  WALDO/MAGIC, INC.  C/P,P,H
HEINLEIN, ROBERT A.  WALDO: GENIUS IN ORBIT  P
HEINLEIN, ROBERT A.  WORLDS OF ROBERT A. HEINLEIN, THE  C/P:2,P
HEINRICH, D. J.  DRAGON’S TOMB, THE [@DUNGEONS & DRAGONS PENHALIGON #2]  P
HEINRICH, D. J.  FALL OF MAGIC, THE [@DUNGEONS & DRAGONS PENHALIGON #3]  C/P,P
HEINRICH, D. J.  TAINTED SWORD, THE [@DUNGEONS & DRAGONS PENHALIGON #1]  P
HELFER, ANDREW  SUPERMAN: THE MAN OF STEEL [DC COMICS]  P
HELFERS, JOHN/LITTLE, DENISE  VALDEMAR COMPANION, THE [VALDEMAR]  R/REF-H
HELLER, CAROL  GATES OF VENSUNOR, THE [@SHUNLAR #1]  P
HELLER, CAROL  SANDS OF KALAVEN, THE [@SHUNLAR 2]  P
HELLER, CAROL  STONES OF STIGA, THE [@SHUNLAR 3]  P
HELMS, RANDEL  TOLKIEN’S WORLD  R/REF-H
HELPRIN, MARK  CITY IN WINTER, A  H
HELPRIN, MARK  WINTER’S TALE  C/P,P
HEMINGWAY, AMANDA  GREENSTONE GRAIL, THE  H
HEMINGWAY, AMANDA  PZYCHE  H
HEMINGWAY, AMANDA  SWORD OF STRAW, THE [SANGREAL 2]  LP
HEMRY, JOHN G.  AGAINST ALL ENEMIES [PAUL SINCLAIR 4]  P
HEMRY, JOHN G.  BURDEN OF PROOF [PAUL SINCLAIR 2]  P
HEMRY, JOHN G.  JUST DETERMINATION, A [PAUL SINCLAIR 1]  P
HEMRY, JOHN G.  RULE OF EVIDENCE [PAUL SINCLAIR 3]  P
HEMRY, JOHN G.  STARK’S COMMAND [ETHAN STARK 2]  P
HEMRY, JOHN G.  STARK’S CRUSADE [ETHAN STARK 3]  P
HEMRY, JOHN G.  STARK’S WAR [ETHAN STARK 1]  C/P,P
HENDEE, BARB/HENDEE, J. C.  DHAMPIR [NOBLE DEAD SERIES 1 1]  P
HENDEE, BARB/HENDEE, J. C.  REBEL FAY [NOBLE DEAD SERIES 1 5]  H:2
HENDEE, BARB/HENDEE, J. C.  SISTER OF THE DEAD [NOBLE DEAD SERIES 1 3]  P
HENDEE, BARB/HENDEE, J. C.  THIEF OF LIVES [NOBLE DEAD SERIES 1 2]  P
HENDEE, BARB/HENDEE, J. C.  TRAITOR TO THE BLOOD [NOBLE DEAD SERIES 1 4]  H
HENDERSON, ALICE  PORTAL THROUGH TIME [@BUFFY]  P
HENDERSON, C. J.  DOUBLE OR NOTHING [@QUANTUM LEAP]  P
HENDERSON, C. J.  OCCULT DETECTIVES OF C.J. HENDERSON, THE  LP

The MITSFS Pinkdex  September 24, 2007  Page 187
AUTHOR  TITLE  CODE
HENDERSON, C. J.  THINGS THAT ARE NOT THERE, THE  LP
HENDERSON, DAN  PARADISE  P
HENDERSON, JAMES  COPPERHEAD  H,P
HENDERSON, JASON  ELEMENT OF FIRE, THE [HIGHLANDER]  P
HENDERSON, JASON  IRON THANE, THE  P,SFWA-P
HENDERSON, JASON  SPAWN OF LOKI, THE  P
HENDERSON, ZENNA  ANYTHING BOX, THE  C/P,P,H
HENDERSON, ZENNA  HOLDING WONDER  C/P,P,H
HENDERSON, ZENNA  INGATHERING  H:2
HENDERSON, ZENNA  PEOPLE: NO DIFFERENT FLESH, THE  C/P:2,P
HENDERSON, ZENNA  PILGRIMAGE  C/P:2,P

HENDRA, TONY/CERF, CHRISTOPHER/ ELBLING, PETER ’80’S: A LOOK BACK AT THE TUMULTUOUS DECADE 1980–1989, THE

SR-VLHA

HENDRICKSON, WALTER B. JR.  CLASS G-ZERO  P
HENDRIX, HOWARD V.  BETTER ANGELS  H,LP
HENDRIX, HOWARD V.  EMPTY CITIES OF THE FULL MOON  LP
HENDRIX, HOWARD V.  LABYRINTH KEY, THE  LP
HENDRIX, HOWARD V.  LIGHTPATHS  P
HENDRIX, HOWARD V.  SPEARS OF GOD  LP
HENDRIX, HOWARD V.  STANDING WAVE  P
HENDRY, ALLAN  UFO HANDBOOK, THE  SFWA
HENIGHAN, TOM  STRANGE ATTRACTORS  P

HENNEBERG, NATHALIE/ HENNEBERG, CHARLES GREEN GODS, THE

C/P,P

HENRICK, RICHARD  VAMPIRE IN MOSCOW [ST. JOHN THE PURSUER]  P
HENSLEY, J. L.  BLACK ROADS, THE  SR-P
HENSLEY, SHANE LACY  DARKFALL [WORLD OF ADEN THUNDERSCAPE]  P
HENSLEY, SHANE LACY  SOLE SURVIVOR [SHATTERZONE 2]  P
HENSTELL, DIANA  NEW MORNING DRAGON  P
HEPPENHEIMER, T. A.  COLONIES IN SPACE  [C/RAN-P]
HERBER, KEITH  DARK PRINCE [WORLD OF DARKNESS VAMPIRE]  C/P,P
HERBER, KEITH  PRINCE OF THE CITY [WORLD OF DARKNESS VAMPIRE]  P
HERBERT, BRIAN/ANDERSON, KEVIN J.  BATTLE OF CORRIN, THE [DUNE JIHAD 3]  H
HERBERT, BRIAN/LANDIS, MARIE  BLOOD ON THE SUN [WORLD OF DARKNESS VAMPIRE]  P
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: HOUSE ATREIDES [DUNE 0.1]  H:2,P
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: HOUSE CORRINO [DUNE 0.3]  H:2,P
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: HOUSE CORRINO (6 CASSETTES, BOXED) [DUNE 0.3]  SR-MM
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: HOUSE HARKONNEN [DUNE 0.2]  H:2,P
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: THE BUTLERIAN JIHAD [DUNE JIHAD 1]  H:2
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: THE MACHINE CRUSADE [DUNE JIHAD 2]  H:2
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: THE MACHINE CRUSADE (PROOF) [DUNE JIHAD 2]  SR-LP
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: WHIPPING MEK [DUNE JIHAD]  B
HERBERT, BRIAN/ANDERSON, KEVIN J.  DUNE: WHIPPING MEK (BOOKLET AND CD) [DUNE JIHAD]  SR-MM:2
HERBERT, BRIAN  GARBAGE CHRONICLES, THE  C/P,P
HERBERT, BRIAN/ANDERSON, KEVIN J.  HUNTERS OF DUNE [DUNE 7]  H
HERBERT, BRIAN/LANDIS, MARIE  MEMORYMAKERS  P
HERBERT, BRIAN  PRISONERS OF ARIONN  C/P,SFWA,P,H
HERBERT, BRIAN  RACE FOR GOD, THE  C/P,SFWA-P,P
HERBERT, BRIAN  SIDNEY’S COMET  C/P:2,P
HERBERT, BRIAN  SUDANNA, SUDANNA  C/P:2,P,H
HERBERT, FRANK/RANSOM, BILL  ASCENSION FACTOR, THE [SHIP 4]  C/P,SFWA-P,P,H
HERBERT, FRANK  BOOK OF FRANK HERBERT, THE  C/P:2,P
HERBERT, FRANK  CHAPTERHOUSE: DUNE [DUNE 6]  C/P,SFWA-P,P,H
HERBERT, FRANK  CHILDREN OF DUNE [DUNE 3]  C/P:2,P,H
HERBERT, FRANK  DESTINATION: VOID [SHIP 1]  C/P:2,P
HERBERT, FRANK  DIRECT DESCENT  C/P:2,P
HERBERT, FRANK  DOSADI EXPERIMENT, THE  C/P,P,H
HERBERT, FRANK  DRAGON IN THE SEA, THE  H:2,P
HERBERT, FRANK  DUNE [DUNE 1]  C/P:3,SFWA-P,P,H
HERBERT, FRANK  DUNE MESSIAH [DUNE 2]  C/P,P,H
HERBERT, FRANK  EYE  H:2,SFWA-P,P
HERBERT, FRANK  EYES OF HEISENBERG, THE  C/P:2,P
HERBERT, FRANK  GOD EMPEROR OF DUNE [DUNE 4]  C/P,P,H,SFWA
HERBERT, FRANK  GODMAKERS, THE  C/P,P,H
HERBERT, FRANK  GREEN BRAIN, THE  C/P:2,P
HERBERT, FRANK  HEAVEN MAKERS, THE  C/P:2,P
HERBERT, FRANK  HEAVEN MAKERS, THE (REVISED)  C/P,P
HERBERT, FRANK  HELLSTROM’S HIVE  C/P,P,H
HERBERT, FRANK  HERETICS OF DUNE [DUNE 5]  C/P,LP,H
HERBERT, FRANK  ILLUSTRATED DUNE, THE [DUNE 1]  LP
HERBERT, FRANK/RANSOM, BILL  JESUS INCIDENT, THE [SHIP 2]  C/P,P,H,SFWA
HERBERT, FRANK/RANSOM, BILL  LAZARUS EFFECT, THE [SHIP 3]  C/P,P,H
HERBERT, FRANK/HERBERT, BRIAN  MAN OF TWO WORLDS  C/P,SFWA-P,P,H
HERBERT, FRANK  NEBULA WINNERS 15  C/PA,PA,SFWA

The MITSFS Pinkdex  September 24, 2007  Page 188
AUTHOR  TITLE  CODE

HERBERT, FRANK/HERBERT, BRIAN/ ANDERSON, KEVIN J.

ROAD TO DUNE, THE [DUNE]

R/REF-H

HERBERT, FRANK  SANTAROGA BARRIER, THE  C/P:2,P
HERBERT, FRANK  SOUL CATCHER  C/P:2,P
HERBERT, FRANK  TWENTY-FIRST CENTURY SUB  C/P:2,P
HERBERT, FRANK  UNDER PRESSURE  C/P:2,P
HERBERT, FRANK  WHIPPING STAR  C/P,P,H
HERBERT, FRANK  WHITE PLAGUE, THE  C/P,P,H
HERBERT, FRANK  WORLDS OF FRANK HERBERT, THE  C/P:2,P
HERBERT, JAMES  DARK, THE  P
HERBERT, JAMES  DOMAIN  P
HERBERT, JAMES  FLUKE  P
HERBERT, JAMES  FOG, THE  P
HERBERT, JAMES  ’48  H
HERBERT, JAMES  GHOSTS OF SLEATH, THE  H,P
HERBERT, JAMES  GHOSTS OF SLEATH, THE (PROOF)  SR-LP
HERBERT, JAMES  LAIR  P
HERBERT, JAMES  MAGIC COTTAGE, THE  H
HERBERT, JAMES  ONCE...  H
HERBERT, JAMES  PORTENT  H,LP
HERBERT, JAMES  SPEAR, THE  P
HERBERT, JAMES  SURVIVOR, THE  P
HERBERT, MARY H.  CITY OF THE LOST [@DRAGONLANCE LINSHA #1]  P
HERBERT, MARY H.  CITY OF THE SORCERERS [DARK HORSE 3]  P
HERBERT, MARY H.  CLANDESTINE CIRCLE, THE [@DRAGONLANCE CHRONICLES]  P
HERBERT, MARY H.  DARK HORSE [DARK HORSE 1]  P
HERBERT, MARY H.  DRAGON’S BLUFF [@DRAGONLANCE CROSSROADS]  P
HERBERT, MARY H.  FLIGHT OF THE FALLEN [@DRAGONLANCE LINSHA #2]  P
HERBERT, MARY H.  LIGHTNING’S DAUGHTER [DARK HORSE 2]  P
HERBERT, MARY H.  RETURN OF THE EXILE [@DRAGONLANCE LINSHA #3]  P
HERBERT, MARY H.  VALORIAN [DARK HORSE 0]  P
HERBERT, MARY H.  VALORIAN’S CHILDREN [DARK HORSE 1,2]  H
HERBERT, MARY H.  VALORIAN’S LEGACY [DARK HORSE 3,4]  H
HERBERT, MARY H.  WINGED MAGIC [DARK HORSE 4]  P
HERLEY, RICHARD  PENAL COLONY, THE  P
HERMAN, J. B.  BLACK SABBAT  P
HERNDON, CORY J.  DISSENSION [@MAGIC: THE GATHERING RAVNICA 3]  P
HERNDON, CORY J.  FIFTH DAWN, THE [@MAGIC: THE GATHERING MIRRODIN #3]  P
HERNDON, CORY J.  GUILDPACT [MAGIC: THE GATHERING RAVNICA 2]  P
HERNDON, CORY J.  RAVNICA: CITY OF GUILDS [MAGIC: THE GATHERING RAVNICA 1]  P
HERNIMAN, MARCUS  SIEGE OF ARRANDIN, THE [@ARRANDIN #1]  P
HERNIMAN, MARCUS  TREASON OF DORTREAN, THE [@ARRANDIN #2]  P
HERSEY, JOHN  CHILD BUYER, THE  C/P,P,H
HERSEY, JOHN  MY PETITION FOR MORE SPACE  C/P,H
HERSEY, JOHN  WHITE LOTUS  H:2

HERSHEY, JENNIFER/DUPRE, TOM/ SILVERSTEIN, JANNA FULL SPECTRUM 5

LPA

HERSHMAN, MORRIS  CRASH OF 2086, THE  P
HERSHMAN, MORRIS  SHAREWORLD  H,SFWA
HERTER, DAVID  CERES STORM  H,P
HERTER, DAVID  EVENING’S EMPIRE  H,LP
HERTZLER, J. G./LANG, JEFFREY  LEFT HAND OF DESTINY BOOK ONE, THE [@STAR TREK DS9]  P
HERTZLER, J. G./LANG, JEFFREY  LEFT HAND OF DESTINY BOOK TWO, THE [@STAR TREK DS9]  P
HERVEY, MICHAEL  FRATERNITY OF THE WEIRD  [C/RAN-P]
HERZOG, ARTHUR  HEAT  P
HERZOG, ARTHUR  SWARM, THE  P
HESSE, HERMAN  STEPPENWOLF  C/P,P
HETLEY, JAMES A.  DRAGON’S EYE [STONEFORT,MAINE 1]  LP
HETLEY, JAMES A.  DRAGON’S TEETH [STONEFORT,MAINE 2]  LP
HETLEY, JAMES A.  SUMMER COUNTRY, THE [MAUREEN PIERCE 1]  LP
HETLEY, JAMES A.  WINTER OAK, THE [MAUREEN PIERCE 2]  LP
HEWSON, DAVID  SOLSTICE  H
HEYDT, DOROTHY J.  POINT OF HONOR, A  C/P:2,P
HEYERDAHL, THOR  AKU-AKU  [C/RAN-P]
HEYWOOD, VICTOR D.  PRISON PLANET  SR-P
HICHENS, ROBERT  DWELLER ON THE THRESHOLD, THE  H
HICHENS, ROBERT  UNEARTHLY, THE  H
HICKAM, HOMER H. JR.  BACK TO THE MOON  P
HICKMAN, STEPHEN F.  LEMURIAN STONE, THE  C/P,P
HICKMAN, TRACY  IMMORTALS, THE  H,P
HICKMAN, TRACY/HICKMAN, LAURA  MYSTIC EMPIRE [@BRONZE CANTICLES #3]  H
HICKMAN, TRACY/HICKMAN, LAURA  MYSTIC QUEST [@BRONZE CANTICLES #2]  H:2
HICKMAN, TRACY/HICKMAN, LAURA  MYSTIC WARRIOR [@BRONZE CANTICLES #1]  P
HICKMAN, TRACY  REQUIEM OF STARS [SONGS OF THE STELLAR WIND 1]  P

AUTHOR  TITLE  CODE
HICKMAN, TRACY  SPEED OF DARKNESS [@STARCRAFT #3]  P
HICKS, CLIFFORD B.  FIRST BOY ON THE MOON  H:2
HICKS, GRANVILLE/BENNETT, RICHARD M.  FIRST TO AWAKEN, THE  H
HIGH, PHILIP E.  INVADER ON MY BACK  DH085
HIGH, PHILIP E.  MAD METROPOLIS, THE  P,DM135
HIGH, PHILIP E.  NO TRUCE WITH TERRA  DF275
HIGH, PHILIP E.  PRODIGAL SUN, THE  C/P:2,P
HIGH, PHILIP E.  REALITY FORBIDDEN  P,DG609
HIGH, PHILIP E.  THESE SAVAGE FUTURIANS  DG623
HIGH, PHILIP E.  TIME MERCENARIES, THE  DH059
HIGH, PHILIP E.  TWIN PLANETS  C/P:2,P
HIGHTMAN, JASON  SAINT OF THE DRAGONS, THE  H
HIGHTOWER, LYNN S.  ALIEN BLUES [ELAKI 1]  P,SFWA-P
HIGHTOWER, LYNN S.  ALIEN EYES [ELAKI 2]  C/P,P
HIGHTOWER, LYNN S.  ALIEN HEAT [ELAKI 3]  P
HIGHTOWER, LYNN S.  ALIEN RITES [ELAKI 4]  P
HILDEBRANDT, BROTHERS/NICHOLS, JERRY URSHURAK  H,SR-LP
HILDICK, E. W.  ACTIVE-ENZYME LEMON-FRESHENED JUNIOR HIGH SCHOOL WITCH, THE  LP
HILGARTNER, BETH  BUSINESS OF FERRETS, A  LP
HILGARTNER, BETH  CATS IN CYBERSPACE  LP
HILL, ALBERT FAY/HILL, DAVID CAMPBELL  INVADER  C/P,P
HILL, CAROL  ELEVEN MILLION MILE HIGH DANCER, THE  H:2,LP
HILL, DAVID CAMPBELL/HILL, ALBERT FAY  DEADLY MESSIAH, THE  C/P,P
HILL, DOUGLAS  ALIEN CITADEL [HUNTSMAN 3]  H
HILL, DOUGLAS  BLADE OF THE POISONER  P
HILL, DOUGLAS  CAVES OF KLYDOR, THE [COLSEC 2]  C/P,P,H
HILL, DOUGLAS  COLSEC REBELLION [COLSEC 3]  H,P
HILL, DOUGLAS  DAY OF THE STARWIND [LAST LEGIONARY 3]  H,P
HILL, DOUGLAS  DEATHWING OVER VEYNAA [LAST LEGIONARY 2]  H,P
HILL, DOUGLAS  DEVIL HIS DUE, THE  C/PA,PA
HILL, DOUGLAS  EXILES OF COLSEC [COLSEC 1]  C/P,P,H
HILL, DOUGLAS  GALACTIC WARLORD [LAST LEGIONARY 1]  H:2,P
HILL, DOUGLAS  HUNTSMAN, THE [HUNTSMAN 1]  H
HILL, DOUGLAS  LIGHTLESS DOME, THE [APOTHEOSIS 1]  P
HILL, DOUGLAS  PLANET OF THE WARLORD [LAST LEGIONARY 4]  H
HILL, DOUGLAS  SHAPE OF SEX TO COME, THE  SR-PA
HILL, DOUGLAS/WILLIAMS, PAT  SUPERNATURAL, THE  [C/RAN-H]
HILL, DOUGLAS  WARRIORS OF THE WASTELAND [HUNTSMAN 2]  H
HILL, DOUGLAS  WINDOW ON THE FUTURE  SR-PA
HILL, DOUGLAS  YOUNG LEGIONARY [LAST LEGIONARY 0]  H
HILL, ERNEST  PITY ABOUT EARTH  DH056
HILL, JOE  HEART-SHAPED BOX  H
HILL, JOHN  HEARTBEEPS  P
HILL, JOHN  LONG SLEEP, THE  P
HILL, RUSSELL  EDGE OF THE EARTH, THE  P
HILLEGAS, MARK R.  FUTURE AS NIGHTMARE, THE  R/REF-H
HILLEGAS, MARK R.  SHADOWS OF IMAGINATION  R/REF-H
HILSON, PHOENIX  WIZARD’S WAR  SR-P
HILTON, JAMES  LOST HORIZON  C/P,SR-P,H
HILTON, JAMES  ORIZZONTE PERDUTO  FP
HILTON, MARGERY  FLOWER OF ETERNITY, THE  P
HINCHCLIFFE, PHILIP  DOCTOR WHO AND THE KEYS OF MARINUS [DOCTOR WHO]  SR-P,C/P:2
HINCHCLIFFE, PHILIP  DOCTOR WHO AND THE MASQUE OF MANDRAGORA [DOCTOR WHO]  C/P,P
HINCHCLIFFE, PHILIP  DOCTOR WHO AND THE SEEDS OF DOOM [DOCTOR WHO]  C/P:2,P
HINE, AL  BEWITCHED  P
HINES, JIM C.  GOBLIN HERO [JIG THE GOBLIN 2]  P
HINES, JIM C.  GOBLINQUEST [JIG THE GOBLIN 1]  H
HINES, LEONARD F.  TIME SNATCHERS  LP
HINKEMEYER, M. THOMAS  CREATOR, THE  P
HINTON, CRAIG  CRYSTAL BUCEPHALUS, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
HINTON, CRAIG  GODENGINE [DOCTOR WHO NEW ADVENTURES]  P
HINTON, CRAIG  MILLENNIAL RITES [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
HINTON, CRAIG  QUANTUM ARCHANGEL, THE [DOCTOR WHO PAST DOCTORS]  P
HINTON, CRAIG  SYNTHESPIANS[TM] [@DOCTOR WHO PAST DOCTORS]  P
HINTZE, NAOMI A.  STONE CARNATION, THE  P
HINZ, CHRISTOPHER  ANACHRONISMS  P
HINZ, CHRISTOPHER  ASH OCK [PARATWA 2]  H,P
HINZ, CHRISTOPHER  LIEGE-KILLER [PARATWA 1]  C/P:2,P,H
HINZ, CHRISTOPHER  PARATWA, THE [PARATWA 3]  SFWA
HIPOLITO, JANE/MCNELLY, WILLIS E.  MARS, WE LOVE YOU  HA,PA,SFWA
HIPPLE, THEODORE W./WRIGHT, ROBERT G.  WORLDS OF SCIENCE FICTION, THE  LPA
HIRSCH, PHIL  SUPERNATURAL  [C/RAN-P]
HIRSCHFELD, BURT  GAS!  C/P,P
HITCHCOCK, ALFRED  ANTI-SOCIAL REGISTER  PA

The MITSFS Pinkdex  September 24, 2007  Page 190
AUTHOR  TITLE  CODE
HITCHCOCK, ALFRED  DOWN BY THE OLD BLOODSTREAM  C/PA,PA
HITCHCOCK, ALFRED  HANGMAN’S DOZEN, A  PA
HITCHCOCK, ALFRED  NOOSE REPORT  PA
HITCHCOCK, ALFRED  SLAY RIDE  C/PA,PA
HITCHCOCK, ALFRED  12 STORIES THEY WOULDN’T LET ME DO ON TV  PA
HITCHCOCK, RAYMOND  VENUS 13  SR-P
HJALMARSSON, KJELL  FORTECKNING PA JULES VERNE MAGASINET  R/IND
HJORTSBERG, WILLIAM  FALLING ANGEL  P
HJORTSBERG, WILLIAM  GRAY MATTERS  H:2,P
HLASKO, MAREK  EIGHTH DAY OF THE WEEK, THE  H
HOARE, ANDY  ROGUE STAR [@WARHAMMER 40K]  P
HOBAN, RUSSELL  LION OF BOAZ-JACHIN AND JACHIN-BOAZ, THE  C/P,P
HOBAN, RUSSELL  MEDUSA FREQUENCY, THE  LP
HOBAN, RUSSELL  MOUSE AND HIS CHILD, THE  P
HOBAN, RUSSELL  PILGERMANN  LP
HOBAN, RUSSELL  RIDDLEY WALKER  H:2,LP
HOBB, ROBIN  ASSASSIN’S APPRENTICE [FARSEER 1]  C/P:3,P
HOBB, ROBIN  ASSASSIN’S QUEST [FARSEER 3]  H,P
HOBB, ROBIN  FOOL’S ERRAND [TAWNY MAN 1]  C/P,P
HOBB, ROBIN  FOOL’S FATE [TAWNY MAN 3]  H
HOBB, ROBIN  FOREST MAGE [SOLDIER SON 2]  H
HOBB, ROBIN  GOLDEN FOOL [TAWNY MAN 2]  H,P
HOBB, ROBIN  MAD SHIP [LIVESHIP TRADERS 2]  C/P,P,H
HOBB, ROBIN  ROYAL ASSASSIN [FARSEER 2]  C/P:2,P
HOBB, ROBIN  SHAMAN’S CROSSING [SOLDIER SON 1]  H
HOBB, ROBIN  SHIP OF DESTINY [LIVESHIP TRADERS 3]  H
HOBB, ROBIN  SHIP OF MAGIC [LIVESHIP TRADERS 1]  H:2,P
HOCH, EDWARD D.  CITY OF BRASS  C/P,P
HOCH, EDWARD D.  DEAR DEAD DAYS  SFWA
HOCH, EDWARD D.  FELLOWSHIP OF THE HAND, THE  H,SFWA,P
HOCH, EDWARD D.  FRANKENSTEIN FACTORY, THE  P
HOCH, EDWARD D.  JUDGES OF HADES, THE  SR-P,C/P
HOCH, EDWARD D.  TRANSVECTION MACHINE, THE  H,SFWA,P
HOCKE, MARTIN  ANCIENT SOLITARY REIGN, THE  H
HOCKING, JOHN C.  CONAN AND THE EMERALD LOTUS [CONAN 38]  LP
HODDER-WILLIAMS, CHRISTOPHER  EGG-SHAPED THING, THE  H:2
HODDER-WILLIAMS, CHRISTOPHER  FISTFUL OF DIGITS, A  SR-P,C/P
HODDER-WILLIAMS, CHRISTOPHER  MAIN EXPERIMENT, THE  H:2,P
HODDER-WILLIAMS, CHRISTOPHER  98.4  SR-P
HODDER-WILLIAMS, CHRISTOPHER  PRAYER MACHINE, THE  H
HODEL, MIKE/GILDEN, MEL  HOUR 25 SHOW, THE (5 CASSETTES)  SR-MM
HODES, RICHARD  PROJECT MEGALON  LP
HODGART, MATTHEW  NEW VOYAGE TO THE COUNTRY OF THE HOUYHNHNMS, A  H
HODGE, BRIAN  OASIS  SFWA-P
HODGELL, P. C.  BLOOD & IVORY [JAMETHIEL DREAM-WEAVER]  LP
HODGELL, P. C.  DARK OF THE MOON [JAMETHIEL DREAM-WEAVER 2]  H:2,SFWA-P,P
HODGELL, P. C.  GOD STALK [JAMETHIEL DREAM-WEAVER 1]  H:2,P
HODGELL, P. C.  SEEKER’S MASK [JAMETHIEL DREAM-WEAVER 3]  SR-H,H:2
HODGELL, P. C.  TO RIDE A RATHORN [JAMETHIEL DREAM-WEAVER 4]  H
HODGMAN, ANN  ADDAMS FAMILY VALUES [ADDAMS FAMILY]  P
HODGMAN, ANN  DARK DREAMS [CHILDREN OF THE NIGHT 1]  P
HODGMAN, ANN  DARK MUSIC [CHILDREN OF THE NIGHT 2]  P

HODGMAN, CHARLES D./ LANGE, NORBERT A.

HANDBOOK OF CHEMISTRY AND PHYSICS (11TH EDITION)

[C/RAN-P]

HODGSON, WILLIAM HOPE  AGUAS PROFUNDAS  FP
HODGSON, WILLIAM HOPE  BOATS OF THE "GLEN CARRIG", THE  C/P:2,LP
HODGSON, WILLIAM HOPE  CARNACKI THE GHOST-FINDER  SR-P,C/P
HODGSON, WILLIAM HOPE  DEEP WATERS  SR-H
HODGSON, WILLIAM HOPE  DREAM OF X, THE  SR-VLH
HODGSON, WILLIAM HOPE  GHOST PIRATES, THE  SR-P,LP
HODGSON, WILLIAM HOPE  GHOST PIRATES AND OTHER REVENANTS OF THE SEA, THE  H
HODGSON, WILLIAM HOPE  HOUSE ON THE BORDERLAND, THE  C/P,LP,H
HODGSON, WILLIAM HOPE  MASTERS OF TERROR I  SR-P
HODGSON, WILLIAM HOPE  NIGHT LAND 1, THE  P
HODGSON, WILLIAM HOPE  NIGHT LAND 2, THE  C/P,P
HODGSON, WILLIAM HOPE  NIGHT LAND AND OTHER PERILOUS ROMANCES, THE  H
HODGSON, WILLIAM HOPE  OUT OF THE STORM  H
HOEG, PETER  SMILLA’S SENSE OF SNOW  C/P,P
HOFFER, ERIC  TRUE BELIEVER, THE  [C/RAN-P:2]
HOFFMAN, ALICE  PRACTICAL MAGIC  P
HOFFMAN, LEE  ALWAYS THE BLACK KNIGHT  C/P:2,P
HOFFMAN, LEE  BRED TO KILL  P
HOFFMAN, LEE  CAVES OF KARST, THE  C/P,P
HOFFMAN, LEE  CHANGE SONG  H,SFWA

AUTHOR  TITLE  CODE
HOFFMAN, LEE  DEAD MAN’S GOLD  P
HOFFMAN, LEE  GUNFIGHT AT LARAMIE  P
HOFFMAN, LEE  IN AND OUT OF QUANDRY  SR-DNESFA2
HOFFMAN, LEE  LEGEND OF BLACKJACK SAM, THE  P
HOFFMAN, LEE  RETURN TO BROKEN CROSSING  P
HOFFMAN, LEE  TELEPOWER  SR-DB0779
HOFFMAN, LEE  VALDEZ HORSES, THE  P
HOFFMAN, LEE  YARBOROUGH BRAND, THE  C/P,P
HOFFMAN, NINA KIRIKI  CATALYST  LP
HOFFMAN, NINA KIRIKI  FISTFUL OF SKY, A  P
HOFFMAN, NINA KIRIKI  PAST THE SIZE OF DREAMING [MATILDA BLACK 2]  P
HOFFMAN, NINA KIRIKI  RED HEART OF MEMORIES, A [MATILDA BLACK]  LP
HOFFMAN, NINA KIRIKI  SILENT STRENGTH OF STONES, THE  P
HOFFMAN, NINA KIRIKI  SPIRITS THAT WALK IN SHADOW  H
HOFFMAN, NINA KIRIKI  THREAD THAT BINDS THE BONES, THE  P,SFWA-P
HOFFMANN, CURTIS H.  PROJECT: MILLENNIUM  C/P,P
HOFFMANN, E. T. A.  BEST TALES OF HOFFMANN, THE  LP
HOGAN, ERNEST  CORTEZ ON JUPITER  P,SFWA-P
HOGAN, ERNEST  HIGH AZTECH  P
HOGAN, JAMES P.  ANGUISHED DAWN, THE [CRADLE OF SATURN 2]  P
HOGAN, JAMES P.  BUG PARK  C/P,P,H
HOGAN, JAMES P.  CODE OF THE LIFEMAKER  H:2,P
HOGAN, JAMES P.  CRADLE OF SATURN [CRADLE OF SATURN 1]  H:2,P
HOGAN, JAMES P.  EHOES OF AN ALIEN SKY  H
HOGAN, JAMES P.  ENDGAME ENIGMA  C/P,P,H
HOGAN, JAMES P.  ENTOVERSE [GIANTS 4]  H:2,P
HOGAN, JAMES P.  GENESIS MACHINE, THE  C/P,SR-P,H
HOGAN, JAMES P.  GENTLE GIANTS OF GANYMEDE, THE [GIANTS 2]  C/P:2,SR-P
HOGAN, JAMES P.  GIANTS NOVELS, THE [GIANTS 1,2,3]  C/P,P,H
HOGAN, JAMES P.  GIANTS’ STAR [GIANTS 3]  C/P:2,P
HOGAN, JAMES P.  IMMORTALITY OPTION, THE  C/P,P,H
HOGAN, JAMES P.  INHERIT THE STARS [GIANTS 1]  C/P:2,SR-P
HOGAN, JAMES P.  LEGEND THAT WAS EARTH, THE  H
HOGAN, JAMES P.  MARTIAN KNIGHTLIFE  H,P
HOGAN, JAMES P.  MINDS, MACHINES AND EVOLUTION  C/P:2,P
HOGAN, JAMES P.  MINERVAN EXPERIMENT, THE [GIANTS 1,2,3]  H:2
HOGAN, JAMES P.  MIRROR MAZE, THE  C/P:2,P
HOGAN, JAMES P.  MISSION TO MINERVA [GIANTS 5]  H
HOGAN, JAMES P.  MULTIPLEX MAN, THE  H:2,P
HOGAN, JAMES P.  MULTIPLEX MAN, THE (PROOF)  SR-LP
HOGAN, JAMES P.  OUT OF TIME  P
HOGAN, JAMES P.  OUTWARD BOUND  P
HOGAN, JAMES P.  PATHS TO OTHERWHERE  H
HOGAN, JAMES P.  PATHS TO OTHERWHERE (PROOF)  SR-LP
HOGAN, JAMES P.  PROTEUS OPERATION, THE  H:2,P,SFWA
HOGAN, JAMES P.  REALTIME INTERRUPT  C/P,P,H
HOGAN, JAMES P.  ROCKETS, REDHEADS & REVOLUTION  C/P:2,P
HOGAN, JAMES P.  STAR CHILD  C/P:2,P
HOGAN, JAMES P.  THRICE UPON A TIME  C/P:2,P
HOGAN, JAMES P.  THRICE UPON A TIME (PROOF)  SR-LP
HOGAN, JAMES P.  TWO FACES OF TOMORROW, THE  C/P:2,P
HOGAN, JAMES P.  VOYAGE FROM YESTERYEAR  C/P,P,H
HOGAN, LEE  BELARUS  P
HOGAN, LEE  ENEMIES  P
HOGAN, ROBERT J.  ACE OF THE WHITE DEATH [@G-8 #3]  SR-P
HOGAN, ROBERT J.  BAT STAFFEL, THE [@G-8 #1]  SR-P
HOGAN, ROBERT J.  BOMBS FROM THE MURDER WOLVES [@G-8 #4]  SR-P
HOGAN, ROBERT J.  FLIGHT FROM THE GRAVE [@G-8 #6]  P
HOGAN, ROBERT J.  MARK OF THE VULTURE, THE [@G-8 #8]  SR-P
HOGAN, ROBERT J.  PURPLE ACES [@G-8 #2]  P
HOGAN, ROBERT J.  SCOURGE OF THE STEEL MASK [@G-8 #1]  P
HOGAN, ROBERT J.  VULTURES OF THE WHITE DEATH [@G-8 #5]  SR-P
HOKE, HELEN  SPOOKS, SPOOKS, SPOOKS  HA
HOLBERG, LUDVIG  JOURNEY OF NIELS KLIM TO THE WORLD UNDERGROUND  LP
HOLDEN, RICHARD CORT  SNOW FURY  H:2,P
HOLDER, NANCY  ANGEL CHRONICLES, THE, VOL. 1 [BUFFY]  C/P,P
HOLDER, NANCY  ANGEL CHRONICLES, THE, VOL. 3 [BUFFY]  C/P,P
HOLDER, NANCY  BLOOD AND FOG [@BUFFY]  P
HOLDER, NANCY  BOOK OF FOURS, THE [@BUFFY]  H:2,P
HOLDER, NANCY/MARIOTTE, JEFF  BURNING, THE [@BUFFY ANGEL UNSEEN 1]  P
HOLDER, NANCY  CARNIVAL OF SOULS [@BUFFY]  P
HOLDER, NANCY  CHOSEN [@BUFFY]  P
HOLDER, NANCY  CITY OF [@ANGEL]  P
HOLDER, NANCY/MARIOTTE, JEFF  DOOR TO ALTERNITY [@BUFFY ANGEL UNSEEN 2]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 192
AUTHOR  TITLE  CODE
HOLDER, NANCY/MARIOTTE, JEFF  ENDANGERED SPECIES [@ANGEL]  P
HOLDER, NANCY  EVIL THAT MEN DO, THE [@BUFFY]  C/P,P
HOLDER, NANCY  HAUNTINGS [@SMALLVILLE]  P
HOLDER, NANCY  HEAT [@BUFFY/ANGEL]  P
HOLDER, NANCY  INVASIONS [@GAMBLER’S STAR #3]  P
HOLDER, NANCY  LEGACIES AND LIES [@GAMBLER’S STAR #2]  P
HOLDER, NANCY/MARIOTTE, JEFF  LONG WAY HOME [@BUFFY ANGEL UNSEEN 3]  P
HOLDER, NANCY  MEASURE OF A MAN [@HIGHLANDER]  P
HOLDER, NANCY  NOT FORGOTTEN [@ANGEL]  P
HOLDER, NANCY  QUEEN OF THE SLAYERS [@BUFFY]  H,LP
HOLDER, NANCY  SCARABIAN NIGHTS [@SABRINA THE TEENAGE WITCH #24]  P
HOLDER, NANCY  SILENCE [@SMALLVILLE]  P
HOLDER, NANCY  SIX FAMILIES, THE [@GAMBLER’S STAR #1]  P
HOLDER, NANCY  SPYING EYES [@SABRINA THE TEENAGE WITCH #14]  P

HOLDSTOCK, ROBERT P./ EDWARDS, MALCOLM ALIEN LANDSCAPES

A

HOLDSTOCK, ROBERT P.  ANCIENT ECHOES  C/P,P
HOLDSTOCK, ROBERT P.  BONE FOREST, THE  P,SFWA-P
HOLDSTOCK, ROBERT P.  BROKEN KINGS, THE [MERLIN CODEX 3]  H
HOLDSTOCK, ROBERT P.  CELTIKA [MERLIN CODEX 1]  P
HOLDSTOCK, ROBERT P.  EARTHWIND  C/P:2,P
HOLDSTOCK, ROBERT P.  ENCYCLOPEDIA OF SCIENCE FICTION  R/REF-H
HOLDSTOCK, ROBERT P.  EYE AMONG THE BLIND  C/P,SFWA,P,H
HOLDSTOCK, ROBERT P.  GATE OF IVORY, GATE OF HORN [RYHOPE WOOD 3]  P
HOLDSTOCK, ROBERT P.  HOLLOWING, THE [RYHOPE WOOD 2]  H,P
HOLDSTOCK, ROBERT P.  IRON GRAIL, THE [MERLIN CODEX 2]  P
HOLDSTOCK, ROBERT P.  LAVONDYSS  H:2,P
HOLDSTOCK, ROBERT P.  MYTHAGO WOOD [RYHOPE WOOD 1]  H:2,SFWA,P
HOLDSTOCK, ROBERT P.  NECROMANCER  P
HOLDSTOCK, ROBERT P.  UNKNOWN REGIONS  P
HOLDSTOCK, ROBERT P.  WHERE TIME WINDS BLOW  H:2,P
HOLLAND, CECELIA  ANGEL AND THE SWORD, THE  LP
HOLLAND, CECELIA  FLOATING WORLDS  C/P,P,H
HOLLAND, CECELIA  PILLAR OF THE SKY  H
HOLLAND, JACK  FIRE QUEEN, THE  P,SFWA-P
HOLLAND, TOM  LORD OF THE DEAD  H,LP
HOLLO, ANSELM  SPACE BALTIC  OVD-2
HOLLOW, JOHN  AGAINST THE NIGHT, THE STARS  R/REF-H
HOLLY, EMMA  CATCHING MIDNIGHT  H
HOLLY, EMMA  DEMON’S DAUGHTER, THE  H
HOLLY, JOAN HUNTER  ASSASSINATION AFFAIR, THE [@MAN FROM UNCLE #10]  C/P:2,P
HOLLY, JOAN HUNTER  DARK PLANET, THE  H,P
HOLLY, JOAN HUNTER  DEATH DOLLS OF LYRA, THE  P
HOLLY, JOAN HUNTER  ENCOUNTER  C/P,P,H
HOLLY, JOAN HUNTER  FLYING EYES, THE  C/P,P
HOLLY, JOAN HUNTER  GRAY ALIENS, THE  H
HOLLY, JOAN HUNTER  GREEN PLANET, THE  C/P:2,P
HOLLY, JOAN HUNTER  KEEPER  SR-P
HOLLY, JOAN HUNTER  MIND TRADERS, THE  C/P,P,H
HOLLY, JOAN HUNTER  RUNNING MAN, THE  C/P:2,P
HOLLY, JOAN HUNTER  SHEPHERD  SR-P
HOLLY, JOAN HUNTER  TIME TWISTERS, THE  C/P,P
HOLMES, BRUCE T.  ANVIL OF THE HEART  P
HOLMES, JOHN ERIC  MAHARS OF PELLUCIDAR  P
HOLMES, JOHN ERIC  MAZE OF PERIL, THE  LP

HOLMES, JOHN ERIC/NIVEN, LARRY/ POURNELLE, JERRY MORDRED [BUCK ROGERS (NIVEN & POURNELLE’S) 2]

C/P,P

HOLMES, ROBERT  TWO DOCTORS, THE [DOCTOR WHO]  SR-P,C/P
HOLMSTEN, BRIAN/LUBERTOZZI, ALEX  COMPLETE WAR OF THE WORLDS, THE  SR-MM
HOLT, CONRAD G.  COSMIC EXODUS  SR-P
HOLT, JOHN R.  WHEN WE DEAD AWAKEN  P
HOLT, JOHN R.  WOLF MOON  P
HOLT, TOM  EXPECTING BEOWULF  LP
HOLT, TOM  EXPECTING SOMEONE TALLER  C/P:3,SFWA-P:2,P
HOLT, TOM  FALLING SIDEWAYS  H
HOLT, TOM  FAUST AMONG EQUALS  P
HOLT, TOM  FLYING DUTCH  H,P
HOLT, TOM  GRAILBLAZERS  P
HOLT, TOM  HERE COMES THE SUN  P
HOLT, TOM  IN YOUR DREAMS [PAUL CARPENTER 2]  H
HOLT, TOM  LITTLE PEOPLE  H
HOLT, TOM  NOTHING BUT BLUE SKIES  P
HOLT, TOM  ODDS AND GODS  P
HOLT, TOM  ONLY HUMAN  P

AUTHOR  TITLE  CODE
HOLT, TOM  OVERTIME  P
HOLT, TOM  PORTABLE DOOR, THE [PAUL CARPENTER 1]  H
HOLT, TOM  SECOND TOM HOLT OMNIBUS, THE  H
HOLT, TOM  SNOW WHITE AND THE SEVEN SAMURAI  C/P,P
HOLT, TOM  VALHALLA  P
HOLT, TOM  WHO’S AFRAID OF BEOWULF?  H:2,P
HOLT, TOM  WISH YOU WERE HERE  P
HOLTON, LEONARD  PACT WITH SATAN, A  P
HOLZER, HANS  ENTRY, THE  P
HOLZER, HANS  PSYCHIC WORLD OF BISHOP PIKE, THE  [C/RAN-P]
HOLZER, HANS  RED CHINDVIT CONSPIRACY, THE  P
HOLZER, HANS  SIGNS OF LOVE AND GLORY  [C/RAN-P]
HOME-GALL, EDWARD R.  HUMAN BAT, THE  SR-P
HOME-GALL, EDWARD R.  HUMAN BAT VS. THE ROBOT GANGSTER, THE  SR-P
HOME, WILLIAM SCOTT/FABIAN, STEPHEN E. HOLLOW FACES, MERCILESS MOONS  VLP
HOMER  ODYSSEY, THE  P
HONE, JOSEPH  SIXTH DIRECTORATE, THE  P
HOOD, DANIEL  BEGGAR’S BANQUET [FANUILH 3]  P
HOOD, DANIEL  FAMILIAR DRAGON, THE [FANUILH 1,2,3]  H
HOOD, DANIEL  FANUILH [FANUILH 1]  C/P,P
HOOD, DANIEL  KING’S CURE [FANUILH 5]  P
HOOD, DANIEL  SCALES OF JUSTICE [FANUILH 4]  P
HOOD, DANIEL  WIZARD’S HEIR [FANUILH 2]  P
HOOD, GWENYTH  COMING OF THE DEMONS, THE  H
HOOD, KEN  DEMON KNIGHT [LONGDIRK 3]  P
HOOD, KEN  DEMON RIDER [LONGDIRK 2]  P
HOOD, KEN  DEMON SWORD [LONGDIRK 1]  P
HOOVER, H. M.  ANOTHER HEAVEN, ANOTHER EARTH  LP
HOOVER, H. M.  CHILDREN OF MORROW  P
HOOVER, H. M.  DELIKON, THE  C/P,P
HOOVER, H. M.  LOST STAR, THE  C/P,P
HOOVER, H. M.  RAINS OF ERIDAN, THE  C/P,P,H
HOOVER, H. M.  RETURN TO EARTH  C/P,P
HOOVER, H. M.  THIS TIME OF DARKNESS  H
HOPE, ANTHONY  PRISONER OF ZENDA, THE [ZENDA 1]  SR-H
HOPE, ANTHONY  RUPERT OF HENTZAU [ZENDA 2]  SR-H
HOPEN, STUART  WARP ANGEL  H,P
HOPKINS, ANDREA  BOOK OF GUINEVERE, THE  H
HOPKINS, BRIAN A.  DIA DE LOS MUERTOS, EL  SR-H
HOPKINS, JACK  SATELLITE NIGHT FEVER [SATELLITE NIGHT 3]  C/P,P
HOPKINS, JACK  SATELLITE NIGHT NEWS [SATELLITE NIGHT 1]  C/P,P
HOPKINS, JACK  SATELLITE NIGHT SPECIAL [SATELLITE NIGHT 2]  C/P,P
HOPKINS, LINTON C.  CANDLE, THE  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 1  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 3  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 4  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 5  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 6  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 8  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 9  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 10  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 11  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 12  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 13  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 14  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 15  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 16  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 17  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 18  SR-L
HOPKINS, MARIANE S.  FANDOM DIRECTORY 19  R/REF-P,C/REF-H
HOPKINSON, NALO  BROWN GIRL IN THE RING  H,LP
HOPKINSON, NALO  MIDNIGHT ROBBER  H,LP
HOPKINSON, NALO  MOJO: CONJURE STORIES (PROOF)  SR-LPA
HOPKINSON, NALO  SKIN FOLK  LP
HOPKINSON, NALO/MEHAN, UPPINDER  SO LONG BEEN DREAMING  LPA
HOPKINSON, NALO/RYMAN, GEOFF  TESSERACTS NINE [TESSERACTS 9]  LPA
HOPKINSON, NALO  WHISPERS FROM THE COTTON TREE ROOT  LP
HOPPE, STEPHANIE T.  WINDRIDER, THE  P
HORAN, ELLAMAY  BALTIMORE CATECHISM 2  [C/RAN-H]
HORLAK, E. E.  STILL LIFE  C/P,P
HORNER, W. H.  CLOAKED IN SHADOW  LPA
HORTON, RICH  FANTASY: THE BEST OF THE YEAR 2006  LPA
HORTON, RICH  FANTASY: THE BEST OF THE YEAR 2007  HA,PA
HORTON, RICH  SCIENCE FICTION: THE BEST OF THE YEAR 2007  LPA

The MITSFS Pinkdex  September 24, 2007  Page 194
AUTHOR  TITLE  CODE
HORVAT, DILWYN  ASSAULT ON OMEGA 4  P
HORVATH, GILLIAN  EVENING AT JOE’S, AN [HIGHLANDER]  LPA
HORWOOD, WILLIAM  DUNCTON WOOD  P
HOSKINS, ROBERT  AGAINST TOMORROW  PA
HOSKINS, ROBERT  EDGE OF NEVER, THE  HA,PA
HOSKINS, ROBERT  FAR-OUT PEOPLE, THE  PA,C/PA
HOSKINS, ROBERT  FIRST STEP OUTWARD  C/PA,PA
HOSKINS, ROBERT  FUTURE NOW, THE  C/PA,PA
HOSKINS, ROBERT  INFINITY ONE  C/PA,PA
HOSKINS, ROBERT  INFINITY TWO  PA,C/PA
HOSKINS, ROBERT  INFINITY THREE  HA,PA
HOSKINS, ROBERT  INFINITY FOUR  PA,C/PA
HOSKINS, ROBERT  INFINITY FIVE  PA
HOSKINS, ROBERT  LIBERATED FUTURE, THE  PA,C/PA
HOSKINS, ROBERT  MASTER OF THE STARS  SR-P
HOSKINS, ROBERT  SHATTERED PEOPLE, THE  H,SFWA
HOSKINS, ROBERT  STARS AROUND US, THE  C/PA,PA
HOSKINS, ROBERT  STRANGE TOMORROWS  C/PA,PA
HOSKINS, ROBERT  SWORDS AGAINST TOMORROW  PA,C/PA
HOSKINS, ROBERT  TO CONTROL THE STARS  C/P:2,P
HOSKINS, ROBERT  TO ESCAPE THE STARS  C/P:2,P
HOSKINS, ROBERT  TOMORROW 1  PA,C/PA
HOSKINS, ROBERT  TOMORROW’S SON  H,SFWA
HOSKINS, ROBERT  WONDERMAKERS  C/PA,PA
HOSKINS, ROBERT  WONDERMAKERS 2  PA
HOSS, NORMAN  HOW AND WHY WONDER BOOK OF STARS  [C/RAN-H]
HOTCHKISS, ROBERT R.  EARTH OF FIRE, SKY OF ICE [SHALRODAN SAGA 1]  LP
HOTTERBEEX, MARCEL  SCIENCE POLITIQUE ET SCIENCE-FICTION (FRENCH)  R/REF-H
HOUARNER, GERARD DANIEL  BARD OF SORCERY, THE  P
HOUGH, S. B.  EXTINCTION BOMBER  H
HOUGHTON MIFFLIN  DILEMMAS  LP
HOUSE, BRANT  STRANGE POWERS OF UNUSUAL PEOPLE  [C/RAN-P]
HOUSEHOLD, GEOFFREY  SPANISH CAVE, THE  P
HOUSEMAN, CLEMENCE  WEREWOLF  VLP
HOUSTON, DAVID  ALIEN PERSPECTIVE  P
HOUSTON, DAVID  GODS IN A VORTEX  P
HOUSTON, DAVID  ICE FROM SPACE [TALES OF TOMORROW 4]  P
HOUSTON, DAVID  INVADERS AT GROUND ZERO [TALES OF TOMORROW 1]  P
HOUSTON, DAVID  RED DUST [TALES OF TOMORROW 2]  P
HOUSTON, DAVID  SUBSTANCE X [TALES OF TOMORROW 3]  P
HOUSTON, DAVID/WEIN, LEN  SWAMP THING  P
HOUSTON, DAVID  WINGMASTER  P
HOVORKA, ROBERT L.  DERELICT  C/P,P
HOWARD, HAYDEN  ESKIMO INVASION, THE  C/P:2,P
HOWARD, IVAN  ESCAPE TO EARTH  PA,C/PA
HOWARD, IVAN  NOVELETS OF SCIENCE FICTION  PA,C/PA
HOWARD, IVAN  RARE SCIENCE FICTION  PA,C/PA
HOWARD, IVAN  6 AND THE SILENT SCREAM  C/PA,PA
HOWARD, IVAN  THINGS  PA,C/PA
HOWARD, IVAN  WAY OUT  PA,C/PA
HOWARD, JOSEPH  DAMIEN — OMEN II [OMEN 2]  P
HOWARD, MADELINE  HIDDEN STARS, THE [RUNE OF UNMAKING 1]  H,LP
HOWARD, ROBERT E.  ALMURIC  C/P:2,P
HOWARD, ROBERT E.  BEYOND THE BORDERS 7  P
HOWARD, ROBERT E.  BLACK CANAAN  P
HOWARD, ROBERT E.  BLACK STRANGER, THE  LP
HOWARD, ROBERT E.  BLACK VULMEA’S VENGEANCE  C/P,VLP
HOWARD, ROBERT E.  BLOODY CROWN OF CONAN, THE [CONAN]  H
HOWARD, ROBERT E.  BOOK OF ROBERT E. HOWARD, THE  P
HOWARD, ROBERT E.  BOOKS OF ROBERT E. HOWARD, THE  P
HOWARD, ROBERT E.  BRAN MAK MORN [BRAN MAK MORN 1]  SR-P,C/P
HOWARD, ROBERT E.  BRAN MAK MORN: THE LAST KING [BRAN MAK MORN]  H
HOWARD, ROBERT E.  COMING OF CONAN THE CIMMERIAN, THE [CONAN]  H,LP

HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/ CARTER, LIN CONAN [CONAN 1]

C/P:2,P

HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/ CARTER, LIN CONAN OF CIMMERIA [CONAN 2]

C/P:2,P

HOWARD, ROBERT E./DE CAMP, L. SPRAGUE CONAN THE ADVENTURER [CONAN 5]  C/P:2,P

HOWARD, ROBERT E./NYBERG, BJORN/ DE CAMP, L. SPRAGUE CONAN THE AVENGER [CONAN 10]

C/P:2,P

HOWARD, ROBERT E.  CONAN THE CONQUEROR [CONAN 9]  C/P:2,SR-DD036,P
HOWARD, ROBERT E./DE CAMP, L. SPRAGUE CONAN: THE FLAME KNIFE [CONAN]  P
HOWARD, ROBERT E./DE CAMP, L. SPRAGUE CONAN THE FREEBOOTER [CONAN 3]  C/P:2,P
HOWARD, ROBERT E./DE CAMP, L. SPRAGUE CONAN THE USURPER [CONAN 8]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 195
AUTHOR  TITLE  CODE

HOWARD, ROBERT E./DE CAMP, L. SPRAGUE/ CARTER, LIN CONAN THE WANDERER [CONAN 4]

C/P,P

HOWARD, ROBERT E./DE CAMP, L. SPRAGUE CONAN THE WARRIOR [CONAN 7]  C/P:2,P
HOWARD, ROBERT E.  CONQUERING SWORD OF CONAN, THE [CONAN]  LP
HOWARD, ROBERT E.  CORMAC MAC ART  P
HOWARD, ROBERT E.  CTHULHU: THE MYTHOS AND KINDRED HORRORS  C/P:2,P
HOWARD, ROBERT E.  DEVIL IN IRON, THE  VLP
HOWARD, ROBERT E.  ECHOES FROM AN IRON HARP  SR-H
HOWARD, ROBERT E.  EONS OF THE NIGHT [@HOWARD LIBRARY #5]  P
HOWARD, ROBERT E.  GARDENS OF FEAR  H
HOWARD, ROBERT E.  GODS OF BAL-SAGOTH, THE  C/P,P
HOWARD, ROBERT E.  HAND OF KANE, THE [SOLOMON KANE 2]  P
HOWARD, ROBERT E.  HOUR OF THE DRAGON, THE [BERKLEY CONAN 1]  H,P
HOWARD, ROBERT E.  HOWARD COLLECTOR, THE  P
HOWARD, ROBERT E.  INCREDIBLE ADVENTURES OF DENNIS DORGAN, THE  SR-H,P
HOWARD, ROBERT E.  IRON MAN, THE  P
HOWARD, ROBERT E./CARTER, LIN  KING KULL [KULL]  C/P,P
HOWARD, ROBERT E.  KULL [KULL]  C/P,P
HOWARD, ROBERT E.  KULL: EXILE OF ATLANTIS [KULL]  H
HOWARD, ROBERT E.  LAST RIDE, THE  P
HOWARD, ROBERT E.  LOST VALLEY OF ISKANDER, THE  SR-H,P
HOWARD, ROBERT E.  MARCHERS OF VALHALLA  C/P:2,P
HOWARD, ROBERT E.  MOON OF SKULLS, THE [SOLOMON KANE 1]  C/P,P
HOWARD, ROBERT E.  PEOPLE OF THE BLACK CIRCLE, THE [BERKLEY CONAN 3]  P
HOWARD, ROBERT E.  PEOPLE OF THE DARK  H
HOWARD, ROBERT E.  PIGEONS FROM HELL  C/P:2,P
HOWARD, ROBERT E.  RED NAILS [BERKLEY CONAN 2]  P
HOWARD, ROBERT E./LUPOFF, RICHARD A.  RETURN OF SKULL-FACE, THE  SR-H
HOWARD, ROBERT E.  ROAD OF AZRAEL, THE  P
HOWARD, ROBERT E.  ROBERT E. HOWARD’S WORLD OF HEROES  P
HOWARD, ROBERT E.  SAVAGE TALES OF SOLOMON KANE, THE [SOLOMON KANE]  H
HOWARD, ROBERT E.  SECOND BOOK OF ROBERT E. HOWARD, THE  C/P,P
HOWARD, ROBERT E.  SHADOW KINGDOMS  H
HOWARD, ROBERT E.  SHE DEVIL, THE  P
HOWARD, ROBERT E.  SKULL-FACE  P
HOWARD, ROBERT E.  SKULL-FACE AND OTHERS  SR-H
HOWARD, ROBERT E.  SKULLS IN THE STARS [SOLOMON KANE 4]  P
HOWARD, ROBERT E.  SOLOMON KANE [SOLOMON KANE 3]  P
HOWARD, ROBERT E.  SOWERS OF THE THUNDER, THE  SR-H,P
HOWARD, ROBERT E.  SWORD WOMAN, THE  C/P,P
HOWARD, ROBERT E.  SWORDS OF SHAHRAZAR  C/P,P
HOWARD, ROBERT E.  THREE-BLADED DOOM  P
HOWARD, ROBERT E.  TIGERS OF THE SEA  SR-H,C/P,P
HOWARD, ROBERT E.  TOWER OF THE ELEPHANT, THE  VLP
HOWARD, ROBERT E.  TRAILS IN DARKNESS  P
HOWARD, ROBERT E.  VALLEY OF THE WORM  H
HOWARD, ROBERT E.  VULTURES OF WHAPETON, THE  P
HOWARD, ROBERT E.  WINGS IN THE NIGHT  LP
HOWARD, ROBERT E.  WORMS OF THE EARTH  SR-H,P
HOWARD, RUSS T.  ULTIMATE HELM, THE [@SPELLJAMMER CLOAKMASTER #6]  C/P,P
HOWARD, STELLA  PROPHECY OF DARKNESS [@XENA]  P
HOWARTH, MATT  CHANGES  CX
HOWATCH, SUSAN  DEVIL ON LAMMAS NIGHT, THE  P
HOWE, DAVID J./STAMMERS, MARK  COMPANIONS [DOCTOR WHO]  R/REF-H

HOWE, DAVID J./STAMMERS, MARK/ WALKER, STEPHEN JAMES EIGHTIES, THE [DOCTOR WHO]

R/REF-H

HOWE, DAVID J./WALKER, STEPHEN JAMES  FIFTH DOCTOR, THE [DOCTOR WHO]  R/REF-P,C/REF-P

HOWE, DAVID J./STAMMERS, MARK/ WALKER, STEPHEN JAMES FIRST DOCTOR, THE [DOCTOR WHO]

R/REF-P,C/REF-P

HOWE, DAVID J./STAMMERS, MARK/ WALKER, STEPHEN JAMES FOURTH DOCTOR, THE [DOCTOR WHO]

R/REF-P

HOWE, DAVID J./STAMMERS, MARK/ WALKER, STEPHEN JAMES SECOND DOCTOR, THE [DOCTOR WHO]

R/REF-P

HOWE, DAVID J./WALKER, STEPHEN JAMES  THIRD DOCTOR, THE [DOCTOR WHO]  R/REF-P
HOWE, DEBORAH/HOWE, JAMES  BUNNICULA [BUNNICULA 1]  LP
HOWE, IMOGEN  FATAL ATTRACTION [TWILIGHT 4]  P
HOWE, JAMES  CELERY STALKS AT MIDNIGHT, THE [BUNNICULA 3]  LP
HOWE, JAMES  HOWLIDAY INN [BUNNICULA 2]  LP
HOYLE, FRED/ELLIOT, JOHN  "A" COMME ANDROMEDE [ANDROMEDA 1]  FP
HOYLE, FRED/ELLIOT, JOHN  A FOR ANDROMEDA [ANDROMEDA 1]  H:2,P
HOYLE, FRED/ELLIOT, JOHN  AMEACA DE ANDROMEDA [ANDROMEDA 1]  FP
HOYLE, FRED/ELLIOT, JOHN  ANDROMEDA BREAKTHROUGH [ANDROMEDA 2]  C/P,P,H
HOYLE, FRED  BLACK CLOUD, THE  H:2,P
HOYLE, FRED  COMET HALLEY  H

AUTHOR  TITLE  CODE
HOYLE, FRED  ELEMENT 79  H:2,P
HOYLE, FRED/HOYLE, GEOFFREY  FIFTH PLANET  H:2,P
HOYLE, FRED/HOYLE, GEOFFREY  INCANDESCENT ONES, THE  P
HOYLE, FRED/HOYLE, GEOFFREY  INFERNO, THE  H:2
HOYLE, FRED/HOYLE, GEOFFREY  INTO DEEPEST SPACE  H:2,P
HOYLE, FRED/HOYLE, GEOFFREY  MOLECULE MEN, THE  H:2
HOYLE, FRED  OCTOBER THE FIRST IS TOO LATE  H:2,P
HOYLE, FRED  OSSIAN’S RIDE  H:2,P
HOYLE, FRED/HOYLE, GEOFFREY  ROCKETS IN URSA MAJOR  H:2,P
HOYLE, FRED/HOYLE, GEOFFREY  SEVEN STEPS TO THE SUN  C/P:2,P
HOYLE, TREVOR  BLAKE’S SEVEN [BLAKE’S 7]  C/P,P
HOYLE, TREVOR  LAST GASP, THE  H,P
HOYLE, TREVOR  PROJECT AVALON [BLAKE’S 7]  C/P,P
HOYLE, TREVOR  SCORPIO ATTACK [BLAKE’S 7]  P
HOYLE, TREVOR  SEEKING THE MYTHICAL FUTURE [Q 1]  C/P,P
HOYLE, TREVOR  THIS SENTIENT EARTH  P
HOYLE, TREVOR  THROUGH THE EYE OF TIME [Q 2]  C/P,P
HOYT, RICHARD  MANNA ENZYME, THE  SFWA-P
HOYT, RICHARD  SIEGE  SFWA-P
HOYT, RICHARD  TROTSKY’S RUN  SFWA-P
HOYT, SARAH A.  ALL NIGHT AWAKE [WILLIAM SHAKESPEARE 2]  P
HOYT, SARAH A.  ANY MAN SO DARING  P
HOYT, SARAH A.  CRAWLING BETWEEN HEAVEN AND EARTH  SR-LP
HOYT, SARAH A.  DRAW ONE IN THE DARK  H
HOYT, SARAH A.  ILL MET BY MOONLIGHT [WILLIAM SHAKESPEARE 1]  P
HPLHS INC.  SHOGGOTH ON THE ROOF, A: THE HPLHS CAST ALBUM  SR-MM
HRUSKA, ALAN  BORROWED TIME  C/P,P
HSIEH, DEAN  ATHENA, ACT 1 [ATHENA 1]  SCX
HSIEH, DEAN  ATHENA, ACT 2 [ATHENA 2]  SCX
HUANZHULOUZHU  BLADES FROM THE WILLOWS  LP
HUBBARD, L. RON  ALIEN AFFAIR, AN [@MISSION EARTH #4]  H,P
HUBBARD, L. RON  BATTLEFIELD EARTH  H,P
HUBBARD, L. RON  BLACK GENESIS [@MISSION EARTH #2]  C/P,SFWA-P,H,P
HUBBARD, L. RON  DEATH QUEST [@MISSION EARTH #6]  H,P
HUBBARD, L. RON  DEATH’S DEPUTY  SR-H,P
HUBBARD, L. RON  DIANETICS  [C/RAN-H]
HUBBARD, L. RON  DISASTER [@MISSION EARTH #8]  C/P,P,H
HUBBARD, L. RON  DOOMED PLANET, THE [@MISSION EARTH #10]  H
HUBBARD, L. RON  ENEMY WITHIN, THE [@MISSION EARTH #3]  H,SFWA-P,P
HUBBARD, L. RON  FEAR  P,GN5.29
HUBBARD, L. RON  FEAR/TYPEWRITER IN THE SKY  P
HUBBARD, L. RON  FEAR/ULTIMATE ADVENTURE, THE  C/P:2,P
HUBBARD, L. RON  FINAL BLACKOUT  SR-H,P
HUBBARD, L. RON  FORTUNE OF FEAR [@MISSION EARTH #5]  H,P
HUBBARD, L. RON  FORTUNE OF FEAR (PROOF) [@MISSION EARTH #5]  SR-LP
HUBBARD, L. RON  INVADERS PLAN, THE [@MISSION EARTH #1]  SFWA-P,H,P
HUBBARD, L. RON  OLE DOC METHUSELAH  C/P:2,P
HUBBARD, L. RON  RETURN TO TOMORROW  C/P:2,P
HUBBARD, L. RON  SEVEN STEPS TO THE ARBITER  P
HUBBARD, L. RON  SLAVES OF SLEEP  C/P:2,P
HUBBARD, L. RON  TYPEWRITER IN THE SKY  SR-P
HUBBARD, L. RON  TYPEWRITER IN THE SKY/FEAR  H:2
HUBBARD, L. RON  VILLAINY VICTORIOUS [@MISSION EARTH #9]  H
HUBBARD, L. RON  VOYAGE OF VENGEANCE [@MISSION EARTH #7]  C/P,P,H
HUBBARD, L. RON  WRITER, THE  R/REF-H
HUBBELL, WILL  CRETACEOUS SEA [CRETACEOUS SEA 1]  P
HUBBELL, WILL  SEA OF TIME [CRETACEOUS SEA 2]  P
HUBSCHMAN, THOMAS  ALPHA-II  P
HUBSCHMAN, THOMAS  SPACE ARK  P
HUCKELBERY, TIMOTHY LEE  WISDOM TEETH [TEETH #1,2,3]  WRD
HUDDY, DELIA  TIME PIPER  P
HUDSON, EDWARD S.  ALIEN DEATH FLEET [@STAR FRONTIER #1]  C/P:2,P
HUDSON, MICHAEL  THIEVES OF LIGHT [PHOTON]  P
HUDSON, W. H.  CRYSTAL AGE, A  H
HUDSON, W. H.  GREEN MANSIONS  H,P
HUFF, TANYA  BETTER PART OF VALOR, THE [TORIN KERR 2]  P
HUFF, TANYA  BLOOD DEBT [VICTORY NELSON 5]  C/P:2,P
HUFF, TANYA  BLOOD LINES [VICTORY NELSON 3]  C/P:2,P
HUFF, TANYA  BLOOD PACT [VICTORY NELSON 4]  C/P:2,P
HUFF, TANYA  BLOOD PRICE [VICTORY NELSON 1]  C/P:2,P
HUFF, TANYA  BLOOD TRAIL [VICTORY NELSON 2]  C/P:2,P
HUFF, TANYA  CHILD OF THE GROVE  C/P,P
HUFF, TANYA  CONFEDERATION OF VALOR, A [TORIN KERR 1,2]  P
HUFF, TANYA  FIFTH QUARTER [QUARTER SINGERS 2]  C/P:2,P

AUTHOR  TITLE  CODE
HUFF, TANYA  FIRE’S STONE, THE [HUFF 4]  C/P,P
HUFF, TANYA  GATE OF DARKNESS, CIRCLE OF LIGHT [HUFF 3]  C/P:2,P
HUFF, TANYA  LAST WIZARD, THE  C/P,P
HUFF, TANYA  LONG HOT SUMMONING [KEEPER’S CHRONICLES 3]  P
HUFF, TANYA  NO QUARTER [QUARTER SINGERS 3]  C/P:2,P
HUFF, TANYA  OF DARKNESS, LIGHT, AND FIRE (OMNIBUS) [HUFF 3,4]  P
HUFF, TANYA  QUARTERED SEA, THE [QUARTER SINGERS 4]  C/P,P
HUFF, TANYA  RELATIVE MAGIC  H,LP
HUFF, TANYA  SCHOLAR OF DECAY [@RAVENLOFT]  P
HUFF, TANYA  SECOND SUMMONING, THE [KEEPER’S CHRONICLES 2]  C/P:2,P
HUFF, TANYA  SING THE FOUR QUARTERS [QUARTER SINGERS 1]  C/P:2,P
HUFF, TANYA  SMOKE AND ASHES [TONY FOSTER 3]  H:2
HUFF, TANYA  SMOKE AND MIRRORS [TONY FOSTER 2]  H
HUFF, TANYA  SMOKE AND SHADOWS [TONY FOSTER 1]  H
HUFF, TANYA  STEALING MAGIC  H,LP
HUFF, TANYA  SUMMON THE KEEPER [KEEPER’S CHRONICLES 1]  C/P:2,P
HUFF, TANYA  VALOR’S CHOICE [TORIN KERR 1]  P
HUFF, TANYA  WHAT HO, MAGIC!  LP
HUGHART, BARRY  BRIDGE OF BIRDS [MASTER LI 1]  C/P,P,H
HUGHART, BARRY  EIGHT SKILLED GENTLEMEN [MASTER LI 3]  C/P,LP,H
HUGHART, BARRY  STORY OF THE STONE, THE [MASTER LI 2]  C/P,P,H
HUGHES, DENIS  EARTH INVASION BATTALION, THE  SR-P
HUGHES, DENIS  MOON WAR  SR-P
HUGHES, DENIS  MURDER BY TELECOPTER  SR-P
HUGHES, DENIS  WAR LORDS OF SPACE  SR-P
HUGHES, DOROTHY B.  DELICATE APE, THE  SR-P,H
HUGHES, EDWARD P.  LONG MYND, THE  C/P,P
HUGHES, EDWARD P.  MASTERS OF THE FIST  C/P,SFWA-P,P
HUGHES, MARIAN  INITIATION  P
HUGHES, MATTHEW  BLACK BRILLION [ARCHONATE 3]  H
HUGHES, MATTHEW  FOOL ME TWICE [ARCHONATE 2]  C/P,P
HUGHES, MATTHEW  FOOLS ERRANT [ARCHONATE 1]  P
HUGHES, MATTHEW  GIST HUNTER, THE [ARCHONATE/HENGHIS HAPTHORN]  H
HUGHES, MATTHEW  GULLIBLE’S TRAVELS [ARCHONATE 1,2]  H
HUGHES, MATTHEW  MAJESTRUM [ARCHONATE/HENGHIS HAPTHORN]  H
HUGHES, MONICA  DEVIL ON MY BACK [ARC ONE 1]  P
HUGHES, MONICA  DREAM CATCHER, THE [ARC ONE 2]  SFWA
HUGHES, MONICA  INVITATION TO THE GAME  LP
HUGHES, MONICA  KEEPER OF THE ISIS LIGHT, THE [ISIS 1]  H,P
HUGHES, MONICA  PROMISE, THE [SANDWRITER 2]  H
HUGHES, RILEY  HILLS WERE LIARS, THE  P
HUGHES, ROBERT DAN  FAITHFUL TRAITOR, THE [WIZARD AND DRAGON 2]  C/P,P
HUGHES, ROBERT DON  FORGING OF THE DRAGON, THE [WIZARD AND DRAGON 1]  C/P:2,P
HUGHES, ROBERT DON  POWER AND THE PROPHET, THE [PELMEN THE POWERSHAPER 3]  C/P:2,P
HUGHES, ROBERT DON  PROPHET OF LAMATH, THE [PELMEN THE POWERSHAPER 1]  C/P:2,P
HUGHES, ROBERT DON  WIZARD IN WAITING, THE [PELMEN THE POWERSHAPER 2]  C/P:2,P
HUGHES, RODNEY  DRAGON KEEPERS, THE  P
HUGHES, RUPERT  DESTINY  H
HUGHES, RYAN  DARKNESS BEFORE THE DAWN, THE [DARK SUN ATHAS #2]  P
HUGHES, RYAN  HARD CRASH [UNREAL 1]  P
HUGHES, WILLIAM  LUST FOR A VAMPIRE  P
HUGHES, ZACH  BOOK OF RACK THE HEALER, THE  C/P:2,P
HUGHES, ZACH  CLOSED SYSTEM  C/P:2,P
HUGHES, ZACH  DARK SIDE, THE  C/P,P
HUGHES, ZACH  DEEP FREEZE  C/P,P
HUGHES, ZACH  FOR TEXAS AND ZED  C/P,P
HUGHES, ZACH  GOLD STAR  C/P,P
HUGHES, ZACH  KILLBIRD  P
HUGHES, ZACH  LEGEND OF MIAREE, THE  C/P,P
HUGHES, ZACH  LIFE FORCE  C/P,P
HUGHES, ZACH  MOTHER LODE  P
HUGHES, ZACH  OMNIFICENCE FACTOR, THE  C/P,P
HUGHES, ZACH  PRESSURE MAN  P
HUGHES, ZACH  SEED OF THE GODS  C/P:2,P
HUGHES, ZACH  STORK FACTOR, THE  C/P:2,P
HUGHES, ZACH  SUNDRINKER  P
HUGHES, ZACH  THUNDERWORLD  P
HUGHES, ZACH  TIDE  H,P
HUGHES, ZACH  TIGER IN THE STARS  SR-P
HULKE, MALCOLM  DOCTOR WHO AND THE CAVE-MONSTERS [DOCTOR WHO]  SR-P,C/P:2
HULKE, MALCOLM  DOCTOR WHO AND THE DINOSAUR INVASION [DOCTOR WHO]  P
HULKE, MALCOLM  DOCTOR WHO AND THE DOOMSDAY WEAPON [DOCTOR WHO]  C/P:2,P
HULKE, MALCOLM  DOCTOR WHO AND THE GREEN DEATH [DOCTOR WHO]  SR-P,C/P:2
HULKE, MALCOLM  DOCTOR WHO AND THE SEA-DEVILS [DOCTOR WHO]  SR-P,C/P:2

AUTHOR  TITLE  CODE
HULKE, MALCOLM  DOCTOR WHO AND THE SPACE WAR [DOCTOR WHO]  SR-P,C/P
HULKE, MALCOLM  DOCTOR WHO AND THE WAR GAMES [DOCTOR WHO]  SR-P,C/P:2
HUNT, ANDREW  CAT’S CRADLE: WITCH MARK [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
HUNT, BARBARA  LITTLE NIGHT MUSIC, A  H
HUNT, BARBARA  SEA CHANGE  H
HUNT, CECIL  WORD ORIGINS  [C/RAN-H]
HUNT, CHARLOTTE  CUP OF THANATOS, THE [PAUL HOLTON 2]  P
HUNT, CHARLOTTE  GILDED SARCOPHAGUS, THE [PAUL HOLTON 1]  P
HUNT, CHARLOTTE  LOTUS VELLUM, THE  P
HUNT, DAVE  ARCHON CONSPIRACY, THE  LP
HUNT, DOUGLAS  EXPLORING THE OCCULT  [C/RAN-P]
HUNT, GILL  ELEKTRON UNION  SR-P
HUNT, GILL  FISSION  SR-P,C/P
HUNT, GILL  GALACTIC STORM  SR-P
HUNT, GILL  HOSTILE WORLDS  SR-P
HUNT, GILL  PLANET X  SR-P
HUNT, GILL  SPACE FLIGHT  SR-P
HUNT, GILL  SPATIAL RAY  SR-P
HUNT, GILL  STATION 7  SR-P
HUNT, GILL  VEGA  SR-P
HUNT, GILL  ZERO FIELD  SR-P
HUNT, WALTER H.  DARK ASCENT, THE  P
HUNT, WALTER H.  DARK CRUSADE, THE [XOR 3]  P
HUNT, WALTER H.  DARK PATH, THE [XOR 2]  H
HUNT, WALTER H.  DARK WING, THE [XOR 1]  C/P,P
HUNTER, ALAN  STORY OF "F", THE  SR-L
HUNTER, ERIN  SUNSET [@WARRIORS NEW PROPHECY #6]  H
HUNTER, EVAN  FIND THE FEATHERED SERPENT  H:2
HUNTER, FAITH  BLOODRING  LP
HUNTER, JUSTIN  FORGED IN BATTLE [@WARHAMMER]  P
HUNTER, KIM  KNIGHT’S DAWN [@RED PAVILIONS #1]  LP
HUNTER, KIM  WIZARD’S FUNERAL [@RED PAVILIONS #2]  P
HUNTER, MOLLIE  THOMAS AND THE WARLOCK  LP
HUNTER, THOMAS O’D.  SOFTLY WALKS THE BEAST  P
HUNTING, GARDNER  VICARION, THE  H
HUNTINGTON, CHARLES  NIGHTMARE ON VEGA 3  C/P:2,P
HUNTINGTON, CHARLES  SOUL STEALERS, THE  C/P,P
HUNTINGTON, GEOFFREY  DEMON WITCH [@RAVENSCLIFF #2]  P
HUNTLEY, TIM  ONE ON ME  C/P,P
HURD, DOUGLAS/OSMOND, ANDREW  SEND HIM VICTORIOUS  P
HURKOS, PETER  PSYCHIC  [C/RAN-P]
HURLEY, MAXWELL  MAX’S BOOK [@PSI PATROL 3]  P
HURLEY, R. J.  BEYOND BELIEF  PA,C/PA
HURST, EARL OLIVER  BIG BOOK OF SPACE, THE  [C/RAN-H]
HURST, FANNIE  PRESIDENT IS BORN, A  H
HURTADO, OSCAR  INTRODUCCION A LA CIENCIA FICCION  FP
HURWOOD, BERNHARDT J.  INVISIBLES, THE  P
HURWOOD, BERNHARDT J.  KINGDOM OF THE SPIDERS  P
HURWOOD, BERNHARDT J.  MIND MASTER, THE  P
HURWOOD, BERNHARDT J.  MONSTERS GALORE  PA
HURWOOD, BERNHARDT J.  STRANGE TALENTS  [C/RAN-P]
HURWOOD, BERNHARDT J.  TERROR BY NIGHT  [C/RAN-P]
HUSS, ROY/ROSS, T. J.  FOCUS ON THE HORROR FILM  R/REF-H
HUSTON, CHARLIE  ALREADY DEAD  LP
HUSTON, CHARLIE  NO DOMINION  LP
HUTIN, SERGE  HISTORY OF ALCHEMY, A  [C/RAN-P]
HUXLEY, ALDOUS  AFTER MANY A SUMMER DIES THE SWAN  SR-P,C/P
HUXLEY, ALDOUS  APE AND ESSENCE  C/P,P,H
HUXLEY, ALDOUS  BRAVE NEW WORLD  H:2,SR-P,P
HUXLEY, ALDOUS  BRAVE NEW WORLD/BRAVE NEW WORLD REVISITED  LP
HUXLEY, ALDOUS  BRAVE NEW WORLD REVISITED  [C/RAN-H,C/RAN-P:2]
HUXLEY, ALDOUS  ISLAND  P
HUXLEY, ALDOUS  POINT COUNTER POINT  P
HUXLEY, ALDOUS  TIME MUST HAVE A STOP  H
HUXLEY, JULIAN  EVOLUTION IN ACTION  [C/RAN-P]
HUYSMANS, J. K.  LA-BAS  LP
HYAMS, EDWARD  998  H
HYAMS, GINA  CAMPFIRE COLLECTION, THE  LPA
HYMAN, JACKIE  SHADOWLIGHT  C/P,P
HYMAN, TOM  JUPITER’S DAUGHTER  H,SFWA-P
HYNE, C. J. CUTCLIFFE  LOST CONTINENT, THE  C/P:2,P

AUTHOR  TITLE  CODE
IAN, JANICE/RESNICK, MIKE  STARS  HA
IBBOTSON, EVA  SECRET OF PLATFORM 13, THE  LP
ICE, KATHY  DISTANT PLANES [MAGIC: THE GATHERING]  HA,PA
ICE, KATHY  TAPESTRIES [MAGIC: THE GATHERING]  LP
IDLE, ERIC  ROAD TO MARS, THE  H
IF EDITORS  BEST FROM IF 1, THE  PA
IF EDITORS  BEST FROM IF 2, THE  C/PA,PA
IGGULDEN, JOHN  BREAKTHROUGH  SR-P
IKIN, VAN  AUSTRALIAN SCIENCE FICTION  HA,LPA
ILES, GREG  FOOTPRINTS OF GOD, THE  H
IMAGE COMICS  FLIGHT, VOLUME ONE  CX
ING, DEAN  ANASAZI  C/P:2,P
ING, DEAN  BIG LIFTERS, THE  C/P,SFWA-P,P,H
ING, DEAN  CATHOUSE [MAN-KZIN WARS]  C/P:2,SFWA-P,P
ING, DEAN  CHERNOBYL SYNDROME, THE  SFWA-P
ING, DEAN  FIREFIGHT 2000  C/P,P
ING, DEAN  HIGH TENSION  P
ING, DEAN  NEMESIS MISSION, THE  H:2
ING, DEAN  PULLING THROUGH  C/P:2,P
ING, DEAN  RANSOM OF BLACK STEALTH ONE, THE  P,SFWA-P
ING, DEAN  SILENT THUNDER  TD31
ING, DEAN  SINGLE COMBAT [SYSTEMIC SHOCK 2]  C/P:2,P
ING, DEAN  SOFT TARGETS  C/P,P,H
ING, DEAN  SYSTEMIC SHOCK [SYSTEMIC SHOCK 1]  C/P:2,P
ING, DEAN  WILD COUNTRY [SYSTEMIC SHOCK 3]  C/P,SFWA-P,P
INGALLS, RACHEL  MRS. CALIBAN  LP
INGERSOLL, BOB/ISABELLA, TONY  CASE OF THE COLONIST’S CORPSE, THE [@STAR TREK SAMUEL COGLEY]  P
INGRID, CHARLES  ALIEN SALUTE [@SAND WARS #4]  C/P,P
INGRID, CHARLES  CELESTIAL HIT LIST [@SAND WARS #3]  C/P,P
INGRID, CHARLES  CHALLENGE MET [@SAND WARS #6]  C/P:2,P
INGRID, CHARLES  DOWNFALL MATRIX, THE [@PATTERNS OF CHAOS #3]  P
INGRID, CHARLES  LASERTOWN BLUES [@SAND WARS #2]  C/P,P
INGRID, CHARLES  LAST RECALL, THE [MARKED MAN 2]  P
INGRID, CHARLES  MARKED MAN, THE [MARKED MAN 1]  C/P:2,P
INGRID, CHARLES  PATH OF FIRE [@PATTERNS OF CHAOS #2]  P
INGRID, CHARLES  PATTERNS OF CHAOS, VOLUME ONE, THE [PATTERNS OF CHAOS 1,2]  P
INGRID, CHARLES  PATTERNS OF CHAOS, VOLUME TWO, THE [PATTERNS OF CHAOS 3,4]  P
INGRID, CHARLES  RADIUS OF DOUBT [@PATTERNS OF CHAOS #1]  P
INGRID, CHARLES  RETURN FIRE [@SAND WARS #5]  C/P,P
INGRID, CHARLES  SOLAR KILL [@SAND WARS #1]  C/P:2,P
INGRID, CHARLES  SOUL FIRE [@PATTERNS OF CHAOS #4]  P
INNES, EVAN  AMERICA 2040 [AMERICA 2040 1]  P
INNES, EVAN  CITY IN THE MIST [AMERICA 2040 3]  P
INNES, EVAN  GOLDEN WORLD, THE [AMERICA 2040 2]  C/P,P
INNES, EVAN  RETURN, THE [AMERICA 2040 4]  P
INNES, HAMMOND  MEDUSA  SFWA-P
INOUYE, JON  NIGHT TIDE, A  SR-P
IPCAR, DAHLOV  DARK HORN BLOWING, A  P
IRELAND, DAVID  WOMAN OF THE FUTURE, A  P
IRISH, WILLIAM  NIGHT HAS A THOUSAND EYES  SR-P
IRVINE, ALEXANDER C.  HAVE ROBOT, WILL TRAVEL [@ASIMOV’S ROBOT MYSTERY]  H
IRVINE, ALEXANDER C.  NARROWS, THE  LP
IRVINE, ALEXANDER C.  ONE KING, ONE SOLDIER  LP
IRVINE, ALEXANDER C.  PICTURES FROM AN EXPEDITION  LP
IRVINE, ALEXANDER C.  SCATTERING OF JADES, A  H,P
IRVINE, IAN  DARK IS THE MOON [@VIEW FROM THE MIRROR #3]  P
IRVINE, IAN  SHADOW ON THE GLASS, A [@VIEW FROM THE MIRROR #1]  C/P:2,P
IRVINE, IAN  TOWER ON THE RIFT, THE [@VIEW FROM THE MIRROR #2]  C/P,P
IRVINE, IAN  WAY BETWEEN THE WORLDS, THE [@VIEW FROM THE MIRROR #4]  P
IRWIN, WALTER/LOVE, G. B.  BEST OF THE BEST OF TREK, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF THE BEST OF TREK II, THE  SFWA,R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #2, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #3, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #4, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #5, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #6, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #7, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #8, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #9, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #10, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #11, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #12, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #13, THE  R/REF-P,C/REF-P

AUTHOR  TITLE  CODE
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #15, THE  R/REF-P,C/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #16, THE  R/REF-P
IRWIN, WALTER/LOVE, G. B.  BEST OF TREK #18, THE  R/REF-P,C/REF-P
ISAAK, ELAINE  EUNUCH’S HEIR, THE [DURHYS FAMILY 2]  LP
ISAAK, ELAINE  SINGER’S CROWN, THE [DURHYS FAMILY 1]  LP
ISHIGURO, KAZUO  NEVER LET ME GO  H
ISIDORE, SARAH  HIDDEN LAND, THE [DAUGHTERS OF BAST 1]  P
ISIDORE, SARAH  SHRINE OF LIGHT [DAUGHTERS OF BAST 2]  P
ISIDORE, SARAH  WORLD TREE, THE [DAUGHTERS OF BAST 3]  P
IVANHOE, MARK  VIRGINTOOTH  P
IVERSON, ERIC  WEREBLOOD  P
IVERSON, ERIC  WERENIGHT  P
IZUBUCHI, YUTAKA/BONES  RAHXEPHON, VOL. 1 [RAHXEPHON 1]  SCX
IZUBUCHI, YUTAKA/BONES  RAHXEPHON, VOL. 2 [RAHXEPHON 2]  SCX

AUTHOR  TITLE  CODE
JABLOKOV, ALEXANDER  BREATH OF SUSPENSION, THE  C/P,P
JABLOKOV, ALEXANDER  CARVE THE SKY  C/P,P,H
JABLOKOV, ALEXANDER  DEEPDRIVE  C/P:2,P
JABLOKOV, ALEXANDER  DEEPER SEA, A  C/P,SFWA,P,H
JABLOKOV, ALEXANDER  NIMBUS  H:2,SFWA,P
JABLOKOV, ALEXANDER  RIVER OF DUST  P
JABLONSKI, CARLA  BINDINGS [@BOOKS OF MAGIC #2]  P
JABLONSKI, CARLA  GYPSY ENCHANTMENT, THE [@CHARMED]  P
JABLONSKI, CARLA  INVITATION, THE [@BOOKS OF MAGIC 1]  P
JABLONSKI, CARLA  RECONINGS [@BOOKS OF MAGIC #6]  P
JABLONSKI, CARLA  SHADOW OF THE SPHINX [@CHARMED]  P
JACCOMA, RICHARD  WEREWOLF’S REVENGE, THE [JAMES UNDERHILL 2]  P
JACCOMA, RICHARD  WEREWOLF’S TALE, THE [JAMES UNDERHILL 1]  P
JACKSON, ANITA  ACTOR, THE  SR-P
JACKSON, ANITA  DEADLY GAME, A  SR-P
JACKSON, ANITA  DREAMS  SR-P
JACKSON, ANITA  EAR, THE  SR-P
JACKSON, ANITA  ’57 T-BIRD  SR-P
JACKSON, ANITA  HOMECOMING  SR-P
JACKSON, ANITA  NO RENT TO PAY  SR-P
JACKSON, ANITA  ZB4  SR-P
JACKSON, FRANCIS/MOORE, PATRICK  LEVEN IN HET HEELAL  FP
JACKSON, SHIRLEY  HAUNTING OF HILL HOUSE, THE  C/P,P
JACKSON, SHIRLEY  SUNDIAL, THE  P
JACKSON, SHIRLEY  WE HAVE ALWAYS LIVED IN THE CASTLE  C/P,P
JACKSON, STEVE  CROWN OF KINGS, THE [@SORCERY! #4]  [C/RAN-P]
JACKSON, STEVE  HOUSE OF HADES  P
JACKSON, STEVE  KHARE — CITYPORT OF TRAPS [@SORCERY! #2]  [C/RAN-P:2]
JACKSON, STEVE  SEVEN SERPENTS, THE [@SORCERY! #3]  [C/RAN-P]
JACKSON, STEVE  SHAMUTANTI HILLS, THE [@SORCERY! #1]  [C/RAN-P:2]
JACKSON, STEVE  SORCERY SPELL BOOK, THE [SORCERY!]  [C/RAN-P]
JACKSON, STEVE/LIVINGSTONE, IAN  TALISMAN OF DEATH  P
JACOBI, CARL  DISCLOSURES IN SCARLET  SR-H
JACOBI, CARL  PORTRAITS IN MOONLIGHT  SR-H
JACOBI, CARL  REVELATIONS IN BLACK  P
JACOBS, FRANK  ALVIN STEADFAST ON VERNACULAR ISLAND  H
JACOBS, JOSEPH  CELTIC FAIRY TALES  LP
JACOBS, MATTHEW  DOCTOR WHO: THE SCRIPT OF THE FILM [DOCTOR WHO]  LP
JACOBSON, KARIE  SIMULATIONS  HA,LPA
JACOBSON, MARK  GOJIRO  P
JACOBY, KATE  BLACK EAGLE RISING [@ELITA #3]  P
JACOBY, KATE  EXILE’S RETURN [@ELITA 1]  P
JACOBY, KATE  REBEL’S CAGE [@ELITA #4]  LP
JACOBY, KATE  VOICE OF THE DEMON [@ELITA #2]  P
JACQUES, BRIAN  ANGEL’S COMMAND, THE [CASTAWAYS OF THE FLYING DUTCHMAN]  P
JACQUES, BRIAN  BELLMAKER, THE [REDWALL]  H,P
JACQUES, BRIAN  HIGH RHULAIN [@REDWALL]  P
JACQUES, BRIAN  LEGEND OF LUKE, THE [REDWALL]  LP
JACQUES, BRIAN  LONG PATROL, THE [REDWALL]  P
JACQUES, BRIAN  LORD BROCKTREE [REDWALL]  H
JACQUES, BRIAN  MARIEL OF REDWALL [REDWALL]  C/P,P,H
JACQUES, BRIAN  MARLFOX [REDWALL]  P
JACQUES, BRIAN  MARTIN THE WARRIOR [REDWALL]  P
JACQUES, BRIAN  MATTIMEO [REDWALL]  P
JACQUES, BRIAN  MOSSFLOWER [REDWALL]  C/P,P
JACQUES, BRIAN  OUTCAST OF REDWALL, THE [REDWALL]  P
JACQUES, BRIAN  PEARLS OF LUTRA [REDWALL]  P
JACQUES, BRIAN  RAKKETY TAM [REDWALL]  LP
JACQUES, BRIAN  REDWALL [REDWALL]  C/P,P
JACQUES, BRIAN  SALAMANDASTRON [REDWALL]  C/P,P
JACQUES, BRIAN  TAGGERUNG [REDWALL]  P
JADE, SYMON  ALTER EVIL [@STARSHIP ORPHEUS #3]  P
JADE, SYMON  COSMIC CARNAGE [@STARSHIP ORPHEUS #2]  P
JADE, SYMON  RETURN FROM THE DEAD [@STARSHIP ORPHEUS #1]  P
JAFFERY, SHELDON  HORRORS AND UNPLEASANTRIES (ARKHAM HOUSE BIBLIOGRAPHY)  R/REF-H
JAFFERY, SHELDON  SENSUOUS SCIENCE FICTION FROM THE WEIRD AND SPICY PULPS  SR-LPA
JAHN, ERNEST T.  SILENT ENEMY, THE  P
JAHN, MICHAEL  ARMADA  C/P:2,P
JAHN, MICHAEL  FRIGHTENERS, THE  P
JAHN, MICHAEL  INTERNATIONAL INCIDENTS [BERKLEY SIX MILLION DOLLAR MAN 2]  P
JAHN, MICHAEL  INVISIBLE MAN, THE  P
JAHN, MICHAEL  RESCUE OF ATHENA ONE, THE [@SIX MILLION DOLLAR MAN #5]  C/P,P
JAHN, MICHAEL  SECRET OF BIGFOOT PASS, THE [BERKLEY SIX MILLION DOLLAR MAN 1]  P
JAHN, MICHAEL  WINE, WOMEN, AND WAR [@SIX MILLION DOLLAR MAN #1]  C/P,P

AUTHOR  TITLE  CODE
JAIVIN, LINDA  ROCK ’N’ ROLL BABES FROM OUTER SPACE  H
JAKES, JOHN  ANNI DELLA TURBINA, GLI  FH
JAKES, JOHN  ASYLUM WORLD, THE  C/P:2,P
JAKES, JOHN  BEST OF JOHN JAKES, THE  C/P,P
JAKES, JOHN  BLACK IN TIME  C/P:2,P
JAKES, JOHN  BRAK THE BARBARIAN [BRAK 1]  P
JAKES, JOHN  BRAK THE BARBARIAN (JAPANESE) [BRAK 1]  FP
JAKES, JOHN  BRAK THE BARBARIAN VERSUS THE MARK OF THE DEMONS [BRAK 2]  C/P:2,P
JAKES, JOHN  BRAK THE BARBARIAN VERSUS THE SORCERESS [BRAK 3]  C/P,P
JAKES, JOHN  CONQUEST OF THE PLANET OF THE APES  C/P:2,P
JAKES, JOHN  FORTUNES OF BRAK, THE [BRAK 5]  P
JAKES, JOHN  HYBRID, THE  C/P:2,P
JAKES, JOHN  LAST MAGICIANS, THE  C/P,P
JAKES, JOHN  MASK OF CHAOS  D78400
JAKES, JOHN  MASTER OF THE DARK GATE  P
JAKES, JOHN  MENTION MY NAME IN ATLANTIS  C/P:2,P
JAKES, JOHN  MONTE CRISTO #99  P
JAKES, JOHN  ON WHEELS  C/P,P
JAKES, JOHN  PLANET WIZARD, THE  C/P:2,P
JAKES, JOHN  ROUES DES TENEBRES, LES  FP
JAKES, JOHN  SECRETS OF STARDEEP  P
JAKES, JOHN  SIX-GUN PLANET  C/P,P
JAKES, JOHN  TIME GATE  C/P,P,H
JAKES, JOHN  TONIGHT WE STEAL THE STARS  D81680
JAKES, JOHN  WHEN THE IDOLS WALKED [BRAK 4]  P
JAKES, JOHN  WHEN THE STAR KINGS DIE  C/P:2,P
JAKES, JOHN  WITCH OF THE DARK GATE  C/P:2,P
JAKOBER, MARIE  BLACK CHALICE, THE  LP
JAKUBOWSKI, MAXIM/CHRISTIAN, M.  MAMMOTH BOOK OF FUTURE COPS, THE  LPA
JAKUBOWSKI, MAXIM/EDWARDS, MALCOLM SF BOOK OF LISTS, THE  R/REF-P
JAKUBOWSKI, MAXIM  TRAVELLING TOWARDS EPSILON  SR-PA
JAMES, CARY  KING & RAVEN  SFWA
JAMES, CARY  KING AND RAVEN  P
JAMES, DAKOTA  MILWAUKEE THE BEAUTIFUL  H
JAMES, DONALD  FALL OF THE RUSSIAN EMPIRE, THE  P
JAMES, EDWIN  VEILED KNOWLEDGE  B
JAMES, HENRY  GHOSTLY TALES OF HENRY JAMES  LP
JAMES, L. DEAN  BOOK OF STONES  C/P,P
JAMES, L. DEAN  KINGSLAYER  P
JAMES, L. DEAN  MOJAVE WELLS  P
JAMES, L. DEAN  SORCERER’S STONE  C/P,P
JAMES, L. DEAN  SUMMERLAND  P
JAMES, LAURENCE  BACKFLASH [@RACK #3]  P
JAMES, LAURENCE  EARTH LIES SLEEPING [@RACK #1]  P
JAMES, LAURENCE  NEW LIFE FOR OLD [RACK 5]  SR-P
JAMES, LAURENCE  PLANET OF THE BLIND [@RACK #4]  P
JAMES, LAURENCE  WAR ON ALEPH [@RACK #2]  P
JAMES, M. R.  BEST GHOST STORIES OF M. R. JAMES  H
JAMES, M. R.  GHOST STORIES OF AN ANTIQUARY  LP
JAMES, P. D.  CHILDREN OF MEN, THE  H,P
JAMES, PETER  SWEET HEART  SFWA
JAMES, ROBY  COMMENCEMENT  P
JAMES, ROBY  COMMITMENT  P
JAMESON, MALCOLM  BULLARD OF THE SPACE PATROL  H:2
JAMESON, MALCOM  TARNISHED UTOPIA  GN5.27
JANIFER, LAURENCE M.  ALIENIST [GERALD KNAVE]  LP
JANIFER, LAURENCE M.  BLOODWORLD  C/P:2,P
JANIFER, LAURENCE M.  COUNTERFEIT HEINLEIN, THE  LP
JANIFER, LAURENCE M.  18 GREATEST SCIENCE FICTION STORIES  PA
JANIFER, LAURENCE M./TREIBICH, S. J.  HIGH HEX, THE  D72400
JANIFER, LAURENCE M.  IMPOSSIBLE?  C/P:2,P
JANIFER, LAURENCE M.  KNAVE AND THE GAME [GERALD KNAVE]  H,SFWA
JANIFER, LAURENCE M.  KNAVE IN HAND [GERALD KNAVE 2]  C/P,P
JANIFER, LAURENCE M.  PIECE OF MARTIN CANN, A  C/P,P
JANIFER, LAURENCE M.  POWER  C/P,P
JANIFER, LAURENCE M.  REEL  H:2,SFWA
JANIFER, LAURENCE M.  SLAVE PLANET  C/P:2,P
JANIFER, LAURENCE M.  SURVIVOR [GERALD KNAVE 1]  C/P:2,P
JANIFER, LAURENCE M./TREIBICH, S. J.  TARGET: TERRA  DH091
JANIFER, LAURENCE M.  TWO [GERALD KNAVE]  LP
JANIFER, LAURENCE M./TREIBICH, S. J.  WAGERED WORLD, THE  D81680
JANIFER, LAURENCE M.  WONDER WAR, THE  C/P:2,P
JANIFER, LAURENCE M.  YOU SANE MEN  C/P:2,P
JANSON, HANK  UNSEEN ASSASSIN, THE  SR-P
JANSSON, TOVE  EXPLOITS OF MOOMINPAPPA, THE  SR-P

AUTHOR  TITLE  CODE
JARMAN, HEATHER  EVOLUTION [@STAR TREK VOYAGER STRING THEORY #3]  P
JARMAN, HEATHER  THIS GRAY SPIRIT [STAR TREK DS9 MISSION GAMMA #2]  C/P,P
JARRETT, DAVID  WITHERWING  C/P:2,P
JARVIS, ROBIN  ALCHYMIST’S CAT, THE [@DEPTFORD HISTORIES #2]  LP
JARVIS, ROBIN  CRYSTAL PRISON, THE [DEPTFORD MICE 2]  LP
JARVIS, ROBIN  OAKEN THRONE, THE [@DEPTFORD HISTORIES #2]  LP
JARVIS, ROBIN  THORN OGRES OF HAGWOOD [HAGWOOD 1]  H
JARVIS, SHARON  INSIDE OUTER SPACE  R/REF-H
JARVIS, SHARON  TRUE TALES OF THE UNKNOWN: THE UNINVITED  PA
JASTROW, ROBERT  JOURNEY TO THE STARS  [C/RAN-H]
JAVOR, FRANK A.  ICE BEAST, THE  C/P,P
JAVOR, FRANK A.  RIM-WORLD LEGACY, THE  C/P:2,P
JAVOR, FRANK A.  RIM-WORLD LEGACY AND BEYOND, THE  P
JAVOR, FRANK A.  SCOR-STING  P
JAY, MEL  ORBIT ONE  P
JEANYNE  MAGICKS AND CEREMONYES  [C/RAN-P]
JEFFERIES, MIKE  CITADEL OF SHADOWS [ELUNDIUM 4]  C/P,P
JEFFERIES, MIKE  GLITTERSPIKE HALL  P
JEFFERIES, MIKE  HALL OF WHISPERS [HEIRS TO GNARLSMYRE]  P
JEFFERIES, MIKE  HIDDEN ECHOES  P
JEFFERIES, MIKE  KNIGHTS OF CAWDOR, THE [ELUNDIUM 5]  P
JEFFERIES, MIKE  PALACE OF KINGS [ELUNDIUM 2]  P
JEFFERIES, MIKE  ROAD TO UNDERFALL, THE [ELUNDIUM 1]  C/P:2,P
JEFFERIES, MIKE  SHADOWLIGHT [ELUNDIUM 3]  P
JEFFERIES, MIKE  SHADOWS IN THE WATCHGATE  P
JEFFERIES, MIKE  STONE ANGELS  P
JEFFERIES, RICHARD  AFTER LONDON  LP
JEFFERIS, B. G./NICHOLS, J. L.  LIGHT ON DARK CORNERS  [C/RAN-H]
JEFFREY, MARK  POCKET AND THE PENDANT, THE  LP
JENKINS, A. M.  REPOSSESSED  SR-LP
JENKINS, HENRY  TEXTUAL POACHERS  R/REF-H
JENKINS, SUSAN/JENKINS, ROBERT  LIFE SIGNS [STAR TREK]  R/REF-H
JENKINS, WILL F.  FORGOTTEN PLANET, THE  H
JENKINS, WILL F.  MURDER OF THE U.S.A., THE  SR-P
JENKINS, WILL F.  SOLDADO ANT, THE  B
JENKS, TUDOR  TALES OF FANTASY  HA
JENNER, JANANN V.  SANDEAGOZU  H
JENNINGS, JAN  VAMPYR  P
JENNINGS, PHILIP S.  MEDIEVAL LEGENDS  HA
JENNINGS, PHILLIP C.  BUG LIFE CHRONICLES, THE  C/P,P
JENNINGS, PHILLIP C.  TOWER TO THE SKY  C/P,SFWA-P,P
JENSEN, JANE  DANTE’S EQUATION  LP
JENSEN, JANE  JUDGMENT DAY  P
JENSEN, JANE  MILLENNIUM RISING  H
JENSEN, JANE  SINS OF THE FATHERS [@GABRIEL KNIGHT]  P
JENSEN, KRIS  FREEMASTER  C/P:2,P
JENSEN, KRIS  HEALER  P
JENSEN, KRIS  MENTOR  P
JENSEN, LIZ  MY DIRTY LITTLE BOOK OF STOLEN TIMES  H
JENSEN, RUBY JEAN  LAKE, THE  SFWA-P
JENSEN, RUBY JEAN  SATAN’S SISTER  P
JENSEN, RUBY JEAN  SUCH A GOOD BABY  SFWA-P
JEPPSON, J. O.  LAST IMMORTAL, THE  C/P,P,H
JEPPSON, J. O.  MYSTERIOUS CURE, THE  SFWA
JEPPSON, J. O.  SECOND EXPERIMENT, THE  P
JERGENS, PHILLIP  BIOTA RISK, THE  P
JERSILD, P. C.  ANIMAL DOCTOR, THE  H,SFWA
JESCHKE, WOLFGANG  LAST DAY OF CREATION, THE  H
JETER, K. W.  AMES DEVOREES, LES  FP
JETER, K. W.  BLOODLETTER [STAR TREK DS9]  C/P:2,P
JETER, K. W.  BOUNTY HUNTER WARS, THE [STAR WARS BOUNTY HUNTER WARS 1,2,3]  H:2
JETER, K. W.  CROSS OF BLOOD [@ALIEN NATION #8]  P
JETER, K. W.  DARK HORIZON [@ALIEN NATION #2]  P
JETER, K. W.  DARK SEEKER  SFWA-P
JETER, K. W.  DEATH ARMS  H
JETER, K. W.  DR. ADDER  H,SFWA,P
JETER, K. W.  DREAMFIELDS, THE  SR-P,C/P
JETER, K. W.  EDGE OF HUMAN, THE [@BLADE RUNNER #2]  P
JETER, K. W.  FAREWELL HORIZONTAL  H:2
JETER, K. W.  GLASS HAMMER, THE  C/P,P,H
JETER, K. W.  HARD MERCHANDISE [@STAR WARS BOUNTY HUNTER WARS 3]  C/P,P
JETER, K. W.  INFERNAL DEVICES  H
JETER, K. W.  MADLANDS  H
JETER, K. W.  MANDALORIAN ARMOR, THE [@STAR WARS BOUNTY HUNTER WARS 1]  C/P,P

AUTHOR  TITLE  CODE
JETER, K. W.  MORLOCK NIGHT  P
JETER, K. W.  NOIR  P
JETER, K. W.  REPLICANT NIGHT [@BLADE RUNNER]  P
JETER, K. W.  SEEKLIGHT  SR-P
JETER, K. W.  SLAVE SHIP [@STAR WARS BOUNTY HUNTER WARS 2]  C/P,P
JETER, K. W.  WARPED [STAR TREK DS9]  H,P
JEURY, MICHAEL  CHRONOLYSIS  H:2
JEURY, MICHEL  L’ORBE ET LA ROUE  FP
JEWETT, ELEANORE M.  HIDDEN TREASURE OF GLASTON, THE  P
JEWISH PUBLICATION SOCIETY  TANAKH  PA
JOELS, KERRY MARK  MARS ONE CREW MANUAL, THE  R/REF-H
JOHANNESSON, OLOF  END OF MAN?, THE  P
JOHANSON, PAULA/TRUDEL, JEAN-LOUIS  TESSERACTS 7  LPA
JOHNS, MARSTON  BEYOND TIME  H
JOHNS, MARSTON  SPACE VOID  H
JOHNSON, ANNABEL/JOHNSON, EDGAR  MEMORY OF DRAGONS, A  H
JOHNSON, ANNABEL/JOHNSON, EDGAR  PRISONER OF PSI  H
JOHNSON, BARBARA FERRY  LIONORS  P
JOHNSON, CHARLES  SORCEROR’S APPRENTICE, THE  H
JOHNSON, CROCKETT  JACKEEN J. O’MALLEY FOR CONGRESS [@BARNABY #3]  SCX
JOHNSON, CROCKETT  MR. O’MALLEY AND THE HAUNTED HOUSE [@BARNABY #2]  SCX
JOHNSON, CROCKETT  MR. O’MALLEY GOES FOR THE GOLD [@BARNABY #4]  SCX
JOHNSON, CROCKETT  MR. O’MALLEY, WIZARD OF WALL STREET [@BARNABY #5]  SCX
JOHNSON, CROCKETT  WANTED: A FAIRY GODFATHER [@BARNABY #1]  SCX
JOHNSON, JALEIGH  HOWLING DELVE, THE [@FORGOTTEN REALMS DUNGEONS]  P
JOHNSON, JAMES B.  DAYSTAR AND SHADOW  P
JOHNSON, JAMES B.  HABU  P
JOHNSON, JAMES B.  MINDHOPPER  C/P,P
JOHNSON, JAMES B.  TREKMASTER  C/P:2,P
JOHNSON, JAMES B.  WORLD LOST, A  P
JOHNSON, KARMELA  EVIDENCE [GARRISON 2]  LP
JOHNSON, KEN  HOUNDS OF DRACULA  P
JOHNSON, KENNETH R.  CHECKLIST OF GERMAN "DIME-NOVELS"  R/IND
JOHNSON, KENNETH R.  INDEX TO MISCELLANEOUS FANTASY MAGAZINES  R/IND
JOHNSON, KENNETH R.  SCIENCE FICTION MAGAZINE CHECKLIST 1  R/IND
JOHNSON, KENNETH R.  SCIENCE FICTION MAGAZINE CHECKLIST 2  R/IND
JOHNSON, KIJ/COX, GREG  DRAGON’S HONOR [STAR TREK TNG]  C/P:2,P
JOHNSON, KIJ  FOX WOMAN, THE  H,LP
JOHNSON, KIJ  FUDOKI  LP
JOHNSON, OLIVER  FORGING OF THE SHADOWS, THE [@LIGHTBRINGER #1]  H,LP
JOHNSON, OLIVER  LAST STAR AT DAWN, THE [@LIGHTBRINGER #3]  H,P
JOHNSON, OLIVER  LORD OF SHADOW KEEP, THE  P
JOHNSON, OLIVER  NATIONS OF THE NIGHT, THE [@LIGHTBRINGER #2]  H,P
JOHNSON, OWEN  COMING OF THE AMAZONS, THE  H
JOHNSON, STANLEY  PRESIDENTIAL PLOT, THE  H
JOHNSON, TODD  SLAMMERS DOWN! [HAMMER’S SLAMMERS]  P,SFWA-P
JOHNSON, WILLIAM  FOCUS ON THE SCIENCE FICTION FILM  R/REF-H
JOHNSTON, JOE  STAR WARS SKETCHBOOK, THE [STAR WARS]  A
JOHNSTON, WILLIAM  AND LOVING IT! [GET SMART 6]  P
JOHNSTON, WILLIAM  ASYLUM  P
JOHNSTON, WILLIAM  BEWITCHED: THE OPPOSITE UNCLE  H
JOHNSTON, WILLIAM  BLOOP BOX, THE [NANNY AND THE PROFESSOR 3]  P
JOHNSTON, WILLIAM  CAPTAIN NICE  P
JOHNSTON, WILLIAM  GET SMART! [GET SMART 1]  P
JOHNSTON, WILLIAM  GET SMART ONCE AGAIN! [GET SMART 3]  C/P,P
JOHNSTON, WILLIAM  LITTLE GREEN MEN, THE [@FLYING NUN #4]  P
JOHNSTON, WILLIAM  LITTLEST REBELS, THE [@FLYING NUN #2]  C/P,P
JOHNSTON, WILLIAM  MAX SMART AND THE GHASTLY GHOST AFFAIR [GET SMART 9]  P
JOHNSTON, WILLIAM  MAX SMART AND THE PERILOUS PELLETS [GET SMART 4]  C/P,P
JOHNSTON, WILLIAM  MAX SMART LOSES CONTROL [GET SMART 8]  P
JOHNSTON, WILLIAM  MAX SMART — THE SPY WHO WENT OUT TO THE COLD [GET SMART 7]  P
JOHNSTON, WILLIAM  MIRACLE AT SAN TANCO [@FLYING NUN #1]  P
JOHNSTON, WILLIAM  MISSED IT BY THAT MUCH! [GET SMART 5]  P
JOHNSTON, WILLIAM  MOTHER OF INVENTION [@FLYING NUN #3]  P
JOHNSTON, WILLIAM  MUNSTERS AND THE GREAT CAMERA CAPER, THE  H
JOHNSTON, WILLIAM  NANNY AND THE PROFESSOR [NANNY AND THE PROFESSOR 1]  P
JOHNSTON, WILLIAM  SAM WESKIT ON THE PLANET FRAMINGHAM  C/P:2,P
JOHNSTON, WILLIAM  SORRY, CHIEF... [GET SMART 2]  P
JOHNSTON, WILLIAM  UNDERGROUND PICNIC, THE [@FLYING NUN #5]  P
JOHNSTONE, WILLIAM W.  ALONE IN THE ASHES [ASHES 5]  P
JOHNSTONE, WILLIAM W.  ANARCHY IN THE ASHES [ASHES 3]  P
JOHNSTONE, WILLIAM W.  BLOOD IN THE ASHES [ASHES 4]  P
JOHNSTONE, WILLIAM W.  CRYING SHAME, A  P
JOHNSTONE, WILLIAM W.  DANGER IN THE ASHES [ASHES 8]  P

AUTHOR  TITLE  CODE
JOHNSTONE, WILLIAM W.  FIRE IN THE ASHES [ASHES 2]  P
JOHNSTONE, WILLIAM W.  OUT OF THE ASHES [ASHES 1]  C/P,P
JOHNSTONE, WILLIAM W.  SMOKE FROM THE ASHES [ASHES 7]  P
JOHNSTONE, WILLIAM W.  TRAPPED IN THE ASHES [ASHES 9]  P
JOHNSTONE, WILLIAM W.  VALOR IN THE ASHES  P
JOHNSTONE, WILLIAM W.  WIND IN THE ASHES [ASHES 6]  C/P,P
JOHNSTONE, WILLIAM W.  WOLFSBANE  P
JONAS, DORIS/JONAS, DAVID  OTHER SENSES, OTHER WORLDS  R/REF-H
JONES, ADRIENNE  MURAL MASTER, THE  H
JONES, BOB  WEIRD MENACE, THE  R/IND
JONES, COURTNEY  IN THE SHADOW OF THE OAK KING  H
JONES, COURTWAY  PRINCE IN CAMELOT, A [DRAGON’S HEIRS 3]  P
JONES, D. F.  COLOSSUS [COLOSSUS 1]  H:2,P
JONES, D. F.  COLOSSUS AND THE CRAB [COLOSSUS 3]  C/P,P
JONES, D. F.  DENVER IS MISSING  C/P,P,H
JONES, D. F.  EARTH HAS BEEN FOUND  C/P,P
JONES, D. F.  FALL OF COLOSSUS, THE [COLOSSUS 2]  H:2,P
JONES, D. F.  FLOATING ZOMBIE, THE  C/P,P
JONES, D. F.  IMPLOSION  H:2,P
JONES, DAVID LEE  MONTEZUMA’S PEARL  P
JONES, DAVID LEE  UNICORN HIGHWAY  P
JONES, DAVID LEE  ZEUS AND CO.  P,SFWA-P
JONES, DENNIS  MASK AND THE SORCERESS, THE [PANDRAGORE 2]  H
JONES, DENNIS  STONE AND THE MAIDEN, THE [PANDRAGORE 1]  C/P,P,H
JONES, DIANA WYNNE  ARCHER’S GOON  P,SFWA-P
JONES, DIANA WYNNE  AUNT MARIA  LP
JONES, DIANA WYNNE  BLACK MARIA  LP
JONES, DIANA WYNNE  CART AND CWIDDER [DALEMARK 1]  C/P:2,LP
JONES, DIANA WYNNE  CASTLE IN THE AIR [CASTLE 2]  H
JONES, DIANA WYNNE  CHANGEOVER  LP
JONES, DIANA WYNNE  CHARMED LIFE [CHRESTOMANCI 1]  C/P:2,P
JONES, DIANA WYNNE  CHRESTOMANCI QUARTET, THE [CHRESTOMANCI 1,2,3,4]  H
JONES, DIANA WYNNE  CHRONICLES OF CHRESTOMANCI, VOLUME II [CHRESTOMANCI 3,2]  P
JONES, DIANA WYNNE  CONRAD’S FATE [CHRESTOMANCI 5]  H
JONES, DIANA WYNNE  CROWN OF DALEMARK, THE [DALEMARK 4]  C/P,P
JONES, DIANA WYNNE  DARK LORD OF DERKHOLM [DERKHOLM 1]  H
JONES, DIANA WYNNE  DEEP SECRET  H
JONES, DIANA WYNNE  DOGSBODY  P
JONES, DIANA WYNNE  DROWNED AMMET [DALEMARK 2]  C/P,P
JONES, DIANA WYNNE  EIGHT DAYS OF LUKE  H
JONES, DIANA WYNNE  EVERARD’S RIDE  SR-H
JONES, DIANA WYNNE  FIRE AND HEMLOCK  H,P
JONES, DIANA WYNNE  HEXWOOD  H
JONES, DIANA WYNNE  HOMEWARD BOUNDERS, THE  H,P
JONES, DIANA WYNNE  HOWL’S MOVING CASTLE [CASTLE 1]  P
JONES, DIANA WYNNE  LIVES OF CHRISTOPHER CHANT, THE [CHRESTOMANCI 4]  LP
JONES, DIANA WYNNE  MAGICIANS OF CAPRONA, THE [CHRESTOMANCI 3]  C/P,P
JONES, DIANA WYNNE  MERLIN CONSPIRACY, THE  H
JONES, DIANA WYNNE  MERLIN CONSPIRACY, THE (PROOF)  H,SR-LP
JONES, DIANA WYNNE  MINOR ARCANA  P
JONES, DIANA WYNNE  MIXED MAGICS [CHRESTOMANCI]  LP
JONES, DIANA WYNNE  OGRE DOWNSTAIRS, THE  H
JONES, DIANA WYNNE  PINHOE EGG, THE [CHRESTOMANCI]  H
JONES, DIANA WYNNE  POWER OF THREE  C/P,P
JONES, DIANA WYNNE  PUSS IN BOOTS  LP
JONES, DIANA WYNNE  SPELLCOATS, THE [DALEMARK 3]  C/P:2,P
JONES, DIANA WYNNE  STOPPING FOR A SPELL  LP
JONES, DIANA WYNNE  SUDDEN WILD MAGIC, A  P,SFWA
JONES, DIANA WYNNE  TALE OF TIME CITY, A  LP
JONES, DIANA WYNNE  TIME OF THE GHOST, THE  C/P,P
JONES, DIANA WYNNE  TOUGH GUIDE TO FANTASYLAND, THE  R/REF-H,C/REF-P
JONES, DIANA WYNNE  UNEXPECTED MAGIC  H
JONES, DIANA WYNNE  WARLOCK AT THE WHEEL  H
JONES, DIANA WYNNE  WILD ROBERT  H
JONES, DIANA WYNNE  WITCH WEEK [CHRESTOMANCI 2]  H,LP
JONES, DIANA WYNNE  WITCH’S BUSINESS  LP
JONES, DIANA WYNNE  YEAR OF THE GRIFFIN [DERKHOLM 2]  P
JONES, DOUGLAS C.  COURT-MARTIAL OF GEORGE ARMSTRONG CUSTER, THE  P
JONES, DOUGLAS C.  ROMAN  SFWA-P
JONES, GONNER  CANCRO 2000  FP
JONES, GUY/JONES, CONSTANCE  THERE WAS A LITTLE MAN  H
JONES, GWYNETH  BOLD AS LOVE [ROCK AND ROLL REICH 1]  LP
JONES, GWYNETH  CASTLES MADE OF SAND [ROCK AND ROLL REICH 2]  LP
JONES, GWYNETH  DIVINE ENDURANCE [DIVINE ENDURANCE 1]  H,SFWA-P,P

The MITSFS Pinkdex  September 24, 2007  Page 206
AUTHOR  TITLE  CODE
JONES, GWYNETH  FLOWERDUST [DIVINE ENDURANCE 2]  LP
JONES, GWYNETH  LIFE  LP
JONES, GWYNETH  NORTH WIND [ALEUTIAN 2]  H,LP
JONES, GWYNETH  PHOENIX CAFE [ALEUTIAN 3]  LP
JONES, GWYNETH  WHITE QUEEN [ALEUTIAN 1]  H
JONES, H. SPENCER  LIFE ON OTHER WORLDS  [C/RAN-P:2]
JONES, J. A.  BLUE LAB  P
JONES, J. V.  BAKER’S BOY, THE [BOOK OF WORDS 1]  C/P,LP
JONES, J. V.  BARBED COIL, THE  H:2,P
JONES, J. V.  CAVERN OF BLACK ICE, A [SWORD OF SHADOWS 1]  H,P
JONES, J. V.  FORTRESS OF GREY ICE, A [SWORD OF SHADOWS 2]  H
JONES, J. V.  MAN BETRAYED, A [BOOK OF WORDS 2]  C/P,LP,H
JONES, J. V.  MASTER AND FOOL [BOOK OF WORDS 3]  C/P:2,LP
JONES, JUDITH KAYE  OBELISK  P
JONES, LANGDON  EYE OF THE LENS, THE  H,P

JONES, LANGDON/MOORCOCK, MICHAEL NEW NATURE OF THE CATASTROPHE, THE (OMNIBUS EDITION)

[CORNELIUS]

PA

JONES, MARGARET  TRANSPLANT  P
JONES, MATTHEW  BAD THERAPY [DOCTOR WHO NEW ADVENTURES]  C/P,P
JONES, MATTHEW  BEYOND THE SUN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
JONES, MCCLURE  CAST DOWN THE STARS  H
JONES, NEIL/PRINGLE, DAVID  DEATHWING [WARHAMMER 40K]  PA
JONES, NEIL R.  DOOMSDAY ON AJIAT [PROFESSOR JAMESON 5]  C/P:2,P
JONES, NEIL R.  PLANET OF THE DOUBLE SUN, THE [PROFESSOR JAMESON 1]  C/P:2,P
JONES, NEIL R.  SPACE WAR [PROFESSOR JAMESON 3]  C/P:2,P
JONES, NEIL R.  SUNLESS WORLD, THE [PROFESSOR JAMESON 2]  C/P:2,P
JONES, NEIL R.  TWIN WORLDS [PROFESSOR JAMESON 4]  C/P:2,P
JONES, RAYMOND F.  ALIEN, THE  C/P:2,GN1.06,P
JONES, RAYMOND F.  CYBERNETIC BRAINS, THE  C/P,P,H
JONES, RAYMOND F.  DEVIATES, THE  SR-P,C/P:2
JONES, RAYMOND F.  GIFT OF THE GODS, THE  B
JONES, RAYMOND F.  KING OF EOLIM, THE  SR-P
JONES, RAYMOND F.  MAN OF TWO WORLDS  C/P:2,P
JONES, RAYMOND F.  MOONBASE ONE  H
JONES, RAYMOND F.  NON-STATISTICAL MAN, THE  C/P:2,P
JONES, RAYMOND F.  PLANET OF LIGHT  H:2
JONES, RAYMOND F.  RENAISSANCE  H
JONES, RAYMOND F.  RENEGADES OF TIME  SR-P,C/P:2
JONES, RAYMOND F.  RIVER AND THE DREAM, THE  SR-P
JONES, RAYMOND F.  SECRET PEOPLE, THE  H
JONES, RAYMOND F.  SON OF THE STARS  H:2
JONES, RAYMOND F.  SYN  C/P:2,P
JONES, RAYMOND F.  THIS ISLAND EARTH  H:2
JONES, RAYMOND F.  TOYMAKER, THE  SR-LP
JONES, RAYMOND F.  VOYAGE TO THE BOTTOM OF THE SEA  H:2
JONES, RAYMOND F./DEL REY, LESTER  WEEPING MAY TARRY  P
JONES, RAYMOND F.  YEAR WHEN STARDUST FELL, THE  H:2
JONES, ROBERT F.  BLOOD SPORT  P
JONES, ROBERT KENNETH  SHUDDER PULPS, THE  R/REF-H
JONES, ROBIN D.  BEGINNING OF UNBELIEF, THE  SFWA
JONES, STEPHEN/SUTTON, DAVID  BEST HORROR FROM FANTASY TALES, THE  LPA
JONES, STEPHEN/CAMPBELL, RAMSEY  BEST NEW HORROR 3  LPA
JONES, STEPHEN  MAMMOTH BOOK OF BEST NEW HORROR, THE  LPA
JONES, STEPHEN  MAMMOTH BOOK OF FRANKENSTEIN, THE  PA
JONES, STEPHEN  MAMMOTH BOOK OF VAMPIRE STORIES BY WOMEN, THE  HA
JONES, STEPHEN  MAMMOTH BOOK OF VAMPIRES, THE  LPA
JONES, TAMARA SILER  GHOSTS IN THE SNOW [DUBRIC BYERLY 1]  C/P,P
JONES, TERRY  DOUGLAS ADAMS’S STARSHIP TITANIC  LP
JONES, TRISTAN  AKA  H,P
JONES, WILLIAM  ARKHAM TALES  LPA
JONES, WILLIAM  ARTIFACTS  LP
JONES, WILLIAM  HORRORS BEYOND  LPA
JORDAN, ANNE DEVEREAUX  FIRES OF THE PAST  SFWA
JORDAN, BRENDA  BRENTWOOD WITCHES, THE  P
JORDAN, ROBERT  CONAN THE DEFENDER [CONAN]  C/P,SFWA-P,LP
JORDAN, ROBERT  CONAN THE DESTROYER [CONAN]  C/P:2,P
JORDAN, ROBERT  CONAN THE INVINCIBLE [CONAN]  C/P:2,P
JORDAN, ROBERT  CONAN THE MAGNIFICENT [CONAN]  C/P:2,P
JORDAN, ROBERT  CONAN THE TRIUMPHANT [CONAN]  C/P:2,SFWA-P,P
JORDAN, ROBERT  CONAN THE UNCONQUERED [CONAN]  C/P:2,SFWA-P,P
JORDAN, ROBERT  CONAN THE VICTORIOUS [CONAN]  C/P,SFWA,SFWA-P:2,P,H
JORDAN, ROBERT  CROSSROADS OF TWILIGHT [WHEEL OF TIME 10]  H
JORDAN, ROBERT  CROWN OF SWORDS, A [WHEEL OF TIME 7]  H:2,P
JORDAN, ROBERT  DRAGON REBORN, THE [WHEEL OF TIME 3]  C/P,P,H:2

The MITSFS Pinkdex  September 24, 2007  Page 207
AUTHOR  TITLE  CODE
JORDAN, ROBERT  DRAGON REBORN, THE (PROOF) [WHEEL OF TIME 3]  SR-LP
JORDAN, ROBERT  EYE OF THE WORLD, THE [WHEEL OF TIME 1]  C/P:2,H,LP,SFWA
JORDAN, ROBERT  EYE OF THE WORLD, THE (CHAPTERS 1-18) [@WHEEL OF TIME 1]  C/P,P
JORDAN, ROBERT  FIRES OF HEAVEN, THE [WHEEL OF TIME 5]  C/P,SFWA,P,H:2
JORDAN, ROBERT  GREAT HUNT, THE [WHEEL OF TIME 2]  H:2,P,SFWA:2
JORDAN, ROBERT  KNIFE OF DREAMS [WHEEL OF TIME 11]  H:2
JORDAN, ROBERT  LORD OF CHAOS [WHEEL OF TIME 6]  H:2,P
JORDAN, ROBERT  NEW SPRING [WHEEL OF TIME 0]  H
JORDAN, ROBERT  PATH OF DAGGERS, THE [WHEEL OF TIME 8]  H:2,P
JORDAN, ROBERT  SHADOW RISING, THE [WHEEL OF TIME 4]  C/P:2,P
JORDAN, ROBERT  WHEEL OF TIME, THE (SYNOPSIS) [WHEEL OF TIME 1,2]  SR-L
JORDAN, ROBERT  WINTER’S HEART [WHEEL OF TIME 9]  H:2,P

JORDAN, ROBERT/PATTERSON, TERESA WORLD OF ROBERT JORDAN’S THE WHEEL OF TIME, THE

[WHEEL OF TIME]

R/REF-H

JORDAN, SHERRYL  WINTER OF FIRE  P
JORGENSEN, IVAR  DEADLY SKY, THE  P
JORGENSEN, IVAR  REST IN AGONY  P
JORGENSEN, IVAR  STARHAVEN  DD351
JORGENSEN, IVAR  TEN FROM INFINITY  C/P,P
JORGENSEN, IVAR  WHOM THE GODS WOULD SLAY  C/P,P
JOSEPH, FRANZ  STAR FLEET TECHNICAL MANUAL  R/REF-H,C/REF-H:3
JOSEPH, FRANZ  STAR TREK BLUEPRINTS  A
JOSEPH, M. K.  HOLE IN THE ZERO, THE  C/P,P,H
JOSHI, S. T.  H.P. LOVECRAFT: A LIFE  R/REF-H
JOSHI, S. T.  INDEX TO THE SELECTED LETTERS OF H. P. LOVECRAFT, AN  R/IND

JOSHI, S. T. =LOVECRAFT, H. P.

INDEX TO THE SELECTED LETTERS OF H. P. LOVECRAFT (SECOND EDITION, KEEP WITH THE SELECTED LETTERS), AN

R/REF-H

JOWETT, SIMON/EVANS, PETER J.  BLACK ATLANTIC [@JUDGE DREDD]  P
JOYCE, GRAHAM  DARK SISTER  H
JOYCE, GRAHAM  DREAMSIDE  H,LP
JOYCE, GRAHAM  INDIGO  P
JOYCE, GRAHAM  LENINGRAD NIGHTS  MB001
JOYCE, GRAHAM  REQUIEM  LP
JOYCE, GRAHAM  TOOTH FAIRY, THE  LP
JOYCE, TOM/STEPHENS, CHRISTOPHER P.  CHECKLIST OF KIM STANLEY ROBINSON, A  R/IND
JUBERT, HERVE  DEVIL’S TANGO [DEVIL DANCES 2]  H
JUDD, CYRIL  GUNNER CADE  H,DD227
JUDD, CYRIL  OUTPOST MARS  C/P:2,P
JUDD, CYRIL  SIN IN SPACE  SR-P,C/P
JUDSON, THEODORE  FITZPATRICK’S WAR  H
JUENGER, ERNST  GLASS BEES, THE  LP
JULES-VERNE, JEAN  JULES VERNE: A BIOGRAPHY  R/REF-H
JULIAN, NANCY R.  PECULIAR MISS PICKETT, THE  LP
JURASIK, PETER/KEITH, WILLIAM H. JR.  DIPLOMATIC ACT  P
JUSTER, NORTON  PHANTOM TOLLBOOTH, THE  H:2,LP

AUTHOR  TITLE  CODE
KAABERSOL, LENE  SHAMER’S SIGNET, THE [SHAMER 2]  H
KADOHATA, CYNTHIA  GLASS MOUNTAINS, THE  P
KADOHATA, CYNTHIA  IN THE HEART OF THE VALLEY OF LOVE  H
KADREY, RICHARD  ANGEL SCENE  ERD2
KADREY, RICHARD  KAMIKAZE L’AMOUR  H
KADREY, RICHARD  METROPHAGE  C/P,SFWA-P,P
KAEWERT, JULIE  AVENGERS, THE  P
KAGAN, DAVID  SUNSTROKE  C/P,P
KAGAN, JANET  HELLSPARK  H:2,SFWA-P,P
KAGAN, JANET  MIRABILE  H
KAGAN, JANET  UHURA’S SONG [STAR TREK]  C/P:3,SFWA-P,P
KAHN, JAMES  POLTERGEIST  H,P
KAHN, JAMES  POLTERGEIST II: THE OTHER SIDE  P
KAHN, JAMES  RETURN OF THE JEDI [STAR WARS MOVIE 6]  C/P,P,H
KAHN, JAMES  TIMEFALL  H,P
KAHN, JAMES  TIME’S DARK LAUGHTER  C/P:2,P
KAHN, JAMES  WORLD ENOUGH, AND TIME  C/P:2,P
KAHN, JOAN  SOME THINGS DARK AND DANGEROUS  HA
KAHN, JOAN  SOME THINGS STRANGE AND SINISTER  HA
KAHN, JOAN  SOME THINGS WEIRD AND WICKED  PA
KAHN, STEVE  MALL, THE  P
KAINEN, RAY  COSMIC GASH, THE  C/P,SR-L
KAINEN, RAY  EARTH STATION SEX  SR-L
KAINEN, RAY  SATYR TREK  SR-L
KALISH, CAROL/BUSIEK, KURT  OPEN SPACE VOL. 1, NO. 1  CX
KAMARCK, LAWRENCE  INFORMED SOURCES  P
KAMERMAN, SYLVIA E.  SPACE AND SCIENCE FICTION PLAYS FOR YOUNG PEOPLE  HA
KAMIDA, VICKI  SABRETOOTH UNLEASHED [X-MEN]  [C/RAN-P]
KAMIN, NICK  EARTHRIM  D66160
KAMIN, NICK  HEROD MEN, THE  D13805
KAMINSKI, RAYMOND  AMAZONS OF SOMELON, THE  P
KANALY, MICHAEL  THOUGHTS OF GOD  H,LP
KANALY, MICHAEL  VIRUS CLANS  P
KANDEL, MICHAEL  CAPTAIN JACK ZODIAC  C/P:2,P
KANDEL, MICHAEL  IN BETWEEN DRAGONS  C/P:2,P
KANDEL, MICHAEL  PANDA RAY  H:2
KANDEL, MICHAEL  STRANGE INVASION  C/P,P
KANER, H.  PEOPLE OF THE TWILIGHT  H:2
KANER, H.  SUN QUEEN, THE  H:2
KANGILASKI, JAAN  HANDS OF GLORY  C/P,P
KANGILASKI, JAAN  SEEKING SWORD, THE  C/P:2,P
KANTOR, MACKINLAY  IF THE SOUTH HAD WON THE CIVIL WAR  P
KAO, TERESA  SHADOW FLOW  LP
KAPLAN, ALINE BOUCHER  KHYREN  C/P:2,SFWA-P,P
KAPP, COLIN  CHAOS WEAPON, THE  C/P:2,P
KAPP, COLIN  ION WAR, THE  C/P:2,P
KAPP, COLIN  LOST WORLDS OF CRONUS, THE [CAGEWORLD 2]  C/P,P
KAPP, COLIN  MANALONE  SR-P
KAPP, COLIN  PATTERNS OF CHAOS  C/P:2,P
KAPP, COLIN  SEARCH FOR THE SUN [CAGEWORLD 1]  C/P:2,P
KAPP, COLIN  STAR-SEARCH  C/P:2,P
KAPP, COLIN  SURVIVAL GAME, THE  C/P,P
KAPP, COLIN  TRANSFINITE MAN  C/P:2,P
KAPP, COLIN  TYRANT OF HADES, THE [CAGEWORLD 3]  C/P,P
KAPP, COLIN  WIZARD OF ANHARITTE, THE  C/P:2,P
KARIG, WALTER  ZOTZ!  H:2
KARINTHY, FRIGYES  VOYAGE TO FAREMIDO/CAPILLARIA  SR-P
KARL, JEAN E.  BUT WE ARE NOT OF EARTH  P
KARL, JEAN E.  TURNING PLACE, THE  H,P
KARLINS, MARVIN/ANDREWS, LEWIS M.  GOMORRAH  H
KARLOFF, BORIS  BORIS KARLOFF’S FAVORITE HORROR STORIES  PA
KARLOFF, BORIS  TALES OF TERROR  HA
KARLOVA, IRINA  DREADFUL HOLLOW  P
KARP, DAVID  ONE  SR-P,H
KARP, MARVIN ALLEN/SETTEL, IRVING  SUDDENLY  C/PA,PA
KARPYSHYN, DREW  PATH OF DESTRUCTION [@STAR WARS DARTH BANE]  H
KARPYSHYN, DREW  REVELATION [@MASS EFFECT]  P
KARPYSHYN, DREW  TEMPLE HILL [FORGOTTEN REALMS CITIES]  P
KARPYSHYN, DREW  THRONE OF BHAAL [@FORGOTTEN REALMS BALDUR’S GATE II]  P
KARR, PHYLLIS ANN  AT AMBERLEAF FAIR  C/P,LP
KARR, PHYLLIS ANN  FROSTFLOWER AND THORN [FROSTFLOWER]  C/P:2,P
KARR, PHYLLIS ANN  FROSTFLOWER AND WINDBOURNE [FROSTFLOWER]  C/P:2,P
KARR, PHYLLIS ANN  IDYLLS OF THE QUEEN, THE  C/P,P
KARR, PHYLLIS ANN  WILDRAITH’S LAST BATTLE  C/P:2,P
KASEM, LOU  DANCE OF DEATH [TWILIGHT 19]  P

The MITSFS Pinkdex  September 24, 2007  Page 210
AUTHOR  TITLE  CODE
KASHINA, ANNA  PRINCESS OF DHAGABAD, THE [SPIRITS OF ANCIENT SANDS 1]  H
KASNER, MICHAEL  KILLING FIELDS [WARKEEP 2030]  P
KASTLE, HERBERT D.  REASSEMBLED MAN, THE  C/P:2,P
KATHRYNS, G. A.  BORDERS OF LIFE, THE  C/P,P
KATO, KEN  RAGE IN HEAVEN, A (COMPLETE) [@YAMATO #1]  H
KATO, KEN  RAGE IN HEAVEN, A (PART 1) [YAMATO]  P
KATO, KEN  RAGE IN HEAVEN, A (PART 2) [@YAMATO #1]  C/P,P
KATO, KEN  WAY OF THE WARRIOR, THE (PART 2) [@YAMATO #2]  P
KATZ, ROBERT  CASSANDRA CROSSING, THE  P
KATZ, ROBERT I.  EDWARD MARET  LP
KATZ, STEVE  SAW  H
KATZ, WELWYN WILTON  WHALE SINGER  SFWA
KATZ, WILLIAM  DEATH DREAMS  P
KATZ, WILLIAM  VISIONS OF TERROR  P
KAUFMAN, DOUGLAS  DARK REALM, THE [@TORG POSSIBILITY WARS #2]  C/P,P
KAVAN, ANNA  ICE  H,P
KAVANAGH, JOHN PATRICK  SIXERS  H
KAVENY, ROZ  FROM ALIEN TO THE MATRIX  R/REF-H
KAVENY, ROZ  MORE TALES FROM THE FORBIDDEN PLANET  HA
KAVENY, ROZ  TALES FROM THE FORBIDDEN PLANET  SR-LPA
KAY, GUY GAVRIEL  DARKEST ROAD, THE [FIONAVAR TAPESTRY 3]  H:2,SFWA,P
KAY, GUY GAVRIEL  LAST LIGHT OF THE SUN, THE  H:2,LP
KAY, GUY GAVRIEL  LIONS OF AL-RASSAN, THE  H,P
KAY, GUY GAVRIEL  LORD OF EMPERORS [SARANTINE MOSAIC 2]  C/P,P,H
KAY, GUY GAVRIEL  SAILING TO SARANTIUM [SARANTINE MOSAIC 1]  H
KAY, GUY GAVRIEL  SAILING TO SARANTIUM (BRITISH) [SARANTINE MOSAIC 1]  C/P,P,H
KAY, GUY GAVRIEL  SONG FOR ARBONNE, A  C/P:2,SFWA-P,P
KAY, GUY GAVRIEL  SUMMER TREE, THE [FIONAVAR TAPESTRY 1]  H:2,SFWA,P
KAY, GUY GAVRIEL  TIGANA  C/P,SFWA-P,P,H:3
KAY, GUY GAVRIEL  WANDERING FIRE, THE [FIONAVAR TAPESTRY 2]  H:2,SFWA,SFWA-P,P
KAY, GUY GAVRIEL  YSABEL  H
KAY, KENNETH/GOLDBERG, MARSHALL  CHENGTU STRAIN, THE  P
KAY, SUSAN  PHANTOM  P
KAYE, MARILYN  CONVERGENCE, THE [@LAST ON EARTH #2]  P
KAYE, MARILYN  MAX ON EARTH [@OUT OF THIS WORLD #1]  P
KAYE, MARILYN  RETURN, THE [@LAST ON EARTH #3]  P
KAYE, MARILYN  VANISHING, THE [@LAST ON EARTH #1]  P
KAYE, MARVIN  AMOROUS UMBRELLA, THE  H,SFWA

KAYE, MARVIN/ BETANCOURT, JOHN GREGORY BEST OF WEIRD TALES: 1923, THE

LPA

KAYE, MARVIN  DEVILS AND DEMONS  HA,SFWA
KAYE, MARVIN  DON’T OPEN THIS BOOK!  HA
KAYE, MARVIN  DRAGON QUINTET, THE  HA
KAYE, MARVIN  FAIR FOLK, THE  HA
KAYE, MARVIN  FIENDS AND CREATURES  C/PA,PA
KAYE, MARVIN  FORBIDDEN PLANETS  HA
KAYE, MARVIN  INCREDIBLE UMBRELLA, THE  P,SFWA
KAYE, MARVIN  LOVERS AND OTHER MONSTERS  HA
KAYE, MARVIN  MASTERPIECES OF TERROR AND THE SUPERNATURAL  SFWA
KAYE, MARVIN/GODWIN, PARKE  MASTERS OF SOLITUDE, THE  H:2,P,SFWA
KAYE, MARVIN  PENGUIN BOOK OF WITCHES AND WARLOCKS, THE  HA,LPA
KAYE, MARVIN  POSSESSION OF IMMANUEL WOLF, THE  H,SFWA
KAYE, MARVIN  ULTIMATE HALLOWEEN, THE  PA
KAYE, MARVIN  VAMPIRE SEXTETTE, THE  HA,PA
KAYE, MARVIN  WEIRD TALES: THE MAGAZINE THAT NEVER DIES  HA
KAYE, MARVIN/GODWIN, PARKE  WINTERMIND  H:2,SFWA,P
KEARNEY, PAUL  FORSAKEN EARTH, THIS [@SEA BEGGARS 2]  LP
KEARNEY, PAUL  HAWKWOD’S VOYAGE [@MONARCHIES OF GOD #1]  P
KEARNEY, PAUL  HERETIC KINGS, THE [@MONARCHIES OF GOD #2]  P
KEARNEY, PAUL  IRON WARS, THE [@MONARCHIES OF GOD #3]  P
KEARNEY, PAUL  MARK OF RAN, THE [@SEA BEGGARS 1]  LP
KEARNEY, PAUL  SECOND EMPIRE, THE [@MONARCHIES OF GOD #4]  P
KEARNEY, PAUL  SHIPS FROM THE WEST [@MONARCHIES OF GOD #5]  C/P,P
KEARY, A./KEARY, E.  HEROES OF ASGARD, THE  SR-H
KECK, DAVID  IN THE EYE OF HEAVEN  H
KEEFE, MATT  OUTLANDER [@NECROMUNDA]  P
KEEL, JOHN A.  FICKLE FINGER OF FATE, THE  P
KEEL, JOHN A.  OUR HAUNTED PLANET  [C/RAN-P:2]
KEENE, BRIAN  BABYLON FALLING  BD00001
KEENE, BRIAN  CONQUEROR WORMS, THE  P
KEENE, DAY/PRUYN, LEONARD G.  WORLD WITHOUT WOMEN  C/P,P
KEEPING, CHARLES  CHARLES KEEPING’S BOOK OF CLASSIC GHOST STORIES  HA
KEITH, ANDREW  BLOOD OF HEROES [@BATTLETECH 11]  C/P,P
KEITH, ANDREW  COHORT OF THE DAMNED [@FIFTH FOREIGN LEGION #3]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 211
AUTHOR  TITLE  CODE
KEITH, ANDREW/KEITH, WILLIAM H. JR.  HONOR AND FIDELITY [@FIFTH FOREIGN LEGION #2]  P
KEITH, ANDREW/KEITH, WILLIAM H. JR.  MARCH OR DIE [@FIFTH FOREIGN LEGION #1]  P
KEITH, BRANDON  AFFAIR OF THE GENTLE SABOTEUR, THE [MAN FROM UNCLE]  H
KEITH, BRANDON  AFFAIR OF THE GUNRUNNER’S GOLD, THE [MAN FROM UNCLE]  H
KEITH, WILLIAM H. JR.  BOLO BRIGADE [BOLOS]  C/P,P
KEITH, WILLIAM H. JR.  BOLO RISING [BOLOS]  C/P,P
KEITH, WILLIAM H. JR.  BOLO STRIKE [BOLOS]  H
KEITH, WILLIAM H. JR.  DECISION AT THUNDER RIFT [@BATTLETECH 6]  P
KEITH, WILLIAM H. JR.  MERCENARY’S STAR [@BATTLETECH 7]  C/P:2,P
KEITH, WILLIAM H. JR.  NOMADS OF THE SKY [BUCK ROGERS INVADERS OF CHARON 2]  P
KEITH, WILLIAM H. JR.  OPERATION EXCALIBUR [@BATTLETECH 27]  P
KEITH, WILLIAM H. JR.  PRICE OF GLORY, THE [@BATTLETECH 8]  C/P:2,P
KEITH, WILLIAM H. JR.  RENEGADE’S HONOR  P
KEITH, WILLIAM H. JR.  RETIEF’S PEACE [RETIEF]  P
KEITH, WILLIAM H. JR.  TACTICS OF DUTY [@BATTLETECH 19]  P
KEITH, WILLIAM H. JR.  TWO OF MINDS [STAR*DRIVE]  P
KEITH, WILLIAM H. JR.  WARSTRIDER [@WARSTRIDER 1]  P,SFWA-P
KEITH, WILLIAM H. JR.  WARSTRIDER: BATTLEMIND [@WARSTRIDER 6]  P
KEITH, WILLIAM H. JR.  WARSTRIDER: JACKERS [@WARSTRIDER 3]  P
KEITH, WILLIAM H. JR.  WARSTRIDER: NETLINK [@WARSTRIDER 5]  C/P,P
KEITH, WILLIAM H. JR.  WARSTRIDER: REBELLION [@WARSTRIDER 2]  P,SFWA-P
KEITH, WILLIAM H. JR.  WARSTRIDER: SYMBIONTS [@WARSTRIDER 4]  P
KELLAR, VON  TRI-PLANET  SR-P
KELLEAM, JOSEPH E.  OVERLORDS FROM SPACE  DD173
KELLEAM, JOSESPH E.  WHEN THE RED KING WOKE  H
KELLEHER, ANNE  SILVER’S BANE [SIDHE AND SHADOWLANDS 2]  LP
KELLEHER, ANNE  SILVER’S LURE  LP
KELLEHER, VICTOR  BEAST OF HEAVEN, THE  H
KELLEHER, VICTOR  BROTHER NIGHT  H
KELLER, DAVID H.  FOLSOM FLINT, THE  SR-H
KELLER, DAVID H.  HUMAN TERMITES, THE  SR-LP
KELLER, DAVID H.  LAST MAGICIAN, THE  SR-LP
KELLER, DAVID H.  LIFE EVERLASTING  SR-H,LP
KELLER, DAVID H., M. D.  HOMMUNCULUS, THE  H
KELLER, MARK M./ANDERSON, SUE  RIVETS!!!  LP
KELLEY, LEO P.  ACCIDENTAL EARTH, THE  C/P:2,P
KELLEY, LEO P.  BACKWARD IN TIME  SR-P
KELLEY, LEO P.  COINS OF MURPH, THE  C/P:2,P
KELLEY, LEO P.  COUNTERFEITS, THE  P
KELLEY, LEO P.  DEAD MOON  SR-P
KELLEY, LEO P.  DEATH SENTENCE  SR-P
KELLEY, LEO P.  EARTH TRIPPER, THE  C/P:2,P
KELLEY, LEO P.  EARTH TWO  SR-P
KELLEY, LEO P.  GOOD-BYE TO EARTH  SR-P
KELLEY, LEO P.  KING OF THE STARS  SR-P
KELLEY, LEO P.  LUKE SUTTON: OUTRIDER  SFWA
KELLEY, LEO P.  MINDMIX  C/P,P
KELLEY, LEO P.  MYTHMASTER  C/P,P
KELLEY, LEO P.  ODYSSEY TO EARTHDEATH  P
KELLEY, LEO P.  ON THE RED WORLD  SR-P
KELLEY, LEO P.  PRISON SATELLITE  SR-P
KELLEY, LEO P.  STAR GOLD  SR-P
KELLEY, LEO P.  SUNWORLD  SR-P
KELLEY, LEO P.  TIME: 110100  H,SFWA
KELLEY, LEO P.  TIME ROGUE  P
KELLEY, LEO P.  TIME TRAP, THE  SR-P
KELLEY, LEO P.  VACATION IN SPACE  SR-P
KELLEY, LEO P.  WHERE NO SUN SHINES  SR-P
KELLEY, LEO P.  WORLDS APART  SR-P
KELLEY, THOMAS P.  I FOUND CLEOPATRA  SR-LP
KELLOGG, MARJORIE BRADLEY  BOOK OF AIR, THE [@DRAGON QUARTET #4]  C/P,P
KELLOGG, MARJORIE BRADLEY  BOOK OF EARTH, THE [@DRAGON QUARTET #1]  C/P,P
KELLOGG, MARJORIE BRADLEY  BOOK OF FIRE, THE [@DRAGON QUARTET #3]  C/P:2,P
KELLOGG, MARJORIE BRADLEY  BOOK OF WATER, THE [@DRAGON QUARTET #2]  C/P,P
KELLOGG, MARJORIE BRADLEY  HARMONY  P,SFWA-P:2

KELLOGG, MARJORIE BRADLEY/ ROSSOW, WILLIAM B.

LEAR’S DAUGHTERS

H

KELLOGG, MARJORIE BRADLEY/ ROSSOW, WILLIAM B.

REIGN OF FIRE [LEAR’S DAUGHTERS 2]

P

KELLOGG, MARJORIE BRADLEY  RUMOR OF ANGELS, A  C/P,P

KELLOGG, MARJORIE BRADLEY/ ROSSOW, WILLIAM B.

WAVE AND THE FLAME, THE [LEAR’S DAUGHTERS 1]

P

KELLY, DAVID J.  BAALBAK QUEST, THE [KILLSTAR 1]  SR-P
KELLY, DAVID J.  TOWER OF DESPAIR [KILLSTAR 2]  SR-P,C/P

AUTHOR  TITLE  CODE
KELLY, ERIN/WIECK, STEWART  CITY OF DARKNESS: UNSEEN [WORLD OF DARKNESS]  C/PA,PA
KELLY, ERIN E.  SPLENDOUR FALLS, THE  PA
KELLY, FRANK K.  STARSHIP INVINCIBLE  LP
KELLY, FRANK K.  STARSHIP INVINCIBLE (PROOF)  SR-VLP
KELLY, JAMES PATRICK  BURN  H
KELLY, JAMES PATRICK/KESSEL, JOHN  FEELING VERY STRANGE  LPA
KELLY, JAMES PATRICK/KESSEL, JOHN  FREEDOM BEACH  SR-H,SFWA-P,LP
KELLY, JAMES PATRICK  HEROINES  SR-H
KELLY, JAMES PATRICK  LOOK INTO THE SUN [MESSENGER 2]  C/P,SFWA-P,P,H
KELLY, JAMES PATRICK  PLANET OF WHISPERS [MESSENGER 1]  C/P,SFWA,SFWA-P,P,H
KELLY, JAMES PATRICK  STRANGE BUT NOT A STRANGER  H
KELLY, JAMES PATRICK  THINK LIKE A DINOSAUR AND OTHER STORIES  H
KELLY, JAMES PATRICK  WILDLIFE  P
KELLY, ROBERT B.  CLOUD PEOPLE, THE  P
KELLY, SHANE  HIDDEN CITY, THE  SR-L
KELLY, TOM/KELLY, ALRYSSA  CAT WHO WALKED THROUGH TIME, THE [DOCTOR WHO]  SR-VLP
KELLY, WALT  TEN EVER-LOVIN’ BLUE-EYED YEARS WITH POGO  CX
KEMENY, JOHN G.  MAN AND THE COMPUTER  [C/RAN-P]
KEMP, EARL  PROCEEDINGS; CHICON III, THE  R/REF-P
KEMP, PAUL  SHADOWBRED [@FORGOTTEN REALMS TWILIGHT WAR #1]  P
KEMP, PAUL S.  DAWN OF NIGHT [@FORGOTTEN REALMS EREVIS CALE #2]  P
KEMP, PAUL S.  MIDNIGHT’S MASK [@FORGOTTEN REALMS EREVIS CALE #3]  P
KEMP, PAUL S.  RESURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #6]  H
KEMP, PAUL S.  SHADOW’S WITNESS [@FORGOTTEN REALMS SEMBIA]  P
KEMP, PAUL S.  TWILIGHT FALLING [@FORGOTTEN REALMS EREVIS CALE #1]  P
KEMPER, WILLIAM E. JR.  ANOTHER MAN’S HELL  SR-P
KENDALL, GORDON  WHITE WING  C/P,SFWA-P,P
KENDRICK, WALTER  FIRE IN THE SKY, A  P
KENNAWAY, JAMES  MIND BENDERS, THE  C/P,P
KENNEALY-MORRISON, PATRICIA  BLACKMANTLE [KELTIAD MISC 1]  H:2,P
KENNEALY-MORRISON, PATRICIA  COPPER CROWN, THE [KELTIAD 1]  C/P:2,SFWA,P
KENNEALY-MORRISON, PATRICIA  DEER’S CRY, THE [KELTIAD]  C/P,P,H
KENNEALY-MORRISON, PATRICIA  HAWK’S GRAY FEATHER, THE [KELTIAD ARTHUR 1]  C/P:2,P
KENNEALY-MORRISON, PATRICIA  HEDGE OF MIST, THE [KELTIAD ARTHUR 3]  H,P
KENNEALY-MORRISON, PATRICIA  OAK ABOVE THE KINGS, THE [KELTIAD ARTHUR 2]  C/P,SFWA,P
KENNEALY-MORRISON, PATRICIA  SILVER BRANCH, THE [KELTIAD 3]  P
KENNEALY-MORRISON, PATRICIA  THRONE OF SCONE, THE [KELTIAD 2]  C/P,P,H
KENNEDY, EDGAR REES  CONQUERORS OF VENUS  SR-P
KENNEDY, EDGAR REES  MYSTERY PLANET, THE  SR-P
KENNEDY, KATHRYNE  BENEATH THE THIRTEEN MOONS  H
KENNEDY, LEIGH  FACES  H,SFWA
KENNEDY, LEIGH  JOURNAL OF NICHOLAS THE AMERICAN, THE  C/P:2,LP,H
KENNEDY, X. J.  OWLSTONE CROWN, THE  P
KENNER, JULIE  CARPE DEMON  LP
KENNETT, RICK  WARRIOR’S STAR, A  SR-P
KENSETT, PERCY F.  AMULET OF TARV, THE  H
KENSON, STEPHEN  BORN TO RUN [@SHADOWRUN KELLAN COLT 1]  P
KENSON, STEPHEN  BURNING TIME, THE [@SHADOWRUN 40]  P
KENSON, STEPHEN  CROSSROADS [SHADOWRUN 36]  P
KENSON, STEPHEN  FALLEN ANGELS [@SHADOWRUN KELLAN COLT 2]  P
KENSON, STEPHEN  GHOST OF WINTER [@MECHWARRIOR]  P
KENSON, STEPHEN  POISON AGENDAS [SHADOWRUN]  P
KENSON, STEPHEN  RAGNAROCK [@SHADOWRUN 38]  P
KENSON, STEPHEN  TECHNOBABEL [@SHADOWRUN 31]  P
KENT, BRAD  BIOLOGY "A"  SR-P
KENT, BRAD  CATALYST  SR-P
KENT, BRAD  FATAL LAW, THE  SR-P
KENT, BRAD  OUT OF THE SILENT PLACES  SR-P
KENT, MELANIE  HEAT WAVE [QUANTUM LEAP]  P
KENT, STEVEN L.  CLONE REPUBLIC, THE [CLONE REPUBLIC 1]  P
KENT, STEVEN L.  ROGUE CLONE [CLONE REPUBLIC 2]  P
KENTON, L. P.  DESTINATION MOON  SR-P
KENYON, ERNEST M.  ROGUE GOLEM  P
KENYON, KAY  BRAIDED WORLD, THE [MAXIMUM ICE 2]  P
KENYON, KAY  BRIGHT OF THE SKY [ENTIRE AND ROSE 1]  H
KENYON, KAY  LEAP POINT  P
KENYON, KAY  MAXIMUM ICE [MAXIMUM ICE 1]  P
KENYON, KAY  RIFT  P
KENYON, KAY  SEEDS OF TIME, THE  C/P,P
KENYON, KAY  TROPIC OF CREATION  P
KENYON, PAUL  OPERATION DOOMSDAY [BARONESS 5]  P
KENYON, SHERRILYN  DANCE WITH THE DEVIL [DARK-HUNTER]  P
KENYON, SHERRILYN  SINS OF THE NIGHT [@DARK-HUNTER]  P

The MITSFS Pinkdex  September 24, 2007  Page 213
AUTHOR  TITLE  CODE

KENYON, SHERRILYN/ASHLEY, AMANDA/ BANKS, L. A./HANDELAND, LORI STROKE OF MIDNIGHT

P

KENYON, SHERRILYN  UNLEASH THE NIGHT [@DARK-HUNTER]  P
KENYON, TIM  ERSATZ NATION  LP
KEPHART, CAROLYN  WYSARD  LP
KERN, GREGORY  BEYOND THE GALACTIC LENS [@CAP KENNEDY #16]  P
KERN, GREGORY  EARTH ENSLAVED [@CAP KENNEDY #9]  P
KERN, GREGORY  EATER OF WORLDS, THE [@CAP KENNEDY #8]  C/P,P
KERN, GREGORY  ENEMY WITHIN THE SKULL [@CAP KENNEDY #4]  P
KERN, GREGORY  GALACTIAD, THE  P
KERN, GREGORY  GALAXY OF THE LOST [@CAP KENNEDY #1]  C/P:2,P
KERN, GREGORY  GENETIC BUCCANEER, THE [@CAP KENNEDY #12]  C/P,P
KERN, GREGORY  GHOLAN GATE, THE [@CAP KENNEDY #7]  P
KERN, GREGORY  GHOSTS OF EPIDORIS, THE [@CAP KENNEDY #14]  C/P,P
KERN, GREGORY  JEWEL OF JARHEN [@CAP KENNEDY #5]  P
KERN, GREGORY  MIMICS OF DEPHENE [@CAP KENNEDY #15]  C/P,P
KERN, GREGORY  MONSTER OF METELAZE [@CAP KENNEDY #3]  P
KERN, GREGORY  PLANET OF DREAD [@CAP KENNEDY #10]  C/P,P
KERN, GREGORY  SEETEE ALERT! [@CAP KENNEDY #6]  C/P,P
KERN, GREGORY  SLAVE SHIP FROM SERGAN [@CAP KENNEDY #2]  C/P,P
KERN, GREGORY  SPAWN OF LABAN [@CAP KENNEDY #11]  C/P,P
KERN, GREGORY  WORLD AFLAME, A [@CAP KENNEDY #13]  P
KERNAGHAN, EILEEN  JOURNEY TO APRILIOTH  C/P:2,P
KERNAGHAN, EILEEN  SARSEN WITCH, THE  P
KERNER, ELIZABETH  LESSER KINDRED, THE [LANEN KAELAR 2]  P
KERNER, ELIZABETH  REDEEMING THE LOST  P
KERNER, ELIZABETH  SONG IN THE SILENCE [LANEN KAELAR 1]  C/P,P
KERNS, DANIEL R.  BORDER DISPUTE [INDIW 2]  P
KERNS, DANIEL R.  HERO  P
KERR, GEOFFREY  UNDER THE INFLUENCE  C/P:2,P
KERR, KATHARINE  BLACK RAVEN, THE [DEVERRY 10/DRAGON MAGE 2]  H,LP
KERR, KATHARINE  BRISTLING WOOD, THE [DEVERRY 3/DEVERRY PROPER 3]  C/P:2,P
KERR, KATHARINE  DAGGERSPELL [DEVERRY 1/DEVERRY PROPER 1]  H:2,SFWA,P
KERR, KATHARINE  DAGGERSPELL (REVISED) [DEVERRY 1/DEVERRY PROPER 1]  C/P:2,P
KERR, KATHARINE  DARKSPELL [DEVERRY 2/DEVERRY PROPER 2]  C/P,SFWA,P,H
KERR, KATHARINE  DARKSPELL (REVISED) [DEVERRY 2/DEVERRY PROPER 2]  C/P,P
KERR, KATHARINE  DAYS OF AIR AND DARKNESS [DEVERRY 8/WESTLANDS 4]  C/P:2,LP
KERR, KATHARINE  DAYS OF BLOOD AND FIRE [DEVERRY 7/WESTLANDS 3]  C/P:2,LP
KERR, KATHARINE  DRAGON REVENANT, THE [DEVERRY 4/DEVERRY PROPER 4]  C/P,LP,H
KERR, KATHARINE/GREENBERG, MARTIN H. ENCHANTED FORESTS  C/PA,PA
KERR, KATHARINE  FIRE DRAGON, THE [DEVERRY 11/DRAGON MAGE 3]  P
KERR, KATHARINE  FREEZE FRAMES  P
KERR, KATHARINE  GOLD FALCON, THE [@SILVER WYRM #1]  H
KERR, KATHARINE/KREIGHBAUM, MARK  PALACE [PINCH 1]  P
KERR, KATHARINE  POLAR CITY BLUES [POLAR CITY 1]  C/P,P
KERR, KATHARINE/DANIEL, KATE  POLAR CITY NIGHTMARE [POLAR CITY 2]  LP
KERR, KATHARINE  RED WYVERN, THE [DEVERRY 9/DRAGON MAGE 1]  C/P,LP
KERR, KATHARINE  RESURRECTION  C/P,P
KERR, KATHARINE  SHIMMERING DOOR, THE  LPA
KERR, KATHARINE  SNARE  H,P
KERR, KATHARINE  SPIRIT STONE, THE [@SILVER WYRM #2]  H
KERR, KATHARINE  TIME OF EXILE, A [DEVERRY 5/WESTLANDS 1]  C/P,P,H
KERR, KATHARINE  TIME OF JUSTICE, A [DEVERRY 7/WESTLANDS 3]  P
KERR, KATHARINE  TIME OF OMENS, A [DEVERRY 6/WESTLANDS 2]  H:2,P
KERR, KATHARINE  TIME OF WAR, A [DEVERRY]  P
KERR, KATHARINE/GREENBERG, MARTIN H. WEIRD TALES FROM SHAKESPEARE  PA,C/PA
KERR, PEG  EMERALD HOUSE RISING  C/P:2,P
KERR, PEG  WILD SWANS, THE  P
KERR, PHILIP  PHILOSOPHICAL INVESTIGATION, A  SFWA
KERR, PHILIP  SECOND ANGEL, THE  H
KERRUISH, JESSIE DOUGLAS  UNDYING MONSTER, THE  P
KERSH, GERALD  MEN WITHOUT BONES  C/P:2,P
KERSH, GERALD  NIGHTSHADE AND DAMNATIONS  C/P:2,P
KERSH, GERALD  ON AN ODD NOTE  C/P,P
KERSH, GERALD  SECRET MASTERS, THE  H:2,P
KESSEL, JOHN  ANOTHER ORPHAN  TD6,SFWA-TD6
KESSEL, JOHN  CORRUPTING DR. NICE  H,LP
KESSEL, JOHN  GOOD NEWS FROM OUTER SPACE  C/P:2,P

KESSEL, JOHN/VAN NAME, MARK L./ BUTNER, RICHARD INTERSECTIONS: THE SYCAMORE HILL ANTHOLOGY

LPA

KESSEL, JOHN  PURE PRODUCT, THE  LP
KESSLER, ANGELA  BEST OF DREAMS OF DECADENCE, THE  C/PA,PA
KESSLER, JACKIE  HELL’S BELLES  LP
KESTERTON, DAVID  DARKLING, THE  H

AUTHOR  TITLE  CODE
KETTERER, DAVID  NEW WORLDS FOR OLD  R/REF-P,C/REF-P
KEY, ALEXANDER  CASE OF THE VANISHING BOY, THE  C/P,P
KEY, ALEXANDER  ESCAPE TO WITCH MOUNTAIN  C/P,P
KEY, ALEXANDER  FLIGHT TO THE LONESOME PLACE  H
KEY, ALEXANDER  FORGOTTEN DOOR, THE  C/P:2,LP
KEY, ALEXANDER  GOLDEN ENEMY, THE  H
KEY, ALEXANDER  INCREDIBLE TIDE, THE  H
KEY, ALEXANDER  RETURN FROM WITCH MOUNTAIN  C/P,P
KEY, TED  CAT FROM OUTER SPACE, THE  C/P:2,P
KEYES, DANIEL  ALGERNON, CHARLIE AND I  R/REF-H
KEYES, DANIEL  FLOWERS FOR ALGERNON  C/P:2,P
KEYES, DANIEL  TOUCH, THE  P
KEYES, J. GREGORY  BLACKGOD, THE [CHILDREN OF THE CHANGELING 2]  H,P
KEYES, J. GREGORY  BLOOD KNIGHT, THE [KINGDOMS OF THORN AND BONE 3]  H
KEYES, J. GREGORY  BRIAR KING, THE [KINGDOMS OF THORN AND BONE 1]  H
KEYES, J. GREGORY  CALCULUS OF ANGELS, A [AGE OF UNREASON 2]  H:2,LP
KEYES, J. GREGORY  CHARNEL PRINCE, THE [KINGDOMS OF THORN AND BONE 2]  H
KEYES, J. GREGORY  CONQUEST [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #1]  P
KEYES, J. GREGORY  DARK GENESIS [@BABYLON 5 PSI CORPS 1]  C/P,P
KEYES, J. GREGORY  DEADLY RELATIONS [@BABYLON 5 PSI CORPS 2]  P
KEYES, J. GREGORY  EDGE OF VICTORY [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY 1,2] H
KEYES, J. GREGORY  EMPIRE OF UNREASON [AGE OF UNREASON 3]  LP
KEYES, J. GREGORY  FINAL PROPHECY, THE [@STAR WARS NEW JEDI ORDER]  C/P,P
KEYES, J. GREGORY  FINAL RECKONING [@BABYLON 5 PSI CORPS 3]  P
KEYES, J. GREGORY  NEWTON’S CANNON [AGE OF UNREASON 1]  H:2,LP
KEYES, J. GREGORY  REBIRTH [@STAR WARS NEW JEDI ORDER EDGE OF VICTORY #2]  P
KEYES, J. GREGORY  SHADOWS OF GOD, THE [AGE OF UNREASON 4]  H,LP
KEYES, J. GREGORY  WATERBORN, THE [CHILDREN OF THE CHANGELING 1]  H,P
KEYES, NOEL  CONTACT  PA,C/PA
KEYES, THOM  BATTLE OF DISNEYLAND, THE  H
KEYHOE, DONALD E.  ALIENS FROM SPACE  SFWA
KEYHOE, DONALD E.  FLYING SAUCER CONSPIRACY, THE  [C/RAN-H]
KIDD, CHICO  PRINTER’S DEVIL, THE  P
KIDD, PAUL  COUNCIL OF BLADES, THE [@FORGOTTEN REALMS NOBLES #6]  P
KIDD, PAUL  DESCENT INTO THE DEPTHS OF THE EARTH [@GREYHAWK]  P
KIDD, PAUL  MUS OF KERBRIDGE  P
KIDD, PAUL  QUEEN OF THE DEMONWEB PITS [@GREYHAWK]  P
KIDD, PAUL  WHITE PLUME MOUNTAIN [@GREYHAWK]  P
KIDD, VIRGINIA  MILLENNIAL WOMEN  HA,PA
KIDDER, TRACY  SOUL OF A NEW MACHINE, THE  [C/RAN-P]
KIEFER, MILDRED S.  PLANET X  H
KIERNAN, CAITLIN R.  DAUGHTER OF HOUNDS  LP
KIERNAN, CAITLIN R.  FIVE OF CUPS  H
KIERNAN, CAITLIN R.  LOW RED MOON  LP
KIERNAN, CAITLIN R.  SILK  C/P,LP
KIERNAN, CAITLIN R.  THRESHOLD  LP
KIERNAN, CAITLIN R.  TO CHARLES FORT, WITH LOVE  H
KIESEL, STANLEY  WAR BETWEEN THE PITIFUL TEACHERS AND THE SPLENDID KIDS, THE  C/P,P
KIESOW, ULRICH  CHARLATAN, THE [RELAMS OF ARKANIA]  P
KIHN, GREG  HORROR SHOW  H
KILIAN, CRAWFORD  BROTHER JONATHAN  P
KILIAN, CRAWFORD  EMPIRE OF TIME, THE [CHRONOPLANE WARS 3]  C/P,P
KILIAN, CRAWFORD  EYAS  C/P:2,P
KILIAN, CRAWFORD  FALL OF THE REPUBLIC, THE [CHRONOPLANE WARS 1]  C/P,P
KILIAN, CRAWFORD  GREENMAGIC [CALINDOR 1]  P
KILIAN, CRAWFORD  GRYPHON  C/P:2,P
KILIAN, CRAWFORD  ICEQUAKE  P
KILIAN, CRAWFORD  LIFTER  P
KILIAN, CRAWFORD  REDMAGIC [CALINDOR 2]  P
KILIAN, CRAWFORD  ROGUE EMPEROR [CHRONOPLANE WARS 2]  C/P,P
KILIAN, CRAWFORD  WRITING SCIENCE FICTION AND FANTASY  R/REF-H
KILLIANY, KEVIN  WOLF HUNTERS [@MECHWARRIOR DARK AGE/BATTLETECH]  P
KILLOUGH, LEE  AVENTINE  C/P:2,P
KILLOUGH, LEE  BLOOD HUNT [GARRETH MIKAELIAN 1]  C/P,SFWA-P,P
KILLOUGH, LEE  BLOODLINKS [GARRETH MIKAELIAN 2]  P,SFWA-P
KILLOUGH, LEE  BLOODWALK [GARRETH MIKAELIAN 1,2]  LP
KILLOUGH, LEE  BRIDLING CHAOS [JANNA BRILL 1,2,3]  LP
KILLOUGH, LEE  DEADLY SILENTS  C/P,P
KILLOUGH, LEE  DOPPELGANGER GAMBIT, THE [JANNA BRILL 1]  C/P,P
KILLOUGH, LEE  DRAGON’S TEETH [JANNA BRILL 3]  C/P:2,P
KILLOUGH, LEE  LEOPARD’S DAUGHTER  C/P:2,P
KILLOUGH, LEE  LIBERTY’S WORLD  C/P,P
KILLOUGH, LEE  MONITOR, THE MINERS AND THE SHREE, THE  C/P:2,P
KILLOUGH, LEE  SPIDER PLAY [JANNA BRILL 2]  P

The MITSFS Pinkdex  September 24, 2007  Page 215
AUTHOR  TITLE  CODE
KILLOUGH, LEE  VOICE OUT OF RAMAH, A  C/P:2,P
KILLUS, JAMES  BOOK OF SHADOWS  C/P,P
KILLUS, JAMES  SUNSMOKE  C/P,P
KILPATRICK, NANCY  BLOODLOVER [POWER OF THE BOOD 4]  LP
KILPATRICK, NANCY/ROCHE, THOMAS S.  GRAVEN IMAGES  LPA
KILPATRICK, NANCY/ROCHE, THOMAS S.  IN THE SHADOW OF THE GARGOYLE  PA
KILWORTH, GARRY  ANGEL  P
KILWORTH, GARRY  FOXES OF FIRSTDARK, THE  H
KILWORTH, GARRY  HOGFOOT RIGHT AND BIRD-HANDS  LP
KILWORTH, GARRY  IN SOLITARY  P
KILWORTH, GARRY  MIDSUMMER’S NIGHTMARE, A  P
KILWORTH, GARRY  NIGHT OF KADAR, THE  C/P,P
KILWORTH, GARRY  SONGBIRDS OF PAIN, THE  P
KILWORTH, GARRY  SPLIT SECOND  P
KILWORTH, GARRY  THEATER OF TIMESMITHS  P
KIMBALL, PROFESSOR JANUS  SCANNERS II: THE NEW ORDER  P
KIMBERLY, GAIL  DRACULA BEGAN  P
KIMBERLY, GAIL  FLYER  C/P,P
KIMBRIEL, KATHERINE ELISKA  FIRE SANCTUARY [NUALA 1]  C/P,P
KIMBRIEL, KATHERINE ELISKA  FIRES OF NUALA [NUALA 2]  C/P:2,P
KIMBRIEL, KATHERINE ELISKA  HIDDEN FIRES [NUALA 3]  C/P:2,P
KIMBRIEL, KATHERINE ELISKA  NIGHT CALLS  P
KIMBRO, JOHN M.  NIGHT OF TEARS  P
KIMBROUGH, KATHERYN  ANN, THE GENTLE [@PHENWICK #20]  P
KIMBROUGH, KATHERYN  AUGUSTA, THE FIRST [@PHENWICK #1]  P
KIMBROUGH, KATHERYN  AUGUSTA, THE SECOND [@PHENWICK #28]  P
KIMBROUGH, KATHERYN  BARBARA, THE VALIANT [@PHENWICK #16]  P
KIMBROUGH, KATHERYN  CAROL, THE PURSUED [@PHENWICK #29]  P
KIMBROUGH, KATHERYN  DOROTHY, THE TERRIFIED [@PHENWICK #19]  P
KIMBROUGH, KATHERYN  EVELYN, THE AMBITIOUS [@PHENWICK #23]  P
KIMBROUGH, KATHERYN  HARRIET, THE HAUNTED [@PHENWICK #10]  P
KIMBROUGH, KATHERYN  ILENE, THE SUPERSTITIOUS [@PHENWICK #14]  P
KIMBROUGH, KATHERYN  IRIS, THE BEWITCHED [@PHENWICK #39]  P
KIMBROUGH, KATHERYN  ISABELLE, THE FRANTIC [@PHENWICK #22]  P
KIMBROUGH, KATHERYN  JANE, THE COURAGEOUS [@PHENWICK #2]  P
KIMBROUGH, KATHERYN  JOANNE, THE UNPREDICTABLE [@PHENWICK #8]  P
KIMBROUGH, KATHERYN  JOYCE, THE BELOVED [@PHENWICK #27]  P
KIMBROUGH, KATHERYN  KATE, THE CURIOUS [@PHENWICK #13]  P
KIMBROUGH, KATHERYN  KATHERINE, THE RETURNED [@PHENWICK #30]  P
KIMBROUGH, KATHERYN  LAURA, THE IMPERILED [@PHENWICK #38]  P
KIMBROUGH, KATHERYN  LETITIA, THE DREAMER [@PHENWICK #35]  P
KIMBROUGH, KATHERYN  MARCIA, THE INNOCENT [@PHENWICK #12]  P
KIMBROUGH, KATHERYN  MARGARET, THE FAITHFUL [@PHENWICK #3]  P
KIMBROUGH, KATHERYN  MILLIJOY, THE DETERMINED [@PHENWICK #15]  P
KIMBROUGH, KATHERYN  NANCY, THE DARING [@PHENWICK #11]  P
KIMBROUGH, KATHERYN  NELLIE, THE OBVIOUS [@PHENWICK #21]  P
KIMBROUGH, KATHERYN  OLGA, THE DISILLUSIONED [@PHENWICK #32]  P
KIMBROUGH, KATHERYN  OLIVIA, THE TORMENTED [@PHENWICK #9]  P
KIMBROUGH, KATHERYN  OPHELIA, THE ANXIOUS [@PHENWICK #18]  P
KIMBROUGH, KATHERYN  PATRICIA, THE BEAUTIFUL [@PHENWICK #4]  P
KIMBROUGH, KATHERYN  PEGGY, THE CONCERNED [@PHENWICK #31]  P
KIMBROUGH, KATHERYN  POLLY, THE WORRIED [@PHENWICK #25]  P
KIMBROUGH, KATHERYN  RACHEL, THE POSESSESED [@PHENWICK #5]  P
KIMBROUGH, KATHERYN  REBECCA, THE MYSTERIOUS [@PHENWICK #7]  P
KIMBROUGH, KATHERYN  RUTH, THE UNSUSPECTING [@PHENWICK #17]  P
KIMBROUGH, KATHERYN  SUSANNAH, THE RIGHTEOUS [@PHENWICK #6]  P
KIMBROUGH, KATHERYN  UNSEEN TORMENT  P
KIMBROUGH, KATHERYN  URSULA, THE PROUD [@PHENWICK #34]  P
KIMBROUGH, KATHERYN  YVONNE, THE CONFIDENT [@PHENWICK #26]  P
KING, BERNARD  STARKADDER [HATHER LAMBISSON 1]  H
KING, BERNARD  VARGR-MOON [HATHER LAMBISSON 2]  H
KING, CYNTHIA  IN THE MORNING OF TIME  H,P
KING, GABRIEL  GOLDEN CAT, THE [WILD ROAD 2]  P
KING, GABRIEL  WILD ROAD, THE [WILD ROAD 1]  C/P,P
KING-HALL, STEPHEN  MOMENT OF NO RETURN  P
KING, HAROLD  PARADIGM RED  H

KING, J. ROBERT ABDUCTION, THE

[@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #1]

P

KING, J. ROBERT  ABYSSAL WARRIORS [@PLANESCAPE BLOOD WARS #2]  P
KING, J. ROBERT  APOCALYPSE [@MAGIC: THE GATHERING INVASION 3]  C/P,P
KING, J. ROBERT  BLOOD HOSTAGES [@PLANESCAPE BLOOD WARS #1]  P
KING, J. ROBERT  CARNIVAL OF FEAR [RAVENLOFT]  P
KING, J. ROBERT  DRAGONS OF MAGIC, THE [@MAGIC: THE GATHERING]  PA
KING, J. ROBERT  HEART OF MIDNIGHT [@RAVENLOFT]  P

AUTHOR  TITLE  CODE
KING, J. ROBERT  INVASION [@MAGIC: THE GATHERING INVASION 1]  C/P:2,P
KING, J. ROBERT  LANCELOT DU LETHE  P
KING, J. ROBERT  LEGIONS [@MAGIC: THE GATHERING ONSLAUGHT 2]  P
KING, J. ROBERT  MAD MERLIN  P
KING, J. ROBERT  MONSTERS OF MAGIC, THE [@MAGIC: THE GATHERING]  PA
KING, J. ROBERT  ONSLAUGHT [@MAGIC: THE GATHERING ONSLAUGHT 1]  P
KING, J. ROBERT  PLANESHIFT [@MAGIC: THE GATHERING INVASION 2]  C/P,P
KING, J. ROBERT  REALMS OF THE UNDERDARK [@FORGOTTEN REALMS UNDERDARK]  PA
KING, J. ROBERT  SCOURGE [@MAGIC: THE GATHERING ONSLAUGHT 3]  P
KING, J. ROBERT  SUMMERHILL HOUNDS  P
KING, J. ROBERT  THRAN, THE [@MAGIC: THE GATHERING]  P
KING, J. ROBERT  TIME STREAMS [@MAGIC: THE GATHERING ARTIFACTS 3]  P
KING, J. ROBERT  VINAS SOLAMNUS [@DRAGONLANCE LOST LEGENDS #1]  P
KING, JAMES B.  MIRTH OF YORE  SCX
KING, PAULA  MAD ROY’S LIGHT  P,SFWA-P
KING, RUFUS  FATAL KISS MYSTERY, THE  SR-P
KING, STEPHEN  BACHMAN BOOKS, THE  C/P,LP,H
KING, STEPHEN  BAG OF BONES  H
KING, STEPHEN  CARRIE  C/P:2,P
KING, STEPHEN  CHRISTINE  P
KING, STEPHEN  CREEPSHOW  CX
KING, STEPHEN  CUJO  C/P:2,P
KING, STEPHEN  CYCLE OF THE WEREWOLF  SR-LP
KING, STEPHEN  DANSE MACABRE  R/REF-P,C/REF-H
KING, STEPHEN  DARK HALF, THE  C/P,P,H
KING, STEPHEN  DARK TOWER, THE [@DARK TOWER #7]  H:2
KING, STEPHEN  DEAD ZONE, THE  H:2,SR-H,P
KING, STEPHEN  DESPERATION  C/P,P
KING, STEPHEN  DIFFERENT SEASONS  C/P:2,P
KING, STEPHEN  DRAWING OF THE THREE, THE [DARK TOWER 2]  C/P:2,@H:2,P
KING, STEPHEN  DREAMCATCHER  H:2
KING, STEPHEN  EVERYTHING’S EVENTUAL  H
KING, STEPHEN  EYES OF THE DRAGON, THE  C/P,P,H
KING, STEPHEN  FIRESTARTER  H:2,P
KING, STEPHEN  FOUR PAST MIDNIGHT  H:2,P
KING, STEPHEN  FROM A BUICK 8  H
KING, STEPHEN  GERALD’S GAME  P
KING, STEPHEN  GIRL WHO LOVED TOM GORDON, THE  P
KING, STEPHEN  GREEN MILE, THE [GREEN MILE 1,2,3,4,5,6]  P
KING, STEPHEN  GUNSLINGER, THE [DARK TOWER 1]  C/P:2,SR-L,LP
KING, STEPHEN  GUNSLINGER, THE (REVISED) [DARK TOWER 1]  H
KING, STEPHEN  INSOMNIA  P
KING, STEPHEN  IT  C/P,P,H
KING, STEPHEN  MISERY  C/P,P,H
KING, STEPHEN  NEEDFUL THINGS  C/P,P,H
KING, STEPHEN  NIGHT SHIFT  C/P,P,H
KING, STEPHEN  NIGHTMARES AND DREAMSCAPES  C/P,P,H
KING, STEPHEN  PET SEMATARY  C/P,SFWA,P,H
KING, STEPHEN  REGULATORS, THE  P
KING, STEPHEN  ROSE MADDER  C/P,P
KING, STEPHEN  RUNNING MAN, THE  P
KING, STEPHEN  ’SALEM’S LOT  C/P,P,H
KING, STEPHEN  SHINING, THE  H,P
KING, STEPHEN  SKELETON CREW  C/P,P,H:2
KING, STEPHEN  SONG OF SUSANNAH [@DARK TOWER #6]  H
KING, STEPHEN  STAND, THE  C/P:2,P
KING, STEPHEN  STAND, THE (UNABRIDGED)  H
KING, STEPHEN/STRAUB, PETER  TALISMAN, THE  C/P,P,H
KING, STEPHEN  THINNER  C/P:2,P
KING, STEPHEN  TOMMYKNOCKERS, THE  H:2,P
KING, STEPHEN  WASTE LANDS, THE [@DARK TOWER #3]  C/P,SFWA,LP,H
KING, STEPHEN  WIZARD AND GLASS [@DARK TOWER #4]  C/P,LP,H
KING, STEPHEN  WOLVES OF THE CALLA [@DARK TOWER #5]  H:2
KING, T. JACKSON  RETREAD SHOP  P
KING, TABITHA  SMALL WORLD  H,P
KING, TAPPAN/MEACHAM, BETH  NIGHTSHADE [WEIRD HEROES 4]  P
KING, VINCENT  ANOTHER END  C/P:2,P
KING, VINCENT  CANDY MAN  C/P:2,P
KING, VINCENT  LIGHT A LAST CANDLE  C/P,P,H
KING, VINCENT  TIME SNAKE AND SUPERCLOWN  SR-P
KING, WILLIAM  BEASTSLAYER [WARHAMMER GOTREK & FELIX 5]  P
KING, WILLIAM  DAEMONSLAYER [WARHAMMER GOTREK & FELIX 3]  P
KING, WILLIAM  DRAGONSLAYER [WARHAMMER GOTREK & FELIX 4]  P
KING, WILLIAM  FARSEER [WARHAMMER 40K ELDAR]  P

AUTHOR  TITLE  CODE
KING, WILLIAM  GIANTSLAYER [WARHAMMER GOTREK & FELIX 7]  P

KING, WILLIAM GOTREK & FELIX: THE FIRST OMNIBUS [@WARHAMMER GOTREK & FELIX 1,2,3]

H

KING, WILLIAM GOTREK & FELIX: THE SECOND OMNIBUS [@WARHAMMER GOTREK & FELIX 4,5,6]

H

KING, WILLIAM  GREY HUNTER [WARHAMMER 40K SPACE WOLF 3]  P
KING, WILLIAM  RAGNAR’S CLAW [WARHAMMER 40K SPACE WOLF 2]  P
KING, WILLIAM  SKAVENSLAYER [WARHAMMER GOTREK & FELIX 2]  P
KING, WILLIAM  SPACE WOLF [WARHAMMER 40K SPACE WOLF 1]  P
KING, WILLIAM  SPACE WOLF OMNIBUS, THE [@WARHAMMER 40K SPACE WOLF 1,2,3]  H
KING, WILLIAM  TROLLSLAYER [WARHAMMER GOTREK & FELIX 1]  P
KING, WILLIAM  VAMPIRESLAYER [WARHAMMER GOTREK & FELIX 6]  P
KING, WILLIAM  WOLFBLADE [@WARHAMMER 40K SPACE WOLF 4]  P
KINGSBURY, DONALD  COURTSHIP RITE  H:2,SFWA,P
KINGSBURY, DONALD  MOON GODDESS AND THE SON, THE  H:2,P
KINGSBURY, DONALD  PSYCHOHISTORICAL CRISIS  H:2,P
KINGSLEY, CHARLES  WATER BABIES, THE  H
KINLEY, GEORGE  FERRY ROCKET  SR-P
KINROSS, ALBERT  FEARSOME ISLAND, THE  SR-P
KINSELLA, W. P.  IOWA BASEBALL CONFEDERACY, THE  H
KINSELLA, W. P.  SHOELESS JOE  P
KIPLING, RUDYARD  KIPLING’S SCIENCE FICTION  H
KIPLING, RUDYARD  PHANTOM RICKSHAW AND OTHER TALES, THE  SR-H
KIPLING, RUDYARD  PUCK OF POOK’S HILL  SR-H,C/P,SR-P
KIPLING, RUDYARD  REWARDS AND FAIRIES [@WORKS OF KIPLING #25]  SR-H,H
KIRALY, MARIE  MINA  H:2
KIRBY, WILLIAM S.  IAPETUS  P
KIRCHOFF, MARY  BLACK WING, THE [@DRAGONLANCE VILLAINS #2]  C/P,P
KIRCHOFF, MARY/NILES, DOUGLAS  FLINT THE KING [@DRAGONLANCE PRELUDES II 2]  C/P,P
KIRCHOFF, MARY  KENDERMORE [@DRAGONLANCE PRELUDES 2]  C/P:2,P
KIRCHOFF, MARY  KNIGHT OF ILLUSION [ENDLESS QUEST 33]  C/P,P
KIRCHOFF, MARY/WARD, JAMES M.  LIGHT ON QUESTS MOUNTAIN  C/P,P
KIRCHOFF, MARY  MEDUSA PLAGUE, THE [@DRAGONLANCE DEFENDERS OF MAGIC #2]  P
KIRCHOFF, MARY  NIGHT OF THE EYE [@DRAGONLANCE DEFENDERS OF MAGIC #1]  C/P,P
KIRCHOFF, MARY  SEVENTH SENTINEL, THE [@DRAGONLANCE DEFENDERS OF MAGIC #3]  P
KIRCHOFF, MARY/WINTER, STEVE  WANDERLUST [@DRAGONLANCE MEETINGS 2]  C/P:2,P
KIRK, RICHARD  LORDS OF THE SHADOWS [@RAVEN #4]  SR-P
KIRK, RICHARD  SWORDSMISTRESS OF CHAOS [@RAVEN #1]  SR-P
KIRK, RICHARD  TIME OF DYING, A [@RAVEN #5]  SR-P,C/P
KIRK, RICHARD  TIME OF GHOSTS, A [@RAVEN 2]  SR-P
KIRK, RUSSELL  PRINCESS OF ALL LANDS, THE  H
KIRST, HANS HELLMUT  SEVENTH DAY, THE  C/P,P
KIRSTEIN, ROSEMARY  OUTSKIRTER’S SECRET, THE [STEERSWOMAN 2]  C/P,P
KIRSTEIN, ROSEMARY  STEERSWOMAN, THE [STEERSWOMAN 1]  C/P:2,P
KITTREDGE, MARY/O’DONNELL, KEVIN JR.  SHELTER, THE  SFWA-P
KJELGAARD, JIM  FIRE-HUNTER  C/P,P
KLAPER, STEVEN  AGENTS OF INSIGHT  P,SFWA-P
KLASKY, MINDY L.  GLASSWRIGHT’S APPRENTICE, THE [GLASSWRIGHT 1]  C/P,P
KLASKY, MINDY L.  GLASSWRIGHT’S JOURNEYMAN, THE [GLASSWRIGHT 3]  P
KLASKY, MINDY L.  GLASSWRIGHT’S PROGRESS, THE [GLASSWRIGHT 2]  P
KLASKY, MINDY L.  SEASON OF SACRIFICE  P
KLASS, JUDY  CRY OF THE ONLIES, THE [STAR TREK]  C/P:2,P
KLATTE, E./GOLZE, G.  LANGENSCHEIDT’S GERMAN-ENGLISH ENGLISH-GERMAN DICTIONARY  DICT
KLAUSE, ANNETTE CURTIS  ALIEN SECRETS  H
KLAVAN, ANDREW  UNCANNY, THE  H
KLEIER, GLENN  LAST DAY, THE  P
KLEIMAN, LOUIS A.  PROJECT ICARUS  [C/RAN-H]
KLEIN, GERARD  DAY BEFORE TOMORROW, THE  C/P:2,P
KLEIN, GERARD  MOTE IN TIME’S EYE, THE  C/P:2,P
KLEIN, GERARD  OVERLORDS OF WAR, THE  H:2,P,SFWA
KLEIN, GERARD  REENCUENTRO  FP
KLEIN, GERARD  REGRESO AL PASADO  FP
KLEIN, GERARD  STARMASTERS’ GAMBIT  C/P:2,P
KLEIN, ROBERT  HALFWAY ACROSS THE GALAXY AND TURN LEFT  P
KLEIN, T. E. D.  CEREMONIES, THE  P
KLEIN, T. E. D.  DARK GODS  H,P
KLEIN, T. E. D.  GREAT STORIES FROM ROD SERLING’S THE TWILIGHT ZONE MAGAZINE  VLPA
KLEINBAUM, N. H.  D.A.R.Y.L.  C/P,P
KLINE, OTIS ADELBERT  JAN IN INDIA  SR-LP
KLINE, OTIS ADELBERT  JAN OF THE JUNGLE  P
KLINE, OTIS ADELBERT  MAZA OF THE MOON  C/P,P,H
KLINE, OTIS ADELBERT  OUTLAWS OF MARS, THE  C/P,P,H
KLINE, OTIS ADELBERT  PLANET OF PERIL  C/P,P,H
KLINE, OTIS ADELBERT  PORT OF PERIL, THE  C/P,P,H

The MITSFS Pinkdex  September 24, 2007  Page 218
AUTHOR  TITLE  CODE
KLINE, OTIS ADELBERT  PRINCE OF PERIL  C/P,P,H
KLINE, OTIS ADELBERT  SWORDSMAN OF MARS, THE  C/P:2,P
KLINE, OTIS ADELBERT  TAM, SON OF THE TIGER  H
KLINKOWITZ, JEROME/SOMER, JOHN  VONNEGUT STATEMENT, THE  R/REF-P
KNAAK, RICHARD A.  BIRTHRIGHT [@DIABLO SIN WAR #1]  P
KNAAK, RICHARD A.  CHILDREN OF THE DRAKE [DRAGONREALM ORIGINS 2]  P
KNAAK, RICHARD A.  CITADEL, THE [@DRAGONLANCE CLASSICS]  P
KNAAK, RICHARD A.  CRYSTAL DRAGON, THE [DRAGONREALM 5]  C/P,P
KNAAK, RICHARD A.  DAY OF THE DRAGON [@WARCRAFT #1]  P
KNAAK, RICHARD A.  DEMON SOUL, THE [@WARCRAFT WAR OF THE ANCIENTS #2]  P
KNAAK, RICHARD A.  DRAGON HUNT [@WARCRAFT SUNWELL 1]  SCX
KNAAK, RICHARD A.  DRAGON TOME [DRAGONREALM ORIGINS 3]  P
KNAAK, RICHARD A.  DUTCHMAN  P
KNAAK, RICHARD A.  EMPIRE OF BLOOD [@DRAGONLANCE MINOTAUR WARS #3]  H
KNAAK, RICHARD A.  EYE OF CHARON, THE [@AGE OF CONAN SOLDIER’S QUEST #2]  P
KNAAK, RICHARD A.  FIREDRAKE [DRAGONREALM 1]  C/P:2,P
KNAAK, RICHARD A.  FROSTWING  C/P,P
KNAAK, RICHARD A.  GOD IN THE MOON, THE [@AGE OF CONAN SOLDIER’S QUEST #1]  P
KNAAK, RICHARD A.  HORSE KING, THE [DRAGONREALM 7]  P
KNAAK, RICHARD A.  ICEDRAGON [DRAGONREALM 2]  C/P,P
KNAAK, RICHARD A.  JANUS MASK, THE  P
KNAAK, RICHARD A.  KAZ THE MINOTAUR [@DRAGONLANCE HEROES II #1]  C/P:2,P
KNAAK, RICHARD A.  KING OF THE GREY  C/P,P
KNAAK, RICHARD A.  KINGDOM OF SHADOW, THE [@DIABLO #3]  C/P,P
KNAAK, RICHARD A.  LAND OF THE MINOTAURS [@DRAGONLANCE LOST HISTORIES #4]  P
KNAAK, RICHARD A.  LEGACY OF BLOOD [@DIABLO #1]  P
KNAAK, RICHARD A.  LEGEND OF HUMA, THE [@DRAGONLANCE HEROES #1]  C/P:2,P
KNAAK, RICHARD A.  MOON OF THE SPIDER [@DIABLO]  P
KNAAK, RICHARD A.  NIGHT OF BLOOD [@DRAGONLANCE MINOTAUR WARS #1]  P
KNAAK, RICHARD A.  REAVERS OF THE BLOOD SEA [@DRAGONLANCE CHAOS WAR 5]  P
KNAAK, RICHARD A.  RUBY FLAMES [SHATTERED LIGHT 1]  P
KNAAK, RICHARD A.  SCALES OF THE SERPENT [@DIABLO SIN WAR #2]  P
KNAAK, RICHARD A.  SHADOW STEED [DRAGONREALM 4]  P
KNAAK, RICHARD A.  SHADOWS OF ICE [@WARCRAFT SUNWELL 2]  SCX
KNAAK, RICHARD A.  SHROUDED REALM, THE [DRAGONREALM ORIGINS 1]  P
KNAAK, RICHARD A.  SILENT ENEMY, THE [@AGE OF CONAN SOLDIER’S QUEST #3]  P
KNAAK, RICHARD A.  SUNDERING, THE [@WARCRAFT WAR OF THE ANCIENTS #3]  P
KNAAK, RICHARD A.  TIDES OF BLOOD [@DRAGONLANCE MINOTAUR WARS #2]  P

KNAAK, RICHARD A./GOLDEN, CHRISTIE/ GRUBB, JEFF/METZEN, CHRIS WARCRAFT ARCHIVE, THE [WARCRAFT]

H

KNAAK, RICHARD A.  WELL OF ETERNITY, THE [@WARCRAFT WAR OF THE ANCIENTS #1]  P
KNAAK, RICHARD A.  WOLFHELM [DRAGONREALM 3]  P
KNEALE, NIGEL  QUATERMASS  SR-P
KNEALE, NIGEL  QUATERMASS AND THE PIT  SR-P
KNEALE, NIGEL  QUATERMASS EXPERIMENT, THE  SR-P
KNEALE, NIGEL  QUATERMASS II  SR-P
KNEBEL, FLETCHER/BAILEY, CHARLES W. II SEVEN DAYS IN MAY  C/P:2,P
KNEIFEL, HANS  AM RAND DES BLAUEN NEBELS [@PERRY RHODAN #8]  FPR
KNEIFEL, HANS  FLUCHT DER ANDROIDEN [@PERRY RHODAN #147]  FPR
KNEIFEL, HANS  HOLLENTANZ DER MARIONETTEN [@PERRY RHODAN #47]  FPR
KNEIFEL, HANS  INSELN DER BEWAHRUNG, DIE [@PERRY RHODAN #130]  FPR
KNEIFEL, HANS  PFLANZEN DES TODES, DIE [@PERRY RHODAN #43]  FPR
KNEIFEL, HANS  SCHATTEN DES KRISTALLENEN TODES, DIE [@PERRY RHODAN #32]  FPR
KNERR, M. E.  SASQUATCH: MONSTER OF THE NORTHWEST WOODS  P
KNERR, MICHAEL  SEX LIFE OF THE GODS, THE  SR-L
KNIGHT, ANGELA  MASTER OF THE NIGHT [MAGEVERSE MASTERS 1]  P
KNIGHT, DAMON  A FOR ANYTHING  C/P,SFWA-P,P,H
KNIGHT, DAMON  ANALOGUE MEN  C/P:2,P
KNIGHT, DAMON  BEST FROM ORBIT, THE  C/PA,PA
KNIGHT, DAMON  BEST OF DAMON KNIGHT, THE  C/P,P,H:2
KNIGHT, DAMON/WILHELM, KATE  BETTER THAN ONE  SR-H
KNIGHT, DAMON  BEYOND THE BARRIER  H:2,P
KNIGHT, DAMON  BEYOND TOMORROW  C/PA,PA
KNIGHT, DAMON  CENTURY OF GREAT SHORT SCIENCE FICTION NOVELS, A  HA,PA
KNIGHT, DAMON  CENTURY OF SCIENCE FICTION, A  C/PA,PA
KNIGHT, DAMON  CITIES OF WONDER  PA,C/PA
KNIGHT, DAMON  CLARION AWARDS, THE  HA,SFWA
KNIGHT, DAMON  CV [CV 1]  C/P,SFWA-P,P,H
KNIGHT, DAMON  DARK SIDE, THE  HA
KNIGHT, DAMON  DIMENSION X  HA
KNIGHT, DAMON  DOUBLE MEANING  TD34
KNIGHT, DAMON  FAR OUT  H:2,P
KNIGHT, DAMON  FIRST CONTACT  C/PA,PA
KNIGHT, DAMON  FIRST FLIGHT  PA,C/PA

The MITSFS Pinkdex  September 24, 2007  Page 219
AUTHOR  TITLE  CODE

KNIGHT, DAMON/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

FIRST VOYAGES

PA

KNIGHT, DAMON  FUTURIANS, THE  R/REF-H
KNIGHT, DAMON  GOLDEN ROAD, THE  HA
KNIGHT, DAMON  HELL’S PAVEMENT  SR-P,C/P:2
KNIGHT, DAMON  HUMPTY DUMPTY: AN OVAL  H,LP
KNIGHT, DAMON  IN DEEP  C/P:2,P
KNIGHT, DAMON  IN SEARCH OF WONDER  R/REF-H,C/REF-H:2
KNIGHT, DAMON  LATE KNIGHT EDITION  R/REF-H
KNIGHT, DAMON  MAN IN THE TREE, THE  P
KNIGHT, DAMON  MASTERS OF EVOLUTION  DD375
KNIGHT, DAMON  METAL SMILE, THE  PA,C/PA
KNIGHT, DAMON  MIND SWITCH  C/P:2,P
KNIGHT, DAMON  MONAD 2  R/REF-P
KNIGHT, DAMON  NEBULA AWARD STORIES  PA,C/PA
KNIGHT, DAMON  NEBULA AWARD STORIES 1965  HA
KNIGHT, DAMON  NOW BEGINS TOMORROW  C/PA,PA
KNIGHT, DAMON  OBSERVERS, THE [CV 2]  H,SFWA-P
KNIGHT, DAMON  OFF CENTER  C/P,DM113,P
KNIGHT, DAMON  ONE SIDE LAUGHING  H
KNIGHT, DAMON  ORBIT 1  C/PA,PA
KNIGHT, DAMON  ORBIT 2  PA,C/PA
KNIGHT, DAMON  ORBIT 3  PA,C/PA
KNIGHT, DAMON  ORBIT 4  C/PA,PA
KNIGHT, DAMON  ORBIT 5  C/PA,PA
KNIGHT, DAMON  ORBIT 6  C/PA,PA
KNIGHT, DAMON  ORBIT 7  C/PA,PA
KNIGHT, DAMON  ORBIT 8  C/PA,PA
KNIGHT, DAMON  ORBIT 9  C/PA,PA
KNIGHT, DAMON  ORBIT 10  C/PA,PA
KNIGHT, DAMON  ORBIT 11  HA,PA
KNIGHT, DAMON  ORBIT 12  C/PA,PA
KNIGHT, DAMON  ORBIT 13  C/PA,PA
KNIGHT, DAMON  ORBIT 14  HA,SFWA
KNIGHT, DAMON  ORBIT 15  HA,SFWA
KNIGHT, DAMON  ORBIT 16  HA,SFWA:2
KNIGHT, DAMON  ORBIT 17  SFWA
KNIGHT, DAMON  ORBIT 18  HA,SFWA
KNIGHT, DAMON  ORBIT 19  HA,SFWA
KNIGHT, DAMON  ORBIT 20  HA,SFWA
KNIGHT, DAMON  ORBIT 21  HA,SFWA
KNIGHT, DAMON  OTHER FOOT, THE  C/P,P
KNIGHT, DAMON  PEOPLE MAKER, THE  SR-P
KNIGHT, DAMON  PERCHANCE TO DREAM  C/PA,PA,SFWA
KNIGHT, DAMON  POCKETFUL OF STARS, A  HA,PA,SFWA
KNIGHT, DAMON  REASONABLE WORLD, A [CV 3]  H:2,SFWA,P
KNIGHT, DAMON  RITHIAN TERROR, THE  C/P,DM113,P
KNIGHT, DAMON  RULE GOLDEN  TD34
KNIGHT, DAMON  RULE GOLDEN AND OTHER STORIES  P
KNIGHT, DAMON  SCIENCE FICTION ARGOSY, A  HA
KNIGHT, DAMON  SCIENCE FICTION INVENTIONS  C/PA,PA
KNIGHT, DAMON  SCIENCE FICTION OF THE THIRTIES  HA
KNIGHT, DAMON  SHAPE OF THINGS, THE  PA,C/PA
KNIGHT, DAMON  SHOCKING THING, A  PA
KNIGHT, DAMON  SUN SABOTEURS, THE  DF108
KNIGHT, DAMON  THIRTEEN FRENCH SCIENCE FICTION STORIES  PA,C/PA
KNIGHT, DAMON  THREE NOVELS  H:2,P
KNIGHT, DAMON  TOMORROW X 4  PA,C/PA
KNIGHT, DAMON  TOWARD INFINITY  HA
KNIGHT, DAMON  TURNING ON  C/P:2,P
KNIGHT, DAMON  TURNING POINTS: ESSAYS ON THE ART OF SCIENCE FICTION  SFWA,R/REF-H
KNIGHT, DAMON  WHY DO BIRDS  H
KNIGHT, DAMON  WORLD AND THORINN, THE  H:2,P,SFWA

KNIGHT, DAMON WORLD WITHOUT CHILDREN/EARTH QUARTER, THE (TWO DAMON KNIGHT SCIENCE FICTION NOVELS)

C/P:2,P

KNIGHT, DAMON  WORLDS TO COME  C/PA,PA
KNIGHT, E. E.  CHOICE OF THE CAT [@VAMPIRE EARTH #2]  P
KNIGHT, E. E.  DRAGON AVENGER [@AGE OF FIRE #2]  LP
KNIGHT, E. E.  DRAGON CHAMPION [@AGE OF FIRE #1]  LP
KNIGHT, E. E.  TALE OF THE THUNDERBOLT [@VAMPIRE EARTH #3]  P
KNIGHT, E. E.  VALENTINE’S EXILE [@VAMPIRE EARTH #5]  H
KNIGHT, E. E.  VALENTINE’S RISING [@VAMPIRE EARTH #4]  P
KNIGHT, E. E.  WAY OF THE WOLF [@VAMPIRE EARTH #1]  P
KNIGHT, ERIC  FLYING YORKSHIREMAN, THE  C/P,P,H
KNIGHT, HARRY ADAM  CARNOSAUR  P

The MITSFS Pinkdex  September 24, 2007  Page 220
AUTHOR  TITLE  CODE
KNIGHT, HARRY ADAM  FUNGUS, THE  H
KNIGHT, MALLORY T.  DIRTY, ROTTEN, DEPRIVING RAY, THE [@MAN FROM TOMCAT #4]  P
KNITTEL, JOHN  NILE GOLD  H
KNOBEL, PHILIP  MR. MOON  P
KNOEPFLMACHER, U. C.  CHRISTMAS CAROL AND OTHER VICTORIAN FAIRY TALES, A  PA
KNOX, CALVIN M.  LEST WE FORGET THEE, EARTH  DD291
KNOX, CALVIN M.  ONE OF OUR ASTEROIDS IS MISSING  DF253
KNOX, CALVIN M.  PLOT AGAINST EARTH, THE  DD358
KNOX, ELIZABETH  DREAMHUNTER [DREAMHUNTER 1]  H
KNOX, ELIZABETH  DREAMQUAKE [DREAMHUNTER 2]  H
KNOX, RICHARD A.  MEMORIES OF THE FUTURE  SR-H
KOCH, HOWARD  ORSON WELLES’ ...WAR OF THE WORLDS (CASSETTE)  SR-MM
KOCH, HOWARD  PANIC BROADCAST, THE  R/REF-P,C/REF-P
KOCH, R.  ERDE GEHT NICHT UNTER, DIE  FH

KOCHER, PAUL H.

MASTER OF MIDDLE-EARTH: THE FICTION OF J. R. R. TOLKIEN R/REF-P,C/REF-H,

C/REF-P

KOCHER, PAUL H.  READER’S GUIDE TO THE SILMARILLION, A  R/REF-H,C/REF-H
KOENIG, WALTER  BUCK ALICE AND THE ACTOR-ROBOT  P
KOENIG, WALTER  CHEKOV’S ENTERPRISE  R/REF-P
KOHLL, MALCOLM  DELTA AND THE BANNERMEN (MISSPELLED ON SPINE) [DOCTOR WHO]  SR-P,C/P
KOKE, JAK  BEYOND THE PALE [SHADOWRUN 30/DRAGON HEART 3]  P
KOKE, JAK  CLOCKWORK ASYLUM [@SHADOWRUN DRAGON HEART 2]  P
KOKE, JAK  DEAD AIR [@SHADOWRUN 22]  P
KOKE, JAK  STRANGER SOULS [SHADOWRUN 26/DRAGON HEART 1]  P
KOLNBERGER, ANTON  AUF UNBEKANNTEM STERN  FH
KOLUPAEV, VICTOR  HERMIT’S SWING  H
KOMAN, VICTOR  JEHOVAH CONTRACT, THE  H,SFWA,P
KOMAN, VICTOR  KINGS OF THE HIGH FRONTIER [HIGH PILGRIMAGE 1]  H
KOMAN, VICTOR  SOLOMON’S KNIFE  H
KOMATSU, SAKYO  JAPAN SINKS  H
KONVITZ, JEFFREY  GUARDIAN, THE  P
KONVITZ, JEFFREY  MONSTER  P
KONWICKI, TADEUSZ  ANTHROPOS-SPECTER BEAST, THE  H
KOOIKER, LEONIE  MAGIC STONE, THE  SFWA
KOOISTRA, JEFFREY D.  DYKSTRA’S WAR  C/P,P
KOONTZ, DEAN R.  ANTI-MAN  P
KOONTZ, DEAN R./KOONTZ, GERDA  APHRODISIAC GIRL  SR-L
KOONTZ, DEAN R.  BEASTCHILD  P
KOONTZ, DEAN R.  BY THE LIGHT OF THE MOON  H
KOONTZ, DEAN R.  COLD FIRE  P
KOONTZ, DEAN R.  CRIMSON WITCH, THE  SR-P,C/P:2
KOONTZ, DEAN R.  DARK OF THE WOODS  D13793
KOONTZ, DEAN R.  DARK SYMPHONY, THE  C/P,P
KOONTZ, DEAN R.  DARKFALL  P
KOONTZ, DEAN R.  DARKNESS IN MY SOUL, A  C/P:2,P
KOONTZ, DEAN R.  DEMON SEED  C/P:2,P
KOONTZ, DEAN R.  DOOR TO DECEMBER, THE  P
KOONTZ, DEAN R.  DRAGON TEARS  H:2,P
KOONTZ, DEAN R.  FACE, THE  SR-LP
KOONTZ, DEAN R.  FACE OF FEAR, THE  C/P,P
KOONTZ, DEAN R.  FALL OF THE DREAM MACHINE, THE  D22600
KOONTZ, DEAN R.  FEAR THAT MAN  D23140
KOONTZ, DEAN R.  FLESH IN THE FURNACE, THE  C/P:2,P
KOONTZ, DEAN R.  HAUNTED EARTH, THE  C/P,P
KOONTZ, DEAN R.  HELL’S GATE  P
KOONTZ, DEAN R.  LIFE EXPECTANCY  H
KOONTZ, DEAN R.  LIGHTNING  C/P,P
KOONTZ, DEAN R.  MASQUE DE L’OUBLI, LE  FP
KOONTZ, DEAN R.  MIDNIGHT  H
KOONTZ, DEAN R.  MIROIRS DE SANG  FP
KOONTZ, DEAN R.  NIGHT CHILLS  H,P
KOONTZ, DEAN R.  NIGHTMARE JOURNEY  H,P
KOONTZ, DEAN R.  ODD THOMAS  H
KOONTZ, DEAN R.  ONE DOOR AWAY FROM HEAVEN  P
KOONTZ, DEAN R.  SOFT COME THE DRAGONS  D13793
KOONTZ, DEAN R.  SOLE SURVIVOR  P
KOONTZ, DEAN R.  STAR QUEST  DH070
KOONTZ, DEAN R.  STARBLOOD  C/P,P
KOONTZ, DEAN R.  STRANGE HIGHWAYS  P
KOONTZ, DEAN R.  STRANGERS  H
KOONTZ, DEAN R.  TAKING, THE  H
KOONTZ, DEAN R.  TIME THIEVES  D00990
KOONTZ, DEAN R.  TWILIGHT EYES  P
KOONTZ, DEAN R.  VISION, THE  H,P
KOONTZ, DEAN R.  WARLOCK  C/P,P

AUTHOR  TITLE  CODE
KOONTZ, DEAN R.  WATCHERS  C/P,SFWA-P,P,H
KOONTZ, DEAN R.  WEREWOLF AMONG US, A  C/P:2,P
KOONTZ, DEAN R.  WHISPERS  P
KORNBLUTH, C. M.  BEST OF C. M. KORNBLUTH, THE  H:2,P
KORNBLUTH, C. M.  EXPLORERS, THE  C/P:2,P
KORNBLUTH, C. M./MERRIL, JUDITH  GUNNER CADE  C/P:2,P
KORNBLUTH, C. M.  HIS SHARE OF GLORY  H:2
KORNBLUTH, C. M.  MARCHING MORONS, THE  C/P:2,P
KORNBLUTH, C. M.  MILE BEYOND THE MOON, A  H:2,P
KORNBLUTH, C. M.  NOT THIS AUGUST  H:2,P
KORNBLUTH, C. M.  NOT THIS AUGUST (REVISED)  P,SFWA-P
KORNBLUTH, C. M.  SYNDIC, THE  H:2,P
KORNBLUTH, C. M.  TAKEOFF  C/P,P,H
KORNBLUTH, C. M.  THIRTEEN O’CLOCK AND OTHER ZERO HOURS  C/P:2,P
KORNBLUTH, MARY  SCIENCE FICTION SHOWCASE  HA,PA
KOSINSKI, JERZY  BEING THERE  C/P,P
KOSINSKI, JERZY  PAINTED BIRD, THE (REVISED)  P
KOSTER, R. M.  CARMICHAEL’S DOG  H
KOTANI, ERIC/ROBERTS, JOHN MADDOX  ACT OF GOD [ISLAND WORLDS 1]  P
KOTANI, ERIC/ROBERTS, JOHN MADDOX  BETWEEN THE STARS [ISLAND WORLDS 3]  C/P:2,SFWA-P,P
KOTANI, ERIC/SMITH, DEAN WESLEY  DEATH OF A NEUTRON STAR [STAR TREK VOYAGER]  C/P:2,P
KOTANI, ERIC/ROBERTS, JOHN MADDOX  DELTA PAVONIS [ISLAND WORLDS 4]  C/P,P
KOTANI, ERIC/ROBERTS, JOHN MADDOX  ISLAND WORLDS, THE [ISLAND WORLDS 2]  C/P,P
KOTZWINKLE, WILLIAM  DOCTOR RAT  H,P
KOTZWINKLE, WILLIAM  E.T.: THE BOOK OF THE GREEN PLANET  C/P:2,P
KOTZWINKLE, WILLIAM  E.T.: THE EXTRA-TERRESTRIAL  C/P,P,H
KOTZWINKLE, WILLIAM  E.T.: THE EXTRA-TERRESTRIAL STORYBOOK  H
KOTZWINKLE, WILLIAM  GREAT WORLD CIRCUS  H
KOTZWINKLE, WILLIAM  JEWEL OF THE MOON  H
KOTZWINKLE, WILLIAM  LEOPARD’S TOOTH, A  P
KOTZWINKLE, WILLIAM  QUEEN OF SWORDS  H,P
KOTZWINKLE, WILLIAM  SUPERMAN III [DC COMICS]  C/P,P
KOVALIC, JOHN  DORK COVENANT  CX
KOZERSKI, PAUL  SKYLOCK  P
KRAMER, DANA  WARHORN  SFWA-P
KRAMER, EDWARD E.  DARK DESTINY III: CHILDREN OF DRACULA  PA
KRAMER, EDWARD E.  TALES OF THE WHITE WOLF [ELRIC]  PA
KRAMER, EDWARD E./CROWTHER, PETER  TOMBS  PA
KRAMER, INA  LIONESS, THE [REALMS OF ARKANIA THALIONMEL 1]  P
KRAMER, INA  SACRIFICE, THE [REALMS OF ARKANIA THALIONMEL 2]  P
KRAMER, NORA  ARROW BOOK OF GHOST STORIES  PA
KRAMER-ROLLS, DANA  HOME IS THE HUNTER [STAR TREK]  C/P:2,P
KRANE, STEVEN  STRANGER INSIDE  P
KRATMAN, TOM  STATE OF DISOBEDIENCE, A (PROOF)  SR-LP
KRAUSE, STALEY/WIECK, STEWART  DEATH AND DAMNATION [WORLD OF DARKNESS WRAITH]  PA
KRAUSE, STALEY/WIECK, STEWART  STRANGE CITY [WORLD OF DARKNESS]  C/PA,PA
KRAUSE, STALEY/WIECK, STEWART  TRUTH UNTIL PARADOX  PA
KRAUSS, LAWRENCE M.  BEYOND STAR TREK  R/REF-P,R/REF-H
KRAUSS, LAWRENCE M.  PHYSICS OF STAR TREK, THE [STAR TREK]  R/REF-P,C/REF-H
KRAUZER, STEVEN M.  BRAINSTORM  P
KREIGHBAUM, MARK  EYES OF GOD, THE [PINCH 2]  P
KRENSKY, STEPHEN  DRAGON CIRCLE, THE  LP
KRENTZ, JAYNE ANN  SWEET STARFIRE  P
KRENTZ, JAYNE ANN  WHITE LIES [ARCANE SOCIETY]  H
KRESS, NANCY  ALIEN LIGHT, AN  H:2
KRESS, NANCY  ALIENS, OF EARTH, THE  H
KRESS, NANCY  BEAKER’S DOZEN  H,LP
KRESS, NANCY  BEGGARS AND CHOOSERS [BEGGARS 2]  H:2,P
KRESS, NANCY  BEGGARS IN SPAIN [BEGGARS 1]  H:2,P
KRESS, NANCY  BEGGARS RIDE [BEGGARS 3]  C/P:2,P
KRESS, NANCY  BRAINROSE  C/P,P
KRESS, NANCY  CROSSFIRE [CROSSFIRE 1]  H
KRESS, NANCY  CRUCIBLE [CROSSFIRE 2]  H
KRESS, NANCY  DANCING ON AIR  LP
KRESS, NANCY  GOLDEN GROVE, THE  C/P,SFWA,P
KRESS, NANCY  MAXIMUM LIGHT  P
KRESS, NANCY  NEBULA AWARDS SHOWCASE 2003  LPA
KRESS, NANCY  NOTHING HUMAN  H
KRESS, NANCY  OATHS AND MIRACLES  C/P,P
KRESS, NANCY  PRICE OF ORANGES, THE  B
KRESS, NANCY  PRINCE OF MORNING BELLS, THE  C/P,P
KRESS, NANCY  PROBABILITY MOON [FALLERS 1]  H:2,P
KRESS, NANCY  PROBABILITY SPACE [FALLERS 3]  H
KRESS, NANCY  PROBABILITY SUN [FALLERS 2]  H

The MITSFS Pinkdex  September 24, 2007  Page 222
AUTHOR  TITLE  CODE
KRESS, NANCY  STINGER  C/P,P
KRESS, NANCY  TRINITY  H
KRESS, NANCY  TRINITY AND OTHER STORIES  P
KRESS, NANCY  WHITE PIPES, THE  C/P,P,H
KRESS, NANCY  YANKED! [@OUT OF TIME #1]  P
KRINARD, SUSAN  SHIELD OF THE SKY  LP
KRING, MICHAEL  CHILDREN OF THE NIGHT [SPACE MAVERICKS 2]  P
KRING, MICHAEL  SPACE MAVERICKS, THE [SPACE MAVERICKS 1]  P
KRITZER, NAOMI  FIRES OF THE FAITHFUL [ELIANA 1]  P
KRITZER, NAOMI  FREEDOM’S APPRENTICE [DEAD RIVERS 2]  P
KRITZER, NAOMI  FREEDOM’S GATE [DEAD RIVERS 1]  C/P,P
KRITZER, NAOMI  TURNING THE STORM [ELIANA 2]  P
KRONENBERGER, LOUIS  CAVALCADE OF COMEDY  HA
KROPP, LLOYD  DRIFT, THE  H,P
KRUCHTEN, MARCIA H.  SKYBORN  P
KRUGER, ROBERT/SWENSON, PATRICK  IMAGINATION FULLY DILATED: SCIENCE FICTION  LPA
KRULIK, NANCY  CORDELIA COLLECTION, THE, VOL. 1 [@BUFFY]  P
KRULIK, NANCY  PASSENGERS, THE [LOST IN SPACE NEW JOURNEYS]  LP
KRULIK, THEODORE  COMPLETE AMBER SOURCEBOOK, THE [AMBER]  R/REF-H
KUBASIK, CHRIS  CHANGELING [@SHADOWRUN #5]  C/P,SFWA-P,P
KUBASIK, CHRIS  IDEAL WAR [@BATTLETECH 9]  C/P,P
KUBASIK, CHRISTOPHER  LONGING RING, THE [@EARTHDAWN]  P
KUBASIK, CHRISTOPHER  POISONED MEMORIES [@EARTHDAWN]  P
KUBE-MCDOWELL, MICHAEL P.  ALTERNITIES  H,SFWA,SFWA-P,P
KUBE-MCDOWELL, MICHAEL P.  BEFORE THE STORM [@STAR WARS BLACK FLEET CRISIS 1]  C/P:2,P
KUBE-MCDOWELL, MICHAEL P.  EMPERY [TRIGON DISUNITY 3]  C/P:2,SFWA-P,P
KUBE-MCDOWELL, MICHAEL P.  EMPRISE [TRIGON DISUNITY 1]  C/P:2,P
KUBE-MCDOWELL, MICHAEL P.  ENIGMA [TRIGON DISUNITY 2]  C/P:2,P
KUBE-MCDOWELL, MICHAEL P.  EXILE  H,SFWA,SFWA-P:2
KUBE-MCDOWELL, MICHAEL P.  ODYSSEY [ROBOT CITY 1]  C/P:2,P
KUBE-MCDOWELL, MICHAEL P.  QUIET POOLS, THE  H:2,SFWA:2,SFWA-P:2,P
KUBE-MCDOWELL, MICHAEL P.  SHIELD OF LIES [STAR WARS BLACK FLEET CRISIS 2]  C/P:2,P
KUBE-MCDOWELL, MICHAEL P.  TYRANT’S TEST [@STAR WARS BLACK FLEET CRISIS 3]  C/P,P
KUBE-MCDOWELL, MICHAEL P.  VECTORS  P
KUBILIUS, WALTER  OTHER SIDE, THE  B
KUBRICK, STANLEY  STANLEY KUBRICK’S A CLOCKWORK ORANGE  LP
KUEBLER, HAROLD W.  TREASURY OF SCIENCE FICTION CLASSICS, THE  SR-HA,HA
KUMMER, FREDERIC ARNOLD  GENTLEMEN IN HADES  H
KUNSTLER, JAMES HOWARD  HUNT, THE  SFWA-P
KUPPERBERG, PAUL  CRIME CAMPAIGN  P
KUPPERBERG, PAUL  MURDERMOON  P
KURLAND, LYNN  STAR OF THE MORNING [NINE KINGDOMS]  LP
KURLAND, MICHAEL/BARTON, S. W.  LAST PRESIDENT, THE  P
KURLAND, MICHAEL  PERCHANCE  P
KURLAND, MICHAEL  PLURIBUS  C/P,SFWA,P,H
KURLAND, MICHAEL  PRINCES OF EARTH, THE  H
KURLAND, MICHAEL  PSI HUNT  C/P:2,P
KURLAND, MICHAEL  STAR GRIFFIN  H,SFWA
KURLAND, MICHAEL  STUDY IN SORCERY, A [LORD DARCY 6]  C/P:2,P
KURLAND, MICHAEL  TEN LITTLE WIZARDS [LORD DARCY 5]  C/P:2,SFWA-P,P
KURLAND, MICHAEL  TOMORROW KNIGHT  C/P,P
KURLAND, MICHAEL  TRANSMISSION ERROR  C/P,P
KURLAND, MICHAEL  UNICORN GIRL, THE  C/P:2,P
KURLAND, MICHAEL  WHENABOUTS OF BURR, THE  C/P,P
KURTEN, BJORN  DANCE OF THE TIGER  C/P:2,P

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER ADEPT, THE [@ADEPT #1]

C/P:2,SFWA-P:2,P

KURTZ, KATHERINE  BASTARD PRINCE, THE [DERYNI HEIRS 3]  H:2,P
KURTZ, KATHERINE  BISHOP’S HEIR, THE [DERYNI KELSON 1]  C/P,P,H,SFWA
KURTZ, KATHERINE  CAMBER OF CULDI [DERYNI CAMBER 1]  C/P,P,H
KURTZ, KATHERINE  CAMBER THE HERETIC [DERYNI CAMBER 3]  C/P:2,P,H
KURTZ, KATHERINE  CHILDE MORGAN [DERYNI]  H
KURTZ, KATHERINE  CHRONICLES OF THE DERYNI, THE [DERYNI CHRONICLES 1,2,3]  H
KURTZ, KATHERINE/REGINALD, ROBERT  CODEX DERYNIANUS II [DERYNI]  R/REF-H
KURTZ, KATHERINE  CRUSADE OF FIRE  PA

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER DAGGER MAGIC [ADEPT 4]

C/P:2,P,H

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER DEATH OF AN ADEPT [ADEPT 5]

H,P

KURTZ, KATHERINE  DERYNI ARCHIVES, THE [DERYNI]  C/P,P,H
KURTZ, KATHERINE  DERYNI CHECKMATE [DERYNI CHRONICLES 2]  C/P:3,P
KURTZ, KATHERINE  DERYNI MAGIC [DERYNI]  C/P:2,P
KURTZ, KATHERINE  DERYNI RISING [DERYNI CHRONICLES 1]  C/P:3,P
KURTZ, KATHERINE  DERYNI TALES [DERYNI]  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 223
AUTHOR  TITLE  CODE
KURTZ, KATHERINE  HARROWING OF GWYNEDD, THE [DERYNI HEIRS 1]  C/P,P,H
KURTZ, KATHERINE  HIGH DERYNI [DERYNI CHRONICLES 3]  C/P:2,SR-P,P
KURTZ, KATHERINE  IN THE KING’S SERVICE [@DERYNI]  P
KURTZ, KATHERINE  KING JAVAN’S YEAR [DERYNI HEIRS 2]  H
KURTZ, KATHERINE  KING KELSON’S BRIDE [DERYNI KELSON 4]  H
KURTZ, KATHERINE  KING’S JUSTICE, THE [DERYNI KELSON 2]  H,P,SFWA
KURTZ, KATHERINE  LAMMAS NIGHT  C/P:2,P
KURTZ, KATHERINE  LEGACY OF LEHR, THE  C/P,P,H

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER LODGE OF THE LYNX, THE [@ADEPT #2]

C/P:2,SFWA-P,P

KURTZ, KATHERINE  ON CRUSADE: MORE TALES OF THE KNIGHTS TEMPLAR  LPA
KURTZ, KATHERINE  QUEST FOR SAINT CAMBER, THE [DERYNI KELSON 3]  C/P,H,P
KURTZ, KATHERINE  SAINT CAMBER [DERYNI CAMBER 2]  C/P,P,H
KURTZ, KATHERINE  ST. PATRICK’S GARGOYLE  H
KURTZ, KATHERINE  TALES OF THE KNIGHTS TEMPLAR  C/PA,PA

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER TEMPLAR TREASURE, THE [@ADEPT #3]

C/P:2,SFWA-P,P

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER TEMPLE AND THE CROWN, THE

P

KURTZ, KATHERINE/ HARRIS, DEBORAH TURNER TEMPLE AND THE STONE, THE

H

KURTZ, KATHERINE  TWO CROWNS FOR AMERICA  H,P
KURTZMAN, HARVEY/ELDER, WILL  PLAYBOY’S LITTLE ANNIE FANNY  CX
KURZAWA, GREG  GIDEON’S WALL  LP
KUSHNER, DONN  BOOK DRAGON, A  C/P:2,P
KUSHNER, ELLEN  BASILISK  C/PA,PA
KUSHNER, ELLEN/SHERMAN, DELIA  FALL OF THE KINGS, THE [SWORDSPOINT 2]  C/P,LP
KUSHNER, ELLEN  PRIVILEGE OF THE SWORD, THE [SWORDSPOINT]  H,LP
KUSHNER, ELLEN  SWORDSPOINT [SWORDSPOINT 1]  C/P,SFWA-P,P,H
KUSHNER, ELLEN  THOMAS THE RHYMER  H:2,P
KUTTNER, HENRY  AHEAD OF TIME  C/P:2,P
KUTTNER, HENRY  BEST OF HENRY KUTTNER, THE  H:2,P
KUTTNER, HENRY  BYPASS TO OTHERNESS  C/P:2,P
KUTTNER, HENRY  CREATURE FROM BEYOND INFINITY, THE  C/P:2,P
KUTTNER, HENRY  DARK WORLD, THE  C/P:2,P
KUTTNER, HENRY/MOORE, C. L.  DEJA DEMAIN  FP
KUTTNER, HENRY  DESTINATION: INFINITY  C/P:2,P
KUTTNER, HENRY  FURY  C/P,P,H
KUTTNER, HENRY  MASK OF CIRCE, THE  C/P:2,P
KUTTNER, HENRY  MURDER OF A MISTRESS  P
KUTTNER, HENRY  MUTANT  C/P:2,P
KUTTNER, HENRY/MOORE, C. L.  NO BOUNDARIES  SR-H,C/P:2,P
KUTTNER, HENRY  RETURN TO OTHERNESS  C/P:2,P
KUTTNER, HENRY  ROBOTS HAVE NO TAILS  C/P:2,P
KUTTNER, HENRY  STARTLING WORLDS OF HENRY KUTTNER, THE  P
KUTTNER, HENRY  TIME AXIS, THE  C/P:2,P
KUTTNER, HENRY/MOORE, C. L.  TWO-HANDED ENGINE  H
KUTTNER, HENRY  VALLEY OF THE FLAME  C/P:2,P
KUTTNER, HENRY  WAY OF THE GODS  B
KUTTNER, HENRY  WELL OF THE WORLDS, THE  C/P:2,P
KYLE, DAVID A.  DRAGON LENSMAN, THE [LENSMEN 7]  C/P:2,P
KYLE, DAVID A.  LENSMAN FROM RIGEL [LENSMEN 8]  C/P:2,P
KYLE, DAVID A.  Z-LENSMAN [LENSMEN 9]  C/P:2,P
KYME, NICK  BACK FROM THE DEAD [@NECROMUNDA]  P

The MITSFS Pinkdex  September 24, 2007  Page 225
AUTHOR  TITLE  CODE
LA PLANTE, RICHARD  TEGNE: SOUL WARRIOR  P
LA SALLE, VICTOR  AFTER THE ATOM  SR-P
LA SALLE, VICTOR  SEVENTH DIMENSION, THE  SR-P
LA SALLE, VICTOR  SUNS IN DUO  SR-P
LABONTE, RICHARD/SCHIMEL, LAWRENCE  FUTURE IS QUEER, THE  LPA
LACKEY, MERCEDES  AERIE [DRAGON JOUSTERS 4]  H
LACKEY, MERCEDES  ALTA [DRAGON JOUSTERS 2]  H,P
LACKEY, MERCEDES  ARROW’S FALL [HERALDS OF VALDEMAR 3]  C/P:2,P
LACKEY, MERCEDES  ARROW’S FLIGHT [HERALDS OF VALDEMAR 2]  C/P:2,P
LACKEY, MERCEDES  ARROWS OF THE QUEEN [HERALDS OF VALDEMAR 1]  C/P:2,P
LACKEY, MERCEDES/GUON, ELLEN  BEDLAM’S BARD [ERIC BANYON 1,2]  H,P
LACKEY, MERCEDES/EDGHILL, ROSEMARY  BEDLAM’S EDGE  HA
LACKEY, MERCEDES/EDGHILL, ROSEMARY  BEYOND WORLD’S END [ERIC BANYON 3]  H:2
LACKEY, MERCEDES/DIXON, LARRY  BLACK GRYPHON, THE [MAGE WARS 1]  C/P,SFWA,P,H
LACKEY, MERCEDES  BLACK SWAN, THE  C/P,P,H
LACKEY, MERCEDES/DIXON, LARRY  BORN TO RUN [SERRATED EDGE 1]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  BRIGHTLY BURNING [VALDEMAR]  H:2,P
LACKEY, MERCEDES  BURNING WATER  C/P:2,P
LACKEY, MERCEDES/GELLIS, ROBERTA  BY SLANDEROUS TONGUES  H
LACKEY, MERCEDES  BY THE SWORD [KEROWYN’S TALE]  C/P:2,P
LACKEY, MERCEDES/SHERMAN, JOSEPHA  CAST OF CORBIES, A [BARDIC CHOICES]  C/P:2,P
LACKEY, MERCEDES/SHERMAN, JOSEPHA  CASTLE OF DECEPTION [BARD’S TALE 1]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  CHILDREN OF THE NIGHT [DIANA TREGARDE]  C/P:2,P
LACKEY, MERCEDES/DIXON, LARRY  CHROME BORNE, THE [SERRATED EDGE 1,4]  C/P,P
LACKEY, MERCEDES/DIXON, LARRY  CHROME CIRCLE [SERRATED EDGE 4]  C/P:2,P
LACKEY, MERCEDES  CROSSROADS [VALDEMAR]  PA
LACKEY, MERCEDES  EAGLE & THE NIGHTINGALES, THE [@BARDIC VOICES #3]  H,P
LACKEY, MERCEDES  EXILE’S HONOR [VALDEMAR]  H,P
LACKEY, MERCEDES  EXILE’S VALOR [VALDEMAR]  H:2
LACKEY, MERCEDES  EXILE’S VALOR (PROOF) [VALDEMAR]  SR-LP
LACKEY, MERCEDES  FAIRY GODMOTHER, THE [FIVE HUNDRED KINGDOMS]  P
LACKEY, MERCEDES  FIDDLER FAIR  C/P,P
LACKEY, MERCEDES  FIRE ROSE, THE  C/P:2,P
LACKEY, MERCEDES  FIREBIRD  C/P,P,H
LACKEY, MERCEDES  FLIGHTS OF FANTASY  C/PA,PA
LACKEY, MERCEDES/EMERSON, RU  FORTRESS OF FROST AND FIRE [BARD’S TALE 2]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  FORTUNE’S FOOL [FIVE HUNDRED KINGDOMS]  H
LACKEY, MERCEDES  FOUR AND TWENTY BLACKBIRDS [BARDIC VOICES 4]  H:2
LACKEY, MERCEDES/GUON, ELLEN  FREEDOM FLIGHT [WING COMMANDER 1]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  GATES OF SLEEP, THE [ELEMENTAL MASTERS 2]  H
LACKEY, MERCEDES/GELLIS, ROBERTA  ILL MET BY MOONLIGHT [SCEPTER’D ISLE 2]  H
LACKEY, MERCEDES/GELLIS, ROBERTA  ILL MET BY MOONLIGHT (PROOF) [SCEPTER’D ISLE 2]  SR-LP
LACKEY, MERCEDES  JINX HIGH [DIANA TREGARDE]  C/P:2,P
LACKEY, MERCEDES  JOUST [DRAGON JOUSTERS 1]  H:2
LACKEY, MERCEDES/GUON, ELLEN  KNIGHT OF GHOSTS AND SHADOWS [ERIC BANYON 1]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  LARK AND THE WREN, THE [BARDIC VOICES 1]  C/P:2,P,H
LACKEY, MERCEDES  LARK AND THE WREN, THE (EXCERPT) [BARDIC VOICES 1]  SR-LP
LACKEY, MERCEDES  LAST HERALD MAGE, THE [LAST HERALD MAGE 1,2,3]  H:2
LACKEY, MERCEDES/EDGHILL, ROSEMARY  MAD MAUDLIN [ERIC BANYON 5]  H:2
LACKEY, MERCEDES  MAGIC’S PAWN [LAST HERALD MAGE 1]  C/P:2,P
LACKEY, MERCEDES  MAGIC’S PRICE [LAST HERALD MAGE 3]  C/P:2,P
LACKEY, MERCEDES  MAGIC’S PROMISE [LAST HERALD MAGE 2]  C/P:2,P
LACKEY, MERCEDES/EDGHILL, ROSEMARY  MUSIC TO MY SORROW [ERIC BANYON 6]  H
LACKEY, MERCEDES  OATHBLOOD [VOWS AND HONOR 3]  C/P:2,P
LACKEY, MERCEDES  OATHBOUND, THE [VOWS AND HONOR 1]  C/P:2,P
LACKEY, MERCEDES  OATHBREAKERS [VOWS AND HONOR 2]  C/P:2,P
LACKEY, MERCEDES  ONE GOOD KNIGHT [FIVE HUNDRED KINGDOMS]  H

LACKEY, MERCEDES/SHEPHERD, MARK/ LISLE, HOLLY OTHERWORLD, THE

P

LACKEY, MERCEDES/MALLORY, JAMES  OUTSTRETCHED SHADOW, THE [OBSIDIAN 1]  H
LACKEY, MERCEDES/DIXON, LARRY  OWLFLIGHT [DARIAN 1]  C/P,P,H
LACKEY, MERCEDES/DIXON, LARRY  OWLKNIGHT [DARIAN 3]  C/P:2,P,H
LACKEY, MERCEDES/DIXON, LARRY  OWLSIGHT [DARIAN 2]  C/P,P,H
LACKEY, MERCEDES  PHOENIX AND ASHES [ELEMENTAL MASTERS 3]  H:2
LACKEY, MERCEDES/SHEPHERD, MARK  PRISON OF SOULS [BARD’S TALE 3]  C/P,SFWA-P,P
LACKEY, MERCEDES  QUEEN’S OWN [HERALDS OF VALDEMAR 1,2,3]  H
LACKEY, MERCEDES  ROBIN AND THE KESTREL, THE [BARDIC VOICES 2]  C/P,@H,P
LACKEY, MERCEDES  SACRED GROUND  C/P,P,H
LACKEY, MERCEDES  SANCTUARY [DRAGON JOUSTERS 3]  H:2
LACKEY, MERCEDES  SERPENT’S SHADOW, THE [ELEMENTAL MASTERS 1]  H

LACKEY, MERCEDES/FLINT, ERIC/ FREER, DAVE SHADOW OF THE LION, THE [SHADOW OF THE LION 1]

H:2,P

LACKEY, MERCEDES/DIXON, LARRY  SILVER GRYPHON, THE [MAGE WARS 3]  C/P,P,H

The MITSFS Pinkdex  September 24, 2007  Page 226
AUTHOR  TITLE  CODE
LACKEY, MERCEDES/EDGHILL, ROSEMARY  SPIRITS WHITE AS LIGHTNING [ERIC BANYON 4]  H
LACKEY, MERCEDES  STORM BREAKING [MAGE STORMS 3]  C/P,P,H:2
LACKEY, MERCEDES  STORM RISING [MAGE STORMS 2]  H:2,SFWA,P
LACKEY, MERCEDES  STORM WARNING [MAGE STORMS 1]  H:2,SFWA-P,P
LACKEY, MERCEDES/GUON, ELLEN  SUMMONED TO TOURNEY [ERIC BANYON 2]  C/P:2,SFWA-P,P
LACKEY, MERCEDES  SUN IN GLORY [VALDEMAR]  PA
LACKEY, MERCEDES  SWORD OF ICE [VALDEMAR]  C/PA,PA
LACKEY, MERCEDES  TAKE A THIEF [VALDEMAR]  H:2

LACKEY, MERCEDES/FLINT, ERIC/ FREER, DAVE THIS ROUGH MAGIC [SHADOW OF THE LION 2]

H,P

LACKEY, MERCEDES/FLINT, ERIC/ FREER, DAVE THIS ROUGH MAGIC (PROOF) [SHADOW OF THE LION 2]

SR-LP

LACKEY, MERCEDES/GELLIS, ROBERTA  THIS SCEPTER’D ISLE [SCEPTER’D ISLE 1]  P
LACKEY, MERCEDES/MALLORY, JAMES  TO LIGHT A CANDLE [OBSIDIAN 2]  H
LACKEY, MERCEDES  WEREHUNTER  C/P:2,P
LACKEY, MERCEDES/SHEPHERD, MARK  WHEELS OF FIRE [SERRATED EDGE 2]  C/P:2,SFWA-P,P
LACKEY, MERCEDES/MALLORY, JAMES  WHEN DARKNESS FALLS [OBSIDIAN 3]  H
LACKEY, MERCEDES/LISLE, HOLLY  WHEN THE BOUGH BREAKS [SERRATED EDGE 3]  C/P:2,SFWA-P,P
LACKEY, MERCEDES/DIXON, LARRY  WHITE GRYPHON, THE [MAGE WARS 2]  C/P,P,H
LACKEY, MERCEDES  WINDS OF CHANGE [MAGE WINDS 2]  C/P,P,H
LACKEY, MERCEDES  WINDS OF FATE [MAGE WINDS 1]  C/P,P,H
LACKEY, MERCEDES  WINDS OF FURY [MAGE WINDS 3]  H:2,SFWA

LACKEY, MERCEDES/LEE, TANITH/ MURPHY, C. E.

WINTER MOON

LP

LACKEY, MERCEDES/FLINT, ERIC/ FREER, DAVE WIZARD OF KARRES, THE [WITCHES OF KARRES 2]

H

LACKEY, MERCEDES  WIZARD OF LONDON, THE [ELEMENTAL MASTERS 4]  H:2
LAFFERTY, R. A.  ANNALS OF KLEPSIS  P
LAFFERTY, R. A.  APOCALYPSES  C/P,P
LAFFERTY, R. A.  ARRIVE AT EASTERWINE  C/P,SFWA,P,H
LAFFERTY, R. A.  DEVIL IS DEAD, THE  C/P:2,P
LAFFERTY, R. A.  DOES ANYONE ELSE HAVE SOMETHING FURTHER TO ADD?  SFWA
LAFFERTY, R. A.  FALL OF ROME, THE  SFWA
LAFFERTY, R. A.  FLAME IS GREEN, THE  H
LAFFERTY, R. A.  FOURTH MANSIONS  C/P:2,P
LAFFERTY, R. A.  NINE HUNDRED GRANDMOTHERS  C/P:2,P
LAFFERTY, R. A.  NOT TO MENTION CAMELS  H
LAFFERTY, R. A.  OKLA HANNALI  P
LAFFERTY, R. A.  PAST MASTER  C/P:2,P
LAFFERTY, R. A.  REEFS OF EARTH, THE  C/P,P
LAFFERTY, R. A.  RINGING CHANGES  C/P,P
LAFFERTY, R. A.  SINDBAD: THE THIRTEENTH VOYAGE  H:2,LP
LAFFERTY, R. A.  SPACE CHANTEY  DH056
LAFFERTY, R. A.  STRANGE DOINGS  H:2,SFWA,P
LAFFERTY, R. A.  THROUGH ELEGANT EYES  SR-H
LAGOWSKI, BARBARA/MUMMA, RICK  TEEN TERMINATORS  P
LAHAYE, TIM/JENKINS, JERRY B.  LEFT BEHIND [@LEFT BEHIND #1]  H:2,LP
LAHAYE, TIM/JENKINS, JERRY B.  NICOLAE [@LEFT BEHIND #3]  LP
LAHAYE, TIM/JENKINS, JERRY B.  SOUL HARVEST [@LEFT BEHIND #4]  LP
LAHAYE, TIM/JENKINS, JERRY B.  TRIBULATION FORCE [@LEFT BEHIND #2]  LP
LAHEY, MICHAEL  QUEST FOR APOLLO  P
LAIDLAW, MARC  DAD’S NUKE  H
LAIDLAW, MARC  NEON LOTUS  C/P,P
LAIDLAW, MARC  THIRD FORCE, THE  LP
LAIN, DOUGLAS  LAST WEEK’S APOCALYPSE  LP
LAIN, T. H.  BLOODY EYE, THE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  CITY OF FIRE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  DEATH RAY, THE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  LIVING DEAD, THE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  OATH OF NERULL [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  PLAGUE OF ICE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  RETURN OF THE DAMNED [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  SAVAGE CAVES, THE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  SUNDERED ARMS, THE [@DUNGEONS & DRAGONS]  P
LAIN, T. H.  TREACHERY’S WAKE [@DUNGEONS & DRAGONS]  P
LAING, ALEXANDER  CADAVER OF GIDEON WYCK, THE  H
LAIRD, CHARLTON  WEBSTER’S NEW WORLD THESAURUS  DICT
LAKE, DAVID J.  FOURTH HEMISPHERE, THE  SR-P
LAKE, DAVID J.  GODS OF XUMA, THE  P
LAKE, DAVID J.  RIGHT HAND OF DEXTRA, THE  C/P,P
LAKE, DAVID J.  RING OF TRUTH, THE  SR-P,C/P
LAKE, DAVID J.  WALKERS ON THE SKY  P
LAKE, DAVID J.  WARLORDS OF XUMA  P
LAKE, DAVID J.  WILDINGS OF WESTRON, THE  C/P,P

AUTHOR  TITLE  CODE
LAKE, JAY  MAINSPRING  H
LAKE, JAY  ROCKET SCIENCE  LP
LAKIN, RITA  DEMON OF THE NIGHT  P
LALLEY, PAUL  COLONY, THE  P
LAMB, HAROLD  MIGHTY MANSLAYER, THE  P
LAMBERT, W. III  ENCORES IN FADE  P
LAMBERT, W. III  MICHAEL: THE MASTER  P
LAMBSHEAD, JOHN  LUCY’S BLADE  H
L’AMOUR, LOUIS  CALIFORNIOS, THE  P
LAMPMAN, EVELYN SIBLEY  SHY STEGOSAURUS OF CRICKET CREEK, THE  C/P,LP,H
LAMPMAN, EVELYN SIBLEY  SHY STEGOSAURUS OF INDIAN SPRINGS, THE  H
LAMPP, JAMES  SANDRA EFFECT, THE  P
LAMPTON, CHRISTOPHER  CROSS OF EMPIRE  SR-P
LAMPTON, CHRISTOPHER  GATEWAY TO LIMBO  H:2,SFWA
LANAGAN, MARGO  BLACK JUICE  H,P
LANAGAN, MARGO  WHITE TIME  H
LANCE, KATHRYN  PANDORA’S CHILDREN [PANDORA 2]  P
LANCE, KATHRYN  PANDORA’S GENES [PANDORA 1]  C/P,P
LANCOUR, GENE  GLOBES OF LLARUM, THE  H,SFWA
LANCOUR, GENE  LERIOS MECCA, THE [DIRSHAN 1]  H,SFWA
LANCOUR, GENE  MAN-EATERS OF CASCALON, THE [DIRSHAN 4]  SFWA
LANCOUR, GENE  SWORD FOR THE EMPIRE [DIRSHAN 3]  H,SFWA
LANCOUR, GENE  WAR MACHINES OF KALINTH, THE [DIRSHAN 2]  H:2,SFWA
LAND, JOHN  DOOMSDAY SPIRAL, THE  P
LAND, JON  VORTEX  P
LANDIS, ARTHUR H.  CAMELOT IN ORBIT [CAMELOT 2]  C/P,P
LANDIS, ARTHUR H.  HOME — TO AVALON [CAMELOT 4]  C/P,P
LANDIS, ARTHUR H.  MAGICK OF CAMELOT, THE [CAMELOT 3]  C/P:2,P
LANDIS, ARTHUR H.  WORLD CALLED CAMELOT, A [CAMELOT 1]  C/P,P
LANDIS, GEOFFREY A.  IMPACT PARAMETER  H:2
LANDIS, GEOFFREY A.  MARS CROSSING  H,P
LANDIS, GEOFFREY A.  MYTHS, LEGENDS, AND TRUE HISTORY  SR-H
LANDON, KRISTIN  HIDDEN WORLDS, THE  P
LANDSBURG, ALAN/LANDSBURG, SALLY  IN SEARCH OF ANCIENT MYSTERIES  [C/RAN-P]
LANDSBURG, ALAN  SECRETS OF THE BERMUDA TRIANGLE  [C/RAN-P]
LANE, AMY  VULNERABLE  LP
LANE, AMY  WOUNDED [LITTLE GODDESS 2]  LP
LANE, ANDY  ALL-CONSUMING FIRE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LANE, ANDY/RICHARDS, JUSTIN  BANQUO LEGACY, THE [DOCTOR WHO EIGHTH DOCTOR]  P
LANE, ANDY/RICHARDS, JUSTIN  DECALOG 3: CONSEQUENCES [DOCTOR WHO]  C/PA,PA
LANE, ANDY/RICHARDS, JUSTIN  DECALOG 4 — RE:GENERATIONS [DOCTOR WHO]  PA
LANE, ANDY  EMPIRE OF GLASS, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
LANE, ANDY/MORTIMORE, JIM  LUCIFER RISING [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LANE, ANDY  ORIGINAL SIN [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LANE, JOHN  MAID OF THURO  H
LANE, JOHN  MAMMALIA  SR-P
LANG, ANDREW  BLUE FAIRY BOOK, THE  SR-HA
LANG, JEFFREY  COHESION [STAR TREK VOYAGER STRING THEORY 1]  P
LANG, JEFFREY  IMMORTAL COIL [STAR TREK TNG]  P
LANG, KING  ASTRO RACE  SR-P
LANG, KING  GYRATOR CONTROL  SR-P
LANG, KING  PROJECTILE WAR  SR-P
LANG, KING  ROCKET INVASION  SR-P
LANG, KING  SPACE LINE  SR-P
LANG, KING  TASK FLIGHT  SR-P
LANG, KING  TERRA!  SR-P
LANG, KING  TRANS-MERCURIAN  SR-P
LANG, SIMON  ALL THE GODS OF EISERNON  C/P:2,P
LANG, SIMON  ELLUVON GIFT, THE  C/P:2,P
LANG, SIMON  HOPESHIP [USS SKIPJACK 3]  P
LANG, SIMON  TIMESLIDE [USS SKIPJACK 2]  P
LANG, SIMON  TRUMPETS OF TAGAN, THE [USS SKIPJACK 1]  P
LANGART, DARREL T.  ANYTHING YOU CAN DO...  H:2
LANGE, JOHN  BINARY  C/P:2,P
LANGE, JOHN  DRUG OF CHOICE  P
LANGE, JOHN  ZERO COOL  C/P,P
LANGFORD, DAVID  LEAKY ESTABLISHMENT, THE  LP
LANGFORD, DAVID  SILENCE OF THE LANGFORD, THE  SR-LP
LANGFORD, DAVID  SPACE EATER, THE  C/P,P
LANGGUTH, A. J.  JESUS CHRISTS  C/P,P
LANGLEY, JOHN PRENTICE  TRAIL BLAZERS OF THE SKIES  H
LANGRISH, KATHERINE  TROLL FELL  H,P
LANGRISH, KATHERINE  TROLL FELL (PROOF)  SR-LP
LANGTON, JANE  MYSTERIOUS CIRCUS, THE (PROOF)  SR-LP

AUTHOR  TITLE  CODE
LANIER, STERLING E.  CURIOUS QUESTS OF BRIGADIER FFELLOWES, THE  SR-H,H
LANIER, STERLING E.  HIERO’S JOURNEY  C/P:2,P
LANIER, STERLING E.  MENACE UNDER MARSWOOD  C/P,P
LANIER, STERLING E.  UNFORSAKEN HIERO, THE  C/P,P,H
LANNING, GEORGE  PEDESTAL, THE  P
LANNING, SERETA  ESCAPE FROM TOMORROW  SR-P
LANSDALE, JOE R.  BESTSELLERS GUARANTEED  P
LANSDALE, JOE R.  BUMPER CROP  H
LANSDALE, JOE R.  DRIVE-IN, THE [@DRIVE-IN #1]  C/P,P
LANSDALE, JOE R.  DRIVE-IN 2, THE [@DRIVE-IN #2]  C/P,P
LANSDALE, JOE R.  NIGHTRUNNERS, THE  SFWA-P
LANSDALE, JOE R.  SHADOWS, KITH AND KIN, THE  SR-H

LANSDALE, JOE R.

STORIES BY MAMA LANSDALE’S YOUNGEST BOY [@AUTHOR’S CHOICE MONTHLY #18]

SR-H

LAPINE, WARREN/PAGEL, STEPHEN  ABSOLUTE MAGNITUDE  HA,LPA
LARBALESTIER, JUSTINE  BATTLE OF THE SEXES IN SCIENCE FICTION, THE  R/REF-H
LARBALESTIER, JUSTINE  DAUGHTERS OF EARTH  LPA
LARKE, GLENDA  AWARE, THE [@ISLES OF GLORY #1]  C/P,P
LARKE, GLENDA  GILFEATHER [@ISLES OF GLORY #2]  P
LARKE, GLENDA  TAINTED, THE [@ISLES OF GLORY #3]  P
LARKIN, DAVID  FANTASTIC KINGDOM, THE  A
LARSEN, J. ANKER  PHILOSOPHER’S STONE, THE  H
LARSON, BOB  OUR TOPIC TODAY IS D&D (RADIO SERMON CASSETTE)  SR-MM
LARSON, GLEN A./THURSTON, ROBERT  APOLLO’S WAR [@BATTLESTAR GALACTICA 1978 #13]  P,SFWA-P
LARSON, GLEN A./THURSTON, ROBERT  BATTLESTAR GALACTICA [BATTLESTAR GALACTICA 1978 #1]  H:2,@P
LARSON, GLEN A./THURSTON, ROBERT  CYLON DEATH MACHINE, THE [@BATTLESTAR GALACTICA 1978 #2]  C/P,P
LARSON, GLEN A./THURSTON, ROBERT  "DIE, CHAMELEON!" [@BATTLESTAR GALACTICA 1978 #12]  P
LARSON, GLEN A./GOULART, RON  EXPERIMENT IN TERRA [@BATTLESTAR GALACTICA 1978 #9]  C/P,P
LARSON, GLEN A./RESNICK, MIKE  GALACTICA DISCOVERS EARTH [@BATTLESTAR GALACTICA 1978 #5]  P
LARSON, GLEN A./GOULART, RON  GREETINGS FROM EARTH [@BATTLESTAR GALACTICA 1978 #8]  P
LARSON, GLEN A./HILL, ROGER  KNIGHT RIDER  P
LARSON, GLEN A./YERMAKOV, NICHOLAS  LIVING LEGEND, THE [@BATTLESTAR GALACTICA 1978 #6]  P
LARSON, GLEN A./GOULART, RON  LONG PATROL, THE [@BATTLESTAR GALACTICA 1978 #10]  P
LARSON, GLEN A./THURSTON, ROBERT  NIGHTMARE MACHINE, THE [@BATTLESTAR GALACTICA 1978 #11]  P
LARSON, GLEN A./THURSTON, ROBERT  SAGA OF A STAR WORLD [@BATTLESTAR GALACTICA 1978]  P
LARSON, GLEN A./THURSTON, ROBERT  SURRENDER THE GALACTICA! [@BATTLESTAR GALACTICA 1978 #14]  P
LARSON, GLEN A./THURSTON, ROBERT  TOMBS OF KOBOL, THE [@BATTLESTAR GALACTICA 1978 #3]  P
LARSON, GLEN A./YERMAKOV, NICHOLAS  WAR OF THE GODS [@BATTLESTAR GALACTICA 1978 #7]  P
LARSON, GLEN A./THURSTON, ROBERT  YOUNG WARRIORS, THE [@BATTLESTAR GALACTICA 1978 #4]  P
LARSON, MAJLISS  PAWNS AND SYMBOLS [STAR TREK]  C/P:2,P
LASKI, MARGHANITA  VICTORIAN CHAISE LONGUE, THE  C/P,P
LASKY, KATHRYN  STAR SPLIT  H
LASSIEUR, A. L.  UNICORN, THE [FIVE RINGS CLAN WAR 2]  P
LASSWITZ, KURD  TWO PLANETS  C/P,P,H
LAST, MARTIN/SEARLES, BAIRD  SCIENCE FICTION QUIZBOOK, THE  [C/RAN-P:2]
LATHAM, PHILIP  FIVE AGAINST VENUS  H:2
LATHAM, PHILIP  MISSING MEN OF SATURN  H:2
LATHAM, WILLIAM  ETERNITY UNBOUND [@SPACE: 1999]  LP
LATHAM, WILLIAM  MARY’S MONSTER  P
LATHAM, WILLIAM  RESURRECTION [@SPACE: 1999]  P
LATNER, ALEXIS GLYNN  HURRICANE MOON  LP
LATTER, SIMON  GLOBAL GLOBULES AFFAIR, THE (BRITISH) [@GIRL FROM UNCLE #1]  SR-P

LATTER, SIMON GOLDEN BOATS OF TARADATA AFFAIR, THE (BRITISH)

[@GIRL FROM UNCLE #3]

SR-P

LAU, MILLER  DARK THANE [@LAST CLANSMAN #2]  P
LAU, MILLER  TALISKER [@LAST CLANSMAN #1]  P
LAUBENTHAL, SANDERS ANNE  EXCALIBUR  C/P,P

LAUGHLIN, CHARLOTTE/ LEVACK, DANIEL J. H.

DE CAMP: AN L. SPRAGUE DE CAMP BIBLIOGRAPHY

R/IND

LAUMER, KEITH  ALIEN MINDS  C/P,P
LAUMER, KEITH  ASSIGNMENT IN NOWHERE [IMPERIUM 3]  C/P:2,P
LAUMER, KEITH  BACK TO THE TIME TRAP  C/P,P
LAUMER, KEITH  BEST OF KEITH LAUMER, THE  C/P:2,P
LAUMER, KEITH  BEYOND THE IMPERIUM [IMPERIUM 2,3]  C/P:2,P
LAUMER, KEITH  BIG SHOW, THE  C/P,P
LAUMER, KEITH  BOLO [BOLOS]  C/P:2,P
LAUMER, KEITH  BREAKING EARTH, THE  P,SFWA-P
LAUMER, KEITH  CATASTROPHE PLANET  C/P:2,P
LAUMER, KEITH  CHRESTOMATHY  C/P:2,P
LAUMER, KEITH  COMPLEAT BOLO, THE [BOLOS]  C/P,SFWA-P,P
LAUMER, KEITH  DAY BEFORE FOREVER, THE/THUNDERHEAD  C/P,P,H
LAUMER, KEITH  DINOSAUR BEACH  H:2,SFWA,P
LAUMER, KEITH  DROWNED QUEEN, THE [AVENGERS 6]  P
LAUMER, KEITH/BROWN, ROSEL GEORGE  EARTHBLOOD  H:2,P

The MITSFS Pinkdex  September 24, 2007  Page 229
AUTHOR  TITLE  CODE
LAUMER, KEITH  EMBASSY [BRION BAYARD 1]  P
LAUMER, KEITH  END AS A HERO  C/P,P
LAUMER, KEITH  ENEMIES FROM BEYOND [INVADERS 2]  P
LAUMER, KEITH  ENVOY TO NEW WORLDS [RETIEF]  C/P:2,DF223,P
LAUMER, KEITH  FAT CHANCE  P
LAUMER, KEITH  GALACTIC DIPLOMAT [RETIEF]  C/P,P
LAUMER, KEITH  GALACTIC ODYSSEY  C/P:2,SFWA-P,P
LAUMER, KEITH  GALAXY BUILDER, THE [LAFAYETTE O’LEARY 4]  C/P:2,P
LAUMER, KEITH  GIORNO PRIMA DELL’ETERNITA, IL  FP
LAUMER, KEITH  GLORY GAME, THE  C/P,P,H,SFWA
LAUMER, KEITH  GOLD BOMB, THE [AVENGERS 7]  C/P,P
LAUMER, KEITH  GREAT TIME MACHINE HOAX, THE  C/P:2,P
LAUMER, KEITH  GREYLORN  C/P:2,P
LAUMER, KEITH  HOUSE IN NOVEMBER, THE  H:2,P
LAUMER, KEITH  INFINITE CAGE, THE  H:2,P
LAUMER, KEITH  INVADERS, THE [INVADERS 1]  C/P:2,P
LAUMER, KEITH  IT’S A MAD, MAD, MAD GALAXY  C/P:2,P
LAUMER, KEITH  JUDSON’S EDEN  C/P:2,P
LAUMER, KEITH  KNIGHT OF DELUSIONS  C/P:2,SFWA-P,P
LAUMER, KEITH  LEGIONS OF SPACE  P
LAUMER, KEITH  LIGHTER SIDE, THE  P
LAUMER, KEITH  LONG TWILIGHT, THE  C/P,P,H
LAUMER, KEITH  MONITORS, THE  C/P:2,P
LAUMER, KEITH  NIGHT OF DELUSIONS  H:2,P
LAUMER, KEITH  NIGHT OF DELUSIONS (COVER FROM GODEY’S THE MAN IN QUESTION)  SR-L
LAUMER, KEITH  NINE BY LAUMER  C/P:2,P
LAUMER, KEITH  ODYSSEY  P
LAUMER, KEITH  ONCE THERE WAS A GIANT  SR-H,H:2,SFWA
LAUMER, KEITH  ONCE THERE WAS A GIANT/NO SHIP BOOTS IN FAIRYLAND  C/P,P
LAUMER, KEITH  OTHER SIDE OF TIME, THE [IMPERIUM 2]  SR-H,C/P:2,P
LAUMER, KEITH  PLAGUE OF DEMONS, A  C/P:2,P
LAUMER, KEITH/DICKSON, GORDON R.  PLANET RUN  C/P:2,P
LAUMER, KEITH/DICKSON, GORDON R.  PLANET RUN AND TWO STORIES  C/P,P
LAUMER, KEITH  RETIEF! [RETIEF]  P
LAUMER, KEITH  RETIEF: AMBASSADOR TO SPACE [RETIEF]  H,P
LAUMER, KEITH  RETIEF AND THE PANGALACTIC PAGEANT OF PULCHRITUDE [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF AND THE RASCALS [RETIEF]  P,SFWA-P
LAUMER, KEITH  RETIEF AND THE WARLORDS [RETIEF]  C/P,P,H
LAUMER, KEITH  RETIEF AT LARGE [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF: DIPLOMAT AT ARMS [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF: EMISSARY TO THE STARS [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF: EMISSARY TO THE STARS (EXPANDED COLLECTION) [RETIEF]  P

LAUMER, KEITH RETIEF: EMISSARY TO THE STARS (MORE EXPANDED COLLECTION)

[RETIEF]

C/P,P

LAUMER, KEITH  RETIEF: ENVOY TO NEW WORLDS [RETIEF]  P
LAUMER, KEITH  RETIEF IN THE RUINS [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF OF THE CDT [RETIEF]  C/P,P,H,SR-H,SFWA
LAUMER, KEITH  RETIEF TO THE RESCUE [RETIEF]  C/P,P,H
LAUMER, KEITH  RETIEF UNBOUND [RETIEF]  C/P:2,P
LAUMER, KEITH  RETIEF’S RANSOM [RETIEF]  C/P,SR-H,P,H
LAUMER, KEITH  RETIEF’S WAR [RETIEF]  C/P:2,P
LAUMER, KEITH  RETURN OF RETIEF, THE [RETIEF]  C/P,P
LAUMER, KEITH  REWARD FOR RETIEF [RETIEF]  C/P:2,SFWA-P,P
LAUMER, KEITH  ROGUE BOLO [BOLOS]  C/P:2,P
LAUMER, KEITH  SHAPE CHANGER, THE [LAFAYETTE O’LEARY 3]  C/P,P,H
LAUMER, KEITH  STAR COLONY  C/P,P,H
LAUMER, KEITH  STAR TREASURE, THE  SR-H,C/P:2,P
LAUMER, KEITH  STARS MUST WAIT, THE [BOLOS]  C/P:2,SFWA-P,P
LAUMER, KEITH  TIME BENDER, THE [LAFAYETTE O’LEARY 1]  C/P:2,P
LAUMER, KEITH  TIME TRAP  SR-H,C/P:2,P
LAUMER, KEITH  TIMETRACKS  C/P:2,P
LAUMER, KEITH  TRACE OF MEMORY, A  C/P:2,SFWA-P,P
LAUMER, KEITH  ULTIMAX MAN, THE  H:2,P
LAUMER, KEITH  UNDEFEATED, THE  C/P:2,P
LAUMER, KEITH  WORLD SHUFFLER, THE [LAFAYETTE O’LEARY 2]  H:2,P
LAUMER, KEITH  WORLDS OF THE IMPERIUM [IMPERIUM 1]  C/P:2,DF127,SFWA-P,P
LAUMER, KEITH  WORLDS OF THE IMPERIUM (EXPANDED) [IMPERIUM 1]  C/P:2,P
LAUMER, KEITH  ZONE YELLOW [IMPERIUM 4]  C/P,P
LAURANCE, ALICE  CASSANDRA RISING  HA,SFWA
LAURENCE, JAMES  FAITH TRIALS, THE, VOL. 1 [BUFFY]  P
LAURENCE, THEODOR  SATAN, SORCERY AND SEX  [C/RAN-H]
LAURENS, MARSHALL  Z EFFECT, THE  C/P:2,P
LAURIA, FRANK  DARK CITY  P
LAURIA, FRANK  DOCTOR ORIENT  P

AUTHOR  TITLE  CODE
LAURIA, FRANK  LADY SATIVA  P
LAURIA, FRANK  PITCH BLACK  P
LAURIA, FRANK  PRIESTESS, THE  P
LAURIA, FRANK  RAGA SIX  P
LAURIA, FRANK  SETH PAPERS, THE  P
LAURIA, VICTORIA  WHAT’S A GHOUL TO DO? [@GHOST HUNTER]  P
LAWHEAD, STEPHEN R.  ARTHUR [@PENDRAGON #3]  P
LAWHEAD, STEPHEN R.  AVALON  P
LAWHEAD, STEPHEN R.  BLACK ROOD, THE [CELTIC CRUSADES 2]  P
LAWHEAD, STEPHEN R.  BYZANTIUM  H,P
LAWHEAD, STEPHEN R.  DREAM THIEF  LP
LAWHEAD, STEPHEN R.  ENDLESS KNOT, THE [SONG OF ALBION #3]  H,@P
LAWHEAD, STEPHEN R.  GRAIL [PENDRAGON 5]  H,P
LAWHEAD, STEPHEN R.  HOOD [KING RAVEN 1]  H
LAWHEAD, STEPHEN R.  IN THE HALL OF THE DRAGON KING [@DRAGON KING #1]  C/P,P
LAWHEAD, STEPHEN R.  IRON LANCE, THE [CELTIC CRUSADES 1]  P
LAWHEAD, STEPHEN R.  MERLIN [@PENDRAGON #2]  P
LAWHEAD, STEPHEN R.  MYSTIC ROSE, THE [@CELTIC CRUSADES 3]  P
LAWHEAD, STEPHEN R.  PARADISE WAR, THE [SONG OF ALBION #1]  H,@P
LAWHEAD, STEPHEN R.  PATRICK SON OF IRELAND  H
LAWHEAD, STEPHEN R.  PENDRAGON [PENDRAGON 4]  C/P,P
LAWHEAD, STEPHEN R.  SEARCH FOR FIERRA, THE [@EMPYRION #1]  LP
LAWHEAD, STEPHEN R.  SIEGE OF DOME, THE [@EMPYRION #2]  LP
LAWHEAD, STEPHEN R.  SILVER HAND, THE [@SONG OF ALBION #2]  P
LAWHEAD, STEPHEN R.  SWORD AND THE FLAME, THE [@DRAGON KING #3]  C/P:2,SFWA-P,P
LAWHEAD, STEPHEN R.  TALIESIN [@PENDRAGON #1]  C/P:2,P
LAWHEAD, STEPHEN R.  WARLORDS OF NIN, THE [@DRAGON KING #2]  C/P,P
LAWRENCE, CONRAD  COUNCIL TO SAVE THE PLANET, THE  LP
LAWRENCE, J. A.  MUDD’S ANGELS [STAR TREK]  C/P:2,P
LAWRENCE, LOUISE  CHILDREN OF THE DUST  H
LAWRENCE, LOUISE  EARTH WITCH, THE  H,P
LAWRENCE, LOUISE  KEEPER OF THE UNIVERSE  H
LAWRENCE, LOUISE  MOONWIND  H
LAWRENCE, LOUISE  POWER OF STARS, THE  H
LAWRENCE, LOUISE  STAR LORD  P
LAWS, ROBIN D.  HONOR OF THE GRAVE [@WARHAMMER ANGELIKA FLEISCHER 1]  P
LAWS, ROBIN D.  LIAR’S PEAK [@WARHAMMER ANGELIKA FLEISCHER 3]  P
LAWS, ROBIN D.  SACRED FLESH [@WARHAMMER ANGELIKA FLEISCHER 2]  P
LAWS, STEPHEN  GHOST TRAIN  SFWA-P
LAWS, STEPHEN  SPECTRE  SFWA-P
LAWSON, ROBERT  MR. WILMER  H
LAWSON, SUSAN  RIDDLE OF THE GRIFFON [@CRIMSON CRYSTAL 1]  P
LAWSON, SYBIL  POSSESSION OF TAMARA, THE  SR-L
LAYMON, RICHARD  CELLAR, THE  P
LAYMON, RICHARD  FLESH  SFWA-P
LAYMON, RICHARD  NIGHT SHOW  SFWA-P
LAYMON, RICHARD  TREAD SOFTLY  SFWA-P
LAZARUS, KEO FELKER  GISMO FROM OUTER SPACE, THE  LP
LAZENBY, NORMAN  TERROR-TRAP!  B
LE FANU, J. SHERIDAN  BEST GHOST STORIES  LPA
LE FANU, J. SHERIDAN  CARMILLA/HAUNTED BARONET, THE  P
LE FANU, J. SHERIDAN  PURCELL PAPERS, THE  SR-H
LE FANU, J. SHERIDAN  ROSE AND THE KEY, THE  LP
LE FANU, J. SHERIDAN  WYLDER’S HAND  LP
LE GUIN, URSULA K.  ALWAYS COMING HOME  C/P:2,SFWA,P
LE GUIN, URSULA K.  ALWAYS COMING HOME (BOXED HARDCOVER AND CASSETTE)  H,SR-MM:3
LE GUIN, URSULA K.  ALWAYS COMING HOME (COMPANION CASSETTE)  SR-MM
LE GUIN, URSULA K.  BEGINNING PLACE, THE  C/P,SFWA,P
LE GUIN, URSULA K.  BIRTHDAY OF THE WORLD, THE  LP
LE GUIN, URSULA K.  BUFFALO GALS  C/P,P
LE GUIN, URSULA K.  CATWINGS  LP
LE GUIN, URSULA K.  CHANGING PLANES  H
LE GUIN, URSULA K.  CITY OF ILLUSIONS [HAINISH 3]  C/P:2,SFWA,P
LE GUIN, URSULA K.  CIUDAD DE ILUSIONES  FP
LE GUIN, URSULA K.  COMPASS ROSE, THE  C/P:2,P,H
LE GUIN, URSULA K.  DANCING AT THE EDGE OF THE WORLD  R/REF-H
LE GUIN, URSULA K.  DISPOSSESSED, THE  C/P,P,H,SFWA
LE GUIN, URSULA K./KIDD, VIRGINIA  EDGES  C/PA,PA
LE GUIN, URSULA K.  EYE OF THE HERON, THE  C/P:2,P
LE GUIN, URSULA K.  FARTHEST SHORE, THE [EARTHSEA 3]  C/P,P,H,SFWA
LE GUIN, URSULA K.  FISHERMAN OF THE INLAND SEA, A  P
LE GUIN, URSULA K.  FISHERMAN OF THE INLAND SEA, A (PRE-PUBLICATION EDITION)  SR-H,H
LE GUIN, URSULA K.  FOUR WAYS TO FORGIVENESS  C/P,P,H
LE GUIN, URSULA K.  FOUR WAYS TO FORGIVENESS (PROOF)  SR-LP

AUTHOR  TITLE  CODE
LE GUIN, URSULA K.  FROM ELFLAND TO POUGHKEEPSIE  R/REF-P
LE GUIN, URSULA K.  GIFTS  P
LE GUIN, URSULA K.  HARD WORDS AND OTHER POEMS  SFWA
LE GUIN, URSULA K./KIDD, VIRGINIA  INTERFACES  HA,PA
LE GUIN, URSULA K.  LANGUAGE OF THE NIGHT, THE  SFWA,R/REF-H
LE GUIN, URSULA K.  LATHE OF HEAVEN, THE  C/P,SFWA,P,H
LE GUIN, URSULA K.  LEFT HAND OF DARKNESS, THE  H:3,P
LE GUIN, URSULA K.  MALAFRENA  H:2,SFWA,P
LE GUIN, URSULA K.  NEBULA AWARD STORIES 11  HA,PA
LE GUIN, URSULA K.  NEW ATLANTIS, THE  SFWA-TD13,TD13
LE GUIN, URSULA K.  NINE LIVES  B
LE GUIN, URSULA K.  ORSINIAN TALES  H:2,P
LE GUIN, URSULA K.  OTHER WIND, THE [EARTHSEA]  LP
LE GUIN, URSULA K.  PLANET OF EXILE [HAINISH 2]  C/P:2,P,SFWA,DG597
LE GUIN, URSULA K.  ROCANNON’S WORLD [HAINISH 1]  C/P:2,SFWA,DG574,P
LE GUIN, URSULA K.  RUEDA DE CIELO, LA  FP
LE GUIN, URSULA K.  SEAROAD  H:2,LP
LE GUIN, URSULA K.  TALES FROM EARTHSEA [EARTHSEA]  H
LE GUIN, URSULA K.  TALES FROM EARTHSEA [EARTHSEA]/OTHER WIND, THE [EARTHSEA]  H
LE GUIN, URSULA K.  TEHANU [EARTHSEA 4]  H:2,P
LE GUIN, URSULA K.  TELLING, THE [HAINISH]  H:2
LE GUIN, URSULA K.  TELLING, THE (PROOF) [HAINISH]  SR-LP
LE GUIN, URSULA K.  THREE HAINISH NOVELS [HAINISH 1,2,3]  H:2
LE GUIN, URSULA K.  THRESHOLD  P
LE GUIN, URSULA K.  TOMBS OF ATUAN, THE [EARTHSEA 2]  C/P,P,H
LE GUIN, URSULA K.  UNLOCKING THE AIR AND OTHER STORIES  LP
LE GUIN, URSULA K.  VERY FAR AWAY FROM ANYWHERE ELSE  P
LE GUIN, URSULA K.  VOICES [WESTERN SHORE 2]  H
LE GUIN, URSULA K.  WIND’S TWELVE QUARTERS, THE  H:2,P,SFWA
LE GUIN, URSULA K.  WIZARD OF EARTHSEA, A [EARTHSEA 1]  C/P,P,H
LE GUIN, URSULA K.  WORD FOR WORLD IS FOREST, THE  C/P:2,P
LE GUIN, URSULA K.  WORLDS OF EXILE AND ILLUSION [HAINISH 1,2,3]  LP
LE MAY, J./LE MAY, D.  CHASSE A L’IMPONDERABLE, LA  FP
LE MAY, J./LE MAY, D.  OENIPS D’ORLON, L  FP
LE PAGE, RAND  "A" MEN  SR-P
LE PAGE, RAND  ASTEROID FORMA  SR-P
LE PAGE, RAND  BEYOND THESE SUNS  SR-P,C/P
LE PAGE, RAND  BLUE ASP  SR-P
LE PAGE, RAND  SATELLITE B. C.  SR-P
LE PAGE, RAND  TIME AND SPACE  SR-P
LE PAGE, RAND  WAR OF ARGOS  SR-P
LE PAGE, RAND  ZERO POINT  SR-P
LE QUEUX, WILLIAM  EYE OF ISTAR, THE  SR-H
LEACOCK, STEPHEN  FRENZIED FICTION  H
LEAR, PETER  GOLDENGIRL  P
LEATH, ROBERT NEAL  OBSIDIAN APE, THE  VLP
LEBBON, TIM  DUSK  LP
LEBEDEV, VLADIMIR/GAGARIN, YURI  SURVIVAL IN SPACE  [C/RAN-P]
LEBLANC, MAURICE  ARSENE LUPIN  H
LEBLANC, MAURICE  SECRET OF SAREK, THE  H
LEBOW, JESS  COLORS OF MAGIC, THE [@MAGIC: THE GATHERING]  PA
LEBOW, JESS  DARKSTEEL EYE, THE [@MAGIC: THE GATHERING MIRRODIN 2]  P
LEBOW, JESS  MASTER OF CHAINS [@FORGOTTEN REALMS FIGHTERS]  P
LEBOW, JESS  MYTHS OF MAGIC ANTHOLOGY, THE [@MAGIC: THE GATHERING]  PA
LEBOW, JESS  SECRETS OF MAGIC, THE [@MAGIC: THE GATHERING]  PA
LEBOW, JESS  WIND OF WAR [FIVE RINGS FOUR WINDS #2]  P
LEE, ADAM  DARK SHORE, THE [@DOMINIONS OF IRTH #1]  C/P,P
LEE, ADAM  OCTOBERLAND [@DOMINIONS OF IRTH #3]  H,P
LEE, ADAM  SHADOW EATER, THE [@DOMINIONS OF IRTH #2]  P
LEE, BRUNO  VERMINTIDE [@WARHAMMER]  P
LEE, CHRISTOPHER/PARRY, MICHEL  FROM THE ARCHIVES OF EVIL #2  PA
LEE, GENTRY  BRIGHT MESSENGERS [RAMA]  C/P,P
LEE, GENTRY  DOUBLE FULL MOON NIGHT [RAMA]  H,P
LEE, GENTRY  TRANQUILITY WARS, THE  H,P
LEE, J. ARDIAN  OUTLAW SWORDS [DYLAN MATHESON 2]  LP
LEE, J. ARDIAN  SON OF THE SWORD [DYLAN MATHESON 1]  LP
LEE, J. ARDIAN  SWORD OF KING JAMES [DYLAN MATHESON 3]  LP
LEE, J. ARDIAN  SWORD OF THE WHITE ROSE  LP
LEE, JOHN/MOORE, B.  MONSTERS AMONG US  [C/RAN-P]
LEE, JOHN  UNICORN DILEMMA, THE [STRAND 2]  P
LEE, JOHN  UNICORN PEACE, THE [STRAND 4]  P
LEE, JOHN  UNICORN QUEST, THE [STRAND 1]  C/P,P
LEE, JOHN  UNICORN SOLUTION, THE [STRAND 3]  C/P:2,SFWA-P,P
LEE, JOHN  UNICORN WAR, THE [STRAND 5]  P

The MITSFS Pinkdex  September 24, 2007  Page 232
AUTHOR  TITLE  CODE
LEE, LISA/MORRIS, TEE  MOREVI [RAFE AND ASKANA]  LP
LEE, RACHEL  SHADOWS OF MYTH  LP
LEE, SAMANTHA  CHILDE ROLANDE  LP

LEE, SHARON/MILLER, STEVE ADVENTURES IN THE LIADEN UNIVERSE 1-5 [LIADEN UNIVERSE 0/

LIADEN ADVENTURES 1,2,3,4,5]

SR-H

LEE, SHARON/MILLER, STEVE  AGENT OF CHANGE [LIADEN UNIVERSE 1]  P
LEE, SHARON/MILLER, STEVE  BALANCE OF TRADE [LIADEN UNIVERSE]  H
LEE, SHARON/MILLER, STEVE  CALAMITY’S CHILD [LIADEN UNIVERSE]  SR-LP
LEE, SHARON/MILLER, STEVE  CARPE DIEM [LIADEN UNIVERSE 3]  C/P,P
LEE, SHARON/MILLER, STEVE  CHANGELING [LIADEN UNIVERSE 0/LIADEN ADVENTURES 6]  SR-LP
LEE, SHARON/MILLER, STEVE  CONFLICT OF HONORS [LIADEN UNIVERSE 2]  C/P,P
LEE, SHARON/MILLER, STEVE  CRYSTAL DRAGON [LIADEN UNIVERSE GREAT MIGRATION 2]  H
LEE, SHARON/MILLER, STEVE  CRYSTAL SOLDIER [LIADEN UNIVERSE GREAT MIGRATION 1]  H
LEE, SHARON/MILLER, STEVE  I DARE [LIADEN UNIVERSE]  LP
LEE, SHARON/MILLER, STEVE  LIADEN UNIVERSE COMPANION: VOLUME I [LIADEN UNIVERSE]  LP
LEE, SHARON/MILLER, STEVE  LOCAL CUSTOM [LIADEN UNIVERSE]  P
LEE, SHARON/MILLER, STEVE  LOOSE CANNON [LIADEN UNIVERSE 0/LIADEN ADVENTURES 7]  SR-LP
LEE, SHARON/MILLER, STEVE  LOW PORT  HA
LEE, SHARON/MILLER, STEVE  MASTER WALK [ADVOCACY 1]  SR-LP
LEE, SHARON/MILLER, STEVE  NAMING OF KINZEL, THE  SR-LP
LEE, SHARON/MILLER, STEVE  PARTNERS IN NECESSITY [LIADEN UNIVERSE 1,2,3]  H
LEE, SHARON/MILLER, STEVE  PILOTS CHOICE [LIADEN UNIVERSE]  H,LP
LEE, SHARON/MILLER, STEVE  PLAN B [LIADEN UNIVERSE 4]  H:2,LP
LEE, SHARON/MILLER, STEVE  QUIET KNIVES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 9]  SR-LP
LEE, SHARON/MILLER, STEVE  QUIET MAGIC  SR-LP
LEE, SHARON/MILLER, STEVE  SCOUT’S PROGRESS [@LIADEN UNIVERSE]  P
LEE, SHARON/MILLER, STEVE  SHADOWS AND SHADES [LIADEN UNIVERSE 0/LIADEN ADVENTURES 8]  SR-LP
LEE, SHARON/MILLER, STEVE  SWORD OF ORION [@BENEATH STRANGE SKIES #1]  LP
LEE, SHARON/MILLER, STEVE  TOMORROW LOG, THE  LP
LEE, SHARON  VARIATIONS THREE:  SR-LP
LEE, STAN/TIMMONS, STAN  ALIEN FACTOR, THE  C/P:2,P
LEE, STAN  AMAZING SPIDER-MAN, THE [MARVEL COMICS]  SCX

LEE, STAN BRING ON THE BAD GUYS: ORIGINS OF THE MARVEL COMICS VILLAINS

[MARVEL COMICS]

CX

LEE, STAN  DOCTOR STRANGE: MASTER OF THE MYSTIC ARTS [MARVEL COMICS]  CX
LEE, STAN  ORIGINS OF MARVEL COMICS [MARVEL COMICS]  CX
LEE, STAN  SON OF ORIGINS OF MARVEL COMICS [MARVEL COMICS]  CX
LEE, STAN/MCCAY, BILL  STAN LEE’S RIFTWORLD ODYSSEY  P
LEE, STAN/DAVID, PETER  ULTIMATE HULK, THE [MARVEL COMICS]  LPA
LEE, STAN  ULTIMATE SILVER SURFER, THE [MARVEL COMICS]  LPA
LEE, STAN  ULTIMATE SPIDER-MAN, THE [MARVEL COMICS]  LPA
LEE, STAN  ULTIMATE SUPER-VILLAINS, THE [MARVEL COMICS]  LPA
LEE, STANLEY R.  DUNN’S CONUNDRUM  C/P,P
LEE, STANLEY R.  GOD PROJECT, THE  H
LEE, TANITH  ANACKIRE [VIS 2]  C/P:2,P
LEE, TANITH  BIRTHGRAVE, THE  C/P:2,P
LEE, TANITH  BITING THE SUN  P
LEE, TANITH  BLACK UNICORN [TANAQUIL 1]  H:2,SFWA,P
LEE, TANITH  BOOK OF THE BEAST, THE [PARADYS 2]  H
LEE, TANITH  BOOK OF THE DAMNED [PARADYS 1]  LP
LEE, TANITH  BOOK OF THE DEAD, THE [PARADYS 3]  H,LP
LEE, TANITH  BOOK OF THE MAD, THE [PARADYS 4]  H,LP
LEE, TANITH  CLAIDI JOURNALS, THE [CLAIDI 1,2,3]  H
LEE, TANITH  COMPANIONS ON THE ROAD  P
LEE, TANITH  COMPANIONS ON THE ROAD/WINTER PLAYERS, THE  H
LEE, TANITH  CYRION  C/P,P
LEE, TANITH  DARK CASTLE, WHITE HORSE  C/P,P
LEE, TANITH  DARK DANCE [BLOOD OPERA 1]  H
LEE, TANITH  DAY BY NIGHT  C/P:2,P
LEE, TANITH  DAYS OF GRASS  P
LEE, TANITH  DEATH’S MASTER [FLAT EARTH #2]  C/P,@P
LEE, TANITH  DELIRIUM’S MISTRESS [@FLAT EARTH #4]  P
LEE, TANITH  DELUSION’S MASTER [@FLAT EARTH #3]  C/P:2,P
LEE, TANITH  DON’T BITE THE SUN  P
LEE, TANITH  DRAGON HOARD, THE  P
LEE, TANITH  DRINKING SAPPHIRE WINE  P
LEE, TANITH  EAST OF MIDNIGHT  C/P,P,H
LEE, TANITH  ELECTRIC FOREST  H:2,P
LEE, TANITH  ELEPHANTASM  P
LEE, TANITH  FACES UNDER WATER [@SECRET BOOKS OF VENUS #1]  H
LEE, TANITH  GOLD UNICORN [TANAQUIL 2]  C/P,P,H
LEE, TANITH  GORGON, THE  P
LEE, TANITH  GORGON AND OTHER BEASTLY TALES, THE  H
LEE, TANITH  HEART BEAST  P

AUTHOR  TITLE  CODE
LEE, TANITH  HEROINE OF THE WORLD, A  C/P,P,H
LEE, TANITH  KILL THE DEAD  C/P,P
LEE, TANITH  LORDS OF DARKNESS, THE [@FLAT EARTH 3,4,5]  H:2
LEE, TANITH  LYCANTHIA, OR, THE CHILDREN OF WOLVES  C/P,P
LEE, TANITH  METALLIC LOVE [SILVER METAL LOVER 2]  P
LEE, TANITH  METALLIC LOVE (PROOF) [SILVER METAL LOVER 2]  SR-P
LEE, TANITH  NIGHT’S DAUGHTER [@FLAT EARTH 1,2]  H:2
LEE, TANITH  NIGHT’S MASTER [@FLAT EARTH #1]  C/P,P
LEE, TANITH  NIGHT’S SORCERIES [@FLAT EARTH #5]  P
LEE, TANITH  NIGHTSHADES  P
LEE, TANITH  PERSONAL DARKNESS [BLOOD OPERA 2]  H
LEE, TANITH  PRINCESS HYNCHATTI AND SOME OTHER SURPRISES  H
LEE, TANITH  QUEST FOR THE WHITE WITCH  P
LEE, TANITH  RED AS BLOOD, OR, TALES FROM THE SISTERS GRIMMER  H:2,P
LEE, TANITH  RED UNICORN [TANAQUIL 3]  H,P
LEE, TANITH  SABELLA, OR, THE BLOOD STONE  C/P,P
LEE, TANITH  SAINT FIRE  H
LEE, TANITH  SECRET BOOKS OF PARADYS I & II, THE [PARADYS 1,2]  H:2
LEE, TANITH  SECRET BOOKS OF PARADYS III & IV, THE [PARADYS 3,4]  H
LEE, TANITH  SECRET BOOKS OF VENUS I & II, THE [VENUS 1,2]  H
LEE, TANITH  SECRET BOOKS OF VENUS III & IV, THE [VENUS 3,4]  H
LEE, TANITH  SILVER METAL LOVER, THE [SILVER METAL LOVER 1]  H:2,P
LEE, TANITH  SOMETIMES, AFTER SUNSET  H
LEE, TANITH  STORM LORD, THE [VIS 1]  C/P,P
LEE, TANITH  SUNG IN SHADOW  C/P:2,P
LEE, TANITH  TAMASTARA  C/P,P
LEE, TANITH  UNSILENT NIGHT  SR-H
LEE, TANITH  VAZKOR, SON OF VAZKOR  C/P,P
LEE, TANITH  VENUS PRESERVED [@VENUS #4]  LP
LEE, TANITH  VOLKHAVAAR  C/P:2,P
LEE, TANITH  WARS OF VIS, THE [VIS 1,2]  H
LEE, TANITH  WHITE AS SNOW  H:2
LEE, TERRENCE  TIME CAMERA  LP
LEE, VERNON  RAVENNA AND HER GHOSTS  SR-P
LEE, WALT  REFERENCE GUIDE TO FANTASTIC FILMS 1 (A–F)  R/IND
LEE, WALT  REFERENCE GUIDE TO FANTASTIC FILMS 2 (G–O)  R/IND
LEE, WALT  REFERENCE GUIDE TO FANTASTIC FILMS 3 (P–Z)  R/IND
LEEDER, MURRAY J. D.  SON OF THUNDER [@FORGOTTEN REALMS FIGHTERS]  P
LEES, DAN  ZODIAC  SFWA
LEFANU, SARAH  IN THE CHINKS OF THE WORLD MACHINE  R/REF-P
LEHMAN, MILTON  THIS HIGH MAN  [C/RAN-P]
LEIBER, FRITZ  BEST OF FRITZ LEIBER, THE  H:2,P
LEIBER, FRITZ  BIG TIME, THE  C/P,DD491,SFWA-P,P,H
LEIBER, FRITZ  BOOK OF FRITZ LEIBER, THE  C/P:2,P
LEIBER, FRITZ  CHANGEWAR  P
LEIBER, FRITZ  CONJURE WIFE  SR-P,C/P:2,TD36
LEIBER, FRITZ  DAY DARK, NIGHT BRIGHT  SR-H
LEIBER, FRITZ  DESTINY TIMES THREE  C/P,GN5.28,BS1
LEIBER, FRITZ  GATHER, DARKNESS!  C/P:2,SFWA,LP
LEIBER, FRITZ  GHOST LIGHT, THE  P,LP
LEIBER, FRITZ  GREEN MILLENNIUM, THE  C/P,H,SR-P,D30300
LEIBER, FRITZ  HEROES AND HORRORS  P
LEIBER, FRITZ  ILL MET IN LANKHMAR [FAFHRD 1,2]  H,SFWA-TD19,TD19
LEIBER, FRITZ  ILL MET IN LANKHMAR (AND OTHER STORIES) [FAFHRD]  P
LEIBER, FRITZ  ILL MET IN LANKHMAR (SHORT STORY) [FAFHRD]  H
LEIBER, FRITZ  KNIGHT AND KNAVE OF SWORDS, THE [FAFHRD]  C/P,SFWA-P:2,P,H
LEIBER, FRITZ  LANKHMAR [FAFHRD]  P
LEIBER, FRITZ  MIND SPIDER AND OTHER STORIES, THE  C/P:2,DD491,P
LEIBER, FRITZ  NIGHT MONSTERS  D30300
LEIBER, FRITZ  NIGHT OF THE WOLF, THE  C/P,P
LEIBER, FRITZ  NIGHT’S BLACK AGENTS (ABRIDGED)  C/P,P
LEIBER, FRITZ  NIGHT’S BLACK AGENTS (REVISED)  C/P:2,P
LEIBER, FRITZ  OUR LADY OF DARKNESS  C/P,TD36,P,H
LEIBER, FRITZ  PAIL OF AIR, A  C/P,P
LEIBER, FRITZ  RIME ISLE  SR-H
LEIBER, FRITZ  SCHWERTER UND EISZAUBER  FP
LEIBER, FRITZ  SECOND BOOK OF FRITZ LEIBER, THE  C/P,P
LEIBER, FRITZ  SECRET SONGS, THE  SR-P
LEIBER, FRITZ  SHADOWS WITH EYES  C/P,P
LEIBER, FRITZ  SHIP OF SHADOWS  TD5
LEIBER, FRITZ  SHIPS TO THE STARS  DF285
LEIBER, FRITZ  SILVER EGGHEADS, THE  C/P:2,P
LEIBER, FRITZ  SINFUL ONES, THE  C/P:2,P
LEIBER, FRITZ  SPECTER IS HAUNTING TEXAS, A  H:2,P,SFWA

AUTHOR  TITLE  CODE
LEIBER, FRITZ  SWORDS AGAINST DEATH [FAFHRD 2]  C/P:2,P
LEIBER, FRITZ  SWORDS AGAINST WIZARDRY [FAFHRD 4]  C/P:2,P
LEIBER, FRITZ  SWORDS AND DEVILTRY [FAFHRD 1]  C/P:2,P
LEIBER, FRITZ  SWORDS AND ICE MAGIC [FAFHRD 6]  C/P:2,P
LEIBER, FRITZ  SWORDS IN THE MIST [FAFHRD 3]  C/P:2,P
LEIBER, FRITZ  SWORDS’ MASTERS [FAFHRD 4,5,6]  H:2
LEIBER, FRITZ  SWORDS OF LANKHMAR, THE [FAFHRD 5]  SR-P,C/P:2
LEIBER, FRITZ  TARZAN AND THE VALLEY OF GOLD [TARZAN]  P
LEIBER, FRITZ  THREE OF SWORDS, THE [FAFHRD 1,2,3]  H
LEIBER, FRITZ  WANDERER, THE  C/P:2,SFWA-P,P
LEIBER, FRITZ  WORLDS OF FRITZ LEIBER, THE  C/P:2,P
LEIBER, FRITZ  YOU’RE ALL ALONE  C/P:2,P
LEIBER, JUSTIN  BEYOND GRAVITY [SALLY FORTH 3]  C/P:2,SFWA-P,P
LEIBER, JUSTIN  BEYOND HUMANITY [SALLY FORTH 2]  C/P:2,P
LEIBER, JUSTIN  BEYOND REJECTION [SALLY FORTH 1]  C/P,P,H
LEIBER, JUSTIN  SWORD AND THE EYE, THE [HOUSE OF EIGIN 1]  C/P,SFWA-P,P
LEIBER, JUSTIN  SWORD AND THE TOWER, THE [HOUSE OF EIGIN 2]  P,SFWA-P
LEIGH, STEPHEN  ABRAXAS MARVEL CIRCUS, THE  C/P:2,P
LEIGH, STEPHEN  ALIEN TONGUE [NEXT WAVE 2]  C/P:2,P
LEIGH, STEPHEN  BONES OF GOD, THE  C/P,P
LEIGH, STEPHEN  CHANGELING [ROBOTS AND ALIENS 1]  C/P:2,P
LEIGH, STEPHEN  CRYSTAL MEMORY, THE  H:2,P
LEIGH, STEPHEN  DANCE OF THE HAG  C/P:2,P
LEIGH, STEPHEN  DARK WATER’S EMBRACE  P
LEIGH, STEPHEN  DINOSAUR CONQUEST [RAY BRADBURY’S DINOSAURS 6]  P
LEIGH, STEPHEN/MILLER, JOHN J.  DINOSAUR EMPIRE [RAY BRADBURY’S DINOSAURS 5]  P
LEIGH, STEPHEN  DINOSAUR PLANET [RAY BRADBURY’S DINOSAURS 2]  P,SFWA-P
LEIGH, STEPHEN/MILLER, JOHN L.  DINOSAUR SAMURAI [RAY BRADBURY’S DINOSAURS 3]  P
LEIGH, STEPHEN  DINOSAUR WARRIORS [RAY BRADBURY’S DINOSAURS 4]  P
LEIGH, STEPHEN  DINOSAUR WORLD [RAY BRADBURY’S DINOSAURS 1]  P
LEIGH, STEPHEN  QUIET OF STONE, A  C/P:2,P
LEIGH, STEPHEN  SECRET OF THE LONA, THE [DR. BONES 1]  C/P,P
LEIGH, STEPHEN  SLOW FALL TO DAWN  C/P:2,P
LEIGH, STEPHEN  SPEAKING STONES  P
LEIMAS, BROOKE  INTRUDER, THE  P
LEIMAS, BROOKE  SUMMER VISITORS, THE  P
LEININGER, ROBERT  BLACK SUN  P
LEINSTER, MURRAY  ALIENS, THE  C/P:2,P
LEINSTER, MURRAY  ASTRONEF PIRATE, L  FP
LEINSTER, MURRAY  BEST OF MURRAY LEINSTER, THE  C/P:2,P
LEINSTER, MURRAY  BLACK GALAXY, THE  C/P,GN4.20
LEINSTER, MURRAY  BRAIN-STEALERS, THE  SR-P,DD079
LEINSTER, MURRAY  CHECKPOINT LAMBDA  C/P,P
LEINSTER, MURRAY  CITY ON THE MOON  DD277,H
LEINSTER, MURRAY  COLONIAL SURVEY  H
LEINSTER, MURRAY  CREATURES OF THE ABYSS  C/P:2,P
LEINSTER, MURRAY  DOCTOR TO THE STARS [MED]  C/P:2,P
LEINSTER, MURRAY  DUPLICATORS, THE  DF275
LEINSTER, MURRAY  FIGHT FOR LIFE  SR-P
LEINSTER, MURRAY  FIRST CONTACTS  H
LEINSTER, MURRAY  FORGOTTEN PLANET, THE  H:2,P,DD146
LEINSTER, MURRAY  FOUR FROM PLANET 5  C/P,P,H
LEINSTER, MURRAY  GATEWAY TO ELSEWHERE  SR-DD053
LEINSTER, MURRAY  GET OFF MY WORLD!  C/P:2,P
LEINSTER, MURRAY  GREKS BRING GIFTS, THE  C/P:2,P
LEINSTER, MURRAY  HOT SPOT, THE [LAND OF THE GIANTS 2]  P
LEINSTER, MURRAY  INVADERS OF SPACE  C/P:2,P
LEINSTER, MURRAY  LAND OF THE GIANTS [LAND OF THE GIANTS 1]  C/P:2,P
LEINSTER, MURRAY  LAST SPACESHIP, THE  GN4.25,H,SR-P
LEINSTER, MURRAY  LONELY PLANET, THE  B
LEINSTER, MURRAY  MED SERIES, THE [MED 1,2,3]  C/P:2,P
LEINSTER, MURRAY  MED SHIP [MED]  C/P,P
LEINSTER, MURRAY  MEN INTO SPACE  C/P:2,P
LEINSTER, MURRAY  MINERS IN THE SKY  C/P:2,P
LEINSTER, MURRAY  MONSTER FROM EARTH’S END, THE  C/P:2,P
LEINSTER, MURRAY  MONSTERS AND SUCH  C/P:2,P
LEINSTER, MURRAY  MUTANT WEAPON, THE [MED 1]  D66525,DD403
LEINSTER, MURRAY  OPERATION: OUTER SPACE  C/P:2,SR-LP,LP
LEINSTER, MURRAY  OPERATION TERROR  C/P:2,P
LEINSTER, MURRAY  OTHER SIDE OF HERE, THE  DD094
LEINSTER, MURRAY  OTHER SIDE OF NOWHERE, THE  C/P:2,P
LEINSTER, MURRAY  PIRATES OF ZAN, THE  D66525,DD403
LEINSTER, MURRAY  PLANET EXPLORER, THE  P
LEINSTER, MURRAY  PLANETS OF ADVENTURE  P

AUTHOR  TITLE  CODE
LEINSTER, MURRAY  QUARANTINE WORLD  P
LEINSTER, MURRAY  SIDEWISE IN TIME  SR-H
LEINSTER, MURRAY  S.O.S. FROM THREE WORLDS [MED 2]  C/P:2,P
LEINSTER, MURRAY  SPACE CAPTAIN  DM135
LEINSTER, MURRAY  SPACE GYPSIES  C/P:2,P
LEINSTER, MURRAY  SPACE PLATFORM  C/P,P,H
LEINSTER, MURRAY  SPACE TUG  C/P:2,P
LEINSTER, MURRAY  TALENTS, INCORPORATED  C/P:2,P
LEINSTER, MURRAY  THIS WORLD IS TABOO [MED 3]  C/P:2,P
LEINSTER, MURRAY  TIME TUNNEL  C/P,P
LEINSTER, MURRAY  TIME TUNNEL, THE [TIME TUNNEL 1]  C/P:2,P
LEINSTER, MURRAY  TIMESLIP! [TIME TUNNEL 2]  C/P:2,P
LEINSTER, MURRAY  TWISTS IN TIME  C/P:2,P
LEINSTER, MURRAY  UNKNOWN, THE  B
LEINSTER, MURRAY  UNKNOWN DANGER [LAND OF THE GIANTS 3]  P
LEINSTER, MURRAY  WAILING ASTEROID, THE  C/P:2,P
LEINSTER, MURRAY  WAR WITH THE GIZMOS  C/P:2,P
LEITH, VALERY  COMPANY OF GLASS, THE [EVERIEN 1]  C/P,LP
LEITH, VALERY  RIDDLED NIGHT, THE [EVERIEN 2]  C/P,LP
LEITH, VALERY  WAY OF THE ROSE, THE [EVERIEN 3]  LP
LEITHAUSER, BRAD  HENCE  H
LEM, STANISLAW  CHAIN OF CHANCE, THE  C/P:2,P,H
LEM, STANISLAW  COSMIC CARNIVAL OF STANISLAW LEM, THE  H,LP
LEM, STANISLAW  CYBERIAD, THE  H:2,SFWA:2,P
LEM, STANISLAW  DZIENNIKI GWIAZDOWE  FP
LEM, STANISLAW  EDEN  FH
LEM, STANISLAW  FIASCO  LP
LEM, STANISLAW  FIEBRE DEL HENO, LA  FP
LEM, STANISLAW  FUTUROLOGICAL CONGRESS, THE  C/P,P,H
LEM, STANISLAW  GLOS PANA  FP
LEM, STANISLAW  HIS MASTER’S VOICE  C/P:2,P
LEM, STANISLAW  IMAGINARY MAGNITUDE  H,P
LEM, STANISLAW  INVESTIGATION, THE  C/P:2,P
LEM, STANISLAW  INVINCIBLE, THE  C/P:3,SFWA,P
LEM, STANISLAW  MEMOIRS FOUND IN A BATHTUB  C/P,SFWA,P,H
LEM, STANISLAW  MEMOIRS OF A SPACE TRAVELER  H:2,P
LEM, STANISLAW  MICROWORLDS  R/REF-H
LEM, STANISLAW  MORE TALES OF PIRX THE PILOT  C/P,P,H
LEM, STANISLAW  MORTAL ENGINES  C/P,P
LEM, STANISLAW  ONE HUMAN MINUTE  H,P
LEM, STANISLAW  OPOWIADANIA WYBRANE  FP
LEM, STANISLAW  PEACE ON EARTH  H
LEM, STANISLAW  PERFECT VACUUM, A  LP
LEM, STANISLAW  RETURN FROM THE STARS  H:2,P
LEM, STANISLAW  SOLARIS  H:2,P
LEM, STANISLAW  STAR DIARIES, THE  C/P:2,P,H
LEM, STANISLAW  TALES OF PIRX THE PILOT  H:2,P
LEM, STANISLAW  WIELKOSC UROJONA  FP
LEMAN, BOB  FEESTERS IN THE LAKE  SR-H
LEMING, R. L.  DAMNATIONS  SR-VLPA
L’ENGLE, MADELEINE  ACCEPTABLE TIME, AN  C/P,P,H
L’ENGLE, MADELEINE  ARM OF THE STARFISH, THE  C/P,P
L’ENGLE, MADELEINE  DRAGONS IN THE WATERS  C/P,P
L’ENGLE, MADELEINE  HOUSE LIKE A LOTUS, A  P
L’ENGLE, MADELEINE  MANY WATERS  C/P,P,H
L’ENGLE, MADELEINE  RING OF ENDLESS LIGHT, A  P
L’ENGLE, MADELEINE  SWIFTLY TILTING PLANET, A [MEG MURRY 3]  C/P,P,H
L’ENGLE, MADELEINE  WIND IN THE DOOR, A [MEG MURRY 2]  C/P,LP,H
L’ENGLE, MADELEINE  WRINKLE IN TIME, A [MEG MURRY 1]  C/P,P,H:2
L’ENGLE, MADELEINE  WRINKLE IN TIME, A (BRITISH) [MEG MURRY 1]  SR-P
L’ENGLE, MADELEINE  YOUNG UNICORNS, THE  P
LENHARD, ELIZABETH  CHARMED AGAIN [@CHARMED]  P
LENHARD, ELIZABETH  DATE WITH DEATH [@CHARMED]  P
LENHARD, ELIZABETH  SOUL OF THE BRIDE [@CHARMED]  P
LEOKUM, LEONARD/POSNICK, PAUL  WEATHER WAR  P
LEON, MARK  GAIA WAR, THE  C/P,P
LEON, MARK  MIND-SURFER  P
LEON, MARK  UNIFIED FIELD, THE  P
LEONARD, GEORGE H.  ALIEN  C/P,P
LEONARD, JONATHAN EDWARD  FLIGHT INTO SPACE  [C/RAN-P:2]
LEONARD, PAUL  DANCING THE CODE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
LEONARD, PAUL/MORTIMORE, JIM  DECALOG 5 — WONDERS [DOCTOR WHO]  PA
LEONARD, PAUL  DREAMSTONE MOON [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
LEONARD, PAUL/WALTERS, NICK  DRY PILGRIMAGE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P

AUTHOR  TITLE  CODE
LEONARD, PAUL  GENOCIDE [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
LEONARD, PAUL  LAST RESORT, THE [DOCTOR WHO EIGHTH DOCTOR]  P
LEONARD, PAUL  REVOLUTION MAN [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
LEONARD, PAUL  SPEED OF FLIGHT [DOCTOR WHO MISSING ADVENTURES]  C/P,P
LEONARD, PAUL  TOY SOLDIERS [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LEONARD, PAUL  TURING TEST, THE [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
LEONARD, PAUL  VENUSIAN LULLABY [DOCTOR WHO MISSING ADVENTURES]  C/P,P
LEONARD, R. CEDRIC  QUEST FOR ATLANTIS  [C/RAN-P]
LEOURIER, CHRISTIAN  MOUNTAINS OF THE SUN, THE  C/P:2,P
LERANGIS, PETER  STAR TREK IV: THE VOYAGE HOME (FOTONOVEL) [STAR TREK]  LP
LERMAN, RHODA  BOOK OF THE NIGHT, THE  P
LERNER, EDWARD M.  MOONSTRUCK  H
LERNER, EDWARD M.  PROBE  C/P:2,P
LERNER, FRED  BOOKMAN’S FANTASY, A  R/REF-P
LERNER, FRED  SILVERLOCK COMPANION, A  R/REF-H
LEROUX, GASTON  FANTOME DE L’OPERA, LE  FP
LEROUX, GASTON  MACHINE TO KILL, THE  H
LEROUX, GASTON  MYSTERY OF THE YELLOW ROOM, THE  H
LEROUX, GASTON  PHANTOM OF THE OPERA, THE  H,P
LESLIE, PETER  AUTUMN ACCELERATOR, THE  P
LESLIE, PETER  DIVING DAMES AFFAIR, THE [@MAN FROM UNCLE #9]  C/P:2,P
LESLIE, PETER  FINGER IN THE SKY AFFAIR, THE [@MAN FROM UNCLE #23]  P
LESLIE, PETER  FINGER IN THE SKY AFFAIR, THE (BRITISH) [@MAN FROM UNCLE #5]  SR-P
LESLIE, PETER  RADIOACTIVE CAMEL AFFAIR, THE [@MAN FROM UNCLE #7]  C/P:2,P
LESLIE, PETER  SPLINTERED SUNGLASSES AFFAIR, THE [@MAN FROM UNCLE #16]  C/P,P
LESLIE, PETER  UNFAIR FARE AFFAIR, THE [@MAN FROM UNCLE #18]  C/P,P
LESSER, MILTON  EARTHBOUND  H:2
LESSER, MILTON  IRRATIONALS, THE  B
LESSER, MILTON  LOOKING FORWARD  HA
LESSER, MILTON  RECRUIT FOR ANDROMEDA  DD358,P
LESSER, MILTON  SECRET OF THE BLACK PLANET  C/P:2,P
LESSER, MILTON  SPACEMEN, GO HOME  H:2
LESSER, MILTON  STADIUM BEYOND THE STARS  H
LESSER, MILTON  STAR SEEKERS, THE  H:2

LESSING, DORIS DOCUMENTS RELATING TO THE SENTIMENTAL AGENTS IN THE VOLYEN

EMPIRE [CANOPUS IN ARGOS 5]

H,LP

LESSING, DORIS MAKING OF THE REPRESENTATIVE FOR PLANET 8, THE

[CANOPUS IN ARGOS 4]

H,LP

LESSING, DORIS MARRIAGES BETWEEN ZONES THREE, FOUR, AND FIVE, THE

[CANOPUS IN ARGOS 2]

H:2,LP

LESSING, DORIS  MEMOIRS OF A SURVIVOR, THE  P
LESSING, DORIS  RE: COLONISED PLANET 5, SHIKASTA [CANOPUS IN ARGOS 1]  H,LP
LESSING, DORIS  SIRIAN EXPERIMENTS, THE [CANOPUS IN ARGOS 3]  H:2,LP
LESTER, COLIN  INTERNATIONAL SCIENCE FICTION YEARBOOK 1979, THE  R/REF-P,C/REF-H
LESTIENNE, VOLDEMAR  FURIOSO  P
LETHEM, JONATHAN  AMNESIA MOON  H
LETHEM, JONATHAN  AS SHE CLIMBED ACROSS THE TABLE  H
LETHEM, JONATHAN  GIRL IN LANDSCAPE  LP
LETHEM, JONATHAN  GUN, WITH OCCASIONAL MUSIC  H:2,LP
LETHEM, JONATHAN  WALL OF THE SKY, THE WALL OF THE EYE, THE  LP
LETTOW, DONNA  ZEALOT [@HIGHLANDER]  P
LETTS, BARRY  DOCTOR WHO AND THE DAEMONS [DOCTOR WHO]  SR-P,C/P:2
LETTS, BARRY  GHOSTS OF N-SPACE, THE [DOCTOR WHO MISSING ADVENTURES]  C/P,P
LETTS, BARRY  ISLAND OF DEATH [DOCTOR WHO PAST DOCTORS]  P
LEVACK, DANIEL J. H.  AMBER DREAMS: A ROGER ZELAZNY BILIOGRAPHY  R/IND
LEVACK, DANIEL J. H./UNDERWOOD, TIM  FANTASMS: A BIBLIOGRAPHY OF THE LITERATURE OF JACK VANCE  R/IND
LEVACK, DANIEL J. H.  PKD: A PHILIP K. DICK BIBLIOGRAPHY  R/IND
LEVEN, JEREMY  CREATOR  C/P,P,H

LEVEN, JEREMY SATAN: HIS PSYCHOTHERAPY AND CURE BY THE UNFORTUNATE DR.

KASSLER, J.S.P.S.

H

LEVENE, MALCOLM  CARDER’S PARADISE  H
LEVENE, REBECCA  BAD TIMING [STRONTIUM DOG]  P
LEVENE, REBECCA/WINSTONE, SIMON  WHERE ANGELS FEAR [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
LEVIE, REX DEAN  INSECT WARRIORS, THE  C/P:2,P
LEVIN, BETTY  SWORD OF CULANN, THE  H
LEVIN, IRA  BOYS FROM BRAZIL, THE  H,SFWA,P
LEVIN, IRA  ROSEMARY’S BABY  C/P,P
LEVIN, IRA  STEPFORD WIVES, THE  P
LEVIN, IRA  THIS PERFECT DAY  C/P,P
LEVINE, BARRY/OKRAND, MARC  POWER KLINGON (CASSETTE) [STAR TREK]  SR-MM
LEVINE, GAIL CARSON  ELLA ENCHANTED  LP
LEVINSON, PAUL  BORROWED TIDES  H
LEVINSON, PAUL  CONSCIOUSNESS PLAGUE, THE [PHIL D’AMATO 2]  H
LEVINSON, PAUL  PIXEL EYE, THE [PHIL D’AMATO 3]  H

AUTHOR  TITLE  CODE
LEVINSON, PAUL  PLOT TO SAVE SOCRATES, THE  H
LEVINSON, PAUL  SILK CODE, THE [PHIL D’AMATO 1]  H:2,P
LEVRERO, MARIO  CIUDAD, LA  FP
LEVY, DAVID  GODS OF FOXCROFT, THE  C/P:2,P
LEVY, EDWARD  BEAST WITHIN, THE  P
LEVY, JOSEPH  GO ASK MALICE [@BUFFY]  H
LEVY, PHIL  ANIMAL FORUM  LP
LEVY, ROGER  DARK HEAVENS  LP
LEVY, ROGER  RECKLESS SLEEP  P
LEWIS, AL  INDEX TO THE SCIENCE FICTION MAGAZINES 1961  SRL/IND
LEWIS, ANTHONY R.  ANNOTATED BIBLIOGRAPHY OF RECURSIVE SCIENCE FICTION, AN  R/IND
LEWIS, ANTHONY R.  CONCORDANCE TO CORDWAINER SMITH (PRELIMINARY EDITION)  R/REF-H
LEWIS, ANTHONY R.  CONCORDANCE TO CORDWAINER SMITH (THIRD EDITION)  R/REF-P
LEWIS, C. S.  BOXEN  H
LEWIS, C. S.  DARK TOWER AND OTHER STORIES, THE  H,LP
LEWIS, C. S.  GREAT DIVORCE, THE  P
LEWIS, C. S.  HORSE AND HIS BOY, THE [NARNIA 5]  C/P:2,P
LEWIS, C. S.  LAST BATTLE, THE [NARNIA 7]  C/P:2,P
LEWIS, C. S.  LION, THE WITCH AND THE WARDROBE, THE [NARNIA 1]  C/P:2,P
LEWIS, C. S.  MAGICIAN’S NEPHEW, THE [NARNIA 6]  C/P:2,P
LEWIS, C. S.  OF OTHER WORLDS  LP
LEWIS, C. S.  OUT OF THE SILENT PLANET [PERELANDRA 1]  C/P,P,H
LEWIS, C. S.  PERELANDRA [PERELANDRA 2]  C/P:2,P,H
LEWIS, C. S.  PILGRIM’S REGRESS, THE  P
LEWIS, C. S.  PRINCE CASPIAN [NARNIA 2]  C/P:2,P
LEWIS, C. S.  PROBLEM OF PAIN, THE  [C/RAN-P]
LEWIS, C. S.  SCREWTAPE LETTERS, THE  C/P:2,P
LEWIS, C. S.  SILVER CHAIR, THE [NARNIA 4]  C/P:2,P
LEWIS, C. S.  THAT HIDEOUS STRENGTH [PERELANDRA 3]  C/P:2,P
LEWIS, C. S.  TILL WE HAVE FACES  LP
LEWIS, C. S.  TORTURED PLANET, THE (PERELANDRA 3 ABRIDGED) [PERELANDRA 3]  C/P:2,P
LEWIS, C. S.  VOYAGE OF THE "DAWN TREADER", THE [NARNIA 3]  C/P:2,P
LEWIS, CHARLES  CAIN FACTOR, THE  H
LEWIS, ELBERT JR.  DAWN OF MAN, THE  LP
LEWIS, IRWIN  DAY NEW YORK TREMBLED, THE  C/P:2,P
LEWIS, IRWIN  DAY THEY INVADED NEW YORK, THE  C/P:2,P
LEWIS, MICK  COMBAT ROCK [DOCTOR WHO PAST DOCTORS]  P
LEWIS, MICK  RAGS [DOCTOR WHO PAST DOCTORS]  P
LEWIS, RICHARD  BLACK HORDE, THE  C/P,P
LEWIS, RICHARD  SPIDERS, THE  P
LEWIS, ROY  EVOLUTION MAN, THE  SR-P
LEWIS, SAM  TALISMAN [@EARTHDAWN]  PA
LEWIS, SINCLAIR  IT CAN’T HAPPEN HERE  C/P,P
LEWIS, TONY  BEST OF ASTOUNDING, THE  LPA
LEWITT, SHARIANN  ANGEL AT APOGEE  C/P:2,SFWA-P,P
LEWITT, SHARIANN  BLIND JUSTICE  P,SFWA-P:2
LEWITT, SHARIANN  CYBERNETIC JUNGLE  C/P,P
LEWITT, SHARIANN  CYBERSONG [STAR TREK VOYAGER]  P
LEWITT, SHARIANN  CYBERSTEALTH [CARGO AND GHOSTER 1]  P
LEWITT, SHARIANN  DANCING VAC [CARGO AND GHOSTER 2]  P
LEWITT, SHARIANN  FIRST AND FINAL RITES  C/P:2,P
LEWITT, SHARIANN  INTERFACE MASQUE  H
LEWITT, SHARIANN  MEMENTO MORI  LP
LEWITT, SHARIANN  REBEL SUTRA  H,P
LEWITT, SHARIANN  SONGS OF CHAOS  C/P,SFWA-P:2,P
LEY, SANDRA  BEYOND TIME  C/PA,PA
LEY, WILLY/BONESTELL, CHESLEY  CONQUEST OF SPACE, THE  R/REF-H
LEY, WILLY  EVENTS IN SPACE  [C/RAN-P:2]
LEY, WILLY/VON BRAUN, WERNHER  EXPLORATION OF MARS, THE  R/REF-H,C/REF-H
LEY, WILLY  MAN-MADE SATELLITES  [C/RAN-H]
LEY, WILLY  MISSILES, MOON PROBES, AND MEGAPARSECS  [C/RAN-P]
LEY, WILLY  RANGER TO THE MOON  [C/RAN-P]
LEY, WILLY  ROCKETS: THE FUTURE OF TRAVEL BEYOND THE STRATOSPHERE  R/REF-H
LEY, WILLY  SATELLITES, ROCKETS AND OUTER SPACE  [C/RAN-P]
LEYNER, MARK  ET TU, BABE  SFWA
LICATA, TONY  GREAT SCIENCE-FICTION  C/PA,PA
LICHTENBERG, JACQUELINE  CITY OF A MILLION LEGENDS [KREN 2]  C/P:2,P
LICHTENBERG, JACQUELINE  DREAMSPY  H
LICHTENBERG, JACQUELINE  DUSHAU [DUSHAU 1]  C/P,P
LICHTENBERG, JACQUELINE  FARFETCH [DUSHAU 2]  C/P,P
LICHTENBERG, JACQUELINE  HOUSE OF ZEOR [SIME/GEN 1]  C/P,P,H,SFWA
LICHTENBERG, JACQUELINE  MAHOGANY TRINROSE [SIME/GEN 4]  C/P:2,SFWA,P
LICHTENBERG, JACQUELINE  MOLT BROTHER [KREN 1]  C/P,P
LICHTENBERG, JACQUELINE  OUTREACH [DUSHAU 3]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 238
AUTHOR  TITLE  CODE
LICHTENBERG, JACQUELINE  RENSIME [SIME/GEN 6]  C/P,SFWA,P,H

LICHTENBERG, JACQUELINE/ MARSHAK, SONDRA/WINSTON, JOAN STAR TREK LIVES!

R/REF-P,C/REF-P:2

LICHTENBERG, JACQUELINE  UNTO ZEOR, FOREVER [SIME/GEN 2]  C/P,P,H,SFWA
LICHTENBERG, JACQUELINE/LORRAH, JEAN ZELEROD’S DOOM [SIME/GEN 8]  C/P:2,P
LICKISS, REBECCA  ECCENTRIC CIRCLES  P
LICKISS, REBECCA  NEVER AFTER  P
LIEBERMAN, ROBERT  BABY  H
LIEBERMAN, ROBERT  PERFECT PEOPLE  P
LIEBSCHER, WALT  ALIEN CARNIVAL  SR-P
LIEF, EVELYN  CLONE REBELLION, THE  P
LIEF, EVELYN  CLONE REBELLION, THE (PROOF)  SR-LP
LIGHTNER, A. M.  DAY OF THE DRONES, THE  C/P:2,P
LIGHTNER, A. M.  GALACTIC TROUBADOURS, THE  H
LIGHTNER, A. M.  GODS OR DEMONS?  H
LIGHTNER, A. M.  SPACE OLYMPICS, THE  P
LIGHTNER, A. M.  STAR CIRCUS  H
LIGHTNER, LEE  SONS OF FENRIS [@WARHAMMER 40K]  P
LILLEY, ERNEST  FUTURE WASHINGTON  LPA
LIMAT, MAURICE  DIEU COULEUR DE NUIT, LE  FP
LIMAT, MAURICE  METRO POUR L’INCONNU  FP
LIMAT, MAURICE  OISEAUX DE VEGA, LES  FP
LIMAT, MAURICE  PLANETOIDE 13  FP
LIMAT, MAURICE  RIEN QU’UNE ETOILE  FP
LIMAT, MAURICE  SOLEIL DE GLACE, LE  FP
LIMAT, MAURICE  TERRE N’EST PAS RONDE, LA  FP
LINAKER, MICHAEL R.  SCORPION  P
LINAWEAVER, BRAD/KRAMER, EDWARD E.  FREE SPACE  LPA
LINAWEAVER, BRAD  MOON OF ICE  H
LINAWEAVER, BRAD  SLIDERS  P
LINDBLAD, JOHN  ZAP DAY  P
LINDGREN, ASTRID  RONIA, THE ROBBER’S DAUGHTER  LP
LINDHOLM, MEGAN  ALIEN EARTH  C/P,P
LINDHOLM, MEGAN  CLOVEN HOOVES  C/P,P
LINDHOLM, MEGAN  HARPY’S FLIGHT [KI AND VANDIEN 1]  C/P,P
LINDHOLM, MEGAN  LIMBRETH GATE, THE [KI AND VANDIEN 3]  C/P,P
LINDHOLM, MEGAN  LUCK OF THE WHEELS [KI AND VANDIEN 4]  C/P,P
LINDHOLM, MEGAN  REINDEER PEOPLE, THE  C/P,SFWA-P,P
LINDHOLM, MEGAN  SAGA OF THE REINDEER PEOPLE, A  H
LINDHOLM, MEGAN  WINDSINGERS, THE [KI AND VANDIEN 2]  C/P,P
LINDHOLM, MEGAN  WIZARD OF THE PIGEONS  C/P,P
LINDHOLM, MEGAN  WOLF’S BROTHER  C/P,P
LINDNER, ROBERT  FIFTY-MINUTE HOUR, THE  [C/RAN-P]
LINDSAY, DAVID  HAUNTED WOMAN, THE  LP
LINDSAY, DAVID  VIOLET APPLE, THE/WITCH, THE  LP
LINDSAY, DAVID  VOYAGE TO ARCTURUS, A  C/P,P,H
LINDSEY, HAL  LATE GREAT PLANET EARTH, THE  [C/RAN-P:2]
LINDSEY, HAL  PROMISE, THE  [C/RAN-P]
LINDSEY, HAL  THERE’S A NEW WORLD COMING  [C/RAN-P:2]
LINDSEY, JOHANNA  WARRIOR’S WOMAN  P
LINDSKOLD, JANE  BROTHER TO DRAGONS, COMPANION TO OWLS  P
LINDSKOLD, JANE  BURIED PYRAMID, THE  P
LINDSKOLD, JANE  CHANGER [ATHANOR 1]  P
LINDSKOLD, JANE  CHILD OF A RAINLESS YEAR  P
LINDSKOLD, JANE  CHRONOMASTER  P
LINDSKOLD, JANE  DRAGON OF DESPAIR, THE [FIREKEEPER 3]  P
LINDSKOLD, JANE  LEGENDS WALKING [ATHANOR 2]  P
LINDSKOLD, JANE  MARKS OF OUR BROTHERS  P
LINDSKOLD, JANE  PIPES OF ORPHEUS, THE  P
LINDSKOLD, JANE  SMOKE AND MIRRORS  P
LINDSKOLD, JANE  THROUGH WOLF’S EYES [FIREKEEPER 1]  P
LINDSKOLD, JANE  WHEN THE GODS ARE SILENT  P
LINDSKOLD, JANE  WOLF CAPTURED [FIREKEEPER 4]  H
LINDSKOLD, JANE  WOLF HUNTING [FIREKEEPER 5]  H
LINDSKOLD, JANE  WOLF’S BLOOD [FIREKEEPER 6]  H
LINDSKOLD, JANE  WOLF’S HEAD, WOLF’S HEART [FIREKEEPER 2]  P
LINDSKOOG, KATHRYN  LION OF JUDAH IN NEVER-NEVER LAND, THE  R/REF-P
LINEHAN, L.  BOHACK: SYMBIOTIC WORLDS  P
LINK, KELLY  MAGIC FOR BEGINNERS  LP
LINK, KELLY  STRANGER THINGS HAPPEN  LP
LINK, KELLY  TRAMPOLINE  LPA
LINKLATER, ERIC  SPELL FOR OLD BONES, A  H
LIONEL, ROBERT  TIME ECHO  C/P:2,P
LIQUORI, SAL  SAL’S BOOK [@PSI PATROL 1]  P

The MITSFS Pinkdex  September 24, 2007  Page 239
AUTHOR  TITLE  CODE
LISLE, HOLLY  BONES OF THE PAST [FAIA RISSEDOTTE 2]  C/P:2,SFWA-P,P
LISLE, HOLLY  COURAGE OF FALCONS [SECRET TEXTS 3]  H,LP
LISLE, HOLLY  CURSE OF THE BLACK HERON [BARD’S TALE 8]  C/P:2,P
LISLE, HOLLY  DEVIL & DAN COOLEY, THE [SYMPATHY FOR THE DEVIL 2]  C/P,P
LISLE, HOLLY  DIPLOMACY OF WOLVES [SECRET TEXTS 1]  C/P:2,LP
LISLE, HOLLY  FIRE IN THE MIST [FAIA RISSEDOTTE 1]  C/P:2,SFWA-P,P
LISLE, HOLLY  GODS OLD AND DARK [WORLD GATES 3]  C/P,P
LISLE, HOLLY/NOLAN, TED  HELL ON HIGH [SYMPATHY FOR THE DEVIL 3]  P
LISLE, HOLLY  HUNTING THE CORRIGAN’S BLOOD  P
LISLE, HOLLY/GUIN, CHRIS  MALL, MAYHEM AND MAGIC  C/P:2,SFWA-P,P
LISLE, HOLLY  MEMORY OF FIRE [WORLD GATES 1]  C/P,P
LISLE, HOLLY  MIND OF THE MAGIC [FAIA RISSEDOTTE 3]  C/P:2,P
LISLE, HOLLY  MINERVA WAKES  C/P:2,SFWA-P,P
LISLE, HOLLY  SYMPATHY FOR THE DEVIL [SYMPATHY FOR THE DEVIL 1]  C/P,P
LISLE, HOLLY  TALYN [KORRE]  H,P
LISLE, HOLLY/ALLSTON, AARON  THUNDER OF THE CAPTAINS [BARD’S TALE 6]  P
LISLE, HOLLY  VENGEANCE OF DRAGONS [SECRET TEXTS 2]  C/P,P
LISLE, HOLLY  VINCALIS THE AGITATOR  C/P,LP
LISLE, HOLLY/ALLSTON, AARON  WRATH OF THE PRINCES [BARD’S TALE 7]  P
LISLE, HOLLY  WRECK OF HEAVEN, THE [WORLD GATES 2]  P
LISTER, MARK  TREKTOONS [STAR TREK]  SCX

LITTELL, J./HESTON, L./HAUPT, H. B./ SOLOTAROFF, S.

FLIGHTS OF FANTASY

HA

LITTELL, JONATHAN  BAD VOLTAGE  P
LITTLE, DENISE  ALIEN PETS  PA
LITTLE, DENISE  CONSTELLATION OF CATS, A  PA
LITTLE, DENISE  COSMIC COCKTAILS  PA
LITTLE, DENISE  CREATURE FANTASTIC  PA
LITTLE, DENISE  DANGEROUS MAGIC, A  PA
LITTLE, DENISE  FAMILIARS  PA
LITTLE, DENISE  HAGS, SIRENS, AND OTHER BAD GIRLS OF FANTASY  PA
LITTLE, DENISE  MAGIC SHOP, THE  PA
LITTLE, DENISE  MAGIC TOYBOX, THE  PA
LITTLE, DENISE  PERCHANCE TO DREAM  PA
LITTLE, DENISE  ROTTEN RELATIONS  PA
LITTLE, DENISE  SORCERER’S ACADEMY, THE  PA
LITTLE, DENISE  TIME AFTER TIME  PA
LITTLE, DENISE  TWICE UPON A TIME  C/PA,PA
LITTLE, DENISE  VENGEANCE FANTASTIC  PA
LITTLE, JANE  PHILOSOPHER’S STONE, THE  H
LIVELY, PENELOPE  WILD HUNT OF THE GHOST HOUNDS, THE  P
LIVINGSTON, BERKELEY  MOVING FINGER, THE  B
LIVINGSTON, HAROLD  CLIMACTICON, THE  C/P:2,P
LIVINGSTONE, IAN  DICING WITH DRAGONS  [C/RAN-P]
LJOKA, DAN  SHELTER  P
LLEWELLYN, ALUN  STRANGE INVADERS, THE  SR-P
LLEWELLYN, EDWARD  BRIGHT COMPANION, THE  P
LLEWELLYN, EDWARD  DOUGLAS CONVOLUTION, THE  P
LLEWELLYN, EDWARD  FUGITIVE IN TRANSIT  C/P,P
LLEWELLYN, EDWARD  PRELUDE TO CHAOS  C/P:2,P
LLEWELLYN, EDWARD  SALVAGE AND DESTROY  C/P:2,P
LLEWELLYN, EDWARD  WORD-BRINGER  P
LLEWELLYN, ROBERT  MAN IN THE RUBBER MASK, THE [RED DWARF]  R/REF-P
LLOYD, JOHN URI  ETIDORHPA  SR-H,P
LLYWELYN, MORGAN  BARD  SFWA-P
LLYWELYN, MORGAN  DRUIDS  P
LLYWELYN, MORGAN  ELEMENTALS, THE  H,P
LLYWELYN, MORGAN/SCOTT, MICHAEL  ETRUSCANS  H,P
LLYWELYN, MORGAN  FINN MAC COOL  H,P
LLYWELYN, MORGAN  ISLES OF THE BLEST, THE  C/P,SFWA-P,P
LLYWELYN, MORGAN  RED BRANCH  P
LLYWELYN, MORGAN/SCOTT, MICHAEL  SILVERHAND [@ARCANA #1]  P
LLYWELYN, MORGAN/SCOTT, MICHAEL  SILVERLIGHT [@ARCANA #2]  P
LOBDELL, JARED  TOLKIEN COMPASS, A  R/REF-P
LOBEL, BRANA  REVENANT, THE  P
LOCK, MASTER  LOCK  WRD
LOCKE, GEORGE  FERRET FANTASY’S CHRISTMAS ANNUAL FOR 1973  R/IND
LOCKE, GEORGE  SCIENCE FICTION FIRST EDITIONS  R/IND
LOCKE, GEORGE  VOYAGES IN SPACE  R/IND
LOCKE, GEORGE  WORLDS APART  HA
LOCKE, JOSEPH  BEN THERE, DONE THAT [@SABRINA THE TEENAGE WITCH #6]  P
LOCKSLEY, REBECCA  THREE SISTERS, THE  P

LOFFICIER, JEAN MARC DOCTOR WHO PROGRAMME GUIDE 1: THE PROGRAMMES, THE

[DOCTOR WHO] R/REF-P,C/REF-P:2

AUTHOR  TITLE  CODE

LOFFICIER, JEAN MARC DOCTOR WHO PROGRAMME GUIDE 2:

WHAT’S WHAT AND WHO’S WHO, THE [DOCTOR WHO]

R/REF-P,C/REF-P:2

LOFFICIER, JEAN MARC  DOCTOR WHO PROGRAMME GUIDE, THE (2ND EDITION) [DOCTOR WHO]  R/REF-P,C/REF-P
LOFFICIER, JEAN MARC  NTH DOCTOR, THE [DOCTOR WHO]  R/REF-P
LOFFICIER, JEAN MARC  TERRESTRIAL INDEX, THE [DOCTOR WHO]  R/REF-P
LOFFICIER, JEAN MARC  UNIVERSAL DATABANK, THE [DOCTOR WHO]  R/REF-P
LOFTING, HUGH  DOCTOR DOLITTLE’S CIRCUS [DOCTOR DOLITTLE]  H,P
LOFTING, HUGH  DOCTOR DOLITTLE’S GARDEN [DOCTOR DOLITTLE]  H
LOFTING, HUGH  DOCTOR DOLITTLE’S POST OFFICE [DOCTOR DOLITTLE]  H,P
LOFTING, HUGH  DOCTOR DOLITTLE’S RETURN [DOCTOR DOLITTLE]  H
LOFTING, HUGH  DOCTOR DOLITTLE’S ZOO [DOCTOR DOLITTLE]  P
LOFTING, HUGH  TWILIGHT OF MAGIC, THE  H
LOFTING, HUGH  VOYAGES OF DOCTOR DOLITTLE, THE [DOCTOR DOLITTLE]  C/P,P,H
LOGAN, CHARLES  SHIPWRECK  P
LOGAN, LES  GAME, THE [DARK FORCES 1]  P
LOGAN, LES  UNNATURAL TALENT [DARK FORCES 7]  P

LOGAN, SIMON DECADENT RETURN OF THE HI-FI QUEEN AND HER EMBRYONIC REPTILE

INFECTION, THE

ERD3

LOGSDON, SYD  FOND FAREWELL TO DYING, A  P
LOGSDON, SYD  JANDRAX  P
LOGSTON, ANNE  DAGGER’S EDGE [SHADOW 4]  P
LOGSTON, ANNE  DAGGER’S POINT [SHADOW 5]  P
LOGSTON, ANNE  EXILE  P
LOGSTON, ANNE  FIREWALK  P
LOGSTON, ANNE  GREENDAUGHTER  P,SFWA-P:2
LOGSTON, ANNE  GUARDIAN’S KEY  P
LOGSTON, ANNE  SHADOW [SHADOW 1]  C/P:2,SFWA-P:2,P
LOGSTON, ANNE  SHADOW DANCE [SHADOW 3]  P
LOGSTON, ANNE  SHADOW HUNT [SHADOW 2]  P
LOGSTON, ANNE  WATERDANCE  P
LOGSTON, ANNE  WILD BLOOD  C/P,P
LONARDO, PAUL  APOSTATE, THE  LP
LONDON, JACK  BEFORE ADAM  H:2,P
LONDON, JACK  IRON HEEL, THE  H,LP
LONDON, JACK  MOON FACE  H:2
LONDON, JACK  SELECTED SCIENCE FICTION AND FANTASY STORIES  H
LONDON, JACK  STAR ROVER, THE  H:2
LONDON, JACK  THIRTEEN TALES OF TERROR  P
LONDON, JACK  WHEN GOD LAUGHS  H
LONG, DUNCAN  ANTI-GRAV UNLIMITED  C/P,P
LONG, FRANK BELKNAP  ...AND OTHERS SHALL BE BORN  SR-DB0809
LONG, FRANK BELKNAP  DARK BEASTS, THE  C/P:2,P
LONG, FRANK BELKNAP  EARLY LONG, THE  H:2,SFWA
LONG, FRANK BELKNAP  HOUNDS OF TINDALOS, THE  C/P,P,H
LONG, FRANK BELKNAP  HOUNDS OF TINDALOS: THE EARLY LONG, THE  P
LONG, FRANK BELKNAP  HOWARD PHILLIPS LOVECRAFT: DREAMER ON THE NIGHTSIDE  R/REF-H
LONG, FRANK BELKNAP  IN MAYAN SPLENDOR  SR-H
LONG, FRANK BELKNAP  IT WAS THE DAY OF THE ROBOT  C/P:2,P
LONG, FRANK BELKNAP  JOHN CARSTAIRS: SPACE DETECTIVE  SR-P,H
LONG, FRANK BELKNAP  JOURNEY INTO DARKNESS  C/P:2,P
LONG, FRANK BELKNAP  LEST EARTH BE CONQUERED  C/P:2,P
LONG, FRANK BELKNAP  MARS IS MY DESTINATION  C/P:2,P
LONG, FRANK BELKNAP  MARTIAN VISITORS, THE  H
LONG, FRANK BELKNAP  MONSTER FROM OUT OF TIME  P
LONG, FRANK BELKNAP  NIGHT FEAR  C/P,P
LONG, FRANK BELKNAP  NIGHT OF THE WOLF, THE  P
LONG, FRANK BELKNAP  ODD SCIENCE FICTION  C/P:2,P
LONG, FRANK BELKNAP  RIM OF THE UNKNOWN, THE  SR-H,P
LONG, FRANK BELKNAP  SO DARK A HERITAGE  P
LONG, FRANK BELKNAP  SPACE STATION #1  DD242,P,C/P
LONG, FRANK BELKNAP  SURVIVAL WORLD  C/P,P
LONG, FRANK BELKNAP  THIS STRANGE TOMORROW  C/P:2,P
LONG, FRANK BELKNAP  THREE FACES OF TIME, THE  P
LONG, FRANK BELKNAP  WOMAN FROM ANOTHER PLANET  SR-L
LONG, JAMES D.  D.R.T. [@BATTLETECH 14]  P
LONG, JEFF  DESCENT, THE  P
LONG, JEFF  YEAR ZERO  P
LONG, JIM  MAIN EVENT [@BATTLETECH 10]  C/P,P
LONG, LYDA BELKNAP  FIRE OF THE WITCHES  P
LONG, LYDA BELKNAP  HOUSE OF THE DEADLY NIGHTSHADE  P
LONG, LYDA BELKNAP  LEMOYNE HERITAGE, THE  P
LONG, LYDA BELKNAP  TO THE DARK TOWER  P
LONG, NATHAN  BROKEN LANCE, THE [@WARHAMMER BLACK HEARTS 2]  P
LONG, NATHAN  VALNIR’S BANE [@WARHAMMER BLACK HEARTS 1]  P

AUTHOR  TITLE  CODE
LONGO, CHRIS  LAST GENE, THE  P
LONGYEAR, BARRY B.  CHANGE, THE [@ALIEN NATION #4]  C/P:2,P
LONGYEAR, BARRY B.  CIRCUS WORLD [CIRCUS WORLD 3]  C/P,P,H
LONGYEAR, BARRY B.  CITY OF BARABOO [CIRCUS WORLD 1]  C/P,SFWA,P,H
LONGYEAR, BARRY B.  ELEPHANT SONG [CIRCUS WORLD 2]  C/P:2,P
LONGYEAR, BARRY B./GERROLD, DAVID  ENEMY MINE  TD6,SFWA-TD6,P
LONGYEAR, BARRY B.  ENEMY PAPERS, THE  H
LONGYEAR, BARRY B.  GOD BOX, THE  C/P:2,P
LONGYEAR, BARRY B.  HOMECOMING, THE  H,SFWA
LONGYEAR, BARRY B.  INFINITY HOLD  C/P:2,P
LONGYEAR, BARRY B.  IT CAME FROM SCHENECTADY  H:2,SFWA,P
LONGYEAR, BARRY B.  LAST ENEMY, THE (DRAFT? MANUSCRIPT)  SR-L
LONGYEAR, BARRY B.  MANIFEST DESTINY  C/P:2,P
LONGYEAR, BARRY B.  NAKED CAME THE ROBOT  C/P:2,P
LONGYEAR, BARRY B.  SCIENCE FICTION WRITER’S WORKSHOP-1  R/REF-P
LONGYEAR, BARRY B.  SEA OF GLASS  H:2,SFWA,P
LONGYEAR, BARRY B.  SLAG LIKE ME [@ALIEN NATION #5]  C/P,P
LONGYEAR, BARRY B.  TOMORROW TESTAMENT, THE  C/P,P
LOOMIS, NOEL  CITY OF GLASS  SR-P
LORD, GLENN  LAST CELT: A BIO-BIBLIOGRAPHY OF ROBERT ERVIN HOWARD, THE  R/IND
LORD, GRAHAM  TIME OUT OF MIND  P
LORD, JEFFREY  BRONZE AXE, THE [@BLADE #1]  C/P:2,P
LORD, JEFFREY  CHAMPION OF THE GODS [@BLADE #21]  P
LORD, JEFFREY  CITY OF THE LIVING DEAD [@BLADE #26]  P
LORD, JEFFREY  CRYSTAL SEAS, THE [@BLADE #16]  P
LORD, JEFFREY  DIMENSION OF DREAMS [@BLADE #11]  C/P,P
LORD, JEFFREY  DIMENSION OF HORROR [@BLADE #30]  P
LORD, JEFFREY  DRAGONS OF ENGLOR, THE [@BLADE #24]  P
LORD, JEFFREY  EMPIRE OF BLOOD [@BLADE #23]  P
LORD, JEFFREY  FORESTS OF GLEOR, THE [@BLADE #22]  P
LORD, JEFFREY  GLADIATORS OF HAPANU [@BLADE #31]  P
LORD, JEFFREY  GOLDEN STEED, THE [@BLADE #13]  P
LORD, JEFFREY  GUARDIANS OF THE CORAL THRONE [@BLADE #20]  C/P,P
LORD, JEFFREY  ICE DRAGON [@BLADE #10]  C/P,P
LORD, JEFFREY  JADE WARRIOR, THE [@BLADE #2]  C/P,P
LORD, JEFFREY  JEWEL OF THARN [@BLADE #3]  C/P:2,P
LORD, JEFFREY  KILLER PLANTS OF BINAARK [@BLADE #33]  P
LORD, JEFFREY  KING OF ZUNGA [@BLADE #12]  P
LORD, JEFFREY  KINGDOM OF ROYTH [@BLADE #9]  P
LORD, JEFFREY  LIBERATOR OF JEDD [@BLADE #5]  C/P:2,P
LORD, JEFFREY  LOOTERS OF THARN [@BLADE #19]  P
LORD, JEFFREY  LORDS OF THE CRIMSON RIVER, THE [@BLADE #35]  P
LORD, JEFFREY  MASTER OF THE HASHOMI [@BLADE #27]  P
LORD, JEFFREY  MONSTER OF THE MAZE [@BLADE #6]  C/P,P
LORD, JEFFREY  MOUNTAINS OF BREGA, THE [@BLADE #17]  P
LORD, JEFFREY  PEARL OF PATMOS [@BLADE #7]  C/P,P
LORD, JEFFREY  PIRATES OF GOHAR [@BLADE #32]  P
LORD, JEFFREY  RETURN TO KALDAK [@BLADE #36]  P
LORD, JEFFREY  RUINS OF KALDAK, THE [@BLADE #34]  P
LORD, JEFFREY  SLAVE OF SARMA [@BLADE #4]  C/P:2,P
LORD, JEFFREY  TEMPLES OF AYOCAN, THE [@BLADE #14]  C/P,P
LORD, JEFFREY  TORIAN PEARLS, THE [@BLADE #25]  P
LORD, JEFFREY  TOWERS OF MELNON, THE [@BLADE #15]  P
LORD, JEFFREY  TREASURE OF THE STARS [@BLADE #29]  P
LORD, JEFFREY  UNDYING WORLD [@BLADE #8]  C/P,P
LORD, JEFFREY  WARLORDS OF GAIKON [@BLADE #18]  P
LORD, JEFFREY  WARRIORS OF LATAN [@BLADE #37]  P
LORD, JEFFREY  WIZARD OF RENTORO [@BLADE #28]  P
LORE, GORDON I. R./DENEAULT, HAROLD H. MYSTERIES OF THE SKIES: UFO’S IN PERSPECTIVE  [C/RAN-H]
LORENZ, J. D. JERRY BROWN: THE MAN ON THE WHITE HORSE [C/RAN-H]

LORENZEN, CORAL E.

FLYING SAUCERS: THE STARTLING EVIDENCE OF THE INVASION FROM OUTER SPACE

[C/RAN-P]

LORNQUEST, OLAF  MOONLOVERS, THE  C/P,P
LORRAH, JEAN  AMBROV KEON [SIME/GEN 7]  C/P:2,P
LORRAH, JEAN  CAPTIVES OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 3]  C/P,P
LORRAH, JEAN/LICHTENBERG, JACQUELINE CHANNEL’S DESTINY [SIME/GEN 5]  C/P:2,SFWA,P
LORRAH, JEAN  DRAGON LORD OF THE SAVAGE EMPIRE [SAVAGE EMPIRE 2]  C/P,P
LORRAH, JEAN  EMPRESS UNBORN [SAVAGE EMPIRE 7]  C/P,P
LORRAH, JEAN/LICHTENBERG, JACQUELINE FIRST CHANNEL [SIME/GEN 3]  C/P,P,H,SFWA
LORRAH, JEAN/HOWLETT, WINSTON A.  FLIGHT TO THE SAVAGE EMPIRE [SAVAGE EMPIRE 4]  C/P,P
LORRAH, JEAN  IDIC EPIDEMIC, THE [STAR TREK]  C/P:2,P
LORRAH, JEAN  METAMORPHOSIS [STAR TREK TNG]  C/P:2,P
LORRAH, JEAN  SAVAGE EMPIRE [SAVAGE EMPIRE 1]  C/P:2,P
LORRAH, JEAN  SORCERERS OF THE FROZEN ISLES [SAVAGE EMPIRE 5]  C/P,P

AUTHOR  TITLE  CODE
LORRAH, JEAN  SURVIVORS [STAR TREK TNG]  C/P:2,P
LORRAH, JEAN  VULCAN ACADEMY MURDERS, THE [STAR TREK]  C/P:2,P
LORRAINE, PAUL  DARK BOUNDARIES  SR-P
LORRAINE, PAUL  TWO WORLDS  SR-P,H
LORRAINE, PAUL  ZENITH-D  SR-P
LORY, ROBERT  CHALLENGE TO DRACULA [@DRACULA #9]  P
LORY, ROBERT  CURSE OF LEO, THE [@HORRORSCOPE #3]  P
LORY, ROBERT  DRACULA RETURNS [@DRACULA #1]  P
LORY, ROBERT  DRACULA’S BROTHERS [@DRACULA #3]  C/P,P
LORY, ROBERT  DRACULA’S DISCIPLE [@DRACULA #8]  C/P,P
LORY, ROBERT  DRACULA’S GOLD [@DRACULA #4]  P
LORY, ROBERT  DRUMS OF DRACULA [@DRACULA #5]  P
LORY, ROBERT  EYES OF BOLSK, THE  D77710,P
LORY, ROBERT  GEMINI SMILE, GEMINI KILL [@HORRORSCOPE #4]  C/P,P
LORY, ROBERT  GREEN FLAMES OF ARIES, THE [@HORRORSCOPE #1]  P
LORY, ROBERT  HAND OF DRACULA, THE [@DRACULA #2]  C/P,P
LORY, ROBERT  HARVEST OF HOODWINKS, A  D52180
LORY, ROBERT  IDENTITY SEVEN  C/P,P
LORY, ROBERT  MASTER OF THE ETRAX  C/P,P
LORY, ROBERT  MASTERS OF THE LAMP  D52180
LORY, ROBERT  REVENGE OF TAURUS, THE [@HORRORSCOPE #2]  P
LORY, ROBERT  THIRTEEN BRACELETS, THE  C/P:2,P
LORY, ROBERT  VEILED WORLD, THE  P,D31755
LORY, ROBERT  WITCHING OF DRACULA, THE [@DRACULA #6]  P
LOTTMAN, EILEEN  EXTRACURRICULAR ACTIVITIES [BIONIC WOMAN 2]  C/P:2,P
LOTTMAN, EILEEN  WELCOME HOME, JAIME [BIONIC WOMAN 1]  P
LOVE, KENNY  MILLENNIUM EVE  H
LOVE, ROSALEEN  TRAVELING TIDE, THE  LP
LOVECRAFT, H. P.  AT THE MOUNTAINS OF MADNESS  H:2,SR-H,P

LOVECRAFT, H. P.

BEST OF H. P. LOVECRAFT: BLOODCURDLING TALES OF HORROR AND THE MACABRE, THE

LP

LOVECRAFT, H. P.  BLACK SEAS OF INFINITY  H:2
LOVECRAFT, H. P.  CASE OF CHARLES DEXTER WARD, THE  C/P:2,P
LOVECRAFT, H. P.  COLOUR OUT OF SPACE AND OTHERS, THE  C/P:3,P
LOVECRAFT, H. P.  CRY HORROR!  C/P,P
LOVECRAFT, H. P.  DAGON AND OTHER MACABRE TALES  SR-H,H
LOVECRAFT, H. P.  DEMONS ET MERVEILLES  FP
LOVECRAFT, H. P.  DOOM THAT CAME TO SARNATH, THE  C/P:2,P
LOVECRAFT, H. P.  DREAM-QUEST OF UNKNOWN KADATH, THE  C/P:2,P
LOVECRAFT, H. P.  DUNWICH HORROR, THE  SR-H,H
LOVECRAFT, H. P.  DUNWICH HORROR, THE (BART)  SR-P
LOVECRAFT, H. P.  DUNWICH HORROR, THE (LANCER)  C/P:2,P
LOVECRAFT, H. P.  FUNGI FROM YUGGOTH AND OTHER POEMS  C/P,P
LOVECRAFT, H. P.  HORROR IN THE MUSEUM AND OTHER REVISIONS, THE  SR-H,C/P,P,H
LOVECRAFT, H. P./CONOVER, WILLIS  LOVECRAFT AT LAST  R/REF-H
LOVECRAFT, H. P./DERLETH, AUGUST  LURKER AT THE THRESHOLD, THE  SR-H,C/P:2,P
LOVECRAFT, H. P.  LURKING FEAR AND OTHER STORIES, THE (AVON)  SR-P,C/P
LOVECRAFT, H. P.  LURKING FEAR AND OTHER STORIES, THE (BALLANTINE)  C/P:2,P
LOVECRAFT, H. P./RAVEN, ANTHONY  OCCULT LOVECRAFT, THE  R/REF-P
LOVECRAFT, H. P.  SELECTED LETTERS I: 1911–1924  R/REF-H
LOVECRAFT, H. P.  SELECTED LETTERS II: 1925–1929  R/REF-H
LOVECRAFT, H. P.  SELECTED LETTERS III: 1929–1931  R/REF-H
LOVECRAFT, H. P.  SELECTED LETTERS IV: 1932–1934  R/REF-H
LOVECRAFT, H. P.  SELECTED LETTERS V: 1934–1937  R/REF-H
LOVECRAFT, H. P.  SHADOW OVER INNSMOUTH AND OTHER STORIES OF HORROR, THE  P
LOVECRAFT, H. P./DERLETH, AUGUST  SHUTTERED ROOM AND OTHER TALES OF HORROR, THE  C/P:2,P
LOVECRAFT, H. P./DERLETH, AUGUST  SURVIVOR AND OTHERS, THE  C/P:2,P
LOVECRAFT, H. P./HENDERSON, C. J.  TALES OF INSPECTOR LEGRASSE, THE  LP
LOVECRAFT, H. P.  TOMB AND OTHER TALES, THE  C/P:2,P
LOVECRAFT, H. P.  TRANSITION OF H.P. LOVECRAFT, THE  LP
LOVECRAFT, H. P.  UNCOLLECTED PROSE AND POETRY  SR-LP
LOVECRAFT, H. P./DERLETH, AUGUST  WATCHERS OUT OF TIME AND OTHERS, THE  SR-H,H
LOVECRAFT, H. P.  WEIRD SHADOW OVER INNSMOUTH, THE  SR-P
LOVECRAFT, H. P.  WINTER WISH AND OTHER POEMS, A  SR-H
LOVEGROVE, J. M. H.  KRILOV CONTINUUM, THE [GUARDIANS 1]  LP
LOVEGROVE, JAMES  HOW THE OTHER HALF LIVES  MB001
LOVEGROVE, JAMES  IMAGINED SLIGHTS  LP
LOVEJOY, JACK  BROTHERHOOD OF DIABLO, THE [@VISION OF BEASTS #3]  P,SFWA-P
LOVEJOY, JACK  CREATION DESCENDING [@VISION OF BEASTS #1]  C/P,P
LOVEJOY, JACK  DEFENDERS OF AR [GUARDIANS OF THE THREE 4]  C/P:2,P
LOVEJOY, JACK  HUNTERS, THE  C/P,P
LOVEJOY, JACK  MAGUS REX  C/P:2,P
LOVEJOY, JACK  OUTWORLD CATS  P
LOVEJOY, JACK  REBEL WITCH, THE  H

The MITSFS Pinkdex  September 24, 2007  Page 243
AUTHOR  TITLE  CODE
LOVEJOY, JACK  SECOND KINGDOM, THE [@VISION OF BEASTS #2]  C/P,SFWA-P,P
LOVEJOY, JACK  STAR GODS  P
LOVELACE, DELOS W.  KING KONG  C/P:2,P
LOVELL, JIM/KLUGER, JEFFREY  APOLLO 13  P
LOVELOCK, JAMES/ALLABY, MICHAEL  GREENING OF MARS, THE  [C/RAN-P]
LOVIN, ROGER  APOSTLE  H,LP
LOVISI, GARY  GARGOYLE, THE  B
LOVISI, GARY  GARGOYLE, THE (REVISED)  LP
LOVISI, GARY  MINESWEEPER  GRD8
LOVISI, GARY  SAGA OF FILSTER STEIN, THE  B
LOW, PROF. A. M.  ADRIFT IN THE STRATOSPHERE  H
LOW, PROF. A. M.  SATELLITE IN SPACE  H
LOWACHEE, KARIN  BURNDIVE  C/P,P
LOWACHEE, KARIN  CAGEBIRD  C/P,P
LOWACHEE, KARIN  WARCHILD  C/P,P
LOWDER, JAMES  CRUSADE [FORGOTTEN REALMS EMPIRES 3]  C/P:2,P
LOWDER, JAMES  KNIGHT OF THE BLACK ROSE [RAVENLOFT]  P
LOWDER, JAMES  PATH OF THE BOLD [SILVER AGE SENTINELS]  LPA
LOWDER, JAMES  PATH OF THE JUST [SILVER AGE SENTINELS]  LPA
LOWDER, JAMES  PRINCE OF LIES [FORGOTTEN REALMS AVATAR 4]  P
LOWDER, JAMES  REALMS OF INFAMY [FORGOTTEN REALMS]  PA
LOWDER, JAMES  REALMS OF VALOR [FORGOTTEN REALMS]  PA
LOWDER, JAMES  RING OF WINTER, THE [FORGOTTEN REALMS HARPERS #5]  P

LOWDER, JAMES/ WHITNEY-ROBINSON, VORONICA SPECTRE OF THE BLACK ROSE [RAVENLOFT]

C/P,P

LOWENTHAL, MARK M.  CRISPAN MAGICKER  C/P,P
LOWERY, LINDA  SECRET SORCERESS  P
LOWERY, LINDA  SPELL OF THE WINTER WIZARD  P
LOWNDES, ROBERT A. W.  BELIEVERS’ WORLD  H
LOWNDES, ROBERT A. W.  MYSTERY OF THE THIRD MINE  H:2
LOWNDES, ROBERT A. W.  PUZZLE PLANET, THE  DD485
LOWNDES, ROBERT A. W.  THREE FACES OF SCIENCE FICTION  R/REF-H
LOWRY, BRIAN  TRUTH IS OUT THERE [X-FILES]  R/REF-P,C/REF-H
LOWRY, LOIS  GIVER, THE  P
LUBAN, MILTON  SPIRIT WAS WILLING, THE  H
LUBAR, DAVID  HIDDEN TALENTS  H:2,LP
LUCAROTTI, JOHN  AZTECS, THE [DOCTOR WHO]  SR-P,C/P:2
LUCAROTTI, JOHN  MARCO POLO [DOCTOR WHO]  SR-P
LUCAROTTI, JOHN  MASSACRE, THE [DOCTOR WHO]  SR-P
LUCAS BOOKS  VECTOR PRIME (CD) [STAR WARS]  SR-MM
LUCAS, F. L.  WOMAN CLOTHED WITH THE SUN AND OTHER STORIES, THE  H
LUCAS, GEORGE  NEW HOPE, A [STAR WARS MOVIE 4]  P
LUCAS, GEORGE/CLAREMONT, CHRIS  SHADOW MOON [SHADOW WAR 1]  C/P,P,H
LUCAS, GEORGE  STAR WARS [STAR WARS]  C/P,H,SR-P
LUCAS, GEORGE  STAR WARS (RADIO SHOW) (7 CASSETTES) [STAR WARS]  SR-MM

LUCAS, GEORGE/BRACKETT, LEIGH/ KASDAN, LAWRENCE STAR WARS: THE SCRIPTS [STAR WARS]

SR-S

LUCASFILM LTD.  ART OF STAR WARS: EPISODE VI, THE [STAR WARS]  A
LUCASFILM LTD.  EMPIRE STRIKES BACK COLLECTORS EDITION  A
LUCASFILM LTD.  RETURN OF THE JEDI SOUVENIR BOOK  A
LUCCHETTE, STEPHEN C.  DOC – FIRST GALACTIC ROAMER  R/REF-H

LUCENO, JAMES AGENTS OF CHAOS

[@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS 1,2]

H

LUCENO, JAMES  BIG EMPTY, THE  C/P,P
LUCENO, JAMES  CLOAK OF DECEPTION [@STAR WARS]  P
LUCENO, JAMES  DARK LORD: THE RISE OF DARTH VADER [STAR WARS]  H
LUCENO, JAMES  DIMENSION X [@WEB WARRIORS #2]  P
LUCENO, JAMES  FEARFUL SYMMETRY, A  C/P,P
LUCENO, JAMES  HERO’S TRIAL [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #1]  P
LUCENO, JAMES  ILLEGAL ALIEN  C/P:2,P
LUCENO, JAMES  JEDI ECLIPSE [@STAR WARS NEW JEDI ORDER AGENTS OF CHAOS #2]  P
LUCENO, JAMES  LABYRINTH OF EVIL [@STAR WARS]  H
LUCENO, JAMES  MEMORIES END [@WEB WARRIORS #1]  P
LUCENO, JAMES  SHADOW, THE  P
LUCENO, JAMES  UNIFYING FORCE, THE [@STAR WARS NEW JEDI ORDER]  P

LUCENO, JAMES UNIFYING FORCE, THE (HARDCOVER AND CD)

[@STAR WARS NEW JEDI ORDER]

H

LUCIE-SMITH, EDWARD  HOLDING YOUR EIGHT HANDS  LPA
LUDLAM, HARRY  COMING OF JONATHAN SMITH, THE  SR-P
LUDWIG, EDWARD W.  MASK OF JON CULON, THE  H
LUGONES, LEOPOLDO  STRANGE FORCES  LP
LUIGI, BELLI  MASTER-MIND MENACE  SR-P
LUKEMAN, TIM  KOREN  H:2,P
LUKEMAN, TIM  RAJAN  H,SFWA,P

AUTHOR  TITLE  CODE
LUKEMAN, TIM  WITCHWOOD  H:2
LUKENS, ADAM  ALIEN WORLD  H
LUKENS, ADAM  CONQUEST OF LIFE  H
LUKENS, ADAM  EEVALU  H
LUKENS, ADAM  GLASS CAGE, THE  H
LUKENS, ADAM  SEA PEOPLE, THE  H
LUKENS, ADAM  SONS OF THE WOLF  H
LUKENS, ADAM  WORLD WITHIN, THE  H
LUKYANENKO, SERGEI  DAY WATCH [NIGHTWATCH 2]  H,LP
LUKYANENKO, SERGEI  NIGHT WATCH [NIGHTWATCH 1]  H,LP
LUKYANENKO, SERGEI  TWILIGHT WATCH [NIGHTWATCH 3]  LP
LUMLEY, BRIAN  BENEATH THE MOORS  SR-H
LUMLEY, BRIAN  BURROWERS BENEATH, THE  C/P,P
LUMLEY, BRIAN  CALLER OF THE BLACK, THE  SR-H
LUMLEY, BRIAN  CLOCK OF DREAMS, THE [TITUS CROW]  P
LUMLEY, BRIAN  COMPLEAT CROW, THE [TITUS CROW]  LP
LUMLEY, BRIAN  DEADSPAWN [NECROSCOPE 5]  P
LUMLEY, BRIAN  DEADSPEAK [NECROSCOPE 4]  P,SFWA-P
LUMLEY, BRIAN  DEMOGORGON  P
LUMLEY, BRIAN  HERO OF DREAMS [DREAMWORLD 1]  LP
LUMLEY, BRIAN  HORROR AT OAKDEENE AND OTHERS  H
LUMLEY, BRIAN  HOUSE OF CTHULHU, THE [PRIMAL LAND 1]  SR-VLP
LUMLEY, BRIAN  HOUSE OF DOORS, THE  C/P,SFWA-P,P
LUMLEY, BRIAN  ICED ON ARAN [DREAMLAND 4]  P
LUMLEY, BRIAN  IN THE MOONS OF BOREA  C/P,P
LUMLEY, BRIAN  KHAI OF ANCIENT KHEM  P
LUMLEY, BRIAN  KHAI OF KHEM  LP
LUMLEY, BRIAN  LAST AERIE, THE [NECROSCOPE 7/VAMPIRE WORLD 2]  H
LUMLEY, BRIAN  MAD MOON OF DREAMS [DREAMLAND 3]  P
LUMLEY, BRIAN  MAZE OF WORLDS  P
LUMLEY, BRIAN  NECROSCOPE [NECROSCOPE 1]  SFWA-P:2
LUMLEY, BRIAN  NECROSCOPE: AVENGERS [NECROSCOPE 13/E-BRANCH 3]  H
LUMLEY, BRIAN  NECROSCOPE: INVADERS [NECROSCOPE 11/E-BRANCH 1]  P
LUMLEY, BRIAN  NECROSCOPE: RESURGENCE [NECROSCOPE 10/LOST YEARS 2]  P
LUMLEY, BRIAN  NECROSCOPE: THE LOST YEARS [NECROSCOPE 9/LOST YEARS 1]  P
LUMLEY, BRIAN  PSYCHAMOK [PSYCHOMECH 3]  C/P,P
LUMLEY, BRIAN  PSYCHOMECH [PSYCHOMECH 1]  C/P,P
LUMLEY, BRIAN  PSYCHOSPHERE [PSYCHOMECH 2]  C/P,P
LUMLEY, BRIAN  SCREAMING SCIENCE FICTION  H
LUMLEY, BRIAN  SHIP OF DREAMS [DREAMWORLD 2]  H
LUMLEY, BRIAN  SORCERY IN SHAD [PRIMAL LAND 3]  H
LUMLEY, BRIAN  SOURCE, THE [NECROSCOPE 3]  C/P,SFWA-P:2,P
LUMLEY, BRIAN  SPAWN OF THE WINDS  C/P,P
LUMLEY, BRIAN  TARRA KHASH: HROSSAK! [PRIMAL LAND 2]  H
LUMLEY, BRIAN  TITUS CROW VOLUME THREE [TITUS CROW 5,6]  LP
LUMLEY, BRIAN  TOUCH, THE [@NECROSCOPE 14]  H
LUMLEY, BRIAN  TRANSITION OF TITUS CROW, THE [TITUS CROW 1]  C/P,P
LUMLEY, BRIAN  VAMPHYRI! [NECROSCOPE 2]  SFWA-P:2
LUMLEY, BRIAN  WHISPERER AND OTHER STORIES, THE  LP
LUNA, KRIS  STELLA RADIUM DISCHARGE  SR-P,H
LUNDWALL, SAM J.  ALICE’S WORLD  D58880
LUNDWALL, SAM J.  BERNHARD THE CONQUEROR  C/P:2,P
LUNDWALL, SAM J.  NO TIME FOR HEROES  D58880
LUNDWALL, SAM J.  SCIENCE FICTION: AN ILLUSTRATED HISTORY  R/REF-H
LUNDWALL, SAM J.  SCIENCE FICTION: WHAT IT’S ALL ABOUT  R/REF-P,C/REF-P:2
LUNDWALL, SAM J.  2018 A.D. OR THE KING KONG BLUES  C/P:2,P
LUPOFF, RICHARD  BEFORE... 12:01 ...AND AFTER  H
LUPOFF, RICHARD  CLAREMONT TALES II  H
LUPOFF, RICHARD  NIGHT OF THE LIVING GATOR! [@MELVINGE 3]  P
LUPOFF, RICHARD A./THOMPSON, DON  ALL IN COLOR FOR A DIME  R/REF-P
LUPOFF, RICHARD A.  BARSOOM: EDGAR RICE BURROUGHS AND THE MARTIAN VISION  R/REF-H
LUPOFF, RICHARD A.  BLACK TOWER, THE [DUNGEON #1]  C/P:2,P
LUPOFF, RICHARD A.  CIRCUMPOLAR!  H,P
LUPOFF, RICHARD A.  COUNTERSOLAR!  H,SFWA,P
LUPOFF, RICHARD A.  CRACK IN THE SKY, THE  P
LUPOFF, RICHARD A.  EDGAR RICE BURROUGHS: MASTER OF ADVENTURE  R/REF-P
LUPOFF, RICHARD A.  FINAL BATTLE, THE [DUNGEON #6]  C/P:2,P
LUPOFF, RICHARD A.  FOREVER CITY, THE  H
LUPOFF, RICHARD A.  GALAXY’S END  P
LUPOFF, RICHARD A.  INTO THE AETHER  C/P:2,P
LUPOFF, RICHARD A.  ONE MILLION CENTURIES  C/P:2,P
LUPOFF, RICHARD A.  SACRED LOCOMOTIVE FLIES  C/P,P
LUPOFF, RICHARD A.  SANDWORLD  C/P,P
LUPOFF, RICHARD A.  SPACE WAR BLUES  P

AUTHOR  TITLE  CODE
LUPOFF, RICHARD A.  SUN’S END  C/P:2,P
LUPOFF, RICHARD A.  SWORD OF THE DEMON  C/P,SFWA,P,H
LUPOFF, RICHARD A.  TRIUNE MAN, THE  C/P,P,H
LUPOFF, RICHARD A.  WHAT IF? VOLUME 1  PA
LUQUIENS, FREDERICK BLISS  SONG OF ROLAND, THE  LP
LUSTBADER, ERIC VAN  BATMAN: THE LAST ANGEL  CX
LUSTBADER, ERIC VAN  BENEATH AN OPAL MOON [SUNSET WARRIOR 4]  H,SFWA,P
LUSTBADER, ERIC VAN  DAI-SAN [SUNSET WARRIOR 3]  C/P,SFWA,P,H
LUSTBADER, ERIC VAN  RING OF FIVE DRAGONS, THE [PEARL 1]  H:2,P
LUSTBADER, ERIC VAN  RING OF FIVE DRAGONS, THE (PROOF) [PEARL 1]  SR-LP
LUSTBADER, ERIC VAN  SHALLOWS OF NIGHT [SUNSET WARRIOR 2]  C/P,SFWA,P,H
LUSTBADER, ERIC VAN  SUNSET WARRIOR, THE [SUNSET WARRIOR 1]  C/P:2,SFWA,P
LUSTBADER, ERIC VAN  VEIL OF A THOUSAND TEARS, THE [PEARL 2]  H
LUSTBADER, ERIC VAN  VEIL OF A THOUSAND TEARS, THE (PROOF) [PEARL 2]  H:2,SR-LP
LUSTBADER, ERIC VAN  WHITE NINJA  C/P,P
LUTHER, RAY  INTERMIND  P
LUTTRELL, WANDA  HOUSE OF ELNORA GARLAND, THE  P
LUTZ, JOHN  RIDE THE LIGHTNING [ALO NUDGER]  SFWA-P
LYDECKER, JOHN  DOCTOR WHO AND WARRIORS’ GATE [DOCTOR WHO]  SR-P,C/P:2
LYDECKER, JOHN  TERMINUS [DOCTOR WHO]  SR-P,C/P:2
LYLE, MEL  MYSTERY OF THE BURNING OCEAN, THE  H
LYLE, PETER  ROLIND OF MERU  C/P,P
LYMINGTON, JOHN  COMING OF THE STRANGERS, THE  P
LYMINGTON, JOHN  FROOMB!  H:2,P
LYMINGTON, JOHN  GREY ONES, THE  C/P,P
LYMINGTON, JOHN  NIGHT OF THE BIG HEAT  C/P,P
LYMINGTON, JOHN  NIGHT SPIDERS, THE  H,P
LYMINGTON, JOHN  NOWHERE PLACE, THE  H:2,SFWA
LYMINGTON, JOHN  SCREAMING FACE, THE  P
LYMINGTON, JOHN  SLEEP EATERS, THE  P
LYMINGTON, JOHN  STAR WITCHES, THE  P
LYMINGTON, JOHN  SWORD ABOVE THE NIGHT, A  P
LYMINGTON, JOHN  TEN MILLION YEARS TO FRIDAY  H:2,P
LYNCH, SCOTT  LIES OF LOCKE LAMORA, THE  H
LYNDON, BARRE/SANGSTER  MAN WHO COULD CHEAT DEATH, THE  P
LYNN, ELIZABETH A.  DANCERS OF ARUN, THE [TORNOR 2]  C/P,P,H,SFWA
LYNN, ELIZABETH A.  DIFFERENT LIGHT, A  C/P:2,P
LYNN, ELIZABETH A.  DRAGON’S TREASURE [DRAGON’S COUNTY 2]  H
LYNN, ELIZABETH A.  DRAGON’S WINTER [DRAGON’S COUNTY 1]  C/P,P
LYNN, ELIZABETH A.  NORTHERN GIRL, THE [TORNOR 3]  C/P:2,SFWA,P
LYNN, ELIZABETH A.  SARDONYX NET, THE  C/P,P,H
LYNN, ELIZABETH A.  SILVER HORSE, THE  SFWA
LYNN, ELIZABETH A.  WATCHTOWER [TORNOR 1]  C/P,P,H,SFWA
LYNN, ELIZABETH A.  WOMAN WHO LOVED THE MOON AND OTHER STORIES, THE  C/P:2,P
LYON, JENE  OUR SUN AND THE WORLDS AROUND IT  [C/RAN-H]
LYONS, STEVE  CONUNDRUM [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LYONS, STEVE  CROOKED WORLD, THE [DOCTOR WHO EIGHTH DOCTOR]  P
LYONS, STEVE  DEATH WORLD [@WARHAMMER 40K]  P
LYONS, STEVE  FINAL SANCTION, THE [DOCTOR WHO PAST DOCTORS]  P
LYONS, STEVE  HEAD GAMES [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
LYONS, STEVE  KILLING GROUND [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
LYONS, STEVE  MICRONAUTS [TIME TRAVELER 3]  P
LYONS, STEVE  MURDER GAME, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
LYONS, STEVE  SALVATION [DOCTOR WHO PAST DOCTORS]  C/P:2,P
LYONS, STEVE  SPACE AGE, THE [DOCTOR WHO EIGHTH DOCTOR]  P
LYONS, STEVE  STEALER OF DREAMS, THE [DOCTOR WHO NINTH DOCTOR]  H
LYONS, STEVE  TIME OF YOUR LIFE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
LYONS, STEVE  WITCH HUNTERS, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
LYRIS, SONIA ORIN  AND PEACE SHALL SLEEP [MAGIC: THE GATHERING]  P
LYTLE, ANDREW  NAME FOR EVIL, A  P

AUTHOR  TITLE  CODE
MACALISTER, KATIE  EVEN VAMPIRES GET THE BLUES  P
MACALISTER, KATIE  FIRE ME UP [AISLING GREY 2]  C/P,P
MACALISTER, KATIE  GIRL’S GUIDE TO VAMPIRES, A  P
MACALISTER, KATIE  LIGHT MY FIRE [AISLING GREY 3]  C/P,P
MACALISTER, KATIE  SEX AND THE SINGLE VAMPIRE  P
MACALISTER, KATIE  SEX, LIES, AND VAMPIRES  P
MACALISTER, KATIE  YOU SLAY ME [AISLING GREY 4]  C/P,P
MACAPP, C. C.  BUMSIDER  C/P:2,P
MACAPP, C. C.  OMHA ABIDES  C/P,P
MACAPP, C. C.  PRISONERS OF THE SKY  C/P,P
MACAPP, C. C.  RECALL NOT EARTH  C/P:2,P
MACAPP, C. C.  SUBB  C/P,P
MACAPP, C. C.  WORLDS OF THE WALL  C/P:2,P
MACARDLE, DOROTHY  UNINVITED, THE  H
MACARTHUR, DAVID S.  THUNDERBOLT MEN, THE  H
MACARTNEY, CLEM  DARK SIDE OF VENUS  SR-P
MACAULAY, DAVID  BAAA  H
MACAULEY, ROBIE  SECRET HISTORY OF TIME TO COME, A  H:2
MACAVOY, R. A.  BELLY OF THE WOLF, THE [@LENS OF THE WORLD #3]  C/P,P
MACAVOY, R. A.  BOOK OF KELLS, THE  C/P:2,P
MACAVOY, R. A.  DAMIANO [DAMIANO 1]  C/P:2,P
MACAVOY, R. A.  DAMIANO’S LUTE [DAMIANO 2]  C/P:2,P
MACAVOY, R. A.  GREY HORSE, THE  C/P:2,P
MACAVOY, R. A.  KING OF THE DEAD [LENS OF THE WORLD 2]  H:2
MACAVOY, R. A.  LENS OF THE WORLD [LENS OF THE WORLD #1]  H:2,@P
MACAVOY, R. A.  RAPHAEL [DAMIANO 3]  C/P:2,P
MACAVOY, R. A.  TEA WITH THE BLACK DRAGON  C/P:2,P
MACAVOY, R. A.  THIRD EAGLE, THE  C/P,P,H
MACAVOY, R. A.  TRIO FOR LUTE, A [DAMIANO 1,2,3]  H:2,P
MACAVOY, R. A.  TWISTING THE ROPE  C/P:2,P
MACCABEE, JOHN  DAY ONE  P
MACCLOUD, MALCOLM  TERA BEYOND, THE  H:2
MACCLURE, VICTOR  ULTIMATUM  H
MACCONNELL, COLUM  TARK AND THE GOLDEN TIDE  P
MACDONALD, GEORGE  AT THE BACK OF THE NORTH WIND  H,P
MACDONALD, GEORGE  EVENOR  C/P,P
MACDONALD, GEORGE  GOLDEN KEY, THE  H
MACDONALD, GEORGE  GRAY WOLF, THE  P
MACDONALD, GEORGE  LILITH  C/P:2,P
MACDONALD, GEORGE  LOST PRINCESS, THE  P
MACDONALD, GEORGE  PHANTASIES AND LILITH  LP
MACDONALD, GEORGE  PHANTASTES  C/P:2,P
MACDONALD, GEORGE  PRINCESS AND CURDIE, THE  SR-P,C/P
MACDONALD, GEORGE  PRINCESS AND THE GOBLIN, THE  SR-P,C/P:2
MACDONALD, GEORGE  PRINCESS AND THE GOBLIN, THE/PRINCESS AND CURDIE, THE  P
MACDONALD, GEORGE  WISE WOMAN, THE  P
MACDONALD, JAMES D.  APOCALYPSE DOOR, THE  H
MACDONALD, JOHN D.  BALLROOM OF THE SKIES  C/P,P,H
MACDONALD, JOHN D.  GIRL, THE GOLD WATCH AND EVERYTHING, THE  C/P,P
MACDONALD, JOHN D.  OTHER TIMES, OTHER WORLDS  C/P:2,P
MACDONALD, JOHN D.  PLANET OF THE DREAMERS  C/P:2,P
MACDONALD, JOHN D.  TIME AND TOMORROW  H:2
MACDONALD, JOHN D.  WINE OF THE DREAMERS  C/P,P,H
MACDOUGALL, W. B.  FALL OF THE NIBELUNGS, THE  A
MACE, DAVID  DEMON-4  C/P,P
MACE, DAVID  FIRE LANCE  C/P,P
MACEK, CARL  ROBOTECH ART 3: THE SENTINELS  A
MACFADDEN BOOKS  FROZEN PLANET, THE  PA,C/PA
MACFADDEN BOOKS  6 FINGERS OF TIME, THE  PA,C/PA
MACGREGOR, ELLEN/PANTELL, DORA  MISS PICKERELL AND THE WEATHER SATELLITE [@MISS PICKERELL #2]  P
MACGREGOR, ELLEN  MISS PICKERELL GOES TO MARS [@MISS PICKERELL]  C/P:2,LP
MACGREGOR, ELLEN/PANTELL, DORA  MISS PICKERELL MEETS MR. H.U.M. [@MISS PICKERELL #3]  P
MACGREGOR, ELLEN/PANTELL, DORA  MISS PICKERELL ON THE MOON [@MISS PICKERELL #1]  H:2,P
MACGREGOR, ELLEN/PANTELL, DORA  MISS PICKERELL TAKES THE BULL BY THE HORNS [@MISS PICKERELL #4]  P
MACGREGOR, KINLEY  KNIGHT OF DARKNESS [LORDS OF AVALON 1]  H
MACGREGOR, KINLEY  SWORD OF DARKNESS [LORDS OF AVALON 2]  H
MACGREGOR, LOREN J.  NET, THE  P,SFWA-P
MACGREGOR, ROB  INDIANA JONES AND THE DANCE OF THE GIANTS [INDIANA JONES]  P
MACGREGOR, ROB  INDIANA JONES AND THE GENESIS DELUGE [INDIANA JONES]  P
MACGREGOR, ROB  INDIANA JONES AND THE INTERIOR WORLD [INDIANA JONES]  C/P,P
MACGREGOR, ROB  INDIANA JONES AND THE LAST CRUSADE [INDIANA JONES]  P
MACGREGOR, ROB  INDIANA JONES AND THE SEVEN VEILS [INDIANA JONES]  P
MACGREGOR, ROB  INDIANA JONES AND THE UNICORN’S LEGACY [INDIANA JONES]  P
MACHEN, ARTHUR  BLACK CRUSADE  P

AUTHOR  TITLE  CODE
MACHEN, ARTHUR  HILL OF DREAMS, THE  H,SR-L
MACHEN, ARTHUR  TALES OF HORROR 1  C/P:2,P
MACHEN, ARTHUR  TALES OF HORROR AND THE SUPERNATURAL 2  P
MACHEN, ARTHUR  TERROR, THE  VLP
MACHEN, ARTHUR  THREE IMPOSTORS, THE  C/P,P
MACINTYRE, F. GWYNPLAINE  WOMAN BETWEEN THE WORLDS, THE  LP
MACK, DAVID  HARBINGER [STAR TREK VANGUARD]  P
MACK, DAVID  REAP THE WHIRLWIND [@STAR TREK VANGUARD]  P
MACK, DAVID  TIME TO HEAL, A [STAR TREK TNG TIME TO 8]  P
MACK, DAVID  TIME TO KILL, A [@STAR TREK TNG TIME TO 7]  P
MACK, DAVID  WARPATH [@STAR TREK DS9]  P
MACK, JOHN E., M. D.  ABDUCTION (CASSETTE)  SR-MM
MACKAY, SCOTT  MEEK, THE  P
MACKAY, SCOTT  OMNIFIX  P
MACKAY, SCOTT  ORBIS  C/P,P
MACKAY, SCOTT  OUTPOST  LP
MACKAY, SCOTT  PHYTOSPHERE  P
MACKAY, SCOTT  TIDES  H
MACKEL, KATHRYN  ALIEN IN A BOTTLE  H
MACKEL, KATHRYN  OUTRIDERS [BIRTHRIGHT PROJECT 1]  LP
MACKEL, KATHRYN  TRACKERS [BIRTHRIGHT PROJECT 2]  LP
MACKELWORTH, R. W.  DIABOLS, THE  C/P:2,P
MACKELWORTH, R. W.  STARFLIGHT 3000  C/P,P
MACKELWORTH, R. W.  TILTANGLE  C/P:2,P
MACKENROTH, NANCY  TREES OF ZHARKA, THE  P
MACKENZIE, ANDREW  FRONTIERS OF THE UNKNOWN  [C/RAN-P]
MACKENZIE, NIGEL  WORLD WITHOUT END  H
MACKEY, MARY  LAST WARRIOR QUEEN, THE  C/P,P,H
MACKLIN, JOHN  BROTHERHOOD OF THE STRANGE  [C/RAN-P]
MACKLIN, JOHN  OUT OF THIS WORLD  [C/RAN-P]
MACKLIN, JOHN  STRANGE AND UNCANNY, THE  [C/RAN-P]
MACKLIN, JOHN  STRANGE DESTINIES  [C/RAN-P]
MACKSEY, KENNETH  INVASION  H
MACLAINE, JENNA  WAGES OF SIN, THE  LP
MACLAY, JOHN  NUKES  SFWA
MACLEAN, ALISTAIR  ICE STATION ZEBRA  P
MACLEAN, ALISTAIR  SATAN BUG, THE  P
MACLEAN, KATHERINE  DIPLOIDS, THE  C/P:2,P
MACLEAN, KATHERINE  MAN IN THE BIRD CAGE, THE  P
MACLEAN, KATHERINE  MISSING MAN  H:2,P
MACLEAN, KATHERINE  TROUBLE WITH YOU EARTH PEOPLE, THE  LP
MACLEISH, RODERICK  PRINCE OMBRA  C/P,P
MACLEOD, CHARLOTTE  CURSE OF THE GIANT HOGWEED, THE [PROFESSOR SHANDY 5]  P
MACLEOD, IAN R.  BREATHMOSS AND OTHER EXHALATIONS  H
MACLEOD, IAN R.  LIGHT AGES, THE  H
MACLEOD, IAN R.  VOYAGES BY STARLIGHT  H
MACLEOD, KEN  CASSINI DIVISION, THE [FALL REVOLUTION]  H,P
MACLEOD, KEN  COSMONAUT KEEP [ENGINES OF LIGHT 1]  H
MACLEOD, KEN  DARK LIGHT [ENGINES OF LIGHT 2]  H
MACLEOD, KEN  ENGINE CITY [ENGINES OF LIGHT 3]  H
MACLEOD, KEN  FALL REVOLUTION, THE [FALL REVOLUTION]  H
MACLEOD, KEN  GIANT LIZARDS FROM ANOTHER STAR  H
MACLEOD, KEN  LEARNING THE WORLD  H
MACLEOD, KEN  NEWTON’S WAKE  H
MACLEOD, KEN  SKY ROAD, THE [FALL REVOLUTION]  H
MACLEOD, KEN  STAR FRACTION, THE [FALL REVOLUTION]  H,LP
MACLEOD, KEN  STONE CANAL, THE [FALL REVOLUTION]  C/P,LP
MACLEOD, SHEILA  XANTHE AND THE ROBOTS  SR-P
MACMILLAN, IAN  BLAKELY’S ARK  P
MACMILLAN PUBLISHING  ALIENS, TRAVELERS, AND OTHER STRANGERS  HA
MACMILLAN PUBLISHING  EARTH AND ELSEWHERE  HA
MACMILLAN PUBLISHING  NEW SOVIET SCIENCE FICTION  HA
MACMILLAN, SCOTT  AT SWORD’S POINT [@KNIGHTS OF THE BLOOD]  C/P,P
MACMILLAN, SCOTT  VAMPYR-SS [@KNIGHTS OF THE BLOOD 1]  C/P,P
MACTYRE, PAUL  DOOMSDAY, 1999  C/P:2,P
MACVEY, JOHN W.  WHISPERS FROM SPACE  SFWA
MADARIAGA, SALVADOR DE  SACRED GIRAFFE, THE  H
MADDEN, TIMOTHY A.  OUTBANKER  C/P,P
MADDOCK, LARRY  EMERALD ELEPHANT GAMBIT, THE [@AGENT OF TERRA #3]  C/P:2,P
MADDOCK, LARRY  FLYING SAUCER GAMBIT, THE [@AGENT OF TERRA #1]  C/P:2,P
MADDOCK, LARRY  GOLDEN GODDESS GAMBIT, THE [@AGENT OF TERRA #2]  C/P:2,P
MADDOCK, LARRY  TIME TRAP GAMBIT, THE [@AGENT OF TERRA #4]  C/P:2,P
MADDOX, TOM  HALO  H:2,P
MADDUX, RACHEL  GREEN KINGDOM, THE  P

AUTHOR  TITLE  CODE
MADISON, J. J.  THING, THE  P
MADSEN, AXEL  UNISAVE  P
MADSEN, DAVID  U.S.S.A.  H:2
MAETERLINCK, MAURICE  UNKNOWN GUEST, THE  H
MAGAZINE OF F & SF  ESPER GREIFEN EIN, DIE  FP
MAGGIN, ELLIOT S.  KINGDOM COME  C/P,P
MAGGIN, ELLIOT S.  SUPERMAN: LAST SON OF KRYPTON [DC COMICS]  C/P:2,P
MAGGIN, ELLIOT S.  SUPERMAN: MIRACLE MONDAY [DC COMICS]  P
MAGIDOFF, ROBERT  RUSSIAN SCIENCE FICTION  HA
MAGIDOFF, ROBERT  RUSSIAN SCIENCE FICTION: 1968  HA
MAGIDOFF, ROBERT  RUSSIAN SCIENCE FICTION: 1969  HA
MAGINN, SIMON  VIRGINS AND MARTYRS  P
MAGLIO, MITCHELL  OFFICIAL STAR TREK QUIZ BOOK, THE [STAR TREK]  [C/RAN-P]
MAGOG, H. -J.  VALLEE ENSEVELIE, LA  FP
MAGROON, VECTOR  BURNING VOID  SR-P
MAGRS, PAUL/HOAD, JEREMY  BLUE ANGEL, THE [DOCTOR WHO EIGHTH DOCTOR]  P
MAGRS, PAUL  MAD DOGS AND ENGLISHMEN [DOCTOR WHO EIGHTH DOCTOR]  P
MAGRS, PAUL  SCARLET EMPRESS, THE [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
MAGRS, PAUL  VERDIGRIS [DOCTOR WHO PAST DOCTORS]  C/P,P
MAGUIRE, GREGORY  SON OF A WITCH  H
MAGUIRE, GREGORY  WICKED  LP
MAHR, KURT  ABTEILUNG PSI [@PERRY RHODAN #131]  FPR
MAHR, KURT  ATOOM-ALARM! (DUTCH) [@PERRY RHODAN #5]  FPR
MAHR, KURT  BEASTS BELOW, THE [PERRY RHODAN 128]  SR-LP
MAHR, KURT  BETWEEN THE GALAXIES [PERRY RHODAN 119]  SR-LP
MAHR, KURT  FORTRESS IN TIME [PERRY RHODAN 123]  SR-LP
MAHR, KURT  PLANETENKONIG, DER [@PERRY RHODAN #61]  FPR
MAHR, KURT  STATION OF THE INVISIBLES [PERRY RHODAN 133]  SR-LP
MAHR, KURT  SUPERNOVA [@PERRY RHODAN #66]  FPR
MAHR, KURT  TOCHTER DES ROBOTERS, DIE [@PERRY RHODAN #6]  FPR
MAHY, MARGARET  CHANGEOVER, THE  C/P,P
MAHY, MARGARET  TRICKSTERS, THE  SFWA
MAILER, NORMAN  ANCIENT EVENINGS  H
MAILER, NORMAN  OF A FIRE ON THE MOON  [C/RAN-P]
MAINE, CHARLES ERIC  ALPH  H:2,P
MAINE, CHARLES ERIC  BATALHA DO VACUO, A  FP
MAINE, CHARLES ERIC  B.E.A.S.T.  C/P:2,P
MAINE, CHARLES ERIC  FIRE PAST THE FUTURE  C/P:2,P
MAINE, CHARLES ERIC  HE OWNED THE WORLD  C/P,P,H
MAINE, CHARLES ERIC  HIGH VACUUM  C/P:2,P
MAINE, CHARLES ERIC  ISOTOPE MAN, THE  H:2
MAINE, CHARLES ERIC  MAN WHO COULDN’T SLEEP, THE  H:2
MAINE, CHARLES ERIC  MIND OF MR. SOAMES, THE  SR-P
MAINE, CHARLES ERIC  SPACEWAYS  SR-P
MAINE, CHARLES ERIC  SURVIVAL MARGIN  C/P:2,P
MAINE, CHARLES ERIC  TIDE WENT OUT, THE  C/P:2,P
MAINE, CHARLES ERIC  TIMELINER  H:2,P
MAINE, CHARLES ERIC  VIRUS DESTRUIDOR, O  FP
MAINE, CHARLES ERIC  WORLD WITHOUT MEN  C/P:2,P
MAJOR, H. M.  ALIEN TRACE, THE [@ALIEN TRACE #1]  P
MAJOR, H. M.  TIME TWISTER [@ALIEN TRACE #2]  P
MAJORS, SIMON  DRUID STONE, THE  C/P:2,P
MALACLYPSE THE YOUNGER  PRINCIPIA DISCORDIA  [C/RAN-H,C/RAN-P]
MALAQUAIS, JEAN  JOKER, THE  P
MALCOLM, DONALD  IRON RAIN, THE  SR-P
MALCOLM, DONALD  UNKNOWN SHORE, THE  SR-P
MALCOLM, SALLY  COST OF HONOR, THE [@STARGATE SG-1 #5]  P
MALCOLM, SALLY  MATTER OF HONOR, A [@STARGATE SG-1 #3]  P
MALCOLM, SALLY  RISING [STARGATE ATLANTIS #1]  P
MALCOLM-SMITH, GEORGE  GRASS IS ALWAYS GREENER, THE  H
MALEC, ALEXANDER  EXTRAPOLASIS  C/P:3,P
MALININA  AL’MANAKH NAUCHNOI FANTASTIKI 1963, VYPUSK 1  FH
MALIS, JODY CAMERON  DARK SHADOWS COOKBOOK, THE  [C/RAN-P]
MALONE, ADRIAN  SECRET, THE  H
MALONEY, GEOFFREY  NOR OF HUMAN...  LPA
MALONEY, MACK  BATTLE AT ZERO POINT [@STARHAWK 4]  P
MALONEY, MACK  CIRCLE WAR [WINGMAN 2]  P
MALONEY, MACK  FINAL STORM, THE [WINGMAN 6]  P
MALONEY, MACK  FOURTH EMPIRE, THE [@STARHAWK 3]  P
MALONEY, MACK  FREEDOM EXPRESS [WINGMAN 7]  P
MALONEY, MACK  LUCIFER CRUSADE, THE [WINGMAN 3]  P
MALONEY, MACK  PLANET AMERICA [@STARHAWK 2]  P
MALONEY, MACK  STARHAWK [@STARHAWK 1]  P
MALONEY, MACK  STORM OVER SATURN [@STARHAWK 5]  P

The MITSFS Pinkdex  September 24, 2007  Page 250
AUTHOR  TITLE  CODE
MALONEY, MACK  TWISTED CROSS, THE [WINGMAN 5]  P
MALONEY, MACK  WINGMAN [WINGMAN 1]  P
MALORY, SIR THOMAS  LE MORTE D’ARTHUR (BAINES RENDITION)  P
MALTZ, MAXWELL, M. D., F. I. C. S.  PSYCHO-CYBERNETICS  [C/RAN-P]
MALZBERG, BARRY N.  BEST OF BARRY N. MALZBERG, THE  C/P:2,P
MALZBERG, BARRY N.  BEST TIME TRAVEL STORIES OF ALL TIME, THE  PA
MALZBERG, BARRY N.  BEYOND APOLLO  H:2,P
MALZBERG, BARRY N.  BREAKFAST IN THE RUINS  R/REF-H
MALZBERG, BARRY N.  CHORALE  H,SFWA
MALZBERG, BARRY N.  CONVERSATIONS  H
MALZBERG, BARRY N.  CROSS OF FIRE, THE  C/P,P
MALZBERG, BARRY N./PRONZINI, BILL  DARK SINS, DARK DREAMS  HA,SFWA
MALZBERG, BARRY N.  DAY OF THE BURNING, THE  C/P,P
MALZBERG, BARRY N.  DESTRUCTION OF THE TEMPLE, THE  C/P,P
MALZBERG, BARRY N.  DOWN HERE IN THE DREAM QUARTER  H,SFWA
MALZBERG, BARRY N./PRONZINI, BILL  END OF SUMMER, THE  C/PA,PA
MALZBERG, BARRY N.  ENGINES OF THE NIGHT, THE  SFWA,R/REF-H,C/REF-H
MALZBERG, BARRY N.  FALLING ASTRONAUTS, THE  P
MALZBERG, BARRY N.  GALAXIES  C/P,P
MALZBERG, BARRY N.  GAMESMAN, THE  C/P,P
MALZBERG, BARRY N.  GUERNICA NIGHT  H:2,P
MALZBERG, BARRY N.  HEROVIT’S WORLD  C/P:2,SFWA,P,H
MALZBERG, BARRY N.  IN MY PARENTS’ BEDROOM  SR-P
MALZBERG, BARRY N.  IN THE ENCLOSURE  P
MALZBERG, BARRY N.  LAST TRANSACTION, THE  P
MALZBERG, BARRY N.  MALZBERG AT LARGE  P
MALZBERG, BARRY N.  MAN WHO LOVED THE MIDNIGHT LADY, THE  SFWA
MALZBERG, BARRY N.  MANY WORLDS OF BARRY MALZBERG, THE  C/P,P
MALZBERG, BARRY N.  MEN INSIDE, THE  P

MALZBERG, BARRY N./ GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

NEGLECTED VISIONS

HA

MALZBERG, BARRY N.  ON A PLANET ALIEN  C/P:2,P
MALZBERG, BARRY N.  OUT FROM GANYMEDE  P
MALZBERG, BARRY N.  OVERLAY  C/P:2,P
MALZBERG, BARRY N.  PASSAGE OF THE LIGHT, THE  LP
MALZBERG, BARRY N.  PHASE IV  P
MALZBERG, BARRY N.  REMAKING OF SIGMUND FREUD, THE  C/P,P
MALZBERG, BARRY N.  REVELATIONS  P
MALZBERG, BARRY N.  SCOP  C/P,P
MALZBERG, BARRY N.  SCREEN  P
MALZBERG, BARRY N.  SODOM AND GOMORRAH BUSINESS, THE  P
MALZBERG, BARRY N.  SPREAD, THE  SR-P,C/P
MALZBERG, BARRY N.  TACTICS OF CONQUEST  C/P:2,P
MAMATAS, NICK  MOVE UNDER GROUND  LP
MANACHINO, ALBERT J.  DARK BIBLE  B
MANADAY, TILPER  TERRAPIN OR  H:2,LP
MANCOUR, T. L.  SPARTACUS [STAR TREK TNG]  C/P:3,P

MANDER, JERRY/DIPPEL, GEORGE/ GOSSAGE, HOWARD GREAT INTERNATIONAL PAPER AIRPLANE BOOK, THE

[C/RAN-H]

MANGELS, ANDY/MARTIN, MICHAEL A.  GOOD THAT MEN DO, THE [STAR TREK ENTERPRISE]  P
MANGELS, ANDY/MARTIN, MICHAEL A.  RED KING, THE [STAR TREK TITAN]  P
MANGELS, ANDY/MARTIN, MICHAEL A.  ROGUE [@STAR TREK TNG SECTION 31]  P
MANGUEL, ALBERTO/GUADALUPI, GIANNI  DICTIONARY OF IMAGINARY PLACES, THE  R/REF-H,C/REF-H
MANGUEL, ALBERTO/GUADALUPI, GIANNI  DICTIONARY OF IMAGINARY PLACES, THE (EXPANDED EDITION)  R/REF-H
MANLEY, SEON/LEWIS, GOGO  SHAPES OF THE SUPERNATURAL  HA
MANN, GEORGE  SEVERED MAN, THE [@TIME HUNTER 5]  LP
MANN, GEORGE  SOLARIS BOOK OF NEW SCIENCE FICTION, THE  PA
MANN, PHILLIP  EYE OF THE QUEEN, THE  H:2,P
MANN, PHILLIP  FALL OF THE FAMILIES, THE [GARDNER 2]  H
MANN, PHILLIP  MASTER OF PAXWAX [GARDNER 1]  H
MANN, PHILLIP  WULFSYARN  C/P,SFWA,SFWA-P,P,H
MANN, THOMAS  TRANSPOSED HEADS, THE  H,P
MANNES, MARYA  THEY  C/P,P
MANNHEIM, KARL  WHEN THE EARTH DIED  P
MANNING, LAURENCE  MAN WHO AWOKE, THE  C/P:2,P
MANNING, P. L.  DESTROYERS, THE  SR-P
MANNING, ROSEMARY  DRAGON IN DANGER  P
MANNION, MICHAEL  DEATH CLOUD  P
MANO, D. KEITH  BRIDGE, THE  C/P,P,H
MANOR BOOKS  4 FUTURES  C/PA,PA
MANSON, CYNTHIA/ARDAI, CHARLES  FUTURE CRIME  HA
MANTLEY, JOHN  TWENTY-SEVENTH DAY, THE  H:2,P
MANVELL, ROGER  DREAMERS, THE  P
MAO TSE-TUNG  QUOTATIONS FROM CHAIRMAN MAO TSE-TUNG  [C/RAN-P]

AUTHOR  TITLE  CODE
MARAS, KARL  ZHORANI: MASTER OF THE UNIVERSE  SR-P
MARCEL, GABRIEL  DECLIN DE LA SAGESSE, LE  FP
MARCELLAS, DIANA  MOTHER OCEAN, DAUGHTER SEA [BRIERLY 1]  P
MARCELLAS, DIANA  SEA LARK’S SONG, THE [BRIERLY 2]  P
MARCH, GENE  SHAKWA, THE  P
MARCO, JOHN  DEVIL’S ARMOR [EYES OF GOD 2]  P
MARCO, JOHN  EYES OF GOD [EYES OF GOD 1]  H,P
MARCO, JOHN  GRAND DESIGN, THE [TYRANTS AND KINGS 2]  C/P,LP
MARCO, JOHN  JACKAL OF NAR, THE [TYRANTS AND KINGS 1]  C/P:2,LP
MARCO, JOHN  SAINTS OF THE SWORD, THE [TYRANTS AND KINGS 3]  H,LP
MARCO, JOHN  SWORD OF ANGELS, THE [EYES OF GOD 3]  H
MARCUS, ROBERT B. JR.  SHADOW ON THE STARS  SR-P
MARDEN, WILLIAM  EXILE OF ELLENDON, THE  H,SFWA
MARGOLIS, JACK S.  LINDA LOVELACE FOR PRESIDENT  SR-P
MARGULIES, LEO  GET OUT OF MY SKY  PA,C/PA:2
MARGULIES, LEO  GHOUL KEEPERS, THE  PA,C/PA
MARGULIES, LEO/FRIEND, OSCAR J.  GIANT ANTHOLOGY OF SCIENCE FICTION, THE  HA
MARGULIES, LEO/FRIEND, OSCAR J.  MY BEST SCIENCE FICTION STORY  C/PA,PA
MARGULIES, LEO/FRIEND, OSCAR J.  RACE TO THE STARS  PA
MARGULIES, LEO  THREE FROM OUT THERE  PA,C/PA
MARGULIES, LEO  THREE IN ONE  PA,C/PA
MARGULIES, LEO  THREE TIMES INFINITY  PA,C/PA
MARGULIES, LEO  UNEXPECTED, THE  C/PA,PA
MARGULIES, LEO  WEIRD TALES  C/PA,PA
MARGULIES, LEO  WORLDS OF WEIRD  C/PA,PA
MARILLIER, JULIET  BLADE OF FORTRIU [@BRIDEI #2]  H
MARILLIER, JULIET  CHILD OF THE PROPHECY [SEVENWATERS 3]  H
MARILLIER, JULIET  DARK MIRROR, THE [@BRIDEI #1]  H
MARILLIER, JULIET  DAUGHTER OF THE FOREST [SEVENWATERS 1]  H
MARILLIER, JULIET  FOXMASK  LP
MARILLIER, JULIET  SON OF THE SHADOWS [SEVENWATERS 2]  H
MARILLIER, JULIET  WELL OF SHADES, THE [BRIDEI 3]  H
MARILLIER, JULIET  WILDWOOD DANCING  H
MARILLIER, JULIET  WOLFSKIN  H
MARINACCIO, DAVE  ALL I REALLY NEED TO KNOW I LEARNED FROM WATCHING STAR TREK  [C/RAN-H:2]
MARIOTTE, JEFF  ATTITUDE OF SILENCE, THE [ANDROMEDA (GENE RODDENBERRY’S) 5]  H
MARIOTTE, JEFF  CLOSE TO THE GROUND [@ANGEL]  C/P,P
MARIOTTE, JEFF  DAWN OF THE ICE BEAR [@AGE OF CONAN MARAUDERS #3]  P
MARIOTTE, JEFF  DENY THY FATHER [@STAR TREK LOST ERA (2355-57 AD)]  P
MARIOTTE, JEFF  GHOST OF THE WALL [@AGE OF CONAN MARAUDERS #1]  P
MARIOTTE, JEFF  HAUNTED [@ANGEL]  P
MARIOTTE, JEFF  HOLLYWOOD NOIR [@ANGEL]  P
MARIOTTE, JEFF  LOVE AND DEATH [@ANGEL]  P
MARIOTTE, JEFF  MIRROR IMAGE [@CHARMED]  P
MARIOTTE, JEFF  SANCTUARY [@ANGEL]  P
MARIOTTE, JEFF  SOLITARY MAN [@ANGEL]  P
MARIOTTE, JEFF  STRANGER TO THE SUN [@ANGEL]  P
MARIOTTE, JEFF  SURVIVAL OF THE FITTEST [@CHARMED]  P
MARIOTTE, JEFF  WINDS OF THE WILD SEA [@AGE OF CONAN MARAUDERS #2]  P
MARIOTTE, JEFF  XANDER YEARS, THE, VOL. 2 [BUFFY]  P
MARK, JAN  AQUARIUS  SFWA
MARK, JAN  ENNEAD, THE  P
MARK, TED  BACK HOME AT THE O.R.G.Y. [MAN FROM ORGY]  SR-P
MARK, TED  COME BE MY O.R.G.Y. [MAN FROM ORGY]  SR-P,C/P
MARKS, ALAN  ANTENNA SYNDROME, THE  C/P,P
MARKS, ANTON  IN THE DAYS OF DREAD  LP
MARKS, LAURIE J.  ARA’S FIELD  P
MARKS, LAURIE J.  DANCING JACK  P
MARKS, LAURIE J.  DELAN THE MISLAID  P
MARKS, LAURIE J.  EARTH LOGIC [ELEMENTAL LOGIC 2]  H
MARKS, LAURIE J.  FIRE LOGIC [ELEMENTAL LOGIC 1]  H
MARKS, LAURIE J.  MOONBANE MAGE, THE  P
MARKS, LAURIE J.  WATCHER’S MASK, THE  P
MARKS, LAURIE J.  WATER LOGIC [ELEMENTAL LOGIC 3]  LP
MARLEY, LOUISE  GLASS HARMONICA, THE  C/P,LP
MARLEY, LOUISE  MAQUISARDE, THE  P
MARLEY, LOUISE  RECEIVE THE GIFT [CANTRIX 3]  P
MARLEY, LOUISE  SING THE LIGHT [CANTRIX 1]  P
MARLEY, LOUISE  SING THE WARMTH [CANTRIX 2]  P
MARLEY, LOUISE  TERRORISTS OF IRUSTAN, THE  C/P:2,LP
MARLEY, STEPHEN  MANAGRA [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
MARLEY, STEPHEN  MORTAL MASK [CHIA BLACK DRAGON 2]  LP
MARLOWE, CHRISTOPHER  DOCTOR FAUSTUS  LP
MARLOWE, STEPHEN  TRANSLATION  P

The MITSFS Pinkdex  September 24, 2007  Page 252
AUTHOR  TITLE  CODE
MARNEY, DEAN  CHRISTMAS TREE THAT ATE MY MOTHER, THE  LP
MARONEY, TIM  BOOK OF DZYAN, THE  R/REF-H
MARQUIS, ROY  MOON MONSTERS, THE  SR-P
MARRIOTT, CRITTENDEN  ISLE OF DEAD SHIPS, THE  H
MARSH, GEOFFREY  FANGS OF THE HOODED DEMON, THE [LINCOLN BLACKTHORNE 4]  SFWA-P
MARSH, GEOFFREY  KING OF SATAN’S EYES, THE [LINCOLN BLACKTHORNE 1]  H,SFWA,SFWA-P
MARSH, GEOFFREY  PATCH OF THE ODIN SOLDIER, THE [LINCOLN BLACKTHORNE 3]  H,SFWA,SFWA-P
MARSH, GEOFFREY  TAIL OF THE ARABIAN, KNIGHT, THE [LINCOLN BLACKTHORNE 2]  P,SFWA,SFWA-P
MARSHACK, ALEXANDER  WORLD IN SPACE, THE  [C/RAN-P:2]
MARSHAK, SONDRA/CULBREATH, MYRNA  FATE OF THE PHOENIX, THE [STAR TREK]  C/P:3,P
MARSHAK, SONDRA/CULBREATH, MYRNA  PRICE OF THE PHOENIX, THE [STAR TREK]  C/P:3,P
MARSHAK, SONDRA/CULBREATH, MYRNA  PROMETHEUS DESIGN, THE [STAR TREK]  C/P:3,P
MARSHAK, SONDRA/CULBREATH, MYRNA  STAR TREK: THE NEW VOYAGES [STAR TREK]  SR-PA,C/PA:2
MARSHAK, SONDRA/CULBREATH, MYRNA  STAR TREK: THE NEW VOYAGES 2 [STAR TREK]  SR-PA,C/PA:2
MARSHAK, SONDRA/CULBREATH, MYRNA  TRIANGLE [STAR TREK]  C/P:3,P
MARSHAL, GENE/WAEDT, CARL F.  INCREDIBLE ADVENTURES #1  SR-LPA
MARSHALL, BRUCE  FATHER MALACHY’S MIRACLE  P
MARSHALL, EDISON  LOST LAND, THE  P
MARSTEN, RICHARD  DANGER: DINOSAURS!  H:2
MARSTEN, RICHARD  DANGER...DINOSAURES!  FP
MARSTEN, RICHARD  ROCKET TO LUNA  H
MARSTON, ANN  BROKEN BLADE [RUNE BLADE 3]  P
MARSTON, ANN  CLOUDBEARER’S SHADOW [SWORD IN EXILE 1]  P
MARSTON, ANN  KING OF SHADOWS [SWORD IN EXILE 2]  P
MARSTON, ANN  KINGMAKER’S SWORD [RUNE BLADE 1]  P
MARSTON, ANN  SWORD AND SHADOW [SWORD IN EXILE 3]  P
MARSTON, ANN  WESTERN KING, THE [RUNE BLADE 2]  P
MARTEL, SUZANNE  CITY UNDER GROUND, THE  H
MARTEN, JACQUELINE  NIGHTMARE IN RED  P
MARTEN, JACQUELINE  VISIONS OF THE DAMNED  P
MARTER, IAN  DOCTOR WHO AND THE ARK IN SPACE [DOCTOR WHO]  SR-P
MARTER, IAN  DOCTOR WHO AND THE ENEMY OF THE WORLD [DOCTOR WHO]  SR-P,C/P
MARTER, IAN  DOCTOR WHO AND THE RIBOS OPERATION [DOCTOR WHO]  SR-P,C/P
MARTER, IAN  DOCTOR WHO AND THE SONTARAN EXPERIMENT [DOCTOR WHO]  SR-P
MARTER, IAN  DOMINATORS, THE [DOCTOR WHO]  SR-P,C/P:2
MARTER, IAN  EARTHSHOCK [DOCTOR WHO]  SR-P,C/P
MARTER, IAN  HARRY SULLIVAN’S WAR [DOCTOR WHO]  SR-P,C/P
MARTER, IAN  INVASION, THE [DOCTOR WHO]  SR-P,P
MARTER, IAN  REIGN OF TERROR, THE [DOCTOR WHO]  SR-P
MARTER, IAN  RESCUE, THE [DOCTOR WHO]  SR-P,C/P
MARTI-IBANEZ, FELIX  ALL THE WONDERS WE SEEK  SR-H
MARTIN, DAVID  SEARCH FOR THE DOCTOR [DOCTOR WHO]  [C/RAN-P]
MARTIN, ED  BUSY BODIES  SR-L
MARTIN, ED  FRANKENSTEIN ’69  SR-L
MARTIN, GAIL Z.  SUMMONER, THE [@NECROMANCER #1]  P
MARTIN, GEORGE R. R.  ACE IN THE HOLE [@WILD CARDS #6]  C/PA,PA
MARTIN, GEORGE R. R.  ACES ABROAD [@WILD CARDS #4]  C/PA,PA
MARTIN, GEORGE R. R.  ACES HIGH [@WILD CARDS #2]  C/PA,PA
MARTIN, GEORGE R. R.  ARMAGEDDON RAG, THE  C/P,P,H

MARTIN, GEORGE R. R.

ARMS OF THE KRAKEN, THE (NOVELLA FROM DRAGON#305) [SONG OF ICE AND FIRE 4.01]

H

MARTIN, GEORGE R. R.  BLACK TRUMP [@WILD CARDS NEW CYCLE 3]  C/PA,PA,SFWA-P
MARTIN, GEORGE R. R.  CARD SHARKS [@WILD CARDS NEW CYCLE 1]  PA,C/PA,SFWA-P
MARTIN, GEORGE R. R.  CLASH OF KINGS, A [SONG OF ICE AND FIRE 2]  C/P,LP,H:2
MARTIN, GEORGE R. R.  CLASH OF KINGS, A (BRITISH) [SONG OF ICE AND FIRE 2]  H
MARTIN, GEORGE R. R./MILLER, JOHN J.  DEAD MAN’S HAND [@WILD CARDS #7]  C/P:2,P
MARTIN, GEORGE R. R.  DEALER’S CHOICE [@WILD CARDS #11]  PA,C/PA
MARTIN, GEORGE R. R.  DOWN AND DIRTY [@WILD CARDS #5]  C/PA,PA
MARTIN, GEORGE R. R.  DYING OF THE LIGHT  H:2,SFWA-P,P
MARTIN, GEORGE R. R.  FEAST FOR CROWS, A [SONG OF ICE AND FIRE 4]  H
MARTIN, GEORGE R. R.  FEVRE DREAM  H:2,P
MARTIN, GEORGE R. R.  GAME OF THRONES, A [SONG OF ICE AND FIRE 1]  C/P:2,P,H
MARTIN, GEORGE R. R.  ICE DRAGON, THE  H
MARTIN, GEORGE R. R.  JOHN W. CAMPBELL AWARDS 5, THE  HA,LPA,SFWA
MARTIN, GEORGE R. R.  JOKERS WILD [@WILD CARDS #3]  C/PA,PA
MARTIN, GEORGE R. R.  JOKERTOWN SHUFFLE [@WILD CARDS #9]  C/PA,PA
MARTIN, GEORGE R. R.  MARKED CARDS [@WILD CARDS NEW CYCLE 2]  PA
MARTIN, GEORGE R. R.  NEW VOICES 1  C/PA,PA
MARTIN, GEORGE R. R.  NEW VOICES 2  C/PA,PA,SFWA-P
MARTIN, GEORGE R. R.  NEW VOICES 3  C/PA,PA
MARTIN, GEORGE R. R.  NEW VOICES 4  C/PA,PA
MARTIN, GEORGE R. R.  NEW VOICES IN SCIENCE FICTION  HA
MARTIN, GEORGE R. R.  NIGHT VISIONS 3  SR-HA
MARTIN, GEORGE R. R.  NIGHTFLYERS  BS5,SFWA-P,LP,C/P:2

The MITSFS Pinkdex  September 24, 2007  Page 253
AUTHOR  TITLE  CODE
MARTIN, GEORGE R. R.  ONE-EYED JACKS [@WILD CARDS #8]  PA,C/PA
MARTIN, GEORGE R. R.  PORTRAITS OF HIS CHILDREN  H,P
MARTIN, GEORGE R. R.  QUARTET  SR-H
MARTIN, GEORGE R. R.  SANDKINGS  C/P:2,P,H
MARTIN, GEORGE R. R.  SONG FOR LYA AND OTHER STORIES, A  C/P:2,P
MARTIN, GEORGE R. R.  SONGS OF STARS AND SHADOWS  C/P,P
MARTIN, GEORGE R. R.  STORM OF SWORDS, A [SONG OF ICE AND FIRE 3]  H:2,P
MARTIN, GEORGE R. R.  STORM OF SWORDS, A (BRITISH) [SONG OF ICE AND FIRE 3]  H
MARTIN, GEORGE R. R.  TUF VOYAGING  C/P,P,H
MARTIN, GEORGE R. R.  WILD CARDS [@WILD CARDS #1]  HA,PA
MARTIN, GEORGE R. R./TUTTLE, LISA  WINDHAVEN  H:2,P
MARTIN, GRAHAM DUNSTAN  GIFTWISH  H
MARTIN, GRAHAM DUNSTAN  TIME-SLIP  P
MARTIN, JACK  HALLOWEEN III: SEASON OF THE WITCH  P
MARTIN, JACK  VIDEODROME  P
MARTIN, JOHN  REVOLT ON JUPITER  P
MARTIN, LORI  DARKLING HILLS, THE  C/P,P
MARTIN, MARCIA  TOM CORBETT: A TRIP TO THE MOON [TOM CORBETT]  BLB
MARTIN, MARK O./BENFORD, GREGORY  DARKER GEOMETRY, A [KNOWN SPACE]  P
MARTIN, MICHAEL A./MANGELS, ANDY  CATHEDRAL [@STAR TREK DS9 MISSION GAMMA #3]  C/P,P
MARTIN, MICHAEL A./MANGELS, ANDY  LAST FULL MEASURE [STAR TREK ENTERPRISE]  P
MARTIN, MICHAEL A./MANGELS, ANDY  SUNDERED, THE [@STAR TREK LOST ERA (2298 AD)]  P
MARTIN, MICHAEL A./MANGELS, ANDY  TAKING WING [@STAR TREK TITAN 1]  P

MARTIN, MICHAEL A./MANGELS, ANDY/ KYM, J. NOAH TRILL: UNJOINED [STAR TREK DS9 WORLDS 2]/

BAJOR: FRAGMENTS AND OMENS [STAR TREK DS9 WORLDS 2]

C/P,P

MARTIN, PETER  SUMMER IN THREE THOUSAND  H
MARTIN, PHILIP  INVASION OF THE ORMAZOIDS [DOCTOR WHO]  [C/RAN-P]
MARTIN, PHILIP  MINDWARP [DOCTOR WHO]  P
MARTIN, PHILIP  MISSION TO MAGNUS [DOCTOR WHO]  P
MARTIN, PHILIP  VENGEANCE ON VAROS [DOCTOR WHO]  SR-P,C/P
MARTIN, RUSS  EDUCATION OF JENNIFER PARRISH, THE  SFWA-P
MARTIN, RUSS  OBSESSION OF SALLY WING, THE  SFWA-P
MARTIN, THOMAS K.  AWAKENING, THE [@MAGELORD 1]  P
MARTIN, THOMAS K.  CALL TO ARMS, A [TWO-EDGED SWORD 3]  P
MARTIN, THOMAS K.  MATTER OF HONOR, A [TWO-EDGED SWORD 2]  P
MARTIN, THOMAS K.  TIME OF MADNESS, THE [@MAGELORD 2]  P
MARTIN, THOMAS K.  TWO-EDGED SWORD, A [TWO-EDGED SWORD 1]  P
MARTIN, VALERIE  MARY REILLY  P
MARTINDALE, LEE  SUCH A PRETTY FACE  HA,LPA
MARTINE-BARNES, ADRIENNE  CRYSTAL SWORD, THE  C/P,P
MARTINE-BARNES, ADRIENNE  DRAGON RISES, THE  P
MARTINE-BARNES, ADRIENNE  FIRE SWORD, THE  C/P:2,P
MARTINE-BARNES, ADRIENNE  RAINBOW SWORD, THE  C/P:2,P
MARTINE-BARNES, ADRIENNE  SEA SWORD, THE  C/P,P
MARTINEZ, A. LEE  GIL’S ALL FRIGHT DINER  LP
MARTINEZ, A. LEE  IN THE COMPANY OF OGRES  LP
MARTINEZ, A. LEE  NAMELESS WITCH, A  LP
MARTINI, VIRGILIO  WORLD WITHOUT WOMEN, THE  H
MARTINSON, HARRY  ANIARA  LP
MARUSEK, DAVID  COUNTING HEADS  H
MARUYAMA, MAGOROH/HARKINS, ARTHUR CULTURES BEYOND THE EARTH [MATT CRUSE 1]  R/REF-P
MARVEL COMICS  DOCTOR WHO 1985 SUMMER SPECIAL CLASSIC [DOCTOR WHO]  CX
MARVEL COMICS  DOCTOR WHO YEARBOOK (1991) [DOCTOR WHO]  HA
MARVEL COMICS  MARK OF THE MANDRAGORA, THE [DOCTOR WHO]  CX
MARVIN, SUSAN  CHALET BOUGY-VILLARS  P
MASEFIELD, JOHN  BOX OF DELIGHTS, THE  P
MASEFIELD, JOHN  MIDNIGHT FOLK, THE  LP
MASELLO, ROBERT  VIGIL  P
MASON, DAVID  DEEP GODS, THE  C/P:2,P
MASON, DAVID  KAVIN’S WORLD  C/P,P
MASON, DAVID  RETURN OF KAVIN, THE  C/P:2,P
MASON, DAVID  SHORES OF TOMORROW, THE  C/P:2,P
MASON, DAVID  SORCERER’S SKULL, THE  P
MASON, DOUGLAS R.  DILATION EFFECT  C/P:2,P
MASON, DOUGLAS R.  EIGHT AGAINST UTOPIA  C/P:2,P
MASON, DOUGLAS R.  END BRINGERS, THE  C/P:2,P
MASON, DOUGLAS R.  HORIZON ALPHA  C/P:2,P
MASON, DOUGLAS R.  MATRIX  C/P,P
MASON, DOUGLAS R.  PHAETON CONDITION, THE  H,P
MASON, DOUGLAS R.  RESURRECTION OF ROGER DIMENT, THE  C/P,P
MASON, DOUGLAS R.  RING OF VIOLENCE  C/P,P
MASON, DOUGLAS R.  SATELLITE 54-ZERO  C/P,P
MASON, LISA  ARACHNE  H,SFWA,P
MASON, LISA  CYBERWEB  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 254
AUTHOR  TITLE  CODE
MASON, LISA  GOLDEN NINETIES, THE  C/P:2,P
MASON, LISA  IMPERIUM AFIRE [@PANGAEA #2]  P
MASON, LISA  IMPERIUM WITHOUT END [@PANGAEA #1]  P
MASON, LISA  SUMMER OF LOVE  LP
MASON, ROBERT  WEAPON  C/P:2,P
MASSA, JACK  MOONCROW  P
MASSIE, ELIZABETH  POWER OF PERSUASION [BUFFY]  P

MASSIE-FERCH, KATHLEEN M./ GREENBERG, MARTIN H./ GILLIAM, RICHARD ANCIENT ENCHANTRESSES

PA

MASSIE-FRENCH, KATHLEEN M./ GREENBERG, MARTIN H.

WARRIOR ENCHANTRESSES

C/PA,PA

MASSON, DAVID I.  CALTRAPS OF TIME, THE  SR-P

MASTER PUBLICATIONS PHANTOM HORDE [PERRY RHODAN 137]/

STAR MAN #1 THROUGH #5 [PERRY RHODAN 137]

SR-LP

MASTERMAN, WALTER S.  GREEN TOAD, THE  H
MASTERS, J. D.  JAGGED STEELE [STEELE 4]  P
MASTERS, J. D.  STEELE [STEELE 1]  P
MASTERTON, GRAHAM  BURNING, THE  SFWA
MASTERTON, GRAHAM  CHARNEL HOUSE  SFWA-P
MASTERTON, GRAHAM  DEATH DREAM  SFWA-P
MASTERTON, GRAHAM  DEVILS OF D-DAY, THE  P
MASTERTON, GRAHAM  DJINN, THE  SFWA-P
MASTERTON, GRAHAM  FAMINE  P
MASTERTON, GRAHAM  MANITOU, THE  P,SFWA-P
MASTERTON, GRAHAM  MASTER OF LIES  H
MASTERTON, GRAHAM  MIRROR  SFWA-P
MASTERTON, GRAHAM  NIGHT PLAGUE  SFWA-P
MASTERTON, GRAHAM  NIGHT WARRIORS  SFWA-P
MASTERTON, GRAHAM  PLAGUE  P
MASTERTON, GRAHAM  REVENGE OF THE MANITOU  P
MASTERTON, GRAHAM  SCARE CARE  C/PA,PA,SFWA-P
MASTERTON, GRAHAM  WALKERS  P
MATHESON, RICHARD  BID TIME RETURN  SR-P
MATHESON, RICHARD  COLLECTED STORIES, VOLUME ONE  LP
MATHESON, RICHARD  COLLECTED STORIES, VOLUME TWO  LP
MATHESON, RICHARD  EARTHBOUND  P
MATHESON, RICHARD  HELL HOUSE  P
MATHESON, RICHARD  I AM LEGEND  H:2,SR-P
MATHESON, RICHARD  INCREDIBLE SHRINKING MAN, THE  P
MATHESON, RICHARD  RIDE THE NIGHTMARE  P
MATHESON, RICHARD/MAINHARDT, RICIA  ROBERT BLOCH: APPRECIATIONS OF THE MASTER  R/REF-P
MATHESON, RICHARD  SHOCK!  C/P:2,P
MATHESON, RICHARD  SHOCK II  C/P:2,P
MATHESON, RICHARD  SHOCK III  C/P:2,P
MATHESON, RICHARD  SHOCK WAVES  C/P,P
MATHESON, RICHARD  SHORES OF SPACE, THE  C/P:2,P
MATHESON, RICHARD  SHRINKING MAN, THE  C/P:2,P
MATHESON, RICHARD  STIR OF ECHOES, A  C/P,P
MATHESON, RICHARD  TERZO DAL SOLE (ITALIAN)  FP
MATHESON, RICHARD  THIRD FROM THE SUN  C/P:2,P
MATHESON, RICHARD  WHAT DREAMS MAY COME  P
MATHESON, RICHARD CHRISTIAN  SCARS AND OTHER DISTINGUISHING MARKS  SFWA-P
MATHEWS, BASIL  WE FIGHT FOR THE FUTURE  [C/RAN-P]
MATHEWS, RICHARD  ALDISS UNBOUND  R/REF-P
MATHEWS, RICHARD  CLOCKWORK UNIVERSE OF ANTHONY BURGESS, THE  R/REF-P

MATHEWS, RICHARD LIGHTNING FROM A CLEAR SKY:

TOLKIEN, THE TRILOGY, AND THE SILMARILLION

R/REF-P

MATHEWS, RICHARD WORLDS BEYOND THE WORLD:

THE FANTASTIC VISION OF WILLIAM MORRIS

R/REF-P

MATHEWS, SUSAN/RUEMMLER, J. D.  MINES OF MORIA  P
MATSON, NORMAN  BATS IN THE BELFRY  H
MATSON, NORMAN  DOCTOR FOGG  H
MATTHEWS, JOHN  BOOK OF ARTHUR, THE  H
MATTHEWS, SUSAN R.  ANGEL OF DESTRUCTION [JUDICIARY]  P
MATTHEWS, SUSAN R.  AVALANCHE SOLDIER  C/P:2,P
MATTHEWS, SUSAN R.  COLONY FLEET  C/P,P
MATTHEWS, SUSAN R.  DEVIL AND DEEP SPACE, THE [JUDICIARY 4]  P
MATTHEWS, SUSAN R.  EXCHANGE OF HOSTAGES, AN [JUDICIARY 1]  C/P,P
MATTHEWS, SUSAN R.  HOUR OF JUDGEMENT [JUDICIARY 3]  C/P,P
MATTHEWS, SUSAN R.  PRISONER OF CONSCIENCE [JUDICIARY 2]  C/P,P
MATURIN, CHARLES ROBERT  MELMOTH THE WANDERER  LP
MATZ, MARC  NOCTURNE FOR A DANGEROUS MAN  P
MATZKIN, M.  OUTER FLEET, THE  C/P,P
MAUGHAM, W. SOMERSET  CATALINA  P

The MITSFS Pinkdex  September 24, 2007  Page 255
AUTHOR  TITLE  CODE
MAX, NICHOLAS  PRESIDENT MCGOVERN’S FIRST TERM  H
MAXEY, JAMES  NOBODY GETS THE GIRL  LP
MAXIM, JOHN  PLATFORMS  P,SFWA-P
MAXIM, JOHN R.  TIME OUT OF MIND  C/P,SFWA-P,P,H
MAXWELL, ANN  CHANGE  C/P,P
MAXWELL, ANN  DANCER’S ILLUSION  C/P:2,P
MAXWELL, ANN  DANCER’S LUCK  C/P:2,P
MAXWELL, ANN  DEAD GOD DANCING, A  C/P,P
MAXWELL, ANN  FIRE DANCER  C/P:2,P
MAXWELL, ANN  JAWS OF MENX, THE  C/P:2,P
MAXWELL, ANN  NAME OF A SHADOW  C/P:2,P
MAXWELL, ANN  SINGER ENIGMA, THE  C/P:2,P
MAXWELL, ANN  TIMESHADOW RIDER  C/P:2,SFWA-P,P
MAXWELL, JOHN C.  WORLD MAKERS, THE  H
MAXWELL, LISA  BERSERKER’S HORSE, THE  P
MAXXE, ROBERT  ARCADE  H
MAY, JULIAN  ADVERSARY, THE [PLIOCENE EXILE 4]  H:2,P
MAY, JULIAN  BLOOD TRILLIUM [TRILLIUM 2]  H:2,P
MAY, JULIAN  CONQUEROR’S MOON [BOREAL MOON 1]  H
MAY, JULIAN  DIAMOND MASK [GALACTIC MILIEU 2]  C/P,P,H
MAY, JULIAN  GOLDEN TORC, THE [PLIOCENE EXILE 2]  C/P,P,H
MAY, JULIAN  INTERVENTION [INTERVENTION 1,2]  H:2,SFWA
MAY, JULIAN  IRONCROWN MOON [BOREAL MOON 2]  H
MAY, JULIAN  JACK THE BODILESS [GALACTIC MILIEU 1]  H:2,P
MAY, JULIAN  MAGNIFICAT [GALACTIC MILIEU 3]  C/P,P,H
MAY, JULIAN  MANY-COLORED LAND, THE [PLIOCENE EXILE 1]  C/P,P,H

MAY, JULIAN MANY-COLORED LAND, THE [PLIOCENE EXILE 1,2]/

GOLDEN TORC, THE [PLIOCENE EXILE 1,2]

H:2

MAY, JULIAN  METACONCERT, THE [@INTERVENTION #2]  C/P:2,P
MAY, JULIAN  NONBORN KING, THE [PLIOCENE EXILE 3]  C/P:2,SFWA,P

MAY, JULIAN NONBORN KING, THE [PLIOCENE EXILE 3,4]/ ADVERSARY, THE [PLIOCENE EXILE 3,4]

H:2

MAY, JULIAN  ORION ARM [RAMPART WORLDS 2]  C/P,P
MAY, JULIAN  PERSEUS SPUR [RAMPART WORLDS 1]  C/P,P

MAY, JULIAN PLIOCENE COMPANION, A R/REF-H,C/REF-H:2,

C/REF-P

MAY, JULIAN  SAGITTARIUS WHORL [RAMPART WORLDS 3]  C/P,P
MAY, JULIAN  SKY TRILLIUM [TRILLIUM]  H,P
MAY, JULIAN  SURVEILLANCE, THE [@INTERVENTION #1]  C/P:2,P
MAYHAR, ARDATH  EXILE ON VLAHIL  H,SFWA
MAYHAR, ARDATH  GOLDEN DREAM: A FUZZY ODYSSEY [FUZZIES]  C/P,LP
MAYHAR, ARDATH  HOW THE GODS WOVE IN KYRANNON  C/P,SFWA,P
MAYHAR, ARDATH  KHI TO FREEDOM  C/P,P
MAYHAR, ARDATH  LORDS OF THE TRIPLE MOONS  C/P:2,SFWA,P,H
MAYHAR, ARDATH  MAKRA CHORIA  SFWA
MAYHAR, ARDATH/FORTIER, RON  MONKEY STATION  P
MAYHAR, ARDATH  RUNES OF THE LYRE  P
MAYHAR, ARDATH  RUNES OF THE LYRE, THE  H
MAYHAR, ARDATH  SAGA OF GRITTEL SUNDOTHA, THE  H
MAYHAR, ARDATH  SEEKERS OF SHAR-NUHN, THE  H,SFWA,P
MAYHAR, ARDATH  SOUL-SINGER OF TYRNOS  C/P,P,H
MAYHAR, ARDATH  SWORD AND THE DAGGER, THE [@BATTLETECH]  P
MAYHAR, ARDATH/FORTIER, RON  TRAIL OF THE SEAHAWKS  P
MAYHAR, ARDATH  WARLOCK’S GIFT  H,SFWA
MAYHAR, ARDATH  WORLD ENDS IN HICKORY HOLLOW, THE  H,SFWA
MAYHEW, VIC/LONG, DOUG  FIREBALL  P
MAYNARD, JEFF  STAR TREK MAPS  A
MAYNARD, RICHARD  QUIET PLACE, THE  P
MAYNARD, RICHARD  RETURN, THE  P
MAYNE, WILLIAM  GHOSTS  HA

MAYO, CLARK KURT VONNEGUT:

THE GOSPEL FROM OUTER SPACE OR YES WE HAVE NO NIRVANAS

R/REF-P,C/REF-P

MAZER, NORMA FOX  SUPERGIRL  C/P,P
MAZZEO, HENRY  HAUNTINGS: TALES OF THE SUPERNATURAL  HA
MCALEER, NEIL  ARTHUR C. CLARKE  R/REF-H
MCALEER, NEIL  EARTHLOVE  LP
MCALLISTER, BRUCE  DREAM BABY  LP
MCALLISTER, BRUCE  HUMANITY PRIME  C/P:2,P
MCALLISTER, M. I.  URCHIN OF THE RIDING STARS [@MISTMANTLE #1]  H
MCALLISTER, P. K.  MAIA’S VEIL [CLOUDSHIPS OF ORION 2]  P
MCALLISTER, P. K.  ORION’S DAGGER [CLOUDSHIPS OF ORION 3]  P
MCALLISTER, P. K.  SIDURI’S NET [CLOUDSHIPS OF ORION 1]  P
MCARTHUR, MAXINE  TIME FUTURE [COMMANDER HALLEY 1]  P
MCARTHUR, MAXINE  TIME PAST [COMMANDER HALLEY 2]  P

The MITSFS Pinkdex  September 24, 2007  Page 256
AUTHOR  TITLE  CODE
MCAULEY, PAUL J.  ANCIENT OF DAYS [@CONFLUENCE #2]  C/P,P,H
MCAULEY, PAUL J.  CHILD OF THE RIVER [@CONFLUENCE #1]  H,P
MCAULEY, PAUL J.  CONFLUENCE [CONFLUENCE 1,2,3]  H
MCAULEY, PAUL J.  ETERNAL LIGHT  C/P,SFWA,P,H
MCAULEY, PAUL J.  EYE OF THE TYGER, THE [DOCTOR WHO]  SR-H
MCAULEY, PAUL J.  FAIRYLAND  C/P,LP,H
MCAULEY, PAUL J.  FOUR HUNDRED BILLION STARS  C/P,P
MCAULEY, PAUL J.  INVISIBLE COUNTRY, THE  P
MCAULEY, PAUL J.  MAKING HISTORY  GZD07306
MCAULEY, PAUL J.  OF THE FALL  P
MCAULEY, PAUL J.  PASQUALE’S ANGEL  P
MCAULEY, PAUL J.  RED DUST  C/P:2,P
MCAULEY, PAUL J.  SECRET OF LIFE, THE  H:2
MCAULEY, PAUL J.  SHRINE OF STARS [@CONFLUENCE #3]  H
MCAULEY, PAUL J.  WHITE DEVILS  P
MCAULEY, PAUL J.  WHOLE WIDE WORLD  H
MCBAIN, GORDON  PATH OF EXOTERRA, THE  P
MCBAIN, GORDON  QUEST OF THE DAWNSTAR  P
MCBIRNIE, W. S.  ANTICHRIST  [C/RAN-P]
MCCAFFREY, ANNE/BALL, MARGARET  ACORNA [ACORNA 1]  H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S PEOPLE [ACORNA 3]

H,P

MCCAFFREY, ANNE/BALL, MARGARET  ACORNA’S QUEST [ACORNA 2]  C/P,P,H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S REBELS [ACORNA 6]

H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S REBELS (PROOF) [ACORNA 6]

SR-LP

MCCAFFREY, ANNE  ACORNA’S SEARCH [ACORNA 5]  H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S TRIUMPH [ACORNA 7]

P

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S TRIUMPH (PROOF) [ACORNA 7]

SR-LP

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN ACORNA’S WORLD [ACORNA 4]

H

MCCAFFREY, ANNE  ALCHEMY AND ACADEME  C/PA,PA
MCCAFFREY, ANNE  ALL THE WEYRS OF PERN [PERN]  H:2,P
MCCAFFREY, ANNE  BLACK HORSES FOR THE KING  H,LP
MCCAFFREY, ANNE  BOOK OF FREEDOMS, THE [BOTANY 1,2,3]  H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN CHANGELINGS [TWINS OF PETAYBEE 1]

H:2

MCCAFFREY, ANNE/STIRLING, S. M.  CITY WHO FOUGHT, THE [SHELLPEOPLE 4]  H:2,P
MCCAFFREY, ANNE  COELURA, THE  C/P,SFWA-P,P,H,SR-H
MCCAFFREY, ANNE/NYE, JODY LYNN  CRISIS ON DOONA [DOONA]  C/P:2,SFWA-P,P
MCCAFFREY, ANNE  CRYSTAL LINE [CRYSTAL SINGER 3]  H:2,P
MCCAFFREY, ANNE  CRYSTAL SINGER [CRYSTAL SINGER 1]  C/P:2,P,H:3
MCCAFFREY, ANNE  DAMIA [ROWAN 2]  C/P,SFWA-P:2,P,H
MCCAFFREY, ANNE  DAMIA’S CHILDREN [ROWAN 3]  C/P,SFWA-P,P,H
MCCAFFREY, ANNE/NYE, JODY LYNN  DEATH OF SLEEP, THE [PLANET PIRATES 2]  C/P,P,H
MCCAFFREY, ANNE  DECISION AT DOONA [DOONA]  C/P:2,P
MCCAFFREY, ANNE  DINOSAUR PLANET [DINOSAUR PLANET 1]  C/P:3,P
MCCAFFREY, ANNE  DINOSAUR PLANET SURVIVORS [DINOSAUR PLANET 2]  C/P:2,P
MCCAFFREY, ANNE  DOLPHIN’S BELL, THE [PERN]  H:2
MCCAFFREY, ANNE  DOLPHINS OF PERN, THE [PERN]  H:2,P
MCCAFFREY, ANNE  DRAGONDRUMS [PERN HARPER HALL 3]  H:2,P
MCCAFFREY, ANNE  DRAGONFLIGHT [PERN DRAGONRIDERS 1]  C/P:2,P
MCCAFFREY, ANNE  DRAGONQUEST [PERN DRAGONRIDERS 2]  C/P:2,P
MCCAFFREY, ANNE  DRAGONRIDERS OF PERN [PERN DRAGONRIDERS 1,2,3]  H:2
MCCAFFREY, ANNE/MCCAFFREY, TODD  DRAGON’S FIRE [PERN]  H
MCCAFFREY, ANNE/MCCAFFREY, TODD  DRAGON’S KIN [PERN]  H
MCCAFFREY, ANNE/MCCAFFREY, TODD  DRAGON’S KIN (PROOF) [PERN]  SR-LP
MCCAFFREY, ANNE  DRAGONSDAWN [PERN]  H:2,P
MCCAFFREY, ANNE  DRAGONSEYE [PERN]  H:2,P
MCCAFFREY, ANNE  DRAGONSINGER [PERN HARPER HALL 2]  H:2,P
MCCAFFREY, ANNE  DRAGONSONG [PERN HARPER HALL 1]  H:2,P
MCCAFFREY, ANNE  EXCHANGE OF GIFTS, AN  H,LP
MCCAFFREY, ANNE  FIRST FALL [PERN]  C/P,P,H:2

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN FIRST WARNING [ACORNA’S CHILDREN 1]

H

MCCAFFREY, ANNE  FREEDOM’S CHALLENGE [BOTANY 3]  H:2,P
MCCAFFREY, ANNE  FREEDOM’S CHOICE [BOTANY 2]  C/P:2,P,H
MCCAFFREY, ANNE  FREEDOM’S LANDING [BOTANY 1]  H:3,P
MCCAFFREY, ANNE  FREEDOM’S RANSOM [BOTANY 4]  H:2
MCCAFFREY, ANNE/MOON, ELIZABETH  GENERATION WARRIORS [PLANET PIRATES 3]  C/P:2,SFWA-P,P
MCCAFFREY, ANNE  GET OFF THE UNICORN  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 257
AUTHOR  TITLE  CODE
MCCAFFREY, ANNE  GIFT OF DRAGONS, A [PERN]  H
MCCAFFREY, ANNE  GIRL WHO HEARD DRAGONS, THE  H:2,SFWA-P,P
MCCAFFREY, ANNE  HARPER HALL OF PERN, THE [PERN HARPER HALL 1,2,3]  H:2
MCCAFFREY, ANNE  IF WISHES WERE HORSES  H
MCCAFFREY, ANNE  IRETA ADVENTURE, THE [DINOSAUR PLANET 1,2]  H
MCCAFFREY, ANNE  KILLASHANDRA [CRYSTAL SINGER 2]  C/P,P,H,SFWA
MCCAFFREY, ANNE  KILTERNAN LEGACY, THE  C/P,P
MCCAFFREY, ANNE  LADY, THE  H,P
MCCAFFREY, ANNE  LYON’S PRIDE  C/P,P,H

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN MAELSTROM [TWINS OF PETAYBEE 2]

H

MCCAFFREY, ANNE  MARK OF MERLIN, THE  C/P,P
MCCAFFREY, ANNE  MASTERHARPER OF PERN, THE [PERN]  H:2,P
MCCAFFREY, ANNE  MORETA: DRAGONLADY OF PERN [PERN]  H:2,P
MCCAFFREY, ANNE  NERILKA’S STORY [PERN]  H:2,P
MCCAFFREY, ANNE  NIMISHA’S SHIP  C/P:2,P,H
MCCAFFREY, ANNE  NO ONE NOTICED THE CAT  H:2
MCCAFFREY, ANNE/BALL, MARGARET  PARTNERSHIP [SHELLPEOPLE 2]  C/P,SFWA-P,P,H
MCCAFFREY, ANNE/BALL, MARGARET  PARTNERSHIP (EXCERPT) [SHELLPEOPLE 2]  SR-LP
MCCAFFREY, ANNE  PEGASUS IN FLIGHT [PEGASUS 2]  H:2,SR-H,P
MCCAFFREY, ANNE  PEGASUS IN SPACE [PEGASUS 3]  H:2,P

MCCAFFREY, ANNE/MOON, ELIZABETH/ NYE, JODY LYNN PLANET PIRATES, THE [PLANET PIRATES 1,2,3]

LP

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN POWER LINES [PETAYBEE 2]

H:2,P

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN POWER PLAY [PETAYBEE 3]

H:2,P

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN POWERS THAT BE [PETAYBEE 1]

H:2,P

MCCAFFREY, ANNE  RENEGADES OF PERN, THE [PERN]  H:2,P
MCCAFFREY, ANNE  RESCUE RUN [PERN]  H:2
MCCAFFREY, ANNE  RESTOREE  C/P:2,P
MCCAFFREY, ANNE  RING OF FEAR  C/P,P
MCCAFFREY, ANNE  ROWAN, THE [ROWAN 1]  C/P,SFWA-P,H:2,P,SFWA
MCCAFFREY, ANNE/MOON, ELIZABETH  SASSINAK [PLANET PIRATES 1]  C/P:2,SFWA-P,P

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN SECOND WAVE [ACORNA’S CHILDREN 2]

H

MCCAFFREY, ANNE  SHIP WHO SANG, THE [SHELLPEOPLE 1]  C/P,P,H
MCCAFFREY, ANNE/LACKEY, MERCEDES  SHIP WHO SEARCHED, THE [SHELLPEOPLE 3]  C/P,P,H
MCCAFFREY, ANNE/NYE, JODY LYNN  SHIP WHO WON, THE [SHELLPEOPLE 5]  C/P,P,H
MCCAFFREY, ANNE  SKIES OF PERN, THE [PERN]  H,P

MCCAFFREY, ANNE/ SCARBOROUGH, ELIZABETH ANN SPACE OPERA

C/PA,PA

MCCAFFREY, ANNE  STITCH IN SNOW  P
MCCAFFREY, ANNE  SURVIVORS, THE [DINOSAUR PLANET 2]  P
MCCAFFREY, ANNE  TIME WHEN, A  SR-H
MCCAFFREY, ANNE  TO RIDE PEGASUS [PEGASUS 1]  C/P:2,P
MCCAFFREY, ANNE  TOWER AND THE HIVE, THE [ROWAN 4]  H,P
MCCAFFREY, ANNE/NYE, JODY LYNN  TREATY AT DOONA [DOONA]  P
MCCAFFREY, ANNE  WHITE DRAGON, THE [PERN DRAGONRIDERS 3]  C/P,P,H
MCCAFFREY, ANNE  WINGS OF PEGASUS, THE [PEGASUS 1,2]  H
MCCAFFREY, ANNE  YEAR OF THE LUCY, THE  P
MCCAFFREY, TODD  DRAGONHOLDER  R/REF-H,C/REF-H
MCCAFFREY, TODD  DRAGONSBLOOD [PERN]  H
MCCALLISTER, BRUCE  DREAM BABY  P
MCCAMMON, ROBERT R.  BAAL  P
MCCAMMON, ROBERT R.  BLUE WORLD  C/P:2,P
MCCAMMON, ROBERT R.  BOY’S LIFE  P
MCCAMMON, ROBERT R.  MYSTERY WALK  C/P,P
MCCAMMON, ROBERT R.  NIGHT BOAT, THE  P
MCCAMMON, ROBERT R.  STINGER  P
MCCAMMON, ROBERT R.  SWAN SONG  C/P,P
MCCAMMON, ROBERT R.  THEY THIRST  C/P,P
MCCAMMON, ROBERT R.  USHER’S PASSING  P
MCCAMMON, ROBERT R.  WOLF’S HOUR, THE  SR-P,C/P:2
MCCANN, EDSON  PREFERRED RISK  H:2,P
MCCARTHY, CORMAC  ROAD, THE  H
MCCARTHY, JUSTIN HUNTLY  DRYAD, THE  H
MCCARTHY, KEVIN/SILVA, DAVID  INTO THE DARKNESS [@FAMILY #2]  P
MCCARTHY, KEVIN/SILVA, DAVID  SPECIAL EFFECTS [@FAMILY #1]  P
MCCARTHY, KEVIN/GORMAN, ED  "THEY’RE HERE..." [INVASION OF THE BODY SNATCHERS]  R/REF-P

MCCARTHY, PATRICK A./ELKINS, CHARLES/ GREENBERG, MARTIN H.

LEGACY OF OLAF STAPLEDON, THE

R/REF-H

MCCARTHY, SHAWNA  ISAAC ASIMOV’S ALIENS AND OUTWORLDERS  C/PA,LPA

The MITSFS Pinkdex  September 24, 2007  Page 258
AUTHOR  TITLE  CODE
MCCARTHY, SHAWNA  ISAAC ASIMOV’S FANTASY!  HA,SFWA
MCCARTHY, SHAWNA  ISAAC ASIMOV’S SPACE OF HER OWN  SFWA
MCCARTHY, SHAWNA  ISAAC ASIMOV’S SPACE OF HER OWN (ABRIDGED)  PA
MCCARTHY, WIL  AGGRESSOR SIX [AGGRESSOR SIX 1]  C/P,P
MCCARTHY, WIL  BLOOM  C/P,P
MCCARTHY, WIL  COLLAPSIUM, THE [QUEENDOM OF SOL 1]  H
MCCARTHY, WIL  FALL OF SIRIUS, THE [AGGRESSOR SIX 2]  P
MCCARTHY, WIL  FLIES FROM THE AMBER  P
MCCARTHY, WIL  LOST IN TRANSMISSION [QUEENDOM OF SOL 3]  C/P,P
MCCARTHY, WIL  MONARCHS OF SOL, THE [QUEENDOM OF SOL 1,2]  H
MCCARTHY, WIL  MURDER IN THE SOLID STATE  C/P:2,P
MCCARTHY, WIL  MURDER IN THE SOLID STATE (PROOF)  SR-LP

MCCARTHY, WIL/GREENBERG, MARTIN H./ HELFERS, JOHN ONCE UPON A GALAXY

PA

MCCARTHY, WIL  TO CRUSH THE MOON [QUEENDOM OF SOL 4]  P
MCCARTHY, WIL  WELLSTONE, THE [QUEENDOM OF SOL 2]  P
MCCARTY, DENNIS  ACROSS THE THLASSA MEY [THLASSA MEY 4]  C/P,P
MCCARTY, DENNIS  BIRTH OF THE BLADE, THE [THLASSA MEY 5]  C/P,P
MCCARTY, DENNIS  FLIGHT TO THLASSA MEY [THLASSA MEY 1]  C/P,P
MCCARTY, DENNIS  LORDS OF THLASSA MEY [THLASSA MEY 3]  C/P,P
MCCARTY, DENNIS  WARRIORS OF THLASSA MEY [THLASSA MEY 2]  P
MCCAULEY, KIRBY  BEYOND MIDNIGHT  PA
MCCAULEY, KIRBY  DARK FORCES  SR-HA,C/PA,PA
MCCAULEY, KIRBY  FRIGHTS  PA
MCCAULEY, KIRBY  NIGHT CHILLS  C/PA,PA
MCCAY, BILL/FLOOD, ELOISE  CHAINS OF COMMAND [STAR TREK TNG]  C/P:3,P
MCCAY, BILL  CROSSOVER [@RIFTWORLD 1]  C/P:2,SFWA-P,P
MCCAY, BILL  REBELLION [STARGATE]  C/P:2,P
MCCAY, BILL  RECONNAISSANCE [STARGATE]  P
MCCAY, BILL  RESISTANCE [STARGATE]  C/P,P
MCCAY, BILL  RETALIATION [STARGATE]  C/P,P
MCCAY, BILL  RETRIBUTION [STARGATE]  P
MCCAY, BILL  VILLAINS [@RIFTWORLD 2]  C/P,P
MCCLARY, THOMAS CALVERT  REBIRTH — WHEN EVERYONE FORGOT  SR-P
MCCLARY, THOMAS CALVERT  THREE THOUSAND YEARS  SR-H,P,DD176
MCCOLLUM, MICHAEL  ANTARES DAWN [ANTARES 1]  C/P:2,P
MCCOLLUM, MICHAEL  ANTARES PASSAGE [ANTARES 2]  C/P,P
MCCOLLUM, MICHAEL  ANTARES VICTORY [ANTARES 3]  LP
MCCOLLUM, MICHAEL  CLOUDS OF SATURN, THE  C/P:2,P
MCCOLLUM, MICHAEL  GIBRALTAR EARTH  LP
MCCOLLUM, MICHAEL  GREATER INFINITY, A  C/P:2,P
MCCOLLUM, MICHAEL  GRIDLOCK AND OTHER STORIES  LP
MCCOLLUM, MICHAEL  LIFE PROBE [LIFE PROBE 1]  C/P:2,P
MCCOLLUM, MICHAEL  PROCYON’S PROMISE [LIFE PROBE 2]  C/P:2,P
MCCOLLUM, MICHAEL  SAILS OF TAU CETI, THE  C/P:2,P
MCCOLLUM, MICHAEL  THUNDER STRIKE!  C/P:2,P
MCCOMAS, ANNETTE PELTZ  EUREKA YEARS, THE  PA
MCCOMAS, J. FRANCIS  SPECIAL WONDER 1  PA
MCCOMAS, J. FRANCIS  SPECIAL WONDER 2  PA,C/PA

MCCOMAS, J. FRANCIS SPECIAL WONDER: THE ANTHONY BOUCHER MEMORIAL ANTHOLOGY OF

FANTASY AND SCIENCE FICTION

HA

MCCONNELL, ASHLEY  BOOK OF THE DEAD [@ANGEL]  P
MCCONNELL, ASHLEY  COURTS OF SORCERY, THE [DEMON WARS 3]  P
MCCONNELL, ASHLEY  FIRST AMENDMENT, THE [@STARGATE SG-1 (MCCONNELL) 3]  P
MCCONNELL, ASHLEY  FOUNTAINS OF MIRLACCA, THE [DEMON WARS 1]  P
MCCONNELL, ASHLEY  ITINERANT EXORCIST, THE [DEMON WARS 2]  P
MCCONNELL, ASHLEY  MORPHEUS FACTOR, THE [@STARGATE SG-1 (MCCONNELL) 4]  P
MCCONNELL, ASHLEY  PRELUDE [QUANTUM LEAP]  C/P,P
MCCONNELL, ASHLEY  PRICE YOU PAY, THE [@STARGATE SG-1 (MCCONNELL) 2]  P
MCCONNELL, ASHLEY  QUANTUM LEAP [QUANTUM LEAP]  C/P,SFWA-P,P
MCCONNELL, ASHLEY  RANDOM MEASURES [@QUANTUM LEAP]  C/P,P
MCCONNELL, ASHLEY  SCIMITAR [HIGHLANDER]  P
MCCONNELL, ASHLEY  STARGATE SG-1 [@STARGATE SG-1 (MCCONNELL) 1]  C/P,P
MCCONNELL, ASHLEY/KOOGLER, DORI  THESE OUR ACTORS [@BUFFY]  C/P,P
MCCONNELL, ASHLEY  TOO CLOSE FOR COMFORT [@QUANTUM LEAP]  C/P:2,P
MCCONNELL, ASHLEY  UNEARTHED  P
MCCONNELL, ASHLEY  WALL, THE [@QUANTUM LEAP]  C/P,SFWA-P:2,P
MCCONNELL, JAMES V.  WORM RE-TURNS, THE  [C/RAN-H]
MCCONNOR, VINCENT  I AM VIDOCQ  SFWA-P

MCCORMACK, UNA/JARMAN, HEATHER CARDASSIA: LOTUS FLOWER, THE [STAR TREK DS9 WORLDS 1]/

ANDOR: PARADIGM [STAR TREK DS9 WORLDS 1]

C/P,P

MCCORMACK, UNA  HOLLOW MEN [@STAR TREK DS9]  P
MCCOY, GLEN  TIMELASH [DOCTOR WHO]  SR-P,C/P
MCCRAY, CHEYENNE  FORBIDDEN MAGIC  P

The MITSFS Pinkdex  September 24, 2007  Page 259
AUTHOR  TITLE  CODE
MCCRUMB, SHARYN  BIMBOS OF THE DEATH SUN  C/P:2,P
MCCRUMB, SHARYN  ZOMBIES OF THE GENE POOL  P
MCCULLOUGH, COLLEEN  CREED FOR THE THIRD MILLENNIUM, A  H,P
MCCULLOUGH, KELLY  WEBMAGE  P
MCCURDY, J. FITSGERALD  BURNING CROWN, THE [SERPENT’S EGG 2]  LP
MCCURDY, J. FITSGERALD  SERPENT’S EGG, THE [SERPENT’S EGG 1]  LP
MCCUTCHEON, GEORGE B.  GRAUSTARK  H:2
MCCUTCHEON, GEORGE B.  PRINCE OF GRAUSTARK, THE  H
MCCUTCHEON, GEORGE B.  TRUXTON KING  H
MCDANIEL, DAVID  ARSENAL OUT OF TIME, THE  C/P:2,P
MCDANIEL, DAVID  DAGGER AFFAIR, THE [@MAN FROM UNCLE #4]  C/P:2,P
MCDANIEL, DAVID  HOLLOW CROWN AFFAIR, THE [@MAN FROM UNCLE #17]  C/P:2,P
MCDANIEL, DAVID  MONSTER WHEEL AFFAIR, THE [@MAN FROM UNCLE #8]  C/P:2,P
MCDANIEL, DAVID  PRISONER #2, THE [PRISONER]  C/P,P
MCDANIEL, DAVID  RAINBOW AFFAIR, THE [@MAN FROM UNCLE #13]  C/P,P
MCDANIEL, DAVID  UTOPIA AFFAIR, THE [@MAN FROM UNCLE #15]  C/P,P
MCDANIEL, DAVID  VAMPIRE AFFAIR, THE [@MAN FROM UNCLE #6]  C/P:2,P
MCDANIEL, LURLENE  BATTLE OF ZORN, THE  P
MCDERMOTT, WILL/RENNIE, GORDON  BLOOD ROYAL [@NECROMUNDA]  P
MCDERMOTT, WILL  CARDINAL CRIMSON [@NECROMUNDA]  P
MCDERMOTT, WILL  JUDGMENT [@MAGIC: THE GATHERING ODYSSEY 3]  P
MCDERMOTT, WILL  LASGUN WEDDING [@NECROMUNDA]  P
MCDERMOTT, WILL  MOONS OF MIRRODIN, THE [@MAGIC: THE GATHERING MIRRODIN 1]  P
MCDEVITT, JACK  ANCIENT SHORES  C/P,P,H:2
MCDEVITT, JACK  CHINDI [PRISCILLA HUTCHINS 3]  H:2
MCDEVITT, JACK  DEEPSIX [PRISCILLA HUTCHINS 2]  H
MCDEVITT, JACK  ENGINES OF GOD, THE [PRISCILLA HUTCHINS 1]  C/P:2,P
MCDEVITT, JACK  ETERNITY ROAD  H
MCDEVITT, JACK  HELLO OUT THERE  LP
MCDEVITT, JACK  HERCULES TEXT, THE  C/P:2,P
MCDEVITT, JACK  INFINITY BEACH  H:2
MCDEVITT, JACK  MOONFALL  C/P,P
MCDEVITT, JACK  ODYSSEY  H
MCDEVITT, JACK  OMEGA [PRISCILLA HUTCHINS 4]  H
MCDEVITT, JACK  OUTBOUND  H
MCDEVITT, JACK  POLARIS [ALEX BENEDICT 2]  H
MCDEVITT, JACK  SEEKER [ALEX BENEDICT 3]  H
MCDEVITT, JACK  STANDARD CANDLES  H
MCDEVITT, JACK  TALENT FOR WAR, A [ALEX BENEDICT 1]  H,SFWA-P,P
MCDONALD, IAN  ARES EXPRESS  P
MCDONALD, IAN  BRASYL  H
MCDONALD, IAN  BROKEN LAND, THE  C/P,LP,P
MCDONALD, IAN  DESOLATION ROAD  C/P:2,P
MCDONALD, IAN  EMPIRE DREAMS  C/P:2,P
MCDONALD, IAN  EVOLUTION’S SHORE [CHAGA 1]  C/P,P,H
MCDONALD, IAN  KING OF MORNING, QUEEN OF DAY  C/P,P
MCDONALD, IAN  KIRINYA [CHAGA 2]  P
MCDONALD, IAN  OUT ON BLUE SIX  P
MCDONALD, IAN  RIVER OF GODS  LP
MCDONALD, IAN  SCISSORS CUT PAPER WRAP STONE  C/P,P
MCDONALD, IAN  SPEAKING IN TONGUES  P
MCDONALD, IAN  TENDELEO’S STORY [CHAGA]  GZD07305,SR-LP
MCDONALD, IAN  TERMINAL CAFE  P
MCDONALD, JOHN  STRATEGY IN POKER, BUSINESS, AND WAR  [C/RAN-P]
MCDONALD, SANDRA  OUTBACK STARS, THE  H
MCDONALD, STEVEN E.  EVENT HORIZON  P
MCDONALD, STEVEN E.  JANUS SYNDROME, THE  C/P,P
MCDONNELL, DAVID  STAR TREK’S GREATEST GUEST STARS  R/REF-P,C/REF-P

MCDONNELL, DAVID/SAYERS, JOHN/ SMITH, TIM L.

STARLOG: SCIENCE FICTION TRIVIA

[C/RAN-P]

MCDONOUGH, ALEX  DRAGON’S BLOOD [SCORPIO 4]  C/P,P
MCDONOUGH, ALEX  DRAGON’S CLAW [SCORPIO 6]  C/P,P
MCDONOUGH, ALEX  DRAGON’S EYE [SCORPIO 5]  P
MCDONOUGH, ALEX  SCORPIO [SCORPIO 1]  C/P:2,P
MCDONOUGH, ALEX  SCORPIO DESCENDING [SCORPIO 3]  C/P,P
MCDONOUGH, ALEX  SCORPIO RISING [SCORPIO 2]  C/P,P
MCDONOUGH, THOMAS R.  ARCHITECTS OF HYPERSPACE, THE  C/P:2,P
MCDONOUGH, THOMAS R.  MISSING MATTER, THE [NEXT WAVE 3]  C/P,P
MCDOWELL, IAN  MERLIN’S GIFT [MORDRED’S CURSE 2]  P
MCDOWELL, IAN  MORDRED’S CURSE [MORDRED’S CURSE 1]  P
MCDOWELL, MICHAEL  ELEMENTALS, THE  P
MCDOWELL, MICHAEL  FORTUNE, THE [@BLACKWATER #5]  P
MCDOWELL, MICHAEL  WAR, THE [@BLACKWATER #6]  P
MCELROY, JOSEPH  PLUS  LP

AUTHOR  TITLE  CODE
MCENROE, RICHARD S.  FLIGHT OF HONOR [FAR STARS AND FUTURE TIMES]  P
MCENROE, RICHARD S.  PROTEUS  C/PA,PA
MCENROE, RICHARD S.  SHATTERED STARS, THE [FAR STARS AND FUTURE TIMES]  C/P:2,P
MCENROE, RICHARD S.  SKINNER [FAR STARS AND FUTURE TIMES]  C/P:2,P
MCENROE, RICHARD S.  WARRIOR’S BLOOD [BUCK ROGERS (NIVEN & POURNELLE’S) 3]  P
MCENROE, RICHARD S.  WARRIOR’S WORLD [BUCK ROGERS (NIVEN & POURNELLE’S) 4]  P
MCEVOY, SETH  BATTERIES NOT INCLUDED [NOT QUITE HUMAN 1]  P
MCEVOY, SETH  DESTINATION: BRAIN  P
MCEVOY, SETH  ROBOT WORLD [INTERPLANETARY SPY 3]  SCX
MCEVOY, SETH  TERROR AT PLAY [NOT QUITE HUMAN 5]  P
MCFADDEN, EDWARD J. III  BEST OF PIRATE WRITINGS 1, THE  LPA
MCFADDEN, EDWARD J. III  2ND COMING, THE  LPA
MCFEDRIES, A. A.  KING OF THE SUN, THE  P
MCGARRY, MARK J.  BLANK SLATE  P
MCGARRY, MARK J.  SUN DOGS  C/P,P
MCGARRY, TERRY  BINDER’S ROAD, THE [ILLUMINATION 2]  P
MCGARRY, TERRY  ILLUMINATION [ILLUMINATION 1]  H
MCGARRY, TERRY  TRIAD [ILLUMINATION 3]  H
MCGEARY, DUNCAN  ICETOWERS  P
MCGEARY, DUNCAN  SNOWCASTLES  P
MCGEARY, DUNCAN  STAR AXE  P
MCGHAN, BARRY  SCIENCE FICTION AND FANTASY PSEUDONYMS  R/IND
MCGILL, GORDON  ABOMINATION, THE [OMEN 5]  C/P,P
MCGILL, GORDON  ARMAGEDDON 2000 [OMEN 4]  P
MCGILL, GORDON  FINAL CONFLICT, THE [OMEN 3]  P
MCGIRT, DAN  DIRTY WORK [JASON COSMO 3]  C/P,P
MCGIRT, DAN  JASON COSMO [JASON COSMO 1]  C/P,P
MCGIRT, DAN  ROYAL CHAOS  C/P,P

MCGIVERN, WILLIAM P./ MCGIVERN, MAUREEN SEEING, THE

P

MCGOUGH, SCOTT  ASSASSIN’S BLADE [@MAGIC: THE GATHERING LEGENDS TWO 1]  P
MCGOUGH, SCOTT  CHAINER’S TORMENT [@MAGIC: THE GATHERING ODYSSEY 2]  P
MCGOUGH, SCOTT  CHAMPION’S TRIAL [@MAGIC: THE GATHERING LEGENDS TWO 3]  P
MCGOUGH, SCOTT  EMPEROR’S FIST [@MAGIC: THE GATHERING LEGENDS TWO 2]  P
MCGOUGH, SCOTT/DELANEY, JOHN  FUTURE SIGHT [@MAGIC: THE GATHERING TIME SPIRAL #3]  P

MCGOUGH, SCOTT GUARDIAN:SAVIORS OF KAMIGAWA

[@MAGIC: THE GATHERING KAMIGAWA 3]

P

MCGOUGH, SCOTT HERITIC: BETRAYERS OF KAMIGAWA

[@MAGIC: THE GATHERING KAMIGAWA 2]

P

MCGOUGH, SCOTT OUTLAW: CHAMPIONS OF KAMIGAWA

[@MAGIC: THE GATHERING KAMIGAWA 1]

C/P,P

MCGOUGH, SCOTT/SANDERS, TIMOTHY  PLANAR CHAOS [@MAGIC: THE GATHERING TIME SPIRAL #2]  P
MCGOUGH, SCOTT  TIME SPIRAL [@MAGIC: THE GATHERING TIME SPIRAL #1]  P
MCGREW, CHANDLER  DARKENING, THE  P
MCGUIRE, PATRICK L.  RED STARS  R/REF-H
MCGUIRE, WILLIAM  FROM TOLKIEN TO OZ  A
MCHUGH, MAUREEN F.  CHINA MOUNTAIN ZHANG  C/P,P,H
MCHUGH, MAUREEN F.  HALF THE DAY IS NIGHT  C/P:2,P,H
MCHUGH, MAUREEN F.  MISSION CHILD  H:2,P
MCHUGH, MAUREEN F.  NEKROPOLIS  H
MCHUGH, VINCENT  I AM THINKING OF MY DARLING  SR-P,H
MCINTEE, DAVID A.  AUTUMN MIST [DOCTOR WHO EIGHTH DOCTOR]  P
MCINTEE, DAVID A.  BULLET TIME [DOCTOR WHO PAST DOCTORS]  P
MCINTEE, DAVID A.  DARK PATH, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
MCINTEE, DAVID A.  ELEVENTH TIGER, THE [DOCTOR WHO PAST DOCTORS]  P
MCINTEE, DAVID A.  FACE OF THE ENEMY, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
MCINTEE, DAVID A.  FIRST FRONTIER [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
MCINTEE, DAVID A.  LORDS OF THE STORM [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
MCINTEE, DAVID A.  MISSION: IMPRACTICAL [DOCTOR WHO PAST DOCTORS]  C/P,P
MCINTEE, DAVID A.  SANCTUARY [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
MCINTEE, DAVID A.  SHADOW OF WENG-CHIANG, THE [DOCTOR WHO MISSING ADVENTURES]  C/P,P
MCINTEE, DAVID A.  WAGES OF SIN, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
MCINTEE, DAVID A.  WHITE DARKNESS [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
MCINTOSH, FIONA  BLOOD AND MEMORY [@QUICKENING #2]  C/P,LP
MCINTOSH, FIONA  BRIDGE OF SOULS [QUICKENING 3]  LP
MCINTOSH, FIONA  MYRREN’S GIFT [QUICKENING 1]  H,LP
MCINTOSH, FIONA  ODALISQUE [PERCHERON 1]  LP
MCINTOSH, J. T.  BORN LEADER  H:2
MCINTOSH, J. T.  FITTEST, THE  H:2
MCINTOSH, J. T.  FLIGHT FROM REBIRTH  C/P:2,P
MCINTOSH, J. T.  MILLION CITIES, THE  C/P:2,P
MCINTOSH, J. T.  NORMAN CONQUEST 2066  SR-P
MCINTOSH, J. T.  ONE IN THREE HUNDRED  DD113,H:2
MCINTOSH, J. T.  PLANET CALLED UTOPIA, A  P

The MITSFS Pinkdex  September 24, 2007  Page 261
AUTHOR  TITLE  CODE
MCINTOSH, J. T.  RULE OF THE PAGBEASTS, THE  SR-P,C/P
MCINTOSH, J. T.  RULER OF THE WORLD  SR-P,C/P:2
MCINTOSH, J. T.  SIX GATES FROM LIMBO  C/P:2,P
MCINTOSH, J. T.  SNOW WHITE AND THE GIANTS  C/P,P
MCINTOSH, J. T.  SUICIDERS, THE  C/P,P
MCINTOSH, J. T.  THIS IS THE WAY THE WORLD BEGINS  P
MCINTOSH, J. T.  TRANSMIGRATION  C/P:2,P
MCINTOSH, J. T.  200 YEARS TO CHRISTMAS  DF113
MCINTOSH, J. T.  WORLD OUT OF MIND  H:2,P
MCINTOSH, J. T.  WORLDS APART  C/P:2,P
MCINTOSH, NEIL  KEEPERS OF THE FLAME [@WARHAMMER STEFAN KUMANSKY 3]  P
MCINTOSH, NEIL  STAR OF ERENGRAD [@WARHAMMER]  P
MCINTOSH, NEIL  TAINT OF EVIL [@WARHAMMER STEFAN KUMANSKY 2]  P

MCINTYRE, VONDA N./ ANDERSON, SUSAN JANICE AURORA: BEYOND EQUALITY

C/PA,PA

MCINTYRE, VONDA N.  BARBARY  C/P,SFWA-P,H,P,SFWA
MCINTYRE, VONDA N.  BRIDE, THE  P
MCINTYRE, VONDA N.  CRYSTAL STAR, THE [@STAR WARS]  C/P:2,P
MCINTYRE, VONDA N.  DREAMSNAKE  C/P,H:2,P
MCINTYRE, VONDA N.  DUTY, HONOR, REDEMPTION [@STAR TREK]  LP
MCINTYRE, VONDA N.  ENTERPRISE: THE FIRST ADVENTURE [STAR TREK]  C/P:3,P
MCINTYRE, VONDA N.  ENTROPY EFFECT, THE [STAR TREK]  C/P:3,P
MCINTYRE, VONDA N.  EXILE WAITING, THE  C/P,P,H
MCINTYRE, VONDA N.  EXILE WAITING, THE (REVISED)  C/P,P
MCINTYRE, VONDA N.  FIREFLOOD AND OTHER STORIES  C/P,P,H:2
MCINTYRE, VONDA N.  METAPHASE [STARFARERS 3]  C/P:2,P
MCINTYRE, VONDA N.  MOON AND THE SUN, THE  H,P
MCINTYRE, VONDA N.  NAUTILUS [STARFARERS 4]  C/P,P
MCINTYRE, VONDA N.  NEBULA AWARDS SHOWCASE 2004  LPA
MCINTYRE, VONDA N.  SCREWTOP  TD7,SFWA-TD7
MCINTYRE, VONDA N.  STAR TREK II: THE WRATH OF KHAN [STAR TREK]  C/P,P,H
MCINTYRE, VONDA N.  STAR TREK III: THE SEARCH FOR SPOCK [STAR TREK]  H:2,P
MCINTYRE, VONDA N.  STAR TREK IV: THE VOYAGE HOME [STAR TREK]  C/P:2,P,H
MCINTYRE, VONDA N.  STARFARERS [STARFARERS 1]  C/P:2,SFWA-P,P
MCINTYRE, VONDA N.  SUPERLUMINAL  C/P:2,SFWA,P,H
MCINTYRE, VONDA N.  TRANSITION [STARFARERS 2]  C/P:2,P
MCKAY, KENNETH  INDECENT RELATIONS  P
MCKENNA, JULIET E.  ASSASSIN’S EDGE, THE [EINARINN 5]  P
MCKENNA, JULIET E.  GAMBLER’S FORTUNE, THE [EINARINN 3]  P
MCKENNA, JULIET E.  NORTHERN STORM [@ALDABRESHIN COMPASS #2]  P
MCKENNA, JULIET E.  SOUTHERN FIRE [ALDABRESHIN COMPASS 1]  H
MCKENNA, JULIET E.  SWORDSMAN’S OATH, THE [EINARINN 2]  H,P
MCKENNA, JULIET E.  THIEF’S GAMBLE, THE [EINARINN 1]  C/P:2,P
MCKENNA, JULIET E.  WARRIOR’S BLOOD, THE [EINARINN 4]  P
MCKENNA, JULIET E.  WESTERN SHORE [@ALDABRESHIN COMPASS #3]  P
MCKENNA, RICHARD  CASEY AGONISTES AND OTHER SCIENCE FICTION AND FANTASY STORIES H,P
MCKENNA, STEPHEN  OLDEST GOD, THE  H
MCKENNEY, KENNETH  MOONCHILD, THE  P
MCKENNEY, KENNETH  PLANTS, THE  P

MCKENZIE, CHUCK/ ROBERTS, TANSY RAYNER AUSTRALIEN ABSURDITIES

LPA

MCKENZIE, ELLEN KINDT  GOLDEN BAND OF EDDRIS, THE  H
MCKENZIE, MELINDA  MAGIC TO DO  P
MCKENZIE, NANCY  CHILD QUEEN, THE [CHILD QUEEN 1]  P
MCKENZIE, NANCY  GRAIL PRINCE  LP
MCKENZIE, NANCY  HIGH QUEEN, THE [CHILD QUEEN 2]  P
MCKEONE, DIXIE LEE  GREATHEART [BIRTHRIGHT 2]  P
MCKEONE, DIXIE LEE  SENTINEL, THE [@WORLD OF ADEN THUNDERSCAPE]  P
MCKEONE, DIXIE LEE  SPIDER’S TEST, THE [@BIRTHRIGHT 5]  P
MCKEONE, DIXIE LEE  TALES OF UNCLE TRAPSPRINGER [@DRAGONLANCE]  P
MCKEONE, LEE  BACKBLAST [CONEK HAYDEN]  P
MCKEONE, LEE  CLONE CRISIS, THE  P
MCKEONE, LEE  STARFIRE DOWN [CONEK HAYDEN]  P
MCKIE, ANGUS  SO BEAUTIFUL AND SO DANGEROUS  CX
MCKIERNAN, DENNIS L.  BREGA PATH, THE [@SILVER CALL #2/MITHGAR 11]  C/P,SFWA,P,H
MCKIERNAN, DENNIS L.  CAVERNS OF SOCRATES  P
MCKIERNAN, DENNIS L.  DARK TIDE, THE [@IRON TOWER #1/MITHGAR 7]  C/P:2,SFWA,P
MCKIERNAN, DENNIS L.  DARKEST DAY, THE [@IRON TOWER #3/MITHGAR 9]  C/P:2,SFWA,P
MCKIERNAN, DENNIS L.  DRAGONDOOM [MITHGAR 5]  C/P:2,P
MCKIERNAN, DENNIS L.  DRAGONSTONE, THE [MITHGAR 1]  P
MCKIERNAN, DENNIS L.  EYE OF THE HUNTER, THE [MITHGAR 12]  C/P,SFWA,LP
MCKIERNAN, DENNIS L.  INTO THE FIRE [HEL’S CRUCIBLE 2/MITHGAR 4]  P
MCKIERNAN, DENNIS L.  INTO THE FORGE [HEL’S CRUCIBLE 1/MITHGAR 3]  P
MCKIERNAN, DENNIS L.  ONCE UPON A WINTER’S NIGHT  P

AUTHOR  TITLE  CODE
MCKIERNAN, DENNIS L.  RED SLIPPERS [@MITHGAR]  H,P
MCKIERNAN, DENNIS L.  SHADOWS OF DOOM [@IRON TOWER #2/MITHGAR 8]  C/P,SFWA,P,H
MCKIERNAN, DENNIS L.  SILVER CALL, THE (OMNIBUS) [MITHGAR 10,11]  P
MCKIERNAN, DENNIS L.  SILVER WOLF, BLACK FALCON [MITHGAR 13]  H
MCKIERNAN, DENNIS L.  TALES OF MITHGAR [MITHGAR 6]  LP
MCKIERNAN, DENNIS L.  TREK TO KRAGGEN-COR [@SILVER CALL #1/MITHGAR 10]  C/P,SFWA,P,H
MCKIERNAN, DENNIS L.  VOYAGE OF THE FOX RIDER [MITHGAR 2]  C/P,P
MCKIERNAN, DENNIS L.  VULGMASTER, THE [@TALES FROM THE ONE-EYED CROW]  C/P:2,SCX,SFWA-P
MCKILLIP, PATRICIA A.  ALPHABET OF THORN  LP
MCKILLIP, PATRICIA A.  BOOK OF ATRIX WOLFE, THE  P
MCKILLIP, PATRICIA A.  CHANGELING SEA, THE  C/P:2,P
MCKILLIP, PATRICIA A.  CYGNET AND THE FIREBIRD, THE [CYGNET 2]  SFWA
MCKILLIP, PATRICIA A.  FOOL’S RUN  C/P,P,H
MCKILLIP, PATRICIA A.  FORGOTTEN BEASTS OF ELD, THE  C/P,P,H
MCKILLIP, PATRICIA A.  HARPIST IN THE WIND  C/P,P,H
MCKILLIP, PATRICIA A.  HARROWING THE DRAGON  H
MCKILLIP, PATRICIA A.  HEIR OF SEA AND FIRE  C/P,P
MCKILLIP, PATRICIA A.  IN THE FORESTS OF SERRE  H
MCKILLIP, PATRICIA A.  MOON AND THE FACE, THE  C/P,P,H
MCKILLIP, PATRICIA A.  MOON-FLASH  C/P,P,H
MCKILLIP, PATRICIA A.  NIGHT GIFT, THE  H,LP
MCKILLIP, PATRICIA A.  OD MAGIC  H:2
MCKILLIP, PATRICIA A.  OMBRIA IN SHADOW  H
MCKILLIP, PATRICIA A.  RIDDLE-MASTER OF HED, THE  C/P,P,H
MCKILLIP, PATRICIA A.  RIDDLE OF STARS  H:2
MCKILLIP, PATRICIA A.  SOMETHING RICH AND STRANGE  LP
MCKILLIP, PATRICIA A.  SORCERESS AND THE CYGNET, THE [CYGNET 1]  C/P:2,SFWA,SFWA-P:2,P
MCKILLIP, PATRICIA A.  STEPPING FROM THE SHADOWS  C/P,P
MCKILLIP, PATRICIA A.  THROME OF THE ERRIL OF SHERILL, THE  C/P,P
MCKILLIP, PATRICIA A.  TOWER AT STONY WOOD, THE  H
MCKILLIP, PATRICIA A.  WINTER ROSE  P
MCKINLEY, ROBIN  BEAUTY  C/P,P,H
MCKINLEY, ROBIN  BLUE SWORD, THE  H:2,P
MCKINLEY, ROBIN  DEERSKIN  P,SFWA
MCKINLEY, ROBIN  DOOR IN THE HEDGE, THE  C/P:2,P
MCKINLEY, ROBIN  HERO AND THE CROWN, THE  H:2,P
MCKINLEY, ROBIN  IMAGINARY LANDS  C/PA,PA
MCKINLEY, ROBIN  KNOT IN THE GRAIN, A  P
MCKINLEY, ROBIN  OUTLAWS OF SHERWOOD, THE  C/P:2,SFWA-P,P
MCKINLEY, ROBIN  ROSE DAUGHTER  P
MCKINLEY, ROBIN  SPINDLE’S END  H:2
MCKINLEY, ROBIN  SUNSHINE  H
MCKINLEY, ROBIN  SUNSHINE (PROOF)  SR-LP
MCKINLEY, ROBIN/DICKINSON, PETER  WATER  LP
MCKINNEY, JACK  ARTIFACT OF THE SYSTEM [@BLACK HOLE TRAVEL AGENCY #2]  C/P:2,P
MCKINNEY, JACK  BATTLE CRY [@ROBOTECH #2]  P
MCKINNEY, JACK  BATTLEHYMN [@ROBOTECH #4]  P
MCKINNEY, JACK  BEFORE THE INVID STORM [@ROBOTECH #21]  C/P,P
MCKINNEY, JACK  DARK POWERS [@SENTINELS #2/ROBOTECH 14]  P
MCKINNEY, JACK  DEVIL’S HAND, THE [@SENTINELS #1/ROBOTECH 13]  C/P,P
MCKINNEY, JACK  DOOMSDAY [@ROBOTECH #6]  P
MCKINNEY, JACK  END OF THE CIRCLE, THE [@ROBOTECH #18]  C/P,P
MCKINNEY, JACK  EVENT HORIZON [@BLACK HOLE TRAVEL AGENCY #1]  C/P:2,P
MCKINNEY, JACK  FINAL NIGHTMARE, THE [@ROBOTECH #9]  C/P,P
MCKINNEY, JACK  FORCE OF ARMS [@ROBOTECH #5]  P
MCKINNEY, JACK  FREE RADICALS [@BLACK HOLE TRAVEL AGENCY #3]  C/P:2,P
MCKINNEY, JACK  GENESIS [@ROBOTECH #1]  C/P:2,P
MCKINNEY, JACK  HOMECOMING [@ROBOTECH #3]  P
MCKINNEY, JACK  HOSTILE TAKEOVER [@BLACK HOLE TRAVEL AGENCY #4]  C/P:2,P
MCKINNEY, JACK  INVID INVASION [@ROBOTECH #10]  C/P,P
MCKINNEY, JACK  KADUNA MEMORIES  C/P:2,P
MCKINNEY, JACK  MASTERS’ GAMBIT, THE [@ROBOTECH #20]  C/P,P
MCKINNEY, JACK  METAL FIRE [@ROBOTECH #8]  C/P,P
MCKINNEY, JACK  METAMORPHOSIS [@ROBOTECH #11]  C/P,P
MCKINNEY, JACK  ROBOTECH [@ROBOTECH #1,2,3]  P
MCKINNEY, JACK  RUBICON [@SENTINELS #5/ROBOTECH 17]  C/P,P
MCKINNEY, JACK  SENTINELS, THE [@SENTINELS #1,2,3/ROBOTECH 13,14,15]  P
MCKINNEY, JACK  SOUTHERN CROSS [@ROBOTECH #7]  C/P,P
MCKINNEY, JACK  SYMPHONY OF LIGHT [@ROBOTECH #12]  P
MCKINNEY, JACK  WORLD KILLERS [@SENTINELS #4/ROBOTECH 16]  C/P,P
MCKINNEY, JACK  ZENTRAEDI REBELLION, THE [@ROBOTECH #19]  P
MCLAREN, TERI  CURSED LAND, THE [MAGIC: THE GATHERING]  P
MCLAREN, TERI  SONG OF TIME [MAGIC: THE GATHERING]  P
MCLAUGHLIN, DEAN  DOME WORLD  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 263
AUTHOR  TITLE  CODE
MCLAUGHLIN, DEAN  FURY FROM EARTH, THE  C/P:2,P
MCLAUGHLIN, DEAN  HAWK AMONG THE SPARROWS  H,SFWA
MCLAUGHLIN, DEAN  MAN WHO WANTED STARS, THE  C/P:2,P
MCLAUGHLIN, IAIN  BLOOD AND HOPE [DOCTOR WHO]  SR-H
MCLAUGHLIN, IAIN/BARTLETT, CLAIRE  ECHOES [@TIME HUNTER 6]  LP

MCLAUGHLIN, MARK/GRAVES, RAIN/ WILSON, DAVID NIALL GOSSAMER EYE, THE

LP

MCLOUGHLIN, JOHN C.  HELIX AND THE SWORD, THE  C/P,P,H
MCLOUGHLIN, JOHN C.  TOOLMAKER KOAN  C/P,SFWA,SFWA-P,P,H
MCMAHON, DONNA  DANCE OF KNIVES  LP
MCMAHON, MICK  HUBSCHMANN EFFECT, THE  P
MCMANUS, JAMES  CHIN MUSIC  H
MCMASTER, MICHAEL  U.S.S. ENTERPRISE BRIDGE BLUEPRINTS  A
MCMASTERS, WILLIAM H.  REVOLT  H
MCMULLEN, SEAN  CENTURION’S EMPIRE, THE  P
MCMULLEN, SEAN  EYES OF THE CALCULOR [GREAT MACHINE 2]  H
MCMULLEN, SEAN  GLASS DRAGONS [MOONWORLDS 2]  P
MCMULLEN, SEAN  MIOCENE ARROW, THE  LP
MCMULLEN, SEAN  SOULS IN THE GREAT MACHINE [GREAT MACHINE 1]  LP
MCMULLEN, SEAN  VOIDFARER [MOONWORLDS]  H
MCMULLEN, SEAN  VOYAGE OF THE SHADOWMOON [MOONWORLDS 1]  H
MCNALLY, CLAIRE  GHOST HOUSE [GHOST HOUSE 1]  P
MCNALLY, CLAIRE  GHOST HOUSE REVENGE [GHOST HOUSE 2]  P
MCNALLY, CLARE  SOMEBODY COME AND PLAY  SFWA-P
MCNALLY, CLARE  STAGE FRIGHT  SFWA-P
MCNALLY, RAYMOND T.  CLUTCH OF VAMPIRES, A  HA,PA
MCNALLY, RAYMOND T./FLORESCU, RADU  IN SEARCH OF DRACULA  R/REF-P,C/REF-P
MCNAUGHTON, BRIAN  EVEN MORE NASTY STORIES  LP
MCNAUGHTON, BRIAN  NASTY STORIES  LP
MCNAUGHTON, BRIAN  SATAN’S LOVE CHILD  SR-P
MCNAUGHTON, BRIAN  SATAN’S MISTRESS  P
MCNAUGHTON, BRIAN  SATAN’S SEDUCTRESS  P
MCNAUGHTON, BRIAN  THRONE OF BONES, THE  H
MCNAUGHTON, BRIAN  WORSE THINGS WAITING  LP
MCNEIL, JOHN  LITTLE BROTHER  H,P
MCNEILL, GRAHAM  AMBASSADOR, THE [@WARHAMMER VON VELTEN 1]  P
MCNEILL, GRAHAM  DEAD SKY, BLACK SUN [@WARHAMMER 40K ULTRAMARINES 3]  P
MCNEILL, GRAHAM  FALSE GODS [@WARHAMMER 40K HORUS HERESY 2]  P
MCNEILL, GRAHAM  GUARDIANS OF THE FOREST [@WARHAMMER]  P
MCNEILL, GRAHAM  NIGHTBRINGER [WARHAMMER 40K ULTRAMARINES 1]  P
MCNEILL, GRAHAM  STORM OF IRON [WARHAMMER 40K]  P

MCNEILL, GRAHAM ULTRAMARINES OMNIBUS, THE

[@WARHAMMER 40K ULTRAMARINES 1,2,3]

LP

MCNEILL, GRAHAM  URSUN’S TEETH [@WARHAMMER VON VELTEN 2]  P
MCNEILL, GRAHAM  WARRIORS OF ULTRAMAR [WARHAMMER 40K ULTRAMARINES 2]  P
MCNEILL, JAMES  SUNKEN CITY, THE  LP
MCNELLY, WILLIS E./STOVER, LEON E.  ABOVE THE HUMAN LANDSCAPE  LPA
MCNELLY, WILLIS E.  DUNE ENCYCLOPEDIA, THE  R/REF-H,C/REF-H
MCPHEE, JAMES  BLOOD QUEST [SURVIVAL 2000 1]  P
MCPHEE, JAMES  FROZEN FIRE [@SURVIVAL 2000 3]  P
MCQUARRIE, RALPH  RETURN OF THE JEDI PORTFOLIO  A
MCQUAY, MIKE  DEADLIEST SHOW IN TOWN, THE [MATHEW SWAIN 3]  P
MCQUAY, MIKE  ESCAPE FROM NEW YORK  P
MCQUAY, MIKE  HOT TIME IN OLD TOWN [MATHEW SWAIN 1]  P
MCQUAY, MIKE  JITTERBUG  P
MCQUAY, MIKE  LIFEKEEPER  C/P:2,P
MCQUAY, MIKE  MEMORIES  C/P:2,P
MCQUAY, MIKE  MOTHER EARTH [MORGAN OF ALB’NY 2]  P
MCQUAY, MIKE  MY SCIENCE PROJECT  C/P,P
MCQUAY, MIKE  NEXUS, THE  C/P,P
MCQUAY, MIKE  PURE BLOOD [MORGAN OF ALB’NY 1]  P
MCQUAY, MIKE  SUSPICION [ROBOT CITY #2]  C/P:2,P
MCQUAY, MIKE  WHEN TROUBLE BECKONS [MATHEW SWAIN 2]  P
MCQUINN, DONALD E.  PRISONER WITHIN, THE [LANNAT 2]  P
MCQUINN, DONALD E.  WANDERER [WARRIOR 2]  P
MCQUINN, DONALD E.  WARRIOR [WARRIOR 1]  C/P,H,LP
MCQUINN, DONALD E.  WITCH [WARRIOR 3]  C/P,LP
MCQUINN, DONALD E.  WITH FULL HONORS [LANNAT 1]  P
MCREYNOLDS, GLENNA  CHALICE AND THE BLADE, THE  P
MCSHERRY, FRANK JR.  FANTASTIC CIVIL WAR, THE  PA,SFWA-P
MCSHERRY, FRANK JR.  FANTASTIC WORLD WAR II, THE  PA,SFWA-P
MCSWEEGAN, EDWARD  ALPHA TRANSIT  LP
MCWHIRTER, NORRIS/MCWHIRTER, ROSS  GUINNESS BOOK OF WORLD RECORDS, THE  [C/RAN-P]
MCWILLIAMS, JUDITH  MAN FROM ATLANTIS, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 264
AUTHOR  TITLE  CODE
MEACHAM, BETH  TERRY’S UNIVERSE  HA,SFWA-P
MEAD, HAROLD  BRIGHT PHOENIX, THE  SR-H,C/P:2,P
MEAD, RICHELLE  SUCCUBUS BLUES  LP
MEAD, SHEPHERD  BIG BALL OF WAX, THE  C/P:2,P
MEAD, SHEPHERD  CAREFULLY CONSIDERED RAPE OF THE WORLD, THE  P
MEADE, RICHARD  EXILE’S QUEST  P
MEADE, RICHARD  SWORD OF MORNING STAR, THE  C/P:2,P
MEANEY, DEE MORRISON  DEATH OF THE RAVEN [LADY BRANWEN 2]  P
MEANEY, DEE MORRISON  ISEULT  P
MEANEY, DEE MORRISON  UNKINDNESS OF RAVENS, AN [LADY BRANWEN 1]  P
MEANEY, JOHN  CONTEXT [NULAPEIRON 2]  H
MEANEY, JOHN  PARADOX [NULAPEIRON 1]  P
MEANEY, JOHN  RESOLUTION [NULAPEIRON 3]  H
MEANEY, JOHN  TO HOLD INFINITY  P
MEDEIROS, TERESA  VAMPIRE WHO LOVED ME, THE  H
MEEK, COLONEL S. P.  DRUMS OF TAPAJOS  H

MEERSON, STEVE/KRIKES, PETER/ BENNETT, HARVE/MEYER, NICHOLAS STAR TREK IV: THE VOYAGE HOME [STAR TREK]

SR-S

MEIER, SHIRLEY  SHADOW’S DAUGHTER [FIFTH MILLENNIUM]  C/P:2,P

MEIER, SHIRLEY/STIRLING, S. M./ WEHRSTEIN, KAREN SHADOW’S SON [FIFTH MILLENNIUM]

C/P,SFWA-P,P

MEIKLE, WILLIAM  WATCHERS: THE COMING OF THE KING  LP
MEIKLEJOHN GRAPHICS  DRAGON CHRONICLES, THE  H:2
MEINHOLD, WILLIAM  SIDONIA THE SORCERESS AND THE AMBER WITCH I  H
MEINHOLD, WILLIAM  SIDONIA THE SORCERESS AND THE AMBER WITCH II  H
MELAMED, LEO  TENTH PLANET, THE  LP
MELDE, G. R.  PACIFIC ADVANCE  SR-P
MELLICK, CARLTON III  STEEL BREAKFAST ERA, THE  ERD3
MELLICK, CARLTON III  SUNSET WITH A BEARD  LP
MELLICK, CARLTON III  TEETH AND TONGUE LANDSCAPE  ERD2
MELUCH, R. M.  CHICAGO RED  C/P:2,P
MELUCH, R. M.  JERUSALEM FIRE  C/P,P
MELUCH, R. M.  MYRIAD, THE [@TOUR OF THE MERRIMACK #1]  H
MELUCH, R. M.  QUEEN’S SQUADRON, THE  SFWA-P
MELUCH, R. M.  SOVEREIGN  C/P,P
MELUCH, R. M.  WIND CHILD  C/P,P
MELUCH, R. M.  WIND DANCERS  C/P,P
MELUCH, R. M.  WOLF STAR [@TOUR OF THE MERRIMACK #2]  H
MELVILLE, HERMAN  MOBY DICK (ABRIDGED)  P
MENARD, ANDREW R.  ABSOLUTE VALUES  H
MENDELSOHN, FELIX JR.  SUPERBABY  C/P,P
MENDELSON, DREW  PILGRIMAGE  C/P,P
MENDLESOHN, FARAH  GLORIFYING TERRORISM  LPA
MENGER, HOWARD  FROM OUTER SPACE  [C/RAN-P]
MENGER, LUCY  THEODORE STURGEON  R/REF-P
MENOTTI, GIAN CARLO/DEAN, LEIGH  HELP, HELP, THE GLOBOLINKS!  H
MENTOR BOOKS  NEW WORLD WRITING 9  PA
MENZEL, DONALD H./TAVES, ERNEST H.  UFO ENIGMA, THE  SFWA
MENZEL, DONALD H./BOYD, LYLE G.  WORLD OF FLYING SAUCERS, THE  [C/RAN-H]
MERAK, A. J.  BARRIER UNKNOWN  SR-P
MERAK, A. J.  DARK ANDROMEDA  SR-P
MERAK, A. J.  DARK MILLENNIUM, THE  C/P:2,P,H
MERAK, A. J.  FROZEN PLANET, THE  P
MERAK, A. J.  HYDROSPHERE  H
MERAK, A. J.  NO DAWN AND NO HORIZON  H
MEREDITH, GEORGE  SHAVING OF SHAGPAT, THE  P
MEREDITH PRESS  THREE FOR TOMORROW  HA
MEREDITH, RICHARD C.  AT THE NARROW PASSAGE [ERIC MATHERS 1]  C/P,P,H
MEREDITH, RICHARD C.  NO BROTHER, NO FRIEND [ERIC MATHERS 2]  C/P,P,H
MEREDITH, RICHARD C.  RUN, COME SEE JERUSALEM!  C/P:2,P
MEREDITH, RICHARD C.  SKY IS FILLED WITH SHIPS, THE  P
MEREDITH, RICHARD C.  VESTIGES OF TIME [ERIC MATHERS 3]  H,SFWA,P
MEREDITH, RICHARD C.  WE ALL DIED AT BREAKAWAY STATION  C/P:2,P
MERETZKY, S. ERIC  CAVERN OF DOOM, THE [ZORK 3]  C/P,P
MERETZKY, S. ERIC  CONQUEST AT QUENDOR [ZORK 4]  SFWA-P
MERETZKY, S. ERIC  FORCES OF KRILL, THE [ZORK 1]  P
MERETZKY, S. ERIC  MALIFESTRO QUEST, THE [ZORK 2]  P
MERLE, ROBERT  DAY OF THE DOLPHIN, THE  C/P,P,H
MERLE, ROBERT  MALEVIL  C/P,P,H
MERLE, ROBERT  VIRILITY FACTOR, THE  H
MERRIAM-WEBSTER  NEW MERRIAM-WEBSTER POCKET DICTIONARY, THE  DICT
MERRIL, JUDITH  BEST OF JUDITH MERRIL, THE  C/P:2,P
MERRIL, JUDITH  BEYOND HUMAN KEN  C/PA,PA
MERRIL, JUDITH  BEYOND THE BARRIERS OF SPACE AND TIME  HA
MERRIL, JUDITH  DAUGHTERS OF EARTH  C/P:2,P

AUTHOR  TITLE  CODE
MERRIL, JUDITH  ENGLAND SWINGS SF  C/PA,PA
MERRIL, JUDITH  GALAXY OF GHOULS  SR-PA,C/PA
MERRIL, JUDITH  HOMECALLING AND OTHER STORIES  H
MERRIL, JUDITH  HUMAN?  SR-PA,C/PA
MERRIL, JUDITH  OFF THE BEATEN ORBIT  PA,C/PA
MERRIL, JUDITH  OUT OF BOUNDS  C/P:2,P
MERRIL, JUDITH  S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY  C/PA,PA
MERRIL, JUDITH  S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 2  PA,C/PA
MERRIL, JUDITH  S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 3  PA,C/PA
MERRIL, JUDITH  S-F: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY 4  PA,C/PA
MERRIL, JUDITH  SF: ’57: THE YEAR’S GREATEST SCIENCE FICTION AND FANTASY  HA
MERRIL, JUDITH  SF: ’59: THE YEAR’S GREATEST SCIENCE-FICTION AND FANTASY  HA
MERRIL, JUDITH  SF: THE BEST OF THE BEST  PA,C/PA
MERRIL, JUDITH  SHADOW ON THE HEARTH  H
MERRIL, JUDITH  SHOT IN THE DARK  SR-PA
MERRIL, JUDITH  SPACE-TIME JOURNAL, THE  PA
MERRIL, JUDITH  SURVIVAL SHIP AND OTHER STORIES  SR-P,C/P
MERRIL, JUDITH  TESSERACTS  LPA
MERRIL, JUDITH  TOMORROW PEOPLE, THE  C/P:2,P
MERRIL, JUDITH  YEAR’S BEST S-F 5, THE  PA,C/PA
MERRIL, JUDITH  YEAR’S BEST S-F 6, THE  PA,C/PA
MERRIL, JUDITH  YEAR’S BEST S-F 7, THE  C/PA,PA
MERRIL, JUDITH  YEAR’S BEST S-F 8, THE  PA,C/PA
MERRIL, JUDITH  YEAR’S BEST S-F 9, THE  C/PA,PA
MERRIL, JUDITH  YEAR’S BEST S-F 10, THE  HA,PA
MERRIL, JUDITH  YEAR’S BEST S-F 11, THE  C/PA,PA
MERRIL, JUDITH  SF 12  PA,C/PA
MERRILL, ALBERT ADAMS  GREAT AWAKENING, THE  SR-H
MERRILL, JEAN  PUSHCART WAR, THE  LP
MERRITT, A./BOK, HANNES  BLACK WHEEL, THE  C/P,P
MERRITT, A.  BURN, WITCH, BURN  SR-LP
MERRITT, A.  CREEP, SHADOW!  H
MERRITT, A.  CREEP, SHADOW CREEP!  SR-P
MERRITT, A.  DWELLERS IN THE MIRAGE  C/P:2,SFWA-P,P
MERRITT, A.  FACE IN THE ABYSS, THE  C/P,SR-LP,P,H
MERRITT, A.  FOX WOMAN AND OTHER STORIES, THE  P
MERRITT, A.  METAL MONSTER, THE  SR-P,C/P:2,SR-LP
MERRITT, A.  MOON POOL, THE  C/P,SR-LP,P,H:2
MERRITT, A.  SEVEN FOOTPRINTS TO SATAN  SR-P,C/P:2
MERRITT, A.  SEVEN FOOTPRINTS TO SATAN/BURN WITCH BURN  H
MERRITT, A.  SHIP OF ISHTAR, THE  SFWA-P,H:2,SR-LP,P
MERRITT, A.  STORY BEHIND THE STORY, THE  [C/RAN-H]
MERWIN, SAM JR.  CHAUVINISTO  P
MERWIN, SAM JR.  HOUSE OF MANY WORLDS, THE  GN2.12,C/P,P,H
MERWIN, SAM JR.  KILLER TO COME  H:2,GN4.22
MERWIN, SAM JR.  SEX WAR, THE (ORIGINALLY THE WHITE WIDOWS)  C/P
MERWIN, SAM JR.  THREE FACES OF TIME  DD121
MERWIN, SAM JR.  TIME SHIFTERS, THE  P
MERWIN, SAM JR.  WHITE WIDOWS, THE  P
MESSINGHAM, SIMON  FACE-EATER, THE [DOCTOR WHO EIGHTH DOCTOR]  P
MESSINGHAM, SIMON  INDESTRUCTIBLE MAN, THE [@DOCTOR WHO PAST DOCTORS]  P
MESSINGHAM, SIMON  INFINITY RACE, THE [DOCTOR WHO EIGHTH DOCTOR]  P
MESSINGHAM, SIMON  STRANGE ENGLAND [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
MESSINGHAM, SIMON  TOMB OF VALDEMAR [DOCTOR WHO PAST DOCTORS]  P
MESSINGHAM, SIMON  ZETA MAJOR [DOCTOR WHO PAST DOCTORS]  C/P,P
MESSMANN, JON  DEADLY DEEP, THE  P
MESTA, GABRIEL  SHADOW OF THE XEL’NAGA [@STARCRAFT #2]  P
METCALF, NORMAN  INDEX OF SCIENCE FICTION MAGAZINES, 1951–1965, THE  R/IND
METZ, DIANA  TALON AND THE DRAGONS OF CRINNELIA  P
METZGER, ARTHUR  INDEX AND SHORT HISTORY OF UNKNOWN, AN  R/IND
METZGER, ROBERT A.  CUSP  H
METZGER, ROBERT A.  PICOVERSE  H
METZGER, ROBERT A.  QUAD WORLD  LP
MEYER, JOHN J.  THIRTEEN SECONDS THAT ROCKED THE WORLD, OR THE MENTATOR  H
MEYER, JOHN J.  TRY ANOTHER WORLD  H
MEYER, KAI  WELLENLAUFER, DIE  FH
MEYER, NICHOLAS  SEVEN-PER-CENT SOLUTION, THE  H,P
MEYERS, RICHARD S.  CRY OF THE BEAST  P
MEYERS, RICHARD S.  DOOM STAR [DOOM STAR 1]  SR-P,C/P:2
MEYERS, RICHARD S.  DOOM STAR NUMBER TWO [DOOM STAR 2]  SR-P
MEYERS, RICHARD S.  RETURN TO DOOMSTAR [DOOM STAR]  C/P,P
MEYERS, RICHARD S.  S-F 2  R/REF-H
MEYERS, ROY  DAUGHTERS OF THE DOLPHIN [DOLPHIN 2]  C/P:2,P
MEYERS, ROY  DESTINY AND THE DOLPHINS [DOLPHIN 3]  C/P:2,P
MEYERS, ROY  DOLPHIN BOY [DOLPHIN 1]  C/P:2,P

AUTHOR  TITLE  CODE
MEZO, FRANCINE  FALL OF WORLDS, THE [@AREIA DARENGA #1]  C/P:2,P
MEZO, FRANCINE  NO EARTHLY SHORE [@AREIA DARENGA #3]  C/P,P
MEZO, FRANCINE  UNLESS SHE BURN [@AREIA DARENGA #2]  P
MEZRICH, BEN  SKIN [@X-FILES]  H
MIALL, ROBERT  FLESH HUNTERS [@UFO #1]  P
MIALL, ROBERT  SPORTING BLOOD [@UFO #2]  P
MICHAELS, BARBARA  BE BURIED IN THE RAIN  P
MICHAELS, BARBARA  GREY BEGINNING, THE  SFWA-P:2
MICHAELS, BARBARA  HERE I STAY  SFWA-P
MICHAELS, BARBARA  HOUSES OF STONE  P
MICHAELS, BARBARA  SEA-KING’S DAUGHTER, THE  P
MICHAELS, BARBARA  WALKER IN SHADOWS, THE  P
MICHAELS, BARBARA  WIZARD’S DAUGHTER, THE  P
MICHAELS, FERN  WITHOUT WARNING  P
MICHAELS, MELISA C.  COLD IRON  P
MICHAELS, MELISA C.  FAR HARBOR  P
MICHAELS, MELISA C.  FIRST BATTLE [SKYRIDER 2]  P,SFWA-P
MICHAELS, MELISA C.  FLOATER FACTOR [SKYRIDER 5]  P,SFWA-P
MICHAELS, MELISA C.  LAST WAR [SKYRIDER 3]  P,SFWA-P
MICHAELS, MELISA C.  PIRATE PRINCE [SKYRIDER 4]  P,SFWA-P
MICHAELS, MELISA C.  SISTER TO THE RAIN  P
MICHAELS, MELISA C.  SKIRMISH [SKYRIDER 1]  C/P,SFWA-P:2,P
MICHALOWSKI, MARK  HALFLIFE [DOCTOR WHO EIGHTH DOCTOR]  P
MICHALOWSKI, MARK  RELATIVE DEMENTIAS [DOCTOR WHO PAST DOCTORS]  P
MICHALSON, KAREN  ENEMY GLORY [ENEMY GLORY 1]  H,P
MICHALSON, KAREN  HECATE’S GLORY [ENEMY GLORY 2]  H
MICHAUD, MARC A.  FIRST WORLD FANTASY CONVENTION: THREE AUTHORS REMEMBER, THE R/REF-P

MICHELINIE, DAVID MAN WHO STOLE TOMORROW, THE

[AVENGERS (MARVEL COMICS CHARACTERS)]

C/P,P

MICHELINIE, DAVID/SMITH, DEAN WESLEY  SPIDER-MAN: CARNAGE IN NEW YORK [SPIDER-MAN]  P
MICHELSON, MIRIAM  AWAKENING OF ZOJAS, THE  H
MICHENER, JAMES A.  SPACE  C/P,P,H
MICKLEM, SARAH  FIRETHORN  LP
MIEDANER, TERREL  SOUL OF ANNA KLANE, THE  C/P,P
MIESEL, SANDRA  AGAINST TIME’S ARROW: THE HIGH CRUSADE OF POUL ANDERSON  R/REF-P
MIESEL, SANDRA  DREAMRIDER  C/P,P
MIESEL, SANDRA  MYTH, SYMBOL AND RELIGION IN THE LORD OF THE RINGS  R/REF-P
MIESEL, SANDRA  SHAMAN  C/P,SFWA-P,P
MIEVILLE, CHINA  IRON COUNCIL [NEW CROBUZON]  H:2
MIEVILLE, CHINA  KING RAT  LP
MIEVILLE, CHINA  LOOKING FOR JAKE  LP
MIEVILLE, CHINA  PERDIDO STREET STATION  C/P,LP,H
MIEVILLE, CHINA  SCAR, THE  H,LP
MIEVILLE, CHINA  UN LUN DUN  H
MIKHAILOVA, S.  FANTASTIKA 73–74  FH
MIKHANOVSKY, VLADIMIR  DOUBLES, THE  P
MIKLOWITZ, GLORIA  AFTER THE BOMB  C/P,P
MILAN, VICTOR  BLACK DRAGON [@BATTLETECH 29]  C/P,P
MILAN, VICTOR  CLD: COLLECTIVE LANDING DETACHMENT  P
MILAN, VICTOR  CLOSE QUARTERS [@BATTLETECH 15]  P
MILAN, VICTOR  CYBERNETIC SAMURAI, THE [CYBERNETIC SAMURAI 1]  H:2,SFWA,P
MILAN, VICTOR  CYBERNETIC SHOGUN, THE [CYBERNETIC SAMURAI 2]  H
MILAN, VICTOR  FLIGHT OF THE FALCON [MECHWARRIOR DARK AGE/BATTLETECH]  P
MILAN, VICTOR  FROM THE DEPTHS [STAR TREK]  C/P:2,P
MILAN, VICTOR  HEARTS OF CHAOS [@BATTLETECH 26]  P
MILAN, VICTOR  RENDING OF FALCONS, A [@MECHWARRIOR DARK AGE/BATTLETECH]  P
MILAN, VICTOR/SNODGRASS, MELINDA  RUNESPEAR  C/P,P
MILAN, VICTOR  TURN OF THE CARDS [WILD CARDS 12]  C/P:2,P
MILAN, VICTOR  WAR IN TETHYR [FORGOTTEN REALMS NOBLES 2]  P
MILES, ELLEN  WIZARDS & WARRIORS [WORLDS OF POWER 5]  P
MILES, KEITH  DRAGON’S TEETH  C/P,P
MILES, LAWRENCE  ADVENTURESS OF HENRIETTA STREET, THE [DOCTOR WHO]  P
MILES, LAWRENCE  ALIEN BODIES [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
MILES, LAWRENCE  BOOK OF THE WAR, THE [@FACTION PARADOX 0]  LPA
MILES, LAWRENCE  CHRISTMAS ON A RATIONAL PLANET [DOCTOR WHO NEW ADVENTURES]  P
MILES, LAWRENCE  DEAD ROMANCE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  P

MILES, LAWRENCE DEAD ROMANCE (WITH ADDITIONAL STORIES)

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

LP

MILES, LAWRENCE  DOWN [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P

MILES, LAWRENCE INTERFERENCE BOOK ONE: SHOCK TACTIC

[DOCTOR WHO PAST DOCTORS]

P

MILES, LAWRENCE INTERFERENCE BOOK TWO: THE HOUR OF THE GEEK

[DOCTOR WHO EIGHTH DOCTOR]

P

MILES, LAWRENCE  THIS TOWN WILL NEVER LET US GO [@FACTION PARADOX 1]  LP

The MITSFS Pinkdex  September 24, 2007  Page 267
AUTHOR  TITLE  CODE
MILES, PATRICIA  GODS IN WINTER, THE  H
MILK, LOUIS G.  DUENO DEL MUNDO  FP
MILLAR, MARTIN  GOOD FAIRIES OF NEW YORK, THE  LP
MILLARD, JOSEPH  GODS HATE KANSAS, THE  C/P:2,P
MILLENNIUM  WEB 2027  PA
MILLER, BENJAMIN E.  ZERO HOUR  P
MILLER, CALVIN  GUARDIANS OF THE SINGREALE  LP
MILLER, CALVIN  SINGER, THE  LP
MILLER, CARL  DRAGONBOUND  P
MILLER, CARL  GOBLIN PLAIN WAR, THE  C/P,P
MILLER, CARL  WARRIOR AND THE WITCH, THE  P
MILLER, ELIZABETH  YOKE, THE  H
MILLER, FAREN  ILLUSIONISTS, THE  P
MILLER, FRANK/GIBBONS, DAVE  GIVE ME LIBERTY  CX
MILLER, FRANK/DARROW, GEOF  HARD BOILED  CX
MILLER, J. P.  SKOOK, THE  H,P
MILLER, JIMMY  BIG WIN, THE  C/P,P
MILLER, JOHN  FIRST POWER PLAY [BUCK ROGERS INNER PLANETS 1]  C/P:2,P
MILLER, JOHN A.  MEN AND VOLTS AT WAR  [C/RAN-P]
MILLER, JOHN J.  DEATH DRAWS FIVE [@WILD CARDS #17]  H
MILLER, JOHN J.  SHADES OF NIGHT, FALLING [@TWILIGHT ZONE 1]  LP
MILLER, JOHN J.  TERRIBLE BEAUTY, A [@WITCHBLADE]  P
MILLER, JON DE BURGH  DEUS LE VOLT [TIME HUNTER 8]  SR-H
MILLER, JON DE BURGH  DYING IN THE SUN [DOCTOR WHO PAST DOCTORS]  P
MILLER, KAREN  ALLIANCES [@STARGATE SG-1 #8]  P
MILLER, KEITH  BOOK OF FLYING, THE  H
MILLER, LINDA LAEL  FOREVER AND THE NIGHT  H
MILLER, MARJORIE M.  ISAAC ASIMOV: A CHECKLIST...  R/IND
MILLER, P. SCHUYLER  TITAN, THE  H:2
MILLER, R. DEWITT/HUNGER, ANNA  MAN WHO LIVED FOREVER, THE  DD162
MILLER, R. DEWITT  REINCARNATION — THE WHOLE STARTLING STORY  [C/RAN-P]
MILLER, R. DEWITT  STRANGER THAN LIFE  [C/RAN-P]

MILLER, RAND/MILLER, ROBYN/ WINGROVE, DAVID MYST: THE BOOK OF ATRUS [@MYST 1]

C/P,P

MILLER, RAND/WINGROVE, DAVID  MYST: THE BOOK OF D’NI [@MYST 2]  P
MILLER, RAND/WINGROVE, DAVID  MYST: THE BOOK OF TI’ANA [@MYST 3]  H,P
MILLER, REX  SLOB  P
MILLER, RICHARD  SNAIL  LP
MILLER, RON  BRADAMANT: THE IRON TEMPEST  H,LP
MILLER, RON  HEARTS AND ARMOR [BRONWYN 3]  P
MILLER, RON  PALACES AND PRISONS  C/P,P
MILLER, RON  PALACES AND PRISONS (REVISED) [BRONWYN 1]  LP
MILLER, RON  SILK AND STEEL [BRONWYN 2]  C/P,P
MILLER, RON  SPACE ART  A
MILLER, RUSTY  JEDI MASTER’S QUIZBOOK, THE  [C/RAN-P]
MILLER, RYAN S. J.  INKARRI  LP
MILLER, SASHA  LADYLORD  C/P,P,H
MILLER, STEPHEN O.  MITHRANDIR  R/REF-P

MILLER, STEPHEN T./ CONTENTO, WILLIAM G.

SCIENCE FICTION, FANTASY, & WEIRD FICTION MAGAZINE INDEX (1890–2003) (CD)

SR-MM

MILLER, STEVE  CHARIOT TO THE STARS  SR-LP
MILLER, WALTER M. JR.  BEST OF WALTER M. MILLER JR., THE  C/P,P

MILLER, WALTER M. JR./ GREENBERG, MARTIN H.

BEYOND ARMAGEDDON

HA,PA,SFWA

MILLER, WALTER M. JR.  CANTICLE FOR LEIBOWITZ, A [LEIBOWITZ 1]  H:2,LP
MILLER, WALTER M. JR.  CONDITIONALLY HUMAN  C/P:2,P
MILLER, WALTER M. JR.  SAINT LEIBOWITZ AND THE WILD HORSE WOMAN [LEIBOWITZ 2]  H:2,LP
MILLER, WALTER M. JR.  VIEW FROM THE STARS, THE  C/P:2,P
MILLER, WARREN  LOOKING FOR THE GENERAL  P
MILLER, WARREN  SIEGE OF HARLEM, THE  P
MILLERS, REINHOLD  TIME EXILE  H
MILLHISER, MARLYS  MIRROR, THE  H
MILLS, C. J.  BRANDER’S BOOK [WINTER WORLD 4]  C/P,P
MILLS, C. J.  EGIL’S BOOK [WINTER WORLD 2]  C/P,P
MILLS, C. J.  KIT’S BOOK [WINTER WORLD 3]  C/P,P
MILLS, C. J.  WINTER WORLD [WINTER WORLD 1]  C/P,P
MILLS, CRAIG  BANE OF LORD CALADON, THE  P
MILLS, CRAIG  DREAMER IN DISCORD, THE  P
MILLS, CRAIG  FLOATING CASTLE, THE [@KING’S QUEST 1]  P
MILLS, CRAIG  SHADOW OF THE CROWN  P
MILLS, JAMES  SEVENTH POWER, THE  P
MILLS, PAT/MITCHELL, ALAN  MEDUSA WAR, THE [@ABC WARRIORS]  P
MILLS, ROBERT E.  STAR FIGHTERS  P
MILLS, ROBERT E.  STAR FORCE  P

AUTHOR  TITLE  CODE
MILLS, ROBERT E.  STAR QUEST  P
MILLS, ROBERT E.  UNDER THE EYE OF NIGHT  P
MILLS, ROBERT P.  BEST FROM FANTASY AND SCIENCE FICTION 9, THE  C/PA,PA
MILLS, ROBERT P.  BEST FROM FANTASY AND SCIENCE FICTION 10, THE  PA,C/PA
MILLS, ROBERT P.  BEST FROM FANTASY AND SCIENCE FICTION 11, THE  PA,C/PA
MILLS, ROBERT P.  DECADE OF FANTASY AND SCIENCE FICTION, A  C/PA,PA
MILLS, ROBERT P.  FLOWERS FOR ALGERNON AND OTHER STORIES  C/PA,PA
MILLS, ROBERT P.  WORLDS OF SCIENCE FICTION, THE  PA,C/PA
MILNE, A. A.  ONCE ON A TIME  C/P,P
MILNE, A. A.  WINNIE ILLE PU  FH
MILNE, A. A.  WINNIE THE POOH  H
MILNE, ROBERT DUNCAN  INTO THE SUN AND OTHER STORIES  SR-H
MINES, SAMUEL  BEST FROM STARTLING STORIES, THE  HA
MINES, SAMUEL  STARTLING STORIES  HA
MINKOV, SVETOSLAV  LADY WITH THE X-RAY EYES, THE  H
MINSKY, MARVIN  ROBOTICS  SFWA
MINTON, PAULA  GIRL FROM NOWHERE, THE  P
MIR PUBLISHERS  EVERYTHING BUT LOVE  SR-PA
MIR PUBLISHERS  JOURNEY ACROSS THREE WORLDS  HA
MIR PUBLISHERS  MOLECULAR CAFE, THE  SR-PA
MIRRLEES, HOPE  LUD-IN-THE-MIST  H,P
MISSION CONTROL/JHO, ZOEV  E.T. 101  H
MISTRAL, BENGO  PIRATES OF CERBERUS  SR-P
MITCHELL, DAVID  CLOUD ATLAS  LP
MITCHELL, EDWARD PAGE  CRYSTAL MAN, THE  H,SFWA
MITCHELL, ELIZABETH  AFTER THE FLAMES [ALIEN STARS 2]  PA
MITCHELL, ELIZABETH  ALIEN STARS [ALIEN STARS 1]  PA
MITCHELL, ELIZABETH  FREE LANCERS [ALIEN STARS 4]  C/PA,PA
MITCHELL, ELIZABETH  UNDER THE WHEEL [ALIEN STARS 3]  PA
MITCHELL, J. A.  DROWSY  H
MITCHELL, J. A.  LAST AMERICAN, THE  H
MITCHELL, J. LESLIE  THREE GO BACK  GN3.15,C/P:2
MITCHELL, KIRK  CRY REPUBLIC [PROCURATOR 3]  C/P,SFWA-P,P
MITCHELL, KIRK  NEVER THE TWAIN  C/P,P
MITCHELL, KIRK  NEW BARBARIANS [PROCURATOR 2]  P
MITCHELL, KIRK  PROCURATOR [PROCURATOR 1]  P
MITCHELL, MARY ANN  CATHERDRAL OF VAMPIRES  P
MITCHELL, MIKE  DEDALUS BOOK OF AUSTRIAN FANTASY: 1890-2000, THE  LPA
MITCHELL, RONALD  DAN OWEN AND THE ANGEL JOE  H
MITCHELL, SANDY  CAVES OF ICE [@WARHAMMER 40K CIAPHAS CAIN 2]  P

MITCHELL, SANDY CIAPHAS CAIN: HERO OF THE IMPERIUM [@WARHAMMER 40K CIAPHAS CAIN 1,2,3]

LP

MITCHELL, SANDY  DEATH OR GLORY [@WARHAMMER 40K CIAPHAS CAIN 4]  P
MITCHELL, SANDY  DEATH’S CITY [@WARHAMMER BLOOD ON THE REIK 2]  P
MITCHELL, SANDY  DEATH’S MESSENGER [@WARHAMMER BLOOD ON THE REIK 1]  C/P,P
MITCHELL, SANDY  DUTY CALLS [@WARHAMMER 40K]  P
MITCHELL, SANDY  FOR THE EMPEROR [@WARHAMMER 40K CIAPHAS CAIN 1]  P
MITCHELL, SANDY  TRAITOR’S HAND, THE [@WARHAMMER 40K CIAPHAS CAIN 3]  P
MITCHELL, SYNE  CHANGELING PLAGUE, THE  P
MITCHELL, SYNE  LAST MORTAL MAN, THE [DEATHLESS 1]  P
MITCHELL, SYNE  MURPHY’S GAMBIT  C/P,P
MITCHELL, SYNE  TECHNOGENESIS  P
MITCHELL, V. E.  ATLANTIS STATION [@STAR TREK TNG STARFLEET ACADEMY #5]  H,LP
MITCHELL, V. E.  ENEMY UNSEEN [STAR TREK]  C/P:3,P
MITCHELL, V. E.  IMBALANCE [STAR TREK TNG]  C/P:2,P
MITCHELL, V. E.  WINDOWS ON A LOST WORLD [STAR TREK]  C/P:2,P
MITCHELSON, AUSTIN/UTECHIN, NICHOLAS EARTHQUAKE MACHINE, THE  P
MITCHISON, NAOMI  MEMOIRS OF A SPACEWOMAN  C/P:2,P
MITCHISON, NAOMI  SOLUTION THREE  H,P
MITROKHINA, S.  LUCHSHII IZ MIROV  FH
MITROKHINA, S./VOSKRESENSKAYA, E.  SOVREMENNAYA ZARUBEZHNAYA FANTASTIKA  FH
MIXON, LAURA J.  ASTROPILOTS  C/P,P
MIXON, LAURA J.  BURNING THE ICE  H
MIXON, LAURA J.  GLASS HOUSES  P
MIXON, LAURA J.  PROXIES  C/P,P,H

MIYAZAKI, HAYAO NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 1

[@NAUSICAA PERFECT #1]

CX

MIYAZAKI, HAYAO NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 2

[@NAUSICAA PERFECT #2]

CX

MIYAZAKI, HAYAO NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 3

[@NAUSICAA PERFECT #3]

CX

MIYAZAKI, HAYAO NAUSICAA OF THE VALLEY OF WIND PERFECT COLLECTION 4

[@NAUSICAA PERFECT #4]

CX

MOBLEY, JANE  PHANTASMAGORIA  PA
MODESITT, L. E. JR.  ADIAMANTE  P

AUTHOR  TITLE  CODE
MODESITT, L. E. JR.  ALECTOR’S CHOICE [COREAN 4]  H
MODESITT, L. E. JR.  ARCHFORM: BEAUTY [ARCHFORM BEAUTY 1]  H
MODESITT, L. E. JR.  CADMIAN’S CHOICE [COREAN 5]  H
MODESITT, L. E. JR.  CHAOS BALANCE, THE [RECLUCE 7]  H:2,P
MODESITT, L. E. JR.  COLORS OF CHAOS [RECLUCE 9]  H:2,P
MODESITT, L. E. JR.  DARKNESSES [COREAN 2]  P
MODESITT, L. E. JR.  DARKSONG RISING [SPELLSONG 3]  C/P,P,H
MODESITT, L. E. JR.  DAWN FOR A DISTANT EARTH [FOREVER HERO 1]  C/P:2,SFWA-P,P
MODESITT, L. E. JR.  DEATH OF CHAOS, THE [RECLUCE 5]  H:2,P
MODESITT, L. E. JR.  ECOLITAN ENIGMA, THE [ECOLITAN 4]  H,P
MODESITT, L. E. JR.  ECOLITAN OPERATION, THE [ECOLITAN 2]  SFWA-P
MODESITT, L. E. JR.  ECOLITAN PRIME [ECOLITAN 1,4]  LP
MODESITT, L. E. JR.  ECOLOGIC ENVOY, THE [ECOLITAN 1]  C/P,P
MODESITT, L. E. JR.  ECOLOGIC SECESSION, THE [ECOLITAN 3]  C/P:2,P
MODESITT, L. E. JR.  ELYSIUM COMMISSION, THE  H
MODESITT, L. E. JR.  EMPIRE & ECOLITAN [ECOLITAN 2,3]  LP
MODESITT, L. E. JR.  ETERNITY ARTIFACT, THE  H
MODESITT, L. E. JR.  ETHOS EFFECT, THE  P
MODESITT, L. E. JR.  FALL OF ANGELS [RECLUCE 6]  H:2,P
MODESITT, L. E. JR.  FIRES OF PARATIME, THE  C/P:2,P,H
MODESITT, L. E. JR.  FLASH [ARCHFORM BEAUTY 2]  H
MODESITT, L. E. JR.  FOREVER HERO, THE [FOREVER HERO 1,2,3]  H
MODESITT, L. E. JR.  GHOST OF THE REVELATOR, THE [JOHAN ESBACH 2]  H,P
MODESITT, L. E. JR.  GHOST OF THE WHITE NIGHTS [JOHAN ESBACH 3]  P
MODESITT, L. E. JR.  GRAVITY DREAMS  H
MODESITT, L. E. JR.  HAMMER OF DARKNESS, THE  C/P,P
MODESITT, L. E. JR.  IN ENDLESS TWILIGHT [FOREVER HERO 3]  C/P,P
MODESITT, L. E. JR.  LEGACIES [COREAN 1]  H
MODESITT, L. E. JR.  LEGACIES (PROOF) [COREAN 1]  SR-LP
MODESITT, L. E. JR.  MAGIC ENGINEER, THE [RECLUCE 3]  P
MODESITT, L. E. JR.  MAGIC OF RECLUCE, THE [RECLUCE 1]  H:2,P
MODESITT, L. E. JR.  MAGI’I OF CYADOR [RECLUCE 10]  H:2,P
MODESITT, L. E. JR.  OCTAGONAL RAVEN, THE  H,P
MODESITT, L. E. JR.  OF TANGIBLE GHOSTS [JOHAN ESBACH 1]  P
MODESITT, L. E. JR.  ORDER WAR, THE [RECLUCE 4]  H,P
MODESITT, L. E. JR.  ORDERMASTER [RECLUCE 13]  H
MODESITT, L. E. JR.  PARAFAITH WAR, THE  P
MODESITT, L. E. JR.  SCEPTERS [COREAN 3]  H
MODESITT, L. E. JR.  SCION OF CYADOR [RECLUCE 11]  H:2
MODESITT, L. E. JR.  SHADOW SORCERESS, THE [SPELLSONG 4]  C/P,P,H
MODESITT, L. E. JR.  SHADOWSINGER [SPELLSONG 5]  H
MODESITT, L. E. JR.  SILENT WARRIOR, THE [FOREVER HERO 2]  C/P,P
MODESITT, L. E. JR.  SOARER’S CHOICE [COREAN 6]  H
MODESITT, L. E. JR.  SOPRANO SORCERESS, THE [SPELLSONG 1]  H,P
MODESITT, L. E. JR.  SPELLSONG WAR, THE [SPELLSONG 2]  H,P
MODESITT, L. E. JR.  TIMEDIVER’S DAWN [TIMEDIVERS 2]  C/P:2,P
MODESITT, L. E. JR.  TIMEGOD, THE [TIMEDIVERS 1]  P
MODESITT, L. E. JR.  TIMEGODS’ WORLD [TIMEDIVERS 1,2]  H
MODESITT, L. E. JR.  TOWERS OF THE SUNSET, THE [RECLUCE 2]  C/P:2,P
MODESITT, L. E. JR.  WELLSPRING OF CHAOS [RECLUCE 12]  H
MODESITT, L. E. JR.  WHITE ORDER, THE [RECLUCE 8]  H:2,P
MOENCH, DOUG  RED RAIN [BATMAN]  CX
MOESTA, REBECCA  LITTLE THINGS [@BUFFY]  P
MOFFETT, JUDITH  PENNTERRA  H,P
MOFFETT, JUDITH  RAGGED WORLD, THE  H,P
MOFFETT, JUDITH  TIME LIKE AN EVER-ROLLING STREAM  H
MOFFITT, DONALD  CRESCENT IN THE SKY [MECHANICAL SKY 1]  C/P,P
MOFFITT, DONALD  GATHERING OF STARS, A [MECHANICAL SKY 2]  P
MOFFITT, DONALD  GENESIS QUEST, THE [GENESIS 1]  C/P:2,P
MOFFITT, DONALD  JUPITER THEFT, THE  C/P:2,P
MOFFITT, DONALD  SECOND GENESIS [GENESIS 2]  C/P:2,P
MOHAN, KIM  MORE AMAZING STORIES  HA
MOHS, MAYO  OTHER WORLDS, OTHER GODS  C/PA,PA
MOLES, DAVID/LAKE, JAY  ALL-STAR ZEPPELIN STORIES  LPA
MOLLOY, J. J.  CACOPHONY (CASSETTE) [STAR TREK]  SR-MM
MOLLOY, PAUL  PENNANT FOR THE KREMLIN, A  P
MOLNAR, FRANZ  LILIOM  H
MOLODAYA GVARDIYA  FANTASTIKA 1991  FH
MOLONEY, JAMES  BOOK OF LIES, THE  H
MOLSTAD, STEPHEN  GODZILLA  H,P
MOLSTAD, STEPHEN  SILENT ZONE [@ID4]  H,P
MOLSTAD, STEPHEN  WAR IN THE DESERT [@ID4]  C/P,P
MONACO, RICHARD  BLOOD AND DREAMS [PARSIVAL 4]  P
MONACO, RICHARD  BROKEN STONE  P

The MITSFS Pinkdex  September 24, 2007  Page 270
AUTHOR  TITLE  CODE
MONACO, RICHARD  FINAL QUEST, THE [PARSIVAL 3]  C/P,SFWA,P
MONACO, RICHARD  GRAIL WAR, THE [PARSIVAL 2]  C/P,P
MONACO, RICHARD  JOURNEY TO THE FLAME  P
MONACO, RICHARD  PARSIVAL OR A KNIGHT’S TALE [PARSIVAL 1]  C/P,P,H
MONACO, RICHARD  RUNES  C/P:2,P
MONACO, RICHARD  UNTO THE BEAST  P
MONETTE, PAUL  NOSFERATU THE VAMPYRE  P
MONETTE, PAUL  PREDATOR [PREDATOR]  C/P,P
MONETTE, SARAH  MELUSINE  H
MONETTE, SARAH  VIRTU, THE  H
MONING, KAREN MARIE  DARKFEVER  H
MONING, KAREN MARIE  SPELL OF THE HIGHLANDER  H
MONTAGUE, C. E.  RIGHT OFF THE MAP  H:2
MONTANA, RON  CATHEDRAL OPTION, THE  P
MONTANARI, GIANNI  NEL NOME DELL’UOMO  FP
MONTANARI, GIANNI  SEPOLTURA, LA  FP

MONTELEONE, ELIZABETH/ MONTELEONE, THOMAS F.

BORDERLANDS 4

C/PA,PA

MONTELEONE, ELIZABETH/ MONTELEONE, THOMAS F.

FROM THE BORDERLANDS

PA

MONTELEONE, THOMAS F.  ARTS AND BEYOND, THE  HA,SFWA
MONTELEONE, THOMAS F.  BLOOD OF THE LAMB, THE  H,P
MONTELEONE, THOMAS F.  BORDERLANDS  PA
MONTELEONE, THOMAS F.  BORDERLANDS 1 (INCLUDES "THE SCENT OF VINEGAR")  C/PA,PA
MONTELEONE, THOMAS F.  BORDERLANDS 2  C/PA,PA
MONTELEONE, THOMAS F.  BORDERLANDS 3  C/PA,PA

MONTELEONE, THOMAS F./ DE CHANCIE, JOHN CROOKED HOUSE

SFWA-P

MONTELEONE, THOMAS F.  DARK STARS AND OTHER ILLUMINATIONS  H,SFWA
MONTELEONE, THOMAS F.  FANTASMA  SFWA-P
MONTELEONE, THOMAS F.  GUARDIAN  C/P,SFWA,P,H
MONTELEONE, THOMAS F.  HORREUR DU METRO, L  FP
MONTELEONE, THOMAS F.  MAGNIFICENT GALLERY, THE  SFWA-P
MONTELEONE, THOMAS F.  NIGHT THINGS  P
MONTELEONE, THOMAS F.  NIGHT-TRAIN  P
MONTELEONE, THOMAS F.  OZYMANDIAS  H
MONTELEONE, THOMAS F.  RANDOM ACCESS MESSAGES OF THE COMPUTER AGE  SR-LPA
MONTELEONE, THOMAS F.  SECRET SEA, THE  P
MONTELEONE, THOMAS F.  SEEDS OF CHANGE  SR-P,C/P:2
MONTELEONE, THOMAS F.  TIME CONNECTION, THE  C/P,P
MONTELEONE, THOMAS F.  TIME-SWEPT CITY, THE  P
MONTGOMERY, FRANCES T.  ON A LARK TO THE PLANETS  H
MONTGOMERY, R. A.  ESCAPE FROM CHINA [TRIO]  P
MONTGOMERY, R. A.  JOURNEY UNDER THE SEA  C/P,P
MONTGOMERY, R. A.  SPACE AND BEYOND  C/P:2,P
MONTGOMERY, RUTH  GIFT OF PROPHECY, A  [C/RAN-P]
MONTY PYTHON  MONTY PYTHON AND THE HOLY GRAIL BOOK  VLP
MOON, ELIZABETH  AGAINST THE ODDS [HERIS SERRANO 7]  H,P
MOON, ELIZABETH  CHANGE OF COMMAND [HERIS SERRANO 6]  H:2,P
MOON, ELIZABETH  COMMAND DECISION [VATTA’S WAR 4]  H:2
MOON, ELIZABETH  DEED OF PAKSENARRION, THE [PAKSENARRION 1,2,3]  H,SFWA
MOON, ELIZABETH  DIVIDED ALLEGIANCE [PAKSENARRION 2]  C/P:3,SFWA-P,P
MOON, ELIZABETH  ENGAGING THE ENEMY [VATTA’S WAR 3]  H:2
MOON, ELIZABETH  HERIS SERRANO [HERIS SERRANO 1,2,3]  H
MOON, ELIZABETH  HUNTING PARTY [HERIS SERRANO 1]  C/P:2,SFWA-P,P
MOON, ELIZABETH  LIAR’S OATH [GIRD 2]  C/P:2,SFWA-P,P
MOON, ELIZABETH  LUNAR ACTIVITY  C/P:2,SFWA-P,P
MOON, ELIZABETH  MARQUE AND REPRISAL [VATTA’S WAR 2]  H,P
MOON, ELIZABETH  OATH OF GOLD [PAKSENARRION 3]  C/P:2,P
MOON, ELIZABETH  ONCE A HERO [HERIS SERRANO 4]  C/P:2,P
MOON, ELIZABETH  PHASES  C/P:2,P
MOON, ELIZABETH  REMNANT POPULATION  C/P,P,H
MOON, ELIZABETH  RULES OF ENGAGEMENT [HERIS SERRANO 5]  C/P,P,H
MOON, ELIZABETH  SHEEPFARMER’S DAUGHTER [PAKSENARRION 1]  C/P:3,SFWA-P:2,P
MOON, ELIZABETH  SPEED OF DARK, THE  H
MOON, ELIZABETH  SPORTING CHANCE [HERIS SERRANO 2]  C/P:2,P
MOON, ELIZABETH  SURRENDER NONE: THE LEGACY OF GIRD [GIRD 1]  C/P:2,SFWA-P,P
MOON, ELIZABETH  TRADING IN DANGER [VATTA’S WAR 1]  H
MOON, ELIZABETH  WINNING COLORS [HERIS SERRANO 3]  C/P:2,SFWA-P,P
MOON, MARGARET/MOON, MAURINE  JUPITER EXPERIMENT, THE  P
MOON, SHELIA  KNEE DEEP IN THE THUNDER  LP

MOONEY, PHILIP F.

SCIENCE FICTION COLLECTIONS IN THE GEORGE ARENTS RESEARCH LIBRARY AT SYRACUSE UNIVERSITY

R/REF-H

MOONEY, RICHARD E.  COLONY: EARTH  [C/RAN-H,C/RAN-P]
MOONEY, RICHARD E.  GODS OF AIR AND DARKNESS  [C/RAN-P]

The MITSFS Pinkdex  September 24, 2007  Page 271
AUTHOR  TITLE  CODE
MOONEY, TED  EASY TRAVEL TO OTHER PLANETS  P
MOORCOCK, MICHAEL  ALIEN HEAT, AN [DANCERS AT THE END OF TIME 1]  H:2,P,SFWA
MOORCOCK, MICHAEL  BANE OF THE BLACK SWORD, THE [ELRIC 5]  C/P:2,P
MOORCOCK, MICHAEL  BARBARIANS OF MARS [MICHAEL KANE 3]  P
MOORCOCK, MICHAEL  BEFORE ARMAGEDDON, VOLUME 1  PA
MOORCOCK, MICHAEL  BEHOLD THE MAN [KARL GLOGAUER 1]  C/P:2,P
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS, THE  PA,C/PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 2, THE  PA,C/PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 3, THE  C/PA,PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 4, THE  PA,C/PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 5, THE  C/PA,PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 6, THE  PA,C/PA
MOORCOCK, MICHAEL  BEST SF STORIES FROM NEW WORLDS 8, THE  SR-PA
MOORCOCK, MICHAEL  BLACK CORRIDOR, THE  H:2,P
MOORCOCK, MICHAEL  BLADES OF MARS [MICHAEL KANE 2]  P
MOORCOCK, MICHAEL  BLOOD [SECOND ETHER 1]  H
MOORCOCK, MICHAEL  BREAKFAST IN THE RUINS [KARL GLOGAUER 2]  H,P
MOORCOCK, MICHAEL  BROTHEL IN ROSENSTRASSE, THE [VON BEK]  LP
MOORCOCK, MICHAEL  BULL AND THE SPEAR, THE [CORUM 4]  C/P:2,P
MOORCOCK, MICHAEL  CHAMPION OF GARATHORM, THE [CASTLE BRASS 2]  C/P,P
MOORCOCK, MICHAEL  CHRONICLES OF CORUM, THE [CORUM 1,2,3]  P
MOORCOCK, MICHAEL  CITY IN THE AUTUMN STARS, THE [VON BEK 2]  C/P,SFWA,SFWA-P,P
MOORCOCK, MICHAEL  CITY OF THE BEAST, THE [MICHAEL KANE 1]  C/P:2,P
MOORCOCK, MICHAEL  CORNELIUS CHRONICLES, THE [CORNELIUS 1,2,3,4]  C/P,P
MOORCOCK, MICHAEL  CORNELIUS CHRONICLES II, THE [CORNELIUS 5,6]  P
MOORCOCK, MICHAEL  CORNELIUS CHRONICLES III, THE [CORNELIUS 7,8]  P
MOORCOCK, MICHAEL  COUNT BRASS [CASTLE BRASS 1]  C/P,P
MOORCOCK, MICHAEL  CURE FOR CANCER, A [CORNELIUS 2]  H
MOORCOCK, MICHAEL/JAMES, PHILIP  DISTANT SUNS, THE  H,SR-LP
MOORCOCK, MICHAEL  DRAGON IN THE SWORD, THE [EREKOSE 3/ETERNAL CHAMPION 3]  C/P,SFWA-P,P,H
MOORCOCK, MICHAEL  DREAMING CITY, THE [ELRIC]  C/P:2,P
MOORCOCK, MICHAEL  DREAMTHIEF’S DAUGHTER, THE  H,P
MOORCOCK, MICHAEL  DYING FOR TOMORROW  C/P,P
MOORCOCK, MICHAEL  EARL AUBEC AND OTHER STORIES [ETERNAL CHAMPION]  P
MOORCOCK, MICHAEL  ELRIC AT THE END OF TIME [ELRIC]  C/P:2,P
MOORCOCK, MICHAEL  ELRIC OF MELNIBONE [ELRIC 1]  P
MOORCOCK, MICHAEL  ELRIC OF MELNIBONE (OMNIBUS EDITION) [ELRIC]  P
MOORCOCK, MICHAEL  ELRIC SAGA PART I, THE [ELRIC 1,2,3]  H:2
MOORCOCK, MICHAEL  ELRIC SAGA PART II, THE [ELRIC 4,5,6]  H:2
MOORCOCK, MICHAEL  ELRIC SAGA PART III, THE [ELRIC]  H
MOORCOCK, MICHAEL  END OF ALL SONGS, THE [DANCERS AT THE END OF TIME 3]  H:2,P,SFWA
MOORCOCK, MICHAEL  ENGLISH ASSASSIN, THE [CORNELIUS 3]  H
MOORCOCK, MICHAEL  ETERNAL CHAMPION, THE [EREKOSE 1/ETERNAL CHAMPION 1]  C/P,SFWA,SFWA-P,P
MOORCOCK, MICHAEL  ETERNAL CHAMPION, THE (OMNIBUS) [ETERNAL CHAMPION]  LP
MOORCOCK, MICHAEL  FABULOUS HARBORS [SECOND ETHER 2]  H,LP
MOORCOCK, MICHAEL  FINAL PROGRAMME, THE [CORNELIUS 1]  C/P:2,P
MOORCOCK, MICHAEL  FIRECLOWN, THE  SR-P,C/P:2
MOORCOCK, MICHAEL  FORTRESS OF THE PEARL, THE [ELRIC]  H:2,SFWA,SFWA-P:2,P
MOORCOCK, MICHAEL  GLORIANA, OR THE UNFULFILL’D QUEEN  H:2,LP
MOORCOCK, MICHAEL  GOLDEN BARGE, THE  C/P,P
MOORCOCK, MICHAEL  HAWKMOON [RUNESTAFF 1,2,3,4]  H
MOORCOCK, MICHAEL  HOLLOW LANDS, THE [DANCERS AT THE END OF TIME 2]  H:2,SFWA,P
MOORCOCK, MICHAEL  ICE SCHOONER, THE  H:2,SFWA,P
MOORCOCK, MICHAEL  JEWEL IN THE SKULL, THE [RUNESTAFF 1]  C/P:2,P
MOORCOCK, MICHAEL  KING OF THE SWORDS, THE [CORUM 3]  C/P:2,P
MOORCOCK, MICHAEL  KNIGHT OF THE SWORDS, THE [CORUM 1]  C/P:2,P
MOORCOCK, MICHAEL  LAND LEVIATHAN, THE [OSWALD BASTABLE 2]  C/P,SFWA,P,H
MOORCOCK, MICHAEL  LAUGHTER OF CARTHAGE, THE [PYAT 2]  H
MOORCOCK, MICHAEL  LEGENDS FROM THE END OF TIME [DANCERS AT THE END OF TIME 4]  H,SFWA,P

MOORCOCK, MICHAEL LIVES AND TIMES OF JERRY CORNELIUS, THE (1979 EDITION)

[CORNELIUS 5]

P

MOORCOCK, MICHAEL LIVES AND TIMES OF JERRY CORNELIUS, THE (2003 EDITION)

[CORNELIUS 5]

LP

MOORCOCK, MICHAEL  LORD OF THE SPIDERS, THE [MICHAEL KANE 2]  C/P,P
MOORCOCK, MICHAEL  MAD GOD’S AMULET, THE [RUNESTAFF 2]  C/P:2,P
MOORCOCK, MICHAEL  MASTERS OF THE PIT OR BARBARIANS OF MARS [MICHAEL KANE 3]  C/P:2,P
MOORCOCK, MICHAEL  MESSIAH AT THE END OF TIME, A [DANCERS AT THE END OF TIME 5]  C/P:2,P
MOORCOCK, MICHAEL  MOTHER LONDON  LP
MOORCOCK, MICHAEL  NEW WORLDS 1  SR-PA
MOORCOCK, MICHAEL  NEW WORLDS 2  SR-PA
MOORCOCK, MICHAEL  NEW WORLDS 3  SR-PA
MOORCOCK, MICHAEL  NEW WORLDS 5  SR-PA
MOORCOCK, MICHAEL/PLATT, CHARLES  NEW WORLDS #5  LPA
MOORCOCK, MICHAEL  NEW WORLDS QUARTERLY #1  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 272
AUTHOR  TITLE  CODE
MOORCOCK, MICHAEL  NEW WORLDS QUARTERLY #2  C/PA,PA
MOORCOCK, MICHAEL  NEW WORLDS QUARTERLY #3  PA,C/PA
MOORCOCK, MICHAEL  NEW WORLDS QUARTERLY #4  C/PA,PA
MOORCOCK, MICHAEL  NOMAD OF TIME, THE  H:2
MOORCOCK, MICHAEL  OAK AND THE RAM, THE [CORUM 5]  C/P,P,H
MOORCOCK, MICHAEL  OPIUM GENERAL, THE  H
MOORCOCK, MICHAEL  PAWN OF CHAOS [ETERNAL CHAMPION]  LPA
MOORCOCK, MICHAEL  PHOENIX IN OBSIDIAN [ETERNAL CHAMPION 2]  SR-P
MOORCOCK, MICHAEL  QUEEN OF THE SWORDS, THE [CORUM 2]  C/P:2,P
MOORCOCK, MICHAEL  QUEST FOR TANELORN, THE [CASTLE BRASS 3]  C/P:2,P
MOORCOCK, MICHAEL  QUEST FOR TANELORN, THE (BRITISH) [CASTLE BRASS 3]  SR-P
MOORCOCK, MICHAEL  REVENGE OF THE ROSE, THE [ELRIC]  H:2,SFWA:2,P
MOORCOCK, MICHAEL  RITUALS OF INFINITY, THE  C/P,P
MOORCOCK, MICHAEL  ROADS BETWEEN THE WORLDS, THE  LP
MOORCOCK, MICHAEL  RUNESTAFF, THE [RUNESTAFF 4]  C/P:2,P
MOORCOCK, MICHAEL  SAILING TO UTOPIA (OMNIBUS EDITION)  P
MOORCOCK, MICHAEL  SAILOR ON THE SEAS OF FATE, THE [ELRIC 2]  C/P:2,P
MOORCOCK, MICHAEL  SECRET OF THE RUNESTAFF, THE [RUNESTAFF 4]  P
MOORCOCK, MICHAEL  SF REPRISE 5  PA
MOORCOCK, MICHAEL  SHORES OF DEATH, THE  C/P,P
MOORCOCK, MICHAEL  SILVER WARRIORS, THE [EREKOSE 2]  C/P:2,P

MOORCOCK, MICHAEL/ CONSTANTINE, STORM SILVERHEART

H

MOORCOCK, MICHAEL  SINGING CITADEL, THE  P
MOORCOCK, MICHAEL  SKRAYLING TREE, THE [ELRIC/VON BEK]  P
MOORCOCK, MICHAEL  SLEEPING SORCERESS, THE [ELRIC]  C/P:2,P
MOORCOCK, MICHAEL  SOJAN  SR-LP
MOORCOCK, MICHAEL  SORCERER’S AMULET [RUNESTAFF 2]  P
MOORCOCK, MICHAEL  STEALER OF SOULS, THE [ELRIC]  SR-H,C/P:2
MOORCOCK, MICHAEL  STEEL TSAR, THE [OSWALD BASTABLE 3]  C/P,P
MOORCOCK, MICHAEL  STORMBRINGER [ELRIC]  SR-H,C/P:2,P
MOORCOCK, MICHAEL  STORMBRINGER (REVISED) [ELRIC 6]  C/P,P
MOORCOCK, MICHAEL  SUNDERED WORLDS, THE  C/P:2,P
MOORCOCK, MICHAEL  SWORD AND THE STALLION, THE [CORUM 6]  C/P,P,H
MOORCOCK, MICHAEL  SWORD OF THE DAWN [RUNESTAFF 3]  C/P:2,P
MOORCOCK, MICHAEL  SWORDS TRILOGY, THE [SWORDS 1,2,3]  C/P,P,H
MOORCOCK, MICHAEL  TIME DWELLER, THE  C/P:2,P

MOORCOCK, MICHAEL/ BUTTERWORTH, MICHAEL TIME OF THE HAWKLORDS, THE

C/P:2,P

MOORCOCK, MICHAEL  TWILIGHT MAN, THE  C/P:2,P
MOORCOCK, MICHAEL  VANISHING TOWER, THE [ELRIC 4]  C/P,P
MOORCOCK, MICHAEL  VELIERO DEI GHIACCI, IL  FP
MOORCOCK, MICHAEL  VON BEK (EXPANDED OMNIBUS EDITION) [VON BEK]  LP
MOORCOCK, MICHAEL  VON BEK (OMNIBUS EDITION) [VON BEK]  P
MOORCOCK, MICHAEL  WAR AMONGST THE ANGELS, THE [SECOND ETHER 3]  H,LP
MOORCOCK, MICHAEL  WAR HOUND AND THE WORLD’S PAIN, THE [VON BEK 1]  C/P,P,H
MOORCOCK, MICHAEL  WARLORD OF THE AIR, THE [OSWALD BASTABLE 1]  C/P:2,P
MOORCOCK, MICHAEL  WARRIORS OF MARS [MICHAEL KANE 1]  P
MOORCOCK, MICHAEL  WEIRD OF THE WHITE WOLF, THE [ELRIC 3]  C/P:2,P
MOORCOCK, MICHAEL  WINDS OF LIMBO, THE  C/P:2,P
MOORCOCK, MICHAEL  WRECKS OF TIME, THE  DH036
MOORE, ALAN/LLOYD, DAVID  V FOR VENDETTA  CX
MOORE, C. L.  BEST OF C.L. MOORE, THE  H:2,P
MOORE, C. L.  BLACK GOD’S SHADOW  SR-H
MOORE, C. L.  DOOMSDAY MORNING  H:2,P
MOORE, C. L./KUTTNER, HENRY  EARTH’S LAST CITADEL  C/P:2,P
MOORE, C. L.  JIREL OF JOIRY  H:2,P
MOORE, C. L.  JUDGMENT NIGHT  C/P,P
MOORE, C. L.  JUDGMENT NIGHT (COLLECTIONS)  C/P:2,P
MOORE, C. L.  NORTHWEST SMITH  C/P,P
MOORE, C. L.  SHAMBLEAU  GN5.31,C/P,SR-P
MOORE, C. L.  SHAMBLEAU AND OTHERS  SR-H,H
MOORE, C. L.  THERE SHALL BE DARKNESS  B
MOORE, C. L.  VINTAGE SEASON [VINTAGE SEASON 1]  SFWA-TD18,TD18
MOORE, CHRISTOPHER  BLOODSUCKING FIENDS: A LOVE STORY  LP
MOORE, CHRISTOPHER  COYOTE BLUE  LP
MOORE, CHRISTOPHER  FLUKE  LP
MOORE, CHRISTOPHER  ISLAND OF THE SEQUINED LOVE NUN  LP
MOORE, CHRISTOPHER  LUST LIZARD OF MELANCHOLY COVE, THE  H,LP
MOORE, CHRISTOPHER  PRACTICAL DEMONKEEPING  H,P
MOORE, DOREEN  BRIDE OF THE DARK CASTLE  P
MOORE, DORINE  BRIDE OF THE DARK CASTLE  P
MOORE, DORINE  CASTLE HOHENFELS  P
MOORE, HARRIS  MARROW EATERS, THE  P

AUTHOR  TITLE  CODE
MOORE, HARRIS  SLATER’S PLANET  C/P:2,P
MOORE, ISABEL  DAY THE COMMUNISTS TOOK OVER AMERICA, THE  SR-P
MOORE, JAMES A.  CHAOS BLEEDS [@BUFFY]  P
MOORE, JAMES A.  FIREWORKS  LP
MOORE, JAMES A.  HELL-STORM [@WORLD OF DARKNESS WEREWOLF]  P
MOORE, JAMES A./MURPHY, KEVIN  HOUSE OF SECRETS [WORLD OF DARKNESS]  P
MOORE, JAMES A.  UNDER THE OVERTREE  LP
MOORE, JOHN  BAD PRINCE CHARLIE  P
MOORE, JOHN  FATE WORSE THAN DRAGONS, A  P
MOORE, JOHN  HEROICS FOR BEGINNERS  C/P,P
MOORE, JOHN  SLAY AND RESCUE  C/P:2,SFWA-P,P
MOORE, JOHN  UNHANDSOME PRINCE, THE  H
MOORE, MOIRA J.  HERO STRIKES BACK, THE [SHIELD AND SOURCE 2]  P
MOORE, MOIRA J.  RESENTING THE HERO [SHIELD AND SOURCE 1]  P
MOORE, PATRICK  EARTH SATELLITES  [C/RAN-H]
MOORE, PATRICK  SPY IN SPACE  P
MOORE, RAYLYN  WHAT HAPPENED TO EMILY GOODE AFTER THE GREAT EXHIBITION  H,LP
MOORE, ROBIN/GRIFFIN, GERALD G.  DEATH DISCIPLE, THE  P
MOORE, ROBIN/GRIFFIN, GERALD G.  LAST COMING, THE  P

MOORE, ROGER E.

ERRAND OF MERCY [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #4]

P

MOORE, ROGER E.  MAELSTROM’S EYE, THE [@SPELLJAMMER CLOAKMASTER #3]  P
MOORE, SEAN A.  CONAN AND THE GRIM GREY GOD [CONAN]  LP
MOORE, SEAN A.  CONAN AND THE SHAMAN’S CURSE [CONAN]  P
MOORE, SHIRLEY  SCIENCE PROJECTS HANDBOOK  [C/RAN-P]
MOORE, SILAS  SCARLET ARENA 30303  H
MOORE, STUART  AMERICAN MEAT [@DARK FUTURE AMERICAN MEAT 1]  P
MOORE, THOMAS  EPICUREAN, THE  H
MOORE, VANCE  ODYSSEY [@MAGIC: THE GATHERING ODYSSEY 1]  P
MOORE, VANCE  PROPHECY [@MAGIC: THE GATHERING MASQUERADE 3]  P
MOORE, WALLACE  BLOOD STONES, THE [BALZAN 1]  P
MOORE, WALLACE  CAVES OF MADNESS, THE [BALZAN 2]  P
MOORE, WALLACE  LIGHTS OF ZETAR, THE [BALZAN 3]  P
MOORE, WARD  BRING THE JUBILEE  H:2,P
MOORE, WARD/BRADFORD, ROBERT  CADUCEUS WILD  C/P,P
MOORE, WARD  GREENER THAN YOU THINK  H:2
MOORE, WARD  GREENER THAN YOU THINK (ABRIDGED)  C/P:2,P
MOORE, WARD/DAVIDSON, AVRAM  JOYLEG  C/P,P,H
MOORE, WARD  LOT & LOT’S DAUGHTER  LP
MOOSER, STEPHEN  MONSTER EXPRESS  P
MOR, CAISEAL  KING OF SLEEP, THE [@WATCHERS #2]  P
MOR, CAISEAL  MEETING OF THE WATERS, THE [@WATCHERS #1]  P
MOR, CAISEAL  RAVEN GAME, THE [@WATCHERS #3]  P
MORAN, DANIEL  FLAME KEY, THE [KEYS TO PARADISE 1]  C/P,P
MORAN, DANIEL KEYS  ARMAGEDDON BLUES, THE  C/P,P
MORAN, DANIEL KEYS  EMERALD EYES [CONTINUING TIME 1]  P
MORAN, DANIEL KEYS  LAST DANCER, THE [CONTINUING TIME 3]  C/P,P
MORAN, DANIEL KEYS  LONG RUN, THE [CONTINUING TIME 2]  C/P,P
MORAN, DANIEL KEYS  RING, THE  H
MORAN, DANIEL KEYS/MORAN, JODI  TERMINAL FREEDOM  SR-H
MORAN, RICHARD  COLD SEA RISING  H
MORAN, RICHARD  EARTH WINTER [NEW ICE AGE 2]  P
MORAN, RICHARD  EMPIRE OF ICE, THE [NEW ICE AGE 1]  H
MORE, SIR THOMAS  UTOPIA  C/P,P
MOREHOUSE, LYDA  APOCALYPSE ARRAY [ARCHANGEL 4]  P
MOREHOUSE, LYDA  ARCHANGEL PROTOCOL [ARCHANGEL 1]  P
MOREHOUSE, LYDA  FALLEN HOST [ARCHANGEL 2]  P
MOREHOUSE, LYDA  MESSIAH NODE [ARCHANGEL 3]  P
MORENO, PEPE  BATMAN: DIGITAL JUSTICE  CX
MORGAN, CHRIS  FRITZ LEIBER: A BIBLIOGRAPHY 1934–1979  R/IND
MORGAN, DAN  CEE-TEE MAN  SR-P
MORGAN, DAN  GUITAR  [C/RAN-P]
MORGAN, DAN  HIGH DESTINY, THE  C/P,P
MORGAN, DAN  INSIDE  C/P:2,P
MORGAN, DAN  MIND TRAP  C/P:2,P
MORGAN, DAN/KIPPAX, JOHN  NEUTRAL STARS, THE  C/P:2,P
MORGAN, DAN  NEW MINDS, THE  C/P,P
MORGAN, DAN/KIPPAX, JOHN  SEED OF STARS  C/P:2,P
MORGAN, DAN  SEVERAL MINDS, THE  SR-P,C/P
MORGAN, DAN/KIPPAX, JOHN  THUNDER OF STARS, A  C/P:2,P
MORGAN, DAREION  SOPHIE’S LEGACY  LP
MORGAN, DE WOLFE  GREAT PYRAMID MYSTERY  P
MORGAN, J. M.  BETWEEN THE DEVIL AND THE DEEP  P
MORGAN, KATHLEEN  GIVER OF ROSES [@GUARDIANS OF GADIEL]  LP

AUTHOR  TITLE  CODE
MORGAN, RICHARD  ALTERED CARBON [TAKESHI KOVACS 1]  H
MORGAN, RICHARD  BROKEN ANGELS [TAKESHI KOVACS 2]  H,LP
MORGAN, RICHARD  MARKET FORCES  H
MORGAN, RICHARD  MARKET FORCES (PROOF)  SR-LP
MORGAN, RICHARD  WOKEN FURIES [TAKESHI KOVACS 3]  H
MORGAN, RICHARD K.  THIRTEEN  H
MORGAN, ROBERT  ALL THINGS UNDER THE MOON [TEDDY LONDON,P.I.]  P
MORGENSTERN, S.  SILENT GONDOLIERS, THE (PROOF)  SR-LP
MORIARTY, CHRIS  SPIN CONTROL  LP
MORIARTY, CHRIS  SPIN STATE  H,LP
MORIARTY, CHRIS  SPIN STATE (PROOF)  SR-LP
MORIOKA, HIROYUKI  PRINCESS OF THE STARS [SEIKAI CREST OF THE STARS #1]  LP
MORIYAMA, DAISUKE  CHRONO CRUSADE, VOL. 1 [CHRONO CRUSADE 1]  SCX
MORLAN, A. R.  AMULET, THE  P
MORLAN, A. R.  DARK JOURNEY  P
MORLAND, ALANNA  LEOPARD LORD  P
MORLAND, ALANNA  SHACKLE AND SWORD  P
MORLEY, HENRY  IDEAL COMMONWEALTHS  SR-HA
MORMON  BOOK OF MORMON, THE  PA
MOROZ, ANNE  NO SAFE PLACE  C/P,P
MORR, KENYON  KINGDOM OF SORROW [@KING’S QUEST 2]  P
MORR, KENYON  SEE NO WEEVIL [@KING’S QUEST 3]  P
MORRELL, DAVID  TOTEM, THE  P
MORRESSY, JOHN  DUDGEON & DRAGONS [@KEDRIGERN CHRONICLES #2]  LP
MORRESSY, JOHN  EXTRATERRITORIAL, THE  SR-P
MORRESSY, JOHN  FROSTWORLD AND DREAMFIRE  C/P,SFWA,P,H:2
MORRESSY, JOHN  GRAYMANTLE [IRON ANGEL 2]  C/P:2,P
MORRESSY, JOHN  HUMANS OF ZIAX II, THE  H
MORRESSY, JOHN  IRONBRAND [IRON ANGEL 1]  C/P,P
MORRESSY, JOHN  KEDRIGERN AND THE CHARMING COUPLE [KEDRIGERN]  P
MORRESSY, JOHN  KEDRIGERN IN WONDERLAND [KEDRIGERN 3]  P
MORRESSY, JOHN  KINGSBANE [IRON ANGEL 3]  P
MORRESSY, JOHN  LAW FOR THE STARS, A  SR-P,C/P
MORRESSY, JOHN  MANSIONS OF SPACE, THE  P
MORRESSY, JOHN  NAIL DOWN THE STARS  H,SFWA
MORRESSY, JOHN  QUESTING OF KEDRIGERN, THE [KEDRIGERN 2]  C/P,P
MORRESSY, JOHN  REMEMBRANCE FOR KEDRIGERN, A [KEDRIGERN]  P
MORRESSY, JOHN  STARBRAT  C/P,P,H
MORRESSY, JOHN  STARDRIFT  C/P,P
MORRESSY, JOHN  TIME OF THE ANNIHILATOR, THE  P
MORRESSY, JOHN  UNDER A CALCULATING STAR  C/P,SFWA,P,H
MORRESSY, JOHN  VOICE FOR PRINCESS, A [KEDRIGERN 1]  C/P,P
MORRIS, HENRY M.  REMARKABLE BIRTH OF PLANET EARTH, THE  [C/RAN-P]
MORRIS, JANET/DRAKE, DAVID  ACTIVE MEASURES  C/P,P
MORRIS, JANET  AFTERWAR  PA
MORRIS, JANET  ANGELS IN HELL [HEROES IN HELL]  PA,SFWA-P
MORRIS, JANET  BEYOND SANCTUARY [THIEVES’ WORLD TEMPUS 1]  H:2,P
MORRIS, JANET  BEYOND THE VEIL [THIEVES’ WORLD TEMPUS 2]  C/P,P,H
MORRIS, JANET  BEYOND WIZARDWALL [THIEVES’ WORLD TEMPUS]  H:2,SFWA-P,P
MORRIS, JANET  CARNELIAN THRONE, THE [SILISTRA 4]  C/P,P
MORRIS, JANET/MORRIS, CHRIS  CITY AT THE EDGE OF TIME [THIEVES’ WORLD TEMPUS 5]  C/P:2,SFWA-P,P
MORRIS, JANET  CRUISER DREAMS [KERRION EMPIRE 2]  H:2,P
MORRIS, JANET  CRUSADERS IN HELL [HEROES IN HELL]  PA
MORRIS, JANET  DREAM DANCER [KERRION EMPIRE 1]  H:2,SFWA,P
MORRIS, JANET  EARTH DREAMS [KERRION EMPIRE 3]  C/P,P,H
MORRIS, JANET/MORRIS, CHRIS  40-MINUTE WAR, THE  H,P
MORRIS, JANET  GOLDEN SWORD, THE [SILISTRA 2]  C/P,P
MORRIS, JANET  HEROES IN HELL [HEROES IN HELL]  PA,C/PA
MORRIS, JANET  HIGH COUCH OF SILISTRA [SILISTRA 1]  C/P:2,P
MORRIS, JANET  I, THE SUN [HATTI EMPIRE]  P
MORRIS, JANET/DRAKE, DAVID  KILL RATIO  P
MORRIS, JANET/MORRIS, CHRIS  LITTLE HELLIAD, THE [HEROES IN HELL]  C/P,SFWA-P,P
MORRIS, JANET  MASTERS IN HELL [HEROES IN HELL]  PA
MORRIS, JANET/MORRIS, CHRIS  MEDUSA  C/P:2,P
MORRIS, JANET/MORRIS, CHRIS  OUTPASSAGE  C/P:2,P
MORRIS, JANET  PROPHETS IN HELL [HEROES IN HELL]  PA,C/PA,SFWA-P
MORRIS, JANET  REBELS IN HELL [HEROES IN HELL]  PA,C/PA
MORRIS, JANET  RETURNING CREATION [SILISTRA (REVISED) 1]  C/P,P
MORRIS, JANET/MORRIS, CHRIS  STALK, THE  P
MORRIS, JANET/MORRIS, CHRIS  STORM SEED [THIEVES’ WORLD TEMPUS 7]  P
MORRIS, JANET/DRAKE, DAVID  TARGET  C/P,SFWA-P,P
MORRIS, JANET  TEMPUS [THIEVES’ WORLD TEMPUS 4]  C/P:2,P
MORRIS, JANET/MORRIS, CHRIS  TEMPUS UNBOUND [THIEVES’ WORLD TEMPUS 6]  C/P:2,P
MORRIS, JANET/MORRIS, CHRIS  THRESHOLD  H:2,SFWA-P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 275
AUTHOR  TITLE  CODE
MORRIS, JANET/MORRIS, CHRIS  TRUST TERRITORY  H:2,P
MORRIS, JANET  WAR IN HELL [HEROES IN HELL]  C/PA,PA,SFWA-P
MORRIS, JANET  WARLORD!  C/P,P
MORRIS, JANET  WIND FROM THE ABYSS [SILISTRA 3]  C/P:3,P
MORRIS, JIM  SHERIFF OF PURGATORY, THE  SFWA
MORRIS, JIM  SPURLOCK: SHERIFF OF PURGATORY  P
MORRIS, JONATHAN  ANACHROPHOBIA [DOCTOR WHO EIGHTH DOCTOR]  P
MORRIS, JONATHAN  FESTIVAL OF DEATH [DOCTOR WHO PAST DOCTORS]  C/P,P
MORRIS, JONATHAN  TOMORROW WINDOWS, THE [DOCTOR WHO EIGHTH DOCTOR]  P
MORRIS, KENNETH  CHALCHIUHITE DRAGON, THE  H
MORRIS, KENNETH  FATES OF THE PRINCES OF DYFED, THE  LP
MORRIS, MARK  BODYSNATCHERS, THE [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
MORRIS, MARK  DEEP BLUE [DOCTOR WHO PAST DOCTORS]  C/P,P
MORRIS, MARK  IMMACULATE, THE  P
MORRIS, TEE  BILLIBUB BADDINGS AND THE CASE OF THE SINGING SWORD  LP
MORRIS, TEE  LEGACY OF MOREVI [@ARATHELLEAN WARS #1]  LP
MORRIS, WILLIAM  CHILD CHRISTOPHER AND GOLDILIND THE FAIR  LP
MORRIS, WILLIAM  GLITTERING PLAIN, THE  H,LP
MORRIS, WILLIAM  GOLDEN WINGS AND OTHER STORIES  LP
MORRIS, WILLIAM  HOUSE OF THE WOLFINGS, THE  LP
MORRIS, WILLIAM  ROOTS OF THE MOUNTAINS, THE  LP
MORRIS, WILLIAM  SUNDERING FLOOD, THE  C/P,P
MORRIS, WILLIAM  WATER OF THE WONDROUS ISLES, THE  C/P:2,P
MORRIS, WILLIAM  WELL AT THE WORLD’S END, THE  P
MORRIS, WILLIAM  WELL AT THE WORLD’S END 1, THE  C/P:2,P
MORRIS, WILLIAM  WELL AT THE WORLD’S END 2, THE  C/P:2,P
MORRIS, WILLIAM  WOOD BEYOND THE WORLD, THE  C/P,P,H
MORRISON, LOUISE D.  J.R R. TOLKIEN’S THE FELLOWSHIP OF THE RING  R/REF-P
MORROUGH, E. R.  TEMPLE SERVANT AND OTHER STORIES, THE  H
MORROW, JAMES  BLAMELESS IN ABADDON [CORPUS DEI 2]  H
MORROW, JAMES  CAT’S PAJAMAS & OTHER STORIES, THE  H
MORROW, JAMES  CITY OF TRUTH  H:2,LP
MORROW, JAMES  CONTINENT OF LIES, THE  C/P,P,H
MORROW, JAMES  ETERNAL FOOTMAN, THE [CORPUS DEI 3]  H
MORROW, JAMES  NEBULA AWARDS 26  HA,LPA,SFWA
MORROW, JAMES  NEBULA AWARDS 27  HA
MORROW, JAMES  NEBULA AWARDS 28  HA,LPA
MORROW, JAMES  ONLY BEGOTTEN DAUGHTER  SFWA-P,H:2,P
MORROW, JAMES/MORROW, KATHRYN  SFWA EUROPEAN HALL OF FAME, THE  HA
MORROW, JAMES  SWATTING AT THE COSMOS  SR-H
MORROW, JAMES  THIS IS THE WAY THE WORLD ENDS  H,SFWA-P
MORROW, JAMES  TOWING JEHOVAH [CORPUS DEI 1]  H:2,LP
MORROW, JAMES  WINE OF VIOLENCE, THE  C/P,P,H
MORSE, A. REYNOLDS  WORKS OF M. P. SHIEL, THE  R/REF-H
MORTIMORE, JIM  BELTEMPEST [DOCTOR WHO EIGHTH DOCTOR]  P
MORTIMORE, JIM  BLOOD HEAT [DOCTOR WHO NEW ADVENTURES]  C/P,P
MORTIMORE, JIM  CAMPAIGN [DOCTOR WHO PAST DOCTORS]  SR-P
MORTIMORE, JIM  CLARK’S LAW [@BABYLON 5 #4]  C/P:2,P
MORTIMORE, JIM  ETERNITY WEEPS [DOCTOR WHO NEW ADVENTURES]  C/P,P
MORTIMORE, JIM  EYE OF HEAVEN [DOCTOR WHO PAST DOCTORS]  C/P,P
MORTIMORE, JIM  PARASITE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P

MORTIMORE, JIM SWORD OF FOREVER, THE

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

C/P,P

MORWOOD, PETER  DEMON LORD, THE [BOOK OF YEARS 2]  C/P,P
MORWOOD, PETER  DRAGON LORD [BOOK OF YEARS 3]  P
MORWOOD, PETER  HORSE LORD, THE [BOOK OF YEARS 1]  P
MORWOOD, PETER/DUANE, DIANE  KEEPER OF THE CITY [@GUARDIANS OF THE THREE 2]  C/P,P
MORWOOD, PETER  RULES OF ENGAGEMENT [STAR TREK]  C/P:2,P
MORWOOD, PETER  WARLORD’S DOMAIN, THE [BOOK OF YEARS 4]  P
MOSCOE, MIKE  FIRST CASUALTY, THE [SOCIETY OF HUMANITY/UNITY PARTY 1]  C/P,P
MOSCOE, MIKE  FIRST DAWN [@LOST MILLENNIUM 1]  C/P,P
MOSCOE, MIKE  LOST DAYS [@LOST MILLENNIUM 3]  C/P,P
MOSCOE, MIKE  PATRIOT’S STAND [@MECHWARRIOR DARK AGE/BATTLETECH]  P
MOSCOE, MIKE  PRICE OF PEACE, THE [SOCIETY OF HUMANITY/UNITY PARTY 2]  C/P,P
MOSCOE, MIKE  SECOND FIRE [@LOST MILLENNIUM 2]  C/P,P
MOSCOE, MIKE  THEY ALSO SERVE [SOCIETY OF HUMANITY/UNITY PARTY 3]  C/P,P
MOSER, NANCY  SECOND TIME AROUND [TIME LOTTERY 2]  LP
MOSER, NANCY  TIME LOTTERY [TIME LOTTERY 1]  LP
MOSKOWITZ, SAM  A. MERRITT: REFLECTIONS IN THE MOON POOL  R/REF-H
MOSKOWITZ, SAM  COMING OF THE ROBOTS, THE  PA,C/PA
MOSKOWITZ, SAM  DOORWAY INTO TIME  PA,C/PA
MOSKOWITZ, SAM  EXPLORERS OF THE INFINITE  R/REF-H
MOSKOWITZ, SAM  EXPLORING OTHER WORLDS  PA,C/PA
MOSKOWITZ, SAM  FUTURES TO INFINITY  PA,C/PA

AUTHOR  TITLE  CODE
MOSKOWITZ, SAM/NORTON, ALDEN H.  HORRORS IN HIDING  PA
MOSKOWITZ, SAM  HORRORS UNKNOWN  HA,PA
MOSKOWITZ, SAM  HORRORS UNSEEN  C/PA,PA
MOSKOWITZ, SAM/ELWOOD, ROGER  HUMAN ZERO, THE  PA,C/PA
MOSKOWITZ, SAM  IMMORTAL STORM, THE  R/REF-H
MOSKOWITZ, SAM  MASTERPIECES OF SCIENCE FICTION  LPA
MOSKOWITZ, SAM  MICROCOSMIC GOD  PA
MOSKOWITZ, SAM  MODERN MASTERPIECES OF SCIENCE FICTION  H,LPA
MOSKOWITZ, SAM  MOON ERA, THE  PA
MOSKOWITZ, SAM/ELWOOD, ROGER  OTHER WORLDS, OTHER TIMES  C/PA,PA

MOSKOWITZ, SAM SCIENCE FICTION BY GASLIGHT: A HISTORY AND ANTHOLOGY OF SCIENCE FICTION IN THE POPULAR MAGAZINES, 1891–1911

HA

MOSKOWITZ, SAM SCIENCE FICTION IN OLD SAN FRANCISCO: HISTORY OF THE MOVEMENT 1854–1890

R/REF-H

MOSKOWITZ, SAM  SEEKERS OF TOMORROW  R/REF-P,C/REF-P:2
MOSKOWITZ, SAM  SENSE OF WONDER, A  PA
MOSKOWITZ, SAM  STRANGE HORIZONS  SFWA,R/REF-H
MOSKOWITZ, SAM/ELWOOD, ROGER  TIME CURVE, THE  PA,C/PA
MOSKOWITZ, SAM  UNDER THE MOONS OF MARS  HA,R/REF-H
MOSKOWITZ, SAM  VORTEX BLASTERS, THE  PA,C/PA
MOSKOWITZ, SAM  WHEN WOMEN RULE  HA
MOSLEY, WALTER  BLUE LIGHT  H
MOSLEY, WALTER  FUTURELAND  H,P
MOSLEY, WALTER  WAVE, THE  H
MOTZ, LLOYD  THIS IS OUTER SPACE  R/REF-P
MOUDY, WALTER  NO MAN ON EARTH  C/P:2,P
MROZEK, SLAWOMIR  ELEPHANT, THE  H
MUDD, STEVE  TANGLED WEBS  P
MUELLER, RICHARD  GHOSTBUSTERS  P
MUELLER, RICHARD  JERNIGAN’S EGG  LP
MUIR, JOHN KENNETH  FORSAKEN, THE [@SPACE: 1999]  P
MUJICA LAINEZ, MANUEL  WANDERING UNICORN, THE  H,P
MULLEN, PATRICIA  STONE MOVERS, THE  P
MULLER, JOHN E.  ALIEN  SR-P
MULLER, JOHN E.  BEYOND THE VOID  SR-P
MULLER, JOHN E.  BEYOND TIME  SR-P
MULLER, JOHN E.  DARK CONTINUUM  SR-P
MULLER, JOHN E.  DAY OF THE BEASTS  C/P,P
MULLER, JOHN E.  DAY THE WORLD DIED, THE  SR-P,C/P
MULLER, JOHN E.  EDGE OF ETERNITY  SR-P
MULLER, JOHN E.  EXORCISTS, THE  SR-P
MULLER, JOHN E.  EYE OF KARNAK, THE  SR-P
MULLER, JOHN E.  IN THE BEGINNING  SR-P
MULLER, JOHN E.  INFINITY MACHINE  SR-P
MULLER, JOHN E.  MAN FROM BEYOND, THE  SR-P
MULLER, JOHN E.  MAN WHO CONQUERED TIME, THE  SR-P
MULLER, JOHN E.  MARK OF THE BEAST  SR-P
MULLER, JOHN E.  MICRO INFINITY  SR-P
MULLER, JOHN E.  MOON ROCKET  H
MULLER, JOHN E.  NEGATIVE ONES, THE  SR-P
MULLER, JOHN E.  NIGHT OF THE BIG FIRE  SR-P
MULLER, JOHN E.  ORBIT ONE  SR-P
MULLER, JOHN E.  OUT OF THE NIGHT  SR-P
MULLER, JOHN E.  PERILOUS GALAXY  SR-P
MULLER, JOHN E.  PHENOMENA X  SR-P
MULLER, JOHN E.  REACTOR XK9  SR-P
MULLER, JOHN E.  RETURN OF ZEUS  SR-P
MULLER, JOHN E.  SPECIAL MISSION  SR-P
MULLER, JOHN E.  SPECTRE OF DARKNESS  SR-P
MULLER, JOHN E.  SURVIVAL PROJECT  SR-P
MULLER, JOHN E.  1000 YEARS ON, A  SR-P
MULLER, JOHN E.  UNINVITED, THE  SR-P
MULLER, JOHN E.  URANIUM 235  SR-P
MULLER, JOHN E.  VENGEANCE OF SIVA  SR-P
MULLER, JOHN E.  VENUS VENTURE, THE  SR-P
MULLER, JOHN E.  X-MACHINE, THE  SR-P
MULOCK, MISS  LITTLE LAME PRINCE, THE/ADVENTURES OF A BROWNIE, THE/POOR PRIN SR-H
MUNDY, TALBOT  CAESAR DIES  C/P,P
MUNDY, TALBOT  DEVIL’S GUARD, THE  P
MUNDY, TALBOT/WILLEFORD, CHARLES  FULL MOON/HIGH PRIEST OF CALIFORNIA  SR-P
MUNDY, TALBOT  HELENE [TROS OF SAMOTHRACE 4]  SR-P
MUNDY, TALBOT  HELMA [TROS OF SAMOTHRACE 2]  SR-P,C/P:2
MUNDY, TALBOT  JIMGRIM  C/P,SR-P,H
MUNDY, TALBOT  KING — OF THE KHYBER RIFLES  H,P
MUNDY, TALBOT  LIAFAIL [TROS OF SAMOTHRACE 3]  SR-P,C/P:2

The MITSFS Pinkdex  September 24, 2007  Page 277
AUTHOR  TITLE  CODE
MUNDY, TALBOT  NINE UNKNOWN, THE  SR-P,C/P,H
MUNDY, TALBOT  OM: THE SECRET OF AHBOR VALLEY  C/P,P
MUNDY, TALBOT  OM: THE SECRET OF AHBOR VALLEY (COVER MISSPELLED)  LP
MUNDY, TALBOT  PURPLE PIRATE, THE  SR-P
MUNDY, TALBOT  QUEEN CLEOPATRA  SR-P

MUNDY, TALBOT TROS OF SAMOTHRACE [TROS OF SAMOTHRACE 1]/ LUD OF LUNDUN [TROS OF SAMOTHRACE 1]

P

MUNN, H. WARNER  IN THE TOMB OF THE BISHOP [@TALES OF THE WEREWOLF CLAN #1]  SR-H
MUNN, H. WARNER  KING OF THE WORLD’S EDGE  C/P:2,P
MUNN, H. WARNER  MASTER GOES HOME, THE [@TALES OF THE WEREWOLF CLAN #2]  SR-H
MUNN, H. WARNER  MERLIN’S GODSON  C/P:2,P
MUNN, H. WARNER  MERLIN’S RING  SR-P,C/P:2
MUNN, H. WARNER  SHIP FROM ATLANTIS, THE  DG618
MUNN, H. WARNER  WEREWOLF OF PONKERT, THE  P
MUNRO, RONA  SURVIVAL [DOCTOR WHO]  SR-P
MUNSEY, TERENCE  FLIGHT OF THE STONEMAN’S SON, THE  P
MUNSON, BRAD  MAD THRONE, THE  P
MURAKAMI, HARUKI  HARD-BOILED WONDERLAND AND THE END OF THE WORLD  H
MURARI, TIMERI  OBLIVION TAPES, THE  P
MURDOCK, M. S.  ARMAGEDDON OFF VESTA [BUCK ROGERS MARTIAN WARS 3]  C/P,P
MURDOCK, M. S.  DYNTERYX  P
MURDOCK, M. S.  HAMMER OF MARS [BUCK ROGERS MARTIAN WARS 2]  C/P,P
MURDOCK, M. S.  PRIME SQUARED [BUCK ROGERS INNER PLANETS 2]  C/P:2,P
MURDOCK, M. S.  REBELLION 2456 [BUCK ROGERS MARTIAN WARS 1]  C/P,P
MURDOCK, M. S.  VENDETTA  P
MURDOCK, M. S.  WEB OF THE ROMULANS [STAR TREK]  C/P:3,P
MURILL, RAY W.  WAR DOGS OF THE GOLDEN HORDE [@MARS ATTACKS]  C/P,P
MURPHY, C. E.  COYOTE DREAMS [@WALKER #3]  LP
MURPHY, C. E.  THUNDERBIRD FALLS [WALKER 2]  H:2,LP
MURPHY, C. E.  URBAN SHAMAN [WALKER 1]  LP

MURPHY, PAT ADVENTURES IN TIME AND SPACE WITH MAX MERRIWELL

[MAX MERRIWELL 4]

H

MURPHY, PAT  CITY, NOT LONG AFTER, THE  C/P,P
MURPHY, PAT  FALLING WOMAN, THE  H,SFWA-P:2,P
MURPHY, PAT  NADYA  P
MURPHY, PAT  POINTS OF DEPARTURE  C/P,P
MURPHY, PAT  RACHEL IN LOVE  B
MURPHY, PAT  SHADOW HUNTER, THE  LP
MURPHY, PAT  THERE AND BACK AGAIN [MAX MERRIWELL 1]  P
MURPHY, PAT  WILD ANGEL [MAX MERRIWELL 2]  P
MURPHY, PATRICK D./HYLES, VERNON  POETIC FANTASTIC, THE  R/REF-H
MURPHY, ROBERT FRANKLIN  GIRL FACTORY, THE [GIRL FACTORY 1]  C/P,P
MURPHY, ROBERT FRANKLIN  KING’S MATE [GIRL FACTORY 2]  P
MURPHY, ROBERT FRANKLIN  MAN-MADE WOMAN, THE [GIRL FACTORY 3]  P
MURPHY, SHIRLEY ROUSSEAU  CASTLE OF HAPE, THE [YNELL 3]  H,P
MURPHY, SHIRLEY ROUSSEAU  CAT ON THE EDGE [JOE GREY 1]  P
MURPHY, SHIRLEY ROUSSEAU  CAT RAISE THE DEAD [JOE GREY 3]  P
MURPHY, SHIRLEY ROUSSEAU  CAT UNDER FIRE [JOE GREY 2]  C/P,P
MURPHY, SHIRLEY ROUSSEAU  CATSWOLD PORTAL, THE  C/P,SFWA:2,P,H
MURPHY, SHIRLEY ROUSSEAU  CAVES OF FIRE AND ICE [YNELL 4]  C/P,P
MURPHY, SHIRLEY ROUSSEAU  DRAGONBARDS, THE  P
MURPHY, SHIRLEY ROUSSEAU  JOINING OF THE STONE, THE [YNELL 5]  H,P
MURPHY, SHIRLEY ROUSSEAU  NIGHTPOOL  H
MURPHY, SHIRLEY ROUSSEAU  RING OF FIRE, THE [YNELL 1]  H,P
MURPHY, SHIRLEY ROUSSEAU  WOLF BELL, THE [YNELL 2]  C/P,P,H
MURPHY, WALTER F.  VICAR OF CHRIST, THE  P
MURPHY, WARREN/SAPIR, RICHARD BEN  REMO  P
MURRAY, DOUG  BLOOD RELATIONS [@WORLD OF DARKNESS VAMPIRE]  P
MURRAY, DOUG  CALL TO BATTLE [WORLD OF DARKNESS RAGE JAY NO-NAME 1]  P
MURRAY, WILL  DOC SAVAGE: REFLECTIONS IN BRONZE  R/REF-P
MURRAY, WILL  DUENDE HISTORY OF THE SHADOW MAGAZINE, THE  R/REF-H
MURRAY, WILL  SECRETS OF DOC SAVAGE  R/REF-P,C/REF-H
MYERS, BARRY  SHANEE, THE  P
MYERS, BARRY  SHANEE WOMAN  P
MYERS, EDWARD  FIRE AND ICE [@MOUNTAIN MADE OF LIGHT #2]  P,SFWA-P
MYERS, EDWARD  MOUNTAIN MADE OF LIGHT, THE [MOUNTAIN MADE OF LIGHT 1]  P
MYERS, EDWARD  SUMMIT, THE  P
MYERS, GARY  HOUSE OF THE WORM, THE  H
MYERS, HENRY  UTMOST ISLAND, THE  H
MYERS, HOWARD L.  CLOUD CHAMBER  P
MYERS, HOWARD L.  CREATURES OF MAN, THE  P
MYERS, JOHN MYERS  HARP AND THE BLADE, THE  P
MYERS, JOHN MYERS  MOON’S FIRE-EATING DAUGHTER, THE  H,LP
MYERS, JOHN MYERS  SILVERLOCK  SR-P,C/P:2

AUTHOR  TITLE  CODE
MYERS, ROBERT E.  JACK WILLIAMSON: A PRIMARY AND SECONDARY BIBLIOGRAPHY  R/IND
MYERS, ROBERT J.  CROSS OF FRANKENSTEIN, THE  H,P
MYERS, ROBERT J.  SLAVE OF FRANKENSTEIN, THE  P
MYERS, ROBERT J.  VIRGIN AND THE VAMPIRE, THE  P
MYERS, WALTER DEAN  BRAINSTORM  P
MYERS, WALTER DEAN  SHADOW OF THE RED MOON  P

AUTHOR  TITLE  CODE
NADER, GEORGE  CHROME  H,P
NADER, RALPH  UNSAFE AT ANY SPEED  [C/RAN-P]
NAGATA, LINDA  BOHR MAKER, THE  P
NAGATA, LINDA  DECEPTION WELL  P
NAGATA, LINDA  LIMIT OF VISION  H
NAGATA, LINDA  MEMORY  H
NAGATA, LINDA  TECH-HEAVEN  P
NAGATA, LINDA  VAST  C/P,P
NAHA, ED  GHOSTBUSTERS 2 [GHOSTBUSTERS]  P
NAHA, ED  PARADISE PLOT, THE  P
NAHA, ED  ROBOCOP  C/P,P
NAHA, ED  SUICIDE PLAGUE, THE  C/P,P
NANOVIC, JOHN L.  DOC SAVAGE, SUPREME ADVENTURER  B
NASA  ELEVENTH SEMIANNUAL REPORT TO CONGRESS 1964  [C/RAN-H]
NASA  MARINER-MARS 1964/FINAL PROJECT REPORT  [C/RAN-H]
NASA  NASA-UNIVERSITY CONFERENCE 1  [C/RAN-H]
NASA  NASA-UNIVERSITY CONFERENCE 2  [C/RAN-H]
NASA  RESULTS OF THE 1ST US MANNED ORBITAL SPACE FLIGHT  [C/RAN-H]
NASA  RESULTS OF THE 2ND US MANNED SUBORBITAL SPACE FLIGHT  [C/RAN-H]

NASA STUDY OF THE PERFORMANCE OF AN ASTRONAUT DURING INGRESS AND EGRESS MANEUVERS THROUGH AIRLOCKS AND PASSAGEWAYS, A

[C/RAN-H]

NASH, MARY  MRS. COVERLET’S MAGICIANS  P
NASIR, JAMIL  DISTANCE HAZE  P
NASIR, JAMIL  HIGHER SPACE, THE  P
NASIR, JAMIL  QUASAR  C/P,P
NASIR, JAMIL  TOWER OF DREAMS  P
NATAS, ETH  EXORCISM  P
NATHAN, ROBERT  BARLY FIELDS, THE  H
NATHAN, ROBERT  BUT GENTLY DAY  H
NATHAN, ROBERT  PORTRAIT OF JENNIE  P
NATHAN, ROBERT  RIVER JOURNEY, THE  H
NATHAN, ROBERT  SIR HENRY  H,LP
NATHAN, ROBERT  STONECLIFF  P
NATHAN, ROBERT  WEANS, THE  H
NATHENSON, JOSEPH  DEEP, VERY DEEP SPACE  C/P,P
NATIONAL COMMISSION ON SPACE  PIONEERING THE SPACE FRONTIER  R/REF-H
NAVARRO, YVONNE  AFTERAGE  P
NAVARRO, YVONNE  BROKEN SUNRISE [@BUFFY WICKED WILLOW #3]  C/P,P
NAVARRO, YVONNE  DARKENING, THE [@BUFFY WICKED WILLOW #1]  C/P,P
NAVARRO, YVONNE  FINAL IMPACT  C/P,P
NAVARRO, YVONNE  HELLBOY  P
NAVARRO, YVONNE  MUSIC OF THE SPEARS [@ALIENS 8]  P
NAVARRO, YVONNE  PALEO [@BUFFY]  C/P,P
NAVARRO, YVONNE  RED SHADOWS  P
NAVARRO, YVONNE  SHATTERED TWILIGHT [@BUFFY WICKED WILLOW #2]  P
NAVARRO, YVONNE  SPECIES  P
NAVARRO, YVONNE  TEMPTED CHAMPIONS [@BUFFY]  C/P,P
NAVARRO, YVONNE  WILLOW FILES, THE, VOL. 1 [@BUFFY]  P
NAVARRO, YVONNE  WILLOW FILES, THE, VOL. 2 [@BUFFY]  P
NAYLOR, GRANT  BETTER THAN LIFE [@RED DWARF 2]  C/P:2,P
NAYLOR, GRANT  INFINITY WELCOMES CAREFUL DRIVERS [@RED DWARF 1]  C/P:2,SFWA-P,P
NAYLOR, GRANT  RED DWARF [RED DWARF 1,2]  H
NEARING, H. JR.  SINISTER RESEARCHES OF C. P. RANSOM, THE  P
NEASON, REBECCA  GUISES OF THE MIND [STAR TREK TNG]  C/P:2,P
NEASON, REBECCA  PATH, THE [@HIGHLANDER]  P
NEASON, REBECCA  SHADOW OF OBSESSION [@HIGHLANDER]  P
NEASON, REBECCA  THIRTEENTH SCROLL, THE [AGHAMORE 2]  P
NEASON, REBECCA  TRUEST POWER, THE [AGHAMORE 1]  P
NEBEL, LONG JOHN  WAY OUT WORLD, THE  [C/RAN-P]
NECKER, CLAIRE  SUPERNATURAL CATS  PA,SFWA
NEEPER, CARY  PLACE BEYOND MAN, A  C/P,SFWA,P,H
NEIDERMAN, ANDREW  BABY SQUAD, THE  P
NEIDERMAN, ANDREW  LOVE CHILD  P,SFWA-P
NEILL, A. S.  LAST MAN ALIVE, THE  H,P
NEILL, ROBERT  WITCH BANE  P
NELSON DOUBLEDAY, INC.  SCIENCE SERVICE SCIENCE PROGRAM (4 BOOKLETS, BOXED)  [C/RAN-H]
NELSON, RAY  VIRTUAL ZEN  P
NELSON, RAY FARADAY  BLAKE’S PROGRESS  SR-P
NELSON, RAY FARADAY  ECOLOG, THE  SR-P
NELSON, RAY FARADAY  PROMETHEUS MAN, THE  LP
NELSON, RAY FARADAY  THEN BEGGARS COULD RIDE  SR-P,C/P
NELSON, RAY FARADAY  TIMEQUEST  P,SFWA-P
NEO-OPSIS SCIENCE FICTION MAGAZINE  CD OF AMAZINGLY NEAT STUFF! (CD)  SR-MM
NEPHEW, WILLIAM/CHESTER, MICHAEL  MOON BASE  [C/RAN-P]

AUTHOR  TITLE  CODE
NEPHEW, WILLIAM/CHESTER, MICHAEL  MOON TRIP  [C/RAN-P]
NESBIT, E.  FIVE CHILDREN AND IT  C/P:2,P
NESBIT, E.  LAST OF THE DRAGONS AND SOME OTHERS, THE  P
NESBIT, E.  PHOENIX AND THE CARPET, THE  H
NESBIT, E.  STORY OF THE AMULET, THE  C/P,SR-P,H
NESFA  DIRECTORY OF FANDOM (1969)  SR-L
NESFA  INDEX TO PERRY RHODAN — US EDITION 1–25  R/IND-F
NESFA  INDEX TO PERRY RHODAN — US EDITION 26–50  R/IND-F
NESFA  INDEX TO SHORT SCIENCE FICTION 1987  R/IND
NESFA  INDEX TO SHORT SCIENCE FICTION FOR 1988  R/IND
NESFA  INDEX TO SHORT SCIENCE FICTION FOR 1989  R/IND
NESFA  INDEX TO THE SCIENCE FICTION MAGAZINES 1966  SRL/IND
NESFA  INDEX TO THE SCIENCE FICTION MAGAZINES 1966–1970  R/IND:2
NESFA  INDEX TO THE SCIENCE FICTION MAGAZINES 1967  SRL/IND
NESFA  INDEX TO THE SCIENCE FICTION MAGAZINES 1968  SRL/IND
NESFA  INDEX TO THE SCIENCE FICTION MAGAZINES 1969  SRL/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1971–1972

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1973

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1974

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1975

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1976

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1977–1978

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1979–1980

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1981

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1982

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1983

R/IND

NESFA INDEX TO THE SCIENCE FICTION MAGAZINES AND ORIGINAL

ANTHOLOGIES 1986

R/IND

NESFA  NESFA HYMNAL, THE, VOLUME 1  [C/RAN-H]
NESFA  NESFA HYMNAL, THE, VOLUME 1, 2ND EDITION  [C/RAN-H:2]
NESFA  NESFA HYMNAL, THE, VOLUME 2  [C/RAN-H:2]
NESVADBA, JOSEF  LOST FACE, THE  H
NETTER, SUSAN  STORM CHILD [TWILIGHT 17]  P
NEUFELD, JOHN  SLEEP TWO, THREE, FOUR!  C/P,P
NEUGROSCHEL, JOACHIM  GREAT TALES OF JEWISH FANTASY AND THE OCCULT  HA,LPA
NEVILLE, KATHERINE  EIGHT, THE  P
NEVILLE, KRIS  BETTYANN  P
NEVILLE, KRIS  INVADERS ON THE MOON  C/P:2,P
NEVILLE, KRIS  MISSION: MANSTOP  SR-P
NEVILLE, KRIS  MUTANTS, THE  C/P,P
NEVILLE, KRIS  PERIL OF THE STARMEN  SR-DB0759
NEVILLE, KRIS  SPECIAL DELIVERY  SR-DB0788
NEVILLE, KRIS  UNEARTH PEOPLE, THE  C/P:2,P
NEW ENGLISH LIBRARY  STRANGE WORLD OF SCIENCE FICTION  SR-VLHA
NEW VICTORIA PUBLISHERS  WOMAN SPACE  SR-LPA
NEWBOLT, SIR HENRY  ALADORE  LP
NEWCOMB, AMBROSE  WINGS OVER THE ROCKIES  SR-H
NEWCOMB, ROBERT  FIFTH SORCERESS, THE [BLOOD AND STONE 1]  H
NEWCOMB, ROBERT  GATES OF DAWN, THE [@BLOOD AND STONE #2]  C/P,P,H
NEWCOMB, ROBERT  GATES OF DAWN, THE (PROOF) [BLOOD AND STONE 2]  H,SR-LP
NEWCOMB, ROBERT  MARCH INTO DARKNESS, A [DESTINIES OF BLOOD AND STONE 2]  H
NEWCOMB, ROBERT  SAVAGE MESSIAH [DESTINIES OF BLOOD AND STONE 1]  H
NEWCOMB, ROBERT  SCROLLS OF THE ANCIENTS, THE [BLOOD AND STONE 3]  H,P
NEWCOMER, ALAN BARD  SPELL SINGERS  PA,C/PA
NEWELL, NEIL K.  RELUCTANT WIZARD, THE  P
NEWLOVE, DONALD  ETERNAL LIFE  LP
NEWMAN, KIM  ANDY WARHOL’S DRACULA  MB002
NEWMAN, KIM  ANNO DRACULA [ANNO DRACULA 1]  H,P
NEWMAN, KIM  BAD DREAMS  H
NEWMAN, KIM  BLOODY RED BARON, THE [ANNO DRACULA 2]  P
NEWMAN, KIM  NIGHT MAYOR, THE  P
NEWMAN, KIM  TIME AND RELATIVE [DOCTOR WHO]  SR-H
NEWMAN, PAUL S.  INVADERS: ALIEN MISSILE THREAT, THE  C/P,BLB
NEWMAN, PAUL S.  POPEYE IN GHOST SHIP TO TREASURE ISLAND  BLB
NEWMAN, RICHARD LOUIS  SIEGE OF ORBITOR  P
NEWMAN, ROBERT  BOY WHO COULD FLY, THE  P
NEWMAN, ROBERT  MERLIN’S MISTAKE  LP
NEWMAN, SHARAN  CHESSBOARD QUEEN, THE [GUINEVERE 2]  LP

The MITSFS Pinkdex  September 24, 2007  Page 281
AUTHOR  TITLE  CODE
NEWMAN, SHARAN  GUINEVERE [GUINEVERE 1]  P
NEWMAN, SHARAN  GUINEVERE EVERMORE [GUINEVERE 3]  H
NEWTON, JULIUS P.  FORGOTTEN RACE, THE  H

NEYLAND, JAMES OFFICIAL BATTLESTAR GALACTICA SCRAPBOOK, THE

[BATTLESTAR GALACTICA 1978]

R/REF-H

NIALL, IAN  BOY WHO SAW TOMORROW, THE  H
NICHOLLS, PETER  SCIENCE FICTION AT LARGE  R/REF-H,C/REF-H
NICHOLLS, PETER  SCIENCE FICTION ENCYCLOPEDIA, THE  R/REF-H,C/REF-H
NICHOLLS, STAN  BODYGUARD OF LIGHTNING [ORCS FIRST BLOOD 1]  LP
NICHOLLS, STAN  COVENANT RISING, THE [DREAMTIME 1]  H,LP
NICHOLLS, STAN  FIRST BLOOD [ORCS FIRST BLOOD 1,2,3]  H
NICHOLLS, STAN  LEGION OF THUNDER [ORCS FIRST BLOOD 2]  LP
NICHOLS, ADAM  PATHLESS WAY, THE  H
NICHOLS, BEVERLY  POWERS THAT BE  [C/RAN-P:2]
NICHOLS, LEIGH  EYES OF DARKNESS, THE  P
NICHOLS, LEIGH  SHADOWFIRES  P
NICHOLS, NICHELLE  BEYOND UHURA [STAR TREK]  R/REF-H

NICHOLS, NICHELLE/ BONANNO, MARGARET WANDER SATURN’S CHILD

H

NICHOLS, RUTH  SONG OF THE PEARL  P
NICHOLS, RUTH  WALK OUT OF THE WORLD, A  P
NICHOLSON, SAM  CAPTAIN EMPIRICAL  C/P,P
NICHOLSON, SAM  LIGHT BEARER, THE  P
NICHOLSON, WILLIAM  WIND SINGER, THE [WIND ON FIRE 1]  H
NIELSEN HAYDEN, PATRICK  NEW SKIES  PA
NIELSEN HAYDEN, PATRICK  STARLIGHT 1  LPA
NIELSEN HAYDEN, PATRICK  STARLIGHT 2  HA,LPA
NIELSEN HAYDEN, PATRICK  STARLIGHT 3  HA
NIELSEN HAYDEN, TERESA  MAKING BOOK  SR-LP
NIESEWAND, PETER  FALLBACK  H,P
NIFFENEGGER, AUDREY  TIME TRAVELER’S WIFE, THE  H,LP
NIGHBERT, DAVID E.  CLOUDS OF MAGELLAN [TIMELAPSE 2]  H,SFWA
NILES, DOUGLAS  BLACK WIZARDS [@FORGOTTEN REALMS MOONSHAE #2]  C/P,P
NILES, DOUGLAS  BREACH IN THE WATERSHED, A [WATERSHED 1]  H,LP
NILES, DOUGLAS  CIRCLE AT CENTER [SEVEN CIRCLES 1]  LP
NILES, DOUGLAS  CORAL KINGDOM, THE [@FORGOTTEN REALMS DRUIDHOME #2]  C/P,P
NILES, DOUGLAS  CROWN AND THE SWORD, THE [@DRAGONLANCE RISE OF SOLAMNIA #2]  P
NILES, DOUGLAS  DARKENHEIGHT [WATERSHED 2]  P
NILES, DOUGLAS  DARKWALKER ON MOONSHAE [@FORGOTTEN REALMS MOONSHAE #1]  C/P:2,P
NILES, DOUGLAS  DARKWELL [@FORGOTTEN REALMS MOONSHAE #3]  C/P:2,P
NILES, DOUGLAS  DRAGONS, THE [@DRAGONLANCE LOST HISTORIES #6]  P
NILES, DOUGLAS  DRUID QUEEN, THE [@FORGOTTEN REALMS DRUIDHOME #3]  C/P,P
NILES, DOUGLAS  EMPEROR OF ANSALON [@DRAGONLANCE VILLAINS #3]  P
NILES, DOUGLAS  FEATHERED DRAGON, THE [@FORGOTTEN REALMS MAZTICA #3]  P
NILES, DOUGLAS  FISTANDANTILUS REBORN [@DRAGONLANCE LOST LEGENDS #2]  P
NILES, DOUGLAS  GODDESS WORLDWEAVER, THE [@SEVEN CIRCLES #3]  LP
NILES, DOUGLAS  GOLDEN ORB, THE [@DRAGONLANCE ICEWALL #2]  P
NILES, DOUGLAS  IMMORTAL GAME [QUEST TRIAD 3]  P
NILES, DOUGLAS  IRONHELM [@FORGOTTEN REALMS MAZTICA #1]  P
NILES, DOUGLAS  KAGONESTI, THE [@DRAGONLANCE LOST HISTORIES #1]  P
NILES, DOUGLAS  KINSLAYER WARS, THE [@DRAGONLANCE ELVEN NATIONS #2]  P
NILES, DOUGLAS  LAST THANE, THE [@DRAGONLANCE CHAOS WAR 2]  P
NILES, DOUGLAS  LORD OF THE ROSE [@DRAGONLANCE RISE OF SOLAMNIA #1]  P

NILES, DOUGLAS MEASURE AND THE TRUTH, THE

[@DRAGONLANCE RISE OF SOLAMNIA #3]

P

NILES, DOUGLAS  MESSENGER, THE [@DRAGONLANCE ICEWALL #1]  P
NILES, DOUGLAS  PROPHET OF MOONSHAE [@FORGOTTEN REALMS DRUIDHOME #1]  C/P,P
NILES, DOUGLAS  PUPPET KING, THE [@DRAGONLANCE CHAOS WAR 4]  P
NILES, DOUGLAS  SUITORS DUEL [QUEST TRIAD 2]  P
NILES, DOUGLAS  TARZAN AND THE WELL OF SLAVES [TARZAN]  P
NILES, DOUGLAS  VIPERHAND [@FORGOTTEN REALMS MAZTICA #2]  C/P:2,P
NILES, DOUGLAS  WAR OF THREE WATERS [WATERSHED 3]  H,LP
NILES, DOUGLAS  WINTERHEIM [@DRAGONLANCE ICEWALL #3]  P
NILES, DOUGLAS  WIZARDS’ CONCLAVE [@DRAGONLANCE AGE OF MORTALS]  P
NILES, DOUGLAS  WORLD FALL [SEVEN CIRCLES 2]  LP
NILES, STEVE  GUNS, DRUGS, AND MONSTERS [CAL MCDONALD 2]  P
NILES, STEVE  WORDS WITHOUT PICTURES  LPA
NILSSON, INGEMAR  INDEX TILL SVENSKA GALAXY  R/IND
NIM, P. S.  DOUBLE MOBIUS SPHERE  C/P:2,P
NIMOY, LEONARD  I AM NOT SPOCK  R/REF-P,C/REF-P
NIMOY, LEONARD  I AM SPOCK  R/REF-H,C/REF-H:2
NIMOY, LEONARD  LEONARD NIMOY SCIENCE FICTION (DEMO CD)  SR-MM
NIMOY, LEONARD  WILL I THINK OF YOU?  [C/RAN-P]
NIMOY, LEONARD  YOU AND I  [C/RAN-P]

The MITSFS Pinkdex  September 24, 2007  Page 282
AUTHOR  TITLE  CODE
NIVEN, LARRY/BARNES, STEVEN  ACHILLES’ CHOICE  H:2,P,SFWA
NIVEN, LARRY  ALL THE MYRIAD WAYS  C/P:2,P
NIVEN, LARRY/BARNES, STEVEN  BARSOOM PROJECT, THE [DREAM PARK]  C/P,SFWA-P,P,H

NIVEN, LARRY/POURNELLE, JERRY/ BARNES, STEVEN BEOWULF’S CHILDREN [HEOROT 2]

H:2,P

NIVEN, LARRY  BEST OF ALL POSSIBLE WARS [@MAN-KZIN WARS]  PA
NIVEN, LARRY/COOPER, BRENDA  BUILDING HARLEQUIN’S MOON  H
NIVEN, LARRY/POURNELLE, JERRY  BURNING CITY, THE [MAGIC 4]  H:2,P
NIVEN, LARRY/POURNELLE, JERRY  BURNING TOWER [MAGIC 5]  H:2
NIVEN, LARRY/BARNES, STEVEN  CALIFORNIA VOODOO GAME, THE [DREAM PARK]  C/P,P,H
NIVEN, LARRY  CHOOSING NAMES [@MAN-KZIN WARS #8]  C/PA,PA
NIVEN, LARRY  CONVERGENT SERIES  C/P:2,P
NIVEN, LARRY  CRASHLANDER [KNOWN SPACE]  C/P:2,P
NIVEN, LARRY/BARNES, STEVEN  DESCENT OF ANANSI, THE  C/P:2,P
NIVEN, LARRY  DESTINY’S ROAD  C/P,P,H
NIVEN, LARRY  DOWN IN FLAMES (OUTLINE) [KNOWN SPACE]  [C/RAN-H:3]
NIVEN, LARRY  DRACO TAVERN, THE  H:2
NIVEN, LARRY/BARNES, STEVEN  DREAM PARK [DREAM PARK]  H:3,P

NIVEN, LARRY/POURNELLE, JERRY/ FLYNN, MICHAEL FALLEN ANGELS

C/P,P,H

NIVEN, LARRY  FLATLANDER [KNOWN SPACE]  P
NIVEN, LARRY  FLIGHT OF THE HORSE, THE  C/P:2,P
NIVEN, LARRY/POURNELLE, JERRY  FOOTFALL  C/P:2,P,H,SFWA
NIVEN, LARRY  GIFT FROM EARTH, A [KNOWN SPACE]  C/P:2,P
NIVEN, LARRY/POURNELLE, JERRY  GRIPPING HAND, THE [MOTE IN GOD’S EYE 2]  H:2,P
NIVEN, LARRY  HOLE IN SPACE, A  C/P:2,P
NIVEN, LARRY  INCONSTANT MOON (BRITISH)  SR-P,C/P
NIVEN, LARRY/POURNELLE, JERRY  INFERNO  C/P:2,P
NIVEN, LARRY  INTEGRAL TREES, THE [INTEGRAL TREES 1]  H:2,P,SFWA

NIVEN, LARRY/POURNELLE, JERRY/ BARNES, STEVEN LEGACY OF HEOROT, THE [HEOROT 1]

H:2,P

NIVEN, LARRY  LIMITS  H:2,P
NIVEN, LARRY  LONG ARM OF GIL HAMILTON, THE [KNOWN SPACE]  SR-P,C/P:2
NIVEN, LARRY/POURNELLE, JERRY  LUCIFER’S HAMMER  H:2,P
NIVEN, LARRY  MAGIC GOES AWAY, THE [MAGIC 1]  H:2,LP
NIVEN, LARRY  MAGIC MAY RETURN, THE [MAGIC 2]  C/PA,LPA
NIVEN, LARRY  MAN-KZIN WARS, THE [@MAN-KZIN WARS #1]  PA,C/PA,SFWA-P
NIVEN, LARRY  MAN-KZIN WARS 2 [@MAN-KZIN WARS #2]  PA,C/PA,SFWA-P:2
NIVEN, LARRY  MAN-KZIN WARS 3 [@MAN-KZIN WARS #3]  PA,C/PA,SFWA-P
NIVEN, LARRY  MAN-KZIN WARS 4 [@MAN-KZIN WARS #4]  PA,C/PA
NIVEN, LARRY  MAN-KZIN WARS 5 [@MAN-KZIN WARS #5]  PA,C/PA,SFWA-P
NIVEN, LARRY  MAN-KZIN WARS 6 [@MAN-KZIN WARS #6]  C/PA,PA
NIVEN, LARRY  MAN-KZIN WARS 7 [@MAN-KZIN WARS #7]  PA,C/PA
NIVEN, LARRY  MAN-KZIN WARS 9 [@MAN-KZIN WARS #9]  C/PA,PA
NIVEN, LARRY  MAN-KZIN WARS 11 [@MAN-KZIN WARS #11]  HA
NIVEN, LARRY  MORE MAGIC [MAGIC 3]  LP
NIVEN, LARRY/POURNELLE, JERRY  MOTE IN GOD’S EYE, THE [MOTE IN GOD’S EYE 1]  C/P,P,H:2,SFWA
NIVEN, LARRY  N-SPACE  H:2,P
NIVEN, LARRY  NEUTRON STAR [KNOWN SPACE]  C/P:2,SR-P
NIVEN, LARRY  NIVEN’S LAWS  SR-H
NIVEN, LARRY/POURNELLE, JERRY  OATH OF FEALTY  C/P,P,H:2
NIVEN, LARRY  PATCHWORK GIRL, THE [KNOWN SPACE]  C/P,LP,H
NIVEN, LARRY  PLAYGROUNDS OF THE MIND  H,P
NIVEN, LARRY  PROTECTOR [KNOWN SPACE]  C/P:2,P
NIVEN, LARRY  RAINBOW MARS  H:2,P
NIVEN, LARRY  RINGWORLD (ORIGINAL) [KNOWN SPACE RINGWORLD 1]  SR-P
NIVEN, LARRY  RINGWORLD (REVISED) [KNOWN SPACE RINGWORLD 1]  C/P,P
NIVEN, LARRY  RINGWORLD ENGINEERS, THE [KNOWN SPACE RINGWORLD 2]  H:2,P
NIVEN, LARRY  RINGWORLD THRONE, THE [KNOWN SPACE RINGWORLD 3]  C/P:2,P
NIVEN, LARRY  RINGWORLD’S CHILDREN [KNOWN SPACE RINGWORLD 4]  H:2
NIVEN, LARRY/BARNES, STEVEN  SATURN’S RACE  H:2,P
NIVEN, LARRY  SCATTERBRAIN  H
NIVEN, LARRY  SHAPE OF SPACE, THE  C/P:2,P
NIVEN, LARRY  SMOKE RING, THE [INTEGRAL TREES 2]  H:2,P,SFWA
NIVEN, LARRY  TALES OF KNOWN SPACE [KNOWN SPACE]  C/P:2,P
NIVEN, LARRY  THREE BOOKS OF KNOWN SPACE [KNOWN SPACE]  LP
NIVEN, LARRY  WORLD OF PTAVVS [KNOWN SPACE]  C/P:2,P
NIVEN, LARRY  WORLD OUT OF TIME, A  C/P,P,H
NIX, GARTH  ABHORSEN TRILOGY, THE [ABHORSEN 1,2,3]  H
NIX, GARTH  ACROSS THE WALL [ABHORSEN]  H
NIX, GARTH  DROWNED WEDNESDAY [@KEYS TO THE KINGDOM #3]  LP
NIX, GARTH  GRIM TUESDAY [@KEYS TO THE KINGDOM #2]  LP
NIX, GARTH  MISTER MONDAY [@KEYS TO THE KINGDOM #1]  LP
NIX, GARTH  RAGWITCH, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 283
AUTHOR  TITLE  CODE
NIX, GARTH  SHADE’S CHILDREN  P
NIX, GARTH  SIR THURSDAY [@KEYS TO THE KINGDOM #4]  LP
NOEL, ATANIELLE ANNYN  DUCHESS OF KNEEDEEP, THE  C/P:2,P
NOEL, ATANIELLE ANNYN  MURDER ON USHER’S PLANET  P
NOEL, ATANIELLE ANNYN  SPEAKER TO HEAVEN  H,SFWA-P
NOEL, RUTH S.  LANGUAGES OF MIDDLE EARTH, THE  R/REF-P,C/REF-P
NOEL, RUTH S.  LANGUAGES OF TOLKIEN’S MIDDLE-EARTH, THE  R/REF-P
NOEL, STERLING  WE WHO SURVIVED THE FIFTH ICE AGE  C/P:2,P

NOJD, RUBEN/TORNBERG, ASTRID/ ANGSTROM, MARGARITA MCKAY’S MODERN ENGLISH-SWEDISH SWEDISH-ENGLISH DICTIONARY

DICT

NOLAN, WILLIAM F.  ALIEN HORIZONS  C/P:2,P
NOLAN, WILLIAM F./GREENBERG, MARTIN H. BRADBURY CHRONICLES, THE  SFWA:2
NOLAN, WILLIAM F./SCHAFER, WILLIAM  CALIFORNIA SORCERY  C/PA,PA
NOLAN, WILLIAM F.  FUTURE IS NOW, THE  HA,PA
NOLAN, WILLIAM F.  HUMAN EQUATION, THE  HA
NOLAN, WILLIAM F.  IMPACT-20  C/P:2,P

NOLAN, WILLIAM F./ JOHNSON, GEORGE CLAYTON LOGAN [LOGAN 1,2,3]

H,SFWA,LP

NOLAN, WILLIAM F./ JOHNSON, GEORGE CLAYTON LOGAN’S RUN [LOGAN 1]

C/P:2,P

NOLAN, WILLIAM F.  LOGAN’S SEARCH [LOGAN 3]  C/P,P
NOLAN, WILLIAM F.  LOGAN’S WORLD [LOGAN 2]  C/P,P
NOLAN, WILLIAM F.  MAN AGAINST TOMORROW  PA,C/PA
NOLAN, WILLIAM F.  PSEUDO-PEOPLE, THE  HA,PA
NOLAN, WILLIAM F./GREENBERG, MARTIN H. SCIENCE FICTION ORIGINS  PA
NOLAN, WILLIAM F.  SEA OF SPACE, A  PA,C/PA
NOLAN, WILLIAM F.  SPACE FOR HIRE [SAM SPACE]  C/P:2,P
NOLAN, WILLIAM F.  THINGS BEYOND MIDNIGHT  H
NOLAN, WILLIAM F.  3 FOR SPACE [SAM SPACE]  LP
NOLAN, WILLIAM F.  3 TO THE HIGHEST POWER  PA,C/PA
NOLAN, WILLIAM F.  WILD GALAXY  H
NOLAN, WILLIAM F.  WILDERNESS OF STARS, A  PA,C/PA
NOLANE, RICHARD D.  TERRA SF  PA
NOLANE, RICHARD D.  TERRA SF 2  C/PA,PA
NOON, JEFF  AUTOMATED ALICE  P
NOON, JEFF  NYMPHOMATION  P
NOON, JEFF  PIXEL JUICE  P
NOON, JEFF  POLLEN  LP
NOON, JEFF  VURT  LP
NOONAN, ROSILAND  WHISPERS FROM THE PAST [@CHARMED]  P
NOONE, EDWINA  CRAGHOLD LEGACY, THE  P
NORDEN, ERIC  STARSONGS AND UNICORNS  C/P,P
NORDEN, ERIC  ULTIMATE SOLUTION, THE  P
NORMAN, ERIC  UNDER-PEOPLE, THE  [C/RAN-P]
NORMAN, JOHN  ASSASSIN OF GOR [GOR 5]  C/P:2,P
NORMAN, JOHN  BEASTS OF GOR [GOR 12]  C/P:2,P
NORMAN, JOHN  BLOOD BROTHERS OF GOR [GOR 18]  C/P,P
NORMAN, JOHN  CAPTAIN, THE [@TELNARIAN 2]  C/P:2,P
NORMAN, JOHN  CAPTIVE OF GOR [GOR 7]  C/P:2,P
NORMAN, JOHN  CHIEFTAIN, THE [@TELNARIAN 1]  C/P,P
NORMAN, JOHN  DANCER OF GOR [GOR 22]  P
NORMAN, JOHN  EXPLORERS OF GOR [GOR 13]  C/P:2,P
NORMAN, JOHN  FIGHTING SLAVE OF GOR [GOR 14]  C/P:2,P
NORMAN, JOHN  GHOST DANCE  P
NORMAN, JOHN  GUARDSMAN OF GOR [GOR 16]  P
NORMAN, JOHN  HUNTERS OF GOR [GOR 8]  C/P:2,P
NORMAN, JOHN  IMAGINATIVE SEX  [C/RAN-P:2]
NORMAN, JOHN  KAJIRA OF GOR [GOR 19]  P
NORMAN, JOHN  KING, THE [@TELNARIAN 3]  P
NORMAN, JOHN  MAGICIANS OF GOR [GOR 25]  C/P:2,P
NORMAN, JOHN  MARAUDERS OF GOR [GOR 9]  C/P:2,P
NORMAN, JOHN  MERCENARIES OF GOR [GOR 21]  C/P,P
NORMAN, JOHN  NOMADS OF GOR [GOR 4]  C/P:2,P
NORMAN, JOHN  OUTLAW OF GOR [GOR 2]  C/P:2,P
NORMAN, JOHN  PLAYERS OF GOR [GOR 20]  C/P:2,P
NORMAN, JOHN  PRIEST-KINGS OF GOR [GOR 3]  C/P:2,P
NORMAN, JOHN  RAIDERS OF GOR [GOR 6]  C/P:2,P
NORMAN, JOHN  RENEGADES OF GOR [GOR 23]  P
NORMAN, JOHN  ROGUE OF GOR [GOR 15]  C/P:2,P
NORMAN, JOHN  SAVAGES OF GOR [GOR 17]  C/P,P
NORMAN, JOHN  SLAVE GIRL OF GOR [GOR 11]  C/P,P
NORMAN, JOHN  TARNSMAN OF GOR [GOR 1]  C/P:2,P
NORMAN, JOHN  TIME SLAVE  C/P:2,P
NORMAN, JOHN  TRIBESMEN OF GOR [GOR 10]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 284
AUTHOR  TITLE  CODE
NORMAN, JOHN  VAGABONDS OF GOR [GOR 24]  P
NORMAN, JOHN  WITNESS OF GOR [GOR 26]  H
NORMAN, LISANNE  BETWEEN DARKNESS AND LIGHT [SHOLAN ALLIANCE 7]  P
NORMAN, LISANNE  BETWEEN DARKNESS AND LIGHT (PROOF) [SHOLAN ALLIANCE 7]  SR-LP
NORMAN, LISANNE  DARK NADIR [SHOLAN ALLIANCE 5]  P
NORMAN, LISANNE  FIRE MARGINS [SHOLAN ALLIANCE 3]  P
NORMAN, LISANNE  FORTUNE’S WHEEL [SHOLAN ALLIANCE 2]  P
NORMAN, LISANNE  RAZOR’S EDGE [SHOLAN ALLIANCE 4]  P
NORMAN, LISANNE  STRONGHOLD RISING [SHOLAN ALLIANCE 6]  P
NORMAN, LISANNE  TURNING POINT [SHOLAN ALLIANCE 1]  C/P,P
NORRIS, FRANK  NUTRO 29  H
NORTH, ANDREW  PLAGUE SHIP [SOLAR QUEEN 2]  DD345
NORTH, ANDREW  SARGASSO OF SPACE [SOLAR QUEEN 1]  DD249
NORTH, ANDREW  VOODOO PLANET  DD345
NORTH, DAVID  FORBIDDEN REGION [@TIME WARRIORS #2]  C/P,P
NORTH, DAVID  FUSE POINT [TIME WARRIORS #1]  C/P,P
NORTH, EDMUND H./COEN, FRANKLIN  METEOR  C/P:2,P
NORTH, ERIC  ANT MEN, THE  H,P
NORTH, STERLING/BOUTELL, C. B.  SPEAK OF THE DEVIL  HA
NORTON, ALDEN H.  AWARD SCIENCE FICTION READER  PA,C/PA
NORTON, ALDEN H.  FUTURES UNLIMITED  C/PA,PA
NORTON, ALDEN H./MOSKOWITZ, SAM  GREAT UNTOLD STORIES OF FANTASY AND HORROR  C/PA,PA
NORTON, ALDEN H.  HAUNTINGS AND HORRORS  PA
NORTON, ALDEN H.  HORROR TIMES TEN  PA
NORTON, ALDEN H.  MASTERS OF HORROR  PA
NORTON, ALDEN H./MOSKOWITZ, SAM  SPACE MAGICIANS, THE  C/PA,PA
NORTON, ANDRE  ANDROID AT ARMS  C/P:2,P
NORTON, ANDRE  AT SWORDS’ POINTS  H:2
NORTON, ANDRE/SMITH, SHERWOOD  ATLANTIS ENDGAME [TIME TRADERS]  P
NORTON, ANDRE  BEAST MASTER, THE [HOSTEEM STORM 1]  C/P,P,H,DD509
NORTON, ANDRE/MCCONCHIE, LYN  BEAST MASTER TEAM [HOSTEEM STORM 3,4]  H
NORTON, ANDRE/MCCONCHIE, LYN  BEAST MASTER’S ARK [HOSTEEM STORM 3]  P
NORTON, ANDRE/MCCONCHIE, LYN  BEAST MASTER’S CIRCUS [HOSTEEM STORM 4]  P
NORTON, ANDRE/MCCONCHIE, LYN  BEAST MASTER’S QUEST [HOSTEEM STORM 5]  H
NORTON, ANDRE  BOOK OF ANDRE NORTON, THE  C/P:2,P
NORTON, ANDRE  BREED TO COME  C/P,P,H
NORTON, ANDRE  BROTHER TO SHADOWS  C/P,P,H
NORTON, ANDRE/CUSHING, ENID  CAROLINE  P
NORTON, ANDRE/GREENBERG, MARTIN H.  CATFANTASTIC  C/PA,PA
NORTON, ANDRE/GREENBERG, MARTIN H.  CATFANTASTIC II  C/PA,PA
NORTON, ANDRE/GREENBERG, MARTIN H.  CATFANTASTIC III  C/PA,PA
NORTON, ANDRE/GREENBERG, MARTIN H.  CATFANTASTIC IV  C/PA,PA
NORTON, ANDRE/GREENBERG, MARTIN H.  CATFANTASTIC V  PA
NORTON, ANDRE  CATSEYE [DIPPLE 1]  C/P,P,H
NORTON, ANDRE/MCCONCHIE, LYN  CIARA’S SONG [WITCH WORLD]  P
NORTON, ANDRE  CROSSROADS OF TIME, THE [CROSSTIME]  C/P:2,DD164,P
NORTON, ANDRE/MILLER, SASHA  CROWN DISOWNED, A [OAK,YEW,ASH,AND ROWAN 3]  H
NORTON, ANDRE  CRYSTAL GRYPHON, THE [WITCH WORLD HIGH HALLACK GRYPHON]  C/P:2,SFWA-P,P
NORTON, ANDRE  DARE TO GO A-HUNTING [MOON SINGER 4]  H:2,SFWA-P,P
NORTON, ANDRE  DARK COMPANION  H
NORTON, ANDRE  DARK PIPER  C/P:2,P,H
NORTON, ANDRE  DARKNESS AND DAWN  P
NORTON, ANDRE/GILBERT, MICHAEL  DAY OF THE NESS, THE  H,P
NORTON, ANDRE  DAYBREAK — 2250 A.D.  C/P:2,SR-DD069,P
NORTON, ANDRE  DEFIANT AGENTS, THE [TIME TRADERS 3]  C/P,P,H

NORTON, ANDRE DEFIANT AGENTS, THE (REVISED) [TIME TRADERS 3,4]/ KEY OUT OF TIME (REVISED) [TIME TRADERS 3,4]

P

NORTON, ANDRE/SMITH, SHERWOOD  DERELICT FOR TRADE [SOLAR QUEEN 5]  C/P,P
NORTON, ANDRE/MILLER, SASHA  DRAGON BLADE [OAK,YEW,ASH,AND ROWAN 4]  H
NORTON, ANDRE  DRAGON MAGIC  C/P,P,H
NORTON, ANDRE  DREAD COMPANION  H:2,P
NORTON, ANDRE/MCCONCHIE, LYN  DUKE’S BALLAD, THE  H
NORTON, ANDRE/SMITH, SHERWOOD  ECHOES IN TIME [TIME TRADERS 6]  H
NORTON, ANDRE/LACKEY, MERCEDES  ELVENBANE, THE [HALFBLOOD 1]  H:2,P
NORTON, ANDRE/LACKEY, MERCEDES  ELVENBLOOD [HALFBLOOD 2]  H:2
NORTON, ANDRE/LACKEY, MERCEDES  ELVENBORN [HALFBLOOD 3]  H:2
NORTON, ANDRE/SHWARTZ, SUSAN  EMPIRE OF THE EAGLE  P
NORTON, ANDRE  EXILES OF THE STARS [MOON SINGER 2]  C/P,P,H
NORTON, ANDRE  EYE OF THE MONSTER  DF147,C/P
NORTON, ANDRE/GRIFFIN, P. M.  FIREHAND [TIME TRADERS 5]  H:2
NORTON, ANDRE  FLIGHT IN YIKTOR [MOON SINGER 3]  H:2,P

NORTON, ANDRE/GRIFFIN, P. M./ SCHAUB, MARY H.

FLIGHT OF VENGEANCE [WITCH WORLD TURNING 2]

P

NORTON, ANDRE  FOLLOW THE DRUM  P

The MITSFS Pinkdex  September 24, 2007  Page 285
AUTHOR  TITLE  CODE
NORTON, ANDRE  FORERUNNER [FORERUNNER 4/DIPPLE 4]  C/P:2,SFWA-P,P
NORTON, ANDRE  FORERUNNER FORAY [FORERUNNER 3]  H:2,P
NORTON, ANDRE  FORERUNNER: THE SECOND VENTURE [FORERUNNER 5]  H:2,SFWA,SFWA-P,P
NORTON, ANDRE  FOUR FROM THE WITCH WORLD  C/PA,PA,SFWA-P
NORTON, ANDRE  FROM THE SEA TO THE STARS  LP
NORTON, ANDRE  FUR MAGIC  C/P:2,P
NORTON, ANDRE  GALACTIC DERELICT [TIME TRADERS 2]  C/P:2,P
NORTON, ANDRE  GARAN THE ETERNAL  C/P,P,H
NORTON, ANDRE  GATE OF THE CAT, THE [WITCH WORLD ESTCARP]  C/P,SFWA,P,H
NORTON, ANDRE/DONALDY, ERNESTINE  GATES TO TOMORROW  HA
NORTON, ANDRE  GODS AND ANDROIDS  P
NORTON, ANDRE  GOLDEN TRILLIUM [TRILLIUM 3]  H:2,P
NORTON, ANDRE  GRAND MASTERS’ CHOICE  PA
NORTON, ANDRE  GRAY MAGIC  LP
NORTON, ANDRE  GRYPHON IN GLORY [WITCH WORLD HIGH HALLACK GRYPHON]  H:2,P
NORTON, ANDRE/CRISPIN, A. C.  GRYPHON’S EYRIE [WITCH WORLD HIGH HALLACK GRYPHON]  H:2,SFWA-P:2,P
NORTON, ANDRE  HANDS OF LYR, THE  C/P,P,H
NORTON, ANDRE  HERE ABIDE MONSTERS  C/P,P,H
NORTON, ANDRE  HIGH SORCERY  C/P:2,P
NORTON, ANDRE  HORN CROWN [WITCH WORLD HIGH HALLACK]  H:2,P
NORTON, ANDRE/MILLER, PHYLLIS  HOUSE OF SHADOWS  C/P:2,SFWA-P,P
NORTON, ANDRE  HUON OF THE HORN  C/P:2,P
NORTON, ANDRE  ICE CROWN  H:2,P
NORTON, ANDRE/SHWARTZ, SUSAN  IMPERIAL LADY  C/P,SFWA,SFWA-P,P
NORTON, ANDRE  IRON BUTTERFLIES  P
NORTON, ANDRE  IRON CAGE  H:2,P
NORTON, ANDRE  JARGOON PARD, THE [WITCH WORLD HIGH HALLACK]  C/P,P,H
NORTON, ANDRE  JUDGMENT ON JANUS [JANUS 1/DIPPLE 2]  H:2,P
NORTON, ANDRE/MCCONCHIE, LYN  KEY OF THE KEPLAN, THE [WITCH WORLD SECRETS 1]  P
NORTON, ANDRE  KEY OUT OF TIME [TIME TRADERS 4]  C/P:2,P
NORTON, ANDRE  KNAVE OF DREAMS  C/P,P,H
NORTON, ANDRE/MILLER, SASHA  KNIGHT OR KNAVE [OAK,YEW,ASH,AND ROWAN 2]  H
NORTON, ANDRE  LAST PLANET, THE [STAR SOLDIERS 2]  DD096,C/P:2,P
NORTON, ANDRE  LAVENDER-GREEN MAGIC  C/P,P,H
NORTON, ANDRE/EDGHILL, ROSEMARY  LEOPARD IN EXILE [CAROLUS REX 2]  H,P
NORTON, ANDRE  LORD OF THUNDER [HOSTEEM STORM 2]  C/P:2,P
NORTON, ANDRE  LORE OF THE WITCH WORLD [WITCH WORLD]  C/P:2,P
NORTON, ANDRE  LOST LANDS OF THE WITCH WORLD [WITCH WORLD ESTCARP 1,2,3]  H
NORTON, ANDRE/SCHAUB, MARY H.  MAGESTONE, THE [WITCH WORLD SECRETS]  C/P,P
NORTON, ANDRE  MAGIC BOOKS, THE  P
NORTON, ANDRE/ADAMS, ROBERT  MAGIC IN ITHKAR 1  C/PA,LPA,SFWA
NORTON, ANDRE/ADAMS, ROBERT  MAGIC IN ITHKAR 2  SFWA,SFWA-P,LP
NORTON, ANDRE/ADAMS, ROBERT  MAGIC IN ITHKAR 3  PA,SFWA-P
NORTON, ANDRE/ADAMS, ROBERT  MAGIC IN ITHKAR 4  C/PA,PA,SFWA-P
NORTON, ANDRE  MANY WORLDS OF ANDRE NORTON, THE  H
NORTON, ANDRE  MARK OF THE CAT, THE  C/P,SFWA,SFWA-P,P
NORTON, ANDRE  MASKS OF THE OUTCASTS [DIPPLE 1,3]  H
NORTON, ANDRE  MERLIN’S MIRROR  C/P:2,P
NORTON, ANDRE/SMITH, SHERWOOD  MIND FOR TRADE, A [SOLAR QUEEN 6]  P
NORTON, ANDRE  MIRROR OF DESTINY  P
NORTON, ANDRE  MOON CALLED  C/P,LP,P
NORTON, ANDRE  MOON MIRROR  H:2,SFWA-P,P
NORTON, ANDRE  MOON OF THREE RINGS [MOON SINGER 1]  C/P:2,P
NORTON, ANDRE  MOONSINGER [MOON SINGER 1,2]  H
NORTON, ANDRE  NIGHT OF MASKS [DIPPLE 3]  C/P,P,H
NORTON, ANDRE  NO NIGHT WITHOUT STARS  H:2,P
NORTON, ANDRE  OCTAGON MAGIC  H:2

NORTON, ANDRE/MATHEWS, PATRICIA/ MILLER, SASHA ON WINGS OF MAGIC [WITCH WORLD TURNING 3]

H:2

NORTON, ANDRE  OPAL-EYED FAN, THE  H:2,P
NORTON, ANDRE  OPERATION TIME SEARCH  C/P:2,P
NORTON, ANDRE  ORDEAL IN OTHERWHERE [FORERUNNER 2]  C/P:2,P
NORTON, ANDRE  OUTSIDE  C/P,P,H
NORTON, ANDRE  PERILOUS DREAMS  C/P:2,P
NORTON, ANDRE  PLAGUE SHIP [SOLAR QUEEN 2]  C/P,P,H
NORTON, ANDRE  PLANETA DE LAS BRUJAS, EL  FP
NORTON, ANDRE  POSTMARKED THE STARS [SOLAR QUEEN 3]  C/P,P,H
NORTON, ANDRE  PRINCE COMMANDS, THE  P
NORTON, ANDRE  QUAG KEEP [QUAG KEEP 1]  H:2,P
NORTON, ANDRE  QUEST CROSSTIME [CROSSTIME]  H:2,P
NORTON, ANDRE  RALESTONE LUCK  C/P,SFWA-P,P
NORTON, ANDRE  RED HART MAGIC  C/P,P,H
NORTON, ANDRE/GRIFFIN, P. M.  REDLINE THE STARS  H
NORTON, ANDRE/RABE, JEAN  RENAISSANCE FAIRE  PA

The MITSFS Pinkdex  September 24, 2007  Page 286
AUTHOR  TITLE  CODE
NORTON, ANDRE/RABE, JEAN  RETURN TO QUAG KEEP [QUAG KEEP 2]  H:2
NORTON, ANDRE/MILLER, PHYLLIS  RIDE THE GREEN DRAGON  H
NORTON, ANDRE  SARGASSO OF SPACE [SOLAR QUEEN 1]  C/P:2,P
NORTON, ANDRE  SCARFACE  P
NORTON, ANDRE  SCENT OF MAGIC  C/P,P,H
NORTON, ANDRE  SEA SIEGE  DF147,P,C/P:2
NORTON, ANDRE  SECRET OF THE LOST RACE  C/P:2,DD381,P
NORTON, ANDRE/MILLER, PHYLLIS  SEVEN SPELLS TO SUNDAY  H:2,P
NORTON, ANDRE  SHADOW HAWK  C/P,P,H
NORTON, ANDRE/EDGHILL, ROSEMARY  SHADOW OF ALBION, THE [CAROLUS REX 1]  H:2,P
NORTON, ANDRE/MCCONCHIE, LYN  SILVER MAY TARNISH [WITCH WORLD]  H
NORTON, ANDRE  SIOUX SPACEMAN, THE  C/P:2,P,DD437
NORTON, ANDRE  SMALL SHADOWS CREEP  HA
NORTON, ANDRE/HOGARTH, GRACE ALLEN  SNEEZE ON SUNDAY  H
NORTON, ANDRE  SNOW SHADOW  P
NORTON, ANDRE/CRISPIN, A. C.  SONGSMITH [WITCH WORLD HIGH HALLACK]  H:2,P
NORTON, ANDRE  SORCERESS OF THE WITCH WORLD [WITCH WORLD ESTCARP 3]  C/P:2,P
NORTON, ANDRE  SPACE PIONEERS  HA
NORTON, ANDRE  SPACE SERVICE  HA
NORTON, ANDRE  SPELL OF THE WITCH WORLD [WITCH WORLD HIGH HALLACK]  C/P:2,P
NORTON, ANDRE  STAND & DELIVER  P
NORTON, ANDRE  STAND TO HORSE  P
NORTON, ANDRE  STAR BORN [PAX]  C/P:2,P,DD299
NORTON, ANDRE  STAR GATE  C/P:2,P
NORTON, ANDRE  STAR GUARD [STAR SOLDIERS 1]  C/P,H:2,P,DD199
NORTON, ANDRE  STAR HUNTER  DD509
NORTON, ANDRE  STAR HUNTER AND VOODOO PLANET  C/P:2,P
NORTON, ANDRE/MADLEE, DOROTHY  STAR KA’AT  C/P,P,H
NORTON, ANDRE/MADLEE, DOROTHY  STAR KA’AT WORLD  C/P,P,H
NORTON, ANDRE/MADLEE, DOROTHY  STAR KA’ATS AND THE PLANT PEOPLE  C/P,P,H
NORTON, ANDRE  STAR MAN’S SON  C/P:2,P
NORTON, ANDRE  STAR RANGERS [STAR SOLDIERS 2]  C/P,P,H
NORTON, ANDRE  STAR SOLDIERS [STAR SOLDIERS 1,2]  P
NORTON, ANDRE  STARS ARE OURS!, THE [PAX]  C/P,H,DD121,P
NORTON, ANDRE  STEEL MAGIC  P
NORTON, ANDRE  STORM OVER WARLOCK [FORERUNNER 1]  C/P,P,H
NORTON, ANDRE/GRIFFIN, P. M.  STORMS OF VICTORY [WITCH WORLD TURNING 1]  C/P,SFWA,P
NORTON, ANDRE  SWORD IS DRAWN, THE  H
NORTON, ANDRE  TALES OF THE WITCH WORLD [WITCH WORLD]  SR-HA,HA
NORTON, ANDRE  TALES OF THE WITCH WORLD 2 [WITCH WORLD]  PA,SFWA-P
NORTON, ANDRE  TALES OF THE WITCH WORLD 3 [WITCH WORLD]  C/PA,PA
NORTON, ANDRE/RABE, JEAN  TASTE OF MAGIC, A  H
NORTON, ANDRE  TEN MILE TREASURE  P
NORTON, ANDRE  THREE AGAINST THE WITCH WORLD [WITCH WORLD ESTCARP 1]  C/P:2,P
NORTON, ANDRE  THREE HANDS FOR SCORPIO  H
NORTON, ANDRE  TIME TRADERS, THE [TIME TRADERS 1]  C/P:2,P

NORTON, ANDRE TIME TRADERS (REVISED) [TIME TRADERS 1,2]/

GALACTIC DERELICT (REVISED) [TIME TRADERS 1,2]

P

NORTON, ANDRE/SMITH, SHERWOOD  TIME TRADERS III [TIME TRADERS 3,4]  H
NORTON, ANDRE/MILLER, SASHA  TO THE KING A DAUGHTER [OAK,YEW,ASH,AND ROWAN 1]  H:2
NORTON, ANDRE  TREY OF SWORDS [WITCH WORLD ESTCARP]  C/P,P,H
NORTON, ANDRE  UNCHARTED STARS [JERN MURDOCK 2]  H:2,P
NORTON, ANDRE  VELVET SHADOWS  P
NORTON, ANDRE  VICTORY ON JANUS [JANUS 2]  C/P,P,H
NORTON, ANDRE  VOORLOPER  H:2,P
NORTON, ANDRE  WARDING OF WITCH WORLD, THE [WITCH WORLD SECRETS]  H:2,P
NORTON, ANDRE  ’WARE HAWK [WITCH WORLD ESTCARP]  H:2,P
NORTON, ANDRE  WARLOCK OF THE WITCH WORLD [WITCH WORLD ESTCARP 2]  C/P:2,P
NORTON, ANDRE  WEB OF THE WITCH WORLD [WITCH WORLD ESTCARP]  C/P:2,P
NORTON, ANDRE  WHEEL OF STARS  C/P,SFWA-P,LP,H
NORTON, ANDRE  WHITE JADE FOX, THE  H:2,P
NORTON, ANDRE  WIND IN THE STONE  C/P,P,H
NORTON, ANDRE  WITCH WORLD [WITCH WORLD ESTCARP]  C/P:2,P
NORTON, ANDRE  WIZARDS’ WORLDS  C/P,P,H
NORTON, ANDRE  WRAITHS OF TIME  H:2,P
NORTON, ANDRE  X FACTOR, THE  C/P,P,H
NORTON, ANDRE  YEAR OF THE UNICORN [WITCH WORLD HIGH HALLACK]  C/P:2,P
NORTON, ANDRE  YURTH BURDEN  C/P:2,P
NORTON, ANDRE  ZARSTHOR’S BANE [WITCH WORLD HIGH HALLACK]  C/P:2,P
NORTON, ANDRE  ZERO STONE, THE [JERN MURDOCK 1]  H:2,P
NORTON, MARY  ARE ALL THE GIANTS DEAD?  LP
NORVIL, MANNING  CROWN OF THE SWORD GOD  P
NORVIL, MANNING  DREAM CHARIOTS  C/P,P
NORVIL, MANNING  WHETTED BRONZE  P

The MITSFS Pinkdex  September 24, 2007  Page 287
AUTHOR  TITLE  CODE
NORWOOD, VICTOR  NIGHT OF THE BLACK HORROR  SR-P
NORWOOD, WARREN  FINAL COMMAND [DOUBLE SPIRAL WAR 3]  P
NORWOOD, WARREN  FIZE OF THE GABRIEL RATCHETS [@WINDHOVER TAPES #3]  C/P:2,P
NORWOOD, WARREN  FLEXING THE WARP [@WINDHOVER TAPES #2]  C/P:2,P
NORWOOD, WARREN  IMAGE OF VOICES, AN [@WINDHOVER TAPES #1]  C/P,P
NORWOOD, WARREN  MIDWAY BETWEEN [DOUBLE SPIRAL WAR 1]  C/P:2,P
NORWOOD, WARREN  PLANET OF FLOWERS [@WINDHOVER TAPES #4]  C/P:2,P
NORWOOD, WARREN  POLAR FLEET [DOUBLE SPIRAL WAR 2]  C/P,P
NORWOOD, WARREN/MYLIUS, RALPH  SEREN CENACLES, THE  P
NORWOOD, WARREN  SHUDDERCHILD  P
NORWOOD, WARREN/ODOM, MEL  STRANDED [@TIME POLICE #3]  C/P,P
NORWOOD, WARREN  TRAPPED! [@TIME POLICE #2]  P
NORWOOD, WARREN  TRUE JAGUAR  C/P,P
NORWOOD, WARREN  VANISHED [@TIME POLICE #1]  C/P,P
NOSTRADAMUS  ORACLES OF NOSTRADAMUS  [C/RAN-H]
NOTKIN, DEBBIE/STEWART, ROGER  OUTER LIMITS VOLUME ONE, THE  LP
NOTKIN, DEBBIE  OUTER LIMITS VOLUME TWO, THE  LP
NOURSE, ALAN E.  BLADERUNNER, THE  C/P,P,H
NOURSE, ALAN E.  COUNTERFEIT MAN, THE  H:2,P
NOURSE, ALAN E./MEYER, J. A.  INVADERS ARE COMING, THE  C/P:2,P
NOURSE, ALAN E.  MAN OBSESSED, A  DD096
NOURSE, ALAN E.  MERCY MEN, THE  C/P,P,H
NOURSE, ALAN E.  NINE PLANETS  [C/RAN-P]
NOURSE, ALAN E.  PSI HIGH AND OTHERS  C/P:2,P
NOURSE, ALAN E.  RAIDERS FROM THE RINGS  C/P:2,P
NOURSE, ALAN E.  ROCKET TO LIMBO  DD385
NOURSE, ALAN E.  SCAVENGERS IN SPACE  C/P:2,P
NOURSE, ALAN E.  SO YOU WANT TO BE A DOCTOR  [C/RAN-P]
NOURSE, ALAN E.  STAR SURGEON  C/P,P,H
NOURSE, ALAN E.  TIGER BY THE TAIL AND OTHER SCIENCE FICTION STORIES  H:2,P
NOURSE, ALAN E.  TROUBLE ON TITAN  C/P,P,H
NOURSE, ALAN E.  UNIVERSE BETWEEN, THE  C/P:2,P
NOVA, CRAIG  WETWARE  LP
NOVAK, KATE/GRUBB, JEFF  AZURE BONDS [FORGOTTEN REALMS FINDER’S STONE 1]  C/P:2,P
NOVAK, KATE/GRUBB, JEFF  FINDER’S BANE [@FORGOTTEN REALMS HARPERS]  P
NOVAK, KATE/GRUBB, JEFF  MASQUERADES [FORGOTTEN REALMS HARPERS 10]  P
NOVAK, KATE/GRUBB, JEFF  SONG OF THE SAURIALS [FORGOTTEN REALMS FINDER’S STONE 3]  P
NOVAK, KATE/GRUBB, JEFF  TYMORA’S LUCK [@FORGOTTEN REALMS LOST GODS]  P
NOVAK, KATE/GRUBB, JEFF  WYVERN’S SPUR, THE [FORGOTTEN REALMS FINDER’S STONE 2]  C/P,P
NOVEMBER, SHARON  FIREBIRDS  HA
NOVEMBER, SHARON  FIREBIRDS RISING  HA
NOVIK  BLACK POWDER WAR [TEMERAIRE 3]  C/P:2,P
NOVIK, NAOMI  HIS MAJESTY’S DRAGON [TEMERAIRE 1]  P
NOVIK, NAOMI  THRONE OF JADE [TEMERAIRE 2]  C/P,P
NOWAK, JURGEN  ZWISCHEN SCHOPFUNG UND UNTERGANG  FP
NOWLAN, PHILIP FRANCIS  ARMAGEDDON 2419 A.D. [BUCK ROGERS]  C/P,P,H

NOWLAN, PHILIP FRANCIS ARMAGEDDON 2419 A.D. (REVISED) [BUCK ROGERS/

BUCK ROGERS (NIVEN & POURNELLE’S) 1]

C/P,P

NOWLAN, PHILIP FRANCIS  BUCK ROGERS AND THE DOOM COMET  BLB
NOWLAN, PHILIP FRANCIS  BUCK ROGERS AND THE PLANETOID PLOT  BLB
NOYES, ALFRED  NO OTHER MAN  H
NUETZEL, CHARLES  IF THIS GOES ON  SR-PA,C/PA
NUETZEL, CHARLES  IMAGES OF TOMORROW  SR-P
NUETZEL, CHARLES  JUNGLE JUNGLE  SR-P
NUETZEL, CHARLES  LAST CALL FOR THE STARS  H
NUETZEL, CHARLES  QUEEN OF BLOOD  SR-L
NUETZEL, CHARLES  RAIDERS OF NOOMAS  SR-P
NUETZEL, CHARLES  SWORDSMEN OF VISTAR  SR-P
NUNES, CLAUDE  INHERIT THE EARTH  DG580
NUNES, CLAUDE/NUNES, RHODA  RECOIL  D71082
NYE, JODY LYNN  DON’T FORGET YOUR SPACESUIT, DEAR  C/PA,PA
NYE, JODY LYNN  DRAGONFIRE  C/P,SFWA-P,P
NYE, JODY LYNN  DRAGONHARPER  C/P:2,SFWA-P,P
NYE, JODY LYNN/MCCAFFREY, ANNE  DRAGONLOVER’S GUIDE TO PERN, THE [PERN]  R/REF-H
NYE, JODY LYNN  ENCYCLOPEDIA OF XANTH  P,SFWA-P
NYE, JODY LYNN  GHOST OF A CHANCE  C/P:2,SFWA-P,P
NYE, JODY LYNN  GRAND TOUR, THE  P
NYE, JODY LYNN  HIGHER MYTHOLOGY [MYTHOLOGY 3]  P
NYE, JODY LYNN  LADY AND THE TIGER, THE [TAYLOR’S ARK 3]  C/P,P
NYE, JODY LYNN  MAGIC TOUCH, THE  P
NYE, JODY LYNN  MEDICINE SHOW [TAYLOR’S ARK 2]  P
NYE, JODY LYNN  MYTHOLOGY 101 [MYTHOLOGY 1]  P
NYE, JODY LYNN  MYTHOLOGY ABROAD [MYTHOLOGY 2]  C/P,P
NYE, JODY LYNN  SCHOOL OF LIGHT  P

The MITSFS Pinkdex  September 24, 2007  Page 288
AUTHOR  TITLE  CODE
NYE, JODY LYNN  SHIP ERRANT, THE [SHELLPEOPLE 6]  C/P,P
NYE, JODY LYNN  TAYLOR’S ARK [TAYLOR’S ARK 1]  P
NYE, JODY LYNN  UNEXPECTED APPRENTICE, AN  H
NYE, JODY LYNN  WAKING IN DREAMLAND  P
NYE, ROBERT  MERLIN  C/P,P

NYLUND, ERIC S./LEE, MICHAEL B./ BERMAN, NANCY/TRAUTMANN, ERIC S.

CRIMSON SKIES

PA

NYLUND, ERIC S.  DRY WATER  LP
NYLUND, ERIC S.  GAME OF UNIVERSE, A  P
NYLUND, ERIC S.  GHOSTS OF ONYX [@HALO]  P
NYLUND, ERIC S.  HALO: FIRST STRIKE [@HALO]  P
NYLUND, ERIC S.  HALO: THE FALL OF REACH [@HALO]  P
NYLUND, ERIC S.  PAWN’S DREAM  P
NYLUND, ERIC S.  SIGNAL SHATTERED, A [SIGNAL TO NOISE 2]  H,P
NYLUND, ERIC S.  SIGNAL TO NOISE [SIGNAL TO NOISE 1]  C/P,P
NYSTUL, BRYAN  TEST OF VENGEANCE [@BATTLETECH 51]  P

The MITSFS Pinkdex  September 24, 2007  Page 289
AUTHOR  TITLE  CODE
OATES, JOYCE CAROL  NIGHT-SIDE  P
OBERNDORF, CHARLES  FORAGERS  C/P:2,P
OBERNDORF, CHARLES  SHELTERED LIVES  C/P,P
OBERNDORF, CHARLES  TESTING  P
OBERNDORF, CHARLES  TESTING (PROOF)  SR-LP
OBOLER, ARCH  HOUSE ON FIRE  C/P,P
O’BRIEN, FITZ-JAMES  DREAM STORIES AND FANTASIES  SFWA
O’BRIEN, FITZ-JAMES  MACABRE TALES  SFWA
O’BRIEN, KEVIN L.  ELDRITCH BLUE  LPA
O’BRIEN, ROBERT C.  MRS. FRISBY AND THE RATS OF NIMH [NIMH 1]  P
O’BRIEN, ROBERT C.  REPORT FROM GROUP 17, A  H,P
O’BRIEN, ROBERT C.  SILVER CROWN, THE  C/P,LP
O’BRIEN, ROBERT C.  Z FOR ZACHARIAH  C/P,P
OBRUCHEV, V. A.  PLUTONIA: AN ADVENTURE THROUGH PREHISTORY  H
OBRUCHEV, V. A.  PUTESHESTVIYA V PROSHLOE I BUDUSHCHEE  FH
OBRUCHEV, VLADIMIR  SANNIKOV LAND  SR-H
OBSTFELD, RAYMOND  REINCARNATION OF REECE ERIKSON, THE  P,SFWA-P
OBUKHOVA, LYDIA  DAUGHTER OF NIGHT  H,P
O’CALLAGHAN, MAXINE  BOGEYMAN, THE  SFWA-P
O’CALLAGHAN, MAXINE  DARK VISIONS  SFWA-P
O’CALLAGHAN, MAXINE  SOMETHING’S CALLING ME HOME  P
O’CONNELL, CHARLES C.  VANISHING ISLAND, THE  H
OCTOPUS BOOKS  OVER THE RAINBOW  HA
O’DAY-FLANNERY, CONSTANCE  BEWITCHED  P
O’DAY-FLANNERY, CONSTANCE  TIMELESS PASSION  P
O’DAY-FLANNERY, CONSTANCE  TIMESWEPT LOVERS  P
ODOM, MEL  BLACK ROAD, THE [@DIABLO #2]  C/P,P
ODOM, MEL  BLADE  P
ODOM, MEL  BRUJA [@ANGEL]  P
ODOM, MEL  CROSSINGS [@BUFFY]  P
ODOM, MEL  CURSED [@BUFFY/ANGEL]  C/P,P
ODOM, MEL  DESTRUCTION OF THE BOOKS, THE [ROVER 2]  H
ODOM, MEL  EXODUS [@HELLGATE LONDON #1]  P
ODOM, MEL  F.R.E.E.FALL [F.R.E.E.LANCERS 2]  P
ODOM, MEL  F.R.E.E.LANCERS [F.R.E.E.LANCERS 1]  P
ODOM, MEL  HARVEST MOON [@SABRINA THE TEENAGE WITCH #15]  P
ODOM, MEL  HEADHUNTERS [@SHADOWRUN 27]  P
ODOM, MEL  HUNTERS OF THE DARK SEA  H
ODOM, MEL  I’LL ZAP MANHATTAN [@SABRINA THE TEENAGE WITCH #18]  P
ODOM, MEL  IMAGE [@ANGEL]  P
ODOM, MEL  JEWEL OF TURMISH, THE [@FORGOTTEN REALMS CITIES]  P
ODOM, MEL  LETHAL INTERFACE  P
ODOM, MEL  LORD OF THE LIBRARIES [ROVER 3]  P

ODOM, MEL LOST LIBRARY OF CORMANTHYR, THE [@FORGOTTEN REALMS LOST EMPIRES]

P

ODOM, MEL  PREYING FOR KEEPS [@SHADOWRUN 21]  P
ODOM, MEL  QUEST FOR THE TRILOGY, THE [ROVER 4]  H
ODOM, MEL  REDEMPTION [@ANGEL]  P
ODOM, MEL  REVENANT [BUFFY]  C/P,P
ODOM, MEL  RISING TIDE [@FORGOTTEN REALMS THREAT FROM THE SEA #1]  P
ODOM, MEL  ROVER, THE [ROVER 1]  P
ODOM, MEL  RUN HARD, DIE FAST [@SHADOWRUN 35]  P
ODOM, MEL  SABRINA GOES TO ROME [@SABRINA THE TEENAGE WITCH]  P

ODOM, MEL SEA DEVIL’S EYE, THE

[@FORGOTTEN REALMS THREAT FROM THE SEA #3]

P

ODOM, MEL  SEA OF MIST, THE [MIGHT AND MAGIC]  P
ODOM, MEL  STALKER ANALOG  P

ODOM, MEL UNDER FALLEN STARS

[@FORGOTTEN REALMS THREAT FROM THE SEA #2]

P

ODOM, MEL  UNNATURAL SELECTION [BUFFY]  P
O’DONNELL, K. M.  DWELLERS OF THE DEEP  D27400
O’DONNELL, K. M.  EMPTY PEOPLE, THE  C/P:2,P
O’DONNELL, K. M.  FINAL WAR AND OTHER FANTASIES  D23775
O’DONNELL, K. M.  GATHER IN THE HALL OF THE PLANETS  D27415
O’DONNELL, K. M.  GUERRA FINALE  FP
O’DONNELL, K. M.  IN THE POCKET AND OTHER S-F STORIES  D27415
O’DONNELL, K. M.  UNIVERSE DAY  C/P:2,P
O’DONNELL, KEVIN JR.  BANDER SNATCH  P
O’DONNELL, KEVIN JR.  CAVERNS [@JOURNEYS OF MCGILL FEIGHAN #1]  C/P:2,P
O’DONNELL, KEVIN JR.  CLIFFS [@JOURNEYS OF MCGILL FEIGHAN #4]  C/P,P
O’DONNELL, KEVIN JR.  FIRE ON THE BORDER  C/P,P
O’DONNELL, KEVIN JR.  LAVA [@JOURNEYS OF MCGILL FEIGHAN #3]  C/P:2,P
O’DONNELL, KEVIN JR.  MAYFLIES  C/P:2,P
O’DONNELL, KEVIN JR.  ORACLE  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 290
AUTHOR  TITLE  CODE
O’DONNELL, KEVIN JR.  REEFS [@JOURNEYS OF MCGILL FEIGHAN #2]  C/P,P
O’DONNELL, KEVIN JR.  WAR OF OMISSION  C/P:2,P
O’DONNELL, PETER  LAST DAY IN LIMBO [MODESTY BLAISE]  SFWA-P
O’DONNELL, PETER  NIGHT OF MORNINGSTAR, THE [MODESTY BLAISE]  SFWA-P
O’DONNELL, PETER  SILVER MISTRESS, THE [MODESTY BLAISE]  SFWA-P
O’DONNELL, PETER  XANADU TALISMAN, THE [@MODESTY BLAISE]  P,SFWA-P
O’DONOGHUE, MICHAEL/SPRINGER, FRANK ADVENTURES OF PHOEBE ZEIT-GEIST, THE  CX
O’DONOHOE, NICK  GNOMEWRENCH IN THE DWARFWORKS, THE [GNOMEWRENCH 1]  C/P,P
O’DONOHOE, NICK  GNOMEWRENCH IN THE PEOPLEWORKS, THE [GNOMEWRENCH 2]  P
O’DONOHOE, NICK  HEALING OF CROSSROADS, THE [BJ VAUGHAN 3]  P
O’DONOHOE, NICK  MAGIC AND THE HEALING, THE [BJ VAUGHAN 1]  P,SFWA-P
O’DONOHOE, NICK  TOO TOO SOLID FLESH  P
O’DONOHOE, NICK  UNDER THE HEALING SIGN [BJ VAUGHAN 2]  P
ODYSSEY PUBLICATIONS  ACTION STORIES  SR-LPA
ODYSSEY PUBLICATIONS  GOLDEN FLEECE, THE  SR-LPA
ODYSSEY PUBLICATIONS  MAGIC CARPET MAGAZINE, THE  SR-LPA
ODYSSEY PUBLICATIONS  ORIENTAL STORIES  SR-LPA
ODYSSEY PUBLICATIONS  STRANGE TALES  SR-LPA
OFFUTT, ANDREW J.  ARDOR ON AROS  C/P,P
OFFUTT, ANDREW J.  CASTLE KEEPS, THE  C/P,P
OFFUTT, ANDREW J.  CHIEFTAIN OF ANDOR  C/P:2,P
OFFUTT, ANDREW J.  CONAN AND THE SORCERER [CONAN]  C/P,LP
OFFUTT, ANDREW J.  CONAN THE MERCENARY [CONAN]  P
OFFUTT, ANDREW J.  DEATHKNIGHT  P
OFFUTT, ANDREW J./LYON, RICHARD K.  DEMON IN THE MIRROR [WAR OF THE WIZARDS 1]  C/P,P
OFFUTT, ANDREW J.  EVIL IS LIVE SPELLED BACKWARDS  P
OFFUTT, ANDREW J./LYON, RICHARD K.  EYES OF SARSIS, THE [WAR OF THE WIZARDS 2]  P
OFFUTT, ANDREW J.  GALACTIC REJECTS, THE  H,P
OFFUTT, ANDREW J./BERRY, D. BRUCE  GENETIC BOMB  C/P,P
OFFUTT, ANDREW J.  GREAT TWENTY-FOUR HOUR "THING", THE  SR-L
OFFUTT, ANDREW J.  IRON LORDS, THE [WAR OF THE GODS ON EARTH 1]  C/P:2,P
OFFUTT, ANDREW J.  KING DRAGON  C/P,P
OFFUTT, ANDREW J.  LADY OF THE SNOWMIST, THE [WAR OF THE GODS ON EARTH 3]  C/P:2,P
OFFUTT, ANDREW J.  MESSENGER OF ZHUVASTOU  C/P:2,P
OFFUTT, ANDREW J.  MISTS OF DOOM, THE [CORMAC 1]  C/P,P
OFFUTT, ANDREW J.  MY LORD BARBARIAN  C/P,P
OFFUTT, ANDREW J.  SHADOW OF SORCERY, THE [THIEVES’ WORLD SHADOWSPAWN]  P,SFWA-P:2
OFFUTT, ANDREW J.  SHADOWS OUT OF HELL [WAR OF THE GODS ON EARTH 2]  C/P,P
OFFUTT, ANDREW J.  SHADOWSPAWN [THIEVES’ WORLD SHADOWSPAWN]  H:2,P
OFFUTT, ANDREW J.  SIGN OF THE MOONBOW, THE [CORMAC 6]  P
OFFUTT, ANDREW J.  SWORD OF SKELOS, THE [BANTAM CONAN 3]  C/P,P
OFFUTT, ANDREW J.  SWORD OF THE GAEL [CORMAC 4]  C/P:2,P
OFFUTT, ANDREW J.  SWORDS AGAINST DARKNESS I  C/PA,PA
OFFUTT, ANDREW J.  SWORDS AGAINST DARKNESS II  C/PA,PA
OFFUTT, ANDREW J.  SWORDS AGAINST DARKNESS III  C/PA,PA
OFFUTT, ANDREW J.  SWORDS AGAINST DARKNESS IV  C/PA,PA
OFFUTT, ANDREW J.  SWORDS AGAINST DARKNESS V  PA
OFFUTT, ANDREW J./TAYLOR, KEITH  TOWER OF DEATH, THE [CORMAC 2]  P
OFFUTT, ANDREW J.  UNDYING WIZARD, THE  C/P:2,P
OFFUTT, ANDREW J./LYON, RICHARD K.  WEB OF THE SPIDER [WAR OF THE WIZARDS 3]  P
OFFUTT, ANDREW J./TAYLOR, KEITH  WHEN DEATH BIRDS FLY [CORMAC 3]  C/P,P
OHLANDER, BEN/DRAKE, DAVID  ENEMY OF MY ENEMY [TERRA NOVA]  C/P,P
OHLE, DAVID  AGE OF SINATRA, THE  LP
O’KEEFE, CLAUDIA  BLACK SNOW DAYS  C/P,P
OKRAND, MARC  CONVERSATIONAL KLINGON (CASSETTE) [STAR TREK]  SR-MM
OKRAND, MARC  KLINGON DICTIONARY, THE [STAR TREK]  R/REF-P
OKRAND, MARC  KLINGON DICTIONARY, THE (WITH 1992 ADDENDUM) [STAR TREK]  R/REF-P,C/REF-H

OKUDA, MICHAEL/OKUDA, DENISE/ MIREK, DEBBIE STAR TREK ENCYCLOPEDIA, THE

R/REF-H,C/REF-H

OKUN, LAWRENCE  ON THE EIGHTH DAY  P
OLAN, SUSAN TORIAN  EARTH REMEMBERS, THE  C/P,P

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

ARTHUR C. CLARKE

R/REF-H

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

ISAAC ASIMOV

R/REF-H

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

PHILIP K. DICK

R/REF-H

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

RAY BRADBURY

R/REF-H

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

TIME OF PASSAGE

HA

OLANDER, JOSEPH D./ GREENBERG, MARTIN H.

URSULA K. LE GUIN

R/REF-H

OLDEN, MARC  BOOK OF SHADOWS  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 291
AUTHOR  TITLE  CODE
OLDEN, MARC  DAI-SHO  H
OLDEN, MARC  POE MUST DIE  P
O’LEARY, PATRICK  DOOR NUMBER THREE  H,LP
O’LEARY, PATRICK  GIFT, THE  H,LP
O’LEARY, PATRICK  IMPOSSIBLE BIRD, THE  LP
OLECK, JACK  TALES FROM THE CRYPT  C/P,P
OLECK, JACK  TALES FROM THE HOUSE OF MYSTERY #1  C/P,P
OLECK, JACK  TALES FROM THE HOUSE OF MYSTERY #2  C/P,P
OLECK, JACK  VAULT OF HORROR, THE  P
OLEMY, P. T.  CLONES, THE  P
OLEXA, KEITH  ABSOLUTELY BRILLIANT IN CHROME  LPA
OLIVASTRO, DOMINIC  ANCIENT PUZZLES  [C/RAN-H]
OLIVER, CHAD  ANOTHER KIND  SR-H,C/P:2,P
OLIVER, CHAD  EDGE OF FOREVER, THE  H
OLIVER, CHAD  FAR FROM THIS EARTH  H
OLIVER, CHAD  FIRES OF FOREVER  B
OLIVER, CHAD  GIANTS IN THE DUST  C/P,P
OLIVER, CHAD  SHADOWS IN THE SUN  H:2,P
OLIVER, CHAD  SHORES OF ANOTHER SEA, THE  C/P,P
OLIVER, CHAD  STAR ABOVE IT, A  H
OLIVER, CHAD  UNEARTHLY NEIGHBORS  H:2,P
OLIVER, CHAD  WINDS OF TIME, THE  H:2,P
OLIVER, STANLEY  AC3  H
OLSA, JAROSLAV JR.  SRDCE TECHNOPOPU  FP
OLSEN, BOB  RHYTHM RIDES THE ROCKET  B
OLSEN, LANCE  TIME FAMINE  LP
OLSEN, LANCE  TONGUING THE ZEITGEIST  LP
OLSHAKER, MARK  EINSTEIN’S BRAIN  P
OLSON, MIKE  ALL ABOARD  LPA
OLTION, JERRY  ABANDON IN PLACE  LP
OLTION, JERRY  ALLIANCE [ROBOTS AND ALIENS 4]  P,SFWA-P:2
OLTION, JERRY  ANYWHERE BUT HERE  H
OLTION, JERRY  FRAME OF REFERENCE  C/P,P
OLTION, JERRY  GETAWAY SPECIAL, THE  H
OLTION, JERRY  HUMANITY [ROBOTS AND ALIENS 6]  P
OLTION, JERRY  MUDD IN YOUR EYE [STAR TREK]  C/P:2,P
OLTION, JERRY  TWILIGHT’S END [STAR TREK]  C/P,P
OLTION, JERRY  WHERE SEA MEETS SKY [STAR TREK CAPTAIN’S TABLE 6]  C/P:2,P
OLTION, KATHY/OLTION, JERRY  FLAMING ARROW, THE [@STAR TREK NEW EARTH #4]  C/P,P
O’MAHONY, DANIEL  CABINET OF LIGHT, THE [DOCTOR WHO/TIME HUNTER 0]  SR-H
O’MAHONY, DANIEL  FALLS THE SHADOW [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
O’MAHONY, DANIEL  MAN IN THE VELVET MASK, THE [DOCTOR WHO MISSING ADVENTURES]  C/P,P
ON SPEC EDITORIAL COLLECTIVE  ON SPEC: THE FIRST FIVE YEARS  PA
O’NEAL, KATHLEEN M.  ABYSS OF LIGHT, AN [GAMANT 1]  P
O’NEAL, KATHLEEN M.  REDEMPTION OF LIGHT [GAMANT 3]  P
O’NEAL, KATHLEEN N.  TREASURE OF LIGHT [GAMANT 2]  P

O’NEIL, DENNIS/ERWIN, STEVE/ GARCIA-LOPEZ, JOSE LUIS BATMAN RETURNS

CX

O’NEIL, DENNIS  BITE OF MONSTERS, THE  C/P:2,P
O’NEIL, DENNIS  HERO’S QUEST [@JUSTICE LEAGUE OF AMERICA GREEN LANTERN]  P
O’NEILL, GERARD K.  HIGH FRONTIER, THE  [C/RAN-P:2]
O’NEILL, JOSEPH  LAND UNDER ENGLAND  SR-P,H:2
O’NEILL, PAMELA/MOHAN, KIM  CHASE INTO SPACE [@CYBORG COMMANDO #2]  P
O’NEILL, PAMELA/MOHAN, KIM  PLANET IN PERIL [@CYBORG COMMANDO #1]  C/P:2,P
O’NEILL, PAMELA/MOHAN, KIM  ULTIMATE PRIZE, THE [@CYBORG COMMANDO #3]  P
O’NEILL, SCOTT  MARTIAN SEXPOT  SR-L
O’NEILL, T. WARREN  REFUTATION OF DARWINISM, THE  R/REF-H

O’NEILL, TIMOTHY R.

INDIVIDUATED HOBBIT:

JUNG, TOLKIEN AND THE ARCHETYPES OF MIDDLE-EARTH, THE

R/REF-H

ONIONS, OLIVER  COLLECTED GHOST STORIES OF OLIVER ONIONS, THE  LP
ONLEY, DAVID C.  SHUTTLE  P
ONOPA, ROBERT  PLEASURE TUBE, THE  C/P:2,P
OPAR PRESS  GHOST STORIES  R/IND
OPIK, ERNST J.  OSCILLATING UNIVERSE, THE  [C/RAN-P]
OPPEL, KENNETH  AIRBORN [MATT CRUSE 1]  H,P
OPPEL, KENNETH  SKYBREAKER [MATT CRUSE 2]  H
OPPENHEIM, E. PHILLIPS  DUMB GODS SPEAK, THE  H
OPPENHEIM, E. PHILLIPS  GREAT PRINCE SHAN, THE  H
OPPENHEIM, E. PHILLIPS  MR. MIRAKEL  H
OPPENHEIM, SHULAMITH  WORLD INVISIBLE, THE  C/P:2,P
ORAM, JOHN  COPENHAGEN AFFAIR, THE [@MAN FROM UNCLE #3]  C/P:2,P

ORAM, JOHN STONE-COLD DEAD IN THE MARKET AFFAIR, THE

[@MAN FROM UNCLE #22]

C/P,P

ORAM, JOHN STONE-COLD DEAD IN THE MARKET AFFAIR, THE (BRITISH)

[@MAN FROM UNCLE #4] SR-P

The MITSFS Pinkdex  September 24, 2007  Page 292
AUTHOR  TITLE  CODE

ORDWAY, FREDERICK I. III/ WAKEFORD, RONALD C.

CONQUERING THE SUN’S EMPIRE

[C/RAN-H]

ORE, REBECCA  ALIEN BOOTLEGGER  LP
ORE, REBECCA  ALIEN BOOTLEGGER AND OTHER STORIES  H
ORE, REBECCA  BECOMING ALIEN [BECOMING ALIEN 1]  H:2,P
ORE, REBECCA  BEING ALIEN [BECOMING ALIEN 2]  H:2,SFWA-P
ORE, REBECCA  GAIA’S TOYS  H:2,SFWA,LP
ORE, REBECCA  HUMAN TO HUMAN [BECOMING ALIEN 3]  H:2,SFWA-P,P
ORE, REBECCA  ILLEGAL REBIRTH OF BILLY THE KID, THE  H,P
ORE, REBECCA  OUTLAW SCHOOL  H,LP
ORE, REBECCA  SLOW FUNERAL  H,P
ORE, REBECCA  TIME’S CHILD  LP
ORE, REBECCA  TIME’S CHILD (PROOF)  SR-LP
O’REILLY, TIMOTHY  FRANK HERBERT  R/REF-H
O’REILLY, TIMOTHY  MAKER OF DUNE, THE  SFWA
ORGILL, DOUGLAS/GRIBBIN, JOHN  BROTHER ESAU  SFWA-P
ORGILL, DOUGLAS/GRIBBIN, JOHN  SIXTH WINTER, THE  H
O’RIORDAN, ROBERT  CADRE LUCIFER [CADRE 2]  C/P,SFWA-P,P
O’RIORDAN, ROBERT  CADRE MESSIAH [CADRE 3]  C/P:2,P
O’RIORDAN, ROBERT  CADRE ONE [CADRE 1]  C/P:2,P
ORKOW, BEN  WHEN TIME STOOD STILL  P
ORLOV, VLADIMIR  DANILOV THE VIOLIST  H
ORMAN, KATE  BLUE BOX [DOCTOR WHO PAST DOCTORS]  P
ORMAN, KATE  LEFT-HANDED HUMMINGBIRD, THE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
ORMAN, KATE  RETURN OF THE LIVING DAD [DOCTOR WHO NEW ADVENTURES]  C/P,P
ORMAN, KATE  ROOM WITH NO DOORS, THE [DOCTOR WHO NEW ADVENTURES]  C/P,P
ORMAN, KATE  SET PIECE [DOCTOR WHO NEW ADVENTURES]  C/P,P
ORMAN, KATE  SLEEPY [DOCTOR WHO NEW ADVENTURES]  C/P:2,P

ORMAN, KATE WALKING TO BABYLON

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

P,PA

ORMAN, KATE  YEAR OF INTELLIGENT TIGERS, THE [DOCTOR WHO EIGHTH DOCTOR]  P
ORMONDROYD, EDWARD  ALL IN GOOD TIME  P
ORMONDROYD, EDWARD  TIME AT THE TOP  P
ORMSBY, ALAN  MOVIE MONSTERS  R/REF-P
O’ROURKE, FRANK  INSTANT GOLD  H:2
ORR, A.  IN THE ICE KING’S PALACE [WORLD IN AMBER 2]  H,SFWA
ORR, A.  WORLD IN AMBER, THE [WORLD IN AMBER 1]  H:2,SFWA-P,P
ORTIZ, LUIS  EMSHWILLER: INFINITY X TWO  R/REF-H
ORWELL, GEORGE  ANIMAL FARM  C/P,P,H
ORWELL, GEORGE  NINETEEN EIGHTY-FOUR  C/P,P,H
OSBORN, DAVID  HEADS (PROOF)  SR-P
OSBORN, E. M.  SHORT VISIT TO ERGON  LP
OSBORNE, CARY  DARKLOOM  C/P,P
OSBORNE, CARY  DEATHWEAVE  C/P,P
OSBORNE, CARY  GLAIVE, THE [IROSHI 2]  P
OSBORNE, CARY  IROSHI [IROSHI 1]  C/P,P
OSBORNE, CARY  PERSEA [IROSHI 3]  P
OSBORNE, DAVID  ALIENS FROM SPACE  H
OSBORNE, DAVID  INVISIBLE BARRIERS  H:2
OSBORNE, DUFFIELD  SECRET OF THE CRATER, THE  H
OSBOURNE, LLOYD  ADVENTURER, THE  H
O’SHEA, PAT  HOUNDS OF THE MORRIGAN, THE  C/P,P
O’SHEA, PATTI  RAVYN’S FLIGHT [JARVED 9 1]  P
O’SHEA, SEAN  OPERATION BOUDOIR  P
OSIER, JOHN  RANKIN: ENEMY OF THE STATE  P
OSTOW, MICOL  CHANGELING PLACES [@CHARMED]  P
OSTOW, MICOL  HOUSE OF SHARDS [@CHARMED]  P
OTFINOSKI, STEVEN  VILLAGE OF VAMPIRES  SR-P
O’TOOLE, CHRISTOPHER  ALIEN EMPIRE  [C/RAN-H]
OTTEN, CHARLOTTE F.  LITERARY WEREWOLF, THE  LPA
OTTUM, BOB  ALL RIGHT, EVERYBODY OFF THE PLANET!  C/P,P,H
OUELLETTE, PIERRE  DEUS MACHINE, THE  C/P,P,H
OVENDEN, MICHAEL W.  LIFE IN THE UNIVERSE  [C/RAN-P]
OVERGARD, WILLIAM  DIVIDE, THE  P
OWEN, BETTY M  ELEVEN GREAT HORROR STORIES  PA
OWEN, BETTY M.  NINE STRANGE STORIES  C/PA,PA
OWEN, BETTY M.  STARSTREAK  PA
OWEN, BETTY M.  STORIES OF THE SUPERNATURAL  PA
OWEN, DEAN  BRIDES OF DRACULA, THE  SR-P
OWEN, DEAN  END OF THE WORLD  P
OWEN, DEAN  KONGA  SR-P
OWEN, FRANK  HUSBAND FOR KUTANI, A  SR-H
OWEN, JAMES A.  HERE, THERE BE DRAGONS  H
OWEN, WALTER  ’MORE THINGS IN HEAVEN...’  H

AUTHOR  TITLE  CODE
OWENS, EVERETT  REGENERATION [@X-FILES]  H
OWENS, MILES  DAUGHTER OF PROPHECY  LP
OWENS, ROBIN D.  GUARDIAN OF HONOR [LLADRANA 1]  LP
OWENS, ROBIN D.  PROTECTOR OF THE FLIGHT [LLADRANA 3]  LP
OWENS, ROBIN D.  SORCERESS OF FAITH [LLADRANA 2]  LP
OWENS, TED  HOW TO CONTACT SPACE PEOPLE  [C/RAN-H]
OWINGS, MARK  JAMES H. SCHMITZ: A BIBLIOGRAPHY  R/IND
OWINGS, MARK/CHALKER, JACK L.  REVISED H. P. LOVECRAFT BIBLIOGRAPHY, THE  R/IND
OWSTON, C. E.  SCARLET SKULL, THE  C/P,P
OZAKI, YEI THEODORA  JAPANESE FAIRY TALES  HA

The MITSFS Pinkdex  September 24, 2007  Page 295
AUTHOR  TITLE  CODE
PABEL MOEWIG KG  ALTMUTANTEN, DIE [@PERRY RHODAN SILBERBANDE #65]  FPR
PABEL MOEWIG KG  ANTI-UNIVERSUM [@PERRY RHODAN SILBERBANDE #68]  FPR
PABEL MOEWIG KG  APHILE [@PERRY RHODAN SILBERBANDE #81]  FPR
PABEL MOEWIG KG  DRITTE MACHT, DIE [@PERRY RHODAN SILBERBANDE #1]  FPR
PABEL MOEWIG KG  ERBE DER YULOCS, DAS [@PERRY RHODAN SILBERBANDE #71]  FPR
PABEL MOEWIG KG  GEHIRN IN FESSELN [@PERRY RHODAN SILBERBANDE #70]  FPR
PABEL MOEWIG KG  HYPERSEUCHE, DIE [@PERRY RHODAN SILBERBANDE #69]  FPR
PABEL MOEWIG KG  IM MAHLSTROM DER STERNE [@PERRY RHODAN SILBERBANDE #77]  FPR
PABEL MOEWIG KG  KONTAKTE MIT DER EWIGKEIT [@PERRY RHODAN SILBERBANDE #72]  FPR
PABEL MOEWIG KG  KONZIL DER SIEBEN [@PERRY RHODAN SILBERBANDE #74]  FPR
PABEL MOEWIG KG  LAREN, DIE [@PERRY RHODAN SILBERBANDE #75]  FPR
PABEL MOEWIG KG  MENSCHHEIT AM SCHEIDEWEG [@PERRY RHODAN SILBERBANDE #80]  FPR

PABEL MOEWIG KG PERRY RHODAN: DIE GROESSTE SCIENCE FICTION-SERIE DER WELT (CD)

[PERRY RHODAN]

SR-MM

PABEL MOEWIG KG  RAUMSCHIFF ERDE [@PERRY RHODAN SILBERBANDE #76]  FPR
PABEL MOEWIG KG  SCHACH DER FINSTERNIS [@PERRY RHODAN SILBERBANDE #73]  FPR
PABEL MOEWIG KG  SUCHE NACH DER ERDE [@PERRY RHODAN SILBERBANDE #78]  FPR
PACKARD, EDWARD  ALIEN INVADERS [@SPACE HAWKS #2]  P
PACKARD, EDWARD  CAVE OF TIME, THE  C/P,P
PACKARD, EDWARD  FASTER THAN LIGHT [@SPACE HAWKS #1]  P
PACKARD, EDWARD  NIGHT OF THE WEREWOLF  LP
PACKARD, EDWARD  THIRD PLANET FROM ALTAIR, THE  P
PACKER, PETER  BLAST OFF INTO SPACE [LOST IN SPACE]  SR-S
PADGETT, LEWIS/MOORE, C. L.  BEYOND EARTH’S GATES  SR-DD069
PADGETT, LEWIS  CHESSBOARD PLANET  GN5.26
PADGETT, LEWIS  DAY HE DIED, THE  SR-P
PADGETT, LEWIS  GNOME THERE WAS, A  H
PADGETT, LEWIS  LINE TO TOMORROW  C/P:2,P
PADGETT, LEWIS  MUTANT  H
PADGETT, LEWIS  WELL OF THE WORLDS  C/P:2,GN3.17
PAGE, GERALD W./REINHARDT, HANK  HEROIC FANTASY  C/PA,PA
PAGE, GERALD W.  NAMELESS PLACES  SR-HA,HA
PAGE, GERALD W.  YEAR’S BEST HORROR STORIES V, THE  PA
PAGE, GERALD W.  YEAR’S BEST HORROR STORIES VI, THE  PA
PAGE, GERALD W.  YEAR’S BEST HORROR STORIES VII, THE  PA
PAGE, IAN/DEVER, JOE  FORBIDDEN CITY, THE [@LONE WOLF #2]  P
PAGE, IAN/DEVER, JOE  GREY STAR THE WIZARD [@LONE WOLF #1]  P
PAGE, IAN/DEVER, JOE  WAR OF THE WIZARDS [@LONE WOLF #4]  P
PAGE, JAKE  OPERATION SHATTERHAND  P
PAGE, NORVELL W.  FLAME WINDS  C/P:2,P
PAGE, NORVELL W.  SONS OF THE BEAR-GOD  P
PAGE, THOMAS  HEPHAESTUS PLAGUE, THE  C/P,P,H
PAGE, THOMAS  MAN WHO WOULD NOT DIE, THE  P

PAINE, DORIS M./MARTINEZ, DIANA GUIDE TO SCIENCE FICTION:

EXPLORING POSSIBILITIES AND ALTERNATIVES

R/REF-P

PAL, GEORGE/MORHAIM, JOE  TIME MACHINE II  C/P,P
PALESTINE, EILEEN  STAR FLEET MEDICAL REFERENCE MANUAL  R/REF-H,C/REF-H
PALMATIER, JOSHUA  SKEWED THRONE, THE [VARIS 1]  H
PALMER, DAVID R.  EMERGENCE  C/P:2,P
PALMER, DAVID R.  THRESHOLD  C/P:2,P
PALMER, JANE  DRUNE, THE  LP
PALMER, JANE  PLANET DWELLER, THE [PLANET DWELLER 1]  P
PALMER, JANE  WATCHER, THE  SR-P
PALMER, JOSEPH  DRAGON STONES  P
PALMER, RAYMOND A.  REAL UFO INVASION, THE  [C/RAN-P]
PALMER, STEPHEN  FLOWERCRASH  LP
PALMER, THOMAS  DREAM SCIENCE  H
PALMIERI, MARCO  CONSTELLATIONS [STAR TREK]  LPA
PALMIERI, MARCO  DISTANT SHORES [STAR TREK VOYAGER]  LPA
PALMIERI, MARCO  LIVES OF DAX, THE [@STAR TREK]  LPA
PALMIERI, MARCO  PROPHECY AND CHANGE [STAR TREK DS9]  LPA
PALTOCK, ROBERT  LIFE & ADVENTURES OF PETER WILKINS, THE  H
PALUMBO, DENNIS  CITY WARS  C/P,P
PALWICK, SUSAN  FATE OF MICE, THE  LP
PALWICK, SUSAN  FLYING IN PLACE  H
PALWICK, SUSAN  NECESSARY BEGGAR, THE  P
PALWICK, SUSAN  SHELTER  LP
PANATI, CHARLES  LINKS  P
PANATI, CHARLES  PLEASURING OF RORY MALONE, THE  P
PANGBORN, EDGAR  COMPANY OF GLORY, THE  P
PANGBORN, EDGAR  DAVY  C/P:2,SFWA-P,P
PANGBORN, EDGAR  GOOD NEIGHBORS AND OTHER STRANGERS  H:2,P
PANGBORN, EDGAR  JUDGMENT OF EVE, THE  C/P:2,P
PANGBORN, EDGAR  MIRROR FOR OBSERVERS, A  H:2,LP,SFWA

AUTHOR  TITLE  CODE
PANGBORN, EDGAR  STILL I PERSIST IN WONDERING  C/P:2,P
PANGBORN, EDGAR  WEST OF THE SUN  H:2,P
PANSHIN, ALEXEI/PANSHIN, CORY  EARTH MAGIC  C/P:2,P
PANSHIN, ALEXEI  FAREWELL TO YESTERDAY’S TOMORROW  H:2,SFWA,P
PANSHIN, ALEXEI  HEINLEIN IN DIMENSION  R/REF-H,C/REF-H
PANSHIN, ALEXEI  MASQUE WORLD [ANTHONY VILLIERS 3]  C/P:2,P
PANSHIN, ALEXEI  RITE OF PASSAGE  C/P:2,P
PANSHIN, ALEXEI/PANSHIN, CORY  SF IN DIMENSION  R/REF-H
PANSHIN, ALEXEI  STAR WELL [ANTHONY VILLIERS 1]  SR-P,C/P:2,P
PANSHIN, ALEXEI  THURB REVOLUTION, THE [ANTHONY VILLIERS 2]  SR-P,C/P:2
PANSHIN, ALEXEI  TRANSMUTATIONS  SR-LP
PANSHIN, ALEXEI/PANSHIN, CORY  WORLD BEYOND THE HILL, THE  R/REF-H
PAOLINI, CHRISTOPHER  ELDEST [INHERITANCE 2]  H
PAOLINI, CHRISTOPHER  ERAGON [INHERITANCE 1]  H
PAPERBACK LIBRARY  BARNABAS COLLINS IN A FUNNY VEIN  [C/RAN-P]
PAPERBACK LIBRARY  INNER LANDSCAPE, THE  C/PA,PA
PAPSON, STEPHEN  ...IN A WORLD NOT OF HIS OWN MAKING  LP
PARAMOUNT  STAR TREK IV, THE VOYAGE HOME  A
PARAMOUNT PICTURES  STAR TREK QUOTE JOURNAL [STAR TREK]  [C/RAN-H]
PARAMOUNT PICTURES  WHAT LAY BEYOND [STAR TREK GATEWAYS #7]  HA,PA
PARDOE, BLAINE LEE  CALL OF DUTY [BATTLETECH]  P
PARDOE, BLAINE LEE  EXODUS ROAD [@BATTLETECH #33/TWILIGHT OF THE CLANS 1]  C/P,P
PARDOE, BLAINE LEE  HIGHLANDER GAMBIT [@BATTLETECH 18]  P
PARDOE, BLAINE LEE  IMPETUS OF WAR [@BATTLETECH 30]  P
PARDOE, BLAINE LEE  MEASURE OF A HERO [@BATTLETECH 48]  P
PARDOE, BLAINE LEE  OPERATION AUDACITY [BATTLETECH]  P
PARDOE, BLAINE LEE  ROAR OF HONOR [@MECHWARRIOR]  P
PARDOE, BLAINE LEE  SURRENDER YOUR DREAMS [@MECHWARRIOR DARK AGE/BATTLETECH]  P
PARDOE, BLAINE LEE  TARGET OF OPPORTUNITY [@MECHWARRIOR DARK AGE/BATTLETECH]  P
PARENTEAU, SHIRLEY  TALKING COFFINS OF CRYO-CITY, THE  H
PARET, JOHN  TEN THOUSAND YEARS IN A BLOCK OF ICE  H
PARIS, MATTHEW  HOLY CITY, THE  LP
PARK, PAUL  CELESTIS  LP
PARK, PAUL  CULT OF LOVING KINDNESS, THE [STARBRIDGE CHRONICLES 3]  H,P
PARK, PAUL  PRINCESS OF ROUMANIA, A [ROUMANIA 1]  P
PARK, PAUL  SOLDIERS OF PARADISE [STARBRIDGE CHRONICLES 1]  C/P,SFWA,P,H
PARK, PAUL  SUGAR FESTIVAL, THE [STARBRIDGE CHRONICLES 1,2]  H:2
PARK, PAUL  SUGAR RAIN [STARBRIDGE CHRONICLES 2]  C/P,P,H
PARK, PAUL  TOURMALINE, THE [ROUMANIA 2]  H
PARK, PAUL  WHITE TYGER, THE  H
PARK, SEVERNA  ANNUNCIATE, THE  H
PARK, SEVERNA  HAND OF PROPHECY  C/P,P
PARK, SEVERNA  SPEAKING DREAMS  P
PARKER, BRANT/HART, JOHNNY  KING IS A FINK, THE  [C/RAN-P]
PARKER, EDGAR  ENCHANTRESS, THE  H
PARKER, JOY  BLUE LIGHT RED LIGHT VOL. 2 #1  SFWA
PARKER, K. J.  BELLY OF THE BOW, THE [FENCER 2]  LP
PARKER, K. J.  COLOURS IN THE STEEL [FENCER 1]  LP
PARKER, K. J.  MEMORY [@SCAVENGER #3]  LP
PARKER, K. J.  PATTERN [@SCAVENGER #2]  LP
PARKER, K. J.  PROOF HOUSE, THE [FENCER 3]  LP
PARKER, K. J.  SHADOW [@SCAVENGER #1]  LP
PARKER, STEVE  REBEL WINTER [WARHAMMER 40K]  P
PARKIN, LANCE/CLAPHAM, MARK  BEIGE PLANET MARS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
PARKIN, LANCE  BIG HUNT, THE [BERNICE SUMMERFIELD]  H
PARKIN, LANCE  COLD FUSION [DOCTOR WHO MISSING ADVENTURES]  C/P,P
PARKIN, LANCE  DYING DAYS, THE [DOCTOR WHO NEW ADVENTURES]  C/P,P
PARKIN, LANCE  FATHER TIME [DOCTOR WHO EIGHTH DOCTOR]  P
PARKIN, LANCE  GALLIFREY CHRONICLES, THE [DOCTOR WHO EIGHTH DOCTOR]  P
PARKIN, LANCE  HISTORY OF THE UNIVERSE, A [DOCTOR WHO]  R/REF-P,C/REF-P
PARKIN, LANCE  INFINITY DOCTORS, THE [DOCTOR WHO EIGHTH DOCTOR]  P
PARKIN, LANCE  JUST WAR [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
PARKIN, LANCE  TRADING FUTURES [DOCTOR WHO EIGHTH DOCTOR]  P
PARKIN, LANCE  WARLORDS OF UTOPIA [@FACTION PARADOX 3]  LP
PARKIN, LANCE  WINNING SIDE, THE [@TIME HUNTER 1]  LP
PARKINSON, DAN  BLOOD TIES [@TIMECOP 3]  P
PARKINSON, DAN  COVENANT OF THE FORGE, THE [@DRAGONLANCE DWARVEN NATIONS #1] C/P,P
PARKINSON, DAN  FACES OF INFINITY [GATES OF TIME 2]  P
PARKINSON, DAN  GATES OF THORBARDIN, THE [@DRAGONLANCE HEROES II #2]  C/P:2,P
PARKINSON, DAN  GULLY DWARVES, THE [@DRAGONLANCE LOST HISTORIES #5]  P
PARKINSON, DAN  HAMMER AND AXE [@DRAGONLANCE DWARVEN NATIONS #2]  P
PARKINSON, DAN  PARADOX GATE [GATES OF TIME 3]  P
PARKINSON, DAN  SCAVENGER, THE [@TIMECOP 2]  P
PARKINSON, DAN  STARSONG  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 297
AUTHOR  TITLE  CODE
PARKINSON, DAN  SWORDSHEATH SCROLL, THE [@DRAGONLANCE DWARVEN NATIONS #3]  P
PARKINSON, DAN  VIPER’S SPAWN [@TIMECOP 1]  P
PARKINSON, DAN  WHISPERS, THE [GATES OF TIME 1]  P
PARKINSON, KEITH  FANTASY ART OF KEITH PARKINSON CALENDAR, THE (2005)  CAL
PARLIER, CAP  PHOENIX SEDUCTION, THE  P
PARNELL, FRANK H./ASHLEY, MIKE  MONTHLY TERRORS  R/IND
PARRISH, GLENN  INVASORES INVADIDOS, LOS  FP
PARRY, DAVID M.  SCARLET EMPIRE, THE  H
PARRY, DENNIS  SURVIVOR, THE  H
PARRY, MICHEL  CHARIOTS OF FIRE  P
PARRY, MICHEL  COUNTESS DRACULA  P
PARRY, MICHEL  DEVIL’S CHILDREN, THE  PA
PARRY, MICHEL  REIGN OF TERROR  PA
PARRY, MICHEL  RIVALS OF FRANKENSTEIN  PA
PARRY, MICHEL  SAVAGE HEROES  HA
PARRY, MICHEL  STRANGE ECSTASIES  PA
PARRY, MICHEL  SUPERHEROES  SR-PA

PARTRIDGE, NORMAN/ GREENBERG, MARTIN H.

IT CAME FROM THE DRIVE-IN

PA

PARTRIDGE, NORMAN  MR. FOX AND OTHER FERAL TALES  SR-LP
PASCALL, JEREMY  GOD: THE ULTIMATE AUTOBIOGRAPHY  H:2
PASSARELLA, J. G.  WITHER  H
PASSARELLA, JOHN  AVATAR [@ANGEL]  P
PASSARELLA, JOHN  GHOUL TROUBLE [BUFFY]  P
PASSARELLA, JOHN  MONOLITH [@ANGEL]  C/P,P
PATROUCH, JOSEPH F. JR.  SCIENCE FICTION OF ISAAC ASIMOV, THE  R/REF-H,C/REF-H
PATTERSON, JAMES  VIRGIN  H,P
PATTERSON, PETER  FAL THE DRAGONHARPER  LP
PATTON, CLIFF  OMNI STRAIN, THE  P
PATTON, FIONA  GOLDEN SWORD, THE  P
PATTON, FIONA  GRANITE SHIELD, THE [LIVING FLAME 3]  P
PATTON, FIONA  PAINTER KNIGHT, THE [LIVING FLAME 2]  P
PATTON, FIONA  SILVER LAKE, THE [@WARRIORS OF ESTAVIA #1]  H
PATTON, FIONA  STONE PRINCE, THE [LIVING FLAME 1]  C/P:2,P
PATTOU, EDITH  HERO’S SONG  H
PATZER, SIDNEY  SHIP FROM NOWHERE, THE (DOUBLE WITH THE MOON MIRAGE)  B
PAUL, BARBARA  BIBBLINGS  C/P,P
PAUL, BARBARA  EXERCISE FOR MADMEN, AN  C/P,P
PAUL, BARBARA  PILLARS OF SALT  C/P,P
PAUL, BARBARA  THREE-MINUTE UNIVERSE, THE [STAR TREK]  C/P:2,P
PAUL, BARBARA  UNDER THE CANOPY  C/P,P
PAUL, DONITA K.  DRAGONKNIGHT [DRAGONSPELL 3]  LP
PAUL, DONITA K.  DRAGONQUEST [DRAGONSPELL 2]  LP
PAUL, DONITA K.  DRAGONSPELL [DRAGONSPELL 1]  LP
PAUL, HUGO  MASTER OF THE UNDEAD  P
PAULSEN, GARY  COMPKILL  C/P,P
PAULSEN, GARY  IMPLOSION EFFECT, THE  P
PAUWELS, LOUIS/BERGIER, JACQUES  IMPOSSIBLE POSSIBILITIES  [C/RAN-P]
PAUWELS, LOUIS/BERGIER, JACQUES  MORNING OF THE MAGICIANS, THE  [C/RAN-P]
PAVER, MICHELLE  WOLF BROTHER [ANCIENT DARKNESS 1]  H
PAVIC, MILORAD  DICTIONARY OF THE KHAZARS (MALE EDITION)  LP
PAVLAT, BOB/EVANS, BILL  FANZINE INDEX  R/IND
PAVLOU, STEL  DECIPHER  P
PAXSON, DIANA L.  BOOK OF THE CAULDRON, THE [@HALLOWED ISLE #3]  LP
PAXSON, DIANA L.  BOOK OF THE SPEAR, THE [@HALLOWED ISLE #2]  LP
PAXSON, DIANA L.  BOOK OF THE STONE, THE [@HALLOWED ISLE #4]  H,LP
PAXSON, DIANA L.  BOOK OF THE SWORD, THE [@HALLOWED ISLE #1]  LP
PAXSON, DIANA L.  BRISINGAMEN  C/P,P
PAXSON, DIANA L.  DRAGONS OF THE RHINE, THE [WODAN’S CHILDREN 2]  C/P,P
PAXSON, DIANA L.  EARTHSTONE, THE [WESTRIA 4]  P,SFWA-P
PAXSON, DIANA L.  HALLOWED ISLE, THE [HALLOWED ISLE 1,2]  P
PAXSON, DIANA L.  JEWEL OF FIRE, THE [WESTRIA 7]  C/P,P
PAXSON, DIANA L.  LADY OF DARKNESS [WESTRIA 2]  C/P,P
PAXSON, DIANA L.  LADY OF LIGHT [WESTRIA 1]  P
PAXSON, DIANA L.  LORD OF HORSES, THE [WODAN’S CHILDREN 3]  P

PAXSON, DIANA L./ MARTINE-BARNES, ADRIENNE MASTER OF EARTH AND WATER [FIONN MAC CUMHAL 1]

H,P

PAXSON, DIANA L.  MISTRESS OF THE JEWELS, THE [WESTRIA 1,2]  C/P,P
PAXSON, DIANA L.  PARADISE TREE, THE  P
PAXSON, DIANA L.  SEA STAR, THE [WESTRIA 5]  P,SFWA-P
PAXSON, DIANA L.  SERPENT’S TOOTH, THE  H,P

PAXSON, DIANA L./ MARTINE-BARNES, ADRIENNE SHIELD BETWEEN THE WORLDS, THE [FIONN MAC CUMHAL 2]

H,P

PAXSON, DIANA L.  SILVERHAIR THE WANDERER [WESTRIA 3]  P

The MITSFS Pinkdex  September 24, 2007  Page 298
AUTHOR  TITLE  CODE
PAXSON, DIANA L.  SWORD AND SORCERESS XXI  PA

PAXSON, DIANA L./ MARTINE-BARNES, ADRIENNE SWORD OF FIRE AND SHADOW [FIONN MAC CUMHAL 3]

P

PAXSON, DIANA L.  WHITE MARE, RED STALLION  P
PAXSON, DIANA L.  WHITE RAVEN, THE  P
PAXSON, DIANA L.  WIND CRYSTAL, THE [WESTRIA 6]  C/P,SFWA-P,P
PAXSON, DIANA L.  WOLF AND THE RAVEN, THE [WODAN’S CHILDREN 1]  H,P
PAYER, SUE  SECOND BODY  P
PAYNE, MICHAEL H.  BLOOD JAGUAR, THE  H,P
PEAK, MICHAEL  CAT HOUSE  P
PEAK, MICHAEL  CATAMOUNT  P
PEAKE, MERVYN  GORMENGHAST [GORMENGHAST 2]  C/P:2,P,H
PEAKE, MERVYN  MR. PYE  SR-P
PEAKE, MERVYN  TITUS ALONE [GORMENGHAST 3]  C/P,P,H
PEAKE, MERVYN  TITUS GROAN [GORMENGHAST 1]  C/P,P,H
PEARL, JACK  DAM OF DEATH [INVADERS]  H
PEARL, JACK  SPACE EAGLE: OPERATION DOOMSDAY, THE  H
PEARL, JACK  SPACE EAGLE: OPERATION STAR VOYAGE, THE  H
PEARLMAN, DANIEL  MEMINI  H
PEARLMAN, GILBERT  YOUNG FRANKENSTEIN  C/P:2,P
PEARSON, DURK/SHAW, SANDY  LIFE EXTENSION  [C/RAN-H:3]
PEARSON, EDWARD  CHAMIEL  C/P,P
PEARSON, RYNE DOUGLAS  OCTOBER’S GHOST  P
PEARY, DANNY  OMNI’S SCREEN FLIGHTS/SCREEN FANTASIES  SFWA
PEAT, F. DAVID  ARTIFICIAL INTELLIGENCE: HOW MACHINES THINK  SFWA-P
PECCEI, AURELIO  ONE HUNDRED PAGES FOR THE FUTURE  [C/RAN-P]
PECK, CLAUDIA  SPIRIT CROSSINGS  P
PECK, IRA  TREASURY OF GREAT GHOST STORIES, A  PA
PECK, RICHARD  BLOSSOM CULP AND THE SLEEP OF DEATH [BLOSSOM CULP 4]  P
PECK, RICHARD  DREADFUL FUTURE OF BLOSSOM CULP, THE [BLOSSOM CULP 3]  C/P,P
PECK, RICHARD  GHOST BELONGED TO ME, THE [BLOSSOM CULP 1]  P
PECK, RICHARD  GHOSTS I HAVE BEEN [BLOSSOM CULP 2]  C/P:2,P
PECK, RICHARD E.  FINAL SOLUTION  H,SFWA
PECK, VERNON  SUPER SEX CAPTAIN  SR-L
PEDERSEN, TED  SPACE CAMP [STAR TREK DS9]  LP
PEDLER, KIT/DAVIS, GERRY  BRAINRACK  P
PEDLER, KIT/DAVIS, GERRY  DYNOSTAR MENACE, THE  H
PEDLER, KIT/DAVIS, GERRY  MUTANT 59: THE PLASTIC-EATERS  H:2,P
PEECHER, JOHN PHILLIP  MAKING OF STAR WARS: THE RETURN OF THE JEDI, THE  R/REF-P
PEEL, JOHN  BEWARE THE METAL CHILDREN [@OUTER LIMITS]  LP
PEEL, JOHN  DEATH OF PRINCES, THE [STAR TREK TNG]  C/P:2,P
PEEL, JOHN  DOCTOR WHO: THE CHASE [DOCTOR WHO]  P

PEEL, JOHN DOCTOR WHO: THE DALEK’S MASTERPLAN - PART I: MISSION TO THE UNKNOWN [DOCTOR WHO]

P

PEEL, JOHN  EVIL OF THE DALEKS [DOCTOR WHO]  P
PEEL, JOHN  EVOLUTION [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
PEEL, JOHN  HERE THERE BE DRAGONS [STAR TREK TNG]  C/P:3,P
PEEL, JOHN  INDEPENDENCE [QUANTUM LEAP]  P
PEEL, JOHN  INNOCENT, THE [@OUTER LIMITS]  LP
PEEL, JOHN  INVADERS, THE [@OUTER LIMITS]  LP
PEEL, JOHN  LEGACY OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
PEEL, JOHN  LOST, THE [@OUTER LIMITS]  LP
PEEL, JOHN  MELTDOWN [@TWENTY NINETY-NINE #5]  LP
PEEL, JOHN  OBJECTIVE: BAJOR [STAR TREK DS9]  C/P,P
PEEL, JOHN/NATION, TERRY  OFFICIAL DOCTOR WHO AND THE DALEKS BOOK, THE  R/REF-H
PEEL, JOHN  POWER OF THE DALEKS, THE [DOCTOR WHO]  C/P,P
PEEL, JOHN  PRISONERS OF PEACE [@STAR TREK DS9]  H,LP
PEEL, JOHN  SECRET OF DRAGONHOME, THE  P
PEEL, JOHN  TIMEWYRM: GENESYS [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
PEEL, JOHN  UPTIME, DOWNTIME  H
PEEL, JOHN  WAR OF THE DALEKS [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
PEEL, JOHN  ZANTI MISFITS, THE [@OUTER LIMITS]  LP
PEEPLES, SAMUEL A.  WHERE NO MAN HAS GONE BEFORE [STAR TREK]  SR-S
PEERY, NORRIS RAY  AUDIO EXTRACTS (CD)  SR-MM
PEIRCE, HAYFORD  BLACK HOLE PLANET  H
PEIRCE, HAYFORD  BURR IN THE GARDEN OF EDEN, THE [MACNAIR OF MACNAIR 2]  LP
PEIRCE, HAYFORD  DINOSAUR PARK  P
PEIRCE, HAYFORD  NAPOLEON DISENTIMED [MACNAIR OF MACNAIR 1]  C/P:2,SFWA-P,P
PEIRCE, HAYFORD  PHYLUM MONSTERS  C/P,SFWA-P,P
PEIRCE, HAYFORD  THIRTEENTH MAJESTRAL, THE  C/P:2,P
PELAN, JOHN  ANTIQUE VINTAGE, AN  B
PELAN, JOHN/ADAMS, BENJAMIN  CHILDREN OF CTHULHU, THE  SR-LPA
PELLEGRINO, CHARLES  DUST  H

The MITSFS Pinkdex  September 24, 2007  Page 299
AUTHOR  TITLE  CODE

PELLEGRINO, CHARLES/ ZEBROWSKI, GEORGE DYSON SPHERE [@STAR TREK TNG]

C/P:2,P

PELLEGRINO, CHARLES  FLYING TO VALHALLA  C/P,P,H

PELLEGRINO, CHARLES/ ZEBROWSKI, GEORGE KILLING STAR, THE

C/P:2,P

PELOSO, DOMINIC  CITY OF PILLARS  LP
PELZ, BRUCE  FANTASY SHOWCASE TAROT, THE [TAROT DECK]  SR-PA
PEMBERTON, VICTOR  FURY FROM THE DEEP [DOCTOR WHO]  SR-P
PEMBERTON, VICTOR  PESCATONS, THE [@DOCTOR WHO]  P
PEMBROKE, LYNNE  CATERPILLARIAN CHRONICLES, THE  P
PENDLETON, DON  CATACLYSM  C/P:2,P
PENDLETON, DON  EYE TO EYE  H,P
PENDLETON, DON  GODMAKERS, THE  C/P:2,P
PENDLETON, DON  GUNS OF TERRA 10, THE  P
PENDLETON, DON  NINETEEN EIGHTY-NINE: POPULATION DOOMSDAY  P
PENDLETON, DON  REVOLT!  SR-P,C/P
PENDLETON, DON  TIME TO TIME [ASHTON FORD 6]  P
PENSWICK, NEIL  PIT, THE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
PEOPLE’S LITERATURE  VISITOR FROM 50,000 YEARS AGO, THE (CHINESE)  FP
PERCY, WALDER  THANATOS SYNDROME, THE  H,P
PERDUE, LEWIS  DA VINCI LEGACY, THE  P
PEREZ, LANA  FIGURE IN THE FROST [@KNIGHTS OF THE SILVER DRAGON #5]  LP
PERKINS, SHELDON  POLARIS  P
PERMAN, DONALD/LEIGHT  WORLD’S WORST WASP JOKES, THE  [C/RAN-P]
PERRET, PATTI  FACES OF FANTASY, THE  R/REF-H
PERRET, PATTI  FACES OF SCIENCE FICTION, THE  R/REF-H
PERRIMAN, COLE  TERMINAL GAMES  P
PERRIN, DON/WEIS, MARGARET  DRACONIAN MEASURES [@DRAGONLANCE KANG’S REGIMENT #2]  P
PERRIN, DON  THEROS IRONFELD [@DRAGONLANCE WARRIORS #4]  P
PERRY, ANNE  COME ARMAGEDDON  H
PERRY, MARK C.  MORIGU: THE DEAD [MORIGU 1]  P
PERRY, MARK C.  MORIGU: THE DESECRATION [MORIGU 2]  P
PERRY, MARK C./PERRY, MAGAHN  WARLOCK’S BLADE, A  P,SFWA-P
PERRY, ROBERT/TUCKER, MIKE  COMPANION PIECE [DOCTOR WHO]  SR-H
PERRY, ROBERT/TUCKER, MIKE  MATRIX [DOCTOR WHO PAST DOCTORS]  C/P,P
PERRY, ROBERT/TUCKER, MIKE  STORM HARVEST [DOCTOR WHO PAST DOCTORS]  C/P,P
PERRY, S. D.  AVATAR, BOOK ONE [@STAR TREK DS9]  P
PERRY, S. D.  AVATAR, BOOK TWO [@STAR TREK DS9]  P
PERRY, S. D.  BERSERKER [@ALIENS 9]  P
PERRY, S. D.  CALIBAN COVE [@RESIDENT EVIL #2]  P
PERRY, S. D.  CITY OF THE DEAD [@RESIDENT EVIL #3]  P
PERRY, S. D.  CLOAK [STAR TREK SECTION 31]  C/P,P
PERRY, S. D.  CODE: VERONICA [@RESIDENT EVIL #6]  P
PERRY, S. D.  LABYRINTH [@ALIENS 7]  P
PERRY, S. D.  NEMESIS [@RESIDENT EVIL #5]  P
PERRY, S. D.  RISING SON [@STAR TREK DS9]  P
PERRY, S. D.  UMBRELLA CONSPIRACY, THE [@RESIDENT EVIL #1]  P
PERRY, S. D.  UNDERWORLD [@RESIDENT EVIL #4]  P
PERRY, S. D.  UNITY [@STAR TREK DS9]  P
PERRY, S. D.  VIRUS  P
PERRY, S. D.  WAR [@ALIENS VS. PREDATOR 3]  C/P,P
PERRY, S. D.  ZERO HOUR [@RESIDENT EVIL #0]  P
PERRY, STEVE  ALBINO KNIFE, THE [MATADOR 5]  C/P:2,SFWA-P:2,P
PERRY, STEVE  BLACK STEEL [MATADOR 6]  C/P,SFWA-P:2,P
PERRY, STEVE  BROTHER DEATH [MATADOR 7]  P,SFWA-P:2
PERRY, STEVE  CIVIL WAR SECRET AGENT  P
PERRY, STEVE  CONAN THE DEFIANT [CONAN]  SFWA-P,LP
PERRY, STEVE  CONAN THE FEARLESS [CONAN]  H,SFWA-P
PERRY, STEVE  CONAN THE FORMIDABLE [CONAN]  C/P,SFWA-P,LP,P
PERRY, STEVE  CONAN THE FREELANCE [CONAN]  P,SFWA-P
PERRY, STEVE  CONAN THE INDOMITABLE [CONAN]  C/P,SFWA,SFWA-P,LP
PERRY, STEVE  DIGITAL EFFECT, THE  C/P,P
PERRY, STEVE  EARTH HIVE [@ALIENS #1]  C/P,P
PERRY, STEVE/PERRY, STEPHANI  FEMALE WAR, THE [@ALIENS #3]  C/P,P
PERRY, STEVE  FOREVER DRUG, THE  C/P:2,P
PERRY, STEVE/PERRY, DAL  GANGSTER CONSPIRACY, THE [@STAR RISK,LTD.]  P
PERRY, STEVE  LONE STAR [STELLAR RANGER 2]  P
PERRY, STEVE  MACHIAVELLI INTERFACE, THE [MATADOR 3]  C/P:2,P
PERRY, STEVE  MAN WHO NEVER MISSED, THE [MATADOR 1]  C/P,P
PERRY, STEVE  MATADORA [MATADOR 2]  C/P:2,P
PERRY, STEVE/SOLOMON, ED  MEN IN BLACK [MEN IN BLACK]  C/P:2,P
PERRY, STEVE  MUSASHI FLEX, THE  P
PERRY, STEVE  NIGHTMARE ASYLUM [@ALIENS #2]  P
PERRY, STEVE  97TH STEP, THE [MATADOR 0]  C/P:2,SFWA-P,P

AUTHOR  TITLE  CODE
PERRY, STEVE/REAVES, MICHAEL  OMEGA CAGE, THE [MATADOR 4]  C/P,SFWA-P,P
PERRY, STEVE/PERRY, STEPHANI  PREY [@ALIENS VS. PREDATOR 1]  P
PERRY, STEVE  SHADOWS OF THE EMPIRE [@STAR WARS]  H:2,P
PERRY, STEVE  SPINDOC  C/P:2,SFWA-P,P
PERRY, STEVE  STELLAR RANGER [STELLAR RANGER 1]  C/P,P
PERRY, STEVE  TARGET EARTH [@PRIMORTALS]  P
PERRY, STEVE/PERRY, DAL  TITAN A.E.  P
PERRY, STEVE  TRINITY VECTOR, THE  C/P:2,P
PERUTZ, LEO  MASTER OF THE DAY OF JUDGMENT, THE  SR-P
PESEK, LUDEK  EARTH IS NEAR, THE  C/P,P
PESEK, LUDEK  LOG OF A MOON EXPEDITION  H
PETAJA, EMIL  ALPHA YES, TERRA NO!  DM121
PETAJA, EMIL  AS DREAM AND SHADOW  SR-H
PETAJA, EMIL  CAVES OF MARS, THE  DM133
PETAJA, EMIL  DOOM OF THE GREEN PLANET [GREEN PLANET 2]  DH070
PETAJA, EMIL  LORD OF THE GREEN PLANET [GREEN PLANET 1]  DH022
PETAJA, EMIL  NETS OF SPACE, THE  C/P,P
PETAJA, EMIL  PATH BEYOND THE STARS, THE  C/P:2,P
PETAJA, EMIL  PRISM, THE  DH051
PETAJA, EMIL  SAGA OF LOST EARTHS [KALEVALA 1]  C/P:2,P
PETAJA, EMIL  SEED OF THE DREAMERS  D06707
PETAJA, EMIL  SHOWCASE OF FANTASY ART  A
PETAJA, EMIL  STAR MILL, THE [KALEVALA 2]  C/P:2,P
PETAJA, EMIL  STARDRIFT  H
PETAJA, EMIL  STOLEN SUN, THE [KALEVALA 3]  DG618
PETAJA, EMIL  TIME TWISTER, THE  C/P:2,P
PETAJA, EMIL  TRAMONTANE [KALEVALA 4]  DH036
PETERKIEWICZ, JERZY  INNER CIRCLE  H
PETERS, DAVID  EXILE [@PHOTON #5]  P
PETERS, DAVID  FOR THE GLORY [@PHOTON #1]  P
PETERS, DAVID  HIGH STAKES [@PHOTON #2]  P
PETERS, DAVID  IN SEARCH OF MOM [@PHOTON #3]  P
PETERS, DAVID  SKIN DEEP [@PHOTON #6]  P
PETERS, DAVID  THIS IS YOUR LIFE, BHODILI [@PHOTON #4]  P
PETERS, ELIZABETH  CAMELOT CAPER, THE  SFWA-P
PETERS, ELIZABETH  DEVIL-MAY-CARE  SFWA-P
PETERS, ELIZABETH  DIE FOR LOVE  SFWA-P
PETERS, ELIZABETH  JACKAL’S HEAD, THE  SFWA-P
PETERS, ELIZABETH  LION IN THE VALLEY  SFWA-P
PETERS, ELIZABETH  NIGHT OF FOUR HUNDRED RABBITS, THE  SFWA-P
PETERS, ELIZABETH  TROJAN GOLD  SFWA-P:2
PETERS, ELLIS  MORBID TASTE FOR BONES, A  P
PETERS, L. T.  11TH PLAGUE, THE  C/P,P,H
PETERS, OTHELLO  WHISPERS FROM THE DARK SIDE OF TOMORROW  P
PETERS, RALPH  RED ARMY  H
PETERS, RALPH  WAR IN 2020, THE  H:2,P
PETERSEN, GAIL  MAKING OF A MONSTER, THE  P
PETERSON, JOHN VICTOR  ROCK THE BIG ROCK  P
PETREY, SUSAN  GIFTS OF BLOOD  P,SFWA-P
PETRONIUS ARBITER, TITUS  SATYRICON  P
PETRUCHA, STEFAN  ASSAMITE [@WORLD OF DARKNESS DARK AGES CLAN #2]  P
PETRUCHA, STEFAN/PENDLETON, THOMAS  LURKER  LP
PETRUCHA, STEFAN/PENDLETON, THOMAS  2 L8 4 U (PROOF) [@WICKED DEAD #1]  SR-LP
PETRUCHA, STEFAN  TUNNEL AT THE END OF THE LIGHT, THE [@TIME HUNTER 2]  LP
PETTERSON, VICKI  SCENT OF SHADOWS, THE [ZODIAC 1]  P
PETTERSON, VICKI  TASTE OF NIGHT, THE [ZODIAC 2]  P
PETYO, ROBERT  INSTITUTE, THE  C/P,P
PEYRAMAURE, MICHEL  VALLEE DES MAMMOUTHS, LA  FH
PEZET, A. WASHINGTON  ARISTOKIA  SR-H
PFEFFER, SUSAN BETH  FUTURE FORWARD [REWIND TO YESTERDAY 2]  H
PFEFFERLE, SETH  STICKMAN  SFWA-P
PFEIL, DONALD J.  LOOK BACK TO EARTH  P
PFEIL, DONALD J.  THROUGH THE REALITY WARP  C/P,P
PFEIL, DONALD J.  VOYAGE TO A FORGOTTEN SUN  C/P:2,P
PHILBRICK, RODMAN/HARNETT, LYNN  ABDUCTION  P
PHILLIES, GEORGE  MINUTEGIRLS, THE  LP
PHILLIES, GEORGE  NINE GEES  LP
PHILLIES, GEORGE  THIS SHINING SEA  LP
PHILLIFENT, JOHN T.  CORFU AFFAIR, THE [@MAN FROM UNCLE #20]  P
PHILLIFENT, JOHN T.  GENIUS UNLIMITED  C/P:2,P
PHILLIFENT, JOHN T.  HIERARCHIES  D53415
PHILLIFENT, JOHN T.  KING OF ARGENT  C/P:2,P
PHILLIFENT, JOHN T.  LIFE WITH LANCELOT  D48245
PHILLIFENT, JOHN T.  MAD SCIENTIST AFFAIR, THE [@MAN FROM UNCLE #5]  C/P:2,P

AUTHOR  TITLE  CODE
PHILLIFENT, JOHN T.  POWER CUBE AFFAIR, THE [@MAN FROM UNCLE #19]  P
PHILLIPS, ALEXANDER M.  DEATH OF THE MOON  B
PHILLIPS, ALEXANDER M.  MISLAID CHARM, THE  SR-H
PHILLIPS, DONALD G.  STAR LORD [@BATTLETECH 23]  P
PHILLIPS, GENE D.  STANLEY KUBRICK: A FILM ODYSSEY  R/REF-P
PHILLIPS, GORDON  HIGH EXPLOSIVE  H
PHILLIPS, MARK  BRAIN TWISTER  C/P:2,P
PHILLIPS, MARK  IMPOSSIBLES, THE  C/P:2,P
PHILLIPS, MARK  SUPERMIND  C/P:2,P
PHILLIPS, ROG  TIME TRAP  SR-P
PHILLIPS, ROG  WORLD OF IF  SR-P
PHILLIPS, ROG  WORLDS WITHIN  SR-P
PHILLIPS, TERRY  SOULFORGE, THE  P
PHILLPOTTS, EDEN  ADDRESS UNKNOWN  H
PHILLPOTTS, EDEN  GREY ROOM, THE  H
PHILMUS, ROBERT M.  INTO THE UNKNOWN  R/REF-H
PHINNEY, WILLIAM C., ET AL.  LUNAR RESOURCES AND THEIR UTILIZATION  [C/RAN-H:2]
PHIPPEN, MARK/LYON, SHAUN  MISSING PIECES [DOCTOR WHO]  VLPA
PHIPPEN, MARK/FAYLE, HELEN  PERFECT TIMING [DOCTOR WHO]  SR-LPA
PHIPSON, JOAN  CATS, THE  P
PHYLOS  DWELLER ON TWO PLANETS, A  H
PICARD, BARBARA LEONIE  FAUN AND THE WOODCUTTER’S DAUGHTER, THE  H
PICARD, BARBARA LEONIE  HERO TALES FROM THE BRITISH ISLES  P
PICCIRILLI, TOM  CHOIR OF ILL CHILDREN, A  P
PICCIRILLI, TOM  CHOIR OF ILL CHILDREN, A (PROOF)  SR-P
PIDDOCK, CHARLES  MONSTER FLY, THE  P
PIERCE, J. CALVIN  DOOR TO AMBERMERE, THE [AMBERMERE 1]  C/P,P
PIERCE, J. CALVIN  SORCERESS OF AMBERMERE, THE [AMBERMERE 2]  P,SFWA-P
PIERCE, J. CALVIN  WIZARD OF AMBERMERE, THE [AMBERMERE 3]  P
PIERCE, MEREDITH ANN  DARKANGEL, THE [DARKANGEL 1]  C/P,P
PIERCE, MEREDITH ANN  DARKANGEL TRILOGY, THE [DARKANGEL 1,2,3]  H
PIERCE, MEREDITH ANN  FIREBRINGER TRILOGY, THE [FIREBRINGER 1,2,3]  H
PIERCE, MEREDITH ANN  GATHERING OF GARGOYLES, A [DARKANGEL 2]  P,SFWA-P
PIERCE, MEREDITH ANN  PEARL OF THE SOUL OF THE WORLD, THE [DARKANGEL 3]  P
PIERCE, MEREDITH ANN  SON OF SUMMER STARS, THE [FIREBRINGER 3]  P
PIERCE, MEREDITH ANN  WOMAN WHO LOVED REINDEER, THE  C/P,SFWA-P,P
PIERCE, TAMORA  ALANNA [SONG OF THE LIONESS 1]  P
PIERCE, TAMORA  BRIAR’S BOOK [@CIRCLE OF MAGIC #4]  C/P,P
PIERCE, TAMORA  DAJA’S BOOK [@CIRCLE OF MAGIC #3]  P
PIERCE, TAMORA  EMPEROR MAGE [@IMMORTALS (PIERCE) #3]  P
PIERCE, TAMORA  FIRST TEST [@PROTECTOR OF THE SMALL 1]  H
PIERCE, TAMORA  IMMORTALS, THE [IMMORTALS (PIERCE) 1,2,3,4]  H
PIERCE, TAMORA  IN THE HAND OF THE GODDESS [SONG OF THE LIONESS 2]  P
PIERCE, TAMORA  LADY KNIGHT [@PROTECTOR OF THE SMALL 4]  H
PIERCE, TAMORA  LIONESS RAMPANT [SONG OF THE LIONESS 4]  H,P
PIERCE, TAMORA  PAGE [@PROTECTOR OF THE SMALL 2]  H
PIERCE, TAMORA  REALMS OF THE GODS, THE [@IMMORTALS (PIERCE) #4]  P
PIERCE, TAMORA  SANDRY’S BOOK [@CIRCLE OF MAGIC #1]  P
PIERCE, TAMORA  SQUIRE [@PROTECTOR OF THE SMALL 3]  H:2
PIERCE, TAMORA  TERRIER [@BEKA COOPER #1]  H
PIERCE, TAMORA  TRICKSTERS [DAUGHTER OF THE LIONESS 1,2]  H
PIERCE, TAMORA  TRIS’S BOOK [@CIRCLE OF MAGIC #2]  P
PIERCE, TAMORA  WILD MAGIC [@IMMORTALS (PIERCE) #1]  P
PIERCE, TAMORA  WILL OF THE EMPRESS, THE [CIRCLE OF MAGIC]  H
PIERCE, TAMORA  WOLF-SPEAKER [@IMMORTALS (PIERCE) #2]  P
PIERCE, TAMORA  WOMAN WHO RIDES LIKE A MAN, THE [SONG OF THE LIONESS 3]  H,P
PIERCY, MARGE  HE, SHE AND IT  H,P
PIERCY, MARGE  WOMAN ON THE EDGE OF TIME  C/P,P
PIERSON, CHRIS  BLADES OF THE TIGER [@DRAGONLANCE TALADAS #1]  P
PIERSON, CHRIS  CHOSEN OF THE GODS [@DRAGONLANCE KINGPRIEST #1]  P
PIERSON, CHRIS  DEZRA’S QUEST [@DRAGONLANCE BRIDGES OF TIME]  P
PIERSON, CHRIS  DIVINE HAMMER [@DRAGONLANCE KINGPRIEST #2]  C/P,P
PIERSON, CHRIS  SACRED FIRE [@DRAGONLANCE KINGPRIEST #3]  P
PIERSON, CHRIS  SHADOW OF THE FLAME [@DRAGONLANCE TALADAS #3]  P
PIERSON, CHRIS  TRAIL OF THE BLACK WYRM [@DRAGONLANCE TALADAS #2]  P
PIERSON, CLARA D.  AMONG THE FOREST PEOPLE  H
PIKE, CHRISTOPHER  LISTENERS, THE  P
PIKE, CHRISTOPHER  RED DICE [@LAST VAMPIRE #3]  P
PIKE, CHRISTOPHER  SATI  P
PIKE, CHRISTOPHER  SEASON OF PASSAGE, THE  H
PIKE, CHRISTOPHER  STARLIGHT CRYSTAL, THE  P
PIKE, CHRISTOPHER  TACHYON WEB, THE  P
PIKE, E. ROYSTON  STRANGE WAYS OF MAN, THE  [C/RAN-P]
PINCHER, CHAPMAN  NOT WITH A BANG  H:3,P

The MITSFS Pinkdex  September 24, 2007  Page 302
AUTHOR  TITLE  CODE
PINCKNEY, JOSEPHINE  GREAT MISCHIEF  H,P
PINI, RICHARD  AGAINST THE WIND [@BLOOD OF TEN CHIEFS #4]  C/PA,LPA,SFWA

PINI, RICHARD/ASPRIN, ROBERT LYNN/ ABBEY, LYNN BLOOD OF TEN CHIEFS, THE [@BLOOD OF TEN CHIEFS #1]

SFWA-P,C/PA,PA

PINI, RICHARD  DARK HOURS [@BLOOD OF TEN CHIEFS #5]  LPA
PINI, RICHARD  RETURN TO CENTAUR [XANTH]  CX
PINI, RICHARD  WINDS OF CHANGE [@BLOOD OF TEN CHIEFS #3]  PA,C/PA,SFWA-P,SFWA

PINI, RICHARD/ASPRIN, ROBERT LYNN/ ABBEY, LYNN WOLFSONG [@BLOOD OF TEN CHIEFS #2]

SFWA-P,C/PA,LPA

PINI, WENDY/PINI, RICHARD  CAPTIVES OF BLUE MOUNTAIN [@COMPLETE ELFQUEST #3]  CX
PINI, WENDY/PINI, RICHARD  ELFQUEST 1  CX
PINI, WENDY/PINI, RICHARD  ELFQUEST 1 (FIRST HALF)  SCX
PINI, WENDY/PINI, RICHARD  ELFQUEST 2  CX
PINI, WENDY/PINI, RICHARD  ELFQUEST 3  CX
PINI, WENDY/PINI, RICHARD  ELFQUEST 4  CX
PINI, WENDY/PINI, RICHARD  ELFQUEST: JOURNEY TO SORROW’S END  C/P:2,LP
PINI, WENDY/PINI, RICHARD  FIRE AND FLIGHT [@COMPLETE ELFQUEST #1]  CX
PINI, WENDY/PINI, RICHARD  FORBIDDEN GROVE, THE [@COMPLETE ELFQUEST #2]  CX
PINI, WENDY/PINI, RICHARD  QUEST BEGINS, THE [COMPLETE ELFQUEST]  LP
PINI, WENDY/PINI, RICHARD  QUEST’S END [@COMPLETE ELFQUEST #4]  CX
PINI, WENDY/PINI, RICHARD  SECRET OF TWO-EDGE, THE [@COMPLETE ELFQUEST #6]  CX
PINI, WENDY/PINI, RICHARD  SIEGE AT BLUE MOUNTAIN [@COMPLETE ELFQUEST #5]  CX
PINKWATER, DANIEL  BORGEL  SFWA
PINKWATER, DANIEL  MAGIC MOSCOW, THE  SFWA-P
PINKWATER, DANIEL MANUS  GUYS FROM SPACE  SFWA-P
PINKWATER, DANIEL MANUS  LIZARD MUSIC  P
PINKWATER, DANIEL MANUS  SNARKOUT BOYS AND THE AVOCADO OF DEATH, THE  P
PINKWATER, DANIEL MANUS  YOUNG ADULTS  LP
PINTO, RICARDO  CHOSEN, THE [STONE DANCE OF THE CHAMELEON 1]  C/P,P,H
PINTO, RICARDO  STANDING DEAD, THE [STONE DANCE OF THE CHAMELEON 2]  H
PIOMBO, AKBAR DEL  IS THAT YOU SIMON?  SR-P
PIPER, H. BEAM  COMPLETE PARATIME, THE [PARATIME POLICE]  H,LP
PIPER, H. BEAM  COSMIC COMPUTER, THE  C/P:2,P
PIPER, H. BEAM/MCGUIRE, JOHN J.  CRISIS IN 2140  DD227
PIPER, H. BEAM  EMPIRE  C/P:2,P
PIPER, H. BEAM  FEDERATION  C/P,LP,H
PIPER, H. BEAM/KURLAND, MICHAEL  FIRST CYCLE  C/P:2,P
PIPER, H. BEAM  FOUR-DAY PLANET  H
PIPER, H. BEAM  FOUR-DAY PLANET/LONE STAR PLANET  C/P:2,P
PIPER, H. BEAM  FUZZIES AND OTHER PEOPLE [FUZZIES 3]  H,P
PIPER, H. BEAM  FUZZY PAPERS, THE [FUZZIES 1,2]  H:2,P
PIPER, H. BEAM  FUZZY SAPIENS [FUZZIES 2]  P
PIPER, H. BEAM  GUNPOWDER GOD [PARATIME POLICE]  P
PIPER, H. BEAM  LITTLE FUZZY [FUZZIES 1]  C/P:2,P
PIPER, H. BEAM  LORD KALVAN OF OTHERWHEN [PARATIME POLICE]  C/P,P,H
PIPER, H. BEAM  OTHER HUMAN RACE, THE [FUZZIES 2]  C/P:2,P
PIPER, H. BEAM  PARATIME [PARATIME POLICE]  C/P:2,P
PIPER, H. BEAM/MCGUIRE, JOHN J.  PLANET FOR TEXANS, A  DD299
PIPER, H. BEAM  SPACE VIKING  C/P:2,P
PIPER, H. BEAM  ULLER UPRISING  C/P:2,P
PIPER, H. BEAM  WORLDS OF H. BEAM PIPER, THE  C/P,P
PIPER, WALTER  SPACE SWAP 1984–8612  SR-L
PISERCHIA, DORIS  BILLION DAYS OF EARTH, A  C/P:2,P
PISERCHIA, DORIS  DEADLY SKY, THE  P
PISERCHIA, DORIS  DIMENSIONEERS, THE  C/P,P
PISERCHIA, DORIS  DOOMTIME  P
PISERCHIA, DORIS  EARTH IN TWILIGHT  P
PISERCHIA, DORIS  EARTHCHILD  C/P,P
PISERCHIA, DORIS  FLUGER, THE  P
PISERCHIA, DORIS  MISTER JUSTICE  H,D53415
PISERCHIA, DORIS  SPACELING  H:2,P
PISERCHIA, DORIS  SPINNER, THE  H:2,P
PISERCHIA, DORIS  STAR RIDER  C/P:2,P
PITMAN, NORMAN H.  CHINESE FAIRY STORIES  H
PIZIKS, STEVEN  CORPORATE MENTALITY  C/P,P
PIZIKS, STEVEN  IN THE COMPANY OF MIND  C/P:2,P
PIZIKS, STEVEN  NANOTECH WAR, THE [STAR TREK VOYAGER]  P
PIZOR, FAITH K./COMP, T. ALLAN  MAN IN THE MOONE AND OTHER LUNAR FANTASIES, THE  HA
PLANCHAT, HENRY-LUC  DEDALE 2  FP
PLANTE, EDMUND  NEW NEIGHBORS, THE  P
PLATT, CHARLES  ASTEROIDE DEI PARIA, L  FP
PLATT, CHARLES  FREE ZONE  C/P,P
PLATT, CHARLES  GARBAGE WORLD  C/P:2,P
PLATT, CHARLES/BAILEY, HILARY  NEW WORLDS #6  LPA

AUTHOR  TITLE  CODE
PLATT, CHARLES  PLANET OF THE VOLES  H,P
PLATT, CHARLES  PLASM [WORLDS OF CHTHON]  C/P,P
PLATT, CHARLES  PROTEKTOR  C/P,P
PLATT, CHARLES  SILICON MAN, THE  C/P:2,P
PLATT, CHARLES  TWILIGHT OF THE CITY  H,P
PLATT, MARC  BATTLEFIELD [DOCTOR WHO]  C/P,P
PLATT, MARC  CAT’S CRADLE: TIME’S CRUCIBLE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
PLATT, MARC  DOWNTIME [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
PLATT, MARC  LUNGBARROW [DOCTOR WHO NEW ADVENTURES]  P
PLATT, RANDALL BETH  OUT OF A FOREST CLEARING  LP
PLAYBOY BOOKS  DEAD ASTRONAUT, THE  PA,C/PA
PLAYBOY BOOKS  FIEND, THE  C/PA,PA
PLAYBOY BOOKS  FROM THE "S" FILE  PA,C/PA
PLAYBOY BOOKS  FULLY AUTOMATED LOVE LIFE OF HENRY KEANRIDGE, THE  C/PA,PA
PLAYBOY BOOKS  LAST TRAIN TO LIMBO  C/PA,PA
PLAYBOY BOOKS  MASKS  C/PA:2,PA
PLAYBOY BOOKS  PLAYBOY BOOK OF HORROR AND THE SUPERNATURAL, THE  C/PA,PA
PLAYBOY BOOKS  PLAYBOY BOOK OF SCIENCE FICTION AND FANTASY, THE  PA,C/PA
PLAYBOY BOOKS  PLAYBOY’S STORIES OF THE SINISTER AND STRANGE  PA
PLAYBOY BOOKS  TRANSIT OF EARTH  PA,C/PA
PLAYBOY BOOKS  WEIRD SHOW  C/PA,PA
PLOWRIGHT, TERESA  DREAMS OF AN UNSEEN PLANET  H,SFWA
PLUMMER, CATHARINE  ALL ABOUT BROTHER BIRD  P

PLUMMER, RACHEL MYSTERY OF THE WIZARD’S TOMB

[@KNIGHTS OF THE SILVER DRAGON #11]

LP

POCKET BOOKS  AFTERMATH [STAR TREK S.C.E. 8]  LPA
POCKET BOOKS  EIGHT SPELLS A WEEK [SABRINA THE TEENAGE WITCH]  PA
POCKET BOOKS  GLASS EMPIRES [STAR TREK MIRROR]  LPA
POCKET BOOKS  OBSIDIAN ALLIANCES [STAR TREK MIRROR]  LPA
POCKET BOOKS  SAND AND STARS [STAR TREK]  LPA
POCKET BOOKS  TWIST OF FAITH [STAR TREK DS9]  LPA
POCKET PULSE  TALES OF THE SLAYER, VOL. 1 [BUFFY]  LP
PODRUG, JUNIUS  DARK PASSAGE  P
POE, EDGAR ALLAN  COMPLETE STORIES AND POEMS OF EDGAR ALLAN POE  H
POE, EDGAR ALLAN  GREAT TALES OF HORROR  P
POE, EDGAR ALLAN/VERNE, JULES  MYSTERY OF ARTHUR GORDON PYM, THE  H
POE, EDGAR ALLAN  PORTABLE EDGAR ALLAN POE, THE  P
POE, EDGAR ALLAN  SCIENCE FICTION OF EDGAR ALLAN POE, THE  C/P,P
POE, EDGAR ALLAN  TALES OF MYSTERY  P
POE, EDGAR ALLAN  WORKS OF EDGAR ALLAN POE 1, THE  SR-H
POE, EDGAR ALLAN  WORKS OF EDGAR ALLAN POE 2, THE  SR-H
POE, EDGAR ALLAN  WORKS OF EDGAR ALLAN POE 3, THE  SR-H
POE, EDGAR ALLAN  WORKS OF EDGAR ALLAN POE 4, THE  SR-H
POE, EDGAR ALLAN  WORKS OF EDGAR ALLAN POE 5, THE  SR-H
POGUE, CHARLES EDWARD  DRAGONHEART  P
POGUE, DAVID  HARD DRIVE  C/P,P
POGUE, WILLIAM R.  HOW DO YOU GO TO THE BATHROOM IN SPACE?  R/REF-P
POHL, CAROL/POHL, FREDERIK  JUPITER  C/PA,PA
POHL, CAROL/POHL, FREDERIK  SCIENCE FICTION DISCOVERIES  PA,C/PA
POHL, CAROL/POHL, FREDERIK  SCIENCE FICTION: THE GREAT YEARS  C/PA,PA
POHL, CAROL/POHL, FREDERIK  SCIENCE FICTION: THE GREAT YEARS II  C/PA,PA
POHL, FREDERIK  ABOMINABLE EARTHMAN, THE  C/P:2,P
POHL, FREDERIK  AGE OF THE PUSSYFOOT, THE  C/P,P,H
POHL, FREDERIK  ALTERNATING CURRENTS  C/P:2,P
POHL, FREDERIK  ANNALS OF THE HEECHEE, THE [HEECHEE 4]  H:2,P,SFWA
POHL, FREDERIK  ASSIGNMENT IN TOMORROW  HA,PA
POHL, FREDERIK/KORNBLUTH, C. M.  BEFORE THE UNIVERSE  C/P:2,P
POHL, FREDERIK  BEST OF FREDERIK POHL, THE  H:2,P
POHL, FREDERIK  BEST SCIENCE FICTION FOR 1972  PA,C/PA
POHL, FREDERIK  BEST SCIENCE FICTION FROM IF  SR-PA,C/PA
POHL, FREDERIK  BEST SCIENCE FICTION FROM WORLDS OF TOMORROW  SR-PA,C/PA
POHL, FREDERIK  BEYOND THE BLUE EVENT HORIZON [HEECHEE 2]  H:2,P,SFWA
POHL, FREDERIK  BEYOND THE END OF TIME  PA,C/PA
POHL, FREDERIK  BIPOHL  C/P:2,P
POHL, FREDERIK  BLACK STAR RISING  C/P,P,H
POHL, FREDERIK  BOY WHO WOULD LIVE FOREVER, THE [HEECHEE 6]  H
POHL, FREDERIK  CASE AGAINST TOMORROW, THE  C/P:2,P
POHL, FREDERIK  CHERNOBYL  C/P:2,P
POHL, FREDERIK  COMING OF THE QUANTUM CATS, THE  C/P,P,H
POHL, FREDERIK  COOL WAR, THE  C/P,P,H
POHL, FREDERIK/KORNBLUTH, C. M.  CRITICAL MASS  C/P:2,P
POHL, FREDERIK  DAY MILLION  C/P:2,P
POHL, FREDERIK  DAY THE MARTIANS CAME, THE  H:2,P
POHL, FREDERIK  DEMON IN THE SKULL  C/P:2,P
POHL, FREDERIK  DIGITS AND DASTARDS  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 304
AUTHOR  TITLE  CODE
POHL, FREDERIK  DOOR TO ANYWHERE  PA
POHL, FREDERIK  DRUNKARD’S WALK  C/P:2,P
POHL, FREDERIK  EARLY POHL, THE  H:2,SFWA
POHL, FREDERIK  EXPERT DREAMERS, THE  C/PA,PA
POHL, FREDERIK  FAR SHORE OF TIME, THE [ESCHATON 3]  C/P,P,H
POHL, FREDERIK/WILLIAMSON, JACK  FARTHEST STAR  C/P:2,P
POHL, FREDERIK  FINAL ENCOUNTER  C/PA,PA
POHL, FREDERIK  SEVENTH GALAXY READER, THE  HA
POHL, FREDERIK  NINTH GALAXY READER, THE  PA,C/PA
POHL, FREDERIK  ELEVENTH GALAXY READER, THE  HA

POHL, FREDERIK/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

GALAXY: THIRTY YEARS OF INNOVATIVE SCIENCE FICTION

SR-HA,HA

POHL, FREDERIK/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

GALAXY: VOLUME 2

PA

POHL, FREDERIK  GATEWAY [HEECHEE 1]  C/P,P,H
POHL, FREDERIK  GATEWAY TRIP, THE [HEECHEE 5]  C/P,P,H
POHL, FREDERIK/KORNBLUTH, C. M.  GLADIATOR-AT-LAW  C/P:2,P
POHL, FREDERIK  GOLD AT THE STARBOW’S END, THE  C/P,P,H

POHL, FREDERIK/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

GREAT SCIENCE FICTION SERIES, THE

HA

POHL, FREDERIK  HEECHEE RENDEZVOUS [HEECHEE 3]  C/P,P,H
POHL, FREDERIK  HOMEGOING  H:2,P
POHL, FREDERIK  IF READER OF SCIENCE FICTION, THE  C/PA,PA
POHL, FREDERIK  SECOND IF READER OF SCIENCE FICTION, THE  C/PA,PA
POHL, FREDERIK  IN THE PROBLEM PIT  P
POHL, FREDERIK  JEM  C/P,P,H
POHL, FREDERIK/WILLIAMSON, JACK  LAND’S END  H,SFWA-P,P
POHL, FREDERIK  MAN PLUS [MAN PLUS 1]  C/P,P,H,SFWA
POHL, FREDERIK  MAN WHO ATE THE WORLD, THE  C/P:2,P
POHL, FREDERIK/THOMAS, THOMAS T.  MARS PLUS [MAN PLUS 2]  H,P
POHL, FREDERIK  MERCHANTS’ WAR, THE [SPACE MERCHANTS 2]  C/P,P,H
POHL, FREDERIK  MIDAS WORLD  C/P,SFWA-P,H:2,P
POHL, FREDERIK  MINING THE OORT  C/P,P,H
POHL, FREDERIK  NARABEDLA, LTD.  C/P,P,H
POHL, FREDERIK  NEBULA WINNERS 14  C/PA,PA
POHL, FREDERIK  NIGHTMARE AGE  C/PA,PA
POHL, FREDERIK  O PIONEER!  H
POHL, FREDERIK  OTHER END OF TIME, THE [ESCHATON 1]  C/P:2,P
POHL, FREDERIK/KORNBLUTH, C. M.  OUR BEST  C/P,P
POHL, FREDERIK  OUTNUMBERING THE DEAD  H
POHL, FREDERIK  PLAGUE OF PYTHONS, A  C/P:2,P
POHL, FREDERIK  PLANETS THREE  C/P,P
POHL, FREDERIK  PLATINUM POHL  H:2
POHL, FREDERIK  POHLSTARS  C/P:2,P
POHL, FREDERIK/DEL REY, LESTER  PREFERRED RISK  C/P,P
POHL, FREDERIK/KORNBLUTH, C. M.  PRESIDENTIAL YEAR  P
POHL, FREDERIK/WILLIAMSON, JACK  REEFS OF SPACE, THE  C/P:2,P
POHL, FREDERIK/WILLIAMSON, JACK  ROGUE STAR  C/P:2,P
POHL, FREDERIK/WILLIAMSON, JACK  SAGA OF CUCKOO, THE  H:2

POHL, FREDERIK/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

SCIENCE FICTION OF THE 40’S

LPA

POHL, FREDERIK  SCIENCE FICTION ROLL OF HONOR, THE  HA,SFWA
POHL, FREDERIK/POHL, FREDERIK IV  SCIENCE FICTION STUDIES IN FILM  R/REF-H
POHL, FREDERIK/KORNBLUTH, C. M.  SEARCH THE SKY  C/P:2,P
POHL, FREDERIK/KORNBLUTH, C. M.  SEARCH THE SKY (REVISED)  C/P,P
POHL, FREDERIK  SFWA GRAND MASTERS 1, THE  LPA
POHL, FREDERIK  SFWA GRAND MASTERS 2, THE  HA,LPA
POHL, FREDERIK  SFWA GRAND MASTERS 3, THE  HA
POHL, FREDERIK  SHADOW OF TOMORROW  SR-PA,C/PA
POHL, FREDERIK  SIEGE OF ETERNITY, THE [ESCHATON 2]  C/P,P,H
POHL, FREDERIK/WILLIAMSON, JACK  SINGERS OF TIME, THE  C/P,LP,H
POHL, FREDERIK  SLAVE SHIP  C/P:2,P
POHL, FREDERIK/KORNBLUTH, C. M.  SPACE MERCHANTS, THE [SPACE MERCHANTS 1]  C/P:2,P
POHL, FREDERIK  STAR FOURTEEN  PA
POHL, FREDERIK  STAR OF STARS  C/PA,PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES  SR-PA,C/PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES 2  SR-PA,C/PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES 3  SR-PA,C/PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES 4  SR-PA,C/PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES 5  SR-PA,C/PA
POHL, FREDERIK  STAR SCIENCE FICTION STORIES 6  SR-PA,C/PA
POHL, FREDERIK  STAR SHORT NOVELS  SR-PA,C/PA
POHL, FREDERIK  STARBURST  C/P,P,H
POHL, FREDERIK/WILLIAMSON, JACK  STARCHILD  C/P,P,H
POHL, FREDERIK/WILLIAMSON, JACK  STARCHILD TRILOGY, THE  H:2,P

The MITSFS Pinkdex  September 24, 2007  Page 305
AUTHOR  TITLE  CODE
POHL, FREDERIK  STOPPING AT SLOWYEAR  C/P:2,P
POHL, FREDERIK  SYZYGY  C/P:2,P
POHL, FREDERIK/HULL, ELIZABETH ANNE  TALES FROM THE PLANET EARTH  SR-HA,C/PA,PA
POHL, FREDERIK  TERROR  C/P:2,P
POHL, FREDERIK  TOMORROW TIMES SEVEN  C/P:3,P
POHL, FREDERIK  TURN LEFT AT THURSDAY  C/P:2,P
POHL, FREDERIK/WILLIAMSON, JACK  UNDERSEA CITY [UNDERSEA 3]  H:2,P
POHL, FREDERIK/WILLIAMSON, JACK  UNDERSEA FLEET [UNDERSEA 2]  C/P:2,P
POHL, FREDERIK/WILLIAMSON, JACK  UNDERSEA QUEST [UNDERSEA 1]  C/P:2,P
POHL, FREDERIK/WILLIAMSON, JACK  UNDERSEA TRILOGY, THE [UNDERSEA 1,2,3]  SFWA-P
POHL, FREDERIK/KORNBLUTH, C. M.  VENUS, INC.  H:2
POHL, FREDERIK  VOICES OF HEAVEN, THE  H:2,P
POHL, FREDERIK/WILLIAMSON, JACK  WALL AROUND A STAR  C/P,P

POHL, FREDERIK WAY THE FUTURE WAS, THE R/REF-P,C/REF-H:2,

C/REF-P

POHL, FREDERIK/KORNBLUTH, C. M.  WOLFBANE  C/P:2,P
POHL, FREDERIK/KORNBLUTH, C. M.  WOLFBANE (REVISED)  P
POHL, FREDERIK/KORNBLUTH, C. M.  WOLFSKLAUW  FP
POHL, FREDERIK/KORNBLUTH, C. M.  WONDER EFFECT, THE  C/P:2,P
POHL, FREDERIK  WORLD AT THE END OF TIME, THE  H:2,LP

POHL, FREDERIK/GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

WORLDS OF IF

HA

POHL, FREDERIK  YEARS OF THE CITY, THE  C/P,P,H
POHL, FREDERIK  YESTERDAY’S TOMORROWS  HA,LPA
POHLE, ROBERT  DOOM OF THREE PLANETS  P
POHLMAN, EDWARD  GOD OF PLANET 607, THE  LP
POLIKARPUS, VIIDO/KING, TAPPAN  DOWN TOWN  P,SFWA,SFWA-P
POLIVKA, GEORGE BRYAN  LEGEND OF THE FIREFISH, THE [TROPHY CHASE 1]  LP
POLLACK, JACK HARRISON  CROISET THE CLAIRVOYANT  [C/RAN-P]
POLLACK, RACHEL  ALQUA DREAMS  H
POLLACK, RACHEL  GOLDEN VANITY  C/P,P
POLLACK, RACHEL/MATTHEWS, CAITLIN  TAROT TALES  PA
POLLACK, RACHEL  TEMPORARY AGENCY [LIVING WORLD 2]  H:2,LP
POLLACK, RACHEL  UNQUENCHABLE FIRE [LIVING WORLD 1]  H,LP
POLLOCK, DALE  SKYWALKING  R/REF-P,C/REF-P
POLLOTTA, NICK  BUREAU 13 [BUREAU 13 1]  P
POLLOTTA, NICK  DAMNED NATION [@BUREAU 13]  LP
POLLOTTA, NICK  DOOMSDAY EXAM [BUREAU 13 2]  C/P,P
POLLOTTA, NICK  FULL MOONSTER [BUREAU 13 3]  P
POLLOTTA, NICK/FOGLIO, PHIL  ILLEGAL ALIENS  C/P:2,P
POLLOTTA, NICK  JUDGMENT NIGHT [@BUREAU 13]  LP
POLLOTTA, NICK  SHADOWBOXER [SHADOWRUN 25]  P
POLLOTTA, NICK  TEQUILA MOCKINGBIRD  LP
POLLOTTA, NICK/CLAY, JAMES  THAT DARN SQUID GOD  H
POMEGRANATE  SPACE (1999)  CAL
POPE, ELIZABETH MARIE  PERILOUS GARD, THE  C/P,P
POPE, ELIZABETH MARIE  SHERWOOD RING, THE  P
POPE, GUSTAVUS W.  JOURNEY TO MARS  LP
POPE, NICK  OPERATION LIGHTNING STRIKE  P
POPESCU, PETRU  IN HOT BLOOD  P
POPKES, STEVEN  CALIBAN LANDING  H,P

POPKES, STEVEN SLOW LIGHTNING TD51170,

SFWA-TD51170

POPP, ROBIN T.  DARKENING, THE [@IMMORTALS]  P
POPULAR LIBRARY  DR. CYCLOPS  PA
PORATH, ELLEN  STEEL AND STONE [@DRAGONLANCE MEETINGS 5]  C/P,P
PORGES, IRWIN  EDGAR RICE BURROUGHS: THE MAN WHO CREATED TARZAN  R/REF-H
PORTER, ANDREW  BOOK OF ELLISON, THE  R/REF-P
PORTNOY, HOWARD N.  HOT RAIN  P
POSSONY, STEFAN T./POURNELLE, JERRY  STRATEGY OF TECHNOLOGY, THE  R/REF-H
POST, J. B.  ATLAS OF FANTASY, AN  A
POTTER, ROBERT R.  BEYOND TIME AND SPACE  LP
POTTS, CHARLES  LOADING LAS VEGAS  LP
POUNS, BRAUNA E.  AMERIKA  P
POURNELLE, JERRY  ARMAGEDDON! [@THERE WILL BE WAR #8]  PA,SFWA-P
POURNELLE, JERRY  BIRTH OF FIRE  C/P:2,P
POURNELLE, JERRY  BLACK HOLES  PA,C/PA
POURNELLE, JERRY  BLOOD AND IRON [@THERE WILL BE WAR #3]  C/PA,PA
POURNELLE, JERRY  BURNING EYE, THE [@WAR WORLD #1]  PA,C/PA,SFWA-P
POURNELLE, JERRY  CALL TO BATTLE! [@THERE WILL BE WAR #7]  PA,SFWA-P
POURNELLE, JERRY/STIRLING, S. M.  CHILDREN’S HOUR, THE [MAN-KZIN WARS]  C/P:2,SFWA-P,P
POURNELLE, JERRY/CARR, JOHN F.  CITIES IN SPACE [ENDLESS FRONTIER 3]  PA
POURNELLE, JERRY/GREEN, ROLAND  CLAN AND CROWN [@JANISSARIES 2]  C/P,LP,H
POURNELLE, JERRY  CODOMINIUM: REVOLT ON WAR WORLD  C/PA,PA,SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 306
AUTHOR  TITLE  CODE
POURNELLE, JERRY  CRASH OF EMPIRE, THE [@IMPERIAL STARS #3]  SFWA-P
POURNELLE, JERRY  DAY OF THE TYRANT [@THERE WILL BE WAR #4]  PA,C/PA,SFWA-P
POURNELLE, JERRY  DEATH’S HEAD REBELLION [@WAR WORLD #2]  PA,C/PA
POURNELLE, JERRY  ENDLESS FRONTIER, THE  PA
POURNELLE, JERRY/CARR, JOHN F.  ENDLESS FRONTIER II, THE  C/PA,PA
POURNELLE, JERRY  ESCAPE FROM THE PLANET OF THE APES  C/P:2,P
POURNELLE, JERRY  EXILES TO GLORY  C/P:2,SFWA-P,P
POURNELLE, JERRY  EXILES TO GLORY (REVISED)  P
POURNELLE, JERRY  FALKENBERG’S LEGION [FALKENBERG’S LEGION]  C/P,P
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS  PA,C/PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS II  PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS III  PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS IV  PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS V  PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS VI  PA
POURNELLE, JERRY/BAEN, JIM  FAR FRONTIERS VII  PA
POURNELLE, JERRY/STIRLING, S. M.  GO TELL THE SPARTANS [FALKENBERG’S LEGION]  C/P:2,SFWA-P,P
POURNELLE, JERRY  GUNS OF DARKNESS [@THERE WILL BE WAR #6]  PA,C/PA,SFWA-P
POURNELLE, JERRY  HIGH JUSTICE  C/P:2,P

POURNELLE, JERRY/STIRLING, S. M./ ING, DEAN HOUSES OF THE KZINTI [MAN-KZIN WARS]

LP

POURNELLE, JERRY  INVASION [@WAR WORLD #4]  PA
POURNELLE, JERRY  JANISSARIES [@JANISSARIES 1]  C/P,LP,H
POURNELLE, JERRY  KING DAVID’S SPACESHIP  C/P,P,H
POURNELLE, JERRY  MEN OF WAR [@THERE WILL BE WAR #2]  C/PA,PA
POURNELLE, JERRY  MERCENARY, THE  C/P:2,P
POURNELLE, JERRY  NEBULA AWARD STORIES 16  C/PA,PA
POURNELLE, JERRY  PRINCE OF MERCENARIES [FALKENBERG’S LEGION]  C/P:2,SFWA-P,P
POURNELLE, JERRY/STIRLING, S. M.  PRINCE OF SPARTA [FALKENBERG’S LEGION]  C/P:2,SFWA-P,P
POURNELLE, JERRY  RED HEROIN  P
POURNELLE, JERRY  REPUBLIC AND EMPIRE [@IMPERIAL STARS #2]  C/PA,PA,SFWA-P
POURNELLE, JERRY  SAURON DOMINION [@WAR WORLD #3]  SFWA-P,C/PA,PA
POURNELLE, JERRY  SCIENCE FICTION YEARBOOK, THE  HA
POURNELLE, JERRY  SPACESHIP FOR THE KING, A  C/P,P
POURNELLE, JERRY  STARS AT WAR, THE [@IMPERIAL STARS #1]  C/PA,PA
POURNELLE, JERRY  STARSWARM  H,P
POURNELLE, JERRY  STEP FARTHER OUT, A  [C/RAN-P:2]
POURNELLE, JERRY/GREEN, ROLAND  STORMS OF VICTORY [@JANISSARIES 3]  C/P,SFWA,P,H
POURNELLE, JERRY/CARR, TERRY  SURVIVAL OF FREEDOM, THE  C/PA,PA
POURNELLE, JERRY  THERE WILL BE WAR [@THERE WILL BE WAR #1]  PA,C/PA
POURNELLE, JERRY/GREEN, ROLAND  TRAN [JANISSARIES]  P
POURNELLE, JERRY  2020 VISION  PA,C/PA
POURNELLE, JERRY  WARRIOR! [@THERE WILL BE WAR #5]  C/PA,PA,SFWA-P
POURNELLE, JERRY  WEST OF HONOR  P
POURNELLE, JERRY  WEST OF HONOR (REVISED)  C/P:2,P
POWAGA, WIESIEK  DEDALUS BOOK OF POLISH FANTASY, THE  LPA
POWE, BRUCE  KILLING GROUND  P
POWELL, F. INGLIS  SNAKE, THE  H
POWERS, J. L.  BLACK ABYSS  SR-P
POWERS, RICHARD  GALATEA 2.2  LP
POWERS, ROBERT M.  SHUTTLE  [C/RAN-P]
POWERS, TIM  ANUBIS GATES, THE  C/P:2,P
POWERS, TIM  DECLARE  H,P
POWERS, TIM  DINNER AT DEVIANT’S PALACE  C/P,P,H
POWERS, TIM  DRAWING OF THE DARK, THE  P
POWERS, TIM  EARTHQUAKE WEATHER [FAULT LINES 2]  P
POWERS, TIM  EPITAPH IN RUST  SR-P
POWERS, TIM  EPITAPH IN RUST, AN  SR-H
POWERS, TIM  EXPIRATION DATE [FAULT LINES 1]  H,P
POWERS, TIM  FAULT LINES [FAULT LINES 1,2]  H
POWERS, TIM  FORSAKE THE SKY  C/P:2,P
POWERS, TIM  LAST CALL  H:2,P
POWERS, TIM  LAST CALL (PROOF)  SR-P
POWERS, TIM  ON STRANGER TIDES  H:2,SFWA-P,P
POWERS, TIM  POWERS OF TWO  H:2
POWERS, TIM  SKIES DISCROWNED, THE  SR-P
POWERS, TIM  STRANGE ITINERARIES  LP
POWERS, TIM  STRESS OF HER REGARD, THE  H,SFWA,SFWA-P
POWERS, TIM  THREE DAYS TO NEVER  H
POYER, D. C.  STEPFATHER BANK  H:2,P
POYER, JOE  CHINESE AGENDA, THE  H
POYER, JOE  NORTH CAPE  H
POYER, JOE  OPERATION MALACCA  H,P
POYER, JOE  SHOOTING OF THE GREEN, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 307
AUTHOR  TITLE  CODE
PRABHUPADA, SWAMI  PERFECTION OF YOGA, THE  [C/RAN-P]
PRAGER, J. SIMON  NEWMAN FACTOR, THE  P
PRAGNELL, FESTUS  GREEN MAN OF GRAYPEC, THE  H
PRATCHETT, TERRY  AMAZING MAURICE AND HIS EDUCATED RODENTS, THE [DISCWORLD 27]  H
PRATCHETT, TERRY  BROMELIAD TRILOGY, THE [BROMELIAD 1,2,3]  H
PRATCHETT, TERRY  CARPE JUGULUM [DISCWORLD 23/DISCWORLD WITCHES 6]  H:2,P
PRATCHETT, TERRY  COLOUR OF MAGIC, THE [DISCWORLD 1/DISCWORLD RINCEWIND 1]  C/P,P,H:2
PRATCHETT, TERRY  DARK SIDE OF THE SUN, THE  H

PRATCHETT, TERRY/STEWART, IAN/ COHEN, JACK DARWIN’S WATCH [@SCIENCE OF DISCWORLD #3/DISCWORLD]

H,P

PRATCHETT, TERRY/BRIGGS, STEPHEN  DISCWORLD COMPANION, THE [DISCWORLD]  R/REF-P
PRATCHETT, TERRY  EQUAL RITES [DISCWORLD 3/DISCWORLD WITCHES 1]  C/P:2,P,H
PRATCHETT, TERRY  ERIC [DISCWORLD 9/DISCWORLD RINCEWIND 4]  C/P:3,SFWA-P,VLP
PRATCHETT, TERRY  FEET OF CLAY [DISCWORLD 19/DISCWORLD WATCHMEN 3]  H:2,P
PRATCHETT, TERRY  FIFTH ELEPHANT, THE [DISCWORLD 24/DISCWORLD WATCHMEN 5]  H:2,P

PRATCHETT, TERRY/STEWART, IAN/ COHEN, JACK GLOBE, THE [@SCIENCE OF DISCWORLD #2]

P

PRATCHETT, TERRY  GOING POSTAL [DISCWORLD 32]  C/P,P,H
PRATCHETT, TERRY  GUARDS! GUARDS! [DISCWORLD 8/DISCWORLD WATCHMEN 1]  H:2,P
PRATCHETT, TERRY  HAT FULL OF SKY, A [DISCWORLD 31/DISCWORLD TIFFANY ACHING 2]  H
PRATCHETT, TERRY  HOGFATHER [DISCWORLD 20/DISCWORLD DEATH 4]  H,P
PRATCHETT, TERRY  INTERESTING TIMES [DISCWORLD 17/DISCWORLD RINCEWIND 5]  C/P,P,H
PRATCHETT, TERRY  JINGO [DISCWORLD 21/DISCWORLD WATCHMEN 4]  H:2,P
PRATCHETT, TERRY  JOHNNY AND THE DEAD [JOHNNY MAXWELL 2]  H
PRATCHETT, TERRY  JOHNNY MAXWELL TRILOGY, THE [JOHNNY MAXWELL 1,2,3]  H
PRATCHETT, TERRY  LAST CONTINENT, THE [DISCWORLD 22/DISCWORLD RINCEWIND 6]  C/P,P
PRATCHETT, TERRY  LAST HERO, THE [DISCWORLD]  VLP
PRATCHETT, TERRY  LIGHT FANTASTIC, THE [DISCWORLD 2/DISCWORLD RINCEWIND 2]  C/P:2,P,H
PRATCHETT, TERRY  LORDS AND LADIES [DISCWORLD 14/DISCWORLD WITCHES 4]  C/P:2,LP,P
PRATCHETT, TERRY  LORDS AND LADIES (PROOF) [DISCWORLD 14/DISCWORLD WITCHES 4]  SR-LP
PRATCHETT, TERRY  MASKERADE [DISCWORLD 18/DISCWORLD WITCHES 5]  C/P,P,H
PRATCHETT, TERRY  MEN AT ARMS [DISCWORLD 15/DISCWORLD WATCHMEN 2]  C/P,P,H
PRATCHETT, TERRY  MONSTROUS REGIMENT [DISCWORLD 30]  H:2,P
PRATCHETT, TERRY  MORT [DISCWORLD 4/DISCWORLD DEATH 1]  C/P,H,P
PRATCHETT, TERRY  MOVING PICTURES [DISCWORLD 10]  C/P:3,P,H
PRATCHETT, TERRY  NIGHT WATCH [DISCWORLD 28/DISCWORLD WATCHMEN 6]  H:2
PRATCHETT, TERRY  ONLY YOU CAN SAVE MANKIND [JOHNNY MAXWELL 1]  H
PRATCHETT, TERRY  PYRAMIDS [DISCWORLD 7]  H:2,P
PRATCHETT, TERRY  REAPER MAN [DISCWORLD 11/DISCWORLD DEATH 2]  H:2,SFWA-P,P

PRATCHETT, TERRY RINCEWIND THE WIZARD [DISCWORLD 1,2,5,9/

DISCWORLD RINCEWIND 1,2,3,4]

H

PRATCHETT, TERRY/STEWART, IAN/ COHEN, JACK SCIENCE OF DISCWORLD, THE (REVISED) [SCIENCE OF DISCWORLD 1/

DISCWORLD]

P

PRATCHETT, TERRY  SMALL GODS [DISCWORLD 13]  H:2,P
PRATCHETT, TERRY  SOUL MUSIC [DISCWORLD 16/DISCWORLD DEATH 3]  C/P:2,P
PRATCHETT, TERRY  SOURCERY [DISCWORLD 5/DISCWORLD RINCEWIND 3]  C/P,P,H
PRATCHETT, TERRY  STRATA  H:2
PRATCHETT, TERRY  THIEF OF TIME [DISCWORLD 26]  H:2,P
PRATCHETT, TERRY  THUD! [DISCWORLD 33/DISCWORLD WATCHMEN 7]  H,P
PRATCHETT, TERRY  TRUTH, THE [DISCWORLD 25]  H:2,P
PRATCHETT, TERRY  WEE FREE MEN, THE [DISCWORLD 29/DISCWORLD TIFFANY ACHING 1]  H,P
PRATCHETT, TERRY  WHERE’S MY COW? [DISCWORLD]  H
PRATCHETT, TERRY  WINTERSMITH [DISCWORLD TIFFANY ACHING 3]  H:2
PRATCHETT, TERRY  WITCHES ABROAD [DISCWORLD 12/DISCWORLD WITCHES 3]  C/P,SFWA-P,P
PRATCHETT, TERRY  WYRD SISTERS [DISCWORLD 6/DISCWORLD WITCHES 2]  C/P,P,H
PRATHER, RICHARD S.  AMBER EFFECT, THE  SFWA-P
PRATHER, RICHARD S.  CHEIM MANUSCRIPT, THE  SFWA-P
PRATHER, RICHARD S.  KUBLA KHAN CAPER, THE  SFWA-P
PRATHER, RICHARD S.  SWEET RIDE, THE  SFWA-P
PRATT, FLETCHER  ALIEN PLANET  C/P,P,H
PRATT, FLETCHER  ALL ABOUT ROCKETS AND JETS  [C/RAN-H]
PRATT, FLETCHER  BLUE STAR, THE  C/P:2,P
PRATT, FLETCHER  DOUBLE IN SPACE  H:2
PRATT, FLETCHER  DOUBLE JEOPARDY  C/P,H,GN5.30
PRATT, FLETCHER  INVADERS FROM RIGEL  C/P:2,P
PRATT, FLETCHER/DE CAMP, L. SPRAGUE  LAND OF UNREASON  C/P:2,P
PRATT, FLETCHER  UNDYING FIRE, THE  C/P,P,H
PRATT, FLETCHER  WELL OF THE UNICORN, THE  C/P:2,P
PRATT, FLETCHER  WORLD OF WONDER  HA
PRATT, THEODORE  MR. LIMPET  H
PRATT, TIM  LITTLE GODS  LP
PRATT, TIM  STRANGE ADVENTURES OF RANGERGIRL, THE  LP
PREHODA, ROBERT W.  YOUR NEXT FIFTY YEARS  R/REF-H,C/REF-P
PREISLER, JEROME  ANNIHILATION [MORTAL KOMBAT 2]  P

AUTHOR  TITLE  CODE
PREISS, BYRON/REAVES, J. MICHAEL  DRAGONWORLD  H:2,P
PREISS, BYRON  GUTS  C/P:2,P
PREISS, BYRON  MICROVERSE, THE  SR-VLHA
PREISS, BYRON  PLANETS, THE  SR-VLHA
PREISS, BYRON/FABIAN, STEPHEN  STARFAWN  C/P:2,SCX
PREISS, BYRON  ULTIMATE ALIEN, THE  LPA
PREISS, BYRON/SILVERBERG, ROBERT  ULTIMATE DINOSAUR, THE  SR-VLHA
PREISS, BYRON  ULTIMATE DRACULA, THE  HA,LPA
PREISS, BYRON  ULTIMATE DRAGON, THE  LPA
PREISS, BYRON  ULTIMATE DRAGON, THE (PROOF)  SR-LPA
PREISS, BYRON  ULTIMATE FRANKENSTEIN, THE  HA,LPA
PREISS, BYRON  ULTIMATE WEREWOLF, THE  HA,LPA
PREISS, BYRON  UNIVERSE, THE  SR-VLHA
PREISS, BYRON  WEIRD HEROES 1  C/PA,PA
PREISS, BYRON  WEIRD HEROES 2  PA
PREISS, BYRON  WEIRD HEROES 6  PA
PREISS, BYRON  WEIRD HEROES 8  C/PA,PA

PRESIDENTIAL COMMISSION REPORT OF THE PRESIDENTIAL COMMISSION ON THE SPACE SHUTTLE

CHALLENGER ACCIDENT

R/REF-H

PRESTON, DOUGLAS/CHILD, LINCOLN  RELIC [RELIC 1]  P
PRESTON, DOUGLAS/CHILD, LINCOLN  RELIQUARY [RELIC 2]  H
PRESTON, RICHARD  HOT ZONE, THE  R/REF-P
PREUSS, PAUL  BREAKING STRAIN [@VENUS PRIME #1]  C/P:2,P
PREUSS, PAUL  BROKEN SYMMETRIES  C/P,P,H
PREUSS, PAUL  BROKEN SYMMETRIES (PROOF)  SR-LP
PREUSS, PAUL  CORE  C/P:2,P
PREUSS, PAUL  DIAMOND MOON, THE [@VENUS PRIME #5]  C/P:2,P
PREUSS, PAUL  GATES OF HEAVEN, THE  C/P:2,P
PREUSS, PAUL  HIDE AND SEEK [@VENUS PRIME #3]  C/P:2,P
PREUSS, PAUL  HUMAN ERROR  H:2
PREUSS, PAUL  MAELSTROM [@VENUS PRIME #2]  C/P:2,P
PREUSS, PAUL  MEDUSA ENCOUNTER, THE [@VENUS PRIME #4]  C/P:3,P
PREUSS, PAUL  RE-ENTRY  C/P:2,P
PREUSS, PAUL  SECRET PASSAGES  H,P
PREUSS, PAUL  SHINING ONES, THE [@VENUS PRIME #6]  C/P,P,H
PREUSS, PAUL  STARFIRE  C/P:2,SFWA-P,P,H
PREUSSLER, OTFRIED  SATANIC MILL, THE  SFWA-P
PRICE, E. HOFFMANN  DEVIL WIVES OF LI FONG, THE  C/P:2,P
PRICE, E. HOFFMANN  JADE ENCHANTRESS, THE  C/P:2,P
PRICE, E. HOFFMANN  OPERATION EXILE  P
PRICE, E. HOFFMANN  OPERATION ISIS  P
PRICE, E. HOFFMANN  OPERATION LONGLIFE  C/P:2,P
PRICE, E. HOFFMANN  OPERATION MISFIT  C/P:2,P
PRICE, E. HOFFMANN  STRANGE GATEWAYS  SR-H
PRICE, JOHN-ALLEN  FRENZY [@MUTANT CHRONICLES APOSTLE OF INSANITY 2]  P
PRICE, JOHN-ALLEN  PURSUIT OF THE PHOENIX, THE  P
PRICE, JORDAN CASTILLO  PSYCOP PARTNERS  LP
PRICE, ROBERT M.  ANTARKTOS CYCLE, THE  LP
PRICE, ROBERT M.  BOOK OF EIBON, THE  LPA
PRICE, ROBERT M.  HASTUR CYCLE, THE (2ND EDITION)  LPA
PRICE, ROBERT M.  LIN CARTER’S ANTON ZARNAK SUPERNATURAL SLEUTH  LPA
PRICE, ROBERT M.  NEW LOVECRAFT CIRCLE, THE  LPA
PRICE, ROBERT M.  NYARLATHOTEP CYCLE, THE  LPA
PRICE, ROBERT M.  TALES OF THE LOVECRAFT MYTHOS  HA
PRICE, ROBERT M.  TALES OUT OF DUNWICH  LPA
PRICE, ROBERT M.  TALES OUT OF INNSMOUTH  LPA
PRICE, ROBERT M.  TSATHOGGUA CYCLE, THE  LPA
PRICE, SUSAN  STERKARM HANDSHAKE, THE  H
PRIEST, CHERIE  FOUR AND TWENTY BLACKBIRDS [EDEN MOORE]  LP
PRIEST, CHERIE  WINGS TO THE KINGDOM [EDEN MOORE]  LP
PRIEST, CHRISTOPHER  AFFIRMATION, THE  SFWA
PRIEST, CHRISTOPHER  ANTICIPATIONS  C/PA,SFWA
PRIEST, CHRISTOPHER  DARKENING ISLAND  H,P
PRIEST, CHRISTOPHER  DREAM OF WESSEX, A  P
PRIEST, CHRISTOPHER  GLAMOUR, THE  H:2,SFWA
PRIEST, CHRISTOPHER  INDOCTRINAIRE  C/P,P,H
PRIEST, CHRISTOPHER  INFINITE SUMMER, AN  C/P,P,H
PRIEST, CHRISTOPHER  INVERTED WORLD, THE  H:2,SFWA,P
PRIEST, CHRISTOPHER  INVERTED WORLD/FUGUE FOR A DARKENING ISLAND  P
PRIEST, CHRISTOPHER  PERFECT LOVER, THE  C/P,P,H
PRIEST, CHRISTOPHER  PRESTIGE, THE  H,LP
PRIEST, CHRISTOPHER  SPACE MACHINE, THE  H,SFWA,P
PRIEST, CHRISTOPHER  SPACE MACHINE, THE/DREAM OF WESSEX, A  P
PRIESTLEY, J. B.  DOOMSDAY MEN, THE  H,P
PRIESTLEY, J. B.  OTHER PLACE, THE  H

AUTHOR  TITLE  CODE
PRIESTLEY, J. B.  THIRTY-FIRST OF JUNE, THE  H
PRIESTLEY, J. B.  THREE TIME-PLAYS  SR-P
PRIESTLEY, LEE  ROCKET MOUSE  P
PRINGLE, DAVID  ANT-MEN OF TIBET AND OTHER STORIES, THE  LPA
PRINGLE, DAVID  BEST OF INTERZONE, THE  HA

PRINGLE, DAVID EARTH IS THE ALIEN PLANET:

J. G. BALLARD’S FOUR-DIMENSIONAL NIGHTMARE

R/REF-P

PRINGLE, DAVID  IGNORANT ARMIES [WARHAMMER]  SR-LP
PRINGLE, DAVID  LAUGHTER OF DARK GODS, THE [WARHAMMER]  PA
PRINGLE, DAVID  MODERN FANTASY  R/REF-P
PRINGLE, DAVID  RED THIRST [WARHAMMER]  LP
PRINGLE, DAVID  ROUTE 666 [DARK FUTURE]  LP
PRINGLE, DAVID  ULTIMATE ENCYCLOPEDIA OF SCIENCE FICTION, THE  R/REF-H
PRINGLE, DAVID  ULTIMATE GUIDE TO SCIENCE FICTION, THE  SFWA
PRINGLE, DAVID  WOLF RIDERS [WARHAMMER]  SR-LP
PRINGLE, ERIC  AWAKENING, THE [DOCTOR WHO]  SR-P,C/P
PROCTOR, GEO. W.  CHICAGO CONVERSION, THE [V]  C/P,P
PROCTOR, GEO. W.  ESPER TRANSFER, THE  P
PROCTOR, GEO. W.  FIRE AT THE CENTER  C/P,P
PROCTOR, GEO. W./UTLEY, STEVEN  LONE STAR UNIVERSE  SR-HA,HA
PROCTOR, GEO. W.  SHADOWMAN  P
PROCTOR, GEO. W.  STARWINGS  P
PROCTOR, GEO. W.  STELLAR FIST  P
PROCTOR, GEO. W.  TEXAS RUN, THE [V]  C/P,P
PRONZINI, BILL/MALZBERG, BARRY N.  BUG-EYED MONSTERS  HA,LP
PRONZINI, BILL  CREATURE!  LP
PRONZINI, BILL  MASQUES  P
PRONZINI, BILL/MALZBERG, BARRY N.  PROSE BOWL  H
PRONZINI, BILL/MALZBERG, BARRY N.  SHARED TOMORROWS: SCIENCE FICTION IN COLLABORATION  HA
PRONZINI, BILL  WEREWOLF!  HA,PA
PROSPERI, PIERO  AUTOCRISI  FP
PRUCHER, JEFF  BRAVE NEW WORDS  R/REF-H
PTACEK, KATHRYN  BLOOD AUTUMN  SFWA-P
PTACEK, KATHRYN  GHOST DANCE  SFWA-P
PTACEK, KATHRYN  HUNTED, THE  SFWA
PTACEK, KATHRYN  IN SILENCE SEALED  SFWA-P
PTACEK, KATHRYN  SHADOWEYES  P
PTACEK, KATHRYN  WOMEN OF DARKNESS  HA,PA,SFWA-P
PTACEK, KATHRYN  WOMEN OF DARKNESS II  SR-HA
PUCCETTI, ROLAND  DEATH OF THE FUHRER, THE  P
PUDNEY, JOHN  EDNA’S FRUIT HAT AND OTHER STORIES  H
PULLMAN, PHILIP  AMBER SPYGLASS, THE [HIS DARK MATERIALS 3]  C/P,P,H

PULLMAN, PHILIP AMBER SPYGLASS, THE (LIMITED EDITION) (CD)

[HIS DARK MATERIALS 1,2,3]

SR-MM

PULLMAN, PHILIP  GOLDEN COMPASS, THE [HIS DARK MATERIALS 1]  C/P:2,P,H
PULLMAN, PHILIP  SUBTLE KNIFE, THE [HIS DARK MATERIALS 2]  C/P,P,H
PULVER, JOSEPH S., SR.  NIGHTMARE’S DISCIPLE  LP
PUNTER, DAVID  LITERATURE OF TERROR, THE  R/REF-H
PURDOM, TOM  BARONS OF BEHAVIOR, THE  C/P:2,P
PURDOM, TOM  FIVE AGAINST ARLANE  DH022
PURDOM, TOM  I WANT THE STARS  DF289
PURDOM, TOM  REDUCTION IN ARMS  C/P:2,P
PURDOM, TOM  TREE LORD OF IMETEN, THE  DM139
PURSER-HALLARD, PHILIP  OF THE CITY OF THE SAVED... [@FACTION PARADOX 2]  LP
PURSER-HALLARD, PHILIP  PECULIAR LIVES [@TIME HUNTER 7]  LP
PURTILL, RICHARD  GOLDEN GRYPHON FEATHER, THE [KAPHTU 1]  C/P:2,P
PURTILL, RICHARD  MIRROR OF HELEN, THE [KAPHTU 3]  P
PURTILL, RICHARD  MURDERCON  SFWA
PURTILL, RICHARD  PARALLEL MAN, THE  C/P:2,P
PURTILL, RICHARD  STOLEN GODDESS, THE [KAPHTU 2]  C/P:2,P
PUTNEY, SUSAN K.  AGAINST ARCTURUS  D00990
PYGASTER, CAL I.  ZERO GRAVITY SWAP  SR-L
PYKE, MAGNUS  UITERSTE MOGELIJKHEDEN VAN DE NATUURWETENSCHAP, DE  FP
PYNCHON, THOMAS  CRYING OF LOT 49, THE  C/P:2,P
PYNCHON, THOMAS  GRAVITY’S RAINBOW  C/P:2,LP

AUTHOR  TITLE  CODE
QUARK  LEGENDS OF THE FERENGI [STAR TREK DS9]  LP
QUICK, W. T.  COLONY [@PLANET OF THE APES]  P
QUICK, W. T.  DREAMS OF FLESH AND SAND  C/P,P
QUICK, W. T.  DREAMS OF GODS AND MEN  P
QUICK, W. T.  FALL, THE [@PLANET OF THE APES]  P
QUICK, W. T.  PLANET OF THE APES  H
QUICK, W. T.  SINGULARITIES  P
QUICK, W. T.  STAR CONTROL INTERBELLUM [@STAR CONTROL]  P
QUICK, W. T.  SYSTEMS  C/P:2,P
QUICK, W. T.  YESTERDAY’S PAWN  P
QUINN, DANIEL  DREAMER  SFWA-P
QUINN, DANIEL  ISHMAEL  H
QUINN, JAMES/WULFF, EVE  FIRST WORLD OF IF, THE  SR-PA
QUINN, JAMES/WULFF, EVE  SECOND WORLD OF IF, THE  SR-PA
QUINN, SEABURY  ADVENTURES OF JULES DE GRANDIN, THE  C/P,P
QUINN, SEABURY  ALIEN FLESH  SR-H
QUINN, SEABURY  CASEBOOK OF JULES DE GRANDIN, THE  P
QUINN, SEABURY  DEVIL’S BRIDE, THE  P
QUINN, SEABURY  HELLFIRE FILES OF JULES DE GRANDIN, THE  P
QUINN, SEABURY  HORROR CHAMBERS OF JULES DE GRANDIN, THE  P
QUINN, SEABURY  SKELETON CLOSET OF JULES DE GRANDIN, THE  C/P:2,P

AUTHOR  TITLE  CODE
RAABE  BIBLIOHOLISM  [C/RAN-P]
RABE, JEAN/HELFERS, JOHN  AFTERSHOCK [@SHADOWRUN]  P
RABE, JEAN  BETRAYAL [@DRAGONLANCE DHAMON #2]  C/P,P
RABE, JEAN  DAWNING OF A NEW AGE, THE [@DRAGONLANCE FIFTH AGE 1]  P
RABE, JEAN  DAY OF THE TEMPEST, THE [@DRAGONLANCE FIFTH AGE]  P
RABE, JEAN  DOWNFALL [@DRAGONLANCE DHAMON #1]  H,P
RABE, JEAN  EVE OF THE MAELSTROM, THE [@DRAGONLANCE FIFTH AGE]  P
RABE, JEAN  FINEST CHOICE, THE  H
RABE, JEAN  FINEST CREATION, THE  H
RABE, JEAN/THOMSEN, BRIAN M.  FURRY FANTASTIC  PA
RABE, JEAN  LAKE OF DEATH, THE [@DRAGONLANCE AGE OF MORTALS]  P
RABE, JEAN  RED MAGIC [@FORGOTTEN REALMS HARPERS #3]  P
RABE, JEAN  REDEMPTION [@DRAGONLANCE DHAMON #3]  H
RABE, JEAN  SILVER STAIR, THE [@DRAGONLANCE BRIDGES OF TIME]  P
RABE, JEAN/GREENBERG, MARTIN H.  SOL’S CHILDREN  PA
RABE, JEAN/GREENBERG, MARTIN H.  TIME TWISTERS  PA
RABELAIS  GARGANTUA AND PANTAGRUEL  P
RABEN, RICHARD/COHEN, HIYAGUHA  BOLDLY LIVE AS YOU’VE NEVER LIVED BEFORE  R/REF-H
RABKIN, ERIC S.  ARTHUR C. CLARKE  R/REF-P
RABKIN, ERIC S.  FANTASTIC WORLDS  LPA
RABKIN, ERIC S.  SCIENCE FICTION: A HISTORICAL ANTHOLOGY  LPA
RACINA, THOM  GREAT LOS ANGELES BLIZZARD, THE  P
RACKHAM, JOHN  ALIEN SEA  DH040
RACKHAM, JOHN  ANYTHING TREE, THE  D89250
RACKHAM, JOHN  BEANSTALK  C/P:2,P
RACKHAM, JOHN  BEASTS OF KOHL, THE  DG592
RACKHAM, JOHN  BEYOND CAPELLA  D05595
RACKHAM, JOHN  DANGER FROM VEGA  DG576
RACKHAM, JOHN  DARK PLANET  D13805
RACKHAM, JOHN  DOUBLE INVADERS, THE  DG623
RACKHAM, JOHN  EARTHSTRINGS  D10293
RACKHAM, JOHN  FLOWER OF DORADIL  D24100
RACKHAM, JOHN  IPOMOEA  D37250
RACKHAM, JOHN  PROXIMA PROJECT, THE  DH091
RACKHAM, JOHN  TIME TO LIVE  DG606
RACKHAM, JOHN  TREASURE OF TAU CETI  D23775
RACKHAM, JOHN  WE, THE VENUSIANS  DM127
RADCLIFFE, GARNETT  LADY FROM VENUS, THE  H
RADFORD, IRENE  DRAGON CIRCLE, THE [@STARGODS #2]  P
RADFORD, IRENE  DRAGON’S TOUCHSTONE, THE [DRAGON NIMBUS HISTORY 1]  C/P,P
RADFORD, IRENE  GLASS DRAGON, THE [DRAGON NIMBUS 1]  C/P:2,P
RADFORD, IRENE  GUARDIAN OF THE BALANCE [@MERLIN’S DESCENDANTS #1]  H,P
RADFORD, IRENE  GUARDIAN OF THE PROMISE [@MERLIN’S DESCENDANTS #4]  H,P
RADFORD, IRENE  GUARDIAN OF THE TRUST [@MERLIN’S DESCENDANTS #2]  P
RADFORD, IRENE  GUARDIAN OF THE VISION [@MERLIN’S DESCENDANTS #3]  P
RADFORD, IRENE  LAST BATTLEMAGE, THE [DRAGON NIMBUS HISTORY 2]  P
RADFORD, IRENE  LONELIEST MAGICIAN, THE [DRAGON NIMBUS 3]  P
RADFORD, IRENE  PERFECT PRINCESS, THE [DRAGON NIMBUS 2]  C/P:2,P
RADFORD, IRENE  RENEGADE DRAGON, THE [DRAGON NIMBUS HISTORY 3]  P
RADFORD, IRENE  WIZARD’S TREASURE, THE [DRAGON NIMBUS 4]  P
RADIO SPIRITS, INC.  SUPERMAN ON RADIO (4 CASSETTES, BOXED)  SR-MM
RAE, HUGH C.  HARKFAST: THE MAKING OF THE KING  P
RAE, HUGH C.  SKINNER  P
RAE, HUGH C.  TRAVELING SOUL, THE  P
RAHMAN, GLENN  HEIR OF DARKNESS  P
RAINEY, STEPHEN MARK  BALAK  LP
RAINEY, STEPHEN MARK  SONG OF CTHULHU  LPA
RALSTON, GILBERT A.  BEN  P
RAMBELLI, ROBERTA  MINISTERO DELLA FELICITA, IL  FP
RAME, DAVID  TUNNEL FROM CALAIS  H
RAMON Y CAJAL, SANTIAGO  VACATION STORIES  H
RAMPTON, ANTHONY  PANTHER BOOK OF HORROR, THE  SR-PA
RAMSLAND, KATHERINE  ANNE RICE TRIVIA BOOK, THE  R/REF-P,C/REF-P
RAMSLAND, KATHERINE  WITCH’S COMPANION, THE [MAYFAIR WITCHES]  R/REF-H
RAND, AYN  ANTHEM  C/P:2,P
RAND, AYN  ATLAS SHRUGGED  C/P:2,P
RAND, AYN  FOUNTAINHEAD, THE  P
RANDA, PETER  COMMANDO DU NON-RETOUR  FP
RANDA, PETER  GRANDE DERIVE, LA  FP
RANDA, PETER  OBJECTIF TAMAX  FP
RANDA, PETER  SURVIVANTS DE KOR, LES  FP
RANDA, PETER  VAGUES D’INVASION  FP
RANDALL, BOB  NEXT, THE  P
RANDALL, FLORENCE ENGEL  WATCHER IN THE WOODS, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 314
AUTHOR  TITLE  CODE
RANDALL, MARTA  CITY IN THE NORTH, A  C/P:2,P
RANDALL, MARTA  DANGEROUS GAMES  C/P:2,P
RANDALL, MARTA  ISLANDS  C/P:2,P
RANDALL, MARTA  JOURNEY  C/P:2,P
RANDALL, MARTA  NEBULA AWARDS 19, THE  HA
RANDALL, MARTA/SILVERBERG, ROBERT  NEW DIMENSIONS 12  PA,C/PA
RANDALL, MARTA  SWORD OF WINTER, THE  H:2,P
RANDALL, MARTA  THOSE WHO FAVOR FIRE  C/P,P
RANDALL, NEIL  SEVEN NO-TRUMP [AMBER]  SFWA-P
RANDALL, NEIL  STORM OF DUST  P,SFWA-P
RANDALL, ROBERT  DAWNING LIGHT, THE  C/P,P,H
RANDALL, ROBERT  SHROUDED PLANET, THE  SR-P,C/P:2,H
RANDLE, KEVIN/CORNETT, ROBERT  ALDEBARAN CAMPAIGN, THE [SEEDS OF WAR 2]  C/P,P
RANDLE, KEVIN/CORNETT, ROBERT  AQUARIAN ATTACK, THE [SEEDS OF WAR 3]  C/P:2,P
RANDLE, KEVIN  CHAIN OF COMMAND [JEFFERSON’S WAR]  P
RANDLE, KEVIN  CITADEL, THE [GALACTIC MI]  P
RANDLE, KEVIN  DEATH OF A REGIMENT [JEFFERSON’S WAR 5]  C/P,P
RANDLE, KEVIN  F.T.L. [@EXPLORATION #3]  P
RANDLE, KEVIN  GALACTIC MI [GALACTIC MI]  P
RANDLE, KEVIN  GALACTIC SILVER STAR, THE [JEFFERSON’S WAR 1]  P
RANDLE, KEVIN  GATE, THE [@EXPLORATION #4]  P
RANDLE, KEVIN/DRISCOLL, RICHARD  INSIDE JOB [@STAR PRECINCT #3]  C/P,P
RANDLE, KEVIN  JANUARY PLATOON, THE [JEFFERSON’S WAR 4]  C/P,P
RANDLE, KEVIN  LOST COLONY, THE [JEFFERSON’S WAR 3]  C/P,P
RANDLE, KEVIN/DRUSCOLL, RICHARD  MINDSLAYER [@STAR PRECINCT #2]  C/P,P
RANDLE, KEVIN  OCTOBER SCENARIO, THE  [C/RAN-P]
RANDLE, KEVIN  PRICE OF COMMAND, THE [JEFFERSON’S WAR 2]  C/P,P
RANDLE, KEVIN  RAT TRAP, THE [GALACTIC MI 2]  C/P,P
RANDLE, KEVIN/CORNETT, ROBERT  REMEMBER GETTYSBURG! [REMEMBER! 2]  P
RANDLE, KEVIN/CORNETT, ROBERT  REMEMBER THE ALAMO! [REMEMBER! 1]  P
RANDLE, KEVIN/CORNETT, ROBERT  REMEMBER THE LITTLE BIGHORN! [REMEMBER! 3]  P
RANDLE, KEVIN/CORNETT, ROBERT  SEEDS OF WAR [SEEDS OF WAR 1]  P
RANDLE, KEVIN  SIGNALS [@EXPLORATION #1]  C/P,P
RANDLE, KEVIN/DRISCOLL, RICHARD  STAR PRECINCT [@STAR PRECINCT #1]  P
RANDLE, KEVIN  STARSHIP [@EXPLORATION #2]  C/P,P
RANDLE, KEVIN D.  HISTORY OF UFO CRASHES, A  [C/RAN-P]
RANDLE, KEVIN D.  OPERATION ROSWELL  H
RANDLE, KEVIN D./SCHMITT, DONALD R.  UFO CRASH AT ROSWELL  [C/RAN-P]
RANDOLPHE, ARABELLA  VAMPIRE TAPES, THE  P
RANDOM HOUSE  THREE: 1971  HA
RANKIN, ROBERT  ANTIPOPE, THE [BRENTFORD 1]  P
RANKIN, ROBERT  APOCALYPSO  P
RANKIN, ROBERT  ARMAGEDDON: THE MUSICAL [ARMAGEDDON 1]  LP
RANKIN, ROBERT  ARMAGEDDON: THE MUSICAL (PROOF) [ARMAGEDDON 1]  SR-LP
RANKIN, ROBERT  BOOK OF ULTIMATE TRUTHS, THE [CORNELIUS MURPHY 1]  P
RANKIN, ROBERT  BRENTFORD TRIANGLE, THE [BRENTFORD 2]  P
RANKIN, ROBERT  BRIGHTONOMICON, THE  LP

RANKIN, ROBERT DANCE OF THE VOODOO HANDBAG, THE

[TRILOGY THAT DARE NOT SPEAK ITS NAME 2]

P

RANKIN, ROBERT  DOG CALLED DEMOLITION, A  P
RANKIN, ROBERT  EAST OF EALING [BRENTFORD 3]  P
RANKIN, ROBERT  EDDIE BEAR, PRIVATE DETECTIVE [EDDIE BEAR 1,2]  H
RANKIN, ROBERT  FANDOM OF THE OPERATOR, THE  P
RANKIN, ROBERT  GARDEN OF UNEARTHLY DELIGHTS, THE  P

RANKIN, ROBERT HOLLOW CHOCOLATE BUNNIES OF THE APOCALYPSE, THE

[EDDIE BEAR 1]

H

RANKIN, ROBERT  KNEES UP MOTHER EARTH [BRENTFORD 7/WITCHES OF CHISWICK 2]  H
RANKIN, ROBERT  MOST AMAZING MAN WHO EVER LIVED, THE [CORNELIUS MURPHY 3]  P
RANKIN, ROBERT  NOSTRADAMUS ATE MY HAMSTER  P
RANKIN, ROBERT  RAIDERS OF THE LOST CAR PARK [CORNELIUS MURPHY 2]  P
RANKIN, ROBERT  SNUFF FICTION  P
RANKIN, ROBERT  SPROUT MASK REPLICA [TRILOGY THAT DARE NOT SPEAK ITS NAME 1]  P
RANKIN, ROBERT  SPROUTS OF WRATH, THE [BRENTFORD 4]  P
RANKIN, ROBERT  SUBURBAN BOOK OF THE DEAD, THE [ARMAGEDDON 3]  P
RANKIN, ROBERT  WAITING FOR GODALMING [TRILOGY THAT DARE NOT SPEAK ITS NAME 3] P
RANKIN, ROBERT  WEB SITE STORY  P
RANKIN, ROBERT  WITCHES OF CHISWICK, THE [WITCHES OF CHISWICK 1]  P
RANKINE, JOHN  ANDROID PLANET [SPACE: 1999 8]  C/P,P
RANKINE, JOHN  ASTRAL QUEST [SPACE: 1999 6]  P
RANKINE, JOHN  BROMIUS PHENOMENON, THE  P
RANKINE, JOHN  INTERSTELLAR TWO FIVE  P
RANKINE, JOHN  LUNAR ATTACK [SPACE: 1999 5]  C/P,P
RANKINE, JOHN  MOON ODYSSEY [SPACE: 1999 2]  C/P:2,P
RANKINE, JOHN  MOONS OF TRIOPUS  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 315
AUTHOR  TITLE  CODE
RANKINE, JOHN  OPERATION UMANAQ  C/P:2,P
RANKINE, JOHN  PHOENIX OF MEGARON [SPACE: 1999 10]  P
RANSOM, BILL  BURN  C/P,P
RANSOM, BILL  JAGUAR  C/P:2,SFWA-P:2,P
RANSOM, BILL  VIRAVAX  P,SFWA
RANSOM, CANDICE  KEY TO THE GRIFFON’S LAIR [@KNIGHTS OF THE SILVER DRAGON #9]  LP
RANSOM, DANIEL  FUGITIVE STARS, THE  P
RANSOM, DANIEL  ZONE SOLDIERS  C/P,P
RAPHAEL, RICK  CODE THREE  C/P:2,P
RAPHAEL, RICK  THIRST QUENCHERS, THE  H
RASCOVICH, MARK  BEDFORD INCIDENT, THE  P
RASMUSSEN, ALIS A.  LABYRINTH GATE, THE  C/P:2,SFWA-P,P
RASMUSSEN, ALIS A.  PASSAGE OF STARS, A [@HIGHROAD #1]  C/P:2,P
RASMUSSEN, ALIS A.  PRICE OF RANSOM, THE [@HIGHROAD #3]  C/P:2,P
RASMUSSEN, ALIS A.  REVOLUTION’S SHORE [@HIGHROAD #2]  P
RASPAIL, JEAN  CAMP OF THE SAINTS, THE  P

RASPE, R. E.

SINGULAR TRAVELS, CAMPAIGNS AND ADVENTURES OF BARON MUNCHAUSEN

LP

RATHBONE, BASIL  BASIL RATHBONE SELECTS STRANGE TALES  PA
RATHJEN, CARL HENRY  FLIGHT OF FEAR [LAND OF THE GIANTS]  H
RAUCH, EARL MAC  BUCKAROO BANZAI  C/P:2,P
RAVENSWOOD, FRITZEN  SPAWNING, THE  P
RAWN, MELANIE  DRAGON PRINCE [DRAGON PRINCE 1]  C/P,P
RAWN, MELANIE  DRAGON TOKEN, THE [DRAGON STAR 2]  H:2,P

RAWN, MELANIE/ROBERSON, JENNIFER/ ELLIOTT, KATE GOLDEN KEY, THE

C/P:2,P

RAWN, MELANIE  KNIGHTS OF THE MORNINGSTAR [QUANTUM LEAP]  C/P:2,P
RAWN, MELANIE  MAGEBORN TRAITOR, THE [@EXILES #2]  C/P,P,H
RAWN, MELANIE  RUINS OF AMBRAI, THE [@EXILES #1]  C/P,P,H
RAWN, MELANIE  SKYBOWL [DRAGON STAR 3]  C/P,SFWA-P,H,P,SFWA
RAWN, MELANIE  SPELLBINDER  H
RAWN, MELANIE  STAR SCROLL, THE [DRAGON PRINCE 2]  C/P:2,P
RAWN, MELANIE  STRONGHOLD [DRAGON STAR 1]  C/P,P,H
RAWN, MELANIE  SUN-RUNNER’S FIRE [DRAGON PRINCE 3]  C/P:2,P
RAY, JEAN  GHOULS IN MY GRAVE  SR-P
RAY, ROBERT  NO STARS FOR US  SR-P
RAYER, FRANCIS G.  JOURNEY TO THE STARS  H
RAYER, FRANCIS G.  STAR SEEKERS, THE  SR-P
RAYJEAN, MAX-ANDRE  ALIEN (CHINESE)  FP
RAYJEAN, MAX-ANDRE  ANTI-HOMMES, LES  FP
RAYJEAN, MAX-ANDRE  CLES DE L’UNIVERS, LES  FP
RAYJEAN, MAX-ANDRE  PLAN S.03  FP
RAYJEAN, MAX-ANDRE  PROJET "KOZNA"  FP
RAYJEAN, MAX-ANDRE  PUISSANCE: FACTEUR 3  FP
RAYJEAN, MAX-ANDRE  SEPTIEME CONTINENT, LE  FP
RAYJEAN, MAX-ANDRE  ZOO DES ASTORS, LE  FP
RAYNER, JACQUELINE  EARTHWORLD [DOCTOR WHO EIGHTH DOCTOR]  P

RAYNER, JACQUELINE PROFESSOR BERNICE SUMMERFIELD AND THE GLASS PRISON

[BERNICE SUMMERFIELD 5]

P

RAYNER, JACQUELINE PROFESSOR BERNICE SUMMERFIELD AND THE SQUIRE’S CRYSTAL

[BERNICE SUMMERFIELD 3]

P

RAYNER, JACQUELINE  SHORT TRIPS: COMPANIONS [DOCTOR WHO]  HA
RAYNER, JACQUELINE  SHORT TRIPS: THE MUSES [DOCTOR WHO]  HA
RAYNER, JACQUELINE  SHORT TRIPS: ZODIAC [DOCTOR WHO]  HA
RAYNER, JACQUELINE  STONE ROSE, THE [@DOCTOR WHO]  H
RAYNER, JACQUELINE  WINNER TAKES ALL [DOCTOR WHO NINTH DOCTOR]  H
RAYNER, JACQUELINE  WOLFSBANE [DOCTOR WHO PAST DOCTORS]  P
RAZZI, JAMES  STAR TREK PUZZLE MANUAL  [C/RAN-P:2]
READ, HERBERT  GREEN CHILD, THE  H
READE, QUINN  QUEST OF THE DARK LADY  P
READY, WILLIAM  TOLKIEN RELATION, THE  R/REF-H
READY, WILLIAM  UNDERSTANDING TOLKIEN AND THE LORD OF THE RINGS  R/REF-P,C/REF-P
REAGEN, JAMES E.  LEAGUE OF THE CRIMSON CRESCENT, THE [WIZARDRY]  P
REAMY, TOM  BLIND VOICES  H:2,P
REAMY, TOM  SAN DIEGO LIGHTFOOT SUE  C/P:2,P
REASONER, JAMES M.  TERRAN GIRLS MAKE WONDERFUL WIVES  GRD8
REAVES, MICHAEL/PERRY, STEVE  BATTLE SURGEONS [@STAR WARS CLONE WARS MEDSTAR 1]  P
REAVES, MICHAEL  BURNING REALM, THE  C/P:2,P
REAVES, MICHAEL  DARKWORLD DETECTIVE  C/P:2,P
REAVES, MICHAEL  DARTH MAUL: SHADOW HUNTER [STAR WARS]  P
REAVES, MICHAEL/PERRY, STEVE  DOME  P,SFWA-P
REAVES, MICHAEL  HELL ON EARTH  H
REAVES, MICHAEL/PERRY, STEVE  HELLSTAR  C/P:2,P
REAVES, MICHAEL  I — ALIEN  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 316
AUTHOR  TITLE  CODE
REAVES, MICHAEL/PERRY, STEVE  MEDSTAR [@STAR WARS CLONE WARS MEDSTAR 1,2]  H
REAVES, MICHAEL  NIGHT HUNTER  SFWA
REAVES, MICHAEL/PELAN, JOHN  SHADOWS OVER BAKER STREET  HA
REAVES, MICHAEL  SHATTERED WORLD, THE  C/P,LP,H
REAVES, MICHAEL  STREET MAGIC  H
REAVES, MICHAEL/PERRY, STEVE  SWORD OF THE SAMURAI  P
REAVES, MICHAEL  VOODOO CHILD  H
REDFINN, MICHAEL  BEING  P
REDMOND, JEFFREY  CHRONICLES OF ER-DA: BOOK ONE [ER-DA 1]  LP
REDMOND, JEFFREY  CHRONICLES OF ER-DA: BOOK TWO [ER-DA 2]  LP
REDMOND, JEFFREY  ER-DAN STORIES, THE [ER-DA]  LP
REED, ALISON  LINEUP, THE  SR-L
REED, DAVID V.  MURDER IN SPACE  C/P:2,GN4.23
REED, DAVID V.  WHISPERING GORILLA, THE  SR-P
REED, KIT  ARMED CAMPS  C/P,P
REED, KIT  DOGS OF TRUTH  LP
REED, KIT  FORT PRIVILEGE  P,SFWA
REED, KIT  MAGIC TIME  H,SFWA,P
REED, KIT  MISTER DA V.  C/P:2,P
REED, KIT  OTHER STORIES AND THE ATTACK OF THE GIANT BABY  P
REED, KIT  REVENGE OF THE SENIOR CITIZENS ** PLUS, THE  H
REED, PETER J.  KURT VONNEGUT, JR.  R/REF-P
REED, RICK R.  FACE WITHOUT A HEART, A  LP
REED, ROBERT  BENEATH THE GATED SKY  H
REED, ROBERT  BEYOND THE VEIL OF STARS  C/P,P,H
REED, ROBERT  BLACK MILK  C/P:2,P
REED, ROBERT  CUCKOO’S BOYS, THE  H
REED, ROBERT  DOWN THE BRIGHT WAY  C/P:2,P
REED, ROBERT  DRAGONS OF SPRINGSPACE, THE  H
REED, ROBERT  EXHALTATION OF LARKS, AN  LP
REED, ROBERT  HORMONE JUNGLE, THE  H,P
REED, ROBERT  LEESHORE, THE  H:2,P
REED, ROBERT  MARROW [MARROW 1]  H:2,P
REED, ROBERT  REMARKABLES, THE  C/P,P
REED, ROBERT  SISTER ALICE  H
REED, ROBERT  WELL OF STARS, THE [MARROW 2]  H
REEVE, ARTHUR B.  DREAM DOCTOR, THE  H
REEVE, PHILIP  DARKLING PLAIN, A [HUNGRY CITY 4]  H
REEVE, PHILIP  DARKLING PLAIN, A (PROOF) [@HUNGRY CITY 4]  SR-LP
REEVE, PHILIP  INFERNAL DEVICES [HUNGRY CITY 3]  H
REEVE, PHILIP  INFERNAL DEVICES (PROOF) [@HUNGRY CITY 3]  SR-LP
REEVE, PHILIP  LARKLIGHT  H
REEVE, PHILIP  MORTAL ENGINES [@HUNGRY CITY 1]  H
REEVE, PHILIP  MORTAL ENGINES (PROOF) [@HUNGRY CITY 1]  SR-LP
REEVE, PHILIP  PREDATOR’S GOLD [HUNGRY CITY 2]  H
REEVES-STEVENS, GARFIELD  DARK MATTER  LP

REEVES-STEVENS, GARFIELD/ REEVES-STEVENS, JUDITH MEMORY PRIME [STAR TREK]

C/P:3,P

REEVES-STEVENS, GARFIELD  NIGHTEYES  H:2,SR-LP

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD DAY OF DESCENT, THE [@ALIEN NATION #1]

C/P,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD FALL OF TEROK NOR [@STAR TREK DS9 MILLENIUM #1]

C/P,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD FEDERATION [STAR TREK]

C/P,P,H

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD INFERNO [@STAR TREK DS9 MILLENIUM #3]

C/P,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD MAKING OF STAR TREK DEEP SPACE NINE, THE [STAR TREK]

R/REF-H

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD NIGHTFEEDER [GALEN SWORD 2]

P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD PRIME DIRECTIVE [STAR TREK]

H:2,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD SHIFTER [@GALEN SWORD #1]

C/P:2,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD WAR OF THE PROPHETS, THE [@STAR TREK DS9 MILLENIUM #2]

C/P,P

REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD WORLDS IN COLLISION [STAR TREK]

LP

REGAN, DIAN CURTIS  PRINCESS NEVERMORE  P

REGINALD, R.

SCIENCE FICTION AND FANTASY LITERATURE 1:

INDEXES TO THE LITERATURE

R/IND

REGINALD, R.

SCIENCE FICTION AND FANTASY LITERATURE 2: CONTEMPORARY SCIENCE FICTION AUTHORS

R/IND

REGINALD, ROBERT  SCIENCE FICTION AND FANTASY LITERATURE, 1975–1991  R/IND

The MITSFS Pinkdex  September 24, 2007  Page 317
AUTHOR  TITLE  CODE
REICHERT, MICKEY ZUCKER  BEASTS OF BARAKHAI [@BARAKHAI #1]  H,P
REICHERT, MICKEY ZUCKER  BEYOND RAGNAROK [RENSHAI CHRONICLES 1]  P,SFWA
REICHERT, MICKEY ZUCKER  BY CHAOS CURSED [BIFROST GUARDIANS 5]  P
REICHERT, MICKEY ZUCKER  CHILD OF THUNDER [RENSHAI 3]  C/P,P
REICHERT, MICKEY ZUCKER  CHILDREN OF WRATH, THE [RENSHAI CHRONICLES 3]  P
REICHERT, MICKEY ZUCKER  DRAGONRANK MASTER [BIFROST GUARDIANS 3]  C/P,P
REICHERT, MICKEY ZUCKER  FLIGHTLESS FALCON  H
REICHERT, MICKEY ZUCKER  GODSLAYER [BIFROST GUARDIANS 1]  C/P:2,P
REICHERT, MICKEY ZUCKER  LAST OF THE RENSHAI, THE [RENSHAI 1]  C/P:2,P
REICHERT, MICKEY ZUCKER  LEGEND OF NIGHTFALL, THE [NIGHTFALL 1]  C/P,P
REICHERT, MICKEY ZUCKER  LOST DRAGONS OF BARAKHAI, THE [@BARAKHAI #2]  H,P
REICHERT, MICKEY ZUCKER  PRINCE OF DEMONS [RENSHAI CHRONICLES 2]  P
REICHERT, MICKEY ZUCKER  SHADOW CLIMBER [BIFROST GUARDIANS 2]  C/P,P
REICHERT, MICKEY ZUCKER  SHADOW’S REALM [BIFROST GUARDIANS 4]  P

REICHERT, MICKEY ZUCKER/ WINGERT, JENNIFER SPIRIT FOX

C/P,P

REICHERT, MICKEY ZUCKER  UNKNOWN SOLDIER, THE  C/P,SFWA-P,P
REICHERT, MICKEY ZUCKER  WESTERN WIZARD, THE [RENSHAI 2]  C/P,P
REID, C. LESTOCK  SONS OF SOLOMON  H

REID, THOMAS M.

EMERALD SCEPTER, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #3]

P

REID, THOMAS M.  GOSSAMER PLAIN, THE [@FORGOTTEN REALMS EMPYREAN ODYSSEY #1]  P
REID, THOMAS M.  GRIDRUNNER [STAR*DRIVE]  C/P,P
REID, THOMAS M.  INSURRECTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #2]  H
REID, THOMAS M.  RUBY GUARDIAN, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #2]  P

REID, THOMAS M.

SAPPHIRE CRESCENT, THE [@FORGOTTEN REALMS SCIONS OF ARRABAR #1]

P

REID, THOMAS M.  TEMPLE OF ELEMENTAL EVIL, THE [@GREYHAWK]  P
REIDA, ALVAH  FAULT LINES  P
REIMANN, KATYA  PRINCE OF FIRE AND ASHES [TIELMARAN 3]  P
REIMANN, KATYA  TREMOR IN THE BITTER EARTH, A [TIELMARAN 2]  H,P
REIMANN, KATYA  WIND FROM A FOREIGN SKY [TIELMARAN 1]  H,P
REIN, HAROLD  FEW WERE LEFT  H
REINIUS, TRISH  PLANET OF TEARS, THE  C/P:2,P
REINSMITH, RICHARD  SAVAGE STARS, THE  P
REIS, RICHARD H.  GEORGE MACDONALD  R/REF-H
REISFELD, RANDI  ALL YOU NEED IS A LOVE SPELL [@SABRINA THE TEENAGE WITCH #7]  P
RELLING, WILLIAM JR.  BRUJO  SFWA-P
RELLING, WILLIAM JR.  NEW MOON  P
RELLING, WILLIAM JR.  SILENT MOON  SFWA-P
RENARD, JOSEPH  MONODYNE CATASTROPHE, THE  P
RENAULT, MARY  BULL FROM THE SEA, THE  P
RENAULT, MARY  KING MUST DIE, THE  P
RENAULT, MARY  KING MUST DIE, THE/BULL FROM THE SEA, THE  H
RENNIE, GORDON  CRUCIBLE [@ROGUE TROOPER]  P
RENNIE, GORDON  EXECUTION HOUR [WARHAMMER 40K]  P
RENNIE, GORDON/MACNEIL, COLIN  IN THE EYE OF TERROR [WARHAMMER BLOODQUEST 2]  CX
RENNIE, GORDON/KOPINSKI, KARL  KAL JERICO [WARHAMMER]  CX
RENNIE, GORDON  SHADOW POINT [@WARHAMMER 40K]  P
RENNIE, GORDON  ZAVANT [WARHAMMER ZAVANT KONNIGER]  P
RENO, KELLY  MISADVENTURES & MERFOLK  LP
RENWICK, BRETT L.  WIRED FOR CHAOS  LP
REPP, ED EARL  RADIUM POOL, THE  SR-H,H
REPP, ED EARL  STELLAR MISSILES, THE  H,SR-LP
RESNICK, LAURA  DESTROYER GODDESS, THE [IN FIRE FORGED 2]  P
RESNICK, LAURA  DISAPPEARING NIGHTLY [ESTHER DIAMOND 1]  LP
RESNICK, LAURA  IN LEGEND BORN  H:2
RESNICK, LAURA  WHITE DRAGON, THE [IN FIRE FORGED 1]  H,P
RESNICK, MIKE  ADVENTURES [LUCIFER JONES 1]  P
RESNICK, MIKE/GREENBERG, MARTIN H.  ALADDIN: MASTER OF THE LAMP  C/PA,PA
RESNICK, MIKE  ALTERNATE KENNEDYS  PA
RESNICK, MIKE  ALTERNATE OUTLAWS  PA
RESNICK, MIKE  ALTERNATE PRESIDENTS  C/PA,PA
RESNICK, MIKE/HAYDEN, PATRICK NIELSEN ALTERNATE SKIFFY  LP
RESNICK, MIKE  ALTERNATE TYRANTS  LPA
RESNICK, MIKE  ALTERNATE WARRIORS  PA
RESNICK, MIKE  ALTERNATE WORLDCONS AND AGAIN, ALTERNATE WORLDCONS  SR-LPA

RESNICK, MIKE BEST ROOTIN’ TOOTIN’ SHOOTIN’ GUNSLINGER IN THE WHOLE DAMNED

GALAXY, THE [GALACTIC MIDWAY 4]

C/P:2,P

RESNICK, MIKE  BIRTHRIGHT: THE BOOK OF MAN  C/P:2,P
RESNICK, MIKE  BRANCH, THE  P
RESNICK, MIKE  BWANA  TD33
RESNICK, MIKE  BWANA/BULLY!  C/P:3,P
RESNICK, MIKE/GREENBERG, MARTIN H.  BY ANY OTHER FAME  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 318
AUTHOR  TITLE  CODE
RESNICK, MIKE/GREENBERG, MARTIN H.  CHRISTMAS GHOSTS  PA
RESNICK, MIKE  DARK LADY, THE  C/P:2,SFWA-P,P

RESNICK, MIKE/GREENBERG, MARTIN H./ ESTLEMAN, LOREN D.

DEALS WITH THE DEVIL

PA,C/PA

RESNICK, MIKE/GREENBERG, MARTIN H.  DINOSAUR FANTASTIC  HA,PA
RESNICK, MIKE  DOWN THESE DARK SPACEWAYS  HA
RESNICK, MIKE  DRAGON AMERICA  LP
RESNICK, MIKE  EROS ASCENDING [VELVET COMET #1]  C/P,@P,H
RESNICK, MIKE  EROS AT NADIR [VELVET COMET #4]  C/P,@P
RESNICK, MIKE  EROS AT ZENITH [VELVET COMET #2]  C/P:2,@P,H
RESNICK, MIKE  EROS DESCENDING [VELVET COMET #3]  C/P,@P
RESNICK, MIKE  GATHERING OF WIDOWMAKERS, A [WIDOWMAKER 4]  H
RESNICK, MIKE  GODDESS OF GANYMEDE, THE  SR-H,P
RESNICK, MIKE  HUNGER IN THE SOUL, A  LP
RESNICK, MIKE  I, ALIEN  PA
RESNICK, MIKE  INFERNO  H
RESNICK, MIKE  INSIDE THE FUNHOUSE  PA,SFWA-P
RESNICK, MIKE  IVORY  C/P:2,SFWA-P,P
RESNICK, MIKE  KIRINYAGA (NOVEL)  H
RESNICK, MIKE  KIRINYAGA (SHORT STORY)  B
RESNICK, MIKE  LUCIFER JONES [LUCIFER JONES 2]  C/P,P
RESNICK, MIKE  MEN WRITING SCIENCE FICTION AS WOMEN  PA
RESNICK, MIKE  MIRACLE OF RARE DESIGN, A  H:2,P
RESNICK, MIKE  MORE WHATDUNITS  PA
RESNICK, MIKE  NEBULA AWARDS SHOWCASE 2007  LPA
RESNICK, MIKE  NEW DREAMS FOR OLD  LP
RESNICK, MIKE  NEW VOICES IN SCIENCE FICTION  PA
RESNICK, MIKE  OFFICIAL GUIDE TO FANTASTIC LITERATURE, THE  R/REF-P
RESNICK, MIKE  ORACLE  C/P,SFWA-P:2,P
RESNICK, MIKE  OUTPOST, THE  H
RESNICK, MIKE  PARADISE  C/P,SFWA-P,P,H
RESNICK, MIKE  PROPHET  C/P,SFWA-P:2,P
RESNICK, MIKE  PURGATORY  C/P:2,P
RESNICK, MIKE  PURSUIT ON GANYMEDE  P
RESNICK, MIKE  REDBEARD  C/P:2,P
RESNICK, MIKE  RETURN OF SANTIAGO, THE [SANTIAGO 2]  H
RESNICK, MIKE/GREENBERG, MARTIN H.  RETURN OF THE DINOSAURS  C/PA,PA
RESNICK, MIKE  SAFARI OF THE MIND, A  LP
RESNICK, MIKE  SANTIAGO [SANTIAGO 1]  C/P:2,SFWA-P,P
RESNICK, MIKE  SECOND CONTACT  C/P:2,SFWA-P,P
RESNICK, MIKE/GREENBERG, MARTIN H.  SHERLOCK HOLMES IN ORBIT  C/PA,PA
RESNICK, MIKE  SIDESHOW [GALACTIC MIDWAY 1]  C/P,P
RESNICK, MIKE  SOLO FLIGHTS THROUGH SHARED WORLDS  LP
RESNICK, MIKE  SOOTHSAYER  C/P:2,SFWA-P:2,P
RESNICK, MIKE  SOUL EATER, THE  C/P,P
RESNICK, MIKE  STALKING THE UNICORN  C/P:2,SFWA-P,P
RESNICK, MIKE  STALKING THE WILD RESNICK  SR-H
RESNICK, MIKE  STARSHIP: MUTINY [STARSHIP THEODORE ROOSEVELT 1]  H
RESNICK, MIKE  STARSHIP:PIRATE [STARSHIP 2]  H
RESNICK, MIKE  THIS IS MY FUNNIEST  LPA
RESNICK, MIKE  THREE-LEGGED HOOTCH DANCER, THE [GALACTIC MIDWAY 2]  C/P,P
RESNICK, MIKE  UNAUTHORIZED AUTOBIOGRAPHIES  B
RESNICK, MIKE/DOZOIS, GARDNER  UNDER AFRICAN SKIES [@FUTURE EARTHS]  C/PA,PA
RESNICK, MIKE/DOZOIS, GARDNER  UNDER SOUTH AMERICAN SKIES [@FUTURE EARTHS]  PA
RESNICK, MIKE  WALPURGIS III  P
RESNICK, MIKE  WHATDUNITS  SFWA-P
RESNICK, MIKE  WIDOWMAKER, THE [WIDOWMAKER 1]  C/P,P
RESNICK, MIKE  WIDOWMAKER REBORN, THE [WIDOWMAKER 2]  C/P,P
RESNICK, MIKE  WIDOWMAKER UNLEASHED, THE [WIDOWMAKER 3]  P
RESNICK, MIKE  WIDOWMAKERS, THE [WIDOWMAKER 1,2,3]  H:2
RESNICK, MIKE  WILD ALIEN TAMER, THE [GALACTIC MIDWAY 3]  C/P,P

RESNICK, MIKE WILL THE LAST PERSON TO LEAVE THE PLANET PLEASE SHUT OFF THE

SUN

H

RESNICK, MIKE/GREENBERG, MARTIN H.  WITCH FANTASTIC  C/PA,PA
RESNICK, MIKE  WOMEN WRITING SCIENCE FICTION AS MEN  PA
RESNICK, MIKE/SICLARI, JOE  WORLDCON GUEST OF HONOR SPEECHES  R/REF-H
REWOLINSKI, LEAH  GEEK SPACE NINE [@STAR WRECK #6]  P
REWOLINSKI, LEAH  GENERATION GAP, THE [@STAR WRECK #1]  C/P,P
REWOLINSKI, LEAH  SPACE, THE FIDO FRONTIER [@STAR WRECK #7]  P
REWOLINSKI, LEAH  UNDISCOVERED NURSING HOME, THE [@STAR WRECK #5]  P
REY, RAINER  REPLICATOR RUN  P
REY, RUSSELL  VALLEY OF TERROR  SR-P
REYNARD, ELIZABETH  MUTINOUS WIND: A SORCERER’S TALE, THE  H
REYNOLD, PAMELA  EARTH TIMES TWO  H

AUTHOR  TITLE  CODE
REYNOLDS, ALASTAIR  ABSOLUTION GAP [REVELATION SPACE 3]  H
REYNOLDS, ALASTAIR  CENTURY RAIN [REVELATION SPACE]  H:2
REYNOLDS, ALASTAIR  CHASM CITY [REVELATION SPACE]  H
REYNOLDS, ALASTAIR  DIAMOND DOGS, TURQUOISE DAYS [REVELATION SPACE]  H
REYNOLDS, ALASTAIR  GALACTIC NORTH  H
REYNOLDS, ALASTAIR  PUSHING ICE  LP
REYNOLDS, ALASTAIR  REDEMPTION ARK [REVELATION SPACE 2]  H:2
REYNOLDS, ALASTAIR  REVELATION SPACE [REVELATION SPACE 1]  C/P,P,H
REYNOLDS, ALASTAIR  ZIMA BLUE  H
REYNOLDS, ALFRED  KITEMAN  P
REYNOLDS, ALFRED  KITEMAN OF KARANGA  H
REYNOLDS, ANTHONY  MARK OF CHAOS [@WARHAMMER MARK OF CHAOS]  P
REYNOLDS, DAVID WEST  STAR WARS EPISODE I: INCREDIBLE CROSS-SECTIONS  [C/RAN-H]
REYNOLDS, DAVID WEST  STAR WARS EPISODE I: THE VISUAL DICTIONARY  [C/RAN-H]
REYNOLDS, ERIC T.  GOLDEN AGE SF  HA
REYNOLDS, G. W. M.  WAGNER, THE WEHR-WOLF  VLP
REYNOLDS, JAMES  GHOSTS IN IRISH HOUSES  [C/RAN-P]
REYNOLDS, JOHN MURRAY  PRIVATE LIFE OF HENRY PERKINS, THE  H
REYNOLDS, KAY  ROBOTECH ART 2  A
REYNOLDS, MACK  ABILITY QUOTIENT  C/P:2,P
REYNOLDS, MACK  AFTER SOME TOMORROW  C/P:2,P
REYNOLDS, MACK  AFTER UTOPIA  C/P,P
REYNOLDS, MACK  AGENTE PLANETARIO X  FP
REYNOLDS, MACK  AMAZON PLANET  C/P:2,P
REYNOLDS, MACK  BEST OF MACK REYNOLDS, THE  C/P:2,P
REYNOLDS, MACK  BEST YE BREED, THE  C/P:2,P
REYNOLDS, MACK  BLACKMAN’S BURDEN  D06612
REYNOLDS, MACK  BORDER, BREED NOR BIRTH  D06612
REYNOLDS, MACK  BRAIN WORLD  C/P,P
REYNOLDS, MACK  CASE OF THE LITTLE GREEN MEN, THE  H
REYNOLDS, MACK  CHAOS IN LAGRANGIA  C/P,P
REYNOLDS, MACK  CODE DUELLO  D11650,DH103
REYNOLDS, MACK  COMMUNE 2000 A. D.  C/P:2,P
REYNOLDS, MACK  COMPOUNDED INTERESTS  SR-H
REYNOLDS, MACK  COMPUTER WAR  DH034,D11650
REYNOLDS, MACK  COSMIC EYE, THE  C/P:2,P
REYNOLDS, MACK  DAWNMAN PLANET  DG580
REYNOLDS, MACK  DAY AFTER TOMORROW  C/P,P
REYNOLDS, MACK/ING, DEAN  DEATHWISH WORLD  C/P:2,P
REYNOLDS, MACK  DEPRESSION OR BUST/DAWNMAN PLANET  C/P:2,P
REYNOLDS, MACK  EARTH UNAWARE  C/P:2,P
REYNOLDS, MACK  EARTH WAR, THE  C/P:2,P
REYNOLDS, MACK  EQUALITY: IN THE YEAR 2000  C/P,P
REYNOLDS, MACK/ING, DEAN  ETERNITY  C/P:2,P
REYNOLDS, MACK  FIVE WAY SECRET AGENT, THE/MERCENARY FROM TOMORROW  C/P:2,P
REYNOLDS, MACK  FRACAS FACTOR, THE  C/P,P
REYNOLDS, MACK  GALACTIC MEDAL OF HONOR  C/P:2,P
REYNOLDS, MACK/ING, DEAN  HOME SWEET HOME: 2010 A. D.  P
REYNOLDS, MACK/BANKS, MICHAEL  JOE MAUSER, MERCENARY FROM TOMORROW  C/P:2,P
REYNOLDS, MACK  KISS BEFORE LOVING, A  P
REYNOLDS, MACK  LAGRANGE FIVE  C/P:2,P
REYNOLDS, MACK  LAGRANGISTS, THE  P
REYNOLDS, MACK  LOOKING BACKWARD FROM THE YEAR 2000  C/P:2,P
REYNOLDS, MACK  MERCENARY FROM TOMORROW  DH065
REYNOLDS, MACK  MISSION TO HORATIUS [STAR TREK]  H:2
REYNOLDS, MACK  OF GODLIKE POWER  C/P:2,P
REYNOLDS, MACK  ONCE DEPARTED  C/P,P
REYNOLDS, MACK/ING, DEAN  OTHER TIME, THE  C/P,P
REYNOLDS, MACK  PERCHANCE TO DREAM  C/P:2,P
REYNOLDS, MACK  PLANETARY AGENT X  DM131
REYNOLDS, MACK  POLICE PATROL: 2000 A.D.  C/P:2,P
REYNOLDS, MACK  RIVAL RIGELIANS, THE  DG632
REYNOLDS, MACK  RIVAL RIGELIANS, THE/PLANETARY AGENT X  P
REYNOLDS, MACK  ROLLTOWN  C/P:2,P
REYNOLDS, MACK  SATELLITE CITY  C/P:2,P
REYNOLDS, MACK  SECTION G: UNITED PLANETS  C/P:2,P
REYNOLDS, MACK  SPACE BARBARIANS, THE  D77710
REYNOLDS, MACK  SPACE SEARCH  C/P:2,P
REYNOLDS, MACK  SPACE VISITOR  C/P:2,P
REYNOLDS, MACK/BANKS, MICHAEL  SWEET DREAMS, SWEET PRINCES  C/P:2,P
REYNOLDS, MACK  TIME GLADIATOR  C/P,P
REYNOLDS, MACK  TOMORROW MIGHT BE DIFFERENT  C/P:2,P
REYNOLDS, MACK  TOWERS OF UTOPIA, THE  C/P:2,P
REYNOLDS, MACK  TRAMPLE AN EMPIRE DOWN  C/P,P
REYNOLDS, MACK/ING, DEAN  TROJAN ORBIT  C/P,P

AUTHOR  TITLE  CODE
REYNOLDS, TED  TIDES OF GOD, THE  C/P,P
RHEINGOLD, HOWARD  SAVAGE REPORT: 1994, THE  SR-P,C/P
RHEINGOLD, HOWARD  WAR OF THE GURUS  SR-P
RHINE, J. B.  NEW FRONTIERS OF THE MIND  [C/RAN-H]
RHODES, APOLLONIUS OF  VOYAGE OF ARGO, THE  P
RHODES, DANIEL  ADVERSARY  SFWA-P
RHODES, DANIEL  NEXT, AFTER LUCIFER  SFWA-P
RHODES, JENNA  FOUR FORGES, THE [ELVEN WAYS 1]  H
RHODES, NATASHA  BLADE TRINITY  P
RICCI, BARBARA GUIGNON  YEAR OF THE RATS, THE  H
RICE, ANNE  BLACKWOOD FARM [LESTAT]  H:2
RICE, ANNE  BLOOD AND GOLD [LESTAT]  H
RICE, ANNE  BLOOD CANTICLE [LESTAT]  H:2
RICE, ANNE  INTERVIEW WITH THE VAMPIRE [LESTAT 1]  C/P:2,P
RICE, ANNE  LASHER [MAYFAIR WITCHES 2]  C/P,P,H
RICE, ANNE  MEMNOCH THE DEVIL [LESTAT 5]  C/P,P,H
RICE, ANNE  MERRICK [MAYFAIR WITCHES/LESTAT]  H:2
RICE, ANNE  MUMMY, THE  C/P,P,H
RICE, ANNE  PANDORA [NEW TALES 1]  C/P,P,H
RICE, ANNE  QUEEN OF THE DAMNED, THE [LESTAT 3]  H:2,P
RICE, ANNE  TALE OF THE BODY THIEF, THE [LESTAT 4]  H:2,P
RICE, ANNE  TALTOS [MAYFAIR WITCHES 3]  H:2,P
RICE, ANNE  VAMPIRE ARMAND, THE [LESTAT 6]  H,LP
RICE, ANNE  VAMPIRE LESTAT, THE [LESTAT 2]  C/P,P,H
RICE, ANNE/PEROZICH, FAYE  VAMPIRE LESTAT, THE (GRAPHIC NOVEL) [LESTAT 2]  CX
RICE, ANNE  VITTORIO THE VAMPIRE [NEW TALES 2]  H:2,P
RICE, ANNE  WITCHING HOUR, THE [MAYFAIR WITCHES 1]  H:2,P
RICE, JANE  IDOL OF THE FLIES, THE  SR-H
RICE, JEFF  NIGHT STALKER, THE  P
RICE, JEFF  NIGHT STRANGLER, THE  C/P,P
RICE, PETER L.  FAR COUNTRY [@BATTLETECH 13]  C/P,P
RICHARD-BESSIERE, F.  AGENT SPATIAL NO. 1  FP
RICHARD-BESSIERE, F.  CERVEAUX SOUS CONTROLE  FP
RICHARD-BESSIERE, F.  CETTE LUEUR QUI VENAIT DES TENEBRES  FP
RICHARD-BESSIERE, F.  INVERSIA  FP
RICHARD-BESSIERE, F.  MAGES DE DEREB, LES  FP
RICHARD-BESSIERE, F.  PAS DE GONIA POUR LES GHARKANDES  FP
RICHARD-BESSIERE, F.  PELOS CAMINHOS DO ESPACO  FP
RICHARDS, EVAN  SOLID GOLD KIDNAPPING [@SIX MILLION DOLLAR MAN #2]  P
RICHARDS, HENRY  HOUR OF THE PHOENIX, THE  H
RICHARDS, JOEL  PINDHAREE  P,SFWA-P
RICHARDS, JUSTIN  ALIENS AND ENEMIES [DOCTOR WHO]  R/REF-H
RICHARDS, JUSTIN  BURNING, THE [DOCTOR WHO EIGHTH DOCTOR]  P
RICHARDS, JUSTIN  CLOCKWISE MAN, THE [DOCTOR WHO NINTH DOCTOR]  H
RICHARDS, JUSTIN  CREATURES AND DEMONS [DOCTOR WHO]  R/REF-H
RICHARDS, JUSTIN  DEMONTAGE [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
RICHARDS, JUSTIN  DEVIANT STRAIN, THE [DOCTOR WHO NINTH DOCTOR]  H
RICHARDS, JUSTIN  DOCTOR WHO: THE LEGEND CONTINUES [DOCTOR WHO]  R/REF-H
RICHARDS, JUSTIN  DRAGONS’ WRATH [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
RICHARDS, JUSTIN  DREAMS OF EMPIRE [DOCTOR WHO PAST DOCTORS]  P
RICHARDS, JUSTIN  GRAVE MATTER [DOCTOR WHO PAST DOCTORS]  P
RICHARDS, JUSTIN  JOY DEVICE, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  P
RICHARDS, JUSTIN  MEDUSA EFFECT, THE [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  P,PA
RICHARDS, JUSTIN  MILLENNIUM SHOCK [DOCTOR WHO PAST DOCTORS]  P
RICHARDS, JUSTIN  OPTION LOCK [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P

RICHARDS, JUSTIN PROFESSOR BERNICE SUMMERFIELD AND THE DOOMSDAY MANUSCRIPT

[BERNICE SUMMERFIELD 1]

P

RICHARDS, JUSTIN  RESURRECTION CASKET, THE [DOCTOR WHO TENTH DOCTOR]  H
RICHARDS, JUSTIN  SANDS OF TIME, THE [DOCTOR WHO MISSING ADVENTURES]  C/P,P
RICHARDS, JUSTIN/COLE, STEPHEN  SHADOW IN THE GLASS, THE [DOCTOR WHO PAST DOCTORS]  P
RICHARDS, JUSTIN  SOMETIME NEVER... [DOCTOR WHO EIGHTH DOCTOR]  P
RICHARDS, JUSTIN  SYSTEM SHOCK [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P

RICHARDS, JUSTIN TEARS OF THE ORACLE

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

P

RICHARDS, JUSTIN  THEATRE OF WAR [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
RICHARDS, JUSTIN  TIME ZERO [DOCTOR WHO EIGHTH DOCTOR]  P
RICHARDS, LEIGH  CALIFIA’S DAUGHTERS  P
RICHARDS, PAUL  OUR SPACE CRAFT IS MISSING!  P
RICHARDS, PAUL  UNBLESSED, THE  P
RICHARDS, R. P. J.  BLONDE GODDESS, THE  H
RICHARDS, TONY  HARVEST BRIDE, THE  SFWA-P
RICHARDSON, KIT  GREYWALKER  LP
RICHARDSON, ROBERT  SECOND SATELLITE  H
RICHMOND, WALT/RICHMOND, LEIGH  CHALLENGE THE HELLMAKER  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 321
AUTHOR  TITLE  CODE
RICHMOND, WALT/RICHMOND, LEIGH  GALLAGHER’S GLACIER  P,D27235
RICHMOND, WALT/RICHMOND, LEIGH  LOST MILLENNIUM, THE  DH029
RICHMOND, WALT/RICHMOND, LEIGH  PHASE TWO  C/P,P
RICHMOND, WALT/RICHMOND, LEIGH  PHOENIX SHIP  D66160
RICHMOND, WALT/RICHMOND, LEIGH  POSITIVE CHARGE  D27235
RICHMOND, WALT/RICHMOND, LEIGH  PROBABILITY CORNER, THE  C/P:2,P
RICHMOND, WALT/RICHMOND, LEIGH  SHOCK WAVE  DG614
RICHMOND, WALT/RICHMOND, LEIGH  SIVA!  C/P,P
RICKERT, M.  MAP OF DREAMS  H
RIDDELL, J. H.  WEIRD STORIES  H
RIDDLE, CHARLES LEE  1953 CHECKDEX, THE  SRL/IND
RIDGEWAY, JAMES  CLOSED CORPORATION, THE  [C/RAN-P]
RIDLEY, JOHN  THOSE WHO WALK IN DARKNESS  H
RIDLEY, R. W.  DELON CITY [@OZ CHRONICLES #2]  LP
RIDLEY, R. W.  TAKERS, THE [@OZ CHRONICLES #1]  LP
RIDPATH, IAN  WORLDS BEYOND  SFWA-VLP
RIEDMANN, AGNES  STORY OF ADAMSVILLE, THE  LP
RIENOW, LEONA  BEWITCHED CAVERNS, THE  H
RIENOW, LEONA TRAIN/RIENOW, ROBERT  YEAR OF THE LAST EAGLE, THE  C/P,P
RILEY, LOUISE  SPELL AT SCOGGIN’S CROSSING, A  H
RIMEL, DUANE W.  TO YITH AND BEYOND  SR-VLH
RIMMER, ROBERT H.  HARRAD EXPERIMENT, THE  C/P,P
RIMMER, ROBERT H.  ZOLOTOV AFFAIR, THE  C/P,P
RIMMER, STEVEN WILLIAM  COVEN  P
RINEHART, MARY ROBERTS  RED LAMP, THE  H,P
RINGO, JOHN  AGAINST THE TIDE [DRAGONS 3]  H
RINGO, JOHN/COCHRANE, JULIE  CALLY’S WAR  H
RINGO, JOHN  EAST OF THE SUN, WEST OF THE MOON [DRAGONS 4]  H
RINGO, JOHN  EMERALD SEA [DRAGONS 2]  H
RINGO, JOHN  GUST FRONT [POSLEEN 2]  P
RINGO, JOHN  HELL’S FAIRE (HARDCOVER AND CD) [POSLEEN 4]  H:2
RINGO, JOHN  HELL’S FAIRE (PROOF) [POSLEEN 4]  H,SR-LP
RINGO, JOHN/WILLIAMSON, MICHAEL Z.  HERO, THE  H
RINGO, JOHN  HYMN BEFORE BATTLE, A [POSLEEN 1]  C/P,P
RINGO, JOHN  INTO THE LOOKING GLASS [LOOKING GLASS 1]  H
RINGO, JOHN  PRINCESS OF WANDS  H
RINGO, JOHN/EVANS, LINDA  ROAD TO DAMASCUS, THE [BOLOS]  H
RINGO, JOHN/EVANS, LINDA  ROAD TO DAMASCUS, THE (PROOF) [BOLOS]  SR-LP
RINGO, JOHN  THERE WILL BE DRAGONS (HARDCOVER AND CD) [DRAGONS 1]  H:2
RINGO, JOHN  THERE WILL BE DRAGONS (PROOF) [DRAGONS 1]  SR-LP
RINGO, JOHN/TAYLOR, TRAVIS S.  VON NEUMANN’S WAR  H
RINGO, JOHN/KRATMAN, TOM  WATCH ON THE RHINE [POSLEEN]  H
RINGO, JOHN  WHEN THE DEVIL DANCES [POSLEEN 3]  H
RINGO, JOHN/KRATMAN, TOM  YELLOW EYES [POSLEEN]  H
RIORDAN, RICK  LIGHTNING THIEF, THE [@PERCY JACKSON & THE OLYMPIANS #1]  H
RIOS, TERE  FIFTEENTH PELICAN, THE  P
RIPLEY, KAREN  ALCHEMIST OF TIME, THE [@SLOW WORLD #3]  P
RIPLEY, KAREN  PERSISTENCE OF MEMORY, THE [@SLOW WORLD #1]  P
RIPLEY, KAREN  PRISONER OF DREAMS  C/P,P
RIPLEY, KAREN  TENTH CLASS, THE  C/P:2,P
RIPLEY, KAREN  WARDEN OF HORSES, THE [@SLOW WORLD #2]  P
RITCHIE, RITA  GRIP OF FEAR  P
RIVERE, ALEC  LOST CITY OF THE DAMNED  SR-L
RIVES, AMELIE  GHOST GARDEN, THE  H
RIVKIN, J. F.  DREAMSTONE, THE [RUNESWORD 3]  C/P,P
RIVKIN, J. F.  MISTRESS OF AMBIGUITIES [SILVERGLASS 4]  P
RIVKIN, J. F.  SILVERGLASS [SILVERGLASS 1]  C/P:2,P
RIVKIN, J. F.  TYRANNOSAURUS REX [AGE OF DINOSAURS 1]  P,SFWA-P
RIVKIN, J. F.  WEB OF WIND [SILVERGLASS 2]  P
RIVKIN, J. F.  WITCH OF RHOSTSHYL [SILVERGLASS 3]  C/P,P
ROBB, J. D.  BORN IN DEATH [EVE DALLAS]  H
ROBB, J. D.  DIVIDED IN DEATH [EVE DALLAS]  H
ROBB, J. D.  IMMORTAL IN DEATH [EVE DALLAS]  P
ROBB, J. D.  JUDGMENT IN DEATH [EVE DALLAS]  P
ROBB, J. D.  ORIGIN IN DEATH [EVE DALLAS]  H

ROBB, J. D./HAMILTON, LAURELL K./ KRINARD, SUSAN/SHAYNE, MAGGIE OUT OF THIS WORLD [EVE DALLAS]

C/P,P

ROBB, J. D.  PURITY IN DEATH [EVE DALLAS]  P
ROBB, J. D.  RAPTURE IN DEATH [EVE DALLAS]  P
ROBB, J. D.  SURVIVOR IN DEATH [EVE DALLAS]  H
ROBB, J. D.  VISIONS IN DEATH [EVE DALLAS]  H
ROBB, JOHN  SPACE BEAM  SR-P
ROBBINS, DAVID  FOX RUN, THE [@ENDWORLD 1]  P
ROBBINS, DAVID  TWIN CITIES RUN [@ENDWORLD #3]  P

AUTHOR  TITLE  CODE
ROBBINS, TOM  JITTERBUG PERFUME  H,P
ROBBINS, TRINA  ETERNALLY BAD  H
ROBERSON, CHRIS  HERE, THERE & EVERYWHERE  H
ROBERSON, CHRIS  PARAGAEA  LP
ROBERSON, JENNIFER  CHILDREN OF THE LION [@CHEYSULI #5,6]  P
ROBERSON, JENNIFER  DAUGHTER OF THE LION [CHEYSULI 6]  C/P,P
ROBERSON, JENNIFER  DEEPWOOD [KARAVANS 2]  H
ROBERSON, JENNIFER  FLIGHT OF THE RAVEN [CHEYSULI 7]  C/P:2,P
ROBERSON, JENNIFER  HIGHWAYMEN: ROBBERS & ROGUES  C/PA,PA
ROBERSON, JENNIFER  KARAVANS [KARAVANS 1]  H
ROBERSON, JENNIFER  LADY OF SHERWOOD  LP
ROBERSON, JENNIFER  LADY OF THE FOREST  P
ROBERSON, JENNIFER  LEGACY OF THE SWORD [CHEYSULI 3]  C/P:2,P
ROBERSON, JENNIFER  LEGACY OF THE WOLF [@CHEYSULI #3,4]  P
ROBERSON, JENNIFER  LION THRONE, THE [@CHEYSULI #7,8]  P
ROBERSON, JENNIFER  OUT OF AVALON  PA
ROBERSON, JENNIFER  PRIDE OF PRINCES, A [CHEYSULI 5]  C/P:2,P
ROBERSON, JENNIFER  RETURN TO AVALON  C/PA,PA
ROBERSON, JENNIFER  SCOTLAND THE BRAVE [@HIGHLANDER]  P
ROBERSON, JENNIFER  SHAPECHANGERS [CHEYSULI 1]  C/P:2,P
ROBERSON, JENNIFER  SHAPECHANGER’S SONG (INCLUDES "KINSPIRIT") [@CHEYSULI #1,2,3]  P
ROBERSON, JENNIFER  SONG OF HOMANA, THE [CHEYSULI 2]  C/P:2,P
ROBERSON, JENNIFER  SWORD-BORN [TIGER 5]  P
ROBERSON, JENNIFER  SWORD-BREAKER [TIGER 4]  C/P:2,SFWA-P:2,P
ROBERSON, JENNIFER  SWORD-DANCER [TIGER 1]  C/P:2,P
ROBERSON, JENNIFER  SWORD-MAKER [TIGER 3]  C/P:2,P
ROBERSON, JENNIFER  SWORD-SINGER [TIGER 2]  C/P:2,P
ROBERSON, JENNIFER  SWORD-SWORN [TIGER 6]  H,P
ROBERSON, JENNIFER  TAPESTRY OF LIONS, A [CHEYSULI 8]  C/P,P
ROBERSON, JENNIFER  TRACK OF THE WHITE WOLF [CHEYSULI 4]  C/P,P
ROBERTS, A. R. R. R.  KLEINE HOBBNIX, DER  FP
ROBERTS, A3R  STAR WARPED  H
ROBERTS, ADAM  GRADISIL  LP
ROBERTS, ADAM  JUPITER MAGNIFIED  SR-LP
ROBERTS, ADAM  ON  H,LP
ROBERTS, ADAM  POLYSTOM  H
ROBERTS, ADAM  SALT  H,LP
ROBERTS, ADAM  SNOW, THE  LP
ROBERTS, ADAM  STONE  C/P,P,H
ROBERTS, BETTE B.  ANNE RICE  R/REF-H
ROBERTS, BURGERT  SPACEWALKS  H
ROBERTS, DOROTHY JAMES  ENCHANTED CUP, THE  H
ROBERTS, DOROTHY JAMES  RETURN OF THE STRANGER  H
ROBERTS, GARETH  ENGLISH WAY OF DEATH, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
ROBERTS, GARETH  HIGHEST SCIENCE, THE [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
ROBERTS, GARETH  I AM A DALEK [DOCTOR WHO]  LP
ROBERTS, GARETH  ONLY HUMAN [DOCTOR WHO NINTH DOCTOR]  H
ROBERTS, GARETH  PLOTTERS, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
ROBERTS, GARETH  ROMANCE OF CRIME, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
ROBERTS, GARETH  TRAGEDY DAY [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
ROBERTS, GARETH  WELL-MANNERED WAR, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
ROBERTS, GARETH  ZAMPER [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
ROBERTS, JANE  EDUCATION OF OVERSOUL 7, THE  C/P,P,H
ROBERTS, JANE  REBELLERS, THE  DF215
ROBERTS, JOHN MADDOX  BLACK SHIELDS, THE [STORMLANDS 2]  C/P:2,P
ROBERTS, JOHN MADDOX  CESTUS DEI  P
ROBERTS, JOHN MADDOX  CLOAK OF ILLUSION [CINGULUM 2]  P,SFWA-P
ROBERTS, JOHN MADDOX  CONAN AND THE AMAZON [CONAN]  P
ROBERTS, JOHN MADDOX  CONAN AND THE MANHUNTERS [CONAN]  P
ROBERTS, JOHN MADDOX  CONAN AND THE TREASURE OF PYTHON [CONAN]  LP
ROBERTS, JOHN MADDOX  CONAN THE BOLD [CONAN]  P,SFWA-P
ROBERTS, JOHN MADDOX  CONAN THE CHAMPION [CONAN]  P
ROBERTS, JOHN MADDOX  CONAN THE MARAUDER [CONAN]  P,SFWA-P
ROBERTS, JOHN MADDOX  CONAN THE ROGUE [CONAN]  H,LP
ROBERTS, JOHN MADDOX  CONAN THE VALOROUS [CONAN]  SFWA,LP
ROBERTS, JOHN MADDOX  ENIGMA VARIATIONS, THE  P
ROBERTS, JOHN MADDOX  ISLANDER, THE [STORMLANDS 1]  C/P:2,SFWA-P,P
ROBERTS, JOHN MADDOX  KING OF THE WOOD  H,SFWA,SFWA-P,P
ROBERTS, JOHN MADDOX  MURDER IN TARSIS [@DRAGONLANCE CLASSICS]  P
ROBERTS, JOHN MADDOX  POISONED LANDS, THE [STORMLANDS 3]  C/P,P
ROBERTS, JOHN MADDOX  QUEENS OF LAND AND SEA [STORMLANDS 5]  P
ROBERTS, JOHN MADDOX  SPACE ANGEL [SPACE ANGEL 1]  C/P:2,P
ROBERTS, JOHN MADDOX  SPACER: WINDOW OF THE MIND [SPACE ANGEL 2]  P
ROBERTS, JOHN MADDOX  STEEL KINGS, THE [STORMLANDS 4]  P

The MITSFS Pinkdex  September 24, 2007  Page 323
AUTHOR  TITLE  CODE
ROBERTS, JOHN MADDOX  STRAYED SHEEP OF CHARUN, THE  H,SFWA
ROBERTS, JOHN MADDOX  SWORD, THE JEWEL AND THE MIRROR, THE [CINGULUM 3]  P,SFWA-P
ROBERTS, KEITH  ANITA  P
ROBERTS, KEITH  CHALK GIANTS, THE  C/P,P
ROBERTS, KEITH  FURIES, THE  C/P:2,P
ROBERTS, KEITH  GRAIN KINGS, THE  P
ROBERTS, KEITH  INNER WHEEL, THE  H,P
ROBERTS, KEITH  KAETI & COMPANY  LP
ROBERTS, KEITH  KITEWORLD  C/P,SFWA,P,H
ROBERTS, KEITH  MACHINES AND MEN  SR-P
ROBERTS, KEITH  MOLLY ZERO  SR-P
ROBERTS, KEITH  PASSING OF THE DRAGONS, THE  C/P,P
ROBERTS, KEITH  PAVANE  C/P:2,P
ROBERTS, KERRY DANIEL  DAGGER OF DOOM, THE [@KNIGHTS OF THE SILVER DRAGON #6]  LP
ROBERTS, KERRY DANIELS  MARK OF THE YUAN-TI [@KNIGHTS OF THE SILVER DRAGON #12]  LP
ROBERTS, LIONEL  FACE OF X, THE  SR-P
ROBERTS, LIONEL  IN-WORLD, THE  SR-P
ROBERTS, MURRAY  EARTHQUAKE-MAKER, THE  B
ROBERTS, MURRAY  SECRET KINGDOM, THE  B
ROBERTS, MURRAY  WORLD IN DARKNESS, THE  B
ROBERTS, NORA  DANCE OF THE GODS [CIRCLE 2]  H
ROBERTS, NORA  MORRIGAN’S CROSS [CIRCLE 1]  H
ROBERTS, NORA  VALLEY OF SILENCE [CIRCLE 3]  H
ROBERTS, PETER  COROBITE MINES, THE  P
ROBERTS, STEVE  MAX HEADROOM: 20 MINUTES INTO THE FUTURE  VLP
ROBERTS, THOM  MAGICAL MIND ADVENTURE OF HANNAH AND COLDY COLDY, THE  H
ROBERTS, WESS/ROSS, BILL  MAKE IT SO [STAR TREK TNG]  H
ROBERTS, WILLO DAVIS  GHOSTS OF HARREL, THE  P
ROBERTS, WILLO DAVIS  GIRL WITH THE SILVER EYES, THE  P
ROBERTSON, MORGAN  OVER THE BORDER  H
ROBERTSON, R. GARCIA  MOON MAID, THE  H
ROBESON, KENNETH  ANGRY GHOST, THE [DOC SAVAGE #86]  @P
ROBESON, KENNETH  ANNIHILIST [DOC SAVAGE #31]  C/P:2,@P
ROBESON, KENNETH  AWFUL EGG, THE [DOC SAVAGE #92]  @P
ROBESON, KENNETH  BLACK CHARIOTS [@AVENGER #30]  C/P:2,P
ROBESON, KENNETH  BLACK DEATH, THE [@AVENGER #22]  C/P,P
ROBESON, KENNETH  BLACK SPOT, THE [DOC SAVAGE #76]  C/P:2,@P
ROBESON, KENNETH  BLOOD COUNTESS, THE [@AVENGER #33]  C/P:2,P
ROBESON, KENNETH  BLOOD RING, THE [@AVENGER #6]  C/P:2,P
ROBESON, KENNETH  BOSS OF TERROR, THE [DOC SAVAGE #85]  @P
ROBESON, KENNETH  BRAND OF THE WEREWOLF [DOC SAVAGE #5]  C/P:2,@P
ROBESON, KENNETH  CARTOON CRIMES, THE [@AVENGER #31]  C/P,P
ROBESON, KENNETH  COLD DEATH [DOC SAVAGE #21]  C/P,@P
ROBESON, KENNETH  CRIMSON SERPENT, THE [DOC SAVAGE #78]  C/P:2,@P
ROBESON, KENNETH  CZAR OF FEAR, THE [DOC SAVAGE #22]  C/P:2,@P
ROBESON, KENNETH  DAGGER IN THE SKY, THE [DOC SAVAGE #40]  C/P:2,@P
ROBESON, KENNETH  DEADLY DWARF, THE [DOC SAVAGE #28]  C/P:2,@P
ROBESON, KENNETH  DEATH IN SILVER [DOC SAVAGE #26]  C/P:2,@P
ROBESON, KENNETH  DEATH IN SLOW MOTION [@AVENGER #18]  C/P:2,P
ROBESON, KENNETH  DEATH MACHINE, THE [@AVENGER #32]  P
ROBESON, KENNETH  DEMON ISLAND [@AVENGER #36]  C/P:2,P
ROBESON, KENNETH  DERRICK DEVIL, THE [DOC SAVAGE #74]  C/P:2,@P
ROBESON, KENNETH  DEVIL GENGHIS, THE [DOC SAVAGE #79]  C/P:2,@P
ROBESON, KENNETH  DEVIL ON THE MOON [DOC SAVAGE #50]  C/P:2,@P
ROBESON, KENNETH  DEVIL’S HORNS, THE [@AVENGER #4]  C/P:2,P

ROBESON, KENNETH DEVILS OF THE DEEP [DOC SAVAGE #123]/ HEADLESS MEN, THE [DOC SAVAGE 124]

@P

ROBESON, KENNETH  DEVIL’S PLAYGROUND, THE [DOC SAVAGE #25]  C/P:2,@P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 1 [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 3 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 4 [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 5 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 6 [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 7 [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 8 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 9 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 10 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 11 [DOC SAVAGE]  P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 12 [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  DOC SAVAGE OMNIBUS 13 [DOC SAVAGE]  P
ROBESON, KENNETH  DOCTOR TIME [@AVENGER #28]  C/P,P
ROBESON, KENNETH  DUST OF DEATH [DOC SAVAGE #32]  C/P:2,@P
ROBESON, KENNETH  EVIL GNOME, THE [DOC SAVAGE #82]  C/P,@P
ROBESON, KENNETH  FANTASTIC ISLAND, THE [DOC SAVAGE #14]  C/P,@P

The MITSFS Pinkdex  September 24, 2007  Page 324
AUTHOR  TITLE  CODE
ROBESON, KENNETH  FEAR CAY [DOC SAVAGE #11]  C/P:2,@P
ROBESON, KENNETH  FEATHERED OCTOPUS, THE [DOC SAVAGE #48]  C/P:2,@P
ROBESON, KENNETH  FLAME BREATHERS, THE [@AVENGER #12]  C/P:2,P
ROBESON, KENNETH  FLAMING FALCONS, THE [DOC SAVAGE #30]  C/P:2,@P
ROBESON, KENNETH  FLIGHT INTO FEAR [DOC SAVAGE]  P
ROBESON, KENNETH  FLYING GOBLIN, THE [DOC SAVAGE #90]  C/P,@P
ROBESON, KENNETH  FORGOTTEN REALM, THE [DOC SAVAGE]  P
ROBESON, KENNETH  FORTRESS OF SOLITUDE [DOC SAVAGE #23]  @P
ROBESON, KENNETH  FRECKLED SHARK, THE [DOC SAVAGE #67]  C/P:2,@P
ROBESON, KENNETH  FRIGHTENED FISH, THE [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  FROSTED DEATH, THE [@AVENGER #5]  C/P:2,P
ROBESON, KENNETH  GHOST LEGION, THE [DOC SAVAGE 3]  H
ROBESON, KENNETH  GIGGLING GHOSTS, THE [DOC SAVAGE #56]  C/P:2,@P
ROBESON, KENNETH  GLASS MAN, THE [@AVENGER #34]  C/P,P
ROBESON, KENNETH  GLASS MOUNTAIN, THE [@AVENGER #8]  C/P:2,P

ROBESON, KENNETH GOBLINS, THE [DOC SAVAGE #125]/

SECRET OF THE SU, THE [DOC SAVAGE 126]

@P

ROBESON, KENNETH  GOLD OGRE, THE [DOC SAVAGE #42]  C/P,@P

ROBESON, KENNETH GOLDEN MAN, THE [DOC SAVAGE #117]/ PERIL IN THE NORTH [DOC SAVAGE 118]

@P

ROBESON, KENNETH  GOLDEN PERIL, THE [DOC SAVAGE #55]  C/P,@P
ROBESON, KENNETH  GREEN DEATH, THE [DOC SAVAGE #65]  C/P:2,@P
ROBESON, KENNETH  GREEN EAGLE, THE [DOC SAVAGE #24]  C/P,@P
ROBESON, KENNETH  GREEN KILLER, THE [@AVENGER #20]  C/P,P
ROBESON, KENNETH  HAPPY KILLERS, THE [@AVENGER #21]  C/P:2,P
ROBESON, KENNETH  HATE GENIUS, THE [DOC SAVAGE #94]  @P
ROBESON, KENNETH  HATE MASTER, THE [@AVENGER #16]  C/P:2,P
ROBESON, KENNETH  HAUNTED OCEAN [DOC SAVAGE #51]  C/P,@P
ROBESON, KENNETH  HE COULD STOP THE WORLD [DOC SAVAGE #54]  C/P,@P
ROBESON, KENNETH  HELL BELOW [DOC SAVAGE #99]/LOST GIANT, THE [DOC SAVAGE 100]  C/P,@P
ROBESON, KENNETH  HEX [DOC SAVAGE #37]  C/P:2,@P
ROBESON, KENNETH  HOUSE OF DEATH [@AVENGER #15]  C/P,P
ROBESON, KENNETH  INCREDIBLE RADIO EXPLOITS OF DOC SAVAGE 1, THE  VLP
ROBESON, KENNETH  IRON SKULL, THE [@AVENGER #35]  P
ROBESON, KENNETH  JADE OGRE, THE [DOC SAVAGE]  C/P:2,P
ROBESON, KENNETH  JIU SAN [DOC SAVAGE #107]/BLACK, BLACK WITCH, THE [DOC SAVAGE 108] @P
ROBESON, KENNETH  JUSTICE INC. [@AVENGER #1]  C/P:2,P
ROBESON, KENNETH  KING MAKER, THE [DOC SAVAGE #80]  C/P:2,@P
ROBESON, KENNETH  LAND OF ALWAYS-NIGHT [DOC SAVAGE #13]  C/P:2,@P
ROBESON, KENNETH  LAND OF FEAR, THE [DOC SAVAGE #75]  C/P:2,@P
ROBESON, KENNETH  LAND OF LONG JUJU [DOC SAVAGE #47]  C/P,@P
ROBESON, KENNETH  LAND OF TERROR, THE [DOC SAVAGE #8]  C/P,@P

ROBESON, KENNETH LAUGH OF DEATH, THE [DOC SAVAGE #119]/ KING OF TERROR, THE [DOC SAVAGE 120]

@P

ROBESON, KENNETH  LIVING FIRE MENACE, THE [DOC SAVAGE #61]  C/P:2,@P
ROBESON, KENNETH  LOST OASIS, THE [DOC SAVAGE #6]  C/P:2,@P
ROBESON, KENNETH  MAD EYES [DOC SAVAGE #34]  C/P,@P
ROBESON, KENNETH  MAD MESA [DOC SAVAGE #66]  C/P:2,@P
ROBESON, KENNETH  MAGIC ISLAND, THE [DOC SAVAGE #89]  C/P,@P
ROBESON, KENNETH  MAJII, THE [DOC SAVAGE #60]  C/P,@P
ROBESON, KENNETH  MAN FROM ATLANTIS, THE [@AVENGER #25]  C/P:2,P
ROBESON, KENNETH  MAN OF BRONZE, THE [DOC SAVAGE #1]  C/P:2,@P
ROBESON, KENNETH  MAN WHO SHOOK THE EARTH, THE [DOC SAVAGE #43]  C/P:2,@P
ROBESON, KENNETH  MEN WHO SMILED NO MORE, THE [DOC SAVAGE #45]  C/P:2,@P
ROBESON, KENNETH  MENTAL WIZARD, THE [DOC SAVAGE #53]  C/P,@P
ROBESON, KENNETH  MERCHANTS OF DISASTER [DOC SAVAGE #41]  C/P:2,@P
ROBESON, KENNETH  METAL MASTER, THE [DOC SAVAGE #72]  C/P:2,@P
ROBESON, KENNETH  METEOR MENACE [DOC SAVAGE #3]  C/P:2,@P
ROBESON, KENNETH  MIDAS MAN, THE [DOC SAVAGE #46]  C/P:2,@P
ROBESON, KENNETH  MIDNIGHT MURDER [@AVENGER #24]  C/P:2,P
ROBESON, KENNETH  MONSTERS, THE [DOC SAVAGE #7]  C/P:2,@P
ROBESON, KENNETH  MOTION MENACE [DOC SAVAGE #64]  C/P,@P
ROBESON, KENNETH  MOUNTAIN MONSTER, THE [DOC SAVAGE #84]  C/P:2,@P
ROBESON, KENNETH  MUNITIONS MASTER, THE [DOC SAVAGE #58]  C/P:2,@P
ROBESON, KENNETH  MURDER MELODY [DOC SAVAGE #15]  C/P:2,@P
ROBESON, KENNETH  MURDER MIRAGE [DOC SAVAGE #71]  C/P:2,@P
ROBESON, KENNETH  MURDER ON WHEELS [@AVENGER #13]  C/P:2,P
ROBESON, KENNETH  MYSTERY ON HAPPY BONES, THE [DOC SAVAGE #96]  @P
ROBESON, KENNETH  MYSTERY ON THE SNOW, THE [DOC SAVAGE #69]  C/P:2,@P
ROBESON, KENNETH  MYSTERY UNDER THE SEA [DOC SAVAGE #27]  C/P,@P
ROBESON, KENNETH  MYSTIC MULLAH, THE [DOC SAVAGE #9]  @P
ROBESON, KENNETH  NEVLO [@AVENGER #17]  C/P,P
ROBESON, KENNETH  NIGHTWITCH DEVIL, THE [@AVENGER #29]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 325
AUTHOR  TITLE  CODE

ROBESON, KENNETH ONE-EYED MYSTIC [DOC SAVAGE #111]/

MAN WHO FELL UP, THE [DOC SAVAGE 112]

@P

ROBESON, KENNETH  OTHER WORLD, THE [DOC SAVAGE #29]  @P
ROBESON, KENNETH  PHANTOM CITY, THE [DOC SAVAGE #10]  C/P,@P

ROBESON, KENNETH PHAROAH’S GHOST, THE [DOC SAVAGE #101]/ TIME TERROR, THE [DOC SAVAGE 102]

@P

ROBESON, KENNETH  PICTURES OF DEATH [@AVENGER #19]  C/P:2,P
ROBESON, KENNETH  PIRATE ISLE [DOC SAVAGE #115]/SPEAKING STONE, THE [DOC SAVAGE 116] @P
ROBESON, KENNETH  PIRATE OF THE PACIFIC [DOC SAVAGE #19]  C/P,@P
ROBESON, KENNETH  PIRATE’S GHOST, THE [DOC SAVAGE #62]  C/P:2,@P
ROBESON, KENNETH  POISON ISLAND [DOC SAVAGE #57]  C/P,@P
ROBESON, KENNETH  POLAR TREASURE, THE [DOC SAVAGE #4]  C/P,@P
ROBESON, KENNETH  PURPLE DRAGON, THE [DOC SAVAGE #91]  @P
ROBESON, KENNETH  PURPLE ZOMBIE, THE [@AVENGER #27]  C/P:2,P
ROBESON, KENNETH  PYTHON ISLE [DOC SAVAGE]  C/P:2,P
ROBESON, KENNETH  QUEST OF QUI [DOC SAVAGE #12]  C/P:2,@P
ROBESON, KENNETH  QUEST OF THE SPIDER [DOC SAVAGE #68]  C/P:2,@P
ROBESON, KENNETH  RED MOON [@AVENGER #26]  C/P:2,P
ROBESON, KENNETH  RED SKULL, THE [DOC SAVAGE #17]  C/P:2,@P
ROBESON, KENNETH  RED SNOW [DOC SAVAGE #38]  C/P,@P
ROBESON, KENNETH  RED SPIDER, THE [DOC SAVAGE #95]  @P
ROBESON, KENNETH  RED TERRORS, THE [DOC SAVAGE #83]  C/P,@P
ROBESON, KENNETH  RESURRECTION DAY [DOC SAVAGE #36]  C/P:2,@P
ROBESON, KENNETH  RIVER OF ICE [@AVENGER #11]  C/P:2,P
ROBESON, KENNETH  ROAR DEVIL, THE [DOC SAVAGE #88]  C/P,@P
ROBESON, KENNETH  SARGASSO OGRE, THE [DOC SAVAGE #18]  H,@P
ROBESON, KENNETH  SATAN BLACK [DOC SAVAGE #97]/CARGO UNKNOWN [DOC SAVAGE 98]  @P
ROBESON, KENNETH  SEA ANGEL, THE [DOC SAVAGE #49]  C/P:2,@P
ROBESON, KENNETH  SEA MAGICIAN, THE [DOC SAVAGE #44]  C/P:2,@P
ROBESON, KENNETH  SECRET IN THE SKY, THE [DOC SAVAGE #20]  C/P:2,@P
ROBESON, KENNETH  SEVEN AGATE DEVILS, THE [DOC SAVAGE #73]  C/P:2,@P

ROBESON, KENNETH SHAPE OF TERROR, THE [DOC SAVAGE #109]/

DEATH HAD YELLOW EYES [DOC SAVAGE 110]

@P

ROBESON, KENNETH  SKY WALKER, THE [@AVENGER #3]  C/P:2,P
ROBESON, KENNETH  SMILING DOGS, THE [@AVENGER #10]  C/P:2,P
ROBESON, KENNETH  SOUTH POLE TERROR, THE [DOC SAVAGE #77]  C/P,@P
ROBESON, KENNETH  SPOOK HOLE [DOC SAVAGE #70]  C/P:2,@P
ROBESON, KENNETH  SPOOK LEGION, THE [DOC SAVAGE #16]  C/P:2,@P
ROBESON, KENNETH  SPOTTED MEN, THE [DOC SAVAGE #87]  @P
ROBESON, KENNETH  SQUEAKING GOBLIN, THE [DOC SAVAGE #35]  C/P:2,@P
ROBESON, KENNETH  STOCKHOLDERS IN DEATH [@AVENGER #7]  C/P:2,P
ROBESON, KENNETH  STONE MAN, THE [DOC SAVAGE #81]  @P
ROBESON, KENNETH  SUBMARINE MYSTERY, THE [DOC SAVAGE #63]  C/P:2,@P

ROBESON, KENNETH TALKING DEVIL, THE [DOC SAVAGE #113]/ TEN TON SNAKES, THE [DOC SAVAGE 114]

@P

ROBESON, KENNETH  TERROR IN THE NAVY, THE [DOC SAVAGE #33]  C/P:2,@P

ROBESON, KENNETH THEY DIED TWICE [DOC SAVAGE #105]/

SCREAMING MAN, THE [DOC SAVAGE 106]

@P

ROBESON, KENNETH  THOUSAND-HEADED MAN, THE [DOC SAVAGE #2]  C/P:2,@P
ROBESON, KENNETH  THREE GOLD CROWNS [@AVENGER #14]  C/P:2,P

ROBESON, KENNETH THREE WILD MEN, THE [DOC SAVAGE #121]/ FIERY MENACE, THE [DOC SAVAGE 122]

@P

ROBESON, KENNETH  TUNED FOR MURDER [@AVENGER #9]  C/P:2,P
ROBESON, KENNETH  TUNNEL TERROR [DOC SAVAGE #93]  @P
ROBESON, KENNETH  VANISHER, THE [DOC SAVAGE #52]  C/P:2,@P

ROBESON, KENNETH WHISKER OF HERCULES, THE [DOC SAVAGE #103]/ MAN WHO WAS SCARED, THE [DOC SAVAGE 104]

@P

ROBESON, KENNETH  WHISTLING WRAITH, THE [DOC SAVAGE]  C/P,P
ROBESON, KENNETH  WHITE EYES [DOC SAVAGE]  P
ROBESON, KENNETH  WILDER CURSE, THE [@AVENGER #23]  C/P,P
ROBESON, KENNETH  WORLD’S FAIR GOBLIN [DOC SAVAGE #39]  C/P,@P
ROBESON, KENNETH  YELLOW CLOUD, THE [DOC SAVAGE #59]  C/P:2,@P
ROBESON, KENNETH  YELLOW HOARD, THE [@AVENGER #2]  C/P:2,P
ROBINETT, STEPHEN  MAN RESPONSIBLE, THE  C/P:2,P
ROBINETT, STEPHEN  PROJECTIONS  C/P,P
ROBINETT, STEPHEN  STARGATE  C/P,P
ROBINS, LANE  MALEDICTE  LP
ROBINS, MADELEINE E.  STONE WAR, THE  C/P,P
ROBINSON, ANDREW J.  STITCH IN TIME, A [STAR TREK DS9]  C/P,P
ROBINSON, ELEANOR  CHRYSALIS OF DEATH  P
ROBINSON, ELEANOR  SILVERLEAF SYNDROME, THE  P
ROBINSON, FRANK M.  DARK BEYOND THE STARS, THE  H:2,P
ROBINSON, FRANK M.  LIFE IN THE DAY OF...AND OTHER SHORT STORIES, A  P

The MITSFS Pinkdex  September 24, 2007  Page 326
AUTHOR  TITLE  CODE
ROBINSON, FRANK M.  POWER, THE  H:2,P
ROBINSON, FRANK S.  CHILDREN OF THE DRAGON  P
ROBINSON, G. P.  DEBT, THE  H
ROBINSON, JEREMY  RAISING THE PAST  LP
ROBINSON, KIM STANLEY  ANARCTICA  P
ROBINSON, KIM STANLEY  BLIND GEOMETER, THE  SFWA-TD13,TD13
ROBINSON, KIM STANLEY  BLUE MARS [MARS 3]  C/P,P,H
ROBINSON, KIM STANLEY  ESCAPE FROM KATHMANDU  C/P,SFWA-P,P,H
ROBINSON, KIM STANLEY  FIFTY DEGREES BELOW [FORTY SIGNS OF RAIN 2]  H:2
ROBINSON, KIM STANLEY  FORTY SIGNS OF RAIN [FORTY SIGNS OF RAIN 1]  P
ROBINSON, KIM STANLEY  FUTURE PRIMITIVE  LPA
ROBINSON, KIM STANLEY  GOLD COAST, THE [THREE CALIFORNIAS]  H,SFWA-P,P
ROBINSON, KIM STANLEY  GREEN MARS [MARS 2]  H:2,LP
ROBINSON, KIM STANLEY  GREEN MARS (NOVELLA) [MARS]  TD1
ROBINSON, KIM STANLEY  ICEHENGE  C/P:2,SFWA-P,LP
ROBINSON, KIM STANLEY  MARTIANS, THE [MARS]  H
ROBINSON, KIM STANLEY  MEMORY OF WHITENESS, THE  SFWA-P,H:2,P,SFWA
ROBINSON, KIM STANLEY  NEBULA AWARDS SHOWCASE 2002  LPA
ROBINSON, KIM STANLEY  PACIFIC EDGE [THREE CALIFORNIAS]  H,P
ROBINSON, KIM STANLEY  PLANET ON THE TABLE, THE  H,SFWA-P,P
ROBINSON, KIM STANLEY  RED MARS [MARS 1]  H:2,P
ROBINSON, KIM STANLEY  REMAKING HISTORY  H,LP

ROBINSON, KIM STANLEY SENSITIVE DEPENDENCE ON INITIAL CONDITIONS, A

[@AUTHOR’S CHOICE MONTHLY #20]

SR-LP

ROBINSON, KIM STANLEY  SHORT, SHARP SHOCK, A  H,SFWA-TD28,P
ROBINSON, KIM STANLEY  SIXTY DAYS AND COUNTING [FORTY SIGNS OF RAIN 3]  H
ROBINSON, KIM STANLEY  WILD SHORE, THE [THREE CALIFORNIAS]  C/P:2,LP
ROBINSON, KIM STANLEY  YEARS OF RICE AND SALT, THE  H,P
ROBINSON, NIGEL  BIRTHRIGHT [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
ROBINSON, NIGEL  DOCTOR WHO CROSSWORD BOOK, THE [DOCTOR WHO]  [C/RAN-P]
ROBINSON, NIGEL  EDGE OF DESTRUCTION, THE [DOCTOR WHO]  SR-P,C/P
ROBINSON, NIGEL  SENSORITES, THE [DOCTOR WHO]  SR-P
ROBINSON, NIGEL  TIME MEDDLER, THE [DOCTOR WHO]  SR-P,P
ROBINSON, NIGEL  TIMEWYRM: APOCALYPSE [DOCTOR WHO NEW ADVENTURES]  C/P,P
ROBINSON, NIGEL  UNDERWATER MENACE, THE [DOCTOR WHO]  SR-P,C/P
ROBINSON, SPIDER  ANTINOMY  C/P:2,P
ROBINSON, SPIDER  BEST OF ALL POSSIBLE WORLDS, THE  C/PA,PA
ROBINSON, SPIDER  BY ANY OTHER NAME  P
ROBINSON, SPIDER  CALLAHAN TOUCH, THE [CALLAHAN’S CROSSTIME SALOON]  C/P,P,H
ROBINSON, SPIDER  CALLAHAN’S CON [CALLAHAN’S CROSSTIME SALOON]  H
ROBINSON, SPIDER  CALLAHAN’S CROSSTIME SALOON [CALLAHAN’S CROSSTIME SALOON]  C/P:2,P
ROBINSON, SPIDER  CALLAHAN’S KEY [CALLAHAN’S CROSSTIME SALOON]  H,P
ROBINSON, SPIDER  CALLAHAN’S LADY [CALLAHAN’S CROSSTIME SALOON]  C/P,SFWA-P:2,P,H,SFWA
ROBINSON, SPIDER  CALLAHAN’S LEGACY [CALLAHAN’S CROSSTIME SALOON]  H,P
ROBINSON, SPIDER  CALLAHAN’S SECRET [CALLAHAN’S CROSSTIME SALOON]  C/P:2,P
ROBINSON, SPIDER  DEATHKILLER [MINDKILLER 1,2]  C/P,P
ROBINSON, SPIDER  FREE LUNCH, THE  H,P
ROBINSON, SPIDER  KILL THE EDITOR  SR-LP
ROBINSON, SPIDER  LADY SLINGS THE BOOZE [CALLAHAN’S CROSSTIME SALOON]  C/P,P,H
ROBINSON, SPIDER  LIFEHOUSE  C/P:2,P
ROBINSON, SPIDER  MELANCHOLY ELEPHANTS  C/P:2,SFWA-P,P
ROBINSON, SPIDER  MINDKILLER [MINDKILLER 1]  C/P,P,H
ROBINSON, SPIDER  NIGHT OF POWER  C/P,P,H
ROBINSON, SPIDER  OFF THE WALL AT CALLAHAN’S [CALLAHAN’S CROSSTIME SALOON]  LP
ROBINSON, SPIDER/ROBINSON, JEANNE  STAR DANCERS, THE [STARDANCE 1,2]  P
ROBINSON, SPIDER/ROBINSON, JEANNE  STARDANCE [STARDANCE 1]  H:2,P
ROBINSON, SPIDER/ROBINSON, JEANNE  STARDANCE (NOVELLA) [STARDANCE 1]  H
ROBINSON, SPIDER/ROBINSON, JEANNE  STARDANCE (PROOF) [STARDANCE 1]  SR-LP
ROBINSON, SPIDER/ROBINSON, JEANNE  STARMIND [STARDANCE 3]  C/P,P,H:2
ROBINSON, SPIDER/ROBINSON, JEANNE  STARSEED [STARDANCE 2]  C/P,SFWA-P:2,P,H
ROBINSON, SPIDER  TELEMPATH  C/P:2,SFWA-P,P
ROBINSON, SPIDER  TIME PRESSURE [MINDKILLER 2]  C/P,SFWA-P,H,P,SFWA
ROBINSON, SPIDER  TIME TRAVELERS STRICTLY CASH [CALLAHAN’S CROSSTIME SALOON]  C/P:2,P
ROBINSON, SPIDER  USER FRIENDLY  C/P:2,P
ROBSON, JUSTINA  KEEPING IT REAL [@QUANTUM GRAVITY #1]  LP
ROBSON, JUSTINA  LIVING NEXT DOOR TO THE GOD OF LOVE  LP
ROBSON, JUSTINA  MAPPA MUNDI  LP
ROBSON, JUSTINA  NATURAL HISTORY  LP
ROBSON, JUSTINA  SILVER SCREEN  P
ROC BOOKS  TASTE OF THE FUTURE, A: FREE SAMPLER  PA
ROCHE, JAMES JEFFREY  SORROWS OF SAP’ED, THE  H
ROCHELLE, WARREN  HARVEST OF CHANGELINGS  H
ROCHELLE, WARREN  WILD BOY, THE  H
ROCKLYNNE, ROSS  MEN AND THE MIRROR, THE  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 327
AUTHOR  TITLE  CODE
ROCKLYNNE, ROSS  SUN DESTROYERS, THE  D93900
ROCKOW, KAREN  FUNERAL CUSTOMS IN TOLKIEN’S TRILOGY  R/REF-P
ROCKWELL, CAREY  DANGER IN DEEP SPACE [TOM CORBETT 2]  H
ROCKWELL, CAREY  ON THE TRAIL OF THE SPACE PIRATES [TOM CORBETT 3]  H:2
ROCKWELL, CAREY  REVOLT ON VENUS, THE [TOM CORBETT 5]  H:2
ROCKWELL, CAREY  ROBOT ROCKET, THE [TOM CORBETT 8]  H:2
ROCKWELL, CAREY  SABOTAGE IN SPACE [TOM CORBETT 7]  H
ROCKWELL, CAREY  SPACE PIONEERS, THE [TOM CORBETT 4]  H
ROCKWELL, CAREY  STAND BY FOR MARS! [TOM CORBETT 1]  H
ROCKWELL, CAREY  TREACHERY IN OUTER SPACE [TOM CORBETT 6]  H
ROCKWELL, MARSHEILA  LEGACY OF WOLVES [@EBERRON INQUISITIVES]  P
ROCKWOOD, ROY  BY AIR EXPRESS TO VENUS  H
ROCKWOOD, ROY  BY SPACE SHIP TO SATURN  H
ROCKWOOD, ROY  CITY BEYOND THE CLOUDS, THE  H
ROCKWOOD, ROY  FIVE THOUSAND MILES UNDERGROUND  H
ROCKWOOD, ROY  LOST ON THE MOON  H
ROCKWOOD, ROY  ON A TORN-AWAY WORLD  H
ROCKWOOD, ROY  THROUGH SPACE TO MARS  H
ROCKWOOD, ROY  UNDER THE OCEAN TO THE SOUTH POLE  H
RODDA, EMILY  WATER SPRITES, THE [FAIRY REALM 8]  H
RODDEN, MARK  STAR BEGOTTEN (TWO VINYL RECORDS)  SR-MM
RODDENBERRY, GENE  CITY ON THE EDGE OF FOREVER (FOTONOVEL) [STAR TREK]  [C/RAN-P]
RODDENBERRY, GENE  INTERGALACTIC TRAVEL (CHINESE)  FP
RODDENBERRY, GENE  MENAGERIE, THE [STAR TREK]  SR-S
RODDENBERRY, GENE/KANDEL, S.  MUDD’S WOMEN [STAR TREK]  SR-S
RODDENBERRY, GENE  STAR TREK GUIDE, THE [STAR TREK]  SR-S

RODDENBERRY, GENE/ZACCARA, NEVIO/ GIOLITTI, ALBERTO STAR TREK: THE KEY COLLECTION [STAR TREK KEY COLLECTION 1]

CX

RODDENBERRY, GENE/GIOLITTI, ALBERTO STAR TREK: THE KEY COLLECTION: VOLUME 2

[STAR TREK KEY COLLECTION 2]

CX

RODDENBERRY, GENE/ZACCARA, NEVIO/ GIOLITTI, ALBERTO STAR TREK: THE KEY COLLECTION: VOLUME 4

[STAR TREK KEY COLLECTION 4]

CX

RODDENBERRY, GENE/MCWILLIAMS, AL/ ZACCARA, NEVIO/GIOLITTI, ALBERTO STAR TREK: THE KEY COLLECTION: VOLUME 5

[STAR TREK KEY COLLECTION 5]

CX

RODDENBERRY, GENE  STAR TREK: THE MOTION PICTURE [STAR TREK]  C/P,P,H
RODDENBERRY, GENE/LIVINGSTON, HAROLD STAR TREK: THE MOTION PICTURE [STAR TREK]  SR-S
RODGERS, ALAN  ALIEN LOVE  LP
RODGERS, ALAN  FIRE  P
RODGERS, ALAN  NIGHT  P
RODGERS, ALAN  RIVER OF OUR DESTINY, THE  LP
RODITI, EDOUARD  DELIGHTS OF TURKEY, THE  LP
ROEHM, BOB  THOMAS BURNETT SWANN BIBLIOGRAPHY, A  R/IND-F
ROESSNER, MICHAELA  STARS COMPEL, THE [TOMMASO 2]  P
ROESSNER, MICHAELA  STARS DISPOSE, THE [TOMMASO 1]  C/P,P
ROESSNER, MICHAELA  VANISHING POINT  H:2,P
ROESSNER, MICHAELA  WALKABOUT WOMAN  C/P,P
ROGERS, ALVA  REQUIEM FOR ASTOUNDING, A  R/REF-H,C/REF-H
ROGERS, BARBARA  PROJECT WEB  H
ROGERS, BRUCE HOLLAND  BONES OF THE WORLD  LPA
ROGERS, CAMERON  MUSIC OF RAZORS, THE  LP
ROGERS, MARK E.  ADVENTURES OF SAMURAI CAT, THE [SAMURAI CAT 1]  SFWA,LP
ROGERS, MARK E.  DEVOURING VOID, THE [@BLOOD OF THE LAMB #2]  P
ROGERS, MARK E.  EXPECTED ONE, THE [BLOOD OF THE LAMB 1]  P
ROGERS, MARK E.  MORE ADVENTURES OF SAMURAI CAT [SAMURAI CAT 2]  SFWA
ROGERS, MARK E.  NIGHTMARE OF GOD, THE [ZORACHUS 2]  P
ROGERS, MARK E.  RIDDLED MAN, THE [@BLOOD OF THE LAMB #3]  P
ROGERS, MARK E.  SAMURAI CAT GOES TO HELL [SAMURAI CAT 5]  LP
ROGERS, MARK E.  SAMURAI CAT IN THE REAL WORLD [SAMURAI CAT 3]  VLP
ROGERS, MARK E.  SWORD OF SAMURAI CAT, THE [SAMURAI CAT 4]  LP
ROGERS, MARK E.  ZORACHUS [ZORACHUS 1]  P
ROGERS, MICHAEL  MINDFOGGER  C/P,P,H
ROGERSOHN, WILLIAM  NORTH DIMENSION  SR-P
ROGO, D. SCOTT/CLARK, JEROME  EARTH’S SECRET INHABITANTS  [C/RAN-P]
ROGO, D. SCOTT  UFO ABDUCTIONS  [C/RAN-P]
ROGOW, ROBERTA  SING A SONG OF TREKKIN’  [C/RAN-H]
ROHAN, MICHAEL SCOTT  ANVIL OF ICE, THE [@WINTER OF THE WORLD #1]  C/P,P
ROHAN, MICHAEL SCOTT  CHASE THE MORNING [STEPHEN FISHER 1]  P,SFWA
ROHAN, MICHAEL SCOTT  CLOUD CASTLES [STEPHEN FISHER 3]  P
ROHAN, MICHAEL SCOTT  FORGE IN THE FOREST, THE [@WINTER OF THE WORLD #2]  P
ROHAN, MICHAEL SCOTT  GATES OF NOON, THE [STEPHEN FISHER 2]  C/P,SFWA,P
ROHAN, MICHAEL SCOTT  HAMMER OF THE SUN, THE [@WINTER OF THE WORLD #3]  H,P
ROHAN, MICHAEL SCOTT/SCOTT, ALLAN  ICE KING, THE  P
ROHAN, MICHAEL SCOTT  MAXIE’S DEMON  P
ROHAN, MICHAEL SCOTT  RUN TO THE STARS  P

AUTHOR  TITLE  CODE
ROHAN, MICHAEL SCOTT  SINGER AND THE SEA, THE [WINTER OF THE WORLD]  P
ROHMER, RICHARD  BALLS!  P
ROHMER, RICHARD  EXXONERATION  H
ROHMER, RICHARD  RETALIATION  H,P
ROHMER, RICHARD  ULTIMATUM  P
ROHMER, SAX  BRIDE OF FU MANCHU, THE  C/P,P
ROHMER, SAX  BROOD OF THE WITCH-QUEEN  C/P,P
ROHMER, SAX  DAUGHTER OF FU MANCHU  H:2,P
ROHMER, SAX  DAY THE WORLD ENDED, THE  C/P:2,P
ROHMER, SAX  DRUMS OF FU MANCHU, THE  C/P,P,H
ROHMER, SAX  EMPEROR FU MANCHU  P
ROHMER, SAX  EMPEROR OF AMERICA, THE  H
ROHMER, SAX  GOLDEN SCORPION, THE  P
ROHMER, SAX  GREEN EYES OF BAST, THE  H,P
ROHMER, SAX  HAND OF FU-MANCHU, THE  H,P
ROHMER, SAX  INSIDIOUS DR. FU MANCHU, THE  C/P,P,H
ROHMER, SAX  ISLAND OF FU MANCHU, THE  C/P:2,P
ROHMER, SAX  MASK OF FU MANCHU, THE  C/P,P
ROHMER, SAX  MOON IS RED, THE  H
ROHMER, SAX  PRESIDENT FU MANCHU  C/P,P
ROHMER, SAX  RE-ENTER FU MANCHU  C/P:2,P
ROHMER, SAX  RETURN OF DR. FU-MANCHU, THE  C/P,P
ROHMER, SAX  RETURN OF SUMURU  P
ROHMER, SAX  SALUTE TO BAZARADA AND OTHER STORIES  H
ROHMER, SAX  SECRET OF HOLM PEEL, THE  P
ROHMER, SAX  SEVEN SINS  H
ROHMER, SAX  SHADOW OF FU MANCHU, THE  C/P:2,P
ROHMER, SAX  SUMURU  P
ROHMER, SAX  TRAIL OF FU MANCHU, THE  C/P,P
ROHMER, SAX  VIRGIN IN FLAMES  H
ROHMER, SAX  WRATH OF FU MANCHU, THE  C/P,P
ROHMER, SAX  YELLOW CLAW, THE  C/P,P
ROLAND, HOWELL  LIVING STONE, THE  P
ROLLESTON, T. W.  CELTIC MYTHS AND LEGENDS  R/REF-H
ROMANO, DEANE  FLIGHT FROM TIME ONE  H,SFWA
ROMANS, R. H.  MOON CONQUERORS, THE  SR-P
ROME, DAVID  SQUAT  SR-P
ROMERO, GEORGE A./SPARROW, SUSANNA  DAWN OF THE DEAD  H,LP
ROMILLY, ERIC  BLEEDING FROM THE ROMAN  H
ROMILUS, ARN  BRAIN PALAEO  SR-P
ROMILUS, ARN  ORGANIC DESTINY  SR-P
ROMINE, ADEN FOSTER/ROMINE, MARY COX ETERNITY STONE, THE  P
ROMKEY, MICHAEL  I, VAMPIRE  P
ROMKEY, MICHAEL  VAMPIRE PAPERS, THE  P
ROMKEY, MICHAEL  VAMPIRE PRINCESS, THE  P
RONALD, BRUCE W.  OUR MAN IN SPACE  DM117
ROPP, THEODORE  WAR IN THE MODERN WORLD  [C/RAN-P]
ROQUELAURE, A. N.  BEAUTY’S PUNISHMENT [BEAUTY 2]  SR-L
ROQUELAURE, A. N.  BEAUTY’S RELEASE [BEAUTY 3]  H,SR-L
ROQUELAURE, A. N.  CLAIMING OF SLEEPING BEAUTY, THE [BEAUTY 1]  H,SR-L
ROSE, COLIN  ACCELERATED LEARNING  [C/RAN-H]
ROSE, JOEL  KILL THE POOR  SFWA
ROSE, LAWRENCE F.  HELL FRUIT, THE  SR-P
ROSE, LLOYD  ALGEBRA OF ICE, THE [DOCTOR WHO PAST DOCTORS]  P
ROSE, LLOYD  CAMERA OBSCURA [DOCTOR WHO EIGHTH DOCTOR]  P
ROSE, LLOYD  CITY OF THE DEAD, THE [DOCTOR WHO EIGHTH DOCTOR]  P
ROSE, LOIS/ROSE, STEPHEN  SHATTERED RING, THE  R/REF-H
ROSE, MARK  ALIEN ENCOUNTERS  R/REF-P
ROSE, MARK  SCIENCE FICTION  R/REF-P
ROSELLE, DANIEL  TRANSFORMATIONS  PA
ROSELLE, DANIEL  TRANSFORMATIONS II  PA,C/PA
ROSEN, SELINA  CHAINS OF DESTRUCTION [CHAINS 2]  LP
ROSEN, SELINA  CHAINS OF FREEDOM [CHAINS 1]  LP
ROSEN, SELINA  QUEEN OF DENIAL [DREWCILA QWAH 1]  LP
ROSEN, SELINA  RECYCLED [DREWCILA QWAH 2]  LP
ROSENBERG, AARON  QUEEN OF BLADES [@STARCRAFT]  P
ROSENBERG, JOEL  CRIMSON SKY, THE [@KEEPERS OF THE HIDDEN WAYS #3]  P
ROSENBERG, JOEL  D’SHAI [D’SHAI 1]  C/P,SFWA-P:2,P,H
ROSENBERG, JOEL  EMILE AND THE DUTCHMAN  C/P,P
ROSENBERG, JOEL  FIRE DUKE, THE [@KEEPERS OF THE HIDDEN WAYS 1]  P
ROSENBERG, JOEL  HEIR APPARENT, THE [@GUARDIANS OF THE FLAME #4]  C/P,P
ROSENBERG, JOEL  HERO  C/P,P
ROSENBERG, JOEL  HEROES, THE [GUARDIANS OF THE FLAME 4,5]  H:2
ROSENBERG, JOEL  HOUR OF THE OCTOPUS [D’SHAI 2]  P

AUTHOR  TITLE  CODE
ROSENBERG, JOEL  KNIGHT MOVES [@PALADINS #2]  H

ROSENBERG, JOEL NOT EXACTLY THE THREE MUSKETEERS

[GUARDIANS OF THE FLAME CASTLE CULLINANE 1]

H

ROSENBERG, JOEL  NOT FOR GLORY  H:2,SFWA-P,P

ROSENBERG, JOEL NOT QUITE SCARAMOUCHE

[GUARDIANS OF THE FLAME CASTLE CULLINANE 2]

H

ROSENBERG, JOEL NOT REALLY THE PRISONER OF ZENDA

[GUARDIANS OF THE FLAME CASTLE CULLINANE 3]

H

ROSENBERG, JOEL  PALADINS [PALADINS 1]  P
ROSENBERG, JOEL  ROAD HOME, THE [GUARDIANS OF THE FLAME]  P
ROSENBERG, JOEL  ROAD TO EHVENOR, THE [GUARDIANS OF THE FLAME 6]  H,P
ROSENBERG, JOEL  SILVER CROWN, THE [@GUARDIANS OF THE FLAME #3]  C/P:2,P
ROSENBERG, JOEL  SILVER STONE, THE [@KEEPERS OF THE HIDDEN WAYS #2]  C/P,P
ROSENBERG, JOEL  SLEEPING DRAGON, THE [@GUARDIANS OF THE FLAME #1]  C/P:2,P
ROSENBERG, JOEL  SWORD AND THE CHAIN, THE [@GUARDIANS OF THE FLAME #2]  C/P:2,P
ROSENBERG, JOEL  TIES OF BLOOD AND SILVER  C/P:2,P
ROSENBERG, JOEL  WARRIOR LIVES, THE [GUARDIANS OF THE FLAME 5]  C/P,P,H
ROSENBERG, JOEL  WARRIORS, THE [GUARDIANS OF THE FLAME 1,2,3]  H:2

ROSENBERG, TEYA/HIXON, MARTHA P./ SCAPPLE, SHARON M./WHITE, DONNA R.

DIANA WYNNE JONES

R/REF-H

ROSENBLUM, MARY  CHIMERA  P
ROSENBLUM, MARY  DRYLANDS, THE  P
ROSENBLUM, MARY  HORIZONS  H
ROSENBLUM, MARY  STONE GARDEN, THE  C/P,P
ROSENBLUM, MARY  SYNTHESIS & OTHER VIRTUAL REALITIES  H
ROSENTHAL, MONROE/MUNSON, DONALD  PRESIDENT KISSINGER  P
ROSHEIM, DAVID L.  GALAXY MAGAZINE: THE DARK AND THE LIGHT YEARS  R/REF-H
ROSHWALD, MORDECAI  AMOTINADOS DO POLAR LION, OS  FP
ROSHWALD, MORDECAI  LEVEL 7  C/P,P,H
ROSHWALD, MORDECAI  SMALL ARMAGEDDON, A  P
ROSNER, SARA  OTHER DIMENSION, THE  PA
ROSNY, J. H.  IRONCASTLE  C/P,P
ROSNY, J. H.  QUEST FOR FIRE, THE  H
ROSNY, J. H.  QUEST OF THE DAWN MAN  P
ROSS, DAVID D.  ARGUS GAMBIT, THE [DREAMERS OF THE DAY 1]  H,SFWA
ROSS, DAVID D.  EIGHTH RANK, THE [DREAMERS OF THE DAY 2]  H,SFWA
ROSS, IAN  AMAZONS [MIND MASTERS 4]  C/P,P
ROSS, IAN  RECYCLED SOULS [MIND MASTERS 5]  P
ROSS, JOSEPH  BEST OF AMAZING, THE  PA
ROSS, MARILYN  BARNABAS COLLINS  C/P,P
ROSS, MARILYN  BARNABAS COLLINS AND QUENTIN’S DEMON  C/P,P
ROSS, MARILYN  BARNABAS COLLINS AND THE GYPSY WITCH  P
ROSS, MARILYN  BARNABAS COLLINS AND THE MYSTERIOUS GHOST  P
ROSS, MARILYN  BARNABAS COLLINS VS. THE WARLOCK  C/P:2,P
ROSS, MARILYN  BARNABAS, QUENTIN AND DOCTOR JEKYLL’S SON  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE AVENGING GHOST  C/P,P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE BODY SNATCHERS  C/P,P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE CRYSTAL COFFIN  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE FRIGHTENED BRIDE  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE GRAVE ROBBERS  P
ROSS, MARILYN  BARNABAS, QUENTIN, AND THE HAUNTED CAVE  C/P:2,P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE MAD MAGICIAN  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE MAGIC POTION  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE MUMMY’S CURSE  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE NIGHTMARE ASSASSIN  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE SCORPIO CURSE  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE SEA GHOST  C/P:2,P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE SERPENT  P
ROSS, MARILYN  BARNABAS, QUENTIN AND THE WITCH’S CURSE  P
ROSS, MARILYN  CURSE OF COLLINWOOD, THE  C/P,P
ROSS, MARILYN  DARK SHADOWS  P
ROSS, MARILYN  DEMON OF BARNABAS COLLINS, THE  C/P,P
ROSS, MARILYN  FOE OF BARNABAS COLLINS, THE  P
ROSS, MARILYN  HOUSE OF DARK SHADOWS  P
ROSS, MARILYN  MYSTERY OF COLLINWOOD, THE  P
ROSS, MARILYN  PERIL OF BARNABAS COLLINS, THE  P
ROSS, MARILYN  PHANTOM AND BARNABAS COLLINS, THE  P
ROSS, MARILYN  SECRET OF BARNABAS COLLINS, THE  P
ROSS, MARILYN  STRANGERS AT COLLINS HOUSE  C/P,P
ROSS, MARILYN  VICTORIA WINTERS  P
ROSS, PHILIP  TALLEY’S TRUTH  SFWA
ROSSITER, OSCAR  TETRASOMY TWO  C/P,SFWA,P,H
ROSSMANN, JOHN F.  MIND MASTERS, THE [@MIND MASTERS #1]  P
ROSSMANN, JOHN F.  SHAMBALLAH [@MIND MASTERS #2]  P

The MITSFS Pinkdex  September 24, 2007  Page 330
AUTHOR  TITLE  CODE
ROSZAK, THEODORE  BUGS  H:2,P
ROSZAK, THEODORE  DREAMWATCHER  SFWA
ROTH, KURT  AGE OF REASON, THE  SR-HA
ROTHBERG, ABRAHAM  SWORD OF THE GOLEM, THE  C/P,P,H
ROTHFUSS, PATRICK  NAME OF THE WIND, THE [KINGKILLER 1]  H
ROTHMAN, CHUCK  STAROAMER’S FATE  C/P:2,P
ROTHMAN, TONY  WORLD IS ROUND, THE  C/P:2,P
ROTSLER, WILLIAM  AND CALL MY KILLER...MODOK! [IRON MAN]  P
ROTSLER, WILLIAM  FAR FRONTIER, THE  P
ROTSLER, WILLIAM  HIDDEN WORLDS OF ZANDRA, THE  H,SFWA
ROTSLER, WILLIAM  MISTER MERLIN 1  C/P:2,LP
ROTSLER, WILLIAM  MISTER MERLIN 2  C/P:2,LP
ROTSLER, WILLIAM  NIGHTMARE [DOCTOR STRANGE]  C/P,P
ROTSLER, WILLIAM  PATRON OF THE ARTS  P
ROTSLER, WILLIAM  STAR TREK II SHORT STORIES [STAR TREK]  P
ROTSLER, WILLIAM  STAR TREK III SHORT STORIES [STAR TREK]  P
ROTSLER, WILLIAM  TO THE LAND OF THE ELECTRIC ANGEL  P
ROTSLER, WILLIAM  ZANDRA  H,SFWA
ROTTENSTEINER, FRANZ  FANTASY BOOK, THE  R/REF-H
ROTTENSTEINER, FRANZ  SCIENCE FICTION BOOK, THE  R/REF-H
ROTTENSTEINER, FRANZ  VIEW FROM ANOTHER SHORE  C/PA,PA,SFWA
ROUCH, JAMES  BLIND FIRE [@ZONE #2]  P
ROUSSEAU, VICTOR  MESSIAH OF THE CYLINDER, THE  H,LP
ROUTLEY, JANE  ARAMAYA  P
ROUTLEY, JANE  FIRE ANGELS [DION HOLYHANDS 1]  H,LP
ROUTLEY, JANE  MAGE HEART [DION HOLYHANDS 2]  C/P,P
ROVIN, JEFF  MORTAL KOMBAT [MORTAL KOMBAT 0]  P
ROVIN, JEFF  PICTORIAL HISTORY OF SCIENCE FICTION FILMS, A  R/REF-H
ROVIN, JEFF  RE-ANIMATOR  P
ROVIN, JEFF  TRANSGALACTIC GUIDE TO SOLAR SYSTEM M-17, THE  H,LP
ROWE, KEN  TRENCH AND BEYOND, THE  LP
ROWE, STEPHANIE  DATE ME, BABY, ONE MORE TIME  P
ROWE, STEPHANIE  MUST LOVE DRAGONS  P
ROWLEY, CHRISTOPHER  ANCIENT ENEMY, THE [ARNA 1]  P
ROWLEY, CHRISTOPHER  BATTLEDRAGON [BAZIL BROKETAIL 4]  P
ROWLEY, CHRISTOPHER  BAZIL BROKETAIL [BAZIL BROKETAIL 1]  C/P,SFWA-P,P
ROWLEY, CHRISTOPHER  BLACK SHIP, THE  C/P:2,P
ROWLEY, CHRISTOPHER  DOOM’S BREAK [@ARNA #3]  P
ROWLEY, CHRISTOPHER  DRAGON AT WORLD’S END, A [BAZIL BROKETAIL 5]  P
ROWLEY, CHRISTOPHER  DRAGON ULTIMATE [BAZIL BROKETAIL 7]  P
ROWLEY, CHRISTOPHER  DRAGONS OF ARGONATH [BAZIL BROKETAIL 6]  P
ROWLEY, CHRISTOPHER  DRAGONS OF WAR [BAZIL BROKETAIL 3]  C/P,P
ROWLEY, CHRISTOPHER  FOUNDER, THE  C/P,P
ROWLEY, CHRISTOPHER  GOLDEN SUNLANDS  C/P,P
ROWLEY, CHRISTOPHER  SHAST WAR, THE [@ARNA 2]  P
ROWLEY, CHRISTOPHER  STARHAMMER  C/P,P
ROWLEY, CHRISTOPHER  SWORD FOR A DRAGON, A [BAZIL BROKETAIL 2]  C/P:2,SFWA-P,P
ROWLEY, CHRISTOPHER  TO A HIGHLAND NATION  C/P,P
ROWLEY, CHRISTOPHER  VANG: THE BATTLEMASTER, THE [VANG]  P
ROWLEY, CHRISTOPHER  VANG: THE MILITARY FORM, THE [VANG]  C/P,P
ROWLEY, CHRISTOPHER  WAR FOR ETERNITY, THE  C/P:2,P
ROWLEY, CHRISTOPHER  WIZARD AND THE FLOATING CITY, THE  P

ROWLING, J. K.

HARRY POTTER AND THE CHAMBER OF SECRETS (AMERICAN) [HARRY POTTER 2]

H:2

ROWLING, J. K.  HARRY POTTER AND THE DEATHLY HALLOWS [HARRY POTTER 7]  H

ROWLING, J. K.

HARRY POTTER AND THE GOBLET OF FIRE (AMERICAN) [HARRY POTTER 4]

H:2,LP

ROWLING, J. K.  HARRY POTTER AND THE GOBLET OF FIRE (BRITISH) [HARRY POTTER 4]  H
ROWLING, J. K.  HARRY POTTER AND THE HALF-BLOOD PRINCE [HARRY POTTER 6]  H:2

ROWLING, J. K.

HARRY POTTER AND THE ORDER OF THE PHOENIX (AMERICAN) [HARRY POTTER 5]

H

ROWLING, J. K.

HARRY POTTER AND THE PHILOSOPHER’S STONE (BRITISH) [HARRY POTTER 1]

LP

ROWLING, J. K.

HARRY POTTER AND THE PRISONER OF AZKABAN (AMERICAN) [HARRY POTTER 3]

H

ROWLING, J. K.

HARRY POTTER AND THE SORCERER’S STONE (AMERICAN) [HARRY POTTER 1]

C/P,LP,H

ROWLING, J. K.  HARRY POTTER UND DER STEIN DER WEISEN [HARRY POTTER 1]  FH
ROXBURY, KYLE  GOLDEN GIRL OF HOCKEINBECK, THE  SR-L
ROY, JOHN FLINT  GUIDE TO BARSOOM, A  R/REF-P,C/REF-P:2
ROYLE, NICHOLAS  COUNTERPARTS  P
RUBEN, WILLIAM S.  DIONYSUS: THE ULTIMATE EXPERIMENT  P
RUBY, LAURA  CHAOS KING, THE [GEORGIE AND BUG 2]  SR-LP
RUCK, BERTA  IMMORTAL GIRL, THE  H

The MITSFS Pinkdex  September 24, 2007  Page 331
AUTHOR  TITLE  CODE
RUCKA, GREG  GRENDEL: PAST PRIME  LP
RUCKER, RUDY  ALL THE VISIONS  OVD-2
RUCKER, RUDY  57TH FRANZ KAFKA, THE  P
RUCKER, RUDY  FOURTH DIMENSION, THE  [C/RAN-H]
RUCKER, RUDY  FREEWARE [WARE 3]  LP
RUCKER, RUDY  FREK AND THE ELIXER  LP
RUCKER, RUDY  GNARL!  H,LP
RUCKER, RUDY  HACKER AND THE ANTS, THE  C/P,P,H
RUCKER, RUDY  HOLLOW EARTH, THE  H,SFWA
RUCKER, RUDY  LIVE ROBOTS [COBB ANDERSON 1,2]  P
RUCKER, RUDY  MAD PROFESSOR  LP
RUCKER, RUDY  MASTER OF SPACE AND TIME  C/P:2,SFWA,P,H
RUCKER, RUDY  MATHEMATICIANS IN LOVE  H
RUCKER, RUDY  MATHENAUTS  HA,SFWA
RUCKER, RUDY  MOLDIES & MEATBOPS [WARE 1,2,3]  H
RUCKER, RUDY  REALWARE [WARE 4]  LP
RUCKER, RUDY  SAUCER WISDOM  LP
RUCKER, RUDY  SECRET OF LIFE, THE  H

RUCKER, RUDY/WILSON, PETER LAMBORN/ WILSON, ROBERT ANTON SEMIOTEXT[E] SF 14

HA,SR-LPA

RUCKER, RUDY  SEX SPHERE, THE  P
RUCKER, RUDY  SOFTWARE [WARE 1]  C/P:2,P
RUCKER, RUDY  SPACELAND  H
RUCKER, RUDY  SPACETIME DONUTS  P
RUCKER, RUDY  WETWARE [WARE 2]  C/P:2,P
RUCKER, RUDY  WHITE LIGHT  P
RUDHYAR, DANE  RANIA  P
RUDITIS, PAUL  AS PUCK WOULD HAVE IT [@CHARMED]  P
RUDITIS, PAUL  BREWING STORM, THE [@CHARMED]  P
RUDITIS, PAUL  SHOCKWAVE [STAR TREK ENTERPRISE]  H,@P
RUDORFF, RAYMOND  DRACULA ARCHIVES, THE  C/P,P
RUESCH, HANS  SLAUGHTER OF THE INNOCENT  [C/RAN-P]
RUFF, MATT  FOOL ON THE HILL  P,SFWA
RUFF, MATT  SEWER, GAS & ELECTRIC  P
RUGGIERO, TONY  NOVEL SAMPLER (CD)  SR-MM:2
RUNTE, ROBERT/MEYNARD, YVES  TESSERACTS 5  C/PA,PA
RUNYAN, C. F.  FLIGHT, THE  LP
RUNYON, CHARLES W.  AMES HOLBROOK, DEITY  C/P,P
RUNYON, CHARLES W.  I, WEAPON  C/P,SFWA,P,H
RUNYON, CHARLES W.  PIG WORLD  H,SFWA,P
RUNYON, CHARLES W.  SOULMATE  P
RUPPELT, EDWARD J.  REPORT ON UNIDENTIFIED FLYING OBJECTS, THE  [C/RAN-H:3,C/RAN-P]

RUSCH, KRISTINE KATHRYN/ ANDERSON, KEVIN J.

AFTERIMAGE

P

RUSCH, KRISTINE KATHRYN/ ANDERSON, KEVIN J.

AFTERIMAGE AFTERSHOCK

LP

RUSCH, KRISTINE KATHRYN  ALIEN INFLUENCES  P
RUSCH, KRISTINE KATHRYN  BEST OF PULPHOUSE, THE  HA
RUSCH, KRISTINE KATHRYN  BLACK KING, THE [BLACK THRONE 2]  P
RUSCH, KRISTINE KATHRYN  BLACK QUEEN, THE [BLACK THRONE 1]  P
RUSCH, KRISTINE KATHRYN  BURIED DEEP [RETRIEVAL ARTIST 4]  C/P,P
RUSCH, KRISTINE KATHRYN  CHANGELING, THE [@FEY 2]  C/P,P
RUSCH, KRISTINE KATHRYN  CONSEQUENCES [RETRIEVAL ARTIST 3]  P
RUSCH, KRISTINE KATHRYN  DISAPPEARED, THE [RETRIEVAL ARTIST 1]  P
RUSCH, KRISTINE KATHRYN  EXTREMES [RETRIEVAL ARTIST 2]  P
RUSCH, KRISTINE KATHRYN  FANTASY LIFE  P
RUSCH, KRISTINE KATHRYN  HEART READERS  C/P,P
RUSCH, KRISTINE KATHRYN  NEW REBELLION, THE [@STAR WARS]  C/P,P
RUSCH, KRISTINE KATHRYN  PALOMA [RETRIEVAL ARTIST 5]  P
RUSCH, KRISTINE KATHRYN  PULPHOUSE 1  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 2  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 3  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 4  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 5  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 6  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 7  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 8  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 9  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 10  SR-HA
RUSCH, KRISTINE KATHRYN  PULPHOUSE 11  SR-HA
RUSCH, KRISTINE KATHRYN  RESISTANCE, THE [@FEY 4]  C/P,P
RUSCH, KRISTINE KATHRYN  RIVAL, THE [@FEY 3]  P
RUSCH, KRISTINE KATHRYN  SACRIFICE, THE [@FEY 1]  C/P,P
RUSCH, KRISTINE KATHRYN  SINS OF THE BLOOD  P

The MITSFS Pinkdex  September 24, 2007  Page 332
AUTHOR  TITLE  CODE
RUSCH, KRISTINE KATHRYN  STORIES FOR AN ENCHANTED AFTERNOON  H

RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY THIN AIR [@STAR TREK NEW EARTH #5]

C/P:2,P

RUSCH, KRISTINE KATHRYN  VICTORY [@FEY 5]  P
RUSCH, KRISTINE KATHRYN  WHITE MISTS OF POWER, THE  C/P:2,P

RUSCH, KRISTINE KATHRYN/ SMITH, DEAN WESLEY X-MEN [X-MEN]

P

RUSE, GARY ALAN  DEATH HUNT ON A DYING PLANET  C/P,P
RUSE, GARY ALAN  GODS OF CERUS MAJOR, THE  H,SFWA
RUSE, GARY ALAN  MORLAC: THE QUEST OF THE GREEN MAGICIAN  P
RUSH, ALISON  LAST OF THE DANU’S CHILDREN, THE  P
RUSHDIE, SALMAN  GRIMUS  SR-P,SFWA
RUSHDIE, SALMAN  SATANIC VERSES, THE  H
RUSK, JAMES JR.  SPACE SLAVES  C/P,P
RUSK, JAMES JR.  TUG OF THE DWARF STAR  P
RUSOFF, GARRY  SPEAR OF FIRE  P
RUSS, JOANNA  ADVENTURES OF ALYX, THE  C/P:2,P
RUSS, JOANNA  AND CHAOS DIED  C/P:2,P
RUSS, JOANNA  EXTRA(ORDINARY) PEOPLE  H,LP
RUSS, JOANNA  FEMALE MAN, THE  C/P:2,P
RUSS, JOANNA  HIDDEN SIDE OF THE MOON, THE  P
RUSS, JOANNA  KITTATINNY  VLP
RUSS, JOANNA  PICNIC ON PARADISE  C/P:2,P
RUSS, JOANNA  SOULS  SFWA-TD11,TD11
RUSS, JOANNA  TWO OF THEM, THE  H,P
RUSS, JOANNA  WE WHO ARE ABOUT TO...  C/P,P
RUSS, JOANNA  ZANZIBAR CAT, THE  C/P,P
RUSSELL, ALAN  SCIENCE FICTION BY THE RIVALS OF H. G. WELLS  HA
RUSSELL, BERTRAND  BASIC WRITINGS OF BERTRAND RUSSELL, THE  [C/RAN-H]
RUSSELL, BERTRAND  NIGHTMARES OF EMINENT PERSONS  H
RUSSELL, ERIC FRANK  BEST OF ERIC FRANK RUSSELL, THE  C/P:2,P
RUSSELL, ERIC FRANK  DEEP SPACE  C/P,P,H
RUSSELL, ERIC FRANK  DREADFUL SANCTUARY (REVISED)  C/P:2,P
RUSSELL, ERIC FRANK  ENTITIES  H
RUSSELL, ERIC FRANK  GREAT EXPLOSION, THE  H:2,P
RUSSELL, ERIC FRANK  MAJOR INGREDIENTS  H
RUSSELL, ERIC FRANK  MEN, MARTIANS AND MACHINES  C/P,P,H
RUSSELL, ERIC FRANK  MINDWARPERS, THE  C/P:2,P
RUSSELL, ERIC FRANK  NEXT OF KIN  C/P:2,P
RUSSELL, ERIC FRANK  SENTINELS FROM SPACE  C/P:2,P
RUSSELL, ERIC FRANK  SENTINELS OF SPACE  SR-DD044
RUSSELL, ERIC FRANK  SINISTER BARRIER  C/P:2,P,GN1.01
RUSSELL, ERIC FRANK  SIX WORLDS YONDER  DD315,D77785
RUSSELL, ERIC FRANK  SOMEWHERE A VOICE  C/P:2,P
RUSSELL, ERIC FRANK  SPACE WILLIES, THE  DD315,D77785
RUSSELL, ERIC FRANK  THREE TO CONQUER  DD215,P,C/P
RUSSELL, ERIC FRANK  WASP  H:2,P
RUSSELL, ERIC FRANK  WASP (UNABRIDGED)  C/P:2,P
RUSSELL, ERIC FRANK  WITH A STRANGE DEVICE  P
RUSSELL, GARY  BUSINESS UNUSUAL [DOCTOR WHO PAST DOCTORS]  C/P:2,P
RUSSELL, GARY  DEADFALL [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
RUSSELL, GARY  DIVIDED LOYALTIES [DOCTOR WHO PAST DOCTORS]  P
RUSSELL, GARY  DOCTOR WHO [DOCTOR WHO]  C/P,P
RUSSELL, GARY  DOCTOR WHO: THE INSIDE STORY [DOCTOR WHO]  R/REF-H
RUSSELL, GARY  INSTRUMENTS OF DARKNESS [DOCTOR WHO]  P
RUSSELL, GARY  INVASION OF THE CAT-PEOPLE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
RUSSELL, GARY  LEGACY [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
RUSSELL, GARY  PLACEBO EFFECT [DOCTOR WHO EIGHTH DOCTOR]  C/P:2,P
RUSSELL, GARY  SCALES OF INJUSTICE, THE [DOCTOR WHO MISSING ADVENTURES]  C/P:2,P
RUSSELL, GARY  SHORT TRIPS: REPERCUSSIONS [DOCTOR WHO]  HA
RUSSELL, GARY  SHORT TRIPS: THE SOLAR SYSTEM [DOCTOR WHO]  HA
RUSSELL, GARY  SPIRAL SCRATCH [DOCTOR WHO PAST DOCTORS]  P
RUSSELL, JAY  BURNING BRIGHT  LP
RUSSELL, JOHN  IN DARK PLACES  H
RUSSELL, JOHN  WHERE THE PAVEMENT ENDS  H
RUSSELL, JOHN ROBERT  CABU  C/P:2,P
RUSSELL, JOHN ROBERT  SAR  C/P:2,P
RUSSELL, JOHN ROBERT  TA  C/P,P
RUSSELL, MARY DORIA  CHILDREN OF GOD  LP
RUSSELL, MARY DORIA  SPARROW, THE  H:2,LP
RUSSELL, RAY  INCUBUS  P
RUSSELL, RAY  PRINCE OF DARKNESS  SR-P
RUSSELL, RAY  SAGITTARIUS  C/P:2,P
RUSSELL, RAY  SARDONICUS AND OTHER STORIES  P

The MITSFS Pinkdex  September 24, 2007  Page 333
AUTHOR  TITLE  CODE
RUSSELL, RAY  UNHOLY TRINITY  P
RUSSELL, SEAN  BENEATH THE VAULTED HILLS [RIVER INTO DARKNESS 1]  C/P,P
RUSSELL, SEAN  COMPASS OF THE SOUL, THE [RIVER INTO DARKNESS 2]  P
RUSSELL, SEAN  GATHERER OF CLOUDS [INITIATE BROTHER 2]  C/P:2,SFWA-P,P
RUSSELL, SEAN  INITIATE BROTHER, THE [INITIATE BROTHER 1]  C/P:2,P
RUSSELL, SEAN  ISLE OF BATTLE, THE [@SWANS’ WAR #2]  H
RUSSELL, SEAN  ONE KINGDOM, THE [@SWANS’ WAR #1]  H
RUSSELL, SEAN  SEA WITHOUT A SHORE [MOONTIDE AND MAGIC RISE 2]  P
RUSSELL, SEAN  SHADOW ROADS, THE [SWANS’ WAR 3]  H
RUSSELL, SEAN  WORLD WITHOUT END [MOONTIDE AND MAGIC RISE 1]  C/P,P
RUSSELL, W. CLARK  FROZEN PIRATE, THE  H
RUSSO, JOHN  BLACK CAT  P
RUSSO, JOHN  LIVING THINGS  P
RUSSO, JOHN  MIDNIGHT  P
RUSSO, JOHN  NIGHT OF THE LIVING DEAD  P
RUSSO, RICHARD PAUL  CARLUCCI’S EDGE [FRANK CARLUCCI 2]  P
RUSSO, RICHARD PAUL  CARLUCCI’S HEART [FRANK CARLUCCI 3]  P
RUSSO, RICHARD PAUL  DESTROYING ANGEL [FRANK CARLUCCI 1]  C/P,P
RUSSO, RICHARD PAUL  INNER ECLIPSE  P,SFWA-P
RUSSO, RICHARD PAUL  SHIP OF FOOLS  C/P,LP,H
RUSSO, RICHARD PAUL  SUBTERRANEAN GALLERY  P,SFWA-P
RUSSON, PENNI  BREATHE  H
RUTHERFORD, MICHAEL  INFINITE KINGDOMS, THE  LP
RUTHERFORD, MICHAEL  TALE AND ITS MASTER, THE  SFWA
RUTMAN, LEO  CLASH OF EAGLES  P
RUUTH, MARIANNE  OUTBREAK  P
RYAN, ALAN  BONES WIZARD, THE  H,SFWA
RYAN, ALAN  CAST A COLD EYE  P
RYAN, ALAN  HALLOWEEN HORRORS  PA,SFWA
RYAN, ALAN  NIGHT VISIONS 1  SR-HA
RYAN, ALAN  PENGUIN BOOK OF VAMPIRE STORIES, THE  LPA
RYAN, ALAN  PERPETUAL LIGHT  HA,PA
RYAN, ALAN  QUADRIPHOBIA  H,SFWA
RYAN, ALAN  VAMPIRES  HA
RYAN, C. J.  DEXTA [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 1]  P

RYAN, C. J.

FIFTH QUADRANT, THE [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS #2]

P

RYAN, C. J.  GLORIOUS TREASON [DEPARTMENT OF EXTRATERRESTRIAL AFFAIRS 2]  P
RYAN, CHARLES C.  STARRY MESSENGER: THE BEST OF GALILEO  HA
RYAN, FERGUS  SIEGE OF MT. NEVERMIND, THE [@DRAGONLANCE CHAOS WAR 6]  P
RYAN, KATHLEEN  RAVNOS [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
RYAN, KATHLEEN  SETITE [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
RYAN, KEVIN  DEMANDS OF HONOR [@STAR TREK ERRAND OF FURY #2]  P
RYAN, KEVIN  EDGE OF THE SWORD, THE [@STAR TREK ERRAND OF VENGEANCE #1]  P
RYAN, KEVIN  KILLING BLOW [@STAR TREK ERRAND OF VENGEANCE #2]  P
RYAN, KEVIN  RIVER OF BLOOD [@STAR TREK ERRAND OF VENGEANCE #3]  P
RYAN, KEVIN  SEEDS OF RAGE [@STAR TREK ERRAND OF FURY #1]  P
RYAN, KEVIN  VAN HELSING  P
RYAN, THOMAS J.  ADOLESCENCE OF P-1, THE  H:2,LP
RYDILL, JESSICA  CHILDREN OF THE SHAMAN  P
RYMAN, GEOFF  AIR  LP
RYMAN, GEOFF  CHILD GARDEN, THE  H
RYMAN, GEOFF  UNCONQUERED COUNTRIES  H
RYMAN, GEOFF  UNCONQUERED COUNTRY, THE  P
RYMAN, GEOFF  WARRIOR WHO CARRIED LIFE, THE  P
RYMER, JAMES  BRIDE OF THE VAMPYRE [@VARNEY #4]  LP
RYMER, JAMES  FATE OF THE VAMPYRE, THE [@VARNEY THE VAMPIRE #5]  LP
RYMER, JAMES  VARNEY THE VAMPYRE I [VARNEY #1]  LP
RYMER, JAMES  VARNEY THE VAMPYRE II [VARNEY #2]  LP
RYPEL, T. C.  GONJI [GONJI]  P

The MITSFS Pinkdex  September 24, 2007  Page 335
AUTHOR  TITLE  CODE
S. C. ENTERPRISES  STAR TREK — AN ANALYSIS OF A PHENOMENON IN SCIENCE FICTION  R/REF-H
SABEN, LIONEL  REPLICA  P
SABERHAGEN, FRED  AFTER THE FACT [PILGRIM 2]  C/P,SFWA-P,P
SABERHAGEN, FRED  ARDNEH’S WORLD [EMPIRE OF THE EAST 3]  P,SFWA-P
SABERHAGEN, FRED  ARIADNE’S WEB [GODS 2]  P
SABERHAGEN, FRED  ARMORY OF SWORDS, AN [SWORDS]  PA
SABERHAGEN, FRED  ARMS OF HERCULES, THE [GODS 3]  C/P,LP
SABERHAGEN, FRED  BERSERKER [BERSERKER 1]  C/P:2,P
SABERHAGEN, FRED  BERSERKER ATTACK, THE  C/P:2,P

SABERHAGEN, FRED BERSERKER BASE [BERSERKER 2] SFWA,SFWA-P,HA:2,

LPA

SABERHAGEN, FRED  BERSERKER: BLUE DEATH [BERSERKER 3]  C/P:2,SFWA,SFWA-P,LP
SABERHAGEN, FRED  BERSERKER FURY [BERSERKER 7]  H,P
SABERHAGEN, FRED  BERSERKER KILL [BERSERKER 6]  H:2
SABERHAGEN, FRED  BERSERKER LIES [BERSERKER]  C/P,P
SABERHAGEN, FRED  BERSERKER MAN [BERSERKER]  C/P:2,P
SABERHAGEN, FRED  BERSERKER PRIME [BERSERKER]  P
SABERHAGEN, FRED  BERSERKER THRONE, THE [BERSERKER 4]  H:2,SFWA-P,LP
SABERHAGEN, FRED  BERSERKER WARS, THE [BERSERKER 1]  C/P,SFWA-P,P
SABERHAGEN, FRED  BERSERKER’S PLANET [BERSERKER 5]  C/P:2,P
SABERHAGEN, FRED  BERSERKER’S STAR [BERSERKER]  P
SABERHAGEN, FRED  BERSERKERS: THE BEGINNING [BERSERKER]  C/P,P
SABERHAGEN, FRED  BERSERKERS: THE ULTIMATE ENEMY [BERSERKER 4]  C/P,P
SABERHAGEN, FRED  BLACK MOUNTAINS, THE [EMPIRE OF THE EAST 2]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  BOOK OF SABERHAGEN, THE  C/P:2,P
SABERHAGEN, FRED  BOOKS OF THE GODS, THE, PART ONE [GODS 1,2,3]  H
SABERHAGEN, FRED  BROKEN LANDS, THE [EMPIRE OF THE EAST 1]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  BROTHER ASSASSIN [BERSERKER 2]  C/P:2,P
SABERHAGEN, FRED  BROTHER BERSERKER [BERSERKER 2]  H
SABERHAGEN, FRED  CENTURY OF PROGRESS, A  C/P:2,SFWA-P,P
SABERHAGEN, FRED  CHANGELING EARTH [EMPIRE OF THE EAST 3]  C/P:2,P
SABERHAGEN, FRED/ZELAZNY, ROGER  COILS  H:2
SABERHAGEN, FRED  COINSPINNER’S STORY [@LOST SWORDS #5]  C/P,SFWA-P,P,H
SABERHAGEN, FRED  COLDNESS IN THE BLOOD, A [DRACULA]  H
SABERHAGEN, FRED  COMPLETE BOOK OF SWORDS, THE [@SWORDS #1,2,3]  H:2
SABERHAGEN, FRED  DANCING BEARS  H:2,P
SABERHAGEN, FRED  DOMINION [DRACULA]  C/P:2,P
SABERHAGEN, FRED  DRACULA TAPE, THE [DRACULA]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  EARTH DESCENDED  C/P:2,SFWA-P,P
SABERHAGEN, FRED  EMPIRE OF THE EAST [EMPIRE OF THE EAST 1,2,3]  H:2,SFWA-P,P
SABERHAGEN, FRED  ENDGAME [@LOST SWORDS #7,8]  H:2
SABERHAGEN, FRED  ERSTE BUCH DER SCHWERTER, DAS  FP
SABERHAGEN, FRED  FACE OF APOLLO, THE [GODS 1]  C/P,LP
SABERHAGEN, FRED  FARSLAYER’S STORY [@LOST SWORDS #4]  H:2,SFWA-P,P
SABERHAGEN, FRED  FIRST BOOK OF SWORDS, THE [@SWORDS #1]  C/P,LP,H
SABERHAGEN, FRED  FIRST TRIAD, THE [@LOST SWORDS #1,2,3]  H:2
SABERHAGEN, FRED  FRANKENSTEIN PAPERS, THE  C/P:2,P
SABERHAGEN, FRED  GOD OF THE GOLDEN FLEECE [GODS 4]  H
SABERHAGEN, FRED  GODS OF FIRE AND THUNDER [GODS 5]  H
SABERHAGEN, FRED  GOLDEN PEOPLE, THE  C/P:2,DM103,P
SABERHAGEN, FRED  HOLMES-DRACULA FILE, THE [DRACULA]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  LOVE CONQUERS ALL  C/P:2,P

SABERHAGEN, FRED/ GREENBERG, MARTIN H.

MACHINES THAT KILL

C/PA,PA

SABERHAGEN, FRED  MASK OF THE SUN, THE  C/P:2,SFWA-P,P
SABERHAGEN, FRED  MATTER OF TASTE, A [DRACULA]  C/P,P,H
SABERHAGEN, FRED  MERLIN’S BONES  C/P,P,H
SABERHAGEN, FRED  MINDSWORD’S STORY [@LOST SWORDS #6]  C/P,SFWA,P,H
SABERHAGEN, FRED  OCTAGON  C/P:2,P
SABERHAGEN, FRED  OLD FRIEND OF THE FAMILY, AN [DRACULA]  C/P:2,SFWA-P,P
SABERHAGEN, FRED/SABERHAGEN, JOAN  PAWN TO INFINITY  C/PA,PA
SABERHAGEN, FRED  PILGRIM [PILGRIM 1,2]  P
SABERHAGEN, FRED  PYRAMIDS [PILGRIM 1]  C/P:2,P
SABERHAGEN, FRED  QUESTION OF TIME, A [DRACULA]  C/P,P,H
SABERHAGEN, FRED  ROGUE BERSERKER [BERSERKER]  H
SABERHAGEN, FRED  SABERHAGEN: MY BEST  C/P,P
SABERHAGEN, FRED  SEANCE FOR A VAMPIRE [DRACULA]  H,P
SABERHAGEN, FRED  SECOND BOOK OF SWORDS, THE [@SWORDS #2]  C/P,P,LP,H
SABERHAGEN, FRED  SECOND TRIAD, THE [@LOST SWORDS #4,5,6]  H:2
SABERHAGEN, FRED  SHARPNESS ON THE NECK, A [DRACULA]  H,P
SABERHAGEN, FRED  SHIELDBREAKER’S STORY [@LOST SWORDS #8]  C/P,P,H
SABERHAGEN, FRED  SHIVA IN STEEL [BERSERKER 8]  C/P,P,H
SABERHAGEN, FRED  SIGHTBLINDER’S STORY [@LOST SWORDS #2]  H:2,P
SABERHAGEN, FRED  SPADEFUL OF SPACETIME, A  C/PA,PA

The MITSFS Pinkdex  September 24, 2007  Page 336
AUTHOR  TITLE  CODE
SABERHAGEN, FRED  SPECIMENS  C/P:2,P
SABERHAGEN, FRED  STONECUTTER’S STORY [@LOST SWORDS #3]  H:2,SFWA-P,P
SABERHAGEN, FRED  THIRD BOOK OF SWORDS, THE [@SWORDS #3]  C/P,SFWA-P,P,H
SABERHAGEN, FRED  THORN [DRACULA]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  ULTIMATE ENEMY, THE [BERSERKER 4]  C/P:2,SFWA-P,P
SABERHAGEN, FRED  VEILS OF AZLAROC, THE  C/P:2,SFWA-P,P
SABERHAGEN, FRED  WATER OF THOUGHT, THE  DM127
SABERHAGEN, FRED  WATER OF THOUGHT, THE (REVISED)  C/P:2,SFWA-P,P
SABERHAGEN, FRED  WAYFINDER’S STORY [@LOST SWORDS #7]  H:2
SABERHAGEN, FRED  WHITE BULL, THE  C/P:2,SFWA-P,P
SABERHAGEN, FRED  WOUNDHEALER’S STORY [@LOST SWORDS #1]  H:2,P
SABINE, TED  SOULSUCKER, THE  C/P:2,P
SACKERMAN, HENRY  LOVE BOMB, THE  P
SACKETT, JEFFREY  BLOOD OF THE IMPALER  P,SFWA-P
SACKETT, JEFFREY  DEMON, THE  P
SACKETT, SUSAN  LETTERS TO STAR TREK  R/REF-P,C/REF-P:2
SACKETT, SUSAN/RODDENBERRY, GENE  MAKING OF STAR TREK: THE MOTION PICTURE, THE [STAR TREK]  R/REF-P,C/REF-H:2

SACKETT, SUSAN/GOLSTEIN, FRED/ GOLDSTEIN, STAN STAR TREK SPEAKS

[C/RAN-H]

SADLER, BARRY  AFRICAN MERCENARY, THE [@CASCA #12]  P
SADLER, BARRY  ASSASSIN, THE [@CASCA #13]  P
SADLER, BARRY  BARBARIAN, THE [@CASCA #5]  P
SADLER, BARRY  CONQUISTADOR, THE [@CASCA #10]  P
SADLER, BARRY  CURSED, THE [@CASCA #18]  P
SADLER, BARRY  DAMNED, THE [@CASCA #7]  P
SADLER, BARRY  DESERT MERCENARY [@CASCA #16]  P
SADLER, BARRY  ETERNAL MERCENARY, THE [@CASCA #1]  C/P,P
SADLER, BARRY  GOD OF DEATH [@CASCA #2]  P
SADLER, BARRY  LEGIONNAIRE, THE [@CASCA #11]  P
SADLER, BARRY  MONGOL, THE [@CASCA #22]  P
SADLER, BARRY  PANZER SOLDIER [@CASCA #4]  C/P,P
SADLER, BARRY  PERSIAN, THE [@CASCA #6]  P
SADLER, BARRY  PHOENIX, THE [@CASCA #14]  P
SADLER, BARRY  PIRATE, THE [@CASCA #15]  P
SADLER, BARRY  SAMURAI, THE [@CASCA #19]  P
SADLER, BARRY  SENTINEL, THE [@CASCA #9]  P
SADLER, BARRY  SOLDIER OF FORTUNE [@CASCA #8]  P
SADLER, BARRY  SOLDIER OF GIDEON [@CASCA #20]  P
SADLER, BARRY  TRENCH SOLDIER, THE [@CASCA #21]  P
SADLER, BARRY  WARLORD, THE [@CASCA #3]  C/P,P
SADLER, BARRY  WARRIOR, THE [@CASCA #17]  P
SADOUL, JACQUES  HIER, L’AN 2000  A
SADOUL, JACQUES  UNIVERS 01  FP
SADOUL, JACQUES  UNIVERS 15  FP
SAGAN, CARL  BROCA’S BRAIN  R/REF-P
SAGAN, CARL/DRUYAN, ANN  COMET  [C/RAN-H]
SAGAN, CARL  COMMUNICATION WITH EXTRATERRESTRIAL INTELLIGENCE  [C/RAN-P]
SAGAN, CARL  CONTACT  C/P,P,H:2
SAGAN, CARL  COSMIC CONNECTION, THE  [C/RAN-P:2]
SAGAN, CARL  DRAGONS OF EDEN, THE  [C/RAN-P:2]
SAGAN, CARL  MURMURS OF EARTH  SFWA
SAGAN, CARL  OTHER WORLDS  [C/RAN-P]
SAGAN, NICK  EDENBORN  LP
SAGAN, NICK  EVERFREE [IDLEWILD 3]  H
SAGAN, NICK  IDLEWILD [IDLEWILD 1]  H
SAGAN, NICK  IDLEWILD (PROOF) [IDLEWILD 1]  SR-LP
SAGARA, MICHELLE  CAST IN COURTLIGHT [ELANTRA 2]  LP
SAGARA, MICHELLE  CAST IN SECRET [ELANTRA 3]  LP
SAGARA, MICHELLE  CAST IN SHADOW [ELANTRA 1]  P
SAGARA, MICHELLE  CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  P
SAGARA, MICHELLE  CHILDREN OF THE BLOOD [SUNDERED 2]  P,LP
SAGARA, MICHELLE  INTO THE DARK LANDS [SUNDERED 1]  P,LP
SAGARA, MICHELLE  LADY OF MERCY [SUNDERED 3]  P
SAGE, ANGIE  MAGYK [SEPTIMUS HEAP 1]  H
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 7, THE  PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 8, THE  C/PA,PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 9, THE  PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 10, THE  PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 11, THE  PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 12, THE  PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 13, THE  HA,PA
SAHA, ARTHUR W.  YEAR’S BEST FANTASY STORIES 14, THE  C/PA,PA
SAID, S. F.  VARJAK PAW  LP
SAILOR, CHARLES  SECOND SON, THE  H

AUTHOR  TITLE  CODE
SAINT ALCORN, LLOYD  HALBERD, DREAM WARRIOR [DREAMQUEST 1]  P
SAINT ALCORN, LLOYD  ON THE SHOULDERS OF GIANTS [DREAMQUEST 2]  P
SAINT ARMAND, BARTON LEVI  ROOTS OF HORROR IN THE FICTION OF H. P. LOVECRAFT, THE  R/REF-H
SAINT CLAIR, MARGARET  AGENT OF THE UNKNOWN  DD150
SAINT CLAIR, MARGARET  CHANGE THE SKY AND OTHER STORIES  C/P,P
SAINT CLAIR, MARGARET  DANCERS OF NOYO, THE  C/P,P
SAINT CLAIR, MARGARET  DOLPHINS OF ALTAIR, THE  C/P:2,P
SAINT CLAIR, MARGARET  GAMES OF NEITH, THE  DD453
SAINT CLAIR, MARGARET  GREEN QUEEN, THE  DD176
SAINT CLAIR, MARGARET  MESSAGE FROM THE EOCENE  DM105
SAINT CLAIR, MARGARET  SHADOW PEOPLE, THE  C/P:2,P
SAINT CLAIR, MARGARET  SIGN OF THE LABRYS  C/P:2,P
SAINT CLAIR, MARGARET  THREE WORLDS OF FUTURITY  DM105
SAINT EXUPERY, ANTOINE DE  LITTLE PRINCE, THE  P
SAINT, H. F.  MEMOIRS OF AN INVISIBLE MAN  H,SFWA
SAINT JOHN, DAVID  COVEN, THE  H
SAINT JOHN, DAVID  DIABOLUS  P
SAINT JOHN, DAVID  VENUS PROBE, THE  P
SAINT JOHN, PHILIP  ROCKET JOCKEY  H
SAINT JOHN, PHILIP  ROCKET PILOT  H
SAINT JOHN, PHILIP  ROCKETS TO NOWHERE  H:2
SAINT, PAUL  SUNS OF CARESH, THE [DOCTOR WHO PAST DOCTORS]  P
SAINTCROW, LILITH  CLOUD WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  LP
SAINTCROW, LILITH  DARK WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  LP
SAINTCROW, LILITH  DEAD MAN RISING [DANTE VALENTINE 2]  P
SAINTCROW, LILITH  FIRE WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  LP
SAINTCROW, LILITH  SOCIETY, THE  LP
SAINTCROW, LILITH  STORM WATCHER [WATCHERS (CIRCLE LIGHTFALL)]  LP
SAINTCROW, LILITH  WORKING FOR THE DEVIL [DANTE VALENTINE 1]  P
SAKERS, DON  CARMEN MIRANDA’S GHOST IS HAUNTING SPACE STATION THREE  PA,C/PA,SFWA-P
SAKERS, DON  LEAVES OF OCTOBER, THE  C/P,SFWA-P,P
SAKI  BEST OF SAKI, THE  LP
SAKI  HUMOR HORROR AND THE SUPERNATURAL  P
SAKI  INCREDIBLE TALES  P
SALE, RICHARD  NOT TOO NARROW...NOT TOO DEEP  SR-P
SALISBURY, JOHN W.  INTRODUCTION TO THE MOON, AN  [C/RAN-H]
SALLIS, JAMES  ASH OF STARS  R/REF-P
SALLIS, JAMES  FEW LAST WORDS, A  H
SALLIS, JAMES  SHORES BENEATH, THE  PA,C/PA
SALMONSON, JESSICA AMANDA  AMAZONS!  C/PA,PA
SALMONSON, JESSICA AMANDA  AMAZONS II  PA,C/PA
SALMONSON, JESSICA AMANDA  GOLDEN NAGINATA, THE [TOMOE GOZEN 2]  C/P:2,P
SALMONSON, JESSICA AMANDA  HEROIC VISIONS  PA,C/PA
SALMONSON, JESSICA AMANDA  HEROIC VISIONS II  PA
SALMONSON, JESSICA AMANDA  OU LU KHEN AND THE BEAUTIFUL MADWOMAN  C/P:2,P
SALMONSON, JESSICA AMANDA  SWORDSWOMAN, THE  C/P,SFWA-P,H:2,P
SALMONSON, JESSICA AMANDA  TALES BY MOONLIGHT  SR-HA,PA,SFWA-P
SALMONSON, JESSICA AMANDA  TALES BY MOONLIGHT II  PA,SFWA-P
SALMONSON, JESSICA AMANDA  THOUSAND SHRINE WARRIOR [TOMOE GOZEN 3]  C/P:2,P
SALMONSON, JESSICA AMANDA  TOMOE GOZEN [TOMOE GOZEN 1]  C/P:2,P
SALMONSON, JESSICA AMANDA  WHAT DID MISS DARLINGTON SEE?  LP
SALSITZ, R. A. V.  DAUGHTER OF DESTINY  C/P,P
SALSITZ, R. A. V.  NIGHT OF DRAGONS  C/P,P
SALSITZ, R. A. V.  UNICORN DANCER, THE  P
SALSITZ, R. A. V.  WHERE DRAGONS LIE  P
SALTERBERG, B. J.  CAPTIVITY [OUTLANDER (SALTERBERG) 1]  LP
SALVATORE, R. A.  ASCENDANCE [DEMONWARS II 1]  H,P
SALVATORE, R. A.  ATTACK OF THE CLONES [STAR WARS MOVIE 2]  H
SALVATORE, R. A.  BASTION OF DARKNESS [YNIS AIELLE 3]  P
SALVATORE, R. A.  CANTICLE [@FORGOTTEN REALMS CLERIC #1]  P
SALVATORE, R. A.  CHAOS CURSE, THE [@FORGOTTEN REALMS CLERIC #5]  C/P,P
SALVATORE, R. A.  CRIMSON SHADOW, THE [CRIMSON SHADOW 1,2,3]  LP
SALVATORE, R. A.  CRYSTAL SHARD, THE [@FORGOTTEN REALMS ICEWIND DALE #1]  C/P:2,P
SALVATORE, R. A.  DARK ELF TRILOGY, THE [@FORGOTTEN REALMS DARK ELF]  LP
SALVATORE, R. A.  DEMON APOSTLE, THE [DEMONWARS 3]  C/P,P,H
SALVATORE, R. A.  DEMON AWAKENS, THE [DEMONWARS 1]  C/P,SR-LP,H
SALVATORE, R. A.  DEMON SPIRIT, THE [DEMONWARS 2]  C/P,P
SALVATORE, R. A.  DRAGON KING, THE [CRIMSON SHADOW 3]  P
SALVATORE, R. A.  DRAGON’S DAGGER, THE [SPEARWIELDER 2]  C/P,P
SALVATORE, R. A.  DRAGONSLAYER’S RETURN [SPEARWIELDER 3]  P
SALVATORE, R. A.  ECHOES OF THE FOURTH MAGIC [YNIS AIELLE 1]  P
SALVATORE, R. A.  EXILE [@FORGOTTEN REALMS DARK ELF #2]  C/P,P
SALVATORE, R. A.  FALLEN FORTRESS, THE [@FORGOTTEN REALMS CLERIC #4]  P
SALVATORE, R. A.  HALFLING’S GEM, THE [@FORGOTTEN REALMS ICEWIND DALE #3]  P

The MITSFS Pinkdex  September 24, 2007  Page 338
AUTHOR  TITLE  CODE
SALVATORE, R. A.  HIGHWAYMAN, THE  P
SALVATORE, R. A.  HOMELAND [@FORGOTTEN REALMS DARK ELF #1]  C/P:2,P
SALVATORE, R. A.  IMMORTALIS [DEMONWARS II 3]  H
SALVATORE, R. A.  IN SYLVAN SHADOWS [@FORGOTTEN REALMS CLERIC #2]  P
SALVATORE, R. A.  LEGACY, THE [FORGOTTEN REALMS LEGACY OF THE DROW 1]  C/P,P,H
SALVATORE, R. A.  LONE DROW, THE [@FORGOTTEN REALMS HUNTER’S BLADES #2]  P
SALVATORE, R. A.  LUTHIEN’S GAMBLE [CRIMSON SHADOW 2]  P
SALVATORE, R. A.  MORTALIS [DEMONWARS 4]  H
SALVATORE, R. A.  NIGHT MASKS [@FORGOTTEN REALMS CLERIC #3]  P
SALVATORE, R. A.  PASSAGE TO DAWN [FORGOTTEN REALMS LEGACY OF THE DROW 4]  P
SALVATORE, R. A.  PROMISE OF THE WITCHKING [@FORGOTTEN REALMS SELLSWORDS #2]  P
SALVATORE, R. A.  ROAD OF THE PATRIARCH [@FORGOTTEN REALMS SELLSWORDS #3]  H
SALVATORE, R. A.  SEA OF SWORDS [@FORGOTTEN REALMS]  P
SALVATORE, R. A.  SERVANT OF THE SHARD [@FORGOTTEN REALMS SELLSWORDS #1]  P
SALVATORE, R. A.  SIEGE OF DARKNESS [FORGOTTEN REALMS LEGACY OF THE DROW 3]  P
SALVATORE, R. A.  SILENT BLADE, THE [FORGOTTEN REALMS PATHS OF DARKNESS 1]  H
SALVATORE, R. A.  SOJOURN [@FORGOTTEN REALMS DARK ELF 3]  C/P,P
SALVATORE, R. A.  SPINE OF THE WORLD, THE [@FORGOTTEN REALMS]  P
SALVATORE, R. A.  STARLESS NIGHT [FORGOTTEN REALMS LEGACY OF THE DROW 2]  H,P
SALVATORE, R. A.  STREAMS OF SILVER [@FORGOTTEN REALMS ICEWIND DALE #2]  C/P:2,P
SALVATORE, R. A.  SWORD OF BEDWYR, THE [CRIMSON SHADOW 1]  P
SALVATORE, R. A.  TARZAN: THE EPIC ADVENTURES [TARZAN]  P
SALVATORE, R. A.  THOUSAND ORCS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #1]  P
SALVATORE, R. A.  TRANSCENDENCE [DEMONWARS II 2]  H
SALVATORE, R. A.  TWO SWORDS, THE [@FORGOTTEN REALMS HUNTER’S BLADES #3]  H
SALVATORE, R. A.  VECTOR PRIME [@STAR WARS NEW JEDI ORDER]  P
SALVATORE, R. A.  WITCH’S DAUGHTER, THE [YNIS AIELLE 2]  P
SALVATORE, R. A.  WOODS OUT BACK, THE [SPEARWIELDER 1]  C/P:2,SFWA-P,P
SAMBROT, WILLIAM  ISLAND OF FEAR AND OTHER SCIENCE FICTION STORIES  C/P:2,P
SAMMS, TARA  FRAYED [DOCTOR WHO]  SR-H

SAMPSON, JEFF DRAGON SPELL [@DRAGONLANCE NEW ADVENTURES #8/

DRAGONLANCE DRAGON QUARTET 4]

LP

SAMPSON, JEFF  EBONY EYE, THE [@DRAGONLANCE NEW ADVENTURES SUNCATCHER #2]  LP

SAMPSON, JEFF WAYWARD WIZARD, THE

[@DRAGONLANCE NEW ADVENTURES SUNCATCHER #1]

LP

SAMPSON, JEFF WIZARD’S BETRAYAL

[@DRAGONLANCE NEW ADVENTURES TRINISTYR #2]

LP

SAMUELS, VICTOR  VAMPIRE WOMEN, THE  C/P,P
SAN SOUCCI, ROBERT D.  MORE SHORT & SHIVERY  H
SAN SOUCI, ROBERT D.  EMERGENCE  P
SANBORN, ROBIN  BOOK OF STIER, THE  C/P:2,P
SANDER, GORDON F.  SERLING  R/REF-H
SANDERS, FRANKLIN  HEILAND  P
SANDERS, LAWRENCE  PASSION OF MOLLY T., THE  P
SANDERS, LAWRENCE  TOMORROW FILE, THE  C/P:2,P
SANDERS, N. K.  EPIC OF GILGAMESH, THE  P
SANDERS, SCOTT RUSSELL  INVISIBLE COMPANY, THE  P,SFWA-P
SANDERS, SCOTT RUSSELL  TERRARIUM  C/P,SFWA-P,P
SANDERS, WILLIAM  ARE WE HAVING FUN YET?  LP
SANDERS, WILLIAM  BALLAD OF BILLY BADASS AND THE ROSE OF TURKESTAN, THE  LP
SANDERS, WILLIAM  JOURNEY TO FUSANG  P
SANDERS, WILLIAM  WILD BLUE AND THE GRAY, THE  P
SANDERSON, BRANDON  ELANTRIS  H
SANDERSON, BRANDON  MISTBORN [MISTBORN 1]  H
SANDERSON, IVAN T.  INVISIBLE RESIDENTS  [C/RAN-P]
SANSWEET, STEPEHN J.  QUOTABLE STAR WARS, THE  R/REF-P
SANT, MONTE  BOOK OF THE DRAGON, THE  A
SANTESSON, HANS STEFAN  CRIME PREVENTION IN THE 30TH CENTURY  HA
SANTESSON, HANS STEFAN  DAYS AFTER TOMORROW, THE  HA
SANTESSON, HANS STEFAN  FANTASTIC UNIVERSE OMNIBUS, THE  C/PA,PA
SANTESSON, HANS STEFAN  FLYING SAUCERS IN FACT AND FICTION  PA
SANTESSON, HANS STEFAN  GENTLE INVADERS  C/PA,PA
SANTESSON, HANS STEFAN  GODS FOR TOMORROW  PA
SANTESSON, HANS STEFAN  MIGHTY BARBARIANS, THE  PA
SANTESSON, HANS STEFAN  MIGHTY SWORDSMEN, THE  PA
SANTESSON, HANS STEFAN  RULERS OF MEN  PA,C/PA
SANTESSON, HANS STEFAN  UNDERSTANDING MU  [C/RAN-P:2]
SAPERSTEIN, DAVID  COCOON [COCOON 1]  C/P,P
SAPERSTEIN, DAVID  METAMORPHOSIS [COCOON 2]  P
SAPIR, RICHARD BEN/MURPHY, WARREN  DR. QUAKE [DESTROYER 5]  P
SAPIR, RICHARD BEN  FAR ARENA, THE  P
SARABONDE, WILLIAM  EDGE OF THE WORLD, THE [FIRST AMERICANS]  P
SARAC, ROGER  THROWBACKS, THE  P
SARBAN  DOLL MAKER, THE  P

AUTHOR  TITLE  CODE
SARBAN  RINGSTONES  P
SARBAN  SOUND OF HIS HORN, THE  C/P:2,P
SARGENT, CARL/GASCOIGNE, MARC  BLACK MADONNA [@SHADOWRUN 20]  P
SARGENT, CARL/GASCOIGNE, MARC  NOSFERATU [SHADOWRUN 14]  C/P,P
SARGENT, CARL/GASCOIGNE, MARC  STREETS OF BLOOD [@SHADOWRUN 8]  C/P,P
SARGENT, CRAIG  CUTTHROAT CANNIBALS, THE [@LAST RANGER #8]  P
SARGENT, CRAIG  DAMNED DISCIPLES, THE [@LAST RANGER #9]  P
SARGENT, CRAIG  VILE VILLAGE, THE [@LAST RANGER #7]  P
SARGENT, PAMELA/ZEBROWSKI, GEORGE  ACROSS THE UNIVERSE [STAR TREK]  C/P,P
SARGENT, PAMELA/WATSON, IAN  AFTERLIVES  PA
SARGENT, PAMELA  ALIEN UPSTAIRS, THE  P,SFWA
SARGENT, PAMELA  BEST OF PAMELA SARGENT, THE  P
SARGENT, PAMELA  BIO-FUTURES  C/PA:2,PA
SARGENT, PAMELA  CHILD OF VENUS [VENUS OF DREAMS 3]  P
SARGENT, PAMELA  CLIMB THE WIND  P
SARGENT, PAMELA  CLONED LIVES  C/P:2,P
SARGENT, PAMELA  CONQUEROR FANTASTIC  PA
SARGENT, PAMELA  EARTHSEED  C/P:2,LP
SARGENT, PAMELA/ZEBROWSKI, GEORGE  FURY SCORNED, A [STAR TREK TNG]  C/P:2,P
SARGENT, PAMELA/ZEBROWSKI, GEORGE  GARTH OF IZAR [STAR TREK]  P
SARGENT, PAMELA  GOLDEN SPACE, THE  H:2,SFWA,P
SARGENT, PAMELA/ZEBROWSKI, GEORGE  HEART OF THE SUN [STAR TREK]  C/P,P
SARGENT, PAMELA  HOMESMIND  LP
SARGENT, PAMELA  MORE WOMEN OF WONDER  C/PA,PA
SARGENT, PAMELA  NEBULA AWARDS 29  LPA
SARGENT, PAMELA  NEBULA AWARDS 30  LPA
SARGENT, PAMELA  NEBULA AWARDS 31  LPA
SARGENT, PAMELA  NEW WOMEN OF WONDER, THE  C/PA,PA
SARGENT, PAMELA  SHORE OF WOMEN, THE  C/P,SFWA,P,H:2
SARGENT, PAMELA  STARSHADOWS  C/P:2,P
SARGENT, PAMELA  SUDDEN STAR, THE  C/P,P
SARGENT, PAMELA  THUMBPRINTS  H
SARGENT, PAMELA  VENUS OF DREAMS [VENUS OF DREAMS 1]  C/P:2,P
SARGENT, PAMELA  VENUS OF SHADOWS [VENUS OF DREAMS 2]  C/P,P,H
SARGENT, PAMELA  WATCHSTAR  H:2,P
SARGENT, PAMELA  WOMEN OF WONDER  C/PA:2,PA
SARGENT, PAMELA  WOMEN OF WONDER: THE CLASSIC YEARS  LPA
SARGENT, PAMELA  WOMEN OF WONDER: THE CONTEMPORARY YEARS  HA,LPA
SARGENT, SARAH  JONAS MCFEE, A.T.P.  SFWA
SARGENT, SARAH  WEIRD HENRY BERG  H
SARGENT, STANLEY C.  ANCIENT EXHUMATIONS +2  LP
SARGENT, STANLEY C.  TAINT OF LOVECRAFT, THE  LP
SARRANTONIO, AL  CAMPBELL WOOD  H,SFWA
SARRANTONIO, AL  EXILE [@FIVE WORLDS #1]  P
SARRANTONIO, AL  FLIGHTS  HA
SARRANTONIO, AL  HAYDN OF MARS [CATS OF MARS 1]  P
SARRANTONIO, AL  HOUSE HAUNTED  P
SARRANTONIO, AL  JOURNEY [@FIVE WORLDS #2]  P
SARRANTONIO, AL  MOONBANE  C/P,SFWA-P,P
SARRANTONIO, AL/GREENBERG, MARTIN H. 100 HAIR-RAISING LITTLE HORROR STORIES  HA
SARRANTONIO, AL  PERSONAL AGENDAS [@BABYLON 5 #8]  C/P,P
SARRANTONIO, AL  QUEEN OF MARS [CATS OF MARS 3]  P
SARRANTONIO, AL  REDSHIFT  HA
SARRANTONIO, AL  RETURN [@FIVE WORLDS #3]  P
SARRANTONIO, AL  SEBASTIAN OF MARS [CATS OF MARS 2]  P
SARRANTONIO, AL  WORMS, THE  SFWA
SARTI, RON  CHRONICLES OF SCAR, THE [SCAR 1]  C/P,P
SARTI, RON  LANTERNS OF GOD, THE [@SCAR #3]  P
SARTI, RON  LEGACY OF THE ANCIENTS [@SCAR #2]  P
SATURDAY EVENING POST  POST READER OF FANTASY AND SCIENCE FICTION, THE  C/PA,PA
SATURDAY EVENING POST  SATURDAY EVENING POST STORIES 1953, THE  HA
SAUER, ROB  VOYAGES  C/PA,PA
SAUL, GEORGE BRANDON  OWLS’ WATCH  C/PA,PA
SAUL, JOHN  ASHES TO ASHES: THE DRAGON’S FLAME [@BLACKSTONE #3]  P
SAUL, JOHN  ASYLUM [@BLACKSTONE #6]  P
SAUL, JOHN  BLACK LIGHTNING  P
SAUL, JOHN  BRAINCHILD  C/P,P
SAUL, JOHN  COMES THE BLIND FURY  C/P,P,H
SAUL, JOHN  CREATURE  P
SAUL, JOHN  CRY FOR THE STRANGERS  P
SAUL, JOHN  DAY OF RECKONING: THE STEREOSCOPE [@BLACKSTONE #5]  C/P,P
SAUL, JOHN  EYE FOR AN EYE: THE DOLL, AN [@BLACKSTONE #1]  P
SAUL, JOHN  GOD PROJECT, THE  H:2
SAUL, JOHN  GUARDIAN  H

AUTHOR  TITLE  CODE
SAUL, JOHN  HELLFIRE  P
SAUL, JOHN  IN THE SHADOW OF EVIL: THE HANDKERCHIEF [@BLACKSTONE #4]  C/P,P
SAUL, JOHN  NATHANIEL  C/P,P
SAUL, JOHN  PUNISH THE SINNERS  P
SAUL, JOHN  SECOND CHILD  P
SAUL, JOHN  SHADOWS  P
SAUL, JOHN  SLEEPWALK  P
SAUL, JOHN  SUFFER THE CHILDREN  P
SAUL, JOHN  TWIST OF FATE: THE LOCKET [@BLACKSTONE #2]  P
SAUL, JOHN  UNLOVED, THE  P
SAUL, JOHN  WHEN THE WIND BLOWS  H,P
SAUNDERS, CHARLES R.  IMARO [IMARO 1]  C/P,P
SAUNDERS, CHARLES R.  QUEST FOR CUSH, THE [IMARO 2]  C/P:2,P
SAUNDERS, CHARLES R.  TRAIL OF BOHU, THE [IMARO 3]  P
SAUNDERS, ELIZABETH A.  WHEN THE BLACK LOTUS BLOOMS  SR-LPA
SAUNDERS, GEORGE  CIVILWARLAND IN BAD DECLINE  LP
SAUNDERS, JAKE/WALDROP, HOWARD  TEXAS-ISRAELI WAR: 1999, THE  C/P:2,P
SAVAGE, BLAKE  ASSIGNMENT IN SPACE WITH RIP FOSTER  H
SAVAGE, BLAKE  RIDE THE GRAY PLANET  C/P,P
SAVAGE, FELICITY  DELTA CITY [HUMILITY GARDEN 2]  P
SAVAGE, FELICITY  EVER: THE WAR IN THE WASTE [EVER 1]  LP
SAVAGE, FELICITY  GARDEN OF SALT [HUMILITY GARDEN 1,2]  H
SAVAGE, FELICITY  HUMILITY GARDEN [HUMILITY GARDEN 1]  C/P,P
SAVAGE, HARDLEY  JETMAN MEETS THE MAD MADAM  SR-L
SAVAGE, MARY  LIKENESS TO VOICES, A  P
SAVAGE, RICHARD  WHEN THE MOON DIED  SR-P
SAVARIN, JULIAN JAY  ARCHIVES OF HAVEN, THE [LEMMUS #3]  @SR-P,H
SAVARIN, JULIAN JAY  BEYOND THE OUTER MIRR [LEMMUS #2]  @SR-P,H
SAVARIN, JULIAN JAY  WAITERS ON THE DANCE [LEMMUS 1]  H
SAVCHENKO, VLADIMIR  SELF-DISCOVERY  LP
SAVILE, STEVEN  DOMINION [WARHAMMER]  P
SAVILE, STEVEN/ONTIS, ALTHEA  ELEMENTAL  HA
SAVILE, STEVEN  RETRIBUTION [@WARHAMMER VON CARSTEIN 3]  P
SAWARD, ERIC  DOCTOR WHO AND THE VISITATION [DOCTOR WHO]  SR-P
SAWARD, ERIC  DOCTOR WHO: ATTACK OF THE CYBERMEN [DOCTOR WHO]  P
SAWARD, ERIC  SLIPBACK [DOCTOR WHO]  SR-P,C/P
SAWARD, ERIC  TWIN DILEMMA, THE [DOCTOR WHO]  SR-P,C/P
SAWYER, ANDY/SEED, DAVID  SPEAKING SCIENCE FICTION  R/REF-P
SAWYER, ROBERT J.  CALCULATING GOD  H,P
SAWYER, ROBERT J.  END OF AN ERA  C/P:2,LP
SAWYER, ROBERT J.  FACTORING HUMANITY  H,P
SAWYER, ROBERT J.  FAR-SEER [DINOSAUR EMPIRE 1]  C/P:2,SFWA-P:2,P
SAWYER, ROBERT J.  FLASHFORWARD  C/P,P,H
SAWYER, ROBERT J.  FOREIGNER [DINOSAUR EMPIRE 3]  P
SAWYER, ROBERT J.  FOSSIL HUNTER [DINOSAUR EMPIRE 2]  P,SFWA-P
SAWYER, ROBERT J.  FRAMESHIFT  H,P
SAWYER, ROBERT J.  GOLDEN FLEECE  H,P
SAWYER, ROBERT J.  HOMINIDS [NEANDERTHAL PARALLAX 1]  H,P
SAWYER, ROBERT J.  HUMANS [NEANDERTHAL PARALLAX 2]  C/P,P
SAWYER, ROBERT J.  HYBRIDS [NEANDERTHAL PARALLAX 3]  H
SAWYER, ROBERT J.  ILLEGAL ALIEN  C/P,P
SAWYER, ROBERT J.  ITERATIONS  LP
SAWYER, ROBERT J.  MINDSCAN  H
SAWYER, ROBERT J.  RELATIVITY  H
SAWYER, ROBERT J.  ROLLBACK  H:2
SAWYER, ROBERT J.  STARPLEX  C/P,P,H
SAWYER, ROBERT J.  TERMINAL EXPERIMENT, THE  H:2,P
SAWYER, ROBERT J./CLINK, CLAYTON  TESSERACTS 6  LPA
SAXON, PETER  CURSE OF RATHLAW, THE [GUARDIANS (SAXON) 0]  C/P,P
SAXON, PETER  DARKEST NIGHT, THE  P
SAXON, PETER  HAUNTING OF ALAN MAIS, THE [GUARDIANS (SAXON) 3]  C/P,P
SAXON, PETER  KILLING BONE, THE [GUARDIANS (SAXON) 1]  P
SAXON, PETER  THROUGH THE DARK CURTAIN [GUARDIANS (SAXON) 4]  P
SAXON, PETER  TORTURER, THE  P
SAXON, PETER  VAMPIRE’S MOON  P
SAXON, PETER  VAMPIRES OF FINISTERE, THE [GUARDIANS (SAXON) 4]  P
SAXON, RICHARD  COSMIC CRUSADE  H
SAXON, RICHARD  FUTURE FOR SALE  SR-P
SAXON, RICHARD  STARS CAME DOWN, THE  H
SAXTON, JOSEPHINE  GROUP FEAST  H:2,SFWA
SAXTON, JOSEPHINE  HIEROS GAMOS OF SAM AND AN SMITH, THE  C/P,P
SAXTON, JOSEPHINE  POWER OF TIME, THE  LP
SAXTON, MARK  HAVOC IN ISLANDIA [ISLANDIA]  H
SAXTON, MARK  ISLAR, THE  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 341
AUTHOR  TITLE  CODE
SAXTON, MARK  TWO KINGDOMS, THE [ISLANDIA]  H:2
SAYERS, DOROTHY L.  HUMAN AND INHUMAN STORIES  PA
SCAEVOLA, PETER  ’68  H
SCALZI, JOHN  AGENT TO THE STARS  SR-H
SCALZI, JOHN  ANDROID’S DREAM, THE  H
SCALZI, JOHN  GHOST BRIGADES, THE [OLD MAN’S WAR 2]  H
SCALZI, JOHN  LAST COLONY, THE [OLD MAN’S WAR 3]  H
SCALZI, JOHN  OLD MAN’S WAR [OLD MAN’S WAR 1]  H:2
SCALZI, JOHN  SAGAN DIARY, THE [OLD MAN’S WAR]  H
SCANLON, MITCHEL  FIFTEEN HOURS [@WARHAMMER 40K]  P
SCARBOROUGH, ELIZABETH ANN  BRONWYN’S BANE [ARGONIA 3]  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  CHANNELING CLEOPATRA  H
SCARBOROUGH, ELIZABETH ANN  CHRISTENING QUEST, THE [ARGONIA 4]  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  DRASTIC DRAGON OF DRACO, TEXAS, THE  P
SCARBOROUGH, ELIZABETH ANN  GODMOTHER, THE [GODMOTHER 1]  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  GODMOTHER’S APPRENTICE, THE [GODMOTHER 2]  C/P,P
SCARBOROUGH, ELIZABETH ANN  GODMOTHER’S WEB, THE [GODMOTHER 3]  H
SCARBOROUGH, ELIZABETH ANN  GOLDCAMP VAMPIRE, THE  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  GOLDCAMP VAMPIRE, THE (PROOF)  SR-LP
SCARBOROUGH, ELIZABETH ANN  HAREM OF AMAN AKBAR, THE  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  HEALER’S WAR, THE  C/P,SFWA-P,P,H
SCARBOROUGH, ELIZABETH ANN  LADY IN THE LOCH, THE  P
SCARBOROUGH, ELIZABETH ANN  LAST REFUGE  C/P,P,H
SCARBOROUGH, ELIZABETH ANN  NOTHING SACRED  H:2,LP,P
SCARBOROUGH, ELIZABETH ANN  PAST LIVES, PRESENT TENSE  LPA
SCARBOROUGH, ELIZABETH ANN  PHANTOM BANJO [@SONGKILLER #1]  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  PICKING THE BALLAD’S BONES [@SONGKILLER #2]  C/P,P
SCARBOROUGH, ELIZABETH ANN  SONG OF SORCERY [ARGONIA 1]  C/P:2,P
SCARBOROUGH, ELIZABETH ANN  SONGS FROM THE SEASHELL ARCHIVES [ARGONIA 1,2]  P
SCARBOROUGH, ELIZABETH ANN  STRUM AGAIN? [@SONGKILLER #3]  P
SCARBOROUGH, ELIZABETH ANN  UNICORN CREED, THE [ARGONIA 2]  C/P:2,P

SCARBOROUGH, ELIZABETH ANN/ GREENBERG, MARTIN H.

WARRIOR PRINCESSES

C/PA,PA

SCHACHNER, NAT  SPACE LAWYER  H
SCHAEFER, FRANK  WHOSE SONG IS SUNG  H,P
SCHAEFFER, SUSAN FROMBERG  DRAGONS OF NORTH CHITTENDON, THE  H,LP
SCHARTZ, VIJAYA  ANAZ-VOOHRI [OPERATION PLEIADES]  P
SCHARTZ, VIJAYA  LOCKDOWN [GARRISON 1]  LP
SCHEALER, JOHN M.  ZIP-ZIP AND HIS FLYING SAUCER  H
SCHEER, K. H./DARLTON, CLARK  AARDE VALT AAN!, DE (DUTCH) [@PERRY RHODAN #4]  FPR
SCHEER, K. H.  BASES SUR VENUS (FRENCH) [@PERRY RHODAN #4]  FPR
SCHEER, K. H./DARLTON, CLARK  FORTERESSE DES SIX LUNES, LA (FRENCH) [@PERRY RHODAN #7]  FPR
SCHEER, K. H.  GUNS OF EVERBLACK, THE [PERRY RHODAN 126]  SR-LP
SCHEER, K. H.  KODEZEICHEN GROSSER BAR  FH
SCHEER, K. H.  KORPS DER VERZWEIFELTEN  FP
SCHEER, K. H.  MILICE DES MUTANTS, LA (FRENCH) [@PERRY RHODAN #3]  FPR
SCHEER, K. H.  OPERATION "ASTREE" (FRENCH) [@PERRY RHODAN #1]  FPR
SCHEER, K. H./DARLTON, CLARK  QUETE COSMIQUE, LA (FRENCH) [@PERRY RHODAN #11]  FPR
SCHEER, K. H.  ROBOT INVITATION, THE [PERRY RHODAN 136]  SR-LP
SCHEER, K. H.  TERRE A PEUR, LA (FRENCH) [@PERRY RHODAN #2]  FPR
SCHEER, K. H.  UBER UNS DAS NICHTS  FP
SCHEER, K. H.  UNENDLICHE RAUM, DER  FP
SCHEER, K. H./DARLTON, CLARK  VAINQUEURS DE VEGA, LES (FRENCH) [@PERRY RHODAN #5]  FPR
SCHENCK, HILBERT  AT THE EYE OF THE OCEAN  P
SCHENCK, HILBERT  CHRONOSEQUENCE  C/P,SFWA-P,P,H
SCHENCK, HILBERT  ROSE FOR ARMAGEDDON, A  P,SFWA-P
SCHENCK, HILBERT  STEAM BIRD  P,SFWA-P
SCHENCK, HILBERT  WAVE RIDER  P
SCHENCK, HILBERT  WAVE RIDER (PROOF)  SR-LP
SCHEND, STEVEN E  BLACKSTAFF [FORGOTTEN REALMS WIZARDS]  P
SCHEURER, CHRISTIAN LORENZ  ENTROPIA: A COLLECTION OF UNUSUALLY RARE STAMPS  A
SCHIFF, STUART DAVID  DEATH  PA
SCHIFF, STUART DAVID  MAD SCIENTISTS  HA
SCHIFF, STUART DAVID  WHISPERS  HA,PA,SFWA
SCHIFF, STUART DAVID  WHISPERS II  HA,SFWA
SCHIFF, STUART DAVID  WHISPERS III  HA,SFWA
SCHIFF, STUART DAVID  WHISPERS IV  HA,SFWA
SCHIFF, STUART DAVID  WHISPERS V  HA,SFWA
SCHIFF, STUART DAVID  WHISPERS VI  HA,SFWA
SCHIFF, STUART DAVID/LEIBER, FRITZ  WORLD FANTASY AWARDS 2, THE  SFWA

SCHIMEL, LAWRENCE/ GREENBERG, MARTIN H.

CAMELOT FANTASTIC

PA

SCHIMEL, LAWRENCE  DRAG QUEEN OF ELFLAND  P

The MITSFS Pinkdex  September 24, 2007  Page 342
AUTHOR  TITLE  CODE

SCHIMEL, LAWRENCE/ GREENBERG, MARTIN H.

FORTUNE TELLER, THE

PA

SCHIMEL, LAWRENCE  THINGS INVISIBLE TO SEE  LPA
SCHLOBIN, ROGER C./HARRISON, IRENE R.  ANDRE NORTON: A PRIMARY AND SECONDARY BIBLIOGRAPHY  R/IND
SCHLOSSER, S. E.  SPOOKY SOUTHWEST  LP
SCHMIDT, DENNIS  CITY OF CRYSTAL SHADOW [QUESTIONER 2]  P
SCHMIDT, DENNIS  DARK PARADISE [QUESTIONER 3]  P
SCHMIDT, DENNIS  FIRST NAME, THE [@TWILIGHT OF THE GODS #1]  C/P:2,P
SCHMIDT, DENNIS  GROA’S OTHER EYE [@TWILIGHT OF THE GODS #2]  C/P:2,P
SCHMIDT, DENNIS  KENSHO  C/P:2,P
SCHMIDT, DENNIS  LABYRINTH [QUESTIONER 1]  P
SCHMIDT, DENNIS  SATORI  C/P,P
SCHMIDT, DENNIS  THREE TRUMPS SOUNDING [@TWILIGHT OF THE GODS #3]  C/P,SFWA-P,P
SCHMIDT, DENNIS  WANDERER  C/P,P
SCHMIDT, DENNIS  WAY-FARER  C/P:2,P
SCHMIDT, STANLEY  ALIENS FROM ANALOG  HA,SFWA
SCHMIDT, STANLEY  ANALOG: WRITER’S CHOICE II  SFWA
SCHMIDT, STANLEY  ANALOG YEARBOOK II  C/PA,PA
SCHMIDT, STANLEY  ARGONAUT  H
SCHMIDT, STANLEY  FROM MIND TO MIND  SFWA
SCHMIDT, STANLEY  LIFEBOAT EARTH  C/P,P
SCHMIDT, STANLEY  NEWTON AND THE QUASI-APPLE  C/P,P,H
SCHMIDT, STANLEY  READERS’ FAVORITES FROM ANALOG 1980 TO 1990  B
SCHMIDT, STANLEY  SINS OF THE FATHERS, THE  P
SCHMIDT, STANLEY  TWEEDLIOOP  H,SFWA,SFWA-P,P
SCHMIDT, STANLEY  UNKNOWN  PA,SFWA-P
SCHMIDT, STANLEY/GREENBERG, MARTIN H. UNKNOWN WORLDS  HA,LPA
SCHMIDT, STANLEY  WHICH WAY TO THE FUTURE?  R/REF-H
SCHMITZ, JAMES H.  AGENT OF VEGA  C/P,P,H
SCHMITZ, JAMES H.  AGENT OF VEGA AND OTHER STORIES  P
SCHMITZ, JAMES H.  BEST OF JAMES H. SCHMITZ, THE  SR-H,H
SCHMITZ, JAMES H.  DEMON BREED, THE  C/P:2,P,H
SCHMITZ, JAMES H.  ETERNAL FRONTIER  LP
SCHMITZ, JAMES H.  ETERNAL FRONTIERS, THE  H:2,P

SCHMITZ, JAMES H./FLINT, ERIC/ GORDON, GUY HUB: DANGEROUS TERRITORY, THE [TELZEY]

C/P,P

SCHMITZ, JAMES H.  LEGACY  C/P:2,P
SCHMITZ, JAMES H.  LION GAME, THE [TELZEY 3]  C/P:2,P
SCHMITZ, JAMES H.  PRIDE OF MONSTERS, A  H:2,P
SCHMITZ, JAMES H.  TALE OF TWO CLOCKS, A  H:2,P
SCHMITZ, JAMES H./FLINT, ERIC  TELZEY AMBERDON [TELZEY]  P
SCHMITZ, JAMES H.  TELZEY TOY, THE [TELZEY 2]  C/P:2,P

SCHMITZ, JAMES H./FLINT, ERIC/ GORDON, GUY T’NT: TELZEY & TRIGGER [TELZEY]

C/P,P

SCHMITZ, JAMES H./FLINT, ERIC  TRIGGER & FRIENDS [TELZEY]  C/P,P
SCHMITZ, JAMES H.  UNIVERSE AGAINST HER, THE [TELZEY 1]  C/P:2,P
SCHMITZ, JAMES H.  WITCHES OF KARRES, THE [WITCHES OF KARRES 1]  H:2,P
SCHNEIDER, JOHN G.  GOLDEN KAZOO, THE  P
SCHNEIDER, PAUL  BALANCE OF TERROR [STAR TREK]  SR-S
SCHOCHET, VICTORIA/SILBERSACK, JOHN  BERKLEY SHOWCASE 1, THE  PA
SCHOCHET, VICTORIA/SILBERSACK, JOHN  BERKLEY SHOWCASE 2, THE  PA
SCHOCHET, VICTORIA/SILBERSACK, JOHN  BERKLEY SHOWCASE 3, THE  C/PA,PA
SCHOCHET, VICTORIA/SILBERSACK, JOHN  BERKLEY SHOWCASE 4, THE  PA
SCHOCHET, VICTORIA/SINGER, MELISSA  BERKLEY SHOWCASE 5, THE  C/PA,PA
SCHOFIELD, SANDY  BIG GAME [PREDATOR]  P
SCHOFIELD, SANDY  BIG GAME, THE [STAR TREK DS9]  C/P:2,P
SCHOFIELD, SANDY  BIG GAME, THE (PROOF) [STAR TREK DS9]  SR-VLP
SCHOFIELD, SANDY  ROGUE [@ALIENS 6]  C/P,P
SCHOLASTIC BOOKS  STORIES OF SUSPENSE  PA
SCHOLES, ROBERT/RABKIN, ERIC S.  SCIENCE FICTION: HISTORY — SCIENCE — VISION  R/REF-P,C/REF-H
SCHOLES, ROBERT  STRUCTURAL FABULATION  R/REF-P
SCHOLZ, CARTER  AMOUNT TO CARRY, THE  LP
SCHOLZ, CARTER/HARCOURT, GLENN  PALIMPSESTS  P
SCHOLZ, CARTER  RADIANCE  H,LP
SCHOONOVER, LAWRENCE  CENTRAL PASSAGE  H:2,P
SCHOW, DAVID J.  KILL RIFF, THE  SR-LP,SFWA-P
SCHOW, DAVID J./FRENTZEN, JEFFREY  OUTER LIMITS: THE OFFICIAL COMPANION, THE  R/REF-H
SCHOW, DAVID J.  SEEING RED  SFWA-P
SCHOW, DAVID J.  SILVER SCREAM  SFWA-P
SCHREIBER, HARVEY K.  EAGLE AND THE SWORD, THE  P
SCHROEDER, KARL  ENGINE OF RECALL, THE  H
SCHROEDER, KARL  LADY OF MAZES  H:2
SCHROEDER, KARL  PERMANENCE  H
SCHROEDER, KARL  SUN OF SUNS [VIRGA 1]  H
SCHROEDER, KARL  VENTUS  H,P

The MITSFS Pinkdex  September 24, 2007  Page 343
AUTHOR  TITLE  CODE
SCHUITEN, FRANCOIS/PEETERS, BENOIT  FEVER IN URBICAND  CX
SCHULMAN, J. NEIL  ALONGSIDE NIGHT  C/P:2,P
SCHULMAN, J. NEIL  RAINBOW CADENZA, THE  C/P,SFWA,P,H
SCHUSTER, HAL  MAKING OF QUANTUM LEAP, THE [QUANTUM LEAP]  R/REF-P
SCHUSTER, HAL/RATHBONE, WENDY  TREK: THE UNAUTHORIZED A–Z [STAR TREK]  R/REF-P,C/REF-P
SCHUTTE, JAMES E.  BUNYIP ARCHIVES, THE  H
SCHWADER, ANN K.  STRANGE STARS & ALIEN SHADOWS  H
SCHWARTZ, ALAN  WANDERING TELLURIAN, THE  DH020
SCHWARTZ, JULIUS/THOMSEN, BRIAN M.  MAN OF TWO WORLDS  R/REF-P
SCHWARTZ, SHEILA  EARTH IN TRANSIT  C/PA,PA
SCHWEIGHOFER, PETER  TALES FROM THE EMPIRE [@STAR WARS]  PA
SCHWEIGHOFER, PETER/CAREY, CRAIG  TALES FROM THE NEW REPUBLIC [@STAR WARS]  PA
SCHWEITZER, DARRELL  CONAN’S WORLD AND ROBERT E. HOWARD  R/REF-P
SCHWEITZER, DARRELL  EXPLORING FANTASY WORLDS  R/REF-P
SCHWEITZER, DARRELL  MASK OF THE SORCERER, THE  H,P
SCHWEITZER, DARRELL  MEANING OF LIFE, THE  B

SCHWEITZER, DARRELL/ VAN HOLLANDER, JASON NECROMANCIES AND NETHERWORLDS

SR-H

SCHWEITZER, DARRELL  SF VOICES  R/REF-P
SCHWEITZER, DARRELL  SHATTERED GODDESS, THE  H,LP
SCHWEITZER, DARRELL  TOM O’BEDLAM’S NIGHT OUT  H
SCHWEITZER, DARRELL  WE ARE ALL LEGENDS  H,LP

SCIENCE AND TECHNOLOGY INFORMATION CENTER, NATIONAL SCIENCE COUNCIL, REPUBLIC OF CHINA SCIENTIFIC RESEARCH ABSTRACTS IN REPUBLIC OF CHINA I: ABSTRACTS

[C/RAN-H]

SCIENCE AND TECHNOLOGY INFORMATION CENTER, NATIONAL SCIENCE COUNCIL, REPUBLIC OF CHINA SCIENTIFIC RESEARCH ABSTRACTS IN REPUBLIC OF CHINA II: INDEXES

[C/RAN-H]

SCIENCE APPLICATIONS INTERNATIONAL CORP.

BEYOND (2003 CALENDAR)

CAL

SCIENCE FICTION BOOK CLUB  VISIONS 2002 CALENDAR  CAL
SCITHERS, GEORGE/SCHWEITZER, DARRELL ANOTHER ROUND AT THE SPACEPORT BAR  PA,C/PA
SCITHERS, GEORGE  ASIMOV’S CHOICE: ASTRONAUTS AND ANDROIDS  PA
SCITHERS, GEORGE  ASIMOV’S CHOICE: BLACK HOLES AND BUG-EYED MONSTERS  PA
SCITHERS, GEORGE  ASIMOV’S CHOICE: COMETS AND COMPUTERS  PA
SCITHERS, GEORGE  ASIMOV’S CHOICE: DARK STARS AND DRAGONS  PA
SCITHERS, GEORGE  ASIMOV’S CHOICE: EXTRATERRESTRIALS AND ECLIPSES  PA
SCITHERS, GEORGE  ISAAC ASIMOV’S ADVENTURES OF SCIENCE FICTION  HA
SCITHERS, GEORGE  ISAAC ASIMOV’S MASTERS OF SCIENCE FICTION  HA
SCITHERS, GEORGE  ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 1  LPA
SCITHERS, GEORGE  ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 2  PA
SCITHERS, GEORGE  ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 3  PA
SCITHERS, GEORGE  ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 4  C/PA,PA
SCITHERS, GEORGE  ISAAC ASIMOV’S SCIENCE FICTION ANTHOLOGY 5  PA
SCITHERS, GEORGE/SCHWEITZER, DARRELL TALES FROM THE SPACEPORT BAR  PA,C/PA
SCLIAR, MOACYR  CENTAUR IN THE GARDEN, THE  LP
SCORTIA, THOMAS N.  ARTERY OF FIRE  H:2,SFWA,P
SCORTIA, THOMAS N.  BEST OF THOMAS N. SCORTIA, THE  H,SFWA
SCORTIA, THOMAS N.  CAUTION! INFLAMMABLE!  H,SFWA,P
SCORTIA, THOMAS N.  EARTHWRECK!  C/P:2,P
SCORTIA, THOMAS N./ZEBROWSKI, GEORGE HUMAN-MACHINES  PA,C/PA
SCORTIA, THOMAS N./ROBINSON, FRANK M. NIGHTMARE FACTOR, THE  P
SCORTIA, THOMAS N./ROBINSON, FRANK M. PROMETHEUS CRISIS, THE  H:2,P
SCORTIA, THOMAS N. STRANGE BEDFELLOWS C/PA,PA

SCORTIA, THOMAS N./ YARBRO, CHELSEA QUINN TWO VIEWS OF WONDER

C/PA,PA

SCOTCH, CHERI  WEREWOLF’S KISS, THE [MARLEY FAMILY WEREWOLF 1]  P
SCOTCH, CHERI  WEREWOLF’S TOUCH, THE [MARLEY FAMILY WEREWOLF 2]  P
SCOTT, ALAN  ANTHRAX MUTATION, THE  C/P:2,P
SCOTT, HOLLY/DUNCAN, JAMES  SIREN SONG [@STARGATE SG-1 #6]  P
SCOTT, JODY  I, VAMPIRE  C/P:2,P
SCOTT, JODY  PASSING FOR HUMAN  C/P,P
SCOTT, MARTIN  DEATH AND THRAXAS [THRAXAS]  P
SCOTT, MARTIN  THRAXAS [THRAXAS]  SR-P,C/P
SCOTT, MELISSA/BARNETT, LISA A.  ARMOR OF LIGHT, THE  C/P,SFWA-P,P,H
SCOTT, MELISSA  BURNING BRIGHT  H:2,P
SCOTT, MELISSA  CHOICE OF DESTINIES, A  C/P,P
SCOTT, MELISSA  DREAMING METAL  LP
SCOTT, MELISSA  DREAMSHIPS  H:2,SFWA,P
SCOTT, MELISSA  EMPRESS OF EARTH, THE [SILENCE LEIGH 3]  C/P:2,SFWA-P,P
SCOTT, MELISSA  FIVE-TWELFTHS OF HEAVEN [SILENCE LEIGH 1]  C/P:2,P
SCOTT, MELISSA  GAME BEYOND, THE  C/P,P
SCOTT, MELISSA  GARDEN, THE [STAR TREK VOYAGER]  P
SCOTT, MELISSA  JAZZ, THE  H
SCOTT, MELISSA  KINDLY ONES, THE  H:2,SFWA-P,P

The MITSFS Pinkdex  September 24, 2007  Page 344
AUTHOR  TITLE  CODE
SCOTT, MELISSA  MIGHTY GOOD ROAD  C/P:2,SFWA-P,P
SCOTT, MELISSA  NIGHT SKY MINE  H,LP
SCOTT, MELISSA/BARNETT, LISA A.  POINT OF DREAMS [NICOLAS RATHE 2]  H
SCOTT, MELISSA/BARNETT, LISA A.  POINT OF HOPES [NICOLAS RATHE 1]  H
SCOTT, MELISSA  PROUD HELIOS [STAR TREK DS9]  C/P:3,P
SCOTT, MELISSA  ROADS OF HEAVEN, THE [SILENCE LEIGH 1,2,3]  H:2
SCOTT, MELISSA  SHADOW MAN  H,SFWA,LP
SCOTT, MELISSA  SHAPES OF THEIR HEARTS, THE  H
SCOTT, MELISSA  SILENCE IN SOLITUDE [SILENCE LEIGH 2]  C/P:2,P
SCOTT, MELISSA  TROUBLE AND HER FRIENDS  H,P
SCOTT, MICHAEL  CULAI HERITAGE, THE [TALES OF THE BARD 1,2,3]  LP
SCOTT, MICHAEL  DEATH’S LAW [TALES OF THE BARD #3]  LP
SCOTT, R. C.  BLOOD SPORT  P
SCOTT, WALTER  IVANHOE  H
SCOTTEN, CORDELL  RENEGADE [ROBOTS AND ALIENS 2]  C/P:3,SFWA-P,P
SCULLY, FRANK  BEHIND THE FLYING SAUCERS  [C/RAN-H]
SEABROCK, WILLIAM  WITCHCRAFT  [C/RAN-P]
SEARLES, BAIRD/THOMSEN, BRIAN M.  HALFLINGS, HOBBITS, WARROWS & WEEFOLK  HA,LPA

SEARLES, BAIRD/MEACHAM, BETH/ FRANKLIN, MICHAEL READER’S GUIDE TO FANTASY, A

R/REF-P,C/REF-P

SEARLES, BAIRD/LAST, MARTIN/ MEACHAM, BETH/FRANKLIN, MICHAEL READER’S GUIDE TO SCIENCE FICTION, A

R/REF-P,C/REF-P:2

SEARLS, HANK  ASTRONAUT, THE  C/P:2,P
SEARLS, HANK  BIG X, THE  H,P
SEARLS, HANK  PILGRIM PROJECT, THE  C/P:2,P
SEARLS, HANK  SOUNDING  LP
SEDGWICK, MARCUS  BOOK OF DEAD DAYS, THE  H
SEGRIFF, LARRY  ALIEN DREAMS  P
SEGRIFF, LARRY  SPACER DREAMS  P
SEHESTEDT, MARK  FROSTFELL [@FORGOTTEN REALMS WIZARDS]  P
SEHESTEDT, MARK  TALES OF THE LAST WAR, THE [@EBERRON]  PA
SEIDMAN, DAVID L.  FIRST CASUALTY, THE [@HERCULES THE LEGENDARY JOURNEYS]  P
SEIDMAN, ERIC  WANTED BY THE INTERGALACTIC SECURITY BUREAU  A
SELBY, CURT  BLOOD COUNTY  P
SELBY, CURT  I, ZOMBIE  C/P,P
SELLERS, CON  MR. TOMORROW  P
SELLINGS, ARTHUR  POWER OF X, THE  C/P:3,P
SELLINGS, ARTHUR  QUY EFFECT, THE  C/P:2,P
SELLINGS, ARTHUR  TELEPATH  C/P:2,P
SELLINGS, ARTHUR  TIME TRANSFER  SR-P
SELLINGS, ARTHUR  UNCENSORED MAN, THE  C/P:2,P
SELTZER, DAVID  OMEN, THE [OMEN 1]  P
SELTZER, DAVID  PROPHECY  P
SEMPLE, LORENZO JR.  DINO DE LAURENTIIS PRODUCTION OF KING KONG, THE  P
SENN, STEVE  CIRCLE IN THE SEA, A  H
SENN, STEVE  DOUBLE DISAPPEARANCE OF WALTER FOZBEK, THE  LP
SENN, STEVE  SPACEBREAD  H
SERLING, CAROL  ADVENTURES IN THE TWILIGHT ZONE  PA
SERLING, CAROL  JOURNEYS TO THE TWILIGHT ZONE  HA,SFWA-P
SERLING, CAROL  RETURN TO THE TWILIGHT ZONE  PA
SERLING, ROBERT  AIRFORCE ONE IS HAUNTED  H,P
SERLING, ROBERT  PRESIDENT’S PLANE IS MISSING, THE  P
SERLING, ROD  FROM THE TWILIGHT ZONE  H
SERLING, ROD  MORE STORIES FROM THE TWILIGHT ZONE  C/P:2,P
SERLING, ROD  NEW STORIES FROM THE TWILIGHT ZONE  C/P:2,P
SERLING, ROD  NIGHT GALLERY  C/P:2,P
SERLING, ROD  NIGHT GALLERY 2  P
SERLING, ROD  ROD SERLING’S DEVILS AND DEMONS  C/PA,PA
SERLING, ROD  ROD SERLING’S OTHER WORLDS  PA
SERLING, ROD  ROD SERLING’S TRIPLE W  PA,C/PA
SERLING, ROD  STORIES FROM THE TWILIGHT ZONE  C/P:2,P
SERRAILLER, IAN  GORGON’S HEAD, THE  P
SERVICE, PAMELA  BEING OF TWO MINDS  SFWA
SERVICE, PAMELA F.  QUESTION OF DESTINY, A  H
SERVICE, PAMELA F.  RELUCTANT GOD, THE  C/P,P
SERVICE, PAMELA F.  STINKER FROM SPACE [STINKER 1]  P
SERVICE, PAMELA F.  STINKER’S RETURN [STINKER 2]  SFWA
SERVICE, PAMELA F.  TOMORROW’S MAGIC [MAGIC’S RETURN 2]  H,P
SERVICE, PAMELA F.  UNDER ALIEN STARS  C/P,SFWA,P
SERVICE, PAMELA F.  VISION QUEST  C/P,P
SERVICE, PAMELA F.  WINTER OF MAGIC’S RETURN [MAGIC’S RETURN 1]  H,P
SERVICE, ROBERT W.  MASTER OF THE MICROBE, THE  H
SERVISS, GARRETT P.  COLUMBUS OF SPACE, A  LP
SERVISS, GARRETT P.  EDISON’S CONQUEST OF MARS  SR-H

The MITSFS Pinkdex  September 24, 2007  Page 345
AUTHOR  TITLE  CODE
SERVISS, GARRETT P.  INVASION OF MARS  SR-P
SERVISS, GARRETT P.  MOON METAL, THE  H
SERVISS, GARRETT P.  SECOND DELUGE, THE  LP
SETH, MARIE  DREAM OF THE DEAD  SR-P
SETH, RONALD  SPY AND THE ATOM GUN, THE  H
SETLOWE, RICHARD  EXPERIMENT, THE  H:2
SETO, ANDY  CROUCHING TIGER HIDDEN DRAGON, VOL. 5  SCX
SETON, ANYA  TURQUOISE, THE  P
SEVARANCE, CAROL  SORCEROUS SEA [ISLAND WARRIOR 3]  P
SEVERANCE, CAROL  DEMON DRUMS  C/P,P
SEVERANCE, CAROL  REEFSONG  C/P,P
SEVERANCE, CAROL  STORM CALLER [ISLAND WARRIOR 2]  P
SEVERN, DAVID  FUTURE TOOK US, THE  SR-P,C/P:2
SEVERSON, ELLEN DODGE  HEDERICK THE THEOCRAT [@DRAGONLANCE VILLAINS #4]  C/P,P
SEYMOUR, ALAN  COMING SELF-DESTRUCTION OF THE UNITED STATES OF AMERICA, THE  SR-P
SFAR, JOANN  RABBI’S CAT, THE  CX
SFBC  TALES OF THE SLAYER: VOLUMES 2 & 3 [BUFFY]  HA
SHAARA, MICHAEL  HERALD, THE  P
SHAARA, MICHAEL  SOLDIER BOY  C/P,P
SHACKLEFORD, JACK D.  EVE OF MIDSUMMER, THE  SR-P
SHACKLEFORD, JACK D.  STRICKLAND DEMON, THE  SR-P
SHACKLEFORD, JACK D.  TANITH  SR-P
SHAFER, ROBERT  CONQUERED PLACE, THE  H
SHAFFER, EUGENE CARL  CLONES, THE  P
SHAFFER, EUGENE CARL  LAST BREATH, THE  SR-P
SHAFFER, EUGENE CARL  PANIC 7  P
SHANKS, EDWARD  PEOPLE OF THE RUINS, THE  H
SHANNON, FRED  WEIGHTLESS IN GAZA  P
SHANNON, LORELEI  RAGS AND OLD IRON  H
SHAPIRO, ERIC  IT’S ONLY TEMPORARY  LP
SHAPIRO, NEIL  MIND CALL  P
SHAPIRO, NEIL  PLANET WITHOUT A NAME  P
SHAPLEY, HARLOW  VIEW FROM A DISTANT STAR, THE  [C/RAN-P:2]
SHAREE, KEITH  GULLIVER’S FUGITIVES [STAR TREK TNG]  C/P:2,P
SHARKEY, JACK  ADDAMS FAMILY, THE [ADDAMS FAMILY]  C/P:2,P
SHARKEY, JACK  SECRET MARTIANS, THE  DD471
SHARKEY, JACK  ULTIMATUM IN 2050 A.D.  DM117
SHARMAN, NICK  JUDGMENT DAY  P
SHARMAN, NICK  SURROGATE, THE  P
SHARMAT, MITCHELL  GIRL OF MANY PARTS, A  P
SHARP, MARGERY  STONE OF CHASTITY, THE  H
SHARPE, JOHN RUFUS III  HOGAR, LORD OF THE ASYR  P

SHATNER, LISABETH WILLIAM SHATNER’S PERSONAL ACCOUNT OF THE MAKING OF STAR

TREK 5 [STAR TREK]

R/REF-P,C/REF-H

SHATNER, WILLIAM  ASHES OF EDEN, THE [STAR TREK ODYSSEY 1]  H,P
SHATNER, WILLIAM  AVENGER [STAR TREK ODYSSEY 3]  P
SHATNER, WILLIAM/TOBIAS, MICHAEL  BELIEVE.  C/P,SFWA-P:2,P
SHATNER, WILLIAM  BEYOND THE STARS [QUEST FOR TOMORROW 4]  H

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD CAPTAIN’S BLOOD [STAR TREK]

H

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD CAPTAIN’S GLORY [STAR TREK]

H

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD CAPTAIN’S PERIL [@STAR TREK]

P

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD DARK VICTORY [STAR TREK MIRROR 2]

P

SHATNER, WILLIAM  DELTA SEARCH [@QUEST FOR TOMORROW 1]  P
SHATNER, WILLIAM  IN ALIEN HANDS [@QUEST FOR TOMORROW 2]  P
SHATNER, WILLIAM  LAW OF WAR, THE [BENTON HAWKES 2]  H,P
SHATNER, WILLIAM  MAN O’ WAR [BENTON HAWKES 1]  H,P

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD PRESERVER [STAR TREK MIRROR 3]

P

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD RETURN, THE [STAR TREK ODYSSEY 2]

P

SHATNER, WILLIAM  SHADOW PLANET [QUEST FOR TOMORROW 5]  H

SHATNER, WILLIAM/ REEVES-STEVENS, JUDITH/ REEVES-STEVENS, GARFIELD SPECTRE [STAR TREK MIRROR 1]

C/P,P

SHATNER, WILLIAM/KRESKI, CHRIS  STAR TREK MEMORIES  R/REF-H

AUTHOR  TITLE  CODE
SHATNER, WILLIAM  STAR TREK MOVIE MEMORIES  R/REF-H,C/REF-P
SHATNER, WILLIAM  STEP INTO CHAOS [@QUEST FOR TOMORROW 3]  P
SHATNER, WILLIAM  TEK KILL [TEK 8]  P
SHATNER, WILLIAM  TEK MONEY [TEK 7]  H,P
SHATNER, WILLIAM  TEK NET [TEK 9]  H,P
SHATNER, WILLIAM  TEK POWER [TEK 6]  P
SHATNER, WILLIAM  TEK SECRET [TEK 5]  P
SHATNER, WILLIAM  TEK VENGEANCE [TEK 4]  H,SFWA-P:2,P
SHATNER, WILLIAM  TEKLAB [TEK 3]  C/P,SFWA-P,P
SHATNER, WILLIAM  TEKLORDS [TEK 2]  C/P:2,SFWA-P:2,P
SHATNER, WILLIAM  TEKWAR [TEK 1]  C/P,SFWA-P:2,P,H
SHATTUCK, RICHARD  HALF-HAUNTED SALOON, THE  H
SHAUGHNESSY, P. J.  TRAUMA  P
SHAW, BOB  BEST OF THE BUSHEL, THE  R/REF-P
SHAW, BOB  BETTER MANTRAP, A  P
SHAW, BOB  CERES SOLUTION, THE  C/P,P
SHAW, BOB  COSMIC KALEIDOSCOPE  C/P,SFWA,P
SHAW, BOB  DAGGER OF THE MIND  C/P,P
SHAW, BOB  DIMENSIONS  P
SHAW, BOB  EASTERCON SPEECHES, THE  R/REF-P
SHAW, BOB  FIRE PATTERN  C/P,P
SHAW, BOB  FUGITIVE WORLDS, THE [TOLLER MARAQUINE 3]  P
SHAW, BOB  GROUND ZERO MAN  C/P:3,P
SHAW, BOB  MEDUSA’S CHILDREN  SFWA-P,H:2,P,SFWA
SHAW, BOB  MESSAGES FOUND IN AN OXYGEN BOTTLE  SR-DNESFA1
SHAW, BOB  NIGHT WALK  SR-P,C/P:2
SHAW, BOB  ONE MILLION TOMORROWS  C/P:2,P
SHAW, BOB  ORBITSVILLE  C/P:2,SFWA-P,P
SHAW, BOB  ORBITSVILLE DEPARTURE  C/P:2,SFWA-P,P
SHAW, BOB  OTHER DAYS, OTHER EYES  C/P:2,P
SHAW, BOB  PALACE OF ETERNITY, THE  C/P:2,P
SHAW, BOB  RAGGED ASTRONAUTS, THE [TOLLER MARAQUINE 1]  H:2,SFWA,P
SHAW, BOB  SHADOW OF HEAVEN  C/P:2,P
SHAW, BOB  SHIP OF STRANGERS  C/P,P
SHAW, BOB  TOMORROW LIES IN AMBUSH  C/P:2,P
SHAW, BOB  TWO-TIMERS, THE  C/P:2,P
SHAW, BOB  VERTIGO  C/P,P
SHAW, BOB  WHO GOES HERE?  C/P,P
SHAW, BOB  WOODEN SPACESHIPS, THE [TOLLER MARAQUINE 2]  C/P:2,SFWA-P,P
SHAW, BOB  WREATH OF STARS, A  C/P,SFWA,SFWA-P,P,H
SHAW, BRIAN  ARGENTIS  SR-P
SHAW, BRIAN  SHIPS OF VERO  SR-P,H
SHAW, DAVID  LABORATORY X  SR-P
SHAW, DAVID  PLANET FEDERATION  SR-P
SHAW, DAVID  SPACE MEN  SR-P
SHAW, FREDERICK L. JR.  ENVOY TO THE DOG STAR  DG614
SHAW, LARRY T.  GREAT SCIENCE FICTION ADVENTURES  PA,C/PA
SHAW, LARRY T.  TERROR!  C/PA,PA
SHAWN, WILLIAM S.  GOLDEN CIRCLE, THE [PHANTOM 5]  P
SHAWN, WILLIAM S.  HYDRA MONSTER, THE [PHANTOM 8]  P
SHAWN, WILLIAM S.  MYSTERY OF THE SEA HORSE, THE [PHANTOM 7]  P
SHAWN, WILLIAM S.  STORY OF THE VEILED LADY, THE [PHANTOM 4]  P
SHCREIBER, ELLEN  DANCE WITH A VAMPIRE (PROOF) [@VAMPIRE KISSES #4]  SR-LP
SHEA, GEORGE  I DIED HERE  SR-P
SHEA, MICHAEL  COLOR OUT OF TIME, THE  C/P,P
SHEA, MICHAEL  I, SAID THE FLY  LP
SHEA, MICHAEL  INCOMPLEAT NIFFT, THE  P
SHEA, MICHAEL  MINES OF BEHEMOTH, THE [NIFFT 2]  P
SHEA, MICHAEL  NIFFT THE LEAN [NIFFT 1]  C/P,P
SHEA, MICHAEL  QUEST FOR SIMBILIS, A [DYING EARTH]  C/P,P
SHEA, ROBERT/WILSON, ROBERT ANTON  EYE IN THE PYRAMID, THE [ILLUMINATUS! 1]  C/P,P
SHEA, ROBERT/WILSON, ROBERT ANTON  GOLDEN APPLE, THE [ILLUMINATUS! 2]  C/P:2,P
SHEA, ROBERT/WILSON, ROBERT ANTON  ILLUMINATUS! TRILOGY, THE [ILLUMINATUS! 1,2,3]  H:3,LP
SHEA, ROBERT/WILSON, ROBERT ANTON  LEVIATHAN [ILLUMINATUS! 3]  C/P:2,P
SHEAR, DAVID  CLONING  C/P,SFWA,P,H
SHEARIN, LISA  MAGIC LOST, TROUBLE FOUND  P
SHECKLEY, ROBERT  AFTER THE FALL  PA
SHECKLEY, ROBERT  ALCHEMICAL MARRIAGE OF ALISTAIR CROMPTON, THE  P
SHECKLEY, ROBERT  ALIEN HARVEST [ALIENS 5]  C/P,P
SHECKLEY, ROBERT  CALL TO ARMS, A [@BABYLON 5]  C/P,P
SHECKLEY, ROBERT  CAN YOU FEEL ANYTHING WHEN I DO THIS?  H:2,SFWA,P
SHECKLEY, ROBERT  CITIZEN IN SPACE  C/P,P,H
SHECKLEY, ROBERT  CROMPTON DIVIDED  C/P,P,H
SHECKLEY, ROBERT  DIMENSION OF MIRACLES  C/P:3,P

AUTHOR  TITLE  CODE
SHECKLEY, ROBERT  DIMENSIONS OF SHECKLEY  H
SHECKLEY, ROBERT  DRAMOCLES  H:2
SHECKLEY, ROBERT  GAME OF X, THE  P
SHECKLEY, ROBERT  GODSHOME  H
SHECKLEY, ROBERT  HUNTER/VICTIM  P
SHECKLEY, ROBERT  IMMORTALITY, INC.  C/P:2,P
SHECKLEY, ROBERT  JOURNEY BEYOND TOMORROW  C/P:2,P
SHECKLEY, ROBERT  LAERTIAN GAMBLE, THE [STAR TREK DS9]  P
SHECKLEY, ROBERT  MASQUE OF MANANA, THE  H
SHECKLEY, ROBERT  MINDSWAP  H:2,P
SHECKLEY, ROBERT  MINOTAUR MAZE  SR-H
SHECKLEY, ROBERT  NOTIONS: UNLIMITED  C/P:2,P
SHECKLEY, ROBERT  OPTIONS  C/P:2,P
SHECKLEY, ROBERT  PEOPLE TRAP, THE  C/P:2,P
SHECKLEY, ROBERT  PEOPLE TRAP, THE/MINDSWAP  C/P,P
SHECKLEY, ROBERT  PILGRIMAGE TO EARTH  C/P:2,P
SHECKLEY, ROBERT  ROBOT WHO LOOKED LIKE ME, THE  C/P:2,P
SHECKLEY, ROBERT  ROBOT WHO LOOKED LIKE ME, THE (BRITISH)  SR-P
SHECKLEY, ROBERT  SHARDS OF SPACE  C/P:2,P
SHECKLEY, ROBERT  STATUS CIVILIZATION, THE  C/P:2,P
SHECKLEY, ROBERT  STATUS CIVILIZATION, THE/NOTIONS: UNLIMITED  C/P:2,P
SHECKLEY, ROBERT  STORE OF INFINITY  C/P:2,P
SHECKLEY, ROBERT  10TH VICTIM, THE  C/P,P
SHECKLEY, ROBERT  UNTOUCHED BY HUMAN HANDS  C/P:2,P
SHECKLEY, ROBERT  VICTIM PRIME  C/P,P
SHECKLEY, ROBERT  WONDERFUL WORLD OF ROBERT SHECKLEY, THE  P
SHEDLEY, ETHAN I.  EARTH SHIP AND STAR SONG  C/P,P
SHEEHAN, CAROLYN/SHEEHAN, EDMUND  MAGNIFI-CAT  SFWA
SHEEHAN, PERLEY POORE  ABYSS OF WONDERS, THE  SR-H
SHEEHAN, PERLEY POORE  ONE GIFT, THE  SR-LP
SHEFFIELD, CHARLES  AFTERMATH  C/P:2,P
SHEFFIELD, CHARLES  AMAZING DR. DARWIN, THE  H
SHEFFIELD, CHARLES  BETWEEN THE STROKES OF NIGHT  C/P:2,P
SHEFFIELD, CHARLES  BETWEEN THE STROKES OF NIGHT (REVISED)  P
SHEFFIELD, CHARLES  BILLION DOLLAR BOY, THE  C/P:2,P
SHEFFIELD, CHARLES  BORDERLANDS OF SCIENCE  R/REF-H
SHEFFIELD, CHARLES  BROTHER TO DRAGONS  C/P,P
SHEFFIELD, CHARLES  COLD AS ICE [BATTACHARIYA 1]  C/P:2,P,H
SHEFFIELD, CHARLES  CONVERGENCE [HERITAGE UNIVERSE 4]  H,P
SHEFFIELD, CHARLES  CONVERGENT SERIES [HERITAGE UNIVERSE (REVISED) 1,2]  P
SHEFFIELD, CHARLES  CYBORG FROM EARTH, THE  C/P,P,H
SHEFFIELD, CHARLES  DANCING WITH MYSELF  C/P:2,P
SHEFFIELD, CHARLES  DARK AS DAY [BATTACHARIYA 3]  H
SHEFFIELD, CHARLES  DIVERGENCE [HERITAGE UNIVERSE 2]  H,P
SHEFFIELD, CHARLES  ERASMUS MAGISTER  C/P,P
SHEFFIELD, CHARLES  GANYMEDE CLUB, THE [BATTACHARIYA 2]  H,P
SHEFFIELD, CHARLES  GEORGIA ON MY MIND AND OTHER PLACES  LP
SHEFFIELD, CHARLES  GODSPEED  C/P,P,H
SHEFFIELD, CHARLES  HERITAGE UNIVERSE, THE [HERITAGE UNIVERSE 1,2,3]  H:2
SHEFFIELD, CHARLES  HIDDEN VARIABLES  C/P,P
SHEFFIELD, CHARLES/POURNELLE, JERRY  HIGHER EDUCATION  C/P:2,P,H
SHEFFIELD, CHARLES  HOW TO SAVE THE WORLD  SFWA,LP
SHEFFIELD, CHARLES  MCANDREW CHRONICLES, THE  C/P:2,P
SHEFFIELD, CHARLES  MIND POOL, THE  C/P:2,SFWA-P,P
SHEFFIELD, CHARLES  MY BROTHER’S KEEPER  C/P:2,P
SHEFFIELD, CHARLES  NIMROD HUNT, THE  C/P:2,P
SHEFFIELD, CHARLES  ONE MAN’S UNIVERSE  LP
SHEFFIELD, CHARLES  PROTEUS COMBINED [PROTEUS]  P
SHEFFIELD, CHARLES  PROTEUS IN THE UNDERWORLD [PROTEUS 3]  C/P,P,H
SHEFFIELD, CHARLES  PROTEUS MANIFEST [PROTEUS 1,2]  H
SHEFFIELD, CHARLES  PROTEUS UNBOUND [PROTEUS 2]  C/P,P
SHEFFIELD, CHARLES  PUTTING UP ROOTS  C/P,P,H
SHEFFIELD, CHARLES  RESURGENCE [HERITAGE UNIVERSE 5]  H:2
SHEFFIELD, CHARLES/BISCHOFF, DAVID F.  SELKIE, THE  H,P
SHEFFIELD, CHARLES  SIGHT OF PROTEUS [PROTEUS 1]  C/P:2,P
SHEFFIELD, CHARLES  SPACE SUITS  LP
SHEFFIELD, CHARLES  SPHERES OF HEAVEN, THE  H:2,P
SHEFFIELD, CHARLES  STARFIRE  C/P,LP,H
SHEFFIELD, CHARLES  SUMMERTIDE [HERITAGE UNIVERSE 1]  C/P,P,H
SHEFFIELD, CHARLES  TOMORROW AND TOMORROW  C/P,P,H
SHEFFIELD, CHARLES  TRADER’S WORLD  C/P:2,P
SHEFFIELD, CHARLES  TRANSCENDENCE [HERITAGE UNIVERSE 3]  H:2,P
SHEFFIELD, CHARLES  TRANSVERGENCE [HERITAGE UNIVERSE (REVISED) 3,4]  C/P,P
SHEFFIELD, CHARLES  VECTORS  C/P,P

AUTHOR  TITLE  CODE
SHEFFIELD, CHARLES  WEB BETWEEN THE WORLDS, THE  C/P:2,LP
SHEFNER, VADIM  UNMAN, THE/KOVRIGIN’S CHRONICLES  H,LP
SHEFNER, VADIM  UNMAN, THE/KOVRIGIN’S CHRONICLES (PROOF)  SR-LP
SHELDON, ANNE  ADVENTURES OF THE FAITHFUL COUNSELOR, THE  LP
SHELDON, LEE  DOOMED PLANET  H
SHELDON, ROY  ATOMS IN ACTION  SR-P,H
SHELDON, ROY  HOUSE OF ENTROPY  SR-P,H
SHELDON, ROY  MENACING SLEEP, THE  SR-P
SHELDON, ROY  METAL EATER, THE  SR-P
SHELDON, ROY  MOMENT OUT OF TIME  SR-P
SHELDON, ROY  TWO DAYS OF TERROR  SR-P
SHELLEY, MARY  FRANKENSTEIN  C/P,P,H:2
SHELLEY, MARY  MORTAL IMMORTAL, THE  SR-LP
SHELLEY, RICK  BUCHANAN CAMPAIGN, THE [SECOND COMMONWEALTH 1]  P
SHELLEY, RICK  CAPTAIN [DIRIGENT MERCENARY CORPS 3]  C/P,P
SHELLEY, RICK  COLONEL [DIRIGENT MERCENARY CORPS 6]  C/P,P
SHELLEY, RICK  DEEP STRIKE [SPEC OPS SQUAD 2]  C/P,P
SHELLEY, RICK  FIRES OF COVENTRY, THE [SECOND COMMONWEALTH 2]  P
SHELLEY, RICK  HERO KING, THE [@VARAYAN #3]  P
SHELLEY, RICK  HERO OF VARAY, THE [@VARAYAN #2]  P
SHELLEY, RICK  HOLDING THE LINE [SPEC OPS SQUAD 1]  P
SHELLEY, RICK  JUMP PAY [LUCKY 13TH]  P
SHELLEY, RICK  LIEUTENANT [DIRIGENT MERCENARY CORPS 2]  C/P,P
SHELLEY, RICK  LIEUTENANT COLONEL [DIRIGENT MERCENARY CORPS 5]  C/P,P
SHELLEY, RICK  MAJOR [DIRIGENT MERCENARY CORPS 4]  C/P,P
SHELLEY, RICK  OFFICER-CADET [DIRIGENT MERCENARY CORPS 1]  C/P,P
SHELLEY, RICK  RETURN TO CAMEREIN [SECOND COMMONWEALTH 3]  P
SHELLEY, RICK  SIDE SHOW [LUCKY 13TH]  P
SHELLEY, RICK  SON OF THE HERO [@VARAYAN #1]  P
SHELLEY, RICK  SUCKER PUNCH [@SPEC OPS SQUAD 3]  P
SHELLEY, RICK  UNTIL RELIEVED [LUCKY 13TH]  P
SHELLEY, RICK  WIZARD AT HOME, THE [@SEVEN TOWERS]  P
SHELLEY, RICK  WIZARD AT MECQ, THE [@SEVEN TOWERS]  P
SHELTON, WILLIAM R.  STOWAWAY TO THE MOON  H
SHEPARD, LESLIE  DRACULA BOOK OF GREAT VAMPIRE STORIES, THE  C/PA,LPA
SHEPARD, LUCIUS  BARNACLE BILL THE SPACER AND OTHER STORIES  LP
SHEPARD, LUCIUS  ENDS OF THE EARTH, THE  P
SHEPARD, LUCIUS  GOLDEN, THE  P
SHEPARD, LUCIUS  GOLDEN, THE (PROOF)  SR-LP
SHEPARD, LUCIUS  GREEN EYES  C/P:2,P
SHEPARD, LUCIUS  JAGUAR HUNTER, THE  H,LP
SHEPARD, LUCIUS  KALIMANTAN  LP
SHEPARD, LUCIUS  LIAR’S HOUSE  SR-H
SHEPARD, LUCIUS  LIFE DURING WARTIME  C/P,P
SHEPARD, LUCIUS  SCALEHUNTER’S BEAUTIFUL DAUGHTER, THE  H
SHEPARD, LUCIUS  VIATOR  H
SHEPHERD, JOEL  BREAKAWAY [CASSANDRA KRESNOV 2]  LP
SHEPHERD, JOEL  CROSSOVER [CASSANDRA KRESNOV 1]  LP
SHEPHERD, MARK  ELVENDUDE  P
SHEPHERD, MARK  ESCAPE FROM ROKSAMUR [BARD’S TALE 4]  P
SHEPHERD, MARK  LAZERWARZ  P
SHEPHERD, MARK  SPIRITRIDE  P
SHEPHERD, MIKE  KRIS LONGKNIFE: DEFIANT [KRIS LONGKNIFE 3]  P
SHEPHERD, MIKE  KRIS LONGKNIFE: DESERTER [KRIS LONGKNIFE 2]  P
SHEPHERD, MIKE  KRIS LONGKNIFE: MUTINEER [KRIS LONGKNIFE 1]  P
SHEPHERD, MIKE  KRIS LONGKNIFE: RESOLUTE [KRIS LONGKNIFE 4]  P
SHERBANIUK, RICHARD  FIFTH HORSEMAN, THE  H
SHERMAN, DAVID/CRAGG, DAN  BACKSHOT [@STARFIST FORCE RECON 1]  P
SHERMAN, DAVID/CRAGG, DAN  BLOOD CONTACT [@STARFIST 4]  P
SHERMAN, DAVID/CRAGG, DAN  FIRESTORM [@STARFIST 12]  H
SHERMAN, DAVID/CRAGG, DAN  FIRST TO FIGHT [@STARFIST 1]  C/P:2,P
SHERMAN, DAVID/CRAGG, DAN  FLASHFIRE [@STARFIST 11]  H
SHERMAN, DAVID  GULF RUN [@DEMONTECH 3]  P
SHERMAN, DAVID/CRAGG, DAN  HANGFIRE [@STARFIST 6]  C/P,P
SHERMAN, DAVID/CRAGG, DAN  JEDI TRIAL [@STAR WARS CLONE WARS]  H
SHERMAN, DAVID/CRAGG, DAN  KINGDOM’S FURY [@STARFIST 8]  P
SHERMAN, DAVID/CRAGG, DAN  KINGDOM’S SWORDS [@STARFIST 7]  C/P,P
SHERMAN, DAVID/CRAGG, DAN  LAZARUS RISING [@STARFIST 9]  H,P
SHERMAN, DAVID  ONSLAUGHT [@DEMONTECH 1]  P
SHERMAN, DAVID/CRAGG, DAN  POINTBLANK [@STARFIST FORCE RECON 2]  P
SHERMAN, DAVID  RALLY POINT [@DEMONTECH 2]  P
SHERMAN, DAVID/CRAGG, DAN  SCHOOL OF FIRE [@STARFIST 2]  P
SHERMAN, DAVID/CRAGG, DAN  STEEL GAUNTLET [@STARFIST 3]  P
SHERMAN, DAVID/CRAGG, DAN  TECHNOKILL [@STARFIST 5]  P

The MITSFS Pinkdex  September 24, 2007  Page 349
AUTHOR  TITLE  CODE
SHERMAN, DAVID/CRAGG, DAN  WORLD OF HURT, A [@STARFIST 10]  P
SHERMAN, DAVID/CRAGG, DAN  WORLD OF HURT, A (PROOF) [@STARFIST 10]  SR-LP
SHERMAN, DELIA  CHANGELING  H
SHERMAN, DELIA/GOSS, THEODORA  INTERFICTIONS  LPA
SHERMAN, DELIA  THROUGH A BRAZEN MIRROR  P
SHERMAN, HAROLD M.  GREEN MAN, THE  SR-P
SHERMAN, HAROLD M.  TAHARA BOY MYSTIC OF INDIA  H
SHERMAN, HAROLD M.  TAHARA IN THE LAND OF YUCATAN  H
SHERMAN, JOEL HENRY  CORPSEMAN  P
SHERMAN, JOEL HENRY  RANDOM FACTOR  P
SHERMAN, JORY  BAMBOO DEMONS, THE [@CHILL #3]  P
SHERMAN, JORY  CHILL [@CHILL #2]  P
SHERMAN, JORY  HOUSE OF SCORPIONS [@CHILL #6]  P
SHERMAN, JORY  PHOENIX MAN, THE [@CHILL #5]  P
SHERMAN, JORY  SATAN’S SEED [@CHILL #1]  P
SHERMAN, JORY  SHADOWS [@CHILL #7]  C/P,P
SHERMAN, JORY  VEGAS VAMPIRE [@CHILL #4]  P
SHERMAN, JOSEPHA  CAPTIVE SOUL, THE [@HIGHLANDER]  P
SHERMAN, JOSEPHA  CHAOS GATE, THE [BARD’S TALE 5]  P
SHERMAN, JOSEPHA/SHWARTZ, SUSAN  EPIPHANY [@STAR TREK VULCAN’S SOUL #3]  H
SHERMAN, JOSEPHA/SHWARTZ, SUSAN  EXILES [STAR TREK VULCAN’S SOUL 2]  H
SHERMAN, JOSEPHA/SHWARTZ, SUSAN  EXODUS [@STAR TREK VULCAN’S SOUL #1]  P
SHERMAN, JOSEPHA  FORGING THE RUNES [PRINCE OF THE SIDHE 2]  P
SHERMAN, JOSEPHA  HORSE OF FLAME, THE  P
SHERMAN, JOSEPHA  KING’S SON, MAGIC’S SON  P
SHERMAN, JOSEPHA  LAMMAS NIGHT  PA,C/PA
SHERMAN, JOSEPHA  ONCE UPON A GALAXY  R/REF-P
SHERMAN, JOSEPHA  SHATTERED OATH, THE [PRINCE OF THE SIDHE 1]  P
SHERMAN, JOSEPHA  SHINING FALCON, THE  C/P:2,P
SHERMAN, JOSEPHA  SON OF DARKNESS  C/P,P,H
SHERMAN, JOSEPHA  STRANGE AND ANCIENT NAME, A  C/P,P,H

SHERMAN, JOSEPHA THROUGH THE LOOKING GLASS

[@ANDROMEDA (GENE RODDENBERRY’S) #4]

P

SHERMAN, JOSEPHA/ DECANDIDO, KEITH R. A.

URBAN NIGHTMARES

PA

SHERMAN, JOSEPHA/SHWARTZ, SUSAN  VULCAN’S FORGE [@STAR TREK]  P
SHERMAN, JOSEPHA/SHWARTZ, SUSAN  VULCAN’S HEART [STAR TREK]  P
SHERRED, T. L.  ALIEN ISLAND  C/P:2,P
SHERRED, T. L./BIGGLE, LLOYD JR.  ALIEN MAIN  H:2,SFWA
SHERRED, T. L.  FIRST PERSON, PECULIAR  C/P:2,P
SHERRELL, CARL  ARCANE  C/P:2,P
SHERRELL, CARL  RAUM  C/P,P
SHERRELL, CARL  SKRAELINGS  P
SHERRELL, CARL  SPACE PRODIGAL, THE  C/P:2,P
SHERRIFF, R. C.  CATACLYSM, THE  SR-P
SHERRIFF, R. C.  HOPKINS MANUSCRIPT, THE  H:2
SHERWOOD, MARTIN  MAXWELL’S DEMON  SR-P
SHERWOOD, MARTIN  SURVIVAL  P
SHETTERLY, WILL  CATS HAVE NO LORD  C/P:2,P
SHETTERLY, WILL  CHIMERA  H
SHETTERLY, WILL  DOGLAND  H
SHETTERLY, WILL  ELSEWHERE [BORDERLANDS 5]  H:2,P
SHETTERLY, WILL/BULL, EMMA  LIAVEK [LIAVEK 1]  C/PA,PA
SHETTERLY, WILL/BULL, EMMA  LIAVEK: FESTIVAL WEEK [LIAVEK 5]  PA
SHETTERLY, WILL/BULL, EMMA  LIAVEK: PLAYERS OF LUCK, THE [LIAVEK 2]  C/PA,PA
SHETTERLY, WILL/BULL, EMMA  LIAVEK: SPELLS OF BINDING [LIAVEK 4]  C/PA,PA
SHETTERLY, WILL/BULL, EMMA  LIAVEK: WIZARD’S ROW [LIAVEK 3]  C/PA,PA
SHETTERLY, WILL  NEVERNEVER [BORDERLANDS 6]  C/P,P,H
SHETTERLY, WILL  TANGLED LANDS, THE  C/P,P
SHETTERLY, WILL  WITCHBLOOD  C/P,P
SHETTLE, ANDREA  FLUTE SONG MAGIC  P
SHIEL, M. P.  BLACK BOX, THE  H
SHIEL, M. P.  CHILDREN OF THE WIND  H
SHIEL, M. P.  COLD STEEL  H
SHIEL, M. P.  DR. KRASINSKI’S SECRET  H
SHIEL, M. P.  LORD OF THE SEA  H,LP
SHIEL, M. P.  LOST VIOL, THE  H
SHIEL, M. P.  PRINCE ZALESKI/CUMMINGS KING MONK  SR-H
SHIEL, M. P.  PURPLE CLOUD, THE  C/P,P,H
SHIEL, M. P.  YELLOW PERIL, THE  H
SHIMERMAN, ARMIN  CAPITAL OFFENSE [@MERCHANT PRINCE #3]  P
SHIMERMAN, ARMIN/SCOTT, MICHAEL  MERCHANT PRINCE, THE [MERCHANT PRINCE 1]  P

SHIMERMAN, ARMIN/ YARBRO, CHELSEA QUINN OUTRAGEOUS FORTUNE [@MERCHANT PRINCE #2] P

AUTHOR  TITLE  CODE
SHIMERMAN, ARMIN/GEORGE III, DAVID R.  34TH RULE, THE [STAR TREK DS9]  C/P,P
SHINER, LEWIS  DESERTED CITIES OF THE HEART  C/P,P,H
SHINER, LEWIS  FRONTERA  P
SHINER, LEWIS  GLIMPSES  P
SHINER, LEWIS  WHEN THE MUSIC’S OVER  PA
SHINEROCK, SIMON M.  DARK LAGOON, THE  P
SHINGLER, WILLIAM G. JR.  CROOKED COMPUTER, THE  P
SHINN, SHARON  ALLELUIA FILES, THE [SAMARIA]  LP
SHINN, SHARON  ANGEL-SEEKER [SAMARIA]  H,P
SHINN, SHARON  ANGELICA [SAMARIA]  H
SHINN, SHARON  ARCHANGEL [SAMARIA 1]  C/P:2,P
SHINN, SHARON  DARK MOON DEFENDER [TWELVE HOUSES]  H
SHINN, SHARON  HEART OF GOLD  C/P,LP,P
SHINN, SHARON  JENNA STARBORN  H,LP
SHINN, SHARON  JOVAH’S ANGEL [SAMARIA 2]  H,P
SHINN, SHARON  MYSTIC AND RIDER [TWELVE HOUSES]  H
SHINN, SHARON  SHAPE-CHANGER’S WIFE, THE  C/P,LP
SHINN, SHARON  SUMMERS AT CASTLE AUBURN  H,LP
SHINN, SHARON  THIRTEENTH HOUSE, THE [TWELVE HOUSES]  H
SHINN, SHARON  WRAPT IN CRYSTAL  H,P
SHINSEI, SHINSEI  STAR TREK: THE MAGNA [STAR TREK]  SCX
SHIPLETT, JUNE LUND  JOURNEY TO YESTERDAY [STACY GORDON 1]  P
SHIPLETT, JUNE LUND  RETURN TO YESTERDAY [STACY GORDON 2]  P
SHIPPEY, TOM  OXFORD BOOK OF FANTASY STORIES, THE  PA
SHIPPEY, TOM  OXFORD BOOK OF SCIENCE FICTION STORIES, THE  HA:2
SHIRAS, WILMAR H.  CHILDREN OF THE ATOM  H:2,P
SHIRLEY, JOHN  BLACK HOLE OF CARCOSA, THE [KAMUS OF KADIZHAR]  C/P,P
SHIRLEY, JOHN  CITY COME A-WALKIN’  P
SHIRLEY, JOHN  CONSTANTINE  P
SHIRLEY, JOHN  DRACULA IN LOVE  C/P,P
SHIRLEY, JOHN  ECLIPSE [SONG CALLED YOUTH 1]  C/P,LP
SHIRLEY, JOHN  ECLIPSE CORONA [SONG CALLED YOUTH 3]  C/P,P
SHIRLEY, JOHN  ECLIPSE PENUMBRA [SONG CALLED YOUTH 2]  P
SHIRLEY, JOHN  SPLENDID CHAOS, A  H
SHIRLEY, JOHN  THREE-RING PSYCHUS  P
SHIRLEY, ROBERT  TEENOCRACY  P
SHIROW, MASAMUNE  GHOST IN THE SHELL  CX
SHIROW, MASAMUNE  PROMETHEAN BALANCE, THE [@APPLESEED #4]  CX
SHIROW, MASAMUNE  PROMETHEAN CHALLENGE, THE [@APPLESEED #1]  CX
SHIROW, MASAMUNE  PROMETHEUS UNBOUND [@APPLESEED #2]  CX
SHIROW, MASAMUNE  SCALES OF PROMETHEUS, THE [@APPLESEED #3]  CX
SHOBIN, DAVID  SEEDING, THE  H
SHOBIN, DAVID  UNBORN, THE  H,P
SHOCK, JULIAN  EXTRATERRESTRIAL  C/P,P
SHOLS, W. W.  NEUN-PLANETEN-SPIEL, DAS  FP
SHOOTER, JIM  SECRET WARS [MARVEL’S FINEST]  CX
SHORTER, PHILIP  HANDFUL OF SILVER, A  P
SHOWALTER, GENA  AWAKEN ME DARKLY [ALIEN HUNTRESS 1]  LP
SHOWALTER, GENA  HEART OF THE DRAGON  P
SHRADER, ALAN ROSS  SATAN’S CHANCE  P
SHRYACK, DENNIS/BUTLER, MICHAEL  CAR, THE  P
SHUPP, MIKE  DEATH’S GRAY LAND [@DESTINY MAKERS #4]  C/P,P
SHUPP, MIKE  LAST RECKONING, THE [@DESTINY MAKERS #5]  C/P,P
SHUPP, MIKE  MORNING OF CREATION [@DESTINY MAKERS #2]  C/P:2,P
SHUPP, MIKE  SOLDIER OF ANOTHER FORTUNE [@DESTINY MAKERS #3]  P
SHUPP, MIKE  WITH FATE CONSPIRE [@DESTINY MAKERS 1]  C/P:2,P
SHUSTERMAN, NEAL  EYES OF KID MIDAS, THE  P
SHUSTERMAN, NEAL  SCORPION SHARDS [STAR SHARDS 1]  P
SHUSTERMAN, NEAL  SHATTERED SKY [STAR SHARDS 3]  P
SHUSTERMAN, NEAL  THIEF OF SOULS [STAR SHARDS 2]  P
SHUTE, NEVIL  IN THE WET  H,P
SHUTE, NEVIL  ON THE BEACH  C/P,P,H
SHUTE, NEVIL  ORDEAL  P
SHWARTZ, SUSAN  ARABESQUES  PA,C/PA
SHWARTZ, SUSAN  ARABESQUES 2  C/PA,PA
SHWARTZ, SUSAN  BYZANTIUM’S CROWN [HEIRS TO BYZANTIUM 1]  C/P,P
SHWARTZ, SUSAN  CROSS AND CRESCENT  H,LP
SHWARTZ, SUSAN  GRAIL OF HEARTS, THE  C/P,P,H
SHWARTZ, SUSAN  HABITATS  C/PA,PA
SHWARTZ, SUSAN  HECATE’S CAULDRON  C/PA,PA
SHWARTZ, SUSAN  HERITAGE OF FLIGHT  C/P:2,SFWA-P,P
SHWARTZ, SUSAN  MOONSINGER’S FRIENDS  HA,PA,SFWA-P
SHWARTZ, SUSAN  QUEENSBLADE  C/P,P
SHWARTZ, SUSAN  SECOND CHANCES  H

The MITSFS Pinkdex  September 24, 2007  Page 351
AUTHOR  TITLE  CODE
SHWARTZ, SUSAN  SHARDS OF EMPIRE  H,P
SHWARTZ, SUSAN  SILK ROADS AND SHADOWS  C/P:2,SFWA-P,P
SHWARTZ, SUSAN/GREENBERG, MARTIN H.  SISTERS IN FANTASY  PA
SHWARTZ, SUSAN/GREENBERG, MARTIN H.  SISTERS IN FANTASY 2  PA
SHWARTZ, SUSAN  WOMAN OF FLOWERS, THE [HEIRS TO BYZANTIUM 2]  C/P,P
SHY, TIMOTHY/SEARLE, RONALD  TERROR OF SAINT TRINIAN’S, THE  H
SICILIANO, SAM  BLOOD FEUD  P
SIDDONS, ANNE RIVERS  HOUSE NEXT DOOR, THE  P
SIDNEY-FRYER, DONALD  EMPEROR OF DREAMS: A CLARK ASHTON SMITH BIBLIOGRAPHY  R/IND
SIDNEY, KATHLEEN M.  MICHAEL AND THE MAGIC MAN  C/P,P
SIEGEL, BARBARA/SIEGEL, SCOTT  BURNING LAND, THE [@FIREBRATS #1]  P
SIEGEL, BARBARA/SIEGEL, SCOTT  SHOCKWAVE [@FIREBRATS #4]  P
SIEGEL, BARBARA/SIEGEL, SCOTT  SURVIVORS [@FIREBRATS #2]  P
SIEGEL, BARBARA/SIEGEL, SCOTT  TANIS, THE SHADOW YEARS [DRAGONLANCE PRELUDES II 3]  C/P:2,P
SIEGEL, BARBARA/SIEGEL, SCOTT  THUNDER MOUNTAIN [@FIREBRATS #3]  P
SIEGEL, JAN  DRAGON CHARMER, THE [FERN CHAPEL 2]  P
SIEGEL, JAN  PROSPERO’S CHILDREN [FERN CHAPEL 1]  P
SIEGEL, JAN  WITCH QUEEN, THE [FERN CHAPEL 3]  H
SIEGEL, MARTIN  AGENT OF ENTROPY  P
SIEGEL, MARTIN  UNREAL PEOPLE, THE  C/P,P
SIEGEL, ROBERT  ALPHA CENTAURI  LP
SIEGEL, ROBERT  WHALESONG  C/P,LP
SIEGEL, SCOTT  TALES FROM TETHEDRIL  PA
SIERRA STUDIOS  RAYMOND E. FEIST’S WORLD OF KRONDOR (CD) [RIFTWAR]  SR-MM
SIEVERT, JOHN  C.A.D.S. [@C.A.D.S. #1]  P
SIEVERT, JOHN  DOOM COMMANDER [@C.A.D.S. #7]  P
SIEVERT, JOHN  TECH BATTLEGROUND [@C.A.D.S. #2]  P
SIEVERT, JOHN  TECH COMMANDO [@C.A.D.S. #3]  P
SIEVERT, JOHN  TECH SATAN [@C.A.D.S. #5]  P
SIEVERT, JOHN  TECH STRIKE FORCE [@C.A.D.S. #4]  P
SIGLER, SCOTT  EARTHCORE  LP
SIGNET ECLIPSE  DRAGON LOVERS  LPA
SILAS, A. E.  PANORAMA EGG, THE  P

SILBERSACK, JOHN/ SCHELLING, CHRISTOPHER MAGIC OF CHRISTMAS, THE

PA

SILBERSACK, JOHN  ROGERS’ RANGERS [BUCK ROGERS (NIVEN & POURNELLE’S) 5]  P
SILENT, WILLIAM T.  LORD OF THE RED SUN  H:2
SILHOUETTE BOOKS  CHARMED DESTINIES  PA
SILKE, JAMES  LORDS OF DESTRUCTION [DEATH DEALER 2]  P,SFWA-P
SILKE, JAMES  PLAGUE OF KNIVES [DEATH DEALER 4]  C/P,SFWA-P,P
SILKE, JAMES  PRISONER OF THE HORNED HELMET [DEATH DEALER 1]  C/P:2,SFWA-P,P
SILKE, JAMES  TOOTH AND CLAW [DEATH DEALER 3]  C/P,SFWA-P,P
SILVA, DAVID B.  BEST OF THE HORROR SHOW  LPA
SILVA, JOSEPH  HOLOCAUST FOR HIRE  P
SILVA, JOSEPH  ISLAND OF DOCTOR MOREAU  P
SILVER, STEVEN H./GREENBERG, MARTIN H. MAGICAL BEGINNINGS  PA
SILVER, STEVEN H./GREENBERG, MARTIN H. WONDROUS BEGINNINGS  C/PA,PA
SILVERBERG, ROBERT  ACROSS A BILLION YEARS  C/P,SFWA-P,P,H
SILVERBERG, ROBERT  ALIEN YEARS, THE  H:2,P
SILVERBERG, ROBERT  ALPHA 1  C/PA,PA
SILVERBERG, ROBERT  ALPHA 2  C/PA,PA
SILVERBERG, ROBERT  ALPHA 3  C/PA,PA
SILVERBERG, ROBERT  ALPHA 4  C/PA,PA
SILVERBERG, ROBERT  ALPHA 5  C/PA,PA
SILVERBERG, ROBERT  ALPHA 6  C/PA,PA
SILVERBERG, ROBERT  ALPHA 7  C/PA,PA
SILVERBERG, ROBERT  ALPHA 8  PA,C/PA
SILVERBERG, ROBERT  ALPHA 9  C/PA,PA
SILVERBERG, ROBERT  ANDROIDS ARE COMING, THE  LPA

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

ARBOR HOUSE TREASURY OF GREAT SCIENCE FICTION SHORT NOVELS, THE

HA

SILVERBERG, ROBERT  AT WINTER’S END [WINTER 1]  H:2,P
SILVERBERG, ROBERT  BEST OF NEW DIMENSIONS, THE  PA
SILVERBERG, ROBERT  BEST OF ROBERT SILVERBERG, THE  C/P:3,P
SILVERBERG, ROBERT  BETWEEN WORLDS  HA
SILVERBERG, ROBERT  BEYOND CONTROL  HA,PA

SILVERBERG, ROBERT/BRUNNER, JOHN/ YARBRO, CHELSEA QUINN BEYOND THE GATE OF WORLDS

PA,SFWA-P

SILVERBERG, ROBERT  BEYOND THE SAFE ZONE  C/P:2,P
SILVERBERG, ROBERT  BOOK OF SKULLS, THE  H:2,P

SILVERBERG, ROBERT BORN WITH THE DEAD C/P,SFWA-TD3,H,P,TD3,

SFWA

SILVERBERG, ROBERT  CALIBRATED ALLIGATOR AND OTHER SCIENCE FICTION STORIES, THE  H
SILVERBERG, ROBERT  CAPRICORN GAMES  H

The MITSFS Pinkdex  September 24, 2007  Page 352
AUTHOR  TITLE  CODE

SILVERBERG, ROBERT/ GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

CAR SINISTER

PA

SILVERBERG, ROBERT  CENTURY OF FANTASY: 1980-1989, A  HA
SILVERBERG, ROBERT  CENTURY OF SCIENCE FICTION: 1950-1959, A  HA
SILVERBERG, ROBERT  CHAINS OF THE SEA  HA,PA
SILVERBERG, ROBERT  CHALLENGE OF CLIMATE, THE  [C/RAN-H]
SILVERBERG, ROBERT  COLLISION COURSE  DF123,C/P:2,SFWA-P
SILVERBERG, ROBERT  CONGLOMEROID COCKTAIL PARTY, THE  C/P:2,P,H
SILVERBERG, ROBERT  CONQUERORS FROM THE DARKNESS  C/P:2,P
SILVERBERG, ROBERT  CRYSTAL SHIP, THE  C/PA,PA
SILVERBERG, ROBERT  CUBE ROOT OF UNCERTAINTY, THE  C/P,P,H
SILVERBERG, ROBERT/FAWCETT, BILL  DANGEROUS INTERFACES [@TIME GATE #2]  PA,C/PA,SFWA-P
SILVERBERG, ROBERT  DARK STARS  PA,C/PA
SILVERBERG, ROBERT  DEEP SPACE  HA,PA
SILVERBERG, ROBERT  DIMENSION THIRTEEN  C/P:2,P
SILVERBERG, ROBERT  DOWNWARD TO THE EARTH  H:2,P
SILVERBERG, ROBERT  DYING INSIDE  C/P,SFWA,P,H
SILVERBERG, ROBERT  EARTH IS THE STRANGEST PLANET  HA
SILVERBERG, ROBERT  EARTHMEN AND STRANGERS  PA,C/PA
SILVERBERG, ROBERT  EARTH’S OTHER SHADOW  C/P,P,H
SILVERBERG, ROBERT  ENDS OF TIME, THE  C/PA,PA
SILVERBERG, ROBERT  EXPLORERS OF SPACE  HA
SILVERBERG, ROBERT  FACE OF THE WATERS, THE  H:2,P

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

FANTASY HALL OF FAME, THE (SILVERBERG & GREENBERG)

HA

SILVERBERG, ROBERT  FANTASY HALL OF FAME, THE (SILVERBERG ONLY)  HA,LPA
SILVERBERG, ROBERT/HABER, KAREN  FANTASY THE BEST OF 2001  PA
SILVERBERG, ROBERT  FAR HORIZONS  C/PA,PA
SILVERBERG, ROBERT  FEAST OF ST. DIONYSUS  H:2,SFWA,P
SILVERBERG, ROBERT  FIRST AMERICAN INTO SPACE  [C/RAN-P]
SILVERBERG, ROBERT  GALACTIC DREAMERS  SFWA
SILVERBERG, ROBERT  GATE OF WORLDS, THE  C/P:2,P
SILVERBERG, ROBERT  GILGAMESH THE KING [GILGAMESH]  C/P,P,H:2
SILVERBERG, ROBERT  GODLING, GO HOME!  P

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

GREAT SCIENCE FICTION OF THE 20TH CENTURY

HA

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

GREAT SF STORIES 1964, THE

HA

SILVERBERG, ROBERT  GREAT SHORT NOVELS OF SCIENCE FICTION  C/PA,PA
SILVERBERG, ROBERT  HAWKSBILL STATION  SFWA-TD26
SILVERBERG, ROBERT  HAWKSBILL STATION (EXPANDED)  H:2,P
SILVERBERG, ROBERT  HOT SKY AT MIDNIGHT  H:2,P
SILVERBERG, ROBERT  IN ANOTHER COUNTRY [VINTAGE SEASON 2]  SFWA-TD18,TD18
SILVERBERG, ROBERT  IN THE BEGINNING  SR-HA
SILVERBERG, ROBERT  INFINITE JESTS: THE LIGHTER SIDE OF SCIENCE FICTION  HA
SILVERBERG, ROBERT  INFINITE WEB, THE  HA
SILVERBERG, ROBERT  INVADERS FROM EARTH  DD286,P,C/P:2
SILVERBERG, ROBERT  INVADERS FROM EARTH/TO WORLDS BEYOND  P
SILVERBERG, ROBERT  INVADERS FROM SPACE  HA
SILVERBERG, ROBERT  KING OF DREAMS, THE [MAJIPOOR PRESTIMION 3]  H,P
SILVERBERG, ROBERT  KINGDOMS OF THE WALL  H:2,P
SILVERBERG, ROBERT  KINGDOMS OF THE WALL (PROOF)  SR-LP
SILVERBERG, ROBERT  LEGENDS (I) (VOL. 2)  C/PA,PA
SILVERBERG, ROBERT  LEGENDS (I) (VOL.S 1-3)  HA
SILVERBERG, ROBERT  LEGENDS II  HA
SILVERBERG, ROBERT  LETTERS FROM ATLANTIS  P,SFWA
SILVERBERG, ROBERT  LION TIME IN TIMBUCTOO  SR-H
SILVERBERG, ROBERT  LONGEST WAY HOME, THE  H
SILVERBERG, ROBERT  LORD OF DARKNESS  H
SILVERBERG, ROBERT  LORD PRESTIMION [MAJIPOOR PRESTIMION 2]  H:2,P
SILVERBERG, ROBERT  LORD VALENTINE’S CASTLE [MAJIPOOR 1]  H:2,SFWA,P
SILVERBERG, ROBERT  LORD VALENTINE’S CASTLE (PROOF) [MAJIPOOR 1]  SR-LP
SILVERBERG, ROBERT  LOST CITIES AND VANISHED CIVILIZATIONS  [C/RAN-P]
SILVERBERG, ROBERT  LOST RACE OF MARS  C/P:2,P
SILVERBERG, ROBERT  LOST WORLDS, UNKNOWN HORIZONS  HA
SILVERBERG, ROBERT  MAJIPOOR CHRONICLES [MAJIPOOR 2]  H:2,P
SILVERBERG, ROBERT  MAN IN THE MAZE, THE  C/P:2,P
SILVERBERG, ROBERT  MASKS OF TIME, THE  C/P:2,P
SILVERBERG, ROBERT  MASKS OF TIME, THE/BORN WITH THE DEAD/DYING INSIDE  P
SILVERBERG, ROBERT  MASTER OF LIFE AND DEATH  SFWA-P,DD237,P,C/P:2
SILVERBERG, ROBERT  MEN AND MACHINES  C/PA,PA
SILVERBERG, ROBERT  MIND TO MIND  HA,PA
SILVERBERG, ROBERT  MIRROR OF INFINITY, THE  PA
SILVERBERG, ROBERT  MOONFERNS AND STARSONGS  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 353
AUTHOR  TITLE  CODE
SILVERBERG, ROBERT  MOUNTAINS OF MAJIPOOR, THE  H:2
SILVERBERG, ROBERT  MURASAKI  SR-HA,C/PA,PA
SILVERBERG, ROBERT  MUTANTS  HA

SILVERBERG, ROBERT/ GREENBERG, MARTIN H./ WAUGH, CHARLES G.

NEANDERTHALS

C/PA,PA

SILVERBERG, ROBERT  NEBULA AWARDS 18, THE  C/PA,PA
SILVERBERG, ROBERT  NEBULA AWARDS SHOWCASE 2001  LPA
SILVERBERG, ROBERT  NEEDLE IN A TIMESTACK  C/P:2,P
SILVERBERG, ROBERT  NEW ATLANTIS, THE  C/PA,PA
SILVERBERG, ROBERT  NEW DIMENSIONS 1  C/PA,PA
SILVERBERG, ROBERT  NEW DIMENSIONS 2  C/PA,PA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 3  C/PA,PA
SILVERBERG, ROBERT  NEW DIMENSIONS 4  C/PA,PA
SILVERBERG, ROBERT  NEW DIMENSIONS 5  HA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 6  HA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 7  HA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 8  HA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 9  HA,SFWA
SILVERBERG, ROBERT  NEW DIMENSIONS 9 (PROOF)  SR-LPA
SILVERBERG, ROBERT  NEW DIMENSIONS 10  HA
SILVERBERG, ROBERT/RANDALL, MARTA  NEW DIMENSIONS 11  C/PA,PA
SILVERBERG, ROBERT  NEW SPRINGTIME, THE [WINTER 2]  C/P,P,H:2
SILVERBERG, ROBERT  NEXT STOP THE STARS  C/P,DF145,SFWA-P

SILVERBERG, ROBERT NIGHTWINGS C/P:2,TD15,SFWA-TD15,

P

SILVERBERG, ROBERT  NO MIND OF MAN  C/PA,PA
SILVERBERG, ROBERT  OTHER DIMENSIONS  HA,PA
SILVERBERG, ROBERT  PAIR FROM SPACE: WE, THE MARAUDERS, A  SR-DB0612
SILVERBERG, ROBERT  PARSECS AND PARABLES  H
SILVERBERG, ROBERT  PLANET KILLERS, THE  DD407
SILVERBERG, ROBERT  PROJECT PENDULUM  C/P,SFWA-P,P,H
SILVERBERG, ROBERT  QUEEN OF SPRINGTIME, THE [WINTER 2]  LP
SILVERBERG, ROBERT  REALITY TRIP AND OTHER IMPLAUSIBILITIES, THE  C/P:2,P
SILVERBERG, ROBERT  RECALLED TO LIFE  C/P:2,SFWA,P
SILVERBERG, ROBERT  REGAN’S PLANET  C/P,P
SILVERBERG, ROBERT  REVOLT ON ALPHA C  C/P,P,H
SILVERBERG, ROBERT  ROBERT SILVERBERG OMNIBUS, A  H,SFWA
SILVERBERG, ROBERT  ROBERT SILVERBERG’S WORLDS OF WONDER  LPA
SILVERBERG, ROBERT  ROMA ETERNA  H
SILVERBERG, ROBERT  SAILING TO BYZANTIUM  SFWA-TD10,TD10
SILVERBERG, ROBERT  SAILING TO BYZANTIUM (COLLECTION)  LP
SILVERBERG, ROBERT  SCIENCE FICTION BESTIARY, THE  HA,PA
SILVERBERG, ROBERT  SCIENCE FICTION HALL OF FAME  C/PA,PA
SILVERBERG, ROBERT/HABER, KAREN  SCIENCE FICTION THE BEST OF 2001  C/PA,PA
SILVERBERG, ROBERT/HABER, KAREN  SCIENCE FICTION THE BEST OF 2002  PA

SILVERBERG, ROBERT/WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SCIENCE FICTIONAL DINOSAUR, THE

PA

SILVERBERG, ROBERT  SECOND TRIP, THE  H:2,P
SILVERBERG, ROBERT  SECRET SHARER, THE  H
SILVERBERG, ROBERT  SEED OF EARTH, THE  P,DF145
SILVERBERG, ROBERT  SHADRACH IN THE FURNACE  C/P,P,H
SILVERBERG, ROBERT  SHORES OF TOMORROW, THE  H
SILVERBERG, ROBERT  SILENT INVADERS, THE  C/P,DF195,SFWA-P,P
SILVERBERG, ROBERT  SON OF MAN  C/P:2,P
SILVERBERG, ROBERT  SORCERERS OF MAJIPOOR [MAJIPOOR PRESTIMION 1]  H:2,P
SILVERBERG, ROBERT  STAR OF GYPSIES  H:2,P
SILVERBERG, ROBERT  STARBORNE  C/P,P,H
SILVERBERG, ROBERT  STARMAN’S QUEST  H
SILVERBERG, ROBERT  STEPSONS OF TERRA  C/P,DD311
SILVERBERG, ROBERT  STOCHASTIC MAN, THE  C/P:2,P,H
SILVERBERG, ROBERT  STRANGE GIFTS  HA
SILVERBERG, ROBERT  SUNDANCE AND OTHER SCIENCE FICTION STORIES  H
SILVERBERG, ROBERT  SUNRISE ON MERCURY  H
SILVERBERG, ROBERT  THEBES OF THE HUNDRED GATES  C/P,P
SILVERBERG, ROBERT  13TH IMMORTAL, THE  DD223
SILVERBERG, ROBERT  THORNS  C/P:2,P
SILVERBERG, ROBERT  THOSE WHO WATCH  C/P:2,P
SILVERBERG, ROBERT  THREADS OF TIME  SR-HA,C/PA,PA
SILVERBERG, ROBERT  THREE NOVELS  P
SILVERBERG, ROBERT  THREE SURVIVED  H
SILVERBERG, ROBERT/FAWCETT, BILL  TIME GATE [@TIME GATE #1]  PA,C/PA
SILVERBERG, ROBERT  TIME HOPPERS, THE  H:2,P
SILVERBERG, ROBERT  TIME OF CHANGES, A  H:2,P
SILVERBERG, ROBERT  TIME OF THE GREAT FREEZE  C/P:2,SFWA-P,P

The MITSFS Pinkdex  September 24, 2007  Page 354
AUTHOR  TITLE  CODE

SILVERBERG, ROBERT/ GREENBERG, MARTIN H.

TIME TRAVELERS, THE

C/PA,PA

SILVERBERG, ROBERT  TO BE CONTINUED  H
SILVERBERG, ROBERT  TO LIVE AGAIN  H:2,P
SILVERBERG, ROBERT  TO OPEN THE SKY  C/P:2,P
SILVERBERG, ROBERT  TO THE LAND OF THE LIVING [GILGAMESH]  C/P,P
SILVERBERG, ROBERT  TO THE STARS  HA
SILVERBERG, ROBERT  TOM O’BEDLAM  C/P,SFWA,P,H
SILVERBERG, ROBERT  TOMORROW’S WORLDS  HA,PA
SILVERBERG, ROBERT  TOWER OF GLASS  C/P,P,H
SILVERBERG, ROBERT  TRIAX  C/PA,PA
SILVERBERG, ROBERT  TRIPS IN TIME  HA
SILVERBERG, ROBERT  UNFAMILIAR TERRITORY  C/P,SFWA,P,H
SILVERBERG, ROBERT/HABER, KAREN  UNIVERSE 1 (SECOND SERIES)  PA
SILVERBERG, ROBERT/HABER, KAREN  UNIVERSE 2 (SECOND SERIES)  HA,LPA
SILVERBERG, ROBERT/HABER, KAREN  UNIVERSE 3 (SECOND SERIES)  C/PA,PA
SILVERBERG, ROBERT  UP THE LINE  C/P:2,P
SILVERBERG, ROBERT  VALENTINE PONTIFEX [MAJIPOOR 3]  H:2,P
SILVERBERG, ROBERT  VALLEY BEYOND TIME  C/P:2,P
SILVERBERG, ROBERT  VOYAGERS IN TIME  PA,C/PA
SILVERBERG, ROBERT  WINDOWS INTO TOMORROW  HA,PA
SILVERBERG, ROBERT  WORLD INSIDE, THE  H:2,SFWA,P
SILVERBERG, ROBERT  WORLD OF A THOUSAND COLORS  H:2,P
SILVERBERG, ROBERT  WORLD’S FAIR, 1992  C/P,HA,P
SILVERBERG, ROBERT  WORLDS OF MAYBE  PA
SIMAK, CLIFFORD D.  ALL FLESH IS GRASS  C/P,P,H
SIMAK, CLIFFORD D.  ALL THE TRAPS OF EARTH AND OTHER STORIES  H:2,P
SIMAK, CLIFFORD D.  BEST OF CLIFFORD D. SIMAK, THE  SR-P
SIMAK, CLIFFORD D.  BEST SCIENCE FICTION STORIES OF CLIFFORD D. SIMAK  H:2,P
SIMAK, CLIFFORD D.  CEMETERY WORLD  H:2,P
SIMAK, CLIFFORD D.  CHOICE OF GODS, A  C/P,P,H
SIMAK, CLIFFORD D.  CITY  C/P,SFWA-P,P,H
SIMAK, CLIFFORD D.  CITY (INCLUDES "EPILOG")  P
SIMAK, CLIFFORD D.  COSMIC ENGINEERS  SR-H,C/P:2,P
SIMAK, CLIFFORD D.  DESTINY DOLL  C/P:2,P
SIMAK, CLIFFORD D.  EMPIRE  C/P:2,GN2.07
SIMAK, CLIFFORD D.  ENCHANTED PILGRIMAGE  H:2,P
SIMAK, CLIFFORD D.  FELLOWSHIP OF THE TALISMAN, THE  H:2,P
SIMAK, CLIFFORD D.  FIRST HE DIED  C/P:2,P
SIMAK, CLIFFORD D.  GOBLIN RESERVATION, THE  C/P,P,H
SIMAK, CLIFFORD D.  HERITAGE OF STARS, A  H:2,P
SIMAK, CLIFFORD D.  HIGHWAY OF ETERNITY  C/P,SFWA,P,H
SIMAK, CLIFFORD D.  MARATHON PHOTOGRAPH, THE  P
SIMAK, CLIFFORD D.  MASTODONIA  H:2,P
SIMAK, CLIFFORD D.  NEBULA AWARD STORIES 6  C/PA,PA,SFWA
SIMAK, CLIFFORD D.  OTHER WORLDS OF CLIFFORD SIMAK  C/P:2,P
SIMAK, CLIFFORD D.  OUR CHILDREN’S CHILDREN  C/P,P,H
SIMAK, CLIFFORD D.  OUT OF THEIR MINDS  C/P:2,P
SIMAK, CLIFFORD D.  OVER THE RIVER AND THROUGH THE WOODS  H
SIMAK, CLIFFORD D.  PROJECT POPE  H:2,P
SIMAK, CLIFFORD D.  RING AROUND THE SUN  H:2,P,SR-DD061
SIMAK, CLIFFORD D.  SHAKESPEARE’S PLANET  H:2,P
SIMAK, CLIFFORD D.  SKIRMISH  C/P,P,H
SIMAK, CLIFFORD D.  SO BRIGHT THE VISION  C/P,DH095
SIMAK, CLIFFORD D.  SPECIAL DELIVERANCE  C/P,P,H:2
SIMAK, CLIFFORD D.  STRANGERS IN THE UNIVERSE  H:2,P
SIMAK, CLIFFORD D.  THEY WALKED LIKE MEN  C/P:2,P
SIMAK, CLIFFORD D.  TIME AND AGAIN  C/P,SFWA,P,H
SIMAK, CLIFFORD D.  TIME IS THE SIMPLEST THING  H:2,SFWA,P
SIMAK, CLIFFORD D.  TROUBLE WITH TYCHO, THE  DD517,C/P:2,P
SIMAK, CLIFFORD D.  VISITORS, THE  H:2,P
SIMAK, CLIFFORD D.  WAY STATION  H:2,P,SFWA
SIMAK, CLIFFORD D.  WEREWOLF PRINCIPLE, THE  H:2,P
SIMAK, CLIFFORD D.  WHERE THE EVIL DWELLS  H:2,P
SIMAK, CLIFFORD D.  WHY CALL THEM BACK FROM HEAVEN?  C/P,P,H
SIMAK, CLIFFORD D.  WORLDS OF CLIFFORD SIMAK, THE  H:2,P
SIMAK, CLIFFORD D.  WORLDS WITHOUT END  C/P:2,P
SIMMONS, DAN  CARRION COMFORT  C/P,P
SIMMONS, DAN  CHILDREN OF THE NIGHT  C/P,P,H
SIMMONS, DAN  CROOK FACTORY, THE  P
SIMMONS, DAN  DARWIN’S BLADE  H
SIMMONS, DAN  ENDYMION [HYPERION 3]  H:2,P
SIMMONS, DAN  FALL OF HYPERION, THE [HYPERION 2]  C/P:2,P,H
SIMMONS, DAN  FIRES OF EDEN  C/P,P,H
SIMMONS, DAN  HOLLOW MAN, THE  C/P,P,H

AUTHOR  TITLE  CODE
SIMMONS, DAN  HYPERION [HYPERION 1]  C/P:2,P
SIMMONS, DAN  HYPERION CANTOS [HYPERION 1,2]  H:3
SIMMONS, DAN  ILIUM [ILIUM 1]  H
SIMMONS, DAN  LOVEDEATH  C/P,P
SIMMONS, DAN  LOVEDEATH (PROOF)  SR-LP
SIMMONS, DAN  OLYMPOS [ILIUM 2]  H
SIMMONS, DAN  PHASES OF GRAVITY  C/P:2,P
SIMMONS, DAN  PRAYERS TO BROKEN STONES  C/P:2,P
SIMMONS, DAN  RISE OF ENDYMION, THE [HYPERION 4]  H:2,P
SIMMONS, DAN  SONG OF KALI  C/P,SFWA-P,P,H
SIMMONS, DAN  SUMMER OF NIGHT  C/P,P
SIMMONS, DAN  TERROR, THE  H
SIMMONS, DAN  WORLDS ENOUGH & TIME  H,LP
SIMMONS, GEOFFREY  ADAM EXPERIMENT, THE  P
SIMMONS, GEOFFREY  MURDOCK  H
SIMMONS, GEOFFREY  PANDEMIC  P
SIMMONS, WM. MARK  DEAD EASY [HALFLIFE 4]  H
SIMMONS, WM. MARK  DEAD ON MY FEET [HALFLIFE 2]  H
SIMMONS, WM. MARK  DREAMLAND CHRONICLES [DREAMLAND 1,2,3]  LP
SIMMONS, WM. MARK  IN THE NET OF DREAMS [DREAMLAND 1]  P
SIMMONS, WM. MARK  ONE FOOT IN THE GRAVE [HALFLIFE 1]  C/P:2,P
SIMON, JO ANN  HOLD FAST TO LOVE  P
SIMON, JO ANN  LOVE ONCE AGAIN  P
SIMON, JO ANN  LOVE ONCE IN PASSING  P
SIMON, LEONARD  REBORN  C/P,P
SIMON, MORRIS  BLACK VESSEL, THE [MYSTARA]  P
SIMON SPOTLIGHT ENTERTAINMENT  TALES OF THE SLAYER, VOL. 4 [@BUFFY]  LPA
SIMON SPOTLIGHT ENTERTAINMENT  WARREN WITCHES, THE [CHARMED]  LPA
SIMON, TOBY  SEARCH FOR PLANET X, THE  [C/RAN-P]
SIMONSON, LOUISE  SUPERMAN: DOOMSDAY & BEYOND [DC COMICS]  LP
SIMPSON, GEORGE E./BURGER, NEAL R.  GHOSTBOAT  P
SIMPSON, GEORGE GAYLORD  DECHRONIZATION OF SAM MAGRUDER, THE  H
SIMPSON, MARGARET  STRANGE ORBIT  LP
SIMPSON, PAUL/HUGHES, DAVID  FARSCAPE: THE ILLUSTRATED COMPANION [FARSCAPE]  R/REF-P
SIMPSON, ROBERT  LESSER EVIL [STAR TREK DS9 MISSION GAMMA #4]  C/P,P
SIMS, MARGOT  ON THE NECESSITY OF BESTIALIZING THE HUMAN FEMALE  [C/RAN-P]
SINCLAIR, ALISON  BLUEHEART  P
SINCLAIR, ALISON  LEGACIES  C/P,P
SINCLAIR, ANDREW  GOG  P
SINCLAIR, DAVID  GILEAD BOMB, THE  C/P,P
SINCLAIR, LINNEA  ACCIDENTAL GODDESS, AN  P
SINCLAIR, LINNEA  FINDERS KEEPERS  P
SINCLAIR, LINNEA  GABRIEL’S GHOST  P
SINCLAIR, LINNEA  GAMES OF COMMAND  P
SINCLAIR, LINNEA  WINTERTIDE  P
SINCLAIR, UPTON  GNOMOBILE, THE  P
SINCLAIR, UPTON  WORLD’S END  H
SINGER, ISAAC BASHEVIS  SATAN IN GORAY  LP
SINGER, JUDITH  THRESHOLD  P
SINGER, KURT  KURT SINGER’S GHOST OMNIBUS  SR-PA
SINGER, KURT  TALES FROM THE UNKNOWN  PA
SINGER, KURT  UNEARTHLY, THE  [C/RAN-P]
SINGER, MARILYN  HORSEMASTER  H
SINGLETON, JACQUI  HEARTSTONE & SABER  LP
SINISALO, JOHANNA  DEDALUS BOOK OF FINNISH FANTASY, THE  LPA
SIODMAK, CURT  CITY IN THE SKY  H:2,P
SIODMAK, CURT  DONOVAN’S BRAIN  C/P:2,P
SIODMAK, CURT  HAUSER’S MEMORY  H:2,P
SIODMAK, CURT  RIDERS TO THE STARS  C/P:2,P
SIODMAK, CURT  SKYPORT  C/P,P,H
SIODMAK, CURT  THIRD EAR, THE  C/P,P,H:2
SIROIS, A. L.  BEGINNINGS OF FOREVER, THE  LP
SIROIS, A. L.  BLIND AMBITIONS  LP
SIROIS, A. L.  BLOOD RELATIONS  LP
SIROTA, MIKE  BERBORA [@BERBORA #1]  C/P,P

SIROTA, MIKE BICYCLING THROUGH SPACE AND TIME

[BICYCLING THROUGH SPACE AND TIME 1]

C/P,SFWA-P:2,P

SIROTA, MIKE  CAVES OF REGLATHIUM, THE [@DANNUS #3]  C/P,P
SIROTA, MIKE  CONQUERORS OF REGLATHIUM, THE [@DANNUS #2]  P
SIROTA, MIKE  DARK STRAITS OF REGLATHIUM, THE [@DANNUS #4]  P
SIROTA, MIKE  DEMON SHADOWS  P
SIROTA, MIKE  DEMONS OF ZAMMAR, THE [@RO-LAN #4]  P
SIROTA, MIKE  FLIGHT FROM BERBORA [@BERBORA #2]  P
SIROTA, MIKE  GOLDEN HAWK OF ZANDRAYA, THE  P

AUTHOR  TITLE  CODE
SIROTA, MIKE  JOURNEY TO MESHARRA [@RO-LAN #3]  P
SIROTA, MIKE  MASTER OF BORANGA [@RO-LAN #1]  P
SIROTA, MIKE  PRISONER OF REGLATHIUM, THE [@DANNUS #1]  P
SIROTA, MIKE  SHROUDED WALKS OF BORANGA, THE [@RO-LAN #2]  P
SIROTA, MIKE  SLAVES OF REGLATHIUM, THE [@DANNUS #5]  P
SIROTA, MIKE  TWENTIETH SON OF ORNON, THE  C/P,P
SIROTA, MIKE  ULTIMATE BIKE PATH, THE [BICYCLING THROUGH SPACE AND TIME 2]  C/P,SFWA-P,P
SIROTA, MIKE  WELL, THE  P
SITCHIN, ZECHARIA  DIVINE ENCOUNTERS  R/REF-P
SITCHIN, ZECHARIA  12TH PLANET, THE  [C/RAN-P]
SIXBURY, GLENN R.  LEGACY [EARTH FINAL CONFLICT]  LP
SIZEMORE, SUSAN  COMPANIONS [@LAWS OF THE BLOOD 3]  C/P,P
SIZEMORE, SUSAN  DECEPTIONS [@LAWS OF THE BLOOD 4]  C/P:2,P
SIZEMORE, SUSAN  HEROES [@LAWS OF THE BLOOD 5]  P
SIZEMORE, SUSAN  HUNT, THE [@LAWS OF THE BLOOD 1]  C/P,P
SIZEMORE, SUSAN  PARTNERS [@LAWS OF THE BLOOD 2]  C/P,P
SKAL, DAVID J.  ANTIBODIES  H,P
SKAL, DAVID J.  SCAVENGERS  P
SKAL, DAVID J.  WHEN WE WERE GOOD  P
SKALDASPILLIR, SIGFRIOUR  WITCH’S WELCOME, A [ERIC BRIGHTEYES 2]  P
SKELTON, MATTHEW  ENDYMION SPRING  H
SKELTON, RED  RED SKELTON’S FAVORITE GHOST STORIES  PA
SKETCHLEY, MARTIN  AFFINITY TRAP, THE [STRUCTURE 1]  LP
SKETCHLEY, MARTIN  DESTINY MASK, THE [STRUCTURE 2]  LP
SKIMIN, ROBERT  GRAY VICTORY  H
SKINKLE, DOROTHY  STAR GIANT  P
SKINNER, AINSLIE  HARROWING, THE  P
SKINNER, B. F.  WALDEN TWO  C/P,P,H
SKINNER, HUBERT M.  READINGS IN FOLK-LORE  [C/RAN-H]
SKIPP, JOHN/SPECTOR, CRAIG  ANIMALS  C/P:2,P
SKIPP, JOHN/SPECTOR, CRAIG  BOOK OF THE DEAD  PA
SKIPP, JOHN/SPECTOR, CRAIG  BRIDGE, THE  P
SKIPP, JOHN/SPECTOR, CRAIG  DEAD LINES  P
SKIPP, JOHN/SPECTOR, CRAIG  FRIGHT NIGHT  P
SKIPP, JOHN/SPECTOR, CRAIG  LIGHT AT THE END, THE  C/P,P
SKOLSKY, SID  AFFECTIONISM SOCIETY, THE  SR-P
SKY, KATHLEEN  BIRTHRIGHT  SR-P,C/P
SKY, KATHLEEN  DEATH’S ANGEL [STAR TREK]  C/P:2,P
SKY, KATHLEEN  ICE PRISON  SR-P
SKY, KATHLEEN  VULCAN! [STAR TREK]  C/P,P
SKY, KATHLEEN  WITCHDAME  C/P:2,P
SLADEK, JOHN  BEST OF JOHN SLADEK, THE  P
SLADEK, JOHN  BLACK AURA  P
SLADEK, JOHN  KEEP THE GIRAFFE BURNING  SR-P
SLADEK, JOHN  MAPS  LP
SLADEK, JOHN  MECHASM  C/P:2,P
SLADEK, JOHN  MULLER-FOKKER EFFECT, THE  C/P:2,P,H
SLADEK, JOHN  REPRODUCTIVE SYSTEM, THE  LP
SLADEK, JOHN  RODERICK [RODERICK THE ROBOT 1]  C/P:2,SFWA-P,P
SLADEK, JOHN  RODERICK AT RANDOM [RODERICK THE ROBOT 2]  P
SLADEK, JOHN  STEAM-DRIVEN BOY, THE  SR-P,C/P
SLADEK, JOHN  TIK-TOK  C/P,P
SLATER, HENRY J.  SHIP OF DESTINY  H
SLATER, HENRY J.  SMASHED WORLD, THE  H
SLATER, JIM  BOY WHO SAVED EARTH, THE  H,P
SLATER, PHILIP  EARTHWALK  [C/RAN-P]
SLAUGHTER, FRANK G.  DEVIL’S GAMBLE  H,P
SLAUGHTER, FRANK G.  EPIDEMIC!  P
SLAVICSEK, BILL/TRAMONTANA, C. J.  STORM KNIGHTS [@TORG POSSIBILITY WARS #1]  P
SLEATOR, WILLIAM  BOY WHO REVERSED HIMSELF, THE  P
SLEATOR, WILLIAM  DUPLICATE, THE  H,P
SLEATOR, WILLIAM  GREEN FUTURES OF TYCHO, THE  C/P,P
SLEATOR, WILLIAM  HOUSE OF STAIRS  C/P:2,P
SLEATOR, WILLIAM  INTERSTELLAR PIG  H,P
SLEATOR, WILLIAM  INTO THE DREAM  C/P,P
SLEATOR, WILLIAM  MARCO’S MILLIONS [MARCO 1]  H,LP
SLEATOR, WILLIAM  SINGULARITY  C/P:2,P
SLESAR, HENRY  20 MILLION MILES TO EARTH  SR-P,C/P
SLOANE, BEN  BLOWN DEAD [HORN]  P
SLOANE, BEN  HOT ZONE [HORN]  P
SLOANE, BEN  OUTLAND STRIP [HORN]  P
SLOANE, WILLIAM  EDGE OF RUNNING WATER, THE  H:2,SR-L,P
SLOANE, WILLIAM  SPACE, SPACE, SPACE  HA
SLOANE, WILLIAM  TO WALK THE NIGHT  C/P,P,H

AUTHOR  TITLE  CODE
SLOBODKIN, LOUIS  SPACE SHIP UNDER THE APPLE TREE, THE  C/P,P
SLOBODKIN, LOUIS  THREE-SEATED SPACE SHIP, THE  C/P,P,H
SLONAKER, LARRY  VOICE OF THE VISITOR  C/P:2,P
SLONCZEWSKI, JOAN  BRAIN PLAGUE [ELYSIUM]  H
SLONCZEWSKI, JOAN  CHILDREN STAR, THE [ELYSIUM]  H:2,P
SLONCZEWSKI, JOAN  DAUGHTER OF ELYSIUM [ELYSIUM]  C/P,P
SLONCZEWSKI, JOAN  DOOR INTO OCEAN, A [ELYSIUM]  H:2,P,SFWA
SLONCZEWSKI, JOAN  STILL FORMS ON FOXFIELD  C/P,P
SLONCZEWSKI, JOAN  WALL AROUND EDEN, THE  C/P,P,H
SLOTE, ALFRED  MY ROBOT BUDDY  C/P,LP
SLOTE, ALFRED  MY TRIP TO ALPHA 1  LP
SLUNG, MICHELE  I SHUDDER AT YOUR TOUCH  C/PA,PA
SLUSSER, GEORGE EDGAR  DELANY INTERSECTION, THE  R/REF-P,C/REF-P
SLUSSER, GEORGE EDGAR  FARTHEST SHORES OF URSULA K. LE GUIN, THE  R/REF-P
SLUSSER, GEORGE EDGAR  HARLAN ELLISON: UNREPENTANT HARLEQUIN  R/REF-P
SLUSSER, GEORGE EDGAR  ROBERT A. HEINLEIN: STRANGER IN HIS OWN LAND  R/REF-P
SLUSSER, GEORGE EDGAR  SPACE ODYSSEYS OF ARTHUR C. CLARKE, THE  R/REF-P
SLUYSER, M.  VAN MOLECUUL TOT MENS  FH
SMALL, AUSTIN J.  AVENGING RAY, THE  H
SMEDMAN, LISA  APPARITION TRAIL, THE  LP
SMEDMAN, LISA  BLOOD SPORT [SHADOWRUN 29]  P
SMEDMAN, LISA  EXTINCTION [@FORGOTTEN REALMS WAR OF THE SPIDER QUEEN #4]  P
SMEDMAN, LISA  FOREVER DRUG, THE [SHADOWRUN 37]  P
SMEDMAN, LISA  HEIRS OF PROPHECY [@FORGOTTEN REALMS SEMBIA]  P
SMEDMAN, LISA  LUCIFER DECK, THE [SHADOWRUN 23]  P
SMEDMAN, LISA  PSYCHOTROPE [@SHADOWRUN 33]  P
SMEDMAN, LISA  SACRIFICE OF THE WIDOW [@FORGOTTEN REALMS LADY PENITENT #1]  C/P,P
SMEDMAN, LISA  TAILS YOU LOSE [@SHADOWRUN 39]  P
SMEDMAN, LISA  VANITY’S BROOD [@FORGOTTEN REALMS HOUSE OF SERPENTS #3]  P
SMEDMAN, LISA  VENOM’S TASTE [@FORGOTTEN REALMS HOUSE OF SERPENTS #1]  P
SMEDMAN, LISA  VIPER’S KISS [@FORGOTTEN REALMS HOUSE OF SERPENTS #2]  P
SMEDS, DAVE  LAW OF THE JUNGLE [@X-MEN]  P
SMEDS, DAVE  SCHEMES OF DRAGONS, THE [TANAGARAN 2]  P
SMEDS, DAVE  SORCERY WITHIN, THE [TANAGARAN 1]  C/P,P
SMETHURST, WILLIAM  PASIPHAE  H
SMIRNOV, G.  FANTASTIKA 1964 GOD  FH
SMITH, A. C. H.  DARK CRYSTAL, THE  C/P,P
SMITH, A. C. H.  LABYRINTH  P
SMITH, A. D.  CORRIDOR, THE  LP
SMITH, ALISON  COME AWAY HOME  SFWA
SMITH, ALSON J.  IMMORTALITY: THE SCIENTIFIC EVIDENCE  [C/RAN-P]
SMITH, ANDREW  FULL CIRCLE [DOCTOR WHO]  SR-P,C/P
SMITH, ARTHUR D. HOWDEN  GREY MAIDEN  C/P:2,P
SMITH, CANON BASIL A.  SCALLION STONE, THE  H
SMITH, CLARK ASHTON  ABOMINATIONS OF YONDO, THE  SR-P
SMITH, CLARK ASHTON  BLACK DIAMONDS, THE  LP
SMITH, CLARK ASHTON  BOOK OF HYPERBOREA, THE  SR-P
SMITH, CLARK ASHTON  CITY OF THE SINGING FLAME, THE  C/P,P
SMITH, CLARK ASHTON  DOOR TO SATURN, THE  H
SMITH, CLARK ASHTON  DOUBLE SHADOW, THE  LP
SMITH, CLARK ASHTON  END OF THE STORY, THE  H
SMITH, CLARK ASHTON  GENIUS LOCI  H
SMITH, CLARK ASHTON  GENIUS LOCI AND OTHER TALES  SR-P
SMITH, CLARK ASHTON  HYPERBOREA  SR-P,C/P:2
SMITH, CLARK ASHTON  LAST INCANTATION, THE  C/P,P
SMITH, CLARK ASHTON  LOST WORLDS  H
SMITH, CLARK ASHTON  LOST WORLDS I  SR-P
SMITH, CLARK ASHTON  LOST WORLDS II  SR-P
SMITH, CLARK ASHTON  MAKER OF GARGOYLES, THE  LP
SMITH, CLARK ASHTON  MONSTER OF THE PROPHECY, THE  P
SMITH, CLARK ASHTON  OTHER DIMENSIONS  SR-H
SMITH, CLARK ASHTON  OUT OF SPACE AND TIME I  SR-P
SMITH, CLARK ASHTON  OUT OF SPACE AND TIME II  SR-P
SMITH, CLARK ASHTON  PLANETS AND DIMENSIONS  R/REF-H
SMITH, CLARK ASHTON  POSEIDONIS  SR-P,C/P
SMITH, CLARK ASHTON  RED WORLD OF POLARIS  H
SMITH, CLARK ASHTON  SELECTED POEMS  SR-H
SMITH, CLARK ASHTON  STRANGE SHADOWS  SR-H
SMITH, CLARK ASHTON  SWORD OF ZAGAN, THE  LP
SMITH, CLARK ASHTON  TALES OF SCIENCE AND SORCERY  SR-P
SMITH, CLARK ASHTON  TALES OF ZOTHIQUE [ZOTHIQUE]  LP
SMITH, CLARK ASHTON  WHITE SYBIL AND OTHER STORIES, THE  H
SMITH, CLARK ASHTON  XICCARPH  SR-P,C/P:2
SMITH, CLARK ASHTON  ZOTHIQUE [ZOTHIQUE]  SR-P,C/P:2

The MITSFS Pinkdex  September 24, 2007  Page 358
AUTHOR  TITLE  CODE
SMITH, CORDELIA TITCOMB  GREAT SCIENCE FICTION STORIES  PA,C/PA
SMITH, CORDWAINER  BEST OF CORDWAINER SMITH, THE  H:2,P
SMITH, CORDWAINER  INSTRUMENTALITY OF MANKIND, THE [INSTRUMENTALITY OF MANKIND] C/P:2,P
SMITH, CORDWAINER NORSTRILIA [INSTRUMENTALITY OF MANKIND] C/P:2,P

SMITH, CORDWAINER NORSTRILIA (COMPLETE AND CORRECTED) [INSTRUMENTALITY OF MANKIND]

SR-H

SMITH, CORDWAINER  PLANET BUYER, THE [INSTRUMENTALITY OF MANKIND]  C/P:2,P
SMITH, CORDWAINER  QUEST OF THE THREE WORLDS  C/P:2,P
SMITH, CORDWAINER  REDISCOVERY OF MAN, THE [INSTRUMENTALITY OF MANKIND]  SR-H
SMITH, CORDWAINER  SPACE LORDS  C/P:2,P
SMITH, CORDWAINER  STARDREAMER  C/P:2,P
SMITH, CORDWAINER  UNDER OLD EARTH  P
SMITH, CORDWAINER  UNDERPEOPLE, THE [INSTRUMENTALITY OF MANKIND]  C/P,P
SMITH, CORDWAINER  YOU WILL NEVER BE THE SAME  C/P:2,P
SMITH, DALE  ALBINO’S DANCER, THE [TIME HUNTER 9]  LP
SMITH, DALE  HERITAGE [DOCTOR WHO PAST DOCTORS]  P
SMITH, DAVID ALEXANDER  FUTURE BOSTON  HA
SMITH, DAVID ALEXANDER  IN THE CUBE  C/P,P,H
SMITH, DAVID ALEXANDER  MARATHON [MARATHON 1]  P
SMITH, DAVID ALEXANDER  RENDEZVOUS [MARATHON 2]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  AGAINST THE PRINCE OF HELL [@RED SONJA #5]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  DEMON NIGHT [@RED SONJA #2]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  ENDITHOR’S DAUGHTER [@RED SONJA #4]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  FOR THE WITCH OF THE MISTS [BRAN MAK MORN 3]  C/P,P
SMITH, DAVID C.  GHOST ARMY, THE [@ORON #5]  P
SMITH, DAVID C.  MASTER OF EVIL, THE [FALL OF THE FIRST WORLD 1]  P
SMITH, DAVID C.  MOSUTHA’S MAGIC [@ORON #3]  P
SMITH, DAVID C.  ORON [@ORON #1]  C/P,P
SMITH, DAVID C.  PASSING OF THE GODS, THE [FALL OF THE FIRST WORLD 3]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  RING OF IKRIBU, THE [@RED SONJA #1]  C/P,P
SMITH, DAVID C.  SORCERER’S SHADOW, THE [@ORON #2]  P
SMITH, DAVID C.  SORROWING VENGEANCE [FALL OF THE FIRST WORLD 2]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  STAR OF DOOM [@RED SONJA #6]  P
SMITH, DAVID C.  VALLEY OF OGRUM, THE [@ORON #4]  P
SMITH, DAVID C./TIERNEY, RICHARD L.  WHEN HELL LAUGHS [@RED SONJA #3]  P
SMITH, DAVID C.  WITCH OF THE INDIES, THE [BLACK VULMEA]  P
SMITH, DEAN WESLEY/CAREY, DIANE  BELLE TERRE [@STAR TREK NEW EARTH #2]  C/P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN BY THE BOOK [STAR TREK ENTERPRISE]

P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN/ HOFFMAN, NINA KIRIKI ECHOES [@STAR TREK VOYAGER]

C/P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN ESCAPE, THE [STAR TREK VOYAGER]

C/P,P

SMITH, DEAN WESLEY  GRAZER CONSPIRACY, THE [MEN IN BLACK]  P
SMITH, DEAN WESLEY  GREEN SALIVA BLUES, THE [MEN IN BLACK]  P
SMITH, DEAN WESLEY  HARD RAIN, A [STAR TREK TNG]  C/P,P
SMITH, DEAN WESLEY  LAYING THE MUSIC TO REST  C/P,SFWA-P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN LONG NIGHT, THE [STAR TREK DS9]

C/P:2,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN MIST, THE [STAR TREK CAPTAIN’S TABLE 3/STAR TREK DS9]

C/P:2,P

SMITH, DEAN WESLEY  PROPHET’S POWER [UNREAL 2]  P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN RINGS OF TAUTEE, THE [STAR TREK]

C/P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN SHADOW [@STAR TREK VOYAGER SECTION 31]

P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN SOLDIERS OF FEAR, THE [STAR TREK INVASION! 2/STAR TREK TNG]

C/P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN STAR TREK KLINGON [@STAR TREK TNG]

C/P,P

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS [STAR TREK]

HA:2,LPA

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS II [STAR TREK]

C/PA,PA

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS III [STAR TREK]

LPA

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS IV [STAR TREK]

LPA

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS V [STAR TREK]

LPA

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS VI [STAR TREK] LPA

The MITSFS Pinkdex  September 24, 2007  Page 359
AUTHOR  TITLE  CODE

SMITH, DEAN WESLEY/ORDOVER, JOHN J./ KASSIN, ELISA J./BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS VII [STAR TREK]

LPA

SMITH, DEAN WESLEY/KASSIN, ELISA J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS 8 [STAR TREK]

LPA

SMITH, DEAN WESLEY/KASSIN, ELISA J./ BLOCK, PAULA M.

STAR TREK: STRANGE NEW WORLDS 9 [STAR TREK]

LPA

SMITH, DEAN WESLEY/BLOCK, PAULA M.  STAR TREK: STRANGE NEW WORLDS 10 [STAR TREK]  LPA

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN TENTH PLANET, THE [@TENTH PLANET 1]

P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN TENTH PLANET: FINAL ASSAULT, THE [@TENTH PLANET 3]

P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN TENTH PLANET: OBLIVION, THE [@TENTH PLANET 2]

P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN TREATY’S LAW [STAR TREK DAY OF HONOR 4]

C/P,P

SMITH, DEAN WESLEY/ RUSCH, KRISTINE KATHRYN VECTORS [@STAR TREK TNG DOUBLE HELIX 2]

P

SMITH, DEAN WESLEY  WHODUNNIT [@SMALLVILLE]  P
SMITH, E. E.  ALIEN REALMS [LORD TEDRIC 4]  SR-P
SMITH, E. E./GOLDIN, STEPHEN  APPOINTMENT AT BLOODSTAR [FAMILY D’ALEMBERT #5]  C/P:2,@P
SMITH, E. E.  BEST OF E. E. "DOC" SMITH, THE  SR-P,C/P
SMITH, E. E.  CHILDREN OF THE LENS [LENSMEN 6]  C/P,P,H
SMITH, E. E./GOLDIN, STEPHEN  CLOCKWORK TRAITOR, THE [FAMILY D’ALEMBERT #3]  C/P:2,@P
SMITH, E. E./GOLDIN, STEPHEN  ECLIPSING BINARIES [FAMILY D’ALEMBERT #8]  C/P:2,@P
SMITH, E. E.  FIRST LENSMAN [LENSMEN 2]  H:2,P
SMITH, E. E.  GALACTIC PATROL [LENSMEN 3]  C/P,P,H
SMITH, E. E.  GALAXY PRIMES, THE  C/P:2,P
SMITH, E. E./GOLDIN, STEPHEN  GETAWAY WORLD [FAMILY D’ALEMBERT #4]  C/P:2,@P
SMITH, E. E.  GRAY LENSMAN [LENSMEN 4]  H:2,P
SMITH, E. E.  HAVE TRENCHCOAT – WILL TRAVEL  SR-H
SMITH, E. E./GOLDIN, STEPHEN  IMPERIAL STARS [FAMILY D’ALEMBERT #1]  C/P:2,@P
SMITH, E. E.  MASTERS OF SPACE  C/P:2,P
SMITH, E. E.  MASTERS OF THE VORTEX  C/P:2,P
SMITH, E. E./GOLDIN, STEPHEN  OMICRON INVASION, THE [FAMILY D’ALEMBERT #9]  @P
SMITH, E. E./GOLDIN, STEPHEN  PLANET OF TREACHERY [FAMILY D’ALEMBERT #7]  C/P:2,@P
SMITH, E. E./EKLUND, GORDON  PURITY PLOT, THE [FAMILY D’ALEMBERT #6]  C/P,@P
SMITH, E. E./GOLDIN, STEPHEN  REVOLT OF THE GALAXY [FAMILY D’ALEMBERT #10]  C/P:2,@P
SMITH, E. E.  SECOND STAGE LENSMEN [LENSMEN 5]  C/P,P,H
SMITH, E. E.  SKYLARK DUQUESNE [SKYLARK 4]  C/P:2,P
SMITH, E. E.  SKYLARK OF SPACE, THE [SKYLARK 1]  C/P:2,P,H
SMITH, E. E.  SKYLARK OF VALERON [SKYLARK 3]  C/P,P,H:2
SMITH, E. E.  SKYLARK THREE [SKYLARK 2]  C/P:2,P,H
SMITH, E. E.  SPACEHOUNDS OF IPC  C/P,P,H
SMITH, E. E./GOLDIN, STEPHEN  STRANGLER’S MOON [FAMILY D’ALEMBERT #2]  C/P:2,@P
SMITH, E. E.  SUBSPACE ENCOUNTER  C/P:2,P
SMITH, E. E.  SUBSPACE EXPLORERS  C/P,P,H
SMITH, E. E.  TRIPLANETARY [LENSMEN 1]  H:2,P
SMITH, E. E.  VORTEX BLASTER, THE  H:2
SMITH, ELLA  TRANSFERENCE, THE  P
SMITH, EVELYN E.  COPY SHOP, THE  SFWA
SMITH, EVELYN E.  PERFECT PLANET, THE  C/P,P,H
SMITH, EVELYN E.  UNPOPULAR PLANET  C/P:2,P
SMITH, GARRETT  BETWEEN WORLDS  SR-LP
SMITH, GEORGE H.  COMING OF THE RATS, THE  SR-P
SMITH, GEORGE H.  DOOMSDAY WING  C/P:2,P
SMITH, GEORGE H.  FOUR DAY WEEKEND, THE  C/P:2,P
SMITH, GEORGE H.  ISLAND SNATCHERS, THE [ARWIN 5]  C/P,P
SMITH, GEORGE H.  KAR KABALLA [ARWIN 3]  D42900
SMITH, GEORGE H.  SECOND WAR OF THE WORLDS, THE [ARWIN 4]  P
SMITH, GEORGE H.  UNENDING NIGHT, THE  C/P:2,P
SMITH, GEORGE H.  WITCH QUEEN OF LOCHLANN [ARWIN 2]  C/P,P
SMITH, GEORGE O.  BRAIN MACHINE, THE  C/P:2,P
SMITH, GEORGE O.  COMPLETE VENUS EQUILATERAL, THE  C/P:2,P
SMITH, GEORGE O.  FIRE IN THE HEAVENS  DD375
SMITH, GEORGE O.  FOURTH ’R’, THE  C/P:2,P
SMITH, GEORGE O.  HELLFLOWER  C/P,P
SMITH, GEORGE O.  HIGHWAYS IN HIDING  H,P
SMITH, GEORGE O.  LOST IN SPACE  DD431
SMITH, GEORGE O.  OPERATION INTERSTELLAR  SR-P
SMITH, GEORGE O.  PATH OF UNREASON, THE  C/P,P,H
SMITH, GEORGE O.  PATTERN FOR CONQUEST  H
SMITH, GEORGE O.  SPACE PLAGUE  C/P:2,P
SMITH, GEORGE O.  TROUBLED STAR  SR-P,C/P
SMITH, GEORGE O.  VENUS EQUILATERAL  C/P,SR-H,P,H

The MITSFS Pinkdex  September 24, 2007  Page 360
AUTHOR  TITLE  CODE
SMITH, GEORGE O.  WORLDS OF GEORGE O.  C/P,P
SMITH, GUS  FEATHER & BONE  LP
SMITH, GUY N.  BATS OUT OF HELL  P
SMITH, GUY N.  CRABS’ MOON [CRABS 2]  P
SMITH, GUY N.  CRABS ON THE RAMPAGE [CRABS 3]  C/P,P
SMITH, GUY N.  DEATHBELL  P
SMITH, GUY N.  ENTOMBED  P
SMITH, GUY N.  KILLER CRABS  P
SMITH, GUY N.  NIGHT OF THE CRABS [CRABS]  P
SMITH, GUY N.  ORIGIN OF THE CRABS, THE [CRABS 1]  P
SMITH, H. ALLEN  AGE OF THE TAIL, THE  H:2,P
SMITH, HOMER W.  KAMONGO  H
SMITH, JAMES V. JR.  BEASTMAKER  P
SMITH, JULIE DEAN  CALL OF MADNESS [CAITHAN CRUSADE 1]  C/P,P
SMITH, JULIE DEAN  MISSION OF MAGIC [CAITHAN CRUSADE 2]  P
SMITH, JULIE DEAN  SAGE OF SARE [CAITHAN CRUSADE 3]  C/P,P
SMITH, JULIE DEAN  WIZARD KING, THE [CAITHAN CRUSADE 4]  P
SMITH, KRISTINE  CODE OF CONDUCT [JANI KILIAN 1]  C/P,P,H
SMITH, KRISTINE  CONTACT IMMINENT [JANI KILIAN 4]  C/P,P
SMITH, KRISTINE  CONTACT IMMINENT (PROOF) [JANI KILIAN 4]  SR-P
SMITH, KRISTINE  ENDGAME (PROOF) [JANI KILIAN 5]  SR-P
SMITH, KRISTINE  LAW OF SURVIVAL [JANI KILIAN 3]  C/P:2,P
SMITH, KRISTINE  RULES OF CONFLICT [JANI KILIAN 2]  C/P,P
SMITH, L. J.  HEART OF VALOR [NIGHT OF THE SOLSTICE 2]  H
SMITH, L. J.  INITIATION, THE [SECRET CIRCLE 1]  P
SMITH, L. NEIL  AMERICAN ZONE, THE [NORTH AMERICAN CONFEDERACY]  LP
SMITH, L. NEIL  BRETTA MARTYN [HENRY MARTYN 2]  H,P
SMITH, L. NEIL  BRIGHTSUIT MACBEAR  C/P,P
SMITH, L. NEIL  CONTACT AND COMMUNE [ERIS 1]  P
SMITH, L. NEIL  CONVERSE AND CONFLICT [ERIS 2]  P
SMITH, L. NEIL  CRYSTAL EMPIRE, THE  H,P
SMITH, L. NEIL  FORGE OF THE ELDERS [ERIS 1,2]  P
SMITH, L. NEIL  GALLATIN DIVERGENCE, THE [NORTH AMERICAN CONFEDERACY]  C/P,P
SMITH, L. NEIL  HENRY MARTYN [HENRY MARTYN 1]  C/P,P
SMITH, L. NEIL  LANDO CALRISSIAN ADVENTURES, THE [STAR WARS]  P
SMITH, L. NEIL  LANDO CALRISSIAN AND THE FLAMEWIND OF OSEON [STAR WARS]  P
SMITH, L. NEIL  LANDO CALRISSIAN AND THE MINDHARP OF SHARU [STAR WARS]  C/P,P
SMITH, L. NEIL  LANDO CALRISSIAN AND THE STARCAVE OF THONBOKA [STAR WARS]  P
SMITH, L. NEIL  NAGASAKI VECTOR, THE [NORTH AMERICAN CONFEDERACY]  C/P:2,P
SMITH, L. NEIL  PALLAS  P
SMITH, L. NEIL  PROBABILITY BROACH, THE [NORTH AMERICAN CONFEDERACY]  C/P:2,P

SMITH, L. NEIL PROBABILITY BROACH, THE (ORIGINAL) [NORTH AMERICAN CONFEDERACY]

H,LP

SMITH, L. NEIL  TAFLAK LYSANDRA [NORTH AMERICAN CONFEDERACY]  P
SMITH, L. NEIL  THEIR MAJESTIES’ BUCKETEERS  P
SMITH, L. NEIL  TOM PAINE MARU [NORTH AMERICAN CONFEDERACY]  C/P:2,P
SMITH, L. NEIL  VENUS BELT, THE [NORTH AMERICAN CONFEDERACY]  C/P,P
SMITH, L. NEIL  WARDOVE, THE  C/P,P
SMITH, MARK  MOON LAMP, THE  P
SMITH, MARK/SMITH, JULIA  SHADOW-MAZE  P
SMITH, MARTIN CRUZ  INDIANS WON, THE  P
SMITH, MARTIN CRUZ  NIGHTWING  C/P:2,P
SMITH, MATTHEW  FINAL CUT, THE [@JUDGE DREDD]  P
SMITH, MICHAEL MARSHALL  MORE TOMORROW & OTHER STORIES  H
SMITH, MICHAEL MARSHALL  ONE OF US  H,LP
SMITH, MICHAEL MARSHALL  ONLY FORWARD  C/P,P,H
SMITH, MICHAEL MARSHALL  SPARES  P
SMITH, MICHAEL MARSHALL  VACCINATOR, THE  MB002
SMITH, MICHAEL MARSHALL  WHAT YOU MAKE IT  LP
SMITH, MITCHELL  KINGDOM RIVER [SNOWFALL 2]  P
SMITH, MITCHELL  MOONRISE [SNOWFALL 3]  P
SMITH, MITCHELL  SNOWFALL [SNOWFALL 1]  P
SMITH, NYX  FADE TO BLACK [@SHADOWRUN 13]  P
SMITH, NYX  STEEL RAIN [@SHADOWRUN 24]  P
SMITH, NYX  STRIPER ASSASSIN [SHADOWRUN 11]  C/P,P
SMITH, NYX  WHO HUNTS THE HUNTER [@SHADOWRUN 16]  C/P,P
SMITH, PERRY MICHAEL  LAST RITES  H:2

SMITH, R. DIXON LOST IN THE RENTHARPIAN HILLS:

SPANNING THE DECADES WITH CARL JACOBI

R/REF-P

SMITH-READY, JERI  EYES OF CROW  LP
SMITH, ROBERT ARTHUR  PREY, THE  P
SMITH, SHERWOOD  AUGUR’S TEACHER [EARTH FINAL CONFLICT]  H,LP
SMITH, SHERWOOD  BEYOND THE LAST STAR  LPA
SMITH, SHERWOOD  CROWN DUEL  P

The MITSFS Pinkdex  September 24, 2007  Page 361
AUTHOR  TITLE  CODE
SMITH, SHERWOOD  INDA  H
SMITH, SHERWOOD  PARADISE DRIFT [ANDROMEDA (GENE RODDENBERRY’S)]  H
SMITH, SHERWOOD/TROWBRIDGE, DAVE  PHOENIX IN FLIGHT, THE [EXORDIUM 1]  P
SMITH, SHERWOOD/TROWBRIDGE, DAVE  PRISON UNSOUGHT, A [EXORDIUM 3]  P
SMITH, SHERWOOD/TROWBRIDGE, DAVE  RIFTER’S COVENANT, THE [EXORDIUM 4]  P
SMITH, SHERWOOD/TROWBRIDGE, DAVE  RULER OF NAUGHT [EXORDIUM 2]  P
SMITH, SHERWOOD/TROWBRIDGE, DAVE  THRONES OF KRONOS, THE [EXORDIUM 5]  P
SMITH, STEPHANIE A.  BOY WHO WAS THROWN AWAY, THE [SNOW-EYES 2]  C/P,P
SMITH, STEPHANIE A.  OTHER NATURE  LP
SMITH, STEPHANIE A.  SNOW-EYES [SNOW-EYES 1]  C/P,P,H
SMITH, SUSAN  CONFESSIONS OF A TEENAGE FROG [@SAMANTHA SLADE #2]  P
SMITH, SUSAN  OUR FRIEND: PUBLIC NUISANCE #1 [@SAMANTHA SLADE #3]  P
SMITH, SUSAN  SAMANTHA SLADE: MONSTER-SITTER [@SAMANTHA SLADE #1]  P
SMITH, SUSAN  TERRORS OF ROCK AND ROLL, THE [@SAMANTHA SLADE #4]  P
SMITH, TERRENCE LORE  YOURS TRULY, FROM HELL  P
SMITH, THORNE  BISHOP’S JAEGERS, THE  H
SMITH, THORNE  DID SHE FALL?  H
SMITH, THORNE  GLORIOUS POOL, THE  P
SMITH, THORNE  NIGHT LIFE OF THE GODS, THE  C/P,P,H
SMITH, THORNE  PASSIONATE WITCH, THE  C/P,P
SMITH, THORNE  RAIN IN THE DOORWAY  C/P:2,P
SMITH, THORNE  SKIN AND BONES  H:2,P
SMITH, THORNE  STRAY LAMB, THE  C/P,P,H
SMITH, THORNE  THORNE SMITH 3-DECKER, THE  H
SMITH, THORNE  TOPPER  C/P,P,H
SMITH, THORNE  TOPPER TAKES A TRIP  C/P,P,H
SMITH, THORNE  TURNABOUT  C/P:2,P
SMITH, WILBUR  SUNBIRD, THE  P
SMITH, WORTH  MIRACLE OF THE AGES: THE GREAT PYRAMID OF GIZEH  [C/RAN-P]
SMYTH, CLIFFORD  GILDED MAN, THE  H
SMYTHE  VERY SNEAKY, ANDY CAPP  [C/RAN-P]
SNELL, EDMUND  Z RAY, THE  H
SNELL, ROY J.  CRIMSON FLASH, THE  H
SNIEGOSKI, THOMAS E.  SOUL TRADE [@ANGEL]  P
SNODGRASS, MELINDA M.  CIRCUIT [CIRCUIT 1]  C/P:2,P
SNODGRASS, MELINDA M.  CIRCUIT BREAKER [CIRCUIT 2]  P,SFWA-P
SNODGRASS, MELINDA M.  DOUBLE SOLITAIRE [@WILD CARDS #10]  C/P:2,P
SNODGRASS, MELINDA M.  FINAL CIRCUIT [CIRCUIT 3]  P
SNODGRASS, MELINDA M.  QUEEN’S GAMBIT DECLINED  P,SFWA-P
SNODGRASS, MELINDA M.  TEARS OF THE SINGERS, THE [STAR TREK]  C/P:3,P
SNOW, JACK  DARK MUSIC AND OTHER SPECTRAL TALES  SR-H
SNYDER, CECIL  HAWKS OF ARCTURUS, THE  C/P,P
SNYDER, GENE  MIND WAR  P
SNYDER, GENE  OGDEN ENIGMA, THE  C/P,P
SNYDER, GUY  TESTAMENT XXI  P
SNYDER, MARIA V.  MAGIC STUDY [YELINA 2]  H
SNYDER, MARIA V.  POISON STUDY [YELINA 1]  H
SNYDER, MIDORI  BELDAN’S FIRE [QUEEN’S QUARTER 3]  C/P:2,P
SNYDER, MIDORI  FLIGHT OF MICHAEL MCBRIDE, THE  C/P,P
SNYDER, MIDORI  INNAMORATI, THE  H
SNYDER, MIDORI  NEW MOON [QUEEN’S QUARTER 1]  C/P:2,P
SNYDER, MIDORI  SADAR’S KEEP [QUEEN’S QUARTER 2]  C/P:2,SFWA-P,P
SNYDER, MIDORI  SOULSTRING  C/P:2,P
SNYDER, ZILPHA KEATLEY  AND ALL BETWEEN [GREEN-SKY 2]  C/P,SFWA-P,P
SNYDER, ZILPHA KEATLEY  BELOW THE ROOT [GREEN-SKY 1]  P,SFWA-P
SNYDER, ZILPHA KEATLEY  BLACK AND BLUE MAGIC  LP
SNYDER, ZILPHA KEATLEY  CHANGING MAZE, THE  SFWA
SNYDER, ZILPHA KEATLEY  UNTIL THE CELEBRATION [GREEN-SKY 3]  P
SOBCHACK, VIVIAN CAROL  LIMITS OF INFINITY: THE AMERICAN SCIENCE FICTION FILM, THE  R/REF-H
SOBEL, ROBERT  FOR WANT OF A NAIL  H
SOBOL, DONALD J.  AMAZING POWER OF ASHUR FINE, THE  H
SOESBEE, REE  CITY OF FORTUNE [@DRAGONLANCE NEW ADVENTURES ELIDOR #3]  LP
SOESBEE, REE  CRANE, THE [FIVE RINGS CLAN WAR 3]  P
SOESBEE, REE  CROWN OF THIEVES [@DRAGONLANCE NEW ADVENTURES ELIDOR #1]  LP
SOESBEE, REE  CRYSTAL CHALICE, THE [@DRAGONLANCE NEW ADVENTURES ELIDOR #2]  LP
SOESBEE, REE  DRAGON, THE [FIVE RINGS CLAN WAR 6]  P

SOESBEE, REE DRAGON SWORD [@DRAGONLANCE NEW ADVENTURES #5/

DRAGONLANCE DRAGON QUARTET 1]

LP

SOESBEE, REE PILLAR OF FLAME, THE

[@DRAGONLANCE NEW ADVENTURES ELEMENTS #1]

LP

SOESBEE, REE  RIDDLE IN THE STONE [@KNIGHTS OF THE SILVER DRAGON #2]  LP
SOESBEE, REE  WIND OF HONOR [FIVE RINGS FOUR WINDS #1]  P
SOESBEE, REE  WIND OF TRUTH [@FIVE RINGS FOUR WINDS #4]  P
SOHL, JERRY  ALTERED EGO, THE  H:2,P

AUTHOR  TITLE  CODE
SOHL, JERRY  ANOMALY, THE  C/P,P
SOHL, JERRY  COSTIGAN’S NEEDLE  H:2,P
SOHL, JERRY  HAPLOIDS, THE  SR-P,H
SOHL, JERRY  I, ALEPPO  SR-P
SOHL, JERRY  MARS MONOPOLY, THE  DD162
SOHL, JERRY  NIGHT SLAVES  C/P:2,P
SOHL, JERRY  ODIOUS ONES, THE  H
SOHL, JERRY  ONE AGAINST HERCULUM  DD381
SOHL, JERRY  POINT ULTIMATE  H:2,P
SOHL, JERRY  TIME DISSOLVER, THE  C/P:2,P
SOHL, JERRY  TRANSCENDENT MAN, THE  C/P:2,P
SOHL, JERRY  UNO CONTRO HERCULUM  FP
SOLOTAROFF, THEODORE  NEW AMERICAN REVIEW 15  PA
SOMERS, BART  ABANDON GALAXY!  C/P:2,P
SOMERS, BART  BEYOND THE BLACK ENIGMA  C/P:2,P
SOMERS, JAMES  HEIR TO THE KING [@SOONE 1]  LP
SOMTOW, S. P.  AQUILA AND THE IRON HORSE [@AQUILIAD #2]  P
SOMTOW, S. P.  AQUILA AND THE SPHINX [@AQUILIAD #3]  P
SOMTOW, S. P.  AQUILA IN THE NEW WORLD [@AQUILIAD #1]  C/P,P
SOMTOW, S. P.  DO COMETS DREAM? [@STAR TREK TNG]  P
SOMTOW, S. P.  FOREST OF THE NIGHT [RIVERRUN 2]  P
SOMTOW, S. P.  FORGETTING PLACES  H
SOMTOW, S. P.  RIVERRUN [RIVERRUN 1]  P
SOMTOW, S. P.  RIVERRUN TRILOGY [RIVERRUN 1,2,3]  LP
SOMTOW, S. P.  SHATTERED HORSE, THE  H,SFWA-P,P
SOMTOW, S. P.  STARSHIP AND HAIKU  C/P:2,P
SOMTOW, S. P.  TEMPLE OF NIGHT [@CROW]  P
SOMTOW, S. P.  ULTIMATE MALLWORLD, THE [MALLWORLD]  LP
SOMTOW, S. P.  VAMPIRE JUNCTION [TIMOTHY VALENTINE 1]  C/P:2,P
SOMTOW, S. P.  VANITAS [TIMOTHY VALENTINE 3]  H
SOPHOCLES  OEDIPUS CYCLE, THE  LP
SOREL, EDWARD  MOON MISSING  LP
SOROKA, CYNTHIA  BEGINNING, THE [@DARK CHRONICLES #1]  P
SOSNA, SHARON  IN THIS AGE OF STONE  C/P,P
SOUKUP, MARTHA  ROSEMARY’S BRAIN  P
SOULBAN, LUCIEN  ALIEN SEA, THE [@DRAGONLANCE CHAMPIONS]  P
SOULBAN, LUCIEN  FLESHWORKS [@NECROMUNDA]  P
SOULE, GEORGE  SHAPE OF TOMORROW, THE  [C/RAN-P:2]
SOUTHWELL, SAMUEL  IF ALL THE REBELS DIE  P
SOUTO, MARCIAL  LLEGAN LOS DRAGONES  FP
SOUZA, MARCIO  END OF THE THIRD WORLD, THE [LOST WORLD 2]  H,LP
SOUZA, MARCIO  ORDER OF THE DAY, THE  P
SOWDEN, LEWIS  TOMORROW’S COMET  H
SPARGER, REX  BARGAIN, THE [DARK FORCES 5]  C/P,P
SPARGER, REX  DOLL, THE [DARK FORCES 3]  P
SPARKS, CAT  AGOG! FANTASTIC FICTION  LPA
SPARKS, CAT  AGOG! SMASHING STORIES  LPA
SPARKS, CAT  AGOG! TERRIFIC TALES  HA
SPARKS, KERRELYN  VAMPS AND THE CITY  P
SPEARING, JUDITH  GHOSTS WHO WENT TO SCHOOL  LP
SPECTOR, CAROLINE  WORLDS WITHOUT END [@SHADOWRUN 18/IMMORTALS 3]  C/P,P
SPECTOR, ROBERT DONALD  SEVEN MASTERPIECES OF GOTHIC HORROR  PA
SPECULATOR  WHOLE (ALMOST) PROZINE CATALOG  SRL/IND
SPEER, JACK  FANCESTRAL VOICES  R/REF-H
SPEER, JACK  FANTASY FICTION DECIMAL CLASSIFICATION  R/IND-F
SPENCER, DAVID  PASSING FANCY [ALIEN NATION 6]  C/P,P
SPENCER, JOHN  ELECTRONIC LULLABY MEAT MARKET, THE  P
SPENCER, JOHN  LIMBO OF THE LOST  [C/RAN-P]
SPENCER, JOHN/EVANS, HILARY  PHENOMENON: FORTY YEARS OF FLYING SAUCERS  [C/RAN-P]
SPENCER, JOHN R.  ICEQUAKE  P
SPENCER, SCOTT  LAST NIGHT AT THE BRAIN THIEVES BALL  C/P,P
SPENCER, WEN  ALIEN TASTE [UKIAH OREGON 1]  C/P,P
SPENCER, WEN  BITTER WATERS [UKIAH OREGON 3]  C/P,P
SPENCER, WEN  BROTHER’S PRICE, A  P
SPENCER, WEN  DOG WARRIOR [UKIAH OREGON 4]  C/P:2,P
SPENCER, WEN  TAINTED TRAIL [UKIAH OREGON 2]  C/P,P
SPENCER, WEN  TINKER [TINKER 1]  P
SPENCER, WEN  WOLF WHO RULES [TINKER 2]  H:2
SPENCER, WILLIAM BROWNING  IRRATIONAL FEARS  H
SPENCER, WILLIAM BROWNING  OCEAN IN ALL ITS DEVICES, THE  SR-H
SPENCER, WILLIAM BROWNING  RESUME WITH MONSTERS  P
SPENCER, WILLIAM BROWNING  RETURN OF COUNT ELECTRIC, THE  P
SPENCER, WILLIAM BROWNING  ZOD WALLOP  P
SPENSER, EDMUND  FAERIE QUEENE I, THE  SR-P

AUTHOR  TITLE  CODE
SPENSER, EDMUND  FAERIE QUEENE II, THE  SR-P
SPERRY, RALPH A.  STATUS QUOTIENT: THE CARRIER  C/P:2,P
SPICER, DOROTHY  CRYSTAL BALL, THE  P
SPIEGL, WALTER  SCIENCE FICTION STORIES 4  FP
SPIEGL, WALTER  SCIENCE FICTION STORIES 23  FP
SPIELBERG, STEVEN  CLOSE ENCOUNTERS OF THE THIRD KIND  H:2,P
SPIELBERG, STEVEN  CLOSE ENCOUNTERS OF THE THIRD KIND: THE SPECIAL EDITION  C/P,P
SPIELMAN, BRENDA GATES  STALKER LORD  P
SPINE-TINGLING PRESS  AUDIO SAMPLER 1 (CASSETTE)  SR-MM
SPINRAD, NORMAN  AGENT OF CHAOS  C/P,P,H
SPINRAD, NORMAN  BUG JACK BARRON  C/P:2,P
SPINRAD, NORMAN  CHILD OF FORTUNE  H:2,P
SPINRAD, NORMAN  DEUS X  P
SPINRAD, NORMAN  GREENHOUSE SUMMER  H
SPINRAD, NORMAN  HE WALKED AMONG US  LP
SPINRAD, NORMAN  IRON DREAM, THE  C/P:2,P
SPINRAD, NORMAN  JOURNALS OF THE PLAGUE YEARS  LP
SPINRAD, NORMAN  LA CIVILTA DEI SOLARI  FP
SPINRAD, NORMAN  LAST HURRAH OF THE GOLDEN HORDE, THE  H:2,P
SPINRAD, NORMAN  LITTLE HEROES  H:2,P
SPINRAD, NORMAN  MEN IN THE JUNGLE, THE  C/P:2,P
SPINRAD, NORMAN  MIND GAME, THE  P
SPINRAD, NORMAN  MODERN SCIENCE FICTION  C/PA,PA
SPINRAD, NORMAN  NEW TOMORROWS, THE  PA
SPINRAD, NORMAN  NO DIRECTION HOME  C/P:2,P
SPINRAD, NORMAN  OTHER AMERICAS  C/P,P
SPINRAD, NORMAN  PASSING THROUGH THE FLAME  P
SPINRAD, NORMAN  PICTURES AT 11  C/P,P
SPINRAD, NORMAN  RIDING THE TORCH  TD23,BS1
SPINRAD, NORMAN  RUSSIAN SPRING  C/P,P,H
SPINRAD, NORMAN  SOLARIANS, THE  C/P:2,P
SPINRAD, NORMAN  SONGS FROM THE STARS  H:2,P
SPINRAD, NORMAN  SONGS FROM THE STARS (PROOF)  SR-LP
SPINRAD, NORMAN  STAR-SPANGLED FUTURE, THE  C/P:2,P
SPINRAD, NORMAN  VOID CAPTAIN’S TALE, THE  H:2,P
SPINRAD, NORMAN  WORLD BETWEEN, A  C/P:2,P
SPOFFORD, HARRIET PRESCOTT  AMBER GODS, THE  LP
SPOOR, RYK E.  DIGITAL KNIGHT  C/P,P
SPRIGGS, ROBIN  WONDROUS STRANGE  LP
SPRINGER, NANCY  APOCALYPSE  C/P:2,SFWA-P,P
SPRINGER, NANCY  BLACK BEAST, THE [BOOKS OF THE ISLE 4]  C/P,P
SPRINGER, NANCY  BOOK OF SUNS, THE  C/P:2,P
SPRINGER, NANCY  CHAINS OF GOLD  H:2,SFWA,P
SPRINGER, NANCY  CHANCE AND OTHER GESTURES OF THE HAND OF FATE  C/P:2,SFWA-P,P
SPRINGER, NANCY  DAMNBANNA  SR-H
SPRINGER, NANCY  FAIR PERIL  C/P,P
SPRINGER, NANCY  GODBOND [SEA KING 3]  P,SFWA-P
SPRINGER, NANCY  GOLDEN SWAN, THE [BOOKS OF THE ISLE 5]  C/P:2,P
SPRINGER, NANCY  HEX WITCH OF SELDOM, THE  C/P,P
SPRINGER, NANCY  I AM MORDRED  P
SPRINGER, NANCY  LARQUE ON THE WING  P
SPRINGER, NANCY  MADBOND [SEA KING 1]  P,SFWA-P
SPRINGER, NANCY  METAL ANGEL  P,SFWA-P
SPRINGER, NANCY  MINDBOND [SEA KING 2]  P
SPRINGER, NANCY  PROM NIGHT  PA,C/PA
SPRINGER, NANCY  SABLE MOON, THE [BOOKS OF THE ISLE 3]  C/P:2,P
SPRINGER, NANCY  SILVER SUN, THE [BOOKS OF THE ISLE 2]  C/P:2,P
SPRINGER, NANCY  WHITE HART, THE [BOOKS OF THE ISLE 1]  H:2,P
SPRINGER, NANCY  WINGS OF FLAME  C/P,P,H
SPRUILL, STEVEN G.  GENESIS SHIELD, THE  C/P,P
SPRUILL, STEVEN G.  HELLSTONE  P
SPRUILL, STEVEN G.  IMPERATOR PLOT, THE  P,SFWA,SFWA-P
SPRUILL, STEVEN G.  JANUS EQUATION, THE  BS4
SPRUILL, STEVEN G.  KEEPERS OF THE GATE  C/P,SFWA,P,H
SPRUILL, STEVEN G.  PARADOX PLANET, THE  C/P,P,H
SPRUILL, STEVEN G.  PSYCHOPATH PLAGUE, THE  C/P,SFWA,SFWA-P,P
SPURRIER, SIMON  FIRE WARRIOR [@WARHAMMER 40K]  P
SPURRIER, SIMON  LORD OF THE NIGHT [@WARHAMMER 40K]  P
ST. JUDE/SIRIUS, R. U./NAGEL, BART  CYBERPUNK HANDBOOK  [C/RAN-P]
STABENOW, DANA  HANDFUL OF STARS, A  P
STABENOW, DANA  POWERS OF DETECTION  LPA
STABENOW, DANA  RED PLANET RUN  P
STABENOW, DANA  SECOND STAR  P
STABLEFORD, BRIAN M.  ALGEBRAIC FANTASIES AND REALISTIC ROMANCES  R/REF-P
STABLEFORD, BRIAN M.  ANGEL OF PAIN, THE [DAVID LYDYARD 2]  H

AUTHOR  TITLE  CODE
STABLEFORD, BRIAN M.  ARCHITECTS OF EMORTALITY [EMORTALITY]  H
STABLEFORD, BRIAN M.  BALANCE OF POWER [DAEDALUS 5]  C/P,P
STABLEFORD, BRIAN M.  BLIND WORM, THE  D06707
STABLEFORD, BRIAN M.  CARNIVAL OF DESTRUCTION, THE [DAVID LYDYARD 3]  P
STABLEFORD, BRIAN M.  CASSANDRA COMPLEX, THE [EMORTALITY]  H
STABLEFORD, BRIAN M.  CASTAWAYS OF TANAGAR, THE  C/P,P
STABLEFORD, BRIAN M.  CITY OF THE SUN [DAEDALUS 4]  C/P,P
STABLEFORD, BRIAN M.  CLASH OF SYMBOLS: THE TRIUMPH OF JAMES BLISH, A  R/REF-P
STABLEFORD, BRIAN M.  CRADLE OF THE SUN  D12140
STABLEFORD, BRIAN M.  CRITICAL THRESHOLD [DAEDALUS 2]  C/P:2,P
STABLEFORD, BRIAN M.  DARK ARARAT [EMORTALITY]  H
STABLEFORD, BRIAN M.  DAY OF WRATH [DIES IRAE 3]  C/P:2,P
STABLEFORD, BRIAN M.  DAYS OF GLORY, THE [DIES IRAE 1]  C/P:2,P
STABLEFORD, BRIAN M.  EMPIRE OF FEAR, THE  C/P,P
STABLEFORD, BRIAN M.  FENRIS DEVICE, THE [GRAINGER 5]  C/P,P
STABLEFORD, BRIAN M.  FIREFLY  LP
STABLEFORD, BRIAN M.  FLORIANS, THE [DAEDALUS 1]  C/P,P
STABLEFORD, BRIAN M.  FOUNTAINS OF YOUTH, THE [EMORTALITY]  H,LP
STABLEFORD, BRIAN M.  GATES OF EDEN, THE  C/P,P
STABLEFORD, BRIAN M.  HALCYON DRIFT, THE [GRAINGER 1]  C/P:2,P
STABLEFORD, BRIAN M.  IN THE KINGDOM OF THE BEASTS [DIES IRAE 2]  C/P:2,P
STABLEFORD, BRIAN M.  INHERIT THE EARTH [EMORTALITY]  H
STABLEFORD, BRIAN M.  INVADERS FROM THE CENTRE [ASGARD 2]  P
STABLEFORD, BRIAN M.  JOURNEY TO THE CENTER [ASGARD 1]  H,P
STABLEFORD, BRIAN M.  LAST DAYS OF THE EDGE OF THE WORLD, THE  P
STABLEFORD, BRIAN M.  MAN IN A CAGE  H
STABLEFORD, BRIAN M.  MIND-RIDERS, THE  P
STABLEFORD, BRIAN M.  OMEGA EXPEDITION, THE [EMORTALITY]  H
STABLEFORD, BRIAN M.  OPTIMAN  C/P:2,P
STABLEFORD, BRIAN M.  PARADISE GAME, THE [GRAINGER 4]  C/P,P
STABLEFORD, BRIAN M.  PARADOX OF THE SETS, THE [DAEDALUS 6]  C/P:2,P
STABLEFORD, BRIAN M.  PROMISED LAND [GRAINGER 3]  H:2,P
STABLEFORD, BRIAN M.  REALMS OF TARTARUS, THE  C/P,P
STABLEFORD, BRIAN M.  RHAPSODY IN BLACK [GRAINGER 2]  C/P:2,P
STABLEFORD, BRIAN M.  SEXUAL CHEMISTRY  P
STABLEFORD, BRIAN M.  SWAN SONG [GRAINGER 6]  P
STABLEFORD, BRIAN M.  SWAN SONGS [GRAINGER 1,2,3,4,5,6]  LP
STABLEFORD, BRIAN M./LANGFORD, DAVID THIRD MILLENNIUM, THE  R/REF-H
STABLEFORD, BRIAN M.  TO CHALLENGE CHAOS [CHALLENGE CHAOS 1]  C/P:2,P
STABLEFORD, BRIAN M.  WALKING SHADOW, THE  P
STABLEFORD, BRIAN M.  WEREWOLVES OF LONDON, THE [DAVID LYDYARD 1]  C/P,P,H
STABLEFORD, BRIAN M.  WILDEBLOOD’S EMPIRE [DAEDALUS 3]  C/P,P
STACKPOLE, MICHAEL A.  ASSUMPTION OF RISK [@BATTLETECH 12]  C/P,P
STACKPOLE, MICHAEL A.  BACTRA WAR, THE [@STAR WARS X-WING 4]  P
STACKPOLE, MICHAEL A.  BLOOD LEGACY [@BATTLETECH #21/BLOOD OF KERENSKY 2]  P
STACKPOLE, MICHAEL A.  BRED FOR WAR [@BATTLETECH 16]  P
STACKPOLE, MICHAEL A.  CARTOMANCY [AGE OF DISCOVERY 2]  LP
STACKPOLE, MICHAEL A.  DARK GLORY WAR, THE [DRAGONCROWN WAR]  P
STACKPOLE, MICHAEL A.  DARK TIDE [@STAR WARS NEW JEDI ORDER DARK TIDE 1,2]  H
STACKPOLE, MICHAEL A.  DEMENTIA [@MUTANT CHRONICLES APOSTLE OF INSANITY 3]  P
STACKPOLE, MICHAEL A.  ENEMY REBORN, AN [REALMS OF CHAOS 2]  P
STACKPOLE, MICHAEL A.  EVIL ASCENDING [DARK CONSPIRACY 2]  P
STACKPOLE, MICHAEL A.  EVIL TRIUMPHANT [DARK CONSPIRACY 3]  P
STACKPOLE, MICHAEL A.  EYES OF SILVER  C/P,P
STACKPOLE, MICHAEL A.  FORTRESS DRACONIS [DRAGONCROWN WAR 1]  LP
STACKPOLE, MICHAEL A.  GATHERING EVIL, A [DARK CONSPIRACY 1]  P
STACKPOLE, MICHAEL A.  GHOST WAR [@MECHWARRIOR DARK AGE 1/BATTLETECH]  P
STACKPOLE, MICHAEL A.  GRAND CRUSADE, THE [DRAGONCROWN WAR 3]  LP
STACKPOLE, MICHAEL A.  GRAND CRUSADE, THE (PROOF) [DRAGONCROWN WAR 3]  SR-LP
STACKPOLE, MICHAEL A.  GRAVE COVENANT [@BATTLETECH #34/TWILIGHT OF THE CLANS 2]  C/P,P
STACKPOLE, MICHAEL A.  HERO BORN, A [REALMS OF CHAOS 1]  P
STACKPOLE, MICHAEL A.  I, JEDI [@STAR WARS]  C/P,P,H
STACKPOLE, MICHAEL A.  ISARD’S REVENGE [@STAR WARS X-WING 8]  C/P,P
STACKPOLE, MICHAEL A.  KRYTOS TRAP, THE [@STAR WARS X-WING 3]  C/P,P
STACKPOLE, MICHAEL A.  LETHAL HERITAGE [@BATTLETECH #20/BLOOD OF KERENSKY 1]  P,SFWA-P
STACKPOLE, MICHAEL A.  LOST DESTINY [@BATTLETECH #22/BLOOD OF KERENSKY 3]  P
STACKPOLE, MICHAEL A.  MALICIOUS INTENT [@BATTLETECH 24]  P
STACKPOLE, MICHAEL A.  MASTERS OF WAR [@MECHWARRIOR DARK AGE]  P
STACKPOLE, MICHAEL A.  NATURAL SELECTION [@BATTLETECH #5]  C/P,P
STACKPOLE, MICHAEL A.  ONCE A HERO  C/P:2,P
STACKPOLE, MICHAEL A.  ONSLAUGHT [@STAR WARS NEW JEDI ORDER DARK TIDE #1]  P
STACKPOLE, MICHAEL A.  PERCHANCE TO DREAM  H
STACKPOLE, MICHAEL A.  PRINCE OF HAVOC [@BATTLETECH #39/TWILIGHT OF THE CLANS 7]  P
STACKPOLE, MICHAEL A.  ROGUE SQUADRON [@STAR WARS X-WING 1]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 365
AUTHOR  TITLE  CODE
STACKPOLE, MICHAEL A.  RUIN [@STAR WARS NEW JEDI ORDER DARK TIDE #2]  P
STACKPOLE, MICHAEL A.  SECRET ATLAS, A [AGE OF DISCOVERY 1]  H
STACKPOLE, MICHAEL A.  TALION REVENANT  P
STACKPOLE, MICHAEL A.  WARRIOR: COUPE [@BATTLETECH #43/WARRIOR 3]  C/P:2,P
STACKPOLE, MICHAEL A.  WARRIOR: EN GARDE [@BATTLETECH #41/WARRIOR 1]  C/P,P
STACKPOLE, MICHAEL A.  WARRIOR: RIPOSTE [@BATTLETECH #42/WARRIOR 2]  C/P:2,P
STACKPOLE, MICHAEL A.  WEDGE’S GAMBLE [@STAR WARS X-WING 2]  C/P,P
STACKPOLE, MICHAEL A.  WHEN DRAGONS RAGE [DRAGONCROWN WAR 2]  C/P,LP
STACKPOLE, MICHAEL A.  WOLF AND RAVEN [@SHADOWRUN 32]  P
STACY, RYDER  AMERICAN DEFIANCE [@DOOMSDAY WARRIOR #7]  P
STACY, RYDER  AMERICAN EDEN [@DOOMSDAY WARRIOR #11]  P
STACY, RYDER  AMERICAN GLORY [@DOOMSDAY WARRIOR #8]  P
STACY, RYDER  AMERICAN REBELLION [@DOOMSDAY WARRIOR #6]  P
STACY, RYDER  AMERICA’S LAST DECLARATION [@DOOMSDAY WARRIOR #5]  P
STACY, RYDER  AMERICA’S ZERO HOUR [@DOOMSDAY WARRIOR #9]  P
STACY, RYDER  BLOODY AMERICA [@DOOMSDAY WARRIOR #4]  P
STACY, RYDER  DEATH AMERICAN STYLE [@DOOMSDAY WARRIOR #12]  P
STACY, RYDER  DOOMSDAY WARRIOR [@DOOMSDAY WARRIOR #1]  P
STACY, RYDER  LAST AMERICAN, THE [@DOOMSDAY WARRIOR #3]  P
STACY, RYDER  RED AMERICA [@DOOMSDAY WARRIOR #2]  P
STADLER, JOHN  ECO-FICTION  PA
STAICAR, TOM  FEMININE EYE, THE  R/REF-H
STALLMAN, ROBERT  BEAST, THE  C/P,P
STALLMAN, ROBERT  CAPTIVE, THE  C/P:2,P
STALLMAN, ROBERT  ORPHAN, THE  C/P,P
STAMEY, SARA  DOUBLE BLIND [KURTIS:P385XL47:RUTH 3]  P
STAMEY, SARA  WILD CARD RUN [KURTIS:P385XL47:RUTH 1]  P,SFWA-P
STAMEY, SARA  WIN, LOSE, DRAW [KURTIS:P385XL47:RUTH 2]  P

STAMMERS, MARK/ WALKER, STEPHEN JAMES DECALOG [DOCTOR WHO]

C/PA,PA

STAMMERS, MARK/ WALKER, STEPHEN JAMES DECALOG 2: LOST PROPERTY [DOCTOR WHO]

C/PA:2,PA

STANEK, ROBERT  ELF QUEEN & THE KING [RUIN MIST]  LP
STANEK, ROBERT  ELF QUEEN & THE KING II [RUIN MIST]  LP
STANEK, ROBERT  ELF QUEEN’S QUEST [RUIN MIST DARK PATH 1]  LP
STANEK, ROBERT  FIELDS OF HONOR [RUIN MIST 3]  LP
STANEK, ROBERT  IN THE SERVICE OF DRAGONS [KEEPER MARTIN 5]  LP
STANEK, ROBERT  IN THE SERVICE OF DRAGONS II [KEEPER MARTIN 6]  LP
STANEK, ROBERT  IN THE SERVICE OF DRAGONS III [KEEPER MARTIN 7]  LP
STANEK, ROBERT  IN THE SERVICE OF DRAGONS IV [KEEPER MARTIN 8]  LP
STANEK, ROBERT  JOURNEY BEYOND THE BEYOND [MAGIC LANDS]  LP
STANEK, ROBERT  KEEPER MARTIN’S TALE [RUIN MIST 1]  LP
STANEK, ROBERT  KINGDOM ALLIANCE [RUIN MIST 2]  LP
STANEK, ROBERT  KINGDOMS AND ELVES OF THE REACHES, THE [KEEPER MARTIN 1]  LP
STANEK, ROBERT  KINGDOMS AND ELVES OF THE REACHES II, THE [KEEPER MARTIN 2]  LP
STANEK, ROBERT  KINGDOMS AND ELVES OF THE REACHES III, THE [KEEPER MARTIN 3]  LP
STANEK, ROBERT  KINGDOMS AND ELVES OF THE REACHES IV, THE [KEEPER MARTIN 4]  LP
STANEK, ROBERT  MARK OF THE DRAGON [RUIN MIST 4]  LP
STANFORD, RAY  WHAT YOUR AURA TELLS ME  SFWA
STANLEY, JOHN  CREATURE FEATURES, UPDATED EDITION  R/REF-H
STANLEY, JOHN  WORLD WAR III  P
STANTON, FRANK  MASS MEDIA AND MASS CULTURE  [C/RAN-H]
STANTON, JEROME  FLYING SAUCERS: HOAX OR REALITY?  [C/RAN-P]
STANTON, KEN  WHIRLWIND BENEATH THE SEA [AQUANAUTS 6]  P
STANTON, MARY  HEAVENLY HORSE FROM THE OUTERMOST WEST, THE  C/P:2,P
STANTON, MARY  PIPER AT THE GATE  C/P,SFWA-P,P
STANTON, PAUL  VILLAGE OF STARS  C/P,P,H
STANWOOD, BROOKS  GLOW, THE  H,P
STAPLEDON, OLAF  DARKNESS AND THE LIGHT  H,LP
STAPLEDON, OLAF  FAR FUTURE CALLING  SR-H
STAPLEDON, OLAF  FLAMES, THE  SR-H
STAPLEDON, OLAF  LAST AND FIRST MEN  SR-H,LP
STAPLEDON, OLAF  LAST AND FIRST MEN/LAST MEN IN LONDON  SR-P,C/P
STAPLEDON, OLAF  LAST AND FIRST MEN/STAR MAKER  H:2,LP
STAPLEDON, OLAF  LAST MEN IN LONDON  P
STAPLEDON, OLAF  NEBULA MAKER/FOUR ENCOUNTERS  LP
STAPLEDON, OLAF  ODD JOHN  C/P:2,P,GN2.08
STAPLEDON, OLAF  ODD JOHN AND SIRIUS  H,LP
STAPLEDON, OLAF  PHILOSOPHY AND LIVING 2  [C/RAN-P]
STAPLEDON, OLAF  SIRIUS  SR-P,C/P
STAPLEDON, OLAF  STAR MAKER  C/P,P,LP,H
STAPP, ROBERT  MORE PERFECT UNION, A  H,P
STAR TREK BOOKS  AMAZING STORIES, THE [STAR TREK]  LPA
STARBUCK, KATHLYN S.  INDIA’S STORY  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 366
AUTHOR  TITLE  CODE
STARBUCK, KATHLYN S.  TIME IN MIND  P
STARHAWK  FIFTH SACRED THING, THE  H
STARK, ED  BEYOND THE ’ZONE [SHATTERZONE 1]  P
STARK, ED  FIFTH HORSEMAN, THE [BLOODSHADOWS]  LP
STARLIN, JIM/GRAZIUNAS, DAINA  AMONG MADMEN  C/P,P
STARLIN, JIM/GRAZIUNAS, DIANA  LADY EL  C/P,SFWA-P,P
STARMONT HOUSE, INC.  RED TWILIGHT/WORLD’S END  SR-LPA
STARMONT HOUSE, INC.  WORLDS WITHIN WORLDS  SR-LPA
STARR, BILL  TREASURE OF WONDERWHAT, THE [FARSTAR & SON 2]  C/P:2,P
STARR, BILL  WAY TO DAWNWORLD, THE [FARSTAR & SON 1]  C/P:2,P
STARR, ROLAND  OPERATION OMINA  H
STASHEFF, CHRISTOPHER/FAWCETT, BILL  BLESSINGS AND CURSES [@CRAFTERS #2]  C/PA,PA,SFWA-P:2
STASHEFF, CHRISTOPHER  COMPANY OF STARS, A [@STARSHIP TROUPERS #1]  C/P,P,H
STASHEFF, CHRISTOPHER/FAWCETT, BILL  CRAFTERS, THE  PA,C/PA,SFWA-P
STASHEFF, CHRISTOPHER  CRUSADING WIZARD, THE [WIZARD IN RHYME 7]  C/P,P
STASHEFF, CHRISTOPHER  DAY THE MAGIC STOPPED, THE  PA
STASHEFF, CHRISTOPHER  DRAGON’S EYE  PA

STASHEFF, CHRISTOPHER/ FORSTCHEN, WILLIAM R.

END RUN [@WING COMMANDER 2]

P,SFWA-P

STASHEFF, CHRISTOPHER  ESCAPE VELOCITY [WARLOCK 0]  C/P:2,P
STASHEFF, CHRISTOPHER  FELINE WIZARD, THE [WIZARD IN RHYME 8]  C/P,P
STASHEFF, CHRISTOPHER  HAUNTED WIZARD, THE [WIZARD IN RHYME 6]  P
STASHEFF, CHRISTOPHER  HER MAJESTY’S WIZARD [WIZARD IN RHYME 1]  C/P,P,H
STASHEFF, CHRISTOPHER  HERE BE MONSTERS [WARLOCK’S HEIRS 5]  P
STASHEFF, CHRISTOPHER  KING KOBOLD [WARLOCK 2]  C/P:2,P
STASHEFF, CHRISTOPHER  KING KOBOLD REVIVED [WARLOCK 2]  C/P:2,P
STASHEFF, CHRISTOPHER  M’LADY WITCH [WARLOCK’S HEIRS 2]  C/P,P
STASHEFF, CHRISTOPHER  MY SON, THE WIZARD [WIZARD IN RHYME 5]  C/P,LP
STASHEFF, CHRISTOPHER  OATHBOUND WIZARD, THE [WIZARD IN RHYME 2]  C/P,P,H
STASHEFF, CHRISTOPHER  ODD WARLOCK OUT [WARLOCK 7,8,9]  H
STASHEFF, CHRISTOPHER  QUICKSILVER’S KNIGHT [WARLOCK’S HEIRS 3]  P
STASHEFF, CHRISTOPHER  SAGE, THE [@STAR STONE #2]  H,P
STASHEFF, CHRISTOPHER  SAINT VIDICON TO THE RESCUE  P
STASHEFF, CHRISTOPHER  SECULAR WIZARD, THE [WIZARD IN RHYME 4]  C/P,P,H
STASHEFF, CHRISTOPHER  SHAMAN, THE [@STAR STONE #1]  P
STASHEFF, CHRISTOPHER  SLIGHT DETOUR, A [@STARSHIP TROUPERS #3]  C/P:2,P
STASHEFF, CHRISTOPHER  SPELL-BOUND SCHOLAR, THE [WARLOCK’S HEIRS]  P
STASHEFF, CHRISTOPHER  TO THE MAGIC BORN [WARLOCK 0,1]  H:2
STASHEFF, CHRISTOPHER  WARLOCK AND SON [WARLOCK 11]  C/P,SFWA-P:2
STASHEFF, CHRISTOPHER  WARLOCK ENLARGED, THE [WARLOCK 2,3,4]  H:2
STASHEFF, CHRISTOPHER  WARLOCK ENRAGED, THE [WARLOCK 4]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK HERETICAL, THE [WARLOCK 7]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK IN SPITE OF HIMSELF, THE [WARLOCK 1]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK INSANE, THE [WARLOCK 9]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK IS MISSING, THE [WARLOCK 6]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK ROCK, THE [WARLOCK 10]  C/P,SFWA-P:2,P
STASHEFF, CHRISTOPHER  WARLOCK UNLOCKED, THE [WARLOCK 3]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK WANDERING, THE [WARLOCK 5]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK’S COMPANION, THE [WARLOCK 8]  C/P:2,P
STASHEFF, CHRISTOPHER  WARLOCK’S LAST RIDE, THE [WARLOCK 12]  P
STASHEFF, CHRISTOPHER  WARLOCK’S NIGHT OUT, THE [WARLOCK 5,6]  H:2
STASHEFF, CHRISTOPHER  WE OPEN ON VENUS [@STARSHIP TROUPERS #2]  C/P:2,P
STASHEFF, CHRISTOPHER  WITCH DOCTOR, THE [WIZARD IN RHYME 3]  C/P,P,H
STASHEFF, CHRISTOPHER  WIZARD AND A WARLORD, A [ROGUE WIZARD 7]  H
STASHEFF, CHRISTOPHER  WIZARD IN A FEUD, A [ROGUE WIZARD 9]  H,P
STASHEFF, CHRISTOPHER  WIZARD IN ABSENTIA, A [WARLOCK’S HEIRS 1]  C/P:2,SFWA-P:2,P
STASHEFF, CHRISTOPHER  WIZARD IN BEDLAM, A [ROGUE WIZARD 2]  H:2,SFWA,P
STASHEFF, CHRISTOPHER  WIZARD IN CHAOS, A [ROGUE WIZARD 5]  H
STASHEFF, CHRISTOPHER  WIZARD IN MIDGARD, A [ROGUE WIZARD 6]  H
STASHEFF, CHRISTOPHER  WIZARD IN MIND, A [ROGUE WIZARD 1]  P
STASHEFF, CHRISTOPHER  WIZARD IN PEACE, A [ROGUE WIZARD 4]  P
STASHEFF, CHRISTOPHER  WIZARD IN THE WAY, A [ROGUE WIZARD 8]  H
STASHEFF, CHRISTOPHER  WIZARD IN WAR, A [ROGUE WIZARD 3]  H,P
STATEHAM, B. R.  BANNERS OF THE SA’YEN  P
STATON, MARY  FROM THE LEGEND OF BIEL  C/P:2,P
STATTEN, VARGO  ACROSS THE AGES  SR-P
STATTEN, VARGO  ANNIHILATION!  SR-P
STATTEN, VARGO  AVENGING MARTIAN, THE  SR-P
STATTEN, VARGO  BLACK AVENGERS, THE  SR-P
STATTEN, VARGO  BLACK BARGAIN  SR-P
STATTEN, VARGO  BLACK WING OF MARS  SR-P
STATTEN, VARGO  BORN OF LUNA  SR-P
STATTEN, VARGO  CATACLYSM  SR-P
STATTEN, VARGO  CATALYST, THE  SR-P

AUTHOR  TITLE  CODE
STATTEN, VARGO  COSMIC FLAME, THE  SR-P
STATTEN, VARGO  DEADLINE TO PLUTO  SR-P
STATTEN, VARGO  DECREATION  SR-P
STATTEN, VARGO  DEVOURING FIRE  SR-P
STATTEN, VARGO  DUST DESTROYERS, THE  SR-P
STATTEN, VARGO  ECLIPSE EXPRESS, THE  SR-P
STATTEN, VARGO  G-BOMB, THE  SR-P
STATTEN, VARGO  GRAND ILLUSION, THE  SR-P
STATTEN, VARGO  I SPY  SR-P
STATTEN, VARGO  INFERNO!  SR-P
STATTEN, VARGO  INNER COSMOS, THE  SR-P
STATTEN, VARGO  LAST MARTIAN, THE  SR-P
STATTEN, VARGO  LAUGHTER IN SPACE  SR-P
STATTEN, VARGO  LIE DESTROYER, THE  SR-P
STATTEN, VARGO  MAN FROM TOMORROW, THE  SR-P
STATTEN, VARGO  MAN IN DUPLICATE  SR-P
STATTEN, VARGO  MAN OF TWO WORLDS  SR-P
STATTEN, VARGO  MICRO MEN, THE  SR-P
STATTEN, VARGO  NEBULA-X  SR-P
STATTEN, VARGO  NEW SATELLITE, THE  SR-P
STATTEN, VARGO  1000-YEAR VOYAGE  SR-P
STATTEN, VARGO  PETRIFIED PLANET, THE  SR-P
STATTEN, VARGO  PIONEER, 1990  SR-P
STATTEN, VARGO  RED INSECTS, THE  SR-P
STATTEN, VARGO  RENEGADE STAR, THE  SR-P
STATTEN, VARGO  SCIENCE METROPOLIS  SR-P
STATTEN, VARGO  SPACE WARP, THE  SR-P
STATTEN, VARGO  SUN-MAKERS, THE  SR-P
STATTEN, VARGO  TIME APPOINTED, A  SR-P
STATTEN, VARGO  TIME BRIDGE, THE  SR-P,C/P
STATTEN, VARGO  TIME TRAP, THE  SR-P
STATTEN, VARGO  TO THE ULTIMATE  SR-P
STATTEN, VARGO  2000 YEARS ON  SR-P
STATTEN, VARGO  ULTRA SPECTRUM  SR-P
STATTEN, VARGO  WANDERER OF SPACE  SR-P
STATTEN, VARGO  WEALTH OF THE VOID  SR-P
STATTEN, VARGO  WORLDS TO CONQUER  SR-P
STATTEN, VARGO  ZERO HOUR  SR-P
STAUB, WENDY CORSI  VOODOO MOON [@CHARMED]  P
STAUDINGER, MICHAEL C.  FALCON RISES, THE  P
STAUFFER, DONALD A.  SAINT AND THE HUNCHBACK, THE  H
STAVIS, R. H.  DANIEL’S VEIL  P
STEAKLEY, JOHN  ARMOR  C/P:2,P
STEAKLEY, JOHN  VAMPIRE$  P,LP
STEARN, JESS  MIRACLE WORKERS, THE  [C/RAN-P]
STEEL, MARK  TROUBLE PLANET  SR-P
STEELE, ADDISON E.  BUCK ROGERS IN THE 25TH CENTURY [@BUCK ROGERS #1]  P
STEELE, ADDISON E.  THAT MAN ON BETA [@BUCK ROGERS #2]  C/P,P
STEELE, ALEXANDER  UNLEASHED IN SPACE [WISHBONE 3]  LP
STEELE, ALLEN  ALL-AMERICAN ALIEN BOY  C/P,P
STEELE, ALLEN  CHRONOSPACE  H
STEELE, ALLEN  CLARKE COUNTY, SPACE  C/P:2,SFWA-P:2,P
STEELE, ALLEN  COYOTE [COYOTE 1]  H
STEELE, ALLEN  COYOTE FRONTIER [COYOTE 3]  H
STEELE, ALLEN  COYOTE RISING [COYOTE 2]  H
STEELE, ALLEN  JERICHO ITERATION, THE  P
STEELE, ALLEN  KING OF INFINITE SPACE, A [NEAR SPACE]  H:2,P
STEELE, ALLEN  LABYRINTH OF NIGHT  C/P:2,SFWA-P,P
STEELE, ALLEN  LUNAR DESCENT  C/P:2,SFWA-P:2,P
STEELE, ALLEN  OCEANSPACE  H:2,P
STEELE, ALLEN  ORBITAL DECAY  SR-P,C/P:2,SFWA-P,P
STEELE, ALLEN  RUDE ASTRONAUTS  C/P,P
STEELE, ALLEN  SPINDRIFT  H
STEELE, ALLEN  TRANQUILITY ALTERNATIVE, THE  P
STEELE, CURTIS  CAVERN OF THE DAMNED  SR-LPA
STEELE, CURTIS  INVISIBLE EMPIRE, THE [OPERATOR 5 #2]  SR-P,C/P
STEELE, CURTIS  LEGIONS OF STARVATION  SR-LPA
STEELE, CURTIS  MASKED INVASION, THE [OPERATOR 5 #1]  SR-P
STEELE, CURTIS  YELLOW SCOURGE, THE [OPERATOR 5 #3]  SR-P
STEELE, JOHN H.  TOWER OF BABEL [MAGE ASCENSION]  P
STEELE, LINDA  IBIS  C/P,P
STEELE, MARY Q.  JOURNEY OUTSIDE  H
STEELE, MARY Q.  TRUE MEN, THE  H
STEELE, MAX  WHERE SHE BRUSHED HER HAIR AND OTHER STORIES  H

AUTHOR  TITLE  CODE
STEEN, MARGUERITE  UNQUIET SPIRIT, THE  C/P,P
STEFFANSON, CON  LION MEN OF MONGO [FLASH GORDON 1]  P
STEFFANSON, CON  PLAGUE OF SOUND, THE [FLASH GORDON 2]  P
STEFFANSON, CON  SPACE CIRCUS, THE [FLASH GORDON 3]  C/P,P
STEFFANSON, CON  TIME TRAP OF MING XIII, THE [FLASH GORDON 4]  C/P,P
STEIGER, BRAD  ALIEN MEETINGS  [C/RAN-P]
STEIGER, BRAD  ATLANTIS RISING  [C/RAN-P]
STEIGER, BRAD  ESP: YOUR SIXTH SENSE  [C/RAN-P:2]
STEIGER, BRAD/WHITE, JOHN  OTHER WORLDS, OTHER UNIVERSES  SFWA
STEIGER, BRAD  STRANGE ENCOUNTERS WITH GHOSTS  [C/RAN-P]
STEIGER, BRAD  STRANGE POWERS OF HEALING  [C/RAN-P]
STEIGER, BRAD  STRANGERS FROM THE SKIES  [C/RAN-P]
STEIN, JEANNE C.  BECOMING, THE [ANNA STRONG 1]  P
STEIN, JEANNE C.  BLOOD DRIVE [ANNA STRONG 2]  P
STEIN, KEVIN  BROTHERS MAJERE [DRAGONLANCE PRELUDES 3]  C/P:2,P
STEIN, KEVIN  GUIDE TO LARRY NIVEN’S RINGWORLD, THE  R/REF-H
STEIN, KEVIN  TWISTED DRAGON [ELFWOOD]  P
STEINBECK, JOHN  ACTS OF KING ARTHUR AND HIS NOBLE KNIGHTS, THE  C/P:2,LP
STEINBECK, JOHN  SHORT REIGN OF PIPPIN IV, THE  H,P
STEINBRUNNER, CHRIS/GOLDBLATT, BURT  CINEMA OF THE FANTASTIC  R/REF-H
STEINER, KURT  OCEANS DU CIEL, LES  FP
STEMBER, NICHOLAS S.  TWILIGHT  LP
STENEMAN, SHEP  EMPIRE STRIKES BACK STORYBOOK  SR-VLH
STEPHENS, BRYNNE  DREAM PALACE, THE  P
STEPHENS, JAMES  CROCK OF GOLD, THE  H,LP
STEPHENS, REED  MAN WHO KILLED HIS BROTHER, THE  C/P,P
STEPHENSON, ANDREW M.  NIGHTWATCH  C/P,P
STEPHENSON, ANDREW M.  WALL OF YEARS, THE  C/P,P
STEPHENSON, NEAL  BIG U, THE  H,LP
STEPHENSON, NEAL/GEORGE, J. FREDERICK COBWEB, THE  LP
STEPHENSON, NEAL  CONFUSION, THE [BAROQUE CYCLE 2]  H,LP
STEPHENSON, NEAL  CRYPTONOMICON  H:3,LP
STEPHENSON, NEAL  DIAMOND AGE, THE  C/P,P,H
STEPHENSON, NEAL  KING OF THE VAGABONDS [BAROQUE CYCLE 1B]  P
STEPHENSON, NEAL  QUICKSILVER [BAROQUE CYCLE 1]  H:3,LP
STEPHENSON, NEAL  SNOW CRASH  C/P:4,P,H
STEPHENSON, NEAL  SYSTEM OF THE WORLD, THE [BAROQUE CYCLE 3]  H
STEPHENSON, NEAL  ZODIAC  C/P:2,LP
STERLING, BRETT  DANGER PLANET  C/P:2,P
STERLING, BRETT  TENTH PLANET, THE  P
STERLING, BRUCE  ARTIFICIAL KID, THE  C/P,SFWA-P,H,P,SFWA
STERLING, BRUCE  CRYSTAL EXPRESS  C/P:2,SFWA-P:2,P
STERLING, BRUCE  DISTRACTION  H,P
STERLING, BRUCE  GLOBALHEAD  C/P:2,P
STERLING, BRUCE  GOOD OLD-FASHIONED FUTURE, A  P
STERLING, BRUCE  HACKER CRACKDOWN, THE  R/REF-P
STERLING, BRUCE  HEAVY WEATHER  C/P,P,H
STERLING, BRUCE  HOLY FIRE  H:2,SR-P
STERLING, BRUCE  INVOLUTION OCEAN  C/P,P
STERLING, BRUCE  ISLANDS IN THE NET  C/P:2,SFWA-P,P
STERLING, BRUCE  MIRRORSHADES  PA,SFWA
STERLING, BRUCE  SCHISMATRIX  C/P,SFWA,P,H
STERLING, BRUCE  VISIONARY IN RESIDENCE  LP
STERLING, BRUCE  ZEITGEIST  H
STERLING, BRUCE  ZENITH ANGLE, THE  H,P
STERN, DANIEL  SUICIDE ACADEMY  P
STERN, DAVE  DAEDALUS’S CHILDREN, PART TWO [STAR TREK ENTERPRISE]  P
STERN, DAVE  DEADALUS [STAR TREK ENTERPRISE]  P
STERN, DAVE  LARA CROFT TOMB RAIDER  P
STERN, DAVE  ROSETTA [STAR TREK ENTERPRISE]  P
STERN, DAVE  TRANSFORMATIONS (CASSETTE) [STAR TREK]  SR-MM
STERN, DAVE  WHAT PRICE HONOR? [STAR TREK ENTERPRISE]  P
STERN, DAVID  FRANCIS GOES TO WASHINGTON  H
STERN, DAVID  NIGHTMARE WORLD [DR. BONES]  P
STERN, PHILIP VAN DOREN  GREAT GHOST STORIES  PA
STERN, PHILIP VAN DOREN  GREAT TALES OF FANTASY AND IMAGINATION  C/PA,PA
STERN, PHILIP VAN DOREN  MIDNIGHT READER  SR-PA

STERN, PHILIP VAN DOREN MOONLIGHT TRAVELER:

GREAT TALES OF FANTASY AND IMAGINATION, THE

HA

STERN, PHILIP VAN DOREN  OTHER SIDE OF THE CLOCK, THE  PA
STERN, PHILIP VAN DOREN  STRANGE BEASTS AND UNNATURAL MONSTERS  PA
STERN, PHILIP VAN DOREN  TRAVELERS IN TIME  HA
STERN, ROGER  DEATH AND LIFE OF SUPERMAN, THE [DC COMICS]  C/P,P,H
STERN, ROGER  STRANGE VISITORS [@SMALLVILLE]  P

AUTHOR  TITLE  CODE
STERN, STEVE  HARRY KAPLAN’S ADVENTURES UNDERGROUND  H
STERNAU, CYNTHIA/GREENBERG, MARTIN H. SECRET PROPHECIES OF NOSTRADAMUS, THE  C/PA,PA

STERNBACH, RICK/OKUDA, MICHAEL STAR TREK THE NEXT GENERATION TECHNICAL MANUAL

[STAR TREK TNG]

R/REF-H,C/REF-H

STERNBACH, RICK/OKUDA, MICHAEL STAR TREK THE NEXT GENERATION TECHNICAL MANUAL

(FIFTH SEASON EDITION) [STAR TREK TNG]

R/REF-H

STERNBERG, JACQUES  FUTURE WITHOUT FUTURE  H,SFWA
STERNBERG, JACQUES  SEXUALIS ’95  P
STETTEN, GEORGE  WEISSENBAUM’S EYE  C/P,P
STEUSSY, MARTI  DREAMS OF DAWN  C/P,P,H
STEUSSY, MARTI  FOREST OF THE NIGHT  C/P:2,P
STEVENS, FRANCIS  CITADEL OF FEAR, THE  C/P:2,P
STEVENS, FRANCIS  CLAIMED  P
STEVENS, FRANCIS  HEADS OF CERBERUS, THE  SR-H,H
STEVENS, JAMES/BISHOP, DAVID  WHO KILLED KENNEDY [DOCTOR WHO]  C/P,P
STEVENSON, E.  AVENGING SPIRIT, THE [TWILIGHT 10]  P
STEVENSON, JENNIFER  TRASH SEX MAGIC  H
STEVENSON, LAURA C.  ISLAND AND THE RING, THE  P
STEVENSON, ROBERT LOUIS  DOCTOR JEKYLL AND MR. HYDE  C/P:2,P
STEVENSON, ROBERT LOUIS  DOCTOR JEKYLL AND MR. HYDE AND OTHER STORIES (POCKET)  P
STEVENSON, ROBERT LOUIS  DOCTOR JEKYLL AND MR. HYDE AND OTHER STORIES (SCHOLASTIC)  P
STEVENSON, ROBERT LOUIS  MERRY MEN, THE/STRANGE CASE OF DR. JEKYLL AND MR. HYDE, THE  H
STEVENSON, ROBERT LOUIS  STRANGE CASE OF DR. JEKYLL AND MR. HYDE AND OTHER STORIES, THE P
STEVER, DAVID/WHYTE, ANDREW ADAMS  COLLECTOR’S POUL ANDERSON, THE  R/IND:2
STEVER, DAVID/BOYAJIAN, JERRY  JOHN SCHOENHERR SF CHECKLIST, A  R/IND:2
STEVERMER, CAROLINE  COLLEGE OF MAGICS, A [COLLEGE OF MAGICS 1]  H
STEVERMER, CAROLINE  SCHOLAR OF MAGICS, A [COLLEGE OF MAGICS 2]  H
STEVERMER, CAROLINE  SERPENT’S EGG, THE  P
STEVERMER, CAROLINE  WHEN THE KING COMES HOME  P
STEWART, DANIEL BLAIR  AKHUNATON  LP
STEWART, FRED MUSTARD  MEPHISTO WALTZ, THE  C/P:2,P
STEWART, FRED MUSTARD  METHUSELAH ENZYME, THE  C/P,P
STEWART, FRED MUSTARD  STAR CHILD  H,P
STEWART, GEORGE R.  EARTH ABIDES  H:2,P
STEWART, IAN  FLATTERLAND  H
STEWART, IAN/COHEN, JACK  HEAVEN  H
STEWART, IAN/COHEN, JAMES  WHEELERS  H:2,P
STEWART, JEAN  ISIS RISING [ISIS 2]  LP
STEWART, JEAN  WARRIORS OF ISIS [ISIS 3]  LP
STEWART, JEAN  WINGED ISIS [ISIS 4]  LP
STEWART, JEAN  WIZARD OF ISIS [ISIS 5]  LP
STEWART, KERRY  RUBY  P
STEWART, MARY  CRYSTAL CAVE, THE [ARTHURIAN 1]  C/P:2,P,LP
STEWART, MARY  HOLLOW HILLS, THE [ARTHURIAN 2]  C/P,P,LP,H
STEWART, MARY  LAST ENCHANTMENT, THE [ARTHURIAN 3]  H:2,LP,P
STEWART, MARY  THORNYHOLD  P
STEWART, MARY  TOUCH NOT THE CAT  P
STEWART, MARY  WALK IN WOLF WOOD, A  P
STEWART, MARY  WICKED DAY, THE [ARTHURIAN 4]  H:2,LP,P
STEWART, MICHAEL  BLINDSIGHT  H
STEWART, MICHAEL  GRACE  H
STEWART, MICHAEL  MONKEY SHINES  H,P
STEWART, SEAN  CLOUDS END  C/P,P,H
STEWART, SEAN  DARK RENDEZVOUS [@STAR WARS CLONE WARS YODA]  P
STEWART, SEAN  GALVESTON [RESURRECTION MAN 2]  LP
STEWART, SEAN  NIGHT WATCH, THE [RESURRECTION MAN 3]  C/P:2,P
STEWART, SEAN  NOBODY’S SON  P
STEWART, SEAN  PASSION PLAY  C/P:2,SFWA-P,P
STEWART, SEAN  PERFECT CIRCLE  H
STEWART, SEAN  RESURRECTION MAN [RESURRECTION MAN 1]  C/P,LP
STEWART, WILL  SEETEE SHIP  SR-H
STEWART, WILL  SEETEE SHOCK  H
STICKGOLD, BOB  CALIFORNIA COVEN PROJECT, THE  P
STICKGOLD, BOB/NOBLE, MARK  GLORYHITS  C/P,P,H
STIEGLER, MARC  DAVID’S SLING  C/P,SFWA-P,P
STIEGLER, MARC  EARTHWEB  C/P,P
STIEGLER, MARC  GENTLE SEDUCTION, THE  P,SFWA-P
STILLMAN, RON  TRACKER  P
STILSON, CHARLES B.  MINOS OF SARDANES  H
STILSON, CHARLES B.  POLARIS AND THE IMMORTALS  H
STILSON, CHARLES B.  POLARIS OF THE SNOWS  H
STINE, G. HARRY  BASTAARD REBELLION, THE [@WARBOTS #3]  C/P:2,P
STINE, G. HARRY  EARTH SATELLITES  [C/RAN-P]
STINE, G. HARRY  FIRST ACTION [@STARSEA INVADERS 1]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 370
AUTHOR  TITLE  CODE
STINE, G. HARRY  FORCE OF ARMS [@WARBOTS #8]  P
STINE, G. HARRY  OPERATION IRON FIST [@WARBOTS #7]  P
STINE, G. HARRY  OPERATION STEEL BAND [@WARBOTS #2]  C/P,P
STINE, G. HARRY  SECOND CONTACT [@STARSEA INVADERS 2]  C/P,P
STINE, G. HARRY  SHUTTLE INTO SPACE  [C/RAN-H]
STINE, G. HARRY  SIERRA MADRE [@WARBOTS #4]  P
STINE, G. HARRY  SPACE ENTERPRISE, THE  [C/RAN-P]
STINE, G. HARRY  SPACE POWER  [C/RAN-P]
STINE, G. HARRY  THIRD ENCOUNTER [@STARSEA INVADERS 3]  P
STINE, G. HARRY  THIRD INDUSTRIAL REVOLUTION, THE  [C/RAN-P:2]
STINE, G. HARRY  WARBOTS [@WARBOTS #1]  C/P,P
STINE, HANK  PRISONER #3, THE [PRISONER]  C/P:2,P
STINE, R. L.  ATTACK OF THE MUTANT [GOOSEBUMPS]  LP
STINE, R. L.  STAY OUT OF THE BASEMENT [GOOSEBUMPS]  LP
STIRLING, S. M.  AGAINST THE TIDE OF YEARS [NANTUCKET 2]  P
STIRLING, S. M./DRAKE, DAVID  ANVIL, THE [GENERAL 3]  C/P:2,SFWA-P,P
STIRLING, S. M.  BETRAYALS [@BABYLON 5 #6]  C/P,P

STIRLING, S. M./TARR, JUDITH/ SHWARTZ, SUSAN/TURTLEDOVE, HARRY BLOOD FEUDS [@WAR WORLD]

C/P,SFWA-P,P

STIRLING, S. M./TARR, JUDITH/ SHWARTZ, SUSAN/TURTLEDOVE, HARRY BLOOD VENGEANCE [@WAR WORLD]

P,SFWA-P

STIRLING, S. M./MEIER, SHIRLEY  CAGE, THE [FIFTH MILLENNIUM]  C/P,SFWA-P:2,P
STIRLING, S. M./DRAKE, DAVID  CHOSEN, THE [GENERAL]  C/P:2,P
STIRLING, S. M./DRAKE, DAVID  CONQUEROR (PROOF) [GENERAL 3,4,5]  SR-LP
STIRLING, S. M.  CONQUISTADOR  H
STIRLING, S. M.  DIES THE FIRE [EMBERVERSE 1]  H
STIRLING, S. M.  DOMINATION, THE [DRAKA (REVISED) 1,2,3]  H
STIRLING, S. M.  DRAKAS! [DRAKA]  C/PA,PA
STIRLING, S. M.  DRAKON [DRAKA 4]  C/P:2,P
STIRLING, S. M./DRAKE, DAVID  FORGE, THE [@GENERAL #1]  C/P:2,P
STIRLING, S. M./DRAKE, DAVID  HAMMER, THE [@GENERAL #2]  C/P:2,SFWA-P,P
STIRLING, S. M.  ISLAND IN THE SEA OF TIME [NANTUCKET 1]  C/P,P
STIRLING, S. M.  MARCHING THROUGH GEORGIA [DRAKA 1]  C/P:2,SFWA-P,P
STIRLING, S. M.  MEETING AT CORVALLIS, A [EMBERVERSE 3]  H
STIRLING, S. M.  ON THE OCEANS OF ETERNITY [NANTUCKET 3]  C/P:2,P
STIRLING, S. M.  PESHAWAR LANCERS, THE  H,P
STIRLING, S. M.  POWER  C/PA,PA,SFWA-P
STIRLING, S. M.  PROTECTOR’S WAR, THE [EMBERVERSE 2]  H
STIRLING, S. M./DRAKE, DAVID  REFORMER, THE [GENERAL]  H,P
STIRLING, S. M./LISLE, HOLLY  ROSE SEA, THE  P
STIRLING, S. M./MEIER, SHIRLEY  SABER AND SHADOW [FIFTH MILLENNIUM]  C/P,P
STIRLING, S. M./MEIER, SHIRLEY  SHARPEST EDGE, THE [FIFTH MILLENNIUM]  C/P,P
STIRLING, S. M.  SHIP AVENGED, THE [SHELLPEOPLE 7]  C/P,P,H
STIRLING, S. M.  SKY PEOPLE, THE  H
STIRLING, S. M.  SNOWBROTHER [FIFTH MILLENNIUM]  C/P:2,P
STIRLING, S. M./DRAKE, DAVID  STEEL, THE [@GENERAL #4]  P,SFWA-P
STIRLING, S. M.  STONE DOGS, THE [DRAKA 3]  C/P,P
STIRLING, S. M./DRAKE, DAVID  SWORD, THE [@GENERAL #5]  C/P:2,P
STIRLING, S. M.  T2: INFILTRATOR [TERMINATOR T 2 1]  H
STIRLING, S. M.  T2: RISING STORM [TERMINATOR T 2 2]  P
STIRLING, S. M.  T2: THE FUTURE WAR [TERMINATOR T 2 3]  P
STIRLING, S. M.  UNDER THE YOKE [DRAKA 2]  C/P,SFWA-P,P
STIRLING, S. M./DRAKE, DAVID  WARLORD  H
STITH, JOHN E.  DEATH TOLLS  P
STITH, JOHN E.  DEEP QUARRY  P
STITH, JOHN E.  MANHATTAN TRANSFER  H,P
STITH, JOHN E.  MEMORY BLANK  C/P:2,P
STITH, JOHN E.  RECKONING INFINITY  P
STITH, JOHN E.  REDSHIFT RENDEZVOUS  C/P,P
STITH, JOHN E.  REUNION ON NEVEREND  P
STITH, JOHN E.  SCAPESCOPE  P
STOCKBRIDGE, GRANT  CITY DESTROYER, THE [@SPIDER #3]  P
STOCKBRIDGE, GRANT  DEATH AND THE SPIDER [@SPIDER #4]  P
STOCKBRIDGE, GRANT  DEATH REIGN OF THE VAMPIRE KING [@SPIDER #1]  P
STOCKBRIDGE, GRANT  MASTER OF MEN! [@SPIDER #2]  P
STOCKTON, FRANK R.  GREAT STONE OF SARDIS, THE  H,P
STOCKTON, FRANK R.  JOHN GAYTHER’S GARDEN AND THE STORIES TOLD THEREIN  H
STOCKTON, FRANK R.  LADY OR THE TIGER AND OTHER STORIES, THE  H,P
STOCKTON, FRANK R.  VIZIER OF THE TWO-HORNED ALEXANDER, THE  H
STODDARD, JAMES  FALSE HOUSE, THE [EVENMERE 2]  P
STODDARD, JAMES  HIGH HOUSE, THE [EVENMERE 1]  H,P
STODDARD, JAMES  HIGH HOUSE, THE (PROOF) [EVENMERE 1]  SR-P
STOKER, BRAM  ANNOTATED DRACULA, THE  SR-VLH
STOKER, BRAM  BRAM STOKER BEDSIDE COMPANION, THE  H

AUTHOR  TITLE  CODE
STOKER, BRAM  DRACULA  H:2,P
STOKER, BRAM  GARDEN OF EVIL, THE  P
STOKER, BRAM  JEWEL OF SEVEN STARS, THE  C/P,VLP
STOKER, BRAM  LADY OF THE SHROUD, THE  P
STOKES, NAOMI M.  TREE PEOPLE, THE  P
STOLBOV, BRUCE  LAST FALL  H,SFWA

STOLZE, GREG ASHES AND ANGEL WINGS

[@WORLD OF DARKNESS DEMON THE FALLEN #1]

P

STOLZE, GREG WRECKAGE OF PARADISE, THE

[@WORLD OF DARKNESS DEMON THE FALLEN #3]

P

STONE, ANDY  SONG OF THE KINGDOM  P
STONE, CHARLOTTE  CHEON OF WELTANLAND I: THE FOUR WISHES  C/P,P
STONE, DAVE  BURNING HEART [DOCTOR WHO MISSING ADVENTURES]  C/P,P
STONE, DAVE  CITADEL OF DREAMS [DOCTOR WHO]  SR-H
STONE, DAVE  DEATH AND DIPLOMACY [DOCTOR WHO NEW ADVENTURES]  C/P,P
STONE, DAVE  DEATHMASQUES [JUDGE DREDD]  P
STONE, DAVE  GOLGOTHA RUN [@DARK FUTURE EDDIE KALISH 1]  P
STONE, DAVE  HEART OF TARDIS [DOCTOR WHO PAST DOCTORS]  P

STONE, DAVE/SUTTON, PAUL/ LIDSTER, JOSEPH LIFE IN PIECES, A [BERNICE SUMMERFIELD]

H

STONE, DAVE MARY-SUE EXTRUSION, THE

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

P

STONE, DAVE  OBLIVION [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P

STONE, DAVE PROFESSOR BERNICE SUMMERFIELD AND THE INFERNAL NEXUS

[BERNICE SUMMERFIELD 4]

P

STONE, DAVE RETURN TO THE FRACTURED PLANET

[DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]

P

STONE, DAVE  SHIP OF FOOLS [DOCTOR WHO NEW ADVENTURES (DOCTORLESS)]  C/P,P
STONE, DAVE  SKY PIRATES! [DOCTOR WHO NEW ADVENTURES]  C/P:2,P
STONE, DAVE  SLOW EMPIRE, THE [DOCTOR WHO EIGHTH DOCTOR]  P
STONE, GEORGE  BLIZZARD  H,P
STONE, GRAHAM  AUSTRALIAN SCIENCE FICTION INDEX 1939–1962  R/IND:2
STONE, GRAHAM  INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 1  R/IND
STONE, GRAHAM  INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 2  R/IND
STONE, GRAHAM  INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 4  R/IND
STONE, GRAHAM  INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 5  R/IND
STONE, GRAHAM  INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 PART 6  R/IND

STONE, GRAHAM INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 VOLUME 2

(AUTHORS)

R/IND

STONE, GRAHAM INDEX TO BRITISH SCIENCE FICTION MAGAZINES 1934–1953 VOLUME 3

(TITLES)

R/IND

STONE, IDELLA PURNELL  14 GREAT TALES OF ESP  PA
STONE, IDELLA PURNELL  NEVER IN THIS WORLD  C/PA,PA
STONE, JOSEPHINE RECTOR  PRAISE ALL THE MOONS OF MORNING  H
STONE, JOSEPHINE RECTOR  THOSE WHO FALL FROM THE SUN  H
STONE, LESLIE F.  OUT OF THE VOID  H
STONE, ROBERT S.  DARK WATERS [CHRONICLES UNBINDING 2]  P
STONE, ROBERT S.  HAZARD’S PRICE [CHRONICLES UNBINDING 1]  P
STONG, PHIL  OTHER WORLDS  HA
STOPY  VESMIRNE HRY  FH
STOREY, MARGARET  DRAGON’S SISTER, THE/TIMOTHY TRAVELS  C/P,LP
STOREY, MARGARET  STONE WIZARD, THE  P
STORM, BRIAN  RED STORM  SR-P
STORM, L. ELIZABETH  ANGELS UNAWARE [QUANTUM LEAP]  P
STORM, L. ELIZABETH  PULITZER [QUANTUM LEAP]  C/P,P
STORME, PETER  THING IN THE BROOK, THE  H
STORMONT, LAN  TAN MING  H
STOUGHTON, RICHARD P.  ULTIMA ONE  P
STOUT, AMY  ROYAL FOUR, THE [ONE LAND 2]  P
STOUT, AMY  SACRED SEVEN, THE  P
STOUT, REX  UNDER THE ANDES  C/P,P,H
STOVER, LEON E./HARRISON, HARRY  APEMAN, SPACEMAN  PA
STOVER, MATTHEW WOODRING  BLADE OF TYSHALE [HARI MICHAELSON 2]  P
STOVER, MATTHEW WOODRING  HEROES DIE [HARI MICHAELSON 1]  H,LP
STOVER, MATTHEW WOODRING  IRON DAWN [BARRA THE PICT 1]  C/P,P
STOVER, MATTHEW WOODRING  JERICHO MOON [BARRA THE PICT 2]  LP
STOVER, MATTHEW WOODRING  REVENGE OF THE SITH [@STAR WARS]  H:2
STOVER, MATTHEW WOODRING  SHATTERPOINT [@STAR WARS CLONE WARS]  P
STOVER, MATTHEW WOODRING  TRAITOR [@STAR WARS NEW JEDI ORDER]  P
STRACHEY, BARBARA  JOURNEYS OF FRODO  R/REF-P
STRACZYNSKI, J. MICHAEL  RISING STARS: BORN IN FIRE [RISING STARS]  CX
STRACZYNSKI, J. MICHAEL  RISING STARS: FIRE AND ASH [@RISING STARS #3]  CX
STRACZYNSKI, J. MICHAEL  RISING STARS: POWER [RISING STARS]  CX
STRACZYNSKI, J. MICHAEL  STRACZYNSKI UNPLUGGED  LP
STRACZYNSKI, J. MICHAEL  TALES FROM THE NEW TWILIGHT ZONE  P,SFWA-P

AUTHOR  TITLE  CODE
STRAHAN, JONATHAN  BEST SCIENCE FICTION AND FANTASY OF THE YEAR VOLUME 1, THE  LPA
STRAHAN, JONATHAN  BEST SHORT NOVELS 2004  HA
STRAHAN, JONATHAN  BEST SHORT NOVELS 2005  HA
STRAHAN, JONATHAN  BEST SHORT NOVELS 2006  HA
STRAHAN, JONATHAN  MODERN GREATS OF SCIENCE FICTION  LPA
STRAHAN, JONATHAN  SCIENCE FICTION: THE VERY BEST OF 2005  LPA
STRATTON, THOMAS  INVISIBILITY AFFAIR, THE [@MAN FROM UNCLE #11]  C/P,P
STRATTON, THOMAS  MIND TWISTER AFFAIR, THE [@MAN FROM UNCLE #12]  C/P:2,P
STRAUB, PETER  FLOATING DRAGON  H,P
STRAUB, PETER  GHOST STORY  P
STRAUB, PETER  HOUSES WITHOUT DOORS  P
STRAUB, PETER  IF YOU COULD SEE ME NOW  C/P,P
STRAUB, PETER  IN THE NIGHT ROOM  H
STRAUB, PETER  SHADOWLAND  H:2,P
STRAUB, PETER  WILD ANIMALS  H
STRAUSS, ERWIN S.  INDEX TO THE S-F MAGAZINES 1966, THE  R/IND:2
STRAUSS, ERWIN S.  INDEX TO THE SF MAGAZINES, 1951–1965  R/IND:2,SRL/IND
STRAUSS, VICTORIA  ARM OF THE STONE, THE  P
STRAUSS, VICTORIA  BURNING LAND, THE  P
STRAUSS, VICTORIA  BURNING LAND, THE (PROOF)  SR-LP
STRAUSS, VICTORIA  GARDEN OF THE STONE, THE  C/P,P
STRAUSS, VICTORIA  WORLDSTONE  C/P,P
STRETE, CRAIG  DEATH CHANTS  SFWA
STRETE, CRAIG  DEATH IN THE SPIRIT HOUSE  H
STRETE, CRAIG  DREAMS THAT BURN IN THE NIGHT  H,SFWA
STRETE, CRAIG  IF ALL ELSE FAILS...  SFWA
STRICK, PHILIP  ANTIGRAV: COSMIC COMEDIES BY SF MASTERS  HA
STRICKLAND, BRAD/STRICKLAND, BARBARA CRISIS ON VULCAN [@STAR TREK STARFLEET ACADEMY #1]  LP
STRICKLAND, BRAD MOON DREAMS C/P:2,P
STRICKLAND, BRAD/STRICKLAND, BARBARA NOVA COMMAND [@STAR TREK TNG STARFLEET ACADEMY #9]  LP
STRICKLAND, BRAD  NUL’S QUEST  P
STRICKLAND, BRAD  STAR GHOST, THE [@STAR TREK DS9]  LP
STRICKLAND, BRAD/STRICKLAND, BARBARA STARFALL [@STAR TREK TNG STARFLEET ACADEMY #8]  LP
STRICKLAND, BRAD  STOWAWAYS [STAR TREK DS9]  H,LP
STRICKLAND, BRAD  TO STAND BENEATH THE SUN  P
STRICKLAND, BRAD  WIZARD’S MOLE  P
STRICKLAND, BRAD/STRICKLAND, BARBARA YOU’RE HISTORY [@SABRINA THE TEENAGE WITCH SALEM #3]  P
STRIEBER, WHITLEY  CAT MAGIC  SFWA-P
STRIEBER, WHITLEY  FORBIDDEN ZONE, THE  H
STRIEBER, WHITLEY  HUNGER, THE  C/P,P,H
STRIEBER, WHITLEY  MAJESTIC  H
STRIEBER, WHITLEY/KUNETKA, JAMES  NATURE’S END  H,P
STRIEBER, WHITLEY  NIGHT CHURCH, THE  H
STRIEBER, WHITLEY  TRANSFORMATION: THE BREAKTHROUGH  [C/RAN-P]
STRIEBER, WHITLEY/KUNETKA, JAMES  WARDAY  C/P:2,P
STRIEBER, WHITLEY  WHITLEY STRIEBER’S ALIENS  PA
STRIEBER, WHITLEY  WILD, THE  SFWA-P
STRIEBER, WHITLEY  WOLF OF SHADOWS  C/P,P,H
STRIEBER, WHITLEY  WOLFEN, THE  P
STRIKE, JEREMY  PROMISING PLANET, A  D24100
STRIKER, FRAN  SECRET OF THE LOST MESA, THE  H
STRINGFIELD, LEONARD H.  SITUATION RED: THE UFO SIEGE  SFWA
STROHM, KEITH FRANCIS  BLADESINGER [@FORGOTTEN REALMS FIGHTERS]  P
STROHM, KEITH FRANCIS  TOMB OF HORRORS, THE [@GREYHAWK]  P
STROSS, CHARLES  ACCELERANDO  H
STROSS, CHARLES  ATROCITY ARCHIVES, THE [ATROCITY ARCHIVES 1]  H
STROSS, CHARLES  CLAN CORPORATE, THE [MERCHANT PRINCES 3]  H:2
STROSS, CHARLES  FAMILY TRADE, THE [MERCHANT PRINCES 1]  H,P
STROSS, CHARLES  GLASSHOUSE  H
STROSS, CHARLES  HIDDEN FAMILY, THE [MERCHANT PRINCES 2]  H
STROSS, CHARLES  IRON SUNRISE [SINGULARITY SKY 2]  H
STROSS, CHARLES  JENNIFER MORGUE, THE [ATROCITY ARCHIVES 2]  H:2
STROSS, CHARLES  MISSILE GAP  H
STROSS, CHARLES  SCRATCH MONKEY  H
STROSS, CHARLES  SINGULARITY SKY [SINGULARITY SKY 1]  C/P,P,H
STROSS, CHARLES  TIMELIKE DIPLOMACY [SINGULARITY SKY 1,2]  H
STROSS, CHARLES  TOAST  H
STROUD, JONATHAN  AMULET OF SAMARKAND, THE [@BARTIMAEUS #1]  H
STROUD, JONATHAN  GOLEM’S EYE, THE [@BARTIMAEUS #2]  H
STROUD, JONATHAN  PTOLEMY’S GATE [@BARTIMAEUS #3]  H
STRUGATSKY, ARKADY/STRUGATSKY, BORIS BEETLE IN THE ANTHILL  H
STRUGATSKY, ARKADY/STRUGATSKY, BORIS DEFINITELY MAYBE  C/P:2,P,H
STRUGATSKY, ARKADY FANTASTIKA 1965 VYPUSK II FH
STRUGATSKY, ARKADY/STRUGATSKY, BORIS FAR RAINBOW/SECOND INVASION FROM MARS, THE  C/P,LP,H

The MITSFS Pinkdex  September 24, 2007  Page 373
AUTHOR  TITLE  CODE
STRUGATSKY, ARKADY/STRUGATSKY, BORIS FINAL CIRCLE OF PARADISE, THE  C/P,P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS HARD TO BE A GOD  C/P,SFWA,P,H
STRUGATSKY, ARKADY/STRUGATSKY, BORIS MONDAY BEGINS ON SATURDAY  P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS NOON: 22ND CENTURY  H,LP
STRUGATSKY, ARKADY/STRUGATSKY, BORIS PRISONERS OF POWER  C/P,P,H
STRUGATSKY, ARKADY/STRUGATSKY, BORIS ROADSIDE PICNIC  C/P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS ROADSIDE PICNIC/TALE OF THE TROIKA  H:2,P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS SNAIL ON THE SLOPE, THE  C/P,P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS SPACE APPRENTICE  H:2
STRUGATSKY, ARKADY/STRUGATSKY, BORIS TIME WANDERERS, THE  H:2,P
STRUGATSKY, ARKADY/STRUGATSKY, BORIS UGLY SWANS, THE  H:2,LP
STRUTTON, BILL  DOCTOR WHO AND THE ZARBI [DOCTOR WHO]  SR-P,C/P:2
STRYKER, HAL  NYPD 2025  C/P,P
STUART, IAN  SATAN BUG, THE  H
STUART, L. T.  HOUSE OF THE LIONS, THE  P
STUART, SIDNEY  BEAST WITH THE RED HANDS, THE  P
STUART, W. J.  FORBIDDEN PLANET  C/P:2,P
STURGEON, THEODORE  ALIENS FOUR  C/P:2,P
STURGEON, THEODORE  AMOK TIME [STAR TREK]  SR-S
STURGEON, THEODORE  AND NOW THE NEWS... [COMPLETE STORIES OF THEODORE STURGEON 9] H
STURGEON, THEODORE  BABY IS THREE [COMPLETE STORIES OF THEODORE STURGEON 6]  H
STURGEON, THEODORE  BABY IS THREE/AND MY FEAR IS GREAT  SR-P,C/P
STURGEON, THEODORE  BEYOND  C/P:2,P
STURGEON, THEODORE  CAMINO A CASA, UN  FP
STURGEON, THEODORE  CASE AND THE DREAMER  H:2,P
STURGEON, THEODORE  CAVIAR  SR-H,C/P:2,P
STURGEON, THEODORE  COSMIC RAPE, THE  C/P:2,P
STURGEON, THEODORE  DREAMING JEWELS, THE  SR-H,C/P,LP,H
STURGEON, THEODORE  DREAMING JEWELS, THE/COSMIC RAPE, THE/VENUS PLUS X  H
STURGEON, THEODORE  E PLURIBUS UNICORN  C/P:2,P
STURGEON, THEODORE  GODBODY  H:2,P
STURGEON, THEODORE  GOLDEN HELIX, THE  H:2,P
STURGEON, THEODORE  KILLDOZER! [COMPLETE STORIES OF THEODORE STURGEON 3]  H

STURGEON, THEODORE MAN WHO LOST THE SEA, THE

[COMPLETE STORIES OF THEODORE STURGEON 10]

H

STURGEON, THEODORE  MICROCOSMIC GOD [@COMPLETE STORIES OF THEODORE STURGEON 2]  LP
STURGEON, THEODORE  MORE THAN HUMAN  H:2,SR-P

STURGEON, THEODORE/NINO, ALEX/ MOENCH, DOUG MORE THAN HUMAN: THE GRAPHIC STORY VERSION

CX

STURGEON, THEODORE NAIL AND THE ORACLE, THE

[COMPLETE STORIES OF THEODORE STURGEON 11]

H

STURGEON, THEODORE  NOT WITHOUT SORCERY  C/P:2,P
STURGEON, THEODORE  PERFECT HOST, THE [COMPLETE STORIES OF THEODORE STURGEON 5]  H
STURGEON, THEODORE  RARE BREED, THE  P

STURGEON, THEODORE SAUCER OF LONELINESS, A

[COMPLETE STORIES OF THEODORE STURGEON 7]

H

STURGEON, THEODORE  SELECTED STORIES  LP
STURGEON, THEODORE  SHORE LEAVE [STAR TREK]  SR-S
STURGEON, THEODORE  SOME OF YOUR BLOOD  C/P:2,P
STURGEON, THEODORE  STARS ARE THE STYX, THE  C/P,P,H
STURGEON, THEODORE  STARSHINE  C/P:2,P
STURGEON, THEODORE  STURGEON IN ORBIT  C/P:2,P
STURGEON, THEODORE  STURGEON IS ALIVE AND WELL...  SR-H,H:2,P
STURGEON, THEODORE  SYNTHETIC MAN, THE  C/P:2,P
STURGEON, THEODORE  THUNDER AND ROSES [COMPLETE STORIES OF THEODORE STURGEON 4]  H
STURGEON, THEODORE  TO HERE AND THE EASEL  C/P,P
STURGEON, THEODORE  TO MARRY MEDUSA  H,SFWA-P,P
STURGEON, THEODORE  TOUCH OF STRANGE, A  C/P,P,H

STURGEON, THEODORE ULTIMATE EGOIST, THE

[@COMPLETE STORIES OF THEODORE STURGEON 1]

LP

STURGEON, THEODORE  VENUS PLUS X  C/P:2,P
STURGEON, THEODORE  VISIONS AND VENTURERS  C/P:2,P
STURGEON, THEODORE  VOYAGE TO THE BOTTOM OF THE SEA  C/P:2,P
STURGEON, THEODORE  WAY HOME, A  C/P:2,P
STURGEON, THEODORE  WIDGET, THE WADGET, AND BOFF, THE  SFWA-TD9,TD9
STURGEON, THEODORE  WORLDS OF THEODORE STURGEON, THE  C/P:2,P
STURGIS, SUSANNA J.  MEMORIES AND VISIONS  LPA
STURGIS, SUSANNA J.  WOMEN WHO WALK THROUGH FIRE, THE  LPA

SUBGENIUS FOUNDATION BOOK OF THE SUBGENIUS:

LUNATIC PROPHECIES FOR THE COMING WEIRD TIMES, THE

VLP

SUBIAGA, ROBERT JR.  EYES  LP
SUBOND, VALERIE  HEIGHTS OF HAVENREST, THE  P
SUCHARITKUL, SOMTOW  ALIEN SWORDMASTER, THE [V]  C/P:2,SFWA-P,P
SUCHARITKUL, SOMTOW  AQUILIAD, THE  C/P:2,P
SUCHARITKUL, SOMTOW  DARKLING WIND, THE [INQUESTOR 4]  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 374
AUTHOR  TITLE  CODE

SUCHARITKUL, SOMTOW DAWNING SHADOWS: LIGHT ON THE SOUND, THE (REVISED)

[INQUESTOR 1]

C/P:2,P

SUCHARITKUL, SOMTOW DAWNING SHADOWS: THRONE OF MADNESS, THE (REVISED)

[INQUESTOR 2]

P

SUCHARITKUL, SOMTOW  FALLEN COUNTRY, THE  C/P,P
SUCHARITKUL, SOMTOW  FIRE FROM THE WINE-DARK SEA  LP
SUCHARITKUL, SOMTOW  LIGHT ON THE SOUND [INQUESTOR 1]  C/P,P
SUCHARITKUL, SOMTOW  MALLWORLD [MALLWORLD]  H:2,P
SUCHARITKUL, SOMTOW  STARSHIP AND HAIKU  C/P
SUCHARITKUL, SOMTOW  SYMPHONY OF TERROR [V]  SFWA-P
SUCHARITKUL, SOMTOW  THRONE OF MADNESS, THE [INQUESTOR 2]  C/P:2,P
SUCHARITKUL, SOMTOW  UTOPIA HUNTERS [INQUESTOR 3]  C/P:2,P
SUDAK, EUDICE  X  C/P:2,P
SUDDABY, DONALD  STAR RAIDERS, THE  SR-P
SUGAR, ANDREW  BIO BLITZ [@ENFORCER #5]  P
SUGAR, ANDREW  CALLING DOCTOR KILL! [@ENFORCER #2]  P
SUGAR, ANDREW  ENFORCER, THE [@ENFORCER #1]  P
SUGAR, ANDREW  KILL CITY [@ENFORCER #3]  P
SUGAR, ANDREW  KILL DEADLINE [@ENFORCER #4]  P
SUGAR, ANDREW  STEEL TRAP [@ENFORCER #6]  P
SULLIVAN, C. W. III  WELSH CELTIC MYTH IN MODERN FANTASY  R/REF-H
SULLIVAN, ELEANOR  ELLERY QUEEN’S LOST MEN  SFWA
SULLIVAN, FAITH  MRS. DEMMING AND THE MYTHICAL BEAST  P,SFWA-P

SULLIVAN, JACK ELEGANT NIGHTMARES:

THE ENGLISH GHOST STORY FROM LE FANU TO BLACKWOOD

R/REF-H

SULLIVAN, MIKE  STATION ZERO-ZERO  P
SULLIVAN, SHEILA  CALLING OF BARA, THE  C/P,P
SULLIVAN, STEPHEN D.  DRAGON ISLES, THE [@DRAGONLANCE CROSSROADS]  P

SULLIVAN, STEPHEN D.

DYING KINGDOM, THE [@DRAGONLANCE NEW ADVENTURES #2/ DRAGONLANCE SPELLBINDER QUARTET 2]

LP

SULLIVAN, STEPHEN D.  LION, THE [FIVE RINGS CLAN WAR 7]  P
SULLIVAN, STEPHEN D.  PHOENIX, THE [FIVE RINGS CLAN WAR 4]  P
SULLIVAN, STEPHEN D.  SCORPION, THE [FIVE RINGS CLAN WAR 1]  P
SULLIVAN, STEPHEN D.  WARRIOR’S HEART [@DRAGONLANCE GOODLUND #1]  LP
SULLIVAN, THOMAS  DIAPASON  P
SULLIVAN, TIM  DESTINY’S END  C/P:2,P
SULLIVAN, TIM  FLORIDA PROJECT, THE [V]  C/P,P
SULLIVAN, TIM  LORDS OF CREATION  C/P,P
SULLIVAN, TIM  MARTIAN VIKING, THE  C/P,P
SULLIVAN, TIM  NEW ENGLAND RESISTANCE, THE [V]  C/P,P
SULLIVAN, TIM  PARASITE WAR, THE  C/P,P
SULLIVAN, TIM  TO CONQUER THE THRONE  P,SFWA-P
SULLIVAN, TRICIA  DREAMING IN SMOKE  P
SULLIVAN, TRICIA  LETHE  C/P,P
SULLIVAN, TRICIA  SOMEONE TO WATCH OVER ME  P
SULLIVAN, WALTER  WE ARE NOT ALONE  [C/RAN-P:2]
SUMMERS, IAN  ART OF THE BROTHERS HILDEBRANDT, THE  A
SUMMERS, IAN  TOMORROW AND BEYOND  A
SUMMERS, MONTAGUE  SUPERNATURAL OMNIBUS: BEST GHOST STORIES, THE  HA
SUMNER, M. C.  SUBSTITUTE, THE  P
SUMNER, MARK  DEVIL’S ENGINE [DEVIL’S WEST 2]  P
SUMNER, MARK  DEVIL’S TOWER [DEVIL’S WEST 1]  C/P,P
SUMNER, MARK  INSANITY, ILLINOIS [@NEWS FROM THE EDGE #2]  P
SUMNER, MARK  MONSTER OF MINNESOTA, THE [@NEWS FROM THE EDGE 1]  P
SUMNER, MARK  PRODIGAL SORCERER, THE [MAGIC: THE GATHERING]  P
SUMNER, MARK  VAMPIRES OF VERMONT [@NEWS FROM THE EDGE 3]  P
SUNBORD, TERRY  NEBULA EXPRESS  LP
SUNBORD, TERRY  ROBINSON CRUSOE 1,000,000 A.D.  LP
SUNDERMAN, JAMES F.  AIR ESCAPE AND EVASION  [C/RAN-H]
SUNDMAN, JOHN COMPTON  CHEAP COMPLEX DEVICES  HA,LPA
SUNDMAN, JOHN F. X.  ACTS OF THE APOSTLES  H:2,LP
SUNDOWN, WILL  POCKETS OF RESISTANCE  P
SUNNY  MONA LISA AWAKENING [MONERE]  LP
SUSANN, JACQUELINE  YARGO  H,P
SUSKIND, PATRICK  PERFUME  P
SUSSEX, LUCY  ABSOLUTE UNCERTAINTY  LP
SUTCLIFF, ROSEMARY  SWORD AT SUNSET  SFWA-P
SUTHERLAND, JAMES  STORMTRACK  C/P:2,P
SUTHERLAND, TUI T.  SO THIS IS HOW IT ENDS [@AVATARS #1]  H
SUTTON, DAVID/JONES, STEPHEN  DARK VOICES 4  SR-PA
SUTTON, HENRY  SACRIFICE, THE  P
SUTTON, HENRY  VECTOR  C/P,P
SUTTON, JEAN/SUTTON, JEFF  LORD OF THE STARS  H
SUTTON, JEFF  ALTON’S UNGUESSABLE  D76096

The MITSFS Pinkdex  September 24, 2007  Page 375
AUTHOR  TITLE  CODE
SUTTON, JEFF  APOLLO AT GO  H:2,P
SUTTON, JEFF  ATOM CONSPIRACY, THE  C/P,P,H
SUTTON, JEFF  BOMBS IN ORBIT  C/P:2,P
SUTTON, JEFF  FIRST ON THE MOON  C/P:2,P
SUTTON, JEFF  H-BOMB OVER AMERICA  P
SUTTON, JEFF  MAN WHO SAW TOMORROW, THE  DH095
SUTTON, JEFF  MINDBLOCKED MAN, THE  C/P:2,P
SUTTON, JEFF  MISSILE LORDS, THE  P
SUTTON, JEFF  SPACEHIVE  C/P:2,P
SUTTON, JEFF  WHISPER FROM THE STARS  C/P,P
SUTTON, LEE  VENUS BOY  H
SUTTON, STEPHEN P.  MORE TALES TO TREMBLE BY  HA
SUTTON, STEPHEN P.  TALES TO TREMBLE BY  HA
SUVIN, DARKO  METAMORPHOSES OF SCIENCE FICTION  R/REF-H

SUVIN, DARKO OTHER WORLDS, OTHER SEAS:

SCIENCE FICTION STORIES FROM SOCIALIST COUNTRIES

HA,PA

SUVIN, DARKO  RUSSIAN SCIENCE FICTION 1956–1974: A BIBLIOGRAPHY  R/IND
SUYIN, HAN  ENCHANTRESS, THE  P
SWAIN, DWIGHT V.  TRANSPOSED MAN, THE  DD113
SWAINSTON, STEPH  YEAR OF OUR WAR, THE [FOURLANDS 1]  H
SWALLOW, JAMES  BLOOD RELATIVE [@ROGUE TROOPER]  P
SWALLOW, JAMES  DEUS ENCARMINE [@WARHAMMER 40K BLOOD ANGELS 1]  P
SWALLOW, JAMES  DEUS SANGUINIUS [@WARHAMMER 40K BLOOD ANGELS 2]  P
SWALLOW, JAMES  FAITH & FIRE [@WARHAMMER 40K SISTERS OF BATTLE 1]  P
SWALLOW, JAMES  FLIGHT OF THE EISENSTEIN, THE [@WARHAMMER 40K HORUS HERESY 4]  P
SWANN, INGO  STAR FIRE  C/P,P
SWANN, S. ANDREW  BROKEN CRESCENT  P
SWANN, S. ANDREW  DRAGONS OF THE CUYAHOGA, THE [DRAGONS OF THE CUYAHOGA 1]  C/P,P
SWANN, S. ANDREW  DWARVES OF WHISKEY ISLAND [DRAGONS OF THE CUYAHOGA 2]  P
SWANN, S. ANDREW  EMPERORS OF THE TWILIGHT [MOREAU 2]  C/P:2,P
SWANN, S. ANDREW  FEARFUL SYMMETRIES [MOREAU 4]  P
SWANN, S. ANDREW  FORESTS OF THE NIGHT [MOREAU 1]  C/P,SFWA-P,P
SWANN, S. ANDREW  GOD’S DICE  P
SWANN, S. ANDREW  PARTISAN [@HOSTILE TAKEOVER #2]  P
SWANN, S. ANDREW  PROFITEER [@HOSTILE TAKEOVER #1]  P
SWANN, S. ANDREW  REVOLUTIONARY [@HOSTILE TAKEOVER #3]  C/P,P
SWANN, S. ANDREW  SPECTERS OF THE DAWN [MOREAU 3]  C/P:2,P
SWANN, THOMAS BURNETT  CRY SILVER BELLS  C/P:2,P
SWANN, THOMAS BURNETT  DAY OF THE MINOTAUR  C/P:2,P
SWANN, THOMAS BURNETT  DOLPHIN AND THE DEEP, THE  C/P:2,P
SWANN, THOMAS BURNETT  FOREST OF FOREVER, THE  C/P:2,P
SWANN, THOMAS BURNETT  GOAT WITHOUT HORNS, THE  C/P:2,P
SWANN, THOMAS BURNETT  GODS ABIDE, THE  C/P:2,P
SWANN, THOMAS BURNETT  GREEN PHOENIX  C/P:2,P
SWANN, THOMAS BURNETT  HOW ARE THE MIGHTY FALLEN  C/P:2,P
SWANN, THOMAS BURNETT  LADY OF THE BEES  C/P:2,P
SWANN, THOMAS BURNETT  MINIKINS OF YAM, THE  C/P:2,P
SWANN, THOMAS BURNETT  MOONDUST  C/P:2,P
SWANN, THOMAS BURNETT  NOT-WORLD, THE  C/P:2,P
SWANN, THOMAS BURNETT  QUEENS WALK IN THE DUSK  SR-H
SWANN, THOMAS BURNETT  TOURNAMENT OF THORNS, THE  C/P:2,P
SWANN, THOMAS BURNETT  WEIRWOODS, THE  C/P:2,P
SWANN, THOMAS BURNETT  WHERE IS THE BIRD OF FIRE?  C/P:2,P
SWANN, THOMAS BURNETT  WILL-O-THE-WISP  C/P,P
SWANN, THOMAS BURNETT  WOLFWINTER  C/P:2,P
SWANWICK, MICHAEL  BEING GARDENER DOZOIS  R/REF-H
SWANWICK, MICHAEL  BONES OF THE EARTH  H,P
SWANWICK, MICHAEL  BONES OF THE EARTH (PROOF)  SR-LP
SWANWICK, MICHAEL  GRAVITY’S ANGELS  LP
SWANWICK, MICHAEL  GRIFFIN’S EGG  H:2,LP
SWANWICK, MICHAEL  IN THE DRIFT  C/P:2,P
SWANWICK, MICHAEL  IRON DRAGON’S DAUGHTER, THE  H:2,P
SWANWICK, MICHAEL  JACK FAUST  H:2,LP
SWANWICK, MICHAEL  MOON DOGS  SR-H,H
SWANWICK, MICHAEL  STATIONS OF THE TIDE  C/P,P,H
SWANWICK, MICHAEL  TALES OF OLD EARTH  H
SWANWICK, MICHAEL  VACUUM FLOWERS  C/P,SFWA,P,H:2
SWEENEY, JOYCE  SHADOW  H
SWENDSON, SHANNA  DAMSEL UNDER STRESS [KATIE CHANDLER 3]  LP
SWENDSON, SHANNA  ENCHANTED, INC. [KATIE CHANDLER 1]  LP
SWENDSON, SHANNA  ONCE UPON STILETTOS [KATIE CHANDLER 2]  LP
SWIASKI, PETER  STANISAW LEM READER, A  R/REF-H
SWIFT, JONATHAN  ANNOTATED GULLIVER’S TRAVELS, THE  H
SWIFT, JONATHAN  GULLIVER’S TRAVELS  C/P:2,SR-L,P
SWIGART, LESLIE KAY  HARLAN ELLISON: A BIBLIOGRAPHICAL CHECKLIST  R/IND

AUTHOR  TITLE  CODE
SWIGART, ROB  A.K.A.: A COSMIC FABLE  LP
SWIGART, ROB  BOOK OF REVELATIONS, THE  H
SWIGART, ROB  PORTAL  H
SWIGART, ROB  TIME TRIP, THE  H
SWINFEN, ANN  IN DEFENCE OF FANTASY  R/REF-P
SWINIARSKI, S. A.  BLOOD & RUST [CLEVELAND UNDEAD 1,2]  P
SWINIARSKI, S. A.  RAVEN [CLEVELAND UNDEAD 1]  P
SWYCAFFER, JEFFERSON P.  BECOME THE HUNTED  P
SWYCAFFER, JEFFERSON P.  EMPIRE’S LEGACY, THE [@CONCORDAT #1]  C/P,P
SWYCAFFER, JEFFERSON P.  NOT IN OUR STARS  C/P,P
SWYCAFFER, JEFFERSON P.  PRAESIDIUM OF ARCHIVE, THE  P
SWYCAFFER, JEFFERSON P.  REVOLT AND REBIRTH [@CONCORDAT #3]  P
SWYCAFFER, JEFFERSON P.  UNIVERSAL PREY, THE  P
SWYCAFFER, JEFFERSON P.  VOYAGE OF THE PLANETSLAYER [@CONCORDAT #2]  C/P:2,P
SWYCAFFER, JEFFERSON P.  WARSPRITE  P
SWYCAFFER, JEFFERSON P.  WEB OF FUTURES  C/P,P
SYKES, S. C.  RED GENESIS [NEXT WAVE 1]  C/P:2,P
SYNGE, URSULA  SWAN’S WING  C/P:2,P
SYVERTSEN, RYDER  DARKNESS DESCENDS [@MYSTIC REBEL]  P
SZILARD, LEO  VOICE OF THE DOLPHINS AND OTHER STORIES, THE  LP
SZILARD, LEO  VOICE OF THE DOLPHINS AND OTHER STORIES, THE (BRITISH)  SR-P
SZYDLOW, JARL  ARK, THE  P

AUTHOR  TITLE  CODE
TABLER, JOSEPH  CAPITOL HILL CLONES  P
TABLER, JOSEPH  MICROWAVE CAPER, THE  P
TABLER, MADELINE  FLIP SIDE OF MEMORY  P
TABORI, PAUL  CLEFT, THE  C/P:2,P
TABORI, PAUL  DEMONS OF SANDORRA, THE  P
TABORI, PAUL  GREEN RAIN, THE  C/P:2,P
TABORI, PAUL  TORTURE MACHINE, THE  C/P,P
TAFF, JOHN F. D.  CLOUD GATHERER, THE [HARAN OF OAS 1]  LP
TAFF, JOHN F. D.  HEIR TO THE SAND [HARAN OF OAS 2]  LP
TAGGERT, BRIAN  SANCTION, THE [V]  SR-S
TAINE, JOHN  BEFORE THE DAWN  SR-H
TAINE, JOHN  COSMIC GEOIDS, THE  SR-H
TAINE, JOHN  CRYSTAL HORDE, THE  SR-H
TAINE, JOHN  FORBIDDEN GARDEN, THE  SR-H,H:2
TAINE, JOHN  G.O.G. 666  SR-H
TAINE, JOHN  GOLD TOOTH, THE  SR-H
TAINE, JOHN  GREATEST ADVENTURE, THE  C/P:2,P
TAINE, JOHN  GREEN FIRE  SR-H
TAINE, JOHN  IRON STAR, THE  SR-H,LP
TAINE, JOHN  ORDA DI CRISTALLO, L  FP
TAINE, JOHN  PURPLE SAPPHIRE, THE  SR-H
TAINE, JOHN  SEEDS OF LIFE  SR-H,C/P:2,GN3.13
TAINE, JOHN  SEEDS OF LIFE/WHITE LILY  H:2,LP
TAINE, JOHN  TIME STREAM, THE  SR-H
TAINE, JOHN  TIME STREAM, THE/GREATEST ADVENTURE, THE/PURPLE SAPPHIRE, THE H:2,LP
TAKAHASHI, RUMIKO  INU-YASHA  CX
TAKEI, GEORGE/ASPRIN, ROBERT LYNN  MIRROR FRIEND, MIRROR FOE  SR-P,C/P
TAKEI, GEORGE  TO THE STARS  R/REF-H,C/REF-H
TALBOT, MICHAEL  BOG, THE  H
TALL, STEPHEN  PEOPLE BEYOND THE WALL, THE  C/P,P
TALL, STEPHEN  RAMSGATE PARADOX, THE  C/P,P
TALL, STEPHEN  STARDUST VOYAGES, THE  C/P:2,P
TALLIS, ROBYN  FIRE IN THE SKY [@PLANET BUILDERS #10]  P
TALLIS, ROBYN  MOUNTAIN OF STOLEN DREAMS [@PLANET BUILDERS #1]  C/P,P
TALLIS, ROBYN  NIGHT OF GHOSTS AND LIGHTNING [@PLANET BUILDERS #2]  C/P,P
TALLIS, ROBYN  NIGHT OF TWO NEW MOONS [@PLANET BUILDERS #6]  P
TALLIS, ROBYN  REBEL FROM ALPHORION [@PLANET BUILDERS #3]  P
TALLIS, ROBYN  ZERO-SUM GAMES [@PLANET BUILDERS #5]  P
TALMADGE-BICKMORE, DEBORAH  APPRENTICE, THE  P
TALMADGE-BICKMORE, DEBORAH  HELDAN, THE  P
TAN, CECILIA  FORGED BONDS  SR-L
TAN, CECILIA  GENDERFLEX  LPA
TAN, CECILIA  SEX MAGICK  SR-L
TAN, CECILIA  SEXTOPIA  SR-L
TAN, CECILIA  TECHNOSEX  SR-L
TAN, CECILIA  TELEPATHS DON’T NEED SAFEWORDS  SR-L
TAN, CECILIA  WIRED HARD  SR-L
TAN, CECILIA  WORLDS OF WOMEN  SR-L
TANDEM BOOKS  MINDWORM, THE  SR-PA
TANKERSLEY, RICHIE  ANGEL CHRONICLES, THE, VOL. 2 [BUFFY]  C/P,P
TANNEHILL, JAYNE  OREGON INVASION, THE [V]  C/P,SFWA-P,P
TARABILDA, JAMES  BLACK SUN  P
TARDI, JACQUES  ADELE & THE BEAST  CX
TARDI, JACQUES  DEMON OF THE EIFFEL TOWER, THE  CX
TARR, JUDITH  ALAMUT [ALAMUT 1]  C/P,LP
TARR, JUDITH  ARROWS OF THE SUN [@AVARYAN RISING #4]  H:2,P
TARR, JUDITH  ARS MAGICA  H:2,P
TARR, JUDITH  AVARYAN RESPLENDENT [AVARYAN RISING 4,5,6]  LP
TARR, JUDITH  DAGGER AND THE CROSS, THE [ALAMUT 2]  H,P
TARR, JUDITH  DAUGHTER OF ZIR [HORSE GODDESS PEOPLE 2]  LP
TARR, JUDITH  DEVIL’S BARGAIN  LP
TARR, JUDITH  FALL OF PRINCES, A [@AVARYAN RISING #3]  C/P,SFWA-P,P,H
TARR, JUDITH  GOLDEN HORN, THE [HOUND AND FALCON #2]  C/P:2,@H,SFWA-P,P
TARR, JUDITH  HALL OF THE MOUNTAIN KING, THE [@AVARYAN RISING #1]  H,SFWA,SFWA-P,P
TARR, JUDITH  HOUND AND THE FALCON, THE [HOUND AND FALCON 1,2,3]  H:2
TARR, JUDITH  HOUNDS OF GOD, THE [HOUND AND FALCON #3]  P,@H,SFWA-P
TARR, JUDITH  HOUSE OF WAR  LP
TARR, JUDITH/TURTLEDOVE, HARRY  HOUSEHOLD GODS  C/P,P
TARR, JUDITH  ISLE OF GLASS, THE [HOUND AND FALCON #1]  C/P,@H,SFWA-P,P
TARR, JUDITH  KINGDOM OF THE GRAIL  LP
TARR, JUDITH  KING’S BLOOD  LP
TARR, JUDITH  LADY OF HAN-GILEN, THE [@AVARYAN RISING #2]  H,SFWA-P,P
TARR, JUDITH  LADY OF HORSES [HORSE GODDESS PEOPLE]  LP
TARR, JUDITH  LORD OF THE TWO LANDS  H

The MITSFS Pinkdex  September 24, 2007  Page 378
AUTHOR  TITLE  CODE
TARR, JUDITH  PRIDE OF KINGS  LP
TARR, JUDITH  QUEEN OF SWORDS  H
TARR, JUDITH  QUEEN OF THE AMAZONS  H:2,LP
TARR, JUDITH  RITE OF CONQUEST  P
TARR, JUDITH  SHEPHERD KINGS, THE  H
TARR, JUDITH  SPEAR OF HEAVEN [@AVARYAN RISING #5]  P
TARR, JUDITH  THRONE OF ISIS  P
TARR, JUDITH  WHITE MARE’S DAUGHTER [HORSE GODDESS PEOPLE 1]  H
TARR, JUDITH  WIND IN CAIRO, A  C/P,P
TATE, PETER  COUNTRY LOVE AND POISON RAIN  H:2,SFWA
TATE, PETER  FACES IN THE FLAMES  H,SFWA
TATE, PETER  GARDENS ONE TO FIVE  H,SFWA
TATE, PETER  GREENCOMBER  H,SFWA
TATE, PETER  MOON ON AN IRON MEADOW  H,SFWA
TATE, PETER  SEAGULLS UNDER GLASS  H
TATE, PETER  THINKING SEAT, THE  C/P,P
TAVERNA PRESS  SHIPBUILDING  LPA
TAYLOR, BERNARD  GODSEND, THE  P
TAYLOR, BERNARD  REAPING, THE  P
TAYLOR, FRANK  HOUSE OF THE HUNTER  SR-P
TAYLOR, G. P.  SHADOWMANCER  H
TAYLOR, G. P.  WORMWOOD (PROOF)  SR-LP
TAYLOR, JANELLE  MOONDUST AND MADNESS  P
TAYLOR, JERI  MOSAIC [@STAR TREK VOYAGER]  C/P,P
TAYLOR, JERI  PATHWAYS [@STAR TREK VOYAGER]  C/P,P
TAYLOR, JERI  UNIFICATION [STAR TREK TNG]  C/P:3,P
TAYLOR, JOHN G.  BLACK HOLES: THE END OF THE UNIVERSE?  SFWA
TAYLOR, KAREN E.  BITTER BLOOD [VAMPIRE LEGACY 2]  P
TAYLOR, KAREN E.  BLOOD SECRETS [VAMPIRE LEGACY 1]  P
TAYLOR, KAREN E.  RESURRECTION [VAMPIRE LEGACY]  C/P,P
TAYLOR, KEITH  BARD [@BARD #1]  C/P:2,P
TAYLOR, KEITH  BARD II [@BARD #2]  C/P,P
TAYLOR, KEITH  CAULDRON OF PLENTY, THE [DANANS 2]  C/P,P
TAYLOR, KEITH  LANCES OF NENGESDUL  SR-P
TAYLOR, KEITH  RAVENS’ GATHERING [@BARD #4]  P
TAYLOR, KEITH  SEARCH FOR THE STARBLADE [DANANS 3]  P
TAYLOR, KEITH  SORCERER’S SACRED ISLE, THE [DANANS 1]  C/P,P
TAYLOR, KEITH  WILD SEA, THE [@BARD #3]  P
TAYLOR, L. A.  BLOSSOM OF ERDA, THE  H,SFWA,P
TAYLOR, L. A.  CAT’S PAW  C/P,P
TAYLOR, ROBERT LEWIS  ADRIFT IN A BONEYARD  C/P,P,H
TAYLOR, SAMUEL W.  HEAVEN KNOWS WHY  H
TAYLOR, TRAVIS S.  QUANTUM CONNECTION, THE  H
TAYLOR, TRAVIS S.  WARP SPEED  H

TEDESCHI, ALBERTO/ FANTONETTI, CARLO ROSSI MONDADORI’S POCKET ITALIAN-ENGLISH ENGLISH-ITALIAN DICTIONARY

DICT

TEDFORD, WILLIAM  HYDRABYSS RED [@TIMEQUEST #2]  P
TEDFORD, WILLIAM  LIQUID DIET  P
TEDFORD, WILLIAM  NEMYDIA DEEP [@TIMEQUEST #3]  P
TEDFORD, WILLIAM  RASHANYN DARK [@TIMEQUEST #1]  C/P,P
TEDFORD, WILLIAM  SILENT GALAXY  P
TEIXEIRA, KEVIN  VIRTUAL SOUL, A  P
TELEP, PETER  DEMOLITION WINTER [SPACE: ABOVE AND BEYOND 2]  C/P,P
TELEP, PETER  DESCENT [@DESCENT 1]  P
TELEP, PETER  EQUINOX [@DESCENT 3]  P
TELEP, PETER  PILGRIM STARS [@WING COMMANDER]  P
TELEP, PETER  RED PLANET  P
TELEP, PETER  SPACE: ABOVE AND BEYOND [SPACE: ABOVE AND BEYOND 1]  P
TELEP, PETER  SQUIRE [SQUIRE 1]  C/P,P
TELEP, PETER  SQUIRE’S BLOOD [SQUIRE 2]  P
TELEP, PETER  SQUIRE’S HONOR [SQUIRE 3]  P
TELEP, PETER  STEALING THUNDER [@DESCENT 2]  P
TELEP, PETER  WING COMMANDER [WING COMMANDER]  P
TEM, MELANIE  REVENANT, THE  P
TEM, MELANIE  WILDING, THE (PROOF)  SR-LP
TEM, STEVE RASNIC  HIGH FANTASTIC  HA
TEMPEST BOOKS  WORLDS AT WAR  SR-PA
TEMPLE, WILLIAM F.  AUTOMATED GOLIATH, THE  DF129
TEMPLE, WILLIAM F.  BATTLE ON VENUS  D76380,DF195
TEMPLE, WILLIAM F.  FLESHPOTS OF SANSATO, THE  H
TEMPLE, WILLIAM F.  FOUR-SIDED TRIANGLE  GN2.09,SR-H
TEMPLE, WILLIAM F.  SHOOT AT THE MOON  C/P:2,P,H:2
TEMPLE, WILLIAM F.  THREE SUNS OF AMARA, THE  D76380,DF129
TENN, WILLIAM  CHILDREN OF WONDER  HA

AUTHOR  TITLE  CODE
TENN, WILLIAM  DANCING NAKED  H
TENN, WILLIAM  HERE COMES CIVILIZATION [COMPLETE TENN 2]  H
TENN, WILLIAM  HUMAN ANGLE, THE  C/P:2,P
TENN, WILLIAM  IMMODEST PROPOSALS [COMPLETE TENN 1]  H
TENN, WILLIAM  LAMP FOR MEDUSA, A  SR-DB0077
TENN, WILLIAM  OF ALL POSSIBLE WORLDS  C/P,P,H
TENN, WILLIAM  OF MEN AND MONSTERS  C/P:2,P
TENN, WILLIAM  OUTSIDERS  C/PA,PA
TENN, WILLIAM  SEVEN SEXES, THE  C/P:2,P
TENN, WILLIAM  SQUARE ROOT OF MAN, THE  C/P:2,P
TENN, WILLIAM  TIME IN ADVANCE  C/P:2,P
TENN, WILLIAM  WOODEN STAR, THE  C/P:2,P
TEPPER, SHERI S.  AFTER LONG SILENCE  C/P:2,P,H
TEPPER, SHERI S.  AWAKENERS, THE [AWAKENERS 1,2]  H:2,LP
TEPPER, SHERI S.  BEAUTY  C/P,P,LP,H
TEPPER, SHERI S.  BLOOD HERITAGE  C/P:2,P
TEPPER, SHERI S.  BONES, THE  P,SFWA-P
TEPPER, SHERI S.  COMPANIONS, THE  H,P
TEPPER, SHERI S.  DERVISH DAUGHTER [TRUE GAME 5]  C/P,SFWA-P,P
TEPPER, SHERI S.  END OF THE GAME, THE [TRUE GAME 4,5,6]  H
TEPPER, SHERI S.  FAMILY TREE, THE  C/P:2,LP
TEPPER, SHERI S.  FLIGHT OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 2]  C/P:2,P
TEPPER, SHERI S.  FRESCO, THE  H
TEPPER, SHERI S.  GATE TO WOMEN’S COUNTRY, THE  C/P,SR-LP,P,H:2
TEPPER, SHERI S.  GIBBON’S DECLINE AND FALL  C/P,P,H
TEPPER, SHERI S.  GRASS  C/P,P,H,SFWA
TEPPER, SHERI S.  JINIAN FOOTSEER [TRUE GAME 4]  C/P,SFWA-P,P
TEPPER, SHERI S.  JINIAN STAR-EYE [TRUE GAME 6]  C/P,SFWA-P,P
TEPPER, SHERI S.  KING’S BLOOD FOUR [TRUE GAME 1]  C/P:2,P
TEPPER, SHERI S.  MARGARETS, THE  H
TEPPER, SHERI S.  MARIANNE, THE MADAME, AND THE MOMENTARY GODS [MARIANNE 2]  C/P:2,P
TEPPER, SHERI S.  MARIANNE, THE MAGUS, AND THE MANTICORE [MARIANNE 1]  C/P:2,P
TEPPER, SHERI S.  MARIANNE, THE MATCHBOX, AND THE MALACHITE MOUSE [MARIANNE 3] C/P,P
TEPPER, SHERI S.  NECROMANCER NINE [TRUE GAME 2]  C/P:2,P
TEPPER, SHERI S.  NORTHSHORE [@AWAKENERS #1]  C/P,SFWA,SFWA-P,P
TEPPER, SHERI S.  PLAGUE OF ANGELS, A  C/P:2,P,H
TEPPER, SHERI S.  RAISING THE STONES  H:2,P
TEPPER, SHERI S.  REVENANTS, THE  C/P:2,P
TEPPER, SHERI S.  SEARCH OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 3]  C/P:2,P
TEPPER, SHERI S.  SHADOW’S END  H:2,P
TEPPER, SHERI S.  SIDESHOW  C/P,P,H
TEPPER, SHERI S.  SINGER FROM THE SEA  H
TEPPER, SHERI S.  SIX MOON DANCE  C/P,P,H:2
TEPPER, SHERI S.  SONG OF MAVIN MANYSHAPED, THE [MAVIN MANYSHAPED 1]  C/P:2,P
TEPPER, SHERI S.  SOUTHSHORE [@AWAKENERS #2]  C/P:2,SFWA-P,P,H
TEPPER, SHERI S.  TRUE GAME, THE [TRUE GAME 1,2,3]  H:2
TEPPER, SHERI S.  VISITOR, THE  H
TEPPER, SHERI S.  WIZARD’S ELEVEN [TRUE GAME 3]  C/P,P
TERMAN, DOUGLAS  BY BALLOON TO THE SAHARA  P
TERMAN, DOUGLAS  FREE FLIGHT  P
TERRY, MELVIN R.  GLIMPSE OF THE PLANET ZINN, A  H
TESSIER, THOMAS  NIGHTWALKER, THE  P
TESTA, DOM  COMET’S CURSE, THE [GALAHAD 1]  LP
TEVIS, WALTER  FAR FROM HOME  H
TEVIS, WALTER  MAN WHO FELL TO EARTH, THE  C/P:2,LP,P
TEVIS, WALTER  MOCKINGBIRD  H,P
TEVIS, WALTER  QUEEN’S GAMBIT, THE  LP
TEVIS, WALTER  STEPS OF THE SUN, THE  C/P,SFWA,SFWA-P,P,H
TEY, JOSEPHINE  SINGING SANDS, THE  P
THACKERAY, W. M.  ROSE AND THE RING, THE  SR-P
THANE, ELSWYTH  TRYST  P
THANET, NEIL  BEYOND THE VEIL  SR-P
THANET, NEIL  MAN WHO CAME BACK, THE  SR-P
THAYER, TIFFANY  ONE-MAN SHOW  H
THELS, H. S.  !S.O.S. PLUTON!  FP
THEOBALD, ROBERT/SCOTT, J. M.  TEG’S 1994  C/P,P
THEROUX, PAUL  O-ZONE  H,P
THEYDON, JOHN  STINGRAY  SR-P
THOMAS, CRAIG  FIREFOX  P
THOMAS, CRAIG  FIREFOX DOWN!  P
THOMAS, CRAIG  WILDCAT  P
THOMAS, CRAIG  WINTER HAWK  P
THOMAS, DAN  SEED, THE  C/P:2,P
THOMAS, ELIZABETH MARSHALL  ANIMAL WIFE, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 380
AUTHOR  TITLE  CODE
THOMAS, ELIZABETH MARSHALL  REINDEER MOON  H,P
THOMAS, G. W.  BOOK OF THE BLACK SUN, THE  LP
THOMAS, JEFFREY  DEADSTOCK [PUNKTOWN]  P
THOMAS, JEFFREY  MONSTROCITY  LP
THOMAS, JEFFREY  PUNKTOWN (REVISED) [PUNKTOWN]  H
THOMAS, MAGGY  BROKEN TIME  P
THOMAS, MARTIN  BEYOND THE SPECTRUM  C/P:2,P
THOMAS, QUENTIN  CHAINS OF LIGHT  P,SFWA-P
THOMAS, ROY/CHAYKIN, HOWARD  STAR WARS [STAR WARS]  SCX
THOMAS, SHEREE R.  DARK MATTER  LPA
THOMAS, SUE  CORRESPONDENCE  H,LP
THOMAS, THEODORE L./WILHELM, KATE  CLONE, THE  C/P:2,P
THOMAS, THEODORE L./WILHELM, KATE  YEAR OF THE CLOUD, THE  H,P
THOMAS, THOMAS T.  CRYGENDER  C/P,SFWA-P,P
THOMAS, THOMAS T.  FIRST CITIZEN  C/P,SFWA-P,P
THOMAS, THOMAS T.  ME  C/P:2,P
THOMPSETT, NEIL LEE  BECOMING HUMAN (PROOF)  SR-P
THOMPSON, ALLYN  AZRIEL UPRISING, THE  P
THOMPSON, DONALD  ANCIENT ENEMY, THE  P
THOMPSON, DONALD  NEXT ENCOUNTER, THE  C/P,P
THOMPSON, E. P.  SYKAOS PAPERS, THE  H
THOMPSON, ELDON  CRIMSON SWORD, THE [ASAHIEL 1]  P
THOMPSON, ELDON  OBSIDIAN KEY, THE [ASAHIEL 2]  H
THOMPSON, G. RICHARD  ROMANTIC GOTHIC TALES 1790-1840  PA
THOMPSON, GENE  LUPE  P
THOMPSON, HARLAN  SILENT RUNNING  SR-P
THOMPSON, HARRISON R.  SEVENTH SENSE  P
THOMPSON, JOYCE  BLUE CHAIR, THE  C/P,P
THOMPSON, JOYCE  CONSCIENCE PLACE  P,SFWA
THOMPSON, JOYCE  HARRY AND THE HENDERSONS  P
THOMPSON, PAUL B./COOK, TONYA C.  ALLIANCES [@DRAGONLANCE ELVEN EXILES #2]  P
THOMPSON, PAUL B./COOK, TONYA C.  BROTHER OF THE DRAGON [@DRAGONLANCE BARBARIANS #2]  P
THOMPSON, PAUL B./COOK, TONYA C.  CHILDREN OF THE PLAINS [@DRAGONLANCE BARBARIANS #1]  P
THOMPSON, PAUL B./COOK, TONYA C.  DARGONESTI, THE [@DRAGONLANCE LOST HISTORIES #3]  P
THOMPSON, PAUL B./CARTER, TONYA R.  DARKNESS & LIGHT [DRAGONLANCE PRELUDES 1]  C/P:2,P
THOMPSON, PAUL B./CARTER, TONYA R.  FIRSTBORN [@DRAGONLANCE ELVEN NATIONS #1]  P
THOMPSON, PAUL B./COOK, TONYA C.  HERO’S JUSTICE, A [@DRAGONLANCE ERGOTH #3]  P
THOMPSON, PAUL B.  MIDDLE OF NOWHERE, THE [@DRAGONLANCE CROSSROADS]  P
THOMPSON, PAUL B.  NEMESIS [@MAGIC: THE GATHERING MASQUERADE 2]  P
THOMPSON, PAUL B./CARTER, TONYA R.  QUALINESTI, THE [@DRAGONLANCE ELVEN NATIONS #3]  P
THOMPSON, PAUL B./CARTER, TONYA R.  RED SANDS  P
THOMPSON, PAUL B./CARTER, TONYA R.  RIVERWIND THE PLAINSMAN [DRAGONLANCE PRELUDES II 1]  C/P,P
THOMPSON, PAUL B./COOK, TONYA C.  SANCTUARY [@DRAGONLANCE ELVEN EXILES #1]  P
THOMPSON, PAUL B./COOK, TONYA C.  SISTER OF THE SWORD [@DRAGONLANCE BARBARIANS #3]  P
THOMPSON, PAUL B.  SUNDIPPER  P
THOMPSON, PAUL B.  THORN AND NEEDLE  P
THOMPSON, PAUL B./COOK, TONYA C.  WARRIOR’S JOURNEY, A [@DRAGONLANCE ERGOTH #1]  C/P,P
THOMPSON, PAUL B./COOK, TONYA C.  WIZARD’S FATE, THE [@DRAGONLANCE ERGOTH #2]  P
THOMPSON, RUTH PLUMLY  PURPLE PRINCE OF OZ, THE  H
THOMPSON, V. M.  PROJECT GOD  P
THOMPSON, W. R.  DEBTORS’ PLANET [STAR TREK TNG]  C/P:2,P
THOMPSON, W. R.  INFILTRATOR [STAR TREK TNG]  C/P,P
THOMPSON, W. R.  SIDESHOW  C/P:2,SFWA-P,P
THOMPSON, WILLIAM IRWIN  ISLANDS OUT OF TIME  H,SFWA

THOMSEN, BRIAN M./ GREENBERG, MARTIN H.

ALTERNATE GETTYSBURGS

HA

THOMSEN, BRIAN M.  AMERICAN FANTASY TRADITION, THE  LPA
THOMSEN, BRIAN M.  MAGE IN THE IRON MASK, THE [@FORGOTTEN REALMS NOBLES #4]  P

THOMSEN, BRIAN M./ GREENBERG, MARTIN H.

MOB MAGIC

PA

THOMSEN, BRIAN M.  NOVEL IDEAS: FANTASY  PA

THOMSEN, BRIAN M./ GREENBERG, MARTIN H.

OCEANS OF MAGIC

C/PA,PA

THOMSEN, BRIAN M./ GREENBERG, MARTIN H.

OCEANS OF SPACE

PA

THOMSEN, BRIAN M.  ONCE AROUND THE REALMS [@FORGOTTEN REALMS]  P
THOMSEN, BRIAN M./KING, J. ROBERT  REALMS OF MAGIC [@FORGOTTEN REALMS]  PA
THOMSEN, BRIAN M.  REALMS OF THE ARCANE [@FORGOTTEN REALMS]  PA

THOMSEN, BRIAN M./ GREENBERG, MARTIN H.

REPENTANT, THE

PA

THOMSEN, BRIAN M.  TALES OF RAVENLOFT [RAVENLOFT]  C/PA:2,PA
THOMSEN, BRIAN M.  YULETIDE UNIVERSE, A (PROOF)  SR-LPA
THOMSON, AMY  COLOR OF DISTANCE, THE  LP
THOMSON, AMY  STORYTELLER  LP

AUTHOR  TITLE  CODE
THOMSON, AMY  THROUGH ALIEN EYES  C/P,LP
THOMSON, AMY  VIRTUAL GIRL  C/P,P
THOMSON, K. GRAHAM  PEOPLE OF THE SOUTH POLE  H
THORNE, IAN  GODZILLA  [C/RAN-H]
THORNLEY, DIANN  DOMINION’S REACH [UNIFIED WORLDS 3]  H,P
THORNLEY, DIANN  ECHOES OF ISSEL [UNIFIED WORLDS 2]  P
THORNLEY, DIANN  GANWOLD’S CHILD [UNIFIED WORLDS 1]  H,P
THORPE, GAV  ANGELS OF DARKNESS [@WARHAMMER 40K]  P
THORPE, GAV  ANNIHILATION SQUAD [WARHAMMER 40K LAST CHANCERS 3]  P
THORPE, GAV  BLADES OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 2]  P
THORPE, GAV  CLAWS OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 1]  P
THORPE, GAV  GRUDGE BEARER [@WARHAMMER]  P
THORPE, GAV  HEART OF CHAOS, THE [WARHAMMER SLAVES TO DARKNESS 3]  P
THORPE, GAV  KILL TEAM [WARHAMMER 40K LAST CHANCERS 2]  P
THORPE, GAV  LAST CHANCERS, THE [@WARHAMMER 40K]  LP
THORPE, GAV  13TH LEGION [WARHAMMER 40K LAST CHANCERS 1]  P
THORPE, TREBOR  LIGHTNING WORLD  SR-P
THURMAN, ROB  MOONSHINE [CAL AND NIKO 2]  P
THURMAN, ROB  NIGHTLIFE [CAL AND NIKO 1]  P
THURSTON, ROBERT  ALICIA II  H,P
THURSTON, ROBERT  BLOODNAME [@BATTLETECH #2/JADE PHOENIX 2]  C/P,P
THURSTON, ROBERT  FALCON GUARD [@BATTLETECH #3/JADE PHOENIX 3]  C/P:2,P
THURSTON, ROBERT  FALCON RISING [@BATTLETECH #40/TWILIGHT OF THE CLANS 8]  P
THURSTON, ROBERT  FREEBIRTH [@BATTLETECH #36/TWILIGHT OF THE CLANS 4]  C/P,P
THURSTON, ROBERT  I AM JADE FALCON [@BATTLETECH 17]  P
THURSTON, ROBERT  INTRUDER [ROBOTS AND ALIENS 3]  C/P,SFWA-P:2,P
THURSTON, ROBERT  Q COLONY  P
THURSTON, ROBERT  ROBOT JOX  C/P,P
THURSTON, ROBERT  SET OF WHEELS  P
THURSTON, ROBERT  WAY OF THE CLANS [@BATTLETECH #1/JADE PHOENIX 1]  C/P:2,SFWA-P:2,P
TIEDEMANN, MARK W.  AURORA [ASIMOV’S ROBOT MYSTERY]  P
TIEDEMANN, MARK W.  CHIMERA [@ASIMOV’S ROBOT MYSTERY]  LP
TIEDEMANN, MARK W.  COMPASS REACH [SECANTIS 1]  H,LP
TIEDEMANN, MARK W.  HOUR OF THE WOLF [TERMINATOR 2]  LP
TIEDEMANN, MARK W.  METAL OF NIGHT [SECANTIS 2]  LP
TIEDEMANN, MARK W.  MIRAGE [@ASIMOV’S ROBOT MYSTERY]  LP
TIEDEMANN, MARK W.  OTHER WAYS [SECANTIS]  SR-LP
TIEDEMANN, MARK W.  PEACE & MEMORY [SECANTIS 3]  LP
TIEDEMANN, MARK W.  REALTIME  LP
TIER, MARK/GREENBERG, MARTIN H.  FREEDOM!  LPA
TIER, MARK/GREENBERG, MARTIN H.  VISIONS OF LIBERTY  PA
TIGER, JOHN  MISSION IMPOSSIBLE  P
TIGER, JOHN  WIPEOUT [@I SPY 4]  P
TIGGES, JOHN  KEVIN BROWNE’S NIGHTALES  LP
TILLEY, PATRICK  CLOUD WARRIOR [AMTRAK WARS 1]  C/P:2,P
TILLEY, PATRICK  DEATH-BRINGER [@AMTRAK WARS #5]  P
TILLEY, PATRICK  FADE-OUT  H,P
TILLEY, PATRICK  FIRST FAMILY, THE [AMTRAK WARS 2]  C/P:2,P
TILLEY, PATRICK  IRON MASTER [AMTRAK WARS 3]  C/P,P
TILLEY, PATRICK  MISSION  LP
TILLEY, ROBERT J.  BIG LOSERS, THE  P
TILTON, LOIS  ACCUSATIONS [@BABYLON 5 #2]  C/P,P
TILTON, LOIS  BETRAYAL [STAR TREK DS9]  C/P:2,P
TILTON, LOIS  DARKSPAWN  LP
TILTON, LOIS  WRITTEN IN VENOM  H,LP
TIMLETT, PETER VALENTINE  POWER OF THE SERPENT, THE  C/P,P
TIMLETT, PETER VALENTINE  SEEDBEARERS, THE  P
TIMLETT, PETER VALENTINE  TWILIGHT OF THE SERPENT  C/P,P
TIMMERMAN, JOHN H.  OTHER WORLDS: THE FANTASY GENRE  R/REF-P
TIMMONS, STAN  ALWAYS DARKEST  LP
TIPTREE, JAMES JR.  BRIGHTNESS FALLS FROM THE AIR  H:2,SFWA-P,P
TIPTREE, JAMES JR.  BYTE BEAUTIFUL  H
TIPTREE, JAMES JR.  COLOR OF NEANDERTHAL EYES, THE  SFWA-TD16
TIPTREE, JAMES JR.  CROWN OF STARS  C/P,SFWA-P,P,H
TIPTREE, JAMES JR.  GIRL WHO WAS PLUGGED IN, THE  TD7,SFWA-TD7
TIPTREE, JAMES JR.  HER SMOKE ROSE UP FOREVER  H
TIPTREE, JAMES JR.  HOUSTON, HOUSTON, DO YOU READ?  SFWA-TD11,TD11
TIPTREE, JAMES JR.  MEET ME AT INFINITY  H
TIPTREE, JAMES JR.  OUT OF THE EVERYWHERE  C/P,P
TIPTREE, JAMES JR.  STAR SONGS OF AN OLD PRIMATE  C/P:2,P
TIPTREE, JAMES JR.  STARRY RIFT, THE  H:2,SFWA-P
TIPTREE, JAMES JR.  TALES OF THE QUINTANA ROO  H
TIPTREE, JAMES JR.  TEN THOUSAND LIGHT-YEARS FROM HOME  C/P:2,P
TIPTREE, JAMES JR.  UP THE WALLS OF THE WORLD  C/P,H:2,P

The MITSFS Pinkdex  September 24, 2007  Page 382
AUTHOR  TITLE  CODE
TIPTREE, JAMES JR.  WARM WORLDS AND OTHERWISE  C/P:2,P
TIRPA, G. W.  IN FLUID SILENCE [DARK MATTER 3]  P
TITCHENER, LOUISE  GREENFIRE  P
TITELMAN, CAROL  ART OF STAR WARS, THE [STAR WARS]  A
TOBIAS, MICHAEL  FATAL EXPOSURE  P
TOBIAS, MICHAEL  VOICE OF THE PLANET  C/P:2,P
TOFFLER, ALVIN  FUTURE SHOCK  [C/RAN-P]
TOFTE, ARTHUR  CRASH LANDING ON IDUNA  SR-P,C/P
TOFTE, ARTHUR  GHOSTHUNTERS, THE  P
TOFTE, ARTHUR  WALLS WITHIN WALLS  SR-P,C/P
TOLKIEN, J. R. R.  ADVENTURES OF TOM BOMBADIL, THE  H
TOLKIEN, J. R. R.  ADVENTURES OF TOM BOMBADIL, THE (BRITISH)  SR-H
TOLKIEN, J. R. R./ANDERSON, DOUGLAS A.  ANNOTATED HOBBIT, THE  R/REF-H
TOLKIEN, J. R. R.  BILBO LE HOBBIT  FP
TOLKIEN, J. R. R.  BILBO’S LAST SONG  H:2
TOLKIEN, J. R. R.  BOOK OF LOST TALES 1, THE [HISTORY OF MIDDLE-EARTH 1]  H:2,LP
TOLKIEN, J. R. R.  BOOK OF LOST TALES 2, THE [HISTORY OF MIDDLE-EARTH 2]  H:2,P
TOLKIEN, J. R. R.  FARMER GILES OF HAM  H
TOLKIEN, J. R. R.  FATHER CHRISTMAS LETTERS, THE  VLP
TOLKIEN, J. R. R.  FELLOWSHIP OF THE RING, THE [@LORD OF THE RINGS #1]  SR-H,C/P,P,H
TOLKIEN, J. R. R.  FELLOWSHIP OF THE RING, THE (ACE) [@LORD OF THE RINGS #1]  SR-P
TOLKIEN, J. R. R.  FINN AND HENGEST: THE FRAGMENT AND THE EPISODE  R/REF-H
TOLKIEN, J. R. R.  HOBBIT, THE  H:2,SR-H,P

TOLKIEN, J. R. R./DIXON, CHARLES/ DEMING, SEAN HOBBIT, THE

CX

TOLKIEN, J. R. R.  KLEINE HOBBIT, DER  FH
TOLKIEN, J. R. R.  LAYS OF BELERIAND, THE [HISTORY OF MIDDLE-EARTH 3]  H:2,P
TOLKIEN, J. R. R.  LORD OF THE RINGS, THE (12 CASSETTES, BOXED) [LORD OF THE RINGS]  SR-MM
TOLKIEN, J. R. R.  LORD OF THE RINGS, THE (13 CASSETTES, BOXED) [LORD OF THE RINGS]  SR-MM
TOLKIEN, J. R. R.  LORD OF THE RINGS, THE (COLLECTORS EDITION) [LORD OF THE RINGS]  SR-VLH
TOLKIEN, J. R. R.  LOST ROAD AND OTHER WRITINGS, THE [HISTORY OF MIDDLE-EARTH 5]  H:2,P
TOLKIEN, J. R. R.  MONSTERS AND THE CRITICS AND OTHER ESSAYS  R/REF-H
TOLKIEN, J. R. R.  MORGOTH’S RING [HISTORY OF MIDDLE-EARTH 10]  H
TOLKIEN, J. R. R.  PEOPLES OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 12]  H
TOLKIEN, J. R. R.  RETURN OF THE KING, THE [@LORD OF THE RINGS #3]  SR-H,H:2,P
TOLKIEN, J. R. R.  RETURN OF THE KING, THE (ACE) [@LORD OF THE RINGS #3]  SR-P
TOLKIEN, J. R. R.  RETURN OF THE SHADOW, THE [HISTORY OF MIDDLE-EARTH 6]  H
TOLKIEN, J. R. R./SWANN, DONALD  ROAD GOES EVER ON, THE  R/REF-H,C/REF-H
TOLKIEN, J. R. R.  ROVERANDOM  H
TOLKIEN, J. R. R.  SAURON DEFEATED [HISTORY OF MIDDLE-EARTH 9]  H
TOLKIEN, J. R. R.  SHAPING OF MIDDLE-EARTH, THE [HISTORY OF MIDDLE-EARTH 4]  H,P
TOLKIEN, J. R. R.  SILMARILLION, THE  C/P,P,H
TOLKIEN, J. R. R.  SILMARILLION, VOL. 1, THE (4 CDS) [LORD OF THE RINGS]  SR-MM
TOLKIEN, J. R. R.  SILMARILLION, VOL. 2, THE (4 CDS) [LORD OF THE RINGS]  SR-MM
TOLKIEN, J. R. R.  SILMARILLION, VOL. 3, THE (5 CDS) [LORD OF THE RINGS]  SR-MM
TOLKIEN, J. R. R.  SIR GAWAIN AND THE GREEN KNIGHT/PEARL/SIR ORFEO  HA,PA
TOLKIEN, J. R. R.  SMITH OF WOOTTON MAJOR  H
TOLKIEN, J. R. R.  SMITH OF WOOTTON MAJOR/FARMER GILES OF HAM  C/P,P,H
TOLKIEN, J. R. R.  TALE OF THE CHILDREN OF HURIN, THE [LORD OF THE RINGS]  H:2
TOLKIEN, J. R. R.  TOLKIEN MISCELLANY, A  H
TOLKIEN, J. R. R.  TOLKIEN READER, THE  C/P:2,P
TOLKIEN, J. R. R.  TREASON OF ISENGARD [HISTORY OF MIDDLE-EARTH 7]  H
TOLKIEN, J. R. R.  TREE AND LEAF  R/REF-H,C/REF-H:2
TOLKIEN, J. R. R.  TREE AND LEAF/MYTHOPOEIA  SR-H:2
TOLKIEN, J. R. R.  TWO TOWERS, THE [@LORD OF THE RINGS #2]  H:3,SR-H,P
TOLKIEN, J. R. R.  TWO TOWERS, THE (ACE) [@LORD OF THE RINGS #2]  SR-P
TOLKIEN, J. R. R.  UNFINISHED TALES  H:2,P
TOLKIEN, J. R. R.  WAR OF THE JEWELS, THE [HISTORY OF MIDDLE-EARTH 11]  H:2
TOLKIEN, J. R. R.  WAR OF THE RING, THE [HISTORY OF MIDDLE-EARTH 8]  H:2
TOLNAY, TOM  CELLULOID GANGS  SFWA
TOLSTOI, ALEXEI  AELITA  H
TOLSTOI, ALEXEI  GARIN DEATH RAY, THE  H:2
TOLSTOY, LEO  FABLES AND FAIRY TALES  LP
TOLSTOY, NIKOLAI  COMING OF THE KING, THE  C/P,P,H:2
TOMAS, ANDREW  WE ARE NOT THE FIRST  [C/RAN-P]
TOMINO, YOSHIYUKI  AWAKENING [@GUNDAM MOBILE SUIT #1]  C/P,P
TOMINO, YOSHIYUKI  CONFRONTATION [@GUNDAM MOBILE SUIT #3]  C/P,P
TOMINO, YOSHIYUKI  ESCALATION [@GUNDAM MOBILE SUIT #2]  C/P:2,P
TOMLINSON, THERESA  MOON RIDERS, THE  H
TOMPKINS, PETER/BIRD, CHRISTOPHER  SECRET LIFE OF PLANTS, THE  [C/RAN-P]
TONKS, ANGELA  MIND OUT OF TIME  H
TOOKER, RICHARD  DAY OF THE BROWN HORDE, THE  H
TOOLIS, LORNA/SKEET, MICHAEL  TESSERACTS 4  PA
TOOMBS, JANE  GATHERING DARKNESS [@MOONRUNNER 2]  P

AUTHOR  TITLE  CODE
TOOMBS, JANE  UNDER THE SHADOW [@MOONRUNNER 1]  C/P:2,P
TOOMEY, ROBERT E. JR.  WORLD OF TROUBLE, A  C/P:2,P
TOPOR, ROLAND  TENANT, THE  P
TOPPING, KEITH  BYZANTIUM! [DOCTOR WHO PAST DOCTORS]  P
TOPPING, KEITH/DAY, MARTIN  DEVIL GOBLINS FROM NEPTUNE, THE [DOCTOR WHO PAST DOCTORS]  C/P:2,P
TOPPING, KEITH  GHOST SHIP [DOCTOR WHO]  SR-H
TOPPING, KEITH/DAY, MARTIN  HOLLOW MEN, THE [DOCTOR WHO PAST DOCTORS]  C/P,P
TOPPING, KEITH  KING OF TERROR, THE [DOCTOR WHO PAST DOCTORS]  C/P,P
TOR  TOR HORROR SAMPLER  SR-PA,C/PA
TOR  TOR SF SAMPLER, THE  P
TOR BOOKS  CAFE PURGATORIUM  SR-HA
TORGESON, ROY  CHRYSALIS  PA
TORGESON, ROY  CHRYSALIS 2  PA,C/PA
TORGESON, ROY  CHRYSALIS 3  PA
TORGESON, ROY  CHRYSALIS 4  C/PA,PA
TORGESON, ROY  CHRYSALIS 5  PA
TORGESON, ROY  CHRYSALIS 6  PA
TORGESON, ROY  CHRYSALIS 7  PA
TORGESON, ROY  CHRYSALIS 8  HA,PA
TORGESON, ROY  CHRYSALIS 9  HA,PA,SFWA
TORGESON, ROY  CHRYSALIS 10  HA,SFWA
TORGESON, ROY  OTHER WORLDS 1  PA
TORGESON, ROY  OTHER WORLDS 2  PA
TORRES, RAUL  ANTOLOGIA ESPANOLA DE CIENCIA FICCION 1  FP
TORRO, PEL  BEYOND THE BARRIER OF SPACE  P
TORRO, PEL  EXILED IN SPACE  H
TORRO, PEL  FACE OF FEAR, THE  SR-P
TORRO, PEL  FORCE 97X  SR-P
TORRO, PEL  FORMULA 29X  SR-P
TORRO, PEL  FROZEN PLANET  H
TORRO, PEL  GALAXY 666  SR-P,C/P
TORRO, PEL  LAST ASTRONAUT, THE  SR-P,C/P
TORRO, PEL  LEGION OF THE LOST  SR-P
TORRO, PEL  MAN OF METAL  H
TORRO, PEL  PHANTOM ONES, THE  SR-P
TORRO, PEL  RETURN, THE  SR-P
TORRO, PEL  SPACE NO BARRIER  SR-P
TORRO, PEL  STRANGE ONES, THE  SR-P
TORRO, PEL  THROUGH THE BARRIER  SR-P
TOVMASYAN, G. M.  EXTRATERRESTRIAL CIVILIZATIONS  [C/RAN-H]
TOWER, S. D.  ASSASSINS OF TAMURIN, THE  H
TOWER, S. D.  ASSASSINS OF TAMURIN, THE (PROOF)  SR-LP
TOWNSEND, TOM  PANZER SPIRIT  C/P,P
TRACY, LOUIS  AMERICAN EMPEROR, AN  H
TRACY, LOUIS  FINAL WAR, THE  H
TRACY, LOUIS  KARL GRIER  H
TRAINOR, JOSEPH  FAMILY CRYPT [TWILIGHT 20]  P
TRAINOR, JOSEPH  WATERY GRAVE [TWILIGHT #18]  P
TRALINS, ROBERT  ANDROID ARMAGGEDON  P
TRALINS, ROBERT  COSMOZOIDS, THE  C/P:2,P
TRALINS, ROBERT  GHOUL LOVER  P
TRALINS, ROBERT  HIDDEN SPECTRE, THE  [C/RAN-P]
TRALINS, ROBERT  RING-A-DING U F O’S, THE [MISS FROM SIS]  P
TRANSUE, JACOB  TWILIGHT OF THE BASILISKS  P
TRAVERS, P. L.  MARY POPPINS OPENS THE DOOR  H
TRAVISS, KAREN  ALLY [WESS’HAR WARS 5]  P
TRAVISS, KAREN  BLOODLINES [@STAR WARS LEGACY OF THE FORCE 2]  P
TRAVISS, KAREN  CITY OF PEARL [WESS’HAR WARS 1]  P
TRAVISS, KAREN  CITY OF PEARL (PROOF) [WESS’HAR WARS 1]  SR-P
TRAVISS, KAREN  CROSSING THE LINE [WESS’HAR WARS 2]  P
TRAVISS, KAREN  HARD CONTACT [@STAR WARS REPUBLIC COMMANDO 1]  C/P,P
TRAVISS, KAREN  MATRIARCH [WESS’HAR WARS 4]  P
TRAVISS, KAREN  SACRIFICE [STAR WARS LEGACY OF THE FORCE 5]  H
TRAVISS, KAREN  TRIPLE ZERO [@STAR WARS REPUBLIC COMMANDO 2]  P
TRAVISS, KAREN  WORLD BEFORE, THE [WESS’HAR WARS 3]  P
TREBOR, ROBERT  XT CALLED STANLEY, AN  C/P:2,P
TREECE, HENRY  GREEN MAN, THE  C/P,P
TRELL, MAX  SMALL GODS AND MR. BARNUM, THE  H
TREMAYNE, PETER  BLOODMIST  P,SFWA-P
TREMAYNE, PETER  BLOODRIGHT  H,P
TREMAYNE, PETER  CURSE OF LOCH NESS, THE  SR-P
TREMAYNE, PETER  FIRES OF LAN-KERN, THE  H
TREVISAN, DALTON  VAMPIRE OF CURITIBA AND OTHER STORIES, THE  H
TREVOR, DAN  NIGHT WHISTLERS, THE [@NIGHT WHISTLERS #1]  P

AUTHOR  TITLE  CODE
TREVOR, DAN  TERMINATE [@NIGHT WHISTLERS #3]  P
TREVOR, ELLESTON  SHOOT, THE  H
TREW, ANTHONY  TWO HOURS TO DARKNESS  P
TRIGO, JOAN  DESIERTO DE NIEBLA Y CENIZAS  FP
TRIMBLE, BJO  ON THE GOOD SHIP ENTERPRISE  R/REF-P
TRIMBLE, BJO  STAR TREK CONCORDANCE, THE  R/REF-H
TRIMBLE, LOUIS  ANTHROPOL  DH059
TRIMBLE, LOUIS  BODELAN WAY, THE  C/P,P
TRIMBLE, LOUIS  CITY MACHINE, THE  C/P:2,P
TRIMBLE, LOUIS/TRIMBLE, JACQUELYN  GUARDIANS OF THE GATE  C/P:2,P
TRIMBLE, LOUIS  NOBLEST EXPERIMENT IN THE GALAXY, THE  D11560
TRIMBLE, LOUIS  WANDERING VARIABLES, THE  C/P:2,P
TROOP, ALAN F.  DRAGON DELASANGRE, THE [DRAGON DELASANGRE 1]  P
TROOP, ALAN F.  DRAGON DELASANGRE, THE (PROOF) [DRAGON DELASANGRE 1]  SR-LP
TROOP, ALAN F.  DRAGON MOON [DRAGON DELASANGRE 2]  P
TROOP, ALAN F.  HOST OF DRAGONS, A [DRAGON DELASANGRE 4]  P
TROOP, ALAN F.  SEADRAGON’S DAUGHTER, THE [DRAGON DELASANGRE 3]  P
TROUT, KILGORE  VENUS ON THE HALF-SHELL  C/P:2,P
TRUMBAUER, LISA TRUTKOFF  HIDDEN DRAGON, THE [@KNIGHTS OF THE SILVER DRAGON #7]  LP
TRYON, THOMAS  OTHER, THE  C/P:2,P
TRYON, TOM  NIGHT MAGIC  H
TSR  ARRIVAL [BUCK ROGERS]  PA
TSR  CATACLYSM, THE [@DRAGONLANCE TALES II #2]  C/PA,PA
TSR  REIGN OF ISTAR, THE [@DRAGONLANCE TALES II #1]  C/PA,PA
TSR  WAR OF THE LANCE, THE [@DRAGONLANCE TALES II #3]  C/PA,PA
TSR GAMES  DEITIES AND DEMIGODS  [C/RAN-H]
TSR GAMES  FIEND FOLIO  [C/RAN-H]
TUBB, E. C.  ALIEN DUST  H
TUBB, E. C.  ALIEN LIFE  LP
TUBB, E. C.  ALIEN SEED [SPACE: 1999 7]  C/P:2,P
TUBB, E. C.  ANGADO [DUMAREST 29]  C/P:2,P
TUBB, E. C.  BASE LUNAR  FP
TUBB, E. C.  BEST SCIENCE FICTION OF E. C. TUBB, THE  H
TUBB, E. C.  BREAKAWAY [SPACE: 1999 1]  C/P:2,P
TUBB, E. C.  CENTURY OF THE MANIKIN  C/P,P
TUBB, E. C.  C.O.D. MARS  DH040
TUBB, E. C.  COLLISION COURSE [SPACE: 1999 4]  C/P,P
TUBB, E. C.  COMING EVENT, THE [DUMAREST 26]  C/P,P
TUBB, E. C.  DEATH GOD’S DOOM [MALKAR 1]  LP
TUBB, E. C.  DEATH IS A DREAM  H,DH034
TUBB, E. C.  DERAI [DUMAREST 2]  C/P,DH077,D89301
TUBB, E. C.  EARTH IS HEAVEN [DUMAREST 27]  C/P,P
TUBB, E. C.  EARTH SET FREE  LP
TUBB, E. C.  ELOISE [DUMAREST 12]  C/P:2,P
TUBB, E. C.  ESCAPE INTO SPACE  H
TUBB, E. C.  EXTRA MAN, THE  LP
TUBB, E. C.  EYE OF THE ZODIAC [DUMAREST 13]  C/P,P
TUBB, E. C.  HAVEN OF DARKNESS [DUMAREST 16]  C/P:2,P
TUBB, E. C.  I FIGHT FOR MARS  LP
TUBB, E. C.  IDUNA’S UNIVERSE [DUMAREST 21]  C/P,P
TUBB, E. C.  INCIDENT ON ATH [DUMAREST 18]  C/P:2,P
TUBB, E. C.  JACK OF SWORDS [DUMAREST 14]  C/P:2,P
TUBB, E. C.  JESTER AT SCAR, THE [DUMAREST 5]  C/P,D81610
TUBB, E. C.  JONDELLE [DUMAREST 10]  C/P,P
TUBB, E. C.  KALGAN THE GOLDEN  LP
TUBB, E. C.  KALIN [DUMAREST 4]  C/P,D42800,P
TUBB, E. C.  LALLIA [DUMAREST 6]  D71082
TUBB, E. C.  LUCK MACHINE, THE  LP
TUBB, E. C.  MAYENNE [DUMAREST 9]  C/P:2,P
TUBB, E. C.  MECHANICAL MONARCH, THE  DD266
TUBB, E. C.  MELOME [DUMAREST 28]  C/P,P
TUBB, E. C.  MOON BASE  C/P:2,P
TUBB, E. C.  MURDER IN SPACE  LP
TUBB, E. C.  MUTANTS REBEL, THE  SR-P
TUBB, E. C.  NECTAR OF HEAVEN [DUMAREST 24]  C/P:2,P
TUBB, E. C.  PANDORA’S BOX  LP
TUBB, E. C.  PAWN OF THE OMPHALOS  C/P,P
TUBB, E. C.  PRIMITIVE, THE  SR-P,C/P
TUBB, E. C.  PRISON OF NIGHT [DUMAREST 17]  C/P:2,P
TUBB, E. C.  QUILLIAN SECTOR, THE [DUMAREST 19]  C/P,P
TUBB, E. C.  RETURN, THE [@DUMAREST #32]  LP
TUBB, E. C.  ROGUE PLANET [SPACE: 1999 9]  C/P,P
TUBB, E. C.  SATURN PATROL  LP
TUBB, E. C.  SCATTER OF STARDUST, A  D79975,LP

AUTHOR  TITLE  CODE
TUBB, E. C.  SLEEPING CITY, THE [MALKAR 2]  LP
TUBB, E. C.  SPACE-BORN, THE  DD193
TUBB, E. C.  SPECTRUM OF A FORGOTTEN SUN [DUMAREST 15]  C/P:2,P
TUBB, E. C.  S.T.A.R. FLIGHT  C/P:2,P
TUBB, E. C.  STARDEATH  C/P,P
TUBB, E. C.  SYMBOL OF TERRA [DUMAREST 30]  C/P,P
TUBB, E. C.  TECHNOS [DUMAREST 7]  C/P,D79975
TUBB, E. C.  TEMPLE OF DEATH  LP
TUBB, E. C.  TEMPLE OF TRUTH, THE [DUMAREST 31]  C/P,P
TUBB, E. C.  TERRA DATA, THE [DUMAREST 22]  C/P,P
TUBB, E. C.  TERRIDAE, THE [DUMAREST 25]  C/P,P
TUBB, E. C.  TOMORROW  GRD18
TUBB, E. C.  TOYMAN [DUMAREST 3]  D23140
TUBB, E. C.  VERUCHIA [DUMAREST 8]  C/P:2,P
TUBB, E. C.  WALL, THE  LP
TUBB, E. C.  WEB OF SAND [DUMAREST 20]  C/P:2,P
TUBB, E. C.  WINDS OF GATH, THE [DUMAREST 1]  C/P,DH027,D89301
TUBB, E. C.  WINDS OF GATH, THE (WITH "FRESH GUY") [DUMAREST 1]  LP
TUBB, E. C.  WORLD AT BAY  SR-P
TUBB, E. C.  WORLD OF PROMISE [DUMAREST 23]  C/P,P
TUBB, E. C.  ZENYA [DUMAREST 11]  C/P:2,P

TUCK, DONALD H.

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 1: WHO’S WHO, A–L, THE

R/IND

TUCK, DONALD H.

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 2: WHO’S WHO, M–Z, THE

R/IND

TUCK, DONALD H.

ENCYCLOPEDIA OF SCIENCE FICTION AND FANTASY 3: MISCELLANEOUS, THE

R/IND

TUCKER, MIKE/PERRY, ROBERT  ILLEGAL ALIEN [DOCTOR WHO PAST DOCTORS]  C/P:2,P
TUCKER, MIKE/PERRY, ROBERT  LOVING THE ALIEN [DOCTOR WHO PAST DOCTORS]  P
TUCKER, MIKE  NIGHTMARE OF BLACK ISLAND, THE [@DOCTOR WHO]  H
TUCKER, MIKE  PRIME TIME [DOCTOR WHO PAST DOCTORS]  P
TUCKER, WILSON  BEST OF WILSON TUCKER, THE  C/P,P
TUCKER, WILSON  CITY IN THE SEA, THE  SR-P,C/P,GN2.11
TUCKER, WILSON  ICE AND IRON  H:2,P,SFWA
TUCKER, WILSON  LINCOLN HUNTERS, THE  H,P
TUCKER, WILSON  LONG LOUD SILENCE, THE  H:2,P
TUCKER, WILSON  MAN FROM TOMORROW  C/P:2,P
TUCKER, WILSON/CRANE, ROBERT  MANDEN FRA I GAR/OJNE I VERDENSRUMMET  FP
TUCKER, WILSON  PROCESSION OF THE DAMNED  P
TUCKER, WILSON  RESURRECTION DAYS  C/P:2,P
TUCKER, WILSON  SCIENCE-FICTION SUBTREASURY, THE  H
TUCKER, WILSON  THIS WITCH  H
TUCKER, WILSON  TIME BOMB  H:2
TUCKER, WILSON  TIME MASTERS, THE (ORIGINAL)  C/P:2,P
TUCKER, WILSON  TIME MASTERS, THE (REVISED)  H:2,P
TUCKER, WILSON  TIME: X  C/P:2,P
TUCKER, WILSON  TO THE TOMBAUGH STATION  DD479
TUCKER, WILSON  TOMORROW PLUS X  C/P:2,P
TUCKER, WILSON  WILD TALENT  C/P,P,H
TUCKER, WILSON  YEAR OF THE QUIET SUN, THE  C/P:2,SFWA-P,P
TULLOCH, JOHN/ALVARADO, MANUEL  DOCTOR WHO: THE UNFOLDING TEXT [DOCTOR WHO]  R/REF-P,C/REF-H
TULLOCH, JOHN/JENKINS, HENRY  SCIENCE FICTION AUDIENCES  R/REF-H
TUNING, WILLIAM  FUZZY BONES [FUZZIES]  C/P:2,P
TUNING, WILLIAM  TORNADO ALLEY  C/P,P
TURK, H. C.  BLACK BODY  H
TURK, H. C.  ETHER ORE  P,SFWA-P
TURNBULL, GERRY  STAR TREK CATALOG [STAR TREK]  [C/RAN-H]
TURNER, ALICE K.  PLAYBOY BOOK OF SCIENCE FICTION, THE  HA,LPA
TURNER, DAVID G.  FIRST EDITIONS OF ANDRE NORTON, THE  R/IND
TURNER, DELIA MARSHALL  NAMELESS MAGERY [MENNENKALTENEI 1]  P
TURNER, DELIA MARSHALL  OF SWORDS AND SPELLS [MENNENKALTENEI 2]  P
TURNER, DELIA MARSHALL  WAYS OF MAGIC, THE [MENNENKALTENEI 1,2]  H
TURNER, FREDERICK  DOUBLE SHADOW, A  H,P
TURNER, GARY/HALPERN, MARTY  SILVER GRYPHON, THE  HA
TURNER, GEORGE  BELOVED SON [ETHICAL SOCIETY 1]  C/P,P
TURNER, GEORGE  BRAIN CHILD  SFWA-P,H:2,P
TURNER, GEORGE  DESTINY MAKERS, THE  H:2,SFWA,P
TURNER, GEORGE  DOWN THERE IN DARKNESS  LP
TURNER, GEORGE  DROWNING TOWERS  H
TURNER, GEORGE  GENETIC SOLDIER  H:2,P
TURNER, GEORGE  VANEGLORY [ETHICAL SOCIETY 2]  P
TURNER, GEORGE  VIEW FROM THE EDGE, THE  SR-LPA
TURNER, GEORGE  YESTERDAY’S MEN [ETHICAL SOCIETY 3]  P
TURNER, GEORGE KIBBE  RED FRIDAY  SR-H

The MITSFS Pinkdex  September 24, 2007  Page 386
AUTHOR  TITLE  CODE
TURNER, GERRY  STRANGER FROM THE DEPTHS  C/P:2,P
TURNER, JAMES  FOURTH GHOST BOOK, THE  SR-PA
TURNER, JAMES  UNLIKELY GHOSTS, THE  SR-PA,HA
TURNER, JIM  CTHULHU 2000  HA,LPA
TURNER, JIM  TALES OF THE CTHULHU MYTHOS  HA
TURNER, KARLA  INTO THE FRINGE  SFWA-P
TURNER, MEGAN WHALEN  INSTEAD OF THREE WISHES  LP
TURNER PUBLISHING  LADY COTTINGTON’S PRESSED FAIRY BOOK  A,H
TURTLEDOVE, HARRY  ADVANCE AND RETREAT [WAR BETWEEN THE PROVINCES 3]  H
TURTLEDOVE, HARRY  AGENT OF BYZANTIUM  H,P
TURTLEDOVE, HARRY  AGENT OF BYZANTIUM (REVISED)  C/P,SFWA-P,P
TURTLEDOVE, HARRY  ALTERNATE GENERALS  PA,C/PA
TURTLEDOVE, HARRY  ALTERNATE GENERALS II  HA
TURTLEDOVE, HARRY/GREEN, ROLAND J.  ALTERNATE GENERALS III  HA
TURTLEDOVE, HARRY  AMERICAN FRONT [CONFEDERATE STATES: GREAT WAR 1]  C/P,P

TURTLEDOVE, HARRY/ GREENBERG, MARTIN H.

BEST ALTERNATE HISTORY STORIES OF THE 20TH CENTURY, THE

LPA

TURTLEDOVE, HARRY/ GREENBERG, MARTIN H.

BEST MILITARY SCIENCE FICTION OF THE 20TH CENTURY, THE

HA,LPA

TURTLEDOVE, HARRY/ GREENBERG, MARTIN H.

BEST TIME TRAVEL STORIES OF THE 20TH CENTURY, THE

LPA

TURTLEDOVE, HARRY  BETWEEN THE RIVERS  H,P
TURTLEDOVE, HARRY  BEYOND THE GAP [OPENING OF THE WORLD]  H
TURTLEDOVE, HARRY  BLOOD AND IRON [CONFEDERATE STATES AMERICAN EMPIRE 1]  C/P,P,H
TURTLEDOVE, HARRY  BREAKTHROUGHS [CONFEDERATE STATES: GREAT WAR 3]  H,P
TURTLEDOVE, HARRY  BRIDGE OF THE SEPARATOR [VIDESSOS]  H
TURTLEDOVE, HARRY  CASE OF THE TOXIC SPELL DUMP, THE  C/P:2,SFWA-P,P

TURTLEDOVE, HARRY CENTER CANNOT HOLD, THE

[CONFEDERATE STATES AMERICAN EMPIRE 2]

H

TURTLEDOVE, HARRY  COLONIZATION: AFTERSHOCKS [WORLDWAR 7]  H,P
TURTLEDOVE, HARRY  COLONIZATION: DOWN TO EARTH [WORLDWAR 6]  C/P,P
TURTLEDOVE, HARRY  COLONIZATION: SECOND CONTACT [WORLDWAR 5]  H
TURTLEDOVE, HARRY  CONAN OF VENARIUM [CONAN]  P
TURTLEDOVE, HARRY  COUNTING UP COUNTING DOWN  LP
TURTLEDOVE, HARRY  DARKNESS DESCENDING [INTO THE DARKNESS 2]  C/P,P,H
TURTLEDOVE, HARRY  DEPARTURES  C/P:2,P
TURTLEDOVE, HARRY  DIFFERENT FLESH, A  C/P:2,P
TURTLEDOVE, HARRY  DISUNITED STATES OF AMERICA, THE [@CROSSTIME TRAFFIC #4]  H
TURTLEDOVE, HARRY/DE CAMP, L. SPRAGUE DOWN IN THE BOTTOMLANDS (AND OTHER PLACES)  P
TURTLEDOVE, HARRY  DRIVE TO THE EAST [@SETTLING ACCOUNTS 2]  H
TURTLEDOVE, HARRY  EARTHGRIP  C/P:2,P
TURTLEDOVE, HARRY  EMPEROR FOR THE LEGION, AN [VIDESSOS 2]  C/P:2,P
TURTLEDOVE, HARRY  ENCHANTER COMPLETED, THE  C/PA,PA
TURTLEDOVE, HARRY  END OF THE BEGINNING  H
TURTLEDOVE, HARRY  EVERY INCH A KING  H
TURTLEDOVE, HARRY/DOYLE, NOREEN  FIRST HEROES, THE  LPA
TURTLEDOVE, HARRY  FOX AND EMPIRE [GERIN THE FOX 3]  P
TURTLEDOVE, HARRY  GRAPPLE, THE [@SETTLING ACCOUNTS 3]  H
TURTLEDOVE, HARRY  GUNPOWDER EMPIRE [CROSSTIME TRAFFIC 1]  H
TURTLEDOVE, HARRY  GUNS OF THE SOUTH, THE  C/P,P,H
TURTLEDOVE, HARRY  HAMMER AND ANVIL [TIME OF TROUBLES 2]  P
TURTLEDOVE, HARRY  HOMEWARD BOUND  H
TURTLEDOVE, HARRY  HOW FEW REMAIN [CONFEDERATE STATES]  C/P:2,P
TURTLEDOVE, HARRY  IN HIGH PLACES [CROSSTIME TRAFFIC 3]  H
TURTLEDOVE, HARRY  IN THE PRESENCE OF MINE ENEMIES  H
TURTLEDOVE, HARRY  INTO THE DARKNESS [INTO THE DARKNESS 1]  C/P,P,H
TURTLEDOVE, HARRY  JAWS OF DARKNESS [INTO THE DARKNESS 5]  P
TURTLEDOVE, HARRY  KALEIDOSCOPE  C/P,P
TURTLEDOVE, HARRY  KING OF THE NORTH  P
TURTLEDOVE, HARRY  KRISPOS OF VIDESSOS [KRISPOS 2]  C/P:2,P
TURTLEDOVE, HARRY  KRISPOS RISING [KRISPOS 1]  C/P:2,P
TURTLEDOVE, HARRY  KRISPOS THE EMPEROR [KRISPOS 3]  P
TURTLEDOVE, HARRY  LEGION OF VIDESSOS, THE [VIDESSOS 3]  C/P:2,P
TURTLEDOVE, HARRY  MARCHING THROUGH PEACHTREE [WAR BETWEEN THE PROVINCES 2]  H
TURTLEDOVE, HARRY  MISPLACED LEGION, THE [VIDESSOS 1]  C/P:2,P
TURTLEDOVE, HARRY  NONINTERFERENCE  C/P,P
TURTLEDOVE, HARRY  OUT OF THE DARKNESS [INTO THE DARKNESS 6]  P
TURTLEDOVE, HARRY  PRINCE OF THE NORTH [GERIN THE FOX 2]  C/P,P
TURTLEDOVE, HARRY  PUGNACIOUS PEACEMAKER, THE  TD20,SFWA-TD20
TURTLEDOVE, HARRY  RETURN ENGAGEMENT [SETTLING ACCOUNTS 1]  H
TURTLEDOVE, HARRY  RULED BRITANNIA  H
TURTLEDOVE, HARRY  RULERS OF THE DARKNESS [INTO THE DARKNESS 4]  H
TURTLEDOVE, HARRY  SENTRY PEAK [WAR BETWEEN THE PROVINCES 1]  H,P
TURTLEDOVE, HARRY  STOLEN THRONE, THE [TIME OF TROUBLES 1]  P

The MITSFS Pinkdex  September 24, 2007  Page 387
AUTHOR  TITLE  CODE
TURTLEDOVE, HARRY  SWORDS OF THE LEGION [VIDESSOS 4]  C/P:2,P
TURTLEDOVE, HARRY  THESSALONICA  C/P:2,P
TURTLEDOVE, HARRY  THOUSAND CITIES, THE [TIME OF TROUBLES 3]  P
TURTLEDOVE, HARRY  THROUGH THE DARKNESS [INTO THE DARKNESS 3]  C/P,P,H

TURTLEDOVE, HARRY VICTORIOUS OPPOSITION, THE

[CONFEDERATE STATES AMERICAN EMPIRE 3]

H

TURTLEDOVE, HARRY  VIDESSOS BESIEGED [VIDESSOS 5]  P
TURTLEDOVE, HARRY  WALK IN HELL [CONFEDERATE STATES: GREAT WAR 2]  C/P,P,H
TURTLEDOVE, HARRY  WERENIGHT [GERIN THE FOX 1]  P,SFWA-P
TURTLEDOVE, HARRY  WISDOM OF THE FOX [GERIN THE FOX (REVISED) 1,2]  C/P,P
TURTLEDOVE, HARRY  WORLD OF DIFFERENCE, A  C/P:2,P

TURTLEDOVE, HARRY/STIRLING, S. M./ GENTLE, MARY/WILLIAMS, WALTER JON WORLDS THAT WEREN’T

H,LP

TURTLEDOVE, HARRY  WORLDWAR: IN THE BALANCE [@WORLDWAR 1]  C/P,P,H:2
TURTLEDOVE, HARRY  WORLDWAR: STRIKING THE BALANCE [@WORLDWAR 4]  C/P,P,H
TURTLEDOVE, HARRY  WORLDWAR: TILTING THE BALANCE [@WORLDWAR 2]  H:2,P
TURTLEDOVE, HARRY  WORLDWAR: UPSETTING THE BALANCE [@WORLDWAR 3]  C/P,P
TUTT, C. T.  TEN DRAGON TAILS  H
TUTTLE, LISA  FAMILIAR SPIRIT  SFWA-P
TUTTLE, LISA  GABRIEL  SFWA-P
TUTTLE, LISA  PILLOW FRIEND, THE  H
TUTTLE, LISA  SILVER BOUGH, THE  H
TUTTLE, LISA  SPACESHIP BUILT OF STONE, A  P
TWAIN, MARK  CONNECTICUT YANKEE IN KING ARTHUR’S COURT, A  C/P,P,H
TWAYNE BOOKS  PETRIFIED PLANET, THE  HA
TWELVE HAWKS, JOHN  DARK RIVER, THE [@FOURTH REALM #2]  H
TWELVE HAWKS, JOHN  TRAVELER, THE [FOURTH REALM 1]  LP
TWENTIETH CENTURY FOX  HOW I SURVIVED MY SUMMER VACATION, VOL. 1 [BUFFY]  C/PA,PA
TWENTIETH CENTURY FOX  LONGEST NIGHT 1, THE [@ANGEL]  PA
TWIFORD, WILLIAM RICHARD  SOWN IN THE DARKNESS A.D. 2,000  H
TWITCHELL, PAUL  TALONS OF TIME  P
TYERS, KATHY  BALANCE POINT [@STAR WARS NEW JEDI ORDER]  P
TYERS, KATHY  CROWN OF FIRE [FIREBIRD 3]  LP
TYERS, KATHY  CRYSTAL WITNESS  C/P,P
TYERS, KATHY  FIREBIRD [FIREBIRD 1]  C/P,P
TYERS, KATHY  FIREBIRD TRILOGY, THE [FIREBIRD 1,2,3]  H
TYERS, KATHY  FUSION FIRE [FIREBIRD 2]  C/P,P
TYERS, KATHY  ONE MIND’S EYE  P
TYERS, KATHY  SHIVERING WORLD  C/P:2,P
TYERS, KATHY  TRUCE AT BAKURA, THE [@STAR WARS]  C/P,P,H
TYLER, J. E. A.  NEW TOLKIEN COMPANION, THE  R/REF-H
TYLER, J. E. A.  TOLKIEN COMPANION, THE  R/REF-H
TYLER, THEODORE  MAN WHOSE NAME WOULDN’T FIT, THE  H:2,P
TYMN, MARSHALL B.  AMERICAN FANTASY AND SCIENCE FICTION  SRL/IND
TYMN, MARSHALL B.  HORROR LITERATURE  R/REF-H

TYMN, MARSHALL B./SCHLOBIN, ROGER C./ CURREY, L. W.

RESEARCH GUIDE TO SCIENCE FICTION STUDIES, A

R/IND

TYRON, TOM  NIGHT MAGIC  H
TYSON, DONALD  NECRONOMICON  LP
TYSON, SALINDRA  WHEEL OF DREAMS  C/P,P

AUTHOR  TITLE  CODE
ULLMAN, JAMES R.  SANDS OF KARAKORUM, THE  P
UMPATHY, RANJAN  MASTERPIECE  H
UNDERWOOD, LAURA J.  DRAGON’S TONGUE [DEMON-BOUND 1]  LP
UNDERWOOD, TIM/MILLER, CHUCK  FEAR ITSELF: THE HORROR FICTION OF STEPHEN KING  R/REF-H
UNDERWOOD, TIM/MILLER, CHUCK  JACK VANCE  R/REF-H
UNITED FRUIT COMPANY  PROBLEMS AND PROGRESS IN BANANA DISEASE RESEARCH  [C/RAN-H]
UPDIKE, JOHN  WITCHES OF EASTWICK, THE  C/P,P,H
URBANCIK, JOHN  BABYLON RISING  BD00001
URRUTIA, BENJAMIN  LATTER DAY SCIENCE FICTION 2  SR-LPA
URRUTIA, BENJAMIN  LATTER DAY SCIENCE FICTION 3  SR-LPA
UTLEY, BRIAN W.  MARTYR  C/P,P
UTTLEY, ALISON  TRAVELER IN TIME, A  H
UTTLEY, ALISON  TRAVELLER IN TIME, A  P

AUTHOR  TITLE  CODE
VACCA, ROBERTO  COMING DARK AGE, THE  SFWA
VALDEMI, MARIA  DEMON LOVER, THE  P
VALE, RENA  BEYOND THE SEALED WORLD  C/P:2,P
VALE, RENA  DAY AFTER DOOMSDAY, THE  C/P:2,P
VALE, RENA  TAURUS FOUR  P
VALE, ROBERT B.  EFFICIENCY IN HADES  SR-H
VALENTE, CATHERYNNE M  IN THE NIGHT GARDEN [@ORPHAN’S TALES #1]  LP
VALENTINE, VICTOR  CURE FOR DEATH  SR-P
VALLANCE, KARL  GLOBAL BLACKOUT  SR-P
VALLEE, JACQUES/TORME, TRACY  FASTWALKER  LP
VALLEJO, BORIS/BELL, JULIE  BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2003  CAL
VALLEJO, BORIS/BELL, JULIE  BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2004  CAL
VALLEJO, BORIS/BELL, JULIE  BORIS VALLEJO & JULIE BELL’S FANTASY CALENDAR 2005  CAL
VALLEJO, BORIS  BORIS VALLEJO CLASSIC 1997 CALENDAR  CAL
VALLEJO, BORIS  BORIS VALLEJO’S FANTASY CALENDAR 1995  CAL
VALLEJO, BORIS  BORIS VALLEJO’S FANTASY CALENDAR 1996  CAL
VALLEJO, BORIS  BORIS VALLEJO’S FANTASY CALENDAR 1998  CAL
VALLEJO, BORIS  BORIS VALLEJO’S FANTASY CALENDAR 2000  CAL
VALLEJO, BORIS  BORIS VALLEJO’S FANTASY CALENDAR 2001  CAL
VALLEJO, BORIS  BORIS VALLEJO’S MYTHOLOGY CALENDAR 1989  CAL
VALLEJO, BORIS  BORIS VALLEJO’S MYTHOLOGY CALENDAR 1992  CAL
VALLEJO, BORIS  BORIS VALLEJO’S MYTHOLOGY CALENDAR 1993  CAL
VALLEJO, BORIS  BORIS VALLEJO’S MYTHOLOGY CALENDAR 1994  CAL
VALLEJO, DORIS/VALLEJO, BORIS  ENCHANTMENT  A
VALLEJO, DORIS  WINDSOUND  C/P,P
VALLESE, RAY/VALLESE, VALERIE  TORMENT [@PLANE SCAPE]  P
VAN ARNAM, DAVE  GREYLAND  C/P:2,P
VAN ARNAM, DAVE  LORD OF BLOOD  C/P,P
VAN ARNAM, DAVE/ARCHER, RON  LOST IN SPACE [LOST IN SPACE]  P
VAN ARNAM, DAVE  PLAYERS OF HELL, THE  SR-DB0077
VAN ARNAM, DAVE  STAR BARBARIAN  C/P,P
VAN ARNAM, DAVE  STAR GLADIATOR  SR-DB0788
VAN ARNAM, DAVE  STARMIND  C/P:2,P
VAN ARNAM, DAVE  WIZARD OF STORMS  C/P,P
VAN ASTEN, GAIL  BLIND KNIGHT, THE  P
VAN ASTEN, GAIL  CHARLEMAGNE’S CHAMPION [ROLAND 1]  C/P,P
VAN ASTEN, GAIL  DARK SWORD’S LOVER, THE [ROLAND 2]  C/P,P
VAN BELKOM, EDO  LORD SOTH [@DRAGONLANCE WARRIORS #6]  P
VAN BELKOM, EDO  WYRM WOLF [WORLD OF DARKNESS]  P
VAN GELDER, GORDON  FOURTH PLANET FROM THE SUN  HA,LPA
VAN GELDER, GORDON  IN LANDS THAT NEVER WERE  LPA
VAN GELDER, GORDON  ONE LAMP  LPA
VAN GORES, ALIDA  MERMAID’S SONG  P
VAN GULIK, ROBERT  CELEBRATED CASES OF JUDGE DEE  LP
VAN HAMME, JEAN/ROSINSKI, GRZEGORZ  THORGAL: CHILD OF THE STARS  CX
VAN HAMME, JEAN/ROSINSKI, GRZEGORZ  THORGAL: THE SORCERESS BETRAYED  CX
VAN HERCK, PAUL  WHERE WERE YOU LAST PLUTERDAY?  C/P:2,P
VAN HISE, DELLA  KILLING TIME (ORIGINAL) [STAR TREK]  C/P:2,SR-P
VAN HISE, DELLA  KILLING TIME (REVISED) [STAR TREK]  C/P:3,P
VAN HISE, JAMES  SCI FI TV: FROM THE TWILIGHT ZONE TO DEEP SPACE NINE  R/REF-P
VAN HISE, JAMES  TREK: THE NEXT GENERATION REVISED EDITION [STAR TREK TNG]  R/REF-H
VAN HISE, JAMES  UNAUTHORIZED HISTORY OF TREK, THE  R/REF-P

VAN HISE, JAMES UNAUTHORIZED TREKKERS’ GUIDE TO THE NEXT GENERATION AND DEEP

SPACE NINE, THE

R/REF-P,C/REF-P

VAN LHIN, ERIK  BATTLE ON MERCURY  H:2
VAN LODEN, ERLE  CURSE OF PLANET KUZ  SR-P
VAN LODEN, ERLE  VOYAGE INTO SPACE  SR-P
VAN LOON, HENDRIK WILLEM  INVASION  H
VAN NAME, MARK L.  ONE JUMP AHEAD  H
VAN PALLANDT, NICHOLAS  ANVIL  P
VAN PELT, JAMES  SUMMER OF THE APOCALYPSE  LP
VAN SCYOC, SYDNEY J.  ASSIGNMENT NOR’DYREN  C/P:2,P
VAN SCYOC, SYDNEY J.  BLUESONG  C/P:2,LP
VAN SCYOC, SYDNEY J.  CLOUDCRY  H:2,P
VAN SCYOC, SYDNEY J.  DARKCHILD  C/P:2,LP
VAN SCYOC, SYDNEY J.  DAUGHTERS OF THE SUNSTONE  H:2
VAN SCYOC, SYDNEY J.  DEEPWATER DREAMS  C/P:2,P
VAN SCYOC, SYDNEY J.  DROWNTIDE  C/P,SFWA-P,P
VAN SCYOC, SYDNEY J.  FEATHER STROKE  C/P,P
VAN SCYOC, SYDNEY J.  SALTFLOWER  C/P:2,P
VAN SCYOC, SYDNEY J.  STAR MOTHER  H:2,P
VAN SCYOC, SYDNEY J.  STARSILK  P
VAN SCYOC, SYDNEY J.  SUNWAIFS  H,P
VAN THAL, HERBERT  LIE TEN NIGHTS AWAKE  PA
VAN THAL, HERBERT  PAN BOOK OF HORROR STORIES, THE  PA

AUTHOR  TITLE  CODE
VAN THAL, HERBERT  THIRD PAN BOOK OF HORROR STORIES, THE  C/PA,PA
VAN THAL, HERBERT  FOURTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN THAL, HERBERT  FIFTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN THAL, HERBERT  SIXTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN THAL, HERBERT  EIGHTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN THAL, HERBERT  ELEVENTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN THAL, HERBERT  THIRTEENTH PAN BOOK OF HORROR STORIES, THE  SR-PA
VAN VOGT, A. E.  ANARCHISTIC COLOSSUS, THE  C/P:2,P
VAN VOGT, A. E.  ASTRONAVE PIRATA, A  FP
VAN VOGT, A. E.  AWAY AND BEYOND  C/P:2,P
VAN VOGT, A. E.  BATTLE OF FOREVER, THE  C/P:2,P
VAN VOGT, A. E.  BEAST, THE  C/P,P,H
VAN VOGT, A. E.  BEST OF A. E. VAN VOGT, THE  C/P,P
VAN VOGT, A. E.  BLAL, THE  C/P,P
VAN VOGT, A. E.  BOOK OF PTATH, THE  C/P:2,P,H
VAN VOGT, A. E.  BOOK OF VAN VOGT, THE  C/P:2,P
VAN VOGT, A. E.  CHANGELING, THE  C/P:2,P
VAN VOGT, A. E.  CHILDREN OF TOMORROW  C/P:2,P
VAN VOGT, A. E.  COMPUTERWORLD  C/P:2,P
VAN VOGT, A. E.  COSMIC ENCOUNTER  H,SFWA
VAN VOGT, A. E.  DARKNESS ON DIAMONDIA, THE  C/P:2,P
VAN VOGT, A. E.  DESTINATION: UNIVERSE!  C/P,P,H
VAN VOGT, A. E.  EARTH FACTOR X (ORIGINALLY THE SECRET GALACTICS)  C/P,P
VAN VOGT, A. E.  EARTH’S LAST FORTRESS  DD431
VAN VOGT, A. E.  EMPIRE OF THE ATOM  C/P,P,H,DD242,SFWA
VAN VOGT, A. E.  FAR-OUT WORLDS OF A.E. VAN VOGT, THE  C/P:2,P
VAN VOGT, A. E.  FUTURE GLITTER  C/P:2,P
VAN VOGT, A. E.  FUTURES PAST  LP
VAN VOGT, A. E.  GRYB, THE  C/P,P
VAN VOGT, A. E.  HOUSE THAT STOOD STILL, THE  C/P,P,H
VAN VOGT, A. E.  LOST: FIFTY SUNS  P
VAN VOGT, A. E.  M33 IN ANDROMEDA  C/P:2,P
VAN VOGT, A. E.  MAN WITH A THOUSAND NAMES, THE  C/P:2,P
VAN VOGT, A. E.  MASTERS OF TIME  C/P,P,H
VAN VOGT, A. E.  MIND CAGE, THE  H:2,P
VAN VOGT, A. E.  MISSION: INTERPLANETARY  C/P:2,P
VAN VOGT, A. E.  MISSION TO THE STARS  C/P:2,P
VAN VOGT, A. E.  MIXED MEN, THE  H:2
VAN VOGT, A. E.  MONSTERS  C/P:2,P
VAN VOGT, A. E.  MOONBEAST  P
VAN VOGT, A. E.  MORE THAN SUPERHUMAN  SR-P,C/P
VAN VOGT, A. E.  NULL-A THREE  C/P:2,P
VAN VOGT, A. E.  ONE AGAINST ETERNITY  DD094
VAN VOGT, A. E./HULL, E. MAYNE  OUT OF THE UNKNOWN  SR-H:2,SR-P
VAN VOGT, A. E.  PENDULUM  C/P:2,P
VAN VOGT, A. E./HULL, E. MAYNE  PLANETS FOR SALE  SR-P,C/P:2
VAN VOGT, A. E.  PLAYERS OF NULL-A, THE  C/P:2,P
VAN VOGT, A. E.  PROXY INTELLIGENCE AND OTHER MIND BENDERS, THE  C/P:2,P
VAN VOGT, A. E.  QUEST FOR THE FUTURE  H:2,P
VAN VOGT, A. E.  REFLECTIONS OF A. E. VAN VOGT  R/REF-P
VAN VOGT, A. E.  RENAISSANCE  C/P:2,P
VAN VOGT, A. E.  ROGUE SHIP  H:2,P
VAN VOGT, A. E.  SECRET GALACTICS, THE  H,LP
VAN VOGT, A. E.  SIEGE OF THE UNSEEN  DD391
VAN VOGT, A. E.  SILKIE, THE  C/P:2,P
VAN VOGT, A. E.  SLAN  C/P,P,H
VAN VOGT, A. E./ANDERSON, KEVIN J.  SLAN HUNTER  H
VAN VOGT, A. E.  SORCIER DE LINN, LE  FP
VAN VOGT, A. E.  SUPERMIND  C/P:2,P
VAN VOGT, A. E.  TRANSFINITE: THE ESSENTIAL A.E. VAN VOGT  H
VAN VOGT, A. E.  TRIAD  H:2
VAN VOGT, A. E.  TWISTED MEN, THE  DF253
VAN VOGT, A. E.  TWO HUNDRED MILLION A.D.  C/P:2,P
VAN VOGT, A. E.  TYRANOPOLIS  SR-P
VAN VOGT, A. E.  UNIVERSE MAKER, THE  SR-DD031,P,C/P:2
VAN VOGT, A. E.  VIOLENT MAN, THE  P
VAN VOGT, A. E.  VOYAGE OF THE SPACE BEAGLE, THE  C/P:2,SFWA,P
VAN VOGT, A. E.  WAR AGAINST THE RULL, THE  H:2,P
VAN VOGT, A. E.  WEAPON MAKERS, THE  C/P,P,H
VAN VOGT, A. E.  WEAPON SHOPS OF ISHER, THE  SR-DD053,C/P:2,P
VAN VOGT, A. E./HULL, E. MAYNE  WINGED MAN, THE  C/P,P,H
VAN VOGT, A. E.  WIZARD OF LINN, THE  C/P:2,P
VAN VOGT, A. E.  WORLD OF NULL-A, THE  SR-DD031,P,H,C/P:2
VAN VOGT, A. E.  WORLDS OF A. E. VAN VOGT, THE  C/P:3,P
VANCE, JACK  ALASTOR [ALASTOR]  H

The MITSFS Pinkdex  September 24, 2007  Page 393
AUTHOR  TITLE  CODE
VANCE, JACK  ANOME, THE  C/P:2,P
VANCE, JACK  ARAMINTA STATION [CADWAL 1]  C/P,SFWA-P,P,H
VANCE, JACK  ARAMINTA STATION (VIE) [@VANCE INTEGRAL EDITION #39/CADWAL 1]  SR-H
VANCE, JACK  ASUTRA, THE  C/P:2,P

VANCE, JACK BAD RONALD [@VANCE INTEGRAL EDITION #12]/

DARK OCEAN, THE [@VANCE INTEGRAL EDITION #12]

SR-H

VANCE, JACK  BEST OF JACK VANCE, THE  C/P:2,P
VANCE, JACK  BIG PLANET  DD295,SFWA-P,H:2,P
VANCE, JACK  BLUE WORLD, THE  C/P:2,P
VANCE, JACK  BLUE WORLD (VIE), THE [@VANCE INTEGRAL EDITION #16]  SR-H
VANCE, JACK  BOOK OF DREAMS, THE [DEMON PRINCES 5]  C/P:2,P

VANCE, JACK BOOK OF DREAMS, THE (VIE) [@VANCE INTEGRAL EDITION #26/

DEMON PRINCES 5]

SR-H

VANCE, JACK  BRAINS OF EARTH, THE  DM141
VANCE, JACK  BRAVE FREE MEN, THE  C/P:2,P
VANCE, JACK  CITY OF THE CHASCH  C/P:2,P
VANCE, JACK  COMPLEAT DYING EARTH, THE [DYING EARTH 1,2,3,4]  H:2
VANCE, JACK  CUGEL THE CLEVER [@VANCE INTEGRAL EDITION #15]  SR-H

VANCE, JACK CUGEL: THE SKYBREAK [@VANCE INTEGRAL EDITION #35]/

SPATTERLIGHT [@VANCE INTEGRAL EDITION #35]

SR-H

VANCE, JACK  CUGEL’S SAGA [DYING EARTH 3]  C/P,P,H
VANCE, JACK  DARK SIDE OF THE MOON, THE  H
VANCE, JACK  DEMON PRINCES, THE [DEMON PRINCES 1,2,3,4,5]  H
VANCE, JACK  DIRDIR, THE  C/P:2,P

VANCE, JACK DOGTOWN TOURIST AGENCY, THE [@VANCE INTEGRAL EDITION #32]/

FREITZKE’S TURN [@VANCE INTEGRAL EDITION #32]

SR-H

VANCE, JACK  DOMAINS OF KORYPHON, THE [@VANCE INTEGRAL EDITION #28]  SR-H

VANCE, JACK DRAGON MASTERS, THE C/P,H,DF185,P,D16640,

SFWA-TD28

VANCE, JACK  DURDANE [@VANCE INTEGRAL EDITION #27]  SR-H
VANCE, JACK  DUST OF FAR SUNS  P
VANCE, JACK  DYING EARTH, THE [DYING EARTH 1]  C/P:2,P
VANCE, JACK  ECCE AND OLD EARTH [CADWAL 2]  C/P,P,H
VANCE, JACK  ECCE AND OLD EARTH (VIE) [@VANCE INTEGRAL EDITION #40/CADWAL 2] SR-H
VANCE, JACK  EIGHT FANTASMS AND MAGICS  C/P:2,P
VANCE, JACK  EMPHYRIO  C/P:2,P
VANCE, JACK  EMPHYRIO (VIE) [@VANCE INTEGRAL EDITION #20]  SR-H
VANCE, JACK  EYES OF THE OVERWORLD, THE [DYING EARTH 2]  SR-H,C/P:2,P
VANCE, JACK  FACE, THE [DEMON PRINCES 4]  C/P:2,P
VANCE, JACK  FACE, THE (VIE) [@VANCE INTEGRAL EDITION #25/DEMON PRINCES 4]  SR-H
VANCE, JACK  FACELESS MAN, THE  P
VANCE, JACK  FIVE GOLD BANDS, THE  C/P,DF185,P,D16640

VANCE, JACK FLESH MASK, THE [@VANCE INTEGRAL EDITION #10]/

STRANGE PEOPLE, QUEER NOTIONS [@VANCE INTEGRAL EDITION #10]/ BIRD ISLAND [@VANCE INTEGRAL EDITION #10]

SR-H

VANCE, JACK FOX VALLEY MURDERS, THE [@VANCE INTEGRAL EDITION #13]/

PLEASANT GROVE MURDERS, THE [@VANCE INTEGRAL EDITION #13]

SR-H

VANCE, JACK  FUTURE TENSE  C/P:2,P
VANCE, JACK  GADGET STORIES [@VANCE INTEGRAL EDITION #3]  SR-H
VANCE, JACK  GALACTIC EFFECTUATOR  C/P,P

VANCE, JACK GOLD AND IRON [@VANCE INTEGRAL EDITION #7]/ CLARGES [@VANCE INTEGRAL EDITION #7]/ LANGUAGES OF PAO, THE [@VANCE INTEGRAL EDITION #7]

SR-H

VANCE, JACK  GOLDEN GIRL AND OTHER STORIES [@VANCE INTEGRAL EDITION #6]  SR-H
VANCE, JACK  GRAY PRINCE, THE  H:2,P
VANCE, JACK  GREEN MAGIC  SR-H,SFWA-P
VANCE, JACK  GREEN PEARL, THE [LYONESSE 2]  H:2,SFWA-P,LP
VANCE, JACK  GREEN PEARL, THE (VIE) [@VANCE INTEGRAL EDITION #37/LYONESSE 2]  SR-H

VANCE, JACK HOUSE ON LILY STREET, THE [@VANCE INTEGRAL EDITION #11]/

VIEW FROM CHICKWEE’S WINDOW, THE

[@VANCE INTEGRAL EDITION #11]

SR-H

VANCE, JACK  HOUSES OF ISZM, THE  D77525,DF265

VANCE, JACK HOUSES OF ISZM AND OTHER STORIES, THE

[@VANCE INTEGRAL EDITION #8]

SR-H

VANCE, JACK  JACK VANCE TREASURY, THE  H
VANCE, JACK  KILLING MACHINE, THE [DEMON PRINCES 2]  C/P:2,P

VANCE, JACK KILLING MACHINE (VIE), THE [@VANCE INTEGRAL EDITION #23/

DEMON PRINCES 2]

SR-H

VANCE, JACK  LANGUAGES OF PAO, THE  C/P:3,SFWA-P,P
VANCE, JACK  LAST CASTLE, THE  TD15,DH021, SFWA-TD15
VANCE, JACK  LIGHT FROM A LONE STAR  SR-H
VANCE, JACK  MADOUC [LYONESSE 3]  C/P,SFWA-P:2,LP,H
VANCE, JACK  MADOUC (VIE) [@VANCE INTEGRAL EDITION #38/LYONESSE 3]  SR-H

The MITSFS Pinkdex  September 24, 2007  Page 394
AUTHOR  TITLE  CODE

VANCE, JACK MAGNIFICENT SHOWBOATS OF THE LOWER VISSEL RIVER,

LUNE XXIII SOUTH, BIG PLANET, THE [@VANCE INTEGRAL EDITION #19]

SR-H

VANCE, JACK MAN IN THE CAGE, THE [@VANCE INTEGRAL EDITION #14]/ DEADLY ISLES, THE [@VANCE INTEGRAL EDITION #14]

SR-H

VANCE, JACK  MANY WORLDS OF MAGNUS RIDOLPH, THE  DM141
VANCE, JACK  MANY WORLDS OF MAGNUS RIDOLPH, THE (EXPANDED)  C/P:2,P
VANCE, JACK  MARUNE: ALASTOR 933 [ALASTOR 2]  C/P:2,P

VANCE, JACK MARUNE: ALASTOR 933 (VIE) [@VANCE INTEGRAL EDITION #30/

ALASTOR 2]

SR-H

VANCE, JACK  MASKE: THAERY  H:2,P
VANCE, JACK  MASKE: THAERY (VIE) [@VANCE INTEGRAL EDITION #33]  SR-H

VANCE, JACK MAZIRIAN THE MAGICIAN [@VANCE INTEGRAL EDITION #1/

DYING EARTH 1]

SR-H

VANCE, JACK MIRACLE WORKERS, THE [@VANCE INTEGRAL EDITION #9]/ DRAGON MASTERS, THE [@VANCE INTEGRAL EDITION #9]/ LAST CASTLE, THE [@VANCE INTEGRAL EDITION #9]

SR-H

VANCE, JACK  MONSTERS IN ORBIT  DM125
VANCE, JACK  MOON MOTH AND OTHER STORIES, THE [@VANCE INTEGRAL EDITION #17] SR-H
VANCE, JACK  NARROW LAND, THE  P
VANCE, JACK  NIGHT LAMP  H
VANCE, JACK  NIGHT LAMP (VIE) [@VANCE INTEGRAL EDITION #42]  SR-H
VANCE, JACK  PALACE OF LOVE, THE [DEMON PRINCES 3]  C/P:2,P

VANCE, JACK PALACE OF LOVE (VIE), THE [@VANCE INTEGRAL EDITION #24/

DEMON PRINCES 3]

SR-H

VANCE, JACK  PLANET OF ADVENTURE  H
VANCE, JACK  PNUME, THE  C/P:2,P
VANCE, JACK  PORTS OF CALL [MYRON TANY 1]  H
VANCE, JACK  PORTS OF CALL & LURULU [MYRON TANY 1,2]  H
VANCE, JACK  PORTS OF CALL (VIE) [@VANCE INTEGRAL EDITION #43/MYRON TANY 1]  SR-H

VANCE, JACK RAPPAREE, THE [@VANCE INTEGRAL EDITION #4]/ BIG PLANET [@VANCE INTEGRAL EDITION #4]/ VANDALS OF THE VOID [@VANCE INTEGRAL EDITION #4]

SR-H

VANCE, JACK  RHIALTO THE MARVELLOUS [DYING EARTH 4]  H,P

VANCE, JACK RHIALTO THE MARVELLOUS (VIE) [@VANCE INTEGRAL EDITION #34/

DYING EARTH 4]

SR-H

VANCE, JACK  SERVANTS OF THE WANKH  C/P:2,P
VANCE, JACK  SHOWBOAT WORLD  C/P:2,SFWA-P,P
VANCE, JACK  SLAVES OF THE KLAU  DD295
VANCE, JACK  SON OF THE TREE  D77525,DF265
VANCE, JACK  SON OF THE TREE AND OTHER STORIES [@VANCE INTEGRAL EDITION #5]  SR-H
VANCE, JACK  SPACE OPERA  C/P:2,P
VANCE, JACK  SPACE OPERA (VIE) [@VANCE INTEGRAL EDITION #18]  SR-H
VANCE, JACK  STAR KING, THE [DEMON PRINCES 1]  C/P:2,P

VANCE, JACK STAR KING (VIE), THE [@VANCE INTEGRAL EDITION #22/

DEMON PRINCES 1]

SR-H

VANCE, JACK  SULDRUN’S GARDEN [LYONESSE 1]  C/P,LP,H
VANCE, JACK  SULDRUN’S GARDEN (VIE) [@VANCE INTEGRAL EDITION #36/LYONESSE 1]  SR-H
VANCE, JACK  THROY [CADWAL 3]  H
VANCE, JACK  THROY (VIE) [@VANCE INTEGRAL EDITION #41/CADWAL 3]  SR-H
VANCE, JACK  TO LIVE FOREVER  C/P:2,P
VANCE, JACK  TRULLION: ALASTOR 2262 [ALASTOR 1]  C/P:2,P

VANCE, JACK TRULLION: ALASTOR 2262 (VIE) [@VANCE INTEGRAL EDITION #29/

ALASTOR 1]

SR-H

VANCE, JACK  TSCHAI [@VANCE INTEGRAL EDITION #21]  SR-H
VANCE, JACK  VANDALS OF THE VOID  H:2

VANCE, JACK WILD THYME AND VIOLETS, OTHER UNPUBLISHED WORKS,

AND ADDENDA [@VANCE INTEGRAL EDITION #44]

SR-H

VANCE, JACK  WORLD BETWEEN AND OTHER STORIES, THE  DM125

VANCE, JACK WORLD THINKER AND OTHER STORIES, THE

[@VANCE INTEGRAL EDITION #2]

SR-H

VANCE, JACK  WORLDS OF JACK VANCE, THE  C/P,P
VANCE, JACK  WYST: ALASTOR 1716 [ALASTOR 3]  C/P:2,P
VANCE, JACK  WYST: ALASTOR 1716 (VIE) [@VANCE INTEGRAL EDITION #31/ALASTOR 3]  SR-H
VANCE, JOHN HOLBROOK  BAD RONALD  P
VANCE, STEVE  ALL THE SHATTERED WORLDS  P
VANCE, STEVE  ASGARD RUN, THE  C/P,P
VANCE, STEVE  HYBRID, THE  P
VANCE, STEVE  PLANET OF THE GAWFS  C/P,P
VANDERMEER, JEFF  CITY OF SAINTS AND MADMEN [AMBERGRIS]  LP
VANDERMEER, JEFF  SECRET LIFE  H
VANDERMEER, JEFF  SHRIEK: AN AFTERWORD [AMBERGRIS]  LP

VANDERMEER, JEFF/ROBERTS, MARK THACKERY T. LAMBSHEAD POCKET GUIDE TO ECCENTRIC &

DISCREDITED DISEASES, THE

LPA

VANDERMEER, JEFF  VENISS UNDERGROUND  LP

AUTHOR  TITLE  CODE
VANDERMEER, JEFF  WHY SHOULD I CUT YOUR THROAT?  R/REF-H
VARDEMAN, ROBERT E.  ALIEN WEB, THE [@MASTERS OF SPACE #2]  P
VARDEMAN, ROBERT E.  ANCIENT HEAVENS  P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  BEASTS OF THE MIST, THE [SWORDS OF RAEMLLYN 5]  C/P:2,P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  BLOOD FOUNTAIN [SWORDS OF RAEMLLYN 3]  C/P:2,P
VARDEMAN, ROBERT E.  CENOTAPH ROAD [CENOTAPH ROAD 1]  P
VARDEMAN, ROBERT E./MILAN, VICTOR  CITY IN THE GLACIER, THE [WAR OF POWERS 2]  C/P:2,P
VARDEMAN, ROBERT E.  COLORS OF CHAOS [@WEAPONS OF CHAOS #3]  P
VARDEMAN, ROBERT E.  CRISIS AT STARLIGHT [@BIOWARRIORS #2]  C/P,P
VARDEMAN, ROBERT E.  CRYSTAL CLOUDS, THE [@JADE DEMONS #3]  P
VARDEMAN, ROBERT E.  DARK LEGACY [@MAGIC: THE GATHERING]  P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  DEATH’S ACOLYTE [SWORDS OF RAEMLLYN 4]  P
VARDEMAN, ROBERT E./MILAN, VICTOR  DEMON OF THE DARK ONES [WAR OF POWERS 6]  C/P:2,P
VARDEMAN, ROBERT E./MILAN, VICTOR  DESTINY STONE, THE [WAR OF POWERS 3]  C/P,P
VARDEMAN, ROBERT E.  ECHOES OF CHAOS, THE [@WEAPONS OF CHAOS #1]  C/P:2,SFWA-P,P
VARDEMAN, ROBERT E.  EQUATIONS OF CHAOS [@WEAPONS OF CHAOS #2]  P
VARDEMAN, ROBERT E./MILAN, VICTOR  FALLEN ONES, THE [WAR OF POWERS 4]  C/P,P
VARDEMAN, ROBERT E.  FIRE AND FOG [CENOTAPH ROAD 5]  C/P,P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  FOR CROWN AND KINGDOM [SWORDS OF RAEMLLYN 6]  P
VARDEMAN, ROBERT E.  FROZEN WAVES, THE [@JADE DEMONS #2]  P
VARDEMAN, ROBERT E.  GLASS WARRIOR, THE [@DEMON CROWN #1]  P,SFWA-P
VARDEMAN, ROBERT E.  HELL HEART [@VOR THE MAELSTROM]  P
VARDEMAN, ROBERT E./MILAN, VICTOR  IN THE SHADOW OF OMIZANTRIM [WAR OF POWERS 5]  C/P,P
VARDEMAN, ROBERT E.  INFINITY PLAGUE, THE [@BIOWARRIORS 1]  P
VARDEMAN, ROBERT E.  IRON TONGUE [CENOTAPH ROAD 4]  C/P,P
VARDEMAN, ROBERT E.  KEY OF ICE AND STEEL [KEYS TO PARADISE 3]  P,SFWA-P
VARDEMAN, ROBERT E.  KLINGON GAMBIT, THE [STAR TREK]  C/P:3,P
VARDEMAN, ROBERT E.  MUTINY ON THE ENTERPRISE [STAR TREK]  C/P:3,P
VARDEMAN, ROBERT E.  PHANTOMS ON THE WIND [@DEMON CROWN #2]  P,SFWA-P
VARDEMAN, ROBERT E.  PILLAR OF NIGHT [CENOTAPH ROAD 6]  C/P:2,P
VARDEMAN, ROBERT E.  PLAGUE IN PARADISE, A [@MASTERS OF SPACE #3]  P
VARDEMAN, ROBERT E.  QUAKING LANDS, THE [@JADE DEMONS #1]  P
VARDEMAN, ROBERT E.  RESONANCE OF BLOOD, THE  P
VARDEMAN, ROBERT E.  RUINS OF POWER, THE [@MECHWARRIOR DARK AGE/BATTLETECH]  P
VARDEMAN, ROBERT E.  SANDCATS OF RHYL, THE  P
VARDEMAN, ROBERT E.  SKELETON LORD’S KEY, THE [KEYS TO PARADISE 2]  C/P,SFWA-P,P
VARDEMAN, ROBERT E.  SORCERER’S SKULL, THE [CENOTAPH ROAD 2]  C/P:2,P
VARDEMAN, ROBERT E.  SPACE VECTORS [@BIOWARRIORS #3]  C/P,P
VARDEMAN, ROBERT E.  STELLAR DEATH PLAN, THE [@MASTERS OF SPACE #1]  C/P,P
VARDEMAN, ROBERT E./MILAN, VICTOR  SUNDERED REALM, THE [WAR OF POWERS 1]  C/P:2,P
VARDEMAN, ROBERT E.  SYMPHONY OF STORMS, A [@DEMON CROWN #3]  C/P:2,SFWA-P,P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  TO DEMONS BOUND [SWORDS OF RAEMLLYN 1]  C/P,P
VARDEMAN, ROBERT E.  WHITE FIRE, THE [@JADE DEMONS #4]  P
VARDEMAN, ROBERT E.  WORLD OF MAZES [CENOTAPH ROAD 3]  C/P,P
VARDEMAN, ROBERT E./PROCTOR, GEO. W.  YOKE OF MAGIC, A [SWORDS OF RAEMLLYN 2]  C/P,P
VARE, DANIELE  DOGE’S RING, THE  H
VARE, ETHLIE ANN/MORRIS, DANIEL  BROKEN PLACES, THE [ANDROMEDA (GENE RODDENBERRY’S)]  H
VARIAN, CHRISTOPHER  DR. TIM BOOK ONE  CX
VARLEY, JOHN  BARBIE MURDERS, THE  C/P:2,P
VARLEY, JOHN  BLUE CHAMPAGNE  C/P,P,H
VARLEY, JOHN  DEMON [GAEA 3]  H:2,LP
VARLEY, JOHN  GOLDEN GLOBE, THE  H,P
VARLEY, JOHN  JOHN VARLEY READER, THE  H
VARLEY, JOHN  MAMMOTH  H
VARLEY, JOHN  MILLENNIUM  H:2,LP
VARLEY, JOHN  OPHIUCHI HOTLINE, THE  C/P,P,H
VARLEY, JOHN  PERSISTENCE OF VISION, THE  H:2,P
VARLEY, JOHN  PICNIC ON NEARSIDE  C/P,P
VARLEY, JOHN  PRESS ENTER  H,SFWA-TD26
VARLEY, JOHN  RED LIGHTNING [RED THUNDER 2]  H
VARLEY, JOHN  RED THUNDER [RED THUNDER 1]  H
VARLEY, JOHN  STEEL BEACH  C/P:2,P,H
VARLEY, JOHN/MAINHARDT, RICIA  SUPERHEROES  HA,LPA
VARLEY, JOHN  TANGO CHARLIE AND FOXTROT ROMEO  SFWA-TD4,TD4
VARLEY, JOHN  TITAN [GAEA 1]  H:2,P
VARLEY, JOHN  WIZARD [GAEA 2]  H:2,P,SFWA
VARNEY, ALLEN  CAST OF FATE [DRAGON DICE]  SR-P
VASILIEV, M. P., ET AL.  SALYUT IN ORBIT  FH
VAUGHN, CARRIE  KITTY AND THE MIDNIGHT HOUR [KITTY NORVILLE 1]  P
VAUGHN, CARRIE  KITTY GOES TO WASHINGTON [KITTY NORVILLE 2]  P
VAUGHN, CARRIE  KITTY TAKES A HOLIDAY [KITTY NORVILLE 3]  P
VEJCHAR, ALFRED  STORY INDEX #1 1955–1965  R/IND
VELASCO, RAYMOND L.  GUIDE TO THE STAR WARS UNIVERSE, A  R/REF-P
VELDE, VIVIAN VANDE  CHANGELING PRINCE, THE  C/P,P

AUTHOR  TITLE  CODE
VELDE, VIVIAN VANDE  COMPANIONS OF THE NIGHT  P
VELDE, VIVIAN VANDE  CONJURER PRINCESS, THE  C/P,P
VELDE, VIVIAN VANDE  DRAGON’S BAIT  P
VELDE, VIVIAN VANDE  NEVER TRUST A DEAD MAN  H
VELDE, VIVIAN VANDE  SPELLBOUND  H
VELDE, VIVIAN VANDE  WELL-TIMED ENCHANTMENT, A  H
VELEY, CHARLES  CHILDREN OF THE DARK  P
VELEY, CHARLES  PLAY TO LIVE [TWILIGHT 7]  P
VELIKOVSKY, IMMANUEL  AGES IN CHAOS  [C/RAN-H]
VELIKOVSKY, IMMANUEL  EARTH IN UPHEAVAL  [C/RAN-P:2]
VELIKOVSKY, IMMANUEL  WORLDS IN COLLISION  [C/RAN-H,C/RAN-P]
VELON, ERNEST  MAN FROM HARDIN, THE  LP
VERCORS  ANIMAUX DENATURES, LES  FP
VERCORS  SYLVA  FP
VERCORS  YOU SHALL KNOW THEM  H:2,P
VERLAG VOLK UND WELT  INDEX J-81 ARBEITET FUR MR. FAUST  FH
VERNAM, GLENN R.  POWER OF THE GODS, THE  P
VERNE, JULES  ADRIFT IN THE PACIFIC  H
VERNE, JULES  ADVENTURES OF THE RAT FAMILY  H
VERNE, JULES  AMONG THE CANNIBALS  H
VERNE, JULES  ANNOTATED JULES VERNE: FROM THE EARTH TO THE MOON, THE  SR-VLH

VERNE, JULES ANNOTATED JULES VERNE:

TWENTY THOUSAND LEAGUES UNDER THE SEA, THE

SR-VLP,H

VERNE, JULES  ANOMALOUS PHENOMENA  H
VERNE, JULES  AROUND THE WORLD IN EIGHTY DAYS  C/P:2,P
VERNE, JULES  AT THE NORTH POLE  H
VERNE, JULES  BEGUM’S FORTUNE, THE  H
VERNE, JULES  BEST OF JULES VERNE  H
VERNE, JULES  BLACK DIAMONDS  H
VERNE, JULES  BURBANK THE NORTHERNER  H
VERNE, JULES  CARPATHIAN CASTLE  H,P
VERNE, JULES  CHANCELLOR, THE  H
VERNE, JULES  CINQ SEMAINES AU BALLON  FP
VERNE, JULES  CITY IN THE SAHARA, THE  H,P
VERNE, JULES  CLAIM ON FORTY MILE CREEK, THE  H
VERNE, JULES  CLIPPER OF THE CLOUDS, THE  H
VERNE, JULES  COLLECTED NOVELS  H
VERNE, JULES  CRYPTOGRAM, THE  H
VERNE, JULES  DANUBE PILOT, THE  H
VERNE, JULES  DE LA TERRE A LA LUNE  FP
VERNE, JULES  DEMON OF CAWNPORE, THE  H,P
VERNE, JULES  DOCTOR OX AND OTHER STORIES  H
VERNE, JULES  DOWN THE AMAZON  H
VERNE, JULES  DRAMA IN LIVONIA  H
VERNE, JULES  DROPPED FROM THE CLOUDS  H
VERNE, JULES  EIGHT HUNDRED LEAGUES ON THE AMAZON  H
VERNE, JULES  END OF THE JOURNEY  H
VERNE, JULES  FACING THE FLAG  H
VERNE, JULES  FIVE WEEKS IN A BALLOON  H
VERNE, JULES  FLIGHT TO FRANCE  H
VERNE, JULES  FLOATING CITY, A  H
VERNE, JULES  FLOOD AND FLAME  H
VERNE, JULES  FOR THE FLAG  C/P,P,H
VERNE, JULES  FROM THE EARTH TO THE MOON  C/P,P
VERNE, JULES  FROM THE EARTH TO THE MOON/TOUR OF THE MOON, A  SR-H
VERNE, JULES  FROM THE EARTH TO THE MOON/TRIP AROUND IT, A  C/P,P,H
VERNE, JULES  FUR COUNTRY OR 70 NORTH LATITUDE  SR-H
VERNE, JULES  GREEN RAY, THE/BLOCKADE RUNNERS, THE  H
VERNE, JULES  HECTOR SERVADAC  FP
VERNE, JULES  HOMEWARD BOUND  H
VERNE, JULES  HUNT FOR THE METEOR, THE  H,P
VERNE, JULES  INTO THE ABYSS  H
VERNE, JULES  INTO THE NIGER BEND  C/P,P,H
VERNE, JULES  JOURNEY TO THE CENTER OF THE EARTH, A (BRADLEY TRANSLATION)  C/P,P
VERNE, JULES  JOURNEY TO THE CENTRE OF THE EARTH, A (EVANS TRANSLATION)  H
VERNE, JULES  JOURNEY TO THE CENTRE OF THE EARTH, A (UNCREDITED TRANSLATION) H:2,P
VERNE, JULES  LEADER OF THE RESISTANCE  H
VERNE, JULES  LIGHTHOUSE AT THE END OF THE WORLD, THE  H
VERNE, JULES  LONG VACATION, A  H
VERNE, JULES  MAITRE DU MONDE  FP
VERNE, JULES  MASTER OF THE WORLD  C/P:2,P,H
VERNE, JULES  MASTERLESS MAN, THE  H
VERNE, JULES  MATHIAS SANDORF TOME II  FP
VERNE, JULES  MEASURING A MERIDIAN  H
VERNE, JULES  MYSTERIOUS DOCUMENT, THE  H

The MITSFS Pinkdex  September 24, 2007  Page 397
AUTHOR  TITLE  CODE
VERNE, JULES  MYSTERIOUS ISLAND, THE (ABRIDGED, BAIR TRANSLATION)  C/P,P
VERNE, JULES  MYSTERIOUS ISLAND, THE (KINGSTON TRANSLATION)  H,P
VERNE, JULES  OFF ON A COMET (ACE BOOKS, 1957, ABRIDGED)  C/P:2,P
VERNE, JULES  PACKAGE HOLIDAY  H
VERNE, JULES  PARIS IN THE TWENTIETH CENTURY  H,LP
VERNE, JULES  PROPELLER ISLAND  H
VERNE, JULES  PURCHASE OF THE NORTH POLE, THE  C/P,P,H

VERNE, JULES ROBINSON CRUSOE SCHOOL, THE/STAR OF THE SOUTH, THE/

PURCHASE OF THE NORTH POLE

SR-H

VERNE, JULES  ROBUR LE CONQUERANT  FP
VERNE, JULES  ROUND THE MOON  H
VERNE, JULES  SALVAGE FROM THE CYNTHIA  H
VERNE, JULES  SCHOOL FOR CRUSOES, A  H
VERNE, JULES  SEA SERPENT, THE  H
VERNE, JULES  SECOND YEAR ASHORE  H
VERNE, JULES  SECRET OF THE ISLAND, THE  H
VERNE, JULES  SECRET OF WILHELM STORITZ, THE  H
VERNE, JULES  SHOW ON ICE, THE  H
VERNE, JULES  SOUTHERN STAR MYSTERY, THE  H
VERNE, JULES  SUN IN ECLIPSE, THE  H
VERNE, JULES  TEXAR THE SOUTHERNER  H
VERNE, JULES  THROUGH THE BEHRING STRAIT  H
VERNE, JULES  TIGERS AND TRAITORS  H,P

VERNE, JULES TO THE SUN? (ROTH TRANSLATION)/

OFF ON A COMET! (ROTH TRANSLATION)

LP

VERNE, JULES  TOUR DU MONDE EN 80 JOURS, LE/RAYON-VERT, LE  FH
VERNE, JULES  TOUR DU MONDE EN QUATRE-VINGTS JOURS, LE  FH
VERNE, JULES  TRAVELLING CIRCUS, THE  H
VERNE, JULES  TRIBULATIONS OF A CHINESE GENTLEMAN, THE  H
VERNE, JULES  20,000 LEAGUES UNDER THE SEA/AROUND THE MOON  H
VERNE, JULES  TWENTY THOUSAND LEAGUES UNDER THE SEA  C/P,P,H,LP
VERNE, JULES  TWENTY THOUSAND LEAGUES UNDER THE SEA (ANNOTATED)  LP
VERNE, JULES  TWENTY THOUSAND LEAGUES UNDER THE SEA (BONNER TRANSLATION)  P
VERNE, JULES  UNWILLING DICTATOR, THE  H
VERNE, JULES  VILLAGE IN THE TREETOPS, THE  H
VERNE, JULES  VINGT MILLE LIEUES SOUS LES MERS  FP
VERNE, JULES  VOYAGE AU CENTRE DE LA TERRE  FP
VERNE, JULES  WILDERNESS OF ICE, THE  H
VERNE, JULES  WORKS OF JULES VERNE  H
VERNE, JULES  YESTERDAY AND TOMORROW  H,P
VERNEY, SARAH J.  CAT BY THE TAIL [@SABRINA THE TEENAGE WITCH SALEM #7]  P
VERNON, KAY  ARDREYS, THE  P
VERNON, KAY  DARK SEAS OF MALTERN MANOR, THE  P
VERNON, ROGER LEE  SPACE FRONTIERS, THE  C/P:2,P
VERRILL, A. HYATT  BRIDGE OF LIGHT, THE  SR-H
VERRILL, A. HYATT  STRANGE STORY OF OUR EARTH, THE  [C/RAN-P]
VERSEAU, DOMINIQUE  GIRL FROM EROSPHERE, THE [YOLANDA 1]  SR-L
VERSEAU, DOMINIQUE  SLAVES OF SPACE [YOLANDA 2]  SR-L
VERSINS, PIERRE  ENCYCLOPEDIE DE L’UTOPIE ET DE LA SCIENCE FICTION  R/REF-H
VERSLUIS, ARTHUR  TELOS  P
VESS, CHARLES  HORNS OF ELFLAND, THE  VLP
VESSER, CAROLYN  HELLWALKER  C/P,SFWA-P,P
VIDAL, GORE  DULUTH  H
VIDAL, GORE  KALKI  H,P
VIDAL, GORE  MESSIAH  C/P:2,P
VIDAL, GORE  SMITHSONIAN INSTITUTION, THE  H
VIDAL, GORE  VISIT TO A SMALL PLANET  C/P:2,P
VIEHL, S. L.  AFTERBURN [BIO RESCUE 2]  H
VIEHL, S. L.  BEYOND VARALLAN [STARDOC 2]  C/P,P
VIEHL, S. L.  BIO RESCUE [BIO RESCUE 1]  H
VIEHL, S. L.  BLADE DANCER [STARDOC JORY RASK]  H,P
VIEHL, S. L.  ENDURANCE [STARDOC 3]  C/P,P
VIEHL, S. L.  ETERNITY ROW [@STARDOC]  C/P,P
VIEHL, S. L.  PLAGUE OF MEMORY [@STARDOC]  P
VIEHL, S. L.  REBEL ICE [@STARDOC 5]  P
VIEHL, S. L.  SHOCKBALL [STARDOC 4]  C/P,P
VIEHL, S. L.  STARDOC [STARDOC 1]  C/P,P

VIERECK, GEORGE SYLVESTER/ ELDRIDGE, PAUL MY FIRST TWO THOUSAND YEARS (ABRIDGED)

C/P,P

VIGGIANO, MICHAEL/FRANSON, DONALD  SCIENCE FICTION TITLE CHANGES  R/IND
VIGLIANTE, MARY  COLONY, THE  P
VIGLIANTE, MARY  LAND, THE  P
VIGLIANTE, MARY  SOURCE OF EVIL  P
VIGLIANTE, MARY  WORSHIP THE NIGHT  P

AUTHOR  TITLE  CODE
VIGUIE, DEBBIE  PIED PIPER [@CHARMED]  P
VILOTT, RHONDI  TOWERS OF REXOR, THE [@DRAGONTALES #4]  P
VINCENT, HARL  DOOMSDAY PLANET, THE  C/P:2,P
VINCENT, RACHEL  STRAY  P
VINGE, JOAN D.  ALIEN BLOOD [CAT 1,2]  H:2
VINGE, JOAN D.  CATSPAW [CAT 2]  C/P:2,P
VINGE, JOAN D.  DREAMFALL [CAT 3]  H:2
VINGE, JOAN D.  EYES OF AMBER  C/P:2,P
VINGE, JOAN D.  FIRESHIP  C/P,P,H:2
VINGE, JOAN D.  FIRESHIP/MOTHER AND CHILD  P
VINGE, JOAN D.  HEAVEN CHRONICLES [HEAVEN 1,2]  C/P,P,H
VINGE, JOAN D.  LADYHAWKE  C/P:2,P
VINGE, JOAN D.  LEGACY [HEAVEN 1]  BS4
VINGE, JOAN D.  LOST IN SPACE [LOST IN SPACE]  P
VINGE, JOAN D.  MAD MAX BEYOND THUNDERDOME  P
VINGE, JOAN D.  OUTCASTS OF HEAVEN BELT, THE [HEAVEN 2]  C/P:2,P
VINGE, JOAN D.  PHOENIX IN THE ASHES  H,SFWA-P,P
VINGE, JOAN D.  PSION [CAT 1]  C/P:2,P
VINGE, JOAN D.  RETURN TO OZ [OZ]  C/P,P
VINGE, JOAN D.  SANTA CLAUS: THE MOVIE  P
VINGE, JOAN D.  SNOW QUEEN, THE [SNOW QUEEN 1]  C/P,SFWA-P,H,P,SFWA
VINGE, JOAN D.  SUMMER QUEEN, THE [SNOW QUEEN 3]  H:2,P
VINGE, JOAN D.  TANGLED UP IN BLUE [SNOW QUEEN 4]  H:2
VINGE, JOAN D.  TIN SOLDIER  TD23
VINGE, JOAN D.  WORLD’S END [SNOW QUEEN 2]  SFWA-P,H:2,P,SFWA
VINGE, VERNOR  ACROSS REALTIME  H:2,SFWA-P,P
VINGE, VERNOR  COLLECTED STORIES OF VERNOR VINGE, THE  H
VINGE, VERNOR  DEEPNESS IN THE SKY, A [DEEP 2]  C/P,P,H
VINGE, VERNOR  FIRE UPON THE DEEP, A [DEEP 1]  H:2,P
VINGE, VERNOR  GRIMM’S WORLD  C/P,P,H
VINGE, VERNOR  MAROONED IN REALTIME  C/P,P,H
VINGE, VERNOR  MONDO DI GRIMM, IL (ITALIAN)  FP
VINGE, VERNOR  PEACE WAR, THE  C/P:2,SFWA,P,H:3
VINGE, VERNOR  RAINBOWS END  H
VINGE, VERNOR  TATJA GRIMM’S WORLD  C/P:2,P
VINGE, VERNOR  THREATS ... AND OTHER PROMISES  C/P:2,SFWA-P,P
VINGE, VERNOR  TRUE NAMES  BS5
VINGE, VERNOR  TRUE NAMES ... AND OTHER DANGERS  P,SFWA-P
VINGE, VERNOR  WITLING, THE  C/P:2,P
VINICOFF, ERIC  MAIDEN FLIGHT  C/P:2,SFWA-P,P
VINICOFF, ERIC/MARTIN, MARCIA  SPACING DUTCHMAN  SR-LP
VINICOFF, ERIC/MARTIN, MARCIA  WEIGHER, THE  C/P,P
VIORST, JUDITH  PROJECTS: SPACE  [C/RAN-P]
VIRGIL  AENEID, THE  P
VITARELLI, ROBERT  STRANGE HAPPENINGS  SR-PA
VITOLA, DENISE  HALF-LIGHT  P
VITOLA, DENISE  MANJINN MOON [TY MERRICK 3]  P
VITOLA, DENISE  OPALITE MOON [TY MERRICK 2]  P
VITOLA, DENISE  QUANTUM MOON [TY MERRICK 1]  P
VITOLA, DENISE  RADON FILE, THE [TY MERRICK 5]  P
VITOLA, DENISE  RED SKY FILE, THE [TY MERRICK 4]  P
VIVIAN, E. CHARLES  CITY OF WONDER  P
VIVIAN, E. CHARLES  FIELDS OF SLEEP  SR-H
VIXEN, RICHARD M.  MAGIC CARPET AND THE CEMENT WALL, THE  H,VLP
VLCEK, ERNST  ENDLOSE ALPTRAUM, DER [@PERRY RHODAN #67]  FPR
VLCEK, ERNST  SATANS UNIVERSUM [@PERRY RHODAN #85]  FPR
VOERMANS, PAUL  AND DISREGARDS THE REST  P
VOINOVICH, VLADIMIR  MOSCOW 2042  H
VOISKUNSKY, E./LUKODYANOV, I.  CREW OF THE MEKONG, THE  P
VOLSKY, PAULA  CURSE OF THE WITCH-QUEEN, THE  C/P:2,P
VOLSKY, PAULA  GATES OF TWILIGHT, THE  LP
VOLSKY, PAULA  GRAND ELLIPSE, THE  P
VOLSKY, PAULA  ILLUSION  C/P,P,H
VOLSKY, PAULA  LUCK OF RELIAN KRU, THE  C/P,SFWA-P,P
VOLSKY, PAULA  SORCERER’S CURSE, THE  C/P,P
VOLSKY, PAULA  SORCERER’S HEIR, THE  C/P,P
VOLSKY, PAULA  SORCEROR’S LADY, THE  P
VOLSKY, PAULA  WHITE TRIBUNAL, THE  C/P:2,LP
VOLSKY, PAULA  WOLF OF WINTER, THE  C/P,LP,H
VOLTA, ORNELLA  VAMPIRE, THE  [C/RAN-P]
VOLTAIRE  BEST KNOWN WORKS OF VOLTAIRE, THE  H
VOLTAIRE  PORTABLE VOLTAIRE, THE  P
VOLTZ, WILLIAM  HUMANS KEEP OUT! [PERRY RHODAN 135]  SR-LP
VOLTZ, WILLIAM  KILLERS FROM HYPERSPACE [PERRY RHODAN 120]  SR-LP

The MITSFS Pinkdex  September 24, 2007  Page 399
AUTHOR  TITLE  CODE
VOLTZ, WILLIAM  ROBOTS, BOMBS AND MUTANTS [PERRY RHODAN 125]  SR-LP
VOLTZ, WILLIAM  WORLD WITHOUT MERCY [PERRY RHODAN 131]  SR-LP
VOLTZ, WILLIAM CH.  STERNEN KAMPFER  FH
VON ALLMEN, STEWART  CONSPICUOUS CONSUMPTION [@WORLD OF DARKNESS WEREWOLF 2]  P
VON BRAUN, WERNHER  FIRST MEN TO THE MOON  H
VON DANIKEN, ERICH  CHARIOTS OF THE GODS?  [C/RAN-P:2]
VON DANIKEN, ERICH  GODS FROM OUTER SPACE  [C/RAN-H,C/RAN-P:2]
VON DANIKEN, ERICH  GOLD OF THE GODS, THE  [C/RAN-H,C/RAN-P]
VON GUNDEN, KENNETH  CRY WOLF [@K-9 CORPS 3]  P
VON GUNDEN, KENNETH  K-9 CORPS [K-9 CORPS 1]  P
VON GUNDEN, KENNETH  PALE COMPANION, THE  C/P,P
VON GUNDEN, KENNETH  SOUNDING STILLNESS, THE  C/P,P
VON GUNDEN, KENNETH  STARSPAWN  P
VON GUNDEN, KENNETH  UNDER FIRE [@K-9 CORPS]  C/P,P
VON HARBOU, THEA  METROPOLIS  FP,C/P:2,P
VON HARBOU, THEA  METROPOLIS (BRITISH)  SR-P
VON KORFF, FREYA GRAFIN  REISE NACH ANTARIA, DIE  FH

VON KUEHNELT-LEDDIHN, ERIK R. CHRISTIANE MOSCOW 1979

SR-H

VON SCHMIDT, CARLOS  SCIENCE FICTION: AUTORES SELECIONADOS  FH
VON STASSBURG, GOTTFRIED  TRISTAN  P
VONARBURG, ELISABETH  DREAMS OF THE SEA  LP
VONARBURG, ELISABETH  IN THE MOTHER’S LAND  P
VONARBURG, ELISABETH  RELUCTANT VOYAGERS  P
VONARBURG, ELISABETH  SILENT CITY, THE  C/P,P
VONARBURG, ELISABETH/BRIERLEY, JANE  TESSERACTS Q  C/PA,PA
VONNEGUT, KURT JR.  BAGOMBO SNUFF BOX  H
VONNEGUT, KURT JR.  BETWEEN TIME AND TIMBUKTU, OR PROMETHEUS-5  H,LP
VONNEGUT, KURT JR.  BLUEBEARD  H,P
VONNEGUT, KURT JR.  BREAKFAST OF CHAMPIONS  H:2,P
VONNEGUT, KURT JR.  CAT’S CRADLE  H:2,P
VONNEGUT, KURT JR.  DEADEYE DICK  H,P
VONNEGUT, KURT JR.  GALAPAGOS  H:2,P
VONNEGUT, KURT JR.  GOD BLESS YOU, MR. ROSEWATER, OR PEARLS BEFORE SWINE  H:2,P
VONNEGUT, KURT JR.  HAPPY BIRTHDAY, WANDA JUNE  LP
VONNEGUT, KURT JR.  HOCUS POCUS  P
VONNEGUT, KURT JR.  MOTHER NIGHT  C/P:3,P
VONNEGUT, KURT JR.  PLAYER PIANO  H:2,P
VONNEGUT, KURT JR.  SIRENS OF TITAN, THE  C/P:2,P
VONNEGUT, KURT JR.  SLAPSTICK  P
VONNEGUT, KURT JR.  SLAUGHTERHOUSE-FIVE, OR THE CHILDREN’S CRUSADE  C/P:2,P
VONNEGUT, KURT JR.  TIMEQUAKE  LP
VONNEGUT, KURT JR.  UTOPIA 14  C/P:2,P
VONNEGUT, KURT JR.  WAMPETERS, FOMA, AND GRANFALLOONS (OPINIONS)  R/REF-P
VONNEGUT, KURT JR.  WELCOME TO THE MONKEY HOUSE  C/P:2,P
VORHIES, JOHN  PRE-EMPT  P
VORNHOLT, JOHN  AFTERSHOCK [@STAR TREK STARFLEET ACADEMY #2]  LP
VORNHOLT, JOHN  ANTIMATTER [STAR TREK DS9]  C/P:2,P
VORNHOLT, JOHN  AVATARS, THE [@PRIMAL RAGE]  P
VORNHOLT, JOHN  BEHIND ENEMY LINES [@STAR TREK DOMINION WAR #1/STAR TREK TNG]  P
VORNHOLT, JOHN  BLOOD OATH [@BABYLON 5 #3]  C/P,P
VORNHOLT, JOHN  CAPTURE THE FLAG [@STAR TREK TNG STARFLEET ACADEMY #4]  LP
VORNHOLT, JOHN  CONTAMINATION [STAR TREK TNG]  C/P:2,P
VORNHOLT, JOHN  COYOTE MOON [BUFFY]  P
VORNHOLT, JOHN  CROSSFIRE [@STAR TREK TNG STARFLEET ACADEMY #11]  LP
VORNHOLT, JOHN  FABULIST, THE  P,SFWA-P
VORNHOLT, JOHN  GEMWORLD, BOOK ONE [@STAR TREK TNG]  C/P,P
VORNHOLT, JOHN  GEMWORLD, BOOK TWO [@STAR TREK TNG]  C/P,P
VORNHOLT, JOHN  GENESIS FORCE [STAR TREK TNG]  P
VORNHOLT, JOHN  GENESIS WAVE BOOK ONE, THE [STAR TREK TNG]  P
VORNHOLT, JOHN  GENESIS WAVE BOOK TWO, THE [STAR TREK TNG]  P
VORNHOLT, JOHN  GENESIS WAVE BOOK THREE, THE [STAR TREK TNG]  H,P
VORNHOLT, JOHN  LEATHER WINGS [EARTH 2]  P
VORNHOLT, JOHN  MASKS [STAR TREK TNG]  C/P:2,P
VORNHOLT, JOHN  MIND MELD [STAR TREK]  C/P,P
VORNHOLT, JOHN  NEMESIS [STAR TREK TNG]  LP
VORNHOLT, JOHN  PRISONER OF CABIN 13 [@SABRINA THE TEENAGE WITCH #11]  P
VORNHOLT, JOHN  QUARANTINE [@STAR TREK TNG DOUBLE HELIX 4]  P
VORNHOLT, JOHN  ROGUE SAUCER [STAR TREK TNG]  C/P:2,P
VORNHOLT, JOHN  SANCTUARY [STAR TREK]  C/P:3,P
VORNHOLT, JOHN  SEVEN CROWN [@BUFFY/ANGEL]  C/P,P
VORNHOLT, JOHN  TIME TO BE BORN, A [@STAR TREK TNG TIME TO 1]  P
VORNHOLT, JOHN  TIME TO DIE, A [@STAR TREK TNG TIME TO 2]  P

The MITSFS Pinkdex  September 24, 2007  Page 400
AUTHOR  TITLE  CODE

VORNHOLT, JOHN TUNNEL THROUGH THE STARS [@STAR TREK DOMINION WAR #3/

STAR TREK TNG]

P

VORNHOLT, JOHN  VOICES [@BABYLON 5 #1]  C/P:2,P
VORNHOLT, JOHN  WAR DRUMS [STAR TREK TNG]  C/P:2,P
VORNHOLT, JOHN  WITCHOPOLY [@SABRINA THE TEENAGE WITCH #22]  P
VOSKRESENSKAYA, E./KLYUEVA, V.  FANTASTIKA 1965 VYPUSK I  FH
VUKCEVICH, RAY  MEET ME IN THE MOON ROOM  LP

The MITSFS Pinkdex  September 24, 2007  Page 401
AUTHOR  TITLE  CODE
WADDELL, PATRICIA  TRUE BLOOD  P
WADE, ELIZABETH  ASHES OF TAMAR, THE  P
WADEY, VICTOR  UNITED PLANETS, THE  H
WAGAR, W. WARREN  SHORT HISTORY OF THE FUTURE, A (SECOND ED.)  LP
WAGENKNECHT, EDWARD  SIX NOVELS OF THE SUPERNATURAL  HA
WAGGONER, DIANA  HILLS OF FARAWAY: A GUIDE TO FANTASY, THE  R/REF-H
WAGGONER, TIM  FORGE OF THE MINDSLAYERS [@EBERRON BLADE OF THE FLAME #2]  P

WAGGONER, TIM RETURN OF THE SORCERESS [@DRAGONLANCE NEW ADVENTURES #4/

DRAGONLANCE SPELLBINDER QUARTET 4]

LP

WAGGONER, TIM TEMPLE OF THE DRAGONSLAYER

[@DRAGONLANCE NEW ADVENTURES #1/ DRAGONLANCE SPELLBINDER QUARTET 1]

LP

WAGGONER, TIM  THIEVES OF BLOOD [@EBERRON BLADE OF THE FLAME #1]  P
WAGNER, KARL EDWARD  BLOODSTONE  C/P:2,SFWA-P,P
WAGNER, KARL EDWARD  DARK CRUSADE  C/P:2,SFWA-P,P
WAGNER, KARL EDWARD  DARKNESS WEAVES  P
WAGNER, KARL EDWARD  DARKNESS WEAVES WITH MANY SHADES...  SR-P
WAGNER, KARL EDWARD  DEATH ANGEL’S SHADOW  C/P,P
WAGNER, KARL EDWARD  ECHOES OF VALOR  PA,SFWA-P
WAGNER, KARL EDWARD  ECHOES OF VALOR 2  HA,PA,SFWA-P
WAGNER, KARL EDWARD  ECHOES OF VALOR 3  C/PA,PA
WAGNER, KARL EDWARD  LEGION FROM THE SHADOWS [BRAN MAK MORN 2]  C/P:2,SFWA-P,P
WAGNER, KARL EDWARD  MIDNIGHT SUN  H
WAGNER, KARL EDWARD  NIGHT WINDS  C/P:2,P
WAGNER, KARL EDWARD  ROAD OF KINGS, THE [BANTAM CONAN 4]  P

WAGNER, KARL EDWARD UNTHREATENED BY THE MORNING LIGHT [@AUTHOR’S CHOICE MONTHLY #2]

SR-H

WAGNER, KARL EDWARD  WHY NOT YOU AND I?  SFWA-P
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES VIII, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES IX, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES X, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XI, THE  C/PA,PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XII, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XIII, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XIV, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XVI, THE  PA
WAGNER, KARL EDWARD  YEAR’S BEST HORROR STORIES XVII, THE  PA
WAGNER, ROBIN S.  INCREDIBLE SHRINKING MORK, THE [MORK AND MINDY 2]  C/P,P
WAGNER, SHARON  HOUSE OF DOOM, HOUSE OF DESIRE  P
WAGNER, SHARON  SHADES OF EVIL  P
WAHL, JAN  SOS BOBOMOBILE!  LP
WAHREN, FRIEDEL  ISAAC ASIMOVS SCIENCE FICTION MAGAZIN 25  FP
WAID, MARK/ROSS, ALEX  KINGDOM COME  CX
WAITMAN, KATIE  DIVIDED, THE  LP
WAITMAN, KATIE  MERRO TREE, THE  P
WAKEFIELD, H. R.  CLOCK STRIKES TWELVE, THE  H
WALDEN, MARK  BEYOND THE THIRTEENTH SUN  P
WALDROP, HOWARD  HEART OF WHITENESSE  SR-H
WALDROP, HOWARD  HOWARD WHO?  H,SFWA
WALDROP, HOWARD  NIGHT OF THE COOTERS  H,SFWA-P
WALDROP, HOWARD  STRANGE MONSTERS OF THE RECENT PAST  C/P,P
WALDROP, HOWARD  THEM BONES  C/P,P
WALKER, BARBARA G.  FEMINIST FAIRY TALES  LP
WALKER, DAVID  LORD’S PINK OCEAN, THE  C/P,P,H
WALKER, HUGH  ARMY OF DARKNESS [MAGIRA 2]  C/P,P
WALKER, HUGH  MESSENGERS OF DARKNESS [MAGIRA 3]  P
WALKER, HUGH  WAR-GAMERS’ WORLD [MAGIRA 1]  P
WALKER, IRMA  INHERIT THE EARTH  H
WALKER, IRMA  PORTAL TO E’EWERE  H
WALKER, JERRY  MISSION ACCOMPLISHED  H
WALKER, LARS  BLOOD AND JUDGMENT  P
WALKER, LARS  ERLING’S WORD  P
WALKER, LARS  WOLF TIME  P
WALKER, LARS  YEAR OF THE WARRIOR, THE  P
WALKER, MARY ALEXANDER  SCATHACH AND MAEVE’S DAUGHTERS, THE  SFWA
WALKER, MAX  CODE NAME: JUDAS [@MISSION IMPOSSIBLE #2]  P
WALKER, MAX  CODE NAME: RAPIER [@MISSION IMPOSSIBLE #3]  P
WALKER, PAUL  SPEAKING OF SCIENCE FICTION: THE PAUL WALKER INTERVIEWS  R/REF-H,C/REF-H
WALKER, ROBERT W.  SUB-ZERO  P
WALKER, SAGE  WHITEOUT  H,LP
WALL, MERVYN  UNFORTUNATE FURSEY, THE  H
WALLACE, BARBARA BROOKS  TROUBLE WITH MISS SWITCH, THE  P
WALLACE, EDGAR/COOPER, MERIAN C.  KING KONG  P
WALLACE, F. L.  ADDRESS: CENTAURI  H:2,P
WALLACE, IAN  CROYD  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 402
AUTHOR  TITLE  CODE
WALLACE, IAN  DEATHSTAR VOYAGE  C/P,P,H:2
WALLACE, IAN  DOCTOR ORPHEUS  H:2,P
WALLACE, IAN  HELLER’S LEAP  C/P:2,P
WALLACE, IAN  LUCIFER COMET, THE  C/P,P
WALLACE, IAN  MEGALOMANIA  P
WALLACE, IAN  PAN SAGITTARIUS  C/P,P,H
WALLACE, IAN  PURLOINED PRINCE, THE  H
WALLACE, IAN  RAPE OF THE SUN, THE  C/P,P
WALLACE, IAN  SIGN OF THE MUTE MEDUSA, THE  P
WALLACE, IAN  VOYAGE TO DARI, A  C/P:2,P
WALLACE, IAN  WORLD ASUNDER, THE  C/P,P
WALLACE, IAN  Z-STING  C/P,P
WALLACE, MAY NICKERSON  GHOST OF DIBBLE HOLLOW, THE  P
WALLACE, NICK  FEAR ITSELF [DOCTOR WHO PAST DOCTORS]  P
WALLACE, PATRICIA  TAINT, THE  P
WALLACE, SEAN  BEST NEW FANTASY  LPA

WALLER, MARK AERIAL/CHONG, HEMAN/ COSTINAS, COSMIN/HEATHER, ROSEMARY/ MCKEE, FRANCIS/REINFURT, DAVID/ RUSHTON, STEVE/TANGEN, LEIF MAGNE PHILIP

SR-P,C/P

WALLERSTEIN, JAMES S.  DEMON’S MIRROR, THE  H
WALLEY, CHRIS  DARK FOUNDATIONS, THE [@LAMB AMONG THE STARS #2]  H
WALLEY, CHRIS  SHADOW AND THE NIGHT, THE [@LAMB AMONG THE STARS #1]  H
WALLING, WILLIAM H.  NO ONE GOES THERE NOW  H,SFWA
WALLING, WILLIAM H.  WORLD I LEFT BEHIND ME, THE  H
WALLIS, DAVE  ONLY LOVERS LEFT ALIVE  C/P,P
WALLIS, G. MCDONALD  LEGEND OF LOST EARTH  DF187
WALLIS, G. MCDONALD  LIGHT OF LILITH, THE  DF108
WALLIS, JAMES  MARK OF DAMNATION [@WARHAMMER MARK OF CHAOS]  P
WALLIS, JAMES  MARK OF HERESY [@WARHAMMER]  P
WALLMANN, JEFFREY N.  DEATH TREK  P
WALLOP, DOUGLASS  MERMAID IN THE SWIMMING POOL, THE  H
WALPOLE, HORACE  CASTLE OF OTRANTO, THE  P
WALSH, J. M.  VANGUARD TO NEPTUNE  SR-P
WALSH, JILL PATON  CHANCE CHILD, A  P
WALTER, ELIZABETH  IN THE MIST AND OTHER UNCANNY ENCOUNTERS  H
WALTER, ELIZABETH  SIN EATER, THE  H
WALTER, W. GREY  CURVE OF THE SNOWFLAKE, THE  H
WALTERS, HUGH  FIRST CONTACT?  H
WALTERS, HUGH  FIRST ON THE MOON  C/P:2,P
WALTERS, HUGH  JOURNEY TO JUPITER  H
WALTERS, HUGH  PASSAGE TO PLUTO  H
WALTERS, HUGH  SPACESHIP TO SATURN  H
WALTERS, NICK  DOMINION [DOCTOR WHO EIGHTH DOCTOR]  C/P,P
WALTERS, NICK  FALL OF YQUATINE, THE [DOCTOR WHO EIGHTH DOCTOR]  P
WALTERS, NICK  RECKLESS ENGINEERING [DOCTOR WHO EIGHTH DOCTOR]  P
WALTERS, NICK  SUPERIOR BEINGS [DOCTOR WHO PAST DOCTORS]  P
WALTERS, R. R.  LILY  SFWA-P
WALTERS, R. R.  LUDLOW’S MILL  P
WALTERS, R. R.  RITUAL, THE  P
WALTHER, DANIEL  BOOK OF SHAI, THE  P
WALTHER, DANIEL  SHAI’S DESTINY  P
WALTON, BARBARA E.  ODYSSEY [@QUANTUM LEAP]  P
WALTON, BRYCE  SONS OF THE OCEAN DEEPS  H:2
WALTON, EVANGELINE  CHILDREN OF LLYR, THE [MABINOGI 2]  C/P:2,P
WALTON, EVANGELINE  ISLAND OF THE MIGHTY, THE [MABINOGI 4]  C/P:2,P
WALTON, EVANGELINE  PRINCE OF ANNWN [MABINOGI 1]  C/P:2,P
WALTON, EVANGELINE  SONG OF RHIANNON, THE [MABINOGI 3]  C/P,SFWA,P
WALTON, EVANGELINE  SWORD IS FORGED, THE  P
WALTON, EVANGELINE  WITCH HOUSE  SR-P,C/P,H
WALTON, JO  FARTHING  H
WALTON, JO  KING’S NAME, THE [TIR TANAGIRI 2]  H
WALTON, JO  KING’S PEACE, THE [TIR TANAGIRI 1]  H
WALTON, JO  PRIZE IN THE GAME, THE [TIR TANAGIRI 3]  H,P
WALTON, JO  PRIZE IN THE GAME, THE (PROOF) [TIR TANAGIRI 3]  SR-LP
WALTON, JO  TOOTH AND CLAW  C/P,P
WANDREI, DONALD  STRANGE HARVEST  SR-H
WANGERIN, WALTER JR.  BOOK OF SORROWS, THE [DUN COW 2]  H
WANGERIN, WALTER JR.  BOOK OF THE DUN COW, THE [DUN COW 1]  C/P:2,P
WARD, CHRISTOPHER  GENTLEMAN INTO GOOSE  SR-H
WARD, DAYTON/DILMORE, KEVIN  FOUNDATIONS [@STAR TREK S.C.E. 5]  P
WARD, DAYTON  GENESIS PROTOCOL, THE  LP
WARD, DAYTON  IN THE NAME OF HONOR [STAR TREK]  P
WARD, DAYTON  LAST WORLD WAR, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 403
AUTHOR  TITLE  CODE
WARD, DAYTON/DILMORE, KEVIN  SUMMON THE THUNDER [STAR TREK VANGUARD]  P
WARD, DAYTON/DILMORE, KEVIN  TIME TO HARVEST, A [@STAR TREK TNG TIME TO 4]  P
WARD, DAYTON/DILMORE, KEVIN  TIME TO SOW, A [@STAR TREK TNG TIME TO 3]  P
WARD, DON  BLACK MAGIC  C/PA,PA
WARD, DON  H. RIDER HAGGARD’S SHE  SR-P
WARD, JAMES M.  DRAGONFRIGATE WIZARD HALCYON BLITHE [HALCYON BLITHE 2]  H
WARD, JAMES M./HONG, JANE COOPER  POOL OF RADIANCE [@FORGOTTEN REALMS]  C/P:2,P
WARD, JAMES M./BROWN, ANNE K.  POOL OF TWILIGHT [@FORGOTTEN REALMS]  C/P,P
WARD, JAMES M./BROWN, ANNE K.  POOLS OF DARKNESS [@FORGOTTEN REALMS]  C/P,P
WARD, JAMES S./WISE, DAVID  PALADINS, THE [@FORGOTTEN REALMS DOUBLE DIAMOND TRIANGLE #2]  P
WARNER BOOKS  FIVE FATES  C/PA,PA
WARNER-CROZETTI, R.  WIDDERBURN HORROR, THE  P
WARNER, DOUGLAS  DEATH ON A WARM WIND  C/P,P
WARNER, HARRY JR.  ALL OUR YESTERDAYS  R/REF-H
WARNER, R. D./WARNER, MICHAEL  GALACTIC RIFT  P
WARNER, SYLVIA TOWNSEND  KINGDOMS OF ELFIN  H:2,LP
WARNER, WILLIAM HENRY  BRIDGE OF TIME, THE  H
WARREN, CHRISTINE  DEMON YOU KNOW, THE [OTHERS (WARREN)]  P
WARREN, CHRISTINE  SHE’S NO FAERIE PRINCESS [OTHERS (WARREN) 2]  P
WARREN, CHRISTINE  WOLF AT THE DOOR [OTHERS (WARREN) 1]  P
WARREN, GEORGE  DOMINANT SPECIES  H,LP
WARREN, JAMES  BEST OF CREEPY, THE  SCX
WARRICK, PATRICIA S.  CYBERNETIC IMAGINATION IN SCIENCE FICTION, THE  R/REF-H
WARRICK, PATRICIA S.  MIND IN MOTION  R/REF-H

WARRICK, PATRICIA S./ GREENBERG, MARTIN H.

NEW AWARENESS: RELIGION THROUGH SCIENCE FICTION, THE

HA

WARRICK, PATRICIA S./ GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

SCIENCE FICTION: CONTEMPORARY MYTHOLOGY

HA,SFWA

WARRICK, PATRICIA S./WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SCIENCE FICTION:

THE SCIENCE FICTION RESEARCH ASSOCIATION ANTHOLOGY

LPA

WARRINGTON, FREDA  AMBER CITADEL, THE [JEWELFIRE 1]  P
WARRINGTON, FREDA  BLACKBIRD IN SILVER, A  LP
WARRINGTON, FREDA  OBSIDIAN TOWER, THE [JEWELFIRE 3]  P
WARRINGTON, FREDA  SAPPHIRE THRONE, THE [JEWELFIRE 2]  P
WARSHOFSKY, FRED  TARGET MOON  [C/RAN-H]
WARTOFSKY, VICTOR  PASSAGE, THE  C/P,P
WASHBURN, MARK  OMEGA THREAT, THE  P
WATCHTOWER  DID MAN GET HERE BY EVOLUTION OR BY CREATION?  [C/RAN-H]
WATCHTOWER  EVOLUTION VERSUS THE NEW WORLD  [C/RAN-P]
WATCHTOWER  "THEN IS FINISHED THE MYSTERY OF GOD"  [C/RAN-H]
WATER, SILAS  MAN WITH ABSOLUTE MOTION, THE  SR-P
WATERS, ELISABETH  CHANGING FATE  P
WATERS, T. A.  CENTERFORCE  P
WATERS, T. A.  PROBABILITY PAD, THE  P
WATKIN, LAWRENCE EDWARD  ON BORROWED TIME  H
WATKINS, WILLIAM JOHN  CENTRIFUGAL RICKSHAW DANCER, THE  P
WATKINS, WILLIAM JOHN  CLICKWHISTLE  SFWA
WATKINS, WILLIAM JOHN  COSMIC THUNDER  P

WATKINS, WILLIAM JOHN/ SNYDER, EUGENE V.

ECODEATH

H,SFWA

WATKINS, WILLIAM JOHN  GOD MACHINE, THE  H,SFWA
WATKINS, WILLIAM JOHN  GOING TO SEE THE END OF THE SKY  P
WATKINS, WILLIAM JOHN  LAST DEATHSHIP OFF ANTARES, THE  C/P,P

WATKINS, WILLIAM JOHN/ SNYDER, EUGENE V.

LITANY OF SH’REEV, THE

H,SFWA

WATKINS, WILLIAM JOHN  WHAT ROUGH BEAST  C/P:2,P
WATSIN, JUDE  STAR WARS EPISODE 1 JOURNAL: DARTH MAUL [STAR WARS]  LP

WATSON, DICK, ELLIOTT, JOHN/ STROME, DICK ZACHERLY FOR PRESIDENT

[C/RAN-H]

WATSON, IAN  ALIEN EMBASSY  C/P,P
WATSON, IAN  BOOK OF BEING, THE  P
WATSON, IAN  BOOK OF IAN WATSON, THE  H
WATSON, IAN  BOOK OF THE RIVER, THE  P
WATSON, IAN  BOOK OF THE STARS, THE  P
WATSON, IAN  BOOKS OF THE BLACK CURRENT, THE  H:2
WATSON, IAN  CHAOS CHILD [WARHAMMER INQUISITION WAR 3]  P
WATSON, IAN  CHEKHOV’S JOURNEY  H,P
WATSON, IAN  DEATHHUNTER  C/P,P,H
WATSON, IAN  DRACO [WARHAMMER INQUISITION WAR 1]  P
WATSON, IAN  EMBEDDING, THE  C/P,SFWA,P,H
WATSON, IAN  FLIES OF MEMORY, THE  H
WATSON, IAN  GARDENS OF DELIGHT, THE  C/P:2,P
WATSON, IAN  HARLEQUIN [WARHAMMER INQUISITION WAR 2]  P

The MITSFS Pinkdex  September 24, 2007  Page 404
AUTHOR  TITLE  CODE
WATSON, IAN  INQUISITION WAR, THE [@WARHAMMER INQUISITION WAR 1,2,3]  LP
WATSON, IAN  JONAH KIT, THE  H,P
WATSON, IAN  MARTIAN INCA, THE  C/P,P,H
WATSON, IAN  MIRACLE VISITORS  C/P,P
WATSON, IAN  MOCKYMEN  H
WATSON, IAN  ORACLE  P
WATSON, IAN  PICTURES AT AN EXHIBITION  SR-PA
WATSON, IAN  QUEENMAGIC, KINGMAGIC  H,P
WATSON, IAN  VERY SLOW TIME MACHINE, THE  C/P:2,P
WATSON, JANE WERNER  DINOSAURS  [C/RAN-H]
WATSON, JUDE  RETURN OF THE DARK SIDE [@STAR WARS LAST OF THE JEDI #6]  LP
WATSON, JULES  DAWN STAG, THE [DALRIADA 2]  H
WATSON, JULES  WHITE MARE, THE [DALRIADA 1]  LP
WATSON, PATRICK  ALTER EGO  H
WATT-EVANS, LAWRENCE  BLOOD OF A DRAGON, THE [ETHSHAR]  C/P:2,P
WATT-EVANS, LAWRENCE  BOOK OF SILENCE, THE [LORDS OF DUS 4]  C/P:2,P
WATT-EVANS, LAWRENCE  CELESTIAL DEBRIS  LP
WATT-EVANS, LAWRENCE  CHROMOSOMAL CODE, THE  C/P,P
WATT-EVANS, LAWRENCE  CROSSTIME TRAFFIC  C/P,P
WATT-EVANS, LAWRENCE  CYBORG AND THE SORCERERS, THE [WAR SURPLUS 1]  C/P:2,P
WATT-EVANS, LAWRENCE  DENNER’S WRECK  C/P,P,H:2
WATT-EVANS, LAWRENCE  DRAGON SOCIETY, THE [OBSIDIAN CHRONICLES 2]  H
WATT-EVANS, LAWRENCE  DRAGON VENOM [OBSIDIAN CHRONICLES 3]  H
WATT-EVANS, LAWRENCE  DRAGON WEATHER [OBSIDIAN CHRONICLES 1]  H
WATT-EVANS, LAWRENCE  IN THE EMPIRE OF SHADOW [THREE WORLDS 2]  LP
WATT-EVANS, LAWRENCE  ITHANALIN’S RESTORATION [ETHSHAR]  H
WATT-EVANS, LAWRENCE  LORDS OF DUS, THE [LORDS OF DUS 1,2,3,4]  H
WATT-EVANS, LAWRENCE  LURE OF THE BASILISK, THE [LORDS OF DUS 1]  C/P:2,P
WATT-EVANS, LAWRENCE  MISENCHANTED SWORD, THE [ETHSHAR]  C/P:2,P
WATT-EVANS, LAWRENCE  NEWER YORK  PA
WATT-EVANS, LAWRENCE  NIGHT OF MADNESS [ETHSHAR]  H
WATT-EVANS, LAWRENCE  NIGHTMARE PEOPLE, THE  LP
WATT-EVANS, LAWRENCE  NIGHTSIDE CITY  P
WATT-EVANS, LAWRENCE  NINTH TALISMAN, THE [CHOSEN 2]  H
WATT-EVANS, LAWRENCE  OUT OF THIS WORLD [THREE WORLDS 1]  H:2,P
WATT-EVANS, LAWRENCE  REBIRTH OF WONDER, THE  C/P,P
WATT-EVANS, LAWRENCE  REIGN OF THE BROWN MAGICIAN, THE [THREE WORLDS 3]  P
WATT-EVANS, LAWRENCE  SEVEN ALTARS OF DUSARRA, THE [LORDS OF DUS 2]  C/P:2,P
WATT-EVANS, LAWRENCE  SHINING STEEL  C/P,P

WATT-EVANS, LAWRENCE/ PARLAGRECO, CARL SPARTACUS FILE, THE

LP

WATT-EVANS, LAWRENCE  SPELL OF THE BLACK DAGGER, THE [ETHSHAR]  C/P:2,P

WATT-EVANS, LAWRENCE/ FRIESNER, ESTHER M.

SPLIT HEIRS

C/P,P,H

WATT-EVANS, LAWRENCE  SPRIGGAN MIRROR, THE [ETHSHAR]  LP
WATT-EVANS, LAWRENCE  SWORD OF BHELEU, THE [LORDS OF DUS 3]  C/P:2,P
WATT-EVANS, LAWRENCE  TAKING FLIGHT [ETHSHAR]  P
WATT-EVANS, LAWRENCE  TOUCHED BY THE GODS  H,P
WATT-EVANS, LAWRENCE  UNWILLING WARLORD, THE [ETHSHAR]  P
WATT-EVANS, LAWRENCE  WITH A SINGLE SPELL [ETHSHAR]  C/P:2,P
WATT-EVANS, LAWRENCE  WIZARD AND THE WAR MACHINE, THE [WAR SURPLUS 2]  C/P,P
WATT-EVANS, LAWRENCE  WIZARD LORD, THE [CHOSEN 1]  H
WATT-EVANS, LAWRENCE  WORLDS OF SHADOW [THREE WORLDS 1,2,3]  H
WATTS, PETER  BEHEMOTH: B-MAX [RIFTERS 3.0]  H
WATTS, PETER  BEHEMOTH: SEPPUKU [RIFTERS 3.1]  H
WATTS, PETER  BLINDSIGHT  H:2
WATTS, PETER  MAELSTROM [RIFTERS 2]  H,P
WATTS, PETER  STARFISH [RIFTERS 1]  H,P

WAUGH, CAROL-LYNN ROSSEL/ GREENBERG, MARTIN/ASIMOV, ISAAC PURRFECT CRIME [PURRFECT SERIES]

PA

WAUGH, CAROL-LYNN ROSSEL/ GREENBERG, MARTIN H./ASIMOV, ISAAC SHOW BUSINESS IS MURDER

PA

WAUGH, CAROL-LYNN ROSSEL/ GREENBERG, MARTIN H./ASIMOV, ISAAC 13 HORRORS OF HALLOWEEN

PA

WAUGH, CAROL-LYNN ROSSEL/ GREENBERG, MARTIN H./ASIMOV, ISAAC TWELVE CRIMES OF CHRISTMAS, THE

PA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

BOOTCAMP 3000

PA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

DANGEROUS VEGETABLES

C/PA,PA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

LOVE 3000

HA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

MAMMOTH BOOK OF FANTASTIC SCIENCE FICTION, THE [@MAMMOTH SF 5] HA,LPA

The MITSFS Pinkdex  September 24, 2007  Page 405
AUTHOR  TITLE  CODE

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

MAMMOTH BOOK OF MODERN SCIENCE FICTION, THE [@MAMMOTH SF 6]

LPA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

MAMMOTH BOOK OF VINTAGE SCIENCE FICTION, THE [@MAMMOTH SF 3]

LPA

WAUGH, CHARLES G./ GREENBERG, MARTIN H./ OLANDER, JOSEPH D.

MYSTERIOUS VISIONS

HA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

ROBOT WARRIORS

PA,C/PA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SCI-FI PRIVATE EYE

PA

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

SPACE WARS

PA,SFWA-P

WAUGH, CHARLES G./ GREENBERG, MARTIN H./ MCSHERRY, FRANK D. JR.

STRANGE MAINE

LP

WAUGH, CHARLES G./ GREENBERG, MARTIN H.

TIME WARS

C/PA,PA,SFWA-P

WAUGH, SYLVIA  MENNYMS IN THE WILDERNESS [MENNYMS]  P
WAYMAN, TONY RUSSELL  ADS INFINITUM  P
WAYMAN, TONY RUSSELL  DUNES OF PRADAI  P
WAYMAN, TONY RUSSELL  WORLD OF THE SLEEPER  DH021
WEAVER, BEN  BROTHERS IN ARMS [SCOTT ST. ANDREW 1]  P
WEAVER, BEN  PATRIOTS IN ARMS [SCOTT ST. ANDREW 3]  P
WEAVER, BEN  REBELS IN ARMS [SCOTT ST. ANDREW 2]  P
WEAVER, HENRIETTA  FLAME AND THE SHADOW-EATER  SR-H
WEAVER, MICHAEL D.  BLOODFANG [BLOODFANG 3]  P
WEAVER, MICHAEL D.  MERCEDES NIGHTS  H,P
WEAVER, MICHAEL D.  MY FATHER IMMORTAL  H,P
WEAVER, MICHAEL D.  NIGHTREAVER [BLOODFANG 2]  P
WEAVER, MICHAEL D.  SECOND INFINITY, A  P
WEAVER, MICHAEL D.  WOLF-DREAMS [BLOODFANG 1]  P
WEBB, JEAN FRANCIS  KING SOLOMON’S MINES  SR-P
WEBB, SHARON  ADVENTURES OF TERRA TARKINGTON, THE  C/P:2,P
WEBB, SHARON  EARTH SONG [EARTH SONG 2]  C/P,P,H
WEBB, SHARON  EARTHCHILD [EARTH SONG 1]  C/P,P,H
WEBB, SHARON  PESTIS 18  C/P,P
WEBB, SHARON  RAM SONG [EARTH SONG 3]  C/P,SFWA,P,H
WEBB, STEPHEN  WHERE IS EVERYBODY?  R/REF-H
WEBER, DAVID  APOCALYPSE TROLL, THE  C/P,P,H
WEBER, DAVID  ARMAGEDDON INHERITANCE, THE [COLIN MACINTYRE 2]  C/P,SFWA-P,P
WEBER, DAVID  ASHES OF VICTORY [HONOR HARRINGTON 9]  H:2,P
WEBER, DAVID  AT ALL COSTS [HONOR HARRINGTON 11]  H
WEBER, DAVID  CHANGER OF WORLDS [HONOR HARRINGTON/WORLDS OF HONOR 3]  HA
WEBER, DAVID/FLINT, ERIC  CROWN OF SLAVES [HONOR HARRINGTON]  H
WEBER, DAVID/WHITE, STEVE  CRUSADE [STARFIRE 2]  C/P:2,SFWA-P,P
WEBER, DAVID  ECHOES OF HONOR [HONOR HARRINGTON 8]  H:2,P
WEBER, DAVID  EMPIRE FROM THE ASHES [COLIN MACINTYRE 1,2,3]  H,SR-LP
WEBER, DAVID  EXCALIBUR ALTERNATIVE, THE [RANKS OF BRONZE]  H,P
WEBER, DAVID  FIELD OF DISHONOR [HONOR HARRINGTON 4]  C/P:2,P
WEBER, DAVID  FLAG IN EXILE [HONOR HARRINGTON 5]  C/P,P,H
WEBER, DAVID  HEIRS OF EMPIRE [COLIN MACINTYRE 3]  C/P:2,P
WEBER, DAVID/EVANS, LINDA  HELL HATH NO FURY  H
WEBER, DAVID/EVANS, LINDA  HELL’S GATE [MULTIVERSE 1]  H
WEBER, DAVID  HONOR AMONG ENEMIES [HONOR HARRINGTON 6]  H,P
WEBER, DAVID  HONOR OF THE QUEEN, THE [HONOR HARRINGTON 2]  C/P:2,SFWA-P,P
WEBER, DAVID/WHITE, STEVE  IN DEATH GROUND [STARFIRE 3]  C/P,P
WEBER, DAVID  IN ENEMY HANDS [HONOR HARRINGTON 7]  H:2,P
WEBER, DAVID  IN FURY BORN  H
WEBER, DAVID/WHITE, STEVE  INSURRECTION [STARFIRE 1]  C/P:2,SFWA-P,P
WEBER, DAVID/RINGO, JOHN  MARCH TO THE SEA [MARCH UPCOUNTRY 2]  H,P
WEBER, DAVID/RINGO, JOHN  MARCH TO THE STARS [MARCH UPCOUNTRY 3]  H,P
WEBER, DAVID/RINGO, JOHN  MARCH UPCOUNTRY [MARCH UPCOUNTRY 1]  H
WEBER, DAVID  MORE THAN HONOR [HONOR HARRINGTON/WORLDS OF HONOR 1]  C/PA,PA
WEBER, DAVID  MUTINEERS’ MOON [COLIN MACINTYRE 1]  C/P:2,SFWA-P,P
WEBER, DAVID  OATH OF SWORDS [HRADANI 2]  C/P,P
WEBER, DAVID  OFF ARMAGEDDON REEF  H
WEBER, DAVID  OLD SOLDIERS [BOLOS]  H
WEBER, DAVID  ON BASILISK STATION [HONOR HARRINGTON 1]  C/P:2,SFWA-P,P
WEBER, DAVID  PATH OF THE FURY  C/P:2,SFWA-P,P

WEBER, DAVID SERVICE OF THE SWORD, THE [HONOR HARRINGTON/

WORLDS OF HONOR 4]

HA

WEBER, DAVID  SHADOW OF SAGANAMI, THE [HONOR HARRINGTON SAGANAMI 1]  H
WEBER, DAVID/WHITE, STEVE  SHIVA OPTION, THE [STARFIRE 4]  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 406
AUTHOR  TITLE  CODE
WEBER, DAVID  SHORT VICTORIOUS WAR, THE [HONOR HARRINGTON 3]  C/P,SFWA-P,P
WEBER, DAVID/FLINT, ERIC  1633 [SIXTEEN THIRTY-TWO 2]  H
WEBER, DAVID/EVANS, LINDA  TRIUMPHANT, THE [@BOLOS #3]  P
WEBER, DAVID  WAR GOD’S OWN, THE [HRADANI 1]  C/P,P,H
WEBER, DAVID  WAR OF HONOR (HARDCOVER AND CD) [HONOR HARRINGTON 10]  H:2
WEBER, DAVID/RINGO, JOHN  WE FEW [MARCH UPCOUNTRY 4]  H
WEBER, DAVID  WIND RIDER’S OATH (HARDCOVER AND CD) [HRADANI 3]  H:2
WEBER, DAVID  WORLDS OF HONOR [HONOR HARRINGTON/WORLDS OF HONOR 2]  C/PA,HA,PA
WEBSTER, FRANK V.  BEN HARDY’S FLYING MACHINE  H
WEBSTER, JOSH  BECKONING, THE  P
WEDDLE, DAVID/LANG, JEFFREY  ABYSS [STAR TREK DS9 SECTION 31]  P
WEDEL, E./ROMANOV, A. S.  ROMANOV’S RUSSIAN-ENGLISH/ENGLISH-RUSSIAN DICTIONARY  DICT
WEHRSTEIN, KAREN  LION’S HEART [FIFTH MILLENNIUM]  C/P,P
WEHRSTEIN, KAREN  LION’S SOUL [FIFTH MILLENNIUM]  C/P:2,P
WEIGALL, ARTHUR  GARDEN OF PARADISE, THE  H
WEIN, ELIZABETH E.  WINTER PRINCE, THE  P
WEIN, LEN/WOLFMAN, MARV  MARVEL SUPERHEROES, THE  PA
WEIN, LEN/WOLFMAN, MARV  MAYHEM IN MANHATTAN [SPIDER-MAN]  P
WEIN, LEN/WOLFMAN, MARV/SILVA, JOSEPH STALKER FROM THE STARS [HULK]  C/P,P
WEIN, LEN  UNTOLD LEGEND OF THE BATMAN, THE  SFWA-P
WEINBAUM, STANLEY G.  BEST OF STANLEY G. WEINBAUM, THE  C/P:2,P
WEINBAUM, STANLEY G.  BLACK FLAME, THE  SR-H,C/P:2,P
WEINBAUM, STANLEY G.  DARK OTHER, THE  SR-H
WEINBAUM, STANLEY G.  MARTIAN ODYSSEY, A  SR-H,C/P:2,P
WEINBAUM, STANLEY G.  NEW ADAM, THE  C/P:2,P
WEINBAUM, STANLEY G.  ODYSSEA NA MARSU  FP
WEINBAUM, STANLEY G.  PARASITE PLANET  B
WEINBAUM, STANLEY G.  RED PERI, THE  SR-H,SR-LP
WEINBERG, GEORGE  NUMBERLAND  H
WEINBERG, LARRY  CURSE, THE  P
WEINBERG, LAWRENCE  TRON: THE STORYBOOK  H
WEINBERG, ROBERT  ANNOTATED GUIDE TO ROBERT E. HOWARD’S SWORD AND SORCERY, THE R/REF-H
WEINBERG, ROBERT  ARMAGEDDON BOX, THE  P
WEINBERG, ROBERT  ASCENSION WARRIOR, THE [@HORIZON WAR #2/MAGE: THE ASCENSION]  P

WEINBERG, ROBERT BIOGRAPHICAL DICTIONARY OF SCIENCE FICTION AND FANTASY

ARTISTS, A

R/REF-H

WEINBERG, ROBERT BLOODWAR

[@WORLD OF DARKNESS VAMPIRE MASQUERADE RED DEATH 1]

P

WEINBERG, ROBERT  CALCULATED MAGIC, A [LOGICAL MAGICIAN 2]  C/P,P
WEINBERG, ROBERT  DR. SATAN  SR-LPA
WEINBERG, ROBERT/REIN-HAGEN, MARK  EMBRACE, THE [WORLD OF DARKNESS VAMPIRE DIARY]  SR-H
WEINBERG, ROBERT  FAR BELOW AND OTHER HORRORS  HA
WEINBERG, ROBERT/MCKINSTRY, LOHR  HERO PULP INDEX, THE  SRL/IND
WEINBERG, ROBERT  LOGICAL MAGICIAN, A [LOGICAL MAGICIAN 1]  C/P:2,P
WEINBERG, ROBERT  LOST FANTASIES #4  SR-LPA
WEINBERG, ROBERT  LOST FANTASIES #5  SR-LPA
WEINBERG, ROBERT  LOST FANTASIES #6  SR-LPA
WEINBERG, ROBERT  LOST FANTASIES #8: THE LAKE OF LIFE  SR-LPA

WEINBERG, ROBERT/ DZIEMIANOWICZ, STEFAN R./ GREENBERG, MARTIN H.

MISTS FROM BEYOND, THE

SFWA,SFWA-P

WEINBERG, ROBERT  MYSTERY OF THE DRAGON’S SHADOW, THE  SR-LP

WEINBERG, ROBERT/ DZIEMIANOWICZ, STEFAN R./ GREENBERG, MARTIN H.

100 VICIOUS LITTLE VAMPIRE STORIES

HA

WEINBERG, ROBERT  REVELRY IN HELL  SR-LPA
WEINBERG, ROBERT  ROAD TO HELL, THE [@HORIZON WAR #1/MAGE: THE ASCENSION]  P
WEINBERG, ROBERT  SIN EATERS, THE  SR-LPA
WEINBERG, ROBERT  SLAVES OF THE BLOODWOLVES  SR-LPA
WEINBERG, ROBERT  WAR IN HEAVEN [@HORIZON WAR #3/MAGE: THE ASCENSION]  P
WEINBERG, ROBERT  WEIRD MENACE #1  SR-LPA
WEINBERG, ROBERT  WT50: A TRIBUTE TO WEIRD TALES  R/REF-H
WEINER, ANDREW  STATION GEHENNA  P
WEINER, ELLIS  HOWARD THE DUCK  C/P,P
WEINER, ELLIS  NATIONAL LAMPOON’S DOON  C/P:2,P
WEINSTEIN, HOWARD  BETTER MAN, THE [STAR TREK]  C/P:2,P
WEINSTEIN, HOWARD  COVENANT OF THE CROWN, THE [STAR TREK]  C/P,P,H
WEINSTEIN, HOWARD  DEEP DOMAIN [STAR TREK]  C/P:3,P
WEINSTEIN, HOWARD/CRISPIN, A. C.  EAST COAST CRISIS [V]  C/P,P
WEINSTEIN, HOWARD  EXILES [STAR TREK TNG]  C/P:2,P
WEINSTEIN, HOWARD  PATH TO CONQUEST [V]  C/P,SFWA-P,P
WEINSTEIN, HOWARD  PERCHANCE TO DREAM [STAR TREK TNG]  C/P:2,P
WEINSTEIN, HOWARD  POWER HUNGRY [STAR TREK TNG]  C/P:3,P
WEINSTEIN, HOWARD  PRISONERS AND PAWNS [V]  C/P,SFWA-P,P

AUTHOR  TITLE  CODE
WEINSTEIN, SOL/ALBRECHT, HOWARD  EXERCISER, THE  P
WEINSTEIN, SOL/ALBRECHT, HOWARD  JONATHAN SEGAL CHICKEN  [C/RAN-P]
WEINSTEIN, SOL  LOXFINGER  LP
WEINSTEIN, SOL  ON THE SECRET SERVICE OF HIS MAJESTY, THE QUEEN  LP
WEIS, MARGARET  AMBER AND ASHES [@DRAGONLANCE DARK DISCIPLE #1]  H
WEIS, MARGARET  AMBER AND IRON [@DRAGONLANCE DARK DISCIPLE #2]  H
WEIS, MARGARET/HICKMAN, TRACY  BEST OF TALES, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  BEST OF TALES II, THE [@DRAGONLANCE]  C/PA,PA
WEIS, MARGARET/PERRIN, DON  BROTHERS IN ARMS [@DRAGONLANCE RAISTLIN 2]  P
WEIS, MARGARET/BALDWIN, DAVID  DARK HEART  P
WEIS, MARGARET/PERRIN, DON  DOOM BRIGADE, THE [@DRAGONLANCE CHAOS WAR 1]  P
WEIS, MARGARET/HICKMAN, TRACY  DOOM OF THE DARKSWORD [@DARKSWORD #2]  C/P:2,P
WEIS, MARGARET  DRAGON-LOVER’S TREASURY OF THE FANTASTIC, A  C/PA,LPA
WEIS, MARGARET  DRAGON-LOVER’S TREASURY OF THE FANTASTIC, A (2 CASSETTES)  SR-MM
WEIS, MARGARET/HICKMAN, TRACY  DRAGON WING [@DEATH GATE #1]  H:2,P
WEIS, MARGARET/HICKMAN, TRACY  DRAGONLANCE: THE ANNOTATED CHRONICLES [DRAGONLANCE 1,2,3]  LP

WEIS, MARGARET/HICKMAN, TRACY DRAGONLANCE: THE ANNOTATED LEGENDS [DRAGONLANCE CHRONICLES 1,2,3]

LP

WEIS, MARGARET/HICKMAN, TRACY  DRAGONS AT WAR, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS IN THE ARCHIVES [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF A FALLEN SUN [DRAGONLANCE WAR OF SOULS 1]  H
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF A LOST STAR [DRAGONLANCE WAR OF SOULS 2]  H
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF A VANISHD MOON [DRAGONLANCE WAR OF SOULS 3]  H
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF AUTUMN TWILIGHT [@DRAGONLANCE CHRONICLES #1]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF CHAOS, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF KRYNN, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF SPRING DAWNING [@DRAGONLANCE CHRONICLES #3]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF SUMMER FLAME [DRAGONLANCE CHRONICLES 4]  C/P:2,P

WEIS, MARGARET/HICKMAN, TRACY DRAGONS OF THE DWARVEN DEPTHS

[@DRAGONLANCE LOST CHRONICLES #1]

H

WEIS, MARGARET/HICKMAN, TRACY DRAGONS OF THE HIGHLORD SKIES

[@DRAGONLANCE LOST CHRONICLES #2]

H

WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF TIME [DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  DRAGONS OF WINTER NIGHT [@DRAGONLANCE CHRONICLES #2]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  ELVEN STAR [@DEATH GATE #2]  C/P,P,H
WEIS, MARGARET  FANTASTIC ALICE  LP
WEIS, MARGARET/HICKMAN, TRACY  FIRE SEA [@DEATH GATE #3]  H:2,P
WEIS, MARGARET/HICKMAN, TRACY  FORGING THE DARKSWORD [@DARKSWORD #1]  C/P:2,P
WEIS, MARGARET  GHOST LEGION [@STAR OF THE GUARDIANS #4]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  GUARDIANS OF THE LOST [@SOVEREIGN STONE #2]  H,P
WEIS, MARGARET/HICKMAN, TRACY  HAND OF CHAOS, THE [@DEATH GATE #5]  H:2,P
WEIS, MARGARET/HICKMAN, TRACY  HAND OF CHAOS, THE (PROOF) [DEATH GATE 5]  SR-LP
WEIS, MARGARET/HICKMAN, TRACY  HEROES AND FOOLS [@DRAGONLANCE FIFTH AGE]  C/PA,PA
WEIS, MARGARET/PERRIN, DON  HUNG OUT [STAR OF THE GUARDIANS MAG FORCE 7]  P
WEIS, MARGARET/HICKMAN, TRACY  INTO THE LABYRINTH [@DEATH GATE #6]  H:2,P
WEIS, MARGARET/HICKMAN, TRACY  JOURNEY INTO THE VOID [SOVEREIGN STONE 3]  H
WEIS, MARGARET/HICKMAN, TRACY  JOURNEY INTO THE VOID (PROOF) [SOVEREIGN STONE 3]  SR-LP
WEIS, MARGARET/HICKMAN, TRACY  KENDER, GULLY DWARVES, AND GNOMES [@DRAGONLANCE TALES #2]  PA,C/PA
WEIS, MARGARET  KING’S SACRIFICE [@STAR OF THE GUARDIANS #3]  C/P:2,P
WEIS, MARGARET  KING’S TEST [@STAR OF THE GUARDIANS #2]  C/P:2,P

WEIS, MARGARET/PERRIN, DON KNIGHTS OF THE BLACK EARTH, THE

[STAR OF THE GUARDIANS MAG FORCE 7]

C/P:2,P

WEIS, MARGARET/HICKMAN, TRACY  LEGACY OF THE DARK SWORD [DARKSWORD]  H
WEIS, MARGARET  LEGENDS [ETERNAL ARCHIVES #1]  PA
WEIS, MARGARET  LOST KING, THE [@STAR OF THE GUARDIANS #1]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  LOVE AND WAR [@DRAGONLANCE TALES #3]  PA,C/PA
WEIS, MARGARET  MAGIC-LOVER’S TREASURY OF THE FANTASTIC, A  LPA
WEIS, MARGARET/HICKMAN, TRACY  MAGIC OF KRYNN, THE [@DRAGONLANCE TALES #1]  PA,C/PA
WEIS, MARGARET/HICKMAN, TRACY  MANTLE OF KENDIS-DAI, THE [@STARSHIELD 1]  P
WEIS, MARGARET  MASTER OF DRAGONS [DRAGONVARLD 3]  H
WEIS, MARGARET  MISTRESS OF DRAGONS [DRAGONVARLD 1]  H
WEIS, MARGARET  NEW AMAZONS  C/PA,PA
WEIS, MARGARET/HICKMAN, TRACY  NIGHTSWORD [@STARSHIELD 3]  H
WEIS, MARGARET/HICKMAN, TRACY  PALADIN OF THE NIGHT, THE [@ROSE OF THE PROPHET #2]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  PLAYERS OF GILEAN, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  PROPHET OF AKHRAN, THE [@ROSE OF THE PROPHET #3]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  REBELS AND TYRANTS [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  RELICS AND OMENS [@DRAGONLANCE]  PA
WEIS, MARGARET/PERRIN, DON  ROBOT BLUES [STAR OF THE GUARDIANS MAG FORCE 7]  C/P,P
WEIS, MARGARET/HICKMAN, TRACY  SEARCH FOR MAGIC, THE [DRAGONLANCE WAR OF SOULS]  PA
WEIS, MARGARET  SEARCH FOR POWER, THE [@DRAGONLANCE]  PA
WEIS, MARGARET/HICKMAN, TRACY  SECOND GENERATION, THE [DRAGONLANCE]  C/P,P,H
WEIS, MARGARET/HICKMAN, TRACY  SENTINELS [@STARSHIELD 2]  H:2
WEIS, MARGARET/HICKMAN, TRACY  SERPENT MAGE [@DEATH GATE #4]  C/P,P,H

AUTHOR  TITLE  CODE
WEIS, MARGARET/HICKMAN, TRACY  SEVENTH GATE, THE [@DEATH GATE #7]  H:2,P
WEIS, MARGARET  SOULFORGE, THE [@DRAGONLANCE RAISTLIN 1]  P
WEIS, MARGARET/HICKMAN, TRACY  TEST OF THE TWINS [@DRAGONLANCE LEGENDS #3]  C/P:2,P
WEIS, MARGARET/BALDWIN, DAVID  TESTAMENT OF THE DRAGON  HA
WEIS, MARGARET/HICKMAN, TRACY  TIME OF THE TWINS [@DRAGONLANCE LEGENDS #1]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  TREASURES OF FANTASY  C/PA,LPA
WEIS, MARGARET/HICKMAN, TRACY  TRIUMPH OF THE DARKSWORD [@DARKSWORD #3]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  WAR OF THE TWINS [@DRAGONLANCE LEGENDS #2]  C/P:2,P
WEIS, MARGARET/HICKMAN, TRACY  WELL OF DARKNESS [SOVEREIGN STONE 1]  H
WEIS, MARGARET/HICKMAN, TRACY  WILL OF THE WANDERER, THE [@ROSE OF THE PROPHET #1]  C/P:2,P
WEISMAN, JOSEPH K.  INTO THE SHADOWS [SHADOWRUN 7]  C/PA,PA
WEISS, BOBBI JG/WEISS, DAVID CODY  BREAKAWAY [@STAR TREK TNG STARFLEET ACADEMY #12]  H,P
WEISS, DAVID CODY/WEISS, BOBBI JG  AGE OF AQUARIUMS [@SABRINA THE TEENAGE WITCH #20]  P
WEISS, DAVID CODY/WEISS, BOBBI JG  GO FETCH! [@SABRINA THE TEENAGE WITCH #13]  P
WEISS, DAVID CODY/WEISS, BOBBI JG  GOOD SWITCH, BAD SWITCH [@SABRINA THE TEENAGE WITCH #3]  P
WEISS, DAVID CODY/WEISS, BOBBI JG  PROM TIME [@SABRINA THE TEENAGE WITCH #21]  P
WEISS, DAVID CODY/WEISS, BOBBI JG  SABRINA THE TEENAGE WITCH [@SABRINA THE TEENAGE WITCH #1]  P
WEISS, DAVID CODY/WEISS, BOBBI JG  SALEM ON TRIAL [@SABRINA THE TEENAGE WITCH #8]  P
WEISSKOPF, T. K. F.  COSMIC TALES: ADVENTURES IN SOL SYSTEM  PA
WEIST, JERRY  BRADBURY: AN ILLUSTRATED LIFE  R/REF-H
WELCH, JANE  BROKEN CHALICE, THE [BOOK OF MAN 2]  P
WELCH, JANE  DAWN OF A DARK AGE [BOOK OF MAN 1]  P
WELCH, JANE  LOST RUNES, THE [RUNESPELL 2]  P
WELCH, JANE  RUNES OF SORCERY, THE [RUNESPELL 3]  P
WELCH, JANE  RUNES OF WAR, THE [RUNESPELL 1]  P
WELCH, MICHELLE M.  BRIGHT AND THE DARK, THE [FIVE COUNTRIES]  C/P,P
WELCH, MICHELLE M.  CONFIDENCE GAME [FIVE COUNTRIES]  P
WELCH, MICHELLE M.  CONFIDENCE GAME (PROOF) [FIVE COUNTRIES]  SR-P
WELDON, FAY  CLONING OF JOANNA MAY, THE  H
WELDON, PHAEDRA  WRAITH [@ZOE MARTINIQUE]  LP
WELLEN, EDWARD  HIJACK  C/P:2,P
WELLER, SAM  BRADBURY CHRONICLES, THE  R/REF-H
WELLER, TOM  SCIENCE MADE STUPID  VLP
WELLES, ORSON  INVASION FROM MARS: INTERPLANETARY STORIES  SR-PA,C/PA
WELLMAN, MANLY WADE  AFTER DARK [JOHN THE BALLADEER]  C/P,SFWA,P,H
WELLMAN, MANLY WADE  BEASTS FROM BEYOND, THE  SR-P
WELLMAN, MANLY WADE  BEYONDERS, THE  C/P,P
WELLMAN, MANLY WADE  CAHENA  H,SFWA
WELLMAN, MANLY WADE  DARK DESTROYERS, THE  DD443
WELLMAN, MANLY WADE  DEVIL IS NOT MOCKED, THE  H
WELLMAN, MANLY WADE  FEARFUL ROCK  H
WELLMAN, MANLY WADE  GIANTS FROM ETERNITY  H
WELLMAN, MANLY WADE  HANGING STONES, THE [JOHN THE BALLADEER]  C/P,SFWA,P,H
WELLMAN, MANLY WADE  INVADING ASTEROID, THE  B
WELLMAN, MANLY WADE  ISLAND IN THE SKY  H
WELLMAN, MANLY WADE  JOHN THE BALLADEER [JOHN THE BALLADEER]  P,SFWA-P
WELLMAN, MANLY WADE  LONELY VIGILS  SR-H
WELLMAN, MANLY WADE  LOST AND THE LURKING, THE [JOHN THE BALLADEER]  H:2,SFWA,P
WELLMAN, MANLY WADE  OLD GODS WAKEN, THE [JOHN THE BALLADEER]  C/P,SFWA,P,H
WELLMAN, MANLY WADE  OWLS HOOT IN THE DAYTIME  H
WELLMAN, MANLY WADE  SCHOOL OF DARKNESS, THE [JOHN THUNSTONE]  H,SFWA
WELLMAN, MANLY WADE/WELLMAN, WADE SHERLOCK HOLMES’S WAR OF THE WORLDS  C/P,P
WELLMAN, MANLY WADE  SIN’S DOORWAY  H
WELLMAN, MANLY WADE  SOJARR OF TITAN  SR-P,C/P
WELLMAN, MANLY WADE  SOLAR INVASION, THE  SR-P,C/P:2
WELLMAN, MANLY WADE  STRANGERS ON THE HEIGHTS  H
WELLMAN, MANLY WADE  TWICE IN TIME  SR-P,C/P:2,SFWA-P
WELLMAN, MANLY WADE  VALLEY SO LOW, THE  SFWA
WELLMAN, MANLY WADE  VOICE OF THE MOUNTAIN, THE [JOHN THE BALLADEER]  H,SFWA
WELLMAN, MANLY WADE  WHAT DREAMS MAY COME [JOHN THUNSTONE]  H,SFWA
WELLMAN, MANLY WADE  WHO FEARS THE DEVIL? [JOHN THE BALLADEER]  SR-P,C/P
WELLMAN, MANLY WADE  WORSE THINGS WAITING  SR-H
WELLS, ANGUS  DARK MAGIC [@GODWARS #2]  C/P:2,P
WELLS, ANGUS  EXILE’S CHALLENGE [EXILES 2]  LP
WELLS, ANGUS  EXILE’S CHILDREN [EXILES 1]  LP
WELLS, ANGUS  FORBIDDEN MAGIC [@GODWARS #1]  P
WELLS, ANGUS  GUARDIAN, THE  P
WELLS, ANGUS  LORDS OF THE SKY [LORDS OF THE SKY 1]  P
WELLS, ANGUS  USURPER, THE [@KINGDOMS #2]  C/P:2,P
WELLS, ANGUS  WAY BENEATH, THE [@KINGDOMS #3]  C/P,P
WELLS, ANGUS  WILD MAGIC [@GODWARS #3]  C/P:2,P
WELLS, ANGUS  WRATH OF ASHAR [@KINGDOMS #1]  C/P:2,P
WELLS, ANGUS  YESTERDAY’S KINGS  P
WELLS, BARRY  DAY THE EARTH CAUGHT FIRE, THE  C/P:2,P

The MITSFS Pinkdex  September 24, 2007  Page 409
AUTHOR  TITLE  CODE
WELLS, BASIL  DOORWAYS TO SPACE  SR-LP
WELLS, BASIL  PLANETS OF ADVENTURE  SR-LP
WELLS, CATHERINE  BEYOND THE GATES  C/P,P
WELLS, CATHERINE  EARTH IS ALL THAT LASTS, THE [CHILDREN OF EARTH 1]  P
WELLS, CATHERINE  EARTH SAVER, THE [CHILDREN OF EARTH 3]  P
WELLS, CATHERINE  MOTHER GRIMM  P
WELLS, H. G.  AUTOCRACY OF MR. PARHAM, THE  H
WELLS, H. G.  BEST SCIENCE FICTION STORIES OF H. G. WELLS  H,LP
WELLS, H. G.  BEST STORIES OF H. G. WELLS  C/P:2,P
WELLS, H. G.  BULPINGTON OF BLUP, THE  H
WELLS, H. G.  COLLECTOR’S BOOK OF SCIENCE FICTION BY H.G. WELLS, THE  H:2
WELLS, H. G.  COMPLETE SCIENCE FICTION TREASURY OF H. G. WELLS, THE  H:2
WELLS, H. G.  COUNTRY OF THE BLIND AND OTHER STORIES, THE  P
WELLS, H. G.  CROQUET PLAYER, THE  H:2
WELLS, H. G.  EMPIRE OF THE ANTS (AND OTHER STORIES), THE  P
WELLS, H. G.  EXPERIMENTS IN AUTOBIOGRAPHY  [C/RAN-H]
WELLS, H. G.  FIRST MEN IN THE MOON, THE  C/P,SR-P,H
WELLS, H. G.  FOOD OF THE GODS, THE  C/P,P,H
WELLS, H. G.  H. G. WELLS SCRAPBOOK, THE  R/REF-H,C/REF-H:2
WELLS, H. G.  HISTORY OF MR. POLLY, THE  H
WELLS, H. G.  HOLY TERROR, THE  SR-H
WELLS, H. G.  IN THE DAYS OF THE COMET  C/P,P,H
WELLS, H. G.  INEXPERIENCED GHOST AND NINE OTHER STORIES, THE  C/P,P
WELLS, H. G.  INVISIBLE MAN, THE  SR-P,C/P:2
WELLS, H. G.  INVISIBLE MAN, THE/WAR OF THE WORLDS, THE  C/P:2,P
WELLS, H. G.  ISLAND OF DOCTOR MOREAU, THE  C/P:2,P
WELLS, H. G.  MEN LIKE GODS  H,P
WELLS, H. G.  MODERN UTOPIA, A  LP
WELLS, H. G.  MR. BRITLING SEES IT THROUGH  H
WELLS, H. G.  OUTLINE OF HISTORY, THE: VOLUME 1 [HISTORY 1]  [C/RAN-H]
WELLS, H. G.  OUTLINE OF HISTORY, THE: VOLUME 2 [HISTORY 2]  [C/RAN-H]
WELLS, H. G.  PRIMEROS HOMBRES EN LA LUNA, LOS  FP
WELLS, H. G.  RESEARCH MAGNIFICENT, THE  H:2
WELLS, H. G./HUXLEY, JULIAN/WELLS, G. P. SCIENCE OF LIFE, THE  [C/RAN-H]
WELLS, H. G.  SEVEN FAMOUS NOVELS  H
WELLS, H. G.  SEVEN SCIENCE FICTION NOVELS OF H.G. WELLS  H:2
WELLS, H. G.  SHORT STORIES OF H. G. WELLS, THE  H:2
WELLS, H. G.  STAR-BEGOTTEN  C/P,P,H
WELLS, H. G.  TALES OF WONDER  H
WELLS, H. G.  THINGS TO COME (FILM SCRIPT WITH NOTES)  H
WELLS, H. G.  THIRTY STRANGE STORIES  H:2
WELLS, H. G.  THREE NOVELS OF THE FUTURE  H
WELLS, H. G.  THREE PROPHETIC NOVELS  H:2,LP
WELLS, H. G.  THREE PROPHETIC SCIENCE FICTION NOVELS OF H.G. WELLS  H,LP
WELLS, H. G.  TIME MACHINE, THE  C/P:2,P
WELLS, H. G.  TIME MACHINE, THE/WAR OF THE WORLDS, THE  P
WELLS, H. G.  TIME MACHINE AND OTHER STORIES, THE  P
WELLS, H. G.  TONO-BUNGAY  H
WELLS, H. G.  TRAVELS OF A REPUBLICAN RADICAL IN SEARCH OF HOT WATER  [C/RAN-P]
WELLS, H. G.  TWENTY-EIGHT SCIENCE FICTION STORIES OF H. G. WELLS  H:2
WELLS, H. G.  UNDYING FIRE, THE  H
WELLS, H. G.  VALLEY OF SPIDERS, THE  P
WELLS, H. G.  WAR IN THE AIR, THE  SR-P
WELLS, H. G.  WAR OF THE WORLDS, THE  C/P,SFWA-P,P,H,LP
WELLS, H. G.  WAR OF THE WORLDS, THE/TIME MACHINE, THE  C/P:2,P
WELLS, H. G.  WAR OF THE WORLDS, THE (CASSETTE)  SR-MM

WELLS, H. G.

WAR OF THE WORLDS, THE TIME MACHINE, AND SELECTED SHORT STORIES, THE

H:2

WELLS, H. G.  WHEN THE SLEEPER WAKES  C/P:2,P
WELLS, H. G.  WONDERFUL VISIT, THE  H
WELLS, H. G.  WORKS OF H. G. WELLS  H
WELLS, H. G.  WORLD SET FREE, THE  H,LP
WELLS, MARTHA  CITY OF BONES  H,P
WELLS, MARTHA  DEATH OF THE NECROMANCER, THE  P
WELLS, MARTHA  ELEMENT OF FIRE, THE  H
WELLS, MARTHA  ENTANGLEMENT [@STARGATE ATLANTIS #6]  P
WELLS, MARTHA  GATE OF GODS, THE [FALL OF ILE-RIEN 3]  H
WELLS, MARTHA  RELIQUARY [STARGATE ATLANTIS #2]  P
WELLS, MARTHA  SHIPS OF AIR, THE [FALL OF ILE-RIEN 2]  H
WELLS, MARTHA  WHEEL OF THE INFINITE  H,P
WELLS, MARTHA  WIZARD HUNTERS, THE [FALL OF ILE-RIEN 1]  H:2
WELLS, ROBERT  CANDLE IN THE SUN  C/P:2,P
WELLS, ROBERT  PARASAURIANS, THE  P
WELLS, ROBERT  RIGHT-HANDED WILDERNESS  C/P:2,P
WELLS, ROBERT  SPACEJACKS, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 410
AUTHOR  TITLE  CODE
WELLS, STUART W. III  SCIENCE FICTION AND HEROIC FANTASY AUTHOR INDEX, THE  R/IND
WELLS, WILLIAM K.  EFFIGIES  H
WENTWORTH, K. D.  BLACK ON BLACK  P
WENTWORTH, K. D.  HOUSE OF MOONS [HOUSE OF MOONS 2]  P
WENTWORTH, K. D.  IMPERIUM GAME, THE  P
WENTWORTH, K. D.  MOONSPEAKER [HOUSE OF MOONS 1]  P
WENTWORTH, K. D.  STARS OVER STARS  P
WERBER, BERNARD  EMPIRE OF THE ANTS  P
WERFEL, FRANZ  STAR OF THE UNBORN  C/P:2,P
WERNER, C. L.  BLOOD & STEEL [@WARHAMMER BRUNNER 2]  P
WERNER, C. L.  BLOOD MONEY [@ARHAMMER BRUNNER 1]  P
WERNER, C. L.  BLOOD OF THE DRAGON [@WARHAMMER BRUNNER 3]  P
WERNER, C. L.  WITCH FINDER [@WARHAMMER WITCH HUNTER 1]  P
WERNER, C. L.  WITCH HUNTER [@WARHAMMER]  P
WERNER, C. L.  WITCH KILLER [WARHAMMER]  P
WERPER, BARTON  TARZAN AND THE SILVER GLOBE [TARZAN]  SR-P
WERPER, BARTON  TARZAN AND THE SNAKE PEOPLE [TARZAN]  SR-P
WERTHAM, FREDRIC  WORLD OF FANZINES: A SPECIAL FORM OF COMMUNICATION, THE  R/REF-H
WESLEY, KATHRYN  10TH KINGDOM, THE  C/P,P
WEST, ANTHONY  H. G. WELLS  R/REF-H
WEST, CARL/MACLEAN, KATHERINE  DARK WING  H
WEST, CHARLES  TENANT, THE  H
WEST END BOOKS  DRAGONS OVER ENGLAND [TORG POSSIBILITY WARS]  PA
WEST END BOOKS  STRANGE TALES FROM THE NILE EMPIRE [TORG POSSIBILITY WARS]  PA
WEST, JESSAMYN  CHILEKINGS, THE  C/P:2,P
WEST, LINDSAY  EMPIRE OF THE ANTS  C/P,P
WEST, MICHELLE  BROKEN CROWN, THE [@SUN SWORD #1]  P
WEST, MICHELLE  CHAINS OF DARKNESS, CHAINS OF LIGHT [SUNDERED 4]  LP
WEST, MICHELLE  HUNTER’S DEATH [AVERALAAN 2]  C/P,P
WEST, MICHELLE  HUNTER’S OATH [AVERALAAN 1]  C/P,P
WEST, MICHELLE  LADY OF MERCY [SUNDERED 3]  LP
WEST, MICHELLE  RIVEN SHIELD, THE [@SUN SWORD #5]  P
WEST, MICHELLE  SEA OF SORROWS [@SUN SWORD #4]  P
WEST, MICHELLE  SHINING COURT, THE [@SUN SWORD #3]  C/P,P
WEST, MICHELLE  SUN SWORD, THE [@SUN SWORD #6]  C/P,P
WEST, MICHELLE  UNCROWNED KING, THE [@SUN SWORD #2]  P
WEST, MORRIS L.  SHOES OF THE FISHERMAN, THE  P
WEST, PAMELA  20/20 VISION  P
WEST, RICHARD C.  TOLKIEN CRITICISM: AN ANNOTATED CHECKLIST  R/IND
WEST, WALLACE  BIRD OF TIME, THE  C/P,P,H
WEST, WALLACE  LORDS OF ATLANTIS  C/P:2,P
WEST, WALLACE  MEMORY BANK, THE  C/P,P,H
WEST, WALLACE  OUTPOSTS IN SPACE  H
WEST, WALLACE  RIVER OF TIME  H
WESTALL, ROBERT  CATS OF SEROSTER, THE  H
WESTALL, ROBERT  DEVIL ON THE ROAD, THE  P
WESTCOTT, C. T.  BLOOD AND BONE [EAGLEHEART 3]  C/P:2,P
WESTCOTT, C. T.  SILVER WINGS AND LEATHER JACKETS [EAGLEHEART 1]  P
WESTERFELD, SCOTT  BLUE NOON [@MIDNIGHTERS #3]  H
WESTERFELD, SCOTT  BLUE NOON (PROOF) [@MIDNIGHTERS #3]  SR-LP
WESTERFELD, SCOTT  EVOLUTION’S DARLING  LP
WESTERFELD, SCOTT  FINE PREY  P
WESTERFELD, SCOTT  KILLING OF WORLDS, THE [SUCCESSION 2]  P
WESTERFELD, SCOTT  PEEPS  LP
WESTERFELD, SCOTT  POLYMORPH  P
WESTERFELD, SCOTT  RISEN EMPIRE, THE [SUCCESSION 1]  P
WESTERFELD, SCOTT  SUCCESSION [SUCCESSION 1,2]  H
WESTERN BOOKS  GOLDEN COMICS DIGEST #5  [C/RAN-P]
WESTLAKE, DONALD E.  HOT ROCK, THE  P
WESTLAKE, DONALD E.  HUMANS  P
WESTLAKE, DONALD E.  LIKELY STORY, A  SFWA-P
WESTLAKE, DONALD E.  SMOKE  P
WESTLAKE, DONALD E.  TOMORROW’S CRIMES  H

WESTMORE, MICHAEL/SIMS, ALAN/ LOOK, BRADLEY M./BIRNES, WILLIAM J.

STAR TREK ALIENS & ARTIFACTS [STAR TREK]

R/REF-P

WESTON, PETER  ANDROMEDA 1  SR-PA,HA
WESTON, PETER  ANDROMEDA 2  SR-PA
WESTON, PETER  ANDROMEDA 3  SR-PA
WESTON, PETER  WITH STARS IN MY EYES  R/REF-H
WESTON, SUSAN B.  CHILDREN OF THE LIGHT  H,SFWA,P
WESTWOOD, ALVIN  SINISTER FORCES  SR-P
WETANSON, BURT/HOOBLER, THOMAS  HUNTERS, THE  H,SFWA,P
WETANSON, BURT/HOOBLER, THOMAS  TREASURE HUNTERS, THE  P
WETHERELL, JUNE  BLUEPRINT FOR YESTERDAY  H,SFWA

The MITSFS Pinkdex  September 24, 2007  Page 411
AUTHOR  TITLE  CODE
WEVERKA, ROBERT  SEARCH  P
WEYRICH, BECKY LEE  THROUGH CAVERNS INFINITE  SR-P
WHALEY, STEPHAN V./COOK, STANLEY J.  MAN UNWEPT  LP
WHARTON, KEN  DIVINE INTERVENTION  P
WHEATLEY, DENNIS  DEVIL RIDES OUT, THE  C/P,P
WHEATLEY, DENNIS  GATEWAY TO HELL  P
WHEATLEY, DENNIS  HAUNTING OF TOBY JUGG, THE  C/P,P
WHEATLEY, DENNIS  KA OF GIFFORD HILLARY, THE  C/P,P
WHEATLEY, DENNIS  SATANIST, THE  P
WHEATLEY, DENNIS  STAR OF ILL-OMEN  SR-P
WHEATLEY, DENNIS  STRANGE CONFLICT  SR-P,C/P:2
WHEATLEY, DENNIS  THEY FOUND ATLANTIS  SR-P,C/P,H
WHEATLEY, DENNIS  THEY USED DARK FORCES  SR-P,C/P
WHEATLEY, DENNIS  TO THE DEVIL A DAUGHTER  C/P,P
WHEDON, JOSS  SERENITY [SERENITY]  R/REF-H

WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA SERENITY 1 [SERENITY 1]

CX

WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA SERENITY 2 [SERENITY 2]

CX

WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA SERENITY 3 [SERENITY 3]

CX

WHEDON, JOSS/MATTHEWS, BRETT/ CONRAD, WILL/MARTIN, LAURA SERENITY: THOSE LEFT BEHIND [SERENITY 1,2,3]

H

WHEELER, DEBORAH  JAYDIUM  C/P,P
WHEELER, DEBORAH  NORTHLIGHT  P
WHEELER, J. CRAIG  KRONE EXPERIMENT, THE  P
WHEELER, RAMONA LOUISE  HAVE STARSHIP, WILL TRAVEL  LP
WHEELER, SCOTT  MATTERS OF FORM  P
WHELAN, MICHAEL  ART OF MICHAEL WHELAN, THE  A

WHELAN, MICHAEL DRAGONS & MYSTICS: THE ART OF MICHAEL WHELAN 2005 CALENDAR

(2005)

CAL

WHIP, LEATHER  CAT O’ FIVE TAILS  WRD
WHIP, LEATHER  VICE’S PLEASURE TOY, THE  WRD
WHITAKER, DAVID  DOCTOR WHO AND THE CRUSADERS [DOCTOR WHO]  SR-P,C/P:2
WHITAKER, DAVID  DOCTOR WHO AND THE DALEKS [DOCTOR WHO]  SR-P,P

WHITAKER, DAVID DOCTOR WHO IN AN EXCITING ADVENTURE WITH THE DALEKS

[DOCTOR WHO]

C/P:2,P

WHITBOURN, JOHN  DOWNS-LORD DAWN [DOWNS-LORD 1]  P
WHITBOURN, JOHN  DOWNS-LORD DAY [DOWNS-LORD 2]  P
WHITBOURN, JOHN  POPES AND PHANTOMS  H
WHITE, ANDREA  SURVIVING ANTARCTICA: REALITY TV 2083  H
WHITE, DALE  IS SOMETHING UP THERE?  [C/RAN-P]
WHITE, ELLEN EMERSON  PRESIDENT’S DAUGHTER, THE  P
WHITE, JAMES  ALIEN EMERGENCIES [SECTOR GENERAL 4,5,6]  LP
WHITE, JAMES  ALIENS AMONG US, THE  C/P:2,P
WHITE, JAMES  ALL JUDGMENT FLED  H:2,P
WHITE, JAMES  AMBULANCE SHIP [SECTOR GENERAL 4]  C/P:2,P
WHITE, JAMES  CODE BLUE — EMERGENCY [SECTOR GENERAL 7]  C/P:2,P
WHITE, JAMES  DEADLY LITTER  C/P:2,P
WHITE, JAMES  DOUBLE CONTACT [SECTOR GENERAL 12]  C/P,P
WHITE, JAMES  DREAM MILLENNIUM, THE  C/P:2,P
WHITE, JAMES  ESCAPE ORBIT, THE  C/P:2,P
WHITE, JAMES  FEDERATION WORLD  C/P:2,P
WHITE, JAMES  FINAL DIAGNOSIS [SECTOR GENERAL 10]  H
WHITE, JAMES  FIRST PROTECTOR, THE [EARTH FINAL CONFLICT]  H,P
WHITE, JAMES  FUTURES PAST  C/P:2,P
WHITE, JAMES  GALACTIC GOURMET, THE [SECTOR GENERAL 9]  H,P
WHITE, JAMES  GENOCIDAL HEALER, THE [SECTOR GENERAL 8]  H:2,P
WHITE, JAMES  HOSPITAL STATION [SECTOR GENERAL 1]  C/P:2,P
WHITE, JAMES  LIFEBOAT  C/P:2,P
WHITE, JAMES  MAJOR OPERATION [SECTOR GENERAL 3]  C/P:2,P
WHITE, JAMES  MIND CHANGER [SECTOR GENERAL 11]  H
WHITE, JAMES  MONSTERS AND MEDICS  C/P:2,P
WHITE, JAMES  SECOND ENDING  DF173
WHITE, JAMES  SECRET VISITORS, THE  DD237,P,C/P:2
WHITE, JAMES  SECTOR GENERAL [SECTOR GENERAL 5]  C/P:2,P
WHITE, JAMES  SILENT STARS GO BY, THE  C/P:2,P
WHITE, JAMES  STAR HEALER [SECTOR GENERAL 6]  C/P:2,P
WHITE, JAMES  STAR SURGEON [SECTOR GENERAL 2]  C/P:2,P
WHITE, JAMES  TOMORROW IS TOO FAR  C/P:2,P
WHITE, JAMES  WATCH BELOW, THE  C/P:2,P
WHITE, JAMES  WHITE PAPERS, THE  SR-H
WHITE, JANE  COMET  H,SFWA
WHITE, JOHN  GAAL THE CONQUEROR  LP
WHITE, JON MANCHIP  DEATH BY DREAMING  P

The MITSFS Pinkdex  September 24, 2007  Page 412
AUTHOR  TITLE  CODE
WHITE, MARY ALICE  LAND OF THE POSSIBLE, THE  P
WHITE, MICHAEL  SUPER SCIENCE  [C/RAN-P:2]
WHITE, RICHARD C.  PATHS OF EVIL [@GAUNTLET DARK LEGACY 1]  P
WHITE, STEVE  BLOOD OF THE HEROES  H
WHITE, STEVE  DEBT OF AGES  P
WHITE, STEVE  DEMON’S GATE  P
WHITE, STEVE  DEMON’S GATE (PROOF)  SR-LP
WHITE, STEVE  DISINHERITED, THE [DISINHERITED 1]  P,SFWA-P
WHITE, STEVE  EAGLE AGAINST THE STARS  C/P,P
WHITE, STEVE  EMPEROR OF DAWN [BASIL CASTELLAN 2]  P
WHITE, STEVE/MEIER, SHIRLEY  EXODUS [STARFIRE 5]  H
WHITE, STEVE  FORGE OF THE TITANS  H
WHITE, STEVE  LEGACY [DISINHERITED 2]  P
WHITE, STEVE  PRINCE OF SUNSET [BASIL CASTELLAN 1]  C/P,P
WHITE, STEVE  PROMETHEUS PROJECT, THE  H
WHITE, STEVE  PROMETHEUS PROJECT, THE (PROOF)  SR-LP
WHITE, STEWART EDWARD  SIGN AT SIX, THE  H
WHITE, T. H.  BOOK OF MERLYN, THE  C/P,P,H
WHITE, T. H.  ELEPHANT AND THE KANGAROO, THE  H
WHITE, T. H.  ILL-MADE KNIGHT, THE  H
WHITE, T. H.  MAHARAJAH AND OTHER STORIES, THE  C/P:2,P
WHITE, T. H.  MASTER, THE  C/P,P
WHITE, T. H.  MISTRESS MASHAM’S REPOSE  C/P,P,H
WHITE, T. H.  ONCE AND FUTURE KING, THE  C/P,P,H
WHITE, TED  ANDROID AVENGER  DM123
WHITE, TED  BEST FROM AMAZING STORIES, THE  PA,C/PA
WHITE, TED  BEST FROM FANTASTIC, THE  C/PA,PA
WHITE, TED  BY FURIES POSSESSED  C/P:2,P
WHITE, TED/BISCHOFF, DAVID F.  FORBIDDEN WORLD  C/P:2,P
WHITE, TED  GREAT GOLD STEAL, THE  C/P,P
WHITE, TED  JEWELS OF ELSEWHEN, THE  C/P,P
WHITE, TED  NO TIME LIKE TOMORROW  H
WHITE, TED  PHOENIX PRIME  C/P:2,P
WHITE, TED  SECRET OF THE MARAUDER SATELLITE  H
WHITE, TED/VAN ARNAM, DAVE  SIDESLIP  C/P:2,P
WHITE, TED  SORCERESS OF QAR, THE  C/P:2,P
WHITE, TED  SPAWN OF THE DEATH MACHINE, THE  C/P:2,P
WHITE, TED  STAR WOLF!  C/P:2,P
WHITE, TONY  CHARLIEUNCLENORFOLKTANGO  LP
WHITEFORD, WYNNE  BREATHING SPACE ONLY  SR-P,C/P
WHITEFORD, WYNNE  HYADES CONTACT, THE  C/P,P
WHITEFORD, WYNNE  LAKE OF THE SUN  C/P,P
WHITEFORD, WYNNE  SAPPHIRE ROAD  SR-P,C/P
WHITEFORD, WYNNE  SPECIALIST, THE  C/P,P
WHITEFORD, WYNNE  THOR’S HAMMER  SR-P,C/P
WHITEHALL LINE  SCIENCE FICTION AUTHORS (PLAYING CARDS)  WRD
WHITEHEAD, HENRY S.  JUMBEE AND OTHER UNCANNY TALES  H

WHITELAW, SONNY/ CHRISTENSEN, ELIZABETH CHOSEN, THE [STARGATE ATLANTIS #3]

P

WHITELAW, SONNY  CITY OF THE GODS [@STARGATE SG-1 #4]  P

WHITELAW, SONNY/ CHRISTENSEN, ELIZABETH EXOGENESIS [@STARGATE ATLANTIS #5]

P

WHITELAW, SONNY/FALLON, JENNIFER  ROSWELL [@STARGATE SG-1 #9]  P
WHITELAW, STELLA  CAT THAT WASN’T THERE, THE  P

WHITELAW, STELLA/GARDINER, JUDY/ RONSON, MARK GRIMALKIN’S TALES

H

WHITESON, LEON  SCANNERS  P
WHITFIELD, KIT  BENIGHTED  LP

WHITFIELD, STEPHEN E./ RODDENBERRY, GENE MAKING OF STAR TREK, THE [STAR TREK]

R/REF-P,C/REF-P:2

WHITMAN, J. T.  STR8 BOLT  LP
WHITMAN, JOHN  CITY OF THE DEAD [@STAR WARS GALAXY OF FEAR #2]  P
WHITMAN, JOHN  EATEN ALIVE [@STAR WARS GALAXY OF FEAR #1]  P
WHITMAN, JOHN  GHOST OF THE JEDI [@STAR WARS GALAXY OF FEAR #5]  P
WHITMAN, JOHN  NIGHTMARE MACHINE, THE [@STAR WARS GALAXY OF FEAR #4]  P
WHITMAN, JOHN  PLANET PLAGUE [@STAR WARS GALAXY OF FEAR #3]  P
WHITMORE, ANDREW  FORTRESS OF ETERNITY, THE  C/P,P
WHITMORE, CHARLES  WINTER’S DAUGHTER  P
WHITNEY-ROBINSON, VORONICA  CRIMSON GOLD, THE [@FORGOTTEN REALMS ROGUES]  P

WHITNEY-ROBINSON, VORONICA/ BLACKMAN, HADEN RUINS OF DANTOOINE [@STAR WARS GALAXIES]

P

WHITNEY-ROBINSON, VORONICA  SANDS OF THE SOUL [@FORGOTTEN REALMS SEMBIA]  P
WHITTEMORE, EDWARD  SINAI TAPESTRY  P
WHITTEN, LESLIE H.  ALCHEMIST, THE  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 413
AUTHOR  TITLE  CODE
WHITTEN, LESLIE H.  MOON OF THE WOLF  P
WHITTINGTON, HARRY  DOOMSDAY AFFAIR, THE [@MAN FROM UNCLE #2]  C/P:2,P
WHYTE, ANDREW ADAMS  SF: THE NEW WRITERS 1973 & 1974  R/IND
WHYTE, H. WALTER  DEEP FREEZE  C/P,P
WHYTE, JACK  EAGLE, THE [CAMULOD 9]  H:2
WHYTE, JACK  EAGLES’ BROOD, THE [CAMULOD 3]  P
WHYTE, JACK  FORT AT RIVER’S BEND, THE [CAMULOD 5]  LP
WHYTE, JACK  LANCE THROWER, THE [CAMULOD]  P
WHYTE, JACK  SAXON SHORE, THE [CAMULOD 4]  P
WHYTE, JACK  SKYSTONE, THE [CAMULOD 1]  P
WHYTE, JACK  SORCERER: METAMORPHOSIS, THE [CAMULOD 6]  P
WHYTE, JACK  UTHER [CAMULOD]  P
WIATER, STANLEY  NIGHT VISIONS 7  SR-HA
WIBBERLEY, LEONARD  ENCOUNTER NEAR VENUS  H:2
WIBBERLEY, LEONARD  FEAST OF FREEDOM, A  P
WIBBERLEY, LEONARD  HOMEWARD TO ITHAKA  H
WIBBERLEY, LEONARD  MOUSE ON THE MOON, THE  C/P:2,P
WIBBERLEY, LEONARD  MOUSE ON WALL STREET, THE  P
WIBBERLEY, LEONARD  MOUSE THAT ROARED, THE  C/P,P,H
WIBBERLEY, LEONARD  MOUSE THAT SAVED THE WEST, THE  H
WIBBERLEY, LEONARD  QUEST OF EXCALIBUR, THE  H,LP
WIECK, STEWART  ANTHOLOGY [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  PA
WIECK, STEWART  BEAST WITHIN, THE [WORLD OF DARKNESS VAMPIRE]  PA

WIECK, STEWART/FLEMING, GHERBOD/ GRIFFIN, ERIC BLOODY SEPTEMBER

[@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #3]

LP

WIECK, STEWART/FLEMING, GHERBOD/ GRIFFIN, ERIC END GAMES

[@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #4]

LP

WIECK, STEWART/FLEMING, GHERBOD/ GRIFFIN, ERIC EYE OF GEHENNA, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #2]

LP

WIECK, STEWART/FLEMING, GHERBOD/ GRIFFIN, ERIC FALL OF ATLANTA, THE

[@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN SAGA #1]

LP

WIECK, STEWART  INHERIT THE EARTH [@HUNTER THE RECKONING]  PA
WIECK, STEWART/BRANSCOME, ANNA  QUINTESSENTIAL WORLD OF DARKNESS, THE [WORLD OF DARKNESS]  LPA
WIECK, STEWART  TOREADOR [@WORLD OF DARKNESS VAMPIRE MASQUERADE CLAN]  P
WIECK, STEWART  WHEN WILL YOU RAGE? [@WORLD OF DARKNESS WEREWOLF]  PA
WIENER, NORBERT  GOD AND GOLEM, INC.  [C/RAN-H]
WIGGIN, KATE DOUGLAS/SMITH, NORA A.  ARABIAN NIGHTS, THE  LP
WIGHTMAN, WAYNE  GANGLION AND OTHER STORIES  SR-H
WILBURN, STEPHEN PAUL  AEON  LP
WILDE, KELLEY  SUITING, THE  SFWA-P
WILDE, OSCAR  PICTURE OF DORIAN GRAY, THE  C/P,P,H
WILDER, CHERRY  DEALERS IN LIGHT AND DARKNESS  LP
WILDER, CHERRY  LUCK OF BRIN’S FIVE, THE [TORIN 1]  H,P
WILDER, CHERRY  NEAREST FIRE, THE [TORIN 2]  H,P
WILDER, CHERRY  PRINCESS OF THE CHAMELN, A [RULERS OF HYLOR 1]  H,P
WILDER, CHERRY  SECOND NATURE  P
WILDER, CHERRY  SIGNS OF LIFE  P
WILDER, CHERRY  SUMMER’S KING, THE [RULERS OF HYLOR 3]  H:2,P
WILDER, CHERRY  TAPESTRY WARRIORS, THE [TORIN 3]  H
WILDER, CHERRY  YORATH THE WOLF [RULERS OF HYLOR 2]  H,SFWA,P
WILDING, PHILIP  SHADOW OVER THE EARTH  H
WILDING, PHILIP  SPACEFLIGHT — VENUS  H
WILFORD, JOHN NOBLE  WE REACH THE MOON  [C/RAN-P]
WILHELM, KATE  ABYSS  C/P,SFWA,P,H
WILHELM, KATE  AND THE ANGELS SING  H
WILHELM, KATE  CHILDREN OF THE WIND  H,SFWA-P,LP
WILHELM, KATE  CITY OF CAIN  C/P:2,P
WILHELM, KATE  CLARION SF  C/PA,PA
WILHELM, KATE  CLEWISTON TEST, THE  C/P:2,P
WILHELM, KATE  CRAZY TIME  C/P,P,H
WILHELM, KATE  DARK DOOR, THE  P
WILHELM, KATE  DOWNSTAIRS ROOM, THE  C/P:2,P
WILHELM, KATE  FAULT LINES  P
WILHELM, KATE  HUYSMAN’S PETS  H,P

WILHELM, KATE INFINITY BOX, THE SFWA-TD12,C/P,SFWA,

TD12,P

WILHELM, KATE  JUNIPER TIME  H:2,SFWA,P
WILHELM, KATE  KILLER THING, THE  H:2,P
WILHELM, KATE  LET THE FIRE FALL  C/P,P,H
WILHELM, KATE  LISTEN, LISTEN  C/P,P
WILHELM, KATE  MARGARET AND I  C/P,P
WILHELM, KATE  MILE LONG SPACESHIP, THE  C/P:2,P
WILHELM, KATE  NEBULA AWARD STORIES 9  C/PA,PA
WILHELM, KATE  NEVERMORE AFFAIR, THE  C/P,P

AUTHOR  TITLE  CODE
WILHELM, KATE  OH SUSANNAH!  H
WILHELM, KATE  SENSE OF SHADOW, A  C/P,P,H
WILHELM, KATE  SOMERSET DREAMS AND OTHER FICTIONS  H,SFWA
WILHELM, KATE  STATE OF GRACE  SR-H
WILHELM, KATE  WELCOME, CHAOS  C/P,SFWA,P,H
WILHELM, KATE  WHERE LATE THE SWEET BIRDS SANG  SFWA-P,H:2,P,SFWA
WILKIN, SIR HUBERT/SHERMAN, HAROLD M. THOUGHTS THROUGH SPACE  H
WILKINS, CARY  TREASURY OF FANTASY, A  HA,LPA
WILKINS-FREEMAN, MARY  COLLECTED GHOST STORIES  SR-H
WILKS, EILEEN  BLOOD LINES [LILY YU 3]  P
WILKS, EILEEN  TEMPTING DANGER [LILY YU 1]  P
WILLANS, GEOFFREY/SEARLE, RONALD  MOLESWORTH’S GUIDE TO THE ATOMMIC AGE  [C/RAN-H]
WILLARD, ELIZA  POWER OF THREE, THE [@CHARMED]  P
WILLARD, NANCY  THINGS INVISIBLE TO SEE  C/P,P
WILLEFORD, CHARLES  MACHINE IN WARD ELEVEN, THE  C/P:2,P
WILLETT, JOHN  AUBADE FOR GAMELON  C/P,P
WILLETT, JOHN  SINGER IN THE STONE, THE  H
WILLEY, ELIZABETH  PRICE OF BLOOD AND HONOR, THE [LANDUC 3]  C/P:2,P
WILLEY, ELIZABETH  SORCERER AND A GENTLEMAN, A [LANDUC 2]  C/P:2,SFWA,P
WILLEY, ELIZABETH  SORCERER AND A GENTLEMAN, A (PROOF) [LANDUC 2]  SR-LP
WILLEY, ELIZABETH  WELL-FAVORED MAN, THE [LANDUC 1]  H,P
WILLIAMS, A. SUSAN  LIFTED VEIL, THE  LPA
WILLIAMS, BILLY DEE/MACGREGOR, ROB  PSI/NET  P
WILLIAMS, CHARLES  ALL HALLOW’S EVE  SR-P,C/P,H
WILLIAMS, CHARLES  DESCENT INTO HELL  H,LP
WILLIAMS, CHARLES  GREATER TRUMPS, THE  SR-P,C/P,H
WILLIAMS, CHARLES  MANY DIMENSIONS  H,LP
WILLIAMS, CHARLES  PLACE OF THE LION, THE  LP
WILLIAMS, CHARLES  SHADOWS OF ECSTASY  H,LP
WILLIAMS, CHARLES  WAR IN HEAVEN  H,LP
WILLIAMS, DAVID  SECOND SIGHT  P
WILLIAMS, GORDON  MICROCOLONY, THE [MICRONAUTS 2]  P
WILLIAMS, GORDON  MICRONAUTS, THE [MICRONAUTS 1]  P
WILLIAMS, GORDON  REVOLT OF THE MICRONAUTS [MICRONAUTS 3]  P
WILLIAMS, GRAHAM  NIGHTMARE FAIR, THE [DOCTOR WHO]  SR-P,P
WILLIAMS, GURNEY III  CALLING STATION E-A-R-T-H  LP
WILLIAMS, J. X.  ESP ORGY  SR-L
WILLIAMS, J. X.  WITCH IN HEAT  P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE ANTI-GRAVITY PAINT [DANNY DUNN]  C/P,P,H
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE AUTOMATIC HOUSE [DANNY DUNN]  C/P,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE FOSSIL CAVE [DANNY DUNN]  C/P,P,H
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE HEAT RAY [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE HOMEWORK MACHINE [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE SMALLIFYING MACHINE [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE SWAMP MONSTER [DANNY DUNN]  C/P,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE UNIVERSAL GLUE [DANNY DUNN]  C/P,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE VOICE FROM SPACE [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN AND THE WEATHER MACHINE [DANNY DUNN]  C/P,P,H
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN, INVISIBLE BOY [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN ON A DESERT ISLAND [DANNY DUNN]  C/P,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN ON THE OCEAN FLOOR [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN, SCIENTIFIC DETECTIVE [DANNY DUNN]  C/P:2,P
WILLIAMS, JAY/ABRASHKIN, RAYMOND  DANNY DUNN, TIME TRAVELER [DANNY DUNN]  C/P,P
WILLIAMS, JAY  HERO FROM OTHERWHERE, THE  C/P:2,P
WILLIAMS, JAY  PEOPLE OF THE AX, THE  C/P,P
WILLIAMS, LIZ  DEMON AND THE CITY, THE [DETECTIVE INSPECTOR CHEN 2]  LP
WILLIAMS, LIZ  EMPIRE OF BONES  P
WILLIAMS, LIZ  GHOST SISTER, THE  C/P,P
WILLIAMS, LIZ  NINE LAYERS OF SKY  P
WILLIAMS, LIZ  NINE LAYERS OF SKY (PROOF)  SR-P
WILLIAMS, LIZ  POISON MASTER, THE  C/P,P
WILLIAMS, LIZ  PRECIOUS DRAGON [DETECTIVE INSPECTOR CHEN]  H
WILLIAMS, LORRAINE DILLE  BUCK ROGERS: THE FIRST 60 YEARS IN THE 25TH CENTURY  CX
WILLIAMS, MARGERY  VELVETEEN RABBIT, THE  SFWA-P
WILLIAMS, MICHAEL  ALLAMANDA [ARCADY 2]  H,LP
WILLIAMS, MICHAEL  ARCADY [ARCADY 1]  C/P,P
WILLIAMS, MICHAEL  BALANCE OF POWER, THE  P
WILLIAMS, MICHAEL/WILLIAMS, TERI  BEFORE THE MASK [@DRAGONLANCE VILLIANS #1]  C/P,P
WILLIAMS, MICHAEL/WILLIAMS, TERI  DARK QUEEN, THE [@DRAGONLANCE VILLAINS #6]  P
WILLIAMS, MICHAEL  FOREST LORD, A  P
WILLIAMS, MICHAEL  GALEN BEKNIGHTED [@DRAGONLANCE HEROES II #3]  C/P,P
WILLIAMS, MICHAEL  OATH AND THE MEASURE, THE [@DRAGONLANCE MEETINGS 4]  P
WILLIAMS, MICHAEL  SORCERER’S APPRENTICE, A  P
WILLIAMS, MICHAEL  WEASEL’S LUCK [@DRAGONLANCE HEROES 3]  C/P:2,P

AUTHOR  TITLE  CODE
WILLIAMS, MICHAEL LINDSAY  FTL: FURTHER THAN LIFE  C/P,P
WILLIAMS, MICHAEL LINDSAY  MARTIAN SPRING  C/P,P
WILLIAMS, MONA  MESSENGER, THE  P
WILLIAMS, NICK BODDIE  ATOM CURTAIN, THE  DD139
WILLIAMS, PAUL  ONLY APPARENTLY REAL  SFWA,R/REF-P
WILLIAMS, PAUL O.  AMBUSH OF SHADOWS, AN [PELBAR 5]  C/P:2,P
WILLIAMS, PAUL O.  BREAKING OF NORTHWALL, THE [PELBAR 1]  C/P:2,P
WILLIAMS, PAUL O.  DOME IN THE FOREST, THE [PELBAR 3]  C/P:3,P
WILLIAMS, PAUL O.  ENDS OF THE CIRCLE, THE [PELBAR 2]  C/P:2,P
WILLIAMS, PAUL O.  FALL OF THE SHELL, THE [PELBAR 4]  C/P:2,P
WILLIAMS, PAUL O.  GIFTS OF THE GORBODUC VANDAL, THE  C/P:2,P
WILLIAMS, PAUL O.  SONG OF THE AXE, THE [PELBAR 6]  C/P:2,P
WILLIAMS, PAUL O.  SWORD OF FORBEARANCE, THE [PELBAR 7]  P
WILLIAMS, REGINALD  TIME TRAVELERS ACADEMY, THE  LP
WILLIAMS, ROBERT MOORE  BEACHHEAD PLANET  C/P:2,P
WILLIAMS, ROBERT MOORE  BELL FROM INFINITY, THE  C/P:2,P
WILLIAMS, ROBERT MOORE  BLUE ATOM, THE  DD322
WILLIAMS, ROBERT MOORE  CHAOS FIGHTERS, THE  C/P:2,P
WILLIAMS, ROBERT MOORE  CONQUEST OF THE SPACE SEA  DD099
WILLIAMS, ROBERT MOORE  DARKNESS BEFORE TOMORROW, THE  DF141
WILLIAMS, ROBERT MOORE  DAY THEY H-BOMBED LOS ANGELES, THE  C/P:2,P
WILLIAMS, ROBERT MOORE  DOOMSDAY EVE  DD215
WILLIAMS, ROBERT MOORE  FLIGHT FROM YESTERDAY  DF223
WILLIAMS, ROBERT MOORE  JONGOR FIGHTS BACK  P
WILLIAMS, ROBERT MOORE  JONGOR OF LOST LAND  C/P,P
WILLIAMS, ROBERT MOORE  KING OF THE FOURTH PLANET  DF149
WILLIAMS, ROBERT MOORE  LOVE IS FOREVER — WE ARE FOR TONIGHT  R/REF-P
WILLIAMS, ROBERT MOORE  LUNAR EYE, THE  DF261
WILLIAMS, ROBERT MOORE  NOW COMES TOMORROW  P
WILLIAMS, ROBERT MOORE  REFUGE FOR TONIGHT  B
WILLIAMS, ROBERT MOORE  RETURN OF JONGOR, THE  C/P,P
WILLIAMS, ROBERT MOORE  SECOND ATLANTIS, THE  C/P:2,P
WILLIAMS, ROBERT MOORE  SEVEN TICKETS TO HELL  C/P,P
WILLIAMS, ROBERT MOORE  STAR WASPS, THE  DF177
WILLIAMS, ROBERT MOORE  TO THE END OF TIME AND OTHER STORIES  DD427
WILLIAMS, ROBERT MOORE  VIGILANTE, 21ST CENTURY  C/P:2,P
WILLIAMS, ROBERT MOORE  VOID BEYOND AND OTHER STORIES, THE  DD322
WILLIAMS, ROBERT MOORE  WALK UP THE SKY  H
WILLIAMS, ROBERT MOORE  WHEN TWO WORLDS MEET  P
WILLIAMS, ROBERT MOORE  WORLD OF THE MASTERMINDS  DD427
WILLIAMS, ROBERT MOORE  ZANTHAR AT MOON’S MADNESS  C/P,P
WILLIAMS, ROBERT MOORE  ZANTHAR AT THE EDGE OF NEVER  P
WILLIAMS, ROBERT MOORE  ZANTHAR AT TRIP’S END  P
WILLIAMS, ROBERT MOORE  ZANTHAR OF THE MANY WORLDS  C/P,P
WILLIAMS, SEAN  BLOOD DEBT, THE [@CATACLYSM #2]  H
WILLIAMS, SEAN  CROOKED LETTER, THE [CATACLYSM 1]  H
WILLIAMS, SEAN  DOORWAY TO ETERNITY  SR-LP
WILLIAMS, SEAN/DIX, SHANE  ECHOES OF EARTH [ECHOES OF EARTH 1]  C/P,P
WILLIAMS, SEAN/DIX, SHANE  EVERGENCE [EVERGENCE 1,2,3]  H
WILLIAMS, SEAN/DIX, SHANE  EVERGENCE: A DARK IMBALANCE [EVERGENCE 3]  P
WILLIAMS, SEAN/DIX, SHANE  EVERGENCE: THE DYING LIGHT [EVERGENCE 2]  P
WILLIAMS, SEAN/DIX, SHANE  EVERGENCE: THE PRODIGAL SUN [EVERGENCE 1]  P
WILLIAMS, SEAN/DIX, SHANE  FORCE HERETIC [@STAR WARS NEW JEDI ORDER FORCE HERETIC 1,2,3]  H
WILLIAMS, SEAN/DIX, SHANE  GEODESICA ASCENT [GEODESICA 1]  P
WILLIAMS, SEAN/DIX, SHANE  GEODESICA DESCENT [GEODESICA 2]  P
WILLIAMS, SEAN  HANGING MOUNTAINS [@CATACLYSM #3]  H
WILLIAMS, SEAN/DIX, SHANE  HEIRS OF EARTH [ECHOES OF EARTH 3]  P
WILLIAMS, SEAN  METAL FATIGUE  H
WILLIAMS, SEAN/DIX, SHANE  ORPHANS OF EARTH [ECHOES OF EARTH 2]  P
WILLIAMS, SEAN/DIX, SHANE  REFUGEE [@STAR WARS NEW JEDI ORDER FORCE HERETIC #2]  P
WILLIAMS, SEAN/DIX, SHANE  REMNANT [@STAR WARS NEW JEDI ORDER FORCE HERETIC #1]  P
WILLIAMS, SEAN  RESURRECTED MAN, THE  H
WILLIAMS, SEAN/DIX, SHANE  REUNION [@STAR WARS NEW JEDI ORDER FORCE HERETIC #3]  P
WILLIAMS, SEAN  SATURN RETURNS  P
WILLIAMS, SHEILA  ASIMOV’S SCIENCE FICTION 30TH ANNIVERSARY ANTHOLOGY  LPA
WILLIAMS, SHEILA/MANSON, CYNTHIA  INTERGALACTIC MERCENARIES  PA
WILLIAMS, SHEILA  ROBOTS FROM ASIMOV’S  B
WILLIAMS, T. OWEN  MONTH FOR MANKIND, A  H
WILLIAMS, TAD  CALIBAN’S HOUR  C/P,P,H
WILLIAMS, TAD  CITY OF GOLDEN SHADOW [@OTHERLAND #1]  H,P
WILLIAMS, TAD  DRAGONBONE CHAIR, THE [MEMORY,SORROW AND THORN 1]  C/P,H:2,P
WILLIAMS, TAD  MIRROR WORLD  H
WILLIAMS, TAD  MOUNTAIN OF BLACK GLASS [@OTHERLAND #3]  C/P,P,H
WILLIAMS, TAD  RITE  H

The MITSFS Pinkdex  September 24, 2007  Page 416
AUTHOR  TITLE  CODE
WILLIAMS, TAD  RIVER OF BLUE FIRE [@OTHERLAND #2]  H,P
WILLIAMS, TAD  SEA OF SILVER LIGHT [@OTHERLAND #4]  H,P
WILLIAMS, TAD  SHADOWMARCH [SHADOWMARCH 1]  H

WILLIAMS, TAD SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #2 (EXCERPT, CHAPTERS 3 AND 4) [SHADOWMARCH 1]

SR-LP

WILLIAMS, TAD SHADOWMARCH, VOLUME 1, SPECIAL SNEAK PREVIEW #3 (EXCERPT, CHAPTERS 5 AND 6) [SHADOWMARCH 1]

SR-LP

WILLIAMS, TAD  SHADOWPLAY [SHADOWMARCH 2]  H
WILLIAMS, TAD  STONE OF FAREWELL [MEMORY,SORROW AND THORN 2]  H,P
WILLIAMS, TAD  TAILCHASER’S SONG  C/P,P,H
WILLIAMS, TAD  TO GREEN ANGEL TOWER [MEMORY,SORROW AND THORN 3]  H:2,SFWA:2
WILLIAMS, TAD  TO GREEN ANGEL TOWER (PART ONE) [MEMORY,SORROW AND THORN 3]  C/P,P
WILLIAMS, TAD  TO GREEN ANGEL TOWER (PART TWO) [MEMORY,SORROW AND THORN 3]  C/P:2,P
WILLIAMS, TAD  WAR OF THE FLOWERS, THE  H
WILLIAMS, THOMAS  BRIDE OF STONE, THE  LP
WILLIAMS, URSULA MORAY  MOONBALL, THE  P
WILLIAMS, WALTER JON  AMBASSADOR OF PROGRESS  C/P,SFWA-P,P
WILLIAMS, WALTER JON  ANGEL STATION  C/P,SFWA-P,P,H
WILLIAMS, WALTER JON  ANGEL STATION (ERRATA SHEET INCLUDED)  SR-H,H,SFWA
WILLIAMS, WALTER JON  ARISTOI  P
WILLIAMS, WALTER JON  CITY ON FIRE [AIAH 2]  H,P
WILLIAMS, WALTER JON  CONVENTIONS OF WAR [@DREAD EMPIRE’S FALL 3]  P
WILLIAMS, WALTER JON  CROWN JEWELS, THE [DRAKE MAIJSTRAL 1]  C/P:2,P
WILLIAMS, WALTER JON  DAYS OF ATONEMENT  C/P,P,H,SFWA
WILLIAMS, WALTER JON  DESTINY’S WAY [@STAR WARS NEW JEDI ORDER]  P
WILLIAMS, WALTER JON  ELEGY FOR ANGELS AND DOGS  TD24
WILLIAMS, WALTER JON  FACETS  C/P,SFWA-P,P,H
WILLIAMS, WALTER JON  FRANKENSTEINS AND FOREIGN DEVILS  SR-H
WILLIAMS, WALTER JON  HARDWIRED  C/P,P,H
WILLIAMS, WALTER JON  HOUSE OF SHARDS [DRAKE MAIJSTRAL 2]  P,SFWA-P
WILLIAMS, WALTER JON  KNIGHT MOVES  C/P,SFWA-P,P
WILLIAMS, WALTER JON  METROPOLITAN [AIAH 1]  H,P
WILLIAMS, WALTER JON  PRAXIS, THE [@DREAD EMPIRE’S FALL 1]  P
WILLIAMS, WALTER JON  PRAXIS, THE (PROOF) [@DREAD EMPIRE’S FALL 1]  SR-P
WILLIAMS, WALTER JON  RIFT, THE  P
WILLIAMS, WALTER JON  ROCK OF AGES [DRAKE MAIJSTRAL 3]  H,SFWA
WILLIAMS, WALTER JON  SOLIP:SYSTEM  SR-H
WILLIAMS, WALTER JON  SUNDERING, THE [@DREAD EMPIRE’S FALL 2]  P
WILLIAMS, WALTER JON  VOICE OF THE WHIRLWIND  C/P:2,SFWA-P,P,H
WILLIAMSON, AL  ON THE LOST CONTINENT OF MONGO [FLASH GORDON]  SCX
WILLIAMSON, CHET  ASH WEDNESDAY  SFWA-P
WILLIAMSON, CHET  CITY OF IRON [@SEARCHERS #1]  P
WILLIAMSON, CHET  CLASH BY NIGHT [CROW]  LP
WILLIAMSON, CHET  EMPIRE OF DUST [@SEARCHERS #2]  C/P,P
WILLIAMSON, CHET  HELL  P
WILLIAMSON, CHET  LOWLAND RIDER  SFWA-P
WILLIAMSON, CHET  MORDENHEIM [RAVENLOFT]  P
WILLIAMSON, CHET  MURDER IN CORMYR [@FORGOTTEN REALMS]  P
WILLIAMSON, CHET  SIEGE OF STONE [@SEARCHERS #3]  C/P:2,P
WILLIAMSON, J. N.  BANISHED, THE  P
WILLIAMSON, J. N.  DEATH-COACH  P
WILLIAMSON, J. N.  GHOST MANSION  P
WILLIAMSON, J. N.  HORROR HOUSE  P
WILLIAMSON, J. N.  HOUNGAN, THE  P
WILLIAMSON, J. N.  MASQUES  SR-HA,SFWA
WILLIAMSON, J. N.  MASQUES 2  SR-HA
WILLIAMSON, J. N.  MASQUES 3  SR-HA
WILLIAMSON, J. N.  MASQUES 4  SR-HA
WILLIAMSON, J. N.  NAKED FLESH OF FEELING, THE  SR-H
WILLIAMSON, J. N.  PREMONITION  P
WILLIAMSON, J. N.  TULPA, THE  P
WILLIAMSON, JACK  AFTER WORLD’S END/LEGION OF TIME, THE  SR-P,C/P:2
WILLIAMSON, JACK  BEACHHEAD  H
WILLIAMSON, JACK  BEST OF JACK WILLIAMSON, THE  C/P:2,P
WILLIAMSON, JACK  BLACK SUN, THE  H,P
WILLIAMSON, JACK  BRIGHT NEW UNIVERSE  C/P:2,P
WILLIAMSON, JACK  BROTHER TO DEMONS, BROTHER TO GODS  C/P,P,H
WILLIAMSON, JACK  COMETEERS, THE  SR-H,C/P:2,P
WILLIAMSON, JACK  CRUCIBLE OF POWER, THE [@COLLECTED STORIES JACK WILLIAMSON #5]  H
WILLIAMSON, JACK  DARKER THAN YOU THINK  SR-H,C/P:2,P
WILLIAMSON, JACK  DEMON MOON  H,P
WILLIAMSON, JACK  DOME AROUND AMERICA  DD118
WILLIAMSON, JACK  DRAGON’S ISLAND  C/P:2,P
WILLIAMSON, JACK  EARLY WILLIAMSON, THE  H,SFWA,P

The MITSFS Pinkdex  September 24, 2007  Page 417
AUTHOR  TITLE  CODE
WILLIAMSON, JACK  FIRECHILD  C/P,SFWA-P,P,H
WILLIAMSON, JACK/BREUER, DR. MILES J.  GIRL FROM MARS, THE  B,GRD17
WILLIAMSON, JACK  GOLDEN BLOOD  C/P:2,P
WILLIAMSON, JACK  GREEN GIRL, THE  SR-P
WILLIAMSON, JACK  H. G. WELLS: CRITIC OF PROGRESS  R/REF-H
WILLIAMSON, JACK  HUMANOID TOUCH, THE  C/P,P,H
WILLIAMSON, JACK  HUMANOIDS, THE  C/P,GN4.21,H,SR-P,P
WILLIAMSON, JACK  HUMANOIDS, THE (EXPANDED)  C/P,P,H
WILLIAMSON, JACK  INTO THE EIGHTH DECADE  SR-H
WILLIAMSON, JACK  IRON GOD, THE  GRD18

WILLIAMSON, JACK LEGION OF SPACE, THE C/P,SFWA-P,SR-H,SR-P,P,

H,GN1.02

WILLIAMSON, JACK  LEGION OF TIME, THE  C/P,SR-H,H
WILLIAMSON, JACK  LIFEBURST  C/P,P,H
WILLIAMSON, JACK  MANSEED  H:2,P
WILLIAMSON, JACK  MAZEWAY  H:2,P

WILLIAMSON, JACK METAL MAN AND OTHERS, THE

[@COLLECTED STORIES JACK WILLIAMSON #1]

H

WILLIAMSON, JACK  MOON CHILDREN, THE  C/P,P,H
WILLIAMSON, JACK  NOT-MEN, THE  P
WILLIAMSON, JACK  ONE AGAINST THE LEGION  C/P:2,P
WILLIAMSON, JACK  PANDORA EFFECT, THE  C/P,P
WILLIAMSON, JACK  PEOPLE MACHINES  C/P,P
WILLIAMSON, JACK  POWER OF BLACKNESS, THE  H,P
WILLIAMSON, JACK  PRINCE OF SPACE, THE  GRD17
WILLIAMSON, JACK  QUEEN OF THE LEGION, THE  C/P,P
WILLIAMSON, JACK  REIGN OF WIZARDRY, THE  C/P:2,P
WILLIAMSON, JACK  SEETEE SHIP  P
WILLIAMSON, JACK  SEETEE SHIP/SEETEE SHOCK  C/P:2,P
WILLIAMSON, JACK  SEETEE SHOCK  P
WILLIAMSON, JACK  SEVENTY-FIVE  H
WILLIAMSON, JACK  SILICON DAGGER, THE  H
WILLIAMSON, JACK  SPIDER ISLAND [@COLLECTED STORIES JACK WILLIAMSON #4]  H
WILLIAMSON, JACK/GUNN, JAMES E.  STAR BRIDGE  C/P,P,H
WILLIAMSON, JACK  STONEHENGE GATE, THE  H
WILLIAMSON, JACK  TERRAFORMING EARTH  H
WILLIAMSON, JACK  THREE FROM THE LEGION  H:2,P
WILLIAMSON, JACK  TRANSFORMING EARTH  P
WILLIAMSON, JACK  TRAPPED IN SPACE  C/P,LP
WILLIAMSON, JACK  TRIAL OF TERRA, THE  C/P:2,P
WILLIAMSON, JACK  WIZARD’S ISLE [@COLLECTED STORIES JACK WILLIAMSON #3]  H
WILLIAMSON, JACK  WOLVES OF DARKNESS [@COLLECTED STORIES JACK WILLIAMSON #2]  H
WILLIAMSON, JACK  WONDER’S CHILD: MY LIFE IN SCIENCE FICTION  SFWA
WILLIAMSON, JULIA  STARS THROUGH MAGIC CASEMENTS, THE  [C/RAN-H]
WILLIAMSON, MICHAEL Z.  FREEHOLD [FREEHOLD WAR 1]  P
WILLIAMSON, MICHAEL Z.  WEAPON, THE [FREEHOLD WAR 2]  H
WILLIAMSON, THAMES  BEGINNING AT DUSK: AN INTERLUDE  H
WILLIS, CONNIE  BELLWETHER  C/P,P
WILLIS, CONNIE  D.A.  H
WILLIS, CONNIE  DOOMSDAY BOOK [DUNWORTHY & FINCH 1]  C/P,LP,H:2
WILLIS, CONNIE  FIRE WATCH  C/P:2,SFWA:2,P
WILLIS, CONNIE  FUTURES IMPERFECT  H
WILLIS, CONNIE  IMPOSSIBLE THINGS  C/P,P,H
WILLIS, CONNIE  INSIDE JOB  H
WILLIS, CONNIE/FELICE, CYNTHIA  LIGHT RAID  C/P,SFWA-P:2,H,P,SFWA
WILLIS, CONNIE  LINCOLN’S DREAMS  H:2,P
WILLIS, CONNIE  MIRACLE AND OTHER CHRISTMAS STORIES  C/P,P
WILLIS, CONNIE  NEBULA AWARDS 33  LPA
WILLIS, CONNIE  NEW HUGO WINNERS 3, THE  PA
WILLIS, CONNIE  PASSAGE  H,P
WILLIS, CONNIE/FELICE, CYNTHIA  PROMISED LAND  C/P,P,H
WILLIS, CONNIE  REMAKE  H
WILLIS, CONNIE  TO SAY NOTHING OF THE DOG [DUNWORTHY & FINCH 2]  H:2,P
WILLIS, CONNIE  UNCHARTED TERRITORY  C/P:2,P

WILLIS, DAN DRAGON KNIGHT [@DRAGONLANCE NEW ADVENTURES #7/

DRAGONLANCE DRAGON QUARTET 3]

LP

WILLIS, DAN DRAGON WELL, THE [@DRAGONLANCE NEW ADVENTURES #3/

DRAGONLANCE SPELLBINDER QUARTET 3]

LP

WILLIS, DAN  WIZARD’S RETURN [@DRAGONLANCE NEW ADVENTURES TRINISTYR #3]  LP
WILLIS, JENNIFER SCHWAMM  WIZARDS  LPA
WILLIS, MAUD  DEVIL’S RAIN, THE  P
WILLIS, PAUL J.  NO CLOCK IN THE FOREST  P,SFWA-P
WILLIS, PAUL J.  STOLEN RIVER, THE  P
WILLY, COLETTE  SEPT DIALOGUES DE BETES  FP
WILSON, CLIFFORD  CRASH GO THE CHARIOTS  [C/RAN-P]

AUTHOR  TITLE  CODE
WILSON, COLIN  DELTA, THE [SPIDER WORLD 4]  P
WILSON, COLIN  DESERT, THE [SPIDER WORLD 1]  P
WILSON, COLIN  FORTRESS, THE [SPIDER WORLD 3]  P
WILSON, COLIN  GLASS CAGE, THE  P
WILSON, COLIN  MIND PARASITES, THE  H:2,LP
WILSON, COLIN  PHILOSOPHER’S STONE, THE  C/P,P,H
WILSON, COLIN  SPACE VAMPIRES, THE  C/P,P,H,SFWA
WILSON, COLIN  TOWER, THE [SPIDER WORLD 2]  P
WILSON, COLIN  TREE BY TOLKIEN  R/REF-P
WILSON, D. HARLAN  PSEUDO-CITY  LP
WILSON, DAVID HENRY  COACHMAN RAT, THE  C/P,SFWA-P,P,H
WILSON, DAVID NIALL  CHRYSALIS [@STAR TREK VOYAGER]  P
WILSON, DAVID NIALL  LASOMBRA [@WORLD OF DARKNESS DARK AGES CLAN #5]  P
WILSON, DAVID NIALL  RELIC OF THE DAWN [@EXALTED #2]  P

WILSON, DAVID NIALL TO DREAM OF DREAMERS LOST

[@WORLD OF DARKNESS GRAILS COVENANT #3]

P

WILSON, DAVID NIALL TO SIFT THROUGH BITTER ASHES

[@WORLD OF DARKNESS GRAILS COVENANT #1]

P

WILSON, DAVID NIALL TO SPEAK IN LIFELESS TONGUES

[@WORLD OF DARKNESS GRAILS COVENANT #2]

P

WILSON, F. PAUL  ALL THE RAGE [REPAIRMAN JACK]  H
WILSON, F. PAUL  BARRENS AND OTHERS, THE  LP
WILSON, F. PAUL  BLACK WIND  H,SFWA-P,P
WILSON, F. PAUL  CONSPIRACIES [REPAIRMAN JACK]  H
WILSON, F. PAUL  CRISSCROSS [REPAIRMAN JACK]  P
WILSON, F. PAUL  DYDEETOWN WORLD  C/P:2,SFWA-P,P
WILSON, F. PAUL  ENEMY OF THE STATE, AN [REPAIRMAN JACK]  C/P,SFWA,P,H
WILSON, F. PAUL  HAUNTED AIR, THE [REPAIRMAN JACK]  P
WILSON, F. PAUL  HEALER [REPAIRMAN JACK]  C/P,P,H,SFWA
WILSON, F. PAUL  HOSTS [REPAIRMAN JACK]  P
WILSON, F. PAUL  INFERNAL [REPAIRMAN JACK]  H
WILSON, F. PAUL  KEEP, THE [ADVERSARY 1]  H,P
WILSON, F. PAUL  LANAGUE CHRONICLES, THE  P,SFWA-P
WILSON, F. PAUL  LEGACIES [REPAIRMAN JACK]  P
WILSON, F. PAUL/COSTELLO, MATTHEW J.  MASQUE  P
WILSON, F. PAUL  MIDNIGHT MASS  H
WILSON, F. PAUL  NIGHTWORLD [ADVERSARY 6]  P
WILSON, F. PAUL  PORTERO METHOD, THE [@SIMS #2]  SR-H
WILSON, F. PAUL  REBORN [ADVERSARY 4]  P
WILSON, F. PAUL  REPRISAL [ADVERSARY 5]  P
WILSON, F. PAUL  SIMS  H
WILSON, F. PAUL  SOFT AND OTHERS  H,P
WILSON, F. PAUL  TERY, THE  BS2
WILSON, F. PAUL  TERY, THE (COMPLETE)  C/P,SFWA-P,P
WILSON, F. PAUL  TOMB, THE [ADVERSARY 2/REPAIRMAN JACK]  C/P:2,P
WILSON, F. PAUL  TOUCH, THE [ADVERSARY 3]  P
WILSON, F. PAUL  WHEELS WITHIN WHEELS [REPAIRMAN JACK]  C/P,SFWA,P,H
WILSON, GAHAN  "...AND THEN WE’LL GET HIM!"  CX
WILSON, GAHAN  CLEFT AND OTHER ODD TALES, THE  LP
WILSON, GAHAN  EVERYBODY’S FAVORITE DUCK  C/P,P
WILSON, GAHAN  FIRST WORLD FANTASY AWARDS  SFWA
WILSON, GAHAN  HARRY, THE FAT BEAR SPY  LP
WILSON, GAHAN  PLAYBOY’S GAHAN WILSON  CX
WILSON, JOHN  GUARDIANS OF ALEXANDER [GOLDBANE 1]  LP
WILSON, RICHARD  AND THEN THE TOWN TOOK OFF  DD437
WILSON, RICHARD  GIRLS FROM PLANET 5, THE  SR-H,C/P:2,P
WILSON, RICHARD  30 DAY WONDER  C/P:2,P
WILSON, RICHARD  THOSE IDIOTS FROM EARTH  C/P,P
WILSON, RICHARD  TIME OUT FOR TOMORROW  C/P:2,P
WILSON, ROBERT  TENTACLES OF DAWN, THE  P
WILSON, ROBERT ANTON  COSMIC TRIGGER/THE FINAL SECRET OF THE ILLUMINATI  P
WILSON, ROBERT ANTON  EARTH WILL SHAKE, THE [@HISTORICAL ILLUMINATUS #1]  C/P:2,LP,H
WILSON, ROBERT ANTON  HOMING PIGEONS [@SCHRODINGER’S CAT #3]  C/P,P
WILSON, ROBERT ANTON  ILLUMINATI PAPERS, THE  LP
WILSON, ROBERT ANTON  MASKS OF THE ILLUMINATI  H:2,P
WILSON, ROBERT ANTON  NATURE’S GOD [@HISTORICAL ILLUMINATUS #3]  P
WILSON, ROBERT ANTON  SCHRODINGER’S CAT TRILOGY [SCHRODINGER’S CAT 1,2,3]  H,LP
WILSON, ROBERT ANTON  SEX MAGICIANS, THE  SR-L
WILSON, ROBERT ANTON  TRICK TOP HAT, THE [@SCHRODINGER’S CAT #2]  C/P:2,P
WILSON, ROBERT ANTON  UNIVERSE NEXT DOOR, THE [@SCHRODINGER’S CAT 1]  C/P,P
WILSON, ROBERT ANTON  WIDOW’S SON, THE [@HISTORICAL ILLUMINATUS #2]  C/P,P
WILSON, ROBERT C.  CROOKED TREE  P
WILSON, ROBERT CHARLES  BIOS  H
WILSON, ROBERT CHARLES  BLIND LAKE  H

The MITSFS Pinkdex  September 24, 2007  Page 419
AUTHOR  TITLE  CODE
WILSON, ROBERT CHARLES  BRIDGE OF YEARS, A  C/P,LP,H
WILSON, ROBERT CHARLES  CHRONOLITHS, THE  C/P,P,H
WILSON, ROBERT CHARLES  DARWINIA  C/P,P,H
WILSON, ROBERT CHARLES  DIVIDE, THE  C/P,LP
WILSON, ROBERT CHARLES  GYPSIES  H,SFWA-P,P
WILSON, ROBERT CHARLES  HARVEST, THE  LP
WILSON, ROBERT CHARLES  HARVEST, THE (PROOF)  SR-LP
WILSON, ROBERT CHARLES  HIDDEN PLACE, A  C/P:2,SFWA-P,P
WILSON, ROBERT CHARLES  MEMORY WIRE  C/P:2,P
WILSON, ROBERT CHARLES  MYSTERIUM  H,LP
WILSON, ROBERT CHARLES  PERSEIDS AND OTHER STORIES, THE  H
WILSON, ROBERT CHARLES  SPIN  H
WILSON, ROBIN  PARAGONS  HA
WILSON, ROBIN SCOTT  CLARION  PA,C/PA
WILSON, ROBIN SCOTT  CLARION II  C/PA,PA
WILSON, ROBIN SCOTT  CLARION III  C/PA,PA
WILSON, ROBIN SCOTT  THOSE WHO CAN  C/PA,PA
WILSON, SNOO  SPACEACHE  C/P,P
WILSON, STEVE  LOST TRAVELLER, THE  C/P,P,H
WILSON, WILLIAM  DETOUR  P
WIMBERLY, LOWRY CHARLES  FOLKLORE IN THE ENGLISH AND SCOTTISH BALLADS  [C/RAN-H]
WINDLING, TERRI  ARMLESS MAIDEN, THE  LPA
WINDLING, TERRI/ARNOLD, MARK ALAN  BORDERLAND [BORDERLANDS 1]  PA,C/PA
WINDLING, TERRI/ARNOLD, MARK ALAN  BORDERTOWN [BORDERLANDS 2]  C/PA,PA
WINDLING, TERRI/ARNOLD, MARK ALAN  ELSEWHERE  C/PA,PA
WINDLING, TERRI/ARNOLD, MARK ALAN  ELSEWHERE II  C/PA,PA
WINDLING, TERRI/ARNOLD, MARK ALAN  ELSEWHERE III  C/PA,PA
WINDLING, TERRI/SHERMAN, DELIA  ESSENTIAL BORDERTOWN, THE [BORDERLAND]  LP
WINDLING, TERRI  FAERY!  C/PA,PA
WINDLING, TERRI  LIFE ON THE BORDER [BORDERLANDS 3]  PA,C/PA
WINDLING, TERRI  WOOD WIFE, THE  P
WINGERT, JENNIFER  GRASP THE STARS  P
WINGROVE, DAVID  BENEATH THE TREE OF HEAVEN [@CHUNG KUO #5]  C/P:2,LP
WINGROVE, DAVID  BROKEN WHEEL, THE [@CHUNG KUO #2]  C/P:2,P
WINGROVE, DAVID  DAYS OF BITTER STRENGTH [@CHUNG KUO #7]  C/P,LP,H
WINGROVE, DAVID  MARRIAGE OF THE LIVING DARK, THE [@CHUNG KUO 8]  P
WINGROVE, DAVID  MIDDLE KINGDOM, THE [@CHUNG KUO #1]  C/P,P,H
WINGROVE, DAVID  STONE WITHIN, THE [CHUNG KUO 4]  C/P:2,LP
WINGROVE, DAVID  WHITE MOON, RED DRAGON [@CHUNG KUO #6]  C/P:2,LP
WINGROVE, DAVID  WHITE MOUNTAIN, THE [@CHUNG KUO #3]  C/P,P,H
WINSKI, NORMAN  SWORD AND THE SORCERER, THE  C/P,P
WINSLOW, PAULINE GLEN  I, MARTHA ADAMS  C/P,P
WINSLOW, PAULINE GLEN  JUDGMENT DAY  P
WINSOR, FREDERICK/PARRY, MARIAN  SPACE CHILD’S MOTHER GOOSE, THE  H,LP
WINSOR, G. MCLEOD  STATION X  H

WINSTON, JOAN MAKING OF THE TREK CONVENTIONS, OR HOW TO THROW A PARTY FOR

12,000 OF YOUR MOST INTIMATE FRIENDS, THE

SFWA

WINTER, DOUGLAS E.  BLACK WINE  HA
WINTER, DOUGLAS E.  FACES OF FEAR  R/REF-H
WINTER, DOUGLAS E.  NIGHT VISIONS 5  SR-HA
WINTER, DOUGLAS E.  PRIME EVIL  HA,PA
WINTER, DOUGLAS E.  SKIN TRADE, THE  PA
WINTERBOTHAM, RUSS  JOYCE OF THE SECRET SQUADRON  H
WINTERBOTHAM, RUSS  MEN FROM ARCTURUS, THE  H
WINTERBOTHAM, RUSS  RED PLANET, THE  C/P:2,P
WINTERBOTHAM, RUSS  SPACE EGG, THE  C/P:2,P
WINTERFELD, HENRY  CASTAWAYS IN LILLIPUT  H
WINTERFELD, HENRY  STAR GIRL  C/P,LP
WINTERS, K. C.  CHOSEN/PLACEBO EFFECT, THE  P
WINTERS, LOGAN  HUNT THE BEAST DOWN [@SPECTROS #2]  P
WINTERS, LOGAN  NATCHEZ [@SPECTROS #3]  P
WINTERS, LOGAN  SILVER CANYON, THE [@SPECTROS #4]  P
WINTERS, LOGAN  SILVERADO [@SPECTROS #1]  P
WINTHROP, ELIZABETH  CASTLE IN THE ATTIC, THE  LP
WISE, HERBERT A./FRASER, PHYLLIS  GREAT TALES OF TERROR AND THE SUPERNATURAL  HA
WISEMAN, DAVID  ADAM’S COMMON  H
WISLER, G. CLIFTON  SEER, THE [ANTRIAN MESSENGER 2]  H
WISMAN, KEN  FROST ON THE WINDOW  LP
WISMER, DON  PLANET OF THE DEAD  P,SFWA-P
WISMER, DON  ROIL OF STARS, A  P,SFWA-P
WISMER, DON  WARRIOR PLANET  P
WISMER, DONALD  STARLUCK  P
WITCOVER, PAUL  DRACULA: ASYLUM  P
WITCOVER, PAUL  TUMBLING AFTER  H

The MITSFS Pinkdex  September 24, 2007  Page 420
AUTHOR  TITLE  CODE
WITCOVER, PAUL  WAKING BEAUTY  H
WITTIG, MONIQUE  ACROSS THE ACHERON  P
WIZARDS OF THE COAST  BEST OF THE REALMS BOOK I, THE [@FORGOTTEN REALMS]  PA
WIZARDS OF THE COAST  DRAGONS: WORLD AFIRE  H
WIZARDS OF THE COAST  SNEAK PREVIEW WINTER 2000  SR-P
WOBIG, ELLEN  YOUTH MONOPOLY, THE  DH048
WODEHOUSE, P. G.  LAUGHING GAS  P
WODHAMS, JACK  AUTHENTIC TOUCH, THE  C/P:2,P
WODHAMS, JACK  FUTURE WAR  SR-P
WODHAMS, JACK  LOOKING FOR BLUCHER  SR-P
WODHAMS, JACK  RYN  SR-P
WOLD, ALLEN L.  BELOW THE THRESHOLD [V]  C/P,SFWA-P,P
WOLD, ALLEN L.  CRIVIT EXPERIMENT, THE [V]  C/P,P
WOLD, ALLEN L.  CROWN OF THE SERPENT  C/P,P
WOLD, ALLEN L.  EYE IN THE STONE, THE  P
WOLD, ALLEN L.  JEWELS OF THE DRAGON  P
WOLD, ALLEN L.  LAIR OF THE CYCLOPS  P
WOLD, ALLEN L.  PLANET MASTERS, THE  H
WOLD, ALLEN L.  PLANET MASTERS, THE (PROOF)  SR-LP
WOLD, ALLEN L.  PURSUIT OF DIANA, THE [V]  C/P,SFWA-P,P
WOLD, ALLEN L.  STAR GOD  H,P
WOLD, ALLEN L.  STAR GOD (PROOF)  SR-LP
WOLF, CHRIS L./MAIKOWSKI, MICHAEL F.  FIRE IN THE SKY  C/P,P
WOLF, GARY K.  GENERATION REMOVED, A  H,SFWA
WOLF, GARY K.  KILLERBOWL  SFWA
WOLF, GARY K.  RESURRECTIONIST, THE  H,SFWA
WOLF, GARY K.  WHO CENSORED ROGER RABBIT? [ROGER RABBIT 1]  C/P:2,P
WOLF, GARY K.  WHO P-P-P-PLUGGED ROGER RABBIT? [ROGER RABBIT 2]  H
WOLF, JACK C./WOLF, BARBARA  GHOSTS, CASTLES, AND VICTIMS: TALES OF GOTHIC HORROR  PA
WOLF, JACK C./FITZ GERALD, GREGORY  PAST, PRESENT AND FUTURE PERFECT  PA
WOLF, JACK C./WOLF, BARBARA  TALES OF THE OCCULT  PA
WOLF, LEONARD  DREAM OF DRACULA, A  R/REF-P
WOLFE, AARON  INVASION  SR-P
WOLFE, BERNARD  LIMBO  C/P,P,H
WOLFE, BERNARD  LIMBO ’90  SR-P
WOLFE, CHRIS ANNE  FIRES OF AGGAR [AGGAR 2]  LP
WOLFE, CHRIS ANNE  SHADOWS OF AGGAR [AGGAR 1]  LP
WOLFE, GARY  SCIENCE FICTION DIALOGUES  R/REF-P
WOLFE, GENE  BOOK OF THE NEW SUN, THE [NEW SUN 1,2,3,4]  H
WOLFE, GENE  BOOK OF THE SHORT SUN, THE [BOOK OF THE SHORT SUN 1,2,3]  H:2
WOLFE, GENE  CALDE OF THE LONG SUN [LONG SUN 3]  C/P,P,H
WOLFE, GENE  CASTLE OF DAYS  LP
WOLFE, GENE  CASTLE OF THE OTTER, THE  R/REF-H,C/REF-H
WOLFE, GENE  CASTLEVIEW  H,SFWA-P,P
WOLFE, GENE  CITADEL OF THE AUTARCH, THE [NEW SUN 4]  C/P,P,H
WOLFE, GENE  CLAW OF THE CONCILIATOR, THE [NEW SUN 2]  C/P,P,H
WOLFE, GENE  DEATH OF DOCTOR ISLAND, THE  SFWA-TD25,TD25
WOLFE, GENE  DEVIL IN A FOREST, THE  C/P:2,P
WOLFE, GENE  ENDANGERED SPECIES  C/P:2,SFWA-P,P
WOLFE, GENE  EPIPHANY OF THE LONG SUN [LONG SUN 3,4]  H
WOLFE, GENE  EXODUS FROM THE LONG SUN [LONG SUN 4]  C/P,P,H
WOLFE, GENE  FIFTH HEAD OF CERBERUS, THE  C/P,SFWA,P,H
WOLFE, GENE  FREE LIVE FREE  C/P,SFWA-P,P,H
WOLFE, GENE  GENE WOLFE’S BOOK OF DAYS  H,SFWA
WOLFE, GENE  IN GREEN’S JUNGLES [BOOK OF THE SHORT SUN 2]  LP
WOLFE, GENE  INNOCENTS ABOARD  H

WOLFE, GENE ISLAND OF DOCTOR DEATH AND OTHER STORIES AND OTHER STORIES,

THE

C/P:3,P

WOLFE, GENE  KNIGHT, THE [WIZARD KNIGHT 1]  H
WOLFE, GENE  LAKE OF THE LONG SUN [LONG SUN 2]  H,P
WOLFE, GENE  LITANY OF THE LONG SUN [LONG SUN 1,2]  H,LP
WOLFE, GENE  NIGHTSIDE THE LONG SUN [LONG SUN 1]  H,P
WOLFE, GENE  ON BLUE’S WATERS [BOOK OF THE SHORT SUN 1]  H,LP
WOLFE, GENE  OPERATION ARES  C/P:2,P
WOLFE, GENE  PANDORA BY HOLLY HOLLANDER  H
WOLFE, GENE  PEACE  C/P,SFWA,P,H
WOLFE, GENE  PLAN(E)T ENGINEERING  SR-H
WOLFE, GENE  RETURN TO THE WHORL [BOOK OF THE SHORT SUN 3]  LP
WOLFE, GENE  SEVEN AMERICAN NIGHTS  SFWA-TD10,TD10
WOLFE, GENE  SHADOW OF THE TORTURER, THE [NEW SUN 1]  C/P,P,H
WOLFE, GENE  SOLDIER OF ARETE [SOLDIER OF THE MIST 2]  H:2,SFWA-P,P
WOLFE, GENE  SOLDIER OF SIDON  H
WOLFE, GENE  SOLDIER OF THE MIST [SOLDIER OF THE MIST 1]  H:2,P
WOLFE, GENE  STARWATER STRAINS  H

The MITSFS Pinkdex  September 24, 2007  Page 421
AUTHOR  TITLE  CODE
WOLFE, GENE  STOREYS FROM THE OLD HOTEL  H
WOLFE, GENE  STRANGE TRAVELERS  LP
WOLFE, GENE  SWORD OF THE LICTOR, THE [NEW SUN 3]  C/P,SFWA-P,P,H
WOLFE, GENE  THERE ARE DOORS  H,SFWA-P,P
WOLFE, GENE  URTH OF THE NEW SUN, THE [NEW SUN 5]  C/P,SFWA-P,P,H
WOLFE, GENE  WIZARD, THE [WIZARD KNIGHT 2]  H
WOLFE, L. STEPHEN/WYSACK, ROY L.  HANDBOOK FOR SPACE PIONEERS  R/REF-H
WOLFE, LOUIS  JOURNEY OF THE OCEANAUTS  C/P,P
WOLFF, ROBERT SHERMAN  CAVES OF MARS (PROOF) [FALCON GOLD 1]  SR-LP
WOLFMAN, MARV  DOOMSDAY [FANTASTIC FOUR]  C/P,P
WOLFMAN, MARV  HISTORY OF THE DC UNIVERSE  CX
WOLFMAN, MARV/WHITE, TED  OZ ENCOUNTER, THE [WEIRD HEROES 5]  C/P,P
WOLLHEIM, DONALD A.  ACE SCIENCE FICTION READER  C/PA,PA
WOLLHEIM, DONALD A.  ADVENTURES IN THE FAR FUTURE  DD073
WOLLHEIM, DONALD A.  ADVENTURES ON OTHER PLANETS  PA,C/PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1972 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1973 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1974 ANNUAL WORLD’S BEST SF, THE  PA,HA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1975 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1976 ANNUAL WORLD’S BEST SF, THE  PA,HA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1977 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1978 ANNUAL WORLD’S BEST SF, THE  PA,HA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1979 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1980 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1981 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1982 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1983 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1984 ANNUAL WORLD’S BEST SF, THE  HA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1985 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1986 ANNUAL WORLD’S BEST SF, THE  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1987 ANNUAL WORLD’S BEST SF, THE  HA,PA
WOLLHEIM, DONALD A.  1988 ANNUAL WORLD’S BEST SF, THE  HA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1989 ANNUAL WORLD’S BEST SF, THE  HA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  1990 ANNUAL WORLD’S BEST SF, THE  C/PA,PA

WOLLHEIM, DONALD A./ ERNSBERGER, GEORGE AVON FANTASY READER, THE

PA,C/PA

WOLLHEIM, DONALD A./ ERNSBERGER, GEORGE 2ND AVON FANTASY READER, THE

C/PA,PA

WOLLHEIM, DONALD A.  BEST FROM THE REST OF THE WORLD, THE  HA,PA,SFWA
WOLLHEIM, DONALD A.  DAW SCIENCE FICTION READER, THE  PA,C/PA
WOLLHEIM, DONALD A.  EARTH IN PERIL, THE  DD205
WOLLHEIM, DONALD A.  END OF THE WORLD, THE  PA,C/PA
WOLLHEIM, DONALD A.  FLIGHT INTO SPACE  SR-PA,C/PA
WOLLHEIM, DONALD A.  HIDDEN PLANET, THE  PA,C/PA
WOLLHEIM, DONALD A.  MACABRE READER, THE  PA,C/PA
WOLLHEIM, DONALD A.  MEN FROM ARIEL, THE  SR-H
WOLLHEIM, DONALD A.  MEN ON THE MOON  DD277,PA,C/PA
WOLLHEIM, DONALD A.  MIKE MARS AND THE MYSTERY SATELLITE [@MIKE MARS #7]  P
WOLLHEIM, DONALD A.  MIKE MARS AROUND THE MOON [@MIKE MARS #8]  P
WOLLHEIM, DONALD A.  MIKE MARS ASTRONAUT [@MIKE MARS #1]  P
WOLLHEIM, DONALD A.  MIKE MARS AT CAPE KENNEDY [@MIKE MARS #3]  P
WOLLHEIM, DONALD A.  MIKE MARS FLIES THE DYNA-SOAR [@MIKE MARS #5]  P
WOLLHEIM, DONALD A.  MIKE MARS FLIES THE X-15 [@MIKE MARS #2]  P
WOLLHEIM, DONALD A.  MIKE MARS IN ORBIT [MIKE MARS #4]  H,@P
WOLLHEIM, DONALD A.  MIKE MARS, SOUTH POLE SPACEMAN [@MIKE MARS #6]  P
WOLLHEIM, DONALD A.  MORE ADVENTURES ON OTHER PLANETS  PA,C/PA
WOLLHEIM, DONALD A.  MORE MACABRE  PA
WOLLHEIM, DONALD A.  OPERATION: PHANTASY  SR-H
WOLLHEIM, DONALD A.  POCKET BOOK OF SCIENCE FICTION, THE  SR-PA,C/PA
WOLLHEIM, DONALD A.  PORTABLE NOVELS OF SCIENCE, THE  HA
WOLLHEIM, DONALD A.  SECRET OF SATURN’S RINGS, THE  SR-H,H,P
WOLLHEIM, DONALD A.  SECRET OF THE MARTIAN MOONS, THE  H,SR-H,P
WOLLHEIM, DONALD A.  SECRET OF THE NINTH PLANET, THE  P,H,SR-H
WOLLHEIM, DONALD A.  SWORDSMEN IN THE SKY  PA,C/PA
WOLLHEIM, DONALD A.  TALES OF OUTER SPACE  DD073
WOLLHEIM, DONALD A.  TERROR IN THE MODERN VEIN  HA
WOLLHEIM, DONALD A.  TWO DOZEN DRAGON EGGS  SR-P
WOLLHEIM, DONALD A.  ULTIMATE INVADER AND OTHER SCIENCE FICTION, THE  SR-DD044
WOLLHEIM, DONALD A.  UNIVERSE MAKERS, THE  R/REF-H
WOLLHEIM, DONALD A.  UP THERE AND OTHER STRANGE DIRECTIONS  SR-H
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  WOLLHEIM’S WORLD’S BEST SF SERIES ONE  PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  WOLLHEIM’S WORLD’S BEST SF SERIES FOUR  PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  WOLLHEIM’S WORLD’S BEST SF SERIES NINE  PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1965  PA,C/PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1966  PA,C/PA

The MITSFS Pinkdex  September 24, 2007  Page 422
AUTHOR  TITLE  CODE
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1967  PA,C/PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1968  PA,C/PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1969  HA,PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1970  C/PA,PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION 1971  C/PA,PA
WOLLHEIM, DONALD A./CARR, TERRY  WORLD’S BEST SCIENCE FICTION FIRST SERIES  C/PA,PA
WOLLHEIM, DONALD A./SAHA, ARTHUR W.  WORLD’S BEST SF 7  PA

WOLLHEIM, ELIZABETH R./ GILBERT, SHEILA E.

DAW 30TH ANNIVERSARY FANTASY

PA

WOLLHEIM, ELIZABETH R./ GILBERT, SHEILA E.

DAW 30TH ANNIVERSARY SCIENCE FICTION

HA

WOLVERTON, BASIL  SPACEHAWK  CX
WOLVERTON, DAVE  BEYOND THE GATE [GOLDEN QUEEN 2]  H,SFWA:2,P
WOLVERTON, DAVE  COURTSHIP OF PRINCESS LEIA, THE [STAR WARS]  H:2,P
WOLVERTON, DAVE  GOLDEN QUEEN, THE [GOLDEN QUEEN 1]  P
WOLVERTON, DAVE  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 9  PA
WOLVERTON, DAVE  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 11  PA
WOLVERTON, DAVE  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 13  PA
WOLVERTON, DAVE  L. RON HUBBARD PRESENTS WRITERS OF THE FUTURE 14  PA
WOLVERTON, DAVE  LORDS OF THE SEVENTH SWARM [GOLDEN QUEEN 3]  H,P
WOLVERTON, DAVE  ON MY WAY TO PARADISE  C/P:2,SFWA-P,P
WOLVERTON, DAVE  PATH OF THE HERO  P
WOLVERTON, DAVE  SERPENT CATCH  C/P:2,P
WOLVERTON, DAVE  VERY STRANGE TRIP, THE (PROOF)  SR-LP
WOMACK, JACK  AMBIENT [DRYCO 1]  H,P
WOMACK, JACK  ELVISSEY [DRYCO 4]  H:2,LP
WOMACK, JACK  HEATHERN [DRYCO 3]  P,SFWA
WOMACK, JACK  RANDOM ACTS OF SENSELESS VIOLENCE [DRYCO 5]  H,LP
WOMACK, JACK  TERRAPLANE [DRYCO 2]  C/P,SFWA-P,P,H
WOOD, BARI  TRIBE, THE  P

WOOD, BRIAN ART OF GEORGE R.R. MARTIN’S A SONG OF ICE AND FIRE, THE

[SONG OF ICE AND FIRE]

A

WOOD, BRIDGET  LOST PRINCE, THE [WOLFKINGS]  P
WOOD, BRIDGET  WOLFKING [WOLFKINGS]  C/P,LP
WOOD, CHRISTOPHER  JAMES BOND AND MOONRAKER  P
WOOD, EDWARD  CHECKLIST TO FANTASTIC UNIVERSE  SRL/IND
WOOD, N. LEE  BLOODRIGHTS  P
WOOD, N. LEE  FARADAY’S ORPHANS  C/P,P
WOOD, N. LEE  LOOKING FOR THE MAHDI  C/P,LP
WOOD, N. LEE  MASTER OF NONE  LP
WOODALL, CLIVE  ONE FOR SORROW, TWO FOR JOY  H

WOODARD, EDWIN/ BISCHOFF, HEATHER WOODARD STOREHOUSES OF THE SNOW

P

WOODBURY, FRANCINE G.  SHADE AND SHADOW  P
WOODCOCK, P. G.  SHORT DICTIONARY OF MYTHOLOGY  [C/RAN-H]
WOODCOTT, KEITH  I SPEAK FOR EARTH  DD497
WOODCOTT, KEITH  LADDER IN THE SKY, THE  DF141
WOODCOTT, KEITH  MARTIAN SPHINX, THE  C/P:2,P
WOODCOTT, KEITH  PSIONIC MENACE, THE  DF199
WOODING, CHRIS  WEAVERS OF SARAMYR [BRAIDED PATH 1]  LP
WOODLEY, RICHARD  ARK OF DOOM [MAN FROM ATLANTIS 4]  P
WOODLEY, RICHARD  DEATH SCOUTS [MAN FROM ATLANTIS 2]  P
WOODLEY, RICHARD  IT’S ALIVE!  P
WOODLEY, RICHARD  KILLER SPORES [MAN FROM ATLANTIS 3]  P
WOODLEY, RICHARD  MAN FROM ATLANTIS [MAN FROM ATLANTIS 1]  C/P,P
WOODROFFE, PATRICK  MYTHOPOEIKON  A
WOODROFFE, PATRICK  SECOND EARTH, THE  A
WOODS, CHRISTINA  WIZARD’S CURSE [@DRAGONLANCE NEW ADVENTURES TRINISTYR #1]  LP
WOODWORTH, STEPHEN  THROUGH VIOLET EYES [VIOLETS 1]  H
WOODWORTH, STEPHEN  WITH RED HANDS [VIOLETS 2]  H
WOOLF, VIRGINIA  ORLANDO  P
WOOLFOLK, WILLIAM  ADAM PROJECT, THE  P
WOOLLEY, PERSIA  QUEEN OF THE SUMMER STARS  H
WOOLRICH, CORNELL  DOOM STONE, THE  P
WOOLRICH, CORNELL  FANTASTIC STORIES OF CORNELL WOOLRICH, THE  H
WOOLVERTON, LINDA  STAR WIND  SFWA
WORLD SCIENCE FICTION SOCIETY  INTERSECTION  R/REF-H
WORMSER, RICHARD  PAN SATYRUS  C/P:2,P
WORMSER, RICHARD  WILD WILD WEST, THE  C/P,P
WORNER, HANS  WIR FANDEN MENSCHEN  FH
WORTHY, PETER A.  REHEARSALS FOR OBLIVION: ACT I  LPA
WOUK, HERMAN  "LOMOKOME" PAPERS, THE  C/P,P
WREDE, PATRICIA C.  BOOK OF ENCHANTMENTS  H
WREDE, PATRICIA C.  CALLING ON DRAGONS [ENCHANTED FOREST 3]  P

AUTHOR  TITLE  CODE
WREDE, PATRICIA C.  CAUGHT IN CRYSTAL [LYRA 1]  C/P:2,SFWA-P,P
WREDE, PATRICIA C.  DAUGHTER OF WITCHES [LYRA 4]  C/P:2,P
WREDE, PATRICIA C.  DEALING WITH DRAGONS [ENCHANTED FOREST 1]  P
WREDE, PATRICIA C./STEVERMER, CAROLINE GRAND TOUR, THE [CECILIA & KATHERINE 2]  H
WREDE, PATRICIA C.  HARP OF IMACH THYSSEL, THE [LYRA 5]  C/P:2,P
WREDE, PATRICIA C.  MAGIC & MALICE [MAIRELON AND KIM 1,2]  H
WREDE, PATRICIA C./STEVERMER, CAROLINE MAGICIANS OF QUALITY [CECILIA & KATHERINE 1,2]  H
WREDE, PATRICIA C.  MAGICIAN’S WARD [MAIRELON AND KIM 2]  H
WREDE, PATRICIA C.  MAIRELON THE MAGICIAN [MAIRELON AND KIM 1]  H,P
WREDE, PATRICIA C./STEVERMER, CAROLINE MISLAID MAGICIAN, THE [CECILIA & KATHERINE 3]  H:2
WREDE, PATRICIA C.  RAVEN RING, THE [LYRA 2]  P
WREDE, PATRICIA C.  SEARCHING FOR DRAGONS [ENCHANTED FOREST 2]  P
WREDE, PATRICIA C.  SEVEN TOWERS, THE  C/P:2,P
WREDE, PATRICIA C.  SHADOW MAGIC [LYRA 3]  C/P:2,P
WREDE, PATRICIA C.  SHADOWS OVER LYRA [LYRA 1,2,3]  H,LP
WREDE, PATRICIA C.  SNOW WHITE AND ROSE RED  H:2,P
WREDE, PATRICIA C./STEVERMER, CAROLINE SORCERY AND CECELIA [CECILIA & KATHERINE 1]  C/P,P
WREDE, PATRICIA C.  TALKING TO DRAGONS [ENCHANTED FOREST 4]  P
WREN, M. K.  GIFT UPON THE SHORE, A  P
WREN, M. K.  HOUSE OF THE WOLF [PHOENIX LEGACY 3]  C/P:2,P
WREN, M. K.  SHADOW OF THE SWAN [PHOENIX LEGACY 2]  C/P:2,P
WREN, M. K.  SWORD OF THE LAMB [PHOENIX LEGACY 1]  C/P:2,P
WREN, THOMAS  DOOMSDAY EFFECT, THE  P
WRENCH, SARA J.  DUKE OF SUMAVA, THE  P
WRIGHT, AUSTIN TAPPAN  ISLANDIA  SR-H,C/P,P
WRIGHT, GARY  ROAD WEST, THE  P
WRIGHT, GRAHAME  JOG RUMMAGE  P
WRIGHT, HELEN S.  MATTER OF OATHS, A  C/P,P
WRIGHT, JOHN C.  FUGITIVES OF CHAOS [ORPHANS OF CHAOS 2]  H
WRIGHT, JOHN C.  GOLDEN AGE, THE [GOLDEN AGE 1]  H:2
WRIGHT, JOHN C.  GOLDEN TRANSCENDENCE, THE [GOLDEN AGE 3]  H
WRIGHT, JOHN C.  MISTS OF EVERNESS [WAR OF THE DREAMING 2]  H
WRIGHT, JOHN C.  ORPHANS OF CHAOS [ORPHANS OF CHAOS 1]  H:2
WRIGHT, JOHN C.  PHOENIX EXULTANT, THE [GOLDEN AGE 2]  H
WRIGHT, JOHN C.  TITANS OF CHAOS [ORPHANS OF CHAOS 3]  H
WRIGHT, KENNETH  MYSTERIOUS PLANET, THE  H:2
WRIGHT, LAN  CREEPING SHROUD, THE  SR-P
WRIGHT, LAN  EXILE FROM XANADU  DM103
WRIGHT, LAN  LAST HOPE OF EARTH, THE  C/P:2,P
WRIGHT, LAN  MAN CALLED DESTINY, A  DD311
WRIGHT, LAN  PICTURES OF PAVANNE, THE  DH048
WRIGHT, LAN  QUI PARLE DE CONQUETE  FP
WRIGHT, LAN  WHO SPEAKS OF CONQUEST?  DD205
WRIGHT, LEE/SHEEHAN, RICHARD G.  THESE WILL CHILL YOU  PA
WRIGHT, LEE/SHEEHAN, RICHARD G.  WAKE UP SCREAMING  PA
WRIGHT, RONALD  SCIENTIFIC ROMANCE, A  H,LP
WRIGHT, S. FOWLER  AMPHIBIANS, THE  C/P:2,GN1.04
WRIGHT, S. FOWLER  DAWN  H
WRIGHT, S. FOWLER  DELUGE  H
WRIGHT, S. FOWLER  DELUGE (PROOF)  SR-LP
WRIGHT, S. FOWLER  DREAM, OR THE SIMIAN MAID  H
WRIGHT, S. FOWLER  ELFWIN  H
WRIGHT, S. FOWLER  ISLAND OF CAPTAIN SPARROW, THE  H
WRIGHT, S. FOWLER  SPIDERS’ WAR  H
WRIGHT, S. FOWLER  VENGEANCE OF GWA  H
WRIGHT, S. FOWLER  WAR OF 1938, THE  SR-H
WRIGHT, S. FOWLER  WORLD BELOW, THE  SR-H,GN1.05
WRIGHT, S. FOWLER  WORLD BELOW, THE (WITH "THE AMPHIBIANS")  LP
WRIGHT, SUSAN  BADLANDS 1, THE [STAR TREK]  C/P:2,P
WRIGHT, SUSAN  BADLANDS 2, THE [STAR TREK]  C/P,P
WRIGHT, SUSAN  BEST AND THE BRIGHTEST, THE [STAR TREK]  P
WRIGHT, SUSAN  DARK PASSIONS BOOK ONE [STAR TREK]  C/P,P
WRIGHT, SUSAN  DARK PASSIONS BOOK TWO [STAR TREK]  C/P,P
WRIGHT, SUSAN  ONE SMALL STEP [STAR TREK GATEWAYS #1]  P
WRIGHT, SUSAN  SINS OF COMMISSION [STAR TREK TNG]  C/P:2,P
WRIGHT, SUSAN  SLAVE MASTERS [SLAVE TRADE 2]  P
WRIGHT, SUSAN  SLAVE TRADE [SLAVE TRADE 1]  P
WRIGHT, SUSAN  SLAVES UNCHAINED [SLAVE TRADE 3]  P
WRIGHT, SUSAN  TEMPEST, THE [STAR TREK DS9]  C/P,P
WRIGHT, SUSAN  TO SERVE AND SUBMIT  P
WRIGHT, SUSAN  VIOLATIONS [STAR TREK VOYAGER]  C/P:2,P
WRIGHT, T. M.  ASCENDING, THE  H
WRIGHT, T. M.  BOUNDARIES  SFWA-P
WRIGHT, T. M.  ISLAND, THE  SFWA-P

The MITSFS Pinkdex  September 24, 2007  Page 424
AUTHOR  TITLE  CODE
WRIGHT, T. M.  PEOPLE OF THE DARK, THE  SFWA-P:2
WRIGHT, T. M.  PLACE, THE  SFWA-P
WRIGHT, T. M.  PLAYGROUND, THE  SFWA-P
WRIGHT, T. M.  STRANGE SEED  P,SFWA-P
WRIGHT, T. M.  WAITING ROOM, THE  SFWA-P
WRIGHT, THOMAS  WORSHIP OF THE GENERATIVE POWERS DURING THE MIDDLE AGES, THE  [C/RAN-P]
WRIGHTSON, BERNI  BACK FOR MORE  CX
WRIGHTSON, PATRICIA  DARK BRIGHT WATER, THE  C/P,P
WRIGHTSON, PATRICIA  DOWN TO EARTH  P
WRIGHTSON, PATRICIA  ICE IS COMING, THE  C/P:2,P
WRIGHTSON, PATRICIA  JOURNEY BEHIND THE WIND  C/P,P
WRIGHTSON, PATRICIA  LITTLE FEAR, A  SFWA
WRIGHTSON, PATRICIA  NARGUN AND THE STARS, THE  LP
WRIGHTSON, PATRICIA  OLDER KIND OF MAGIC, AN  P
WSFS  SPLITTING INFINITY  VLP
WU, DINGBO/MURPHY, PATRICK D.  SCIENCE FICTION FROM CHINA  HA
WU, WILLIAM F.  COSMIC BOMBER, THE [DR. BONES]  P
WU, WILLIAM F.  CYBORG [@ROBOT CITY #3]  C/P:2,SFWA-P,P
WU, WILLIAM F.  DICTATOR [@ROBOTS IN TIME 4]  C/P,P
WU, WILLIAM F.  EMPEROR [@ROBOTS IN TIME 5]  P
WU, WILLIAM F.  HONG ON THE RANGE [JACK HONG]  H
WU, WILLIAM F.  IN LUNACY [@MUTANT CHRONICLES APOSTLE OF INSANITY 1]  P
WU, WILLIAM F.  INVADER [@ROBOTS IN TIME 6]  P
WU, WILLIAM F.  MARAUDER [@ROBOTS IN TIME 2]  P,SFWA-P
WU, WILLIAM F.  MASTERPLAY  C/P:2,P
WU, WILLIAM F.  PERIHELION [@ROBOT CITY #6]  C/P:2,SFWA-P,P
WU, WILLIAM F.  PREDATOR [@ROBOTS IN TIME 1]  P,SFWA-P
WU, WILLIAM F.  ROBIN HOOD AMBUSH, THE [TIME TOURS 1]  P
WU, WILLIAM F.  SHADE OF LO MAN GONG, THE [JACK HONG 1]  B
WU, WILLIAM F.  SHAUNESSY FONG [JACK HONG 2]  B
WU, WILLIAM F.  WARRIOR [@ROBOTS IN TIME 3]  P
WU, WILLIAM F.  WONG’S LOST AND FOUND EMPORIUM AND OTHER ODDITIES  SR-H
WUL, STEFAN  ARMADILHA EM ZARKASS  FP
WUL, STEFAN  CATIVEIRO HUMANO  FP
WUL, STEFAN  IMPERIO DOS MUTANTES, O  FP
WUL, STEFAN  TEMPLE OF THE PAST, THE  H,SFWA
WULF, RICH  FLIGHT OF THE DYING SUN [@EBERRON HEIRS OF ASH #2]  P
WULF, RICH  VOYAGE OF THE MOURNING DAWN [@EBERRON HEIRS OF ASH #1]  P
WULFF, RICH  WIND OF JUSTICE [@FIVE RINGS FOUR WINDS #3]  P
WUPI  DESTINATION: TERRE!  CX
WURTS, JANNY  CURSE OF THE MISTWRAITH [WARS OF LIGHT AND SHADOW 1]  C/P:3,SFWA,P
WURTS, JANNY  CYCLE OF FIRE, THE [CYCLE OF FIRE 1,2,3]  LP

WURTS, JANNY FUGITIVE PRINCE [WARS OF LIGHT AND SHADOW 3/

ALLIANCE OF LIGHT 1]

H,P

WURTS, JANNY GRAND CONSPIRACY [WARS OF LIGHT AND SHADOW 4/

ALLIANCE OF LIGHT 2]

H

WURTS, JANNY  KEEPER OF THE KEYS [CYCLE OF FIRE 2]  C/P:2,P
WURTS, JANNY  MASTER OF WHITE STORM, THE  C/P:2,SFWA-P,P
WURTS, JANNY  PERIL’S GATE [WARS OF LIGHT AND SHADOW 5/ALLIANCE OF LIGHT 3]  H
WURTS, JANNY  SHADOWFANE [CYCLE OF FIRE 3]  C/P:2,P

WURTS, JANNY SHIPS OF MERIOR (COMPLETE) [WARS OF LIGHT AND SHADOW 2/

SHIPS OF MERIOR 1,2]

H

WURTS, JANNY SHIPS OF MERIOR (PARTIAL) [WARS OF LIGHT AND SHADOW 2/

SHIPS OF MERIOR 1]

C/P:2,P

WURTS, JANNY  SORCERER’S LEGACY  C/P:2,P
WURTS, JANNY  STORMWARDEN [CYCLE OF FIRE 1]  C/P:2,SFWA-P,P
WURTS, JANNY  THAT WAY LIES CAMELOT  H,P
WURTS, JANNY  TO RIDE HELL’S CHASM  H

WURTS, JANNY WARHOST OF VASTMARK [WARS OF LIGHT AND SHADOW 2/

SHIPS OF MERIOR 2]

C/P,P

WYATT, JAMES  IN THE CLAWS OF THE TIGER [@EBERRON WAR-TORN #3]  P
WYATT, JOAN  MIDDLE-EARTH ALBUM, A  A
WYATT, LEE  FLESH HUNTERS, THE  SR-L
WYATT, STEPHEN  GREATEST SHOW IN THE GALAXY, THE [DOCTOR WHO]  P
WYATT, STEPHEN  PARADISE TOWERS [DOCTOR WHO]  P
WYLDE, THOMAS  GARUKAN BLOOD [DR. BONES]  P
WYLDE, THOMAS  JOURNEY TO RILLA [DR. BONES]  P
WYLDE, THOMAS  PITFALL [ALIEN SPEEDWAY 2]  C/P:2,P
WYLDE, THOMAS  PITFALL (PROOF) [ALIEN SPEEDWAY 2]  SR-LP
WYLDE, THOMAS  WEB, THE [ALIEN SPEEDWAY 3]  C/P:2,P
WYLER, ROSE/AMES, GERALD  GOLDEN BOOK OF ASTRONOMY  [C/RAN-H]
WYLIE, JONATHAN  AGE OF CHAOS, THE [@UNBALANCED EARTH #3]  C/P,P
WYLIE, JONATHAN  CENTER OF THE CIRCLE, THE [SERVANTS OF ARK 2]  C/P,P
WYLIE, JONATHAN  DREAMS OF STONE [@UNBALANCED EARTH #1]  P

AUTHOR  TITLE  CODE
WYLIE, JONATHAN  FIRST NAMED, THE [SERVANTS OF ARK 1]  C/P,P
WYLIE, JONATHAN  LIGHTLESS KINGDOM, THE [@UNBALANCED EARTH #2]  C/P,P
WYLIE, JONATHAN  MAGE-BORN CHILD, THE [SERVANTS OF ARK 3]  C/P:2,P
WYLIE, PHILIP/BALMER, EDWIN  AFTER WORLDS COLLIDE [WHEN WORLDS COLLIDE 2]  C/P,P,H:2
WYLIE, PHILIP  ANSWER, THE  C/P,P
WYLIE, PHILIP  DISAPPEARANCE, THE  C/P,P
WYLIE, PHILIP  END OF THE DREAM, THE  C/P,P,H
WYLIE, PHILIP  GLADIATOR  H,LP,P
WYLIE, PHILIP  LOS ANGELES: A.D. 2017  C/P:2,P
WYLIE, PHILIP  MURDERER INVISIBLE, THE  C/P,P
WYLIE, PHILIP  NINE RITTENHOUSE SQUARE  P
WYLIE, PHILIP  SMUGGLED ATOM BOMB, THE  C/P,P
WYLIE, PHILIP  SPY WHO SPOKE PORPOISE, THE  C/P,P
WYLIE, PHILIP  TOMORROW!  C/P,P,H
WYLIE, PHILIP  TRIUMPH  C/P:2,P
WYLIE, PHILIP/BALMER, EDWIN  WHEN WORLDS COLLIDE [WHEN WORLDS COLLIDE 1]  C/P,P,H:2
WYNDHAM, JOHN  CHOCKY  H:2,P
WYNDHAM, JOHN  CHRYSALIDS, THE  C/P:2,P
WYNDHAM, JOHN  CONSIDER HER WAYS AND OTHERS  SR-P,C/P
WYNDHAM, JOHN  DAY OF THE TRIFFIDS, THE (AMERICAN) [TRIFFIDS 1]  C/P,P,H
WYNDHAM, JOHN  DAY OF THE TRIFFIDS, THE (BRITISH) [TRIFFIDS 1]  SR-P,C/P
WYNDHAM, JOHN  INFINITE MOMENT, THE  C/P:2,P
WYNDHAM, JOHN  JIZZLE  SR-P
WYNDHAM, JOHN  KRAKEN WAKES, THE  SR-P,C/P:2
WYNDHAM, JOHN  MIDWICH CUCKOOS, THE  H:2,SR-H,P
WYNDHAM, JOHN  NO PLACE LIKE EARTH  SR-H
WYNDHAM, JOHN  OUT OF THE DEEPS  C/P:2,P
WYNDHAM, JOHN/PARKES, LUCAS  OUTWARD URGE, THE  C/P:2,P
WYNDHAM, JOHN  RE-BIRTH  C/P:2,P
WYNDHAM, JOHN  REVOLT OF THE TRIFFIDS  SR-P,C/P
WYNDHAM, JOHN  SECRET PEOPLE, THE  C/P,P
WYNDHAM, JOHN  SEEDS OF TIME, THE  SR-P,C/P
WYNDHAM, JOHN  SLEEPERS OF MARS  SR-P,C/P
WYNDHAM, JOHN  STOWAWAY TO MARS  C/P:2,P
WYNDHAM, JOHN  TALES OF GOOSEFLESH AND LAUGHTER  C/P:2,P
WYNDHAM, JOHN  TROUBLE WITH LICHEN (AMERICAN)  C/P,P,H
WYNDHAM, JOHN  TROUBLE WITH LICHEN (BRITISH)  SR-P
WYNDHAM, JOHN  VILLAGE OF THE DAMNED  C/P:2,P
WYNDHAM, JOHN  WANDERERS OF TIME  SR-P
WYNDHAM, JOHN  WEB  SR-P
WYNORSKI, JAMES  THEY CAME FROM OUTER SPACE  SFWA

AUTHOR  TITLE  CODE
X, DOCTOR  INTERN, THE  [C/RAN-P]
XI TSUN TSAI  MONKEY VISITS A BLACK HOLE, THE (CHINESE)  FP
XI TSUN TSAI  SEEDING THE CLOUDS (CHINESE)  FP

AUTHOR  TITLE  CODE
YABLONSKY, LEWIS  ROBOPATHS  [C/RAN-H,C/RAN-P]
YAMADA, MASAKI  AFTER THE LONG GOODBYE [GHOST IN THE SHELL 2]  H
YARBRO, CHELSEA QUINN  ANGRY ANGEL, THE [SISTERS OF THE NIGHT]  H,LP
YARBRO, CHELSEA QUINN  ARIOSTO  C/P:2,P
YARBRO, CHELSEA QUINN  BAD MEDICINE [CHARLIE MOON 1]  P
YARBRO, CHELSEA QUINN  BAROQUE FABLE, A  P
YARBRO, CHELSEA QUINN  BEASTNIGHTS  P
YARBRO, CHELSEA QUINN  BETTER IN THE DARK [SAINT-GERMAIN 8]  H,LP
YARBRO, CHELSEA QUINN  BLOOD GAMES [SAINT-GERMAIN 3]  H:2,SFWA-P,P
YARBRO, CHELSEA QUINN  BLOOD ROSES [SAINT-GERMAIN 12]  LP
YARBRO, CHELSEA QUINN  CANDLE FOR D’ARTAGNAN, A [SAINT-GERMAIN OLIVIA 3]  H:2,LP
YARBRO, CHELSEA QUINN  CAT’S CLAW [CHARLIE MOON 4]  P
YARBRO, CHELSEA QUINN  CAUTIONARY TALES  C/P,SFWA,P
YARBRO, CHELSEA QUINN  COME TWILIGHT [SAINT-GERMAIN 14]  H
YARBRO, CHELSEA QUINN  COMMUNION BLOOD [SAINT-GERMAIN 13]  H
YARBRO, CHELSEA QUINN  CROWN OF EMPIRE [CRISIS OF EMPIRE 4]  P
YARBRO, CHELSEA QUINN  CRUSADER’S TORCH [SAINT-GERMAIN OLIVIA 2]  C/P,SFWA-P,P,H:2
YARBRO, CHELSEA QUINN  DARK LIGHT [SHATTERED LIGHT 2]  P
YARBRO, CHELSEA QUINN  DARK OF THE SUN [SAINT-GERMAIN]  H
YARBRO, CHELSEA QUINN  DARKER JEWELS [SAINT-GERMAIN]  H,LP
YARBRO, CHELSEA QUINN  DEAD AND BURIED  C/P,P,H
YARBRO, CHELSEA QUINN  FALSE DAWN  H:2,SFWA,P
YARBRO, CHELSEA QUINN  FALSE NOTES [CHARLIE MOON 2]  P
YARBRO, CHELSEA QUINN  FEAST IN EXILE, A [SAINT-GERMAIN]  H
YARBRO, CHELSEA QUINN  FIRECODE  C/P,P
YARBRO, CHELSEA QUINN  FLAME IN BYZANTIUM, A [SAINT-GERMAIN OLIVIA 1]  H:2,SFWA-P,P
YARBRO, CHELSEA QUINN  GODFORSAKEN, THE  C/P:2,P
YARBRO, CHELSEA QUINN  HOTEL TRANSYLVANIA [SAINT-GERMAIN 1]  H:2,SFWA-P,P
YARBRO, CHELSEA QUINN  HYACINTHS  H,SFWA
YARBRO, CHELSEA QUINN  MANSIONS OF DARKNESS [SAINT-GERMAIN]  LP
YARBRO, CHELSEA QUINN  MIDNIGHT HARVEST [SAINT-GERMAIN]  H
YARBRO, CHELSEA QUINN  MORTAL GLAMOUR, A  C/P,P
YARBRO, CHELSEA QUINN  NOMADS  C/P,P
YARBRO, CHELSEA QUINN  OGILVIE, TALLANT & MOON  H
YARBRO, CHELSEA QUINN  OUT OF THE HOUSE OF LIFE [SAINT-GERMAIN 7]  H
YARBRO, CHELSEA QUINN  PALACE, THE [SAINT-GERMAIN 2]  H:2,SFWA-P,P
YARBRO, CHELSEA QUINN  PATH OF THE ECLIPSE [SAINT-GERMAIN 4]  H,SFWA,SFWA-P,P
YARBRO, CHELSEA QUINN  POISON FRUIT [CHARLIE MOON 3]  P
YARBRO, CHELSEA QUINN  SAINT-GERMAIN CHRONICLES, THE [SAINT-GERMAIN 6]  H,P
YARBRO, CHELSEA QUINN  SIGNS AND PORTENTS  H,P
YARBRO, CHELSEA QUINN  SINS OF OMISSION  C/P:2,P
YARBRO, CHELSEA QUINN  TAJI’S SYNDROME  C/P:2,P
YARBRO, CHELSEA QUINN  TEMPTING FATE [SAINT-GERMAIN 5]  H:2
YARBRO, CHELSEA QUINN  TIME OF THE FOURTH HORSEMAN  C/P,P,H,SFWA
YARBRO, CHELSEA QUINN  TO THE HIGH REDOUBT  C/P,P
YARBRO, CHELSEA QUINN  WRIT IN BLOOD [SAINT-GERMAIN 11]  H
YARIV, FRAN POKRAS  HALLOWING, THE  P
YATES, ALAN  CORIOLANUS, THE CHARIOT!  P
YATES, W. R.  DIASPORAH  C/P,P
YEFFETH, GLENN  NAVIGATING THE GOLDEN COMPASS  R/REF-H
YEFREMOV, IVAN  ANDROMEDA  SR-P
YEFREMOV, IVAN  LAND OF FOAM, THE  H
YEFREMOV, IVAN  SOCHINENIE (COLLECTED WORKS) I  FH
YEFREMOV, IVAN  SOCHINENIE (COLLECTED WORKS) II  FH
YEFREMOV, IVAN  SOCHINENIE (COLLECTED WORKS) III BOOK 1  FH
YEFREMOV, IVAN  SOCHINENIE (COLLECTED WORKS) III BOOK 2  FH
YEFREMOV, IVAN  STORIES  H
YEOVIL, JACK  BEASTS IN VELVET [WARHAMMER GENEVIEVE 3]  P
YEOVIL, JACK  COMEBACK TOUR [DARK FUTURE]  LP
YEOVIL, JACK  DEMON DOWNLOAD [DARK FUTURE DEMON DOWNLOAD 1]  LP
YEOVIL, JACK  DRACHENFELS [WARHAMMER GENEVIEVE 1]  P
YEOVIL, JACK  GENEVIEVE UNDEAD [WARHAMMER GENEVIEVE 2]  P
YEOVIL, JACK  KROKODIL TEARS [DARK FUTURE]  LP
YEOVIL, JACK  ROUTE 666 [DARK FUTURE DEMON DOWNLOAD 2]  P
YEOVIL, JACK  SILVER NAILS [WARHAMMER GENEVIEVE 4]  P
YEP, LAURENCE  DRAGON CAULDRON [DRAGON OF THE LOST SEA 3]  SFWA
YEP, LAURENCE  MONSTER MAKERS, INC.  P,SFWA
YEP, LAURENCE  SEADEMONS  C/P,SFWA,P,H
YEP, LAURENCE  SHADOW LORD [STAR TREK]  C/P:3,P
YEP, LAURENCE  SWEETWATER  H,LP
YERKA, JACK  MIND FIELDS  A
YERMAKOV, NICHOLAS  CLIQUE  P
YERMAKOV, NICHOLAS  EPIPHANY  C/P,P

The MITSFS Pinkdex  September 24, 2007  Page 430
AUTHOR  TITLE  CODE
YERMAKOV, NICHOLAS  FALL INTO DARKNESS  P
YERMAKOV, NICHOLAS  JEHAD  C/P:2,P
YERMAKOV, NICHOLAS  JOURNEY FROM FLESH  C/P,P
YERMAKOV, NICHOLAS  LAST COMMUNION  C/P:2,P
YI HUANG  COLLECTION OF SCIENCE FICTION (CHINESE)  FP
YOKE, CARL B.  ROGER ZELAZNY  R/REF-P
YOLEN, JANE  BOOKS OF GREAT ALTA, THE [JENNA 1,2]  H
YOLEN, JANE  BRIAR ROSE  C/P,P,H
YOLEN, JANE  CARDS OF GRIEF  C/P,P
YOLEN, JANE  COMMANDER TOAD AND THE PLANET OF THE GRAPES  LP
YOLEN, JANE  DOVE ISABEAU  H
YOLEN, JANE  DRAGONFIELD AND OTHER STORIES  C/P,P
YOLEN, JANE  DRAGON’S BLOOD [PIT DRAGONS 1]  C/P,P
YOLEN, JANE  HEART’S BLOOD [PIT DRAGONS 2]  C/P,P
YOLEN, JANE  HOBBY [YOUNG MERLIN 2]  H
YOLEN, JANE  MAGIC THREE OF SOLATIA, THE  P
YOLEN, JANE  ONE-ARMED QUEEN, THE [ALTA 3]  C/P,P
YOLEN, JANE  PIT DRAGON TRILOGY, THE [PIT DRAGONS 1,2,3]  H
YOLEN, JANE  SENDING OF DRAGONS, A [PIT DRAGONS 3]  H,P
YOLEN, JANE  SISTER EMILY’S LIGHTSHIP AND OTHER STORIES  H
YOLEN, JANE  SISTER LIGHT, SISTER DARK [ALTA 1]  P,SFWA-P
YOLEN, JANE  STORYTELLER  H
YOLEN, JANE  WHITE JENNA [ALTA 2]  C/P,P,H
YOLEN, JANE  XANADU  SR-HA
YOLEN, JANE  XANADU 2  HA,LPA
YOLEN, JANE  XANADU 3  HA,LPA

YOLEN, JANE/HAYDEN, PATRICK NIELSEN YEAR’S BEST SCIENCE FICTION AND FANTASY FOR TEENS,

FIRST ANNUAL COLLECTION, THE

LPA

YOLEN, JANE  ZOO 2000  HA
YORK, J. STEVEN  FORTRESS OF LIES [@MECHWARRIOR DARK AGE/BATTLETECH]  P
YORK, J. STEVEN  SCION OF THE SERPENT [@AGE OF CONAN ANOK #1]  P
YORK, J. STEVEN  TRIAL BY CHAOS [@MECHWARRIOR DARK AGE/BATTLETECH]  P

YORK, REBECCA/LAUREY, ROSEMARY/ THOMAS-SUNDSTROM, LINDA IMMORTAL BAD BOYS

LP

YORKE, CHRISTY  WISHING GARDEN, THE  P
YORKE, PRESTON  SPACE-TIME TASK FORCE  SR-P
YOUMANS, BRIAN  BEST OF THE REST 2  LPA
YOUMANS, BRIAN  BEST OF THE REST 3  LPA
YOUNG, JANINE ELLEN  BRIDGE, THE  P
YOUNG, JANINE ELLEN  CINDERBLOCK  P
YOUNG, JIM  ARMED MEMORY  C/P,P
YOUNG, JIM  FACE OF THE DEEP, THE  P
YOUNG, JIM  FACE OF THE DEEP (PROOF)  SR-LP
YOUNG, RED  SEX LIFE OF THE IMMORTALS  SR-L
YOUNG, ROBERT F.  COPPA PIENA DI STELLE, UNA  FP
YOUNG, ROBERT F.  ERIDAHN  C/P,P
YOUNG, ROBERT F.  GLASS OF STARS, A  H
YOUNG, ROBERT F.  LAST YGGDRASIL, THE  C/P,P
YOUNG, ROBERT F.  STARFINDER  C/P:2,P
YOUNG, ROBERT F.  VIZIER’S SECOND DAUGHTER, THE  P
YOUNG, ROBERT F.  WORLDS OF ROBERT F. YOUNG, THE  SR-P,H
YOUNG, ROY V.  CAPTAINS OUTRAGEOUS [CAPTAINS OUTRAGEOUS 1]  P
YOUNG, ROY V.  YOR’S REVENGE [CAPTAINS OUTRAGEOUS 2]  P
YOUNGER, JACK  DEMON  P
YOUNGER, JACK  REST IN AGONY  P
YOURCENAR, MARGUERITE  ORIENTAL TALES  H
YOUTH PUBLICATIONS  SCIENCE FICTION STORIES  SR-VLPA
YULSMAN, JERRY  ELLEANDER MORNING  C/P,P,H

AUTHOR  TITLE  CODE
ZACHARY, HUGH  GWEN, IN GREEN  P
ZACHERLEY  ZACHERLEY’S MIDNIGHT SNACKS  PA,C/PA
ZACHERLEY  ZACHERLEY’S VULTURE STEW  PA,C/PA
ZACKEY, CHRISTOPHER A.  LOST MYTHOLOGIES  SR-LP
ZAGAT, ARTHUR LEO  SEVEN OUT OF TIME  SR-H
ZAHN, TIMOTHY  ALLEGIANCE [@STAR WARS]  H:3
ZAHN, TIMOTHY  ANGELMASS  H:2,P
ZAHN, TIMOTHY  BACKLASH MISSION, THE [BLACKCOLLAR 2]  C/P,P
ZAHN, TIMOTHY  BLACKCOLLAR, THE [BLACKCOLLAR 1]  C/P:2,P
ZAHN, TIMOTHY  BLACKCOLLAR: THE JUDAS SOLUTION [BLACKCOLLAR 2]  H
ZAHN, TIMOTHY  CASCADE POINT  TD2,SFWA-TD2
ZAHN, TIMOTHY  CASCADE POINT (AND OTHER STORIES)  C/P,P,H
ZAHN, TIMOTHY  COBRA [COBRA 1]  C/P:2,P
ZAHN, TIMOTHY  COBRA BARGAIN [COBRA 3]  C/P:3,P
ZAHN, TIMOTHY  COBRA STRIKE [COBRA 2]  C/P:2,P
ZAHN, TIMOTHY  COBRA TRILOGY, THE [COBRA 1,2,3]  H
ZAHN, TIMOTHY  COBRAS TWO [COBRA 1,2]  C/P:2,SFWA-P,P
ZAHN, TIMOTHY  COMING OF AGE, A  C/P:2,SFWA,P
ZAHN, TIMOTHY  CONQUERORS’ HERITAGE [CONQUERORS 2]  C/P,P,H
ZAHN, TIMOTHY  CONQUERORS’ LEGACY [CONQUERORS 3]  H,P
ZAHN, TIMOTHY  CONQUERORS’ PRIDE [CONQUERORS 1]  C/P,P,H
ZAHN, TIMOTHY  DARK FORCE RISING [STAR WARS]  H:2,P
ZAHN, TIMOTHY  DEADMAN SWITCH  C/P:2,P
ZAHN, TIMOTHY  DISTANT FRIENDS AND OTHERS  C/P,SFWA-P,P
ZAHN, TIMOTHY  DRAGON AND HERDSMAN [DRAGONBACK 4]  H
ZAHN, TIMOTHY  DRAGON AND THIEF [DRAGONBACK 1]  H
ZAHN, TIMOTHY  DRAGONBACK BARGAIN [DRAGONBACK 1,2,3]  H
ZAHN, TIMOTHY  GREEN AND THE GRAY, THE  H
ZAHN, TIMOTHY  HEIR TO THE EMPIRE [STAR WARS]  H:2,P
ZAHN, TIMOTHY  ICARUS HUNT, THE  H,P
ZAHN, TIMOTHY  LAST COMMAND, THE [STAR WARS]  H:2,P
ZAHN, TIMOTHY  MANTA’S GIFT  H
ZAHN, TIMOTHY  NIGHT TRAIN TO RIGEL  H:2
ZAHN, TIMOTHY  OUTBOUND FLIGHT [@STAR WARS]  H
ZAHN, TIMOTHY  SPECTER OF THE PAST [STAR WARS]  P
ZAHN, TIMOTHY  SPINNERET  C/P,P,H
ZAHN, TIMOTHY  SURVIVOR’S QUEST [STAR WARS]  H
ZAHN, TIMOTHY  TIME BOMB  C/P:2,SFWA-P,P
ZAHN, TIMOTHY  TRIPLET  C/P:2,P
ZAHN, TIMOTHY  VISION OF THE FUTURE [STAR WARS]  C/P,P
ZAHN, TIMOTHY  WARHORSE  C/P:2,SFWA-P,P
ZAKOUR, JOHN/GANEM, LAWRENCE  DOOMSDAY BRUNETTE, THE [ZACHARY NIXON JOHNSON 2]  C/P,P
ZAKOUR, JOHN  FROST-HAIRED VIXEN, THE [ZACHARY NIXON JOHNSON 4]  P
ZAKOUR, JOHN/GANEM, LAWRENCE  PLUTONIUM BLONDE, THE [ZACHARY NIXON JOHNSON 1]  C/P,P
ZAKOUR, JOHN/GANEM, LAWRENCE  RADIOACTIVE REDHEAD, THE [ZACHARY NIXON JOHNSON 3]  P
ZAMIATIN, EUGENE  WE (GREGORY ZILBOORG TRANSLATION)  C/P:2,P
ZAMYATIN, YEVGENY  WE (MIRRA GINSBURG TRANSLATION)  C/P,P,H
ZAREM, LEWIS  GREEN MAN FROM SPACE, THE  H
ZAREM, LEWIS/CANTWELL, RAY  SUPERJET  H
ZEBROWSKI, GEORGE  ASHES AND STARS  C/P,P
ZEBROWSKI, GEORGE  BLACK POCKETS  HA
ZEBROWSKI, GEORGE  BRUTE ORBITS  P
ZEBROWSKI, GEORGE  CAVE OF STARS  P
ZEBROWSKI, GEORGE  MACROLIFE  C/P,SFWA,P
ZEBROWSKI, GEORGE  MONADIC UNIVERSE, THE  C/P,P
ZEBROWSKI, GEORGE  NEBULA AWARDS 20  HA,LPA
ZEBROWSKI, GEORGE  NEBULA AWARDS 21  LPA
ZEBROWSKI, GEORGE  NEBULA AWARDS 22  LPA
ZEBROWSKI, GEORGE  OMEGA POINT, THE  C/P:2,P
ZEBROWSKI, GEORGE  OMEGA POINT TRILOGY, THE  P
ZEBROWSKI, GEORGE  STAR WEB, THE  SR-P
ZEBROWSKI, GEORGE  STARS WILL SPEAK, THE [SUNSPACE 2]  H
ZEBROWSKI, GEORGE  STRANGER SUNS  C/P,P
ZEBROWSKI, GEORGE  SUNSPACER [SUNSPACE 1]  LP
ZEBROWSKI, GEORGE  SYNERGY 1  PA
ZEBROWSKI, GEORGE  SYNERGY 2  PA
ZEBROWSKI, GEORGE  SYNERGY 3  PA
ZEBROWSKI, GEORGE  SYNERGY 4  PA
ZEBROWSKI, GEORGE  TOMORROW TODAY  SR-LPA
ZEDDIES, ANN TONSOR  DEATHGIFT  P
ZEDDIES, ANN TONSOR  SKY ROAD  P
ZEDDIES, ANN TONSOR  STEEL HELIX  P
ZEIGFREID, KARL  ANDROID  SR-P
ZEIGFREID, KARL  ATOMIC NEMESIS  SR-P

The MITSFS Pinkdex  September 24, 2007  Page 432
AUTHOR  TITLE  CODE
ZEIGFREID, KARL  BARRIER 346  SR-P,C/P
ZEIGFREID, KARL  CHARIOT INTO TIME  SR-P
ZEIGFREID, KARL  ESCAPE TO INFINITY  SR-P
ZEIGFREID, KARL  GIRL FROM TOMORROW, THE  SR-P
ZEIGFREID, KARL  GODS OF DARKNESS  SR-P
ZEIGFREID, KARL  NO WAY BACK  SR-P
ZEIGFREID, KARL  PROJECTION INFINITY  SR-P,C/P
ZEIGFREID, KARL  RADAR ALERT  SR-P,C/P
ZEIGFREID, KARL  WALK THROUGH TO-MORROW  SR-P,C/P
ZEIGFREID, KARL  WORLD OF TOMORROW  SR-P
ZEIGFREID, KARL  WORLD THAT NEVER WAS, THE  SR-P
ZEIGFREID, KARL  ZERO MINUS X  SR-P
ZELAZNY, ROGER/SABERHAGEN, FRED  BLACK THRONE, THE  C/P,SFWA-P,P,H
ZELAZNY, ROGER  BLOOD OF AMBER [AMBER 7]  C/P,H,@P,SFWA
ZELAZNY, ROGER  BRIDGE OF ASHES  C/P:2,P
ZELAZNY, ROGER/SHECKLEY, ROBERT  BRING ME THE HEAD OF PRINCE CHARMING [AZZIE 1]  H:2,LP,P
ZELAZNY, ROGER  CHANGELING [POL DETSON 1]  C/P,H,LP
ZELAZNY, ROGER  CHANGING LAND, THE  C/P,P,H
ZELAZNY, ROGER  CHRONICLES OF AMBER 1, THE [AMBER 1,2,3]  H:2
ZELAZNY, ROGER  CHRONICLES OF AMBER 2, THE [AMBER 4,5]  H:2
ZELAZNY, ROGER/SABERHAGEN, FRED  COILS  C/P:2,SFWA-P,LP
ZELAZNY, ROGER  COURTS OF CHAOS, THE [AMBER 5]  C/P,@P,H,SFWA
ZELAZNY, ROGER  CREATURES OF LIGHT AND DARKNESS  C/P:3,P
ZELAZNY, ROGER  DAMNATION ALLEY  C/P:2,P
ZELAZNY, ROGER  DARK TRAVELING, A  C/P,P,H
ZELAZNY, ROGER  DILVISH, THE DAMNED  C/P,P,H
ZELAZNY, ROGER/LINDSKOLD, JANE  DONNERJACK  H:2,P

ZELAZNY, ROGER DOORS OF HIS FACE, THE LAMPS OF HIS MOUTH, AND OTHER STORIES,

THE

C/P,SFWA,P,H

ZELAZNY, ROGER  DOORWAYS IN THE SAND  C/P,P,H,SFWA
ZELAZNY, ROGER  DREAM MASTER, THE  C/P:2,SFWA-P,P
ZELAZNY, ROGER  EYE OF CAT  H:2,SFWA,P
ZELAZNY, ROGER/SHECKLEY, ROBERT  FARCE TO BE RECKONED WITH, A [AZZIE 3]  H,LP
ZELAZNY, ROGER/THOMAS, THOMAS T.  FLARE  C/P:2,SFWA-P,P
ZELAZNY, ROGER  FOUR FOR TOMORROW  C/P:2,SFWA-P,P
ZELAZNY, ROGER  FROST AND FIRE  C/P:2,P,H
ZELAZNY, ROGER  GRAVEYARD HEART, THE  TD24
ZELAZNY, ROGER  GREAT BOOK OF AMBER, THE [AMBER 1,2,3,4,5,6,7,8,9,10]  LP
ZELAZNY, ROGER  GUNS OF AVALON, THE [AMBER 2]  C/P:3,@P
ZELAZNY, ROGER  HAND OF OBERON, THE [AMBER 4]  C/P,@P,H,SFWA
ZELAZNY, ROGER  HE WHO SHAPES  SFWA-TD12,TD12
ZELAZNY, ROGER  HOME IS THE HANGMAN  TD21,SFWA-TD21
ZELAZNY, ROGER/SHECKLEY, ROBERT  IF AT FAUST YOU DON’T SUCCEED [AZZIE 2]  P,LP
ZELAZNY, ROGER/SHECKLEY, ROBERT  IF AT FAUST YOU DON’T SUCCEED (PROOF) [AZZIE 2]  SR-LP
ZELAZNY, ROGER/MORROW, GRAY  ILLUSTRATED ROGER ZELAZNY, THE  CX
ZELAZNY, ROGER  ISLE OF THE DEAD  C/P:2,SFWA-P,P
ZELAZNY, ROGER  JACK OF SHADOWS  C/P,P,H,SFWA
ZELAZNY, ROGER  KNIGHT OF SHADOWS [AMBER 9]  C/P,H,@P,SFWA
ZELAZNY, ROGER  LAST DEFENDER OF CAMELOT, THE  C/P,P,H
ZELAZNY, ROGER/LINDSKOLD, JANE  LORD DEMON  H,P
ZELAZNY, ROGER  LORD OF LIGHT  C/P,P,H
ZELAZNY, ROGER  MADWAND [POL DETSON 2]  C/P,H,LP
ZELAZNY, ROGER  MANNA FROM HEAVEN  H
ZELAZNY, ROGER/THOMAS, THOMAS T.  MASK OF LOKI, THE  C/P:2,SFWA-P,P
ZELAZNY, ROGER  MY NAME IS LEGION  C/P:2,P
ZELAZNY, ROGER  NEBULA AWARD STORIES 3  PA,C/PA
ZELAZNY, ROGER  NIGHT IN THE LONESOME OCTOBER, A  C/P,SFWA,P,H
ZELAZNY, ROGER  NINE PRINCES IN AMBER [AMBER 1]  C/P:2,@P
ZELAZNY, ROGER  PRINCE OF CHAOS [AMBER 10]  H:2,@P
ZELAZNY, ROGER  ROADMARKS  C/P:2,P
ZELAZNY, ROGER/RANDALL, NEIL  ROGER ZELAZNY’S VISUAL GUIDE TO CASTLE AMBER  R/REF-P,C/REF-H:3
ZELAZNY, ROGER  ROSE FOR ECCLESIASTES, A  P
ZELAZNY, ROGER  SIGN OF CHAOS [AMBER 8]  C/P,H,@P
ZELAZNY, ROGER  SIGN OF THE UNICORN [AMBER 3]  C/P:2,@P
ZELAZNY, ROGER  THIS IMMORTAL  SFWA-P:2,H:2,P
ZELAZNY, ROGER  TO DIE IN ITALBAR  H:2,P,SFWA
ZELAZNY, ROGER  TODAY WE CHOOSE FACES  C/P:2,P
ZELAZNY, ROGER  TODAY WE CHOOSE FACES/BRIDGE OF ASHES  C/P,P
ZELAZNY, ROGER  TRUMPS OF DOOM [AMBER 6]  C/P,H,@P,SFWA
ZELAZNY, ROGER  UNICORN VARIATIONS  H:2,P
ZELAZNY, ROGER  WARRIORS OF BLOOD AND DREAM  C/PA,PA
ZELAZNY, ROGER  WHEEL OF FORTUNE  C/PA,PA
ZELAZNY, ROGER  WIZARD WORLD [POL DETSON 1,2]  C/P:2,P
ZELENETZ, ALAN/PIERARD, JOHN  STEVEN BRUST’S JHEREG [DRAGAERA VLAD TALTOS]  CX
ZERWICK, CHLOE/BROWN, HARRISON  CASSIOPEIA AFFAIR, THE  P

The MITSFS Pinkdex  September 24, 2007  Page 433
AUTHOR  TITLE  CODE
ZETFORD, TULLY  BOOSTED MAN, THE [@HOOK #2]  C/P,P
ZETFORD, TULLY  STAR CITY [@HOOK #3]  P
ZETFORD, TULLY  VIRILITY GENE, THE [@HOOK #4]  C/P,P
ZETFORD, TULLY  WHIRLPOOL OF STARS [@HOOK #1]  P
ZETTEL, SARAH  FIREBIRD’S VENGEANCE, THE [ISAVALTA 3]  H
ZETTEL, SARAH  FOOL’S WAR  C/P:2,P
ZETTEL, SARAH  FOR CAMELOT’S HONOR  LP
ZETTEL, SARAH  IN CAMELOT’S SHADOW  LP
ZETTEL, SARAH  KINGDOM OF CAGES  H,P
ZETTEL, SARAH  PLAYING GOD  H,P
ZETTEL, SARAH  QUIET INVASION, THE  H:2,P
ZETTEL, SARAH  RECLAMATION  C/P,P,H
ZETTEL, SARAH  SORCERER’S TREASON, A [ISAVALTA 1]  H
ZETTEL, SARAH  SWORD OF THE DECEIVER [ISAVALTA 4]  H
ZETTEL, SARAH  USURPER’S CROWN, THE [ISAVALTA 2]  H
ZICREE, MARC SCOTT/HAMBLY, BARBARA  MAGIC TIME [MAGIC TIME]  H

ZICREE, MARC SCOTT/ BOHNHOFF, MAYA KAATHRYN MAGIC TIME: ANGELFIRE [MAGIC TIME]

H

ZICREE, MARC SCOTT  TWILIGHT ZONE COMPANION, THE (FIRST EDITION)  [C/RAN-H]
ZICREE, MARC SCOTT  TWILIGHT ZONE COMPANION, THE (SECOND EDITION)  R/REF-H
ZIEROLD, NORMAN  SKYSCRAPER DOOM, THE  SFWA
ZIM, HERBERT/BAKER, ROBERT H.  STARS  [C/RAN-P]
ZIMMER, PAUL EDWIN/DE CLES, JON  BLOOD OF THE COLYN MUR  P
ZIMMER, PAUL EDWIN  GATHERING OF HEROES, A [@DARK BORDER]  C/P,P
ZIMMER, PAUL EDWIN  INGULF THE MAD [DARK BORDER]  C/P,SFWA-P,P
ZIMMER, PAUL EDWIN  KING CHONDOS’ RIDE [@DARK BORDER #2]  C/P:2,P
ZIMMER, PAUL EDWIN  LOST PRINCE, THE [@DARK BORDER #1]  C/P:2,P
ZIMPEL, LLOYD  MEETING THE BEAR: JOURNAL OF THE BLACK WARS  C/P,P
ZINDELL, DAVID  BROKEN GOD, THE [REQUIEM FOR HOMO SAPIENS 2]  C/P:2,P
ZINDELL, DAVID  LIGHTSTONE, THE [LIGHTSTONE 1]  H
ZINDELL, DAVID  NEVERNESS [REQUIEM FOR HOMO SAPIENS 1]  C/P:2,P,H
ZINDELL, DAVID  SILVER SWORD, THE [LIGHTSTONE 2]  H
ZINDELL, DAVID  WAR IN HEAVEN [REQUIEM FOR HOMO SAPIENS 4]  P
ZINDELL, DAVID  WILD, THE [REQUIEM FOR HOMO SAPIENS 3]  P
ZIPES, JACK  DON’T BET ON THE PRINCE  LP
ZOLAR  ENCYCLOPEDIA OF ANCIENT AND FORBIDDEN KNOWLEDGE, THE  [C/RAN-P]
ZOSS, JOEL  CHRONICLE  C/P:2,P
ZUBRIN, ROBERT  FIRST LANDING  H
ZUROY, MICHAEL  SECOND DEATH  SFWA